Guile Reference Manual

Edition 2.0.11, revision 1, for use with Guile 2.0.11

The Guile Developers

This manual documents Guile version 2.0.11. Copyright (C) 1996, 1997, 2000, 2001, 2002, 2003, 2004, 2005, 2009, 2010, 2011, 2012, 2013, 2014 Free Software Foundation. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License.”

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Contributors to this Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 The Guile License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Guile and Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Combining with C Code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Guile and the GNU Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Interactive Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Supporting Multiple Languages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 Obtaining and Installing Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.7 Organisation of this Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.8 Typographical Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 Hello Guile! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.1 Running Guile Interactively . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2 Running Guile Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3 Linking Guile into Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4 Writing Guile Extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.5 Using the Guile Module System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2.5.1 Using Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.5.2 Writing new Modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.3 Putting Extensions into Modules . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.6 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3 Hello Scheme!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.1 Data Types, Values and Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.1 Latent Typing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.2 Values and Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.3 Defining and Setting Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3.2 The Representation and Use of Procedures . . . . . . . . . . . . . . . . . . . . 17 3.2.1 Procedures as Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2.2 Simple Procedure Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.2.3 Creating and Using a New Procedure . . . . . . . . . . . . . . . . . . . . . 19 3.2.4 Lambda Alternatives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3.3 Expressions and Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3.3.1 Evaluating Expressions and Executing Programs. . . . . . . . . . 21 3.3.1.1 Evaluating Literal Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.3.1.2 Evaluating a Variable Reference . . . . . . . . . . . . . . . . . . . . . 22 3.3.1.3 Evaluating a Procedure Invocation Expression . . . . . . . 22 3.3.1.4 Evaluating Special Syntactic Expressions . . . . . . . . . . . . 23

3.3.2 Tail calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.3.3 Using the Guile REPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.3.4 Summary of Common Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.4 The Concept of Closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.4.1 Names, Locations, Values and Environments . . . . . . . . . . . . . . 26 3.4.2 Local Variables and Environments . . . . . . . . . . . . . . . . . . . . . . . . 27 3.4.3 Environment Chaining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.4.4 Lexical Scope . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3.4.4.1 An Example of Non-Lexical Scoping . . . . . . . . . . . . . . . . . 28 3.4.5 Closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.4.6 Example 1: A Serial Number Generator . . . . . . . . . . . . . . . . . . 31 3.4.7 Example 2: A Shared Persistent Variable . . . . . . . . . . . . . . . . . 31 3.4.8 Example 3: The Callback Closure Problem . . . . . . . . . . . . . . . 32 3.4.9 Example 4: Object Orientation . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.5 Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4 Programming in Scheme . . . . . . . . . . . . . . . . . . . . . . 35

4.1 Guile’s Implementation of Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2 Invoking Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.2.1 Command-line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2.2 Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

4.3 Guile Scripting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.3.1 The Top of a Script File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.3.2 The Meta Switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.3.3 Command Line Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.3.4 Scripting Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.4 Using Guile Interactively . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.4.1 The Init File, ~/.guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.4.2 Readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.4.3 Value History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.4.4 REPL Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

4.4.4.1 Help Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.4.4.2 Module Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.4.4.3 Language Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.4.4.4 Compile Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.4.4.5 Profile Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.4.4.6 Debug Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.4.4.7 Inspect Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.4.4.8 System Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

4.4.5 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.4.6 Interactive Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

4.5 Using Guile in Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.6 Using Guile Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.7 Installing Site Packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

5 Programming in C . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

5.1 Parallel Installations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.2 Linking Programs With Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

5.2.1 Guile Initialization Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.2.2 A Sample Guile Main Program . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.2.3 Building the Example with Make . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.2.4 Building the Example with Autoconf . . . . . . . . . . . . . . . . . . . . . 59

5.3 Linking Guile with Libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

5.3.1 A Sample Guile Extension. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.4 General concepts for using libguile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

5.4.1 Dynamic Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.4.2 Garbage Collection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 5.4.3 Control Flow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.4.4 Asynchronous Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.4.5 Multi-Threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

5.5 Defining New Types (Smobs). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 5.5.1 Describing a New Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 5.5.2 Creating Smob Instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 5.5.3 Type checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.5.4 Garbage Collecting Smobs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 5.5.5 Remembering During Operations . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.5.6 Double Smobs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.5.7 The Complete Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

5.6 Function Snarfing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 5.7 An Overview of Guile Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

5.7.1 How One Might Extend Dia Using Guile. . . . . . . . . . . . . . . . . . 83 5.7.1.1 Deciding Why You Want to Add Guile . . . . . . . . . . . . . . 84 5.7.1.2 Four Steps Required to Add Guile . . . . . . . . . . . . . . . . . . . 84 5.7.1.3 How to Represent Dia Data in Scheme . . . . . . . . . . . . . . . 85 5.7.1.4 Writing Guile Primitives for Dia . . . . . . . . . . . . . . . . . . . . . 86 5.7.1.5 Providing a Hook for the Evaluation of Scheme Code

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 5.7.1.6 Top-level Structure of Guile-enabled Dia . . . . . . . . . . . . . 87 5.7.1.7 Going Further with Dia and Guile . . . . . . . . . . . . . . . . . . . 88

5.7.2 Why Scheme is More Hackable Than C . . . . . . . . . . . . . . . . . . . 90 5.7.3 Example: Using Guile for an Application Testbed. . . . . . . . . 90 5.7.4 A Choice of Programming Options . . . . . . . . . . . . . . . . . . . . . . . 91

5.7.4.1 What Functionality is Already Available? . . . . . . . . . . . . 91 5.7.4.2 Functional and Performance Constraints . . . . . . . . . . . . . 92 5.7.4.3 Your Preferred Programming Style . . . . . . . . . . . . . . . . . . 92 5.7.4.4 What Controls Program Execution? . . . . . . . . . . . . . . . . . 92

5.7.5 How About Application Users? . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.8 Autoconf Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

5.8.1 Autoconf Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 5.8.2 Autoconf Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 5.8.3 Using Autoconf Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6 API Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.1 Overview of the Guile API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 6.2 Deprecation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.3 The SCM Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.4 Initializing Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.5 Snarfing Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 6.6 Simple Generic Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

6.6.1 Booleans . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 6.6.2 Numerical data types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

6.6.2.1 Scheme’s Numerical “Tower” . . . . . . . . . . . . . . . . . . . . . . . 105 6.6.2.2 Integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.6.2.3 Real and Rational Numbers . . . . . . . . . . . . . . . . . . . . . . . . 110 6.6.2.4 Complex Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 6.6.2.5 Exact and Inexact Numbers . . . . . . . . . . . . . . . . . . . . . . . . 113 6.6.2.6 Read Syntax for Numerical Data . . . . . . . . . . . . . . . . . . . 114 6.6.2.7 Operations on Integer Values . . . . . . . . . . . . . . . . . . . . . . . 116 6.6.2.8 Comparison Predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 6.6.2.9 Converting Numbers To and From Strings . . . . . . . . . . 118 6.6.2.10 Complex Number Operations. . . . . . . . . . . . . . . . . . . . . . 118 6.6.2.11 Arithmetic Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 6.6.2.12 Scientific Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 6.6.2.13 Bitwise Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 6.6.2.14 Random Number Generation . . . . . . . . . . . . . . . . . . . . . . 127

6.6.3 Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.6.4 Character Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

6.6.4.1 Character Set Predicates/Comparison . . . . . . . . . . . . . . 133 6.6.4.2 Iterating Over Character Sets . . . . . . . . . . . . . . . . . . . . . . 134 6.6.4.3 Creating Character Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 6.6.4.4 Querying Character Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 6.6.4.5 Character-Set Algebra. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 6.6.4.6 Standard Character Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

6.6.5 Strings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 6.6.5.1 String Read Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 6.6.5.2 String Predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 6.6.5.3 String Constructors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 6.6.5.4 List/String conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 6.6.5.5 String Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 6.6.5.6 String Modification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 6.6.5.7 String Comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 6.6.5.8 String Searching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 6.6.5.9 Alphabetic Case Mapping . . . . . . . . . . . . . . . . . . . . . . . . . . 154 6.6.5.10 Reversing and Appending Strings. . . . . . . . . . . . . . . . . . 155 6.6.5.11 Mapping, Folding, and Unfolding . . . . . . . . . . . . . . . . . . 156 6.6.5.12 Miscellaneous String Operations . . . . . . . . . . . . . . . . . . . 157 6.6.5.13 Representing Strings as Bytes . . . . . . . . . . . . . . . . . . . . . 158 6.6.5.14 Conversion to/from C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 6.6.5.15 String Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

6.6.6 Bytevectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 6.6.6.1 Endianness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 6.6.6.2 Manipulating Bytevectors . . . . . . . . . . . . . . . . . . . . . . . . . . 164 6.6.6.3 Interpreting Bytevector Contents as Integers . . . . . . . . 165 6.6.6.4 Converting Bytevectors to/from Integer Lists . . . . . . . 168 6.6.6.5 Interpreting Bytevector Contents as Floating Point

Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 6.6.6.6 Interpreting Bytevector Contents as Unicode Strings

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 6.6.6.7 Accessing Bytevectors with the Array API . . . . . . . . . . 170 6.6.6.8 Accessing Bytevectors with the SRFI-4 API . . . . . . . . 170

6.6.7 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 6.6.7.1 Symbols as Discrete Data . . . . . . . . . . . . . . . . . . . . . . . . . . 171 6.6.7.2 Symbols as Lookup Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 6.6.7.3 Symbols as Denoting Variables . . . . . . . . . . . . . . . . . . . . . 173 6.6.7.4 Operations Related to Symbols . . . . . . . . . . . . . . . . . . . . . 173 6.6.7.5 Function Slots and Property Lists . . . . . . . . . . . . . . . . . . 176 6.6.7.6 Extended Read Syntax for Symbols. . . . . . . . . . . . . . . . . 178 6.6.7.7 Uninterned Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

6.6.8 Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 6.6.8.1 Why Use Keywords? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 6.6.8.2 Coding With Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 6.6.8.3 Keyword Read Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 6.6.8.4 Keyword Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

6.6.9 “Functionality-Centric” Data Types . . . . . . . . . . . . . . . . . . . . . 184 6.7 Compound Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

6.7.1 Pairs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 6.7.2 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

6.7.2.1 List Read Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 6.7.2.2 List Predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 6.7.2.3 List Constructors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 6.7.2.4 List Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 6.7.2.5 Append and Reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 6.7.2.6 List Modification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 6.7.2.7 List Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 6.7.2.8 List Mapping. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

6.7.3 Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 6.7.3.1 Read Syntax for Vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 6.7.3.2 Dynamic Vector Creation and Validation . . . . . . . . . . . 194 6.7.3.3 Accessing and Modifying Vector Contents. . . . . . . . . . . 195 6.7.3.4 Vector Accessing from C . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 6.7.3.5 Uniform Numeric Vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . 197

6.7.4 Bit Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 6.7.5 Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200

6.7.5.1 Array Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 6.7.5.2 Array Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 6.7.5.3 Shared Arrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 6.7.5.4 Accessing Arrays from C . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

6.7.6 VLists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 6.7.7 Record Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 6.7.8 SRFI-9 Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

Non-toplevel Record Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 Custom Printers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 Functional “Setters” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

6.7.9 Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 6.7.10 Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

6.7.10.1 Vtables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 6.7.10.2 Structure Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 6.7.10.3 Vtable Contents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 6.7.10.4 Meta-Vtables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 6.7.10.5 Vtable Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 6.7.10.6 Tail Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

6.7.11 Dictionary Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 6.7.12 Association Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

6.7.12.1 Alist Key Equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 6.7.12.2 Adding or Setting Alist Entries . . . . . . . . . . . . . . . . . . . . 228 6.7.12.3 Retrieving Alist Entries . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 6.7.12.4 Removing Alist Entries. . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 6.7.12.5 Sloppy Alist Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 6.7.12.6 Alist Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

6.7.13 VList-Based Hash Lists or “VHashes” . . . . . . . . . . . . . . . . . . 234 6.7.14 Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

6.7.14.1 Hash Table Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 6.7.14.2 Hash Table Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237

6.8 Smobs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 6.9 Procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244

6.9.1 Lambda: Basic Procedure Creation . . . . . . . . . . . . . . . . . . . . . . 244 6.9.2 Primitive Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 6.9.3 Compiled Procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 6.9.4 Optional Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248

6.9.4.1 lambda* and define*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 6.9.4.2 (ice-9 optargs) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250

6.9.5 Case-lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 6.9.6 Higher-Order Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 6.9.7 Procedure Properties and Meta-information . . . . . . . . . . . . . 254 6.9.8 Procedures with Setters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 6.9.9 Inlinable Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256

6.10 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 6.10.1 Defining Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 6.10.2 Syntax-rules Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

6.10.2.1 Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 6.10.2.2 Hygiene . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 6.10.2.3 Shorthands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 6.10.2.4 Reporting Syntax Errors in Macros . . . . . . . . . . . . . . . . 262 6.10.2.5 Specifying a Custom Ellipsis Identifier . . . . . . . . . . . . . 262 6.10.2.6 Further Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263

6.10.3 Support for the syntax-case System. . . . . . . . . . . . . . . . . . . 263 6.10.3.1 Why syntax-case? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 6.10.3.2 Custom Ellipsis Identifiers for syntax-case Macros . . 267

6.10.4 Syntax Transformer Helpers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 6.10.5 Lisp-style Macro Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 6.10.6 Identifier Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 6.10.7 Syntax Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 6.10.8 Eval-when . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 6.10.9 Internal Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275

6.11 General Utility Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 6.11.1 Equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 6.11.2 Object Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 6.11.3 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 6.11.4 Copying Deep Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 6.11.5 General String Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 6.11.6 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281

6.11.6.1 Hook Usage by Example . . . . . . . . . . . . . . . . . . . . . . . . . . 281 6.11.6.2 Hook Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 6.11.6.3 Handling Scheme-level hooks from C code . . . . . . . . . 284 6.11.6.4 Hooks For C Code.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 6.11.6.5 Hooks for Garbage Collection . . . . . . . . . . . . . . . . . . . . . 285 6.11.6.6 Hooks into the Guile REPL . . . . . . . . . . . . . . . . . . . . . . . 286

6.12 Definitions and Variable Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 6.12.1 Top Level Variable Definitions. . . . . . . . . . . . . . . . . . . . . . . . . . 286 6.12.2 Local Variable Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 6.12.3 Internal definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 6.12.4 Querying variable bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 6.12.5 Binding multiple return values . . . . . . . . . . . . . . . . . . . . . . . . . 290

6.13 Controlling the Flow of Program Execution . . . . . . . . . . . . . . . . . 290 6.13.1 Sequencing and Splicing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 6.13.2 Simple Conditional Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . 292 6.13.3 Conditional Evaluation of a Sequence of Expressions. … 293 6.13.4 Iteration mechanisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 6.13.5 Prompts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296

6.13.5.1 Prompt Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 6.13.5.2 Shift, Reset, and All That . . . . . . . . . . . . . . . . . . . . . . . . . 298

6.13.6 Continuations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 6.13.7 Returning and Accepting Multiple Values . . . . . . . . . . . . . . 301 6.13.8 Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303

6.13.8.1 Exception Terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 6.13.8.2 Catching Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 6.13.8.3 Throw Handlers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 6.13.8.4 Throwing Exceptions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 6.13.8.5 How Guile Implements Exceptions. . . . . . . . . . . . . . . . . 308

6.13.9 Procedures for Signaling Errors . . . . . . . . . . . . . . . . . . . . . . . . 309 6.13.10 Dynamic Wind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 6.13.11 How to Handle Errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313

6.13.11.1 C Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 6.13.11.2 Signalling Type Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

6.13.12 Continuation Barriers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 6.14 Input and Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316

6.14.1 Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 6.14.2 Reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 6.14.3 Writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 6.14.4 Closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 6.14.5 Random Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 6.14.6 Line Oriented and Delimited Text . . . . . . . . . . . . . . . . . . . . . . 322 6.14.7 Block reading and writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 6.14.8 Default Ports for Input, Output and Errors. . . . . . . . . . . . . 325 6.14.9 Types of Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326

6.14.9.1 File Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 6.14.9.2 String Ports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 6.14.9.3 Soft Ports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 6.14.9.4 Void Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332

6.14.10 R6RS I/O Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 6.14.10.1 File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 6.14.10.2 File Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 6.14.10.3 Buffer Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 6.14.10.4 Transcoders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 6.14.10.5 The End-of-File Object . . . . . . . . . . . . . . . . . . . . . . . . . . 336 6.14.10.6 Port Manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 6.14.10.7 Input Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 6.14.10.8 Binary Input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 6.14.10.9 Textual Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 6.14.10.10 Output Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 6.14.10.11 Binary Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 6.14.10.12 Textual Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344

6.14.11 Using and Extending Ports in C. . . . . . . . . . . . . . . . . . . . . . . 345 6.14.11.1 C Port Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 6.14.11.2 Port Implementation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346

6.14.12 Handling of Unicode byte order marks. . . . . . . . . . . . . . . . . 348 6.15 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

6.15.1 Regexp Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 350 6.15.2 Match Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 6.15.3 Backslash Escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355

6.16 LALR(1) Parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 6.17 Reading and Evaluating Scheme Code. . . . . . . . . . . . . . . . . . . . . . . 357

6.17.1 Scheme Syntax: Standard and Guile Extensions . . . . . . . . 357 6.17.1.1 Expression Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 6.17.1.2 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358 6.17.1.3 Block Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 6.17.1.4 Case Sensitivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 6.17.1.5 Keyword Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 6.17.1.6 Reader Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360

6.17.2 Reading Scheme Code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 6.17.3 Writing Scheme Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 6.17.4 Procedures for On the Fly Evaluation . . . . . . . . . . . . . . . . . . 362 6.17.5 Compiling Scheme Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 6.17.6 Loading Scheme Code from File . . . . . . . . . . . . . . . . . . . . . . . . 366 6.17.7 Load Paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 6.17.8 Character Encoding of Source Files. . . . . . . . . . . . . . . . . . . . . 370 6.17.9 Delayed Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 6.17.10 Local Evaluation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 6.17.11 Local Inclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 6.17.12 REPL Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 6.17.13 Cooperative REPL Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374

6.18 Memory Management and Garbage Collection . . . . . . . . . . . . . . . 374 6.18.1 Function related to Garbage Collection . . . . . . . . . . . . . . . . . 374 6.18.2 Memory Blocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 376

6.18.2.1 Upgrading from scm must malloc et al. . . . . . . . . . . . 378 6.18.3 Weak References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378

6.18.3.1 Weak hash tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 6.18.3.2 Weak vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379

6.18.4 Guardians . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 6.19 Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381

6.19.1 General Information about Modules . . . . . . . . . . . . . . . . . . . . 381 6.19.2 Using Guile Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 6.19.3 Creating Guile Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 6.19.4 Modules and the File System. . . . . . . . . . . . . . . . . . . . . . . . . . . 386 6.19.5 R6RS Version References. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 6.19.6 R6RS Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 6.19.7 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 6.19.8 Module System Reflection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 6.19.9 Accessing Modules from C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 6.19.10 provide and require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 6.19.11 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395

6.20 Foreign Function Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 6.20.1 Foreign Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 6.20.2 Foreign Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 6.20.3 C Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 6.20.4 Modules and Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 6.20.5 Foreign Pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402

6.20.5.1 Foreign Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 6.20.5.2 Foreign Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 6.20.5.3 Void Pointers and Byte Access . . . . . . . . . . . . . . . . . . . . 404 6.20.5.4 Foreign Structs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406

6.20.6 Dynamic FFI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 6.21 Threads, Mutexes, Asyncs and Dynamic Roots . . . . . . . . . . . . . . 410

6.21.1 Arbiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 6.21.2 Asyncs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411

6.21.2.1 System asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 6.21.2.2 User asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412

6.21.3 Threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 6.21.4 Mutexes and Condition Variables. . . . . . . . . . . . . . . . . . . . . . . 415 6.21.5 Blocking in Guile Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 6.21.6 Critical Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 6.21.7 Fluids and Dynamic States. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 6.21.8 Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 6.21.9 Futures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 424 6.21.10 Parallel forms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425

6.22 Configuration, Features and Runtime Options . . . . . . . . . . . . . . . 427 6.22.1 Configuration, Build and Installation . . . . . . . . . . . . . . . . . . . 427 6.22.2 Feature Tracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429

6.22.2.1 Feature Manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 6.22.2.2 Common Feature Symbols. . . . . . . . . . . . . . . . . . . . . . . . . 429

6.22.3 Runtime Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431

6.22.3.1 Examples of option use. . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 6.23 Support for Other Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 432

6.23.1 Using Other Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 432 6.23.2 Emacs Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433

6.23.2.1 Nil. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 6.23.2.2 Equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 6.23.2.3 Dynamic Binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 6.23.2.4 Other Elisp Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435

6.23.3 ECMAScript. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 6.24 Support for Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436

6.24.1 Internationalization with Guile . . . . . . . . . . . . . . . . . . . . . . . . . 436 6.24.2 Text Collation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 6.24.3 Character Case Mapping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 6.24.4 Number Input and Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 6.24.5 Accessing Locale Information. . . . . . . . . . . . . . . . . . . . . . . . . . . 439 6.24.6 Gettext Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442

6.25 Debugging Infrastructure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444

6.25.1 Evaluation and the Scheme Stack . . . . . . . . . . . . . . . . . . . . . . 445 6.25.1.1 Stack Capture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 6.25.1.2 Stacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 6.25.1.3 Frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446

6.25.2 Source Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 6.25.3 Programmatic Error Handling. . . . . . . . . . . . . . . . . . . . . . . . . . 449

6.25.3.1 Catching Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 6.25.3.2 Capturing the full error stack. . . . . . . . . . . . . . . . . . . . . . 450 6.25.3.3 Pre-Unwind Debugging. . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 6.25.3.4 Debug options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454

6.25.4 Traps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 6.25.4.1 VM Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 6.25.4.2 Trap Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457 6.25.4.3 Low-Level Traps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 6.25.4.4 Tracing Traps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 6.25.4.5 Trap States . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 6.25.4.6 High-Level Traps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462

6.25.5 GDB Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 6.26 Code Coverage Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464

7 Guile Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467

7.1 SLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 7.1.1 SLIB installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 7.1.2 JACAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468

7.2 POSIX System Calls and Networking . . . . . . . . . . . . . . . . . . . . . . . . . 468 7.2.1 POSIX Interface Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 7.2.2 Ports and File Descriptors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 7.2.3 File System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 7.2.4 User Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 7.2.5 Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 7.2.6 Runtime Environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 7.2.7 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 7.2.8 Signals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 7.2.9 Terminals and Ptys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 7.2.10 Pipes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 7.2.11 Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501

7.2.11.1 Network Address Conversion . . . . . . . . . . . . . . . . . . . . . . 501 7.2.11.2 Network Databases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 7.2.11.3 Network Socket Address. . . . . . . . . . . . . . . . . . . . . . . . . . . 509 7.2.11.4 Network Sockets and Communication. . . . . . . . . . . . . . 511 7.2.11.5 Network Socket Examples . . . . . . . . . . . . . . . . . . . . . . . . . 516

7.2.12 System Identification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 7.2.13 Locales . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 7.2.14 Encryption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518

7.3 HTTP, the Web, and All That. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 7.3.1 Types and the Web . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 7.3.2 Universal Resource Identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 7.3.3 The Hyper-Text Transfer Protocol. . . . . . . . . . . . . . . . . . . . . . . 523 7.3.4 HTTP Headers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525

7.3.4.1 HTTP Header Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 7.3.4.2 General Headers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 7.3.4.3 Entity Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 7.3.4.4 Request Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 7.3.4.5 Response Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532

7.3.5 Transfer Codings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 7.3.6 HTTP Requests. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534

7.3.6.1 An Important Note on Character Sets . . . . . . . . . . . . . . 534 7.3.6.2 Request API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534

7.3.7 HTTP Responses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 7.3.8 Web Client. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 7.3.9 Web Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 7.3.10 Web Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542

7.3.10.1 Hello, World!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 543 7.3.10.2 Inspecting the Request . . . . . . . . . . . . . . . . . . . . . . . . . . . . 543 7.3.10.3 Higher-Level Interfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . 544 7.3.10.4 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 545

7.4 The (ice-9 getopt-long) Module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 7.4.1 A Short getopt-long Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 7.4.2 How to Write an Option Specification . . . . . . . . . . . . . . . . . . . 547 7.4.3 Expected Command Line Format. . . . . . . . . . . . . . . . . . . . . . . . 548 7.4.4 Reference Documentation for getopt-long . . . . . . . . . . . . . . 549 7.4.5 Reference Documentation for option-ref . . . . . . . . . . . . . . . 550

7.5 SRFI Support Modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550 7.5.1 About SRFI Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550 7.5.2 SRFI-0 - cond-expand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 7.5.3 SRFI-1 - List library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552

7.5.3.1 Constructors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 7.5.3.2 Predicates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 7.5.3.3 Selectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 7.5.3.4 Length, Append, Concatenate, etc. . . . . . . . . . . . . . . . . . 555 7.5.3.5 Fold, Unfold & Map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 7.5.3.6 Filtering and Partitioning . . . . . . . . . . . . . . . . . . . . . . . . . . 559 7.5.3.7 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 7.5.3.8 Deleting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 7.5.3.9 Association Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 7.5.3.10 Set Operations on Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . 563

7.5.4 SRFI-2 - and-let* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 7.5.5 SRFI-4 - Homogeneous numeric vector datatypes . . . . . . . . 566

7.5.5.1 SRFI-4 - Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 7.5.5.2 SRFI-4 - API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 7.5.5.3 SRFI-4 - Relation to bytevectors. . . . . . . . . . . . . . . . . . . . 574 7.5.5.4 SRFI-4 - Guile extensions . . . . . . . . . . . . . . . . . . . . . . . . . . 574

7.5.6 SRFI-6 - Basic String Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 7.5.7 SRFI-8 - receive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 7.5.8 SRFI-9 - define-record-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 7.5.9 SRFI-10 - Hash-Comma Reader Extension. . . . . . . . . . . . . . . 575 7.5.10 SRFI-11 - let-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 7.5.11 SRFI-13 - String Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 7.5.12 SRFI-14 - Character-set Library. . . . . . . . . . . . . . . . . . . . . . . . 578 7.5.13 SRFI-16 - case-lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 7.5.14 SRFI-17 - Generalized set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 7.5.15 SRFI-18 - Multithreading support . . . . . . . . . . . . . . . . . . . . . . 578

7.5.15.1 SRFI-18 Threads. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 7.5.15.2 SRFI-18 Mutexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 580 7.5.15.3 SRFI-18 Condition variables. . . . . . . . . . . . . . . . . . . . . . . 581 7.5.15.4 SRFI-18 Time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 7.5.15.5 SRFI-18 Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582

7.5.16 SRFI-19 - Time/Date Library . . . . . . . . . . . . . . . . . . . . . . . . . . 583 7.5.16.1 SRFI-19 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 7.5.16.2 SRFI-19 Time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 7.5.16.3 SRFI-19 Date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 7.5.16.4 SRFI-19 Time/Date conversions . . . . . . . . . . . . . . . . . . . 587 7.5.16.5 SRFI-19 Date to string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 7.5.16.6 SRFI-19 String to date . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589

7.5.17 SRFI-23 - Error Reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 7.5.18 SRFI-26 - specializing parameters . . . . . . . . . . . . . . . . . . . . . . 590 7.5.19 SRFI-27 - Sources of Random Bits . . . . . . . . . . . . . . . . . . . . . 591

7.5.19.1 The Default Random Source . . . . . . . . . . . . . . . . . . . . . . 592 7.5.19.2 Random Sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 7.5.19.3 Obtaining random number generator procedures … 593

7.5.20 SRFI-30 - Nested Multi-line Comments. . . . . . . . . . . . . . . . . 593 7.5.21 SRFI-31 - A special form ‘rec’ for recursive evaluation . . 593 7.5.22 SRFI-34 - Exception handling for programs. . . . . . . . . . . . . 594 7.5.23 SRFI-35 - Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 7.5.24 SRFI-37 - args-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 7.5.25 SRFI-38 - External Representation for Data With Shared

Structure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 7.5.26 SRFI-39 - Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599 7.5.27 SRFI-41 - Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599

7.5.27.1 SRFI-41 Stream Fundamentals . . . . . . . . . . . . . . . . . . . . 599 7.5.27.2 SRFI-41 Stream Primitives . . . . . . . . . . . . . . . . . . . . . . . . 599 7.5.27.3 SRFI-41 Stream Library. . . . . . . . . . . . . . . . . . . . . . . . . . . 601

7.5.28 SRFI-42 - Eager Comprehensions. . . . . . . . . . . . . . . . . . . . . . . 608 7.5.29 SRFI-43 - Vector Library. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609

7.5.29.1 SRFI-43 Constructors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 7.5.29.2 SRFI-43 Predicates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 7.5.29.3 SRFI-43 Selectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 7.5.29.4 SRFI-43 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 7.5.29.5 SRFI-43 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 7.5.29.6 SRFI-43 Mutators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 7.5.29.7 SRFI-43 Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613

7.5.30 SRFI-45 - Primitives for Expressing Iterative Lazy

Algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 7.5.31 SRFI-46 Basic syntax-rules Extensions . . . . . . . . . . . . . . . . . 615 7.5.32 SRFI-55 - Requiring Features . . . . . . . . . . . . . . . . . . . . . . . . . . 615 7.5.33 SRFI-60 - Integers as Bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 7.5.34 SRFI-61 - A more general cond clause . . . . . . . . . . . . . . . . . . 617 7.5.35 SRFI-62 - S-expression comments. . . . . . . . . . . . . . . . . . . . . . . 618 7.5.36 SRFI-64 - A Scheme API for test suites. . . . . . . . . . . . . . . . . 618 7.5.37 SRFI-67 - Compare procedures . . . . . . . . . . . . . . . . . . . . . . . . . 618 7.5.38 SRFI-69 - Basic hash tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618

7.5.38.1 Creating hash tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 7.5.38.2 Accessing table items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 7.5.38.3 Table properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 7.5.38.4 Hash table algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620

7.5.39 SRFI-87 => in case clauses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 7.5.40 SRFI-88 Keyword Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 7.5.41 SRFI-98 Accessing environment variables. . . . . . . . . . . . . . . 621 7.5.42 SRFI-105 Curly-infix expressions. . . . . . . . . . . . . . . . . . . . . . . 621 7.5.43 SRFI-111 Boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622

7.6 R6RS Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 7.6.1 Incompatibilities with the R6RS. . . . . . . . . . . . . . . . . . . . . . . . . 622 7.6.2 R6RS Standard Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623

7.6.2.1 Library Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 7.6.2.2 rnrs base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 7.6.2.3 rnrs unicode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 7.6.2.4 rnrs bytevectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 7.6.2.5 rnrs lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 7.6.2.6 rnrs sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 7.6.2.7 rnrs control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 634 7.6.2.8 R6RS Records. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 634 7.6.2.9 rnrs records syntactic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 635 7.6.2.10 rnrs records procedural . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 7.6.2.11 rnrs records inspection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 7.6.2.12 rnrs exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 638 7.6.2.13 rnrs conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 639 7.6.2.14 I/O Conditions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 7.6.2.15 rnrs io ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 7.6.2.16 rnrs io simple . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 7.6.2.17 rnrs files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 7.6.2.18 rnrs programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 7.6.2.19 rnrs arithmetic fixnums . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 7.6.2.20 rnrs arithmetic flonums. . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 7.6.2.21 rnrs arithmetic bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 7.6.2.22 rnrs syntax-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 7.6.2.23 rnrs hashtables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 7.6.2.24 rnrs enums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 654 7.6.2.25 rnrs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 7.6.2.26 rnrs eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 7.6.2.27 rnrs mutable-pairs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 7.6.2.28 rnrs mutable-strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 7.6.2.29 rnrs r5rs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656

7.7 Pattern Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657 7.8 Readline Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660

7.8.1 Loading Readline Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 7.8.2 Readline Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 7.8.3 Readline Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661

7.8.3.1 Readline Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 662 7.8.3.2 Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663

7.9 Pretty Printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 7.10 Formatted Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 7.11 File Tree Walk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 7.12 Queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 7.13 Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 7.14 Buffered Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 7.15 Expect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 7.16 sxml-match: Pattern Matching of SXML. . . . . . . . . . . . . . . . . . . . 687

Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 Matching XML Elements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 Ellipses in Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 Ellipses in Quasiquote’d Output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 Matching Nodesets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 Matching the “Rest” of a Nodeset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 Matching the Unmatched Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 Default Values in Attribute Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 Guards in Patterns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 Catamorphisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 Named-Catamorphisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 692 sxml-match-let and sxml-match-let* . . . . . . . . . . . . . . . . . . . . . . . . 692

7.17 The Scheme shell (scsh). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 7.18 Curried Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 7.19 Statprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 7.20 Implementation notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 7.21 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 7.22 SXML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697

7.22.1 SXML Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 7.22.2 Reading and Writing XML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 7.22.3 SSAX: A Functional XML Parsing Toolkit. . . . . . . . . . . . . . 700

7.22.3.1 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 7.22.3.2 Implementation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 7.22.3.3 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702

7.22.4 Transforming SXML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 7.22.4.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 7.22.4.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704

7.22.5 SXML Tree Fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 7.22.5.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705 7.22.5.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705

7.22.6 SXPath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706 7.22.6.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706 7.22.6.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707

7.22.7 (sxml ssax input-parse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 7.22.7.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 7.22.7.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709

7.22.8 (sxml apply-templates). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 7.22.8.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 7.22.8.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710

7.23 Texinfo Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710

7.23.1 (texinfo) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 7.23.1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 7.23.1.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710

7.23.2 (texinfo docbook) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 7.23.2.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 7.23.2.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711

7.23.3 (texinfo html). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 7.23.3.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 7.23.3.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712

7.23.4 (texinfo indexing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 7.23.4.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 7.23.4.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713

7.23.5 (texinfo string-utils). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 7.23.5.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 7.23.5.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713

7.23.6 (texinfo plain-text). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.6.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.6.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716

7.23.7 (texinfo serialize) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.7.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.7.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716

7.23.8 (texinfo reflection) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.8.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 7.23.8.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716

8 GOOPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719

8.1 Copyright Notice. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 8.2 Class Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 8.3 Instance Creation and Slot Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . 720 8.4 Slot Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 8.5 Illustrating Slot Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724 8.6 Methods and Generic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 726

8.6.1 Accessors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 8.6.2 Extending Primitives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 8.6.3 Merging Generics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 8.6.4 Next-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 8.6.5 Generic Function and Method Examples. . . . . . . . . . . . . . . . . 730 8.6.6 Handling Invocation Errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733

8.7 Inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733 8.7.1 Class Precedence List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 734 8.7.2 Sorting Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 735

8.8 Introspection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 736 8.8.1 Classes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 736 8.8.2 Instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 8.8.3 Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 8.8.4 Generic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 8.8.5 Accessing Slots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740

8.9 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 8.10 GOOPS Object Miscellany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 8.11 The Metaobject Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 743

8.11.1 Metaobjects and the Metaobject Protocol . . . . . . . . . . . . . . 743 8.11.2 Metaclasses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 745 8.11.3 MOP Specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 746 8.11.4 Instance Creation Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 746 8.11.5 Class Definition Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 747 8.11.6 Customizing Class Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . 750 8.11.7 Method Definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 752 8.11.8 Method Definition Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . 752 8.11.9 Generic Function Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753 8.11.10 Generic Function Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . 754

8.12 Redefining a Class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 755 8.12.1 Default Class Redefinition Behaviour . . . . . . . . . . . . . . . . . . . 755 8.12.2 Customizing Class Redefinition. . . . . . . . . . . . . . . . . . . . . . . . . 755

8.13 Changing the Class of an Instance. . . . . . . . . . . . . . . . . . . . . . . . . . . 756

9 Guile Implementation. . . . . . . . . . . . . . . . . . . . . . . . 759

9.1 A Brief History of Guile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 9.1.1 The Emacs Thesis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 9.1.2 Early Days. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 9.1.3 A Scheme of Many Maintainers. . . . . . . . . . . . . . . . . . . . . . . . . . 760 9.1.4 A Timeline of Selected Guile Releases . . . . . . . . . . . . . . . . . . . 761 9.1.5 Status, or: Your Help Needed . . . . . . . . . . . . . . . . . . . . . . . . . . . 762

9.2 Data Representation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 763 9.2.1 A Simple Representation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 764 9.2.2 Faster Integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 764 9.2.3 Cheaper Pairs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 766 9.2.4 Conservative Garbage Collection . . . . . . . . . . . . . . . . . . . . . . . . 767 9.2.5 The SCM Type in Guile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768

9.2.5.1 Relationship between SCM and scm_t_bits. . . . . . . . . . 768 9.2.5.2 Immediate objects. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 9.2.5.3 Non-immediate objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 9.2.5.4 Allocating Cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 9.2.5.5 Heap Cell Type Information . . . . . . . . . . . . . . . . . . . . . . . . 771 9.2.5.6 Accessing Cell Entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771

9.3 A Virtual Machine for Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 772 9.3.1 Why a VM? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 772 9.3.2 VM Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 9.3.3 Stack Layout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 9.3.4 Variables and the VM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 775 9.3.5 Compiled Procedures are VM Programs . . . . . . . . . . . . . . . . . 776 9.3.6 Instruction Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 777

9.3.6.1 Lexical Environment Instructions . . . . . . . . . . . . . . . . . . . 777 9.3.6.2 Top-Level Environment Instructions . . . . . . . . . . . . . . . . 779 9.3.6.3 Procedure Call and Return Instructions. . . . . . . . . . . . . 780 9.3.6.4 Function Prologue Instructions . . . . . . . . . . . . . . . . . . . . . 782 9.3.6.5 Trampoline Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 9.3.6.6 Branch Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 9.3.6.7 Data Constructor Instructions . . . . . . . . . . . . . . . . . . . . . . 785 9.3.6.8 Loading Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786 9.3.6.9 Dynamic Environment Instructions . . . . . . . . . . . . . . . . . 787 9.3.6.10 Miscellaneous Instructions. . . . . . . . . . . . . . . . . . . . . . . . . 788 9.3.6.11 Inlined Scheme Instructions . . . . . . . . . . . . . . . . . . . . . . . 789 9.3.6.12 Inlined Mathematical Instructions . . . . . . . . . . . . . . . . . 790 9.3.6.13 Inlined Bytevector Instructions . . . . . . . . . . . . . . . . . . . . 790

9.4 Compiling to the Virtual Machine . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 9.4.1 Compiler Tower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 792 9.4.2 The Scheme Compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 793 9.4.3 Tree-IL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 795 9.4.4 GLIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 xviii Guile Reference Manual

9.4.5 Assembly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 802 9.4.6 Bytecode and Objcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 803 9.4.7 Writing New High-Level Languages. . . . . . . . . . . . . . . . . . . . . . 804 9.4.8 Extending the Compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 804

Appendix A GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 807 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 815 Procedure Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 821 Variable Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 861 Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 865 R5RS Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 867

Preface

This manual describes how to use Guile, GNU’s Ubiquitous Intelligent Language for Extensions. It relates particularly to Guile version 2.0.11.

Contributors to this Manual

Like Guile itself, the Guile reference manual is a living entity, cared for by many people over a long period of time. As such, it is hard to identify individuals of whom to say “yes, this person, she wrote the manual.”

Still, among the many contributions, some caretakers stand out. First among them is Neil Jerram, who has been working on this document for ten years now. Neil’s attention both to detail and to the big picture have made a real difference in the understanding of a generation of Guile hackers.

Next we should note Marius Vollmer’s effect on this document. Marius maintained Guile during a period in which Guile’s API was clarified—put to the fire, so to speak—and he had the good sense to effect the same change on the manual.

Martin Grabmueller made substantial contributions throughout the manual in preparation for the Guile 1.6 release, including filling out a lot of the documentation of Scheme data types, control mechanisms and procedures. In addition, he wrote the documentation for Guile’s SRFI modules and modules associated with the Guile REPL.

Ludovic Courtès and Andy Wingo, the Guile maintainers at the time of this writing (late 2010), have also made their dent in the manual, writing documentation for new modules and subsystems in Guile 2.0. They are also responsible for ensuring that the existing text retains its relevance as Guile evolves. See Section 2.6 [Reporting Bugs], page 12, for more information on reporting problems in this manual.

The content for the first versions of this manual incorporated and was inspired by documents from Aubrey Jaffer, author of the SCM system on which Guile was based, and from Tom Lord, Guile’s first maintainer. Although most of this text has been rewritten, all of it was important, and some of the structure remains.

The manual for the first versions of Guile were largely written, edited, and compiled by Mark Galassi and Jim Blandy. In particular, Jim wrote the original tutorial on Guile’s data representation and the C API for accessing Guile objects.

Significant portions were also contributed by Thien-Thi Nguyen, Kevin Ryde, Mikael Djurfeldt, Christian Lynbech, Julian Graham, Gary Houston, Tim Pierce, and a few dozen more. You, reader, are most welcome to join their esteemed ranks. Visit Guile’s web site

at http://www.gnu.org/software/guile/ to find out how to get involved.

The Guile License

Guile is Free Software. Guile is copyrighted, not public domain, and there are restrictions on its distribution or redistribution, but these restrictions are designed to permit everything a cooperating person would want to do.

• The Guile library (libguile) and supporting files are published under the terms of the GNU Lesser General Public License version 3 or later. See the files COPYING.LESSER

    and COPYING.

• The Guile readline module is published under the terms of the GNU General Public License version 3 or later. See the file COPYING.

• The manual you’re now reading is published under the terms of the GNU Free Documentation License (see Appendix A [GNU Free Documentation License], page 807).

C code linking to the Guile library is subject to terms of that library. Basically such code may be published on any terms, provided users can re-link against a new or modified version of Guile.

C code linking to the Guile readline module is subject to the terms of that module. Basically such code must be published on Free terms.

Scheme level code written to be run by Guile (but not derived from Guile itself) is not restricted in any way, and may be published on any terms. We encourage authors to publish on Free terms.

You must be aware there is no warranty whatsoever for Guile. This is described in full in the licenses.

1 Introduction

Guile is an implementation of the Scheme programming language. Scheme (http:// schemers.org/) is an elegant and conceptually simple dialect of Lisp, originated by Guy Steele and Gerald Sussman, and since evolved by the series of reports known as RnRS (the Revised^{n} Reports on Scheme).

Unlike, for example, Python or Perl, Scheme has no benevolent dictator. There are many Scheme implementations, with different characteristics and with communities and academic activities around them, and the language develops as a result of the interplay between these. Guile’s particular characteristics are that

• it is easy to combine with other code written in C

• it has a historical and continuing connection with the GNU Project

• it emphasizes interactive and incremental programming

• it actually supports several languages, not just Scheme.

The next few sections explain what we mean by these points. The sections after that cover how you can obtain and install Guile, and the typographical conventions that we use in this manual.

1.1 Guile and Scheme

Guile implements Scheme as described in the Revised^{5} Report on the Algorithmic Language Scheme (usually known as R5RS), providing clean and general data and control structures. Guile goes beyond the rather austere language presented in R5RS, extending it with a module system, full access to POSIX system calls, networking support, multiple threads, dynamic linking, a foreign function call interface, powerful string processing, and many other features needed for programming in the real world.

The Scheme community has recently agreed and published R6RS, the latest installment in the RnRS series. R6RS significantly expands the core Scheme language, and standardises many non-core functions that implementations—including Guile—have previously done in different ways. Guile has been updated to incorporate some of the features of R6RS, and to adjust some existing features to conform to the R6RS specification, but it is by no means a complete R6RS implementation. See Section 7.6 [R6RS Support], page 622.

Between R5RS and R6RS, the SRFI process (http://srfi.schemers.org/) standardised interfaces for many practical needs, such as multithreaded programming and multidimensional arrays. Guile supports many SRFIs, as documented in detail in Section 7.5 [SRFI Support], page 550.

In summary, so far as relationship to the Scheme standards is concerned, Guile is an R5RS implementation with many extensions, some of which conform to SRFIs or to the relevant parts of R6RS.

1.2 Combining with C Code

Like a shell, Guile can run interactively—reading expressions from the user, evaluating them, and displaying the results—or as a script interpreter, reading and executing Scheme code from a file. Guile also provides an object library, libguile, that allows other applications to easily incorporate a complete Scheme interpreter. An application can then use Guile as an extension language, a clean and powerful configuration language, or as multi-purpose “glue”, connecting primitives provided by the application. It is easy to call Scheme code from C code and vice versa, giving the application designer full control of how and when to invoke the interpreter. Applications can add new functions, data types, control structures, and even syntax to Guile, creating a domain-specific language tailored to the task at hand, but based on a robust language design.

This kind of combination is helped by four aspects of Guile’s design and history. First is that Guile has always been targeted as an extension language. Hence its C API has always been of great importance, and has been developed accordingly. Second and third are rather technical points—that Guile uses conservative garbage collection, and that it implements the Scheme concept of continuations by copying and reinstating the C stack—but whose practical consequence is that most existing C code can be glued into Guile as is, without needing modifications to cope with strange Scheme execution flows. Last is the module system, which helps extensions to coexist without stepping on each others’ toes.

Guile’s module system allows one to break up a large program into manageable sections with well-defined interfaces between them. Modules may contain a mixture of interpreted and compiled code; Guile can use either static or dynamic linking to incorporate compiled code. Modules also encourage developers to package up useful collections of routines for general distribution; as of this writing, one can find Emacs interfaces, database access routines, compilers, GUI toolkit interfaces, and HTTP client functions, among others.

1.3 Guile and the GNU Project

Guile was conceived by the GNU Project following the fantastic success of Emacs Lisp as an extension language within Emacs. Just as Emacs Lisp allowed complete and unanticipated applications to be written within the Emacs environment, the idea was that Guile should do the same for other GNU Project applications. This remains true today.

The idea of extensibility is closely related to the GNU project’s primary goal, that of promoting software freedom. Software freedom means that people receiving a software package can modify or enhance it to their own desires, including in ways that may not have occurred at all to the software’s original developers. For programs written in a compiled language like C, this freedom covers modifying and rebuilding the C code; but if the program also provides an extension language, that is usually a much friendlier and lower-barrier-of-entry way for the user to start making their own changes.

Guile is now used by GNU project applications such as AutoGen, Lilypond, Denemo, Mailutils, TeXmacs and Gnucash, and we hope that there will be many more in future.

1.4 Interactive Programming

Non-free software has no interest in its users being able to see how it works. They are supposed to just accept it, or to report problems and hope that the source code owners will choose to work on them.

Free software aims to work reliably just as much as non-free software does, but it should also empower its users by making its workings available. This is useful for many reasons, including education, auditing and enhancements, as well as for debugging problems.

The ideal free software system achieves this by making it easy for interested users to see the source code for a feature that they are using, and to follow through that source code step-by-step, as it runs. In Emacs, good examples of this are the source code hyperlinks in the help system, and edebug. Then, for bonus points and maximising the ability for the user to experiment quickly with code changes, the system should allow parts of the source code to be modified and reloaded into the running program, to take immediate effect.

Guile is designed for this kind of interactive programming, and this distinguishes it from many Scheme implementations that instead prioritise running a fixed Scheme program as fast as possible—because there are tradeoffs between performance and the ability to modify parts of an already running program. There are faster Schemes than Guile, but Guile is a GNU project and so prioritises the GNU vision of programming freedom and experimentation.

1.5 Supporting Multiple Languages

Since the 2.0 release, Guile’s architecture supports compiling any language to its core virtual machine bytecode, and Scheme is just one of the supported languages. Other supported languages are Emacs Lisp, ECMAScript (commonly known as Javascript) and Brainfuck, and work is under discussion for Lua, Ruby and Python.

This means that users can program applications which use Guile in the language of their choice, rather than having the tastes of the application’s author imposed on them.

1.6 Obtaining and Installing Guile

Guile can be obtained from the main GNU archive site ftp://ftp.gnu.org or any of its mirrors. The file will be named guile-version.tar.gz. The current version is 2.0.11, so the file you should grab is:

   ftp://ftp.gnu.org/gnu/guile/guile-2.0.11.tar.gz

To unbundle Guile use the instruction

     zcat guile-2.0.11.tar.gz | tar xvf -

which will create a directory called guile-2.0.11 with all the sources. You can look at the file INSTALL for detailed instructions on how to build and install Guile, but you should be able to just do

     cd guile-2.0.11
     ./configure
     make
     make install

This will install the Guile executable guile, the Guile library libguile and various associated header files and support libraries. It will also install the Guile reference manual.

Since this manual frequently refers to the Scheme “standard”, also known as R5RS, or the “Revised^{5} Report on the Algorithmic Language Scheme”, we have included the report in the Guile distribution; see Section “Introduction” in Revised(5) Report on the Algorithmic Language Scheme. This will also be installed in your info directory.

1.7 Organisation of this Manual

The rest of this manual is organised into the following chapters.

Chapter 2: Hello Guile!

A whirlwind tour shows how Guile can be used interactively and as a script interpreter, how to link Guile into your own applications, and how to write modules of interpreted and compiled code for use with Guile. Everything introduced here is documented again and in full by the later parts of the manual.

Chapter 3: Hello Scheme!

For readers new to Scheme, this chapter provides an introduction to the basic ideas of the Scheme language. This material would apply to any Scheme implementation and so does not make reference to anything Guile-specific.

Chapter 4: Programming in Scheme Provides an overview of programming in Scheme with Guile. It covers how to invoke the guile program from the command-line and how to write scripts in Scheme. It also introduces the extensions that Guile offers beyond standard Scheme.

Chapter 5: Programming in C Provides an overview of how to use Guile in a C program. It discusses the fundamental concepts that you need to understand to access the features of Guile, such as dynamic types and the garbage collector. It explains in a tutorial like manner how to define new data types and functions for the use by Scheme programs.

Chapter 6: Guile API Reference This part of the manual documents the Guile API in functionality-based groups with the Scheme and C interfaces presented side by side.

Chapter 7: Guile Modules Describes some important modules, distributed as part of the Guile distribution, that extend the functionality provided by the Guile Scheme core.

Chapter 8: GOOPS Describes GOOPS, an object oriented extension to Guile that provides classes, multiple inheritance and generic functions.

1.8 Typographical Conventions

In examples and procedure descriptions and all other places where the evaluation of Scheme expression is shown, we use some notation for denoting the output and evaluation results of expressions.

The symbol ‘⇒’ is used to tell which value is returned by an evaluation:

(+ 1 2)
⇒ 3

Some procedures produce some output besides returning a value. This is denoted by the symbol ‘ ⊣ ’.

     (begin (display 1) (newline) 'hooray)
     ⊣ 1
     ⇒ hooray

As you can see, this code prints ‘1’ (denoted by ‘ ⊣ ’), and returns hooray (denoted by ‘⇒’).

2 Hello Guile!

This chapter presents a quick tour of all the ways that Guile can be used. There are additional examples in the examples/ directory in the Guile source distribution. It also explains how best to report any problems that you find.

The following examples assume that Guile has been installed in /usr/local/.

2.1 Running Guile Interactively

In its simplest form, Guile acts as an interactive interpreter for the Scheme programming language, reading and evaluating Scheme expressions the user enters from the terminal. Here is a sample interaction between Guile and a user; the user’s input appears after the $

and scheme@(guile-user)> prompts:
$ guile
scheme@(guile-user)> (+ 1 2 3)                ; add some numbers
$1 = 6
scheme@(guile-user)> (define (factorial n)    ; define a function
                       (if (zero? n) 1 (* n (factorial (- n 1)))))
scheme@(guile-user)> (factorial 20)
$2 = 2432902008176640000
scheme@(guile-user)> (getpwnam "root")        ; look in /etc/passwd
$3 = #("root" "x" 0 0 "root" "/root" "/bin/bash")
scheme@(guile-user)> C-d
$

2.2 Running Guile Scripts

Like AWK, Perl, or any shell, Guile can interpret script files. A Guile script is simply a file of Scheme code with some extra information at the beginning which tells the operating system how to invoke Guile, and then tells Guile how to handle the Scheme code.

Here is a trivial Guile script. See Section 4.3 [Guile Scripting], page 41, for more details.

#!/usr/local/bin/guile -s
!#
(display "Hello, world!")
(newline)

2.3 Linking Guile into Programs

The Guile interpreter is available as an object library, to be linked into applications using Scheme as a configuration or extension language.

Here is simple-guile.c, source code for a program that will produce a complete Guile interpreter. In addition to all usual functions provided by Guile, it will also offer the function

my-hostname.
#include <stdlib.h>
#include <libguile.h>
static SCM

my_hostname (void)
{
  char *s = getenv ("HOSTNAME");
  if (s == NULL)
    return SCM_BOOL_F;
  else
    return scm_from_locale_string (s);
}
static void
inner_main (void *data, int argc, char **argv)
{
  scm_c_define_gsubr ("my-hostname", 0, 0, 0, my_hostname);
  scm_shell (argc, argv);
}
int
main (int argc, char **argv)
{
  scm_boot_guile (argc, argv, inner_main, 0);
  return 0; /* never reached */
}

When Guile is correctly installed on your system, the above program can be compiled and linked like this:

$ gcc -o simple-guile simple-guile.c \
    `pkg-config --cflags --libs guile-2.0`

When it is run, it behaves just like the guile program except that you can also call the new my-hostname function.

$ ./simple-guile
scheme@(guile-user)> (+ 1 2 3)
$1 = 6
scheme@(guile-user)> (my-hostname)
"burns"

2.4 Writing Guile Extensions

You can link Guile into your program and make Scheme available to the users of your program. You can also link your library into Guile and make its functionality available to all users of Guile.

A library that is linked into Guile is called an extension, but it really just is an ordinary object library.

The following example shows how to write a simple extension for Guile that makes the j0 function available to Scheme code.

#include <math.h>
#include <libguile.h>

      SCM
      j0_wrapper (SCM x)
      {
        return scm_from_double (j0 (scm_to_double (x)));
      }
      void
      init_bessel ()
      {
        scm_c_define_gsubr ("j0", 1, 0, 0, j0_wrapper);
      }

This C source file needs to be compiled into a shared library. Here is how to do it on GNU/Linux:

      gcc `pkg-config --cflags guile-2.0` \
        -shared -o libguile-bessel.so -fPIC bessel.c

For creating shared libraries portably, we recommend the use of GNU Libtool (see Section “Introduction” in GNU Libtool).

A shared library can be loaded into a running Guile process with the function loadextension. The j0 is then immediately available:

      $ guile
      scheme@(guile-user)> (load-extension "./libguile-bessel" "init_bessel")
      scheme@(guile-user)> (j0 2)
      $1 = 0.223890779141236

For more on how to install your extension, see Section 4.7 [Installing Site Packages], page 55.

2.5 Using the Guile Module System

Guile has support for dividing a program into modules. By using modules, you can group related code together and manage the composition of complete programs from largely independent parts.

For more details on the module system beyond this introductory material, See Section 6.19 [Modules], page 381.

2.5.1 Using Modules

Guile comes with a lot of useful modules, for example for string processing or command line parsing. Additionally, there exist many Guile modules written by other Guile hackers, but which have to be installed manually.

Here is a sample interactive session that shows how to use the (ice-9 popen) module which provides the means for communicating with other processes over pipes together with the (ice-9 rdelim) module that provides the function read-line.

      $ guile
      scheme@(guile-user)> (use-modules (ice-9 popen))
      scheme@(guile-user)> (use-modules (ice-9 rdelim))
      scheme@(guile-user)> (define p (open-input-pipe "ls -l"))
      scheme@(guile-user)> (read-line p)
      $1 = "total 30"
      scheme@(guile-user)> (read-line p)
      $2 = "drwxr-sr-x    2 mgrabmue mgrabmue     1024 Mar 29 19:57 CVS"

2.5.2 Writing new Modules

You can create new modules using the syntactic form define-module. All definitions following this form until the next define-module are placed into the new module.

One module is usually placed into one file, and that file is installed in a location where Guile can automatically find it. The following session shows a simple example.

      $ cat /usr/local/share/guile/site/foo/bar.scm
      (define-module (foo bar)
        #:export (frob))
      (define (frob x) (* 2 x))
      $ guile
      scheme@(guile-user)> (use-modules (foo bar))
      scheme@(guile-user)> (frob 12)
      $1 = 24

For more on how to install your module, see Section 4.7 [Installing Site Packages], page 55.

2.5.3 Putting Extensions into Modules

In addition to Scheme code you can also put things that are defined in C into a module.

You do this by writing a small Scheme file that defines the module and call loadextension directly in the body of the module.

      $ cat /usr/local/share/guile/site/math/bessel.scm
      (define-module (math bessel)
        #:export (j0))
      (load-extension "libguile-bessel" "init_bessel")
      $ file /usr/local/lib/guile/2.0/extensions/libguile-bessel.so
      ... ELF 32-bit LSB shared object ...
      $ guile
      scheme@(guile-user)> (use-modules (math bessel))
      scheme@(guile-user)> (j0 2)
      $1 = 0.223890779141236

See Section 6.20.4 [Modules and Extensions], page 401, for more information.

2.6 Reporting Bugs

Any problems with the installation should be reported to bug-guile@gnu.org.

If you find a bug in Guile, please report it to the Guile developers, so they can fix it. They may also be able to suggest workarounds when it is not possible for you to apply the bug-fix or install a new version of Guile yourself.

Before sending in bug reports, please check with the following list that you really have found a bug.

• Whenever documentation and actual behavior differ, you have certainly found a bug, either in the documentation or in the program.

• When Guile crashes, it is a bug.

• When Guile hangs or takes forever to complete a task, it is a bug.

• When calculations produce wrong results, it is a bug.

• When Guile signals an error for valid Scheme programs, it is a bug.

• When Guile does not signal an error for invalid Scheme programs, it may be a bug, unless this is explicitly documented.

• When some part of the documentation is not clear and does not make sense to you even after re-reading the section, it is a bug.

Before reporting the bug, check whether any programs you have loaded into Guile, including your .guile file, set any variables that may affect the functioning of Guile. Also, see whether the problem happens in a freshly started Guile without loading your .guile file (start Guile with the -q switch to prevent loading the init file). If the problem does not occur then, you must report the precise contents of any programs that you must load into Guile in order to cause the problem to occur.

When you write a bug report, please make sure to include as much of the information described below in the report. If you can’t figure out some of the items, it is not a problem, but the more information we get, the more likely we can diagnose and fix the bug.

• The version number of Guile. You can get this information from invoking ‘guile --version’ at your shell, or calling (version) from within Guile.

• Your machine type, as determined by the config.guess shell script. If you have a Guile checkout, this file is located in build-aux; otherwise you can fetch

the latest version from http://git.savannah.gnu.org/gitweb/ ?p=config.
git;a=blob_plain;f=config.guess;hb=HEAD.
     $ build-aux/config.guess
     x86_64-unknown-linux-gnu

• If you installed Guile from a binary package, the version of that package. On systems that use RPM, use rpm -qa | grep guile. On systems that use DPKG, dpkg -l |

grep guile.

• If you built Guile yourself, the build configuration that you used:

     $ ./config.status --config
     '--enable-error-on-warning' '--disable-deprecated'...

• A complete description of how to reproduce the bug.

If you have a Scheme program that produces the bug, please include it in the bug report. If your program is too big to include. please try to reduce your code to a minimal test case. If you can reproduce your problem at the REPL, that is best. Give a transcript of the expressions you typed at the REPL.

• A description of the incorrect behavior. For example, "The Guile process gets a fatal signal," or, "The resulting output is as follows, which I think is wrong." If the manifestation of the bug is a Guile error message, it is important to report the precise text of the error message, and a backtrace showing how the Scheme program arrived at the error. This can be done using the ,backtrace command in Guile’s debugger.

If your bug causes Guile to crash, additional information from a low-level debugger such as GDB might be helpful. If you have built Guile yourself, you can run Guile under GDB via the meta/gdb-uninstalled-guile script. Instead of invoking Guile as usual, invoke the wrapper script, type run to start the process, then backtrace when the crash comes. Include that backtrace in your report.

3 Hello Scheme!

In this chapter, we introduce the basic concepts that underpin the elegance and power of the Scheme language.

Readers who already possess a background knowledge of Scheme may happily skip this chapter. For the reader who is new to the language, however, the following discussions on data, procedures, expressions and closure are designed to provide a minimum level of Scheme understanding that is more or less assumed by the chapters that follow.

The style of this introductory material aims about halfway between the terse precision of R5RS and the discursiveness of existing Scheme tutorials. For pointers to useful Scheme resources on the web, please see Section 3.5 [Further Reading], page 34.

3.1 Data Types, Values and Variables

This section discusses the representation of data types and values, what it means for Scheme to be a latently typed language, and the role of variables. We conclude by introducing the Scheme syntaxes for defining a new variable, and for changing the value of an existing variable.

3.1.1 Latent Typing

The term latent typing is used to describe a computer language, such as Scheme, for which you cannot, in general, simply look at a program’s source code and determine what type of data will be associated with a particular variable, or with the result of a particular expression.

Sometimes, of course, you can tell from the code what the type of an expression will be. If you have a line in your program that sets the variable x to the numeric value 1, you can be certain that, immediately after that line has executed (and in the absence of multiple threads), x has the numeric value 1. Or if you write a procedure that is designed to concatenate two strings, it is likely that the rest of your application will always invoke this procedure with two string parameters, and quite probable that the procedure would go wrong in some way if it was ever invoked with parameters that were not both strings.

Nevertheless, the point is that there is nothing in Scheme which requires the procedure parameters always to be strings, or x always to hold a numeric value, and there is no way of declaring in your program that such constraints should always be obeyed. In the same vein, there is no way to declare the expected type of a procedure’s return value.

Instead, the types of variables and expressions are only known – in general – at run time. If you need to check at some point that a value has the expected type, Scheme provides run time procedures that you can invoke to do so. But equally, it can be perfectly valid for two separate invocations of the same procedure to specify arguments with different types, and to return values with different types.

The next subsection explains what this means in practice, for the ways that Scheme programs use data types, values and variables.

3.1.2 Values and Variables

Scheme provides many data types that you can use to represent your data. Primitive types include characters, strings, numbers and procedures. Compound types, which allow a group of primitive and compound values to be stored together, include lists, pairs, vectors and multi-dimensional arrays. In addition, Guile allows applications to define their own data types, with the same status as the built-in standard Scheme types.

As a Scheme program runs, values of all types pop in and out of existence. Sometimes values are stored in variables, but more commonly they pass seamlessly from being the result of one computation to being one of the parameters for the next.

Consider an example. A string value is created because the interpreter reads in a literal string from your program’s source code. Then a numeric value is created as the result of calculating the length of the string. A second numeric value is created by doubling the calculated length. Finally the program creates a list with two elements – the doubled length and the original string itself – and stores this list in a program variable.

All of the values involved here – in fact, all values in Scheme – carry their type with them. In other words, every value “knows,” at runtime, what kind of value it is. A number, a string, a list, whatever.

A variable, on the other hand, has no fixed type. A variable – x, say – is simply the name of a location – a box – in which you can store any kind of Scheme value. So the same variable in a program may hold a number at one moment, a list of procedures the next, and later a pair of strings. The “type” of a variable – insofar as the idea is meaningful at all – is simply the type of whatever value the variable happens to be storing at a particular moment.

3.1.3 Defining and Setting Variables

To define a new variable, you use Scheme’s define syntax like this:

     (define variable-name value)

This makes a new variable called variable-name and stores value in it as the variable’s initial value. For example:

     ;; Make a variable `x' with initial numeric value 1.
     (define x 1)
     ;; Make a variable `organization' with an initial string value.
     (define organization "Free Software Foundation")

(In Scheme, a semicolon marks the beginning of a comment that continues until the end of the line. So the lines beginning ;; are comments.)

Changing the value of an already existing variable is very similar, except that define is replaced by the Scheme syntax set!, like this:

     (set! variable-name new-value)

Remember that variables do not have fixed types, so new-value may have a completely different type from whatever was previously stored in the location named by variable-name. Both of the following examples are therefore correct.

     ;; Change the value of `x' to 5.
     (set! x 5)
     ;; Change the value of `organization' to the FSF's street number.
     (set! organization 545)

In these examples, value and new-value are literal numeric or string values. In general, however, value and new-value can be any Scheme expression. Even though we have not yet covered the forms that Scheme expressions can take (see Section 3.3 [About Expressions], page 20), you can probably guess what the following set! example does…

     (set! x (+ x 1))

(Note: this is not a complete description of define and set!, because we need to introduce some other aspects of Scheme before the missing pieces can be filled in. If, however, you are already familiar with the structure of Scheme, you may like to read about those missing pieces immediately by jumping ahead to the following references.

• Section 3.2.4 [Lambda Alternatives], page 20, to read about an alternative form of the define syntax that can be used when defining new procedures.

• Section 6.9.8 [Procedures with Setters], page 255, to read about an alternative form of the set! syntax that helps with changing a single value in the depths of a compound data structure.)

• See Section 6.12.3 [Internal Definitions], page 289, to read about using define other than at top level in a Scheme program, including a discussion of when it works to use define rather than set! to change the value of an existing variable.

3.2 The Representation and Use of Procedures

This section introduces the basics of using and creating Scheme procedures. It discusses the representation of procedures as just another kind of Scheme value, and shows how procedure invocation expressions are constructed. We then explain how lambda is used to create new procedures, and conclude by presenting the various shorthand forms of define that can be used instead of writing an explicit lambda expression.

3.2.1 Procedures as Values

One of the great simplifications of Scheme is that a procedure is just another type of value, and that procedure values can be passed around and stored in variables in exactly the same way as, for example, strings and lists. When we talk about a built-in standard Scheme procedure such as open-input-file, what we actually mean is that there is a pre-defined top level variable called open-input-file, whose value is a procedure that implements what R5RS says that open-input-file should do.

Note that this is quite different from many dialects of Lisp — including Emacs Lisp — in which a program can use the same name with two quite separate meanings: one meaning identifies a Lisp function, while the other meaning identifies a Lisp variable, whose value need have nothing to do with the function that is associated with the first meaning. In these dialects, functions and variables are said to live in different namespaces.

In Scheme, on the other hand, all names belong to a single unified namespace, and the variables that these names identify can hold any kind of Scheme value, including procedure values.

One consequence of the “procedures as values” idea is that, if you don’t happen to like the standard name for a Scheme procedure, you can change it.

For example, call-with-current-continuation is a very important standard Scheme procedure, but it also has a very long name! So, many programmers use the following definition to assign the same procedure value to the more convenient name call/cc.

     (define call/cc call-with-current-continuation)

Let’s understand exactly how this works. The definition creates a new variable call/cc, and then sets its value to the value of the variable call-with-current-continuation; the latter value is a procedure that implements the behaviour that R5RS specifies under the name “call-with-current-continuation”. So call/cc ends up holding this value as well.

Now that call/cc holds the required procedure value, you could choose to use callwith-current-continuation for a completely different purpose, or just change its value so that you will get an error if you accidentally use call-with-current-continuation as a procedure in your program rather than call/cc. For example:

     (set! call-with-current-continuation "Not a procedure any more!")

Or you could just leave call-with-current-continuation as it was. It’s perfectly fine for more than one variable to hold the same procedure value.

3.2.2 Simple Procedure Invocation

A procedure invocation in Scheme is written like this:

     (procedure [arg1 [arg2 ...]])

In this expression, procedure can be any Scheme expression whose value is a procedure. Most commonly, however, procedure is simply the name of a variable whose value is a procedure.

For example, string-append is a standard Scheme procedure whose behaviour is to concatenate together all the arguments, which are expected to be strings, that it is given. So the expression

     (string-append "/home" "/" "andrew")

is a procedure invocation whose result is the string value "/home/andrew".

Similarly, string-length is a standard Scheme procedure that returns the length of a single string argument, so

     (string-length "abc")

is a procedure invocation whose result is the numeric value 3.

Each of the parameters in a procedure invocation can itself be any Scheme expression. Since a procedure invocation is itself a type of expression, we can put these two examples together to get

     (string-length (string-append "/home" "/" "andrew"))

— a procedure invocation whose result is the numeric value 12.

(You may be wondering what happens if the two examples are combined the other way round. If we do this, we can make a procedure invocation expression that is syntactically correct:

     (string-append "/home" (string-length "abc"))

but when this expression is executed, it will cause an error, because the result of (stringlength "abc") is a numeric value, and string-append is not designed to accept a numeric value as one of its arguments.)

3.2.3 Creating and Using a New Procedure

Scheme has lots of standard procedures, and Guile provides all of these via predefined top level variables. All of these standard procedures are documented in the later chapters of this reference manual.

Before very long, though, you will want to create new procedures that encapsulate aspects of your own applications’ functionality. To do this, you can use the famous lambda syntax.

For example, the value of the following Scheme expression

     (lambda (name address) expression ...)

is a newly created procedure that takes two arguments: name and address. The behaviour of the new procedure is determined by the sequence of expressions in the body of the procedure definition. (Typically, these expressions would use the arguments in some way, or else there wouldn’t be any point in giving them to the procedure.) When invoked, the new procedure returns a value that is the value of the last expression in the procedure body.

To make things more concrete, let’s suppose that the two arguments are both strings, and that the purpose of this procedure is to form a combined string that includes these arguments. Then the full lambda expression might look like this:

     (lambda (name address)
       (string-append "Name=" name ":Address=" address))

We noted in the previous subsection that the procedure part of a procedure invocation expression can be any Scheme expression whose value is a procedure. But that’s exactly what a lambda expression is! So we can use a lambda expression directly in a procedure invocation, like this:

     ((lambda (name address)
        (string-append "Name=" name ":Address=" address))
      "FSF"
      "Cambridge")

This is a valid procedure invocation expression, and its result is the string:

     "Name=FSF:Address=Cambridge"

It is more common, though, to store the procedure value in a variable —

     (define make-combined-string
       (lambda (name address)
         (string-append "Name=" name ":Address=" address)))

— and then to use the variable name in the procedure invocation:

     (make-combined-string "FSF" "Cambridge")

Which has exactly the same result.

It’s important to note that procedures created using lambda have exactly the same status as the standard built in Scheme procedures, and can be invoked, passed around, and stored in variables in exactly the same ways.

3.2.4 Lambda Alternatives

Since it is so common in Scheme programs to want to create a procedure and then store it in a variable, there is an alternative form of the define syntax that allows you to do just that.

A define expression of the form

     (define (name [arg1 [arg2 ...]])
       expression ...)

is exactly equivalent to the longer form

     (define name
       (lambda ([arg1 [arg2 ...]])
         expression ...))

So, for example, the definition of make-combined-string in the previous subsection could equally be written:

     (define (make-combined-string name address)
       (string-append "Name=" name ":Address=" address))

This kind of procedure definition creates a procedure that requires exactly the expected number of arguments. There are two further forms of the lambda expression, which create a procedure that can accept a variable number of arguments:

     (lambda (arg1 ... . args) expression ...)
     (lambda args expression ...)

The corresponding forms of the alternative define syntax are:

     (define (name arg1 ... . args) expression ...)
     (define (name . args) expression ...)

For details on how these forms work, see See Section 6.9.1 [Lambda], page 244.

Prior to Guile 2.0, Guile provided an extension to define syntax that allowed you to nest the previous extension up to an arbitrary depth. These are no longer provided by default, and instead have been moved to Section 7.18 [Curried Definitions], page 693

(It could be argued that the alternative define forms are rather confusing, especially for newcomers to the Scheme language, as they hide both the role of lambda and the fact that procedures are values that are stored in variables in the some way as any other kind of value. On the other hand, they are very convenient, and they are also a good example of another of Scheme’s powerful features: the ability to specify arbitrary syntactic transformations at run time, which can be applied to subsequently read input.)

3.3 Expressions and Evaluation

So far, we have met expressions that do things, such as the define expressions that create and initialize new variables, and we have also talked about expressions that have values, for example the value of the procedure invocation expression:

     (string-append "/home" "/" "andrew")

but we haven’t yet been precise about what causes an expression like this procedure invocation to be reduced to its “value”, or how the processing of such expressions relates to the execution of a Scheme program as a whole.

This section clarifies what we mean by an expression’s value, by introducing the idea of evaluation. It discusses the side effects that evaluation can have, explains how each of the various types of Scheme expression is evaluated, and describes the behaviour and use of the Guile REPL as a mechanism for exploring evaluation. The section concludes with a very brief summary of Scheme’s common syntactic expressions.

3.3.1 Evaluating Expressions and Executing Programs

In Scheme, the process of executing an expression is known as evaluation. Evaluation has two kinds of result:

• the value of the evaluated expression

• the side effects of the evaluation, which consist of any effects of evaluating the expression that are not represented by the value.

Of the expressions that we have met so far, define and set! expressions have side effects — the creation or modification of a variable — but no value; lambda expressions have values — the newly constructed procedures — but no side effects; and procedure invocation expressions, in general, have either values, or side effects, or both.

It is tempting to try to define more intuitively what we mean by “value” and “side effects”, and what the difference between them is. In general, though, this is extremely difficult. It is also unnecessary; instead, we can quite happily define the behaviour of a Scheme program by specifying how Scheme executes a program as a whole, and then by describing the value and side effects of evaluation for each type of expression individually. So, some^{1} definitions…

• A Scheme program consists of a sequence of expressions.

• A Scheme interpreter executes the program by evaluating these expressions in order, one by one.

• An expression can be • a piece of literal data, such as a number 2.3 or a string "Hello world!"

• a variable name

• a procedure invocation expression

• one of Scheme’s special syntactic expressions.

The following subsections describe how each of these types of expression is evaluated.

3.3.1.1 Evaluating Literal Data

When a literal data expression is evaluated, the value of the expression is simply the value that the expression describes. The evaluation of a literal data expression has no side effects. So, for example,

• the value of the expression "abc" is the string value "abc" • the value of the expression 3+4i is the complex number 3 + 4i

1 These definitions are approximate. For the whole and detailed truth, see Section “Formal syntax and

semantics” in The Revised(5) Report on the Algorithmic Language Scheme.

• the value of the expression #(1 2 3) is a three-element vector containing the numeric values 1, 2 and 3.

For any data type which can be expressed literally like this, the syntax of the literal data expression for that data type — in other words, what you need to write in your code to indicate a literal value of that type — is known as the data type’s read syntax. This manual specifies the read syntax for each such data type in the section that describes that data type.

Some data types do not have a read syntax. Procedures, for example, cannot be expressed as literal data; they must be created using a lambda expression (see Section 3.2.3 [Creating a Procedure], page 19) or implicitly using the shorthand form of define (see Section 3.2.4 [Lambda Alternatives], page 20).

3.3.1.2 Evaluating a Variable Reference

When an expression that consists simply of a variable name is evaluated, the value of the expression is the value of the named variable. The evaluation of a variable reference expression has no side effects.

So, after

     (define key "Paul Evans")

the value of the expression key is the string value "Paul Evans". If key is then modified by

     (set! key 3.74)

the value of the expression key is the numeric value 3.74.

If there is no variable with the specified name, evaluation of the variable reference expression signals an error.

3.3.1.3 Evaluating a Procedure Invocation Expression

This is where evaluation starts getting interesting! As already noted, a procedure invocation expression has the form

     (procedure [arg1 [arg2 ...]])

where procedure must be an expression whose value, when evaluated, is a procedure. The evaluation of a procedure invocation expression like this proceeds by

• evaluating individually the expressions procedure, arg1, arg2, and so on

• calling the procedure that is the value of the procedure expression with the list of values obtained from the evaluations of arg1, arg2 etc. as its parameters.

For a procedure defined in Scheme, “calling the procedure with the list of values as its parameters” means binding the values to the procedure’s formal parameters and then evaluating the sequence of expressions that make up the body of the procedure definition. The value of the procedure invocation expression is the value of the last evaluated expression in the procedure body. The side effects of calling the procedure are the combination of the side effects of the sequence of evaluations of expressions in the procedure body.

For a built-in procedure, the value and side-effects of calling the procedure are best described by that procedure’s documentation.

Note that the complete side effects of evaluating a procedure invocation expression consist not only of the side effects of the procedure call, but also of any side effects of the preceding evaluation of the expressions procedure, arg1, arg2, and so on.

To illustrate this, let’s look again at the procedure invocation expression:

     (string-length (string-append "/home" "/" "andrew"))

In the outermost expression, procedure is string-length and arg1 is (string-append

"/home" "/" "andrew").

• Evaluation of string-length, which is a variable, gives a procedure value that implements the expected behaviour for “string-length”.

• Evaluation of (string-append "/home" "/" "andrew"), which is another procedure

invocation expression, means evaluating each of

• string-append, which gives a procedure value that implements the expected behaviour for “string-append”

• "/home", which gives the string value "/home"

• "/", which gives the string value "/"

• "andrew", which gives the string value "andrew"

and then invoking the procedure value with this list of string values as its arguments. The resulting value is a single string value that is the concatenation of all the arguments,

    namely "/home/andrew".

In the evaluation of the outermost expression, the interpreter can now invoke the procedure value obtained from procedure with the value obtained from arg1 as its arguments. The resulting value is a numeric value that is the length of the argument string, which is 12.

3.3.1.4 Evaluating Special Syntactic Expressions

When a procedure invocation expression is evaluated, the procedure and all the argument expressions must be evaluated before the procedure can be invoked. Special syntactic expressions are special because they are able to manipulate their arguments in an unevaluated form, and can choose whether to evaluate any or all of the argument expressions.

Why is this needed? Consider a program fragment that asks the user whether or not to delete a file, and then deletes the file if the user answers yes.

     (if (string=? (read-answer "Should I delete this file?")
                   "yes")
         (delete-file file))

If the outermost (if ...) expression here was a procedure invocation expression, the expression (delete-file file), whose side effect is to actually delete a file, would already have been evaluated before the if procedure even got invoked! Clearly this is no use — the whole point of an if expression is that the consequent expression is only evaluated if the condition of the if expression is “true”.

Therefore if must be special syntax, not a procedure. Other special syntaxes that we have already met are define, set! and lambda. define and set! are syntax because they need to know the variable name that is given as the first argument in a define or set! expression, not that variable’s value. lambda is syntax because it does not immediately evaluate the expressions that define the procedure body; instead it creates a procedure object that incorporates these expressions so that they can be evaluated in the future, when that procedure is invoked.

The rules for evaluating each special syntactic expression are specified individually for each special syntax. For a summary of standard special syntax, see See Section 3.3.4 [Syntax Summary], page 25.

3.3.2 Tail calls

Scheme is “properly tail recursive”, meaning that tail calls or recursions from certain contexts do not consume stack space or other resources and can therefore be used on arbitrarily large data or for an arbitrarily long calculation. Consider for example,

     (define (foo n)
       (display n)
       (newline)
       (foo (1+ n)))
     (foo 1)

⊣

     1
     2
     3
     ...

foo prints numbers infinitely, starting from the given n. It’s implemented by printing n then recursing to itself to print n + 1 and so on. This recursion is a tail call, it’s the last thing done, and in Scheme such tail calls can be made without limit.

Or consider a case where a value is returned, a version of the SRFI-1 last function (see Section 7.5.3.3 [SRFI-1 Selectors], page 554) returning the last element of a list,

     (define (my-last lst)
       (if (null? (cdr lst))
           (car lst)
           (my-last (cdr lst))))
     (my-last '(1 2 3)) ⇒ 3

If the list has more than one element, my-last applies itself to the cdr. This recursion is a tail call, there’s no code after it, and the return value is the return value from that call. In Scheme this can be used on an arbitrarily long list argument.

A proper tail call is only available from certain contexts, namely the following special form positions,

• and — last expression

• begin — last expression

• case — last expression in each clause

• cond — last expression in each clause, and the call to a => procedure is a tail call • do — last result expression • if — “true” and “false” leg expressions

• lambda — last expression in body

  • let, let*, letrec, let-syntax, letrec-syntax — last expression in body

• or — last expression

The following core functions make tail calls, • apply — tail call to given procedure

• call-with-current-continuation — tail call to the procedure receiving the new continuation

• call-with-values — tail call to the values-receiving procedure

• eval — tail call to evaluate the form

• string-any, string-every — tail call to predicate on the last character (if that point is reached)

The above are just core functions and special forms. Tail calls in other modules are described with the relevant documentation, for example SRFI-1 any and every (see Section 7.5.3.7 [SRFI-1 Searching], page 560).

It will be noted there are a lot of places which could potentially be tail calls, for instance the last call in a for-each, but only those explicitly described are guaranteed.

3.3.3 Using the Guile REPL

If you start Guile without specifying a particular program for it to execute, Guile enters its standard Read Evaluate Print Loop — or REPL for short. In this mode, Guile repeatedly reads in the next Scheme expression that the user types, evaluates it, and prints the resulting value.

The REPL is a useful mechanism for exploring the evaluation behaviour described in the previous subsection. If you type string-append, for example, the REPL replies #<primitive-procedure string-append>, illustrating the relationship between the variable string-append and the procedure value stored in that variable.

In this manual, the notation ⇒ is used to mean “evaluates to”. Wherever you see an example of the form

     expression

⇒

     result

feel free to try it out yourself by typing expression into the REPL and checking that it gives the expected result.

3.3.4 Summary of Common Syntax

This subsection lists the most commonly used Scheme syntactic expressions, simply so that you will recognize common special syntax when you see it. For a full description of each of these syntaxes, follow the appropriate reference.

lambda (see Section 6.9.1 [Lambda], page 244) is used to construct procedure objects. define (see Section 6.12.1 [Top Level], page 286) is used to create a new variable and

set its initial value.

set! (see Section 6.12.1 [Top Level], page 286) is used to modify an existing variable’s value.

let, let* and letrec (see Section 6.12.2 [Local Bindings], page 287) create an inner lexical environment for the evaluation of a sequence of expressions, in which a specified set of local variables is bound to the values of a corresponding set of expressions. For an introduction to environments, see See Section 3.4 [About Closure], page 26.

begin (see Section 6.13.1 [begin], page 290) executes a sequence of expressions in order and returns the value of the last expression. Note that this is not the same as a procedure which returns its last argument, because the evaluation of a procedure invocation expression does not guarantee to evaluate the arguments in order.

if and cond (see Section 6.13.2 [Conditionals], page 292) provide conditional evaluation of argument expressions depending on whether one or more conditions evaluate to “true” or “false”.

case (see Section 6.13.2 [Conditionals], page 292) provides conditional evaluation of argument expressions depending on whether a variable has one of a specified group of values.

and (see Section 6.13.3 [and or], page 293) executes a sequence of expressions in order until either there are no expressions left, or one of them evaluates to “false”.

or (see Section 6.13.3 [and or], page 293) executes a sequence of expressions in order until either there are no expressions left, or one of them evaluates to “true”.

3.4 The Concept of Closure

The concept of closure is the idea that a lambda expression “captures” the variable bindings that are in lexical scope at the point where the lambda expression occurs. The procedure created by the lambda expression can refer to and mutate the captured bindings, and the values of those bindings persist between procedure calls.

This section explains and explores the various parts of this idea in more detail.

3.4.1 Names, Locations, Values and Environments

We said earlier that a variable name in a Scheme program is associated with a location in which any kind of Scheme value may be stored. (Incidentally, the term “vcell” is often used in Lisp and Scheme circles as an alternative to “location”.) Thus part of what we mean when we talk about “creating a variable” is in fact establishing an association between a name, or identifier, that is used by the Scheme program code, and the variable location to which that name refers. Although the value that is stored in that location may change, the location to which a given name refers is always the same.

We can illustrate this by breaking down the operation of the define syntax into three

parts: define

• creates a new location

• establishes an association between that location and the name specified as the first argument of the define expression

• stores in that location the value obtained by evaluating the second argument of the define expression.

A collection of associations between names and locations is called an environment. When you create a top level variable in a program using define, the name-location association for that variable is added to the “top level” environment. The “top level” environment also includes name-location associations for all the procedures that are supplied by standard Scheme.

It is also possible to create environments other than the top level one, and to create variable bindings, or name-location associations, in those environments. This ability is a key ingredient in the concept of closure; the next subsection shows how it is done.

3.4.2 Local Variables and Environments

We have seen how to create top level variables using the define syntax (see Section 3.1.3 [Definition], page 16). It is often useful to create variables that are more limited in their scope, typically as part of a procedure body. In Scheme, this is done using the let syntax, or one of its modified forms let* and letrec. These syntaxes are described in full later in the manual (see Section 6.12.2 [Local Bindings], page 287). Here our purpose is to illustrate their use just enough that we can see how local variables work.

For example, the following code uses a local variable s to simplify the computation of the area of a triangle given the lengths of its three sides.

     (define a 5.3)
     (define b 4.7)
     (define c 2.8)
     (define area
       (let ((s (/ (+ a b c) 2)))
         (sqrt (* s (- s a) (- s b) (- s c)))))

The effect of the let expression is to create a new environment and, within this environment, an association between the name s and a new location whose initial value is obtained by evaluating (/ (+ a b c) 2). The expressions in the body of the let, namely (sqrt (* s (- s a) (- s b) (- s c))), are then evaluated in the context of the new environment, and the value of the last expression evaluated becomes the value of the whole let expression, and therefore the value of the variable area.

3.4.3 Environment Chaining

In the example of the previous subsection, we glossed over an important point. The body of the let expression in that example refers not only to the local variable s, but also to the top level variables a, b, c and sqrt. (sqrt is the standard Scheme procedure for calculating a square root.) If the body of the let expression is evaluated in the context of the local let environment, how does the evaluation get at the values of these top level variables?

The answer is that the local environment created by a let expression automatically has a reference to its containing environment — in this case the top level environment — and that the Scheme interpreter automatically looks for a variable binding in the containing environment if it doesn’t find one in the local environment. More generally, every environment except for the top level one has a reference to its containing environment, and the interpreter keeps searching back up the chain of environments — from most local to top level — until it either finds a variable binding for the required identifier or exhausts the chain.

This description also determines what happens when there is more than one variable binding with the same name. Suppose, continuing the example of the previous subsection, that there was also a pre-existing top level variable s created by the expression:

     (define s "Some beans, my lord!")

Then both the top level environment and the local let environment would contain bindings for the name s. When evaluating code within the let body, the interpreter looks first in the local let environment, and so finds the binding for s created by the let syntax. Even though this environment has a reference to the top level environment, which also has a binding for s, the interpreter doesn’t get as far as looking there. When evaluating code outside the let body, the interpreter looks up variable names in the top level environment, so the name s refers to the top level variable.

Within the let body, the binding for s in the local environment is said to shadow the binding for s in the top level environment.

3.4.4 Lexical Scope

The rules that we have just been describing are the details of how Scheme implements “lexical scoping”. This subsection takes a brief diversion to explain what lexical scope means in general and to present an example of non-lexical scoping.

“Lexical scope” in general is the idea that

• an identifier at a particular place in a program always refers to the same variable location — where “always” means “every time that the containing expression is executed”, and that

• the variable location to which it refers can be determined by static examination of the source code context in which that identifier appears, without having to consider the flow of execution through the program as a whole.

In practice, lexical scoping is the norm for most programming languages, and probably corresponds to what you would intuitively consider to be “normal”. You may even be wondering how the situation could possibly — and usefully — be otherwise. To demonstrate that another kind of scoping is possible, therefore, and to compare it against lexical scoping, the following subsection presents an example of non-lexical scoping and examines in detail how its behavior differs from the corresponding lexically scoped code.

3.4.4.1 An Example of Non-Lexical Scoping

To demonstrate that non-lexical scoping does exist and can be useful, we present the following example from Emacs Lisp, which is a “dynamically scoped” language.

     (defvar currency-abbreviation "USD")
     (defun currency-string (units hundredths)
       (concat currency-abbreviation
               (number-to-string units)
               "."
               (number-to-string hundredths)))
     (defun french-currency-string (units hundredths)
       (let ((currency-abbreviation "FRF"))

         (currency-string units hundredths)))

The question to focus on here is: what does the identifier currency-abbreviation refer to in the currency-string function? The answer, in Emacs Lisp, is that all variable bindings go onto a single stack, and that currency-abbreviation refers to the topmost binding from that stack which has the name “currency-abbreviation”. The binding that is created by the defvar form, to the value "USD", is only relevant if none of the code that calls currency-string rebinds the name “currency-abbreviation” in the meanwhile.

The second function french-currency-string works precisely by taking advantage of this behaviour. It creates a new binding for the name “currency-abbreviation” which overrides the one established by the defvar form.

     ;; Note!  This is Emacs Lisp evaluation, not Scheme!
     (french-currency-string 33 44)

⇒

     "FRF33.44"

Now let’s look at the corresponding, lexically scoped Scheme code:

     (define currency-abbreviation "USD")
     (define (currency-string units hundredths)
       (string-append currency-abbreviation
                      (number->string units)
                      "."
                      (number->string hundredths)))
     (define (french-currency-string units hundredths)
       (let ((currency-abbreviation "FRF"))
         (currency-string units hundredths)))

According to the rules of lexical scoping, the currency-abbreviation in currencystring refers to the variable location in the innermost environment at that point in the code which has a binding for currency-abbreviation, which is the variable location in the top level environment created by the preceding (define currency-abbreviation ...) expression.

In Scheme, therefore, the french-currency-string procedure does not work as intended. The variable binding that it creates for “currency-abbreviation” is purely local to the code that forms the body of the let expression. Since this code doesn’t directly use the name “currency-abbreviation” at all, the binding is pointless.

     (french-currency-string 33 44)

⇒

     "USD33.44"

This begs the question of how the Emacs Lisp behaviour can be implemented in Scheme. In general, this is a design question whose answer depends upon the problem that is being addressed. In this case, the best answer may be that currency-string should be redesigned so that it can take an optional third argument. This third argument, if supplied, is interpreted as a currency abbreviation that overrides the default.

It is possible to change french-currency-string so that it mostly works without changing currency-string, but the fix is inelegant, and susceptible to interrupts that could leave the currency-abbreviation variable in the wrong state:

     (define (french-currency-string units hundredths)
       (set! currency-abbreviation "FRF")
       (let ((result (currency-string units hundredths)))
         (set! currency-abbreviation "USD")
         result))

The key point here is that the code does not create any local binding for the identifier currency-abbreviation, so all occurrences of this identifier refer to the top level variable.

3.4.5 Closure

Consider a let expression that doesn’t contain any lambdas:

     (let ((s (/ (+ a b c) 2)))
       (sqrt (* s (- s a) (- s b) (- s c))))

When the Scheme interpreter evaluates this, it

• creates a new environment with a reference to the environment that was current when it encountered the let

• creates a variable binding for s in the new environment, with value given by (/ (+ a b

    c) 2)

• evaluates the expression in the body of the let in the context of the new local environment, and remembers the value V

• forgets the local environment

• continues evaluating the expression that contained the let, using the value V as the value of the let expression, in the context of the containing environment.

After the let expression has been evaluated, the local environment that was created is simply forgotten, and there is no longer any way to access the binding that was created in this environment. If the same code is evaluated again, it will follow the same steps again, creating a second new local environment that has no connection with the first, and then forgetting this one as well.

If the let body contains a lambda expression, however, the local environment is not forgotten. Instead, it becomes associated with the procedure that is created by the lambda expression, and is reinstated every time that that procedure is called. In detail, this works as follows.

• When the Scheme interpreter evaluates a lambda expression, to create a procedure object, it stores the current environment as part of the procedure definition.

• Then, whenever that procedure is called, the interpreter reinstates the environment that is stored in the procedure definition and evaluates the procedure body within the context of that environment.

The result is that the procedure body is always evaluated in the context of the environment that was current when the procedure was created.

This is what is meant by closure. The next few subsections present examples that explore the usefulness of this concept.

3.4.6 Example 1: A Serial Number Generator

This example uses closure to create a procedure with a variable binding that is private to the procedure, like a local variable, but whose value persists between procedure calls.

     (define (make-serial-number-generator)
       (let ((current-serial-number 0))
         (lambda ()
           (set! current-serial-number (+ current-serial-number 1))
           current-serial-number)))
     (define entry-sn-generator (make-serial-number-generator))
     (entry-sn-generator)

⇒

     1
     (entry-sn-generator)

⇒

     2

When make-serial-number-generator is called, it creates a local environment with a binding for current-serial-number whose initial value is 0, then, within this environment, creates a procedure. The local environment is stored within the created procedure object and so persists for the lifetime of the created procedure.

Every time the created procedure is invoked, it increments the value of the currentserial-number binding in the captured environment and then returns the current value.

Note that make-serial-number-generator can be called again to create a second serial number generator that is independent of the first. Every new invocation of make-serial-number-generator creates a new local let environment and returns a new procedure object with an association to this environment.

3.4.7 Example 2: A Shared Persistent Variable

This example uses closure to create two procedures, get-balance and deposit, that both refer to the same captured local environment so that they can both access the balance variable binding inside that environment. The value of this variable binding persists between calls to either procedure.

Note that the captured balance variable binding is private to these two procedures: it is not directly accessible to any other code. It can only be accessed indirectly via get-balance or deposit, as illustrated by the withdraw procedure.

     (define get-balance #f)
     (define deposit #f)
     (let ((balance 0))
       (set! get-balance
             (lambda ()
               balance))
       (set! deposit

             (lambda (amount)
               (set! balance (+ balance amount))
               balance)))
     (define (withdraw amount)
       (deposit (- amount)))
     (get-balance)

⇒

     0
     (deposit 50)

⇒

     50
     (withdraw 75)

⇒

     -25

An important detail here is that the get-balance and deposit variables must be set up by defineing them at top level and then set!ing their values inside the let body. Using define within the let body would not work: this would create variable bindings within the local let environment that would not be accessible at top level.

3.4.8 Example 3: The Callback Closure Problem

A frequently used programming model for library code is to allow an application to register a callback function for the library to call when some particular event occurs. It is often useful for the application to make several such registrations using the same callback function, for example if several similar library events can be handled using the same application code, but the need then arises to distinguish the callback function calls that are associated with one callback registration from those that are associated with different callback registrations.

In languages without the ability to create functions dynamically, this problem is usually solved by passing a user_data parameter on the registration call, and including the value of this parameter as one of the parameters on the callback function. Here is an example of declarations using this solution in C:

     typedef void (event_handler_t) (int event_type,
                                     void *user_data);
     void register_callback (int event_type,
                             event_handler_t *handler,
                             void *user_data);

In Scheme, closure can be used to achieve the same functionality without requiring the library code to store a user-data for each callback registration.

     ;; In the library:
     (define (register-callback event-type handler-proc)
       ...)

     ;; In the application:
     (define (make-handler event-type user-data)
       (lambda ()
         ...
         <code referencing event-type and user-data>
         ...))
     (register-callback event-type
                        (make-handler event-type ...))

As far as the library is concerned, handler-proc is a procedure with no arguments, and all the library has to do is call it when the appropriate event occurs. From the application’s point of view, though, the handler procedure has used closure to capture an environment that includes all the context that the handler code needs — event-type and user-data — to handle the event correctly.

3.4.9 Example 4: Object Orientation

Closure is the capture of an environment, containing persistent variable bindings, within the definition of a procedure or a set of related procedures. This is rather similar to the idea in some object oriented languages of encapsulating a set of related data variables inside an “object”, together with a set of “methods” that operate on the encapsulated data. The following example shows how closure can be used to emulate the ideas of objects, methods and encapsulation in Scheme.

     (define (make-account)
       (let ((balance 0))
         (define (get-balance)
           balance)
         (define (deposit amount)
           (set! balance (+ balance amount))
           balance)
         (define (withdraw amount)
           (deposit (- amount)))
         (lambda args
           (apply
             (case (car args)
               ((get-balance) get-balance)
               ((deposit) deposit)
               ((withdraw) withdraw)
               (else (error "Invalid method!")))
             (cdr args)))))

Each call to make-account creates and returns a new procedure, created by the expression in the example code that begins “(lambda args”.

     (define my-account (make-account))

my-account

⇒

#<procedure args>

This procedure acts as an account object with methods get-balance, deposit and withdraw. To apply one of the methods to the account, you call the procedure with a symbol indicating the required method as the first parameter, followed by any other parameters that are required by that method.

(my-account 'get-balance)

⇒

0
(my-account 'withdraw 5)

⇒

-5
(my-account 'deposit 396)

⇒

391
(my-account 'get-balance)

⇒

391

Note how, in this example, both the current balance and the helper procedures getbalance, deposit and withdraw, used to implement the guts of the account object’s methods, are all stored in variable bindings within the private local environment captured by the lambda expression that creates the account object procedure.

3.5 Further Reading

• The website http://www.schemers.org/ is a good starting point for all things Scheme.

• Dorai Sitaram’s online Scheme tutorial, Teach Yourself Scheme in Fixnum Days, at

http://www.ccs.neu.edu/home/dorai/t-y-scheme/t-y-scheme.html. Includes a

nice explanation of continuations.

• The complete text of Structure and Interpretation of Computer Programs, the classic introduction to computer science and Scheme by Hal Abelson, Jerry Sussman and Julie Sussman, is now available online at http://mitpress.mit.edu/sicp/sicp. html. This site also provides teaching materials related to the book, and all the source code used in the book, in a form suitable for loading and running.

4 Programming in Scheme

Guile’s core language is Scheme, and a lot can be achieved simply by using Guile to write and run Scheme programs — as opposed to having to dive into C code. In this part of the manual, we explain how to use Guile in this mode, and describe the tools that Guile provides to help you with script writing, debugging, and packaging your programs for distribution.

For detailed reference information on the variables, functions, and so on that make up Guile’s application programming interface (API), see Chapter 6 [API Reference], page 99.

4.1 Guile’s Implementation of Scheme

Guile’s core language is Scheme, which is specified and described in the series of reports known as RnRS. RnRS is shorthand for the Revised$^n$ Report on the Algorithmic Language Scheme. Guile complies fully with R5RS (see Section “Introduction” in R5RS), and implements some aspects of R6RS.

Guile also has many extensions that go beyond these reports. Some of the areas where Guile extends R5RS are:

• Guile’s interactive documentation system

• Guile’s support for POSIX-compliant network programming

• GOOPS – Guile’s framework for object oriented programming.

4.2 Invoking Guile

Many features of Guile depend on and can be changed by information that the user provides either before or when Guile is started. Below is a description of what information to provide and how to provide it.

4.2.1 Command-line Options

Here we describe Guile’s command-line processing in detail. Guile processes its arguments from left to right, recognizing the switches described below. For examples, see Section 4.3.4 [Scripting Examples], page 44.

script arg...
-s script arg...

By default, Guile will read a file named on the command line as a script. Any command-line arguments arg... following script become the script’s arguments; the command-line function returns a list of strings of the form (script

          arg...).

It is possible to name a file using a leading hyphen, for example, -myfile.scm. In this case, the file name must be preceded by -s to tell Guile that a (script) file is being named. Scripts are read and evaluated as Scheme source code just as the load function would. After loading script, Guile exits.

-c expr arg...

Evaluate expr as Scheme code, and then exit. Any command-line arguments arg... following expr become command-line arguments; the command-line function returns a list of strings of the form (guile arg...), where guile is the path of the Guile executable.

-- arg... Run interactively, prompting the user for expressions and evaluating them. Any command-line arguments arg... following the -- become command-line arguments for the interactive session; the command-line function returns a list of strings of the form (guile arg...), where guile is the path of the Guile executable.

-L directory

Add directory to the front of Guile’s module load path. The given directories are searched in the order given on the command line and before any directories in the GUILE_LOAD_PATH environment variable. Paths added here are not in effect during execution of the user’s .guile file.

-C directory

Like -L, but adjusts the load path for compiled files.

-x extension

Add extension to the front of Guile’s load extension list (see Section 6.17.7 [Load Paths], page 368). The specified extensions are tried in the order given on the command line, and before the default load extensions. Extensions added here are not in effect during execution of the user’s .guile file.

-l file Load Scheme source code from file, and continue processing the command line.

-e function

Make function the entry point of the script. After loading the script file (with -s) or evaluating the expression (with -c), apply function to a list containing the program name and the command-line arguments—the list provided by the

command-line function.

A -e switch can appear anywhere in the argument list, but Guile always invokes the function as the last action it performs. This is weird, but because of the way script invocation works under POSIX, the -s option must always come last in the list. The function is most often a simple symbol that names a function that is defined in the script. It can also be of the form (@ module-name symbol), and in that case, the symbol is looked up in the module named module-name. For compatibility with some versions of Guile 1.4, you can also use the form (symbol ...) (that is, a list of only symbols that doesn’t start with @), which

is equivalent to (@ (symbol ...) main), or (symbol ...) symbol (that is, a

list of only symbols followed by a symbol), which is equivalent to (@ (symbol ...) symbol). We recommend to use the equivalent forms directly since they correspond to the (@ ...) read syntax that can be used in normal code. See Section 6.19.2 [Using Guile Modules], page 382 and Section 4.3.4 [Scripting Examples], page 44.

-ds Treat a final -s option as if it occurred at this point in the command line; load the script here. This switch is necessary because, although the POSIX script invocation mechanism effectively requires the -s option to appear last, the programmer may well want to run the script before other actions requested on the command line. For examples, see Section 4.3.4 [Scripting Examples], page 44.

\ Read more command-line arguments, starting from the second line of the script file. See Section 4.3.2 [The Meta Switch], page 41.

--use-srfi=list

The option --use-srfi expects a comma-separated list of numbers, each representing a SRFI module to be loaded into the interpreter before evaluating a script file or starting the REPL. Additionally, the feature identifier for the loaded SRFIs is recognized by the procedure cond-expand when this option is used. Here is an example that loads the modules SRFI-8 (’receive’) and SRFI-13 (’string library’) before the GUILE interpreter is started:

     guile --use-srfi=8,13

--debug Start with the debugging virtual machine (VM) engine. Using the debugging VM will enable support for VM hooks, which are needed for tracing, breakpoints, and accurate call counts when profiling. The debugging VM is slower than the regular VM, though, by about ten percent. See Section 6.25.4.1 [VM Hooks], page 455, for more information. By default, the debugging VM engine is only used when entering an interactive session. When executing a script with -s or -c, the normal, faster VM is used by default.

--no-debug

Do not use the debugging VM engine, even when entering an interactive session. Note that, despite the name, Guile running with --no-debug does support the usual debugging facilities, such as printing a detailed backtrace upon error. The only difference with --debug is lack of support for VM hooks and the facilities that build upon it (see above).

-q Do not load the initialization file, .guile. This option only has an effect when running interactively; running scripts does not load the .guile file. See Section 4.4.1 [Init File], page 46.

--listen[=p]

While this program runs, listen on a local port or a path for REPL clients. If p starts with a number, it is assumed to be a local port on which to listen. If it starts with a forward slash, it is assumed to be a path to a UNIX domain socket on which to listen. If p is not given, the default is local port 37146. If you look at it upside down, it almost spells “Guile”. If you have netcat installed, you should be able to nc localhost 37146 and get a Guile prompt. Alternately you can fire up Emacs and connect to the process; see Section 4.5 [Using Guile in Emacs], page 54 for more details. Note that opening a port allows anyone who can connect to that port—in the TCP case, any local user—to do anything Guile can do, as the user that the Guile process is running as. Do not use --listen on multi-user machines. Of course, if you do not pass --listen to Guile, no port will be opened. That said, --listen is great for interactive debugging and development.

--auto-compile

Compile source files automatically (default behavior).

--fresh-auto-compile

Treat the auto-compilation cache as invalid, forcing recompilation.

--no-auto-compile

Disable automatic source file compilation.

--language=lang

For the remainder of the command line arguments, assume that files mentioned with -l and expressions passed with -c are written in lang. lang must be the name of one of the languages supported by the compiler (see Section 9.4.1 [Compiler Tower], page 792). When run interactively, set the REPL’s language to lang (see Section 4.4 [Using Guile Interactively], page 46). The default language is scheme; other interesting values include elisp (for Emacs Lisp), and ecmascript. The example below shows the evaluation of expressions in Scheme, Emacs Lisp, and ECMAScript:

     guile -c "(apply + '(1 2))"
     guile --language=elisp -c "(= (funcall (symbol-function '+) 1 2) 3)"
     guile --language=ecmascript -c '(function (x) { return x * x; })(2);'

To load a file written in Scheme and one written in Emacs Lisp, and then start a Scheme REPL, type:

     guile -l foo.scm --language=elisp -l foo.el --language=scheme
-h, --help

Display help on invoking Guile, and then exit.

-v, --version

Display the current version of Guile, and then exit.

4.2.2 Environment Variables

The environment is a feature of the operating system; it consists of a collection of variables with names and values. Each variable is called an environment variable (or, sometimes, a “shell variable”); environment variable names are case-sensitive, and it is conventional to use upper-case letters only. The values are all text strings, even those that are written as numerals. (Note that here we are referring to names and values that are defined in the operating system shell from which Guile is invoked. This is not the same as a Scheme environment that is defined within a running instance of Guile. For a description of Scheme environments, see Section 3.4.1 [About Environments], page 26.)

How to set environment variables before starting Guile depends on the operating system and, especially, the shell that you are using. For example, here is how to tell Guile to provide detailed warning messages about deprecated features by setting GUILE_WARN_DEPRECATED using Bash:

$ export GUILE_WARN_DEPRECATED="detailed"
$ guile

Or, detailed warnings can be turned on for a single invocation using:

$ env GUILE_WARN_DEPRECATED="detailed" guile

If you wish to retrieve or change the value of the shell environment variables that affect the run-time behavior of Guile from within a running instance of Guile, see Section 7.2.6 [Runtime Environment], page 488.

Here are the environment variables that affect the run-time behavior of Guile:

GUILE_AUTO_COMPILE

This is a flag that can be used to tell Guile whether or not to compile Scheme source files automatically. Starting with Guile 2.0, Scheme source files will be compiled automatically, by default. If a compiled (.go) file corresponding to a .scm file is not found or is not newer than the .scm file, the .scm file will be compiled on the fly, and the resulting .go file stored away. An advisory note will be printed on the console. Compiled files will be stored in the directory $XDG_CACHE_HOME/guile/ccache, where XDG_CACHE_HOME defaults to the directory $HOME/.cache. This directory will be created if it does not already exist. Note that this mechanism depends on the timestamp of the .go file being newer than that of the .scm file; if the .scm or .go files are moved after installation, care should be taken to preserve their original timestamps. Set GUILE_AUTO_COMPILE to zero (0), to prevent Scheme files from being compiled automatically. Set this variable to “fresh” to tell Guile to compile Scheme files whether they are newer than the compiled files or not. See Section 6.17.5 [Compilation], page 364.

GUILE_HISTORY

This variable names the file that holds the Guile REPL command history. You can specify a different history file by setting this environment variable. By default, the history file is $HOME/.guile_history.

GUILE_INSTALL_LOCALE

This is a flag that can be used to tell Guile whether or not to install the current locale at startup, via a call to (setlocale LC_ALL ""). See Section 7.2.13 [Locales], page 518, for more information on locales. You may explicitly indicate that you do not want to install the locale by setting GUILE_INSTALL_LOCALE to 0, or explicitly enable it by setting the variable to

1.

Usually, installing the current locale is the right thing to do. It allows Guile to correctly parse and print strings with non-ASCII characters. However, for compatibility with previous Guile 2.0 releases, this option is off by default. The next stable release series of Guile (the 2.2 series) will install locales by default.

GUILE_STACK_SIZE

Guile currently has a limited stack size for Scheme computations. Attempting to call too many nested functions will signal an error. This is good to detect infinite recursion, but sometimes the limit is reached for normal computations. This environment variable, if set to a positive integer, specifies the number of Scheme value slots to allocate for the stack. In the future we will implement stacks that can grow and shrink, but for now this hack will have to do.

GUILE_LOAD_COMPILED_PATH

This variable may be used to augment the path that is searched for compiled Scheme files (.go files) when loading. Its value should be a colon-separated list of directories. If it contains the special path component ... (ellipsis), then the default path is put in place of the ellipsis, otherwise the default path is placed at the end. The result is stored in %load-compiled-path (see Section 6.17.7 [Load Paths], page 368). Here is an example using the Bash shell that adds the current directory, ., and

the relative directory ../my-library to %load-compiled-path:
     $ export GUILE_LOAD_COMPILED_PATH=".:../my-library"
     $ guile -c '(display %load-compiled-path) (newline)'
     (. ../my-library /usr/local/lib/guile/2.0/ccache)
GUILE_LOAD_PATH

This variable may be used to augment the path that is searched for Scheme files when loading. Its value should be a colon-separated list of directories. If it contains the special path component ... (ellipsis), then the default path is put in place of the ellipsis, otherwise the default path is placed at the end. The result is stored in %load-path (see Section 6.17.7 [Load Paths], page 368). Here is an example using the Bash shell that prepends the current directory to %load-path, and adds the relative directory ../srfi to the end:

     $ env GUILE_LOAD_PATH=".:...:../srfi" \
     guile -c '(display %load-path) (newline)'
     (. /usr/local/share/guile/2.0 \
     /usr/local/share/guile/site/2.0 \
     /usr/local/share/guile/site \
     /usr/local/share/guile \
     ../srfi)

(Note: The line breaks, above, are for documentation purposes only, and not required in the actual example.)

GUILE_WARN_DEPRECATED

As Guile evolves, some features will be eliminated or replaced by newer features. To help users migrate their code as this evolution occurs, Guile will issue warning messages about code that uses features that have been marked for eventual elimination. GUILE_WARN_DEPRECATED can be set to “no” to tell Guile not to display these warning messages, or set to “detailed” to tell Guile to display more lengthy messages describing the warning. See Section 6.2 [Deprecation], page 100.

HOME Guile uses the environment variable HOME, the name of your home directory, to locate various files, such as .guile or .guile_history.

4.3 Guile Scripting

Like AWK, Perl, or any shell, Guile can interpret script files. A Guile script is simply a file of Scheme code with some extra information at the beginning which tells the operating system how to invoke Guile, and then tells Guile how to handle the Scheme code.

4.3.1 The Top of a Script File

The first line of a Guile script must tell the operating system to use Guile to evaluate the script, and then tell Guile how to go about doing that. Here is the simplest case:

• The first two characters of the file must be ‘#!’.

The operating system interprets this to mean that the rest of the line is the name of an executable that can interpret the script. Guile, however, interprets these characters as the beginning of a multi-line comment, terminated by the characters ‘!#’ on a line by themselves. (This is an extension to the syntax described in R5RS, added to support shell scripts.)

• Immediately after those two characters must come the full pathname to the Guile interpreter. On most systems, this would be ‘/usr/local/bin/guile’.

• Then must come a space, followed by a command-line argument to pass to Guile; this should be ‘-s’. This switch tells Guile to run a script, instead of soliciting the user for input from the terminal. There are more elaborate things one can do here; see Section 4.3.2 [The Meta Switch], page 41.

• Follow this with a newline.

• The second line of the script should contain only the characters ‘!#’ — just like the top of the file, but reversed. The operating system never reads this far, but Guile treats this as the end of the comment begun on the first line by the ‘#!’ characters.

• If this source code file is not ASCII or ISO-8859-1 encoded, a coding declaration such as coding: utf-8 should appear in a comment somewhere in the first five lines of the file: see Section 6.17.8 [Character Encoding of Source Files], page 370.

• The rest of the file should be a Scheme program.

Guile reads the program, evaluating expressions in the order that they appear. Upon reaching the end of the file, Guile exits.

4.3.2 The Meta Switch

Guile’s command-line switches allow the programmer to describe reasonably complicated actions in scripts. Unfortunately, the POSIX script invocation mechanism only allows one argument to appear on the ‘#!’ line after the path to the Guile executable, and imposes arbitrary limits on that argument’s length. Suppose you wrote a script starting like this:

 #!/usr/local/bin/guile -e main -s
 !#
 (define (main args)
   (map (lambda (arg) (display arg) (display " "))
        (cdr args))
   (newline))

The intended meaning is clear: load the file, and then call main on the command-line arguments. However, the system will treat everything after the Guile path as a single argument — the string "-e main -s" — which is not what we want.

As a workaround, the meta switch \ allows the Guile programmer to specify an arbitrary number of options without patching the kernel. If the first argument to Guile is \, Guile will open the script file whose name follows the \, parse arguments starting from the file’s second line (according to rules described below), and substitute them for the \ switch.

Working in concert with the meta switch, Guile treats the characters ‘#!’ as the beginning of a comment which extends through the next line containing only the characters ‘!#’. This sort of comment may appear anywhere in a Guile program, but it is most useful at the top of a file, meshing magically with the POSIX script invocation mechanism.

Thus, consider a script named /u/jimb/ekko which starts like this:

     #!/usr/local/bin/guile \
     -e main -s
     !#
     (define (main args)
             (map (lambda (arg) (display arg) (display " "))
                  (cdr args))
             (newline))

Suppose a user invokes this script as follows:

     $ /u/jimb/ekko a b c

Here’s what happens:

• the operating system recognizes the ‘#!’ token at the top of the file, and rewrites the

command line to:

         /usr/local/bin/guile \ /u/jimb/ekko a b c

This is the usual behavior, prescribed by POSIX.

• When Guile sees the first two arguments, \ /u/jimb/ekko, it opens /u/jimb/ekko, parses the three arguments -e, main, and -s from it, and substitutes them for the \ switch. Thus, Guile’s command line now reads:

         /usr/local/bin/guile -e main -s /u/jimb/ekko a b c

• Guile then processes these switches: it loads /u/jimb/ekko as a file of Scheme code (treating the first three lines as a comment), and then performs the application (main

    "/u/jimb/ekko" "a" "b" "c").

When Guile sees the meta switch \, it parses command-line argument from the script file according to the following rules:

• Each space character terminates an argument. This means that two spaces in a row introduce an argument "".

• The tab character is not permitted (unless you quote it with the backslash character, as described below), to avoid confusion.

• The newline character terminates the sequence of arguments, and will also terminate a final non-empty argument. (However, a newline following a space will not introduce a final empty-string argument; it only terminates the argument list.)

• The backslash character is the escape character. It escapes backslash, space, tab, and newline. The ANSI C escape sequences like \n and \t are also supported. These produce argument constituents; the two-character combination \n doesn’t act like a terminating newline. The escape sequence \NNN for exactly three octal digits reads as the character whose ASCII code is NNN. As above, characters produced this way are argument constituents. Backslash followed by other characters is not allowed.

4.3.3 Command Line Handling

The ability to accept and handle command line arguments is very important when writing Guile scripts to solve particular problems, such as extracting information from text files or interfacing with existing command line applications. This chapter describes how Guile makes command line arguments available to a Guile script, and the utilities that Guile provides to help with the processing of command line arguments.

When a Guile script is invoked, Guile makes the command line arguments accessible via the procedure command-line, which returns the arguments as a list of strings.

For example, if the script

     #! /usr/local/bin/guile -s
     !#
     (write (command-line))
     (newline)

is saved in a file cmdline-test.scm and invoked using the command line ./cmdline-

test.scm bar.txt -o foo -frumple grob, the output is
     ("./cmdline-test.scm" "bar.txt" "-o" "foo" "-frumple" "grob")

If the script invocation includes a -e option, specifying a procedure to call after loading the script, Guile will call that procedure with (command-line) as its argument. So a script that uses -e doesn’t need to refer explicitly to command-line in its code. For example, the script above would have identical behaviour if it was written instead like this:

     #! /usr/local/bin/guile \
     -e main -s
     !#
     (define (main args)
       (write args)
       (newline))

(Note the use of the meta switch \ so that the script invocation can include more than one Guile option: See Section 4.3.2 [The Meta Switch], page 41.)

These scripts use the #! POSIX convention so that they can be executed using their own file names directly, as in the example command line ./cmdline-test.scm bar.txt -o foo -frumple grob. But they can also be executed by typing out the implied Guile command line in full, as in:

     $ guile -s ./cmdline-test.scm bar.txt -o foo -frumple grob

or

     $ guile -e main -s ./cmdline-test2.scm bar.txt -o foo -frumple grob

Even when a script is invoked using this longer form, the arguments that the script receives are the same as if it had been invoked using the short form. Guile ensures that the (command-line) or -e arguments are independent of how the script is invoked, by stripping off the arguments that Guile itself processes.

A script is free to parse and handle its command line arguments in any way that it chooses. Where the set of possible options and arguments is complex, however, it can get tricky to extract all the options, check the validity of given arguments, and so on. This task can be greatly simplified by taking advantage of the module (ice-9 getopt-long), which is distributed with Guile, See Section 7.4 [getopt-long], page 546.

4.3.4 Scripting Examples

To start with, here are some examples of invoking Guile directly:

guile -- a b c

Run Guile interactively; (command-line) will return

          ("/usr/local/bin/guile" "a" "b" "c").
guile -s /u/jimb/ex2 a b c
          Load the file /u/jimb/ex2; (command-line) will return
          ("/u/jimb/ex2" "a" "b" "c").
guile -c '(write %load-path) (newline)'

Write the value of the variable %load-path, print a newline, and exit.

guile -e main -s /u/jimb/ex4 foo

Load the file /u/jimb/ex4, and then call the function main, passing it the list

          ("/u/jimb/ex4" "foo").
guile -l first -ds -l last -s script

Load the files first, script, and last, in that order. The -ds switch says when to process the -s switch. For a more motivated example, see the scripts below.

Here is a very simple Guile script:

     #!/usr/local/bin/guile -s
     !#
     (display "Hello, world!")
     (newline)

The first line marks the file as a Guile script. When the user invokes it, the system runs /usr/local/bin/guile to interpret the script, passing -s, the script’s filename, and any arguments given to the script as command-line arguments. When Guile sees -s script, it loads script. Thus, running this program produces the output:

     Hello, world!

Here is a script which prints the factorial of its argument:

     #!/usr/local/bin/guile -s
     !#
     (define (fact n)
       (if (zero? n) 1
         (* n (fact (- n 1)))))
     (display (fact (string->number (cadr (command-line)))))

(newline)

In action:

$ ./fact 5
120
$

However, suppose we want to use the definition of fact in this file from another script. We can’t simply load the script file, and then use fact’s definition, because the script will try to compute and display a factorial when we load it. To avoid this problem, we might write the script this way:

#!/usr/local/bin/guile \
-e main -s
!#
(define (fact n)
  (if (zero? n) 1
    (* n (fact (- n 1)))))
(define (main args)
  (display (fact (string->number (cadr args))))
  (newline))

This version packages the actions the script should perform in a function, main. This allows us to load the file purely for its definitions, without any extraneous computation taking place. Then we used the meta switch \ and the entry point switch -e to tell Guile to call main after loading the script.

$ ./fact 50
30414093201713378043612608166064768844377641568960512000000000000

Suppose that we now want to write a script which computes the choose function: given a set of m distinct objects, (choose n m) is the number of distinct subsets containing n objects each. It’s easy to write choose given fact, so we might write the script this way:

#!/usr/local/bin/guile \
-l fact -e main -s
!#
(define (choose n m)
  (/ (fact m) (* (fact (- m n)) (fact n))))
(define (main args)
  (let ((n (string->number (cadr args)))
        (m (string->number (caddr args))))
    (display (choose n m))
    (newline)))

The command-line arguments here tell Guile to first load the file fact, and then run the script, with main as the entry point. In other words, the choose script can use definitions made in the fact script. Here are some sample runs:

$ ./choose 0 4
1
$ ./choose 1 4

     4
     $ ./choose 2 4
     6
     $ ./choose 3 4
     4
     $ ./choose 4 4
     1
     $ ./choose 50 100
     100891344545564193334812497256

4.4 Using Guile Interactively

When you start up Guile by typing just guile, without a -c argument or the name of a script to execute, you get an interactive interpreter where you can enter Scheme expressions, and Guile will evaluate them and print the results for you. Here are some simple examples.

     scheme@(guile-user)> (+ 3 4 5)
     $1 = 12
     scheme@(guile-user)> (display "Hello world!\n")
     Hello world!
     scheme@(guile-user)> (values 'a 'b)
     $2 = a
     $3 = b

This mode of use is called a REPL, which is short for “Read-Eval-Print Loop”, because the Guile interpreter first reads the expression that you have typed, then evaluates it, and then prints the result.

The prompt shows you what language and module you are in. In this case, the current language is scheme, and the current module is (guile-user). See Section 6.23 [Other Languages], page 432, for more information on Guile’s support for languages other than Scheme.

4.4.1 The Init File, ~/.guile

When run interactively, Guile will load a local initialization file from ~/.guile. This file should contain Scheme expressions for evaluation.

This facility lets the user customize their interactive Guile environment, pulling in extra modules or parameterizing the REPL implementation.

To run Guile without loading the init file, use the -q command-line option.

4.4.2 Readline

To make it easier for you to repeat and vary previously entered expressions, or to edit the expression that you’re typing in, Guile can use the GNU Readline library. This is not enabled by default because of licensing reasons, but all you need to activate Readline is the following pair of lines.

     scheme@(guile-user)> (use-modules (ice-9 readline))
     scheme@(guile-user)> (activate-readline)

It’s a good idea to put these two lines (without the scheme@(guile-user)> prompts) in your .guile file. See Section 4.4.1 [Init File], page 46, for more on .guile.

4.4.3 Value History

Just as Readline helps you to reuse a previous input line, value history allows you to use the result of a previous evaluation in a new expression. When value history is enabled, each evaluation result is automatically assigned to the next in the sequence of variables $1, $2, . . .. You can then use these variables in subsequent expressions.

scheme@(guile-user)> (iota 10)
$1 = (0 1 2 3 4 5 6 7 8 9)
scheme@(guile-user)> (apply * (cdr $1))
$2 = 362880
scheme@(guile-user)> (sqrt $2)
$3 = 602.3952191045344
scheme@(guile-user)> (cons $2 $1)
$4 = (362880 0 1 2 3 4 5 6 7 8 9)

Value history is enabled by default, because Guile’s REPL imports the (ice-9 history) module. Value history may be turned off or on within the repl, using the options interface:

scheme@(guile-user)> ,option value-history #f
scheme@(guile-user)> 'foo
foo
scheme@(guile-user)> ,option value-history #t
scheme@(guile-user)> 'bar
$5 = bar

Note that previously recorded values are still accessible, even if value history is off. In rare cases, these references to past computations can cause Guile to use too much memory. One may clear these values, possibly enabling garbage collection, via the clear-value-history! procedure, described below.

The programmatic interface to value history is in a module:

(use-modules (ice-9 history))
[Scheme Procedure] value-history-enabled?

Return true if value history is enabled, or false otherwise.

[Scheme Procedure] enable-value-history!

Turn on value history, if it was off.

[Scheme Procedure] disable-value-history!

Turn off value history, if it was on.

[Scheme Procedure] clear-value-history!

Clear the value history. If the stored values are not captured by some other data structure or closure, they may then be reclaimed by the garbage collector.

4.4.4 REPL Commands

The REPL exists to read expressions, evaluate them, and then print their results. But sometimes one wants to tell the REPL to evaluate an expression in a different way, or to do something else altogether. A user can affect the way the REPL works with a REPL command.

The previous section had an example of a command, in the form of ,option.

scheme@(guile-user)> ,option value-history #t

Commands are distinguished from expressions by their initial comma (‘,’). Since a comma cannot begin an expression in most languages, it is an effective indicator to the REPL that the following text forms a command, not an expression.

REPL commands are convenient because they are always there. Even if the current module doesn’t have a binding for pretty-print, one can always ,pretty-print.

The following sections document the various commands, grouped together by functionality. Many of the commands have abbreviations; see the online help (,help) for more information.

4.4.4.1 Help Commands

When Guile starts interactively, it notifies the user that help can be had by typing ‘,help’. Indeed, help is a command, and a particularly useful one, as it allows the user to discover the rest of the commands.

[REPL Command] help [all | group | [-c] command]

Show help. With one argument, tries to look up the argument as a group name, giving help on that group if successful. Otherwise tries to look up the argument as a command, giving help on the command. If there is a command whose name is also a group name, use the ‘-c command’ form to give help on the command instead of the group. Without any argument, a list of help commands and command groups are displayed.

[REPL Command] show [topic]

Gives information about Guile. With one argument, tries to show a particular piece of information; currently supported topics are ‘warranty’ (or ‘w’), ‘copying’ (or ‘c’), and ‘version’ (or ‘v’). Without any argument, a list of topics is displayed.

[REPL Command] apropos regexp

Find bindings/modules/packages.

[REPL Command] describe obj

Show description/documentation.

4.4.4.2 Module Commands

[REPL Command] module [module]

Change modules / Show current module.

[REPL Command] import module …

Import modules / List those imported.

[REPL Command] load file

Load a file in the current module.

[REPL Command] reload [module]

Reload the given module, or the current module if none was given.

[REPL Command] binding

List current bindings.

[REPL Command] in module expression [REPL Command] in module command arg …

Evaluate an expression, or alternatively, execute another meta-command in the context of a module. For example, ‘,in (foo bar) ,binding’ will show the bindings in the module (foo bar).

4.4.4.3 Language Commands

[REPL Command] language language

Change languages.

4.4.4.4 Compile Commands

[REPL Command] compile exp

Generate compiled code.

[REPL Command] compile-file file

Compile a file.

[REPL Command] expand exp

Expand any macros in a form.

[REPL Command] optimize exp

Run the optimizer on a piece of code and print the result.

[REPL Command] disassemble exp

Disassemble a compiled procedure.

[REPL Command] disassemble-file file

Disassemble a file.

4.4.4.5 Profile Commands

[REPL Command] time exp

Time execution.

[REPL Command] profile exp

Profile execution.

[REPL Command] trace exp [#:width w] [#:max-indent i] Trace execution. By default, the trace will limit its width to the width of your terminal, or width if specified. Nested procedure invocations will be printed farther to the right, though if the width of the indentation passes the max-indent, the indentation is abbreviated.

4.4.4.6 Debug Commands

These debugging commands are only available within a recursive REPL; they do not work at the top level.

[REPL Command] backtrace [count] [#:width w] [#:full? f] Print a backtrace. Print a backtrace of all stack frames, or innermost count frames. If count is negative, the last count frames will be shown.

[REPL Command] up [count] Select a calling stack frame. Select and print stack frames that called this one. An argument says how many frames up to go.

[REPL Command] down [count]

Select a called stack frame. Select and print stack frames called by this one. An argument says how many frames down to go.

[REPL Command] frame [idx]

Show a frame. Show the selected frame. With an argument, select a frame by index, then show it.

[REPL Command] procedure

Print the procedure for the selected frame.

[REPL Command] locals

Show local variables. Show locally-bound variables in the selected frame.

[REPL Command] error-message

[REPL Command] error

Show error message. Display the message associated with the error that started the current debugging REPL.

[REPL Command] registers

Show the VM registers associated with the current frame. See Section 9.3.3 [Stack Layout], page 774, for more information on VM stack frames.

[REPL Command] width [cols]

Sets the number of display columns in the output of ,backtrace and ,locals to cols. If cols is not given, the width of the terminal is used.

The next 3 commands work at any REPL.

[REPL Command] break proc

Set a breakpoint at proc.

[REPL Command] break-at-source file line

Set a breakpoint at the given source location.

[REPL Command] tracepoint proc

Set a tracepoint on the given procedure. This will cause all calls to the procedure to print out a tracing message. See Section 6.25.4.4 [Tracing Traps], page 460, for more information.

The rest of the commands in this subsection all apply only when the stack is continuable — in other words when it makes sense for the program that the stack comes from to continue running. Usually this means that the program stopped because of a trap or a breakpoint.

[REPL Command] step

Tell the debugged program to step to the next source location.

[REPL Command] next

Tell the debugged program to step to the next source location in the same frame. (See Section 6.25.4 [Traps], page 455 for the details of how this works.)

[REPL Command] finish

Tell the program being debugged to continue running until the completion of the current stack frame, and at that time to print the result and reenter the REPL.

4.4.4.7 Inspect Commands

[REPL Command] inspect exp

Inspect the result(s) of evaluating exp.

[REPL Command] pretty-print exp

Pretty-print the result(s) of evaluating exp.

4.4.4.8 System Commands

[REPL Command] gc

Garbage collection.

[REPL Command] statistics

Display statistics.

[REPL Command] option [name] [exp]

With no arguments, lists all options. With one argument, shows the current value of the name option. With two arguments, sets the name option to the result of evaluating the Scheme expression exp.

[REPL Command] quit

Quit this session.

Current REPL options include:

compile-options

The options used when compiling expressions entered at the REPL. See Section 6.17.5 [Compilation], page 364, for more on compilation options. interp Whether to interpret or compile expressions given at the REPL, if such a choice

is available. Off by default (indicating compilation).

prompt A customized REPL prompt. #f by default, indicating the default prompt.

print A procedure of two arguments used to print the result of evaluating each expression. The arguments are the current REPL and the value to print. By default, #f, to use the default procedure.

value-history

Whether value history is on or not. See Section 4.4.3 [Value History], page 47.

on-error What to do when an error happens. By default, debug, meaning to enter the debugger. Other values include backtrace, to show a backtrace without entering the debugger, or report, to simply show a short error printout.

Default values for REPL options may be set using repl-default-option-set! from

(system repl common):

[Scheme Procedure] repl-default-option-set! key value

Set the default value of a REPL option. This function is particularly useful in a user’s init file. See Section 4.4.1 [Init File], page 46.

4.4.5 Error Handling

When code being evaluated from the REPL hits an error, Guile enters a new prompt, allowing you to inspect the context of the error.

     scheme@(guile-user)> (map string-append '("a" "b") '("c" #\d))
     ERROR: In procedure string-append:
     ERROR: Wrong type (expecting string): #\d
     Entering a new prompt.  Type `,bt' for a backtrace or `,q' to continue.
     scheme@(guile-user) [1]>

The new prompt runs inside the old one, in the dynamic context of the error. It is a recursive REPL, augmented with a reified representation of the stack, ready for debugging.

,backtrace (abbreviated ,bt) displays the Scheme call stack at the point where the error occurred:

     scheme@(guile-user) [1]> ,bt
                1 (map #<procedure string-append _> ("a" "b") ("c" #\d))
                0 (string-append "b" #\d)

In the above example, the backtrace doesn’t have much source information, as map and string-append are both primitives. But in the general case, the space on the left of the backtrace indicates the line and column in which a given procedure calls another.

You can exit a recursive REPL in the same way that you exit any REPL: via ‘(quit)’, ‘,quit’ (abbreviated ‘,q’), or C-d, among other options.

4.4.6 Interactive Debugging

A recursive debugging REPL exposes a number of other meta-commands that inspect the state of the computation at the time of the error. These commands allow you to

• display the Scheme call stack at the point where the error occurred; • move up and down the call stack, to see in detail the expression being evaluated, or

the procedure being applied, in each frame; and

• examine the values of variables and expressions in the context of each frame.

See Section 4.4.4.6 [Debug Commands], page 50, for documentation of the individual commands. This section aims to give more of a walkthrough of a typical debugging session.

First, we’re going to need a good error. Let’s try to macroexpand the expression (unquote foo), outside of a quasiquote form, and see how the macroexpander reports this error.

     scheme@(guile-user)> (macroexpand '(unquote foo))
     ERROR: In procedure macroexpand:
     ERROR: unquote: expression not valid outside of quasiquote in (unquote foo)
     Entering a new prompt.  Type `,bt' for a backtrace or `,q' to continue.
     scheme@(guile-user) [1]>

The backtrace command, which can also be invoked as bt, displays the call stack (aka backtrace) at the point where the debugger was entered:

     scheme@(guile-user) [1]> ,bt
     In ice-9/psyntax.scm:
       1130:21  3 (chi-top (unquote foo) () ((top)) e (eval) (hygiene #))
       1071:30  2 (syntax-type (unquote foo) () ((top)) #f #f (# #) #f)
       1368:28  1 (chi-macro #<procedure de9360 at ice-9/psyntax.scm...> ...)
     In unknown file:
                0 (scm-error syntax-error macroexpand "~a: ~a in ~a" # #f)

A call stack consists of a sequence of stack frames, with each frame describing one procedure which is waiting to do something with the values returned by another. Here we see that there are four frames on the stack.

Note that macroexpand is not on the stack – it must have made a tail call to chi-top, as indeed we would find if we searched ice-9/psyntax.scm for its definition.

When you enter the debugger, the innermost frame is selected, which means that the commands for getting information about the “current” frame, or for evaluating expressions in the context of the current frame, will do so by default with respect to the innermost frame. To select a different frame, so that these operations will apply to it instead, use the up, down and frame commands like this:

     scheme@(guile-user) [1]> ,up
     In ice-9/psyntax.scm:
       1368:28  1 (chi-macro #<procedure de9360 at ice-9/psyntax.scm...> ...)
     scheme@(guile-user) [1]> ,frame 3
     In ice-9/psyntax.scm:
       1130:21  3 (chi-top (unquote foo) () ((top)) e (eval) (hygiene #))
     scheme@(guile-user) [1]> ,down
     In ice-9/psyntax.scm:
       1071:30  2 (syntax-type (unquote foo) () ((top)) #f #f (# #) #f)

Perhaps we’re interested in what’s going on in frame 2, so we take a look at its local variables:

scheme@(guile-user) [1]> ,locals
Local variables:
$1 = e = (unquote foo)
$2 = r = ()
$3 = w = ((top))
$4 = s = #f
$5 = rib = #f
$6 = mod = (hygiene guile-user)
$7 = for-car? = #f
$8 = first = unquote
$9 = ftype = macro
$10 = fval = #<procedure de9360 at ice-9/psyntax.scm:2817:2 (x)>
$11 = fe = unquote
$12 = fw = ((top))
$13 = fs = #f
$14 = fmod = (hygiene guile-user)

All of the values are accessible by their value-history names ($n):

scheme@(guile-user) [1]> $10
$15 = #<procedure de9360 at ice-9/psyntax.scm:2817:2 (x)>

We can even invoke the procedure at the REPL directly:

scheme@(guile-user) [1]> ($10 'not-going-to-work)
ERROR: In procedure macroexpand:
ERROR: source expression failed to match any pattern in not-going-to-work
Entering a new prompt.  Type `,bt' for a backtrace or `,q' to continue.

Well at this point we’ve caused an error within an error. Let’s just quit back to the top level:

scheme@(guile-user) [2]> ,q
scheme@(guile-user) [1]> ,q
scheme@(guile-user)>

Finally, as a word to the wise: hackers close their REPL prompts with C-d.

4.5 Using Guile in Emacs

Any text editor can edit Scheme, but some are better than others. Emacs is the best, of course, and not just because it is a fine text editor. Emacs has good support for Scheme out of the box, with sensible indentation rules, parenthesis-matching, syntax highlighting, and even a set of keybindings for structural editing, allowing navigation, cut-and-paste, and transposition operations that work on balanced S-expressions.

As good as it is, though, two things will vastly improve your experience with Emacs and Guile.

The first is Taylor Campbell’s Paredit (http://www.emacswiki.org/emacs/ParEdit).

You should not code in any dialect of Lisp without Paredit. (They say that unopinionated writing is boring—hence this tone—but it’s the truth, regardless.) Paredit is the bee’s knees.

The second is José Antonio Ortega Ruiz’s Geiser (http://www.nongnu.org/geiser/). Geiser complements Emacs’ scheme-mode with tight integration to running Guile processes via a comint-mode REPL buffer.

Of course there are keybindings to switch to the REPL, and a good REPL environment, but Geiser goes beyond that, providing:

• Form evaluation in the context of the current file’s module.

• Macro expansion.

• File/module loading and/or compilation.

• Namespace-aware identifier completion (including local bindings, names visible in the current module, and module names).

• Autodoc: the echo area shows information about the signature of the procedure/macro around point automatically.

• Jump to definition of identifier at point.

• Access to documentation (including docstrings when the implementation provides it).

• Listings of identifiers exported by a given module.

• Listings of callers/callees of procedures.

• Rudimentary support for debugging and error navigation.

• Support for multiple, simultaneous REPLs.

See Geiser’s web page at http://www.nongnu.org/geiser/, for more information.

4.6 Using Guile Tools

Guile also comes with a growing number of command-line utilities: a compiler, a disassembler, some module inspectors, and in the future, a system to install Guile packages from the internet. These tools may be invoked using the guild program.

     $ guild compile -o foo.go foo.scm
     wrote `foo.go'

This program used to be called guile-tools up to Guile version 2.0.1, and for backward compatibility it still may be called as such. However we changed the name to guild, not only because it is pleasantly shorter and easier to read, but also because this tool will serve to bind Guile wizards together, by allowing hackers to share code with each other using a CPAN-like system.

See Section 6.17.5 [Compilation], page 364, for more on guild compile.

A complete list of guild scripts can be had by invoking guild list, or simply guild.

4.7 Installing Site Packages

At some point, you will probably want to share your code with other people. To do so effectively, it is important to follow a set of common conventions, to make it easy for the user to install and use your package.

The first thing to do is to install your Scheme files where Guile can find them. When Guile goes to find a Scheme file, it will search a load path to find the file: first in Guile’s own path, then in paths for site packages. A site package is any Scheme code that is installed and not part of Guile itself. See Section 6.17.7 [Load Paths], page 368, for more on load paths.

There are several site paths, for historical reasons, but the one that should generally be used can be obtained by invoking the %site-dir procedure. See Section 6.22.1 [Build Config], page 427. If Guile 2.0 is installed on your system in /usr/, then (%site-dir) will be /usr/share/guile/site/2.0. Scheme files should be installed there.

If you do not install compiled .go files, Guile will compile your modules and programs when they are first used, and cache them in the user’s home directory. See Section 6.17.5 [Compilation], page 364, for more on auto-compilation. However, it is better to compile the files before they are installed, and to just copy the files to a place that Guile can find them.

As with Scheme files, Guile searches a path to find compiled .go files, the %loadcompiled-path. By default, this path has two entries: a path for Guile’s files, and a path for site packages. You should install your .go files into the latter directory, whose value is returned by invoking the %site-ccache-dir procedure. As in the previous example, if Guile 2.0 is installed on your system in /usr/, then (%site-ccache-dir) site packages will

be /usr/lib/guile/2.0/site-ccache.

Note that a .go file will only be loaded in preference to a .scm file if it is newer. For that reason, you should install your Scheme files first, and your compiled files second. Load Paths, for more on the loading process.

Finally, although this section is only about Scheme, sometimes you need to install C extensions too. Shared libraries should be installed in the extensions dir. This value can be had from the build config (see Section 6.22.1 [Build Config], page 427). Again, if Guile 2.0 is installed on your system in /usr/, then the extensions dir will be

/usr/lib/guile/2.0/extensions.

5 Programming in C

This part of the manual explains the general concepts that you need to understand when interfacing to Guile from C. You will learn about how the latent typing of Scheme is embedded into the static typing of C, how the garbage collection of Guile is made available to C code, and how continuations influence the control flow in a C program.

This knowledge should make it straightforward to add new functions to Guile that can be called from Scheme. Adding new data types is also possible and is done by defining smobs.

The Section 5.7 [Programming Overview], page 83 section of this part contains general musings and guidelines about programming with Guile. It explores different ways to design a program around Guile, or how to embed Guile into existing programs.

For a pedagogical yet detailed explanation of how the data representation of Guile is implemented, See Section 9.2 [Data Representation], page 763. You don’t need to know the details given there to use Guile from C, but they are useful when you want to modify Guile itself or when you are just curious about how it is all done.

For detailed reference information on the variables, functions etc. that make up Guile’s application programming interface (API), See Chapter 6 [API Reference], page 99.

5.1 Parallel Installations

Guile provides strong API and ABI stability guarantees during stable series, so that if a user writes a program against Guile version 2.0.3, it will be compatible with some future version 2.0.7. We say in this case that 2.0 is the effective version, composed of the major and minor versions, in this case 2 and 0.

Users may install multiple effective versions of Guile, with each version’s headers, libraries, and Scheme files under their own directories. This provides the necessary stability guarantee for users, while also allowing Guile developers to evolve the language and its implementation.

However, parallel installability does have a down-side, in that users need to know which version of Guile to ask for, when they build against Guile. Guile solves this problem by installing a file to be read by the pkg-config utility, a tool to query installed packages by name. Guile encodes the version into its pkg-config name, so that users can ask for

guile-2.0 or guile-2.2, as appropriate.

For effective version 2.0, for example, you would invoke pkg-config --cflags --libs guile-2.0 to get the compilation and linking flags necessary to link to version 2.0 of Guile. You would typically run pkg-config during the configuration phase of your program and use the obtained information in the Makefile.

Guile’s pkg-config file, guile-2.0.pc, defines additional useful variables:

sitedir The default directory where Guile looks for Scheme source and compiled files (see Section 4.7 [Installing Site Packages], page 55). Run pkg-config guile-2.0 --variable=sitedir to see its value. See Section 5.8.2 [Autoconf Macros], page 94, for more on how to use it from Autoconf.

extensiondir

The default directory where Guile looks for extensions—i.e., shared libraries providing additional features (see Section 6.20.4 [Modules and Extensions],

          page 401). Run pkg-config guile-2.0 --variable=extensiondir to see its

value.

See the pkg-config man page, for more information, or its web site, http://pkg-config. freedesktop.org/. See Section 5.8 [Autoconf Support], page 94, for more on checking for Guile from within a configure.ac file.

5.2 Linking Programs With Guile

This section covers the mechanics of linking your program with Guile on a typical POSIX system.

The header file <libguile.h> provides declarations for all of Guile’s functions and constants. You should #include it at the head of any C source file that uses identifiers described in this manual. Once you’ve compiled your source files, you need to link them against the Guile object code library, libguile.

As noted in the previous section, <libguile.h> is not in the default search path for headers. The following command lines give respectively the C compilation and link flags needed to build programs using Guile 2.0:

     pkg-config guile-2.0 --cflags
     pkg-config guile-2.0 --libs

5.2.1 Guile Initialization Functions

To initialize Guile, you can use one of several functions. The first, scm_with_guile, is the most portable way to initialize Guile. It will initialize Guile when necessary and then call a function that you can specify. Multiple threads can call scm_with_guile concurrently and it can also be called more than once in a given thread. The global state of Guile will survive from one call of scm_with_guile to the next. Your function is called from within scm_with_guile since the garbage collector of Guile needs to know where the stack of each thread is.

A second function, scm_init_guile, initializes Guile for the current thread. When it returns, you can use the Guile API in the current thread. This function employs some non-portable magic to learn about stack bounds and might thus not be available on all platforms.

One common way to use Guile is to write a set of C functions which perform some useful task, make them callable from Scheme, and then link the program with Guile. This yields a Scheme interpreter just like guile, but augmented with extra functions for some specific application — a special-purpose scripting language.

In this situation, the application should probably process its command-line arguments in the same manner as the stock Guile interpreter. To make that straightforward, Guile

provides the scm_boot_guile and scm_shell function.

For more about these functions, see Section 6.4 [Initialization], page 101.

5.2.2 A Sample Guile Main Program

Here is simple-guile.c, source code for a main and an inner_main function that will produce a complete Guile interpreter.

/* simple-guile.c --- Start Guile from C.  */
#include <libguile.h>
static void
inner_main (void *closure, int argc, char **argv)
{
  /* preparation */
  scm_shell (argc, argv);
  /* after exit */
}
int
main (int argc, char **argv)
{
  scm_boot_guile (argc, argv, inner_main, 0);
  return 0; /* never reached, see inner_main */
}

The main function calls scm_boot_guile to initialize Guile, passing it inner_main. Once scm_boot_guile is ready, it invokes inner_main, which calls scm_shell to process the command-line arguments in the usual way.

5.2.3 Building the Example with Make

Here is a Makefile which you can use to compile the example program. It uses pkg-config to learn about the necessary compiler and linker flags.

# Use GCC, if you have it installed.
CC=gcc
# Tell the C compiler where to find <libguile.h>
CFLAGS=`pkg-config --cflags guile-2.0`
# Tell the linker what libraries to use and where to find them.
LIBS=`pkg-config --libs guile-2.0`
simple-guile: simple-guile.o
        ${CC} simple-guile.o ${LIBS} -o simple-guile
simple-guile.o: simple-guile.c
        ${CC} -c ${CFLAGS} simple-guile.c

5.2.4 Building the Example with Autoconf

If you are using the GNU Autoconf package to make your application more portable, Autoconf will settle many of the details in the Makefile automatically, making it much simpler and more portable; we recommend using Autoconf with Guile. Here is a configure.ac file for simple-guile that uses the standard PKG_CHECK_MODULES macro to check for Guile. Autoconf will process this file into a configure script. We recommend invoking Autoconf via the autoreconf utility.

AC_INIT(simple-guile.c)
# Find a C compiler.
AC_PROG_CC
# Check for Guile
PKG_CHECK_MODULES([GUILE], [guile-2.0])
# Generate a Makefile, based on the results.
AC_OUTPUT(Makefile)
Run autoreconf -vif to generate configure.

Here is a Makefile.in template, from which the configure script produces a Makefile customized for the host system:

# The configure script fills in these values.
CC=@CC@
CFLAGS=@GUILE_CFLAGS@
LIBS=@GUILE_LIBS@
simple-guile: simple-guile.o
        ${CC} simple-guile.o ${LIBS} -o simple-guile
simple-guile.o: simple-guile.c
        ${CC} -c ${CFLAGS} simple-guile.c

The developer should use Autoconf to generate the configure script from the configure.ac template, and distribute configure with the application. Here’s how a user might go about building the application:

$ ls
Makefile.in     configure*      configure.ac    simple-guile.c
$ ./configure
checking for gcc... ccache gcc
checking whether the C compiler works... yes
checking for C compiler default output file name... a.out
checking for suffix of executables...
checking whether we are cross compiling... no
checking for suffix of object files... o
checking whether we are using the GNU C compiler... yes
checking whether ccache gcc accepts -g... yes
checking for ccache gcc option to accept ISO C89... none needed
checking for pkg-config... /usr/bin/pkg-config
checking pkg-config is at least version 0.9.0... yes
checking for GUILE... yes
configure: creating ./config.status

     config.status: creating Makefile
     $ make
     [...]
     $ ./simple-guile
     guile> (+ 1 2 3)
     6
     guile> (getpwnam "jimb")
     #("jimb" "83Z7d75W2tyJQ" 4008 10 "Jim Blandy" "/u/jimb"
       "/usr/local/bin/bash")
     guile> (exit)
     $

5.3 Linking Guile with Libraries

The previous section has briefly explained how to write programs that make use of an embedded Guile interpreter. But sometimes, all you want to do is make new primitive procedures and data types available to the Scheme programmer. Writing a new version of guile is inconvenient in this case and it would in fact make the life of the users of your new features needlessly hard.

For example, suppose that there is a program guile-db that is a version of Guile with additional features for accessing a database. People who want to write Scheme programs that use these features would have to use guile-db instead of the usual guile program. Now suppose that there is also a program guile-gtk that extends Guile with access to the popular Gtk+ toolkit for graphical user interfaces. People who want to write GUIs in Scheme would have to use guile-gtk. Now, what happens when you want to write a Scheme application that uses a GUI to let the user access a database? You would have to write a third program that incorporates both the database stuff and the GUI stuff. This might not be easy (because guile-gtk might be a quite obscure program, say) and taking this example further makes it easy to see that this approach can not work in practice.

It would have been much better if both the database features and the GUI feature had been provided as libraries that can just be linked with guile. Guile makes it easy to do just this, and we encourage you to make your extensions to Guile available as libraries whenever possible.

You write the new primitive procedures and data types in the normal fashion, and link them into a shared library instead of into a stand-alone program. The shared library can then be loaded dynamically by Guile.

5.3.1 A Sample Guile Extension

This section explains how to make the Bessel functions of the C library available to Scheme. First we need to write the appropriate glue code to convert the arguments and return values of the functions from Scheme to C and back. Additionally, we need a function that will add them to the set of Guile primitives. Because this is just an example, we will only implement this for the j0 function.

Consider the following file bessel.c.

      #include <math.h>
      #include <libguile.h>

      SCM
      j0_wrapper (SCM x)
      {
        return scm_from_double (j0 (scm_to_double (x)));
      }
      void
      init_bessel ()
      {
        scm_c_define_gsubr ("j0", 1, 0, 0, j0_wrapper);
      }

This C source file needs to be compiled into a shared library. Here is how to do it on GNU/Linux:

      gcc `pkg-config --cflags guile-2.0` \
        -shared -o libguile-bessel.so -fPIC bessel.c

For creating shared libraries portably, we recommend the use of GNU Libtool (see Section “Introduction” in GNU Libtool).

A shared library can be loaded into a running Guile process with the function loadextension. In addition to the name of the library to load, this function also expects the name of a function from that library that will be called to initialize it. For our example, we are going to call the function init_bessel which will make j0_wrapper available to Scheme programs with the name j0. Note that we do not specify a filename extension such as .so when invoking load-extension. The right extension for the host platform will be provided automatically.

     (load-extension "libguile-bessel" "init_bessel")
     (j0 2)
     ⇒ 0.223890779141236

For this to work, load-extension must be able to find libguile-bessel, of course. It will look in the places that are usual for your operating system, and it will additionally look into the directories listed in the LTDL_LIBRARY_PATH environment variable.

To see how these Guile extensions via shared libraries relate to the module system, See Section 2.5.3 [Putting Extensions into Modules], page 12.

5.4 General concepts for using libguile

When you want to embed the Guile Scheme interpreter into your program or library, you need to link it against the libguile library (see Section 5.2 [Linking Programs With Guile], page 58). Once you have done this, your C code has access to a number of data types and functions that can be used to invoke the interpreter, or make new functions that you have written in C available to be called from Scheme code, among other things.

Scheme is different from C in a number of significant ways, and Guile tries to make the advantages of Scheme available to C as well. Thus, in addition to a Scheme interpreter, libguile also offers dynamic types, garbage collection, continuations, arithmetic on arbitrary sized numbers, and other things.

The two fundamental concepts are dynamic types and garbage collection. You need to understand how libguile offers them to C programs in order to use the rest of libguile. Also, the more general control flow of Scheme caused by continuations needs to be dealt with.

Running asynchronous signal handlers and multi-threading is known to C code already, but there are of course a few additional rules when using them together with libguile.

5.4.1 Dynamic Types

Scheme is a dynamically-typed language; this means that the system cannot, in general, determine the type of a given expression at compile time. Types only become apparent at run time. Variables do not have fixed types; a variable may hold a pair at one point, an integer at the next, and a thousand-element vector later. Instead, values, not variables, have fixed types.

In order to implement standard Scheme functions like pair? and string? and provide garbage collection, the representation of every value must contain enough information to accurately determine its type at run time. Often, Scheme systems also use this information to determine whether a program has attempted to apply an operation to an inappropriately typed value (such as taking the car of a string).

Because variables, pairs, and vectors may hold values of any type, Scheme implementations use a uniform representation for values — a single type large enough to hold either a complete value or a pointer to a complete value, along with the necessary typing information.

In Guile, this uniform representation of all Scheme values is the C type SCM. This is an opaque type and its size is typically equivalent to that of a pointer to void. Thus, SCM values can be passed around efficiently and they take up reasonably little storage on their own.

The most important rule is: You never access a SCM value directly; you only pass it to functions or macros defined in libguile.

As an obvious example, although a SCM variable can contain integers, you can of course not compute the sum of two SCM values by adding them with the C + operator. You must use the libguile function scm_sum.

Less obvious and therefore more important to keep in mind is that you also cannot directly test SCM values for trueness. In Scheme, the value #f is considered false and of course a SCM variable can represent that value. But there is no guarantee that the SCM representation of #f looks false to C code as well. You need to use scm_is_true or scm_ is_false to test a SCM value for trueness or falseness, respectively.

You also can not directly compare two SCM values to find out whether they are identical (that is, whether they are eq? in Scheme terms). You need to use scm_is_eq for this.

The one exception is that you can directly assign a SCM value to a SCM variable by using the C = operator.

The following (contrived) example shows how to do it right. It implements a function of two arguments (a and flag) that returns a+1 if flag is true, else it returns a unchanged.

     SCM
     my_incrementing_function (SCM a, SCM flag)
     {
       SCM result;
       if (scm_is_true (flag))

         result = scm_sum (a, scm_from_int (1));
       else
         result = a;
       return result;
     }

Often, you need to convert between SCM values and appropriate C values. For example, we needed to convert the integer 1 to its SCM representation in order to add it to a. Libguile provides many function to do these conversions, both from C to SCM and from SCM to C.

The conversion functions follow a common naming pattern: those that make a SCM value from a C value have names of the form scm_from_type (...) and those that convert a SCM value to a C value use the form scm_to_type (...).

However, it is best to avoid converting values when you can. When you must combine C values and SCM values in a computation, it is often better to convert the C values to SCM values and do the computation by using libguile functions than to the other way around (converting SCM to C and doing the computation some other way).

As a simple example, consider this version of my_incrementing_function from above:

     SCM
     my_other_incrementing_function (SCM a, SCM flag)
     {
       int result;
       if (scm_is_true (flag))
         result = scm_to_int (a) + 1;
       else
         result = scm_to_int (a);
       return scm_from_int (result);
     }

This version is much less general than the original one: it will only work for values A that can fit into a int. The original function will work for all values that Guile can represent and that scm_sum can understand, including integers bigger than long long, floating point numbers, complex numbers, and new numerical types that have been added to Guile by third-party libraries.

Also, computing with SCM is not necessarily inefficient. Small integers will be encoded directly in the SCM value, for example, and do not need any additional memory on the heap. See Section 9.2 [Data Representation], page 763 to find out the details.

Some special SCM values are available to C code without needing to convert them from C values:

Scheme value C representation

#f            SCM_BOOL_F
#t            SCM_BOOL_T
()            SCM_EOL

In addition to SCM, Guile also defines the related type scm_t_bits. This is an unsigned integral type of sufficient size to hold all information that is directly contained in a SCM value. The scm_t_bits type is used internally by Guile to do all the bit twiddling explained in Section 9.2 [Data Representation], page 763, but you will encounter it occasionally in low-level user code as well.

5.4.2 Garbage Collection

As explained above, the SCM type can represent all Scheme values. Some values fit entirely into a SCM value (such as small integers), but other values require additional storage in the heap (such as strings and vectors). This additional storage is managed automatically by Guile. You don’t need to explicitly deallocate it when a SCM value is no longer used.

Two things must be guaranteed so that Guile is able to manage the storage automatically: it must know about all blocks of memory that have ever been allocated for Scheme values, and it must know about all Scheme values that are still being used. Given this knowledge, Guile can periodically free all blocks that have been allocated but are not used by any active Scheme values. This activity is called garbage collection.

It is easy for Guile to remember all blocks of memory that it has allocated for use by Scheme values, but you need to help it with finding all Scheme values that are in use by C code.

You do this when writing a SMOB mark function, for example (see Section 5.5.4 [Garbage Collecting Smobs], page 75). By calling this function, the garbage collector learns about all references that your SMOB has to other SCM values.

Other references to SCM objects, such as global variables of type SCM or other random data structures in the heap that contain fields of type SCM, can be made visible to the garbage collector by calling the functions scm_gc_protect or scm_permanent_object. You normally use these functions for long lived objects such as a hash table that is stored in a global variable. For temporary references in local variables or function arguments, using these functions would be too expensive.

These references are handled differently: Local variables (and function arguments) of type SCM are automatically visible to the garbage collector. This works because the collector scans the stack for potential references to SCM objects and considers all referenced objects to be alive. The scanning considers each and every word of the stack, regardless of what it is actually used for, and then decides whether it could possibly be a reference to a SCM object. Thus, the scanning is guaranteed to find all actual references, but it might also find words that only accidentally look like references. These ‘false positives’ might keep SCM objects alive that would otherwise be considered dead. While this might waste memory, keeping an object around longer than it strictly needs to is harmless. This is why this technique is called “conservative garbage collection”. In practice, the wasted memory seems to be no problem.

The stack of every thread is scanned in this way and the registers of the CPU and all other memory locations where local variables or function parameters might show up are included in this scan as well.

The consequence of the conservative scanning is that you can just declare local variables and function parameters of type SCM and be sure that the garbage collector will not free the corresponding objects.

However, a local variable or function parameter is only protected as long as it is really on the stack (or in some register). As an optimization, the C compiler might reuse its location for some other value and the SCM object would no longer be protected. Normally, this leads to exactly the right behavior: the compiler will only overwrite a reference when it is no longer needed and thus the object becomes unprotected precisely when the reference disappears, just as wanted.

There are situations, however, where a SCM object needs to be around longer than its reference from a local variable or function parameter. This happens, for example, when you retrieve some pointer from a smob and work with that pointer directly. The reference to the SCM smob object might be dead after the pointer has been retrieved, but the pointer itself (and the memory pointed to) is still in use and thus the smob object must be protected. The compiler does not know about this connection and might overwrite the SCM reference too early.

To get around this problem, you can use scm_remember_upto_here_1 and its cousins. It will keep the compiler from overwriting the reference. For a typical example of its use, see Section 5.5.5 [Remembering During Operations], page 77.

5.4.3 Control Flow

Scheme has a more general view of program flow than C, both locally and non-locally.

Controlling the local flow of control involves things like gotos, loops, calling functions and returning from them. Non-local control flow refers to situations where the program jumps across one or more levels of function activations without using the normal call or return operations.

The primitive means of C for local control flow is the goto statement, together with if. Loops done with for, while or do could in principle be rewritten with just goto and if. In Scheme, the primitive means for local control flow is the function call (together with if). Thus, the repetition of some computation in a loop is ultimately implemented by a function that calls itself, that is, by recursion.

This approach is theoretically very powerful since it is easier to reason formally about recursion than about gotos. In C, using recursion exclusively would not be practical, though, since it would eat up the stack very quickly. In Scheme, however, it is practical: function calls that appear in a tail position do not use any additional stack space (see Section 3.3.2 [Tail Calls], page 24).

A function call is in a tail position when it is the last thing the calling function does. The value returned by the called function is immediately returned from the calling function. In the following example, the call to bar-1 is in a tail position, while the call to bar-2 is not. (The call to 1- in foo-2 is in a tail position, though.)

     (define (foo-1 x)
       (bar-1 (1- x)))
     (define (foo-2 x)
       (1- (bar-2 x)))

Thus, when you take care to recurse only in tail positions, the recursion will only use constant stack space and will be as good as a loop constructed from gotos.

Scheme offers a few syntactic abstractions (do and named let) that make writing loops slightly easier.

But only Scheme functions can call other functions in a tail position: C functions can not. This matters when you have, say, two functions that call each other recursively to form a common loop. The following (unrealistic) example shows how one might go about determining whether a non-negative integer n is even or odd.

     (define (my-even? n)
       (cond ((zero? n) #t)
             (else (my-odd? (1- n)))))
     (define (my-odd? n)
       (cond ((zero? n) #f)
             (else (my-even? (1- n)))))

Because the calls to my-even? and my-odd? are in tail positions, these two procedures can be applied to arbitrary large integers without overflowing the stack. (They will still take a lot of time, of course.)

However, when one or both of the two procedures would be rewritten in C, it could no longer call its companion in a tail position (since C does not have this concept). You might need to take this consideration into account when deciding which parts of your program to write in Scheme and which in C.

In addition to calling functions and returning from them, a Scheme program can also exit non-locally from a function so that the control flow returns directly to an outer level. This means that some functions might not return at all.

Even more, it is not only possible to jump to some outer level of control, a Scheme program can also jump back into the middle of a function that has already exited. This might cause some functions to return more than once.

In general, these non-local jumps are done by invoking continuations that have previously been captured using call-with-current-continuation. Guile also offers a slightly restricted set of functions, catch and throw, that can only be used for non-local exits. This restriction makes them more efficient. Error reporting (with the function error) is implemented by invoking throw, for example. The functions catch and throw belong to the topic of exceptions.

Since Scheme functions can call C functions and vice versa, C code can experience the more general control flow of Scheme as well. It is possible that a C function will not return at all, or will return more than once. While C does offer setjmp and longjmp for non-local exits, it is still an unusual thing for C code. In contrast, non-local exits are very common in Scheme, mostly to report errors.

You need to be prepared for the non-local jumps in the control flow whenever you use a function from libguile: it is best to assume that any libguile function might signal an error or run a pending signal handler (which in turn can do arbitrary things).

It is often necessary to take cleanup actions when the control leaves a function nonlocally. Also, when the control returns non-locally, some setup actions might be called for. For example, the Scheme function with-output-to-port needs to modify the global state so

that current-output-port returns the port passed to with-output-to-port. The global

output port needs to be reset to its previous value when with-output-to-port returns normally or when it is exited non-locally. Likewise, the port needs to be set again when control enters non-locally.

Scheme code can use the dynamic-wind function to arrange for the setting and resetting of the global state. C code can use the corresponding scm_internal_dynamic_wind function, or a scm_dynwind_begin/scm_dynwind_end pair together with suitable ’dynwind actions’ (see Section 6.13.10 [Dynamic Wind], page 309).

Instead of coping with non-local control flow, you can also prevent it by erecting a continuation barrier, See Section 6.13.12 [Continuation Barriers], page 315. The function scm_c_with_continuation_barrier, for example, is guaranteed to return exactly once.

5.4.4 Asynchronous Signals

You can not call libguile functions from handlers for POSIX signals, but you can register Scheme handlers for POSIX signals such as SIGINT. These handlers do not run during the actual signal delivery. Instead, they are run when the program (more precisely, the thread that the handler has been registered for) reaches the next safe point.

The libguile functions themselves have many such safe points. Consequently, you must be prepared for arbitrary actions anytime you call a libguile function. For example, even scm_cons can contain a safe point and when a signal handler is pending for your thread, calling scm_cons will run this handler and anything might happen, including a non-local exit although scm_cons would not ordinarily do such a thing on its own.

If you do not want to allow the running of asynchronous signal handlers, you can block them temporarily with scm_dynwind_block_asyncs, for example. See See Section 6.21.2.1 [System asyncs], page 411.

Since signal handling in Guile relies on safe points, you need to make sure that your functions do offer enough of them. Normally, calling libguile functions in the normal course of action is all that is needed. But when a thread might spent a long time in a code section that calls no libguile function, it is good to include explicit safe points. This can allow the user to interrupt your code with C-c, for example.

You can do this with the macro SCM_TICK. This macro is syntactically a statement. That is, you could use it like this:

     while (1)
       {
         SCM_TICK;
         do_some_work ();
       }

Frequent execution of a safe point is even more important in multi threaded programs, See Section 5.4.5 [Multi-Threading], page 68.

5.4.5 Multi-Threading

Guile can be used in multi-threaded programs just as well as in single-threaded ones.

Each thread that wants to use functions from libguile must put itself into guile mode and must then follow a few rules. If it doesn’t want to honor these rules in certain situations, a thread can temporarily leave guile mode (but can no longer use libguile functions during that time, of course).

   Threads enter guile mode by calling scm_with_guile, scm_boot_guile, or scm_init_

guile. As explained in the reference documentation for these functions, Guile will then learn about the stack bounds of the thread and can protect the SCM values that are stored in local variables. When a thread puts itself into guile mode for the first time, it gets a Scheme representation and is listed by all-threads, for example.

Threads in guile mode can block (e.g., do blocking I/O) without causing any problems^{1}; temporarily leaving guile mode with scm_without_guile before blocking slightly improves GC performance, though. For some common blocking operations, Guile provides convenience functions. For example, if you want to lock a pthread mutex while in guile mode,

you might want to use scm_pthread_mutex_lock which is just like pthread_mutex_lock

except that it leaves guile mode while blocking.

All libguile functions are (intended to be) robust in the face of multiple threads using them concurrently. This means that there is no risk of the internal data structures of libguile becoming corrupted in such a way that the process crashes.

A program might still produce nonsensical results, though. Taking hashtables as an example, Guile guarantees that you can use them from multiple threads concurrently and a hashtable will always remain a valid hashtable and Guile will not crash when you access it. It does not guarantee, however, that inserting into it concurrently from two threads will give useful results: only one insertion might actually happen, none might happen, or the table might in general be modified in a totally arbitrary manner. (It will still be a valid hashtable, but not the one that you might have expected.) Guile might also signal an error when it detects a harmful race condition.

Thus, you need to put in additional synchronizations when multiple threads want to use a single hashtable, or any other mutable Scheme object.

When writing C code for use with libguile, you should try to make it robust as well. An example that converts a list into a vector will help to illustrate. Here is a correct version:

     SCM
     my_list_to_vector (SCM list)
     {
       SCM vector = scm_make_vector (scm_length (list), SCM_UNDEFINED);
       size_t len, i;
       len = scm_c_vector_length (vector);
       i = 0;
       while (i < len && scm_is_pair (list))
         {
           scm_c_vector_set_x (vector, i, scm_car (list));
           list = scm_cdr (list);
           i++;
         }
       return vector;
     }

The first thing to note is that storing into a SCM location concurrently from multiple threads is guaranteed to be robust: you don’t know which value wins but it will in any case be a valid SCM value.

1 In Guile 1.8, a thread blocking in guile mode would prevent garbage collection to occur. Thus, threads

had to leave guile mode whenever they could block. This is no longer needed with Guile 2.0.

But there is no guarantee that the list referenced by list is not modified in another thread while the loop iterates over it. Thus, while copying its elements into the vector, the list might get longer or shorter. For this reason, the loop must check both that it doesn’t overrun the vector and that it doesn’t overrun the list. Otherwise, scm_c_vector_set_x would raise an error if the index is out of range, and scm_car and scm_cdr would raise an error if the value is not a pair.

It is safe to use scm_car and scm_cdr on the local variable list once it is known that the variable contains a pair. The contents of the pair might change spontaneously, but it will always stay a valid pair (and a local variable will of course not spontaneously point to a different Scheme object).

Likewise, a vector such as the one returned by scm_make_vector is guaranteed to always stay the same length so that it is safe to only use scm c vector length once and store the result. (In the example, vector is safe anyway since it is a fresh object that no other thread can possibly know about until it is returned from my_list_to_vector.)

Of course the behavior of my_list_to_vector is suboptimal when list does indeed get asynchronously lengthened or shortened in another thread. But it is robust: it will always return a valid vector. That vector might be shorter than expected, or its last elements might be unspecified, but it is a valid vector and if a program wants to rule out these cases, it must avoid modifying the list asynchronously.

Here is another version that is also correct:

     SCM
     my_pedantic_list_to_vector (SCM list)
     {
       SCM vector = scm_make_vector (scm_length (list), SCM_UNDEFINED);
       size_t len, i;
       len = scm_c_vector_length (vector);
       i = 0;
       while (i < len)
         {
           scm_c_vector_set_x (vector, i, scm_car (list));
           list = scm_cdr (list);
           i++;
         }
       return vector;
     }

This version relies on the error-checking behavior of scm_car and scm_cdr. When the list is shortened (that is, when list holds a non-pair), scm_car will throw an error. This might be preferable to just returning a half-initialized vector.

The API for accessing vectors and arrays of various kinds from C takes a slightly different approach to thread-robustness. In order to get at the raw memory that stores the elements of an array, you need to reserve that array as long as you need the raw memory. During the time an array is reserved, its elements can still spontaneously change their values, but the memory itself and other things like the size of the array are guaranteed to stay fixed.

Any operation that would change these parameters of an array that is currently reserved will signal an error. In order to avoid these errors, a program should of course put suitable synchronization mechanisms in place. As you can see, Guile itself is again only concerned about robustness, not about correctness: without proper synchronization, your program will likely not be correct, but the worst consequence is an error message.

Real thread-safety often requires that a critical section of code is executed in a certain restricted manner. A common requirement is that the code section is not entered a second time when it is already being executed. Locking a mutex while in that section ensures that no other thread will start executing it, blocking asyncs ensures that no asynchronous code enters the section again from the current thread, and the error checking of Guile mutexes guarantees that an error is signalled when the current thread accidentally reenters the critical section via recursive function calls.

Guile provides two mechanisms to support critical sections as outlined above. You

can either use the macros SCM_CRITICAL_SECTION_START and SCM_CRITICAL_SECTION_END

for very simple sections; or use a dynwind context together with a call to scm_dynwind_

critical_section.

The macros only work reliably for critical sections that are guaranteed to not cause a non-local exit. They also do not detect an accidental reentry by the current thread. Thus, you should probably only use them to delimit critical sections that do not contain calls to libguile functions or to other external functions that might do complicated things.

The function scm_dynwind_critical_section, on the other hand, will correctly deal with non-local exits because it requires a dynwind context. Also, by using a separate mutex for each critical section, it can detect accidental reentries.

5.5 Defining New Types (Smobs)

Smobs are Guile’s mechanism for adding new primitive types to the system. The term “smob” was coined by Aubrey Jaffer, who says it comes from “small object”, referring to the fact that they are quite limited in size: they can hold just one pointer to a larger memory block plus 16 extra bits.

To define a new smob type, the programmer provides Guile with some essential information about the type — how to print it, how to garbage collect it, and so on — and Guile allocates a fresh type tag for it. The programmer can then use scm_c_define_gsubr to make a set of C functions visible to Scheme code that create and operate on these objects.

(You can find a complete version of the example code used in this section in the Guile distribution, in doc/example-smob. That directory includes a makefile and a suitable main function, so you can build a complete interactive Guile shell, extended with the datatypes described here.)

5.5.1 Describing a New Type

To define a new type, the programmer must write two functions to manage instances of the type:

print Guile will apply this function to each instance of the new type to print the value, as for display or write. The default print function prints #<NAME ADDRESS> where NAME is the first argument passed to scm_make_smob_type. equalp If Scheme code asks the equal? function to compare two instances of the same

smob type, Guile calls this function. It should return SCM_BOOL_T if a and b should be considered equal?, or SCM_BOOL_F otherwise. If equalp is NULL, equal? will assume that two instances of this type are never equal? unless they are eq?.

When the only resource associated with a smob is memory managed by the garbage collector—i.e., memory allocated with the scm_gc_malloc functions—this is sufficient. However, when a smob is associated with other kinds of resources, it may be necessary to define one of the following functions, or both:

mark Guile will apply this function to each instance of the new type it encounters during garbage collection. This function is responsible for telling the collector about any other SCM values that the object has stored, and that are in memory regions not already scanned by the garbage collector. See Section 5.5.4 [Garbage Collecting Smobs], page 75, for more details.

free Guile will apply this function to each instance of the new type that is to be deallocated. The function should release all resources held by the object. This is analogous to the Java finalization method—it is invoked at an unspecified time (when garbage collection occurs) after the object is dead. See Section 5.5.4 [Garbage Collecting Smobs], page 75, for more details. This function operates while the heap is in an inconsistent state and must therefore be careful. See Section 6.8 [Smobs], page 241, for details about what this function is allowed to do.

To actually register the new smob type, call scm_make_smob_type. It returns a value of type scm_t_bits which identifies the new smob type.

The four special functions described above are registered by calling one of scm_set_

smob_mark, scm_set_smob_free, scm_set_smob_print, or scm_set_smob_equalp, as ap-

propriate. Each function is intended to be used at most once per type, and the call should be placed immediately following the call to scm_make_smob_type.

There can only be at most 256 different smob types in the system. Instead of registering a huge number of smob types (for example, one for each relevant C struct in your application), it is sometimes better to register just one and implement a second layer of type dispatching on top of it. This second layer might use the 16 extra bits to extend its type, for example.

Here is how one might declare and register a new type representing eight-bit gray-scale images:

     #include <libguile.h>
     struct image {
       int width, height;
       char *pixels;
       /* The name of this image */
       SCM name;
       /* A function to call when this image is

          modified, e.g., to update the screen,
          or SCM_BOOL_F if no action necessary */
       SCM update_func;
     };
     static scm_t_bits image_tag;
     void
     init_image_type (void)
     {
       image_tag = scm_make_smob_type ("image", sizeof (struct image));
       scm_set_smob_mark (image_tag, mark_image);
       scm_set_smob_free (image_tag, free_image);
       scm_set_smob_print (image_tag, print_image);
     }

5.5.2 Creating Smob Instances

Normally, smobs can have one immediate word of data. This word stores either a pointer to an additional memory block that holds the real data, or it might hold the data itself when it fits. The word is large enough for a SCM value, a pointer to void, or an integer that fits

into a size_t or ssize_t.

You can also create smobs that have two or three immediate words, and when these words suffice to store all data, it is more efficient to use these super-sized smobs instead of using a normal smob plus a memory block. See Section 5.5.6 [Double Smobs], page 78, for their discussion.

Guile provides functions for managing memory which are often helpful when implementing smobs. See Section 6.18.2 [Memory Blocks], page 376.

To retrieve the immediate word of a smob, you use the macro SCM_SMOB_DATA. It can be set with SCM_SET_SMOB_DATA. The 16 extra bits can be accessed with SCM_SMOB_FLAGS

and SCM_SET_SMOB_FLAGS.

The two macros SCM_SMOB_DATA and SCM_SET_SMOB_DATA treat the immediate word as if it were of type scm_t_bits, which is an unsigned integer type large enough to hold a pointer to void. Thus you can use these macros to store arbitrary pointers in the smob word.

When you want to store a SCM value directly in the immediate word of a smob, you

should use the macros SCM_SMOB_OBJECT and SCM_SET_SMOB_OBJECT to access it.

Creating a smob instance can be tricky when it consists of multiple steps that allocate resources. Most of the time, this is mainly about allocating memory to hold associated data structures. Using memory managed by the garbage collector simplifies things: the garbage collector will automatically scan those data structures for pointers, and reclaim them when they are no longer referenced.

Continuing the example from above, if the global variable image_tag contains a tag returned by scm_make_smob_type, here is how we could construct a smob whose immediate word contains a pointer to a freshly allocated struct image:

     SCM

make_image (SCM name, SCM s_width, SCM s_height)
{
SCM smob;
struct image *image;
int width = scm_to_int (s_width);
int height = scm_to_int (s_height);
/* Step 1: Allocate the memory block.
 */
image = (struct image *)
   scm_gc_malloc (sizeof (struct image), "image");
/* Step 2: Initialize it with straight code.
 */
image->width = width;
image->height = height;
image->pixels = NULL;
image->name = SCM_BOOL_F;
image->update_func = SCM_BOOL_F;
/* Step 3: Create the smob.
 */
smob = scm_new_smob (image_tag, image);
/* Step 4: Finish the initialization.
 */
image->name = name;
image->pixels =
  scm_gc_malloc_pointerless (width * height, "image pixels");
return smob;
}

We use scm_gc_malloc_pointerless for the pixel buffer to tell the garbage collector

not to scan it for pointers. Calls to scm_gc_malloc, scm_new_smob, and scm_gc_malloc_

pointerless raise an exception in out-of-memory conditions; the garbage collector is able to reclaim previously allocated memory if that happens.

5.5.3 Type checking

Functions that operate on smobs should check that the passed SCM value indeed is a suitable smob before accessing its data. They can do this with scm_assert_smob_type.

For example, here is a simple function that operates on an image smob, and checks the type of its argument.

SCM
clear_image (SCM image_smob)
{
int area;

       struct image *image;
       scm_assert_smob_type (image_tag, image_smob);
       image = (struct image *) SCM_SMOB_DATA (image_smob);
       area = image->width * image->height;
       memset (image->pixels, 0, area);
       /* Invoke the image's update function.
        */
       if (scm_is_true (image->update_func))
         scm_call_0 (image->update_func);
       scm_remember_upto_here_1 (image_smob);
       return SCM_UNSPECIFIED;
     }

See Section 5.5.5 [Remembering During Operations], page 77 for an explanation of the

call to scm_remember_upto_here_1.

5.5.4 Garbage Collecting Smobs

Once a smob has been released to the tender mercies of the Scheme system, it must be prepared to survive garbage collection. In the example above, all the memory associated with the smob is managed by the garbage collector because we used the scm_gc_allocation functions. Thus, no special care must be taken: the garbage collector automatically scans them and reclaims any unused memory.

However, when data associated with a smob is managed in some other way—e.g., malloc’d memory or file descriptors—it is possible to specify a free function to release those resources when the smob is reclaimed, and a mark function to mark Scheme objects otherwise invisible to the garbage collector.

As described in more detail elsewhere (see Section 9.2.4 [Conservative GC], page 767), every object in the Scheme system has a mark bit, which the garbage collector uses to tell live objects from dead ones. When collection starts, every object’s mark bit is clear. The collector traces pointers through the heap, starting from objects known to be live, and sets the mark bit on each object it encounters. When it can find no more unmarked objects, the collector walks all objects, live and dead, frees those whose mark bits are still clear, and clears the mark bit on the others.

The two main portions of the collection are called the mark phase, during which the collector marks live objects, and the sweep phase, during which the collector frees all unmarked objects.

The mark bit of a smob lives in a special memory region. When the collector encounters a smob, it sets the smob’s mark bit, and uses the smob’s type tag to find the appropriate mark function for that smob. It then calls this mark function, passing it the smob as its only argument.

The mark function is responsible for marking any other Scheme objects the smob refers to. If it does not do so, the objects’ mark bits will still be clear when the collector begins to sweep, and the collector will free them. If this occurs, it will probably break, or at least confuse, any code operating on the smob; the smob’s SCM values will have become dangling references.

To mark an arbitrary Scheme object, the mark function calls scm_gc_mark.

Thus, here is how we might write mark_image—again this is not needed in our example since we used the scm_gc_allocation routines, so this is just for the sake of illustration:

     SCM
     mark_image (SCM image_smob)
     {
       /* Mark the image's name and update function.  */
       struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);
       scm_gc_mark (image->name);
       scm_gc_mark (image->update_func);
       return SCM_BOOL_F;
     }

Note that, even though the image’s update_func could be an arbitrarily complex structure (representing a procedure and any values enclosed in its environment), scm_gc_mark will recurse as necessary to mark all its components. Because scm_gc_mark sets an object’s mark bit before it recurses, it is not confused by circular structures.

As an optimization, the collector will mark whatever value is returned by the mark function; this helps limit depth of recursion during the mark phase. Thus, the code above should really be written as:

     SCM
     mark_image (SCM image_smob)
     {
       /* Mark the image's name and update function.  */
       struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);
       scm_gc_mark (image->name);
       return image->update_func;
     }

Finally, when the collector encounters an unmarked smob during the sweep phase, it uses the smob’s tag to find the appropriate free function for the smob. It then calls that function, passing it the smob as its only argument.

The free function must release any resources used by the smob. However, it must not free objects managed by the collector; the collector will take care of them. For historical reasons, the return type of the free function should be size_t, an unsigned integral type; the free function should always return zero.

Here is how we might write the free_image function for the image smob type—again for the sake of illustration, since our example does not need it thanks to the use of the scm_gc_allocation routines:

     size_t
     free_image (SCM image_smob)

{
struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);
scm_gc_free (image->pixels,
             image->width * image->height,
             "image pixels");
scm_gc_free (image, sizeof (struct image), "image");
return 0;
}

During the sweep phase, the garbage collector will clear the mark bits on all live objects. The code which implements a smob need not do this itself.

There is no way for smob code to be notified when collection is complete. It is usually a good idea to minimize the amount of processing done during garbage

collection; keep the mark and free functions very simple. Since collections occur at unpredictable times, it is easy for any unusual activity to interfere with normal code.

5.5.5 Remembering During Operations

It’s important that a smob is visible to the garbage collector whenever its contents are being accessed. Otherwise it could be freed while code is still using it.

For example, consider a procedure to convert image data to a list of pixel values.

SCM
image_to_list (SCM image_smob)
{
struct image *image;
SCM lst;
int i;
scm_assert_smob_type (image_tag, image_smob);
image = (struct image *) SCM_SMOB_DATA (image_smob);
lst = SCM_EOL;
for (i = image->width * image->height - 1; i >= 0; i--)
  lst = scm_cons (scm_from_char (image->pixels[i]), lst);
scm_remember_upto_here_1 (image_smob);
return lst;
}

In the loop, only the image pointer is used and the C compiler has no reason to keep the image_smob value anywhere. If scm_cons results in a garbage collection, image_smob might not be on the stack or anywhere else and could be freed, leaving the loop accessing freed data. The use of scm_remember_upto_here_1 prevents this, by creating a reference to image_smob after all data accesses.

There’s no need to do the same for lst, since that’s the return value and the compiler will certainly keep it in a register or somewhere throughout the routine.

The clear_image example previously shown (see Section 5.5.3 [Type checking], page 74)

also used scm_remember_upto_here_1 for this reason.

It’s only in quite rare circumstances that a missing scm_remember_upto_here_1 will bite, but when it happens the consequences are serious. Fortunately the rule is simple: whenever calling a Guile library function or doing something that might, ensure that the SCM of a smob is referenced past all accesses to its insides. Do this by adding an scm_ remember_upto_here_1 if there are no other references.

In a multi-threaded program, the rule is the same. As far as a given thread is concerned, a garbage collection still only occurs within a Guile library function, not at an arbitrary time. (Guile waits for all threads to reach one of its library functions, and holds them there while the collector runs.)

5.5.6 Double Smobs

Smobs are called smob because they are small: they normally have only room for one void* or SCM value plus 16 bits. The reason for this is that smobs are directly implemented by using the low-level, two-word cells of Guile that are also used to implement pairs, for example. (see Section 9.2 [Data Representation], page 763 for the details.) One word of the two-word cells is used for SCM_SMOB_DATA (or SCM_SMOB_OBJECT), the other contains the 16-bit type tag and the 16 extra bits.

In addition to the fundamental two-word cells, Guile also has four-word cells, which are appropriately called double cells. You can use them for double smobs and get two more immediate words of type scm_t_bits.

A double smob is created with scm_new_double_smob. Its immediate words can be

retrieved as scm_t_bits with SCM_SMOB_DATA_2 and SCM_SMOB_DATA_3 in addition to SCM_

SMOB_DATA. Unsurprisingly, the words can be set to scm_t_bits values with SCM_SET_

SMOB_DATA_2 and SCM_SET_SMOB_DATA_3.
   Of course there are also SCM_SMOB_OBJECT_2, SCM_SMOB_OBJECT_3, SCM_SET_SMOB_
OBJECT_2, and SCM_SET_SMOB_OBJECT_3.

5.5.7 The Complete Example

Here is the complete text of the implementation of the image datatype, as presented in the sections above. We also provide a definition for the smob’s print function, and make some objects and functions static, to clarify exactly what the surrounding code is using.

As mentioned above, you can find this code in the Guile distribution, in doc/examplesmob. That directory includes a makefile and a suitable main function, so you can build a complete interactive Guile shell, extended with the datatypes described here.)

     /* file "image-type.c" */
     #include <stdlib.h>
     #include <libguile.h>
     static scm_t_bits image_tag;
     struct image {
       int width, height;

char *pixels;
/* The name of this image */
SCM name;
/* A function to call when this image is
   modified, e.g., to update the screen,
   or SCM_BOOL_F if no action necessary */
SCM update_func;
};
static SCM
make_image (SCM name, SCM s_width, SCM s_height)
{
SCM smob;
struct image *image;
int width = scm_to_int (s_width);
int height = scm_to_int (s_height);
/* Step 1: Allocate the memory block.
 */
image = (struct image *)
   scm_gc_malloc (sizeof (struct image), "image");
/* Step 2: Initialize it with straight code.
 */
image->width = width;
image->height = height;
image->pixels = NULL;
image->name = SCM_BOOL_F;
image->update_func = SCM_BOOL_F;
/* Step 3: Create the smob.
 */
smob = scm_new_smob (image_tag, image);
/* Step 4: Finish the initialization.
 */
image->name = name;
image->pixels =
   scm_gc_malloc (width * height, "image pixels");
return smob;
}
SCM
clear_image (SCM image_smob)

{
int area;
struct image *image;
scm_assert_smob_type (image_tag, image_smob);
image = (struct image *) SCM_SMOB_DATA (image_smob);
area = image->width * image->height;
memset (image->pixels, 0, area);
/* Invoke the image's update function.
 */
if (scm_is_true (image->update_func))
  scm_call_0 (image->update_func);
scm_remember_upto_here_1 (image_smob);
return SCM_UNSPECIFIED;
}
static SCM
mark_image (SCM image_smob)
{
/* Mark the image's name and update function.  */
struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);
scm_gc_mark (image->name);
return image->update_func;
}
static size_t
free_image (SCM image_smob)
{
struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);
scm_gc_free (image->pixels,
             image->width * image->height,
             "image pixels");
scm_gc_free (image, sizeof (struct image), "image");
return 0;
}
static int
print_image (SCM image_smob, SCM port, scm_print_state *pstate)
{
struct image *image = (struct image *) SCM_SMOB_DATA (image_smob);

  scm_puts ("#<image ", port);
  scm_display (image->name, port);
  scm_puts (">", port);
  /* non-zero means success */
  return 1;
}
void
init_image_type (void)
{
  image_tag = scm_make_smob_type ("image", sizeof (struct image));
  scm_set_smob_mark (image_tag, mark_image);
  scm_set_smob_free (image_tag, free_image);
  scm_set_smob_print (image_tag, print_image);
  scm_c_define_gsubr ("clear-image", 1, 0, 0, clear_image);
  scm_c_define_gsubr ("make-image", 3, 0, 0, make_image);
}

Here is a sample build and interaction with the code from the example-smob directory, on the author’s machine:

zwingli:example-smob$ make CC=gcc
gcc `pkg-config --cflags guile-2.0` -c image-type.c -o image-type.o
gcc `pkg-config --cflags guile-2.0` -c myguile.c -o myguile.o
gcc image-type.o myguile.o `pkg-config --libs guile-2.0` -o myguile
zwingli:example-smob$ ./myguile
guile> make-image
#<primitive-procedure make-image>
guile> (define i (make-image "Whistler's Mother" 100 100))
guile> i
#<image Whistler's Mother>
guile> (clear-image i)
guile> (clear-image 4)
ERROR: In procedure clear-image in expression (clear-image 4):
ERROR: Wrong type (expecting image): 4
ABORT: (wrong-type-arg)
Type "(backtrace)" to get more information.
guile>

5.6 Function Snarfing

When writing C code for use with Guile, you typically define a set of C functions, and then make some of them visible to the Scheme world by calling scm_c_define_gsubr or related functions. If you have many functions to publish, it can sometimes be annoying to keep the list of calls to scm_c_define_gsubr in sync with the list of function definitions.

Guile provides the guile-snarf program to manage this problem. Using this tool, you can keep all the information needed to define the function alongside the function definition itself; guile-snarf will extract this information from your source code, and automatically generate a file of calls to scm_c_define_gsubr which you can #include into an initialization function.

The snarfing mechanism works for many kind of initialization actions, not just for collecting calls to scm_c_define_gsubr. For a full list of what can be done, See Section 6.5 [Snarfing Macros], page 102.

The guile-snarf program is invoked like this:

      guile-snarf [-o outfile] [cpp-args ...]

This command will extract initialization actions to outfile. When no outfile has been specified or when outfile is -, standard output will be used. The C preprocessor is called with cpp-args (which usually include an input file) and the output is filtered to extract the initialization actions.

If there are errors during processing, outfile is deleted and the program exits with nonzero status.

During snarfing, the pre-processor macro SCM_MAGIC_SNARFER is defined. You could use this to avoid including snarfer output files that don’t yet exist by writing code like this:

      #ifndef SCM_MAGIC_SNARFER
      #include "foo.x"
      #endif

Here is how you might define the Scheme function clear-image, implemented by the C

function clear_image:
     #include <libguile.h>
     SCM_DEFINE (clear_image, "clear-image", 1, 0, 0,
                 (SCM image_smob),
                 "Clear the image.")
     {
       /* C code to clear the image in image_smob... */
     }
     void
     init_image_type ()
     {
     #include "image-type.x"
     }

The SCM_DEFINE declaration says that the C function clear_image implements a Scheme function called clear-image, which takes one required argument (of type SCM and named image_smob), no optional arguments, and no rest argument. The string "Clear the image." provides a short help text for the function, it is called a docstring.

SCM_DEFINE macro also defines a static array of characters initialized to the Scheme name of the function. In this case, s_clear_image is set to the C string, "clear-image". You might want to use this symbol when generating error messages.

Assuming the text above lives in a file named image-type.c, you will need to execute the following command to prepare this file for compilation:

     guile-snarf -o image-type.x image-type.c

This scans image-type.c for SCM_DEFINE declarations, and writes to image-type.x the output:

     scm_c_define_gsubr ("clear-image", 1, 0, 0, (SCM (*)() ) clear_image);

When compiled normally, SCM_DEFINE is a macro which expands to the function header

for clear_image.

Note that the output file name matches the #include from the input file. Also, you still need to provide all the same information you would if you were using scm_c_define_gsubr yourself, but you can place the information near the function definition itself, so it is less likely to become incorrect or out-of-date.

If you have many files that guile-snarf must process, you should consider using a fragment like the following in your Makefile:

     snarfcppopts = $(DEFS) $(INCLUDES) $(CPPFLAGS) $(CFLAGS)
     .SUFFIXES: .x
     .c.x:
     guile-snarf -o $@ $< $(snarfcppopts)

This tells make to run guile-snarf to produce each needed .x file from the corresponding .c file.

The program guile-snarf passes its command-line arguments directly to the C preprocessor, which it uses to extract the information it needs from the source code. this means you can pass normal compilation flags to guile-snarf to define preprocessor symbols, add header file directories, and so on.

5.7 An Overview of Guile Programming

Guile is designed as an extension language interpreter that is straightforward to integrate with applications written in C (and C++). The big win here for the application developer is that Guile integration, as the Guile web page says, “lowers your project’s hacktivation energy.” Lowering the hacktivation energy means that you, as the application developer, and your users, reap the benefits that flow from being able to extend the application in a high level extension language rather than in plain old C.

In abstract terms, it’s difficult to explain what this really means and what the integration process involves, so instead let’s begin by jumping straight into an example of how you might integrate Guile into an existing program, and what you could expect to gain by so doing. With that example under our belts, we’ll then return to a more general analysis of the arguments involved and the range of programming options available.

5.7.1 How One Might Extend Dia Using Guile

Dia is a free software program for drawing schematic diagrams like flow charts and floor plans (http://www.gnome.org/projects/dia/). This section conducts the thought experiment of adding Guile to Dia. In so doing, it aims to illustrate several of the steps and considerations involved in adding Guile to applications in general.

5.7.1.1 Deciding Why You Want to Add Guile

First off, you should understand why you want to add Guile to Dia at all, and that means forming a picture of what Dia does and how it does it. So, what are the constituents of the Dia application?

• Most importantly, the application domain objects — in other words, the concepts that differentiate Dia from another application such as a word processor or spreadsheet: shapes, templates, connectors, pages, plus the properties of all these things.

• The code that manages the graphical face of the application, including the layout and display of the objects above.

• The code that handles input events, which indicate that the application user is wanting to do something.

(In other words, a textbook example of the model - view - controller paradigm.)

Next question: how will Dia benefit once the Guile integration is complete? Several (positive!) answers are possible here, and the choice is obviously up to the application developers. Still, one answer is that the main benefit will be the ability to manipulate Dia’s application domain objects from Scheme.

Suppose that Dia made a set of procedures available in Scheme, representing the most basic operations on objects such as shapes, connectors, and so on. Using Scheme, the application user could then write code that builds upon these basic operations to create more complex procedures. For example, given basic procedures to enumerate the objects on a page, to determine whether an object is a square, and to change the fill pattern of a single shape, the user can write a Scheme procedure to change the fill pattern of all squares on the current page:

     (define (change-squares'-fill-pattern new-pattern)
       (for-each-shape current-page
         (lambda (shape)
           (if (square? shape)
               (change-fill-pattern shape new-pattern)))))

5.7.1.2 Four Steps Required to Add Guile

Assuming this objective, four steps are needed to achieve it.

First, you need a way of representing your application-specific objects — such as shape in the previous example — when they are passed into the Scheme world. Unless your objects are so simple that they map naturally into builtin Scheme data types like numbers and strings, you will probably want to use Guile’s SMOB interface to create a new Scheme data type for your objects.

Second, you need to write code for the basic operations like for-each-shape and square? such that they access and manipulate your existing data structures correctly, and then make these operations available as primitives on the Scheme level.

Third, you need to provide some mechanism within the Dia application that a user can hook into to cause arbitrary Scheme code to be evaluated.

Finally, you need to restructure your top-level application C code a little so that it initializes the Guile interpreter correctly and declares your SMOBs and primitives to the Scheme world.

The following subsections expand on these four points in turn.

5.7.1.3 How to Represent Dia Data in Scheme

For all but the most trivial applications, you will probably want to allow some representation of your domain objects to exist on the Scheme level. This is where the idea of SMOBs comes in, and with it issues of lifetime management and garbage collection.

To get more concrete about this, let’s look again at the example we gave earlier of how application users can use Guile to build higher-level functions from the primitives that Dia itself provides.

     (define (change-squares'-fill-pattern new-pattern)
       (for-each-shape current-page
         (lambda (shape)
           (if (square? shape)
               (change-fill-pattern shape new-pattern)))))

Consider what is stored here in the variable shape. For each shape on the current page, the for-each-shape primitive calls (lambda (shape) ...) with an argument representing that shape. Question is: how is that argument represented on the Scheme level? The issues are as follows.

• Whatever the representation, it has to be decodable again by the C code for the square? and change-fill-pattern primitives. In other words, a primitive like square? has somehow to be able to turn the value that it receives back into something that points to the underlying C structure describing a shape.

• The representation must also cope with Scheme code holding on to the value for later use. What happens if the Scheme code stores shape in a global variable, but then that shape is deleted (in a way that the Scheme code is not aware of), and later on some other Scheme code uses that global variable again in a call to, say, square??

• The lifetime and memory allocation of objects that exist only in the Scheme world is managed automatically by Guile’s garbage collector using one simple rule: when there are no remaining references to an object, the object is considered dead and so its memory is freed. But for objects that exist in both C and Scheme, the picture is more complicated; in the case of Dia, where the shape argument passes transiently in and out of the Scheme world, it would be quite wrong the delete the underlying C shape just because the Scheme code has finished evaluation. How do we avoid this happening?

One resolution of these issues is for the Scheme-level representation of a shape to be a new, Scheme-specific C structure wrapped up as a SMOB. The SMOB is what is passed into and out of Scheme code, and the Scheme-specific C structure inside the SMOB points to Dia’s underlying C structure so that the code for primitives like square? can get at it.

To cope with an underlying shape being deleted while Scheme code is still holding onto a Scheme shape value, the underlying C structure should have a new field that points to the Scheme-specific SMOB. When a shape is deleted, the relevant code chains through to the Scheme-specific structure and sets its pointer back to the underlying structure to NULL. Thus the SMOB value for the shape continues to exist, but any primitive code that tries to use it will detect that the underlying shape has been deleted because the underlying structure pointer is NULL.

So, to summarize the steps involved in this resolution of the problem (and assuming that the underlying C structure for a shape is struct dia_shape):

• Define a new Scheme-specific structure that points to the underlying C structure:

         struct dia_guile_shape
         {
           struct dia_shape * c_shape;   /* NULL => deleted */
         }
  • Add a field to struct dia_shape that points to its struct dia_guile_shape if it has

one —

         struct dia_shape
         {
           ...
           struct dia_guile_shape * guile_shape;
         }

— so that C code can set guile_shape->c_shape to NULL when the underlying shape is deleted.

  • Wrap struct dia_guile_shape as a SMOB type.

• Whenever you need to represent a C shape onto the Scheme level, create a SMOB instance for it, and pass that.

• In primitive code that receives a shape SMOB instance, check the c_shape field when decoding it, to find out whether the underlying C shape is still there.

As far as memory management is concerned, the SMOB values and their Scheme-specific structures are under the control of the garbage collector, whereas the underlying C structures are explicitly managed in exactly the same way that Dia managed them before we thought of adding Guile.

When the garbage collector decides to free a shape SMOB value, it calls the SMOB free function that was specified when defining the shape SMOB type. To maintain the correctness of the guile_shape field in the underlying C structure, this function should chain through to the underlying C structure (if it still exists) and set its guile_shape field to NULL.

For full documentation on defining and using SMOB types, see Section 5.5 [Defining New Types (Smobs)], page 71.

5.7.1.4 Writing Guile Primitives for Dia

Once the details of object representation are decided, writing the primitive function code that you need is usually straightforward.

A primitive is simply a C function whose arguments and return value are all of type SCM, and whose body does whatever you want it to do. As an example, here is a possible implementation of the square? primitive:

     static SCM square_p (SCM shape)
     {
       struct dia_guile_shape * guile_shape;
       /* Check that arg is really a shape SMOB. */

       scm_assert_smob_type (shape_tag, shape);
       /* Access Scheme-specific shape structure. */
       guile_shape = SCM_SMOB_DATA (shape);
       /* Find out if underlying shape exists and is a
          square; return answer as a Scheme boolean. */
       return scm_from_bool (guile_shape->c_shape &&
                             (guile_shape->c_shape->type == DIA_SQUARE));
     }

Notice how easy it is to chain through from the SCM shape parameter that square_p receives — which is a SMOB — to the Scheme-specific structure inside the SMOB, and thence to the underlying C structure for the shape.

   In this code, scm_assert_smob_type, SCM_SMOB_DATA, and scm_from_bool are from the

standard Guile API. We assume that shape_tag was given to us when we made the shape

SMOB type, using scm_make_smob_type. The call to scm_assert_smob_type ensures that

shape is indeed a shape. This is needed to guard against Scheme code using the square? procedure incorrectly, as in (square? "hello"); Scheme’s latent typing means that usage errors like this must be caught at run time.

Having written the C code for your primitives, you need to make them available as Scheme procedures by calling the scm_c_define_gsubr function. scm_c_define_gsubr (see Section 6.9.2 [Primitive Procedures], page 245) takes arguments that specify the Scheme-level name for the primitive and how many required, optional and rest arguments it can accept. The square? primitive always requires exactly one argument, so the call to make it available in Scheme reads like this:

     scm_c_define_gsubr ("square?", 1, 0, 0, square_p);

For where to put this call, see the subsection after next on the structure of Guile-enabled code (see Section 5.7.1.6 [Dia Structure], page 87).

5.7.1.5 Providing a Hook for the Evaluation of Scheme Code

To make the Guile integration useful, you have to design some kind of hook into your application that application users can use to cause their Scheme code to be evaluated.

Technically, this is straightforward; you just have to decide on a mechanism that is appropriate for your application. Think of Emacs, for example: when you type ESC :, you get a prompt where you can type in any Elisp code, which Emacs will then evaluate. Or, again like Emacs, you could provide a mechanism (such as an init file) to allow Scheme code to be associated with a particular key sequence, and evaluate the code when that key sequence is entered.

In either case, once you have the Scheme code that you want to evaluate, as a null terminated string, you can tell Guile to evaluate it by calling the scm_c_eval_string function.

5.7.1.6 Top-level Structure of Guile-enabled Dia

Let’s assume that the pre-Guile Dia code looks structurally like this:

  • main ()

• do lots of initialization and setup stuff

• enter Gtk main loop

When you add Guile to a program, one (rather technical) requirement is that Guile’s garbage collector needs to know where the bottom of the C stack is. The easiest way to ensure this is to use scm_boot_guile like this:

  • main ()

• do lots of initialization and setup stuff

     • scm_boot_guile (argc, argv, inner_main, NULL)
  • inner_main ()

• define all SMOB types

• export primitives to Scheme using scm_c_define_gsubr

• enter Gtk main loop

In other words, you move the guts of what was previously in your main function into a new function called inner_main, and then add a scm_boot_guile call, with inner_main as a parameter, to the end of main.

Assuming that you are using SMOBs and have written primitive code as described in the preceding subsections, you also need to insert calls to declare your new SMOBs and export the primitives to Scheme. These declarations must happen inside the dynamic scope of the scm_boot_guile call, but also before any code is run that could possibly use them — the beginning of inner_main is an ideal place for this.

5.7.1.7 Going Further with Dia and Guile

The steps described so far implement an initial Guile integration that already gives a lot of additional power to Dia application users. But there are further steps that you could take, and it’s interesting to consider a few of these.

In general, you could progressively move more of Dia’s source code from C into Scheme. This might make the code more maintainable and extensible, and it could open the door to new programming paradigms that are tricky to effect in C but straightforward in Scheme.

A specific example of this is that you could use the guile-gtk package, which provides Scheme-level procedures for most of the Gtk+ library, to move the code that lays out and displays Dia objects from C to Scheme.

As you follow this path, it naturally becomes less useful to maintain a distinction between Dia’s original non-Guile-related source code, and its later code implementing SMOBs and primitives for the Scheme world.

For example, suppose that the original source code had a dia_change_fill_pattern function:

     void dia_change_fill_pattern (struct dia_shape * shape,
                                   struct dia_pattern * pattern)
     {
       /* real pattern change work */
     }

During initial Guile integration, you add a change_fill_pattern primitive for Scheme purposes, which accesses the underlying structures from its SMOB values and uses dia_

change_fill_pattern to do the real work:

     SCM change_fill_pattern (SCM shape, SCM pattern)
     {
       struct dia_shape * d_shape;
       struct dia_pattern * d_pattern;
       ...
       dia_change_fill_pattern (d_shape, d_pattern);
       return SCM_UNSPECIFIED;
     }
   At this point, it makes sense to keep dia_change_fill_pattern and change_fill_

pattern separate, because dia_change_fill_pattern can also be called without going through Scheme at all, say because the user clicks a button which causes a C-registered Gtk+ callback to be called.

But, if the code for creating buttons and registering their callbacks is moved into Scheme (using guile-gtk), it may become true that dia_change_fill_pattern can no longer be called other than through Scheme. In which case, it makes sense to abolish it and move its contents directly into change_fill_pattern, like this:

     SCM change_fill_pattern (SCM shape, SCM pattern)
     {
       struct dia_shape * d_shape;
       struct dia_pattern * d_pattern;
       ...
       /* real pattern change work */
       return SCM_UNSPECIFIED;
     }

So further Guile integration progressively reduces the amount of functional C code that you have to maintain over the long term.

A similar argument applies to data representation. In the discussion of SMOBs earlier, issues arose because of the different memory management and lifetime models that normally apply to data structures in C and in Scheme. However, with further Guile integration, you can resolve this issue in a more radical way by allowing all your data structures to be under the control of the garbage collector, and kept alive by references from the Scheme world. Instead of maintaining an array or linked list of shapes in C, you would instead maintain a list in Scheme.

   Rather like the coalescing of dia_change_fill_pattern and change_fill_pattern,

the practical upshot of such a change is that you would no longer have to keep the dia_ shape and dia_guile_shape structures separate, and so wouldn’t need to worry about the pointers between them. Instead, you could change the SMOB definition to wrap the dia_shape structure directly, and send dia_guile_shape off to the scrap yard. Cut out the middle man!

Finally, we come to the holy grail of Guile’s free software / extension language approach. Once you have a Scheme representation for interesting Dia data types like shapes, and a handy bunch of primitives for manipulating them, it suddenly becomes clear that you have a bundle of functionality that could have far-ranging use beyond Dia itself. In other words, the data types and primitives could now become a library, and Dia becomes just one of the many possible applications using that library — albeit, at this early stage, a rather important one!

In this model, Guile becomes just the glue that binds everything together. Imagine an application that usefully combined functionality from Dia, Gnumeric and GnuCash — it’s tricky right now, because no such application yet exists; but it’ll happen some day …

5.7.2 Why Scheme is More Hackable Than C

Underlying Guile’s value proposition is the assumption that programming in a high level language, specifically Guile’s implementation of Scheme, is necessarily better in some way than programming in C. What do we mean by this claim, and how can we be so sure?

One class of advantages applies not only to Scheme, but more generally to any interpretable, high level, scripting language, such as Emacs Lisp, Python, Ruby, or TEX’s macro language. Common features of all such languages, when compared to C, are that:

• They lend themselves to rapid and experimental development cycles, owing usually to a combination of their interpretability and the integrated development environment in which they are used.

• They free developers from some of the low level bookkeeping tasks associated with C programming, notably memory management.

• They provide high level features such as container objects and exception handling that make common programming tasks easier.

In the case of Scheme, particular features that make programming easier — and more fun!

— are its powerful mechanisms for abstracting parts of programs (closures — see Section 3.4 [About Closure], page 26) and for iteration (see Section 6.13.4 [while do], page 294).

The evidence in support of this argument is empirical: the huge amount of code that has been written in extension languages for applications that support this mechanism. Most notable are extensions written in Emacs Lisp for GNU Emacs, in TEX’s macro language for TEX, and in Script-Fu for the Gimp, but there is increasingly now a significant code eco-system for Guile-based applications as well, such as Lilypond and GnuCash. It is close to inconceivable that similar amounts of functionality could have been added to these applications just by writing new code in their base implementation languages.

5.7.3 Example: Using Guile for an Application Testbed

As an example of what this means in practice, imagine writing a testbed for an application that is tested by submitting various requests (via a C interface) and validating the output received. Suppose further that the application keeps an idea of its current state, and that the “correct” output for a given request may depend on the current application state. A complete “white box”^{2} test plan for this application would aim to submit all possible requests in each distinguishable state, and validate the output for all request/state combinations.

2 A white box test plan is one that incorporates knowledge of the internal design of the application under

test.

To write all this test code in C would be very tedious. Suppose instead that the testbed code adds a single new C function, to submit an arbitrary request and return the response, and then uses Guile to export this function as a Scheme procedure. The rest of the testbed can then be written in Scheme, and so benefits from all the advantages of programming in Scheme that were described in the previous section.

(In this particular example, there is an additional benefit of writing most of the testbed in Scheme. A common problem for white box testing is that mistakes and mistaken assumptions in the application under test can easily be reproduced in the testbed code. It is more difficult to copy mistakes like this when the testbed is written in a different language from the application.)

5.7.4 A Choice of Programming Options

The preceding arguments and example point to a model of Guile programming that is applicable in many cases. According to this model, Guile programming involves a balance between C and Scheme programming, with the aim being to extract the greatest possible Scheme level benefit from the least amount of C level work.

The C level work required in this model usually consists of packaging and exporting functions and application objects such that they can be seen and manipulated on the Scheme level. To help with this, Guile’s C language interface includes utility features that aim to make this kind of integration very easy for the application developer. These features are documented later in this part of the manual: see REFFIXME.

This model, though, is really just one of a range of possible programming options. If all of the functionality that you need is available from Scheme, you could choose instead to write your whole application in Scheme (or one of the other high level languages that Guile supports through translation), and simply use Guile as an interpreter for Scheme. (In the future, we hope that Guile will also be able to compile Scheme code, so lessening the performance gap between C and Scheme code.) Or, at the other end of the C–Scheme scale, you could write the majority of your application in C, and only call out to Guile occasionally for specific actions such as reading a configuration file or executing a userspecified extension. The choices boil down to two basic questions:

• Which parts of the application do you write in C, and which in Scheme (or another high level translated language)?

• How do you design the interface between the C and Scheme parts of your application?

These are of course design questions, and the right design for any given application will always depend upon the particular requirements that you are trying to meet. In the context of Guile, however, there are some generally applicable considerations that can help you when designing your answers.

5.7.4.1 What Functionality is Already Available?

Suppose, for the sake of argument, that you would prefer to write your whole application in Scheme. Then the API available to you consists of:

• standard Scheme • plus the extensions to standard Scheme provided by Guile in its core distribution

• plus any additional functionality that you or others have packaged so that it can be loaded as a Guile Scheme module.

A module in the last category can either be a pure Scheme module — in other words a collection of utility procedures coded in Scheme — or a module that provides a Scheme interface to an extension library coded in C — in other words a nice package where someone else has done the work of wrapping up some useful C code for you. The set of available modules is growing quickly and already includes such useful examples as (gtk gtk), which makes Gtk+ drawing functions available in Scheme, and (database postgres), which provides SQL access to a Postgres database.

Given the growing collection of pre-existing modules, it is quite feasible that your application could be implemented by combining a selection of these modules together with new application code written in Scheme.

If this approach is not enough, because the functionality that your application needs is not already available in this form, and it is impossible to write the new functionality in Scheme, you will need to write some C code. If the required function is already available in C (e.g. in a library), all you need is a little glue to connect it to the world of Guile. If not, you need both to write the basic code and to plumb it into Guile.

In either case, two general considerations are important. Firstly, what is the interface by which the functionality is presented to the Scheme world? Does the interface consist only of function calls (for example, a simple drawing interface), or does it need to include objects of some kind that can be passed between C and Scheme and manipulated by both worlds. Secondly, how does the lifetime and memory management of objects in the C code relate to the garbage collection governed approach of Scheme objects? In the case where the basic C code is not already written, most of the difficulties of memory management can be avoided by using Guile’s C interface features from the start.

For the full documentation on writing C code for Guile and connecting existing C code to the Guile world, see REFFIXME.

5.7.4.2 Functional and Performance Constraints

5.7.4.3 Your Preferred Programming Style

5.7.4.4 What Controls Program Execution?

5.7.5 How About Application Users?

So far we have considered what Guile programming means for an application developer. But what if you are instead using an existing Guile-based application, and want to know what your options are for programming and extending this application?

The answer to this question varies from one application to another, because the options available depend inevitably on whether the application developer has provided any hooks for you to hang your own code on and, if there are such hooks, what they allow you to do.^{3} For example… • If the application permits you to load and execute any Guile code, the world is your

3 Of course, in the world of free software, you always have the freedom to modify the application’s source

code to your own requirements. Here we are concerned with the extension options that the application

has provided for without your needing to modify its source code.

oyster. You can extend the application in any way that you choose.

• A more cautious application might allow you to load and execute Guile code, but only in a safe environment, where the interface available is restricted by the application from the standard Guile API.

• Or a really fearful application might not provide a hook to really execute user code at all, but just use Scheme syntax as a convenient way for users to specify application data or configuration options.

In the last two cases, what you can do is, by definition, restricted by the application, and you should refer to the application’s own manual to find out your options.

The most well known example of the first case is Emacs, with its extension language Emacs Lisp: as well as being a text editor, Emacs supports the loading and execution of arbitrary Emacs Lisp code. The result of such openness has been dramatic: Emacs now benefits from user-contributed Emacs Lisp libraries that extend the basic editing function to do everything from reading news to psychoanalysis and playing adventure games. The only limitation is that extensions are restricted to the functionality provided by Emacs’s built-in set of primitive operations. For example, you can interact and display data by manipulating the contents of an Emacs buffer, but you can’t pop-up and draw a window with a layout that is totally different to the Emacs standard.

This situation with a Guile application that supports the loading of arbitrary user code is similar, except perhaps even more so, because Guile also supports the loading of extension libraries written in C. This last point enables user code to add new primitive operations to Guile, and so to bypass the limitation present in Emacs Lisp.

At this point, the distinction between an application developer and an application user becomes rather blurred. Instead of seeing yourself as a user extending an application, you could equally well say that you are developing a new application of your own using some of the primitive functionality provided by the original application. As such, all the discussions of the preceding sections of this chapter are relevant to how you can proceed with developing your extension.

5.8 Autoconf Support

Autoconf, a part of the GNU build system, makes it easy for users to build your package. This section documents Guile’s Autoconf support.

5.8.1 Autoconf Background

As explained in the GNU Autoconf Manual, any package needs configuration at build-time (see Section “Introduction” in The GNU Autoconf Manual). If your package uses Guile (or uses a package that in turn uses Guile), you probably need to know what specific Guile features are available and details about them.

The way to do this is to write feature tests and arrange for their execution by the configure script, typically by adding the tests to configure.ac, and running autoconf to create configure. Users of your package then run configure in the normal way.

Macros are a way to make common feature tests easy to express. Autoconf provides a wide range of macros (see Section “Existing Tests” in The GNU Autoconf Manual), and Guile installation provides Guile-specific tests in the areas of: program detection, compilation flags reporting, and Scheme module checks.

5.8.2 Autoconf Macros

As mentioned earlier in this chapter, Guile supports parallel installation, and uses pkgconfig to let the user choose which version of Guile they are interested in. pkg-config has its own set of Autoconf macros that are probably installed on most every development system. The most useful of these macros is PKG_CHECK_MODULES.

     PKG_CHECK_MODULES([GUILE], [guile-2.0])

This example looks for Guile and sets the GUILE_CFLAGS and GUILE_LIBS variables accordingly, or prints an error and exits if Guile was not found.

Guile comes with additional Autoconf macros providing more information, installed as

prefix/share/aclocal/guile.m4. Their names all begin with GUILE_.

[Autoconf Macro] GUILE_PKG [VERSIONS]

This macro runs the pkg-config tool to find development files for an available version of Guile. By default, this macro will search for the latest stable version of Guile (e.g. 2.0), falling back to the previous stable version (e.g. 1.8) if it is available. If no guile-VERSION.pc file is found, an error is signalled. The found version is stored in GUILE EFFECTIVE VERSION. If GUILE_PROGS was already invoked, this macro ensures that the development files have the same effective version as the Guile program. GUILE EFFECTIVE VERSION is marked for substitution, as by AC_SUBST.

[Autoconf Macro] GUILE_FLAGS

This macro runs the pkg-config tool to find out how to compile and link programs against Guile. It sets four variables: GUILE CFLAGS, GUILE LDFLAGS, GUILE LIBS, and GUILE LTLIBS. GUILE CFLAGS: flags to pass to a C or C++ compiler to build code that uses Guile header files. This is almost always just one or more -I flags.

GUILE LDFLAGS: flags to pass to the compiler to link a program against Guile. This includes -lguile-VERSION for the Guile library itself, and may also include one or more -L flag to tell the compiler where to find the libraries. But it does not include flags that influence the program’s runtime search path for libraries, and will therefore lead to a program that fails to start, unless all necessary libraries are installed in a standard location such as /usr/lib. GUILE LIBS and GUILE LTLIBS: flags to pass to the compiler or to libtool, respectively, to link a program against Guile. It includes flags that augment the program’s runtime search path for libraries, so that shared libraries will be found at the location where they were during linking, even in non-standard locations. GUILE LIBS is to be used when linking the program directly with the compiler, whereas GUILE LTLIBS is to be used when linking the program is done through libtool. The variables are marked for substitution, as by AC_SUBST.

[Autoconf Macro] GUILE_SITE_DIR

This looks for Guile’s "site" directory, usually something like PRE-FIX/share/guile/site, and sets var GUILE SITE to the path. Note that the var name is different from the macro name. The variable is marked for substitution, as by AC_SUBST.

[Autoconf Macro] GUILE_PROGS [VERSION]

This macro looks for programs guile and guild, setting variables GUILE and GUILD to their paths, respectively. If guile is not found, signal an error. By default, this macro will search for the latest stable version of Guile (e.g. 2.0). x.y or x.y.z versions can be specified. If an older version is found, the macro will signal an error. The effective version of the found guile is set to GUILE EFFECTIVE VERSION. This macro ensures that the effective version is compatible with the result of a previous invocation of GUILE_FLAGS, if any. As a legacy interface, it also looks for guile-config and guile-tools, setting GUILE CONFIG and GUILE TOOLS. The variables are marked for substitution, as by AC_SUBST.

[Autoconf Macro] GUILE_CHECK_RETVAL var check

var is a shell variable name to be set to the return value. check is a Guile Scheme expression, evaluated with "$GUILE -c", and returning either 0 or non-#f to indicate the check passed. Non-0 number or #f indicates failure. Avoid using the character "#" since that confuses autoconf.

[Autoconf Macro] GUILE_MODULE_CHECK var module featuretest description var is a shell variable name to be set to "yes" or "no". module is a list of symbols, like: (ice-9 common-list). featuretest is an expression acceptable to GUILE CHECK, q.v. description is a present-tense verb phrase (passed to AC MSG CHECKING).

[Autoconf Macro] GUILE_MODULE_AVAILABLE var module

var is a shell variable name to be set to "yes" or "no". module is a list of symbols, like: (ice-9 common-list).

[Autoconf Macro] GUILE_MODULE_REQUIRED symlist

symlist is a list of symbols, WITHOUT surrounding parens, like: ice-9 common-list.

[Autoconf Macro] GUILE_MODULE_EXPORTS var module modvar

var is a shell variable to be set to "yes" or "no". module is a list of symbols, like: (ice-9 common-list). modvar is the Guile Scheme variable to check.

[Autoconf Macro] GUILE_MODULE_REQUIRED_EXPORT module modvar

module is a list of symbols, like: (ice-9 common-list). modvar is the Guile Scheme variable to check.

5.8.3 Using Autoconf Macros

Using the autoconf macros is straightforward: Add the macro "calls" (actually instantiations) to configure.ac, run aclocal, and finally, run autoconf. If your system doesn’t have guile.m4 installed, place the desired macro definitions (AC_DEFUN forms) in acinclude.m4, and aclocal will do the right thing.

Some of the macros can be used inside normal shell constructs: if foo ; then GUILE_ BAZ ; fi, but this is not guaranteed. It’s probably a good idea to instantiate macros at top-level.

We now include two examples, one simple and one complicated.

The first example is for a package that uses libguile, and thus needs to know how to compile and link against it. So we use PKG_CHECK_MODULES to set the vars GUILE_CFLAGS and GUILE_LIBS, which are automatically substituted in the Makefile.

     In configure.ac:
       PKG_CHECK_MODULES([GUILE], [guile-2.0])
     In Makefile.in:
       GUILE_CFLAGS  = @GUILE_CFLAGS@
       GUILE_LIBS = @GUILE_LIBS@
       myprog.o: myprog.c
               $(CC) -o $ $(GUILE_CFLAGS) $<
       myprog: myprog.o
               $(CC) -o $ $< $(GUILE_LIBS)

The second example is for a package of Guile Scheme modules that uses an external program and other Guile Scheme modules (some might call this a "pure scheme" package). So we use the GUILE_SITE_DIR macro, a regular AC_PATH_PROG macro, and the GUILE_

MODULE_AVAILABLE macro.
     In configure.ac:
       GUILE_SITE_DIR
       probably_wont_work=""

# pgtype pgtable
GUILE_MODULE_AVAILABLE(have_guile_pg, (database postgres))
test $have_guile_pg = no &&
probably_wont_work="(my pgtype) (my pgtable) $probably_wont_work"
# gpgutils
AC_PATH_PROG(GNUPG,gpg)
test x"$GNUPG" = x &&
probably_wont_work="(my gpgutils) $probably_wont_work"
if test ! "$probably_wont_work" = "" ; then
p="         ***"
echo
echo "$p"
echo "$p NOTE:"
echo "$p The following modules probably won't work:"
echo "$p   $probably_wont_work"
echo "$p They can be installed anyway, and will work if their"
echo "$p dependencies are installed later.  Please see README."
echo "$p"
echo
fi
In Makefile.in:
instdir = @GUILE_SITE@/my
install:
  $(INSTALL) my/*.scm $(instdir)

6 API Reference

Guile provides an application programming interface (API) to developers in two core languages: Scheme and C. This part of the manual contains reference documentation for all of the functionality that is available through both Scheme and C interfaces.

6.1 Overview of the Guile API

Guile’s application programming interface (API) makes functionality available that an application developer can use in either C or Scheme programming. The interface consists of elements that may be macros, functions or variables in C, and procedures, variables, syntax or other types of object in Scheme.

Many elements are available to both Scheme and C, in a form that is appropriate. For example, the assq Scheme procedure is also available as scm_assq to C code. These elements are documented only once, addressing both the Scheme and C aspects of them.

The Scheme name of an element is related to its C name in a regular way. Also, a C function takes its parameters in a systematic way.

Normally, the name of a C function can be derived given its Scheme name, using some simple textual transformations:

• Replace - (hyphen) with _(underscore).

• Replace ? (question mark) with _p.

• Replace ! (exclamation point) with _x.

• Replace internal -> with _to_.

• Replace <= (less than or equal) with _leq.

• Replace >= (greater than or equal) with _geq.

• Replace < (less than) with _less.

• Replace > (greater than) with _gr.

• Prefix with scm_.

A C function always takes a fixed number of arguments of type SCM, even when the corresponding Scheme function takes a variable number.

For some Scheme functions, some last arguments are optional; the corresponding C function must always be invoked with all optional arguments specified. To get the effect as if an argument has not been specified, pass SCM_UNDEFINED as its value. You can not do this for an argument in the middle; when one argument is SCM_UNDEFINED all the ones following it must be SCM_UNDEFINED as well.

Some Scheme functions take an arbitrary number of rest arguments; the corresponding C function must be invoked with a list of all these arguments. This list is always the last argument of the C function.

These two variants can also be combined.

The type of the return value of a C function that corresponds to a Scheme function is always SCM. In the descriptions below, types are therefore often omitted but for the return value and for the arguments.

6.2 Deprecation

From time to time functions and other features of Guile become obsolete. Guile’s deprecation is a mechanism that can help you cope with this.

When you use a feature that is deprecated, you will likely get a warning message at runtime. Also, if you have a new enough toolchain, using a deprecated function from libguile will cause a link-time warning.

The primary source for information about just what interfaces are deprecated in a given release is the file NEWS. That file also documents what you should use instead of the obsoleted things.

The file README contains instructions on how to control the inclusion or removal of the deprecated features from the public API of Guile, and how to control the deprecation warning messages.

The idea behind this mechanism is that normally all deprecated interfaces are available, but you get feedback when compiling and running code that uses them, so that you can migrate to the newer APIs at your leisure.

6.3 The SCM Type

Guile represents all Scheme values with the single C type SCM. For an introduction to this topic, See Section 5.4.1 [Dynamic Types], page 63.

[C Type] SCM

SCM is the user level abstract C type that is used to represent all of Guile’s Scheme objects, no matter what the Scheme object type is. No C operation except assignment is guaranteed to work with variables of type SCM, so you should only use macros and functions to work with SCM values. Values are converted between C data types and the SCM type with utility functions and macros.

[C Type] scm_t_bits

scm_t_bits is an unsigned integral data type that is guaranteed to be large enough to hold all information that is required to represent any Scheme object. While this data type is mostly used to implement Guile’s internals, the use of this type is also necessary to write certain kinds of extensions to Guile.

[C Type] scm_t_signed_bits

This is a signed integral type of the same size as scm_t_bits.

[C Macro] scm_t_bits SCM_UNPACK (SCM x)

Transforms the SCM value x into its representation as an integral type. Only after applying SCM_UNPACK it is possible to access the bits and contents of the SCM value.

[C Macro] SCM SCM_PACK (scm t bits x)

Takes a valid integral representation of a Scheme object and transforms it into its representation as a SCM value.

6.4 Initializing Guile

Each thread that wants to use functions from the Guile API needs to put itself into guile mode with either scm_with_guile or scm_init_guile. The global state of Guile is initialized automatically when the first thread enters guile mode.

When a thread wants to block outside of a Guile API function, it should leave guile mode temporarily with scm_without_guile, See Section 6.21.5 [Blocking], page 418.

Threads that are created by call-with-new-thread or scm_spawn_thread start out in

guile mode so you don’t need to initialize them.

[C Function] void * scm_with_guile (void *(*func)(void *), void *data)

Call func, passing it data and return what func returns. While func is running, the current thread is in guile mode and can thus use the Guile API. When scm_with_guile is called from guile mode, the thread remains in guile mode

when scm_with_guile returns.

Otherwise, it puts the current thread into guile mode and, if needed, gives it a Scheme representation that is contained in the list returned by all-threads, for example. This Scheme representation is not removed when scm_with_guile returns so that a given thread is always represented by the same Scheme value during its lifetime, if at all. When this is the first thread that enters guile mode, the global state of Guile is initialized before calling func.

The function func is called via scm_with_continuation_barrier; thus, scm_with_

guile returns exactly once. When scm_with_guile returns, the thread is no longer in guile mode (except when scm_with_guile was called from guile mode, see above). Thus, only func can store SCM variables on the stack and be sure that they are protected from the garbage collector. See scm_init_guile for another approach at initializing Guile that does not have this restriction. It is OK to call scm_with_guile while a thread has temporarily left guile mode via scm_without_guile. It will then simply temporarily enter guile mode again.

[C Function] void scm_init_guile ()

Arrange things so that all of the code in the current thread executes as if from within a call to scm_with_guile. That is, all functions called by the current thread can assume that SCM values on their stack frames are protected from the garbage collector (except when the thread has explicitly left guile mode, of course). When scm_init_guile is called from a thread that already has been in guile mode once, nothing happens. This behavior matters when you call scm_init_guile while the thread has only temporarily left guile mode: in that case the thread will not be in guile mode after scm_init_guile returns. Thus, you should not use scm_init_guile in such a scenario. When a uncaught throw happens in a thread that has been put into guile mode via scm_init_guile, a short message is printed to the current error port and the thread is exited via scm_pthread_exit (NULL). No restrictions are placed on continuations.

The function scm_init_guile might not be available on all platforms since it requires some stack-bounds-finding magic that might not have been ported to all platforms that Guile runs on. Thus, if you can, it is better to use scm_with_guile or its variation scm_boot_guile instead of this function.

[C Function] void scm_boot_guile (int argc, char **argv, void (*main_func)

(void *data, int argc, char **argv), void *data) Enter guile mode as with scm_with_guile and call main func, passing it data, argc, and argv as indicated. When main func returns, scm_boot_guile calls exit (0); scm_boot_guile never returns. If you want some other exit value, have main func call exit itself. If you don’t want to exit at all, use scm_with_guile instead of

scm_boot_guile.

The function scm_boot_guile arranges for the Scheme command-line function to return the strings given by argc and argv. If main func modifies argc or argv, it should call scm_set_program_arguments with the final list, so Scheme code will know which arguments have been processed (see Section 7.2.6 [Runtime Environment], page 488).

[C Function] void scm_shell (int argc, char **argv)

Process command-line arguments in the manner of the guile executable. This includes loading the normal Guile initialization files, interacting with the user or running any scripts or expressions specified by -s or -e options, and then exiting. See Section 4.2 [Invoking Guile], page 35, for more details. Since this function does not return, you must do all application-specific initialization before calling this function.

6.5 Snarfing Macros

The following macros do two different things: when compiled normally, they expand in one way; when processed during snarfing, they cause the guile-snarf program to pick up some initialization code, See Section 5.6 [Function Snarfing], page 81.

The descriptions below use the term ‘normally’ to refer to the case when the code is compiled normally, and ‘while snarfing’ when the code is processed by guile-snarf.

[C Macro] SCM_SNARF_INIT (code)

Normally, SCM_SNARF_INIT expands to nothing; while snarfing, it causes code to be included in the initialization action file, followed by a semicolon. This is the fundamental macro for snarfing initialization actions. The more specialized macros below use it internally.

[C Macro] SCM_DEFINE (c name, scheme name, req, opt, var, arglist, docstring)

Normally, this macro expands into

      static const char s_c_name[] = scheme_name;
      SCM
      c_name arglist

While snarfing, it causes

      scm_c_define_gsubr (s_c_name, req, opt, var,
                          c_name);

to be added to the initialization actions. Thus, you can use it to declare a C function named c name that will be made available to Scheme with the name scheme name.

Note that the arglist argument must have parentheses around it.

[C Macro] SCM_SYMBOL (c name, scheme name) [C Macro] SCM_GLOBAL_SYMBOL (c name, scheme name)

Normally, these macros expand into

      static SCM c_name

or

      SCM c_name

respectively. While snarfing, they both expand into the initialization code

      c_name = scm_permanent_object (scm_from_locale_symbol (scheme_name));

Thus, you can use them declare a static or global variable of type SCM that will be initialized to the symbol named scheme name.

[C Macro] SCM_KEYWORD (c name, scheme name)

[C Macro] SCM_GLOBAL_KEYWORD (c name, scheme name)

Normally, these macros expand into

      static SCM c_name

or

      SCM c_name

respectively. While snarfing, they both expand into the initialization code

      c_name = scm_permanent_object (scm_c_make_keyword (scheme_name));

Thus, you can use them declare a static or global variable of type SCM that will be initialized to the keyword named scheme name.

[C Macro] SCM_VARIABLE (c name, scheme name)

[C Macro] SCM_GLOBAL_VARIABLE (c name, scheme name)

These macros are equivalent to SCM_VARIABLE_INIT and SCM_GLOBAL_VARIABLE_

INIT, respectively, with a value of SCM_BOOL_F.

[C Macro] SCM_VARIABLE_INIT (c name, scheme name, value)

[C Macro] SCM_GLOBAL_VARIABLE_INIT (c name, scheme name, value)

Normally, these macros expand into

      static SCM c_name

or

      SCM c_name

respectively. While snarfing, they both expand into the initialization code

      c_name = scm_permanent_object (scm_c_define (scheme_name, value));

Thus, you can use them declare a static or global C variable of type SCM that will be initialized to the object representing the Scheme variable named scheme name in the current module. The variable will be defined when it doesn’t already exist. It is always set to value.

6.6 Simple Generic Data Types

This chapter describes those of Guile’s simple data types which are primarily used for their role as items of generic data. By simple we mean data types that are not primarily used as containers to hold other data — i.e. pairs, lists, vectors and so on. For the documentation of such compound data types, see Section 6.7 [Compound Data Types], page 185.

6.6.1 Booleans

The two boolean values are #t for true and #f for false. They can also be written as #true and #false, as per R7RS.

Boolean values are returned by predicate procedures, such as the general equality predicates eq?, eqv? and equal? (see Section 6.11.1 [Equality], page 276) and numerical and string comparison operators like string=? (see Section 6.6.5.7 [String Comparison], page 148) and <= (see Section 6.6.2.8 [Comparison], page 117).

     (<= 3 8)
     ⇒ #t
     (<= 3 -3)
     ⇒ #f
     (equal? "house" "houses")
     ⇒ #f
     (eq? #f #f)

⇒

     #t

In test condition contexts like if and cond (see Section 6.13.2 [Conditionals], page 292), where a group of subexpressions will be evaluated only if a condition expression evaluates to “true”, “true” means any value at all except #f.

     (if #t "yes" "no")
     ⇒ "yes"
     (if 0 "yes" "no")
     ⇒ "yes"
     (if #f "yes" "no")
     ⇒ "no"

A result of this asymmetry is that typical Scheme source code more often uses #f explicitly than #t: #f is necessary to represent an if or cond false value, whereas #t is not necessary to represent an if or cond true value.

It is important to note that #f is not equivalent to any other Scheme value. In particular, #f is not the same as the number 0 (like in C and C++), and not the same as the “empty list” (like in some Lisp dialects).

In C, the two Scheme boolean values are available as the two constants SCM_BOOL_T for #t and SCM_BOOL_F for #f. Care must be taken with the false value SCM_BOOL_F: it is not false when used in C conditionals. In order to test for it, use scm_is_false or scm_is_true.

[Scheme Procedure] not x [C Function] scm_not (x)

Return #t if x is #f, else return #f.

[Scheme Procedure] boolean? obj [C Function] scm_boolean_p (obj)

Return #t if obj is either #t or #f, else return #f.

[C Macro] SCM SCM_BOOL_T

The SCM representation of the Scheme object #t.

[C Macro] SCM SCM_BOOL_F

The SCM representation of the Scheme object #f.

[C Function] int scm_is_true (SCM obj)

Return 0 if obj is #f, else return 1.

[C Function] int scm_is_false (SCM obj)

Return 1 if obj is #f, else return 0.

[C Function] int scm_is_bool (SCM obj)

Return 1 if obj is either #t or #f, else return 0.

[C Function] SCM scm_from_bool (int val)

Return #f if val is 0, else return #t.

[C Function] int scm_to_bool (SCM val)

Return 1 if val is SCM_BOOL_T, return 0 when val is SCM_BOOL_F, else signal a ‘wrong type’ error. You should probably use scm_is_true instead of this function when you just want to test a SCM value for trueness.

6.6.2 Numerical data types

Guile supports a rich “tower” of numerical types — integer, rational, real and complex — and provides an extensive set of mathematical and scientific functions for operating on numerical data. This section of the manual documents those types and functions.

You may also find it illuminating to read R5RS’s presentation of numbers in Scheme, which is particularly clear and accessible: see Section “Numbers” in R5RS.

6.6.2.1 Scheme’s Numerical “Tower”

Scheme’s numerical “tower” consists of the following categories of numbers:

integers Whole numbers, positive or negative; e.g. –5, 0, 18.

rationals The set of numbers that can be expressed as p/q where p and q are integers; e.g. 9/16 works, but pi (an irrational number) doesn’t. These include integers (n/1).

real numbers The set of numbers that describes all possible positions along a one-dimensional line. This includes rationals as well as irrational numbers. complex numbers

The set of numbers that describes all possible positions in a two dimensional space. This includes real as well as imaginary numbers (a + bi, where a is the real part, b is the imaginary part, and i is the square root of −1.)

It is called a tower because each category “sits on” the one that follows it, in the sense that every integer is also a rational, every rational is also real, and every real number is also a complex number (but with zero imaginary part).

In addition to the classification into integers, rationals, reals and complex numbers, Scheme also distinguishes between whether a number is represented exactly or not. For example, the result of 2 sin(π/4) is exactly \sqrt{2}, but Guile can represent neither π/4 nor \sqrt{2} exactly. Instead, it stores an inexact approximation, using the C type double.

Guile can represent exact rationals of any magnitude, inexact rationals that fit into a C double, and inexact complex numbers with double real and imaginary parts.

The number? predicate may be applied to any Scheme value to discover whether the value is any of the supported numerical types.

[Scheme Procedure] number? obj [C Function] scm_number_p (obj)

Return #t if obj is any kind of number, else #f.

For example:

     (number? 3)
     ⇒ #t
     (number? "hello there!")
     ⇒ #f
     (define pi 3.141592654)
     (number? pi)
     ⇒ #t

[C Function] int scm_is_number (SCM obj)

     This is equivalent to scm_is_true (scm_number_p (obj)).

The next few subsections document each of Guile’s numerical data types in detail.

6.6.2.2 Integers

Integers are whole numbers, that is numbers with no fractional part, such as 2, 83, and −3789.

Integers in Guile can be arbitrarily big, as shown by the following example.

     (define (factorial n)
       (let loop ((n n) (product 1))
         (if (= n 0)
             product
             (loop (- n 1) (* product n)))))

     (factorial 3)
     ⇒ 6
     (factorial 20)
     ⇒ 2432902008176640000
     (- (factorial 45))
     ⇒ -119622220865480194561963161495657715064383733760000000000

Readers whose background is in programming languages where integers are limited by the need to fit into just 4 or 8 bytes of memory may find this surprising, or suspect that Guile’s representation of integers is inefficient. In fact, Guile achieves a near optimal balance of convenience and efficiency by using the host computer’s native representation of integers where possible, and a more general representation where the required number does not fit in the native form. Conversion between these two representations is automatic and completely invisible to the Scheme level programmer.

C has a host of different integer types, and Guile offers a host of functions to convert between them and the SCM representation. For example, a C int can be handled with scm_to_int and scm_from_int. Guile also defines a few C integer types of its own, to help with differences between systems.

C integer types that are not covered can be handled with the generic scm_

to_signed_integer and scm_from_signed_integer for signed types, or with
scm_to_unsigned_integer and scm_from_unsigned_integer for unsigned types.

Scheme integers can be exact and inexact. For example, a number written as 3.0 with an explicit decimal-point is inexact, but it is also an integer. The functions integer? and scm_is_integer report true for such a number, but the functions exact-integer?, scm_

is_exact_integer, scm_is_signed_integer, and scm_is_unsigned_integer only allow

exact integers and thus report false. Likewise, the conversion functions like scm_to_signed_ integer only accept exact integers.

The motivation for this behavior is that the inexactness of a number should not be lost silently. If you want to allow inexact integers, you can explicitly insert a call to inexact->exact or to its C equivalent scm_inexact_to_exact. (Only inexact integers will be converted by this call into exact integers; inexact non-integers will become exact fractions.)

[Scheme Procedure] integer? x [C Function] scm_integer_p (x)

Return #t if x is an exact or inexact integer number, else return #f.

          (integer? 487)
          ⇒ #t
          (integer? 3.0)
          ⇒ #t
          (integer? -3.4)
          ⇒ #f

     (integer? +inf.0)
     ⇒ #f
[C Function] int scm_is_integer (SCM x)
This is equivalent to scm_is_true (scm_integer_p (x)).

[Scheme Procedure] exact-integer? x

[C Function] scm_exact_integer_p (x)

Return #t if x is an exact integer number, else return #f.

     (exact-integer? 37)
     ⇒ #t
     (exact-integer? 3.0)
     ⇒ #f
[C Function] int scm_is_exact_integer (SCM x)
This is equivalent to scm_is_true (scm_exact_integer_p (x)).
[C Type] scm_t_int8
[C Type] scm_t_uint8
[C Type] scm_t_int16
[C Type] scm_t_uint16
[C Type] scm_t_int32
[C Type] scm_t_uint32
[C Type] scm_t_int64
[C Type] scm_t_uint64
[C Type] scm_t_intmax
[C Type] scm_t_uintmax

The C types are equivalent to the corresponding ISO C types but are defined on all platforms, with the exception of scm_t_int64 and scm_t_uint64, which are only defined when a 64-bit type is available. For example, scm_t_int8 is equivalent to

int8_t.

You can regard these definitions as a stop-gap measure until all platforms provide these types. If you know that all the platforms that you are interested in already provide these types, it is better to use them directly instead of the types provided by Guile.

[C Function] int scm_is_signed_integer (SCM x, scm t intmax min,

scm t intmax max)

[C Function] int scm_is_unsigned_integer (SCM x, scm t uintmax min,

scm t uintmax max) Return 1 when x represents an exact integer that is between min and max, inclusive.

These functions can be used to check whether a SCM value will fit into a given range, such as the range of a given C integer type. If you just want to convert a SCM value to a given C integer type, use one of the conversion functions directly.

[C Function] scm_t_intmax scm_to_signed_integer (SCM x, scm t intmax min,

scm t intmax max)

[C Function] scm_t_uintmax scm_to_unsigned_integer (SCM x, scm t uintmax

min, scm t uintmax max) When x represents an exact integer that is between min and max inclusive, return that integer. Else signal an error, either a ‘wrong-type’ error when x is not an exact integer, or an ‘out-of-range’ error when it doesn’t fit the given range.

[C Function] SCM scm_from_signed_integer (scm t intmax x)
[C Function] SCM scm_from_unsigned_integer (scm t uintmax x)

Return the SCM value that represents the integer x. This function will always succeed and will always return an exact number.

[C Function] char scm_to_char (SCM x)

[C Function] signed char scm_to_schar (SCM x)
[C Function] unsigned char scm_to_uchar (SCM x)
[C Function] short scm_to_short (SCM x)
[C Function] unsigned short scm_to_ushort (SCM x)

[C Function] int scm_to_int (SCM x)

[C Function] unsigned int scm_to_uint (SCM x)

[C Function] long scm_to_long (SCM x)

[C Function] unsigned long scm_to_ulong (SCM x)
[C Function] long long scm_to_long_long (SCM x)
[C Function] unsigned long long scm_to_ulong_long (SCM x)
[C Function] size_t scm_to_size_t (SCM x)
[C Function] ssize_t scm_to_ssize_t (SCM x)
[C Function] scm_t_ptrdiff scm_to_ptrdiff_t (SCM x)
[C Function] scm_t_int8 scm_to_int8 (SCM x)
[C Function] scm_t_uint8 scm_to_uint8 (SCM x)
[C Function] scm_t_int16 scm_to_int16 (SCM x)
[C Function] scm_t_uint16 scm_to_uint16 (SCM x)
[C Function] scm_t_int32 scm_to_int32 (SCM x)
[C Function] scm_t_uint32 scm_to_uint32 (SCM x)
[C Function] scm_t_int64 scm_to_int64 (SCM x)
[C Function] scm_t_uint64 scm_to_uint64 (SCM x)
[C Function] scm_t_intmax scm_to_intmax (SCM x)
[C Function] scm_t_uintmax scm_to_uintmax (SCM x)

When x represents an exact integer that fits into the indicated C type, return that integer. Else signal an error, either a ‘wrong-type’ error when x is not an exact integer, or an ‘out-of-range’ error when it doesn’t fit the given range.

     The functions scm_to_long_long, scm_to_ulong_long, scm_to_int64, and scm_to_

uint64 are only available when the corresponding types are.

[C Function] SCM scm_from_char (char x) [C Function] SCM scm_from_schar (signed char x) [C Function] SCM scm_from_uchar (unsigned char x) [C Function] SCM scm_from_short (short x) [C Function] SCM scm_from_ushort (unsigned short x)

[C Function] SCM scm_from_int (int x) [C Function] SCM scm_from_uint (unsigned int x) [C Function] SCM scm_from_long (long x) [C Function] SCM scm_from_ulong (unsigned long x) [C Function] SCM scm_from_long_long (long long x) [C Function] SCM scm_from_ulong_long (unsigned long long x) [C Function] SCM scm_from_size_t (size t x) [C Function] SCM scm_from_ssize_t (ssize t x) [C Function] SCM scm_from_ptrdiff_t (scm t ptrdiff x) [C Function] SCM scm_from_int8 (scm t int8 x) [C Function] SCM scm_from_uint8 (scm t uint8 x) [C Function] SCM scm_from_int16 (scm t int16 x) [C Function] SCM scm_from_uint16 (scm t uint16 x) [C Function] SCM scm_from_int32 (scm t int32 x) [C Function] SCM scm_from_uint32 (scm t uint32 x) [C Function] SCM scm_from_int64 (scm t int64 x) [C Function] SCM scm_from_uint64 (scm t uint64 x) [C Function] SCM scm_from_intmax (scm t intmax x) [C Function] SCM scm_from_uintmax (scm t uintmax x)

Return the SCM value that represents the integer x. These functions will always succeed and will always return an exact number.

[C Function] void scm_to_mpz (SCM val, mpz t rop)

Assign val to the multiple precision integer rop. val must be an exact integer, otherwise an error will be signalled. rop must have been initialized with mpz_init before this function is called. When rop is no longer needed the occupied space must be freed with mpz_clear. See Section “Initializing Integers” in GNU MP Manual, for details.

[C Function] SCM scm_from_mpz (mpz t val)

Return the SCM value that represents val.

6.6.2.3 Real and Rational Numbers

Mathematically, the real numbers are the set of numbers that describe all possible points along a continuous, infinite, one-dimensional line. The rational numbers are the set of all numbers that can be written as fractions p/q, where p and q are integers. All rational numbers are also real, but there are real numbers that are not rational, for example \sqrt{2}, and π.

Guile can represent both exact and inexact rational numbers, but it cannot represent precise finite irrational numbers. Exact rationals are represented by storing the numerator and denominator as two exact integers. Inexact rationals are stored as floating point numbers using the C type double.

Exact rationals are written as a fraction of integers. There must be no whitespace around the slash:

     1/2
     -22/7

Even though the actual encoding of inexact rationals is in binary, it may be helpful to think of it as a decimal number with a limited number of significant figures and a decimal point somewhere, since this corresponds to the standard notation for non-whole numbers. For example:

     0.34
     -0.00000142857931198
     -5648394822220000000000.0
     4.0

The limited precision of Guile’s encoding means that any finite “real” number in Guile can be written in a rational form, by multiplying and then dividing by sufficient powers of 10 (or in fact, 2). For example, ‘-0.00000142857931198’ is the same as −142857931198 divided by 100000000000000000. In Guile’s current incarnation, therefore, the rational? and real? predicates are equivalent for finite numbers.

Dividing by an exact zero leads to a error message, as one might expect. However, dividing by an inexact zero does not produce an error. Instead, the result of the division is either plus or minus infinity, depending on the sign of the divided number and the sign of the zero divisor (some platforms support signed zeroes ‘-0.0’ and ‘+0.0’; ‘0.0’ is the same as ‘+0.0’).

Dividing zero by an inexact zero yields a NaN (‘not a number’) value, although they are actually considered numbers by Scheme. Attempts to compare a NaN value with any number (including itself) using =, <, >, <= or >= always returns #f. Although a NaN value is not = to itself, it is both eqv? and equal? to itself and other NaN values. However, the preferred way to test for them is by using nan?.

The real NaN values and infinities are written ‘+nan.0’, ‘+inf.0’ and ‘-inf.0’. This syntax is also recognized by read as an extension to the usual Scheme syntax. These special values are considered by Scheme to be inexact real numbers but not rational. Note that non-real complex numbers may also contain infinities or NaN values in their real or imaginary parts. To test a real number to see if it is infinite, a NaN value, or neither, use inf?, nan?, or finite?, respectively. Every real number in Scheme belongs to precisely one of those three classes.

On platforms that follow IEEE 754 for their floating point arithmetic, the ‘+inf.0’, ‘-inf.0’, and ‘+nan.0’ values are implemented using the corresponding IEEE 754 values. They behave in arithmetic operations like IEEE 754 describes it, i.e., (= +nan.0 +nan.0)

⇒ #f.

[Scheme Procedure] real? obj [C Function] scm_real_p (obj)

Return #t if obj is a real number, else #f. Note that the sets of integer and rational values form subsets of the set of real numbers, so the predicate will also be fulfilled if obj is an integer number or a rational number.

[Scheme Procedure] rational? x

[C Function] scm_rational_p (x)

Return #t if x is a rational number, #f otherwise. Note that the set of integer values forms a subset of the set of rational numbers, i.e. the predicate will also be fulfilled if x is an integer number.

[Scheme Procedure] rationalize x eps [C Function] scm_rationalize (x, eps)

Returns the simplest rational number differing from x by no more than eps. As required by R5RS, rationalize only returns an exact result when both its arguments are exact. Thus, you might need to use inexact->exact on the arguments.

     (rationalize (inexact->exact 1.2) 1/100)
     ⇒ 6/5

[Scheme Procedure] inf? x [C Function] scm_inf_p (x)

Return #t if the real number x is ‘+inf.0’ or ‘-inf.0’. Otherwise return #f.

[Scheme Procedure] nan? x [C Function] scm_nan_p (x)

Return #t if the real number x is ‘+nan.0’, or #f otherwise.

[Scheme Procedure] finite? x [C Function] scm_finite_p (x)

Return #t if the real number x is neither infinite nor a NaN, #f otherwise.

[Scheme Procedure] nan [C Function] scm_nan ()

Return ‘+nan.0’, a NaN value.

[Scheme Procedure] inf [C Function] scm_inf ()

Return ‘+inf.0’, positive infinity.

[Scheme Procedure] numerator x [C Function] scm_numerator (x)

Return the numerator of the rational number x.

[Scheme Procedure] denominator x

[C Function] scm_denominator (x)

Return the denominator of the rational number x.

[C Function] int scm_is_real (SCM val) [C Function] int scm_is_rational (SCM val)

Equivalent to scm_is_true (scm_real_p (val)) and scm_is_true (scm_rational_

p (val)), respectively.

[C Function] double scm_to_double (SCM val)

Returns the number closest to val that is representable as a double. Returns infinity for a val that is too large in magnitude. The argument val must be a real number.

[C Function] SCM scm_from_double (double val)

Return the SCM value that represents val. The returned value is inexact according to the predicate inexact?, but it will be exactly equal to val.

6.6.2.4 Complex Numbers

Complex numbers are the set of numbers that describe all possible points in a two-dimensional space. The two coordinates of a particular point in this space are known as the real and imaginary parts of the complex number that describes that point.

In Guile, complex numbers are written in rectangular form as the sum of their real and imaginary parts, using the symbol i to indicate the imaginary part.

     3+4i

⇒

     3.0+4.0i
     (* 3-8i 2.3+0.3i)

⇒

     9.3-17.5i

Polar form can also be used, with an ‘@’ between magnitude and angle,

     1@3.141592 ⇒ -1.0      (approx)
     -1@1.57079 ⇒ 0.0-1.0i  (approx)

Guile represents a complex number as a pair of inexact reals, so the real and imaginary parts of a complex number have the same properties of inexactness and limited precision as single inexact real numbers.

Note that each part of a complex number may contain any inexact real value, including the special values ‘+nan.0’, ‘+inf.0’ and ‘-inf.0’, as well as either of the signed zeroes

‘0.0’ or ‘-0.0’.

[Scheme Procedure] complex? z [C Function] scm_complex_p (z)

Return #t if z is a complex number, #f otherwise. Note that the sets of real, rational and integer values form subsets of the set of complex numbers, i.e. the predicate will also be fulfilled if z is a real, rational or integer number.

[C Function] int scm_is_complex (SCM val)

     Equivalent to scm_is_true (scm_complex_p (val)).

6.6.2.5 Exact and Inexact Numbers

R5RS requires that, with few exceptions, a calculation involving inexact numbers always produces an inexact result. To meet this requirement, Guile distinguishes between an exact integer value such as ‘5’ and the corresponding inexact integer value which, to the limited precision available, has no fractional part, and is printed as ‘5.0’. Guile will only convert the latter value to the former when forced to do so by an invocation of the inexact->exact procedure.

The only exception to the above requirement is when the values of the inexact numbers do not affect the result. For example (expt n 0) is ‘1’ for any value of n, therefore (expt 5.0 0) is permitted to return an exact ‘1’.

[Scheme Procedure] exact? z [C Function] scm_exact_p (z)

Return #t if the number z is exact, #f otherwise.

     (exact? 2)
     ⇒ #t
     (exact? 0.5)
     ⇒ #f
     (exact? (/ 2))
     ⇒ #t

[C Function] int scm_is_exact (SCM z)

Return a 1 if the number z is exact, and 0 otherwise. This is equivalent to scm_is_

true (scm_exact_p (z)).

An alternate approch to testing the exactness of a number is to use scm_is_signed_

integer or scm_is_unsigned_integer.

[Scheme Procedure] inexact? z [C Function] scm_inexact_p (z)

Return #t if the number z is inexact, #f else.

[C Function] int scm_is_inexact (SCM z)

Return a 1 if the number z is inexact, and 0 otherwise. This is equivalent to scm_

is_true (scm_inexact_p (z)).

[Scheme Procedure] inexact->exact z

[C Function] scm_inexact_to_exact (z)

Return an exact number that is numerically closest to z, when there is one. For inexact rationals, Guile returns the exact rational that is numerically equal to the inexact rational. Inexact complex numbers with a non-zero imaginary part can not be made exact.

     (inexact->exact 0.5)
     ⇒ 1/2

The following happens because 12/10 is not exactly representable as a double (on most platforms). However, when reading a decimal number that has been marked exact with the “#e” prefix, Guile is able to represent it correctly.

     (inexact->exact 1.2)
     ⇒ 5404319552844595/4503599627370496
     #e1.2
     ⇒ 6/5

[Scheme Procedure] exact->inexact z

[C Function] scm_exact_to_inexact (z)

Convert the number z to its inexact representation.

6.6.2.6 Read Syntax for Numerical Data

The read syntax for integers is a string of digits, optionally preceded by a minus or plus character, a code indicating the base in which the integer is encoded, and a code indicating whether the number is exact or inexact. The supported base codes are:

#b

#B the integer is written in binary (base 2)

#o

#O the integer is written in octal (base 8)

#d

#D the integer is written in decimal (base 10)

#x

#X the integer is written in hexadecimal (base 16)

If the base code is omitted, the integer is assumed to be decimal. The following examples show how these base codes are used.

     -13
     ⇒ -13
     #d-13
     ⇒ -13
     #x-13
     ⇒ -19
     #b+1101
     ⇒ 13
     #o377
     ⇒ 255

The codes for indicating exactness (which can, incidentally, be applied to all numerical values) are:

#e

#E the number is exact

#i

#I the number is inexact.

If the exactness indicator is omitted, the number is exact unless it contains a radix point. Since Guile can not represent exact complex numbers, an error is signalled when asking for them.

     (exact? 1.2)
     ⇒ #f
     (exact? #e1.2)
     ⇒ #t
     (exact? #e+1i)
     ERROR: Wrong type argument

Guile also understands the syntax ‘+inf.0’ and ‘-inf.0’ for plus and minus infinity, respectively. The value must be written exactly as shown, that is, they always must have a sign and exactly one zero digit after the decimal point. It also understands ‘+nan.0’ and ‘-nan.0’ for the special ‘not-a-number’ value. The sign is ignored for ‘not-a-number’ and the value is always printed as ‘+nan.0’.

6.6.2.7 Operations on Integer Values

[Scheme Procedure] odd? n [C Function] scm_odd_p (n)

Return #t if n is an odd number, #f otherwise.

[Scheme Procedure] even? n [C Function] scm_even_p (n)

Return #t if n is an even number, #f otherwise.

[Scheme Procedure] quotient n d [Scheme Procedure] remainder n d [C Function] scm_quotient (n, d) [C Function] scm_remainder (n, d)

Return the quotient or remainder from n divided by d. The quotient is rounded towards zero, and the remainder will have the same sign as n. In all cases quotient and remainder satisfy n = q ∗ d + r.

     (remainder 13 4) ⇒ 1
     (remainder -13 4) ⇒ -1
See also truncate-quotient, truncate-remainder and related operations in Sec-

tion 6.6.2.11 [Arithmetic], page 119.

[Scheme Procedure] modulo n d [C Function] scm_modulo (n, d)

Return the remainder from n divided by d, with the same sign as d.

     (modulo 13 4) ⇒ 1

     (modulo -13 4) ⇒ 3
     (modulo 13 -4) ⇒ -3
     (modulo -13 -4) ⇒ -1

See also floor-quotient, floor-remainder and related operations in Section 6.6.2.11 [Arithmetic], page 119.

[Scheme Procedure] gcd x… [C Function] scm_gcd (x, y)

Return the greatest common divisor of all arguments. If called without arguments, 0 is returned. The C function scm_gcd always takes two arguments, while the Scheme function can take an arbitrary number.

[Scheme Procedure] lcm x… [C Function] scm_lcm (x, y)

Return the least common multiple of the arguments. If called without arguments, 1 is returned. The C function scm_lcm always takes two arguments, while the Scheme function can take an arbitrary number.

[Scheme Procedure] modulo-expt n k m [C Function] scm_modulo_expt (n, k, m)

Return n raised to the integer exponent k, modulo m.

     (modulo-expt 2 3 5)
        ⇒ 3
[Scheme Procedure] exact-integer-sqrt k
[C Function] void scm_exact_integer_sqrt (SCM k, SCM *s, SCM *r)

Return two exact non-negative integers s and r such that k = s^{2} + r and s^{2} <= k < (s + 1)^{2}. An error is raised if k is not an exact non-negative integer.

     (exact-integer-sqrt 10) ⇒ 3 and 1

6.6.2.8 Comparison Predicates

The C comparison functions below always takes two arguments, while the Scheme functions can take an arbitrary number. Also keep in mind that the C functions return one of the Scheme boolean values SCM_BOOL_T or SCM_BOOL_F which are both true as far as C is concerned. Thus, always write scm_is_true (scm_num_eq_p (x, y)) when testing the two Scheme numbers x and y for equality, for example.

[Scheme Procedure] = [C Function] scm_num_eq_p (x, y)

Return #t if all parameters are numerically equal.

[Scheme Procedure] < [C Function] scm_less_p (x, y)

Return #t if the list of parameters is monotonically increasing.

[Scheme Procedure] > [C Function] scm_gr_p (x, y)

Return #t if the list of parameters is monotonically decreasing.

[Scheme Procedure] <= [C Function] scm_leq_p (x, y)

Return #t if the list of parameters is monotonically non-decreasing.

[Scheme Procedure] >= [C Function] scm_geq_p (x, y)

Return #t if the list of parameters is monotonically non-increasing.

[Scheme Procedure] zero? z [C Function] scm_zero_p (z)

Return #t if z is an exact or inexact number equal to zero.

[Scheme Procedure] positive? x

[C Function] scm_positive_p (x)

Return #t if x is an exact or inexact number greater than zero.

[Scheme Procedure] negative? x

[C Function] scm_negative_p (x)

Return #t if x is an exact or inexact number less than zero.

6.6.2.9 Converting Numbers To and From Strings

The following procedures read and write numbers according to their external representation as defined by R5RS (see Section “Lexical structure” in The Revised^5 Report on the Algorithmic Language Scheme). See Section 6.24.4 [Number Input and Output], page 439, for locale-dependent number parsing.

[Scheme Procedure] number->string n [radix]

[C Function] scm_number_to_string (n, radix)

Return a string holding the external representation of the number n in the given radix. If n is inexact, a radix of 10 will be used.

[Scheme Procedure] string->number string [radix] [C Function] scm_string_to_number (string, radix)

Return a number of the maximally precise representation expressed by the given string. radix must be an exact integer, either 2, 8, 10, or 16. If supplied, radix is a default radix that may be overridden by an explicit radix prefix in string (e.g. "#o177"). If radix is not supplied, then the default radix is 10. If string is not a syntactically valid notation for a number, then string->number returns #f.

[C Function] SCM scm_c_locale_stringn_to_number (const char *string, size t

len, unsigned radix) As per string->number above, but taking a C string, as pointer and length. The string characters should be in the current locale encoding (locale in the name refers only to that, there’s no locale-dependent parsing).

6.6.2.10 Complex Number Operations

[Scheme Procedure] make-rectangular real part imaginary part [C Function] scm_make_rectangular (real part, imaginary part)

Return a complex number constructed of the given real-part and imaginary-part parts.

[Scheme Procedure] make-polar mag ang [C Function] scm_make_polar (mag, ang)

Return the complex number mag * e^(i * ang).

[Scheme Procedure] real-part z [C Function] scm_real_part (z)

Return the real part of the number z.

[Scheme Procedure] imag-part z [C Function] scm_imag_part (z)

Return the imaginary part of the number z.

[Scheme Procedure] magnitude z [C Function] scm_magnitude (z)

Return the magnitude of the number z. This is the same as abs for real arguments, but also allows complex numbers.

[Scheme Procedure] angle z [C Function] scm_angle (z)

Return the angle of the complex number z.

[C Function] SCM scm_c_make_rectangular (double re, double im) [C Function] SCM scm_c_make_polar (double x, double y)

Like scm_make_rectangular or scm_make_polar, respectively, but these functions take doubles as their arguments.

[C Function] double scm_c_real_part (z)
[C Function] double scm_c_imag_part (z)

Returns the real or imaginary part of z as a double.

[C Function] double scm_c_magnitude (z)
[C Function] double scm_c_angle (z)

Returns the magnitude or angle of z as a double.

6.6.2.11 Arithmetic Functions

The C arithmetic functions below always takes two arguments, while the Scheme functions can take an arbitrary number. When you need to invoke them with just one argument, for example to compute the equivalent of (- x), pass SCM_UNDEFINED as the second one:

scm_difference (x, SCM_UNDEFINED).

[Scheme Procedure] + z1 …

[C Function] scm_sum (z1, z2)

Return the sum of all parameter values. Return 0 if called without any parameters.

[Scheme Procedure] - z1 z2 …

[C Function] scm_difference (z1, z2)

If called with one argument z1, -z1 is returned. Otherwise the sum of all but the first argument are subtracted from the first argument.

[Scheme Procedure] * z1 …

[C Function] scm_product (z1, z2)

Return the product of all arguments. If called without arguments, 1 is returned.

[Scheme Procedure] / z1 z2 …

[C Function] scm_divide (z1, z2)

Divide the first argument by the product of the remaining arguments. If called with one argument z1, 1/z1 is returned.

[Scheme Procedure] 1+ z [C Function] scm_oneplus (z)

Return z + 1.

[Scheme Procedure] 1- z [C function] scm_oneminus (z)

Return z − 1.

[Scheme Procedure] abs x [C Function] scm_abs (x)

Return the absolute value of x. x must be a number with zero imaginary part. To calculate the magnitude of a complex number, use magnitude instead.

[Scheme Procedure] max x1 x2 …

[C Function] scm_max (x1, x2)

Return the maximum of all parameter values.

[Scheme Procedure] min x1 x2 …

[C Function] scm_min (x1, x2)

Return the minimum of all parameter values.

[Scheme Procedure] truncate x

[C Function] scm_truncate_number (x)

Round the inexact number x towards zero.

[Scheme Procedure] round x

[C Function] scm_round_number (x)

Round the inexact number x to the nearest integer. When exactly halfway between two integers, round to the even one.

[Scheme Procedure] floor x [C Function] scm_floor (x)

Round the number x towards minus infinity.

[Scheme Procedure] ceiling x [C Function] scm_ceiling (x)

Round the number x towards infinity.

[C Function] double scm_c_truncate (double x)

[C Function] double scm_c_round (double x)

Like scm_truncate_number or scm_round_number, respectively, but these functions take and return double values.

[Scheme Procedure] euclidean/ x y

[Scheme Procedure] euclidean-quotient x y
[Scheme Procedure] euclidean-remainder x y

[C Function] void scm_euclidean_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_euclidean_quotient (SCM x, SCM y)
[C Function] SCM scm_euclidean_remainder (SCM x, SCM y)

These procedures accept two real numbers x and y, where the divisor y must be non-

zero. euclidean-quotient returns the integer q and euclidean-remainder returns

the real number r such that x = q ∗ y + r and 0 <= r < |y|. euclidean/ returns both q and r, and is more efficient than computing each separately. Note that when y > 0, euclidean-quotient returns floor(x/y), otherwise it returns ceiling(x/y). Note that these operators are equivalent to the R6RS operators div, mod, and div-

and-mod.
     (euclidean-quotient 123 10) ⇒ 12
     (euclidean-remainder 123 10) ⇒ 3
     (euclidean/ 123 10) ⇒ 12 and 3
     (euclidean/ 123 -10) ⇒ -12 and 3
     (euclidean/ -123 10) ⇒ -13 and 7
     (euclidean/ -123 -10) ⇒ 13 and 7

(euclidean/ -123.2 -63.5) ⇒ 2.0 and 3.8
(euclidean/ 16/3 -10/7) ⇒ -3 and 22/21

[Scheme Procedure] floor/ x y [Scheme Procedure] floor-quotient x y

[Scheme Procedure] floor-remainder x y

[C Function] void scm_floor_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_floor_quotient (x, y)
[C Function] SCM scm_floor_remainder (x, y)

These procedures accept two real numbers x and y, where the divisor y must be non-zero. floor-quotient returns the integer q and floor-remainder returns the real number r such that q = floor(x/y) and x = q ∗ y + r. floor/ returns both q and r, and is more efficient than computing each separately. Note that r, if non-zero, will have the same sign as y. When x and y are integers, floor-remainder is equivalent to the R5RS integer-only operator modulo.

(floor-quotient 123 10) ⇒ 12
(floor-remainder 123 10) ⇒ 3
(floor/ 123 10) ⇒ 12 and 3
(floor/ 123 -10) ⇒ -13 and -7
(floor/ -123 10) ⇒ -13 and 7
(floor/ -123 -10) ⇒ 12 and -3
(floor/ -123.2 -63.5) ⇒ 1.0 and -59.7
(floor/ 16/3 -10/7) ⇒ -4 and -8/21

[Scheme Procedure] ceiling/ x y

[Scheme Procedure] ceiling-quotient x y
[Scheme Procedure] ceiling-remainder x y

[C Function] void scm_ceiling_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_ceiling_quotient (x, y)
[C Function] SCM scm_ceiling_remainder (x, y)

These procedures accept two real numbers x and y, where the divisor y must be non-zero. ceiling-quotient returns the integer q and ceiling-remainder returns the real number r such that q = ceiling(x/y) and x = q ∗ y + r. ceiling/ returns both q and r, and is more efficient than computing each separately. Note that r, if non-zero, will have the opposite sign of y.

(ceiling-quotient 123 10) ⇒ 13
(ceiling-remainder 123 10) ⇒ -7
(ceiling/ 123 10) ⇒ 13 and -7
(ceiling/ 123 -10) ⇒ -12 and 3
(ceiling/ -123 10) ⇒ -12 and -3
(ceiling/ -123 -10) ⇒ 13 and 7
(ceiling/ -123.2 -63.5) ⇒ 2.0 and 3.8
(ceiling/ 16/3 -10/7) ⇒ -3 and 22/21

[Scheme Procedure] truncate/ x y

[Scheme Procedure] truncate-quotient x y

[Scheme Procedure] truncate-remainder x y

[C Function] void scm_truncate_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_truncate_quotient (x, y)
[C Function] SCM scm_truncate_remainder (x, y)

These procedures accept two real numbers x and y, where the divisor y must be nonzero. truncate-quotient returns the integer q and truncate-remainder returns the real number r such that q is x/y rounded toward zero, and x = q ∗ y + r. truncate/ returns both q and r, and is more efficient than computing each separately. Note that r, if non-zero, will have the same sign as x. When x and y are integers, these operators are equivalent to the R5RS integer-only

operators quotient and remainder.
(truncate-quotient 123 10) ⇒ 12
(truncate-remainder 123 10) ⇒ 3
(truncate/ 123 10) ⇒ 12 and 3
(truncate/ 123 -10) ⇒ -12 and 3
(truncate/ -123 10) ⇒ -12 and -3
(truncate/ -123 -10) ⇒ 12 and -3
(truncate/ -123.2 -63.5) ⇒ 1.0 and -59.7
(truncate/ 16/3 -10/7) ⇒ -3 and 22/21

[Scheme Procedure] centered/ x y

[Scheme Procedure] centered-quotient x y
[Scheme Procedure] centered-remainder x y

[C Function] void scm_centered_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_centered_quotient (SCM x, SCM y)
[C Function] SCM scm_centered_remainder (SCM x, SCM y)

These procedures accept two real numbers x and y, where the divisor y must be non-

zero. centered-quotient returns the integer q and centered-remainder returns

the real number r such that x = q ∗ y + r and −|y/2| <= r < |y/2|. centered/ returns both q and r, and is more efficient than computing each separately. Note that centered-quotient returns x/y rounded to the nearest integer. When x/y lies exactly half-way between two integers, the tie is broken according to the sign of y. If y > 0, ties are rounded toward positive infinity, otherwise they are rounded toward negative infinity. This is a consequence of the requirement that −|y/2| <= r < |y/2|. Note that these operators are equivalent to the R6RS operators div0, mod0, and

div0-and-mod0.
(centered-quotient 123 10) ⇒ 12
(centered-remainder 123 10) ⇒ 3
(centered/ 123 10) ⇒ 12 and 3
(centered/ 123 -10) ⇒ -12 and 3
(centered/ -123 10) ⇒ -12 and -3
(centered/ -123 -10) ⇒ 12 and -3
(centered/ 125 10) ⇒ 13 and -5
(centered/ 127 10) ⇒ 13 and -3
(centered/ 135 10) ⇒ 14 and -5
(centered/ -123.2 -63.5) ⇒ 2.0 and 3.8

(centered/ 16/3 -10/7) ⇒ -4 and -8/21

[Scheme Procedure] round/ x y [Scheme Procedure] round-quotient x y

[Scheme Procedure] round-remainder x y

[C Function] void scm_round_divide (SCM x, SCM y, SCM *q, SCM *r)

[C Function] SCM scm_round_quotient (x, y)
[C Function] SCM scm_round_remainder (x, y)

These procedures accept two real numbers x and y, where the divisor y must be non-zero. round-quotient returns the integer q and round-remainder returns the real number r such that x = q ∗ y + r and q is x/y rounded to the nearest integer, with ties going to the nearest even integer. round/ returns both q and r, and is more efficient than computing each separately. Note that round/ and centered/ are almost equivalent, but their behavior differs when x/y lies exactly half-way between two integers. In this case, round/ chooses the nearest even integer, whereas centered/ chooses in such a way to satisfy the constraint −|y/2| <= r < |y/2|, which is stronger than the corresponding constraint for round/, −|y/2| <= r <= |y/2|. In particular, when x and y are integers, the number of possible remainders returned by centered/ is |y|, whereas the number of possible remainders returned by round/ is |y| + 1 when y is even.

(round-quotient 123 10) ⇒ 12
(round-remainder 123 10) ⇒ 3
(round/ 123 10) ⇒ 12 and 3
(round/ 123 -10) ⇒ -12 and 3
(round/ -123 10) ⇒ -12 and -3
(round/ -123 -10) ⇒ 12 and -3
(round/ 125 10) ⇒ 12 and 5
(round/ 127 10) ⇒ 13 and -3
(round/ 135 10) ⇒ 14 and -5
(round/ -123.2 -63.5) ⇒ 2.0 and 3.8
(round/ 16/3 -10/7) ⇒ -4 and -8/21

6.6.2.12 Scientific Functions

The following procedures accept any kind of number as arguments, including complex numbers.

[Scheme Procedure] sqrt z

Return the square root of z. Of the two possible roots (positive and negative), the one with a positive real part is returned, or if that’s zero then a positive imaginary part. Thus,

(sqrt 9.0)  ⇒ 3.0
(sqrt -9.0)  ⇒ 0.0+3.0i
(sqrt 1.0+1.0i) ⇒ 1.09868411346781+0.455089860562227i
(sqrt -1.0-1.0i) ⇒ 0.455089860562227-1.09868411346781i

[Scheme Procedure] expt z1 z2

Return z1 raised to the power of z2.

[Scheme Procedure] sin z

Return the sine of z.

[Scheme Procedure] cos z

Return the cosine of z.

[Scheme Procedure] tan z

Return the tangent of z.

[Scheme Procedure] asin z

Return the arcsine of z.

[Scheme Procedure] acos z

Return the arccosine of z.

[Scheme Procedure] atan z [Scheme Procedure] atan y x

Return the arctangent of z, or of y/x.

[Scheme Procedure] exp z

Return e to the power of z, where e is the base of natural logarithms (2.71828…).

[Scheme Procedure] log z

Return the natural logarithm of z.

[Scheme Procedure] log10 z

Return the base 10 logarithm of z.

[Scheme Procedure] sinh z

Return the hyperbolic sine of z.

[Scheme Procedure] cosh z

Return the hyperbolic cosine of z.

[Scheme Procedure] tanh z

Return the hyperbolic tangent of z.

[Scheme Procedure] asinh z

Return the hyperbolic arcsine of z.

[Scheme Procedure] acosh z

Return the hyperbolic arccosine of z.

[Scheme Procedure] atanh z

Return the hyperbolic arctangent of z.

6.6.2.13 Bitwise Operations

For the following bitwise functions, negative numbers are treated as infinite precision twoscomplements. For instance −6 is bits … 111010, with infinitely many ones on the left. It can be seen that adding 6 (binary 110) to such a bit pattern gives all zeros.

[Scheme Procedure] logand n1 n2 … [C Function] scm_logand (n1, n2)

Return the bitwise and of the integer arguments.

(logand) ⇒ -1
(logand 7) ⇒ 7
(logand #b111 #b011 #b001) ⇒ 1

[Scheme Procedure] logior n1 n2 … [C Function] scm_logior (n1, n2)

Return the bitwise or of the integer arguments.

(logior) ⇒ 0
(logior 7) ⇒ 7
(logior #b000 #b001 #b011) ⇒ 3

[Scheme Procedure] logxor n1 n2 … [C Function] scm_loxor (n1, n2)

Return the bitwise xor of the integer arguments. A bit is set in the result if it is set in an odd number of arguments.

(logxor) ⇒ 0
(logxor 7) ⇒ 7
(logxor #b000 #b001 #b011) ⇒ 2
(logxor #b000 #b001 #b011 #b011) ⇒ 1

[Scheme Procedure] lognot n [C Function] scm_lognot (n)

Return the integer which is the ones-complement of the integer argument, ie. each 0 bit is changed to 1 and each 1 bit to 0.

(number->string (lognot #b10000000) 2)
   ⇒ "-10000001"
(number->string (lognot #b0) 2)
   ⇒ "-1"

[Scheme Procedure] logtest j k [C Function] scm_logtest (j, k)

Test whether j and k have any 1 bits in common. This is equivalent to (not (zero? (logand j k))), but without actually calculating the logand, just testing for nonzero.

(logtest #b0100 #b1011) ⇒ #f
(logtest #b0100 #b0111) ⇒ #t

[Scheme Procedure] logbit? index j [C Function] scm_logbit_p (index, j)

Test whether bit number index in j is set. index starts from 0 for the least significant bit.

(logbit? 0 #b1101) ⇒ #t
(logbit? 1 #b1101) ⇒ #f
(logbit? 2 #b1101) ⇒ #t
(logbit? 3 #b1101) ⇒ #t
(logbit? 4 #b1101) ⇒ #f

[Scheme Procedure] ash n count [C Function] scm_ash (n, count)

Return floor(n ∗ 2^{c}ount). n and count must be exact integers. With n viewed as an infinite-precision twos-complement integer, ash means a left shift introducing zero bits when count is positive, or a right shift dropping bits when count is negative. This is an “arithmetic” shift.

(number->string (ash #b1 3) 2)  ⇒ "1000"
(number->string (ash #b1010 -1) 2) ⇒ "101"
;; -23 is bits ...11101001, -6 is bits ...111010
(ash -23 -2) ⇒ -6

[Scheme Procedure] round-ash n count [C Function] scm_round_ash (n, count)

Return round(n ∗ 2^{c}ount). n and count must be exact integers. With n viewed as an infinite-precision twos-complement integer, round-ash means a left shift introducing zero bits when count is positive, or a right shift rounding to the nearest integer (with ties going to the nearest even integer) when count is negative. This is a rounded “arithmetic” shift.

(number->string (round-ash #b1 3) 2)  ⇒ \"1000\"
(number->string (round-ash #b1010 -1) 2) ⇒ \"101\"
(number->string (round-ash #b1010 -2) 2) ⇒ \"10\"
(number->string (round-ash #b1011 -2) 2) ⇒ \"11\"
(number->string (round-ash #b1101 -2) 2) ⇒ \"11\"
(number->string (round-ash #b1110 -2) 2) ⇒ \"100\"

[Scheme Procedure] logcount n [C Function] scm_logcount (n)

Return the number of bits in integer n. If n is positive, the 1-bits in its binary representation are counted. If negative, the 0-bits in its two’s-complement binary representation are counted. If zero, 0 is returned.

(logcount #b10101010)
   ⇒ 4
(logcount 0)
   ⇒ 0
(logcount -2)
   ⇒ 1

[Scheme Procedure] integer-length n

[C Function] scm_integer_length (n)

Return the number of bits necessary to represent n.

For positive n this is how many bits to the most significant one bit. For negative n it’s how many bits to the most significant zero bit in twos complement form.

(integer-length #b10101010) ⇒ 8
(integer-length #b1111)  ⇒ 4
(integer-length 0)  ⇒ 0
(integer-length -1)  ⇒ 0
(integer-length -256)  ⇒ 8
(integer-length -257)  ⇒ 9

[Scheme Procedure] integer-expt n k

[C Function] scm_integer_expt (n, k)

Return n raised to the power k. k must be an exact integer, n can be any number.

Negative k is supported, and results in 1/n^{|}k| in the usual way. n^{0} is 1, as usual, and that includes 0^{0} is 1.

(integer-expt 2 5) ⇒ 32
(integer-expt -3 3) ⇒ -27
(integer-expt 5 -3) ⇒ 1/125
(integer-expt 0 0) ⇒ 1

[Scheme Procedure] bit-extract n start end [C Function] scm_bit_extract (n, start, end)

Return the integer composed of the start (inclusive) through end (exclusive) bits of n. The startth bit becomes the 0-th bit in the result.

(number->string (bit-extract #b1101101010 0 4) 2)
   ⇒ "1010"
(number->string (bit-extract #b1101101010 4 9) 2)
   ⇒ "10110"

6.6.2.14 Random Number Generation

Pseudo-random numbers are generated from a random state object, which can be created with seed->random-state or datum->random-state. An external representation (i.e. one which can written with write and read with read) of a random state object can be obtained via random-state->datum. The state parameter to the various functions below is optional, it defaults to the state object in the *random-state* variable.

[Scheme Procedure] copy-random-state [state]

[C Function] scm_copy_random_state (state)

Return a copy of the random state state.

[Scheme Procedure] random n [state] [C Function] scm_random (n, state)

Return a number in [0, n).

Accepts a positive integer or real n and returns a number of the same type between zero (inclusive) and n (exclusive). The values returned have a uniform distribution.

[Scheme Procedure] random:exp [state] [C Function] scm_random_exp (state)

Return an inexact real in an exponential distribution with mean 1. For an exponential distribution with mean u use (* u (random:exp)).

[Scheme Procedure] random:hollow-sphere! vect [state]

[C Function] scm_random_hollow_sphere_x (vect, state)

Fills vect with inexact real random numbers the sum of whose squares is equal to 1.0. Thinking of vect as coordinates in space of dimension n = (vector-length vect), the coordinates are uniformly distributed over the surface of the unit n-sphere.

[Scheme Procedure] random:normal [state] [C Function] scm_random_normal (state)

Return an inexact real in a normal distribution. The distribution used has mean 0 and standard deviation 1. For a normal distribution with mean m and standard

deviation d use (+ m (* d (random:normal))).

[Scheme Procedure] random:normal-vector! vect [state]

[C Function] scm_random_normal_vector_x (vect, state)

Fills vect with inexact real random numbers that are independent and standard normally distributed (i.e., with mean 0 and variance 1).

[Scheme Procedure] random:solid-sphere! vect [state]

[C Function] scm_random_solid_sphere_x (vect, state)

Fills vect with inexact real random numbers the sum of whose squares is less than 1.0. Thinking of vect as coordinates in space of dimension n = (vector-length vect), the coordinates are uniformly distributed within the unit n-sphere.

[Scheme Procedure] random:uniform [state]

[C Function] scm_random_uniform (state)

Return a uniformly distributed inexact real random number in [0,1).

[Scheme Procedure] seed->random-state seed

[C Function] scm_seed_to_random_state (seed)

Return a new random state using seed.

[Scheme Procedure] datum->random-state datum

[C Function] scm_datum_to_random_state (datum)

Return a new random state from datum, which should have been obtained by random-

state->datum.

[Scheme Procedure] random-state->datum state

[C Function] scm_random_state_to_datum (state)

Return a datum representation of state that may be written out and read back with the Scheme reader.

[Scheme Procedure] random-state-from-platform
[C Function] scm_random_state_from_platform ()

Construct a new random state seeded from a platform-specific source of entropy, appropriate for use in non-security-critical applications. Currently /dev/urandom is tried first, or else the seed is based on the time, date, process ID, an address from a freshly allocated heap cell, an address from the local stack frame, and a high-resolution timer if available.

[Variable] *random-state*

The global random state used by the above functions when the state parameter is not given.

Note that the initial value of *random-state* is the same every time Guile starts up. Therefore, if you don’t pass a state parameter to the above procedures, and you don’t

set *random-state* to (seed->random-state your-seed), where your-seed is something

that isn’t the same every time, you’ll get the same sequence of “random” numbers on every run.

For example, unless the relevant source code has changed, (map random (cdr (iota 30))), if the first use of random numbers since Guile started up, will always give:

     (map random (cdr (iota 19)))

⇒

     (0 1 1 2 2 2 1 2 6 7 10 0 5 3 12 5 5 12)

To seed the random state in a sensible way for non-security-critical applications, do this during initialization of your program:

     (set! *random-state* (random-state-from-platform))

6.6.3 Characters

In Scheme, there is a data type to describe a single character.

Defining what exactly a character is can be more complicated than it seems. Guile follows the advice of R6RS and uses The Unicode Standard to help define what a character is. So, for Guile, a character is anything in the Unicode Character Database.

The Unicode Character Database is basically a table of characters indexed using integers called ’code points’. Valid code points are in the ranges 0 to #xD7FF inclusive or #xE000 to #x10FFFF inclusive, which is about 1.1 million code points.

Any code point that has been assigned to a character or that has otherwise been given a meaning by Unicode is called a ’designated code point’. Most of the designated code points, about 200,000 of them, indicate characters, accents or other combining marks that modify other characters, symbols, whitespace, and control characters. Some are not characters but indicators that suggest how to format or display neighboring characters.

If a code point is not a designated code point – if it has not been assigned to a character by The Unicode Standard – it is a ’reserved code point’, meaning that they are reserved for future use. Most of the code points, about 800,000, are ’reserved code points’.

By convention, a Unicode code point is written as “U+XXXX” where “XXXX” is a hexadecimal number. Please note that this convenient notation is not valid code. Guile does not interpret “U+XXXX” as a character.

In Scheme, a character literal is written as #\name where name is the name of the character that you want. Printable characters have their usual single character name; for example, #\a is a lower case a.

Some of the code points are ’combining characters’ that are not meant to be printed by themselves but are instead meant to modify the appearance of the previous character. For combining characters, an alternate form of the character literal is #\ followed by U+25CC (a small, dotted circle), followed by the combining character. This allows the combining character to be drawn on the circle, not on the backslash of #\.

Many of the non-printing characters, such as whitespace characters and control characters, also have names.

The most commonly used non-printing characters have long character names, described in the table below.

Character Name  Codepoint

#\nul  U+0000
#\alarm  u+0007
#\backspace U+0008
#\tab  U+0009
#\linefeed U+000A
#\newline  U+000A
#\vtab  U+000B
#\page  U+000C
#\return  U+000D
#\esc  U+001B
#\space  U+0020
#\delete  U+007F

There are also short names for all of the “C0 control characters” (those with code points below 32). The following table lists the short name for each character.

0 = #\nul  1 = #\soh  2 = #\stx  3 = #\etx
4 = #\eot  5 = #\enq  6 = #\ack  7 = #\bel
8 = #\bs  9 = #\ht  10 = #\lf  11 = #\vt
12 = #\ff  13 = #\cr  14 = #\so  15 = #\si
16 = #\dle  17 = #\dc1  18 = #\dc2  19 = #\dc3
20 = #\dc4  21 = #\nak  22 = #\syn  23 = #\etb
24 = #\can  25 = #\em  26 = #\sub  27 = #\esc
28 = #\fs  29 = #\gs  30 = #\rs  31 = #\us

32 = #\sp

The short name for the “delete” character (code point U+007F) is #\del.

The R7RS name for the “escape” character (code point U+001B) is #\escape.

There are also a few alternative names left over for compatibility with previous versions of Guile.

Alternate  Standard

#\nl          #\newline
#\np          #\page
#\null        #\nul

Characters may also be written using their code point values. They can be written with as an octal number, such as #\10 for #\bs or #\177 for #\del.

If one prefers hex to octal, there is an additional syntax for character escapes: #\xHHHH – the letter ’x’ followed by a hexadecimal number of one to eight digits.

[Scheme Procedure] char? x [C Function] scm_char_p (x)

Return #t if x is a character, else #f.

Fundamentally, the character comparison operations below are numeric comparisons of the character’s code points.

[Scheme Procedure] char=? x y

Return #t if code point of x is equal to the code point of y, else #f.

[Scheme Procedure] char<? x y

Return #t if the code point of x is less than the code point of y, else #f.

[Scheme Procedure] char<=? x y

Return #t if the code point of x is less than or equal to the code point of y, else #f.

[Scheme Procedure] char>? x y

Return #t if the code point of x is greater than the code point of y, else #f.

[Scheme Procedure] char>=? x y

Return #t if the code point of x is greater than or equal to the code point of y, else

#f.

Case-insensitive character comparisons use Unicode case folding. In case folding comparisons, if a character is lowercase and has an uppercase form that can be expressed as a single character, it is converted to uppercase before comparison. All other characters undergo no conversion before the comparison occurs. This includes the German sharp S (Eszett) which is not uppercased before conversion because its uppercase form has two characters. Unicode case folding is language independent: it uses rules that are generally true, but, it cannot cover all cases for all languages.

[Scheme Procedure] char-ci=? x y

Return #t if the case-folded code point of x is the same as the case-folded code point of y, else #f.

[Scheme Procedure] char-ci<? x y

Return #t if the case-folded code point of x is less than the case-folded code point of y, else #f.

[Scheme Procedure] char-ci<=? x y

Return #t if the case-folded code point of x is less than or equal to the case-folded code point of y, else #f.

[Scheme Procedure] char-ci>? x y

Return #t if the case-folded code point of x is greater than the case-folded code point of y, else #f.

[Scheme Procedure] char-ci>=? x y

Return #t if the case-folded code point of x is greater than or equal to the case-folded code point of y, else #f.

[Scheme Procedure] char-alphabetic? chr

[C Function] scm_char_alphabetic_p (chr)

Return #t if chr is alphabetic, else #f.

[Scheme Procedure] char-numeric? chr

[C Function] scm_char_numeric_p (chr)

Return #t if chr is numeric, else #f.

[Scheme Procedure] char-whitespace? chr

[C Function] scm_char_whitespace_p (chr)

Return #t if chr is whitespace, else #f.

[Scheme Procedure] char-upper-case? chr

[C Function] scm_char_upper_case_p (chr)

Return #t if chr is uppercase, else #f.

[Scheme Procedure] char-lower-case? chr

[C Function] scm_char_lower_case_p (chr)

Return #t if chr is lowercase, else #f.

[Scheme Procedure] char-is-both? chr

[C Function] scm_char_is_both_p (chr)

Return #t if chr is either uppercase or lowercase, else #f.

[Scheme Procedure] char-general-category chr
[C Function] scm_char_general_category (chr)

Return a symbol giving the two-letter name of the Unicode general category assigned to chr or #f if no named category is assigned. The following table provides a list of category names along with their meanings.

Lu  Uppercase letter  Pf  Final quote punctuation
Ll  Lowercase letter  Po  Other punctuation
Lt  Titlecase letter  Sm  Math symbol
Lm  Modifier letter  Sc  Currency symbol
Lo  Other letter  Sk  Modifier symbol
Mn  Non-spacing mark  So  Other symbol
Mc  Combining spacing mark  Zs  Space separator
Me  Enclosing mark  Zl  Line separator
Nd  Decimal digit number  Zp  Paragraph separator
Nl  Letter number  Cc  Control
No  Other number  Cf  Format
Pc  Connector punctuation  Cs  Surrogate
Pd  Dash punctuation  Co  Private use
Ps  Open punctuation  Cn  Unassigned
Pe  Close punctuation
Pi  Initial quote punctuation

[Scheme Procedure] char->integer chr

[C Function] scm_char_to_integer (chr)

Return the code point of chr.

[Scheme Procedure] integer->char n

[C Function] scm_integer_to_char (n)

Return the character that has code point n. The integer n must be a valid code point. Valid code points are in the ranges 0 to #xD7FF inclusive or #xE000 to #x10FFFF inclusive.

[Scheme Procedure] char-upcase chr [C Function] scm_char_upcase (chr)

Return the uppercase character version of chr.

[Scheme Procedure] char-downcase chr

[C Function] scm_char_downcase (chr)

Return the lowercase character version of chr.

[Scheme Procedure] char-titlecase chr

[C Function] scm_char_titlecase (chr)

Return the titlecase character version of chr if one exists; otherwise return the uppercase version.

For most characters these will be the same, but the Unicode Standard includes certain digraph compatibility characters, such as U+01F3 “dz”, for which the uppercase and titlecase characters are different (U+01F1 “DZ” and U+01F2 “Dz” in this case, respectively).

[C Function] scm_t_wchar scm_c_upcase (scm t wchar c)
[C Function] scm_t_wchar scm_c_downcase (scm t wchar c)
[C Function] scm_t_wchar scm_c_titlecase (scm t wchar c)

These C functions take an integer representation of a Unicode codepoint and return the codepoint corresponding to its uppercase, lowercase, and titlecase forms respectively. The type scm_t_wchar is a signed, 32-bit integer.

6.6.4 Character Sets

The features described in this section correspond directly to SRFI-14.

The data type charset implements sets of characters (see Section 6.6.3 [Characters], page 129). Because the internal representation of character sets is not visible to the user, a lot of procedures for handling them are provided.

Character sets can be created, extended, tested for the membership of a characters and be compared to other character sets.

6.6.4.1 Character Set Predicates/Comparison

Use these procedures for testing whether an object is a character set, or whether several character sets are equal or subsets of each other. char-set-hash can be used for calculating a hash value, maybe for usage in fast lookup procedures.

[Scheme Procedure] char-set? obj [C Function] scm_char_set_p (obj)

Return #t if obj is a character set, #f otherwise.

[Scheme Procedure] char-set= char set … [C Function] scm_char_set_eq (char sets)

Return #t if all given character sets are equal.

[Scheme Procedure] char-set<= char set … [C Function] scm_char_set_leq (char sets)

Return #t if every character set char seti is a subset of character set char seti+1.

[Scheme Procedure] char-set-hash cs [bound] [C Function] scm_char_set_hash (cs, bound)

Compute a hash value for the character set cs. If bound is given and non-zero, it restricts the returned value to the range 0 … bound - 1.

6.6.4.2 Iterating Over Character Sets

Character set cursors are a means for iterating over the members of a character sets. After creating a character set cursor with char-set-cursor, a cursor can be dereferenced with char-set-ref, advanced to the next member with char-set-cursor-next. Whether a cursor has passed past the last element of the set can be checked with end-of-char-set?.

Additionally, mapping and (un-)folding procedures for character sets are provided.

[Scheme Procedure] char-set-cursor cs

[C Function] scm_char_set_cursor (cs)

Return a cursor into the character set cs.

[Scheme Procedure] char-set-ref cs cursor [C Function] scm_char_set_ref (cs, cursor)

Return the character at the current cursor position cursor in the character set cs. It is an error to pass a cursor for which end-of-char-set? returns true.

[Scheme Procedure] char-set-cursor-next cs cursor

[C Function] scm_char_set_cursor_next (cs, cursor)

Advance the character set cursor cursor to the next character in the character set cs. It is an error if the cursor given satisfies end-of-char-set?.

[Scheme Procedure] end-of-char-set? cursor

[C Function] scm_end_of_char_set_p (cursor)

Return #t if cursor has reached the end of a character set, #f otherwise.

[Scheme Procedure] char-set-fold kons knil cs [C Function] scm_char_set_fold (kons, knil, cs)

Fold the procedure kons over the character set cs, initializing it with knil.

[Scheme Procedure] char-set-unfold p f g seed [base cs] [C Function] scm_char_set_unfold (p, f, g, seed, base cs)

This is a fundamental constructor for character sets.

• g is used to generate a series of “seed” values from the initial seed: seed, (g seed), (g^2 seed), (g^3 seed), …

• p tells us when to stop – when it returns true when applied to one of the seed values.

• f maps each seed value to a character. These characters are added to the base character set base cs to form the result; base cs defaults to the empty set.

[Scheme Procedure] char-set-unfold! p f g seed base cs [C Function] scm_char_set_unfold_x (p, f, g, seed, base cs)

This is a fundamental constructor for character sets.

• g is used to generate a series of “seed” values from the initial seed: seed, (g seed), (g^2 seed), (g^3 seed), …

• p tells us when to stop – when it returns true when applied to one of the seed values.

• f maps each seed value to a character. These characters are added to the base character set base cs to form the result; base cs defaults to the empty set.

[Scheme Procedure] char-set-for-each proc cs

[C Function] scm_char_set_for_each (proc, cs)

Apply proc to every character in the character set cs. The return value is not specified.

[Scheme Procedure] char-set-map proc cs [C Function] scm_char_set_map (proc, cs)

Map the procedure proc over every character in cs. proc must be a character -> character procedure.

6.6.4.3 Creating Character Sets

New character sets are produced with these procedures.

[Scheme Procedure] char-set-copy cs

[C Function] scm_char_set_copy (cs)

Return a newly allocated character set containing all characters in cs.

[Scheme Procedure] char-set chr … [C Function] scm_char_set (chrs)

Return a character set containing all given characters.

[Scheme Procedure] list->char-set list [base cs] [C Function] scm_list_to_char_set (list, base cs)

Convert the character list list to a character set. If the character set base cs is given, the character in this set are also included in the result.

[Scheme Procedure] list->char-set! list base cs [C Function] scm_list_to_char_set_x (list, base cs)

Convert the character list list to a character set. The characters are added to base cs and base cs is returned.

[Scheme Procedure] string->char-set str [base cs] [C Function] scm_string_to_char_set (str, base cs)

Convert the string str to a character set. If the character set base cs is given, the characters in this set are also included in the result.

[Scheme Procedure] string->char-set! str base cs

[C Function] scm_string_to_char_set_x (str, base cs)

Convert the string str to a character set. The characters from the string are added to base cs, and base cs is returned.

[Scheme Procedure] char-set-filter pred cs [base cs] [C Function] scm_char_set_filter (pred, cs, base cs)

Return a character set containing every character from cs so that it satisfies pred. If provided, the characters from base cs are added to the result.

[Scheme Procedure] char-set-filter! pred cs base cs [C Function] scm_char_set_filter_x (pred, cs, base cs)

Return a character set containing every character from cs so that it satisfies pred. The characters are added to base cs and base cs is returned.

[Scheme Procedure] ucs-range->char-set lower upper [error [base cs]]

[C Function] scm_ucs_range_to_char_set (lower, upper, error, base cs)

Return a character set containing all characters whose character codes lie in the halfopen range [lower,upper).

If error is a true value, an error is signalled if the specified range contains characters which are not contained in the implemented character range. If error is #f, these characters are silently left out of the resulting character set.

The characters in base cs are added to the result, if given.

[Scheme Procedure] ucs-range->char-set! lower upper error base cs

[C Function] scm_ucs_range_to_char_set_x (lower, upper, error, base cs)

Return a character set containing all characters whose character codes lie in the halfopen range [lower,upper).

If error is a true value, an error is signalled if the specified range contains characters which are not contained in the implemented character range. If error is #f, these characters are silently left out of the resulting character set.

The characters are added to base cs and base cs is returned.

[Scheme Procedure] ->char-set x

[C Function] scm_to_char_set (x)

Coerces x into a char-set. x may be a string, character or char-set. A string is converted to the set of its constituent characters; a character is converted to a singleton set; a char-set is returned as-is.

6.6.4.4 Querying Character Sets

Access the elements and other information of a character set with these procedures.

[Scheme Procedure] %char-set-dump cs

Returns an association list containing debugging information for cs. The association list has the following entries.

char-set The char-set itself

len The number of groups of contiguous code points the char-set contains

ranges A list of lists where each sublist is a range of code points and their associated characters

The return value of this function cannot be relied upon to be consistent between versions of Guile and should not be used in code.

[Scheme Procedure] char-set-size cs

[C Function] scm_char_set_size (cs)

Return the number of elements in character set cs.

[Scheme Procedure] char-set-count pred cs [C Function] scm_char_set_count (pred, cs)

Return the number of the elements int the character set cs which satisfy the predicate pred.

[Scheme Procedure] char-set->list cs

[C Function] scm_char_set_to_list (cs)

Return a list containing the elements of the character set cs.

[Scheme Procedure] char-set->string cs

[C Function] scm_char_set_to_string (cs)

Return a string containing the elements of the character set cs. The order in which the characters are placed in the string is not defined.

[Scheme Procedure] char-set-contains? cs ch

[C Function] scm_char_set_contains_p (cs, ch)

Return #t if the character ch is contained in the character set cs, or #f otherwise.

[Scheme Procedure] char-set-every pred cs [C Function] scm_char_set_every (pred, cs)

Return a true value if every character in the character set cs satisfies the predicate pred.

[Scheme Procedure] char-set-any pred cs [C Function] scm_char_set_any (pred, cs)

Return a true value if any character in the character set cs satisfies the predicate pred.

6.6.4.5 Character-Set Algebra

Character sets can be manipulated with the common set algebra operation, such as union, complement, intersection etc. All of these procedures provide side-effecting variants, which modify their character set argument(s).

[Scheme Procedure] char-set-adjoin cs chr … [C Function] scm_char_set_adjoin (cs, chrs)

Add all character arguments to the first argument, which must be a character set.

[Scheme Procedure] char-set-delete cs chr … [C Function] scm_char_set_delete (cs, chrs)

Delete all character arguments from the first argument, which must be a character set.

[Scheme Procedure] char-set-adjoin! cs chr …

[C Function] scm_char_set_adjoin_x (cs, chrs)

Add all character arguments to the first argument, which must be a character set.

[Scheme Procedure] char-set-delete! cs chr …

[C Function] scm_char_set_delete_x (cs, chrs)

Delete all character arguments from the first argument, which must be a character set.

[Scheme Procedure] char-set-complement cs
[C Function] scm_char_set_complement (cs)

Return the complement of the character set cs.

Note that the complement of a character set is likely to contain many reserved code points (code points that are not associated with characters). It may be helpful to modify the output of char-set-complement by computing its intersection with the set of designated

code points, char-set:designated.

[Scheme Procedure] char-set-union cs … [C Function] scm_char_set_union (char sets)

Return the union of all argument character sets.

[Scheme Procedure] char-set-intersection cs …

[C Function] scm_char_set_intersection (char sets)

Return the intersection of all argument character sets.

[Scheme Procedure] char-set-difference cs1 cs … [C Function] scm_char_set_difference (cs1, char sets)

Return the difference of all argument character sets.

[Scheme Procedure] char-set-xor cs … [C Function] scm_char_set_xor (char sets)

Return the exclusive-or of all argument character sets.

[Scheme Procedure] char-set-diff+intersection cs1 cs . . .
[C Function] scm_char_set_diff_plus_intersection (cs1, char sets)

Return the difference and the intersection of all argument character sets.

[Scheme Procedure] char-set-complement! cs
[C Function] scm_char_set_complement_x (cs)

Return the complement of the character set cs.

[Scheme Procedure] char-set-union! cs1 cs … [C Function] scm_char_set_union_x (cs1, char sets)

Return the union of all argument character sets.

[Scheme Procedure] char-set-intersection! cs1 cs …

[C Function] scm_char_set_intersection_x (cs1, char sets)

Return the intersection of all argument character sets.

[Scheme Procedure] char-set-difference! cs1 cs …

[C Function] scm_char_set_difference_x (cs1, char sets)

Return the difference of all argument character sets.

[Scheme Procedure] char-set-xor! cs1 cs … [C Function] scm_char_set_xor_x (cs1, char sets)

Return the exclusive-or of all argument character sets.

[Scheme Procedure] char-set-diff+intersection! cs1 cs2 cs …

[C Function] scm_char_set_diff_plus_intersection_x (cs1, cs2, char sets)

Return the difference and the intersection of all argument character sets.

6.6.4.6 Standard Character Sets

In order to make the use of the character set data type and procedures useful, several predefined character set variables exist.

These character sets are locale independent and are not recomputed upon a setlocale call. They contain characters from the whole range of Unicode code points. For instance, char-set:letter contains about 100,000 characters.

[Scheme Variable] char-set:lower-case
[C Variable] scm_char_set_lower_case

All lower-case characters.

[Scheme Variable] char-set:upper-case
[C Variable] scm_char_set_upper_case

All upper-case characters.

[Scheme Variable] char-set:title-case
[C Variable] scm_char_set_title_case

All single characters that function as if they were an upper-case letter followed by a lower-case letter.

[Scheme Variable] char-set:letter

[C Variable] scm_char_set_letter
All letters.  This includes char-set:lower-case, char-set:upper-case,

char-set:title-case, and many letters that have no case at all. For example, Chinese and Japanese characters typically have no concept of case.

[Scheme Variable] char-set:digit

[C Variable] scm_char_set_digit

All digits.

[Scheme Variable] char-set:letter+digit
[C Variable] scm_char_set_letter_and_digit
     The union of char-set:letter and char-set:digit.
[Scheme Variable] char-set:graphic
[C Variable] scm_char_set_graphic

All characters which would put ink on the paper.

[Scheme Variable] char-set:printing
[C Variable] scm_char_set_printing
     The union of char-set:graphic and char-set:whitespace.
[Scheme Variable] char-set:whitespace
[C Variable] scm_char_set_whitespace

All whitespace characters.

[Scheme Variable] char-set:blank

[C Variable] scm_char_set_blank

All horizontal whitespace characters, which notably includes #\space and #\tab.

[Scheme Variable] char-set:iso-control
[C Variable] scm_char_set_iso_control

The ISO control characters are the C0 control characters (U+0000 to U+001F), delete (U+007F), and the C1 control characters (U+0080 to U+009F).

[Scheme Variable] char-set:punctuation
[C Variable] scm_char_set_punctuation

All punctuation characters, such as the characters !"#%&'()*,-./:;?@[\\]_{}

[Scheme Variable] char-set:symbol

[C Variable] scm_char_set_symbol

All symbol characters, such as the characters $+<=>^`|~.

[Scheme Variable] char-set:hex-digit
[C Variable] scm_char_set_hex_digit
     The hexadecimal digits 0123456789abcdefABCDEF.

[Scheme Variable] char-set:ascii

[C Variable] scm_char_set_ascii

All ASCII characters.

[Scheme Variable] char-set:empty

[C Variable] scm_char_set_empty

The empty character set.

[Scheme Variable] char-set:designated
[C Variable] scm_char_set_designated

This character set contains all designated code points. This includes all the code points to which Unicode has assigned a character or other meaning.

[Scheme Variable] char-set:full

[C Variable] scm_char_set_full

This character set contains all possible code points. This includes both designated and reserved code points.

6.6.5 Strings

Strings are fixed-length sequences of characters. They can be created by calling constructor procedures, but they can also literally get entered at the REPL or in Scheme source files.

Strings always carry the information about how many characters they are composed of with them, so there is no special end-of-string character, like in C. That means that Scheme strings can contain any character, even the ‘#\nul’ character ‘\0’.

To use strings efficiently, you need to know a bit about how Guile implements them. In Guile, a string consists of two parts, a head and the actual memory where the characters are stored. When a string (or a substring of it) is copied, only a new head gets created, the memory is usually not copied. The two heads start out pointing to the same memory.

When one of these two strings is modified, as with string-set!, their common memory does get copied so that each string has its own memory and modifying one does not accidentally modify the other as well. Thus, Guile’s strings are ‘copy on write’; the actual copying of their memory is delayed until one string is written to.

This implementation makes functions like substring very efficient in the common case that no modifications are done to the involved strings.

If you do know that your strings are getting modified right away, you can use substring/copy instead of substring. This function performs the copy immediately at the time of creation. This is more efficient, especially in a multi-threaded program. Also, substring/copy can avoid the problem that a short substring holds on to the memory of a very large original string that could otherwise be recycled.

If you want to avoid the copy altogether, so that modifications of one string show up in the other, you can use substring/shared. The strings created by this procedure are called mutation sharing substrings since the substring and the original string share modifications to each other.

If you want to prevent modifications, use substring/read-only.

Guile provides all procedures of SRFI-13 and a few more.

6.6.5.1 String Read Syntax

The read syntax for strings is an arbitrarily long sequence of characters enclosed in double quotes (").

Backslash is an escape character and can be used to insert the following special characters. \" and \\ are R5RS standard, \| is R7RS standard, the next seven are R6RS standard — notice they follow C syntax — and the remaining four are Guile extensions.

\\ Backslash character.

\" Double quote character (an unescaped " is otherwise the end of the string).

\| Vertical bar character.

\a Bell character (ASCII 7).

\f Formfeed character (ASCII 12).

\n Newline character (ASCII 10).

\r Carriage return character (ASCII 13).

\t Tab character (ASCII 9).

\v Vertical tab character (ASCII 11).

\b Backspace character (ASCII 8).

\0 NUL character (ASCII 0).

\ followed by newline (ASCII 10)

Nothing. This way if \ is the last character in a line, the string will continue with the first character from the next line, without a line break. If the hungry-eol-escapes reader option is enabled, which is not the case by default, leading whitespace on the next line is discarded.

               "foo\
                 bar"
               ⇒ "foo  bar"
               (read-enable 'hungry-eol-escapes)
               "foo\
                 bar"
               ⇒ "foobar"

\xHH Character code given by two hexadecimal digits. For example \x7f for an ASCII DEL (127).

\uHHHH Character code given by four hexadecimal digits. For example \u0100 for a capital A with macron (U+0100).

\UHHHHHH Character code given by six hexadecimal digits. For example \U010402.

The following are examples of string literals:

     "foo"
     "bar plonk"
     "Hello World"
     "\"Hi\", he said."

The three escape sequences \xHH, \uHHHH and \UHHHHHH were chosen to not break compatibility with code written for previous versions of Guile. The R6RS specification suggests a different, incompatible syntax for hex escapes: \xHHHH; – a character code followed by one to eight hexadecimal digits terminated with a semicolon. If this escape format is desired instead, it can be enabled with the reader option r6rs-hex-escapes.

     (read-enable 'r6rs-hex-escapes)

For more on reader options, See Section 6.17.2 [Scheme Read], page 360.

6.6.5.2 String Predicates

The following procedures can be used to check whether a given string fulfills some specified property.

[Scheme Procedure] string? obj [C Function] scm_string_p (obj)

Return #t if obj is a string, else #f.

[C Function] int scm_is_string (SCM obj)

Returns 1 if obj is a string, 0 otherwise.

[Scheme Procedure] string-null? str

[C Function] scm_string_null_p (str)

Return #t if str’s length is zero, and #f otherwise.

     (string-null? "") ⇒ #t
     y  ⇒ "foo"
     (string-null? y)  ⇒ #f

[Scheme Procedure] string-any char pred s [start [end]] [C Function] scm_string_any (char pred, s, start, end)

Check if char pred is true for any character in string s. char pred can be a character to check for any equal to that, or a character set (see Section 6.6.4 [Character Sets], page 133) to check for any in that set, or a predicate procedure to call. For a procedure, calls (char_pred c) are made successively on the characters from start to end. If char pred returns true (ie. non-#f), string-any stops and that return value is the return from string-any. The call on the last character (ie. at end − 1), if that point is reached, is a tail call. If there are no characters in s (ie. start equals end) then the return is #f.

[Scheme Procedure] string-every char pred s [start [end]] [C Function] scm_string_every (char pred, s, start, end)

Check if char pred is true for every character in string s. char pred can be a character to check for every character equal to that, or a character set (see Section 6.6.4 [Character Sets], page 133) to check for every character being in that set, or a predicate procedure to call. For a procedure, calls (char_pred c) are made successively on the characters from start to end. If char pred returns #f, string-every stops and returns #f. The call on the last character (ie. at end − 1), if that point is reached, is a tail call and the return from that call is the return from string-every. If there are no characters in s (ie. start equals end) then the return is #t.

6.6.5.3 String Constructors

The string constructor procedures create new string objects, possibly initializing them with some specified character data. See also See Section 6.6.5.5 [String Selection], page 145, for ways to create strings from existing strings.

[Scheme Procedure] string char…

Return a newly allocated string made from the given character arguments.

     (string #\x #\y #\z) ⇒ "xyz"
     (string)  ⇒ ""

[Scheme Procedure] list->string lst [C Function] scm_string (lst)

Return a newly allocated string made from a list of characters.

     (list->string '(#\a #\b #\c)) ⇒ "abc"
[Scheme Procedure] reverse-list->string lst
[C Function] scm_reverse_list_to_string (lst)

Return a newly allocated string made from a list of characters, in reverse order.

     (reverse-list->string '(#\a #\B #\c)) ⇒ "cBa"

[Scheme Procedure] make-string k [chr] [C Function] scm_make_string (k, chr)

Return a newly allocated string of length k. If chr is given, then all elements of the string are initialized to chr, otherwise the contents of the string are unspecified.

[C Function] SCM scm_c_make_string (size t len, SCM chr)

Like scm_make_string, but expects the length as a size_t.

[Scheme Procedure] string-tabulate proc len [C Function] scm_string_tabulate (proc, len)

proc is an integer->char procedure. Construct a string of size len by applying proc to each index to produce the corresponding string element. The order in which proc is applied to the indices is not specified.

[Scheme Procedure] string-join ls [delimiter [grammar]] [C Function] scm_string_join (ls, delimiter, grammar)

Append the string in the string list ls, using the string delimiter as a delimiter between the elements of ls. grammar is a symbol which specifies how the delimiter is placed between the strings, and defaults to the symbol infix.

infix Insert the separator between list elements. An empty string will produce an empty list.

strict-infix

Like infix, but will raise an error if given the empty list.

suffix Insert the separator after every list element.

prefix Insert the separator before each list element.

6.6.5.4 List/String conversion

When processing strings, it is often convenient to first convert them into a list representation by using the procedure string->list, work with the resulting list, and then convert it back into a string. These procedures are useful for similar tasks.

[Scheme Procedure] string->list str [start [end]] [C Function] scm_substring_to_list (str, start, end)

[C Function] scm_string_to_list (str)

Convert the string str into a list of characters.

[Scheme Procedure] string-split str char pred [C Function] scm_string_split (str, char pred)

Split the string str into a list of substrings delimited by appearances of characters that

• equal char pred, if it is a character,

• satisfy the predicate char pred, if it is a procedure,

• are in the set char pred, if it is a character set.

Note that an empty substring between separator characters will result in an empty string in the result list.

     (string-split "root:x:0:0:root:/root:/bin/bash" #\:)

⇒

     ("root" "x" "0" "0" "root" "/root" "/bin/bash")
     (string-split "::" #\:)

⇒

     ("" "" "")
     (string-split "" #\:)

⇒

     ("")

6.6.5.5 String Selection

Portions of strings can be extracted by these procedures. string-ref delivers individual characters whereas substring can be used to extract substrings from longer strings.

[Scheme Procedure] string-length string [C Function] scm_string_length (string)

Return the number of characters in string.

[C Function] size_t scm_c_string_length (SCM str)

Return the number of characters in str as a size_t.

[Scheme Procedure] string-ref str k [C Function] scm_string_ref (str, k)

Return character k of str using zero-origin indexing. k must be a valid index of str.

[C Function] SCM scm_c_string_ref (SCM str, size t k)

Return character k of str using zero-origin indexing. k must be a valid index of str.

[Scheme Procedure] string-copy str [start [end]] [C Function] scm_substring_copy (str, start, end) [C Function] scm_string_copy (str)

Return a copy of the given string str.

The returned string shares storage with str initially, but it is copied as soon as one of the two strings is modified.

[Scheme Procedure] substring str start [end]

[C Function] scm_substring (str, start, end)

Return a new string formed from the characters of str beginning with index start (inclusive) and ending with index end (exclusive). str must be a string, start and end must be exact integers satisfying:

0 <= start <= end <= (string-length str).

The returned string shares storage with str initially, but it is copied as soon as one of the two strings is modified.

[Scheme Procedure] substring/shared str start [end] [C Function] scm_substring_shared (str, start, end)

Like substring, but the strings continue to share their storage even if they are modified. Thus, modifications to str show up in the new string, and vice versa.

[Scheme Procedure] substring/copy str start [end] [C Function] scm_substring_copy (str, start, end)

Like substring, but the storage for the new string is copied immediately.

[Scheme Procedure] substring/read-only str start [end] [C Function] scm_substring_read_only (str, start, end)

Like substring, but the resulting string can not be modified.

[C Function] SCM scm_c_substring (SCM str, size t start, size t end) [C Function] SCM scm_c_substring_shared (SCM str, size t start, size t end) [C Function] SCM scm_c_substring_copy (SCM str, size t start, size t end)

[C Function] SCM scm_c_substring_read_only (SCM str, size t start, size t end)

Like scm_substring, etc. but the bounds are given as a size_t.

[Scheme Procedure] string-take s n [C Function] scm_string_take (s, n)

Return the n first characters of s.

[Scheme Procedure] string-drop s n [C Function] scm_string_drop (s, n)

Return all but the first n characters of s.

[Scheme Procedure] string-take-right s n

[C Function] scm_string_take_right (s, n)

Return the n last characters of s.

[Scheme Procedure] string-drop-right s n

[C Function] scm_string_drop_right (s, n)

Return all but the last n characters of s.

[Scheme Procedure] string-pad s len [chr [start [end]]] [Scheme Procedure] string-pad-right s len [chr [start [end]]]

[C Function] scm_string_pad (s, len, chr, start, end) [C Function] scm_string_pad_right (s, len, chr, start, end)

Take characters start to end from the string s and either pad with chr or truncate them to give len characters. string-pad pads or truncates on the left, so for example

     (string-pad "x" 3)  ⇒ "  x"
     (string-pad "abcde" 3) ⇒ "cde"

string-pad-right pads or truncates on the right, so for example

     (string-pad-right "x" 3)  ⇒ "x  "
     (string-pad-right "abcde" 3) ⇒ "abc"

[Scheme Procedure] string-trim s [char pred [start [end]]] [Scheme Procedure] string-trim-right s [char pred [start [end]]] [Scheme Procedure] string-trim-both s [char pred [start [end]]]

[C Function] scm_string_trim (s, char pred, start, end) [C Function] scm_string_trim_right (s, char pred, start, end) [C Function] scm_string_trim_both (s, char pred, start, end)

Trim occurrences of char pred from the ends of s. string-trim trims char pred characters from the left (start) of the string, stringtrim-right trims them from the right (end) of the string, string-trim-both trims from both ends. char pred can be a character, a character set, or a predicate procedure to call on each character. If char pred is not given the default is whitespace as per charset:whitespace (see Section 6.6.4.6 [Standard Character Sets], page 139).

     (string-trim " x ")  ⇒ "x "
     (string-trim-right "banana" #\a) ⇒ "banan"
     (string-trim-both ".,xy:;" char-set:punctuation)
                       ⇒ "xy"
     (string-trim-both "xyzzy" (lambda (c)
                                  (or (eqv? c #\x)
                                      (eqv? c #\y))))
                       ⇒ "zz"

6.6.5.6 String Modification

These procedures are for modifying strings in-place. This means that the result of the operation is not a new string; instead, the original string’s memory representation is modified.

[Scheme Procedure] string-set! str k chr [C Function] scm_string_set_x (str, k, chr)

Store chr in element k of str and return an unspecified value. k must be a valid index of str.

[C Function] void scm_c_string_set_x (SCM str, size t k, SCM chr)

Like scm_string_set_x, but the index is given as a size_t.

[Scheme Procedure] string-fill! str chr [start [end]] [C Function] scm_substring_fill_x (str, chr, start, end) [C Function] scm_string_fill_x (str, chr)

Stores chr in every element of the given str and returns an unspecified value.

[Scheme Procedure] substring-fill! str start end fill [C Function] scm_substring_fill_x (str, start, end, fill)

Change every character in str between start and end to fill.

     (define y (string-copy "abcdefg"))
     (substring-fill! y 1 3 #\r)
     y
     ⇒ "arrdefg"

[Scheme Procedure] substring-move! str1 start1 end1 str2 start2 [C Function] scm_substring_move_x (str1, start1, end1, str2, start2)

Copy the substring of str1 bounded by start1 and end1 into str2 beginning at position start2. str1 and str2 can be the same string.

[Scheme Procedure] string-copy! target tstart s [start [end]] [C Function] scm_string_copy_x (target, tstart, s, start, end)

Copy the sequence of characters from index range [start, end) in string s to string target, beginning at index tstart. The characters are copied left-to-right or right-to-left as needed – the copy is guaranteed to work, even if target and s are the same string. It is an error if the copy operation runs off the end of the target string.

6.6.5.7 String Comparison

The procedures in this section are similar to the character ordering predicates (see Section 6.6.3 [Characters], page 129), but are defined on character sequences.

The first set is specified in R5RS and has names that end in ?. The second set is specified in SRFI-13 and the names have not ending ?.

The predicates ending in -ci ignore the character case when comparing strings. For now, case-insensitive comparison is done using the R5RS rules, where every lower-case character that has a single character upper-case form is converted to uppercase before comparison. See See Section 6.24.2 [Text Collation], page 437, for locale-dependent string comparison.

[Scheme Procedure] string=? s1 s2 s3 …

Lexicographic equality predicate; return #t if all strings are the same length and contain the same characters in the same positions, otherwise return #f. The procedure string-ci=? treats upper and lower case letters as though they were the same character, but string=? treats upper and lower case as distinct characters.

[Scheme Procedure] string<? s1 s2 s3 …

Lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically less than str i+1.

[Scheme Procedure] string<=? s1 s2 s3 …

Lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically less than or equal to str i+1.

[Scheme Procedure] string>? s1 s2 s3 …

Lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically greater than str i+1.

[Scheme Procedure] string>=? s1 s2 s3 …

Lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically greater than or equal to str i+1.

[Scheme Procedure] string-ci=? s1 s2 s3 …

Case-insensitive string equality predicate; return #t if all strings are the same length and their component characters match (ignoring case) at each position; otherwise return #f.

[Scheme Procedure] string-ci<? s1 s2 s3 …

Case insensitive lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically less than str i+1 regardless of case.

[Scheme Procedure] string-ci<=? s1 s2 s3 …

Case insensitive lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically less than or equal to str i+1 regardless of case.

[Scheme Procedure] string-ci>? s1 s2 s3 …

Case insensitive lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically greater than str i+1 regardless of case.

[Scheme Procedure] string-ci>=? s1 s2 s3 …

Case insensitive lexicographic ordering predicate; return #t if, for every pair of consecutive string arguments str i and str i+1, str i is lexicographically greater than or equal to str i+1 regardless of case.

[Scheme Procedure] string-compare s1 s2 proc lt proc eq proc gt [start1 [end1

[start2 [end2]]]]

[C Function] scm_string_compare (s1, s2, proc lt, proc eq, proc gt, start1, end1,

start2, end2) Apply proc lt, proc eq, proc gt to the mismatch index, depending upon whether s1 is less than, equal to, or greater than s2. The mismatch index is the largest index i such that for every 0 <= j < i, s1[j] = s2[j] – that is, i is the first position that does not match.

[Scheme Procedure] string-compare-ci s1 s2 proc lt proc eq proc gt [start1 [end1

[start2 [end2]]]]

[C Function] scm_string_compare_ci (s1, s2, proc lt, proc eq, proc gt, start1,

end1, start2, end2) Apply proc lt, proc eq, proc gt to the mismatch index, depending upon whether s1 is less than, equal to, or greater than s2. The mismatch index is the largest index i such that for every 0 <= j < i, s1[j] = s2[j] – that is, i is the first position where the lowercased letters do not match.

[Scheme Procedure] string= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_eq (s1, s2, start1, end1, start2, end2)

Return #f if s1 and s2 are not equal, a true value otherwise.

[Scheme Procedure] string<> s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_neq (s1, s2, start1, end1, start2, end2)

Return #f if s1 and s2 are equal, a true value otherwise.

[Scheme Procedure] string< s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_lt (s1, s2, start1, end1, start2, end2)

Return #f if s1 is greater or equal to s2, a true value otherwise.

[Scheme Procedure] string> s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_gt (s1, s2, start1, end1, start2, end2)

Return #f if s1 is less or equal to s2, a true value otherwise.

[Scheme Procedure] string<= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_le (s1, s2, start1, end1, start2, end2)

Return #f if s1 is greater to s2, a true value otherwise.

[Scheme Procedure] string>= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ge (s1, s2, start1, end1, start2, end2)

Return #f if s1 is less to s2, a true value otherwise.

[Scheme Procedure] string-ci= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_eq (s1, s2, start1, end1, start2, end2)

Return #f if s1 and s2 are not equal, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-ci<> s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_neq (s1, s2, start1, end1, start2, end2)

Return #f if s1 and s2 are equal, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-ci< s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_lt (s1, s2, start1, end1, start2, end2)

Return #f if s1 is greater or equal to s2, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-ci> s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_gt (s1, s2, start1, end1, start2, end2)

Return #f if s1 is less or equal to s2, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-ci<= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_le (s1, s2, start1, end1, start2, end2)

Return #f if s1 is greater to s2, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-ci>= s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_ci_ge (s1, s2, start1, end1, start2, end2)

Return #f if s1 is less to s2, a true value otherwise. The character comparison is done case-insensitively.

[Scheme Procedure] string-hash s [bound [start [end]]]

[C Function] scm_substring_hash (s, bound, start, end)

Compute a hash value for s. The optional argument bound is a non-negative exact integer specifying the range of the hash function. A positive value restricts the return value to the range [0,bound).

[Scheme Procedure] string-hash-ci s [bound [start [end]]] [C Function] scm_substring_hash_ci (s, bound, start, end)

Compute a hash value for s. The optional argument bound is a non-negative exact integer specifying the range of the hash function. A positive value restricts the return value to the range [0,bound).

Because the same visual appearance of an abstract Unicode character can be obtained via multiple sequences of Unicode characters, even the case-insensitive string comparison functions described above may return #f when presented with strings containing different representations of the same character. For example, the Unicode character “LATIN SMALL LETTER S WITH DOT BELOW AND DOT ABOVE” can be represented with a single character (U+1E69) or by the character “LATIN SMALL LETTER S” (U+0073) followed by the combining marks “COMBINING DOT BELOW” (U+0323) and “COMBINING DOT ABOVE” (U+0307).

For this reason, it is often desirable to ensure that the strings to be compared are using a mutually consistent representation for every character. The Unicode standard defines two methods of normalizing the contents of strings: Decomposition, which breaks composite characters into a set of constituent characters with an ordering defined by the Unicode Standard; and composition, which performs the converse.

There are two decomposition operations. “Canonical decomposition” produces character sequences that share the same visual appearance as the original characters, while “compatibility decomposition” produces ones whose visual appearances may differ from the originals but which represent the same abstract character.

These operations are encapsulated in the following set of normalization forms:

NFD Characters are decomposed to their canonical forms.

NFKD Characters are decomposed to their compatibility forms.

NFC Characters are decomposed to their canonical forms, then composed.

NFKC Characters are decomposed to their compatibility forms, then composed.

The functions below put their arguments into one of the forms described above.

[Scheme Procedure] string-normalize-nfd s
[C Function] scm_string_normalize_nfd (s)

Return the NFD normalized form of s.

[Scheme Procedure] string-normalize-nfkd s
[C Function] scm_string_normalize_nfkd (s)

Return the NFKD normalized form of s.

[Scheme Procedure] string-normalize-nfc s
[C Function] scm_string_normalize_nfc (s)

Return the NFC normalized form of s.

[Scheme Procedure] string-normalize-nfkc s
[C Function] scm_string_normalize_nfkc (s)

Return the NFKC normalized form of s.

6.6.5.8 String Searching

[Scheme Procedure] string-index s char pred [start [end]] [C Function] scm_string_index (s, char pred, start, end)

Search through the string s from left to right, returning the index of the first occurrence of a character which

• equals char pred, if it is character,

• satisfies the predicate char pred, if it is a procedure,

• is in the set char pred, if it is a character set.

Return #f if no match is found.

[Scheme Procedure] string-rindex s char pred [start [end]]

[C Function] scm_string_rindex (s, char pred, start, end)

Search through the string s from right to left, returning the index of the last occurrence of a character which

• equals char pred, if it is character,

• satisfies the predicate char pred, if it is a procedure,

• is in the set if char pred is a character set.

Return #f if no match is found.

[Scheme Procedure] string-prefix-length s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_prefix_length (s1, s2, start1, end1, start2, end2)

Return the length of the longest common prefix of the two strings.

[Scheme Procedure] string-prefix-length-ci s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_prefix_length_ci (s1, s2, start1, end1, start2, end2)

Return the length of the longest common prefix of the two strings, ignoring character case.

[Scheme Procedure] string-suffix-length s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_suffix_length (s1, s2, start1, end1, start2, end2)

Return the length of the longest common suffix of the two strings.

[Scheme Procedure] string-suffix-length-ci s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_suffix_length_ci (s1, s2, start1, end1, start2, end2)

Return the length of the longest common suffix of the two strings, ignoring character case.

[Scheme Procedure] string-prefix? s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_prefix_p (s1, s2, start1, end1, start2, end2)

Is s1 a prefix of s2?

[Scheme Procedure] string-prefix-ci? s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_prefix_ci_p (s1, s2, start1, end1, start2, end2)

Is s1 a prefix of s2, ignoring character case?

[Scheme Procedure] string-suffix? s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_suffix_p (s1, s2, start1, end1, start2, end2)

Is s1 a suffix of s2?

[Scheme Procedure] string-suffix-ci? s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_suffix_ci_p (s1, s2, start1, end1, start2, end2)

Is s1 a suffix of s2, ignoring character case?

[Scheme Procedure] string-index-right s char pred [start [end]] [C Function] scm_string_index_right (s, char pred, start, end)

Search through the string s from right to left, returning the index of the last occurrence of a character which

• equals char pred, if it is character,

• satisfies the predicate char pred, if it is a procedure,

• is in the set if char pred is a character set.

Return #f if no match is found.

[Scheme Procedure] string-skip s char pred [start [end]] [C Function] scm_string_skip (s, char pred, start, end)

Search through the string s from left to right, returning the index of the first occurrence of a character which

• does not equal char pred, if it is character,

• does not satisfy the predicate char pred, if it is a procedure,

• is not in the set if char pred is a character set.

[Scheme Procedure] string-skip-right s char pred [start [end]]

[C Function] scm_string_skip_right (s, char pred, start, end)

Search through the string s from right to left, returning the index of the last occurrence of a character which

• does not equal char pred, if it is character,

• does not satisfy the predicate char pred, if it is a procedure,

• is not in the set if char pred is a character set.

[Scheme Procedure] string-count s char pred [start [end]] [C Function] scm_string_count (s, char pred, start, end)

Return the count of the number of characters in the string s which • equals char pred, if it is character,

• satisfies the predicate char pred, if it is a procedure.

• is in the set char pred, if it is a character set.

[Scheme Procedure] string-contains s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_contains (s1, s2, start1, end1, start2, end2)

Does string s1 contain string s2? Return the index in s1 where s2 occurs as a substring, or false. The optional start/end indices restrict the operation to the indicated substrings.

[Scheme Procedure] string-contains-ci s1 s2 [start1 [end1 [start2 [end2]]]]

[C Function] scm_string_contains_ci (s1, s2, start1, end1, start2, end2)

Does string s1 contain string s2? Return the index in s1 where s2 occurs as a substring, or false. The optional start/end indices restrict the operation to the indicated substrings. Character comparison is done case-insensitively.

6.6.5.9 Alphabetic Case Mapping

These are procedures for mapping strings to their upper- or lower-case equivalents, respectively, or for capitalizing strings.

They use the basic case mapping rules for Unicode characters. No special language or context rules are considered. The resulting strings are guaranteed to be the same length as the input strings.

See Section 6.24.3 [Character Case Mapping], page 438, for locale-dependent case conversions.

[Scheme Procedure] string-upcase str [start [end]] [C Function] scm_substring_upcase (str, start, end)

[C Function] scm_string_upcase (str)

Upcase every character in str.

[Scheme Procedure] string-upcase! str [start [end]] [C Function] scm_substring_upcase_x (str, start, end)

[C Function] scm_string_upcase_x (str)

Destructively upcase every character in str.

     (string-upcase! y)
     ⇒ "ARRDEFG"
     y
     ⇒ "ARRDEFG"

[Scheme Procedure] string-downcase str [start [end]] [C Function] scm_substring_downcase (str, start, end)

[C Function] scm_string_downcase (str)

Downcase every character in str.

[Scheme Procedure] string-downcase! str [start [end]] [C Function] scm_substring_downcase_x (str, start, end)

[C Function] scm_string_downcase_x (str)

Destructively downcase every character in str.

     y
     ⇒ "ARRDEFG"
     (string-downcase! y)
     ⇒ "arrdefg"
     y
     ⇒ "arrdefg"

[Scheme Procedure] string-capitalize str

[C Function] scm_string_capitalize (str)

Return a freshly allocated string with the characters in str, where the first character of every word is capitalized.

[Scheme Procedure] string-capitalize! str

[C Function] scm_string_capitalize_x (str)

Upcase the first character of every word in str destructively and return str.

     y  ⇒ "hello world"
     (string-capitalize! y) ⇒ "Hello World"
     y  ⇒ "Hello World"

[Scheme Procedure] string-titlecase str [start [end]] [C Function] scm_string_titlecase (str, start, end)

Titlecase every first character in a word in str.

[Scheme Procedure] string-titlecase! str [start [end]] [C Function] scm_string_titlecase_x (str, start, end)

Destructively titlecase every first character in a word in str.

6.6.5.10 Reversing and Appending Strings

[Scheme Procedure] string-reverse str [start [end]] [C Function] scm_string_reverse (str, start, end)

Reverse the string str. The optional arguments start and end delimit the region of str to operate on.

[Scheme Procedure] string-reverse! str [start [end]] [C Function] scm_string_reverse_x (str, start, end)

Reverse the string str in-place. The optional arguments start and end delimit the region of str to operate on. The return value is unspecified.

[Scheme Procedure] string-append arg …

[C Function] scm_string_append (args)

Return a newly allocated string whose characters form the concatenation of the given strings, arg . . ..

     (let ((h "hello "))
       (string-append h "world"))
     ⇒ "hello world"

[Scheme Procedure] string-append/shared arg …

[C Function] scm_string_append_shared (args)

Like string-append, but the result may share memory with the argument strings.

[Scheme Procedure] string-concatenate ls

[C Function] scm_string_concatenate (ls)

Append the elements (which must be strings) of ls together into a single string. Guaranteed to return a freshly allocated string.

[Scheme Procedure] string-concatenate-reverse ls [final string [end]]

[C Function] scm_string_concatenate_reverse (ls, final string, end)

Without optional arguments, this procedure is equivalent to

     (string-concatenate (reverse ls))

If the optional argument final string is specified, it is consed onto the beginning to ls before performing the list-reverse and string-concatenate operations. If end is given, only the characters of final string up to index end are used. Guaranteed to return a freshly allocated string.

[Scheme Procedure] string-concatenate/shared ls
[C Function] scm_string_concatenate_shared (ls)

Like string-concatenate, but the result may share memory with the strings in the list ls.

[Scheme Procedure] string-concatenate-reverse/shared ls [final string [end]]

[C Function] scm_string_concatenate_reverse_shared (ls, final string, end)

Like string-concatenate-reverse, but the result may share memory with the strings in the ls arguments.

6.6.5.11 Mapping, Folding, and Unfolding

[Scheme Procedure] string-map proc s [start [end]]

[C Function] scm_string_map (proc, s, start, end)

proc is a char->char procedure, it is mapped over s. The order in which the procedure is applied to the string elements is not specified.

[Scheme Procedure] string-map! proc s [start [end]]

[C Function] scm_string_map_x (proc, s, start, end)

proc is a char->char procedure, it is mapped over s. The order in which the procedure is applied to the string elements is not specified. The string s is modified in-place, the return value is not specified.

[Scheme Procedure] string-for-each proc s [start [end]] [C Function] scm_string_for_each (proc, s, start, end)

proc is mapped over s in left-to-right order. The return value is not specified.

[Scheme Procedure] string-for-each-index proc s [start [end]] [C Function] scm_string_for_each_index (proc, s, start, end)

Call (proc i) for each index i in s, from left to right.

For example, to change characters to alternately upper and lower case,

     (define str (string-copy "studly"))
     (string-for-each-index
         (lambda (i)
           (string-set! str i
             ((if (even? i) char-upcase char-downcase)
              (string-ref str i))))
         str)
     str ⇒ "StUdLy"

[Scheme Procedure] string-fold kons knil s [start [end]] [C Function] scm_string_fold (kons, knil, s, start, end)

Fold kons over the characters of s, with knil as the terminating element, from left to right. kons must expect two arguments: The actual character and the last result of kons’ application.

[Scheme Procedure] string-fold-right kons knil s [start [end]]

[C Function] scm_string_fold_right (kons, knil, s, start, end)

Fold kons over the characters of s, with knil as the terminating element, from right to left. kons must expect two arguments: The actual character and the last result of kons’ application.

[Scheme Procedure] string-unfold p f g seed [base [make final]] [C Function] scm_string_unfold (p, f, g, seed, base, make final)

• g is used to generate a series of seed values from the initial seed: seed, (g seed), (g^2 seed), (g^3 seed), …

• p tells us when to stop – when it returns true when applied to one of these seed values.

• f maps each seed value to the corresponding character in the result string. These chars are assembled into the string in a left-to-right order.

• base is the optional initial/leftmost portion of the constructed string; it default to the empty string.

• make final is applied to the terminal seed value (on which p returns true) to produce the final/rightmost portion of the constructed string. The default is nothing extra.

[Scheme Procedure] string-unfold-right p f g seed [base [make final]]

[C Function] scm_string_unfold_right (p, f, g, seed, base, make final)

• g is used to generate a series of seed values from the initial seed: seed, (g seed), (g^2 seed), (g^3 seed), …

• p tells us when to stop – when it returns true when applied to one of these seed values.

• f maps each seed value to the corresponding character in the result string. These chars are assembled into the string in a right-to-left order.

• base is the optional initial/rightmost portion of the constructed string; it default to the empty string.

• make final is applied to the terminal seed value (on which p returns true) to produce the final/leftmost portion of the constructed string. It defaults to (lambda

(x) ).

6.6.5.12 Miscellaneous String Operations

[Scheme Procedure] xsubstring s from [to [start [end]]]

[C Function] scm_xsubstring (s, from, to, start, end)

This is the extended substring procedure that implements replicated copying of a substring of some string.

s is a string, start and end are optional arguments that demarcate a substring of s, defaulting to 0 and the length of s. Replicate this substring up and down index space, in both the positive and negative directions. xsubstring returns the substring of this string beginning at index from, and ending at to, which defaults to from + (end - start).

[Scheme Procedure] string-xcopy! target tstart s sfrom [sto [start [end]]] [C Function] scm_string_xcopy_x (target, tstart, s, sfrom, sto, start, end)

Exactly the same as xsubstring, but the extracted text is written into the string target starting at index tstart. The operation is not defined if (eq? target s) or these arguments share storage – you cannot copy a string on top of itself.

[Scheme Procedure] string-replace s1 s2 [start1 [end1 [start2 [end2]]]] [C Function] scm_string_replace (s1, s2, start1, end1, start2, end2)

Return the string s1, but with the characters start1 … end1 replaced by the characters start2 … end2 from s2.

[Scheme Procedure] string-tokenize s [token set [start [end]]]

[C Function] scm_string_tokenize (s, token set, start, end)

Split the string s into a list of substrings, where each substring is a maximal nonempty contiguous sequence of characters from the character set token set, which defaults to char-set:graphic. If start or end indices are provided, they restrict string-tokenize to operating on the indicated substring of s.

[Scheme Procedure] string-filter char pred s [start [end]]

[C Function] scm_string_filter (char pred, s, start, end)

Filter the string s, retaining only those characters which satisfy char pred.

If char pred is a procedure, it is applied to each character as a predicate, if it is a character, it is tested for equality and if it is a character set, it is tested for membership.

[Scheme Procedure] string-delete char pred s [start [end]]

[C Function] scm_string_delete (char pred, s, start, end)

Delete characters satisfying char pred from s.

If char pred is a procedure, it is applied to each character as a predicate, if it is a character, it is tested for equality and if it is a character set, it is tested for membership.

6.6.5.13 Representing Strings as Bytes

Out in the cold world outside of Guile, not all strings are treated in the same way. Out there there are only bytes, and there are many ways of representing a strings (sequences of characters) as binary data (sequences of bytes).

As a user, usually you don’t have to think about this very much. When you type on your keyboard, your system encodes your keystrokes as bytes according to the locale that you have configured on your computer. Guile uses the locale to decode those bytes back into characters – hopefully the same characters that you typed in.

All is not so clear when dealing with a system with multiple users, such as a web server. Your web server might get a request from one user for data encoded in the ISO-8859-1 character set, and then another request from a different user for UTF-8 data.

Guile provides an iconv module for converting between strings and sequences of bytes. See Section 6.6.6 [Bytevectors], page 163, for more on how Guile represents raw byte sequences. This module gets its name from the common unix command of the same name.

Note that often it is sufficient to just read and write strings from ports instead of using these functions. To do this, specify the port encoding using set-port-encoding!. See Section 6.14.1 [Ports], page 316, for more on ports and character encodings.

Unlike the rest of the procedures in this section, you have to load the iconv module before having access to these procedures:

(use-modules (ice-9 iconv))

[Scheme Procedure] string->bytevector string encoding [conversion-strategy] Encode string as a sequence of bytes. The string will be encoded in the character set specified by the encoding string. If the string has characters that cannot be represented in the encoding, by default this procedure raises an encoding-error. Pass a conversion-strategy argument to specify other behaviors. The return value is a bytevector. See Section 6.6.6 [Bytevectors], page 163, for more on bytevectors. See Section 6.14.1 [Ports], page 316, for more on character encodings and conversion strategies.

[Scheme Procedure] bytevector->string bytevector encoding [conversion-strategy] Decode bytevector into a string. The bytes will be decoded from the character set by the encoding string. If the bytes do not form a valid encoding, by default this procedure raises an decoding-error. As with string->bytevector, pass the optional conversion-strategy argument to modify this behavior. See Section 6.14.1 [Ports], page 316, for more on character encodings and conversion strategies.

[Scheme Procedure] call-with-output-encoded-string encoding proc

[conversion-strategy] Like call-with-output-string, but instead of returning a string, returns a encoding of the string according to encoding, as a bytevector. This procedure can be more efficient than collecting a string and then converting it via string->bytevector.

6.6.5.14 Conversion to/from C

When creating a Scheme string from a C string or when converting a Scheme string to a C string, the concept of character encoding becomes important.

In C, a string is just a sequence of bytes, and the character encoding describes the relation between these bytes and the actual characters that make up the string. For Scheme strings, character encoding is not an issue (most of the time), since in Scheme you usually treat strings as character sequences, not byte sequences.

Converting to C and converting from C each have their own challenges. When converting from C to Scheme, it is important that the sequence of bytes in the

C string be valid with respect to its encoding. ASCII strings, for example, can’t have any bytes greater than 127. An ASCII byte greater than 127 is considered ill-formed and cannot be converted into a Scheme character.

Problems can occur in the reverse operation as well. Not all character encodings can hold all possible Scheme characters. Some encodings, like ASCII for example, can only describe a small subset of all possible characters. So, when converting to C, one must first decide what to do with Scheme characters that can’t be represented in the C string.

Converting a Scheme string to a C string will often allocate fresh memory to hold the result. You must take care that this memory is properly freed eventually. In many cases, this can be achieved by using scm_dynwind_free inside an appropriate dynwind context, See Section 6.13.10 [Dynamic Wind], page 309.

[C Function] SCM scm_from_locale_string (const char *str) [C Function] SCM scm_from_locale_stringn (const char *str, size t len)

Creates a new Scheme string that has the same contents as str when interpreted in the character encoding of the current locale. For scm_from_locale_string, str must be null-terminated. For scm_from_locale_stringn, len specifies the length of str in bytes, and str does not need to be null-terminated. If len is (size_t)-1, then str does need to be nullterminated and the real length will be found with strlen. If the C string is ill-formed, an error will be raised. Note that these functions should not be used to convert C string constants, because there is no guarantee that the current locale will match that of the execution character set, used for string and character constants. Most modern C compilers use UTF-8 by default, so to convert C string constants we recommend scm_from_utf8_string.

[C Function] SCM scm_take_locale_string (char *str)

[C Function] SCM scm_take_locale_stringn (char *str, size t len)

Like scm_from_locale_string and scm_from_locale_stringn, respectively, but

also frees str with free eventually. Thus, you can use this function when you would free str anyway immediately after creating the Scheme string. In certain cases, Guile can then use str directly as its internal representation.

[C Function] char * scm_to_locale_string (SCM str)

[C Function] char * scm_to_locale_stringn (SCM str, size t *lenp)

Returns a C string with the same contents as str in the character encoding of the current locale. The C string must be freed with free eventually, maybe by using scm_dynwind_free, See Section 6.13.10 [Dynamic Wind], page 309. For scm_to_locale_string, the returned string is null-terminated and an error is signalled when str contains #\nul characters. For scm_to_locale_stringn and lenp not NULL, str might contain #\nul characters and the length of the returned string in bytes is stored in *lenp. The returned string will not be null-terminated in this case. If lenp is NULL, scm_to_locale_stringn

behaves like scm_to_locale_string.

If a character in str cannot be represented in the character encoding of the current locale, the default port conversion strategy is used. See Section 6.14.1 [Ports], page 316, for more on conversion strategies. If the conversion strategy is error, an error will be raised. If it is substitute, a replacement character, such as a question mark, will be inserted in its place. If it is escape, a hex escape will be inserted in its place.

[C Function] size_t scm_to_locale_stringbuf (SCM str, char *buf, size t

max len) Puts str as a C string in the current locale encoding into the memory pointed to by buf. The buffer at buf has room for max len bytes and scm_to_local_stringbuf will never store more than that. No terminating '\0' will be stored. The return value of scm_to_locale_stringbuf is the number of bytes that are needed for all of str, regardless of whether buf was large enough to hold them. Thus, when the return value is larger than max len, only max len bytes have been stored and you probably need to try again with a larger buffer.

For most situations, string conversion should occur using the current locale, such as with the functions above. But there may be cases where one wants to convert strings from a character encoding other than the locale’s character encoding. For these cases, the lowerlevel functions scm_to_stringn and scm_from_stringn are provided. These functions should seldom be necessary if one is properly using locales.

[C Type] scm_t_string_failed_conversion_handler

This is an enumerated type that can take one of three values: SCM_

FAILED_CONVERSION_ERROR,  SCM_FAILED_CONVERSION_QUESTION_MARK,  and
SCM_FAILED_CONVERSION_ESCAPE_SEQUENCE. They are used to indicate a strategy

for handling characters that cannot be converted to or from a given character encoding. SCM_FAILED_CONVERSION_ERROR indicates that a conversion should throw an error if some characters cannot be converted. SCM_FAILED_CONVERSION_ QUESTION_MARK indicates that a conversion should replace unconvertable characters

with the question mark character. And, SCM_FAILED_CONVERSION_ESCAPE_SEQUENCE

requests that a conversion should replace an unconvertable character with an escape sequence. While all three strategies apply when converting Scheme strings to C, only

SCM_FAILED_CONVERSION_ERROR and SCM_FAILED_CONVERSION_QUESTION_MARK can

be used when converting C strings to Scheme.

[C Function] char *scm_to_stringn (SCM str, size t *lenp, const char *encoding,

scm t string failed conversion handler handler) This function returns a newly allocated C string from the Guile string str. The length of the returned string in bytes will be returned in lenp. The character encoding of the C string is passed as the ASCII, null-terminated C string encoding. The handler parameter gives a strategy for dealing with characters that cannot be converted into encoding. If lenp is NULL, this function will return a null-terminated C string. It will throw an error if the string contains a null character. The Scheme interface to this function is string->bytevector, from the ice-9 iconv module. See Section 6.6.5.13 [Representing Strings as Bytes], page 158.

[C Function] SCM scm_from_stringn (const char *str, size t len, const char

*encoding, scm t string failed conversion handler handler) This function returns a scheme string from the C string str. The length in bytes of the C string is input as len. The encoding of the C string is passed as the ASCII, null-terminated C string encoding. The handler parameters suggests a strategy for dealing with unconvertable characters. The Scheme interface to this function is bytevector->string. See Section 6.6.5.13 [Representing Strings as Bytes], page 158.

The following conversion functions are provided as a convenience for the most commonly used encodings.

[C Function] SCM scm_from_latin1_string (const char *str) [C Function] SCM scm_from_utf8_string (const char *str) [C Function] SCM scm_from_utf32_string (const scm t wchar *str)

Return a scheme string from the null-terminated C string str, which is ISO-8859-1-, UTF-8-, or UTF-32-encoded. These functions should be used to convert hard-coded C string constants into Scheme strings.

[C Function] SCM scm_from_latin1_stringn (const char *str, size t len) [C Function] SCM scm_from_utf8_stringn (const char *str, size t len)

[C Function] SCM scm_from_utf32_stringn (const scm t wchar *str, size t len)

Return a scheme string from C string str, which is ISO-8859-1-, UTF-8-, or UTF-32-encoded, of length len. len is the number of bytes pointed to by str for scm_from_

latin1_stringn and scm_from_utf8_stringn; it is the number of elements (code

points) in str in the case of scm_from_utf32_stringn.

[C function] char *scm_to_latin1_stringn (SCM str, size t *lenp) [C function] char *scm_to_utf8_stringn (SCM str, size t *lenp)

[C function] scm_t_wchar *scm_to_utf32_stringn (SCM str, size t *lenp)

Return a newly allocated, ISO-8859-1-, UTF-8-, or UTF-32-encoded C string from Scheme string str. An error is thrown when str cannot be converted to the specified encoding. If lenp is NULL, the returned C string will be null terminated, and an error will be thrown if the C string would otherwise contain null characters. If lenp is not NULL, the string is not null terminated, and the length of the returned string is returned in lenp. The length returned is the number of bytes for scm_to_latin1_ stringn and scm_to_utf8_stringn; it is the number of elements (code points) for

scm_to_utf32_stringn.

6.6.5.15 String Internals

Guile stores each string in memory as a contiguous array of Unicode code points along with an associated set of attributes. If all of the code points of a string have an integer range between 0 and 255 inclusive, the code point array is stored as one byte per code point: it is stored as an ISO-8859-1 (aka Latin-1) string. If any of the code points of the string has an integer value greater that 255, the code point array is stored as four bytes per code point: it is stored as a UTF-32 string.

Conversion between the one-byte-per-code-point and four-bytes-per-code-point representations happens automatically as necessary.

No API is provided to set the internal representation of strings; however, there are pair of procedures available to query it. These are debugging procedures. Using them in production code is discouraged, since the details of Guile’s internal representation of strings may change from release to release.

[Scheme Procedure] string-bytes-per-char str
[C Function] scm_string_bytes_per_char (str)

Return the number of bytes used to encode a Unicode code point in string str. The result is one or four.

[Scheme Procedure] %string-dump str

[C Function] scm_sys_string_dump (str)

Returns an association list containing debugging information for str. The association list has the following entries.

string The string itself.

start The start index of the string into its stringbuf

length The length of the string

shared If this string is a substring, it returns its parent string. Otherwise, it returns #f

read-only

#t if the string is read-only

stringbuf-chars

A new string containing this string’s stringbuf’s characters

stringbuf-length

The number of characters in this stringbuf

stringbuf-shared

#t if this stringbuf is shared

stringbuf-wide

#t if this stringbuf’s characters are stored in a 32-bit buffer, or #f if they are stored in an 8-bit buffer

6.6.6 Bytevectors

A bytevector is a raw bit string. The (rnrs bytevectors) module provides the programming interface specified by the Revised^6 Report on the Algorithmic Language Scheme (R6RS) (http://www.r6rs.org/). It contains procedures to manipulate bytevectors and interpret their contents in a number of ways: bytevector contents can be accessed as signed or unsigned integer of various sizes and endianness, as IEEE-754 floating point numbers, or as strings. It is a useful tool to encode and decode binary data.

The R6RS (Section 4.3.4) specifies an external representation for bytevectors, whereby the octets (integers in the range 0–255) contained in the bytevector are represented as a list prefixed by #vu8:

#vu8(1 53 204)

denotes a 3-byte bytevector containing the octets 1, 53, and 204. Like string literals, booleans, etc., bytevectors are “self-quoting”, i.e., they do not need to be quoted:

#vu8(1 53 204)
⇒ #vu8(1 53 204)

Bytevectors can be used with the binary input/output primitives of the R6RS (see Section 6.14.10 [R6RS I/O Ports], page 332).

6.6.6.1 Endianness

Some of the following procedures take an endianness parameter. The endianness is defined as the order of bytes in multi-byte numbers: numbers encoded in big endian have their most significant bytes written first, whereas numbers encoded in little endian have their least significant bytes first^{1}.

Little-endian is the native endianness of the IA32 architecture and its derivatives, while big-endian is native to SPARC and PowerPC, among others. The native-endianness procedure returns the native endianness of the machine it runs on.

[Scheme Procedure] native-endianness
[C Function] scm_native_endianness ()

Return a value denoting the native endianness of the host machine.

[Scheme Macro] endianness symbol

Return an object denoting the endianness specified by symbol. If symbol is neither big nor little then an error is raised at expand-time.

[C Variable] scm_endianness_big
[C Variable] scm_endianness_little

The objects denoting big- and little-endianness, respectively.

6.6.6.2 Manipulating Bytevectors

Bytevectors can be created, copied, and analyzed with the following procedures and C functions.

[Scheme Procedure] make-bytevector len [fill] [C Function] scm_make_bytevector (len, fill)

[C Function] scm_c_make_bytevector (size t len)

Return a new bytevector of len bytes. Optionally, if fill is given, fill it with fill; fill must be in the range [-128,255].

[Scheme Procedure] bytevector? obj

[C Function] scm_bytevector_p (obj)

Return true if obj is a bytevector.

[C Function] int scm_is_bytevector (SCM obj)
Equivalent to scm_is_true (scm_bytevector_p (obj)).

[Scheme Procedure] bytevector-length bv

[C Function] scm_bytevector_length (bv)

Return the length in bytes of bytevector bv.

[C Function] size_t scm_c_bytevector_length (SCM bv)

Likewise, return the length in bytes of bytevector bv.

1 Big-endian and little-endian are the most common “endiannesses”, but others do exist. For instance,

the GNU MP library allows word order to be specified independently of byte order (see Section “Integer Import and Export” in The GNU Multiple Precision Arithmetic Library Manual).

[Scheme Procedure] bytevector=? bv1 bv2 [C Function] scm_bytevector_eq_p (bv1, bv2)

Return is bv1 equals to bv2—i.e., if they have the same length and contents.

[Scheme Procedure] bytevector-fill! bv fill

[C Function] scm_bytevector_fill_x (bv, fill)

Fill bytevector bv with fill, a byte.

[Scheme Procedure] bytevector-copy! source source-start target target-start len [C Function] scm_bytevector_copy_x (source, source start, target, target start,

len) Copy len bytes from source into target, starting reading from source-start (a positive index within source) and start writing at target-start. It is permitted for the source and target regions to overlap.

[Scheme Procedure] bytevector-copy bv

[C Function] scm_bytevector_copy (bv)

Return a newly allocated copy of bv.

[C Function] scm_t_uint8 scm_c_bytevector_ref (SCM bv, size t index)

Return the byte at index in bytevector bv.

[C Function] void scm_c_bytevector_set_x (SCM bv, size t index, scm t uint8

value) Set the byte at index in bv to value.

Low-level C macros are available. They do not perform any type-checking; as such they should be used with care.

[C Macro] size_t SCM_BYTEVECTOR_LENGTH (bv)

Return the length in bytes of bytevector bv.

[C Macro] signed char * SCM_BYTEVECTOR_CONTENTS (bv)

Return a pointer to the contents of bytevector bv.

6.6.6.3 Interpreting Bytevector Contents as Integers

The contents of a bytevector can be interpreted as a sequence of integers of any given size, sign, and endianness.

(let ((bv (make-bytevector 4)))
  (bytevector-u8-set! bv 0 #x12)
  (bytevector-u8-set! bv 1 #x34)
  (bytevector-u8-set! bv 2 #x56)
  (bytevector-u8-set! bv 3 #x78)
  (map (lambda (number)
         (number->string number 16))
       (list (bytevector-u8-ref bv 0)
             (bytevector-u16-ref bv 0 (endianness big))
             (bytevector-u32-ref bv 0 (endianness little)))))

     ⇒ ("12" "1234" "78563412")

The most generic procedures to interpret bytevector contents as integers are described below.

[Scheme Procedure] bytevector-uint-ref bv index endianness size [C Function] scm_bytevector_uint_ref (bv, index, endianness, size)

Return the size-byte long unsigned integer at index index in bv, decoded according to endianness.

[Scheme Procedure] bytevector-sint-ref bv index endianness size [C Function] scm_bytevector_sint_ref (bv, index, endianness, size)

Return the size-byte long signed integer at index index in bv, decoded according to endianness.

[Scheme Procedure] bytevector-uint-set! bv index value endianness size [C Function] scm_bytevector_uint_set_x (bv, index, value, endianness, size)

Set the size-byte long unsigned integer at index to value, encoded according to endianness.

[Scheme Procedure] bytevector-sint-set! bv index value endianness size [C Function] scm_bytevector_sint_set_x (bv, index, value, endianness, size)

Set the size-byte long signed integer at index to value, encoded according to endianness.

The following procedures are similar to the ones above, but specialized to a given integer size:

[Scheme Procedure] bytevector-u8-ref bv index [Scheme Procedure] bytevector-s8-ref bv index [Scheme Procedure] bytevector-u16-ref bv index endianness [Scheme Procedure] bytevector-s16-ref bv index endianness [Scheme Procedure] bytevector-u32-ref bv index endianness [Scheme Procedure] bytevector-s32-ref bv index endianness [Scheme Procedure] bytevector-u64-ref bv index endianness [Scheme Procedure] bytevector-s64-ref bv index endianness

[C Function] scm_bytevector_u8_ref (bv, index)
[C Function] scm_bytevector_s8_ref (bv, index)

[C Function] scm_bytevector_u16_ref (bv, index, endianness) [C Function] scm_bytevector_s16_ref (bv, index, endianness) [C Function] scm_bytevector_u32_ref (bv, index, endianness) [C Function] scm_bytevector_s32_ref (bv, index, endianness) [C Function] scm_bytevector_u64_ref (bv, index, endianness) [C Function] scm_bytevector_s64_ref (bv, index, endianness)

Return the unsigned n-bit (signed) integer (where n is 8, 16, 32 or 64) from bv at index, decoded according to endianness.

[Scheme Procedure] bytevector-u8-set! bv index value [Scheme Procedure] bytevector-s8-set! bv index value

[Scheme Procedure] bytevector-u16-set! bv index value endianness [Scheme Procedure] bytevector-s16-set! bv index value endianness [Scheme Procedure] bytevector-u32-set! bv index value endianness [Scheme Procedure] bytevector-s32-set! bv index value endianness [Scheme Procedure] bytevector-u64-set! bv index value endianness [Scheme Procedure] bytevector-s64-set! bv index value endianness

[C Function] scm_bytevector_u8_set_x (bv, index, value) [C Function] scm_bytevector_s8_set_x (bv, index, value)

[C Function] scm_bytevector_u16_set_x (bv, index, value, endianness) [C Function] scm_bytevector_s16_set_x (bv, index, value, endianness) [C Function] scm_bytevector_u32_set_x (bv, index, value, endianness) [C Function] scm_bytevector_s32_set_x (bv, index, value, endianness) [C Function] scm_bytevector_u64_set_x (bv, index, value, endianness) [C Function] scm_bytevector_s64_set_x (bv, index, value, endianness)

Store value as an n-bit (signed) integer (where n is 8, 16, 32 or 64) in bv at index, encoded according to endianness.

Finally, a variant specialized for the host’s endianness is available for each of these functions (with the exception of the u8 accessors, for obvious reasons):

[Scheme Procedure] bytevector-u16-native-ref bv index
[Scheme Procedure] bytevector-s16-native-ref bv index
[Scheme Procedure] bytevector-u32-native-ref bv index
[Scheme Procedure] bytevector-s32-native-ref bv index
[Scheme Procedure] bytevector-u64-native-ref bv index
[Scheme Procedure] bytevector-s64-native-ref bv index
[C Function] scm_bytevector_u16_native_ref (bv, index)
[C Function] scm_bytevector_s16_native_ref (bv, index)
[C Function] scm_bytevector_u32_native_ref (bv, index)
[C Function] scm_bytevector_s32_native_ref (bv, index)
[C Function] scm_bytevector_u64_native_ref (bv, index)
[C Function] scm_bytevector_s64_native_ref (bv, index)

Return the unsigned n-bit (signed) integer (where n is 8, 16, 32 or 64) from bv at index, decoded according to the host’s native endianness.

[Scheme Procedure] bytevector-u16-native-set! bv index value [Scheme Procedure] bytevector-s16-native-set! bv index value [Scheme Procedure] bytevector-u32-native-set! bv index value [Scheme Procedure] bytevector-s32-native-set! bv index value [Scheme Procedure] bytevector-u64-native-set! bv index value [Scheme Procedure] bytevector-s64-native-set! bv index value

[C Function] scm_bytevector_u16_native_set_x (bv, index, value)
[C Function] scm_bytevector_s16_native_set_x (bv, index, value)
[C Function] scm_bytevector_u32_native_set_x (bv, index, value)
[C Function] scm_bytevector_s32_native_set_x (bv, index, value)
[C Function] scm_bytevector_u64_native_set_x (bv, index, value)

[C Function] scm_bytevector_s64_native_set_x (bv, index, value)

Store value as an n-bit (signed) integer (where n is 8, 16, 32 or 64) in bv at index, encoded according to the host’s native endianness.

6.6.6.4 Converting Bytevectors to/from Integer Lists

Bytevector contents can readily be converted to/from lists of signed or unsigned integers:

(bytevector->sint-list (u8-list->bytevector (make-list 4 255))
                       (endianness little) 2)
⇒ (-1 -1)
[Scheme Procedure] bytevector->u8-list bv
[C Function] scm_bytevector_to_u8_list (bv)

Return a newly allocated list of unsigned 8-bit integers from the contents of bv.

[Scheme Procedure] u8-list->bytevector lst

[C Function] scm_u8_list_to_bytevector (lst)

Return a newly allocated bytevector consisting of the unsigned 8-bit integers listed in lst.

[Scheme Procedure] bytevector->uint-list bv endianness size [C Function] scm_bytevector_to_uint_list (bv, endianness, size)

Return a list of unsigned integers of size bytes representing the contents of bv, decoded according to endianness.

[Scheme Procedure] bytevector->sint-list bv endianness size [C Function] scm_bytevector_to_sint_list (bv, endianness, size)

Return a list of signed integers of size bytes representing the contents of bv, decoded according to endianness.

[Scheme Procedure] uint-list->bytevector lst endianness size [C Function] scm_uint_list_to_bytevector (lst, endianness, size)

Return a new bytevector containing the unsigned integers listed in lst and encoded on size bytes according to endianness.

[Scheme Procedure] sint-list->bytevector lst endianness size [C Function] scm_sint_list_to_bytevector (lst, endianness, size)

Return a new bytevector containing the signed integers listed in lst and encoded on size bytes according to endianness.

6.6.6.5 Interpreting Bytevector Contents as Floating Point Numbers

Bytevector contents can also be accessed as IEEE-754 single- or double-precision floating point numbers (respectively 32 and 64-bit long) using the procedures described here.

[Scheme Procedure] bytevector-ieee-single-ref bv index endianness [Scheme Procedure] bytevector-ieee-double-ref bv index endianness [C Function] scm_bytevector_ieee_single_ref (bv, index, endianness) [C Function] scm_bytevector_ieee_double_ref (bv, index, endianness)

Return the IEEE-754 single-precision floating point number from bv at index according to endianness.

[Scheme Procedure] bytevector-ieee-single-set! bv index value endianness [Scheme Procedure] bytevector-ieee-double-set! bv index value endianness [C Function] scm_bytevector_ieee_single_set_x (bv, index, value, endianness) [C Function] scm_bytevector_ieee_double_set_x (bv, index, value, endianness)

Store real number value in bv at index according to endianness.

Specialized procedures are also available:

[Scheme Procedure] bytevector-ieee-single-native-ref bv index
[Scheme Procedure] bytevector-ieee-double-native-ref bv index
[C Function] scm_bytevector_ieee_single_native_ref (bv, index)
[C Function] scm_bytevector_ieee_double_native_ref (bv, index)

Return the IEEE-754 single-precision floating point number from bv at index according to the host’s native endianness.

[Scheme Procedure] bytevector-ieee-single-native-set! bv index value
[Scheme Procedure] bytevector-ieee-double-native-set! bv index value
[C Function] scm_bytevector_ieee_single_native_set_x (bv, index, value)
[C Function] scm_bytevector_ieee_double_native_set_x (bv, index, value)

Store real number value in bv at index according to the host’s native endianness.

6.6.6.6 Interpreting Bytevector Contents as Unicode Strings

Bytevector contents can also be interpreted as Unicode strings encoded in one of the most commonly available encoding formats. See Section 6.6.5.13 [Representing Strings as Bytes], page 158, for a more generic interface.

     (utf8->string (u8-list->bytevector '(99 97 102 101)))
     ⇒ "cafe"
     (string->utf8 "café") ;; SMALL LATIN LETTER E WITH ACUTE ACCENT
     ⇒ #vu8(99 97 102 195 169)

[Scheme Procedure] string->utf8 str [Scheme Procedure] string->utf16 str [endianness] [Scheme Procedure] string->utf32 str [endianness]

[C Function] scm_string_to_utf8 (str)

[C Function] scm_string_to_utf16 (str, endianness) [C Function] scm_string_to_utf32 (str, endianness)

Return a newly allocated bytevector that contains the UTF-8, UTF-16, or UTF-32 (aka. UCS-4) encoding of str. For UTF-16 and UTF-32, endianness should be the symbol big or little; when omitted, it defaults to big endian.

[Scheme Procedure] utf8->string utf [Scheme Procedure] utf16->string utf [endianness] [Scheme Procedure] utf32->string utf [endianness]

[C Function] scm_utf8_to_string (utf)

[C Function] scm_utf16_to_string (utf, endianness)

[C Function] scm_utf32_to_string (utf, endianness)

Return a newly allocated string that contains from the UTF-8-, UTF-16-, or UTF-32-decoded contents of bytevector utf. For UTF-16 and UTF-32, endianness should be the symbol big or little; when omitted, it defaults to big endian.

6.6.6.7 Accessing Bytevectors with the Array API

As an extension to the R6RS, Guile allows bytevectors to be manipulated with the array procedures (see Section 6.7.5 [Arrays], page 200). When using these APIs, bytes are accessed one at a time as 8-bit unsigned integers:

     (define bv #vu8(0 1 2 3))
     (array? bv)
     ⇒ #t
     (array-rank bv)
     ⇒ 1
     (array-ref bv 2)
     ⇒ 2
     ;; Note the different argument order on array-set!.
     (array-set! bv 77 2)
     (array-ref bv 2)
     ⇒ 77
     (array-type bv)
     ⇒ vu8

6.6.6.8 Accessing Bytevectors with the SRFI-4 API

Bytevectors may also be accessed with the SRFI-4 API. See Section 7.5.5.3 [SRFI-4 and Bytevectors], page 574, for more information.

6.6.7 Symbols

Symbols in Scheme are widely used in three ways: as items of discrete data, as lookup keys for alists and hash tables, and to denote variable references.

A symbol is similar to a string in that it is defined by a sequence of characters. The sequence of characters is known as the symbol’s name. In the usual case — that is, where the symbol’s name doesn’t include any characters that could be confused with other elements of Scheme syntax — a symbol is written in a Scheme program by writing the sequence of characters that make up the name, without any quotation marks or other special syntax. For example, the symbol whose name is “multiply-by-2” is written, simply:

     multiply-by-2

Notice how this differs from a string with contents “multiply-by-2”, which is written with double quotation marks, like this:

     "multiply-by-2"

Looking beyond how they are written, symbols are different from strings in two important respects.

The first important difference is uniqueness. If the same-looking string is read twice from two different places in a program, the result is two different string objects whose contents just happen to be the same. If, on the other hand, the same-looking symbol is read twice from two different places in a program, the result is the same symbol object both times.

Given two read symbols, you can use eq? to test whether they are the same (that is, have the same name). eq? is the most efficient comparison operator in Scheme, and comparing two symbols like this is as fast as comparing, for example, two numbers. Given two strings, on the other hand, you must use equal? or string=?, which are much slower comparison operators, to determine whether the strings have the same contents.

     (define sym1 (quote hello))
     (define sym2 (quote hello))
     (eq? sym1 sym2) ⇒ #t
     (define str1 "hello")
     (define str2 "hello")
     (eq? str1 str2) ⇒ #f
     (equal? str1 str2) ⇒ #t

The second important difference is that symbols, unlike strings, are not self-evaluating. This is why we need the (quote ...)s in the example above: (quote hello) evaluates to the symbol named "hello" itself, whereas an unquoted hello is read as the symbol named "hello" and evaluated as a variable reference … about which more below (see Section 6.6.7.3 [Symbol Variables], page 173).

6.6.7.1 Symbols as Discrete Data

Numbers and symbols are similar to the extent that they both lend themselves to eq? comparison. But symbols are more descriptive than numbers, because a symbol’s name can be used directly to describe the concept for which that symbol stands.

For example, imagine that you need to represent some colours in a computer program. Using numbers, you would have to choose arbitrarily some mapping between numbers and colours, and then take care to use that mapping consistently:

     ;; 1=red, 2=green, 3=purple
     (if (eq? (colour-of car) 1)
         ...)

You can make the mapping more explicit and the code more readable by defining constants:

     (define red 1)
     (define green 2)
     (define purple 3)
     (if (eq? (colour-of car) red)
         ...)

But the simplest and clearest approach is not to use numbers at all, but symbols whose names specify the colours that they refer to:

     (if (eq? (colour-of car) 'red)
         ...)

The descriptive advantages of symbols over numbers increase as the set of concepts that you want to describe grows. Suppose that a car object can have other properties as well, such as whether it has or uses:

• automatic or manual transmission

• leaded or unleaded fuel

• power steering (or not).

Then a car’s combined property set could be naturally represented and manipulated as a list of symbols:

     (properties-of car1)

⇒

     (red manual unleaded power-steering)
     (if (memq 'power-steering (properties-of car1))
         (display "Unfit people can drive this car.\n")
         (display "You'll need strong arms to drive this car!\n"))

⊣

     Unfit people can drive this car.

Remember, the fundamental property of symbols that we are relying on here is that an occurrence of 'red in one part of a program is an indistinguishable symbol from an occurrence of 'red in another part of a program; this means that symbols can usefully be compared using eq?. At the same time, symbols have naturally descriptive names. This combination of efficiency and descriptive power makes them ideal for use as discrete data.

6.6.7.2 Symbols as Lookup Keys

Given their efficiency and descriptive power, it is natural to use symbols as the keys in an association list or hash table.

To illustrate this, consider a more structured representation of the car properties example from the preceding subsection. Rather than mixing all the properties up together in a flat list, we could use an association list like this:

     (define car1-properties '((colour . red)
                               (transmission . manual)
                               (fuel . unleaded)
                               (steering . power-assisted)))

Notice how this structure is more explicit and extensible than the flat list. For example it makes clear that manual refers to the transmission rather than, say, the windows or the locking of the car. It also allows further properties to use the same symbols among their possible values without becoming ambiguous:

     (define car1-properties '((colour . red)
                               (transmission . manual)
                               (fuel . unleaded)
                               (steering . power-assisted)
                               (seat-colour . red)

                               (locking . manual)))

With a representation like this, it is easy to use the efficient assq-XXX family of procedures (see Section 6.7.12 [Association Lists], page 228) to extract or change individual pieces of information:

     (assq-ref car1-properties 'fuel) ⇒ unleaded
     (assq-ref car1-properties 'transmission) ⇒ manual
     (assq-set! car1-properties 'seat-colour 'black)

⇒

     ((colour . red)
      (transmission . manual)
      (fuel . unleaded)
      (steering . power-assisted)
      (seat-colour . black)
      (locking . manual)))

Hash tables also have keys, and exactly the same arguments apply to the use of symbols in hash tables as in association lists. The hash value that Guile uses to decide where to add a symbol-keyed entry to a hash table can be obtained by calling the symbol-hash procedure:

[Scheme Procedure] symbol-hash symbol [C Function] scm_symbol_hash (symbol)

Return a hash value for symbol.

See Section 6.7.14 [Hash Tables], page 236 for information about hash tables in general, and for why you might choose to use a hash table rather than an association list.

6.6.7.3 Symbols as Denoting Variables

When an unquoted symbol in a Scheme program is evaluated, it is interpreted as a variable reference, and the result of the evaluation is the appropriate variable’s value.

For example, when the expression (string-length "abcd") is read and evaluated, the sequence of characters string-length is read as the symbol whose name is "string-length". This symbol is associated with a variable whose value is the procedure that implements string length calculation. Therefore evaluation of the string-length symbol results in that procedure.

The details of the connection between an unquoted symbol and the variable to which it refers are explained elsewhere. See Section 6.12 [Binding Constructs], page 286, for how associations between symbols and variables are created, and Section 6.19 [Modules], page 381, for how those associations are affected by Guile’s module system.

6.6.7.4 Operations Related to Symbols

Given any Scheme value, you can determine whether it is a symbol using the symbol? primitive:

[Scheme Procedure] symbol? obj [C Function] scm_symbol_p (obj)

Return #t if obj is a symbol, otherwise return #f.

[C Function] int scm_is_symbol (SCM val)

Equivalent to scm_is_true (scm_symbol_p (val)).

Once you know that you have a symbol, you can obtain its name as a string by calling symbol->string. Note that Guile differs by default from R5RS on the details of symbol- >string as regards case-sensitivity:

[Scheme Procedure] symbol->string s

[C Function] scm_symbol_to_string (s)

Return the name of symbol s as a string. By default, Guile reads symbols casesensitively, so the string returned will have the same case variation as the sequence of characters that caused s to be created. If Guile is set to read symbols case-insensitively (as specified by R5RS), and s comes into being as part of a literal expression (see Section “Literal expressions” in The Revised^5 Report on Scheme) or by a call to the read or string-ci->symbol procedures, Guile converts any alphabetic characters in the symbol’s name to lower case before creating the symbol object, so the string returned here will be in lower case. If s was created by string->symbol, the case of characters in the string returned will be the same as that in the string that was passed to string->symbol, regardless of Guile’s case-sensitivity setting at the time s was created. It is an error to apply mutation procedures like string-set! to strings returned by this procedure.

Most symbols are created by writing them literally in code. However it is also possible to create symbols programmatically using the following procedures:

[Scheme Procedure] symbol char…

Return a newly allocated symbol made from the given character arguments.

     (symbol #\x #\y #\z) ⇒ xyz

[Scheme Procedure] list->symbol lst

Return a newly allocated symbol made from a list of characters.

     (list->symbol '(#\a #\b #\c)) ⇒ abc

[Scheme Procedure] symbol-append arg …

Return a newly allocated symbol whose characters form the concatenation of the given symbols, arg . . ..

     (let ((h 'hello))
       (symbol-append h 'world))
     ⇒ helloworld

[Scheme Procedure] string->symbol string

[C Function] scm_string_to_symbol (string)

Return the symbol whose name is string. This procedure can create symbols with names containing special characters or letters in the non-standard case, but it is usually a bad idea to create such symbols because in some implementations of Scheme they cannot be read as themselves.

[Scheme Procedure] string-ci->symbol str

[C Function] scm_string_ci_to_symbol (str)

Return the symbol whose name is str. If Guile is currently reading symbols caseinsensitively, str is converted to lowercase before the returned symbol is looked up or created.

The following examples illustrate Guile’s detailed behaviour as regards the case-sensitivity of symbols:

(read-enable 'case-insensitive)   ; R5RS compliant behaviour
(symbol->string 'flying-fish)  ⇒ "flying-fish"
(symbol->string 'Martin)  ⇒ "martin"
(symbol->string
   (string->symbol "Malvina")) ⇒ "Malvina"
(eq? 'mISSISSIppi 'mississippi) ⇒ #t
(string->symbol "mISSISSIppi") ⇒ mISSISSIppi
(eq? 'bitBlt (string->symbol "bitBlt")) ⇒ #f
(eq? 'LolliPop
  (string->symbol (symbol->string 'LolliPop))) ⇒ #t
(string=? "K. Harper, M.D."
  (symbol->string
    (string->symbol "K. Harper, M.D."))) ⇒ #t
(read-disable 'case-insensitive)   ; Guile default behaviour
(symbol->string 'flying-fish)  ⇒ "flying-fish"
(symbol->string 'Martin)  ⇒ "Martin"
(symbol->string
   (string->symbol "Malvina")) ⇒ "Malvina"
(eq? 'mISSISSIppi 'mississippi) ⇒ #f
(string->symbol "mISSISSIppi") ⇒ mISSISSIppi
(eq? 'bitBlt (string->symbol "bitBlt")) ⇒ #t
(eq? 'LolliPop
  (string->symbol (symbol->string 'LolliPop))) ⇒ #t
(string=? "K. Harper, M.D."
  (symbol->string
    (string->symbol "K. Harper, M.D."))) ⇒ #t

From C, there are lower level functions that construct a Scheme symbol from a C string in the current locale encoding.

When you want to do more from C, you should convert between symbols and strings

using scm_symbol_to_string and scm_string_to_symbol and work with the strings.

[C Function] SCM scm_from_latin1_symbol (const char *name) [C Function] SCM scm_from_utf8_symbol (const char *name)

Construct and return a Scheme symbol whose name is specified by the null-terminated C string name. These are appropriate when the C string is hard-coded in the source code.

[C Function] SCM scm_from_locale_symbol (const char *name) [C Function] SCM scm_from_locale_symboln (const char *name, size t len)

Construct and return a Scheme symbol whose name is specified by name. For scm_

from_locale_symbol, name must be null terminated; for scm_from_locale_symboln

the length of name is specified explicitly by len. Note that these functions should not be used when name is a C string constant, because there is no guarantee that the current locale will match that of the execution character set, used for string and character constants. Most modern C compilers use UTF-8 by default, so in such cases we recommend scm_from_utf8_symbol.

[C Function] SCM scm_take_locale_symbol (char *str)

[C Function] SCM scm_take_locale_symboln (char *str, size t len)

Like scm_from_locale_symbol and scm_from_locale_symboln, respectively, but

also frees str with free eventually. Thus, you can use this function when you would free str anyway immediately after creating the Scheme string. In certain cases, Guile can then use str directly as its internal representation.

The size of a symbol can also be obtained from C:

[C Function] size_t scm_c_symbol_length (SCM sym)

Return the number of characters in sym.

Finally, some applications, especially those that generate new Scheme code dynamically, need to generate symbols for use in the generated code. The gensym primitive meets this need:

[Scheme Procedure] gensym [prefix] [C Function] scm_gensym (prefix)

Create a new symbol with a name constructed from a prefix and a counter value. The string prefix can be specified as an optional argument. Default prefix is ‘ g’. The counter is increased by 1 at each call. There is no provision for resetting the counter.

The symbols generated by gensym are likely to be unique, since their names begin with a space and it is only otherwise possible to generate such symbols if a programmer goes out of their way to do so. Uniqueness can be guaranteed by instead using uninterned symbols (see Section 6.6.7.7 [Symbol Uninterned], page 178), though they can’t be usefully written out and read back in.

6.6.7.5 Function Slots and Property Lists

In traditional Lisp dialects, symbols are often understood as having three kinds of value at once:

• a variable value, which is used when the symbol appears in code in a variable reference context • a function value, which is used when the symbol appears in code in a function name

position (i.e. as the first element in an unquoted list)

• a property list value, which is used when the symbol is given as the first argument to Lisp’s put or get functions.

Although Scheme (as one of its simplifications with respect to Lisp) does away with the distinction between variable and function namespaces, Guile currently retains some elements of the traditional structure in case they turn out to be useful when implementing translators for other languages, in particular Emacs Lisp.

Specifically, Guile symbols have two extra slots, one for a symbol’s property list, and one for its “function value.” The following procedures are provided to access these slots.

[Scheme Procedure] symbol-fref symbol [C Function] scm_symbol_fref (symbol)

Return the contents of symbol’s function slot.

[Scheme Procedure] symbol-fset! symbol value [C Function] scm_symbol_fset_x (symbol, value)

Set the contents of symbol’s function slot to value.

[Scheme Procedure] symbol-pref symbol [C Function] scm_symbol_pref (symbol)

Return the property list currently associated with symbol.

[Scheme Procedure] symbol-pset! symbol value [C Function] scm_symbol_pset_x (symbol, value)

Set symbol’s property list to value.

[Scheme Procedure] symbol-property sym prop

From sym’s property list, return the value for property prop. The assumption is that sym’s property list is an association list whose keys are distinguished from each other using equal?; prop should be one of the keys in that list. If the property list has no entry for prop, symbol-property returns #f.

[Scheme Procedure] set-symbol-property! sym prop val

In sym’s property list, set the value for property prop to val, or add a new entry for prop, with value val, if none already exists. For the structure of the property list, see

symbol-property.

[Scheme Procedure] symbol-property-remove! sym prop

From sym’s property list, remove the entry for property prop, if there is one. For the structure of the property list, see symbol-property.

Support for these extra slots may be removed in a future release, and it is probably better to avoid using them. For a more modern and Schemely approach to properties, see Section 6.11.2 [Object Properties], page 278.

6.6.7.6 Extended Read Syntax for Symbols

The read syntax for a symbol is a sequence of letters, digits, and extended alphabetic characters, beginning with a character that cannot begin a number. In addition, the special cases of +, -, and ... are read as symbols even though numbers can begin with +, - or ..

Extended alphabetic characters may be used within identifiers as if they were letters. The set of extended alphabetic characters is:

     ! $ % & * + - . / : < = > ? @ ^_~

In addition to the standard read syntax defined above (which is taken from R5RS (see Section “Formal syntax” in The Revised^5 Report on Scheme)), Guile provides an extended symbol read syntax that allows the inclusion of unusual characters such as space characters, newlines and parentheses. If (for whatever reason) you need to write a symbol containing characters not mentioned above, you can do so as follows.

• Begin the symbol with the characters #{,

• write the characters of the symbol and

• finish the symbol with the characters }#.

Here are a few examples of this form of read syntax. The first symbol needs to use extended syntax because it contains a space character, the second because it contains a line break, and the last because it looks like a number.

     #{foo bar}#
     #{what
     ever}#
     #{4242}#

Although Guile provides this extended read syntax for symbols, widespread usage of it is discouraged because it is not portable and not very readable.

Alternatively, if you enable the r7rs-symbols read option (see see Section 6.17.2 [Scheme Read], page 360), you can write arbitrary symbols using the same notation used for strings, except delimited by vertical bars instead of double quotes.

     |foo bar|
     |\x3BB; is a greek lambda|
     |\| is a vertical bar|

Note that there’s also an r7rs-symbols print option (see Section 6.17.3 [Scheme Write], page 361). To enable the use of this notation, evaluate one or both of the following expressions:

     (read-enable  'r7rs-symbols)
     (print-enable 'r7rs-symbols)

6.6.7.7 Uninterned Symbols

What makes symbols useful is that they are automatically kept unique. There are no two symbols that are distinct objects but have the same name. But of course, there is no rule without exception. In addition to the normal symbols that have been discussed up to now, you can also create special uninterned symbols that behave slightly differently.

To understand what is different about them and why they might be useful, we look at how normal symbols are actually kept unique.

Whenever Guile wants to find the symbol with a specific name, for example during read or when executing string->symbol, it first looks into a table of all existing symbols to find out whether a symbol with the given name already exists. When this is the case, Guile just returns that symbol. When not, a new symbol with the name is created and entered into the table so that it can be found later.

Sometimes you might want to create a symbol that is guaranteed ‘fresh’, i.e. a symbol that did not exist previously. You might also want to somehow guarantee that no one else will ever unintentionally stumble across your symbol in the future. These properties of a symbol are often needed when generating code during macro expansion. When introducing new temporary variables, you want to guarantee that they don’t conflict with variables in other people’s code.

The simplest way to arrange for this is to create a new symbol but not enter it into the global table of all symbols. That way, no one will ever get access to your symbol by chance. Symbols that are not in the table are called uninterned. Of course, symbols that are in the table are called interned.

You create new uninterned symbols with the function make-symbol. You can test whether a symbol is interned or not with symbol-interned?.

Uninterned symbols break the rule that the name of a symbol uniquely identifies the symbol object. Because of this, they can not be written out and read back in like interned symbols. Currently, Guile has no support for reading uninterned symbols. Note that the function gensym does not return uninterned symbols for this reason.

[Scheme Procedure] make-symbol name [C Function] scm_make_symbol (name)

Return a new uninterned symbol with the name name. The returned symbol is guaranteed to be unique and future calls to string->symbol will not return it.

[Scheme Procedure] symbol-interned? symbol

[C Function] scm_symbol_interned_p (symbol)

Return #t if symbol is interned, otherwise return #f.

For example:

     (define foo-1 (string->symbol "foo"))
     (define foo-2 (string->symbol "foo"))
     (define foo-3 (make-symbol "foo"))
     (define foo-4 (make-symbol "foo"))
     (eq? foo-1 foo-2)
     ⇒ #t
     ; Two interned symbols with the same name are the same object,
     (eq? foo-1 foo-3)
     ⇒ #f
     ; but a call to make-symbol with the same name returns a

     ; distinct object.
     (eq? foo-3 foo-4)
     ⇒ #f
     ; A call to make-symbol always returns a new object, even for
     ; the same name.
     foo-3
     ⇒ #<uninterned-symbol foo 8085290>
     ; Uninterned symbols print differently from interned symbols,
     (symbol? foo-3)
     ⇒ #t
     ; but they are still symbols,
     (symbol-interned? foo-3)
     ⇒ #f
     ; just not interned.

6.6.8 Keywords

Keywords are self-evaluating objects with a convenient read syntax that makes them easy to type.

Guile’s keyword support conforms to R5RS, and adds a (switchable) read syntax extension to permit keywords to begin with : as well as #:, or to end with :.

6.6.8.1 Why Use Keywords?

Keywords are useful in contexts where a program or procedure wants to be able to accept a large number of optional arguments without making its interface unmanageable.

To illustrate this, consider a hypothetical make-window procedure, which creates a new window on the screen for drawing into using some graphical toolkit. There are many parameters that the caller might like to specify, but which could also be sensibly defaulted, for example:

• color depth – Default: the color depth for the screen

• background color – Default: white

• width – Default: 600

• height – Default: 400

If make-window did not use keywords, the caller would have to pass in a value for each possible argument, remembering the correct argument order and using a special value to indicate the default value for that argument:

     (make-window 'default              ;; Color depth
                  'default              ;; Background color
                  800                   ;; Width
                  100                   ;; Height
                  ...)                  ;; More make-window arguments

With keywords, on the other hand, defaulted arguments are omitted, and non-default arguments are clearly tagged by the appropriate keyword. As a result, the invocation becomes much clearer:

     (make-window #:width 800 #:height 100)

On the other hand, for a simpler procedure with few arguments, the use of keywords would be a hindrance rather than a help. The primitive procedure cons, for example, would not be improved if it had to be invoked as

     (cons #:car x #:cdr y)

So the decision whether to use keywords or not is purely pragmatic: use them if they will clarify the procedure invocation at point of call.

6.6.8.2 Coding With Keywords

If a procedure wants to support keywords, it should take a rest argument and then use whatever means is convenient to extract keywords and their corresponding arguments from the contents of that rest argument.

The following example illustrates the principle: the code for make-window uses a helper procedure called get-keyword-value to extract individual keyword arguments from the rest argument.

     (define (get-keyword-value args keyword default)
       (let ((kv (memq keyword args)))
         (if (and kv (>= (length kv) 2))
             (cadr kv)
             default)))
     (define (make-window . args)
       (let ((depth  (get-keyword-value args #:depth  screen-depth))
             (bg     (get-keyword-value args #:bg     "white"))
             (width  (get-keyword-value args #:width  800))
             (height (get-keyword-value args #:height 100))
             ...)
         ...))

But you don’t need to write get-keyword-value. The (ice-9 optargs) module provides a set of powerful macros that you can use to implement keyword-supporting procedures like this:

     (use-modules (ice-9 optargs))
     (define (make-window . args)
       (let-keywords args #f ((depth  screen-depth)
                              (bg     "white")
                              (width  800)
                              (height 100))
         ...))

Or, even more economically, like this:

     (use-modules (ice-9 optargs))

     (define* (make-window #:key (depth  screen-depth)
                                 (bg     "white")
                                 (width  800)
                                 (height 100))
       ...)

For further details on let-keywords, define* and other facilities provided by the (ice-9 optargs) module, see Section 6.9.4 [Optional Arguments], page 248.

To handle keyword arguments from procedures implemented in C, use scm_c_bind_ keyword_arguments (see Section 6.6.8.4 [Keyword Procedures], page 183).

6.6.8.3 Keyword Read Syntax

Guile, by default, only recognizes a keyword syntax that is compatible with R5RS. A token of the form #:NAME, where NAME has the same syntax as a Scheme symbol (see Section 6.6.7.6 [Symbol Read Syntax], page 178), is the external representation of the keyword named NAME. Keyword objects print using this syntax as well, so values containing keyword objects can be read back into Guile. When used in an expression, keywords are self-quoting objects.

If the keyword read option is set to 'prefix, Guile also recognizes the alternative read syntax :NAME. Otherwise, tokens of the form :NAME are read as symbols, as required by R5RS.

If the keyword read option is set to 'postfix, Guile recognizes the SRFI-88 read syntax NAME: (see Section 7.5.40 [SRFI-88], page 620). Otherwise, tokens of this form are read as symbols.

To enable and disable the alternative non-R5RS keyword syntax, you use the read-set! procedure documented Section 6.17.2 [Scheme Read], page 360. Note that the prefix and postfix syntax are mutually exclusive.

     (read-set! keywords 'prefix)
     #:type

⇒

     #:type
     :type

⇒

     #:type
     (read-set! keywords 'postfix)
     type:

⇒

     #:type
     :type

⇒

     :type

(read-set! keywords #f)
#:type

⇒

#:type
:type

⊣

ERROR: In expression :type:
ERROR: Unbound variable: :type
ABORT: (unbound-variable)

6.6.8.4 Keyword Procedures

[Scheme Procedure] keyword? obj [C Function] scm_keyword_p (obj)

Return #t if the argument obj is a keyword, else #f.

[Scheme Procedure] keyword->symbol keyword

[C Function] scm_keyword_to_symbol (keyword)

Return the symbol with the same name as keyword.

[Scheme Procedure] symbol->keyword symbol

[C Function] scm_symbol_to_keyword (symbol)

Return the keyword with the same name as symbol.

[C Function] int scm_is_keyword (SCM obj)

Equivalent to scm_is_true (scm_keyword_p (obj)).

[C Function] SCM scm_from_locale_keyword (const char *name) [C Function] SCM scm_from_locale_keywordn (const char *name, size t len)

Equivalent to scm_symbol_to_keyword (scm_from_locale_symbol (name)) and
scm_symbol_to_keyword (scm_from_locale_symboln (name, len)), respectively.

Note that these functions should not be used when name is a C string constant, because there is no guarantee that the current locale will match that of the execution character set, used for string and character constants. Most modern C compilers use UTF-8 by default, so in such cases we recommend scm_from_utf8_keyword.

[C Function] SCM scm_from_latin1_keyword (const char *name) [C Function] SCM scm_from_utf8_keyword (const char *name)

Equivalent to scm_symbol_to_keyword (scm_from_latin1_symbol (name)) and
scm_symbol_to_keyword (scm_from_utf8_symbol (name)), respectively.

[C Function] void scm_c_bind_keyword_arguments (const char *subr, SCM rest,

scm t keyword arguments flags flags, SCM keyword1, SCM *argp1, …, SCM keywordN, SCM *argpN, SCM_UNDEFINED)

Extract the specified keyword arguments from rest, which is not modified. If the keyword argument keyword1 is present in rest with an associated value, that value is stored in the variable pointed to by argp1, otherwise the variable is left unchanged. Similarly for the other keywords and argument pointers up to keywordN and argpN. The argument list to scm_c_bind_keyword_arguments must be terminated by SCM_

UNDEFINED.

Note that since the variables pointed to by argp1 through argpN are left unchanged if the associated keyword argument is not present, they should be initialized to their default values before calling scm_c_bind_keyword_arguments. Alternatively, you can initialize them to SCM_UNDEFINED before the call, and then use SCM_UNBNDP after the call to see which ones were provided. If an unrecognized keyword argument is present in rest and flags does not contain SCM_ALLOW_OTHER_KEYS, or if non-keyword arguments are present and flags does not contain SCM_ALLOW_NON_KEYWORD_ARGUMENTS, an exception is raised. subr should be the name of the procedure receiving the keyword arguments, for purposes of error reporting. For example:

     SCM k_delimiter;
     SCM k_grammar;
     SCM sym_infix;
     SCM my_string_join (SCM strings, SCM rest)
     {
       SCM delimiter = SCM_UNDEFINED;
       SCM grammar   = sym_infix;
       scm_c_bind_keyword_arguments ("my-string-join", rest, 0,
                                     k_delimiter, &delimiter,
                                     k_grammar, &grammar,
                                     SCM_UNDEFINED);
       if (SCM_UNBNDP (delimiter))
         delimiter = scm_from_utf8_string (" ");
       return scm_string_join (strings, delimiter, grammar);
     }
     void my_init ()
     {
       k_delimiter = scm_from_utf8_keyword ("delimiter");
       k_grammar   = scm_from_utf8_keyword ("grammar");
       sym_infix   = scm_from_utf8_symbol  ("infix");
       scm_c_define_gsubr ("my-string-join", 1, 0, 1, my_string_join);
     }

6.6.9 “Functionality-Centric” Data Types

Procedures and macros are documented in their own sections: see Section 6.9 [Procedures], page 244 and Section 6.10 [Macros], page 257.

Variable objects are documented as part of the description of Guile’s module system: see Section 6.19.7 [Variables], page 390.

Asyncs, dynamic roots and fluids are described in the section on scheduling: see Section 6.21 [Scheduling], page 410.

Hooks are documented in the section on general utility functions: see Section 6.11.6 [Hooks], page 281.

Ports are described in the section on I/O: see Section 6.14 [Input and Output], page 316. Regular expressions are described in their own section: see Section 6.15 [Regular Ex-

pressions], page 349.

6.7 Compound Data Types

This chapter describes Guile’s compound data types. By compound we mean that the primary purpose of these data types is to act as containers for other kinds of data (including other compound objects). For instance, a (non-uniform) vector with length 5 is a container that can hold five arbitrary Scheme objects.

The various kinds of container object differ from each other in how their memory is allocated, how they are indexed, and how particular values can be looked up within them.

6.7.1 Pairs

Pairs are used to combine two Scheme objects into one compound object. Hence the name: A pair stores a pair of objects.

The data type pair is extremely important in Scheme, just like in any other Lisp dialect. The reason is that pairs are not only used to make two values available as one object, but that pairs are used for constructing lists of values. Because lists are so important in Scheme, they are described in a section of their own (see Section 6.7.2 [Lists], page 188).

Pairs can literally get entered in source code or at the REPL, in the so-called dotted list syntax. This syntax consists of an opening parentheses, the first element of the pair, a dot, the second element and a closing parentheses. The following example shows how a pair consisting of the two numbers 1 and 2, and a pair containing the symbols foo and bar can be entered. It is very important to write the whitespace before and after the dot, because otherwise the Scheme parser would not be able to figure out where to split the tokens.

     (1 . 2)
     (foo . bar)

But beware, if you want to try out these examples, you have to quote the expressions. More information about quotation is available in the section Section 6.17.1.1 [Expression Syntax], page 357. The correct way to try these examples is as follows.

     '(1 . 2)

⇒

     (1 . 2)
     '(foo . bar)

⇒

     (foo . bar)

A new pair is made by calling the procedure cons with two arguments. Then the argument values are stored into a newly allocated pair, and the pair is returned. The name cons stands for "construct". Use the procedure pair? to test whether a given Scheme object is a pair or not.

[Scheme Procedure] cons x y [C Function] scm_cons (x, y)

Return a newly allocated pair whose car is x and whose cdr is y. The pair is guaranteed to be different (in the sense of eq?) from every previously existing object.

[Scheme Procedure] pair? x [C Function] scm_pair_p (x)

Return #t if x is a pair; otherwise return #f.

[C Function] int scm_is_pair (SCM x)

Return 1 when x is a pair; otherwise return 0.

The two parts of a pair are traditionally called car and cdr. They can be retrieved with procedures of the same name (car and cdr), and can be modified with the procedures

set-car! and set-cdr!.

Since a very common operation in Scheme programs is to access the car of a car of a pair, or the car of the cdr of a pair, etc., the procedures called caar, cadr and so on are also predefined. However, using these procedures is often detrimental to readability, and error-prone. Thus, accessing the contents of a list is usually better achieved using pattern matching techniques (see Section 7.7 [Pattern Matching], page 657).

[Scheme Procedure] car pair [Scheme Procedure] cdr pair

[C Function] scm_car (pair) [C Function] scm_cdr (pair)

Return the car or the cdr of pair, respectively.

[C Macro] SCM SCM_CAR (SCM pair) [C Macro] SCM SCM_CDR (SCM pair)

These two macros are the fastest way to access the car or cdr of a pair; they can be thought of as compiling into a single memory reference. These macros do no checking at all. The argument pair must be a valid pair.

[Scheme Procedure] cddr pair [Scheme Procedure] cdar pair [Scheme Procedure] cadr pair [Scheme Procedure] caar pair [Scheme Procedure] cdddr pair [Scheme Procedure] cddar pair [Scheme Procedure] cdadr pair [Scheme Procedure] cdaar pair [Scheme Procedure] caddr pair [Scheme Procedure] cadar pair [Scheme Procedure] caadr pair [Scheme Procedure] caaar pair [Scheme Procedure] cddddr pair [Scheme Procedure] cdddar pair [Scheme Procedure] cddadr pair [Scheme Procedure] cddaar pair [Scheme Procedure] cdaddr pair [Scheme Procedure] cdadar pair [Scheme Procedure] cdaadr pair [Scheme Procedure] cdaaar pair [Scheme Procedure] cadddr pair [Scheme Procedure] caddar pair [Scheme Procedure] cadadr pair [Scheme Procedure] cadaar pair [Scheme Procedure] caaddr pair [Scheme Procedure] caadar pair [Scheme Procedure] caaadr pair [Scheme Procedure] caaaar pair [C Function] scm_cddr (pair) [C Function] scm_cdar (pair) [C Function] scm_cadr (pair) [C Function] scm_caar (pair) [C Function] scm_cdddr (pair) [C Function] scm_cddar (pair) [C Function] scm_cdadr (pair) [C Function] scm_cdaar (pair) [C Function] scm_caddr (pair) [C Function] scm_cadar (pair) [C Function] scm_caadr (pair) [C Function] scm_caaar (pair) [C Function] scm_cddddr (pair) [C Function] scm_cdddar (pair) [C Function] scm_cddadr (pair) [C Function] scm_cddaar (pair) [C Function] scm_cdaddr (pair) [C Function] scm_cdadar (pair) [C Function] scm_cdaadr (pair) [C Function] scm_cdaaar (pair) [C Function] scm_cadddr (pair) [C Function] scm_caddar (pair) [C Function] scm_cadadr (pair) [C Function] scm_cadaar (pair) [C Function] scm_caaddr (pair) [C Function] scm_caadar (pair) [C Function] scm_caaadr (pair) [C Function] scm_caaaar (pair)

These procedures are compositions of car and cdr, where for example caddr could be defined by

          (define caddr (lambda (x) (car (cdr (cdr x)))))

cadr, caddr and cadddr pick out the second, third or fourth elements of a list, respectively. SRFI-1 provides the same under the names second, third and fourth (see Section 7.5.3.3 [SRFI-1 Selectors], page 554).

[Scheme Procedure] set-car! pair value

[C Function] scm_set_car_x (pair, value)

Stores value in the car field of pair. The value returned by set-car! is unspecified.

[Scheme Procedure] set-cdr! pair value

[C Function] scm_set_cdr_x (pair, value)

Stores value in the cdr field of pair. The value returned by set-cdr! is unspecified.

6.7.2 Lists

A very important data type in Scheme—as well as in all other Lisp dialects—is the data type list.^{2}

This is the short definition of what a list is:

• Either the empty list (),

• or a pair which has a list in its cdr.

6.7.2.1 List Read Syntax

The syntax for lists is an opening parentheses, then all the elements of the list (separated by whitespace) and finally a closing parentheses.^{3}.

(1 2 3) ; a list of the numbers 1, 2 and 3

("foo" bar 3.1415) ; a string, a symbol and a real number

() ; the empty list The last example needs a bit more explanation. A list with no elements, called the

empty list, is special in some ways. It is used for terminating lists by storing it into the cdr of the last pair that makes up a list. An example will clear that up:

     (car '(1))

⇒

     1
     (cdr '(1))

⇒

     ()

This example also shows that lists have to be quoted when written (see Section 6.17.1.1 [Expression Syntax], page 357), because they would otherwise be mistakingly taken as procedure applications (see Section 3.2.2 [Simple Invocation], page 18).

6.7.2.2 List Predicates

Often it is useful to test whether a given Scheme object is a list or not. List-processing procedures could use this information to test whether their input is valid, or they could do different things depending on the datatype of their arguments.

2 Strictly speaking, Scheme does not have a real datatype list. Lists are made up of chained pairs, and

only exist by definition—a list is a chain of pairs which looks like a list.

3 Note that there is no separation character between the list elements, like a comma or a semicolon.

[Scheme Procedure] list? x [C Function] scm_list_p (x)

Return #t if x is a proper list, else #f.

The predicate null? is often used in list-processing code to tell whether a given list has run out of elements. That is, a loop somehow deals with the elements of a list until the list satisfies null?. Then, the algorithm terminates.

[Scheme Procedure] null? x [C Function] scm_null_p (x)

Return #t if x is the empty list, else #f.

[C Function] int scm_is_null (SCM x)

Return 1 when x is the empty list; otherwise return 0.

6.7.2.3 List Constructors

This section describes the procedures for constructing new lists. list simply returns a list where the elements are the arguments, cons* is similar, but the last argument is stored in the cdr of the last pair of the list.

[Scheme Procedure] list elem … [C Function] scm_list_1 (elem1)

[C Function] scm_list_2 (elem1, elem2) [C Function] scm_list_3 (elem1, elem2, elem3) [C Function] scm_list_4 (elem1, elem2, elem3, elem4) [C Function] scm_list_5 (elem1, elem2, elem3, elem4, elem5)

[C Function] scm_list_n (elem1, …, elemN, SCM_UNDEFINED)

Return a new list containing elements elem . . .. scm_list_n takes a variable number of arguments, terminated by the special SCM_ UNDEFINED. That final SCM_UNDEFINED is not included in the list. None of elem … can themselves be SCM_UNDEFINED, or scm_list_n will terminate at that point.

[Scheme Procedure] cons* arg1 arg2 … Like list, but the last arg provides the tail of the constructed list, returning (cons arg1 (cons arg2 (cons ... argn))). Requires at least one argument. If given one argument, that argument is returned as result. This function is called list* in some other Schemes and in Common LISP.

[Scheme Procedure] list-copy lst [C Function] scm_list_copy (lst)

Return a (newly-created) copy of lst.

[Scheme Procedure] make-list n [init]

Create a list containing of n elements, where each element is initialized to init. init defaults to the empty list () if not given.

Note that list-copy only makes a copy of the pairs which make up the spine of the lists. The list elements are not copied, which means that modifying the elements of the new list also modifies the elements of the old list. On the other hand, applying procedures like set-cdr! or delv! to the new list will not alter the old list. If you also need to copy the list elements (making a deep copy), use the procedure copy-tree (see Section 6.11.4 [Copying], page 280).

6.7.2.4 List Selection

These procedures are used to get some information about a list, or to retrieve one or more elements of a list.

[Scheme Procedure] length lst [C Function] scm_length (lst)

Return the number of elements in list lst.

[Scheme Procedure] last-pair lst [C Function] scm_last_pair (lst)

Return the last pair in lst, signalling an error if lst is circular.

[Scheme Procedure] list-ref list k [C Function] scm_list_ref (list, k)

Return the kth element from list.

[Scheme Procedure] list-tail lst k [Scheme Procedure] list-cdr-ref lst k [C Function] scm_list_tail (lst, k)

Return the "tail" of lst beginning with its kth element. The first element of the list is considered to be element 0. list-tail and list-cdr-ref are identical. It may help to think of list-cdr-ref as accessing the kth cdr of the list, or returning the results of cdring k times down lst.

[Scheme Procedure] list-head lst k [C Function] scm_list_head (lst, k)

Copy the first k elements from lst into a new list, and return it.

6.7.2.5 Append and Reverse

append and append! are used to concatenate two or more lists in order to form a new list. reverse and reverse! return lists with the same elements as their arguments, but in reverse order. The procedure variants with an ! directly modify the pairs which form the list, whereas the other procedures create new pairs. This is why you should be careful when using the side-effecting variants.

[Scheme Procedure] append lst … obj [Scheme Procedure] append [Scheme Procedure] append! lst … obj [Scheme Procedure] append! [C Function] scm_append (lstlst) [C Function] scm_append_x (lstlst)

Return a list comprising all the elements of lists lst … obj. If called with no arguments, return the empty list.

     (append '(x) '(y))  ⇒ (x y)
     (append '(a) '(b c d))  ⇒ (a b c d)
     (append '(a (b)) '((c)))  ⇒ (a (b) (c))

The last argument obj may actually be any object; an improper list results if the last argument is not a proper list.

     (append '(a b) '(c . d))  ⇒ (a b c . d)
     (append '() 'a)  ⇒ a

append doesn’t modify the given lists, but the return may share structure with the final obj. append! is permitted, but not required, to modify the given lists to form its return.

For scm_append and scm_append_x, lstlst is a list of the list operands lst … obj. That lstlst itself is not modified or used in the return.

[Scheme Procedure] reverse lst [Scheme Procedure] reverse! lst [newtail] [C Function] scm_reverse (lst) [C Function] scm_reverse_x (lst, newtail)

Return a list comprising the elements of lst, in reverse order.

reverse constructs a new list. reverse! is permitted, but not required, to modify lst in constructing its return.

For reverse!, the optional newtail is appended to the result. newtail isn’t reversed, it simply becomes the list tail. For scm_reverse_x, the newtail parameter is mandatory, but can be SCM_EOL if no further tail is required.

6.7.2.6 List Modification

The following procedures modify an existing list, either by changing elements of the list, or by changing the list structure itself.

[Scheme Procedure] list-set! list k val [C Function] scm_list_set_x (list, k, val)

Set the kth element of list to val.

[Scheme Procedure] list-cdr-set! list k val [C Function] scm_list_cdr_set_x (list, k, val)

Set the kth cdr of list to val.

[Scheme Procedure] delq item lst [C Function] scm_delq (item, lst)

Return a newly-created copy of lst with elements eq? to item removed. This procedure mirrors memq: delq compares elements of lst against item with eq?.

[Scheme Procedure] delv item lst [C Function] scm_delv (item, lst)

Return a newly-created copy of lst with elements eqv? to item removed. This procedure mirrors memv: delv compares elements of lst against item with eqv?.

[Scheme Procedure] delete item lst

[C Function] scm_delete (item, lst)

Return a newly-created copy of lst with elements equal? to item removed. This procedure mirrors member: delete compares elements of lst against item with equal?. See also SRFI-1 which has an extended delete (Section 7.5.3.8 [SRFI-1 Deleting], page 561), and also an lset-difference which can delete multiple items in one call (Section 7.5.3.10 [SRFI-1 Set Operations], page 563).

[Scheme Procedure] delq! item lst [Scheme Procedure] delv! item lst

[Scheme Procedure] delete! item lst [C Function] scm_delq_x (item, lst) [C Function] scm_delv_x (item, lst) [C Function] scm_delete_x (item, lst)

These procedures are destructive versions of delq, delv and delete: they modify the pointers in the existing lst rather than creating a new list. Caveat evaluator: Like other destructive list functions, these functions cannot modify the binding of lst, and so cannot be used to delete the first element of lst destructively.

[Scheme Procedure] delq1! item lst

[C Function] scm_delq1_x (item, lst)

Like delq!, but only deletes the first occurrence of item from lst. Tests for equality

using eq?. See also delv1! and delete1!.

[Scheme Procedure] delv1! item lst

[C Function] scm_delv1_x (item, lst)

Like delv!, but only deletes the first occurrence of item from lst. Tests for equality

using eqv?. See also delq1! and delete1!.

[Scheme Procedure] delete1! item lst [C Function] scm_delete1_x (item, lst)

Like delete!, but only deletes the first occurrence of item from lst. Tests for equality

using equal?. See also delq1! and delv1!.

[Scheme Procedure] filter pred lst

[Scheme Procedure] filter! pred lst

Return a list containing all elements from lst which satisfy the predicate pred. The elements in the result list have the same order as in lst. The order in which pred is applied to the list elements is not specified. filter does not change lst, but the result may share a tail with it. filter! may modify lst to construct its return.

6.7.2.7 List Searching

The following procedures search lists for particular elements. They use different comparison predicates for comparing list elements with the object to be searched. When they fail, they return #f, otherwise they return the sublist whose car is equal to the search object, where equality depends on the equality predicate used.

[Scheme Procedure] memq x lst [C Function] scm_memq (x, lst)

Return the first sublist of lst whose car is eq? to x where the sublists of lst are the non-empty lists returned by (list-tail lst k) for k less than the length of lst. If x does not occur in lst, then #f (not the empty list) is returned.

[Scheme Procedure] memv x lst [C Function] scm_memv (x, lst)

Return the first sublist of lst whose car is eqv? to x where the sublists of lst are the non-empty lists returned by (list-tail lst k) for k less than the length of lst. If x does not occur in lst, then #f (not the empty list) is returned.

[Scheme Procedure] member x lst [C Function] scm_member (x, lst)

Return the first sublist of lst whose car is equal? to x where the sublists of lst are the non-empty lists returned by (list-tail lst k) for k less than the length of lst. If x does not occur in lst, then #f (not the empty list) is returned. See also SRFI-1 which has an extended member function (Section 7.5.3.7 [SRFI-1 Searching], page 560).

6.7.2.8 List Mapping

List processing is very convenient in Scheme because the process of iterating over the elements of a list can be highly abstracted. The procedures in this section are the most basic iterating procedures for lists. They take a procedure and one or more lists as arguments, and apply the procedure to each element of the list. They differ in their return value.

[Scheme Procedure] map proc arg1 arg2 …

[Scheme Procedure] map-in-order proc arg1 arg2 …

[C Function] scm_map (proc, arg1, args)

Apply proc to each element of the list arg1 (if only two arguments are given), or to the corresponding elements of the argument lists (if more than two arguments are given). The result(s) of the procedure applications are saved and returned in a list. For map, the order of procedure applications is not specified, map-in-order applies the procedure from left to right to the list elements.

[Scheme Procedure] for-each proc arg1 arg2 … Like map, but the procedure is always applied from left to right, and the result(s) of the procedure applications are thrown away. The return value is not specified.

See also SRFI-1 which extends these functions to take lists of unequal lengths (Section 7.5.3.5 [SRFI-1 Fold and Map], page 556).

6.7.3 Vectors

Vectors are sequences of Scheme objects. Unlike lists, the length of a vector, once the vector is created, cannot be changed. The advantage of vectors over lists is that the time required to access one element of a vector given its position (synonymous with index), a zero-origin number, is constant, whereas lists have an access time linear to the position of the accessed element in the list.

Vectors can contain any kind of Scheme object; it is even possible to have different types of objects in the same vector. For vectors containing vectors, you may wish to use arrays, instead. Note, too, that vectors are the special case of one dimensional non-uniform arrays and that most array procedures operate happily on vectors (see Section 6.7.5 [Arrays], page 200).

Also see Section 7.5.29 [SRFI-43], page 609, for a comprehensive vector library.

6.7.3.1 Read Syntax for Vectors

Vectors can literally be entered in source code, just like strings, characters or some of the other data types. The read syntax for vectors is as follows: A sharp sign (#), followed by an opening parentheses, all elements of the vector in their respective read syntax, and finally a closing parentheses. Like strings, vectors do not have to be quoted.

The following are examples of the read syntax for vectors; where the first vector only contains numbers and the second three different object types: a string, a symbol and a number in hexadecimal notation.

     #(1 2 3)
     #("Hello" foo #xdeadbeef)

6.7.3.2 Dynamic Vector Creation and Validation

Instead of creating a vector implicitly by using the read syntax just described, you can create a vector dynamically by calling one of the vector and list->vector primitives with the list of Scheme values that you want to place into a vector. The size of the vector thus created is determined implicitly by the number of arguments given.

[Scheme Procedure] vector arg … [Scheme Procedure] list->vector l [C Function] scm_vector (l)

Return a newly allocated vector composed of the given arguments. Analogous to

     list.
          (vector 'a 'b 'c) ⇒ #(a b c)

The inverse operation is vector->list:

[Scheme Procedure] vector->list v

[C Function] scm_vector_to_list (v)

Return a newly allocated list composed of the elements of v.

          (vector->list #(dah dah didah)) ⇒ (dah dah didah)
          (list->vector '(dididit dah)) ⇒ #(dididit dah)

To allocate a vector with an explicitly specified size, use make-vector. With this primitive you can also specify an initial value for the vector elements (the same value for all elements, that is):

[Scheme Procedure] make-vector len [fill] [C Function] scm_make_vector (len, fill)

Return a newly allocated vector of len elements. If a second argument is given, then each position is initialized to fill. Otherwise the initial contents of each position is unspecified.

[C Function] SCM scm_c_make_vector (size t k, SCM fill)

Like scm_make_vector, but the length is given as a size_t.

To check whether an arbitrary Scheme value is a vector, use the vector? primitive:

[Scheme Procedure] vector? obj [C Function] scm_vector_p (obj)

Return #t if obj is a vector, otherwise return #f.

[C Function] int scm_is_vector (SCM obj)

Return non-zero when obj is a vector, otherwise return zero.

6.7.3.3 Accessing and Modifying Vector Contents

vector-length and vector-ref return information about a given vector, respectively its size and the elements that are contained in the vector.

[Scheme Procedure] vector-length vector [C Function] scm_vector_length (vector)

Return the number of elements in vector as an exact integer.

[C Function] size_t scm_c_vector_length (SCM vec)

Return the number of elements in vec as a size_t.

[Scheme Procedure] vector-ref vec k [C Function] scm_vector_ref (vec, k)

Return the contents of position k of vec. k must be a valid index of vec.

     (vector-ref #(1 1 2 3 5 8 13 21) 5) ⇒ 8
     (vector-ref #(1 1 2 3 5 8 13 21)
         (let ((i (round (* 2 (acos -1)))))
           (if (inexact? i)
             (inexact->exact i)
                i))) ⇒ 13

[C Function] SCM scm_c_vector_ref (SCM vec, size t k)

Return the contents of position k (a size_t) of vec.

A vector created by one of the dynamic vector constructor procedures (see Section 6.7.3.2 [Vector Creation], page 194) can be modified using the following procedures.

NOTE: According to R5RS, it is an error to use any of these procedures on a literally read vector, because such vectors should be considered as constants. Currently, however, Guile does not detect this error.

[Scheme Procedure] vector-set! vec k obj [C Function] scm_vector_set_x (vec, k, obj)

Store obj in position k of vec. k must be a valid index of vec. The value returned by ‘vector-set!’ is unspecified.

     (let ((vec (vector 0 '(2 2 2 2) "Anna")))
       (vector-set! vec 1 '("Sue" "Sue"))
       vec) ⇒ #(0 ("Sue" "Sue") "Anna")

[C Function] void scm_c_vector_set_x (SCM vec, size t k, SCM obj)

Store obj in position k (a size_t) of vec.

[Scheme Procedure] vector-fill! vec fill [C Function] scm_vector_fill_x (vec, fill)

Store fill in every position of vec. The value returned by vector-fill! is unspecified.

[Scheme Procedure] vector-copy vec [C Function] scm_vector_copy (vec)

Return a copy of vec.

[Scheme Procedure] vector-move-left! vec1 start1 end1 vec2 start2 [C Function] scm_vector_move_left_x (vec1, start1, end1, vec2, start2)

Copy elements from vec1, positions start1 to end1, to vec2 starting at position start2. start1 and start2 are inclusive indices; end1 is exclusive. vector-move-left! copies elements in leftmost order. Therefore, in the case where vec1 and vec2 refer to the same vector, vector-move-left! is usually appropriate when start1 is greater than start2.

[Scheme Procedure] vector-move-right! vec1 start1 end1 vec2 start2 [C Function] scm_vector_move_right_x (vec1, start1, end1, vec2, start2)

Copy elements from vec1, positions start1 to end1, to vec2 starting at position start2. start1 and start2 are inclusive indices; end1 is exclusive. vector-move-right! copies elements in rightmost order. Therefore, in the case where vec1 and vec2 refer to the same vector, vector-move-right! is usually appropriate when start1 is less than start2.

6.7.3.4 Vector Accessing from C

A vector can be read and modified from C with the functions scm_c_vector_ref and scm_ c_vector_set_x, for example. In addition to these functions, there are two more ways to access vectors from C that might be more efficient in certain situations: you can restrict yourself to simple vectors and then use the very fast simple vector macros; or you can use the very general framework for accessing all kinds of arrays (see Section 6.7.5.4 [Accessing Arrays from C], page 208), which is more verbose, but can deal efficiently with all kinds of vectors (and arrays). For vectors, you can use the scm_vector_elements and scm_vector_ writable_elements functions as shortcuts.

[C Function] int scm_is_simple_vector (SCM obj)

Return non-zero if obj is a simple vector, else return zero. A simple vector is a vector that can be used with the SCM_SIMPLE_* macros below. The following functions are guaranteed to return simple vectors: scm_make_vector,

scm_c_make_vector, scm_vector, scm_list_to_vector.
[C Macro] size_t SCM_SIMPLE_VECTOR_LENGTH (SCM vec)

Evaluates to the length of the simple vector vec. No type checking is done.

[C Macro] SCM SCM_SIMPLE_VECTOR_REF (SCM vec, size t idx)

Evaluates to the element at position idx in the simple vector vec. No type or range checking is done.

[C Macro] void SCM_SIMPLE_VECTOR_SET (SCM vec, size t idx, SCM val)

Sets the element at position idx in the simple vector vec to val. No type or range checking is done.

[C Function] const SCM * scm_vector_elements (SCM vec, scm t array handle

*handle, size t *lenp, ssize t *incp) Acquire a handle for the vector vec and return a pointer to the elements of it. This pointer can only be used to read the elements of vec. When vec is not a vector, an error is signaled. The handle must eventually be released with scm_array_handle_

release.

The variables pointed to by lenp and incp are filled with the number of elements of the vector and the increment (number of elements) between successive elements, respectively. Successive elements of vec need not be contiguous in their underlying “root vector” returned here; hence the increment is not necessarily equal to 1 and may well be negative too (see Section 6.7.5.3 [Shared Arrays], page 206). The following example shows the typical way to use this function. It creates a list of all elements of vec (in reverse order).

     scm_t_array_handle handle;
     size_t i, len;
     ssize_t inc;
     const SCM *elt;
     SCM list;
     elt = scm_vector_elements (vec, &handle, &len, &inc);
     list = SCM_EOL;
     for (i = 0; i < len; i++, elt += inc)
       list = scm_cons (*elt, list);
     scm_array_handle_release (&handle);
[C Function] SCM * scm_vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp) Like scm_vector_elements but the pointer can be used to modify the vector. The following example shows the typical way to use this function. It fills a vector with #t.

     scm_t_array_handle handle;
     size_t i, len;
     ssize_t inc;
     SCM *elt;
     elt = scm_vector_writable_elements (vec, &handle, &len, &inc);
     for (i = 0; i < len; i++, elt += inc)
       *elt = SCM_BOOL_T;
     scm_array_handle_release (&handle);

6.7.3.5 Uniform Numeric Vectors

A uniform numeric vector is a vector whose elements are all of a single numeric type. Guile offers uniform numeric vectors for signed and unsigned 8-bit, 16-bit, 32-bit, and 64-bit integers, two sizes of floating point values, and complex floating-point numbers of these two sizes. See Section 7.5.5 [SRFI-4], page 566, for more information.

For many purposes, bytevectors work just as well as uniform vectors, and have the advantage that they integrate well with binary input and output. See Section 6.6.6 [Bytevectors], page 163, for more information on bytevectors.

6.7.4 Bit Vectors

Bit vectors are zero-origin, one-dimensional arrays of booleans. They are displayed as a sequence of 0s and 1s prefixed by #*, e.g.,

(make-bitvector 8 #f) ⇒
#*00000000

Bit vectors are the special case of one dimensional bit arrays, and can thus be used with the array procedures, See Section 6.7.5 [Arrays], page 200.

[Scheme Procedure] bitvector? obj [C Function] scm_bitvector_p (obj)

Return #t when obj is a bitvector, else return #f.

[C Function] int scm_is_bitvector (SCM obj)

Return 1 when obj is a bitvector, else return 0.

[Scheme Procedure] make-bitvector len [fill] [C Function] scm_make_bitvector (len, fill)

Create a new bitvector of length len and optionally initialize all elements to fill.

[C Function] SCM scm_c_make_bitvector (size t len, SCM fill)

Like scm_make_bitvector, but the length is given as a size_t.

[Scheme Procedure] bitvector bit … [C Function] scm_bitvector (bits)

Create a new bitvector with the arguments as elements.

[Scheme Procedure] bitvector-length vec

[C Function] scm_bitvector_length (vec)

Return the length of the bitvector vec.

[C Function] size_t scm_c_bitvector_length (SCM vec)

Like scm_bitvector_length, but the length is returned as a size_t.

[Scheme Procedure] bitvector-ref vec idx [C Function] scm_bitvector_ref (vec, idx)

Return the element at index idx of the bitvector vec.

[C Function] SCM scm_c_bitvector_ref (SCM vec, size t idx)

Return the element at index idx of the bitvector vec.

[Scheme Procedure] bitvector-set! vec idx val [C Function] scm_bitvector_set_x (vec, idx, val)

Set the element at index idx of the bitvector vec when val is true, else clear it.

[C Function] SCM scm_c_bitvector_set_x (SCM vec, size t idx, SCM val)

Set the element at index idx of the bitvector vec when val is true, else clear it.

[Scheme Procedure] bitvector-fill! vec val

[C Function] scm_bitvector_fill_x (vec, val)

Set all elements of the bitvector vec when val is true, else clear them.

[Scheme Procedure] list->bitvector list

[C Function] scm_list_to_bitvector (list)

Return a new bitvector initialized with the elements of list.

[Scheme Procedure] bitvector->list vec

[C Function] scm_bitvector_to_list (vec)

Return a new list initialized with the elements of the bitvector vec.

[Scheme Procedure] bit-count bool bitvector [C Function] scm_bit_count (bool, bitvector)

Return a count of how many entries in bitvector are equal to bool. For example,

     (bit-count #f #*000111000) ⇒ 6

[Scheme Procedure] bit-position bool bitvector start [C Function] scm_bit_position (bool, bitvector, start)

Return the index of the first occurrence of bool in bitvector, starting from start. If there is no bool entry between start and the end of bitvector, then return #f. For example,

     (bit-position #t #*000101 0) ⇒ 3
     (bit-position #f #*0001111 3) ⇒ #f

[Scheme Procedure] bit-invert! bitvector [C Function] scm_bit_invert_x (bitvector)

Modify bitvector by replacing each element with its negation.

[Scheme Procedure] bit-set*! bitvector uvec bool [C Function] scm_bit_set_star_x (bitvector, uvec, bool)

Set entries of bitvector to bool, with uvec selecting the entries to change. The return value is unspecified.

If uvec is a bit vector, then those entries where it has #t are the ones in bitvector which are set to bool. uvec and bitvector must be the same length. When bool is #t it’s like uvec is OR’ed into bitvector. Or when bool is #f it can be seen as an ANDNOT.

     (define bv #*01000010)
     (bit-set*! bv #*10010001 #t)
     bv
     ⇒ #*11010011

If uvec is a uniform vector of unsigned long integers, then they’re indexes into bitvector which are set to bool.

          (define bv #*01000010)
          (bit-set*! bv #u(5 2 7) #t)
          bv
          ⇒ #*01100111

[Scheme Procedure] bit-count* bitvector uvec bool [C Function] scm_bit_count_star (bitvector, uvec, bool)

Return a count of how many entries in bitvector are equal to bool, with uvec selecting the entries to consider. uvec is interpreted in the same way as for bit-set*! above. Namely, if uvec is a bit vector then entries which have #t there are considered in bitvector. Or if uvec is a uniform vector of unsigned long integers then it’s the indexes in bitvector to consider. For example,

          (bit-count* #*01110111 #*11001101 #t) ⇒ 3
          (bit-count* #*01110111 #u(7 0 4) #f) ⇒ 2
[C Function] const scm_t_uint32 * scm_bitvector_elements (SCM vec,

scm t array handle *handle, size t *offp, size t *lenp, ssize t *incp) Like scm_vector_elements (see Section 6.7.3.4 [Vector Accessing from C], page 196), but for bitvectors. The variable pointed to by offp is set to the value returned by

     scm_array_handle_bit_elements_offset. See scm_array_handle_bit_elements

for how to use the returned pointer and the offset.

[C Function] scm_t_uint32 * scm_bitvector_writable_elements (SCM vec,

scm t array handle *handle, size t *offp, size t *lenp, ssize t *incp) Like scm_bitvector_elements, but the pointer is good for reading and writing.

6.7.5 Arrays

Arrays are a collection of cells organized into an arbitrary number of dimensions. Each cell can be accessed in constant time by supplying an index for each dimension.

In the current implementation, an array uses a vector of some kind for the actual storage of its elements. Any kind of vector will do, so you can have arrays of uniform numeric values, arrays of characters, arrays of bits, and of course, arrays of arbitrary Scheme values. For example, arrays with an underlying c64vector might be nice for digital signal processing, while arrays made from a u8vector might be used to hold gray-scale images.

The number of dimensions of an array is called its rank. Thus, a matrix is an array of rank 2, while a vector has rank 1. When accessing an array element, you have to specify one exact integer for each dimension. These integers are called the indices of the element. An array specifies the allowed range of indices for each dimension via an inclusive lower and upper bound. These bounds can well be negative, but the upper bound must be greater than or equal to the lower bound minus one. When all lower bounds of an array are zero, it is called a zero-origin array.

Arrays can be of rank 0, which could be interpreted as a scalar. Thus, a zero-rank array can store exactly one object and the list of indices of this element is the empty list.

Arrays contain zero elements when one of their dimensions has a zero length. These empty arrays maintain information about their shape: a matrix with zero columns and 3 rows is different from a matrix with 3 columns and zero rows, which again is different from a vector of length zero.

The array procedures are all polymorphic, treating strings, uniform numeric vectors, bytevectors, bit vectors and ordinary vectors as one dimensional arrays.

6.7.5.1 Array Syntax

An array is displayed as # followed by its rank, followed by a tag that describes the underlying vector, optionally followed by information about its shape, and finally followed by the cells, organized into dimensions using parentheses.

In more words, the array tag is of the form

       #<rank><vectag><@lower><:len><@lower><:len>...

where <rank> is a positive integer in decimal giving the rank of the array. It is omitted when the rank is 1 and the array is non-shared and has zero-origin (see below). For shared arrays and for a non-zero origin, the rank is always printed even when it is 1 to distinguish them from ordinary vectors.

The <vectag> part is the tag for a uniform numeric vector, like u8, s16, etc, b for bitvectors, or a for strings. It is empty for ordinary vectors.

The <@lower> part is a ‘@’ character followed by a signed integer in decimal giving the lower bound of a dimension. There is one <@lower> for each dimension. When all lower bounds are zero, all <@lower> parts are omitted.

The <:len> part is a ‘:’ character followed by an unsigned integer in decimal giving the length of a dimension. Like for the lower bounds, there is one <:len> for each dimension, and the <:len> part always follows the <@lower> part for a dimension. Lengths are only then printed when they can’t be deduced from the nested lists of elements of the array literal, which can happen when at least one length is zero.

As a special case, an array of rank 0 is printed as #0<vectag>(<scalar>), where <scalar> is the result of printing the single element of the array.

Thus,

#(1 2 3) is an ordinary array of rank 1 with lower bound 0 in dimension 0. (I.e., a regular vector.)

#@2(1 2 3)

is an ordinary array of rank 1 with lower bound 2 in dimension 0.

#2((1 2 3) (4 5 6))

is a non-uniform array of rank 2; a 3×3 matrix with index ranges 0..2 and 0..2.

#u32(0 1 2)

is a uniform u8 array of rank 1.

#2u32@2@3((1 2) (2 3))

is a uniform u8 array of rank 2 with index ranges 2..3 and 3..4.

#2() is a two-dimensional array with index ranges 0..-1 and 0..-1, i.e. both dimensions have length zero.

#2:0:2() is a two-dimensional array with index ranges 0..-1 and 0..1, i.e. the first dimension has length zero, but the second has length 2.

#0(12) is a rank-zero array with contents 12.

In addition, bytevectors are also arrays, but use a different syntax (see Section 6.6.6 [Bytevectors], page 163):

#vu8(1 2 3)

is a 3-byte long bytevector, with contents 1, 2, 3.

6.7.5.2 Array Procedures

When an array is created, the range of each dimension must be specified, e.g., to create a 2×3 array with a zero-based index:

(make-array 'ho 2 3) ⇒ #2((ho ho ho) (ho ho ho))

The range of each dimension can also be given explicitly, e.g., another way to create the same array:

(make-array 'ho '(0 1) '(0 2)) ⇒ #2((ho ho ho) (ho ho ho))

The following procedures can be used with arrays (or vectors). An argument shown as idx… means one parameter for each dimension in the array. A idxlist argument means a list of such values, one for each dimension.

[Scheme Procedure] array? obj

[C Function] scm_array_p (obj, unused)

Return #t if the obj is an array, and #f if not. The second argument to scm array p is there for historical reasons, but it is not used. You should always pass SCM_UNDEFINED as its value.

[Scheme Procedure] typed-array? obj type [C Function] scm_typed_array_p (obj, type)

Return #t if the obj is an array of type type, and #f if not.

[C Function] int scm_is_array (SCM obj)

Return 1 if the obj is an array and 0 if not.

[C Function] int scm_is_typed_array (SCM obj, SCM type)

Return 0 if the obj is an array of type type, and 1 if not.

[Scheme Procedure] make-array fill bound … [C Function] scm_make_array (fill, bounds)

Equivalent to (make-typed-array #t fill bound ...).

[Scheme Procedure] make-typed-array type fill bound … [C Function] scm_make_typed_array (type, fill, bounds)

Create and return an array that has as many dimensions as there are bounds and (maybe) fill it with fill. The underlying storage vector is created according to type, which must be a symbol whose name is the ‘vectag’ of the array as explained above, or #t for ordinary, nonspecialized arrays. For example, using the symbol f64 for type will create an array that uses a f64vector for storing its elements, and a will use a string.

When fill is not the special unspecified value, the new array is filled with fill. Otherwise, the initial contents of the array is unspecified. The special unspecified value is stored in the variable *unspecified* so that for example (make-typed-array 'u32 *unspecified* 4) creates a uninitialized u32 vector of length 4.

Each bound may be a positive non-zero integer n, in which case the index for that dimension can range from 0 through n-1; or an explicit index range specifier in the form (LOWER UPPER), where both lower and upper are integers, possibly less than zero, and possibly the same number (however, lower cannot be greater than upper).

[Scheme Procedure] list->array dimspec list

Equivalent to (list->typed-array #t dimspec list).

[Scheme Procedure] list->typed-array type dimspec list [C Function] scm_list_to_typed_array (type, dimspec, list)

Return an array of the type indicated by type with elements the same as those of list.

The argument dimspec determines the number of dimensions of the array and their lower bounds. When dimspec is an exact integer, it gives the number of dimensions directly and all lower bounds are zero. When it is a list of exact integers, then each element is the lower index bound of a dimension, and there will be as many dimensions as elements in the list.

[Scheme Procedure] array-type array [C Function] scm_array_type (array)

Return the type of array. This is the ‘vectag’ used for printing array (or #t for ordinary arrays) and can be used with make-typed-array to create an array of the same kind as array.

[Scheme Procedure] array-ref array idx …

[C Function] scm_array_ref (array, idxlist)

Return the element at (idx ...) in array.

     (define a (make-array 999 '(1 2) '(3 4)))
     (array-ref a 2 4) ⇒ 999

[Scheme Procedure] array-in-bounds? array idx … [C Function] scm_array_in_bounds_p (array, idxlist)

Return #t if the given indices would be acceptable to array-ref.

     (define a (make-array #f '(1 2) '(3 4)))
     (array-in-bounds? a 2 3) ⇒ #t
     (array-in-bounds? a 0 0) ⇒ #f

[Scheme Procedure] array-set! array obj idx … [C Function] scm_array_set_x (array, obj, idxlist)

Set the element at (idx ...) in array to obj. The return value is unspecified.

     (define a (make-array #f '(0 1) '(0 1)))
     (array-set! a #t 1 1)
     a ⇒ #2((#f #f) (#f #t))

[Scheme Procedure] array-shape array [Scheme Procedure] array-dimensions array

[C Function] scm_array_dimensions (array)

Return a list of the bounds for each dimension of array.

array-shape gives (lower upper) for each dimension. array-dimensions instead

returns just upper + 1 for dimensions with a 0 lower bound. Both are suitable as

input to make-array.

For example,

     (define a (make-array 'foo '(-1 3) 5))
     (array-shape a)  ⇒ ((-1 3) (0 4))
     (array-dimensions a) ⇒ ((-1 3) 5)

[Scheme Procedure] array-length array [C Function] scm_array_length (array)

[C Function] size_t scm c array length (array)

Return the length of an array: its first dimension. It is an error to ask for the length of an array of rank 0.

[Scheme Procedure] array-rank array [C Function] scm_array_rank (array)

Return the rank of array.

[C Function] size_t scm_c_array_rank (SCM array)

Return the rank of array as a size_t.

[Scheme Procedure] array->list array [C Function] scm_array_to_list (array)

Return a list consisting of all the elements, in order, of array.

[Scheme Procedure] array-copy! src dst [Scheme Procedure] array-copy-in-order! src dst [C Function] scm_array_copy_x (src, dst)

Copy every element from vector or array src to the corresponding element of dst. dst must have the same rank as src, and be at least as large in each dimension. The return value is unspecified.

[Scheme Procedure] array-fill! array fill [C Function] scm_array_fill_x (array, fill)

Store fill in every element of array. The value returned is unspecified.

[Scheme Procedure] array-equal? array …

Return #t if all arguments are arrays with the same shape, the same type, and have corresponding elements which are either equal? or array-equal?. This function differs from equal? (see Section 6.11.1 [Equality], page 276) in that all arguments must be arrays.

[Scheme Procedure] array-map! dst proc src … [Scheme Procedure] array-map-in-order! dst proc src1 … srcN

[C Function] scm_array_map_x (dst, proc, srclist)

Set each element of the dst array to values obtained from calls to proc. The value returned is unspecified. Each call is (proc elem1 ... elemN), where each elem is from the corresponding src array, at the dst index. array-map-in-order! makes the calls in row-major order, array-map! makes them in an unspecified order. The src arrays must have the same number of dimensions as dst, and must have a range for each dimension which covers the range in dst. This ensures all dst indices are valid in each src.

[Scheme Procedure] array-for-each proc src1 src2 …

[C Function] scm_array_for_each (proc, src1, srclist)

Apply proc to each tuple of elements of src1 src2 …, in row-major order. The value returned is unspecified.

[Scheme Procedure] array-index-map! dst proc

[C Function] scm_array_index_map_x (dst, proc)

Set each element of the dst array to values returned by calls to proc. The value returned is unspecified. Each call is (proc i1 ... iN), where i1 …iN is the destination index, one parameter for each dimension. The order in which the calls are made is unspecified. For example, to create a 4 × 4 matrix representing a cyclic group,

( \begin{matrix}0 & 1 & 2 & 3 \\ 1 & 2 & 3 & 0 \\ 2 & 3 & 0 & 1 \\ 3 & 0 & 1 & 2 \end{matrix} )

     (define a (make-array #f 4 4))
     (array-index-map! a (lambda (i j)
                           (modulo (+ i j) 4)))

[Scheme Procedure] uniform-array-read! ra [port or fd [start [end]]] [C Function] scm_uniform_array_read_x (ra, port or fd, start, end)

Attempt to read all elements of array ra, in lexicographic order, as binary objects from port or fd. If an end of file is encountered, the objects up to that point are put into ra (starting at the beginning) and the remainder of the array is unchanged. The optional arguments start and end allow a specified region of a vector (or linearized array) to be read, leaving the remainder of the vector unchanged. uniform-array-read! returns the number of objects read. port or fd may be omitted, in which case it defaults to the value returned by (current-input-port).

[Scheme Procedure] uniform-array-write ra [port or fd [start [end]]] [C Function] scm_uniform_array_write (ra, port or fd, start, end)

Writes all elements of ra as binary objects to port or fd. The optional arguments start and end allow a specified region of a vector (or linearized array) to be written. The number of objects actually written is returned. port or fd may be omitted, in which case it defaults to the value returned by (current-output-port).

6.7.5.3 Shared Arrays

[Scheme Procedure] make-shared-array oldarray mapfunc bound … [C Function] scm_make_shared_array (oldarray, mapfunc, boundlist)

Return a new array which shares the storage of oldarray. Changes made through either affect the same underlying storage. The bound … arguments are the shape of the new array, the same as make-array (see Section 6.7.5.2 [Array Procedures], page 202). mapfunc translates coordinates from the new array to the oldarray. It’s called as (mapfunc newidx1 ...) with one parameter for each dimension of the new array, and should return a list of indices for oldarray, one for each dimension of oldarray. mapfunc must be affine linear, meaning that each oldarray index must be formed by adding integer multiples (possibly negative) of some or all of newidx1 etc, plus a possible integer offset. The multiples and offset must be the same in each call.

One good use for a shared array is to restrict the range of some dimensions, so as to apply say array-for-each or array-fill! to only part of an array. The plain list function can be used for mapfunc in this case, making no changes to the index values. For example,

     (make-shared-array #2((a b c) (d e f) (g h i)) list 3 2)
     ⇒ #2((a b) (d e) (g h))

The new array can have fewer dimensions than oldarray, for example to take a column from an array.

     (make-shared-array #2((a b c) (d e f) (g h i))
                        (lambda (i) (list i 2))
                        '(0 2))
     ⇒ #1(c f i)

A diagonal can be taken by using the single new array index for both row and column in the old array. For example,

     (make-shared-array #2((a b c) (d e f) (g h i))
                        (lambda (i) (list i i))
                        '(0 2))
     ⇒ #1(a e i)

Dimensions can be increased by for instance considering portions of a one dimensional array as rows in a two dimensional array. (array-contents below can do the opposite, flattening an array.)

     (make-shared-array #1(a b c d e f g h i j k l)
                        (lambda (i j) (list (+ (* i 3) j)))
                        4 3)
     ⇒ #2((a b c) (d e f) (g h i) (j k l))

By negating an index the order that elements appear can be reversed. The following just reverses the column order,

     (make-shared-array #2((a b c) (d e f) (g h i))
                        (lambda (i j) (list i (- 2 j)))

                        3 3)
     ⇒ #2((c b a) (f e d) (i h g))

A fixed offset on indexes allows for instance a change from a 0 based to a 1 based array,

     (define x #2((a b c) (d e f) (g h i)))
     (define y (make-shared-array x
                                  (lambda (i j) (list (1- i) (1- j)))
                                  '(1 3) '(1 3)))
     (array-ref x 0 0) ⇒ a
     (array-ref y 1 1) ⇒ a

A multiple on an index allows every Nth element of an array to be taken. The following is every third element,

     (make-shared-array #1(a b c d e f g h i j k l)
                        (lambda (i) (list (* i 3)))
                        4)
     ⇒ #1(a d g j)

The above examples can be combined to make weird and wonderful selections from an array, but it’s important to note that because mapfunc must be affine linear, arbitrary permutations are not possible. In the current implementation, mapfunc is not called for every access to the new array but only on some sample points to establish a base and stride for new array indices in oldarray data. A few sample points are enough because mapfunc is linear.

[Scheme Procedure] shared-array-increments array
[C Function] scm_shared_array_increments (array)

For each dimension, return the distance between elements in the root vector.

[Scheme Procedure] shared-array-offset array

[C Function] scm_shared_array_offset (array)

Return the root vector index of the first element in the array.

[Scheme Procedure] shared-array-root array

[C Function] scm_shared_array_root (array)

Return the root vector of a shared array.

[Scheme Procedure] array-contents array [strict] [C Function] scm_array_contents (array, strict)

If array may be unrolled into a one dimensional shared array without changing their order (last subscript changing fastest), then array-contents returns that shared array, otherwise it returns #f. All arrays made by make-array and make-typed-array may be unrolled, some arrays made by make-shared-array may not be. If the optional argument strict is provided, a shared array will be returned only if its elements are stored internally contiguous in memory.

[Scheme Procedure] transpose-array array dim1 dim2 … [C Function] scm_transpose_array (array, dimlist)

Return an array sharing contents with array, but with dimensions arranged in a different order. There must be one dim argument for each dimension of array. dim1, dim2, … should be integers between 0 and the rank of the array to be returned. Each integer in that range must appear at least once in the argument list.

The values of dim1, dim2, … correspond to dimensions in the array to be returned, and their positions in the argument list to dimensions of array. Several dims may have the same value, in which case the returned array will have smaller rank than array.

          (transpose-array '#2((a b) (c d)) 1 0) ⇒ #2((a c) (b d))
          (transpose-array '#2((a b) (c d)) 0 0) ⇒ #1(a d)
          (transpose-array '#3(((a b c) (d e f)) ((1 2 3) (4 5 6))) 1 1 0) ⇒
                          #2((a 4) (b 5) (c 6))

6.7.5.4 Accessing Arrays from C

For interworking with external C code, Guile provides an API to allow C code to access the elements of a Scheme array. In particular, for uniform numeric arrays, the API exposes the underlying uniform data as a C array of numbers of the relevant type.

While pointers to the elements of an array are in use, the array itself must be protected so that the pointer remains valid. Such a protected array is said to be reserved. A reserved array can be read but modifications to it that would cause the pointer to its elements to become invalid are prevented. When you attempt such a modification, an error is signalled.

(This is similar to locking the array while it is in use, but without the danger of a deadlock. In a multi-threaded program, you will need additional synchronization to avoid modifying reserved arrays.)

You must take care to always unreserve an array after reserving it, even in the presence of non-local exits. If a non-local exit can happen between these two calls, you should install a dynwind context that releases the array when it is left (see Section 6.13.10 [Dynamic Wind], page 309).

In addition, array reserving and unreserving must be properly paired. For instance, when reserving two or more arrays in a certain order, you need to unreserve them in the opposite order.

Once you have reserved an array and have retrieved the pointer to its elements, you must figure out the layout of the elements in memory. Guile allows slices to be taken out of arrays without actually making a copy, such as making an alias for the diagonal of a matrix that can be treated as a vector. Arrays that result from such an operation are not stored contiguously in memory and when working with their elements directly, you need to take this into account.

The layout of array elements in memory can be defined via a mapping function that computes a scalar position from a vector of indices. The scalar position then is the offset of the element with the given indices from the start of the storage block of the array.

In Guile, this mapping function is restricted to be affine: all mapping functions of Guile

arrays can be written as p = b + c[0]*i[0] + c[1]*i[1] + ... + c[n-1]*i[n-1] where

i[k] is the kth index and n is the rank of the array. For example, a matrix of size 3x3 would have b == 0, c[0] == 3 and c[1] == 1. When you transpose this matrix (with transposearray, say), you will get an array whose mapping function has b == 0, c[0] == 1 and c[1]

== 3.

The function scm_array_handle_dims gives you (indirect) access to the coefficients

c[k].

Note that there are no functions for accessing the elements of a character array yet. Once the string implementation of Guile has been changed to use Unicode, we will provide them.

[C Type] scm_t_array_handle

This is a structure type that holds all information necessary to manage the reservation of arrays as explained above. Structures of this type must be allocated on the stack and must only be accessed by the functions listed below.

[C Function] void scm_array_get_handle (SCM array, scm t array handle

*handle) Reserve array, which must be an array, and prepare handle to be used with the functions below. You must eventually call scm_array_handle_release on handle, and do this in a properly nested fashion, as explained above. The structure pointed to by handle does not need to be initialized before calling this function.

[C Function] void scm_array_handle_release (scm t array handle *handle)

End the array reservation represented by handle. After a call to this function, handle might be used for another reservation.

[C Function] size_t scm_array_handle_rank (scm t array handle *handle)

Return the rank of the array represented by handle.

[C Type] scm_t_array_dim

This structure type holds information about the layout of one dimension of an array. It includes the following fields:

ssize_t lbnd
ssize_t ubnd

The lower and upper bounds (both inclusive) of the permissible index range for the given dimension. Both values can be negative, but lbnd is always less than or equal to ubnd.

ssize_t inc

The distance from one element of this dimension to the next. Note, too, that this can be negative.

[C Function] const scm_t_array_dim * scm_array_handle_dims

(scm t array handle *handle) Return a pointer to a C vector of information about the dimensions of the array represented by handle. This pointer is valid as long as the array remains reserved. As explained above, the scm_t_array_dim structures returned by this function can be used calculate the position of an element in the storage block of the array from its indices. This position can then be used as an index into the C array pointer returned by the

various scm_array_handle_<foo>_elements functions, or with scm_array_handle_
ref and scm_array_handle_set.

Here is how one can compute the position pos of an element given its indices in the vector indices:

     ssize_t indices[RANK];
     scm_t_array_dim *dims;
     ssize_t pos;
     size_t i;
     pos = 0;
     for (i = 0; i < RANK; i++)
       {
         if (indices[i] < dims[i].lbnd || indices[i] > dims[i].ubnd)
           out_of_range ();
         pos += (indices[i] - dims[i].lbnd) * dims[i].inc;
       }

[C Function] ssize_t scm_array_handle_pos (scm t array handle *handle, SCM

indices) Compute the position corresponding to indices, a list of indices. The position is computed as described above for scm_array_handle_dims. The number of the indices and their range is checked and an appropriate error is signalled for invalid indices.

[C Function] SCM scm_array_handle_ref (scm t array handle *handle, ssize t pos)

Return the element at position pos in the storage block of the array represented by handle. Any kind of array is acceptable. No range checking is done on pos.

[C Function] void scm_array_handle_set (scm t array handle *handle, ssize t

pos, SCM val) Set the element at position pos in the storage block of the array represented by handle to val. Any kind of array is acceptable. No range checking is done on pos. An error is signalled when the array can not store val.

[C Function] const SCM * scm_array_handle_elements (scm t array handle

*handle) Return a pointer to the elements of a ordinary array of general Scheme values (i.e., a non-uniform array) for reading. This pointer is valid as long as the array remains reserved.

[C Function] SCM * scm_array_handle_writable_elements (scm t array handle

*handle) Like scm_array_handle_elements, but the pointer is good for reading and writing.

[C Function] const void * scm_array_handle_uniform_elements

(scm t array handle *handle) Return a pointer to the elements of a uniform numeric array for reading. This pointer is valid as long as the array remains reserved. The size of each element is given by

scm_array_handle_uniform_element_size.

[C Function] void * scm_array_handle_uniform_writable_elements

(scm t array handle *handle) Like scm_array_handle_uniform_elements, but the pointer is good reading and writing.

[C Function] size_t scm_array_handle_uniform_element_size

(scm t array handle *handle) Return the size of one element of the uniform numeric array represented by handle.

[C Function] const scm_t_uint8 * scm_array_handle_u8_elements

(scm t array handle *handle)

[C Function] const scm_t_int8 * scm_array_handle_s8_elements

(scm t array handle *handle)

[C Function] const scm_t_uint16 * scm_array_handle_u16_elements

(scm t array handle *handle)

[C Function] const scm_t_int16 * scm_array_handle_s16_elements

(scm t array handle *handle)

[C Function] const scm_t_uint32 * scm_array_handle_u32_elements

(scm t array handle *handle)

[C Function] const scm_t_int32 * scm_array_handle_s32_elements

(scm t array handle *handle)

[C Function] const scm_t_uint64 * scm_array_handle_u64_elements

(scm t array handle *handle)

[C Function] const scm_t_int64 * scm_array_handle_s64_elements

(scm t array handle *handle)

[C Function] const float * scm_array_handle_f32_elements

(scm t array handle *handle)

[C Function] const double * scm_array_handle_f64_elements

(scm t array handle *handle)

[C Function] const float * scm_array_handle_c32_elements

(scm t array handle *handle)

[C Function] const double * scm_array_handle_c64_elements

(scm t array handle *handle) Return a pointer to the elements of a uniform numeric array of the indicated kind for reading. This pointer is valid as long as the array remains reserved.

The pointers for c32 and c64 uniform numeric arrays point to pairs of floating point numbers. The even index holds the real part, the odd index the imaginary part of the complex number.

[C Function] scm_t_uint8 * scm_array_handle_u8_writable_elements

(scm t array handle *handle)

[C Function] scm_t_int8 * scm_array_handle_s8_writable_elements

(scm t array handle *handle)

[C Function] scm_t_uint16 * scm_array_handle_u16_writable_elements

(scm t array handle *handle)

[C Function] scm_t_int16 * scm_array_handle_s16_writable_elements

(scm t array handle *handle)

[C Function] scm_t_uint32 * scm_array_handle_u32_writable_elements

(scm t array handle *handle)

[C Function] scm_t_int32 * scm_array_handle_s32_writable_elements

(scm t array handle *handle)

[C Function] scm_t_uint64 * scm_array_handle_u64_writable_elements

(scm t array handle *handle)

[C Function] scm_t_int64 * scm_array_handle_s64_writable_elements

(scm t array handle *handle)

[C Function] float * scm_array_handle_f32_writable_elements

(scm t array handle *handle)

[C Function] double * scm_array_handle_f64_writable_elements

(scm t array handle *handle)

[C Function] float * scm_array_handle_c32_writable_elements

(scm t array handle *handle)

[C Function] double * scm_array_handle_c64_writable_elements

(scm t array handle *handle) Like scm_array_handle_<kind>_elements, but the pointer is good for reading and writing.

[C Function] const scm_t_uint32 * scm_array_handle_bit_elements

(scm t array handle *handle) Return a pointer to the words that store the bits of the represented array, which must be a bit array. Unlike other arrays, bit arrays have an additional offset that must be figured into index

calculations. That offset is returned by scm_array_handle_bit_elements_offset.

To find a certain bit you first need to calculate its position as explained above for scm_array_handle_dims and then add the offset. This gives the absolute position of the bit, which is always a non-negative integer. Each word of the bit array storage block contains exactly 32 bits, with the least significant bit in that word having the lowest absolute position number. The next word contains the next 32 bits. Thus, the following code can be used to access a bit whose position according to

scm_array_handle_dims is given in pos:
     SCM bit_array;
     scm_t_array_handle handle;
     scm_t_uint32 *bits;
     ssize_t pos;
     size_t abs_pos;
     size_t word_pos, mask;
     scm_array_get_handle (&bit_array, &handle);
     bits = scm_array_handle_bit_elements (&handle);
     pos = ...
     abs_pos = pos + scm_array_handle_bit_elements_offset (&handle);
     word_pos = abs_pos / 32;

          mask = 1L << (abs_pos % 32);
          if (bits[word_pos] & mask)
            /* bit is set. */
          scm_array_handle_release (&handle);
[C Function] scm_t_uint32 * scm_array_handle_bit_writable_elements

(scm t array handle *handle) Like scm_array_handle_bit_elements but the pointer is good for reading and writing. You must take care not to modify bits outside of the allowed index range of the array, even for contiguous arrays.

6.7.6 VLists

The (ice-9 vlist) module provides an implementation of the VList data structure designed by Phil Bagwell in 2002. VLists are immutable lists, which can contain any Scheme object. They improve on standard Scheme linked lists in several areas:

• Random access has typically constant-time complexity.

• Computing the length of a VList has time complexity logarithmic in the number of elements.

• VLists use less storage space than standard lists.

• VList elements are stored in contiguous regions, which improves memory locality and leads to more efficient use of hardware caches.

The idea behind VLists is to store vlist elements in increasingly large contiguous blocks (implemented as vectors here). These blocks are linked to one another using a pointer to the next block and an offset within that block. The size of these blocks form a geometric series with ratio block-growth-factor (2 by default).

The VList structure also serves as the basis for the VList-based hash lists or “vhashes”, an immutable dictionary type (see Section 6.7.13 [VHashes], page 234).

However, the current implementation in (ice-9 vlist) has several noteworthy shortcomings:

• It is not thread-safe. Although operations on vlists are all referentially transparent (i.e., purely functional), adding elements to a vlist with vlist-cons mutates part of its internal structure, which makes it non-thread-safe. This could be fixed, but it would

    slow down vlist-cons.

• vlist-cons always allocates at least as much memory as cons. Again, Phil Bagwell describes how to fix it, but that would require tuning the garbage collector in a way that may not be generally beneficial.

• vlist-cons is a Scheme procedure compiled to bytecode, and it does not compete with the straightforward C implementation of cons, and with the fact that the VM has a special cons instruction.

We hope to address these in the future.

The programming interface exported by (ice-9 vlist) is defined below. Most of it is the same as SRFI-1 with an added vlist- prefix to function names.

[Scheme Procedure] vlist? obj

Return true if obj is a VList.

[Scheme Variable] vlist-null

The empty VList. Note that it’s possible to create an empty VList not eq? to vlistnull; thus, callers should always use vlist-null? when testing whether a VList is empty.

[Scheme Procedure] vlist-null? vlist

Return true if vlist is empty.

[Scheme Procedure] vlist-cons item vlist

Return a new vlist with item as its head and vlist as its tail.

[Scheme Procedure] vlist-head vlist

Return the head of vlist.

[Scheme Procedure] vlist-tail vlist

Return the tail of vlist.

[Scheme Variable] block-growth-factor

A fluid that defines the growth factor of VList blocks, 2 by default.

The functions below provide the usual set of higher-level list operations.

[Scheme Procedure] vlist-fold proc init vlist

[Scheme Procedure] vlist-fold-right proc init vlist

Fold over vlist, calling proc for each element, as for SRFI-1 fold and fold-right (see Section 7.5.3 [SRFI-1], page 552).

[Scheme Procedure] vlist-ref vlist index Return the element at index index in vlist. This is typically a constant-time operation.

[Scheme Procedure] vlist-length vlist

Return the length of vlist. This is typically logarithmic in the number of elements in vlist.

[Scheme Procedure] vlist-reverse vlist

Return a new vlist whose content are those of vlist in reverse order.

[Scheme Procedure] vlist-map proc vlist

Map proc over the elements of vlist and return a new vlist.

[Scheme Procedure] vlist-for-each proc vlist

Call proc on each element of vlist. The result is unspecified.

[Scheme Procedure] vlist-drop vlist count

Return a new vlist that does not contain the count first elements of vlist. This is typically a constant-time operation.

[Scheme Procedure] vlist-take vlist count

Return a new vlist that contains only the count first elements of vlist.

[Scheme Procedure] vlist-filter pred vlist

Return a new vlist containing all the elements from vlist that satisfy pred.

[Scheme Procedure] vlist-delete x vlist [equal?]

Return a new vlist corresponding to vlist without the elements equal? to x.

[Scheme Procedure] vlist-unfold p f g seed [tail-gen]

[Scheme Procedure] vlist-unfold-right p f g seed [tail]

Return a new vlist, as for SRFI-1 unfold and unfold-right (see Section 7.5.3 [SRFI-1], page 552).

[Scheme Procedure] vlist-append vlist …

Append the given vlists and return the resulting vlist.

[Scheme Procedure] list->vlist lst

Return a new vlist whose contents correspond to lst.

[Scheme Procedure] vlist->list vlist

Return a new list whose contents match those of vlist.

6.7.7 Record Overview

Records, also called structures, are Scheme’s primary mechanism to define new disjoint types. A record type defines a list of fields that instances of the type consist of. This is like

C’s struct.

Historically, Guile has offered several different ways to define record types and to create records, offering different features, and making different trade-offs. Over the years, each “standard” has also come with its own new record interface, leading to a maze of record APIs.

At the highest level is SRFI-9, a high-level record interface implemented by most Scheme implementations (see Section 6.7.8 [SRFI-9 Records], page 216). It defines a simple and efficient syntactic abstraction of record types and their associated type predicate, fields, and field accessors. SRFI-9 is suitable for most uses, and this is the recommended way to create record types in Guile. Similar high-level record APIs include SRFI-35 (see Section 7.5.23 [SRFI-35], page 594) and R6RS records (see Section 7.6.2.9 [rnrs records syntactic], page 635).

Then comes Guile’s historical “records” API (see Section 6.7.9 [Records], page 219). Record types defined this way are first-class objects. Introspection facilities are available, allowing users to query the list of fields or the value of a specific field at run-time, without prior knowledge of the type.

Finally, the common denominator of these interfaces is Guile’s structure API (see Section 6.7.10 [Structures], page 220). Guile’s structures are the low-level building block for all other record APIs. Application writers will normally not need to use it.

Records created with these APIs may all be pattern-matched using Guile’s standard pattern matcher (see Section 7.7 [Pattern Matching], page 657).

6.7.8 SRFI-9 Records

SRFI-9 standardizes a syntax for defining new record types and creating predicate, constructor, and field getter and setter functions. In Guile this is the recommended option to create new record types (see Section 6.7.7 [Record Overview], page 215). It can be used with:

(use-modules (srfi srfi-9))
[Scheme Syntax] define-record-type type

(constructor fieldname …)

predicate

(fieldname accessor [modifier]) …

Create a new record type, and make various defines for using it. This syntax can only occur at the top-level, not nested within some other form. type is bound to the record type, which is as per the return from the core makerecord-type. type also provides the name for the record, as per record-type-name. constructor is bound to a function to be called as (constructor fieldval ...) to create a new record of this type. The arguments are initial values for the fields, one argument for each field, in the order they appear in the define-record-type form. The fieldnames provide the names for the record fields, as per the core record-type-fields etc, and are referred to in the subsequent accessor/modifier forms. predicate is bound to a function to be called as (predicate obj). It returns #t or #f according to whether obj is a record of this type. Each accessor is bound to a function to be called (accessor record) to retrieve the respective field from a record. Similarly each modifier is bound to a function to be called (modifier record val) to set the respective field in a record.

An example will illustrate typical usage,

(define-record-type <employee>
  (make-employee name age salary)
  employee?
  (name    employee-name)
  (age     employee-age    set-employee-age!)
  (salary  employee-salary set-employee-salary!))

This creates a new employee data type, with name, age and salary fields. Accessor functions are created for each field, but no modifier function for the name (the intention in this example being that it’s established only when an employee object is created). These can all then be used as for example,

<employee> ⇒ #<record-type <employee>>
(define fred (make-employee "Fred" 45 20000.00))
(employee? fred)  ⇒ #t
(employee-age fred)  ⇒ 45
(set-employee-salary! fred 25000.00)  ;; pay rise

The functions created by define-record-type are ordinary top-level defines. They can be redefined or set! as desired, exported from a module, etc.

Non-toplevel Record Definitions

The SRFI-9 specification explicitly disallows record definitions in a non-toplevel context, such as inside lambda body or inside a let block. However, Guile’s implementation does not enforce that restriction.

Custom Printers

You may use set-record-type-printer! to customize the default printing behavior of records. This is a Guile extension and is not part of SRFI-9. It is located in the (srfi

srfi-9 gnu) module.
[Scheme Syntax] set-record-type-printer! name proc

Where type corresponds to the first argument of define-record-type, and proc is a procedure accepting two arguments, the record to print, and an output port.

This example prints the employee’s name in brackets, for instance [Fred].

     (set-record-type-printer! <employee>
       (lambda (record port)
         (write-char #\[ port)
         (display (employee-name record) port)
         (write-char #\] port)))

Functional “Setters”

When writing code in a functional style, it is desirable to never alter the contents of records. For such code, a simple way to return new record instances based on existing ones is highly desirable.

The (srfi srfi-9 gnu) module extends SRFI-9 with facilities to return new record instances based on existing ones, only with one or more field values changed—functional

setters. First, the define-immutable-record-type works like define-record-type, ex-

cept that fields are immutable and setters are defined as functional setters.

[Scheme Syntax] define-immutable-record-type type

(constructor fieldname …)

predicate

(fieldname accessor [modifier]) … Define type as a new record type, like define-record-type. However, the record type is made immutable (records may not be mutated, even with struct-set!), and any modifier is defined to be a functional setter—a procedure that returns a new record instance with the specified field changed, and leaves the original unchanged (see example below.)

In addition, the generic set-field and set-fields macros may be applied to any SRFI-9 record.

[Scheme Syntax] set-field record (field sub-fields ...) value Return a new record of record’s type whose fields are equal to the corresponding fields of record except for the one specified by field. field must be the name of the getter corresponding to the field of record being “set”. Subsequent sub-fields must be record getters designating sub-fields within that field value to be set (see example below.)

[Scheme Syntax] set-fields record ((field sub-fields ...) value) ...

Like set-field, but can be used to set more than one field at a time. This expands to code that is more efficient than a series of single set-field calls.

To illustrate the use of functional setters, let’s assume these two record type definitions:

(define-record-type <address>
(address street city country)
address?
(street  address-street)
(city    address-city)
(country address-country))
(define-immutable-record-type <person>
(person age email address)
person?
(age     person-age set-person-age)
(email   person-email set-person-email)
(address person-address set-person-address))

First, note that the <person> record type definition introduces named functional setters. These may be used like this:

(define fsf-address
(address "Franklin Street" "Boston" "USA"))
(define rms
(person 30 "rms@gnu.org" fsf-address))
(and (equal? (set-person-age rms 60)
           (person 60 "rms@gnu.org" fsf-address))
   (= (person-age rms) 30))
⇒ #t

Here, the original <person> record, to which rms is bound, is left unchanged.

Now, suppose we want to change both the street and age of rms. This can be achieved

using set-fields:
(set-fields rms
((person-age) 60)
((person-address address-street) "Temple Place"))
⇒ #<<person> age: 60 email: "rms@gnu.org"
address: #<<address> street: "Temple Place" city: "Boston" country: "USA">>

Notice how the above changed two fields of rms, including the street field of its address field, in a concise way. Also note that set-fields works equally well for types defined with

just define-record-type.

6.7.9 Records

A record type is a first class object representing a user-defined data type. A record is an instance of a record type.

Note that in many ways, this interface is too low-level for every-day use. Most uses of records are better served by SRFI-9 records. See Section 6.7.8 [SRFI-9 Records], page 216.

[Scheme Procedure] record? obj

Return #t if obj is a record of any type and #f otherwise. Note that record? may be true of any Scheme value; there is no promise that records are disjoint with other Scheme types.

[Scheme Procedure] make-record-type type-name field-names [print] Create and return a new record-type descriptor. type-name is a string naming the type. Currently it’s only used in the printed representation of records, and in diagnostics. field-names is a list of symbols naming the fields of a record of the type. Duplicates are not allowed among these symbols.

     (make-record-type "employee" '(name age salary))

The optional print argument is a function used by display, write, etc, for printing a record of the new type. It’s called as (print record port) and should look at record and write to port.

[Scheme Procedure] record-constructor rtd [field-names]

Return a procedure for constructing new members of the type represented by rtd. The returned procedure accepts exactly as many arguments as there are symbols in the given list, field-names; these are used, in order, as the initial values of those fields in a new record, which is returned by the constructor procedure. The values of any fields not named in that list are unspecified. The field-names argument defaults to the list of field names in the call to make-record-type that created the type represented by rtd; if the field-names argument is provided, it is an error if it contains any duplicates or any symbols not in the default list.

[Scheme Procedure] record-predicate rtd

Return a procedure for testing membership in the type represented by rtd. The returned procedure accepts exactly one argument and returns a true value if the argument is a member of the indicated record type; it returns a false value otherwise.

[Scheme Procedure] record-accessor rtd field-name

Return a procedure for reading the value of a particular field of a member of the type represented by rtd. The returned procedure accepts exactly one argument which must be a record of the appropriate type; it returns the current value of the field named by the symbol field-name in that record. The symbol field-name must be a member of the list of field-names in the call to make-record-type that created the type represented by rtd.

[Scheme Procedure] record-modifier rtd field-name

Return a procedure for writing the value of a particular field of a member of the type represented by rtd. The returned procedure accepts exactly two arguments: first, a record of the appropriate type, and second, an arbitrary Scheme value; it modifies the field named by the symbol field-name in that record to contain the given value. The returned value of the modifier procedure is unspecified. The symbol field-name must be a member of the list of field-names in the call to make-record-type that created the type represented by rtd.

[Scheme Procedure] record-type-descriptor record

Return a record-type descriptor representing the type of the given record. That is, for example, if the returned descriptor were passed to record-predicate, the resulting predicate would return a true value when passed the given record. Note that it is not necessarily the case that the returned descriptor is the one that was passed to record-constructor in the call that created the constructor procedure that created the given record.

[Scheme Procedure] record-type-name rtd

Return the type-name associated with the type represented by rtd. The returned value is eqv? to the type-name argument given in the call to make-record-type that created the type represented by rtd.

[Scheme Procedure] record-type-fields rtd

Return a list of the symbols naming the fields in members of the type represented by rtd. The returned value is equal? to the field-names argument given in the call to make-record-type that created the type represented by rtd.

6.7.10 Structures

A structure is a first class data type which holds Scheme values or C words in fields numbered 0 upwards. A vtable is a structure that represents a structure type, giving field types and permissions, and an optional print function for write etc.

Structures are lower level than records (see Section 6.7.9 [Records], page 219). Usually, when you need to represent structured data, you just want to use records. But sometimes you need to implement new kinds of structured data abstractions, and for that purpose structures are useful. Indeed, records in Guile are implemented with structures.

6.7.10.1 Vtables

A vtable is a structure type, specifying its layout, and other information. A vtable is actually itself a structure, but there’s no need to worry about that initially (see Section 6.7.10.3 [Vtable Contents], page 222.)

[Scheme Procedure] make-vtable fields [print]

Create a new vtable. fields is a string describing the fields in the structures to be created. Each field is represented by two characters, a type letter and a permissions letter, for example "pw". The types are as follows.

• p – a Scheme value. “p” stands for “protected” meaning it’s protected against garbage collection.

• u – an arbitrary word of data (an scm_t_bits). At the Scheme level it’s read and written as an unsigned integer. “u” stands for “uninterpreted” (it’s not treated as a Scheme value), or “unprotected” (it’s not marked during GC), or “unsigned long” (its size), or all of these things.

• s – a self-reference. Such a field holds the SCM value of the structure itself (a circular reference). This can be useful in C code where you might have a pointer to the data array, and want to get the Scheme SCM handle for the structure. In Scheme code it has no use.

The second letter for each field is a permission code, • w – writable, the field can be read and written.

• r – read-only, the field can be read but not written.

• o – opaque, the field can be neither read nor written at the Scheme level. This can be used for fields which should only be used from C code.

Here are some examples. See Section 6.7.10.6 [Tail Arrays], page 227, for information on the legacy tail array facility.

     (make-vtable "pw")      ;; one writable field
     (make-vtable "prpw")    ;; one read-only and one writable
     (make-vtable "pwuwuw")  ;; one scheme and two uninterpreted

The optional print argument is a function called by display and write (etc) to give a printed representation of a structure created from this vtable. It’s called (print struct port) and should look at struct and write to port. The default print merely gives a form like ‘#<struct ADDR:ADDR>’ with a pair of machine addresses. The following print function for example shows the two fields of its structure.

     (make-vtable "prpw"
                  (lambda (struct port)
                    (format port "#<~a and ~a>"
                            (struct-ref struct 0)
                            (struct-ref struct 1))))

6.7.10.2 Structure Basics

This section describes the basic procedures for working with structures. make-struct creates a structure, and struct-ref and struct-set! access its fields.

[Scheme Procedure] make-struct vtable tail-size init …

[Scheme Procedure] make-struct/no-tail vtable init …

Create a new structure, with layout per the given vtable (see Section 6.7.10.1 [Vtables], page 220). The optional init… arguments are initial values for the fields of the structure. This is the only way to put values in read-only fields. If there are fewer init arguments than fields then the defaults are #f for a Scheme field (type p) or 0 for an uninterpreted field (type u). Structures also have the ability to allocate a variable number of additional cells at the end, at their tails. However, this legacy tail array facilty is confusing and inefficient, and so we do not recommend it. See Section 6.7.10.6 [Tail Arrays], page 227, for more on the legacy tail array interface.

Type s self-reference fields, permission o opaque fields, and the count field of a tail array are all ignored for the init arguments, ie. an argument is not consumed by such a field. An s is always set to the structure itself, an o is always set to #f or 0 (with the intention that C code will do something to it later), and the tail count is always the given tail-size. For example,

     (define v (make-vtable "prpwpw"))
     (define s (make-struct v 0 123 "abc" 456))
     (struct-ref s 0) ⇒ 123
     (struct-ref s 1) ⇒ "abc"

[C Function] SCM scm_make_struct (SCM vtable, SCM tail size, SCM init list) [C Function] SCM scm_c_make_struct (SCM vtable, SCM tail size, SCM init, ...) [C Function] SCM scm_c_make_structv (SCM vtable, SCM tail size, size t n inits,

scm t bits init[]) There are a few ways to make structures from C. scm_make_struct takes a list, scm_ c_make_struct takes variable arguments terminated with SCM UNDEFINED, and

scm_c_make_structv takes a packed array.

[Scheme Procedure] struct? obj [C Function] scm_struct_p (obj)

Return #t if obj is a structure, or #f if not.

[Scheme Procedure] struct-ref struct n [C Function] scm_struct_ref (struct, n)

Return the contents of field number n in struct. The first field is number 0. An error is thrown if n is out of range, or if the field cannot be read because it’s o opaque.

[Scheme Procedure] struct-set! struct n value

[C Function] scm_struct_set_x (struct, n, value)

Set field number n in struct to value. The first field is number 0. An error is thrown if n is out of range, or if the field cannot be written because it’s r read-only or o opaque.

[Scheme Procedure] struct-vtable struct [C Function] scm_struct_vtable (struct)

Return the vtable that describes struct. The vtable is effectively the type of the structure. See Section 6.7.10.3 [Vtable Contents], page 222, for more on vtables.

6.7.10.3 Vtable Contents

A vtable is itself a structure. It has a specific set of fields describing various aspects of its instances: the structures created from a vtable. Some of the fields are internal to Guile, some of them are part of the public interface, and there may be additional fields added on by the user.

Every vtable has a field for the layout of their instances, a field for the procedure used to print its instances, and a field for the name of the vtable itself. Access to the layout and printer is exposed directly via field indexes. Access to the vtable name is exposed via accessor procedures.

[Scheme Variable] vtable-index-layout
[C Macro] scm_vtable_index_layout

The field number of the layout specification in a vtable. The layout specification is a symbol like pwpw formed from the fields string passed to make-vtable, or created by make-struct-layout (see Section 6.7.10.4 [Meta-Vtables], page 223).

     (define v (make-vtable "pwpw" 0))
     (struct-ref v vtable-index-layout) ⇒ pwpw

This field is read-only, since the layout of structures using a vtable cannot be changed.

[Scheme Variable] vtable-index-printer
[C Macro] scm_vtable_index_printer

The field number of the printer function. This field contains #f if the default print function should be used.

     (define (my-print-func struct port)
       ...)
     (define v (make-vtable "pwpw" my-print-func))
     (struct-ref v vtable-index-printer) ⇒ my-print-func

This field is writable, allowing the print function to be changed dynamically.

[Scheme Procedure] struct-vtable-name vtable [Scheme Procedure] set-struct-vtable-name! vtable name

[C Function] scm_struct_vtable_name (vtable)
[C Function] scm_set_struct_vtable_name_x (vtable, name)

Get or set the name of vtable. name is a symbol and is used in the default print function when printing structures created from vtable.

     (define v (make-vtable "pw"))
     (set-struct-vtable-name! v 'my-name)
     (define s (make-struct v 0))
     (display s) ⊣ #<my-name b7ab3ae0:b7ab3730>

6.7.10.4 Meta-Vtables

As a structure, a vtable also has a vtable, which is also a structure. Structures, their vtables, the vtables of the vtables, and so on form a tree of structures. Making a new structure adds a leaf to the tree, and if that structure is a vtable, it may be used to create other leaves.

If you traverse up the tree of vtables, via calling struct-vtable, eventually you reach a root which is the vtable of itself:

scheme@(guile-user)> (current-module)
$1 = #<directory (guile-user) 221b090>
scheme@(guile-user)> (struct-vtable $1)
$2 = #<record-type module>
scheme@(guile-user)> (struct-vtable $2)
$3 = #<<standard-vtable> 12c30a0>

scheme@(guile-user)> (struct-vtable $3)
$4 = #<<standard-vtable> 12c3fa0>
scheme@(guile-user)> (struct-vtable $4)
$5 = #<<standard-vtable> 12c3fa0>
scheme@(guile-user)> <standard-vtable>
$6 = #<<standard-vtable> 12c3fa0>

In this example, we can say that $1 is an instance of $2, $2 is an instance of $3, $3 is an instance of $4, and $4, strangely enough, is an instance of itself. The value bound to $4 in this console session also bound to <standard-vtable> in the default environment.

[Scheme Variable] <standard-vtable>

A meta-vtable, useful for making new vtables.

All of these values are structures. All but $1 are vtables. As $2 is an instance of $3, and $3 is a vtable, we can say that $3 is a meta-vtable: a vtable that can create vtables.

With this definition, we can specify more precisely what a vtable is: a vtable is a structure made from a meta-vtable. Making a structure from a meta-vtable runs some special checks to ensure that the first field of the structure is a valid layout. Additionally, if these checks see that the layout of the child vtable contains all the required fields of a vtable, in the correct order, then the child vtable will also be a meta-table, inheriting a magical bit from the parent.

[Scheme Procedure] struct-vtable? obj

[C Function] scm_struct_vtable_p (obj)

Return #t if obj is a vtable structure: an instance of a meta-vtable.

<standard-vtable> is a root of the vtable tree. (Normally there is only one root in a given Guile process, but due to some legacy interfaces there may be more than one.)

The set of required fields of a vtable is the set of fields in the <standard-vtable>, and is bound to standard-vtable-fields in the default environment. It is possible to create a meta-vtable that with additional fields in its layout, which can be used to create vtables with additional data:

scheme@(guile-user)> (struct-ref $3 vtable-index-layout)
$6 = pruhsruhpwphuhuhprprpw
scheme@(guile-user)> (struct-ref $4 vtable-index-layout)
$7 = pruhsruhpwphuhuh
scheme@(guile-user)> standard-vtable-fields
$8 = "pruhsruhpwphuhuh"
scheme@(guile-user)> (struct-ref $2 vtable-offset-user)
$9 = module

In this continuation of our earlier example, $2 is a vtable that has extra fields, because its vtable, $3, was made from a meta-vtable with an extended layout. vtable-offset-user is a convenient definition that indicates the number of fields in standard-vtable-fields.

[Scheme Variable] standard-vtable-fields

A string containing the orderedq set of fields that a vtable must have.

[Scheme Variable] vtable-offset-user

The first index in a vtable that is available for a user.

[Scheme Procedure] make-struct-layout fields

[C Function] scm_make_struct_layout (fields)

Return a structure layout symbol, from a fields string. fields is as described under make-vtable (see Section 6.7.10.1 [Vtables], page 220). An invalid fields string is an error.

With these definitions, one can define make-vtable in this way:

     (define* (make-vtable fields #:optional printer)
       (make-struct/no-tail <standard-vtable>
         (make-struct-layout fields)
         printer))

6.7.10.5 Vtable Example

Let us bring these points together with an example. Consider a simple object system with single inheritance. Objects will be normal structures, and classes will be vtables with three extra class fields: the name of the class, the parent class, and the list of fields.

So, first we need a meta-vtable that allocates instances with these extra class fields.

     (define <class>
       (make-vtable
        (string-append standard-vtable-fields "pwpwpw")
        (lambda (x port)
          (format port "<<class> ~a>" (class-name x)))))
     (define (class? x)
       (and (struct? x)
            (eq? (struct-vtable x) <class>)))

To make a structure with a specific meta-vtable, we will use make-struct/no-tail, passing it the computed instance layout and printer, as with make-vtable, and additionally the extra three class fields.

     (define (make-class name parent fields)
       (let* ((fields (compute-fields parent fields))
              (layout (compute-layout fields)))
         (make-struct/no-tail <class>
           layout
           (lambda (x port)
             (print-instance x port))
           name
           parent
           fields)))

Instances will store their associated data in slots in the structure: as many slots as there are fields. The compute-layout procedure below can compute a layout, and field-index returns the slot corresponding to a field.

(define-syntax-rule (define-accessor name n)
  (define (name obj)
    (struct-ref obj n)))
;; Accessors for classes
(define-accessor class-name (+ vtable-offset-user 0))
(define-accessor class-parent (+ vtable-offset-user 1))
(define-accessor class-fields (+ vtable-offset-user 2))
(define (compute-fields parent fields)
  (if parent
      (append (class-fields parent) fields)
      fields))
(define (compute-layout fields)
  (make-struct-layout
   (string-concatenate (make-list (length fields) "pw"))))
(define (field-index class field)
  (list-index (class-fields class) field))
(define (print-instance x port)
  (format port "<~a" (class-name (struct-vtable x)))
  (for-each (lambda (field idx)
              (format port " ~a: ~a" field (struct-ref x idx)))
            (class-fields (struct-vtable x))
            (iota (length (class-fields (struct-vtable x)))))
  (format port ">"))

So, at this point we can actually make a few classes:

(define-syntax-rule (define-class name parent field ...)
  (define name (make-class 'name parent '(field ...))))
(define-class <surface> #f
  width height)
(define-class <window> <surface>
  x y)

And finally, make an instance:

(make-struct/no-tail <window> 400 300 10 20)
⇒ <<window> width: 400 height: 300 x: 10 y: 20>

And that’s that. Note that there are many possible optimizations and feature enhancements that can be made to this object system, and the included GOOPS system does make most of them. For more simple use cases, the records facility is usually sufficient. But sometimes you need to make new kinds of data abstractions, and for that purpose, structs are here.

6.7.10.6 Tail Arrays

Guile’s structures have a facility whereby each instance of a vtable can contain a variablelength tail array of values. The length of the tail array is stored in the structure. This facility was originally intended to allow C code to expose raw C structures with word-sized tail arrays to Scheme.

However, the tail array facility is confusing and doesn’t work very well. It is very rarely used, but it insinuates itself into all invocations of make-struct. For this reason the clumsily-named make-struct/no-tail procedure can actually be more elegant in actual use, because it doesn’t have a random 0 argument stuck in the middle.

Tail arrays also inhibit optimization by allowing instances to affect their shapes. In the absence of tail arrays, all instances of a given vtable have the same number and kinds of fields. This uniformity can be exploited by the runtime and the optimizer. The presence of tail arrays make some of these optimizations more difficult.

Finally, the tail array facility is ad-hoc and does not compose with the rest of Guile. If a Guile user wants an array with user-specified length, it’s best to use a vector. It is more clear in the code, and the standard optimization techniques will do a good job with it.

That said, we should mention some details about the interface. A vtable that has tail array has upper-case permission descriptors: W, R or O, correspoding to tail arrays of writable, read-only, or opaque elements. A tail array permission descriptor may only appear in the last element of a vtable layout.

For exampple, ‘pW’ indicates a tail of writable Scheme-valued fields. The ‘pW’ field itself holds the tail size, and the tail fields come after it.

     (define v (make-vtable "prpW")) ;; one fixed then a tail array
     (define s (make-struct v 6 "fixed field" 'x 'y))
     (struct-ref s 0) ⇒ "fixed field"
     (struct-ref s 1) ⇒ 2    ;; tail size
     (struct-ref s 2) ⇒ x    ;; tail array ...
     (struct-ref s 3) ⇒ y
     (struct-ref s 4) ⇒ #f

6.7.11 Dictionary Types

A dictionary object is a data structure used to index information in a user-defined way. In standard Scheme, the main aggregate data types are lists and vectors. Lists are not really indexed at all, and vectors are indexed only by number (e.g. (vector-ref foo 5)). Often you will find it useful to index your data on some other type; for example, in a library catalog you might want to look up a book by the name of its author. Dictionaries are used to help you organize information in such a way.

An association list (or alist for short) is a list of key-value pairs. Each pair represents a single quantity or object; the car of the pair is a key which is used to identify the object, and the cdr is the object’s value.

A hash table also permits you to index objects with arbitrary keys, but in a way that makes looking up any one object extremely fast. A well-designed hash system makes hash table lookups almost as fast as conventional array or vector references.

Alists are popular among Lisp programmers because they use only the language’s primitive operations (lists, car, cdr and the equality primitives). No changes to the language core are necessary. Therefore, with Scheme’s built-in list manipulation facilities, it is very convenient to handle data stored in an association list. Also, alists are highly portable and can be easily implemented on even the most minimal Lisp systems.

However, alists are inefficient, especially for storing large quantities of data. Because we want Guile to be useful for large software systems as well as small ones, Guile provides a rich set of tools for using either association lists or hash tables.

6.7.12 Association Lists

An association list is a conventional data structure that is often used to implement simple key-value databases. It consists of a list of entries in which each entry is a pair. The key of each entry is the car of the pair and the value of each entry is the cdr.

     ASSOCIATION LIST ::=  '( (KEY1 . VALUE1)
                              (KEY2 . VALUE2)
                              (KEY3 . VALUE3)
                              ...
                            )

Association lists are also known, for short, as alists.

The structure of an association list is just one example of the infinite number of possible structures that can be built using pairs and lists. As such, the keys and values in an association list can be manipulated using the general list structure procedures cons, car, cdr, set-car!, set-cdr! and so on. However, because association lists are so useful, Guile also provides specific procedures for manipulating them.

6.7.12.1 Alist Key Equality

All of Guile’s dedicated association list procedures, apart from acons, come in three flavours, depending on the level of equality that is required to decide whether an existing key in the association list is the same as the key that the procedure call uses to identify the required entry.

• Procedures with assq in their name use eq? to determine key equality.

• Procedures with assv in their name use eqv? to determine key equality.

• Procedures with assoc in their name use equal? to determine key equality.

acons is an exception because it is used to build association lists which do not require their entries’ keys to be unique.

6.7.12.2 Adding or Setting Alist Entries

acons adds a new entry to an association list and returns the combined association list. The combined alist is formed by consing the new entry onto the head of the alist specified in the acons procedure call. So the specified alist is not modified, but its contents become shared with the tail of the combined alist that acons returns.

In the most common usage of acons, a variable holding the original association list is updated with the combined alist:

     (set! address-list (acons name address address-list))

In such cases, it doesn’t matter that the old and new values of address-list share some of their contents, since the old value is usually no longer independently accessible.

Note that acons adds the specified new entry regardless of whether the alist may already contain entries with keys that are, in some sense, the same as that of the new entry. Thus acons is ideal for building alists where there is no concept of key uniqueness.

(set! task-list (acons 3 "pay gas bill" '()))
task-list

⇒

((3 . "pay gas bill"))
(set! task-list (acons 3 "tidy bedroom" task-list))
task-list

⇒

((3 . "tidy bedroom") (3 . "pay gas bill"))

assq-set!, assv-set! and assoc-set! are used to add or replace an entry in an association list where there is a concept of key uniqueness. If the specified association list already contains an entry whose key is the same as that specified in the procedure call, the existing entry is replaced by the new one. Otherwise, the new entry is consed onto the head of the old association list to create the combined alist. In all cases, these procedures return the combined alist.

assq-set! and friends may destructively modify the structure of the old association list in such a way that an existing variable is correctly updated without having to set! it to the value returned:

address-list

⇒

(("mary" . "34 Elm Road") ("james" . "16 Bow Street"))
(assoc-set! address-list "james" "1a London Road")

⇒

(("mary" . "34 Elm Road") ("james" . "1a London Road"))
address-list

⇒

(("mary" . "34 Elm Road") ("james" . "1a London Road"))

Or they may not:

(assoc-set! address-list "bob" "11 Newington Avenue")

⇒

(("bob" . "11 Newington Avenue") ("mary" . "34 Elm Road")
 ("james" . "1a London Road"))
address-list

⇒

(("mary" . "34 Elm Road") ("james" . "1a London Road"))

The only safe way to update an association list variable when adding or replacing an entry like this is to set! the variable to the returned value:

(set! address-list
      (assoc-set! address-list "bob" "11 Newington Avenue"))

     address-list

⇒

     (("bob" . "11 Newington Avenue") ("mary" . "34 Elm Road")
      ("james" . "1a London Road"))

Because of this slight inconvenience, you may find it more convenient to use hash tables to store dictionary data. If your application will not be modifying the contents of an alist very often, this may not make much difference to you.

If you need to keep the old value of an association list in a form independent from the list that results from modification by acons, assq-set!, assv-set! or assoc-set!, use list-copy to copy the old association list before modifying it.

[Scheme Procedure] acons key value alist [C Function] scm_acons (key, value, alist)

Add a new key-value pair to alist. A new pair is created whose car is key and whose cdr is value, and the pair is consed onto alist, and the new list is returned. This function is not destructive; alist is not modified.

[Scheme Procedure] assq-set! alist key val [Scheme Procedure] assv-set! alist key value [Scheme Procedure] assoc-set! alist key value [C Function] scm_assq_set_x (alist, key, val) [C Function] scm_assv_set_x (alist, key, val)

[C Function] scm_assoc_set_x (alist, key, val)

Reassociate key in alist with value: find any existing alist entry for key and associate it with the new value. If alist does not contain an entry for key, add a new one. Return the (possibly new) alist. These functions do not attempt to verify the structure of alist, and so may cause unusual results if passed an object that is not an association list.

6.7.12.3 Retrieving Alist Entries

assq, assv and assoc find the entry in an alist for a given key, and return the (key . value) pair. assq-ref, assv-ref and assoc-ref do a similar lookup, but return just the value.

[Scheme Procedure] assq key alist [Scheme Procedure] assv key alist [Scheme Procedure] assoc key alist [C Function] scm_assq (key, alist) [C Function] scm_assv (key, alist) [C Function] scm_assoc (key, alist)

Return the first entry in alist with the given key. The return is the pair (KEY . VALUE) from alist. If there’s no matching entry the return is #f. assq compares keys with eq?, assv uses eqv? and assoc uses equal?. See also SRFI-1 which has an extended assoc (Section 7.5.3.9 [SRFI-1 Association Lists], page 562).

[Scheme Procedure] assq-ref alist key [Scheme Procedure] assv-ref alist key [Scheme Procedure] assoc-ref alist key [C Function] scm_assq_ref (alist, key) [C Function] scm_assv_ref (alist, key) [C Function] scm_assoc_ref (alist, key)

Return the value from the first entry in alist with the given key, or #f if there’s no such entry.

assq-ref compares keys with eq?, assv-ref uses eqv? and assoc-ref uses equal?.

Notice these functions have the key argument last, like other -ref functions, but this is opposite to what assq etc above use. When the return is #f it can be either key not found, or an entry which happens to have value #f in the cdr. Use assq etc above if you need to differentiate these cases.

6.7.12.4 Removing Alist Entries

To remove the element from an association list whose key matches a specified key, use assqremove!, assv-remove! or assoc-remove! (depending, as usual, on the level of equality required between the key that you specify and the keys in the association list).

As with assq-set! and friends, the specified alist may or may not be modified destructively, and the only safe way to update a variable containing the alist is to set! it to the value that assq-remove! and friends return.

address-list

⇒

(("bob" . "11 Newington Avenue") ("mary" . "34 Elm Road")
 ("james" . "1a London Road"))
(set! address-list (assoc-remove! address-list "mary"))
address-list

⇒

(("bob" . "11 Newington Avenue") ("james" . "1a London Road"))

Note that, when assq/v/oc-remove! is used to modify an association list that has been constructed only using the corresponding assq/v/oc-set!, there can be at most one matching entry in the alist, so the question of multiple entries being removed in one go does not arise. If assq/v/oc-remove! is applied to an association list that has been constructed using acons, or an assq/v/oc-set! with a different level of equality, or any mixture of these, it removes only the first matching entry from the alist, even if the alist might contain further matching entries. For example:

(define address-list '())
(set! address-list (assq-set! address-list "mary" "11 Elm Street"))
(set! address-list (assq-set! address-list "mary" "57 Pine Drive"))
address-list

⇒

(("mary" . "57 Pine Drive") ("mary" . "11 Elm Street"))
(set! address-list (assoc-remove! address-list "mary"))
address-list

⇒

     (("mary" . "11 Elm Street"))

In this example, the two instances of the string "mary" are not the same when compared using eq?, so the two assq-set! calls add two distinct entries to address-list. When compared using equal?, both "mary"s in address-list are the same as the "mary" in the assoc-remove! call, but assoc-remove! stops after removing the first matching entry that it finds, and so one of the "mary" entries is left in place.

[Scheme Procedure] assq-remove! alist key [Scheme Procedure] assv-remove! alist key [Scheme Procedure] assoc-remove! alist key [C Function] scm_assq_remove_x (alist, key) [C Function] scm_assv_remove_x (alist, key) [C Function] scm_assoc_remove_x (alist, key)

Delete the first entry in alist associated with key, and return the resulting alist.

6.7.12.5 Sloppy Alist Functions

sloppy-assq, sloppy-assv and sloppy-assoc behave like the corresponding non-sloppy-

procedures, except that they return #f when the specified association list is not well-formed, where the non-sloppy- versions would signal an error.

Specifically, there are two conditions for which the non-sloppy- procedures signal an error, which the sloppy- procedures handle instead by returning #f. Firstly, if the specified alist as a whole is not a proper list:

     (assoc "mary" '((1 . 2) ("key" . "door") . "open sesame"))

⇒

     ERROR: In procedure assoc in expression (assoc "mary" (quote #)):
     ERROR: Wrong type argument in position 2 (expecting
        association list): ((1 . 2) ("key" . "door") . "open sesame")
     (sloppy-assoc "mary" '((1 . 2) ("key" . "door") . "open sesame"))

⇒

     #f

Secondly, if one of the entries in the specified alist is not a pair:

     (assoc 2 '((1 . 1) 2 (3 . 9)))

⇒

     ERROR: In procedure assoc in expression (assoc 2 (quote #)):
     ERROR: Wrong type argument in position 2 (expecting
        association list): ((1 . 1) 2 (3 . 9))
     (sloppy-assoc 2 '((1 . 1) 2 (3 . 9)))

⇒

     #f

Unless you are explicitly working with badly formed association lists, it is much safer to use the non-sloppy- procedures, because they help to highlight coding and data errors that the sloppy- versions would silently cover up.

[Scheme Procedure] sloppy-assq key alist [C Function] scm_sloppy_assq (key, alist)

Behaves like assq but does not do any error checking. Recommended only for use in Guile internals.

[Scheme Procedure] sloppy-assv key alist [C Function] scm_sloppy_assv (key, alist)

Behaves like assv but does not do any error checking. Recommended only for use in Guile internals.

[Scheme Procedure] sloppy-assoc key alist [C Function] scm_sloppy_assoc (key, alist)

Behaves like assoc but does not do any error checking. Recommended only for use in Guile internals.

6.7.12.6 Alist Example

Here is a longer example of how alists may be used in practice.

(define capitals '(("New York" . "Albany")
                   ("Oregon"   . "Salem")
                   ("Florida"  . "Miami")))
;; What's the capital of Oregon?
(assoc "Oregon" capitals)  ⇒ ("Oregon" . "Salem")
(assoc-ref capitals "Oregon") ⇒ "Salem"
;; We left out South Dakota.
(set! capitals
      (assoc-set! capitals "South Dakota" "Pierre"))
capitals
⇒ (("South Dakota" . "Pierre")
    ("New York" . "Albany")
    ("Oregon" . "Salem")
    ("Florida" . "Miami"))
;; And we got Florida wrong.
(set! capitals
      (assoc-set! capitals "Florida" "Tallahassee"))
capitals
⇒ (("South Dakota" . "Pierre")
    ("New York" . "Albany")
    ("Oregon" . "Salem")
    ("Florida" . "Tallahassee"))
;; After Oregon secedes, we can remove it.
(set! capitals
      (assoc-remove! capitals "Oregon"))
capitals

⇒ (("South Dakota" . "Pierre")
    ("New York" . "Albany")
    ("Florida" . "Tallahassee"))

6.7.13 VList-Based Hash Lists or “VHashes”

The (ice-9 vlist) module provides an implementation of VList-based hash lists (see Section 6.7.6 [VLists], page 213). VList-based hash lists, or vhashes, are an immutable dictionary type similar to association lists that maps keys to values. However, unlike association lists, accessing a value given its key is typically a constant-time operation.

The VHash programming interface of (ice-9 vlist) is mostly the same as that of association lists found in SRFI-1, with procedure names prefixed by vhash- instead of alist- (see Section 7.5.3.9 [SRFI-1 Association Lists], page 562).

In addition, vhashes can be manipulated using VList operations:

(vlist-head (vhash-consq 'a 1 vlist-null))
⇒ (a . 1)
(define vh1 (vhash-consq 'b 2 (vhash-consq 'a 1 vlist-null)))
(define vh2 (vhash-consq 'c 3 (vlist-tail vh1)))
(vhash-assq 'a vh2)
⇒ (a . 1)
(vhash-assq 'b vh2)
⇒ #f
(vhash-assq 'c vh2)
⇒ (c . 3)
(vlist->list vh2)
⇒ ((c . 3) (a . 1))

However, keep in mind that procedures that construct new VLists (vlist-map, vlistfilter, etc.) return raw VLists, not vhashes:

(define vh (alist->vhash '((a . 1) (b . 2) (c . 3)) hashq))
(vhash-assq 'a vh)
⇒ (a . 1)
(define vl
  ;; This will create a raw vlist.
  (vlist-filter (lambda (key+value) (odd? (cdr key+value))) vh))
(vhash-assq 'a vl)
⇒ ERROR: Wrong type argument in position 2
(vlist->list vl)
⇒ ((a . 1) (c . 3))

[Scheme Procedure] vhash? obj

Return true if obj is a vhash.

[Scheme Procedure] vhash-cons key value vhash [hash-proc] [Scheme Procedure] vhash-consq key value vhash [Scheme Procedure] vhash-consv key value vhash

Return a new hash list based on vhash where key is associated with value, using hash-proc to compute the hash of key. vhash must be either vlist-null or a vhash returned by a previous call to vhash-cons. hash-proc defaults to hash (see Section 6.7.14.2 [Hash Table Reference], page 237). With vhash-consq, the hashq hash function is used; with vhash-consv the hashv hash function is used. All vhash-cons calls made to construct a vhash should use the same hash-proc. Failing to do that, the result is undefined.

[Scheme Procedure] vhash-assoc key vhash [equal? [hash-proc]]

[Scheme Procedure] vhash-assq key vhash [Scheme Procedure] vhash-assv key vhash

Return the first key/value pair from vhash whose key is equal to key according to the equal? equality predicate (which defaults to equal?), and using hash-proc (which defaults to hash) to compute the hash of key. The second form uses eq? as the equality predicate and hashq as the hash function; the last form uses eqv? and hashv. Note that it is important to consistently use the same hash function for hash-proc as was passed to vhash-cons. Failing to do that, the result is unpredictable.

[Scheme Procedure] vhash-delete key vhash [equal? [hash-proc]]

[Scheme Procedure] vhash-delq key vhash [Scheme Procedure] vhash-delv key vhash

Remove all associations from vhash with key, comparing keys with equal? (which defaults to equal?), and computing the hash of key using hash-proc (which defaults to hash). The second form uses eq? as the equality predicate and hashq as the hash function; the last one uses eqv? and hashv. Again the choice of hash-proc must be consistent with previous calls to vhash-cons.

[Scheme Procedure] vhash-fold proc init vhash

[Scheme Procedure] vhash-fold-right proc init vhash

Fold over the key/value elements of vhash in the given direction, with each call to proc having the form (proc key value result), where result is the result of the previous call to proc and init the value of result for the first call to proc.

[Scheme Procedure] vhash-fold* proc init key vhash [equal? [hash]] [Scheme Procedure] vhash-foldq* proc init key vhash [Scheme Procedure] vhash-foldv* proc init key vhash

Fold over all the values associated with key in vhash, with each call to proc having the form (proc value result), where result is the result of the previous call to proc and init the value of result for the first call to proc. Keys in vhash are hashed using hash are compared using equal?. The second form uses eq? as the equality predicate and hashq as the hash function; the third one uses

eqv? and hashv.

Example:

     (define vh
       (alist->vhash '((a . 1) (a . 2) (z . 0) (a . 3))))

     (vhash-fold* cons '() 'a vh)
     ⇒ (3 2 1)
     (vhash-fold* cons '() 'z vh)
     ⇒ (0)

[Scheme Procedure] alist->vhash alist [hash-proc]

Return the vhash corresponding to alist, an association list, using hash-proc to compute key hashes. When omitted, hash-proc defaults to hash.

6.7.14 Hash Tables

Hash tables are dictionaries which offer similar functionality as association lists: They provide a mapping from keys to values. The difference is that association lists need time linear in the size of elements when searching for entries, whereas hash tables can normally search in constant time. The drawback is that hash tables require a little bit more memory, and that you can not use the normal list procedures (see Section 6.7.2 [Lists], page 188) for working with them.

6.7.14.1 Hash Table Examples

For demonstration purposes, this section gives a few usage examples of some hash table procedures, together with some explanation what they do.

First we start by creating a new hash table with 31 slots, and populate it with two key/value pairs.

(define h (make-hash-table 31))
;; This is an opaque object
h

⇒

#<hash-table 0/31>
;; Inserting into a hash table can be done with hashq-set!
(hashq-set! h 'foo "bar")

⇒

"bar"
(hashq-set! h 'braz "zonk")

⇒

"zonk"
;; Or with hash-create-handle!
(hashq-create-handle! h 'frob #f)

⇒

(frob . #f)

You can get the value for a given key with the procedure hashq-ref, but the problem with this procedure is that you cannot reliably determine whether a key does exists in the table. The reason is that the procedure returns #f if the key is not in the table, but it will return the same value if the key is in the table and just happens to have the value #f, as you can see in the following examples.

(hashq-ref h 'foo)

⇒

"bar"
(hashq-ref h 'frob)

⇒

#f
(hashq-ref h 'not-there)

⇒

#f

Better is to use the procedure hashq-get-handle, which makes a distinction between the two cases. Just like assq, this procedure returns a key/value-pair on success, and #f if the key is not found.

(hashq-get-handle h 'foo)

⇒

(foo . "bar")
(hashq-get-handle h 'not-there)

⇒

#f

Interesting results can be computed by using hash-fold to work through each element. This example will count the total number of elements:

(hash-fold (lambda (key value seed) (+ 1 seed)) 0 h)

⇒

3

The same thing can be done with the procedure hash-count, which can also count the number of elements matching a particular predicate. For example, count the number of elements with string values:

(hash-count (lambda (key value) (string? value)) h)

⇒

2

Counting all the elements is a simple task using const:

(hash-count (const #t) h)

⇒

3

6.7.14.2 Hash Table Reference

Like the association list functions, the hash table functions come in several varieties, according to the equality test used for the keys. Plain hash- functions use equal?, hashqfunctions use eq?, hashv- functions use eqv?, and the hashx- functions use an application supplied test.

A single make-hash-table creates a hash table suitable for use with any set of functions, but it’s imperative that just one set is then used consistently, or results will be unpredictable.

Hash tables are implemented as a vector indexed by a hash value formed from the key, with an association list of key/value pairs for each bucket in case distinct keys hash together. Direct access to the pairs in those lists is provided by the -handle- functions.

When the number of entries in a hash table goes above a threshold, the vector is made larger and the entries are rehashed, to prevent the bucket lists from becoming too long and slowing down accesses. When the number of entries goes below a threshold, the vector is shrunk to save space.

For the hashx- “extended” routines, an application supplies a hash function producing an integer index like hashq etc below, and an assoc alist search function like assq etc (see Section 6.7.12.3 [Retrieving Alist Entries], page 230). Here’s an example of such functions implementing case-insensitive hashing of string keys,

     (use-modules (srfi srfi-1)
                  (srfi srfi-13))
     (define (my-hash str size)
       (remainder (string-hash-ci str) size))
     (define (my-assoc str alist)
       (find (lambda (pair) (string-ci=? str (car pair))) alist))
     (define my-table (make-hash-table))
     (hashx-set! my-hash my-assoc my-table "foo" 123)
     (hashx-ref my-hash my-assoc my-table "FOO")
     ⇒ 123

In a hashx- hash function the aim is to spread keys across the vector, so bucket lists don’t become long. But the actual values are arbitrary as long as they’re in the range 0 to size −1. Helpful functions for forming a hash value, in addition to hashq etc below, include symbolhash (see Section 6.6.7.2 [Symbol Keys], page 172), string-hash and string-hash-ci (see Section 6.6.5.7 [String Comparison], page 148), and char-set-hash (see Section 6.6.4.1 [Character Set Predicates/Comparison], page 133).

[Scheme Procedure] make-hash-table [size]

Create a new hash table object, with an optional minimum vector size.

When size is given, the table vector will still grow and shrink automatically, as described above, but with size as a minimum. If an application knows roughly how many entries the table will hold then it can use size to avoid rehashing when initial entries are added.

[Scheme Procedure] alist->hash-table alist [Scheme Procedure] alist->hashq-table alist [Scheme Procedure] alist->hashv-table alist [Scheme Procedure] alist->hashx-table hash assoc alist

Convert alist into a hash table. When keys are repeated in alist, the leftmost association takes precedence.

          (use-modules (ice-9 hash-table))
          (alist->hash-table '((foo . 1) (bar . 2)))

When converting to an extended hash table, custom hash and assoc procedures must be provided.

          (alist->hashx-table hash assoc '((foo . 1) (bar . 2)))

[Scheme Procedure] hash-table? obj

[C Function] scm_hash_table_p (obj)

Return #t if obj is a abstract hash table object.

[Scheme Procedure] hash-clear! table [C Function] scm_hash_clear_x (table)

Remove all items from table (without triggering a resize).

[Scheme Procedure] hash-ref table key [dflt]

[Scheme Procedure] hashq-ref table key [dflt] [Scheme Procedure] hashv-ref table key [dflt]

[Scheme Procedure] hashx-ref hash assoc table key [dflt] [C Function] scm_hash_ref (table, key, dflt) [C Function] scm_hashq_ref (table, key, dflt) [C Function] scm_hashv_ref (table, key, dflt) [C Function] scm_hashx_ref (hash, assoc, table, key, dflt)

Lookup key in the given hash table, and return the associated value. If key is not found, return dflt, or #f if dflt is not given.

[Scheme Procedure] hash-set! table key val [Scheme Procedure] hashq-set! table key val [Scheme Procedure] hashv-set! table key val [Scheme Procedure] hashx-set! hash assoc table key val [C Function] scm_hash_set_x (table, key, val)

[C Function] scm_hashq_set_x (table, key, val) [C Function] scm_hashv_set_x (table, key, val)

[C Function] scm_hashx_set_x (hash, assoc, table, key, val)

Associate val with key in the given hash table. If key is already present then it’s associated value is changed. If it’s not present then a new entry is created.

[Scheme Procedure] hash-remove! table key [Scheme Procedure] hashq-remove! table key [Scheme Procedure] hashv-remove! table key

[Scheme Procedure] hashx-remove! hash assoc table key

[C Function] scm_hash_remove_x (table, key) [C Function] scm_hashq_remove_x (table, key) [C Function] scm_hashv_remove_x (table, key)

[C Function] scm_hashx_remove_x (hash, assoc, table, key)

Remove any association for key in the given hash table. If key is not in table then nothing is done.

[Scheme Procedure] hash key size [Scheme Procedure] hashq key size [Scheme Procedure] hashv key size [C Function] scm_hash (key, size) [C Function] scm_hashq (key, size) [C Function] scm_hashv (key, size)

Return a hash value for key. This is a number in the range 0 to size − 1, which is suitable for use in a hash table of the given size. Note that hashq and hashv may use internal addresses of objects, so if an object is garbage collected and re-created it can have a different hash value, even when the two are notionally eq?. For instance with symbols,

     (hashq 'something 123) ⇒ 19
     (gc)
     (hashq 'something 123) ⇒ 62

In normal use this is not a problem, since an object entered into a hash table won’t be garbage collected until removed. It’s only if hashing calculations are somehow separated from normal references that its lifetime needs to be considered.

[Scheme Procedure] hash-get-handle table key [Scheme Procedure] hashq-get-handle table key [Scheme Procedure] hashv-get-handle table key

[Scheme Procedure] hashx-get-handle hash assoc table key

[C Function] scm_hash_get_handle (table, key) [C Function] scm_hashq_get_handle (table, key) [C Function] scm_hashv_get_handle (table, key)

[C Function] scm_hashx_get_handle (hash, assoc, table, key)

Return the (key . value) pair for key in the given hash table, or #f if key is not in table.

[Scheme Procedure] hash-create-handle! table key init [Scheme Procedure] hashq-create-handle! table key init [Scheme Procedure] hashv-create-handle! table key init

[Scheme Procedure] hashx-create-handle! hash assoc table key init

[C Function] scm_hash_create_handle_x (table, key, init) [C Function] scm_hashq_create_handle_x (table, key, init) [C Function] scm_hashv_create_handle_x (table, key, init)

[C Function] scm_hashx_create_handle_x (hash, assoc, table, key, init)

Return the (key . value) pair for key in the given hash table. If key is not in table then create an entry for it with init as the value, and return that pair.

[Scheme Procedure] hash-map->list proc table [Scheme Procedure] hash-for-each proc table [C Function] scm_hash_map_to_list (proc, table) [C Function] scm_hash_for_each (proc, table)

Apply proc to the entries in the given hash table. Each call is (proc key value). hash-map->list returns a list of the results from these calls, hash-for-each discards the results and returns an unspecified value. Calls are made over the table entries in an unspecified order, and for hash-map->list the order of the values in the returned list is unspecified. Results will be unpredictable if table is modified while iterating.

For example the following returns a new alist comprising all the entries from mytable, in no particular order.

     (hash-map->list cons mytable)

[Scheme Procedure] hash-for-each-handle proc table

[C Function] scm_hash_for_each_handle (proc, table)

Apply proc to the entries in the given hash table. Each call is (proc handle), where handle is a (key . value) pair. Return an unspecified value. hash-for-each-handle differs from hash-for-each only in the argument list of proc.

[Scheme Procedure] hash-fold proc init table [C Function] scm_hash_fold (proc, init, table)

Accumulate a result by applying proc to the elements of the given hash table. Each call is (proc key value prior-result), where key and value are from the table and prior-result is the return from the previous proc call. For the first call, prior-result is the given init value. Calls are made over the table entries in an unspecified order. Results will be unpredictable if table is modified while hash-fold is running. For example, the following returns a count of how many keys in mytable are strings.

     (hash-fold (lambda (key value prior)
                  (if (string? key) (1+ prior) prior))
                0 mytable)

[Scheme Procedure] hash-count pred table [C Function] scm_hash_count (pred, table)

Return the number of elements in the given hash table that cause (pred key value) to return true. To quickly determine the total number of elements, use (const #t) for pred.

6.8 Smobs

This chapter contains reference information related to defining and working with smobs. See Section 5.5 [Defining New Types (Smobs)], page 71 for a tutorial-like introduction to smobs.

[Function] scm_t_bits scm_make_smob_type (const char *name, size t size)

This function adds a new smob type, named name, with instance size size, to the system. The return value is a tag that is used in creating instances of the type. If size is 0, the default free function will do nothing. If size is not 0, the default free function will deallocate the memory block pointed to by

SCM_SMOB_DATA with scm_gc_free. The what parameter in the call to scm_gc_free

will be name. Default values are provided for the mark, free, print, and equalp functions, as described in Section 5.5 [Defining New Types (Smobs)], page 71. If you want to customize any of these functions, the call to scm_make_smob_type should be immediately followed by

calls to one or several of scm_set_smob_mark, scm_set_smob_free, scm_set_smob_
print, and/or scm_set_smob_equalp.

[C Function] void scm_set_smob_free (scm t bits tc, size t (*free) (SCM obj))

This function sets the smob freeing procedure (sometimes referred to as a finalizer) for the smob type specified by the tag tc. tc is the tag returned by scm_make_smob_type. The free procedure must deallocate all resources that are directly associated with the smob instance obj. It must assume that all SCM values that it references have already been freed and are thus invalid. It must also not call any libguile function or macro except scm_gc_free, SCM_SMOB_

FLAGS, SCM_SMOB_DATA, SCM_SMOB_DATA_2, and SCM_SMOB_DATA_3.

The free procedure must return 0. Note that defining a freeing procedure is not necessary if the resources associated with obj consists only of memory allocated with scm_gc_malloc or scm_gc_malloc_ pointerless because this memory is automatically reclaimed by the garbage collector when it is no longer needed (see Section 6.18.2 [Memory Blocks], page 376).

[C Function] void scm_set_smob_mark (scm t bits tc, SCM (*mark) (SCM obj))

This function sets the smob marking procedure for the smob type specified by the tag tc. tc is the tag returned by scm_make_smob_type. Defining a marking procedure may sometimes be unnecessary because large parts of the process’ memory (with the exception of scm_gc_malloc_pointerless regions, and malloc- or scm_malloc-allocated memory) are scanned for live pointers^{4}. The mark procedure must cause scm_gc_mark to be called for every SCM value that is directly referenced by the smob instance obj. One of these SCM values can be returned from the procedure and Guile will call scm_gc_mark for it. This can be used to avoid deep recursions for smob instances that form a list. It must not call any libguile function or macro except scm_gc_mark, SCM_SMOB_FLAGS,

SCM_SMOB_DATA, SCM_SMOB_DATA_2, and SCM_SMOB_DATA_3.

[C Function] void scm_set_smob_print (scm t bits tc, int (*print) (SCM obj,

SCM port, scm print state* pstate)) This function sets the smob printing procedure for the smob type specified by the tag tc. tc is the tag returned by scm_make_smob_type. The print procedure should output a textual representation of the smob instance obj to port, using information in pstate. The textual representation should be of the form #<name ...>. This ensures that read will not interpret it as some other Scheme value. It is often best to ignore pstate and just print to port with scm_display, scm_write,

scm_simple_format, and scm_puts.

[C Function] void scm_set_smob_equalp (scm t bits tc, SCM (*equalp) (SCM

obj1, SCM obj2)) This function sets the smob equality-testing predicate for the smob type specified by the tag tc. tc is the tag returned by scm_make_smob_type.

4 Conversely, in Guile up to the 1.8 series, the marking procedure was always required. The reason is that

Guile’s GC would only look for pointers in the memory area used for built-in types (the cell heap), not in user-allocated or statically allocated memory. This approach is often referred to as precise marking.

The equalp procedure should return SCM_BOOL_T when obj1 is equal? to obj2. Else it should return SCM_BOOL_F. Both obj1 and obj2 are instances of the smob type tc.

[C Function] void scm_assert_smob_type (scm t bits tag, SCM val)

When val is a smob of the type indicated by tag, do nothing. Else, signal an error.

[C Macro] int SCM_SMOB_PREDICATE (scm t bits tag, SCM exp)

Return true if exp is a smob instance of the type indicated by tag, or false otherwise. The expression exp can be evaluated more than once, so it shouldn’t contain any side effects.

[C Function] SCM scm_new_smob (scm t bits tag, void *data) [C Function] SCM scm_new_double_smob (scm t bits tag, void *data, void *data2,

void *data3) Make a new smob of the type with tag tag and smob data data, data2, and data3, as appropriate. The tag is what has been returned by scm_make_smob_type. The initial values data, data2, and data3 are of type scm_t_bits; when you want to use them for SCM values, these values need to be converted to a scm_t_bits first by using SCM_UNPACK. The flags of the smob instance start out as zero.

[C Macro] scm_t_bits SCM_SMOB_FLAGS (SCM obj)

Return the 16 extra bits of the smob obj. No meaning is predefined for these bits, you can use them freely.

[C Macro] scm_t_bits SCM_SET_SMOB_FLAGS (SCM obj, scm t bits flags)

Set the 16 extra bits of the smob obj to flags. No meaning is predefined for these bits, you can use them freely.

[C Macro] scm_t_bits SCM_SMOB_DATA (SCM obj)
[C Macro] scm_t_bits SCM_SMOB_DATA_2 (SCM obj)
[C Macro] scm_t_bits SCM_SMOB_DATA_3 (SCM obj)

Return the first (second, third) immediate word of the smob obj as a scm_t_bits value. When the word contains a SCM value, use SCM_SMOB_OBJECT (etc.) instead.

[C Macro] void SCM_SET_SMOB_DATA (SCM obj, scm t bits val) [C Macro] void SCM_SET_SMOB_DATA_2 (SCM obj, scm t bits val) [C Macro] void SCM_SET_SMOB_DATA_3 (SCM obj, scm t bits val)

Set the first (second, third) immediate word of the smob obj to val. When the word should be set to a SCM value, use SCM_SMOB_SET_OBJECT (etc.) instead.

[C Macro] SCM SCM_SMOB_OBJECT (SCM obj)
[C Macro] SCM SCM_SMOB_OBJECT_2 (SCM obj)
[C Macro] SCM SCM_SMOB_OBJECT_3 (SCM obj)

Return the first (second, third) immediate word of the smob obj as a SCM value. When the word contains a scm_t_bits value, use SCM_SMOB_DATA (etc.) instead.

[C Macro] void SCM_SET_SMOB_OBJECT (SCM obj, SCM val)
[C Macro] void SCM_SET_SMOB_OBJECT_2 (SCM obj, SCM val)

[C Macro] void SCM_SET_SMOB_OBJECT_3 (SCM obj, SCM val)

Set the first (second, third) immediate word of the smob obj to val. When the word should be set to a scm_t_bits value, use SCM_SMOB_SET_DATA (etc.) instead.

[C Macro] SCM * SCM_SMOB_OBJECT_LOC (SCM obj)
[C Macro] SCM * SCM_SMOB_OBJECT_2_LOC (SCM obj)
[C Macro] SCM * SCM_SMOB_OBJECT_3_LOC (SCM obj)

Return a pointer to the first (second, third) immediate word of the smob obj. Note that this is a pointer to SCM. If you need to work with scm_t_bits values, use

SCM_PACK and SCM_UNPACK, as appropriate.
[Function] SCM scm_markcdr (SCM x)

Mark the references in the smob x, assuming that x’s first data word contains an ordinary Scheme object, and x refers to no other objects. This function simply returns x’s first data word.

6.9 Procedures

6.9.1 Lambda: Basic Procedure Creation

A lambda expression evaluates to a procedure. The environment which is in effect when a lambda expression is evaluated is enclosed in the newly created procedure, this is referred to as a closure (see Section 3.4 [About Closure], page 26).

When a procedure created by lambda is called with some actual arguments, the environment enclosed in the procedure is extended by binding the variables named in the formal argument list to new locations and storing the actual arguments into these locations. Then the body of the lambda expression is evaluated sequentially. The result of the last expression in the procedure body is then the result of the procedure invocation.

The following examples will show how procedures can be created using lambda, and what you can do with these procedures.

(lambda (x) (+ x x))  ⇒ a procedure
((lambda (x) (+ x x)) 4) ⇒ 8

The fact that the environment in effect when creating a procedure is enclosed in the procedure is shown with this example:

(define add4
  (let ((x 4))
    (lambda (y) (+ x y))))
(add4 6)  ⇒ 10

[syntax] lambda formals body formals should be a formal argument list as described in the following table.

(variable1 ...)

The procedure takes a fixed number of arguments; when the procedure is called, the arguments will be stored into the newly created location for the formal variables. variable The procedure takes any number of arguments; when the procedure is

called, the sequence of actual arguments will converted into a list and stored into the newly created location for the formal variable.

(variable1 ... variablen . variablen+1)

If a space-delimited period precedes the last variable, then the procedure takes n or more variables where n is the number of formal arguments before the period. There must be at least one argument before the period. The first n actual arguments will be stored into the newly allocated locations for the first n formal arguments and the sequence of the remaining actual arguments is converted into a list and the stored into the location for the last formal argument. If there are exactly n actual arguments, the empty list is stored into the location of the last formal argument.

The list in variable or variablen+1 is always newly created and the procedure can modify it if desired. This is the case even when the procedure is invoked via apply, the required part of the list argument there will be copied (see Section 6.17.4 [Procedures for On the Fly Evaluation], page 362).

body is a sequence of Scheme expressions which are evaluated in order when the procedure is invoked.

6.9.2 Primitive Procedures

Procedures written in C can be registered for use from Scheme, provided they take only arguments of type SCM and return SCM values. scm_c_define_gsubr is likely to be the most useful mechanism, combining the process of registration (scm_c_make_gsubr) and definition

(scm_define).

[Function] SCM scm_c_make_gsubr (const char *name, int req, int opt, int rst, fcn)

Register a C procedure fcn as a “subr” — a primitive subroutine that can be called from Scheme. It will be associated with the given name but no environment binding will be created. The arguments req, opt and rst specify the number of required, optional and “rest” arguments respectively. The total number of these arguments should match the actual number of arguments to fcn, but may not exceed 10. The number of rest arguments should be 0 or 1. scm_c_make_gsubr returns a value of type SCM which is a “handle” for the procedure.

[Function] SCM scm_c_define_gsubr (const char *name, int req, int opt, int rst,

fcn) Register a C procedure fcn, as for scm_c_make_gsubr above, and additionally create a top-level Scheme binding for the procedure in the “current environment” using scm_ define. scm_c_define_gsubr returns a handle for the procedure in the same way as scm_c_make_gsubr, which is usually not further required.

6.9.3 Compiled Procedures

The evaluation strategy given in Section 6.9.1 [Lambda], page 244 describes how procedures are interpreted. Interpretation operates directly on expanded Scheme source code, recursively calling the evaluator to obtain the value of nested expressions.

Most procedures are compiled, however. This means that Guile has done some precomputation on the procedure, to determine what it will need to do each time the procedure runs. Compiled procedures run faster than interpreted procedures.

Loading files is the normal way that compiled procedures come to being. If Guile sees that a file is uncompiled, or that its compiled file is out of date, it will attempt to compile the file when it is loaded, and save the result to disk. Procedures can be compiled at runtime as well. See Section 6.17 [Read/Load/Eval/Compile], page 357, for more information on runtime compilation.

Compiled procedures, also known as programs, respond all procedures that operate on procedures. In addition, there are a few more accessors for low-level details on programs.

Most people won’t need to use the routines described in this section, but it’s good to have them documented. You’ll have to include the appropriate module first, though:

(use-modules (system vm program))

[Scheme Procedure] program? obj [C Function] scm_program_p (obj)

Returns #t if obj is a compiled procedure, or #f otherwise.

[Scheme Procedure] program-objcode program [C Function] scm_program_objcode (program)

Returns the object code associated with this program. See Section 9.4.6 [Bytecode and Objcode], page 803, for more information.

[Scheme Procedure] program-objects program [C Function] scm_program_objects (program)

Returns the “object table” associated with this program, as a vector. See Section 9.3.5 [VM Programs], page 776, for more information.

[Scheme Procedure] program-module program [C Function] scm_program_module (program)

Returns the module that was current when this program was created. Can return #f if the compiler could determine that this information was unnecessary.

[Scheme Procedure] program-free-variables program

[C Function] scm_program_free_variables (program)

Returns the set of free variables that this program captures in its closure, as a vector. If a closure is code with data, you can get the code from program-objcode, and the

data via program-free-variables.

Some of the values captured are actually in variable “boxes”. See Section 9.3.4 [Variables and the VM], page 775, for more information. Users must not modify the returned value unless they think they’re really clever.

[Scheme Procedure] program-meta program [C Function] scm_program_meta (program)

Return the metadata thunk of program, or #f if it has no metadata. When called, a metadata thunk returns a list of the following form: (bindings sources arities . properties). The format of each of these elements is discussed below.

[Scheme Procedure] program-bindings program

[Scheme Procedure] make-binding name boxed? index start end

[Scheme Procedure] binding:name binding [Scheme Procedure] binding:boxed? binding [Scheme Procedure] binding:index binding [Scheme Procedure] binding:start binding [Scheme Procedure] binding:end binding

Bindings annotations for programs, along with their accessors. Bindings declare names and liveness extents for block-local variables. The best way to see what these are is to play around with them at a REPL. See Section 9.3.2 [VM Concepts], page 773, for more information. Note that bindings information is stored in a program as part of its metadata thunk, so including it in the generated object code does not impose a runtime performance penalty.

[Scheme Procedure] program-sources program [Scheme Procedure] source:addr source [Scheme Procedure] source:line source [Scheme Procedure] source:column source [Scheme Procedure] source:file source

Source location annotations for programs, along with their accessors. Source location information propagates through the compiler and ends up being serialized to the program’s metadata. This information is keyed by the offset of the instruction pointer within the object code of the program. Specifically, it is keyed on the ip just following an instruction, so that backtraces can find the source location of a call that is in progress.

[Scheme Procedure] program-arities program [C Function] scm_program_arities (program) [Scheme Procedure] program-arity program ip [Scheme Procedure] arity:start arity [Scheme Procedure] arity:end arity [Scheme Procedure] arity:nreq arity [Scheme Procedure] arity:nopt arity [Scheme Procedure] arity:rest? arity [Scheme Procedure] arity:kw arity

[Scheme Procedure] arity:allow-other-keys? arity

Accessors for a representation of the “arity” of a program. The normal case is that a procedure has one arity. For example, (lambda (x) x), takes one required argument, and that’s it. One could access that number of re-

quired arguments via (arity:nreq (program-arities (lambda (x) x))). Similarly,

arity:nopt gets the number of optional arguments, and arity:rest? returns a true value if the procedure has a rest arg. arity:kw returns a list of (kw . idx) pairs, if the procedure has keyword arguments. The idx refers to the idxth local variable; See Section 9.3.4 [Variables and the VM], page 775, for more information. Finally arity:allow-other-keys? returns a true value if other keys are allowed. See Section 6.9.4 [Optional Arguments], page 248, for more information. So what about arity:start and arity:end, then? They return the range of bytes in the program’s bytecode for which a given arity is valid. You see, a procedure can actually have more than one arity. The question, “what is a procedure’s arity” only really makes sense at certain points in the program, delimited by these arity:start and arity:end values.

[Scheme Procedure] program-arguments-alist program [ip]

Return an association list describing the arguments that program accepts, or #f if the information cannot be obtained. The alist keys that are currently defined are ‘required’, ‘optional’, ‘keyword’, ‘allowother-keys?’, and ‘rest’. For example:

     (program-arguments-alist
      (lambda* (a b #:optional c #:key (d 1) #:rest e)
        #t)) ⇒
     ((required . (a b))
      (optional . (c))
      (keyword . ((#:d . 4)))
      (allow-other-keys? . #f)
      (rest . d))

[Scheme Procedure] program-lambda-list program [ip]

Return a representation of the arguments of program as a lambda list, or #f if this information is not available. For example:

     (program-lambda-list
      (lambda* (a b #:optional c #:key (d 1) #:rest e)
        #t)) ⇒

6.9.4 Optional Arguments

Scheme procedures, as defined in R5RS, can either handle a fixed number of actual arguments, or a fixed number of actual arguments followed by arbitrarily many additional arguments. Writing procedures of variable arity can be useful, but unfortunately, the syntactic means for handling argument lists of varying length is a bit inconvenient. It is possible to give names to the fixed number of arguments, but the remaining (optional) arguments can be only referenced as a list of values (see Section 6.9.1 [Lambda], page 244).

For this reason, Guile provides an extension to lambda, lambda*, which allows the user to define procedures with optional and keyword arguments. In addition, Guile’s virtual machine has low-level support for optional and keyword argument dispatch. Calls to procedures with optional and keyword arguments can be made cheaply, without allocating a rest list.

6.9.4.1 lambda* and define*.

lambda* is like lambda, except with some extensions to allow optional and keyword arguments.

[library syntax] lambda* ([var…]

[#:optional vardef…]

[#:key vardef… [#:allow-other-keys]] [#:rest var | . var]) body1 body2 …

Create a procedure which takes optional and/or keyword arguments specified with

#:optional and #:key. For example,
     (lambda* (a b #:optional c d . e) '())

is a procedure with fixed arguments a and b, optional arguments c and d, and rest argument e. If the optional arguments are omitted in a call, the variables for them are bound to #f. Likewise, define* is syntactic sugar for defining procedures using lambda*. lambda* can also make procedures with keyword arguments. For example, a procedure defined like this:

     (define* (sir-yes-sir #:key action how-high)
       (list action how-high))
can be called as (sir-yes-sir #:action 'jump), (sir-yes-sir #:how-high
13), (sir-yes-sir #:action 'lay-down #:how-high 0), or just (sir-yes-sir).

Whichever arguments are given as keywords are bound to values (and those not given are #f). Optional and keyword arguments can also have default values to take when not present in a call, by giving a two-element list of variable name and expression. For example in

     (define* (frob foo #:optional (bar 42) #:key (baz 73))
       (list foo bar baz))

foo is a fixed argument, bar is an optional argument with default value 42, and baz is a keyword argument with default value 73. Default value expressions are not evaluated unless they are needed, and until the procedure is called. Normally it’s an error if a call has keywords other than those specified by #:key, but adding #:allow-other-keys to the definition (after the keyword argument declarations) will ignore unknown keywords. If a call has a keyword given twice, the last value is used. For example,

     (define* (flips #:key (heads 0) (tails 0))
       (display (list heads tails)))
     (flips #:heads 37 #:tails 42 #:heads 99)
     ⊣ (99 42)

#:rest is a synonym for the dotted syntax rest argument. The argument lists (a . b) and (a #:rest b) are equivalent in all respects. This is provided for more similarity to DSSSL, MIT-Scheme and Kawa among others, as well as for refugees from other Lisp dialects. When #:key is used together with a rest argument, the keyword parameters in a call all remain in the rest list. This is the same as Common Lisp. For example,

     ((lambda* (#:key (x 0) #:allow-other-keys #:rest r)
        (display r))
      #:x 123 #:y 456)
     ⊣ (#:x 123 #:y 456)

#:optional and #:key establish their bindings successively, from left to right. This means default expressions can refer back to prior parameters, for example

     (lambda* (start #:optional (end (+ 10 start)))
       (do ((i start (1+ i)))
           ((> i end))
         (display i)))

The exception to this left-to-right scoping rule is the rest argument. If there is a rest argument, it is bound after the optional arguments, but before the keyword arguments.

6.9.4.2 (ice-9 optargs)

Before Guile 2.0, lambda* and define* were implemented using macros that processed rest list arguments. This was not optimal, as calling procedures with optional arguments had to allocate rest lists at every procedure invocation. Guile 2.0 improved this situation by bringing optional and keyword arguments into Guile’s core.

However there are occasions in which you have a list and want to parse it for optional or keyword arguments. Guile’s (ice-9 optargs) provides some macros to help with that task.

The syntax let-optional and let-optional* are for destructuring rest argument lists and giving names to the various list elements. let-optional binds all variables simultaneously, while let-optional* binds them sequentially, consistent with let and let* (see Section 6.12.2 [Local Bindings], page 287).

[library syntax] let-optional rest-arg (binding …) body1 body2 … [library syntax] let-optional* rest-arg (binding …) body1 body2 …

These two macros give you an optional argument interface that is very Schemey and introduces no fancy syntax. They are compatible with the scsh macros of the same name, but are slightly extended. Each of binding may be of one of the forms var or (var default-value). rest-arg should be the rest-argument of the procedures these are used from. The items in rest-arg are sequentially bound to the variable names are given. When rest-arg runs out, the remaining vars are bound either to the default values or #f if no default value was specified. rest-arg remains bound to whatever may have been left of rest-arg. After binding the variables, the expressions body1 body2 … are evaluated in order.

Similarly, let-keywords and let-keywords* extract values from keyword style argument lists, binding local variables to those values or to defaults.

[library syntax] let-keywords args allow-other-keys? (binding …) body1 body2 … [library syntax] let-keywords* args allow-other-keys? (binding …) body1 body2

… args is evaluated and should give a list of the form (#:keyword1 value1 #:keyword2 value2 ...). The bindings are variables and default expressions, with the variables to be set (by name) from the keyword values. The body1 body2 … forms are then evaluated and the last is the result. An example will make the syntax clearest,

     (define args '(#:xyzzy "hello" #:foo "world"))
     (let-keywords args #t
           ((foo  "default for foo")
            (bar  (string-append "default" "for" "bar")))
       (display foo)
       (display ", ")
       (display bar))
     ⊣ world, defaultforbar

The binding for foo comes from the #:foo keyword in args. But the binding for bar is the default in the let-keywords, since there’s no #:bar in the args. allow-other-keys? is evaluated and controls whether unknown keywords are allowed in the args list. When true other keys are ignored (such as #:xyzzy in the example), when #f an error is thrown for anything unknown.

(ice-9 optargs) also provides some more define* sugar, which is not so useful with modern Guile coding, but still supported: define*-public is the lambda* version of

define-public; defmacro* and defmacro*-public exist for defining macros with the

improved argument list handling possibilities. The -public versions not only define the procedures/macros, but also export them from the current module.

[library syntax] define*-public formals body1 body2 …

Like a mix of define* and define-public.

[library syntax] defmacro* name formals body1 body2 … [library syntax] defmacro*-public name formals body1 body2 …

These are just like defmacro and defmacro-public except that they take lambda*- style extended parameter lists, where #:optional, #:key, #:allow-other-keys and #:rest are allowed with the usual semantics. Here is an example of a macro with an optional argument:

     (defmacro* transmogrify (a #:optional b)
       (a 1))

6.9.5 Case-lambda

R5RS’s rest arguments are indeed useful and very general, but they often aren’t the most appropriate or efficient means to get the job done. For example, lambda* is a much better solution to the optional argument problem than lambda with rest arguments.

Likewise, case-lambda works well for when you want one procedure to do double duty (or triple, or ...), without the penalty of consing a rest list.

For example:

(define (make-accum n)
  (case-lambda
    (() n)
    ((m) (set! n (+ n m)) n)))

     (define a (make-accum 20))
     (a) ⇒ 20
     (a 10) ⇒ 30
     (a) ⇒ 30

The value returned by a case-lambda form is a procedure which matches the number of actual arguments against the formals in the various clauses, in order. The first matching clause is selected, the corresponding values from the actual parameter list are bound to the variable names in the clauses and the body of the clause is evaluated. If no clause matches, an error is signalled.

The syntax of the case-lambda form is defined in the following EBNF grammar. Formals means a formal argument list just like with lambda (see Section 6.9.1 [Lambda], page 244).

     <case-lambda>
        --> (case-lambda <case-lambda-clause>*)
        --> (case-lambda <docstring> <case-lambda-clause>*)
     <case-lambda-clause>
        --> (<formals> <definition-or-command>*)
     <formals>
        --> (<identifier>*)
          | (<identifier>* . <identifier>)
          | <identifier>

Rest lists can be useful with case-lambda:

     (define plus
       (case-lambda
         "Return the sum of all arguments."
         (() 0)
         ((a) a)
         ((a b) (+ a b))
         ((a b . rest) (apply plus (+ a b) rest))))
     (plus 1 2 3) ⇒ 6

Also, for completeness. Guile defines case-lambda* as well, which is like case-lambda, except with lambda* clauses. A case-lambda* clause matches if the arguments fill the required arguments, but are not too many for the optional and/or rest arguments.

Keyword arguments are possible with case-lambda* as well, but they do not contribute to the “matching” behavior, and their interactions with required, optional, and rest arguments can be surprising.

For the purposes of case-lambda* (and of case-lambda, as a special case), a clause matches if it has enough required arguments, and not too many positional arguments. The required arguments are any arguments before the #:optional, #:key, and #:rest arguments. Positional arguments are the required arguments, together with the optional arguments.

In the absence of #:key or #:rest arguments, it’s easy to see how there could be too many positional arguments: you pass 5 arguments to a function that only takes 4 arguments, including optional arguments. If there is a #:rest argument, there can never be too many positional arguments: any application with enough required arguments for a clause will match that clause, even if there are also #:key arguments.

Otherwise, for applications to a clause with #:key arguments (and without a #:rest argument), a clause will match there only if there are enough required arguments and if the next argument after binding required and optional arguments, if any, is a keyword. For efficiency reasons, Guile is currently unable to include keyword arguments in the matching algorithm. Clauses match on positional arguments only, not by comparing a given keyword to the available set of keyword arguments that a function has.

Some examples follow.

     (define f
       (case-lambda*
         ((a #:optional b) 'clause-1)
         ((a #:optional b #:key c) 'clause-2)
         ((a #:key d) 'clause-3)
         ((#:key e #:rest f) 'clause-4)))
     (f) ⇒ clause-4
     (f 1) ⇒ clause-1
     (f) ⇒ clause-4
     (f #:e 10) clause-1
     (f 1 #:foo) clause-1
     (f 1 #:c 2) clause-2
     (f #:a #:b #:c #:d #:e) clause-4
     ;; clause-2 will match anything that clause-3 would match.
     (f 1 #:d 2) ⇒ error: bad keyword args in clause 2

Don’t forget that the clauses are matched in order, and the first matching clause will be taken. This can result in a keyword being bound to a required argument, as in the case of

f #:e 10.

6.9.6 Higher-Order Functions

As a functional programming language, Scheme allows the definition of higher-order functions, i.e., functions that take functions as arguments and/or return functions. Utilities to derive procedures from other procedures are provided and described below.

[Scheme Procedure] const value

Return a procedure that accepts any number of arguments and returns value.

          (procedure? (const 3))  ⇒ #t
          ((const 'hello))  ⇒ hello
          ((const 'hello) 'world)  ⇒ hello

[Scheme Procedure] negate proc

Return a procedure with the same arity as proc that returns the not of proc’s result.

          (procedure? (negate number?)) ⇒ #t
          ((negate odd?) 2)  ⇒ #t
          ((negate real?) 'dream)  ⇒ #t

          ((negate string-prefix?) "GNU" "GNU Guile")
                                        ⇒ #f
          (filter (negate number?) '(a 2 "b"))
                                        ⇒ (a "b")

[Scheme Procedure] compose proc1 proc2 … Compose proc1 with the procedures proc2 … such that the last proc argument is applied first and proc1 last, and return the resulting procedure. The given procedures must have compatible arity.

          (procedure? (compose 1+ 1-)) ⇒ #t
          ((compose sqrt 1+ 1+) 2)  ⇒ 2.0
          ((compose 1+ sqrt) 3)  ⇒ 2.73205080756888
          (eq? (compose 1+) 1+)  ⇒ #t
          ((compose zip unzip2) '((1 2) (a b)))
                                       ⇒ ((1 2) (a b))

[Scheme Procedure] identity x

Return X.

[Scheme Procedure] and=> value proc When value is #f, return #f. Otherwise, return (proc value).

6.9.7 Procedure Properties and Meta-information

In addition to the information that is strictly necessary to run, procedures may have other associated information. For example, the name of a procedure is information not for the procedure, but about the procedure. This meta-information can be accessed via the procedure properties interface.

The first group of procedures in this meta-interface are predicates to test whether a Scheme object is a procedure, or a special procedure, respectively. procedure? is the most general predicates, it returns #t for any kind of procedure.

[Scheme Procedure] procedure? obj [C Function] scm_procedure_p (obj)

Return #t if obj is a procedure.

[Scheme Procedure] thunk? obj [C Function] scm_thunk_p (obj)

Return #t if obj is a thunk—a procedure that does not accept arguments.

Procedure properties are general properties associated with procedures. These can be the name of a procedure or other relevant information, such as debug hints.

[Scheme Procedure] procedure-name proc

[C Function] scm_procedure_name (proc)

Return the name of the procedure proc

[Scheme Procedure] procedure-source proc

[C Function] scm_procedure_source (proc)

Return the source of the procedure proc. Returns #f if the source code is not available.

[Scheme Procedure] procedure-properties proc

[C Function] scm_procedure_properties (proc)

Return the properties associated with proc, as an association list.

[Scheme Procedure] procedure-property proc key

[C Function] scm_procedure_property (proc, key)

Return the property of proc with name key.

[Scheme Procedure] set-procedure-properties! proc alist

[C Function] scm_set_procedure_properties_x (proc, alist)

Set proc’s property list to alist.

[Scheme Procedure] set-procedure-property! proc key value

[C Function] scm_set_procedure_property_x (proc, key, value)

In proc’s property list, set the property named key to value.

Documentation for a procedure can be accessed with the procedure procedure-

documentation.
[Scheme Procedure] procedure-documentation proc
[C Function] scm_procedure_documentation (proc)

Return the documentation string associated with proc. By convention, if a procedure contains more than one expression and the first expression is a string constant, that string is assumed to contain documentation for that procedure.

6.9.8 Procedures with Setters

A procedure with setter is a special kind of procedure which normally behaves like any accessor procedure, that is a procedure which accesses a data structure. The difference is that this kind of procedure has a so-called setter attached, which is a procedure for storing something into a data structure.

Procedures with setters are treated specially when the procedure appears in the special form set! (REFFIXME). How it works is best shown by example.

Suppose we have a procedure called foo-ref, which accepts two arguments, a value of type foo and an integer. The procedure returns the value stored at the given index in the foo object. Let f be a variable containing such a foo data structure.^{5}

     (foo-ref f 0)  ⇒ bar
     (foo-ref f 1)  ⇒ braz

Also suppose that a corresponding setter procedure called foo-set! does exist.

     (foo-set! f 0 'bla)
     (foo-ref f 0)  ⇒ bla

Now we could create a new procedure called foo, which is a procedure with setter, by calling make-procedure-with-setter with the accessor and setter procedures foo-ref and foo-set!. Let us call this new procedure foo.

5 Working definitions would be:

(define foo-ref vector-ref)

(define foo-set! vector-set!)

(define f (make-vector 2 #f))

     (define foo (make-procedure-with-setter foo-ref foo-set!))

foo can from now an be used to either read from the data structure stored in f, or to write into the structure.

     (set! (foo f 0) 'dum)
     (foo f 0)  ⇒ dum

[Scheme Procedure] make-procedure-with-setter procedure setter

[C Function] scm_make_procedure_with_setter (procedure, setter)

Create a new procedure which behaves like procedure, but with the associated setter setter.

[Scheme Procedure] procedure-with-setter? obj
[C Function] scm_procedure_with_setter_p (obj)

Return #t if obj is a procedure with an associated setter procedure.

[Scheme Procedure] procedure proc [C Function] scm_procedure (proc)

Return the procedure of proc, which must be an applicable struct.

[Scheme Procedure] setter proc

Return the setter of proc, which must be either a procedure with setter or an operator struct.

6.9.9 Inlinable Procedures

You can define an inlinable procedure by using define-inlinable instead of define. An inlinable procedure behaves the same as a regular procedure, but direct calls will result in the procedure body being inlined into the caller.

Bear in mind that starting from version 2.0.3, Guile has a partial evaluator that can inline the body of inner procedures when deemed appropriate:

     scheme@(guile-user)> ,optimize (define (foo x)
                                      (define (bar) (+ x 3))
                                      (* (bar) 2))
     $1 = (define foo
            (lambda (#{x 94}#) (* (+ #{x 94}# 3) 2)))

The partial evaluator does not inline top-level bindings, though, so this is a situation where you may find it interesting to use define-inlinable.

Procedures defined with define-inlinable are always inlined, at all direct call sites. This eliminates function call overhead at the expense of an increase in code size. Additionally, the caller will not transparently use the new definition if the inline procedure is redefined. It is not possible to trace an inlined procedures or install a breakpoint in it (see Section 6.25.4 [Traps], page 455). For these reasons, you should not make a procedure inlinable unless it demonstrably improves performance in a crucial way.

In general, only small procedures should be considered for inlining, as making large procedures inlinable will probably result in an increase in code size. Additionally, the elimination of the call overhead rarely matters for large procedures.

[Scheme Syntax] define-inlinable (name parameter …) body1 body2 … Define name as a procedure with parameters parameters and bodies body1, body2, . . ..

6.10 Macros

At its best, programming in Lisp is an iterative process of building up a language appropriate to the problem at hand, and then solving the problem in that language. Defining new procedures is part of that, but Lisp also allows the user to extend its syntax, with its famous macros.

Macros are syntactic extensions which cause the expression that they appear in to be transformed in some way before being evaluated. In expressions that are intended for macro transformation, the identifier that names the relevant macro must appear as the first element, like this:

     (macro-name macro-args ...)

Macro expansion is a separate phase of evaluation, run before code is interpreted or compiled. A macro is a program that runs on programs, translating an embedded language into core Scheme^{6}.

6.10.1 Defining Macros

A macro is a binding between a keyword and a syntax transformer. Since it’s difficult to discuss define-syntax without discussing the format of transformers, consider the following example macro definition:

     (define-syntax when
       (syntax-rules ()
         ((when condition exp ...)
          (if condition
              (begin exp ...)))))
     (when #t
       (display "hey ho\n")
       (display "let's go\n"))
     ⊣ hey ho
     ⊣ let's go

In this example, the when binding is bound with define-syntax. Syntax transformers are discussed in more depth in Section 6.10.2 [Syntax Rules], page 258 and Section 6.10.3 [Syntax Case], page 263.

[Syntax] define-syntax keyword transformer Bind keyword to the syntax transformer obtained by evaluating transformer. After a macro has been defined, further instances of keyword in Scheme source code will invoke the syntax transformer defined by transformer.

One can also establish local syntactic bindings with let-syntax.

6 These days such embedded languages are often referred to as embedded domain-specific languages, or

EDSLs.

[Syntax] let-syntax ((keyword transformer) …) exp1 exp2 … Bind each keyword to its corresponding transformer while expanding exp1 exp2 . . .. A let-syntax binding only exists at expansion-time.

          (let-syntax ((unless
                        (syntax-rules ()
                          ((unless condition exp ...)
                           (if (not condition)
                               (begin exp ...))))))
            (unless #t
              (primitive-exit 1))
            "rock rock rock")
          ⇒ "rock rock rock"

A define-syntax form is valid anywhere a definition may appear: at the top-level, or locally. Just as a local define expands out to an instance of letrec, a local define-syntax

expands out to letrec-syntax.

[Syntax] letrec-syntax ((keyword transformer) …) exp1 exp2 … Bind each keyword to its corresponding transformer while expanding exp1 exp2 . . .. In the spirit of letrec versus let, an expansion produced by transformer may reference a keyword bound by the same letrec-syntax.

          (letrec-syntax ((my-or
                           (syntax-rules ()
                             ((my-or)
                              #t)
                             ((my-or exp)
                              exp)
                             ((my-or exp rest ...)
                              (let ((t exp))
                                (if t
                                    t
                                    (my-or rest ...)))))))
            (my-or #f "rockaway beach"))
          ⇒ "rockaway beach"

6.10.2 Syntax-rules Macros

syntax-rules macros are simple, pattern-driven syntax transformers, with a beauty worthy of Scheme.

[Syntax] syntax-rules literals (pattern template) … Create a syntax transformer that will rewrite an expression using the rules embodied in the pattern and template clauses.

A syntax-rules macro consists of three parts: the literals (if any), the patterns, and as many templates as there are patterns.

When the syntax expander sees the invocation of a syntax-rules macro, it matches the expression against the patterns, in order, and rewrites the expression using the template from the first matching pattern. If no pattern matches, a syntax error is signalled.

6.10.2.1 Patterns

We have already seen some examples of patterns in the previous section: (unless condition exp ...), (my-or exp), and so on. A pattern is structured like the expression that it is to match. It can have nested structure as well, like (let ((var val) ...) exp exp* ...). Broadly speaking, patterns are made of lists, improper lists, vectors, identifiers, and datums. Users can match a sequence of patterns using the ellipsis (...).

Identifiers in a pattern are called literals if they are present in the syntax-rules literals list, and pattern variables otherwise. When building up the macro output, the expander replaces instances of a pattern variable in the template with the matched subexpression.

     (define-syntax kwote
       (syntax-rules ()
         ((kwote exp)
          (quote exp))))
     (kwote (foo . bar))
     ⇒ (foo . bar)

An improper list of patterns matches as rest arguments do:

     (define-syntax let1
       (syntax-rules ()
         ((_(var val) . exps)
          (let ((var val)) . exps))))

However this definition of let1 probably isn’t what you want, as the tail pattern exps will match non-lists, like (let1 (foo 'bar) . baz). So often instead of using improper lists as patterns, ellipsized patterns are better. Instances of a pattern variable in the template must be followed by an ellipsis.

     (define-syntax let1
       (syntax-rules ()
         ((_(var val) exp ...)
          (let ((var val)) exp ...))))

This let1 probably still doesn’t do what we want, because the body matches sequences of zero expressions, like (let1 (foo 'bar)). In this case we need to assert we have at least one body expression. A common idiom for this is to name the ellipsized pattern variable with an asterisk:

     (define-syntax let1
       (syntax-rules ()
         ((_(var val) exp exp* ...)
          (let ((var val)) exp exp* ...))))

A vector of patterns matches a vector whose contents match the patterns, including ellipsizing and tail patterns.

     (define-syntax letv
       (syntax-rules ()
         ((_#((var val) ...) exp exp* ...)
          (let ((var val) ...) exp exp* ...))))
     (letv #((foo 'bar)) foo)
     ⇒ bar

Literals are used to match specific datums in an expression, like the use of => and else in cond expressions.

(define-syntax cond1
  (syntax-rules (=> else)
    ((cond1 test => fun)
     (let ((exp test))
       (if exp (fun exp) #f)))
    ((cond1 test exp exp* ...)
     (if test (begin exp exp* ...)))
    ((cond1 else exp exp* ...)
     (begin exp exp* ...))))
(define (square x) (* x x))
(cond1 10 => square)
⇒ 100
(let ((=> #t))
  (cond1 10 => square))
⇒ #<procedure square (x)>

A literal matches an input expression if the input expression is an identifier with the same name as the literal, and both are unbound^{7}.

If a pattern is not a list, vector, or an identifier, it matches as a literal, with equal?.

(define-syntax define-matcher-macro
  (syntax-rules ()
    ((_name lit)
     (define-syntax name
       (syntax-rules ()
        ((_lit) #t)
        ((_else) #f))))))
(define-matcher-macro is-literal-foo? "foo")
(is-literal-foo? "foo")
⇒ #t
(is-literal-foo? "bar")
⇒ #f
(let ((foo "foo"))
  (is-literal-foo? foo))
⇒ #f

The last example indicates that matching happens at expansion-time, not at run-time. Syntax-rules macros are always used as (macro . args), and the macro will always be a

symbol. Correspondingly, a syntax-rules pattern must be a list (proper or improper), and the first pattern in that list must be an identifier. Incidentally it can be any identifier – it doesn’t have to actually be the name of the macro. Thus the following three are equivalent:

7 Language lawyers probably see the need here for use of literal-identifier=? rather than free-

identifier=?, and would probably be correct. Patches accepted.

     (define-syntax when
       (syntax-rules ()
         ((when c e ...)
          (if c (begin e ...)))))
     (define-syntax when
       (syntax-rules ()
         ((_c e ...)
          (if c (begin e ...)))))
     (define-syntax when
       (syntax-rules ()
         ((something-else-entirely c e ...)
          (if c (begin e ...)))))

For clarity, use one of the first two variants. Also note that since the pattern variable will always match the macro itself (e.g., cond1), it is actually left unbound in the template.

6.10.2.2 Hygiene

syntax-rules macros have a magical property: they preserve referential transparency. When you read a macro definition, any free bindings in that macro are resolved relative to the macro definition; and when you read a macro instantiation, all free bindings in that expression are resolved relative to the expression.

This property is sometimes known as hygiene, and it does aid in code cleanliness. In your macro definitions, you can feel free to introduce temporary variables, without worrying about inadvertently introducing bindings into the macro expansion.

Consider the definition of my-or from the previous section:

     (define-syntax my-or
       (syntax-rules ()
         ((my-or)
          #t)
         ((my-or exp)
          exp)
         ((my-or exp rest ...)
          (let ((t exp))
            (if t
                t
                (my-or rest ...))))))

A naive expansion of (let ((t #t)) (my-or #f t)) would yield:

     (let ((t #t))
       (let ((t #f))
         (if t t t)))
     ⇒ #f

Which clearly is not what we want. Somehow the t in the definition is distinct from the t at the site of use; and it is indeed this distinction that is maintained by the syntax expander, when expanding hygienic macros.

This discussion is mostly relevant in the context of traditional Lisp macros (see Section 6.10.5 [Defmacros], page 270), which do not preserve referential transparency. Hygiene adds to the expressive power of Scheme.

6.10.2.3 Shorthands

One often ends up writing simple one-clause syntax-rules macros. There is a convenient shorthand for this idiom, in the form of define-syntax-rule.

[Syntax] define-syntax-rule (keyword . pattern) [docstring] template Define keyword as a new syntax-rules macro with one clause.

Cast into this form, our when example is significantly shorter:

     (define-syntax-rule (when c e ...)
       (if c (begin e ...)))

6.10.2.4 Reporting Syntax Errors in Macros

[Syntax] syntax-error message [arg ...]

Report an error at macro-expansion time. message must be a string literal, and the optional arg operands can be arbitrary expressions providing additional information.

syntax-error is intended to be used within syntax-rules templates. For example:

     (define-syntax simple-let
       (syntax-rules ()
         ((_(head ... ((x . y) val) . tail)
             body1 body2 ...)
          (syntax-error
           "expected an identifier but got"
           (x . y)))
         ((_((name val) ...) body1 body2 ...)
          ((lambda (name ...) body1 body2 ...)
           val ...))))

6.10.2.5 Specifying a Custom Ellipsis Identifier

When writing macros that generate macro definitions, it is convenient to use a different ellipsis identifier at each level. Guile allows the desired ellipsis identifier to be specified as the first operand to syntax-rules, as specified by SRFI-46 and R7RS. For example:

     (define-syntax define-quotation-macros
       (syntax-rules ()
         ((_(macro-name head-symbol) ...)
          (begin (define-syntax macro-name
                   (syntax-rules ::: ()
                     ((_x :::)
                      (quote (head-symbol x :::)))))
                 ...))))
     (define-quotation-macros (quote-a a) (quote-b b) (quote-c c))
     (quote-a 1 2 3) ⇒ (a 1 2 3)

6.10.2.6 Further Information

For a formal definition of syntax-rules and its pattern language, see See Section “Macros” in Revised(5) Report on the Algorithmic Language Scheme.

syntax-rules macros are simple and clean, but do they have limitations. They do not lend themselves to expressive error messages: patterns either match or they don’t. Their ability to generate code is limited to template-driven expansion; often one needs to define a number of helper macros to get real work done. Sometimes one wants to introduce a binding into the lexical context of the generated code; this is impossible with syntaxrules. Relatedly, they cannot programmatically generate identifiers.

The solution to all of these problems is to use syntax-case if you need its features. But if for some reason you’re stuck with syntax-rules, you might enjoy Joe Marshall’s

syntax-rules Primer for the Merely Eccentric (http://sites.google.com/site/
evalapply/eccentric.txt).

6.10.3 Support for the syntax-case System

syntax-case macros are procedural syntax transformers, with a power worthy of Scheme.

[Syntax] syntax-case syntax literals (pattern [guard] exp) … Match the syntax object syntax against the given patterns, in order. If a pattern matches, return the result of evaluating the associated exp.

Compare the following definitions of when:

     (define-syntax when
       (syntax-rules ()
         ((_test e e* ...)
          (if test (begin e e* ...)))))
     (define-syntax when
       (lambda (x)
         (syntax-case x ()
           ((_test e e* ...)
            #'(if test (begin e e* ...))))))

Clearly, the syntax-case definition is similar to its syntax-rules counterpart, and equally clearly there are some differences. The syntax-case definition is wrapped in a lambda, a function of one argument; that argument is passed to the syntax-case invocation; and the “return value” of the macro has a #' prefix.

All of these differences stem from the fact that syntax-case does not define a syntax transformer itself – instead, syntax-case expressions provide a way to destructure a syntax object, and to rebuild syntax objects as output.

So the lambda wrapper is simply a leaky implementation detail, that syntax transformers are just functions that transform syntax to syntax. This should not be surprising, given that we have already described macros as “programs that write programs”. syntax-case is simply a way to take apart and put together program text, and to be a valid syntax transformer it needs to be wrapped in a procedure.

Unlike traditional Lisp macros (see Section 6.10.5 [Defmacros], page 270), syntax-case macros transform syntax objects, not raw Scheme forms. Recall the naive expansion of my-or given in the previous section:

     (let ((t #t))
       (my-or #f t))
     ;; naive expansion:
     (let ((t #t))
       (let ((t #f))
         (if t t t)))

Raw Scheme forms simply don’t have enough information to distinguish the first two t instances in (if t t t) from the third t. So instead of representing identifiers as symbols, the syntax expander represents identifiers as annotated syntax objects, attaching such information to those syntax objects as is needed to maintain referential transparency.

[Syntax] syntax form

Create a syntax object wrapping form within the current lexical context.

Syntax objects are typically created internally to the process of expansion, but it is possible to create them outside of syntax expansion:

     (syntax (foo bar baz))
     ⇒ #<some representation of that syntax>

However it is more common, and useful, to create syntax objects when building output from a syntax-case expression.

     (define-syntax add1
       (lambda (x)
         (syntax-case x ()
           ((_exp)
            (syntax (+ exp 1))))))

It is not strictly necessary for a syntax-case expression to return a syntax object, because syntax-case expressions can be used in helper functions, or otherwise used outside of syntax expansion itself. However a syntax transformer procedure must return a syntax object, so most uses of syntax-case do end up returning syntax objects.

Here in this case, the form that built the return value was (syntax (+ exp 1)). The interesting thing about this is that within a syntax expression, any appearance of a pattern variable is substituted into the resulting syntax object, carrying with it all relevant metadata from the source expression, such as lexical identity and source location.

Indeed, a pattern variable may only be referenced from inside a syntax form. The syntax expander would raise an error when defining add1 if it found exp referenced outside

a syntax form.

Since syntax appears frequently in macro-heavy code, it has a special reader macro: #'. #'foo is transformed by the reader into (syntax foo), just as 'foo is transformed into

(quote foo).

The pattern language used by syntax-case is conveniently the same language used by syntax-rules. Given this, Guile actually defines syntax-rules in terms of syntax-case:

     (define-syntax syntax-rules
       (lambda (x)
         (syntax-case x ()
           ((_(k ...) ((keyword . pattern) template) ...)
            #'(lambda (x)
                (syntax-case x (k ...)
                  ((dummy . pattern) #'template)
                  ...))))))

And that’s that.

6.10.3.1 Why syntax-case?

The examples we have shown thus far could just as well have been expressed with syntaxrules, and have just shown that syntax-case is more verbose, which is true. But there is a difference: syntax-case creates procedural macros, giving the full power of Scheme to the macro expander. This has many practical applications.

A common desire is to be able to match a form only if it is an identifier. This is impossible with syntax-rules, given the datum matching forms. But with syntax-case it is easy:

[Scheme Procedure] identifier? syntax-object Returns #t if syntax-object is an identifier, or #f otherwise.

     ;; relying on previous add1 definition
     (define-syntax add1!
       (lambda (x)
         (syntax-case x ()
           ((_var) (identifier? #'var)
            #'(set! var (add1 var))))))
     (define foo 0)
     (add1! foo)
     foo ⇒ 1
     (add1! "not-an-identifier") ⇒ error
   With syntax-rules, the error for (add1! "not-an-identifier") would be something

like “invalid set!”. With syntax-case, it will say something like “invalid add1!”, because we attach the guard clause to the pattern: (identifier? #'var). This becomes more important with more complicated macros. It is necessary to use identifier?, because to the expander, an identifier is more than a bare symbol.

Note that even in the guard clause, we reference the var pattern variable within a syntax form, via #'var.

Another common desire is to introduce bindings into the lexical context of the output expression. One example would be in the so-called “anaphoric macros”, like aif. Anaphoric macros bind some expression to a well-known identifier, often it, within their bodies. For example, in (aif (foo) (bar it)), it would be bound to the result of (foo).

To begin with, we should mention a solution that doesn’t work:

     ;; doesn't work
     (define-syntax aif

       (lambda (x)
         (syntax-case x ()
           ((_test then else)
            #'(let ((it test))
                (if it then else))))))

The reason that this doesn’t work is that, by default, the expander will preserve referential transparency; the then and else expressions won’t have access to the binding of

it.

But they can, if we explicitly introduce a binding via datum->syntax.

[Scheme Procedure] datum->syntax for-syntax datum Create a syntax object that wraps datum, within the lexical context corresponding to the syntax object for-syntax.

For completeness, we should mention that it is possible to strip the metadata from a syntax object, returning a raw Scheme datum:

[Scheme Procedure] syntax->datum syntax-object

Strip the metadata from syntax-object, returning its contents as a raw Scheme datum.

In this case we want to introduce it in the context of the whole expression, so we can create a syntax object as (datum->syntax x 'it), where x is the whole expression, as passed to the transformer procedure.

Here’s another solution that doesn’t work:

     ;; doesn't work either
     (define-syntax aif
       (lambda (x)
         (syntax-case x ()
           ((_test then else)
            (let ((it (datum->syntax x 'it)))
              #'(let ((it test))
                  (if it then else)))))))

The reason that this one doesn’t work is that there are really two environments at work here – the environment of pattern variables, as bound by syntax-case, and the environment of lexical variables, as bound by normal Scheme. The outer let form establishes a binding in the environment of lexical variables, but the inner let form is inside a syntax form, where only pattern variables will be substituted. Here we need to introduce a piece of the lexical environment into the pattern variable environment, and we can do so using syntax-case itself:

     ;; works, but is obtuse
     (define-syntax aif
       (lambda (x)
         (syntax-case x ()
           ((_test then else)
            ;; invoking syntax-case on the generated
            ;; syntax object to expose it to `syntax'
            (syntax-case (datum->syntax x 'it) ()

              (it
                #'(let ((it test))
                    (if it then else))))))))
     (aif (getuid) (display it) (display "none")) (newline)
     ⊣ 500

However there are easier ways to write this. with-syntax is often convenient:

[Syntax] with-syntax ((pat val) …) exp … Bind patterns pat from their corresponding values val, within the lexical context of exp . . ..

          ;; better
          (define-syntax aif
            (lambda (x)
              (syntax-case x ()
                ((_test then else)
                 (with-syntax ((it (datum->syntax x 'it)))
                   #'(let ((it test))
                       (if it then else)))))))

As you might imagine, with-syntax is defined in terms of syntax-case. But even that might be off-putting to you if you are an old Lisp macro hacker, used to building macro output with quasiquote. The issue is that with-syntax creates a separation between the point of definition of a value and its point of substitution.

So for cases in which a quasiquote style makes more sense, syntax-case also de-

fines quasisyntax, and the related unsyntax and unsyntax-splicing, abbreviated by

the reader as #`, #,, and #,@, respectively.

For example, to define a macro that inserts a compile-time timestamp into a source file, one may write:

     (define-syntax display-compile-timestamp
       (lambda (x)
         (syntax-case x ()
           ((_)
            #`(begin
               (display "The compile timestamp was: ")
               (display #,(current-time))
               (newline))))))

Readers interested in further information on syntax-case macros should see R. Kent Dybvig’s excellent The Scheme Programming Language, either edition 3 or 4, in the chapter on syntax. Dybvig was the primary author of the syntax-case system. The book itself is

available online at http://scheme.com/tspl4/.

6.10.3.2 Custom Ellipsis Identifiers for syntax-case Macros

When writing procedural macros that generate macro definitions, it is convenient to use a different ellipsis identifier at each level. Guile supports this for procedural macros using the

with-ellipsis special form:

[Syntax] with-ellipsis ellipsis body …

ellipsis must be an identifier. Evaluate body in a special lexical environment such that all macro patterns and templates within body will use ellipsis as the ellipsis identifier instead of the usual three dots (...).

For example:

(define-syntax define-quotation-macros
  (lambda (x)
    (syntax-case x ()
      ((_(macro-name head-symbol) ...)
       #'(begin (define-syntax macro-name
                  (lambda (x)
                    (with-ellipsis :::
                      (syntax-case x ()
                        ((_x :::)
                         #'(quote (head-symbol x :::)))))))
                ...)))))
(define-quotation-macros (quote-a a) (quote-b b) (quote-c c))
(quote-a 1 2 3) ⇒ (a 1 2 3)

Note that with-ellipsis does not affect the ellipsis identifier of the generated code, unless with-ellipsis is included around the generated code.

6.10.4 Syntax Transformer Helpers

As noted in the previous section, Guile’s syntax expander operates on syntax objects. Procedural macros consume and produce syntax objects. This section describes some of the auxiliary helpers that procedural macros can use to compare, generate, and query objects of this data type.

[Scheme Procedure] bound-identifier=? a b

Return #t if the syntax objects a and b refer to the same lexically-bound identifier, or #f otherwise.

[Scheme Procedure] free-identifier=? a b

Return #t if the syntax objects a and b refer to the same free identifier, or #f otherwise.

[Scheme Procedure] generate-temporaries ls

Return a list of temporary identifiers as long as ls is long.

[Scheme Procedure] syntax-source x

Return the source properties that correspond to the syntax object x. See Section 6.25.2 [Source Properties], page 447, for more information.

Guile also offers some more experimental interfaces in a separate module. As was the case with the Large Hadron Collider, it is unclear to our senior macrologists whether adding these interfaces will result in awesomeness or in the destruction of Guile via the creation of a singularity. We will preserve their functionality through the 2.0 series, but we reserve the right to modify them in a future stable series, to a more than usual degree.

(use-modules (system syntax))

[Scheme Procedure] syntax-module id

Return the name of the module whose source contains the identifier id.

[Scheme Procedure] syntax-local-binding id

Resolve the identifer id, a syntax object, within the current lexical environment, and return two values, the binding type and a binding value. The binding type is a symbol, which may be one of the following:

lexical A lexically-bound variable. The value is a unique token (in the sense of eq?) identifying this binding.

macro A syntax transformer, either local or global. The value is the transformer procedure.

pattern-variable

A pattern variable, bound via syntax-case. The value is an opaque object, internal to the expander.

ellipsis An internal binding, bound via with-ellipsis. The value is the (antimarked) local ellipsis identifier.

displaced-lexical

A lexical variable that has gone out of scope. This can happen if a badly-written procedural macro saves a syntax object, then attempts to introduce it in a context in which it is unbound. The value is #f.

global A global binding. The value is a pair, whose head is the symbol, and whose tail is the name of the module in which to resolve the symbol.

other Some other binding, like lambda or other core bindings. The value is #f.

This is a very low-level procedure, with limited uses. One case in which it is useful is to build abstractions that associate auxiliary information with macros:

     (define aux-property (make-object-property))
     (define-syntax-rule (with-aux aux value)
       (let ((trans value))
         (set! (aux-property trans) aux)
         trans))
     (define-syntax retrieve-aux
       (lambda (x)
         (syntax-case x ()
           ((x id)
            (call-with-values (lambda () (syntax-local-binding #'id))
              (lambda (type val)
                (with-syntax ((aux (datum->syntax #'here
                                                  (and (eq? type 'macro)
                                                       (aux-property val)))))
                  #''aux)))))))
     (define-syntax foo
       (with-aux 'bar
         (syntax-rules () ((_) 'foo))))

     (foo)
     ⇒ foo
     (retrieve-aux foo)
     ⇒ bar

syntax-local-binding must be called within the dynamic extent of a syntax transformer; to call it otherwise will signal an error.

[Scheme Procedure] syntax-locally-bound-identifiers id

Return a list of identifiers that were visible lexically when the identifier id was created, in order from outermost to innermost. This procedure is intended to be used in specialized procedural macros, to provide a macro with the set of bound identifiers that the macro can reference. As a technical implementation detail, the identifiers returned by syntax-locally-bound-identifiers will be anti-marked, like the syntax object that is given as input to a macro. This is to signal to the macro expander that these bindings were present in the original source, and do not need to be hygienically renamed, as would be the case with other introduced identifiers. See the discussion of hygiene in section 12.1 of the R6RS, for more information on marks.

     (define (local-lexicals id)
       (filter (lambda (x)
                 (eq? (syntax-local-binding x) 'lexical))
               (syntax-locally-bound-identifiers id)))
     (define-syntax lexicals
       (lambda (x)
         (syntax-case x ()
           ((lexicals) #'(lexicals lexicals))
           ((lexicals scope)
            (with-syntax (((id ...) (local-lexicals #'scope)))
              #'(list (cons 'id id) ...))))))
     (let* ((x 10) (x 20)) (lexicals))
     ⇒ ((x . 10) (x . 20))

6.10.5 Lisp-style Macro Definitions

The traditional way to define macros in Lisp is very similar to procedure definitions. The key differences are that the macro definition body should return a list that describes the transformed expression, and that the definition is marked as a macro definition (rather than a procedure definition) by the use of a different definition keyword: in Lisp, defmacro rather than defun, and in Scheme, define-macro rather than define.

Guile supports this style of macro definition using both defmacro and define-macro. The only difference between them is how the macro name and arguments are grouped together in the definition:

(defmacro name (args ...) body ...)

is the same as

(define-macro (name args ...) body ...)

The difference is analogous to the corresponding difference between Lisp’s defun and Scheme’s define.

Having read the previous section on syntax-case, it’s probably clear that Guile actually implements defmacros in terms of syntax-case, applying the transformer on the expression between invocations of syntax->datum and datum->syntax. This realization leads us to the problem with defmacros, that they do not preserve referential transparency. One can be careful to not introduce bindings into expanded code, via liberal use of gensym, but there is no getting around the lack of referential transparency for free bindings in the macro itself.

Even a macro as simple as our when from before is difficult to get right:

     (define-macro (when cond exp . rest)
       `(if ,cond
            (begin ,exp . ,rest)))
     (when #f (display "Launching missiles!\n"))
     ⇒ #f
     (let ((if list))
       (when #f (display "Launching missiles!\n")))
     ⊣ Launching missiles!
     ⇒ (#f #<unspecified>)

Guile’s perspective is that defmacros have had a good run, but that modern macros should be written with syntax-rules or syntax-case. There are still many uses of defmacros within Guile itself, but we will be phasing them out over time. Of course we won’t take away defmacro or define-macro themselves, as there is lots of code out there that uses them.

6.10.6 Identifier Macros

When the syntax expander sees a form in which the first element is a macro, the whole form gets passed to the macro’s syntax transformer. One may visualize this as:

     (define-syntax foo foo-transformer)
     (foo arg...)
     ;; expands via
     (foo-transformer #'(foo arg...))

If, on the other hand, a macro is referenced in some other part of a form, the syntax transformer is invoked with only the macro reference, not the whole form.

     (define-syntax foo foo-transformer)
     foo
     ;; expands via
     (foo-transformer #'foo)

This allows bare identifier references to be replaced programmatically via a macro. syntax-rules provides some syntax to effect this transformation more easily.

[Syntax] identifier-syntax exp

Returns a macro transformer that will replace occurrences of the macro with exp.

For example, if you are importing external code written in terms of fx+, the fixnum addition operator, but Guile doesn’t have fx+, you may use the following to replace fx+ with +:

(define-syntax fx+ (identifier-syntax +))

There is also special support for recognizing identifiers on the left-hand side of a set! expression, as in the following:

(define-syntax foo foo-transformer)
(set! foo val)
;; expands via
(foo-transformer #'(set! foo val))
;; if foo-transformer is a "variable transformer"

As the example notes, the transformer procedure must be explicitly marked as being a “variable transformer”, as most macros aren’t written to discriminate on the form in the operator position.

[Scheme Procedure] make-variable-transformer transformer

Mark the transformer procedure as being a “variable transformer”. In practice this means that, when bound to a syntactic keyword, it may detect references to that keyword on the left-hand-side of a set!.

     (define bar 10)
     (define-syntax bar-alias
       (make-variable-transformer
        (lambda (x)
          (syntax-case x (set!)
            ((set! var val) #'(set! bar val))
            ((var arg ...) #'(bar arg ...))
            (var (identifier? #'var) #'bar)))))
     bar-alias ⇒ 10
     (set! bar-alias 20)
     bar ⇒ 20
     (set! bar 30)
     bar-alias ⇒ 30

There is an extension to identifier-syntax which allows it to handle the set! case as well:

[Syntax] identifier-syntax (var exp1) ((set! var val) exp2)

Create a variable transformer. The first clause is used for references to the variable in operator or operand position, and the second for appearances of the variable on the left-hand-side of an assignment. For example, the previous bar-alias example could be expressed more succinctly like this:

     (define-syntax bar-alias
       (identifier-syntax
         (var bar)
         ((set! var val) (set! bar val))))

As before, the templates in identifier-syntax forms do not need wrapping in #' syntax forms.

6.10.7 Syntax Parameters

Syntax parameters^{8} are a mechanism for rebinding a macro definition within the dynamic extent of a macro expansion. This provides a convenient solution to one of the most common types of unhygienic macro: those that introduce a unhygienic binding each time the macro is used. Examples include a lambda form with a return keyword, or class macros that introduce a special self binding.

With syntax parameters, instead of introducing the binding unhygienically each time, we instead create one binding for the keyword, which we can then adjust later when we want the keyword to have a different meaning. As no new bindings are introduced, hygiene is preserved. This is similar to the dynamic binding mechanisms we have at run-time (see Section 7.5.26 [SRFI-39], page 599), except that the dynamic binding only occurs during macro expansion. The code after macro expansion remains lexically scoped.

[Syntax] define-syntax-parameter keyword transformer

Binds keyword to the value obtained by evaluating transformer. The transformer provides the default expansion for the syntax parameter, and in the absence of syntaxparameterize, is functionally equivalent to define-syntax. Usually, you will just want to have the transformer throw a syntax error indicating that the keyword is supposed to be used in conjunction with another macro, for example:

     (define-syntax-parameter return
       (lambda (stx)
         (syntax-violation 'return "return used outside of a lambda^" stx)))

[Syntax] syntax-parameterize ((keyword transformer) …) exp … Adjusts keyword … to use the values obtained by evaluating their transformer …, in the expansion of the exp … forms. Each keyword must be bound to a syntaxparameter. syntax-parameterize differs from let-syntax, in that the binding is not shadowed, but adjusted, and so uses of the keyword in the expansion of exp … use the new transformers. This is somewhat similar to how parameterize adjusts the values of regular parameters, rather than creating new bindings.

     (define-syntax lambda^
       (syntax-rules ()
         [(lambda^argument-list body body* ...)
          (lambda argument-list
            (call-with-current-continuation
             (lambda (escape)
               ;; In the body we adjust the 'return' keyword so that calls
               ;; to 'return' are replaced with calls to the escape
               ;; continuation.
               (syntax-parameterize ([return (syntax-rules ()
                                               [(return vals (... ...))
                                                (escape vals (... ...))])])

8 Described in the paper Keeping it Clean with Syntax Parameters by Barzilay, Culpepper and Flatt.

                      body body* ...))))]))
          ;; Now we can write functions that return early.  Here, 'product' will
          ;; return immediately if it sees any 0 element.
          (define product
            (lambda^(list)
                     (fold (lambda (n o)
                             (if (zero? n)
                                 (return 0)
                                 (* n o)))
                           1
                           list)))

6.10.8 Eval-when

As syntax-case macros have the whole power of Scheme available to them, they present a problem regarding time: when a macro runs, what parts of the program are available for the macro to use?

The default answer to this question is that when you import a module (via definemodule or use-modules), that module will be loaded up at expansion-time, as well as at run-time. Additionally, top-level syntactic definitions within one compilation unit made by define-syntax are also evaluated at expansion time, in the order that they appear in the compilation unit (file).

But if a syntactic definition needs to call out to a normal procedure at expansion-time, it might well need need special declarations to indicate that the procedure should be made available at expansion-time.

For example, the following code will work at a REPL, but not in a file:

     ;; incorrect
     (use-modules (srfi srfi-19))
     (define (date) (date->string (current-date)))
     (define-syntax %date (identifier-syntax (date)))
     (define *compilation-date* %date)

It works at a REPL because the expressions are evaluated one-by-one, in order, but if placed in a file, the expressions are expanded one-by-one, but not evaluated until the compiled file is loaded.

The fix is to use eval-when.

     ;; correct: using eval-when
     (use-modules (srfi srfi-19))
     (eval-when (expand load eval)
       (define (date) (date->string (current-date))))
     (define-syntax %date (identifier-syntax (date)))
     (define *compilation-date* %date)

[Syntax] eval-when conditions exp...

Evaluate exp... under the given conditions. Valid conditions include:

expand Evaluate during macro expansion, whether compiling or not. load Evaluate during the evaluation phase of compiled code, e.g. when loading

a compiled module or running compiled code at the REPL.

eval Evaluate during the evaluation phase of non-compiled code.

compile Evaluate during macro expansion, but only when compiling.

In other words, when using the primitive evaluator, eval-when expressions with expand are run during macro expansion, and those with eval are run during the evaluation phase.

When using the compiler, eval-when expressions with either expand or compile are run during macro expansion, and those with load are run during the evaluation phase.

When in doubt, use the three conditions (expand load eval), as in the example above. Other uses of eval-when may void your warranty or poison your cat.

6.10.9 Internal Macros

[Scheme Procedure] make-syntax-transformer name type binding

Construct a syntax transformer object. This is part of Guile’s low-level support for syntax-case.

[Scheme Procedure] macro? obj [C Function] scm_macro_p (obj)

Return #t if obj is a syntax transformer, or #f otherwise.

Note that it’s a bit difficult to actually get a macro as a first-class object; simply naming it (like case) will produce a syntax error. But it is possible to get these objects using module-ref:

     (macro? (module-ref (current-module) 'case))
     ⇒ #t

[Scheme Procedure] macro-type m

[C Function] scm_macro_type (m)

Return the type that was given when m was constructed, via make-syntax-

transformer.

[Scheme Procedure] macro-name m

[C Function] scm_macro_name (m)

Return the name of the macro m.

[Scheme Procedure] macro-binding m

[C Function] scm_macro_binding (m)

Return the binding of the macro m.

[Scheme Procedure] macro-transformer m

[C Function] scm_macro_transformer (m)

Return the transformer of the macro m. This will return a procedure, for which one may ask the docstring. That’s the whole reason this section is documented. Actually a part of the result of macro-binding.

6.11 General Utility Functions

This chapter contains information about procedures which are not cleanly tied to a specific data type. Because of their wide range of applications, they are collected in a utility chapter.

6.11.1 Equality

There are three kinds of core equality predicates in Scheme, described below. The same kinds of comparisons arise in other functions, like memq and friends (see Section 6.7.2.7 [List Searching], page 192).

For all three tests, objects of different types are never equal. So for instance a list and a vector are not equal?, even if their contents are the same. Exact and inexact numbers are considered different types too, and are hence not equal even if their values are the same.

eq? tests just for the same object (essentially a pointer comparison). This is fast, and can be used when searching for a particular object, or when working with symbols or keywords (which are always unique objects).

eqv? extends eq? to look at the value of numbers and characters. It can for instance be used somewhat like = (see Section 6.6.2.8 [Comparison], page 117) but without an error if one operand isn’t a number.

equal? goes further, it looks (recursively) into the contents of lists, vectors, etc. This is good for instance on lists that have been read or calculated in various places and are the same, just not made up of the same pairs. Such lists look the same (when printed), and equal? will consider them the same.

[Scheme Procedure] eq? x y [C Function] scm_eq_p (x, y)

Return #t if x and y are the same object, except for numbers and characters. For example,

          (define x (vector 1 2 3))
          (define y (vector 1 2 3))
          (eq? x x) ⇒ #t
          (eq? x y) ⇒ #f

Numbers and characters are not equal to any other object, but the problem is they’re not necessarily eq? to themselves either. This is even so when the number comes directly from a variable,

          (let ((n (+ 2 3)))
            (eq? n n))  ⇒ *unspecified*

Generally eqv? below should be used when comparing numbers or characters. = (see Section 6.6.2.8 [Comparison], page 117) or char=? (see Section 6.6.3 [Characters], page 129) can be used too.

It’s worth noting that end-of-list (), #t, #f, a symbol of a given name, and a keyword of a given name, are unique objects. There’s just one of each, so for instance no matter how () arises in a program, it’s the same object and can be compared with

     eq?,

     (define x (cdr '(123)))
     (define y (cdr '(456)))
     (eq? x y) ⇒ #t
     (define x (string->symbol "foo"))
     (eq? x 'foo) ⇒ #t

[C Function] int scm_is_eq (SCM x, SCM y)

Return 1 when x and y are equal in the sense of eq?, otherwise return 0. The == operator should not be used on SCM values, an SCM is a C type which cannot necessarily be compared using == (see Section 6.3 [The SCM Type], page 100).

[Scheme Procedure] eqv? x y [C Function] scm_eqv_p (x, y)

Return #t if x and y are the same object, or for characters and numbers the same value. On objects except characters and numbers, eqv? is the same as eq? above, it’s true if x and y are the same object. If x and y are numbers or characters, eqv? compares their type and value. An exact number is not eqv? to an inexact number (even if their value is the same).

     (eqv? 3 (+ 1 2)) ⇒ #t

     (eqv? 1 1.0)  ⇒ #f

[Scheme Procedure] equal? x y [C Function] scm_equal_p (x, y)

Return #t if x and y are the same type, and their contents or value are equal. For a pair, string, vector, array or structure, equal? compares the contents, and does so using the same equal? recursively, so a deep structure can be traversed.

     (equal? (list 1 2 3) (list 1 2 3)) ⇒ #t
     (equal? (list 1 2 3) (vector 1 2 3)) ⇒ #f

For other objects, equal? compares as per eqv? above, which means characters and numbers are compared by type and value (and like eqv?, exact and inexact numbers are not equal?, even if their value is the same).

     (equal? 3 (+ 1 2)) ⇒ #t

     (equal? 1 1.0)  ⇒ #f

Hash tables are currently only compared as per eq?, so two different tables are not equal?, even if their contents are the same. equal? does not support circular data structures, it may go into an infinite loop if asked to compare two circular lists or similar. New application-defined object types (see Section 5.5 [Defining New Types (Smobs)], page 71) have an equalp handler which is called by equal?. This lets an application traverse the contents or control what is considered equal? for two objects of such a type. If there’s no such handler, the default is to just compare as per eq?.

6.11.2 Object Properties

It’s often useful to associate a piece of additional information with a Scheme object even though that object does not have a dedicated slot available in which the additional information could be stored. Object properties allow you to do just that.

Guile’s representation of an object property is a procedure-with-setter (see Section 6.9.8 [Procedures with Setters], page 255) that can be used with the generalized form of set! (REFFIXME) to set and retrieve that property for any Scheme object. So, setting a property looks like this:

     (set! (my-property obj1) value-for-obj1)
     (set! (my-property obj2) value-for-obj2)

And retrieving values of the same property looks like this:

     (my-property obj1)

⇒

     value-for-obj1
     (my-property obj2)

⇒

     value-for-obj2

To create an object property in the first place, use the make-object-property procedure:

     (define my-property (make-object-property))
[Scheme Procedure] make-object-property

Create and return an object property. An object property is a procedure-with-setter that can be called in two ways. (set! (property obj) val) sets obj’s property to val. (property obj) returns the current setting of obj’s property.

A single object property created by make-object-property can associate distinct property values with all Scheme values that are distinguishable by eq? (including, for example, integers).

Internally, object properties are implemented using a weak key hash table. This means that, as long as a Scheme value with property values is protected from garbage collection, its property values are also protected. When the Scheme value is collected, its entry in the property table is removed and so the (ex-) property values are no longer protected by the table.

Guile also implements a more traditional Lispy interface to properties, in which each object has an list of key-value pairs associated with it. Properties in that list are keyed by symbols. This is a legacy interface; you should use weak hash tables or object properties instead.

[Scheme Procedure] object-properties obj

[C Function] scm_object_properties (obj)

Return obj’s property list.

[Scheme Procedure] set-object-properties! obj alist

[C Function] scm_set_object_properties_x (obj, alist)

Set obj’s property list to alist.

[Scheme Procedure] object-property obj key [C Function] scm_object_property (obj, key)

Return the property of obj with name key.

[Scheme Procedure] set-object-property! obj key value [C Function] scm_set_object_property_x (obj, key, value)

In obj’s property list, set the property named key to value.

6.11.3 Sorting

Sorting is very important in computer programs. Therefore, Guile comes with several sorting procedures built-in. As always, procedures with names ending in ! are side-effecting, that means that they may modify their parameters in order to produce their results.

The first group of procedures can be used to merge two lists (which must be already sorted on their own) and produce sorted lists containing all elements of the input lists.

[Scheme Procedure] merge alist blist less [C Function] scm_merge (alist, blist, less)

Merge two already sorted lists into one. Given two lists alist and blist, such that

     (sorted? alist less?) and (sorted? blist less?), return a new list in which the

elements of alist and blist have been stably interleaved so that (sorted? (merge alist blist less?) less?). Note: this does not accept vectors.

[Scheme Procedure] merge! alist blist less [C Function] scm_merge_x (alist, blist, less)

Takes two lists alist and blist such that (sorted? alist less?) and (sorted? blist less?) and returns a new list in which the elements of alist and blist have been

     stably interleaved so that (sorted? (merge alist blist less?) less?). This is

the destructive variant of merge Note: this does not accept vectors.

The following procedures can operate on sequences which are either vectors or list. According to the given arguments, they return sorted vectors or lists, respectively. The first of the following procedures determines whether a sequence is already sorted, the other sort a given sequence. The variants with names starting with stable- are special in that they maintain a special property of the input sequences: If two or more elements are the same according to the comparison predicate, they are left in the same order as they appeared in the input.

[Scheme Procedure] sorted? items less

[C Function] scm_sorted_p (items, less)

Return #t if items is a list or vector such that, for each element x and the next element y of items, (less y x) returns #f. Otherwise return #f.

[Scheme Procedure] sort items less [C Function] scm_sort (items, less)

Sort the sequence items, which may be a list or a vector. less is used for comparing the sequence elements. This is not a stable sort.

[Scheme Procedure] sort! items less [C Function] scm_sort_x (items, less)

Sort the sequence items, which may be a list or a vector. less is used for comparing the sequence elements. The sorting is destructive, that means that the input sequence is modified to produce the sorted result. This is not a stable sort.

[Scheme Procedure] stable-sort items less [C Function] scm_stable_sort (items, less)

Sort the sequence items, which may be a list or a vector. less is used for comparing the sequence elements. This is a stable sort.

[Scheme Procedure] stable-sort! items less [C Function] scm_stable_sort_x (items, less)

Sort the sequence items, which may be a list or a vector. less is used for comparing the sequence elements. The sorting is destructive, that means that the input sequence is modified to produce the sorted result. This is a stable sort.

The procedures in the last group only accept lists or vectors as input, as their names indicate.

[Scheme Procedure] sort-list items less [C Function] scm_sort_list (items, less)

Sort the list items, using less for comparing the list elements. This is a stable sort.

[Scheme Procedure] sort-list! items less [C Function] scm_sort_list_x (items, less)

Sort the list items, using less for comparing the list elements. The sorting is destructive, that means that the input list is modified to produce the sorted result. This is a stable sort.

[Scheme Procedure] restricted-vector-sort! vec less startpos endpos [C Function] scm_restricted_vector_sort_x (vec, less, startpos, endpos)

Sort the vector vec, using less for comparing the vector elements. startpos (inclusively) and endpos (exclusively) delimit the range of the vector which gets sorted. The return value is not specified.

6.11.4 Copying Deep Structures

The procedures for copying lists (see Section 6.7.2 [Lists], page 188) only produce a flat copy of the input list, and currently Guile does not even contain procedures for copying vectors. copy-tree can be used for these application, as it does not only copy the spine of a list, but also copies any pairs in the cars of the input lists.

[Scheme Procedure] copy-tree obj [C Function] scm_copy_tree (obj)

Recursively copy the data tree that is bound to obj, and return the new data structure. copy-tree recurses down the contents of both pairs and vectors (since both cons cells and vector cells may point to arbitrary objects), and stops recursing when it hits any other object.

6.11.5 General String Conversion

When debugging Scheme programs, but also for providing a human-friendly interface, a procedure for converting any Scheme object into string format is very useful. Conversion from/to strings can of course be done with specialized procedures when the data type of the object to convert is known, but with this procedure, it is often more comfortable.

object->string converts an object by using a print procedure for writing to a string port, and then returning the resulting string. Converting an object back from the string is only possible if the object type has a read syntax and the read syntax is preserved by the printing procedure.

[Scheme Procedure] object->string obj [printer] [C Function] scm_object_to_string (obj, printer)

Return a Scheme string obtained by printing obj. Printing function can be specified by the optional second argument printer (default: write).

6.11.6 Hooks

A hook is a list of procedures to be called at well defined points in time. Typically, an application provides a hook h and promises its users that it will call all of the procedures in h at a defined point in the application’s processing. By adding its own procedure to h, an application user can tap into or even influence the progress of the application.

Guile itself provides several such hooks for debugging and customization purposes: these are listed in a subsection below.

When an application first creates a hook, it needs to know how many arguments will be passed to the hook’s procedures when the hook is run. The chosen number of arguments (which may be none) is declared when the hook is created, and all the procedures that are added to that hook must be capable of accepting that number of arguments.

A hook is created using make-hook. A procedure can be added to or removed from a hook using add-hook! or remove-hook!, and all of a hook’s procedures can be removed together using reset-hook!. When an application wants to run a hook, it does so using

run-hook.

6.11.6.1 Hook Usage by Example

Hook usage is shown by some examples in this section. First, we will define a hook of arity 2 — that is, the procedures stored in the hook will have to accept two arguments.

     (define hook (make-hook 2))
     hook
     ⇒ #<hook 2 40286c90>

Now we are ready to add some procedures to the newly created hook with add-hook!. In the following example, two procedures are added, which print different messages and do different things with their arguments.

     (add-hook! hook (lambda (x y)
                         (display "Foo: ")
                         (display (+ x y))
                         (newline)))
     (add-hook! hook (lambda (x y)

                         (display "Bar: ")
                         (display (* x y))
                         (newline)))

Once the procedures have been added, we can invoke the hook using run-hook.

     (run-hook hook 3 4)
     ⊣ Bar: 12
     ⊣ Foo: 7

Note that the procedures are called in the reverse of the order with which they were added. This is because the default behaviour of add-hook! is to add its procedure to the front of the hook’s procedure list. You can force add-hook! to add its procedure to the end of the list instead by providing a third #t argument on the second call to add-hook!.

     (add-hook! hook (lambda (x y)
                         (display "Foo: ")
                         (display (+ x y))
                         (newline)))
     (add-hook! hook (lambda (x y)
                         (display "Bar: ")
                         (display (* x y))
                         (newline))

#t) ; <- Change here!

     (run-hook hook 3 4)
     ⊣ Foo: 7
     ⊣ Bar: 12

6.11.6.2 Hook Reference

When you create a hook with make-hook, you must specify the arity of the procedures which can be added to the hook. If the arity is not given explicitly as an argument to make-hook, it defaults to zero. All procedures of a given hook must have the same arity, and when the procedures are invoked using run-hook, the number of arguments passed must match the arity specified at hook creation time.

The order in which procedures are added to a hook matters. If the third parameter to add-hook! is omitted or is equal to #f, the procedure is added in front of the procedures which might already be on that hook, otherwise the procedure is added at the end. The procedures are always called from the front to the end of the list when they are invoked via

run-hook.

The ordering of the list of procedures returned by hook->list matches the order in which those procedures would be called if the hook was run using run-hook.

Note that the C functions in the following entries are for handling Scheme-level hooks in C. There are also C-level hooks which have their own interface (see Section 6.11.6.4 [C Hooks], page 284).

[Scheme Procedure] make-hook [n args] [C Function] scm_make_hook (n args)

Create a hook for storing procedure of arity n args. n args defaults to zero. The returned value is a hook object to be used with the other hook procedures.

[Scheme Procedure] hook? x [C Function] scm_hook_p (x)

Return #t if x is a hook, #f otherwise.

[Scheme Procedure] hook-empty? hook [C Function] scm_hook_empty_p (hook)

Return #t if hook is an empty hook, #f otherwise.

[Scheme Procedure] add-hook! hook proc [append p] [C Function] scm_add_hook_x (hook, proc, append p)

Add the procedure proc to the hook hook. The procedure is added to the end if append p is true, otherwise it is added to the front. The return value of this procedure is not specified.

[Scheme Procedure] remove-hook! hook proc [C Function] scm_remove_hook_x (hook, proc)

Remove the procedure proc from the hook hook. The return value of this procedure is not specified.

[Scheme Procedure] reset-hook! hook [C Function] scm_reset_hook_x (hook)

Remove all procedures from the hook hook. The return value of this procedure is not specified.

[Scheme Procedure] hook->list hook [C Function] scm_hook_to_list (hook)

Convert the procedure list of hook to a list.

[Scheme Procedure] run-hook hook arg … [C Function] scm_run_hook (hook, args)

Apply all procedures from the hook hook to the arguments arg . . .. The order of the procedure application is first to last. The return value of this procedure is not specified.

If, in C code, you are certain that you have a hook object and well formed argument list for that hook, you can also use scm_c_run_hook, which is identical to scm_run_hook but does no type checking.

[C Function] void scm_c_run_hook (SCM hook, SCM args)

The same as scm_run_hook but without any type checking to confirm that hook is actually a hook object and that args is a well-formed list matching the arity of the hook.

For C code, SCM_HOOKP is a faster alternative to scm_hook_p:

[C Macro] int SCM_HOOKP (x)

Return 1 if x is a Scheme-level hook, 0 otherwise.

6.11.6.3 Handling Scheme-level hooks from C code

Here is an example of how to handle Scheme-level hooks from C code using the above functions.

     if (scm_is_true (scm_hook_p (obj)))
       /* handle Scheme-level hook using C functions */
       scm_reset_hook_x (obj);
     else
       /* do something else (obj is not a hook) */

6.11.6.4 Hooks For C Code.

The hooks already described are intended to be populated by Scheme-level procedures. In addition to this, the Guile library provides an independent set of interfaces for the creation and manipulation of hooks that are designed to be populated by functions implemented in C.

The original motivation here was to provide a kind of hook that could safely be invoked at various points during garbage collection. Scheme-level hooks are unsuitable for this purpose as running them could itself require memory allocation, which would then invoke garbage collection recursively … However, it is also the case that these hooks are easier to work with than the Scheme-level ones if you only want to register C functions with them. So if that is mainly what your code needs to do, you may prefer to use this interface.

To create a C hook, you should allocate storage for a structure of type scm_t_c_hook and then initialize it using scm_c_hook_init.

[C Type] scm_t_c_hook

Data type for a C hook. The internals of this type should be treated as opaque.

[C Enum] scm_t_c_hook_type

Enumeration of possible hook types, which are:

     SCM_C_HOOK_NORMAL

Type of hook for which all the registered functions will always be called.

     SCM_C_HOOK_OR

Type of hook for which the sequence of registered functions will be called only until one of them returns C true (a non-NULL pointer).

     SCM_C_HOOK_AND

Type of hook for which the sequence of registered functions will be called only until one of them returns C false (a NULL pointer).

[C Function] void scm_c_hook_init (scm t c hook *hook, void *hook data,

scm t c hook type type) Initialize the C hook at memory pointed to by hook. type should be one of the values of the scm_t_c_hook_type enumeration, and controls how the hook functions will be called. hook data is a closure parameter that will be passed to all registered hook functions when they are called.

To add or remove a C function from a C hook, use scm_c_hook_add or scm_c_hook_ remove. A hook function must expect three void * parameters which are, respectively: hook data The hook closure data that was specified at the time the hook was initialized

     by scm_c_hook_init.

func data The function closure data that was specified at the time that that function was registered with the hook by scm_c_hook_add.

data The call closure data specified by the scm_c_hook_run call that runs the hook.

[C Type] scm_t_c_hook_function

Function type for a C hook function: takes three void * parameters and returns a void * result.

[C Function] void scm_c_hook_add (scm t c hook *hook, scm t c hook function

func, void *func data, int appendp) Add function func, with function closure data func data, to the C hook hook. The new function is appended to the hook’s list of functions if appendp is non-zero, otherwise prepended.

[C Function] void scm_c_hook_remove (scm t c hook *hook,

scm t c hook function func, void *func data) Remove function func, with function closure data func data, from the C hook hook. scm_c_hook_remove checks both func and func data so as to allow for the same func being registered multiple times with different closure data.

Finally, to invoke a C hook, call the scm_c_hook_run function specifying the hook and the call closure data for this run:

[C Function] void * scm_c_hook_run (scm t c hook *hook, void *data)

Run the C hook hook will call closure data data. Subject to the variations for hook

types SCM_C_HOOK_OR and SCM_C_HOOK_AND, scm_c_hook_run calls hook’s registered

functions in turn, passing them the hook’s closure data, each function’s closure data, and the call closure data. scm_c_hook_run’s return value is the return value of the last function to be called.

6.11.6.5 Hooks for Garbage Collection

Whenever Guile performs a garbage collection, it calls the following hooks in the order shown.

[C Hook] scm_before_gc_c_hook

C hook called at the very start of a garbage collection, after setting scm_gc_running_ p to 1, but before entering the GC critical section. If garbage collection is blocked because scm_block_gc is non-zero, GC exits early soon after calling this hook, and no further hooks will be called.

[C Hook] scm_before_mark_c_hook

C hook called before beginning the mark phase of garbage collection, after the GC thread has entered a critical section.

[C Hook] scm_before_sweep_c_hook

C hook called before beginning the sweep phase of garbage collection. This is the same as at the end of the mark phase, since nothing else happens between marking and sweeping.

[C Hook] scm_after_sweep_c_hook

C hook called after the end of the sweep phase of garbage collection, but while the GC thread is still inside its critical section.

[C Hook] scm_after_gc_c_hook

C hook called at the very end of a garbage collection, after the GC thread has left its critical section.

[Scheme Hook] after-gc-hook

Scheme hook with arity 0. This hook is run asynchronously (see Section 6.21.2 [Asyncs], page 411) soon after the GC has completed and any other events that were deferred during garbage collection have been processed. (Also accessible from C

     with the name scm_after_gc_hook.)

All the C hooks listed here have type SCM_C_HOOK_NORMAL, are initialized with hook closure data NULL, are invoked by scm_c_hook_run with call closure data NULL.

The Scheme hook after-gc-hook is particularly useful in conjunction with guardians (see Section 6.18.4 [Guardians], page 380). Typically, if you are using a guardian, you want to call the guardian after garbage collection to see if any of the objects added to the guardian have been collected. By adding a thunk that performs this call to after-gc-hook, you can ensure that your guardian is tested after every garbage collection cycle.

6.11.6.6 Hooks into the Guile REPL

6.12 Definitions and Variable Bindings

Scheme supports the definition of variables in different contexts. Variables can be defined at the top level, so that they are visible in the entire program, and variables can be defined locally to procedures and expressions. This is important for modularity and data abstraction.

6.12.1 Top Level Variable Definitions

At the top level of a program (i.e., not nested within any other expression), a definition of the form

     (define a value)

defines a variable called a and sets it to the value value.

If the variable already exists in the current module, because it has already been created by a previous define expression with the same name, its value is simply changed to the new value. In this case, then, the above form is completely equivalent to

     (set! a value)

This equivalence means that define can be used interchangeably with set! to change the value of variables at the top level of the REPL or a Scheme source file. It is useful during interactive development when reloading a Scheme file that you have modified, because it allows the define expressions in that file to work as expected both the first time that the file is loaded and on subsequent occasions.

Note, though, that define and set! are not always equivalent. For example, a set! is not allowed if the named variable does not already exist, and the two expressions can behave differently in the case where there are imported variables visible from another module.

[Scheme Syntax] define name value Create a top level variable named name with value value. If the named variable already exists, just change its value. The return value of a define expression is unspecified.

The C API equivalents of define are scm_define and scm_c_define, which differ from each other in whether the variable name is specified as a SCM symbol or as a null-terminated C string.

[C Function] scm_define (sym, value) [C Function] scm_c_define (const char *name, value)

C equivalents of define, with variable name specified either by sym, a symbol, or by name, a null-terminated C string. Both variants return the new or preexisting variable object.

define (when it occurs at top level), scm_define and scm_c_define all create or set the value of a variable in the top level environment of the current module. If there was not already a variable with the specified name belonging to the current module, but a similarly named variable from another module was visible through having been imported, the newly created variable in the current module will shadow the imported variable, such that the imported variable is no longer visible.

Attention: Scheme definitions inside local binding constructs (see Section 6.12.2 [Local Bindings], page 287) act differently (see Section 6.12.3 [Internal Definitions], page 289).

Many people end up in a development style of adding and changing definitions at runtime, building out their program without restarting it. (You can do this using reload-module, the reload REPL command, the load procedure, or even just pasting code into a REPL.) If you are one of these people, you will find that sometimes you there are some variables that you don’t want to redefine all the time. For these, use define-once.

[Scheme Syntax] define-once name value

Create a top level variable named name with value value, but only if name is not already bound in the current module.

Old Lispers probably know define-once under its Lisp name, defvar.

6.12.2 Local Variable Bindings

As opposed to definitions at the top level, which creates bindings that are visible to all code in a module, it is also possible to define variables which are only visible in a well-defined part of the program. Normally, this part of a program will be a procedure or a subexpression of a procedure.

With the constructs for local binding (let, let*, letrec, and letrec*), the Scheme language has a block structure like most other programming languages since the days of Algol 60. Readers familiar to languages like C or Java should already be used to this concept, but the family of let expressions has a few properties which are well worth knowing.

The most basic local binding construct is let.

[syntax] let bindings body bindings has the form

     ((variable1 init1) ...)

that is zero or more two-element lists of a variable and an arbitrary expression each. All variable names must be distinct. A let expression is evaluated as follows.

• All init expressions are evaluated.

• New storage is allocated for the variables.

• The values of the init expressions are stored into the variables.

• The expressions in body are evaluated in order, and the value of the last expression is returned as the value of the let expression.

The init expressions are not allowed to refer to any of the variables.

The other binding constructs are variations on the same theme: making new values, binding them to variables, and executing a body in that new, extended lexical context.

[syntax] let* bindings body Similar to let, but the variable bindings are performed sequentially, that means that all init expression are allowed to use the variables defined on their left in the binding list. A let* expression can always be expressed with nested let expressions.

     (let* ((a 1) (b a))
        b)

≡

     (let ((a 1))
       (let ((b a))
         b))

[syntax] letrec bindings body Similar to let, but it is possible to refer to the variable from lambda expression created in any of the inits. That is, procedures created in the init expression can recursively refer to the defined variables.

     (letrec ((even? (lambda (n)
                       (if (zero? n)
                           #t
                           (odd? (- n 1)))))
              (odd? (lambda (n)
                       (if (zero? n)
                           #f
                           (even? (- n 1))))))
       (even? 88))

⇒

     #t

Note that while the init expressions may refer to the new variables, they may not access their values. For example, making the even? function above creates a closure (see Section 3.4 [About Closure], page 26) referencing the odd? variable. But odd? can’t be called until after execution has entered the body.

[syntax] letrec* bindings body Similar to letrec, except the init expressions are bound to their variables in order. letrec* thus relaxes the letrec restriction, in that later init expressions may refer to the values of previously bound variables.

          (letrec ((a 42)
                   (b (+ a 10)))  ;; Illegal access
            (* a b))
          ;; The behavior of the expression above is unspecified
          (letrec* ((a 42)
                    (b (+ a 10)))
            (* a b))
          ⇒ 2184

There is also an alternative form of the let form, which is used for expressing iteration. Because of the use as a looping construct, this form (the named let) is documented in the section about iteration (see Section 6.13.4 [while do], page 294)

6.12.3 Internal definitions

A define form which appears inside the body of a lambda, let, let*, letrec, letrec* or equivalent expression is called an internal definition. An internal definition differs from a top level definition (see Section 6.12.1 [Top Level], page 286), because the definition is only visible inside the complete body of the enclosing form. Let us examine the following example.

     (let ((frumble "froz"))
        (define banana (lambda () (apple 'peach)))
        (define apple (lambda (x) x))
        (banana))

⇒

     peach

Here the enclosing form is a let, so the defines in the let-body are internal definitions. Because the scope of the internal definitions is the complete body of the let-expression, the lambda-expression which gets bound to the variable banana may refer to the variable apple, even though its definition appears lexically after the definition of banana. This is because a sequence of internal definition acts as if it were a letrec* expression.

     (let ()
       (define a 1)
       (define b 2)
       (+ a b))

is equivalent to

     (let ()
       (letrec* ((a 1) (b 2))
         (+ a b)))

Internal definitions are only allowed at the beginning of the body of an enclosing expression. They may not be mixed with other expressions.

Another noteworthy difference to top level definitions is that within one group of internal definitions all variable names must be distinct. That means where on the top level a second define for a given variable acts like a set!, an exception is thrown for internal definitions with duplicate bindings.

As a historical note, it used to be that internal bindings were expanded in terms of letrec, not letrec*. This was the situation for the R5RS report and before. However with the R6RS, it was recognized that sequential definition was a more intuitive expansion, as in the following case:

     (let ()
       (define a 1)
       (define b (+ a a))
       (+ a b))

Guile decided to follow the R6RS in this regard, and now expands internal definitions using

letrec*.

6.12.4 Querying variable bindings

Guile provides a procedure for checking whether a symbol is bound in the top level environment.

[Scheme Procedure] defined? sym [module] [C Function] scm_defined_p (sym, module)

Return #t if sym is defined in the module module or the current module when module is not specified; otherwise return #f.

6.12.5 Binding multiple return values

[Syntax] define-values formals expression The expression is evaluated, and the formals are bound to the return values in the same way that the formals in a lambda expression are matched to the arguments in a procedure call.

     (define-values (q r) (floor/ 10 3))
     (list q r) ⇒ (3 1)
     (define-values (x . y) (values 1 2 3))
     x ⇒ 1
     y ⇒ (2 3)
     (define-values x (values 1 2 3))
     x ⇒ (1 2 3)

6.13 Controlling the Flow of Program Execution

See Section 5.4.3 [Control Flow], page 66 for a discussion of how the more general control flow of Scheme affects C code.

6.13.1 Sequencing and Splicing

As an expression, the begin syntax is used to evaluate a sequence of sub-expressions in order. Consider the conditional expression below:

     (if (> x 0)
         (begin (display "greater") (newline)))

If the test is true, we want to display “greater” to the current output port, then display a newline. We use begin to form a compound expression out of this sequence of subexpressions.

[syntax] begin expr …

The expression(s) are evaluated in left-to-right order and the value of the last expression is returned as the value of the begin-expression. This expression type is used when the expressions before the last one are evaluated for their side effects.

The begin syntax has another role in definition context (see Section 6.12.3 [Internal Definitions], page 289). A begin form in a definition context splices its subforms into its place. For example, consider the following procedure:

     (define (make-seal)
       (define-sealant seal open)
       (values seal open))

Let us assume the existence of a define-sealant macro that expands out to some definitions wrapped in a begin, like so:

     (define (make-seal)
       (begin
         (define seal-tag
           (list 'seal))
         (define (seal x)
           (cons seal-tag x))
         (define (sealed? x)
           (and (pair? x) (eq? (car x) seal-tag)))
         (define (open x)
           (if (sealed? x)
               (cdr x)
               (error "Expected a sealed value:" x))))
       (values seal open))

Here, because the begin is in definition context, its subforms are spliced into the place of the begin. This allows the definitions created by the macro to be visible to the following expression, the values form.

It is a fine point, but splicing and sequencing are different. It can make sense to splice zero forms, because it can make sense to have zero internal definitions before the expressions in a procedure or lexical binding form. However it does not make sense to have a sequence of zero expressions, because in that case it would not be clear what the value of the sequence would be, because in a sequence of zero expressions, there can be no last value. Sequencing zero expressions is an error.

It would be more elegant in some ways to eliminate splicing from the Scheme language, and without macros (see Section 6.10 [Macros], page 257), that would be a good idea. But it is useful to be able to write macros that expand out to multiple definitions, as in define-sealant above, so Scheme abuses the begin form for these two tasks.

6.13.2 Simple Conditional Evaluation

Guile provides three syntactic constructs for conditional evaluation. if is the normal ifthen-else expression (with an optional else branch), cond is a conditional expression with multiple branches and case branches if an expression has one of a set of constant values.

[syntax] if test consequent [alternate] All arguments may be arbitrary expressions. First, test is evaluated. If it returns a true value, the expression consequent is evaluated and alternate is ignored. If test evaluates to #f, alternate is evaluated instead. The values of the evaluated branch (consequent or alternate) are returned as the values of the if expression. When alternate is omitted and the test evaluates to #f, the value of the expression is not specified.

When you go to write an if without an alternate (a one-armed if), part of what you are expressing is that you don’t care about the return value (or values) of the expression. As such, you are more interested in the effect of evaluating the consequent expression. (By convention, we use the word statement to refer to an expression that is evaluated for effect, not for value).

In such a case, it is considered more clear to express these intentions with these special forms, when and unless. As an added bonus, these forms accept multiple statements to evaluate, which are implicitly wrapped in a begin.

[Scheme Syntax] when test statement1 statement2 ... [Scheme Syntax] unless test statement1 statement2 ...

The actual definitions of these forms are in many ways their most clear documentation:

     (define-syntax-rule (when test stmt stmt* ...)
       (if test (begin stmt stmt* ...)))
     (define-syntax-rule (unless condition stmt stmt* ...)
       (if (not test) (begin stmt stmt* ...)))

That is to say, when evaluates its consequent statements in order if test is true. unless is the opposite: it evaluates the statements if test is false.

[syntax] cond clause1 clause2 …

Each cond-clause must look like this:

     (test expression ...)

where test and expression are arbitrary expression, or like this

     (test => expression)

where expression must evaluate to a procedure. The tests of the clauses are evaluated in order and as soon as one of them evaluates to a true values, the corresponding expressions are evaluated in order and the last value is returned as the value of the cond-expression. For the => clause type, expression is evaluated and the resulting procedure is applied to the value of test. The result of this procedure application is then the result of the cond-expression. One additional cond-clause is available as an extension to standard Scheme:

     (test guard => expression)

where guard and expression must evaluate to procedures. For this clause type, test may return multiple values, and cond ignores its boolean state; instead, cond evaluates guard and applies the resulting procedure to the value(s) of test, as if guard were the consumer argument of call-with-values. If the result of that procedure call is a true value, it evaluates expression and applies the resulting procedure to the value(s) of test, in the same manner as the guard was called. The test of the last clause may be the symbol else. Then, if none of the preceding tests is true, the expressions following the else are evaluated to produce the result of the cond-expression.

[syntax] case key clause1 clause2 …

key may be any expression, and the clauses must have the form

     ((datum1 ...) expr1 expr2 ...)

or

     ((datum1 ...) => expression)

and the last clause may have the form

     (else expr1 expr2 ...)

or

     (else => expression)

All datums must be distinct. First, key is evaluated. The result of this evaluation is compared against all datum values using eqv?. When this comparison succeeds, the expression(s) following the datum are evaluated from left to right, returning the value of the last expression as the result of the case expression. If the key matches no datum and there is an else-clause, the expressions following the else are evaluated. If there is no such clause, the result of the expression is unspecified. For the => clause types, expression is evaluated and the resulting procedure is applied to the value of key. The result of this procedure application is then the result of the case-expression.

6.13.3 Conditional Evaluation of a Sequence of Expressions

and and or evaluate all their arguments in order, similar to begin, but evaluation stops as soon as one of the expressions evaluates to false or true, respectively.

[syntax] and expr … Evaluate the exprs from left to right and stop evaluation as soon as one expression evaluates to #f; the remaining expressions are not evaluated. The value of the last evaluated expression is returned. If no expression evaluates to #f, the value of the last expression is returned. If used without expressions, #t is returned.

[syntax] or expr … Evaluate the exprs from left to right and stop evaluation as soon as one expression evaluates to a true value (that is, a value different from #f); the remaining expressions are not evaluated. The value of the last evaluated expression is returned. If all expressions evaluate to #f, #f is returned. If used without expressions, #f is returned.

6.13.4 Iteration mechanisms

Scheme has only few iteration mechanisms, mainly because iteration in Scheme programs is normally expressed using recursion. Nevertheless, R5RS defines a construct for programming loops, calling do. In addition, Guile has an explicit looping syntax called while.

[syntax] do ((variable init [step]) …) (test expr …) body … Bind variables and evaluate body until test is true. The return value is the last expr after test, if given. A simple example will illustrate the basic form,

     (do ((i 1 (1+ i)))
         ((> i 4))
       (display i))
     ⊣ 1234

Or with two variables and a final return value,

     (do ((i 1 (1+ i))
          (p 3 (* 3 p)))
         ((> i 4)
          p)
       (format #t "3**~s is ~s\n" i p))

⊣

     3**1 is 3
     3**2 is 9
     3**3 is 27
     3**4 is 81

⇒

     789

The variable bindings are established like a let, in that the expressions are all evaluated and then all bindings made. When iterating, the optional step expressions are evaluated with the previous bindings in scope, then new bindings all made. The test expression is a termination condition. Looping stops when the test is true. It’s evaluated before running the body each time, so if it’s true the first time then body is not run at all. The optional exprs after the test are evaluated at the end of looping, with the final variable bindings available. The last expr gives the return value, or if there are no exprs the return value is unspecified. Each iteration establishes bindings to fresh locations for the variables, like a new let for each iteration. This is done for variables without step expressions too. The following illustrates this, showing how a new i is captured by the lambda in each iteration (see Section 3.4 [The Concept of Closure], page 26).

     (define lst '())
     (do ((i 1 (1+ i)))
         ((> i 4))

       (set! lst (cons (lambda () i) lst)))
     (map (lambda (proc) (proc)) lst)

⇒

     (4 3 2 1)

[syntax] while cond body … Run a loop executing the body forms while cond is true. cond is tested at the start of each iteration, so if it’s #f the first time then body is not executed at all. Within while, two extra bindings are provided, they can be used from both cond and body.

[Scheme Procedure] break break-arg … Break out of the while form.

[Scheme Procedure] continue

Abandon the current iteration, go back to the start and test cond again, etc.

If the loop terminates normally, by the cond evaluating to #f, then the while expression as a whole evaluates to #f. If it terminates by a call to break with some number of arguments, those arguments are returned from the while expression, as multiple values. Otherwise if it terminates by a call to break with no arguments, then return value is #t.

     (while #f (error "not reached")) ⇒ #f
     (while #t (break)) ⇒ #t
     (while #t (break 1 2 3)) ⇒ 1 2 3

Each while form gets its own break and continue procedures, operating on that while. This means when loops are nested the outer break can be used to escape all the way out. For example,

     (while (test1)
       (let ((outer-break break))
         (while (test2)
           (if (something)
             (outer-break #f))
           ...)))

Note that each break and continue procedure can only be used within the dynamic extent of its while. Outside the while their behaviour is unspecified.

Another very common way of expressing iteration in Scheme programs is the use of the so-called named let.

Named let is a variant of let which creates a procedure and calls it in one step. Because of the newly created procedure, named let is more powerful than do–it can be used for iteration, but also for arbitrary recursion.

[syntax] let variable bindings body For the definition of bindings see the documentation about let (see Section 6.12.2 [Local Bindings], page 287). Named let works as follows:

• A new procedure which accepts as many arguments as are in bindings is created and bound locally (using let) to variable. The new procedure’s formal argument names are the name of the variables.

• The body expressions are inserted into the newly created procedure.

• The procedure is called with the init expressions as the formal arguments.

The next example implements a loop which iterates (by recursion) 1000 times.

     (let lp ((x 1000))
       (if (positive? x)
           (lp (- x 1))
           x))

⇒

     0

6.13.5 Prompts

Prompts are control-flow barriers between different parts of a program. In the same way that a user sees a shell prompt (e.g., the Bash prompt) as a barrier between the operating system and her programs, Scheme prompts allow the Scheme programmer to treat parts of programs as if they were running in different operating systems.

We use this roundabout explanation because, unless you’re a functional programming junkie, you probably haven’t heard the term, “delimited, composable continuation”. That’s OK; it’s a relatively recent topic, but a very useful one to know about.

6.13.5.1 Prompt Primitives

Guile’s primitive delimited control operators are call-with-prompt and abort-to-prompt.

[Scheme Procedure] call-with-prompt tag thunk handler

Set up a prompt, and call thunk within that prompt. During the dynamic extent of the call to thunk, a prompt named tag will be present in the dynamic context, such that if a user calls abort-to-prompt (see below) with that tag, control rewinds back to the prompt, and the handler is run. handler must be a procedure. The first argument to handler will be the state of the computation begun when thunk was called, and ending with the call to abort-to-prompt. The remaining arguments to handler are those passed to abort-to-prompt.

[Scheme Procedure] make-prompt-tag [stem]

Make a new prompt tag. Currently prompt tags are generated symbols. This may change in some future Guile version.

[Scheme Procedure] default-prompt-tag

Return the default prompt tag. Having a distinguished default prompt tag allows some useful prompt and abort idioms, discussed in the next section.

[Scheme Procedure] abort-to-prompt tag val1 val2 …

Unwind the dynamic and control context to the nearest prompt named tag, also passing the given values.

C programmers may recognize call-with-prompt and abort-to-prompt as a fancy kind of setjmp and longjmp, respectively. Prompts are indeed quite useful as non-local escape mechanisms. Guile’s catch and throw are implemented in terms of prompts. Prompts are more convenient than longjmp, in that one has the opportunity to pass multiple values to the jump target.

Also unlike longjmp, the prompt handler is given the full state of the process that was aborted, as the first argument to the prompt’s handler. That state is the continuation of the computation wrapped by the prompt. It is a delimited continuation, because it is not the whole continuation of the program; rather, just the computation initiated by the call

to call-with-prompt.

The continuation is a procedure, and may be reinstated simply by invoking it, with any number of values. Here’s where things get interesting, and complicated as well. Besides being described as delimited, continuations reified by prompts are also composable, because invoking a prompt-saved continuation composes that continuation with the current one.

Imagine you have saved a continuation via call-with-prompt:

     (define cont
       (call-with-prompt
        ;; tag
        'foo
        ;; thunk
        (lambda ()
          (+ 34 (abort-to-prompt 'foo)))
        ;; handler
        (lambda (k) k)))

The resulting continuation is the addition of 34. It’s as if you had written:

     (define cont
       (lambda (x)
         (+ 34 x)))

So, if we call cont with one numeric value, we get that number, incremented by 34:

     (cont 8)
     ⇒ 42
     (* 2 (cont 8))
     ⇒ 84

The last example illustrates what we mean when we say, "composes with the current continuation". We mean that there is a current continuation – some remaining things to compute, like (lambda (x) (* x 2)) – and that calling the saved continuation doesn’t wipe out the current continuation, it composes the saved continuation with the current one.

We’re belaboring the point here because traditional Scheme continuations, as discussed in the next section, aren’t composable, and are actually less expressive than continuations captured by prompts. But there’s a place for them both.

Before moving on, we should mention that if the handler of a prompt is a lambda expression, and the first argument isn’t referenced, an abort to that prompt will not cause a continuation to be reified. This can be an important efficiency consideration to keep in mind.

One example where this optimization matters is escape continuations. Escape continuations are delimited continuations whose only use is to make a non-local exit—i.e., to escape from the current continuation. Such continuations are invoked only once, and for this reason they are sometimes called one-shot continuations. A common use of escape continuations is when throwing an exception (see Section 6.13.8 [Exceptions], page 303).

The constructs below are syntactic sugar atop prompts to simplify the use of escape continuations.

[Scheme Procedure] call-with-escape-continuation proc

[Scheme Procedure] call/ec proc

Call proc with an escape continuation.

In the example below, the return continuation is used to escape the continuation of the call to fold.

          (use-modules (ice-9 control)
                       (srfi srfi-1))
          (define (prefix x lst)
            ;; Return all the elements before the first occurrence
            ;; of X in LST.
            (call/ec
              (lambda (return)
                (fold (lambda (element prefix)
                        (if (equal? element x)
                            (return (reverse prefix))  ; escape `fold'
                            (cons element prefix)))
                      '()
                      lst))))
          (prefix 'a '(0 1 2 a 3 4 5))
          ⇒ (0 1 2)
[Scheme Syntax] let-escape-continuation k body . . .

[Scheme Syntax] let/ec k body …

Bind k within body to an escape continuation.

     This is equivalent to (call/ec (lambda (k) body ...)).

6.13.5.2 Shift, Reset, and All That

There is a whole zoo of delimited control operators, and as it does not seem to be a bounded set, Guile implements support for them in a separate module:

     (use-modules (ice-9 control))

Firstly, we have a helpful abbreviation for the call-with-prompt operator.

[Scheme Syntax] % expr [Scheme Syntax] % expr handler [Scheme Syntax] % tag expr handler

Evaluate expr in a prompt, optionally specifying a tag and a handler. If no tag is given, the default prompt tag is used.

If no handler is given, a default handler is installed. The default handler accepts a procedure of one argument, which will called on the captured continuation, within a prompt.

Sometimes it’s easier just to show code, as in this case:

          (define (default-prompt-handler k proc)
            (% (default-prompt-tag)
               (proc k)
               default-prompt-handler))

The % symbol is chosen because it looks like a prompt.

Likewise there is an abbreviation for abort-to-prompt, which assumes the default prompt tag:

[Scheme Procedure] abort val1 val2 … Abort to the default prompt tag, passing val1 val2 … to the handler.

As mentioned before, (ice-9 control) also provides other delimited control operators. This section is a bit technical, and first-time users of delimited continuations should probably come back to it after some practice with %.

Still here? So, when one implements a delimited control operator like call-with-prompt, one needs to make two decisions. Firstly, does the handler run within or outside the prompt? Having the handler run within the prompt allows an abort inside the handler to return to the same prompt handler, which is often useful. However it prevents tail calls from the handler, so it is less general.

Similarly, does invoking a captured continuation reinstate a prompt? Again we have the tradeoff of convenience versus proper tail calls.

These decisions are captured in the Felleisen F operator. If neither the continuations nor the handlers implicitly add a prompt, the operator is known as –F–. This is the case

for Guile’s call-with-prompt and abort-to-prompt.

If both continuation and handler implicitly add prompts, then the operator is +F+. shift and reset are such operators.

[Scheme Syntax] reset body1 body2 … Establish a prompt, and evaluate body1 body2 … within that prompt.

The prompt handler is designed to work with shift, described below.

[Scheme Syntax] shift cont body1 body2 … Abort to the nearest reset, and evaluate body1 body2 … in a context in which the captured continuation is bound to cont.

As mentioned above, taken together, the body1 body2 … expressions and the invocations of cont implicitly establish a prompt.

Interested readers are invited to explore Oleg Kiselyov’s wonderful web site at http:// okmij.org/ftp/, for more information on these operators.

6.13.6 Continuations

A “continuation” is the code that will execute when a given function or expression returns. For example, consider

(define (foo)
  (display "hello\n")
  (display (bar)) (newline)
  (exit))

The continuation from the call to bar comprises a display of the value returned, a newline and an exit. This can be expressed as a function of one argument.

(lambda (r)
  (display r) (newline)
  (exit))

In Scheme, continuations are represented as special procedures just like this. The special property is that when a continuation is called it abandons the current program location and jumps directly to that represented by the continuation.

A continuation is like a dynamic label, capturing at run-time a point in program execution, including all the nested calls that have lead to it (or rather the code that will execute when those calls return).

Continuations are created with the following functions.

[Scheme Procedure] call-with-current-continuation proc

[Scheme Procedure] call/cc proc

Capture the current continuation and call (proc cont) with it. The return value is the value returned by proc, or when (cont value) is later invoked, the return is the value passed. Normally cont should be called with one argument, but when the location resumed is expecting multiple values (see Section 6.13.7 [Multiple Values], page 301) then they should be passed as multiple arguments, for instance (cont x y z). cont may only be used from the same side of a continuation barrier as it was created (see Section 6.13.12 [Continuation Barriers], page 315), and in a multi-threaded program only from the thread in which it was created. The call to proc is not part of the continuation captured, it runs only when the continuation is created. Often a program will want to store cont somewhere for later use; this can be done in proc.

The call in the name call-with-current-continuation refers to the way a call to

proc gives the newly created continuation. It’s not related to the way a call is used later to invoke that continuation.

call/cc is an alias for call-with-current-continuation. This is in common use

since the latter is rather long.

Here is a simple example,

(define kont #f)
(format #t "the return is ~a\n"

             (call/cc (lambda (k)
                        (set! kont k)
                        1)))
     ⇒ the return is 1
     (kont 2)
     ⇒ the return is 2

call/cc captures a continuation in which the value returned is going to be displayed by format. The lambda stores this in kont and gives an initial return 1 which is displayed. The later invocation of kont resumes the captured point, but this time returning 2, which is displayed.

When Guile is run interactively, a call to format like this has an implicit return back to the read-eval-print loop. call/cc captures that like any other return, which is why interactively kont will come back to read more input.

C programmers may note that call/cc is like setjmp in the way it records at runtime a point in program execution. A call to a continuation is like a longjmp in that it abandons the present location and goes to the recorded one. Like longjmp, the value passed to the continuation is the value returned by call/cc on resuming there. However longjmp can only go up the program stack, but the continuation mechanism can go anywhere.

When a continuation is invoked, call/cc and subsequent code effectively “returns” a second time. It can be confusing to imagine a function returning more times than it was called. It may help instead to think of it being stealthily re-entered and then program flow going on as normal.

dynamic-wind (see Section 6.13.10 [Dynamic Wind], page 309) can be used to ensure setup and cleanup code is run when a program locus is resumed or abandoned through the continuation mechanism.

Continuations are a powerful mechanism, and can be used to implement almost any sort of control structure, such as loops, coroutines, or exception handlers.

However the implementation of continuations in Guile is not as efficient as one might hope, because Guile is designed to cooperate with programs written in other languages, such as C, which do not know about continuations. Basically continuations are captured by a block copy of the stack, and resumed by copying back.

For this reason, continuations captured by call/cc should be used only when there is no other simple way to achieve the desired result, or when the elegance of the continuation mechanism outweighs the need for performance.

Escapes upwards from loops or nested functions are generally best handled with prompts (see Section 6.13.5 [Prompts], page 296). Coroutines can be efficiently implemented with cooperating threads (a thread holds a full program stack but doesn’t copy it around the way continuations do).

6.13.7 Returning and Accepting Multiple Values

Scheme allows a procedure to return more than one value to its caller. This is quite different to other languages which only allow single-value returns. Returning multiple values is different from returning a list (or pair or vector) of values to the caller, because conceptually not one compound object is returned, but several distinct values.

The primitive procedures for handling multiple values are values and call-with-values. values is used for returning multiple values from a procedure. This is done by placing a call to values with zero or more arguments in tail position in a procedure body. call-with-values combines a procedure returning multiple values with a procedure which accepts these values as parameters.

[Scheme Procedure] values arg … [C Function] scm_values (args)

Delivers all of its arguments to its continuation. Except for continuations created by the call-with-values procedure, all continuations take exactly one value. The effect of passing no value or more than one value to continuations that were not created by

call-with-values is unspecified.

For scm_values, args is a list of arguments and the return is a multiple-values object which the caller can return. In the current implementation that object shares structure with args, so args should not be modified subsequently.

[C Function] SCM scm_c_values (SCM *base, size t n)

scm_c_values is an alternative to scm_values. It creates a new values object, and copies into it the n values starting from base. Currently this creates a list and passes it to scm_values, but we expect that in the future we will be able to use more a efficient representation.

[C Function] size_t scm_c_nvalues (SCM obj)

If obj is a multiple-values object, returns the number of values it contains. Otherwise returns 1.

[C Function] SCM scm_c_value_ref (SCM obj, size t idx)

Returns the value at the position specified by idx in obj. Note that obj will ordinarily be a multiple-values object, but it need not be. Any other object represents a single value (itself), and is handled appropriately.

[Scheme Procedure] call-with-values producer consumer

Calls its producer argument with no values and a continuation that, when passed some values, calls the consumer procedure with those values as arguments. The continuation for the call to consumer is the continuation of the call to call-with-

values.
     (call-with-values (lambda () (values 4 5))
                       (lambda (a b) b))
     ⇒ 5
     (call-with-values * -)
     ⇒ -1

In addition to the fundamental procedures described above, Guile has a module which exports a syntax called receive, which is much more convenient. This is in the (ice-9 receive) and is the same as specified by SRFI-8 (see Section 7.5.7 [SRFI-8], page 575).

     (use-modules (ice-9 receive))

[library syntax] receive formals expr body … Evaluate the expression expr, and bind the result values (zero or more) to the formal arguments in formals. formals is a list of symbols, like the argument list in a lambda (see Section 6.9.1 [Lambda], page 244). After binding the variables, the expressions in body … are evaluated in order, the return value is the result from the last expression.

For example getting results from partition in SRFI-1 (see Section 7.5.3 [SRFI-1], page 552),

          (receive (odds evens)
              (partition odd? '(7 4 2 8 3))
            (display odds)
            (display " and ")
            (display evens))
          ⊣ (7 3) and (4 2 8)

6.13.8 Exceptions

A common requirement in applications is to want to jump non-locally from the depths of a computation back to, say, the application’s main processing loop. Usually, the place that is the target of the jump is somewhere in the calling stack of procedures that called the procedure that wants to jump back. For example, typical logic for a key press driven application might look something like this:

     main-loop:
       read the next key press and call dispatch-key
     dispatch-key:
       lookup the key in a keymap and call an appropriate procedure,
       say find-file
     find-file:
       interactively read the required file name, then call
       find-specified-file
     find-specified-file:
       check whether file exists; if not, jump back to main-loop
       ...

The jump back to main-loop could be achieved by returning through the stack one procedure at a time, using the return value of each procedure to indicate the error condition, but Guile (like most modern programming languages) provides an additional mechanism called exception handling that can be used to implement such jumps much more conveniently.

6.13.8.1 Exception Terminology

There are several variations on the terminology for dealing with non-local jumps. It is useful to be aware of them, and to realize that they all refer to the same basic mechanism.

• Actually making a non-local jump may be called raising an exception, raising a signal, throwing an exception or doing a long jump. When the jump indicates an error condition, people may talk about signalling, raising or throwing an error.

• Handling the jump at its target may be referred to as catching or handling the exception, signal or, where an error condition is involved, error.

Where signal and signalling are used, special care is needed to avoid the risk of confusion with POSIX signals.

This manual prefers to speak of throwing and catching exceptions, since this terminology matches the corresponding Guile primitives.

The exception mechanism described in this section has connections with delimited continuations (see Section 6.13.5 [Prompts], page 296). In particular, throwing an exception is akin to invoking an escape continuation (see Section 6.13.5.1 [Prompt Primitives], page 296).

6.13.8.2 Catching Exceptions

catch is used to set up a target for a possible non-local jump. The arguments of a catch expression are a key, which restricts the set of exceptions to which this catch applies, a thunk that specifies the code to execute and one or two handler procedures that say what to do if an exception is thrown while executing the code. If the execution thunk executes normally, which means without throwing any exceptions, the handler procedures are not called at all.

When an exception is thrown using the throw function, the first argument of the throw is a symbol that indicates the type of the exception. For example, Guile throws an exception using the symbol numerical-overflow to indicate numerical overflow errors such as division by zero:

     (/ 1 0)

⇒

     ABORT: (numerical-overflow)

The key argument in a catch expression corresponds to this symbol. key may be a specific symbol, such as numerical-overflow, in which case the catch applies specifically to exceptions of that type; or it may be #t, which means that the catch applies to all exceptions, irrespective of their type.

The second argument of a catch expression should be a thunk (i.e. a procedure that accepts no arguments) that specifies the normal case code. The catch is active for the execution of this thunk, including any code called directly or indirectly by the thunk’s body. Evaluation of the catch expression activates the catch and then calls this thunk.

The third argument of a catch expression is a handler procedure. If an exception is thrown, this procedure is called with exactly the arguments specified by the throw. Therefore, the handler procedure must be designed to accept a number of arguments that corresponds to the number of arguments in all throw expressions that can be caught by

this catch.

The fourth, optional argument of a catch expression is another handler procedure, called the pre-unwind handler. It differs from the third argument in that if an exception is thrown, it is called, before the third argument handler, in exactly the dynamic context of the throw expression that threw the exception. This means that it is useful for capturing or displaying the stack at the point of the throw, or for examining other aspects of the dynamic context, such as fluid values, before the context is unwound back to that of the prevailing catch.

[Scheme Procedure] catch key thunk handler [pre-unwind-handler]

[C Function] scm_catch_with_pre_unwind_handler (key, thunk, handler,

pre unwind handler)

[C Function] scm_catch (key, thunk, handler)

Invoke thunk in the dynamic context of handler for exceptions matching key. If thunk throws to the symbol key, then handler is invoked this way:

     (handler key args ...)

key is a symbol or #t. thunk takes no arguments. If thunk returns normally, that is the return value of

catch.

Handler is invoked outside the scope of its own catch. If handler again throws to the same key, a new handler from further up the call chain is invoked. If the key is #t, then a throw to any symbol will match this call to catch. If a pre-unwind-handler is given and thunk throws an exception that matches key, Guile calls the pre-unwind-handler before unwinding the dynamic state and invoking the main handler. pre-unwind-handler should be a procedure with the same signature as handler, that is (lambda (key . args)). It is typically used to save the stack at the point where the exception occurred, but can also query other parts of the dynamic state at that point, such as fluid values. A pre-unwind-handler can exit either normally or non-locally. If it exits normally, Guile unwinds the stack and dynamic context and then calls the normal (third argument) handler. If it exits non-locally, that exit determines the continuation.

If a handler procedure needs to match a variety of throw expressions with varying numbers of arguments, you should write it like this:

(lambda (key . args)
  ...)

The key argument is guaranteed always to be present, because a throw without a key is not valid. The number and interpretation of the args varies from one type of exception to another, but should be specified by the documentation for each exception type.

Note that, once the normal (post-unwind) handler procedure is invoked, the catch that led to the handler procedure being called is no longer active. Therefore, if the handler procedure itself throws an exception, that exception can only be caught by another active catch higher up the call stack, if there is one.

[C Function] SCM scm_c_catch (SCM tag, scm t catch body body, void *body data, scm t catch handler handler, void *handler data, scm t catch handler pre unwind handler, void *pre unwind handler data)

[C Function] SCM scm_internal_catch (SCM tag, scm t catch body body, void

*body data, scm t catch handler handler, void *handler data)

The above scm_catch_with_pre_unwind_handler and scm_catch take Scheme pro-

cedures as body and handler arguments. scm_c_catch and scm_internal_catch are equivalents taking C functions. body is called as body (body_data) with a catch on exceptions of the given tag type. If an exception is caught, pre unwind handler and handler are called as handler (handler_data, key, args). key and args are the SCM key and argument list from

the throw.

body and handler should have the following prototypes. scm_t_catch_body and scm_t_catch_handler are pointer typedefs for these.

     SCM body (void *data);
     SCM handler (void *data, SCM key, SCM args);

The body data and handler data parameters are passed to the respective calls so an application can communicate extra information to those functions. If the data consists of an SCM object, care should be taken that it isn’t garbage collected while still required. If the SCM is a local C variable, one way to protect it is to pass a pointer to that variable as the data parameter, since the C compiler will then know the value must be held on the stack. Another way is to use scm_remember_upto_here_1 (see Section 5.5.5 [Remembering During Operations], page 77).

6.13.8.3 Throw Handlers

It’s sometimes useful to be able to intercept an exception that is being thrown before the stack is unwound. This could be to clean up some related state, to print a backtrace, or to pass information about the exception to a debugger, for example. The with-throw-handler procedure provides a way to do this.

[Scheme Procedure] with-throw-handler key thunk handler [C Function] scm_with_throw_handler (key, thunk, handler)

Add handler to the dynamic context as a throw handler for key key, then invoke thunk. This behaves exactly like catch, except that it does not unwind the stack before invoking handler. If the handler procedure returns normally, Guile rethrows the same exception again to the next innermost catch or throw handler. handler may exit nonlocally, of course, via an explicit throw or via invoking a continuation.

Typically handler is used to display a backtrace of the stack at the point where the corresponding throw occurred, or to save off this information for possible display later.

Not unwinding the stack means that throwing an exception that is handled via a throw handler is equivalent to calling the throw handler handler inline instead of each throw, and then omitting the surrounding with-throw-handler. In other words,

(with-throw-handler 'key
  (lambda () ... (throw 'key args ...) ...)
  handler)

is mostly equivalent to

((lambda () ... (handler 'key args ...) ...))

In particular, the dynamic context when handler is invoked is that of the site where throw is called. The examples are not quite equivalent, because the body of a with-throw-handler is not in tail position with respect to the with-throw-handler, and if handler exits normally, Guile arranges to rethrow the error, but hopefully the intention is clear.

(For an introduction to what is meant by dynamic context, See Section 6.13.10 [Dynamic Wind], page 309.)

[C Function] SCM scm_c_with_throw_handler (SCM tag, scm t catch body body,

void *body data, scm t catch handler handler, void *handler data, int lazy catch p)

The above scm_with_throw_handler takes Scheme procedures as body (thunk) and handler arguments. scm_c_with_throw_handler is an equivalent taking C functions. See scm_c_catch (see Section 6.13.8.2 [Catch], page 304) for a description of the parameters, the behaviour however of course follows with-throw-handler.

If thunk throws an exception, Guile handles that exception by invoking the innermost catch or throw handler whose key matches that of the exception. When the innermost thing is a throw handler, Guile calls the specified handler procedure using (apply handler key args). The handler procedure may either return normally or exit non-locally. If it returns normally, Guile passes the exception on to the next innermost catch or throw handler. If it exits non-locally, that exit determines the continuation.

The behaviour of a throw handler is very similar to that of a catch expression’s optional pre-unwind handler. In particular, a throw handler’s handler procedure is invoked in the exact dynamic context of the throw expression, just as a pre-unwind handler is. withthrow-handler may be seen as a half-catch: it does everything that a catch would do until the point where catch would start unwinding the stack and dynamic context, but then it rethrows to the next innermost catch or throw handler instead.

Note also that since the dynamic context is not unwound, if a with-throw-handler handler throws to a key that does not match the with-throw-handler expression’s key, the new throw may be handled by a catch or throw handler that is closer to the throw

than the first with-throw-handler.

Here is an example to illustrate this behavior:

     (catch 'a
       (lambda ()
         (with-throw-handler 'b
           (lambda ()
             (catch 'a
               (lambda ()
                 (throw 'b))
               inner-handler))
           (lambda (key . args)
             (throw 'a))))
       outer-handler)

This code will call inner-handler and then continue with the continuation of the inner

catch.

6.13.8.4 Throwing Exceptions

The throw primitive is used to throw an exception. One argument, the key, is mandatory, and must be a symbol; it indicates the type of exception that is being thrown. Following the key, throw accepts any number of additional arguments, whose meaning depends on the exception type. The documentation for each possible type of exception should specify the additional arguments that are expected for that kind of exception.

[Scheme Procedure] throw key arg … [C Function] scm_throw (key, args)

Invoke the catch form matching key, passing arg … to the handler. key is a symbol. It will match catches of the same symbol or of #t. If there is no handler at all, Guile prints an error and then exits.

When an exception is thrown, it will be caught by the innermost catch or throw handler that applies to the type of the thrown exception; in other words, whose key is either #t or the same symbol as that used in the throw expression. Once Guile has identified the appropriate catch or throw handler, it handles the exception by applying the relevant handler procedure(s) to the arguments of the throw.

If there is no appropriate catch or throw handler for a thrown exception, Guile prints an error to the current error port indicating an uncaught exception, and then exits. In practice, it is quite difficult to observe this behaviour, because Guile when used interactively installs a top level catch handler that will catch all exceptions and print an appropriate error message without exiting. For example, this is what happens if you try to throw an unhandled exception in the standard Guile REPL; note that Guile’s command loop continues after the error message:

     guile> (throw 'badex)
     <unnamed port>:3:1: In procedure gsubr-apply ...
     <unnamed port>:3:1: unhandled-exception: badex
     ABORT: (misc-error)
     guile>

The default uncaught exception behaviour can be observed by evaluating a throw expression from the shell command line:

     $ guile -c "(begin (throw 'badex) (display \"here\\n\"))"
     guile: uncaught throw to badex: ()
     $

That Guile exits immediately following the uncaught exception is shown by the absence of any output from the display expression, because Guile never gets to the point of evaluating that expression.

6.13.8.5 How Guile Implements Exceptions

It is traditional in Scheme to implement exception systems using call-with-current-continuation. Continuations (see Section 6.13.6 [Continuations], page 300) are such a powerful concept that any other control mechanism — including catch and throw — can be implemented in terms of them.

Guile does not implement catch and throw like this, though. Why not? Because Guile is specifically designed to be easy to integrate with applications written in C. In a mixed Scheme/C environment, the concept of continuation must logically include “what happens next” in the C parts of the application as well as the Scheme parts, and it turns out that the only reasonable way of implementing continuations like this is to save and restore the complete C stack.

So Guile’s implementation of call-with-current-continuation is a stack copying one. This allows it to interact well with ordinary C code, but means that creating and calling a continuation is slowed down by the time that it takes to copy the C stack.

The more targeted mechanism provided by catch and throw does not need to save and restore the C stack because the throw always jumps to a location higher up the stack of the code that executes the throw. Therefore Guile implements the catch and throw primitives independently of call-with-current-continuation, in a way that takes advantage of this upwards only nature of exceptions.

6.13.9 Procedures for Signaling Errors

Guile provides a set of convenience procedures for signaling error conditions that are implemented on top of the exception primitives just described.

[Scheme Procedure] error msg arg … Raise an error with key misc-error and a message constructed by displaying msg and writing arg . . ..

[Scheme Procedure] scm-error key subr message args data [C Function] scm_error_scm (key, subr, message, args, data)

Raise an error with key key. subr can be a string naming the procedure associated with the error, or #f. message is the error message string, possibly containing ~S and ~A escapes. When an error is reported, these are replaced by formatting the corresponding members of args: ~A (was %s in older versions of Guile) formats using display and ~S (was %S) formats using write. data is a list or #f depending on key: if key is system-error then it should be a list containing the Unix errno value; If key is signal then it should be a list containing the Unix signal number; If key is

     out-of-range, wrong-type-arg, or keyword-argument-error, it is a list containing

the bad value; otherwise it will usually be #f.

[Scheme Procedure] strerror err [C Function] scm_strerror (err)

Return the Unix error message corresponding to err, an integer errno value.

When setlocale has been called (see Section 7.2.13 [Locales], page 518), the message is in the language and charset of LC_MESSAGES. (This is done by the C library.)

[syntax] false-if-exception expr

Returns the result of evaluating its argument; however if an exception occurs then #f is returned instead.

6.13.10 Dynamic Wind

For Scheme code, the fundamental procedure to react to non-local entry and exits of dynamic

contexts is dynamic-wind. C code could use scm_internal_dynamic_wind, but since C

does not allow the convenient construction of anonymous procedures that close over lexical variables, this will be, well, inconvenient.

Therefore, Guile offers the functions scm_dynwind_begin and scm_dynwind_end to delimit a dynamic extent. Within this dynamic extent, which is called a dynwind context, you can perform various dynwind actions that control what happens when the dynwind context is entered or left. For example, you can register a cleanup routine with scm_dynwind_ unwind_handler that is executed when the context is left. There are several other more specialized dynwind actions as well, for example to temporarily block the execution of asyncs or to temporarily change the current output port. They are described elsewhere in this manual.

Here is an example that shows how to prevent memory leaks.

/* Suppose there is a function called FOO in some library that you
 would like to make available to Scheme code (or to C code that
 follows the Scheme conventions).
 FOO takes two C strings and returns a new string.  When an error has
 occurred in FOO, it returns NULL.
*/
char *foo (char *s1, char *s2);
/* SCM_FOO interfaces the C function FOO to the Scheme way of life.
 It takes care to free up all temporary strings in the case of
 non-local exits.
*/
SCM
scm_foo (SCM s1, SCM s2)
{
char *c_s1, *c_s2, *c_res;
scm_dynwind_begin (0);
c_s1 = scm_to_locale_string (s1);
/* Call 'free (c_s1)' when the dynwind context is left.
*/
scm_dynwind_unwind_handler (free, c_s1, SCM_F_WIND_EXPLICITLY);
c_s2 = scm_to_locale_string (s2);
/* Same as above, but more concisely.
*/
scm_dynwind_free (c_s2);
c_res = foo (c_s1, c_s2);
if (c_res == NULL)
  scm_memory_error ("foo");
scm_dynwind_end ();

return scm_take_locale_string (res);
}

[Scheme Procedure] dynamic-wind in guard thunk out guard [C Function] scm_dynamic_wind (in guard, thunk, out guard)

All three arguments must be 0-argument procedures. in guard is called, then thunk, then out guard. If, any time during the execution of thunk, the dynamic extent of the dynamic-wind expression is escaped non-locally, out guard is called. If the dynamic extent of the dynamic-wind is re-entered, in guard is called. Thus in guard and out guard may be called any number of times.

(define x 'normal-binding)
⇒ x
(define a-cont
  (call-with-current-continuation
   (lambda (escape)
     (let ((old-x x))
       (dynamic-wind
           ;; in-guard:
           ;;
           (lambda () (set! x 'special-binding))
           ;; thunk
           ;;
           (lambda () (display x) (newline)
                      (call-with-current-continuation escape)
                      (display x) (newline)
                      x)
           ;; out-guard:
           ;;
           (lambda () (set! x old-x)))))))
;; Prints:
special-binding
;; Evaluates to:
⇒ a-cont
x
⇒ normal-binding
(a-cont #f)
;; Prints:
special-binding
;; Evaluates to:
⇒ a-cont  ;; the value of the (define a-cont...)
x
⇒ normal-binding
a-cont

     ⇒ special-binding
[C Type] scm_t_dynwind_flags

This is an enumeration of several flags that modify the behavior of scm_dynwind_ begin. The flags are listed in the following table.

SCM_F_DYNWIND_REWINDABLE

The dynamic context is rewindable. This means that it can be reentered non-locally (via the invocation of a continuation). The default is that a dynwind context can not be reentered non-locally.

[C Function] void scm_dynwind_begin (scm t dynwind flags flags)

The function scm_dynwind_begin starts a new dynamic context and makes it the ‘current’ one. The flags argument determines the default behavior of the context. Normally, use 0. This will result in a context that can not be reentered with a captured continuation. When you are prepared to handle reentries, include SCM_F_DYNWIND_REWINDABLE in flags. Being prepared for reentry means that the effects of unwind handlers can be undone on reentry. In the example above, we want to prevent a memory leak on non-local exit and thus register an unwind handler that frees the memory. But once the memory is freed, we can not get it back on reentry. Thus reentry can not be allowed. The consequence is that continuations become less useful when non-reentrant contexts are captured, but you don’t need to worry about that too much. The context is ended either implicitly when a non-local exit happens, or explicitly with scm_dynwind_end. You must make sure that a dynwind context is indeed ended properly. If you fail to call scm_dynwind_end for each scm_dynwind_begin, the behavior is undefined.

[C Function] void scm_dynwind_end ()

End the current dynamic context explicitly and make the previous one current.

[C Type] scm_t_wind_flags

This is an enumeration of several flags that modify the behavior of scm_dynwind_

unwind_handler and scm_dynwind_rewind_handler. The flags are listed in the fol-

lowing table.

SCM_F_WIND_EXPLICITLY

The registered action is also carried out when the dynwind context is entered or left locally.

[C Function] void scm_dynwind_unwind_handler (void (*func)(void *), void

*data, scm t wind flags flags)

[C Function] void scm_dynwind_unwind_handler_with_scm (void (*func)(SCM),

SCM data, scm t wind flags flags) Arranges for func to be called with data as its arguments when the current context ends implicitly. If flags contains SCM_F_WIND_EXPLICITLY, func is also called when the context ends explicitly with scm_dynwind_end.

The function scm_dynwind_unwind_handler_with_scm takes care that data is pro-

tected from garbage collection.

[C Function] void scm_dynwind_rewind_handler (void (*func)(void *), void

*data, scm t wind flags flags)

[C Function] void scm_dynwind_rewind_handler_with_scm (void (*func)(SCM),

SCM data, scm t wind flags flags) Arrange for func to be called with data as its argument when the current context is restarted by rewinding the stack. When flags contains SCM_F_WIND_EXPLICITLY, func is called immediately as well.

The function scm_dynwind_rewind_handler_with_scm takes care that data is pro-

tected from garbage collection.

[C Function] void scm_dynwind_free (void *mem)

Arrange for mem to be freed automatically whenever the current context is exited, whether normally or non-locally. scm_dynwind_free (mem) is an equivalent short-

hand for scm_dynwind_unwind_handler (free, mem, SCM_F_WIND_EXPLICITLY).

6.13.11 How to Handle Errors

Error handling is based on catch and throw. Errors are always thrown with a key and four arguments:

• key: a symbol which indicates the type of error. The symbols used by libguile are listed below.

• subr: the name of the procedure from which the error is thrown, or #f.

• message: a string (possibly language and system dependent) describing the error. The tokens ~A and ~S can be embedded within the message: they will be replaced with members of the args list when the message is printed. ~A indicates an argument printed using display, while ~S indicates an argument printed using write. message can also be #f, to allow it to be derived from the key by the error handler (may be useful if the key is to be thrown from both C and Scheme).

• args: a list of arguments to be used to expand ~A and ~S tokens in message. Can also be #f if no arguments are required.

• rest: a list of any additional objects required. e.g., when the key is 'system-error, this contains the C errno value. Can also be #f if no additional objects are required.

In addition to catch and throw, the following Scheme facilities are available:

[Scheme Procedure] display-error frame port subr message args rest [C Function] scm_display_error (frame, port, subr, message, args, rest)

Display an error message to the output port port. frame is the frame in which the error occurred, subr is the name of the procedure in which the error occurred and message is the actual error message, which may contain formatting instructions. These will format the arguments in the list args accordingly. rest is currently ignored.

The following are the error keys defined by libguile and the situations in which they are used:

• error-signal: thrown after receiving an unhandled fatal signal such as SIGSEGV, SIGBUS, SIGFPE etc. The rest argument in the throw contains the coded signal number (at present this is not the same as the usual Unix signal number).

• system-error: thrown after the operating system indicates an error condition. The rest argument in the throw contains the errno value.

• numerical-overflow: numerical overflow.

• out-of-range: the arguments to a procedure do not fall within the accepted domain.

• wrong-type-arg: an argument to a procedure has the wrong type.

• wrong-number-of-args: a procedure was called with the wrong number of arguments.

• memory-allocation-error: memory allocation error.

• stack-overflow: stack overflow error.

• regular-expression-syntax: errors generated by the regular expression library.

• misc-error: other errors.

6.13.11.1 C Support

In the following C functions, SUBR and MESSAGE parameters can be NULL to give the effect of #f described above.

[C Function] SCM scm_error (SCM key, char *subr, char *message, SCM args,

   SCM rest)

Throw an error, as per scm-error (see Section 6.13.9 [Error Reporting], page 309).

[C Function] void scm_syserror (char *subr)
[C Function] void scm_syserror_msg (char *subr, char *message, SCM args)

Throw an error with key system-error and supply errno in the rest argument. For scm_syserror the message is generated using strerror. Care should be taken that any code in between the failing operation and the call to these routines doesn’t change errno.

[C Function] void scm_num_overflow (char *subr)
[C Function] void scm_out_of_range (char *subr, SCM bad_value)
[C Function] void scm_wrong_num_args (SCM proc)
[C Function] void scm_wrong_type_arg (char *subr, int argnum, SCM
   bad_value)
[C Function] void scm_wrong_type_arg_msg (char *subr, int argnum, SCM
   bad_value, const char *expected)
[C Function] void scm_memory_error (char *subr)

[C Function] void scm_misc_error (const char *subr, const char *message, SCM

   args)

Throw an error with the various keys described above. In scm_wrong_num_args, proc should be a Scheme symbol which is the name of the procedure incorrectly invoked. The other routines take the name of the invoked procedure as a C string. In scm_wrong_type_arg_msg, expected is a C string describing the type of argument that was expected. In scm_misc_error, message is the error message string, possibly containing simpleformat escapes (see Section 6.14.3 [Writing], page 320), and the corresponding arguments in the args list.

6.13.11.2 Signalling Type Errors

Every function visible at the Scheme level should aggressively check the types of its arguments, to avoid misinterpreting a value, and perhaps causing a segmentation fault. Guile provides some macros to make this easier.

[Macro] void SCM_ASSERT (int test, SCM obj, unsigned int position, const char

   *subr)
[Macro] void SCM_ASSERT_TYPE (int test, SCM obj, unsigned int position,

const char *subr, const char *expected) If test is zero, signal a “wrong type argument” error, attributed to the subroutine named subr, operating on the value obj, which is the position’th argument of subr.

In SCM_ASSERT_TYPE, expected is a C string describing the type of argument that was expected.

[Macro] int SCM_ARG1
[Macro] int SCM_ARG2
[Macro] int SCM_ARG3
[Macro] int SCM_ARG4
[Macro] int SCM_ARG5
[Macro] int SCM_ARG6
[Macro] int SCM_ARG7

One of the above values can be used for position to indicate the number of the argument of subr which is being checked. Alternatively, a positive integer number can be used, which allows to check arguments after the seventh. However, for parameter numbers up to seven it is preferable to use SCM_ARGN instead of the corresponding raw number, since it will make the code easier to understand.

[Macro] int SCM_ARGn

Passing a value of zero or SCM_ARGn for position allows to leave it unspecified which argument’s type is incorrect. Again, SCM_ARGn should be preferred over a raw zero constant.

6.13.12 Continuation Barriers

The non-local flow of control caused by continuations might sometimes not be wanted. You can use with-continuation-barrier to erect fences that continuations can not pass.

[Scheme Procedure] with-continuation-barrier proc
[C Function] scm_with_continuation_barrier (proc)

Call proc and return its result. Do not allow the invocation of continuations that would leave or enter the dynamic extent of the call to with-continuation-barrier. Such an attempt causes an error to be signaled.

Throws (such as errors) that are not caught from within proc are caught by withcontinuation-barrier. In that case, a short message is printed to the current error port and #f is returned.

Thus, with-continuation-barrier returns exactly once.

[C Function] void * scm_c_with_continuation_barrier (void *(*func) (void

*), void *data) Like scm_with_continuation_barrier but call func on data. When an error is caught, NULL is returned.

6.14 Input and Output

6.14.1 Ports

Sequential input/output in Scheme is represented by operations on a port. This chapter explains the operations that Guile provides for working with ports.

Ports are created by opening, for instance open-file for a file (see Section 6.14.9.1 [File Ports], page 326). Characters can be read from an input port and written to an output port, or both on an input/output port. A port can be closed (see Section 6.14.4 [Closing], page 321) when no longer required, after which any attempt to read or write is an error.

The formal definition of a port is very generic: an input port is simply “an object which can deliver characters on demand,” and an output port is “an object which can accept characters.” Because this definition is so loose, it is easy to write functions that simulate ports in software. Soft ports and string ports are two interesting and powerful examples of this technique. (see Section 6.14.9.3 [Soft Ports], page 331, and Section 6.14.9.2 [String Ports], page 329.)

Ports are garbage collected in the usual way (see Section 6.18 [Memory Management], page 374), and will be closed at that time if not already closed. In this case any errors occurring in the close will not be reported. Usually a program will want to explicitly close so as to be sure all its operations have been successful. Of course if a program has abandoned something due to an error or other condition then closing problems are probably not of interest.

It is strongly recommended that file ports be closed explicitly when no longer required. Most systems have limits on how many files can be open, both on a per-process and a system-wide basis. A program that uses many files should take care not to hit those limits. The same applies to similar system resources such as pipes and sockets.

Note that automatic garbage collection is triggered only by memory consumption, not by file or other resource usage, so a program cannot rely on that to keep it away from system limits. An explicit call to gc can of course be relied on to pick up unreferenced ports. If program flow makes it hard to be certain when to close then this may be an acceptable way to control resource usage.

All file access uses the “LFS” large file support functions when available, so files bigger than 2 Gbytes (2^{3}1 bytes) can be read and written on a 32-bit system.

Each port has an associated character encoding that controls how bytes read from the port are converted to characters and string and controls how characters and strings written to the port are converted to bytes. When ports are created, they inherit their character encoding from the current locale, but, that can be modified after the port is created.

Currently, the ports only work with non-modal encodings. Most encodings are nonmodal, meaning that the conversion of bytes to a string doesn’t depend on its context: the same byte sequence will always return the same string. A couple of modal encodings are in common use, like ISO-2022-JP and ISO-2022-KR, and they are not yet supported.

Each port also has an associated conversion strategy: what to do when a Guile character can’t be converted to the port’s encoded character representation for output. There are three possible strategies: to raise an error, to replace the character with a hex escape, or to replace the character with a substitute character.

[Scheme Procedure] input-port? x

[C Function] scm_input_port_p (x)

Return #t if x is an input port, otherwise return #f. Any object satisfying this predicate also satisfies port?.

[Scheme Procedure] output-port? x

[C Function] scm_output_port_p (x)

Return #t if x is an output port, otherwise return #f. Any object satisfying this predicate also satisfies port?.

[Scheme Procedure] port? x [C Function] scm_port_p (x)

Return a boolean indicating whether x is a port. Equivalent to (or (input-port?

x) (output-port? x)).

[Scheme Procedure] set-port-encoding! port enc

[C Function] scm_set_port_encoding_x (port, enc)

Sets the character encoding that will be used to interpret all port I/O. enc is a string containing the name of an encoding. Valid encoding names are those defined by IANA

(http://www.iana.org/assignments/character-sets).
[Scheme Variable] %default-port-encoding

A fluid containing #f or the name of the encoding to be used by default for newly created ports (see Section 6.21.7 [Fluids and Dynamic States], page 419). The value #f is equivalent to "ISO-8859-1". New ports are created with the encoding appropriate for the current locale if setlocale has been called or the value specified by this fluid otherwise.

[Scheme Procedure] port-encoding port

[C Function] scm_port_encoding (port)

Returns, as a string, the character encoding that port uses to interpret its input and output. The value #f is equivalent to "ISO-8859-1".

[Scheme Procedure] set-port-conversion-strategy! port sym
[C Function] scm_set_port_conversion_strategy_x (port, sym)

Sets the behavior of the interpreter when outputting a character that is not representable in the port’s current encoding. sym can be either 'error, 'substitute, or 'escape. If it is 'error, an error will be thrown when an nonconvertible character is encountered. If it is 'substitute, then nonconvertible characters will be replaced with approximate characters, or with question marks if no approximately correct character is available. If it is 'escape, it will appear as a hex escape when output. If port is an open port, the conversion error behavior is set for that port. If it is #f, it is set as the default behavior for any future ports that get created in this thread.

[Scheme Procedure] port-conversion-strategy port
[C Function] scm_port_conversion_strategy (port)

Returns the behavior of the port when outputting a character that is not representable in the port’s current encoding. It returns the symbol error if unrepresentable characters should cause exceptions, substitute if the port should try to replace unrepresentable characters with question marks or approximate characters, or escape if unrepresentable characters should be converted to string escapes. If port is #f, then the current default behavior will be returned. New ports will have this default behavior when they are created.

[Scheme Variable] %default-port-conversion-strategy

The fluid that defines the conversion strategy for newly created ports, and for other

conversion routines such as scm_to_stringn, scm_from_stringn, string->pointer,
and pointer->string.

Its value must be one of the symbols described above, with the same semantics:

'error, 'substitute, or 'escape.

When Guile starts, its value is 'substitute.

Note that (set-port-conversion-strategy! #f sym) is equivalent to (fluid-set!
%default-port-conversion-strategy sym).

6.14.2 Reading

[Generic procedures for reading from ports.]

These procedures pertain to reading characters and strings from ports. To read general S-expressions from ports, See Section 6.17.2 [Scheme Read], page 360.

[Scheme Procedure] eof-object? x

[C Function] scm_eof_object_p (x)

Return #t if x is an end-of-file object; otherwise return #f.

[Scheme Procedure] char-ready? [port] [C Function] scm_char_ready_p (port)

Return #t if a character is ready on input port and return #f otherwise. If charready? returns #t then the next read-char operation on port is guaranteed not to hang. If port is a file port at end of file then char-ready? returns #t. char-ready? exists to make it possible for a program to accept characters from interactive ports without getting stuck waiting for input. Any input editors associated with such ports must make sure that characters whose existence has been asserted by char-ready? cannot be rubbed out. If char-ready? were to return #f at end of file, a port at end of file would be indistinguishable from an interactive port that has no ready characters.

[Scheme Procedure] read-char [port] [C Function] scm_read_char (port)

Return the next character available from port, updating port to point to the following character. If no more characters are available, the end-of-file object is returned. When port’s data cannot be decoded according to its character encoding, a decodingerror is raised and port points past the erroneous byte sequence.

[C Function] size_t scm_c_read (SCM port, void *buffer, size t size)

Read up to size bytes from port and store them in buffer. The return value is the number of bytes actually read, which can be less than size if end-of-file has been reached. Note that this function does not update port-line and port-column below.

[Scheme Procedure] peek-char [port] [C Function] scm_peek_char (port)

Return the next character available from port, without updating port to point to the following character. If no more characters are available, the end-of-file object is returned. The value returned by a call to peek-char is the same as the value that would have been returned by a call to read-char on the same port. The only difference is that the very next call to read-char or peek-char on that port will return the value returned by the preceding call to peek-char. In particular, a call to peek-char on an interactive port will hang waiting for input whenever a call to read-char would have hung. As for read-char, a decoding-error may be raised if such a situation occurs. However, unlike with read-char, port still points at the beginning of the erroneous byte sequence when the error is raised.

[Scheme Procedure] unread-char cobj [port] [C Function] scm_unread_char (cobj, port)

Place character cobj in port so that it will be read by the next read operation. If called multiple times, the unread characters will be read again in last-in first-out order. If port is not supplied, the current input port is used.

[Scheme Procedure] unread-string str port [C Function] scm_unread_string (str, port)

Place the string str in port so that its characters will be read from left-to-right as the next characters from port during subsequent read operations. If called multiple times, the unread characters will be read again in last-in first-out order. If port is not supplied, the current-input-port is used.

[Scheme Procedure] drain-input port [C Function] scm_drain_input (port)

This procedure clears a port’s input buffers, similar to the way that force-output clears the output buffer. The contents of the buffers are returned as a single string, e.g.,

     (define p (open-input-file ...))
     (drain-input p) => empty string, nothing buffered yet.
     (unread-char (read-char p) p)
     (drain-input p) => initial chars from p, up to the buffer size.

Draining the buffers may be useful for cleanly finishing buffered I/O so that the file descriptor can be used directly for further input.

[Scheme Procedure] port-column port [Scheme Procedure] port-line port [C Function] scm_port_column (port) [C Function] scm_port_line (port)

Return the current column number or line number of port. If the number is unknown, the result is #f. Otherwise, the result is a 0-origin integer - i.e. the first character of the first line is line 0, column 0. (However, when you display a file position, for example in an error message, we recommend you add 1 to get 1-origin integers. This is because lines and column numbers traditionally start with 1, and that is what non-programmers will find most natural.)

[Scheme Procedure] set-port-column! port column [Scheme Procedure] set-port-line! port line [C Function] scm_set_port_column_x (port, column) [C Function] scm_set_port_line_x (port, line)

Set the current column or line number of port.

6.14.3 Writing

[Generic procedures for writing to ports.]

These procedures are for writing characters and strings to ports. For more information on writing arbitrary Scheme objects to ports, See Section 6.17.3 [Scheme Write], page 361.

[Scheme Procedure] get-print-state port

[C Function] scm_get_print_state (port)

Return the print state of the port port. If port has no associated print state, #f is returned.

[Scheme Procedure] newline [port] [C Function] scm_newline (port)

Send a newline to port. If port is omitted, send to the current output port.

[Scheme Procedure] port-with-print-state port [pstate]

[C Function] scm_port_with_print_state (port, pstate)

Create a new port which behaves like port, but with an included print state pstate. pstate is optional. If pstate isn’t supplied and port already has a print state, the old print state is reused.

[Scheme Procedure] simple-format destination message . args [C Function] scm_simple_format (destination, message, args)

Write message to destination, defaulting to the current output port. message can contain ~A (was %s) and ~S (was %S) escapes. When printed, the escapes are replaced with corresponding members of args: ~A formats using display and ~S formats using write. If destination is #t, then use the current output port, if destination is #f, then return a string containing the formatted text. Does not add a trailing newline.

[Scheme Procedure] write-char chr [port] [C Function] scm_write_char (chr, port)

Send character chr to port.

[C Function] void scm_c_write (SCM port, const void *buffer, size t size)

Write size bytes at buffer to port.

Note that this function does not update port-line and port-column (see Section 6.14.2 [Reading], page 318).

[Scheme Procedure] force-output [port] [C Function] scm_force_output (port)

Flush the specified output port, or the current output port if port is omitted. The current output buffer contents are passed to the underlying port implementation (e.g., in the case of fports, the data will be written to the file and the output buffer will be cleared.) It has no effect on an unbuffered port. The return value is unspecified.

[Scheme Procedure] flush-all-ports

[C Function] scm_flush_all_ports ()

Equivalent to calling force-output on all open output ports. The return value is unspecified.

6.14.4 Closing

[Scheme Procedure] close-port port [C Function] scm_close_port (port)

Close the specified port object. Return #t if it successfully closes a port or #f if it was already closed. An exception may be raised if an error occurs, for example when flushing buffered output. See also Section 7.2.2 [Ports and File Descriptors], page 469, for a procedure which can close file descriptors.

[Scheme Procedure] close-input-port port [Scheme Procedure] close-output-port port

[C Function] scm_close_input_port (port)
[C Function] scm_close_output_port (port)

Close the specified input or output port. An exception may be raised if an error occurs while closing. If port is already closed, nothing is done. The return value is unspecified. See also Section 7.2.2 [Ports and File Descriptors], page 469, for a procedure which can close file descriptors.

[Scheme Procedure] port-closed? port

[C Function] scm_port_closed_p (port)

Return #t if port is closed or #f if it is open.

6.14.5 Random Access

[Scheme Procedure] seek fd port offset whence [C Function] scm_seek (fd port, offset, whence)

Sets the current position of fd port to the integer offset, which is interpreted according to the value of whence. One of the following variables should be supplied for whence:

[Variable] SEEK_SET

Seek from the beginning of the file.

[Variable] SEEK_CUR

Seek from the current position.

[Variable] SEEK_END

Seek from the end of the file.

If fd port is a file descriptor, the underlying system call is lseek. port may be a string port. The value returned is the new position in the file. This means that the current position of a port can be obtained using:

     (seek port 0 SEEK_CUR)

[Scheme Procedure] ftell fd port

[C Function] scm_ftell (fd port)

Return an integer representing the current position of fd port, measured from the beginning. Equivalent to:

     (seek port 0 SEEK_CUR)

[Scheme Procedure] truncate-file file [length] [C Function] scm_truncate_file (file, length)

Truncate file to length bytes. file can be a filename string, a port object, or an integer file descriptor. The return value is unspecified. For a port or file descriptor length can be omitted, in which case the file is truncated at the current position (per ftell above). On most systems a file can be extended by giving a length greater than the current size, but this is not mandatory in the POSIX standard.

6.14.6 Line Oriented and Delimited Text

The delimited-I/O module can be accessed with:

(use-modules (ice-9 rdelim))

It can be used to read or write lines of text, or read text delimited by a specified set of characters. It’s similar to the (scsh rdelim) module from guile-scsh, but does not use multiple values or character sets and has an extra procedure write-line.

[Scheme Procedure] read-line [port] [handle-delim] Return a line of text from port if specified, otherwise from the value returned by (current-input-port). Under Unix, a line of text is terminated by the first end-of-line character or by end-of-file. If handle-delim is specified, it should be one of the following symbols:

trim Discard the terminating delimiter. This is the default, but it will be impossible to tell whether the read terminated with a delimiter or endof-file.

concat Append the terminating delimiter (if any) to the returned string.

peek Push the terminating delimiter (if any) back on to the port. split Return a pair containing the string read from the port and the terminating

delimiter or end-of-file object.

Like read-char, this procedure can throw to decoding-error (see Section 6.14.2 [Reading], page 318).

[Scheme Procedure] read-line! buf [port]

Read a line of text into the supplied string buf and return the number of characters added to buf. If buf is filled, then #f is returned. Read from port if specified, otherwise from the value returned by (current-input-port).

[Scheme Procedure] read-delimited delims [port] [handle-delim] Read text until one of the characters in the string delims is found or end-of-file is reached. Read from port if supplied, otherwise from the value returned by (currentinput-port). handle-delim takes the same values as described for read-line.

[Scheme Procedure] read-delimited! delims buf [port] [handle-delim] [start] [end] Read text into the supplied string buf. If a delimiter was found, return the number of characters written, except if handledelim is split, in which case the return value is a pair, as noted above. As a special case, if port was already at end-of-stream, the EOF object is returned. Also, if no characters were written because the buffer was full, #f is returned. It’s something of a wacky interface, to be honest.

[Scheme Procedure] write-line obj [port] [C Function] scm_write_line (obj, port)

Display obj and a newline character to port. If port is not specified, (currentoutput-port) is used. This function is equivalent to:

     (display obj [port])
     (newline [port])

In the past, Guile did not have a procedure that would just read out all of the characters from a port. As a workaround, many people just called read-delimited with no delimiters, knowing that would produce the behavior they wanted. This prompted Guile developers to add some routines that would read all characters from a port. So it is that (ice-9 rdelim) is also the home for procedures that can reading undelimited text:

[Scheme Procedure] read-string [port] [count]

Read all of the characters out of port and return them as a string. If the count is present, treat it as a limit to the number of characters to read. By default, read from the current input port, with no size limit on the result. This procedure always returns a string, even if no characters were read.

[Scheme Procedure] read-string! buf [port] [start] [end]

Fill buf with characters read from port, defaulting to the current input port. Return the number of characters read. If start or end are specified, store data only into the substring of str bounded by start and end (which default to the beginning and end of the string, respectively).

Some of the aforementioned I/O functions rely on the following C primitives. These will mainly be of interest to people hacking Guile internals.

[Scheme Procedure] %read-delimited! delims str gobble [port [start [end]]] [C Function] scm_read_delimited_x (delims, str, gobble, port, start, end)

Read characters from port into str until one of the characters in the delims string is encountered. If gobble is true, discard the delimiter character; otherwise, leave it in the input stream for the next read. If port is not specified, use the value of (currentinput-port). If start or end are specified, store data only into the substring of str bounded by start and end (which default to the beginning and end of the string, respectively). Return a pair consisting of the delimiter that terminated the string and the number of characters read. If reading stopped at the end of file, the delimiter returned is the eof-object; if the string was filled without encountering a delimiter, this value is #f.

[Scheme Procedure] %read-line [port] [C Function] scm_read_line (port)

Read a newline-terminated line from port, allocating storage as necessary. The newline terminator (if any) is removed from the string, and a pair consisting of the line and its delimiter is returned. The delimiter may be either a newline or the eof-object; if %read-line is called at the end of file, it returns the pair (#<eof> . #<eof>).

6.14.7 Block reading and writing

The Block-string-I/O module can be accessed with:

(use-modules (ice-9 rw))

It currently contains procedures that help to implement the (scsh rw) module in guilescsh.

[Scheme Procedure] read-string!/partial str [port or fdes [start [end]]]

[C Function] scm_read_string_x_partial (str, port or fdes, start, end)

Read characters from a port or file descriptor into a string str. A port must have an underlying file descriptor — a so-called fport. This procedure is scsh-compatible and can efficiently read large strings. It will:

• attempt to fill the entire string, unless the start and/or end arguments are supplied. i.e., start defaults to 0 and end defaults to (string-length str)

• use the current input port if port or fdes is not supplied.

• return fewer than the requested number of characters in some cases, e.g., on end of file, if interrupted by a signal, or if not all the characters are immediately available.

• wait indefinitely for some input if no characters are currently available, unless the port is in non-blocking mode.

• read characters from the port’s input buffers if available, instead from the underlying file descriptor.

• return #f if end-of-file is encountered before reading any characters, otherwise return the number of characters read.

• return 0 if the port is in non-blocking mode and no characters are immediately available.

• return 0 if the request is for 0 bytes, with no end-of-file check.

[Scheme Procedure] write-string/partial str [port or fdes [start [end]]]

[C Function] scm_write_string_partial (str, port or fdes, start, end)

Write characters from a string str to a port or file descriptor. A port must have an underlying file descriptor — a so-called fport. This procedure is scsh-compatible and can efficiently write large strings. It will:

• attempt to write the entire string, unless the start and/or end arguments are supplied. i.e., start defaults to 0 and end defaults to (string-length str)

• use the current output port if port of fdes is not supplied.

• in the case of a buffered port, store the characters in the port’s output buffer, if all will fit. If they will not fit then any existing buffered characters will be flushed before attempting to write the new characters directly to the underlying file descriptor. If the port is in non-blocking mode and buffered characters can not be flushed immediately, then an EAGAIN system-error exception will be raised (Note: scsh does not support the use of non-blocking buffered ports.)

• write fewer than the requested number of characters in some cases, e.g., if interrupted by a signal or if not all of the output can be accepted immediately.

• wait indefinitely for at least one character from str to be accepted by the port, unless the port is in non-blocking mode.

• return the number of characters accepted by the port.

• return 0 if the port is in non-blocking mode and can not accept at least one character from str immediately

• return 0 immediately if the request size is 0 bytes.

6.14.8 Default Ports for Input, Output and Errors

[Scheme Procedure] current-input-port
[C Function] scm_current_input_port ()

Return the current input port. This is the default port used by many input procedures. Initially this is the standard input in Unix and C terminology. When the standard input is a tty the port is unbuffered, otherwise it’s fully buffered. Unbuffered input is good if an application runs an interactive subprocess, since any type-ahead input won’t go into Guile’s buffer and be unavailable to the subprocess. Note that Guile buffering is completely separate from the tty “line discipline”. In the usual cooked mode on a tty Guile only sees a line of input once the user presses

Return.
[Scheme Procedure] current-output-port
[C Function] scm_current_output_port ()

Return the current output port. This is the default port used by many output procedures.

Initially this is the standard output in Unix and C terminology. When the standard output is a tty this port is unbuffered, otherwise it’s fully buffered.

Unbuffered output to a tty is good for ensuring progress output or a prompt is seen. But an application which always prints whole lines could change to line buffered, or an application with a lot of output could go fully buffered and perhaps make explicit force-output calls (see Section 6.14.3 [Writing], page 320) at selected points.

[Scheme Procedure] current-error-port
[C Function] scm_current_error_port ()

Return the port to which errors and warnings should be sent.

Initially this is the standard error in Unix and C terminology. When the standard error is a tty this port is unbuffered, otherwise it’s fully buffered.

[Scheme Procedure] set-current-input-port port
[Scheme Procedure] set-current-output-port port
[Scheme Procedure] set-current-error-port port
[C Function] scm_set_current_input_port (port)
[C Function] scm_set_current_output_port (port)
[C Function] scm_set_current_error_port (port)
Change the ports returned by current-input-port, current-output-port and

current-error-port, respectively, so that they use the supplied port for input or output.

[C Function] void scm_dynwind_current_input_port (SCM port)
[C Function] void scm_dynwind_current_output_port (SCM port)
[C Function] void scm_dynwind_current_error_port (SCM port)

These functions must be used inside a pair of calls to scm_dynwind_begin and scm_ dynwind_end (see Section 6.13.10 [Dynamic Wind], page 309). During the dynwind context, the indicated port is set to port.

More precisely, the current port is swapped with a ‘backup’ value whenever the dynwind context is entered or left. The backup value is initialized with the port argument.

6.14.9 Types of Port

[Types of port; how to make them.]

6.14.9.1 File Ports

The following procedures are used to open file ports. See also Section 7.2.2 [Ports and File Descriptors], page 469, for an interface to the Unix open system call.

Most systems have limits on how many files can be open, so it’s strongly recommended that file ports be closed explicitly when no longer required (see Section 6.14.1 [Ports], page 316).

[Scheme Procedure] open-file filename mode [#:guess-encoding=#f]

[#:encoding=#f]

[C Function] scm_open_file_with_encoding (filename, mode, guess encoding, encoding)

[C Function] scm_open_file (filename, mode)

Open the file whose name is filename, and return a port representing that file. The attributes of the port are determined by the mode string. The way in which this is interpreted is similar to C stdio. The first character must be one of the following:

‘r’ Open an existing file for input.

‘w’ Open a file for output, creating it if it doesn’t already exist or removing its contents if it does.

‘a’ Open a file for output, creating it if it doesn’t already exist. All writes to the port will go to the end of the file. The "append mode" can be turned off while the port is in use see Section 7.2.2 [Ports and File Descriptors], page 469

The following additional characters can be appended:

‘+’ Open the port for both input and output. E.g., r+: open an existing file for both input and output.

‘0’ Create an "unbuffered" port. In this case input and output operations are passed directly to the underlying port implementation without additional buffering. This is likely to slow down I/O operations. The buffering mode can be changed while a port is in use see Section 7.2.2 [Ports and File Descriptors], page 469

‘l’ Add line-buffering to the port. The port output buffer will be automatically flushed whenever a newline character is written.

‘b’ Use binary mode, ensuring that each byte in the file will be read as one Scheme character. To provide this property, the file will be opened with the 8-bit character encoding "ISO-8859-1", ignoring the default port encoding. See Section 6.14.1 [Ports], page 316, for more information on port encodings. Note that while it is possible to read and write binary data as characters or strings, it is usually better to treat bytes as octets, and byte sequences as bytevectors. See Section 6.14.10.8 [R6RS Binary Input], page 339, and Section 6.14.10.11 [R6RS Binary Output], page 344, for more. This option had another historical meaning, for DOS compatibility: in the default (textual) mode, DOS reads a CR-LF sequence as one LF byte. The b flag prevents this from happening, adding O_BINARY to the underlying open call. Still, the flag is generally useful because of its port encoding ramifications.

Unless binary mode is requested, the character encoding of the new port is determined as follows: First, if guess-encoding is true, the file-encoding procedure is used to guess the encoding of the file (see Section 6.17.8 [Character Encoding of Source Files], page 370). If guess-encoding is false or if file-encoding fails, encoding is used unless it is also false. As a last resort, the default port encoding is used. See Section 6.14.1 [Ports], page 316, for more information on port encodings. It is an error to pass a non-false guess-encoding or encoding if binary mode is requested.

If a file cannot be opened with the access requested, open-file throws an exception. When the file is opened, its encoding is set to the current %default-port-encoding, unless the b flag was supplied. Sometimes it is desirable to honor Emacs-style coding declarations in files^{9}. When that is the case, the file-encoding procedure can be used as follows (see Section 6.17.8 [Character Encoding of Source Files], page 370):

     (let* ((port     (open-input-file file))
            (encoding (file-encoding port)))
       (set-port-encoding! port (or encoding (port-encoding port))))

In theory we could create read/write ports which were buffered in one direction only. However this isn’t included in the current interfaces.

[Scheme Procedure] open-input-file filename [#:guess-encoding=#f]

[#:encoding=#f] [#:binary=#f] Open filename for input. If binary is true, open the port in binary mode, otherwise use text mode. encoding and guess-encoding determine the character encoding as described above for open-file. Equivalent to

     (open-file filename
                (if binary "rb" "r")
                #:guess-encoding guess-encoding
                #:encoding encoding)

[Scheme Procedure] open-output-file filename [#:encoding=#f] [#:binary=#f] Open filename for output. If binary is true, open the port in binary mode, otherwise use text mode. encoding specifies the character encoding as described above for open-file. Equivalent to

     (open-file filename
                (if binary "wb" "w")
                #:encoding encoding)

[Scheme Procedure] call-with-input-file filename proc [#:guess-encoding=#f]

[#:encoding=#f] [#:binary=#f]

[Scheme Procedure] call-with-output-file filename proc [#:encoding=#f]

[#:binary=#f] Open filename for input or output, and call (proc port) with the resulting port. Return the value returned by proc. filename is opened as per open-input-file or open-output-file respectively, and an error is signaled if it cannot be opened. When proc returns, the port is closed. If proc does not return (e.g. if it throws an error), then the port might not be closed automatically, though it will be garbage collected in the usual way if not otherwise referenced.

[Scheme Procedure] with-input-from-file filename thunk [#:guess-encoding=#f]

[#:encoding=#f] [#:binary=#f]

[Scheme Procedure] with-output-to-file filename thunk [#:encoding=#f] [#:binary=#f] [Scheme Procedure] with-error-to-file filename thunk [#:encoding=#f]

9 Guile 2.0.0 to 2.0.7 would do this by default. This behavior was deemed inappropriate and disabled

starting from Guile 2.0.8.

[#:binary=#f] Open filename and call (thunk) with the new port setup as respectively the current-

input-port, current-output-port, or current-error-port. Return the value re-

turned by thunk. filename is opened as per open-input-file or open-output-file respectively, and an error is signaled if it cannot be opened.

When thunk returns, the port is closed and the previous setting of the respective current port is restored.

The current port setting is managed with dynamic-wind, so the previous value is restored no matter how thunk exits (eg. an exception), and if thunk is re-entered (via a captured continuation) then it’s set again to the filename port.

The port is closed when thunk returns normally, but not when exited via an exception or new continuation. This ensures it’s still ready for use if thunk is re-entered by a captured continuation. Of course the port is always garbage collected and closed in the usual way when no longer referenced anywhere.

[Scheme Procedure] port-mode port [C Function] scm_port_mode (port)

Return the port modes associated with the open port port. These will not necessarily be identical to the modes used when the port was opened, since modes such as "append" which are used only during port creation are not retained.

[Scheme Procedure] port-filename port

[C Function] scm_port_filename (port)

Return the filename associated with port, or #f if no filename is associated with the port.

port must be open, port-filename cannot be used once the port is closed.

[Scheme Procedure] set-port-filename! port filename [C Function] scm_set_port_filename_x (port, filename)

Change the filename associated with port, using the current input port if none is specified. Note that this does not change the port’s source of data, but only the value that is returned by port-filename and reported in diagnostic output.

[Scheme Procedure] file-port? obj [C Function] scm_file_port_p (obj)

Determine whether obj is a port that is related to a file.

6.14.9.2 String Ports

The following allow string ports to be opened by analogy to R4RS file port facilities:

With string ports, the port-encoding is treated differently than other types of ports. When string ports are created, they do not inherit a character encoding from the current locale. They are given a default locale that allows them to handle all valid string characters. Typically one should not modify a string port’s character encoding away from its default.

[Scheme Procedure] call-with-output-string proc
[C Function] scm_call_with_output_string (proc)

Calls the one-argument procedure proc with a newly created output port. When the function returns, the string composed of the characters written into the port is returned. proc should not close the port. Note that which characters can be written to a string port depend on the port’s encoding. The default encoding of string ports is specified by the %default-port-encoding fluid (see Section 6.14.1 [Ports], page 316). For instance, it is an error to write Greek letter alpha to an ISO-8859-1-encoded string port since this character cannot be represented with ISO-8859-1:

     (define alpha (integer->char #x03b1)) ; GREEK SMALL LETTER ALPHA
     (with-fluids ((%default-port-encoding "ISO-8859-1"))
       (call-with-output-string
         (lambda (p)
           (display alpha p))))

⇒

     Throw to key `encoding-error'

Changing the string port’s encoding to a Unicode-capable encoding such as UTF-8 solves the problem.

[Scheme Procedure] call-with-input-string string proc

[C Function] scm_call_with_input_string (string, proc)

Calls the one-argument procedure proc with a newly created input port from which string’s contents may be read. The value yielded by the proc is returned.

[Scheme Procedure] with-output-to-string thunk

Calls the zero-argument procedure thunk with the current output port set temporarily to a new string port. It returns a string composed of the characters written to the current output. See call-with-output-string above for character encoding considerations.

[Scheme Procedure] with-input-from-string string thunk

Calls the zero-argument procedure thunk with the current input port set temporarily to a string port opened on the specified string. The value yielded by thunk is returned.

[Scheme Procedure] open-input-string str

[C Function] scm_open_input_string (str)

Take a string and return an input port that delivers characters from the string. The port can be closed by close-input-port, though its storage will be reclaimed by the garbage collector if it becomes inaccessible.

[Scheme Procedure] open-output-string
[C Function] scm_open_output_string ()

Return an output port that will accumulate characters for retrieval by get-output-string. The port can be closed by the procedure close-output-port, though its storage will be reclaimed by the garbage collector if it becomes inaccessible.

[Scheme Procedure] get-output-string port

[C Function] scm_get_output_string (port)

Given an output port created by open-output-string, return a string consisting of the characters that have been output to the port so far.

get-output-string must be used before closing port, once closed the string cannot be obtained.

A string port can be used in many procedures which accept a port but which are not dependent on implementation details of fports. E.g., seeking and truncating will work on a string port, but trying to extract the file descriptor number will fail.

6.14.9.3 Soft Ports

A soft-port is a port based on a vector of procedures capable of accepting or delivering characters. It allows emulation of I/O ports.

[Scheme Procedure] make-soft-port pv modes [C Function] scm_make_soft_port (pv, modes)

Return a port capable of receiving or delivering characters as specified by the modes string (see Section 6.14.9.1 [File Ports], page 326). pv must be a vector of length 5 or 6. Its components are as follows:

0. procedure accepting one character for output

1. procedure accepting a string for output

2. thunk for flushing output

3. thunk for getting one character

4. thunk for closing port (not by garbage collection)

5. (if present and not #f) thunk for computing the number of characters that can be read from the port without blocking.

For an output-only port only elements 0, 1, 2, and 4 need be procedures. For an input-only port only elements 3 and 4 need be procedures. Thunks 2 and 4 can instead be #f if there is no useful operation for them to perform.

If thunk 3 returns #f or an eof-object (see Section “Input” in The Revised^5 Report on Scheme) it indicates that the port has reached end-of-file. For example:

     (define stdout (current-output-port))
     (define p (make-soft-port
                (vector
                 (lambda (c) (write c stdout))
                 (lambda (s) (display s stdout))
                 (lambda () (display "." stdout))
                 (lambda () (char-upcase (read-char)))
                 (lambda () (display "@" stdout)))
                "rw"))
     (write p p) ⇒ #<input-output: soft 8081e20>

6.14.9.4 Void Ports

This kind of port causes any data to be discarded when written to, and always returns the end-of-file object when read from.

[Scheme Procedure] %make-void-port mode

[C Function] scm_sys_make_void_port (mode)

Create and return a new void port. A void port acts like /dev/null. The mode argument specifies the input/output modes for this port: see the documentation for open-file in Section 6.14.9.1 [File Ports], page 326.

6.14.10 R6RS I/O Ports

The I/O port API of the Revised Report^6 on the Algorithmic Language Scheme (R6RS) (http://www.r6rs.org/) is provided by the (rnrs io ports) module. It provides features, such as binary I/O and Unicode string I/O, that complement or refine Guile’s historical port API presented above (see Section 6.14 [Input and Output], page 316). Note that R6RS ports are not disjoint from Guile’s native ports, so Guile-specific procedures will work on ports created using the R6RS API, and vice versa.

The text in this section is taken from the R6RS standard libraries document, with only minor adaptions for inclusion in this manual. The Guile developers offer their thanks to the R6RS editors for having provided the report’s text under permissive conditions making this possible.

Note: The implementation of this R6RS API is not complete yet.

A subset of the (rnrs io ports) module, plus one non-standard procedure ungetbytevector (see Section 6.14.10.8 [R6RS Binary Input], page 339), is provided by the (ice-9 binary-ports) module. It contains binary input/output procedures and does not rely on R6RS support.

6.14.10.1 File Names

Some of the procedures described in this chapter accept a file name as an argument. Valid values for such a file name include strings that name a file using the native notation of file system paths on an implementation’s underlying operating system, and may include implementation-dependent values as well.

A filename parameter name means that the corresponding argument must be a file name.

6.14.10.2 File Options

When opening a file, the various procedures in this library accept a file-options object that encapsulates flags to specify how the file is to be opened. A file-options object is an enum-set (see Section 7.6.2.24 [rnrs enums], page 654) over the symbols constituting valid file options.

A file-options parameter name means that the corresponding argument must be a fileoptions object.

[Scheme Syntax] file-options file-options-symbol ...

Each file-options-symbol must be a symbol.

The file-options syntax returns a file-options object that encapsulates the specified options.

When supplied to an operation that opens a file for output, the file-options object returned by (file-options) specifies that the file is created if it does not exist and an exception with condition type &i/o-file-already-exists is raised if it does exist. The following standard options can be included to modify the default behavior.

     no-create

If the file does not already exist, it is not created; instead, an exception with condition type &i/o-file-does-not-exist is raised. If the file already exists, the exception with condition type &i/o-file-already-exists is not raised and the file is truncated to zero length.

no-fail If the file already exists, the exception with condition type &i/o-file-already-exists is not raised, even if no-create is not included, and the file is truncated to zero length.

     no-truncate

If the file already exists and the exception with condition type &i/ofile-already-exists has been inhibited by inclusion of no-create or no-fail, the file is not truncated, but the port’s current position is still set to the beginning of the file.

These options have no effect when a file is opened only for input. Symbols other than those listed above may be used as file-options-symbols; they have implementationspecific meaning, if any.

Note: Only the name of file-options-symbol is significant.

6.14.10.3 Buffer Modes

Each port has an associated buffer mode. For an output port, the buffer mode defines when an output operation flushes the buffer associated with the output port. For an input port, the buffer mode defines how much data will be read to satisfy read operations. The possible buffer modes are the symbols none for no buffering, line for flushing upon line endings and reading up to line endings, or other implementation-dependent behavior, and block for arbitrary buffering. This section uses the parameter name buffer-mode for arguments that must be buffer-mode symbols.

If two ports are connected to the same mutable source, both ports are unbuffered, and reading a byte or character from that shared source via one of the two ports would change the bytes or characters seen via the other port, a lookahead operation on one port will render the peeked byte or character inaccessible via the other port, while a subsequent read operation on the peeked port will see the peeked byte or character even though the port is otherwise unbuffered.

In other words, the semantics of buffering is defined in terms of side effects on shared mutable sources, and a lookahead operation has the same side effect on the shared source as a read operation.

[Scheme Syntax] buffer-mode buffer-mode-symbol

buffer-mode-symbol must be a symbol whose name is one of none, line, and block. The result is the corresponding symbol, and specifies the associated buffer mode.

Note: Only the name of buffer-mode-symbol is significant.

[Scheme Procedure] buffer-mode? obj

Returns #t if the argument is a valid buffer-mode symbol, and returns #f otherwise.

6.14.10.4 Transcoders

Several different Unicode encoding schemes describe standard ways to encode characters and strings as byte sequences and to decode those sequences. Within this document, a codec is an immutable Scheme object that represents a Unicode or similar encoding scheme.

An end-of-line style is a symbol that, if it is not none, describes how a textual port transcodes representations of line endings.

A transcoder is an immutable Scheme object that combines a codec with an end-of-line style and a method for handling decoding errors. Each transcoder represents some specific bidirectional (but not necessarily lossless), possibly stateful translation between byte sequences and Unicode characters and strings. Every transcoder can operate in the input direction (bytes to characters) or in the output direction (characters to bytes). A transcoder parameter name means that the corresponding argument must be a transcoder.

A binary port is a port that supports binary I/O, does not have an associated transcoder and does not support textual I/O. A textual port is a port that supports textual I/O, and does not support binary I/O. A textual port may or may not have an associated transcoder.

[Scheme Procedure] latin-1-codec [Scheme Procedure] utf-8-codec [Scheme Procedure] utf-16-codec

These are predefined codecs for the ISO 8859-1, UTF-8, and UTF-16 encoding schemes. A call to any of these procedures returns a value that is equal in the sense of eqv? to the result of any other call to the same procedure.

[Scheme Syntax] eol-style eol-style-symbol

eol-style-symbol should be a symbol whose name is one of lf, cr, crlf, nel, crnel,

ls, and none.

The form evaluates to the corresponding symbol. If the name of eol-style-symbol is not one of these symbols, the effect and result are implementation-dependent; in particular, the result may be an eol-style symbol acceptable as an eol-style argument to make-transcoder. Otherwise, an exception is raised. All eol-style symbols except none describe a specific line-ending encoding:

lf linefeed

cr carriage return

crlf carriage return, linefeed

nel next line

crnel carriage return, next line

ls line separator

For a textual port with a transcoder, and whose transcoder has an eol-style symbol none, no conversion occurs. For a textual input port, any eol-style symbol other than none means that all of the above line-ending encodings are recognized and are translated into a single linefeed. For a textual output port, none and lf are equivalent. Linefeed characters are encoded according to the specified eol-style symbol, and all other characters that participate in possible line endings are encoded as is.

Note: Only the name of eol-style-symbol is significant.

[Scheme Procedure] native-eol-style

Returns the default end-of-line style of the underlying platform, e.g., lf on Unix and crlf on Windows.

[Condition Type] &i/o-decoding

[Scheme Procedure] make-i/o-decoding-error port

[Scheme Procedure] i/o-decoding-error? obj

This condition type could be defined by

     (define-condition-type &i/o-decoding &i/o-port
       make-i/o-decoding-error i/o-decoding-error?)

An exception with this type is raised when one of the operations for textual input from a port encounters a sequence of bytes that cannot be translated into a character or string by the input direction of the port’s transcoder. When such an exception is raised, the port’s position is past the invalid encoding.

[Condition Type] &i/o-encoding [Scheme Procedure] make-i/o-encoding-error port char [Scheme Procedure] i/o-encoding-error? obj [Scheme Procedure] i/o-encoding-error-char condition

This condition type could be defined by

     (define-condition-type &i/o-encoding &i/o-port
       make-i/o-encoding-error i/o-encoding-error?
       (char i/o-encoding-error-char))

An exception with this type is raised when one of the operations for textual output to a port encounters a character that cannot be translated into bytes by the output direction of the port’s transcoder. char is the character that could not be encoded.

[Scheme Syntax] error-handling-mode error-handling-mode-symbol

error-handling-mode-symbol should be a symbol whose name is one of ignore, raise, and replace. The form evaluates to the corresponding symbol. If error-handling-mode-symbol is not one of these identifiers, effect and result are implementation-dependent: The result may be an error-handling-mode symbol acceptable as a handling-mode argument to make-transcoder. If it is not acceptable as a handling-mode argument to make-transcoder, an exception is raised.

Note: Only the name of error-handling-mode-symbol is significant. The error-handling mode of a transcoder specifies the behavior of textual I/O operations in the presence of encoding or decoding errors. If a textual input operation encounters an invalid or incomplete character encoding, and the error-handling mode is ignore, an appropriate number of bytes of the invalid encoding are ignored and decoding continues with the following bytes.

If the error-handling mode is replace, the replacement character U+FFFD is injected into the data stream, an appropriate number of bytes are ignored, and decoding continues with the following bytes. If the error-handling mode is raise, an exception with condition type &i/o-decoding is raised. If a textual output operation encounters a character it cannot encode, and the errorhandling mode is ignore, the character is ignored and encoding continues with the next character. If the error-handling mode is replace, a codec-specific replacement character is emitted by the transcoder, and encoding continues with the next character. The replacement character is U+FFFD for transcoders whose codec is one of the Unicode encodings, but is the ? character for the Latin-1 encoding. If the error-handling mode is raise, an exception with condition type &i/o-encoding is raised.

[Scheme Procedure] make-transcoder codec [Scheme Procedure] make-transcoder codec eol-style

[Scheme Procedure] make-transcoder codec eol-style handling-mode codec must be a codec; eol-style, if present, an eol-style symbol; and handling-mode, if present, an error-handling-mode symbol. eol-style may be omitted, in which case it defaults to the native end-of-line style of the underlying platform. handling-mode may be omitted, in which case it defaults to replace. The result is a transcoder with the behavior specified by its arguments.

[Scheme procedure] native-transcoder

Returns an implementation-dependent transcoder that represents a possibly localedependent “native” transcoding.

[Scheme Procedure] transcoder-codec transcoder [Scheme Procedure] transcoder-eol-style transcoder

[Scheme Procedure] transcoder-error-handling-mode transcoder

These are accessors for transcoder objects; when applied to a transcoder returned by make-transcoder, they return the codec, eol-style, and handling-mode arguments, respectively.

[Scheme Procedure] bytevector->string bytevector transcoder Returns the string that results from transcoding the bytevector according to the input direction of the transcoder.

[Scheme Procedure] string->bytevector string transcoder

Returns the bytevector that results from transcoding the string according to the output direction of the transcoder.

6.14.10.5 The End-of-File Object

R5RS’ eof-object? procedure is provided by the (rnrs io ports) module:

[Scheme Procedure] eof-object? obj

[C Function] scm_eof_object_p (obj)

Return true if obj is the end-of-file (EOF) object.

In addition, the following procedure is provided:

[Scheme Procedure] eof-object

[C Function] scm_eof_object ()

Return the end-of-file (EOF) object.

     (eof-object? (eof-object))
     ⇒ #t

6.14.10.6 Port Manipulation

The procedures listed below operate on any kind of R6RS I/O port.

[Scheme Procedure] port? obj

Returns #t if the argument is a port, and returns #f otherwise.

[Scheme Procedure] port-transcoder port

Returns the transcoder associated with port if port is textual and has an associated transcoder, and returns #f if port is binary or does not have an associated transcoder.

[Scheme Procedure] binary-port? port

Return #t if port is a binary port, suitable for binary data input/output.

Note that internally Guile does not differentiate between binary and textual ports, unlike the R6RS. Thus, this procedure returns true when port does not have an associated encoding—i.e., when (port-encoding port) is #f (see Section 6.14.1 [Ports], page 316). This is the case for ports returned by R6RS procedures such as open-

bytevector-input-port and make-custom-binary-output-port.

However, Guile currently does not prevent use of textual I/O procedures such as display or read-char with binary ports. Doing so “upgrades” the port from binary to textual, under the ISO-8859-1 encoding. Likewise, Guile does not prevent use of set-port-encoding! on a binary port, which also turns it into a “textual” port.

[Scheme Procedure] textual-port? port

Always return #t, as all ports can be used for textual I/O in Guile.

[Scheme Procedure] transcoded-port binary-port transcoder The transcoded-port procedure returns a new textual port with the specified transcoder. Otherwise the new textual port’s state is largely the same as that of binary-port. If binary-port is an input port, the new textual port will be an input port and will transcode the bytes that have not yet been read from binary-port. If binary-port is an output port, the new textual port will be an output port and will transcode output characters into bytes that are written to the byte sink represented by binary-port.

As a side effect, however, transcoded-port closes binary-port in a special way that allows the new textual port to continue to use the byte source or sink represented by binary-port, even though binary-port itself is closed and cannot be used by the input and output operations described in this chapter.

[Scheme Procedure] port-position port

If port supports it (see below), return the offset (an integer) indicating where the next octet will be read from/written to in port. If port does not support this operation, an error condition is raised. This is similar to Guile’s seek procedure with the SEEK_CUR argument (see Section 6.14.5 [Random Access], page 321).

[Scheme Procedure] port-has-port-position? port

Return #t is port supports port-position.

[Scheme Procedure] set-port-position! port offset

If port supports it (see below), set the position where the next octet will be read from/written to port to offset (an integer). If port does not support this operation, an error condition is raised. This is similar to Guile’s seek procedure with the SEEK_SET argument (see Section 6.14.5 [Random Access], page 321).

[Scheme Procedure] port-has-set-port-position!? port

Return #t is port supports set-port-position!.

[Scheme Procedure] call-with-port port proc

Call proc, passing it port and closing port upon exit of proc. Return the return values of proc.

6.14.10.7 Input Ports

[Scheme Procedure] input-port? obj

Returns #t if the argument is an input port (or a combined input and output port), and returns #f otherwise.

[Scheme Procedure] port-eof? input-port Returns #t if the lookahead-u8 procedure (if input-port is a binary port) or the lookahead-char procedure (if input-port is a textual port) would return the endof-file object, and #f otherwise. The operation may block indefinitely if no data is available but the port cannot be determined to be at end of file.

[Scheme Procedure] open-file-input-port filename [Scheme Procedure] open-file-input-port filename file-options

[Scheme Procedure] open-file-input-port filename file-options buffer-mode [Scheme Procedure] open-file-input-port filename file-options buffer-mode

maybe-transcoder maybe-transcoder must be either a transcoder or #f. The open-file-input-port procedure returns an input port for the named file. The file-options and maybe-transcoder arguments are optional. The file-options argument, which may determine various aspects of the returned port (see Section 6.14.10.2 [R6RS File Options], page 332), defaults to the value of (file-

options).

The buffer-mode argument, if supplied, must be one of the symbols that name a buffer mode. The buffer-mode argument defaults to block.

If maybe-transcoder is a transcoder, it becomes the transcoder associated with the returned port. If maybe-transcoder is #f or absent, the port will be a binary port and will support the port-position and set-port-position! operations. Otherwise the port will be a textual port, and whether it supports the port-position and set-port-position! operations is implementation-dependent (and possibly transcoder-dependent).

[Scheme Procedure] standard-input-port

Returns a fresh binary input port connected to standard input. Whether the port supports the port-position and set-port-position! operations is implementationdependent.

[Scheme Procedure] current-input-port

This returns a default textual port for input. Normally, this default port is associated with standard input, but can be dynamically re-assigned using the with-input-from-file procedure from the io simple (6) library (see Section 7.6.2.16 [rnrs io simple], page 643). The port may or may not have an associated transcoder; if it does, the transcoder is implementation-dependent.

6.14.10.8 Binary Input

R6RS binary input ports can be created with the procedures described below.

[Scheme Procedure] open-bytevector-input-port bv [transcoder]

[C Function] scm_open_bytevector_input_port (bv, transcoder)

Return an input port whose contents are drawn from bytevector bv (see Section 6.6.6 [Bytevectors], page 163). The transcoder argument is currently not supported.

[Scheme Procedure] make-custom-binary-input-port id read! get-position

set-position! close

[C Function] scm_make_custom_binary_input_port (id, read!, get-position,

set-position!, close) Return a new custom binary input port^{10} named id (a string) whose input is drained by invoking read! and passing it a bytevector, an index where bytes should be written, and the number of bytes to read. The read! procedure must return an integer indicating the number of bytes read, or 0 to indicate the end-of-file. Optionally, if get-position is not #f, it must be a thunk that will be called when port-position is invoked on the custom binary port and should return an integer indicating the position within the underlying data stream; if get-position was not supplied, the returned port does not support port-position. Likewise, if set-position! is not #f, it should be a one-argument procedure. When set-port-position! is invoked on the custom binary input port, set-position! is passed an integer indicating the position of the next byte is to read. Finally, if close is not #f, it must be a thunk. It is invoked when the custom binary input port is closed.

10 This is similar in spirit to Guile’s soft ports (see Section 6.14.9.3 [Soft Ports], page 331).

The returned port is fully buffered by default, but its buffering mode can be changed using setvbuf (see Section 7.2.2 [Ports and File Descriptors], page 469). Using a custom binary input port, the open-bytevector-input-port procedure could be implemented as follows:

     (define (open-bytevector-input-port source)
       (define position 0)
       (define length (bytevector-length source))
       (define (read! bv start count)
         (let ((count (min count (- length position))))
           (bytevector-copy! source position
                             bv start count)
           (set! position (+ position count))
           count))
       (define (get-position) position)
       (define (set-position! new-position)
         (set! position new-position))
       (make-custom-binary-input-port "the port" read!
                                       get-position
                                       set-position!))
     (read (open-bytevector-input-port (string->utf8 "hello")))
     ⇒ hello

Binary input is achieved using the procedures below:

[Scheme Procedure] get-u8 port [C Function] scm_get_u8 (port)

Return an octet read from port, a binary input port, blocking as necessary, or the end-of-file object.

[Scheme Procedure] lookahead-u8 port [C Function] scm_lookahead_u8 (port)

Like get-u8 but does not update port’s position to point past the octet.

[Scheme Procedure] get-bytevector-n port count [C Function] scm_get_bytevector_n (port, count)

Read count octets from port, blocking as necessary and return a bytevector containing the octets read. If fewer bytes are available, a bytevector smaller than count is returned.

[Scheme Procedure] get-bytevector-n! port bv start count [C Function] scm_get_bytevector_n_x (port, bv, start, count)

Read count bytes from port and store them in bv starting at index start. Return either the number of bytes actually read or the end-of-file object.

[Scheme Procedure] get-bytevector-some port

[C Function] scm_get_bytevector_some (port)

Read from port, blocking as necessary, until bytes are available or an end-of-file is reached. Return either the end-of-file object or a new bytevector containing some of the available bytes (at least one), and update the port position to point just past these bytes.

[Scheme Procedure] get-bytevector-all port

[C Function] scm_get_bytevector_all (port)

Read from port, blocking as necessary, until the end-of-file is reached. Return either a new bytevector containing the data read or the end-of-file object (if no data were available).

The (ice-9 binary-ports) module provides the following procedure as an extension to

(rnrs io ports):

[Scheme Procedure] unget-bytevector port bv [start [count]]

[C Function] scm_unget_bytevector (port, bv, start, count)

Place the contents of bv in port, optionally starting at index start and limiting to count octets, so that its bytes will be read from left-to-right as the next bytes from port during subsequent read operations. If called multiple times, the unread bytes will be read again in last-in first-out order.

6.14.10.9 Textual Input

[Scheme Procedure] get-char textual-input-port Reads from textual-input-port, blocking as necessary, until a complete character is available from textual-input-port, or until an end of file is reached.

If a complete character is available before the next end of file, get-char returns that character and updates the input port to point past the character. If an end of file is reached before any character is read, get-char returns the end-of-file object.

[Scheme Procedure] lookahead-char textual-input-port The lookahead-char procedure is like get-char, but it does not update textualinput-port to point past the character.

[Scheme Procedure] get-string-n textual-input-port count count must be an exact, non-negative integer object, representing the number of characters to be read.

The get-string-n procedure reads from textual-input-port, blocking as necessary, until count characters are available, or until an end of file is reached.

If count characters are available before end of file, get-string-n returns a string consisting of those count characters. If fewer characters are available before an end of file, but one or more characters can be read, get-string-n returns a string containing those characters. In either case, the input port is updated to point just past the characters read. If no characters can be read before an end of file, the end-of-file object is returned.

[Scheme Procedure] get-string-n! textual-input-port string start count start and count must be exact, non-negative integer objects, with count representing the number of characters to be read. string must be a string with at least $start + count$ characters. The get-string-n! procedure reads from textual-input-port in the same manner as get-string-n. If count characters are available before an end of file, they are written into string starting at index start, and count is returned. If fewer characters are available before an end of file, but one or more can be read, those characters are written into string starting at index start and the number of characters actually read is returned as an exact integer object. If no characters can be read before an end of file, the end-of-file object is returned.

[Scheme Procedure] get-string-all textual-input-port Reads from textual-input-port until an end of file, decoding characters in the same

manner as get-string-n and get-string-n!.

If characters are available before the end of file, a string containing all the characters decoded from that data are returned. If no character precedes the end of file, the end-of-file object is returned.

[Scheme Procedure] get-line textual-input-port Reads from textual-input-port up to and including the linefeed character or end of file, decoding characters in the same manner as get-string-n and get-string-n!. If a linefeed character is read, a string containing all of the text up to (but not including) the linefeed character is returned, and the port is updated to point just past the linefeed character. If an end of file is encountered before any linefeed character is read, but some characters have been read and decoded as characters, a string containing those characters is returned. If an end of file is encountered before any characters are read, the end-of-file object is returned.

Note: The end-of-line style, if not none, will cause all line endings to be read as linefeed characters. See Section 6.14.10.4 [R6RS Transcoders], page 334.

[Scheme Procedure] get-datum textual-input-port count Reads an external representation from textual-input-port and returns the datum it represents. The get-datum procedure returns the next datum that can be parsed from the given textual-input-port, updating textual-input-port to point exactly past the end of the external representation of the object. Any interlexeme space (comment or whitespace, see Section 6.17.1 [Scheme Syntax], page 357) in the input is first skipped. If an end of file occurs after the interlexeme space, the end-of-file object (see Section 6.14.10.5 [R6RS End-of-File], page 336) is returned. If a character inconsistent with an external representation is encountered in the input, an exception with condition types &lexical and &i/o-read is raised. Also, if the end of file is encountered after the beginning of an external representation, but the external representation is incomplete and therefore cannot be parsed, an exception with condition types &lexical and &i/o-read is raised.

6.14.10.10 Output Ports

[Scheme Procedure] output-port? obj

Returns #t if the argument is an output port (or a combined input and output port), #f otherwise.

[Scheme Procedure] flush-output-port port

Flushes any buffered output from the buffer of output-port to the underlying file, device, or object. The flush-output-port procedure returns an unspecified values.

[Scheme Procedure] open-file-output-port filename [Scheme Procedure] open-file-output-port filename file-options

[Scheme Procedure] open-file-output-port filename file-options buffer-mode [Scheme Procedure] open-file-output-port filename file-options buffer-mode

maybe-transcoder maybe-transcoder must be either a transcoder or #f.

The open-file-output-port procedure returns an output port for the named file.

The file-options argument, which may determine various aspects of the returned port (see Section 6.14.10.2 [R6RS File Options], page 332), defaults to the value of (file-

options).

The buffer-mode argument, if supplied, must be one of the symbols that name a buffer mode. The buffer-mode argument defaults to block.

If maybe-transcoder is a transcoder, it becomes the transcoder associated with the port.

If maybe-transcoder is #f or absent, the port will be a binary port and will support the port-position and set-port-position! operations. Otherwise the port will be a textual port, and whether it supports the port-position and set-port-position! operations is implementation-dependent (and possibly transcoder-dependent).

[Scheme Procedure] standard-output-port
[Scheme Procedure] standard-error-port

Returns a fresh binary output port connected to the standard output or standard error respectively. Whether the port supports the port-position and set-port-position! operations is implementation-dependent.

[Scheme Procedure] current-output-port
[Scheme Procedure] current-error-port

These return default textual ports for regular output and error output. Normally, these default ports are associated with standard output, and standard error, respectively. The return value of current-output-port can be dynamically re-assigned using the with-output-to-file procedure from the io simple (6) library (see Section 7.6.2.16 [rnrs io simple], page 643). A port returned by one of these procedures may or may not have an associated transcoder; if it does, the transcoder is implementation-dependent.

6.14.10.11 Binary Output

Binary output ports can be created with the procedures below.

[Scheme Procedure] open-bytevector-output-port [transcoder]

[C Function] scm_open_bytevector_output_port (transcoder)

Return two values: a binary output port and a procedure. The latter should be called with zero arguments to obtain a bytevector containing the data accumulated by the port, as illustrated below.

     (call-with-values
       (lambda ()
         (open-bytevector-output-port))
       (lambda (port get-bytevector)
         (display "hello" port)
         (get-bytevector)))

     ⇒ #vu8(104 101 108 108 111)

The transcoder argument is currently not supported.

[Scheme Procedure] make-custom-binary-output-port id write! get-position

set-position! close

[C Function] scm_make_custom_binary_output_port (id, write!, get-position,

set-position!, close) Return a new custom binary output port named id (a string) whose output is sunk by invoking write! and passing it a bytevector, an index where bytes should be read from this bytevector, and the number of bytes to be “written”. The write! procedure must return an integer indicating the number of bytes actually written; when it is passed 0 as the number of bytes to write, it should behave as though an end-of-file was sent to the byte sink.

The other arguments are as for make-custom-binary-input-port (see

Section 6.14.10.8 [R6RS Binary Input], page 339).

Writing to a binary output port can be done using the following procedures:

[Scheme Procedure] put-u8 port octet [C Function] scm_put_u8 (port, octet)

Write octet, an integer in the 0–255 range, to port, a binary output port.

[Scheme Procedure] put-bytevector port bv [start [count]]

[C Function] scm_put_bytevector (port, bv, start, count)

Write the contents of bv to port, optionally starting at index start and limiting to count octets.

6.14.10.12 Textual Output

[Scheme Procedure] put-char port char

Writes char to the port. The put-char procedure returns an unspecified value.

[Scheme Procedure] put-string port string

[Scheme Procedure] put-string port string start [Scheme Procedure] put-string port string start count

start and count must be non-negative exact integer objects. string must have a length of at least start + count. start defaults to 0. count defaults to (string − lengthstring) − start$. The put-string procedure writes the count characters of string starting at index start to the port. The put-string procedure returns an unspecified value.

[Scheme Procedure] put-datum textual-output-port datum datum should be a datum value. The put-datum procedure writes an external representation of datum to textual-output-port. The specific external representation is implementation-dependent. However, whenever possible, an implementation should produce a representation for which get-datum, when reading the representation, will return an object equal (in the sense of equal?) to datum.

Note: Not all datums may allow producing an external representation for which get-datum will produce an object that is equal to the original. Specifically, NaNs contained in datum may make this impossible. Note: The put-datum procedure merely writes the external representation, but no trailing delimiter. If put-datum is used to write several subsequent external representations to an output port, care should be taken to delimit them properly so they can be read back in by subsequent calls

          to get-datum.

6.14.11 Using and Extending Ports in C

6.14.11.1 C Port Interface

This section describes how to use Scheme ports from C.

Port basics

There are two main data structures. A port type object (ptob) is of type scm_ptob_ descriptor. A port instance is of type scm_port. Given an SCM variable which points to a port, the corresponding C port object can be obtained using the SCM_PTAB_ENTRY macro. The ptob can be obtained by using SCM_PTOBNUM to give an index into the scm_ptobs global array.

Port buffers

An input port always has a read buffer and an output port always has a write buffer. However the size of these buffers is not guaranteed to be more than one byte (e.g., the shortbuf field in scm_port which is used when no other buffer is allocated). The way in which the buffers are allocated depends on the implementation of the ptob. For example in the case of an fport, buffers may be allocated with malloc when the port is created, but in the case of an strport the underlying string is used as the buffer.

The rw_random flag

Special treatment is required for ports which can be seeked at random. Before various operations, such as seeking the port or changing from input to output on a bidirectional port or vice versa, the port implementation must be given a chance to update its state. The write buffer is updated by calling the flush ptob procedure and the input buffer is updated by calling the end_input ptob procedure. In the case of an fport, flush causes buffered output to be written to the file descriptor, while end_input causes the descriptor position to be adjusted to account for buffered input which was never read.

The special treatment must be performed if the rw_random flag in the port is non-zero.

The rw_active variable

The rw_active variable in the port is only used if rw_random is set. It’s defined as an enum with the following values:

SCM_PORT_READ

the read buffer may have unread data.

SCM_PORT_WRITE

the write buffer may have unwritten data.

SCM_PORT_NEITHER

neither the write nor the read buffer has data.

Reading from a port.

To read from a port, it’s possible to either call existing libguile procedures such as scm_getc and scm_read_line or to read data from the read buffer directly. Reading from the buffer involves the following steps:

1. Flush output on the port, if rw_active is SCM_PORT_WRITE.

2. Fill the read buffer, if it’s empty, using scm_fill_input.

3. Read the data from the buffer and update the read position in the buffer. Steps 2) and 3) may be repeated as many times as required.

4. Set rw active to SCM_PORT_READ if rw_random is set.

5. update the port’s line and column counts.

Writing to a port.

To write data to a port, calling scm_lfwrite should be sufficient for most purposes. This takes care of the following steps:

1. End input on the port, if rw_active is SCM_PORT_READ.

2. Pass the data to the ptob implementation using the write ptob procedure. The advantage of using the ptob write instead of manipulating the write buffer directly is that it allows the data to be written in one operation even if the port is using the single-byte

    shortbuf.
 3. Set rw_active to SCM_PORT_WRITE if rw_random is set.

6.14.11.2 Port Implementation

This section describes how to implement a new port type in C.

As described in the previous section, a port type object (ptob) is a structure of type

scm_ptob_descriptor. A ptob is created by calling scm_make_port_type.

[Function] scm_t_bits scm_make_port_type (char *name, int (*fill input) (SCM

port), void (*write) (SCM port, const void *data, size t size)) Return a new port type object. The name, fill input and write parameters are initial values for those port type fields, as described below. The other fields are initialized with default values and can be changed later.

All of the elements of the ptob, apart from name, are procedures which collectively implement the port behaviour. Creating a new port type mostly involves writing these procedures.

name A pointer to a NUL terminated string: the name of the port type. This is the only element of scm_ptob_descriptor which is not a procedure. Set via the

first argument to scm_make_port_type.

mark Called during garbage collection to mark any SCM objects that a port object may contain. It doesn’t need to be set unless the port has SCM components. Set using

[Function] void scm_set_port_mark (scm t bits tc, SCM (*mark)

(SCM port))

free Called when the port is collected during gc. It should free any resources used by the port. Set using

[Function] void scm_set_port_free (scm t bits tc, size t (*free)

(SCM port))

print Called when write is called on the port object, to print a port description. E.g., for an fport it may produce something like: #<input: /etc/passwd 3>. Set using

[Function] void scm_set_port_print (scm t bits tc, int (*print) (SCM

port, SCM dest port, scm print state *pstate)) The first argument port is the object being printed, the second argument dest port is where its description should go.

equalp Not used at present. Set using

[Function] void scm_set_port_equalp (scm t bits tc, SCM (*equalp)

(SCM, SCM))

close Called when the port is closed, unless it was collected during gc. It should free any resources used by the port. Set using

[Function] void scm_set_port_close (scm t bits tc, int (*close) (SCM

port))

write Accept data which is to be written using the port. The port implementation may choose to buffer the data instead of processing it directly. Set via the third

argument to scm_make_port_type.

flush Complete the processing of buffered output data. Reset the value of rw_active

to SCM_PORT_NEITHER. Set using

[Function] void scm_set_port_flush (scm t bits tc, void (*flush)

(SCM port))

end_input

Perform any synchronization required when switching from input to output on the port. Reset the value of rw_active to SCM_PORT_NEITHER. Set using

[Function] void scm_set_port_end_input (scm t bits tc, void

(*end input) (SCM port, int offset))

fill_input

Read new data into the read buffer and return the first character. It can be assumed that the read buffer is empty when this procedure is called. Set via the second argument to scm_make_port_type.

input_waiting

Return a lower bound on the number of bytes that could be read from the port without blocking. It can be assumed that the current state of rw_active is

SCM_PORT_NEITHER. Set using
[Function] void scm_set_port_input_waiting (scm t bits tc, int

(*input waiting) (SCM port))

seek Set the current position of the port. The procedure can not make any assumptions about the value of rw_active when it’s called. It can reset the buffers first if desired by using something like:

     if (pt->rw_active == SCM_PORT_READ)
       scm_end_input (port);
     else if (pt->rw_active == SCM_PORT_WRITE)
       ptob->flush (port);

However note that this will have the side effect of discarding any data in the unread-char buffer, in addition to any side effects from the end_input and flush ptob procedures. This is undesirable when seek is called to measure the current position of the port, i.e., (seek p 0 SEEK_CUR). The libguile fport and string port implementations take care to avoid this problem. The procedure is set using

[Function] void scm_set_port_seek (scm t bits tc, scm t off (*seek)

(SCM port, scm t off offset, int whence))

truncate Truncate the port data to be specified length. It can be assumed that the

current state of rw_active is SCM_PORT_NEITHER. Set using

[Function] void scm_set_port_truncate (scm t bits tc, void

(*truncate) (SCM port, scm t off length))

6.14.12 Handling of Unicode byte order marks.

This section documents the finer points of Guile’s handling of Unicode byte order marks (BOMs). A byte order mark (U+FEFF) is typically found at the start of a UTF-16 or UTF-32 stream, to allow readers to reliably determine the byte order. Occasionally, a BOM is found at the start of a UTF-8 stream, but this is much less common and not generally recommended.

Guile attempts to handle BOMs automatically, and in accordance with the recommendations of the Unicode Standard, when the port encoding is set to UTF-8, UTF-16, or UTF-32. In brief, Guile automatically writes a BOM at the start of a UTF-16 or UTF-32 stream, and automatically consumes one from the start of a UTF-8, UTF-16, or UTF-32 stream.

As specified in the Unicode Standard, a BOM is only handled specially at the start of a stream, and only if the port encoding is set to UTF-8, UTF-16 or UTF-32. If the port encoding is set to UTF-16BE, UTF-16LE, UTF-32BE, or UTF-32LE, then BOMs are not handled specially, and none of the special handling described in this section applies.

• To ensure that Guile will properly detect the byte order of a UTF-16 or UTF-32 stream, you must perform a textual read before any writes, seeks, or binary I/O. Guile will not attempt to read a BOM unless a read is explicitly requested at the start of the stream.

• If a textual write is performed before the first read, then an arbitrary byte order will be chosen. Currently, big endian is the default on all platforms, but that may change in the future. If you wish to explicitly control the byte order of an output stream, set the port encoding to UTF-16BE, UTF-16LE, UTF-32BE, or UTF-32LE, and explicitly write a BOM (#\xFEFF) if desired.

• If set-port-encoding! is called in the middle of a stream, Guile treats this as a new logical “start of stream” for purposes of BOM handling, and will forget about any BOMs that had previously been seen. Therefore, it may choose a different byte order than had been used previously. This is intended to support multiple logical text streams embedded within a larger binary stream.

• Binary I/O operations are not guaranteed to update Guile’s notion of whether the port is at the “start of the stream”, nor are they guaranteed to produce or consume BOMs.

• For ports that support seeking (e.g. normal files), the input and output streams are considered linked: if the user reads first, then a BOM will be consumed (if appropriate), but later writes will not produce a BOM. Similarly, if the user writes first, then later reads will not consume a BOM.

• For ports that do not support seeking (e.g. pipes, sockets, and terminals), the input and output streams are considered independent for purposes of BOM handling: the first read will consume a BOM (if appropriate), and the first write will also produce a BOM (if appropriate). However, the input and output streams will always use the same byte order.

• Seeks to the beginning of a file will set the “start of stream” flags. Therefore, a subsequent textual read or write will consume or produce a BOM. However, unlike set-port-encoding!, if a byte order had already been chosen for the port, it will remain in effect after a seek, and cannot be changed by the presence of a BOM. Seeks anywhere other than the beginning of a file clear the “start of stream” flags.

6.15 Regular Expressions

A regular expression (or regexp) is a pattern that describes a whole class of strings. A full description of regular expressions and their syntax is beyond the scope of this manual; an introduction can be found in the Emacs manual (see Section “Syntax of Regular Expressions” in The GNU Emacs Manual), or in many general Unix reference books.

If your system does not include a POSIX regular expression library, and you have not linked Guile with a third-party regexp library such as Rx, these functions will not be available. You can tell whether your Guile installation includes regular expression support by checking whether (provided? 'regex) returns true.

The following regexp and string matching features are provided by the (ice-9 regex) module. Before using the described functions, you should load this module by executing

(use-modules (ice-9 regex)).

6.15.1 Regexp Functions

By default, Guile supports POSIX extended regular expressions. That means that the characters ‘(’, ‘)’, ‘+’ and ‘?’ are special, and must be escaped if you wish to match the literal characters.

This regular expression interface was modeled after that implemented by SCSH, the Scheme Shell. It is intended to be upwardly compatible with SCSH regular expressions.

Zero bytes (#\nul) cannot be used in regex patterns or input strings, since the underlying C functions treat that as the end of string. If there’s a zero byte an error is thrown.

Internally, patterns and input strings are converted to the current locale’s encoding, and then passed to the C library’s regular expression routines (see Section “Regular Expressions” in The GNU C Library Reference Manual). The returned match structures always point to characters in the strings, not to individual bytes, even in the case of multi-byte encodings.

[Scheme Procedure] string-match pattern str [start] Compile the string pattern into a regular expression and compare it with str. The optional numeric argument start specifies the position of str at which to begin matching. string-match returns a match structure which describes what, if anything, was matched by the regular expression. See Section 6.15.2 [Match Structures], page 354. If str does not match pattern at all, string-match returns #f.

Two examples of a match follow. In the first example, the pattern matches the four digits in the match string. In the second, the pattern matches nothing.

     (string-match "[0-9][0-9][0-9][0-9]" "blah2002")
     ⇒ #("blah2002" (4 . 8))
     (string-match "[A-Za-z]" "123456")
     ⇒ #f

Each time string-match is called, it must compile its pattern argument into a regular expression structure. This operation is expensive, which makes string-match inefficient if the same regular expression is used several times (for example, in a loop). For better performance, you can compile a regular expression in advance and then match strings against the compiled regexp.

[Scheme Procedure] make-regexp pat flag… [C Function] scm_make_regexp (pat, flaglst)

Compile the regular expression described by pat, and return the compiled regexp structure. If pat does not describe a legal regular expression, make-regexp throws a

     regular-expression-syntax error.

The flag arguments change the behavior of the compiled regular expression. The following values may be supplied:

[Variable] regexp/icase

Consider uppercase and lowercase letters to be the same when matching.

[Variable] regexp/newline

If a newline appears in the target string, then permit the ‘^’ and ‘$’ operators to match immediately after or immediately before the newline, respectively. Also, the ‘.’ and ‘[^...]’ operators will never match a newline character. The intent of this flag is to treat the target string as a buffer containing many lines of text, and the regular expression as a pattern that may match a single one of those lines.

[Variable] regexp/basic

Compile a basic (“obsolete”) regexp instead of the extended (“modern”) regexps that are the default. Basic regexps do not consider ‘|’, ‘+’ or ‘?’ to be special characters, and require the ‘{...}’ and ‘(...)’ metacharacters to be backslashescaped (see Section 6.15.3 [Backslash Escapes], page 355). There are several other differences between basic and extended regular expressions, but these are the most significant.

[Variable] regexp/extended

Compile an extended regular expression rather than a basic regexp. This is the default behavior; this flag will not usually be needed. If a call to make-regexp includes both regexp/basic and regexp/extended flags, the one which comes last will override the earlier one.

[Scheme Procedure] regexp-exec rx str [start [flags]]

[C Function] scm_regexp_exec (rx, str, start, flags)

Match the compiled regular expression rx against str. If the optional integer start argument is provided, begin matching from that position in the string. Return a match structure describing the results of the match, or #f if no match could be found.

The flags argument changes the matching behavior. The following flag values may be supplied, use logior (see Section 6.6.2.13 [Bitwise Operations], page 125) to combine them,

[Variable] regexp/notbol

Consider that the start offset into str is not the beginning of a line and should not match operator ‘^’.

If rx was created with the regexp/newline option above, ‘^’ will still match after a newline in str.

[Variable] regexp/noteol

Consider that the end of str is not the end of a line and should not match operator ‘$’.

If rx was created with the regexp/newline option above, ‘$’ will still match before a newline in str.

;; Regexp to match uppercase letters
(define r (make-regexp "[A-Z]*"))
;; Regexp to match letters, ignoring case
(define ri (make-regexp "[A-Z]*" regexp/icase))
;; Search for bob using regexp r
(match:substring (regexp-exec r "bob"))
⇒ ""                  ; no match
;; Search for bob using regexp ri
(match:substring (regexp-exec ri "Bob"))
⇒ "Bob"               ; matched case insensitive

[Scheme Procedure] regexp? obj [C Function] scm_regexp_p (obj)

Return #t if obj is a compiled regular expression, or #f otherwise.

[Scheme Procedure] list-matches regexp str [flags]

Return a list of match structures which are the non-overlapping matches of regexp in str. regexp can be either a pattern string or a compiled regexp. The flags argument is as per regexp-exec above.

     (map match:substring (list-matches "[a-z]+" "abc 42 def 78"))
     ⇒ ("abc" "def")

[Scheme Procedure] fold-matches regexp str init proc [flags] Apply proc to the non-overlapping matches of regexp in str, to build a result. regexp can be either a pattern string or a compiled regexp. The flags argument is as per

regexp-exec above.

proc is called as (proc match prev) where match is a match structure and prev is the previous return from proc. For the first call prev is the given init parameter. fold-matches returns the final value from proc. For example to count matches,

     (fold-matches "[a-z][0-9]" "abc x1 def y2" 0
                   (lambda (match count)
                     (1+ count)))
     ⇒ 2

Regular expressions are commonly used to find patterns in one string and replace them with the contents of another string. The following functions are convenient ways to do this.

[Scheme Procedure] regexp-substitute port match item …

Write to port selected parts of the match structure match. Or if port is #f then form a string from those parts and return that. Each item specifies a part to be written, and may be one of the following, • A string. String arguments are written out verbatim.

• An integer. The submatch with that number is written (match:substring). Zero is the entire match.

• The symbol ‘pre’. The portion of the matched string preceding the regexp match

    is written (match:prefix).

• The symbol ‘post’. The portion of the matched string following the regexp match

    is written (match:suffix).

For example, changing a match and retaining the text before and after,

     (regexp-substitute #f (string-match "[0-9]+" "number 25 is good")
                        'pre "37" 'post)
     ⇒ "number 37 is good"

Or matching a yyyymmdd format date such as ‘20020828’ and re-ordering and hyphenating the fields.

     (define date-regex
        "([0-9][0-9][0-9][0-9])([0-9][0-9])([0-9][0-9])")
     (define s "Date 20020429 12am.")
     (regexp-substitute #f (string-match date-regex s)
                        'pre 2 "-" 3 "-" 1 'post " (" 0 ")")
     ⇒ "Date 04-29-2002 12am. (20020429)"

[Scheme Procedure] regexp-substitute/global port regexp target item…

Write to port selected parts of matches of regexp in target. If port is #f then form a string from those parts and return that. regexp can be a string or a compiled regex. This is similar to regexp-substitute, but allows global substitutions on target. Each item behaves as per regexp-substitute, with the following differences,

• A function. Called as (item match) with the match structure for the regexp match, it should return a string to be written to port.

• The symbol ‘post’. This doesn’t output anything, but instead causes regexpsubstitute/global to recurse on the unmatched portion of target. This must be supplied to perform a global search and replace on target; without it regexp-substitute/global returns after a single match and output.

For example, to collapse runs of tabs and spaces to a single hyphen each,

     (regexp-substitute/global #f "[ \t]+"  "this   is   the text"
                               'pre "-" 'post)
     ⇒ "this-is-the-text"

Or using a function to reverse the letters in each word,

     (regexp-substitute/global #f "[a-z]+"  "to do and not-do"
       'pre (lambda (m) (string-reverse (match:substring m))) 'post)
     ⇒ "ot od dna ton-od"

Without the post symbol, just one regexp match is made. For example the following is the date example from regexp-substitute above, without the need for the separate

string-match call.

     (define date-regex
        "([0-9][0-9][0-9][0-9])([0-9][0-9])([0-9][0-9])")
     (define s "Date 20020429 12am.")
     (regexp-substitute/global #f date-regex s
                               'pre 2 "-" 3 "-" 1 'post " (" 0 ")")

     ⇒ "Date 04-29-2002 12am. (20020429)"

6.15.2 Match Structures

A match structure is the object returned by string-match and regexp-exec. It describes which portion of a string, if any, matched the given regular expression. Match structures include: a reference to the string that was checked for matches; the starting and ending positions of the regexp match; and, if the regexp included any parenthesized subexpressions, the starting and ending positions of each submatch.

In each of the regexp match functions described below, the match argument must be a match structure returned by a previous call to string-match or regexp-exec. Most of these functions return some information about the original target string that was matched against a regular expression; we will call that string target for easy reference.

[Scheme Procedure] regexp-match? obj

Return #t if obj is a match structure returned by a previous call to regexp-exec, or #f otherwise.

[Scheme Procedure] match:substring match [n]

Return the portion of target matched by subexpression number n. Submatch 0 (the default) represents the entire regexp match. If the regular expression as a whole matched, but the subexpression number n did not match, return #f.

(define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
(match:substring s)
⇒ "2002"
;; match starting at offset 6 in the string
(match:substring
  (string-match "[0-9][0-9][0-9][0-9]" "blah987654" 6))
⇒ "7654"

[Scheme Procedure] match:start match [n]

Return the starting position of submatch number n.

In the following example, the result is 4, since the match starts at character index 4:

(define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
(match:start s)
⇒ 4

[Scheme Procedure] match:end match [n]

Return the ending position of submatch number n.

In the following example, the result is 8, since the match runs between characters 4 and 8 (i.e. the “2002”).

     (define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
     (match:end s)
     ⇒ 8

[Scheme Procedure] match:prefix match

Return the unmatched portion of target preceding the regexp match.

          (define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
          (match:prefix s)
          ⇒ "blah"

[Scheme Procedure] match:suffix match

Return the unmatched portion of target following the regexp match.

     (define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
     (match:suffix s)
     ⇒ "foo"

[Scheme Procedure] match:count match

Return the number of parenthesized subexpressions from match. Note that the entire regular expression match itself counts as a subexpression, and failed submatches are included in the count.

[Scheme Procedure] match:string match

Return the original target string.

     (define s (string-match "[0-9][0-9][0-9][0-9]" "blah2002foo"))
     (match:string s)
     ⇒ "blah2002foo"

6.15.3 Backslash Escapes

Sometimes you will want a regexp to match characters like ‘*’ or ‘$’ exactly. For example, to check whether a particular string represents a menu entry from an Info node, it would be useful to match it against a regexp like ‘^* [^:]*::’. However, this won’t work; because the asterisk is a metacharacter, it won’t match the ‘*’ at the beginning of the string. In this case, we want to make the first asterisk un-magic.

You can do this by preceding the metacharacter with a backslash character ‘\’. (This is also called quoting the metacharacter, and is known as a backslash escape.) When Guile sees a backslash in a regular expression, it considers the following glyph to be an ordinary character, no matter what special meaning it would ordinarily have. Therefore, we can make the above example work by changing the regexp to ‘^\* [^:]*::’. The ‘\*’ sequence tells the regular expression engine to match only a single asterisk in the target string.

Since the backslash is itself a metacharacter, you may force a regexp to match a backslash in the target string by preceding the backslash with itself. For example, to find variable references in a TEX program, you might want to find occurrences of the string ‘\let\’ followed by any number of alphabetic characters. The regular expression ‘\\let\\[A-Za-z]*’ would do this: the double backslashes in the regexp each match a single backslash in the target string.

[Scheme Procedure] regexp-quote str

Quote each special character found in str with a backslash, and return the resulting string.

Very important: Using backslash escapes in Guile source code (as in Emacs Lisp or C) can be tricky, because the backslash character has special meaning for the Guile reader. For example, if Guile encounters the character sequence ‘\n’ in the middle of a string while processing Scheme code, it replaces those characters with a newline character. Similarly, the character sequence ‘\t’ is replaced by a horizontal tab. Several of these escape sequences are processed by the Guile reader before your code is executed. Unrecognized escape sequences are ignored: if the characters ‘\*’ appear in a string, they will be translated to the single character ‘*’.

This translation is obviously undesirable for regular expressions, since we want to be able to include backslashes in a string in order to escape regexp metacharacters. Therefore, to make sure that a backslash is preserved in a string in your Guile program, you must use two consecutive backslashes:

     (define Info-menu-entry-pattern (make-regexp "^\\* [^:]*"))

The string in this example is preprocessed by the Guile reader before any code is executed. The resulting argument to make-regexp is the string ‘^\* [^:]*’, which is what we really want.

This also means that in order to write a regular expression that matches a single backslash character, the regular expression string in the source code must include four backslashes. Each consecutive pair of backslashes gets translated by the Guile reader to a single backslash, and the resulting double-backslash is interpreted by the regexp engine as matching a single backslash character. Hence:

     (define tex-variable-pattern (make-regexp "\\\\let\\\\=[A-Za-z]*"))

The reason for the unwieldiness of this syntax is historical. Both regular expression pattern matchers and Unix string processing systems have traditionally used backslashes with the special meanings described above. The POSIX regular expression specification and ANSI C standard both require these semantics. Attempting to abandon either convention would cause other kinds of compatibility problems, possibly more severe ones. Therefore, without extending the Scheme reader to support strings with different quoting conventions (an ungainly and confusing extension when implemented in other languages), we must adhere to this cumbersome escape syntax.

6.16 LALR(1) Parsing

The (system base lalr) module provides the lalr-scm LALR(1) parser generator by Do-

minique Boucher (http://code.google.com/p/lalr-scm/). lalr-scm uses the same

algorithm as GNU Bison (see Section “Introduction” in Bison, The Yacc-compatible Parser Generator). Parsers are defined using the lalr-parser macro.

[Scheme Syntax] lalr-parser [options] tokens rules...

Generate an LALR(1) syntax analyzer. tokens is a list of symbols representing the terminal symbols of the grammar. rules are the grammar production rules.

Each rule has the form (non-terminal (rhs ...) : action ...), where

non-terminal is the name of the rule, rhs are the right-hand sides, i.e., the production rule, and action is a semantic action associated with the rule. The generated parser is a two-argument procedure that takes a tokenizer and a syntax error procedure. The tokenizer should be a thunk that returns lexical tokens as produced by make-lexical-token. The syntax error procedure may be called with at least an error message (a string), and optionally the lexical token that caused the error.

Please refer to the lalr-scm documentation for details.

6.17 Reading and Evaluating Scheme Code

This chapter describes Guile functions that are concerned with reading, loading, evaluating, and compiling Scheme code at run time.

6.17.1 Scheme Syntax: Standard and Guile Extensions

6.17.1.1 Expression Syntax

An expression to be evaluated takes one of the following forms.

symbol A symbol is evaluated by dereferencing. A binding of that symbol is sought

and the value there used. For example,

     (define x 123)
     x ⇒ 123
(proc args...)

A parenthesised expression is a function call. proc and each argument are evaluated, then the function (which proc evaluated to) is called with those arguments. The order in which proc and the arguments are evaluated is unspecified, so be careful when using expressions with side effects.

     (max 1 2 3) ⇒ 3
     (define (get-some-proc)  min)
     ((get-some-proc) 1 2 3) ⇒ 1

The same sort of parenthesised form is used for a macro invocation, but in that case the arguments are not evaluated. See the descriptions of macros for more on this (see Section 6.10 [Macros], page 257, and see Section 6.10.2 [Syntax Rules], page 258).

constant Number, string, character and boolean constants evaluate “to themselves”, so

can appear as literals.

     123  ⇒ 123
     99.9  ⇒ 99.9
     "hello" ⇒ "hello"
     #\z  ⇒ #\z
     #t  ⇒ #t

Note that an application must not attempt to modify literal strings, since they may be in read-only memory.

(quote data)

'data Quoting is used to obtain a literal symbol (instead of a variable reference), a literal list (instead of a function call), or a literal vector. ' is simply a shorthand for a quote form. For example,

     'x  ⇒ x

     '(1 2 3)  ⇒ (1 2 3)
     '#(1 (2 3) 4)  ⇒ #(1 (2 3) 4)
     (quote x)  ⇒ x
     (quote (1 2 3))  ⇒ (1 2 3)
     (quote #(1 (2 3) 4)) ⇒ #(1 (2 3) 4)

Note that an application must not attempt to modify literal lists or vectors obtained from a quote form, since they may be in read-only memory.

(quasiquote data)

`data Backquote quasi-quotation is like quote, but selected sub-expressions are evaluated. This is a convenient way to construct a list or vector structure most of which is constant, but at certain points should have expressions substituted. The same effect can always be had with suitable list, cons or vector calls, but quasi-quoting is often easier.

(unquote expr)

,expr Within the quasiquote data, unquote or , indicates an expression to be evaluated and inserted. The comma syntax , is simply a shorthand for an unquote form. For example,

               `(1 2 ,(* 9 9) 3 4)  ⇒ (1 2 81 3 4)

               `(1 (unquote (+ 1 1)) 3) ⇒ (1 2 3)
               `#(1 ,(/ 12 2))  ⇒ #(1 6)

(unquote-splicing expr)

,@expr Within the quasiquote data, unquote-splicing or ,@ indicates an expression to be evaluated and the elements of the returned list inserted. expr must evaluate to a list. The “comma-at” syntax ,@ is simply a shorthand for an unquote-splicing form.

               (define x '(2 3))
               `(1 ,@x 4)  ⇒ (1 2 3 4)
               `(1 (unquote-splicing (map 1+ x))) ⇒ (1 3 4)
               `#(9 ,@x 9)  ⇒ #(9 2 3 9)

Notice ,@ differs from plain , in the way one level of nesting is stripped. For ,@ the elements of a returned list are inserted, whereas with , it would be the list itself inserted.

6.17.1.2 Comments

Comments in Scheme source files are written by starting them with a semicolon character (;). The comment then reaches up to the end of the line. Comments can begin at any column, and the may be inserted on the same line as Scheme code.

     ; Comment
     ;; Comment too
     (define x 1)        ; Comment after expression
     (let ((y 1))
       ;; Display something.
       (display y)
     ;;; Comment at left margin.
       (display (+ y 1)))

It is common to use a single semicolon for comments following expressions on a line, to use two semicolons for comments which are indented like code, and three semicolons for comments which start at column 0, even if they are inside an indented code block. This convention is used when indenting code in Emacs’ Scheme mode.

6.17.1.3 Block Comments

In addition to the standard line comments defined by R5RS, Guile has another comment type for multiline comments, called block comments. This type of comment begins with the character sequence #! and ends with the characters !#, which must appear on a line of their own. These comments are compatible with the block comments in the Scheme Shell scsh (see Section 7.17 [The Scheme shell (scsh)], page 693). The characters #! were chosen because they are the magic characters used in shell scripts for indicating that the name of the program for executing the script follows on the same line.

Thus a Guile script often starts like this.

     #! /usr/local/bin/guile -s
     !#

More details on Guile scripting can be found in the scripting section (see Section 4.3 [Guile Scripting], page 41).

Similarly, Guile (starting from version 2.0) supports nested block comments as specified

by R6RS and SRFI-30 (http://srfi.schemers.org/srfi-30/srfi-30.html):
     (+ 1 #| this is a #| nested |# block comment |# 2)
     ⇒ 3

For backward compatibility, this syntax can be overridden with read-hash-extend (see Section 6.17.1.6 [Reader Extensions], page 360).

There is one special case where the contents of a comment can actually affect the interpretation of code. When a character encoding declaration, such as coding: utf-8 appears in one of the first few lines of a source file, it indicates to Guile’s default reader that this source code file is not ASCII. For details see Section 6.17.8 [Character Encoding of Source Files], page 370.

6.17.1.4 Case Sensitivity

Scheme as defined in R5RS is not case sensitive when reading symbols. Guile, on the contrary is case sensitive by default, so the identifiers

     guile-whuzzy
     Guile-Whuzzy

are the same in R5RS Scheme, but are different in Guile.

It is possible to turn off case sensitivity in Guile by setting the reader option caseinsensitive. For more information on reader options, See Section 6.17.2 [Scheme Read], page 360.

(read-enable 'case-insensitive)

It is also possible to disable (or enable) case sensitivity within a single file by placing the reader directives #!fold-case (or #!no-fold-case) within the file itself.

6.17.1.5 Keyword Syntax

6.17.1.6 Reader Extensions

[Scheme Procedure] read-hash-extend chr proc [C Function] scm_read_hash_extend (chr, proc)

Install the procedure proc for reading expressions starting with the character sequence # and chr. proc will be called with two arguments: the character chr and the port to read further data from. The object returned will be the return value of read. Passing #f for proc will remove a previous setting.

6.17.2 Reading Scheme Code

[Scheme Procedure] read [port] [C Function] scm_read (port)

Read an s-expression from the input port port, or from the current input port if port is not specified. Any whitespace before the next token is discarded.

The behaviour of Guile’s Scheme reader can be modified by manipulating its read options.

[Scheme Procedure] read-options [setting]

Display the current settings of the global read options. If setting is omitted, only a short form of the current read options is printed. Otherwise if setting is the symbol help, a complete options description is displayed.

The set of available options, and their default values, may be had by invoking readoptions at the prompt.

scheme@(guile-user)> (read-options)
(square-brackets keywords #f positions)
scheme@(guile-user)> (read-options 'help)
copy              no    Copy source code expressions.
positions         yes   Record positions of source code expressions.
case-insensitive  no    Convert symbols to lower case.
keywords          #f    Style of keyword recognition: #f, 'prefix or 'postfix.
r6rs-hex-escapes  no    Use R6RS variable-length character and string hex escapes.
square-brackets   yes   Treat `[' and `]' as parentheses, for R6RS compatibility.
hungry-eol-escapes no   In strings, consume leading whitespace after an
                        escaped end-of-line.
curly-infix       no    Support SRFI-105 curly infix expressions.
r7rs-symbols      no    Support R7RS |...| symbol notation.

Note that Guile also includes a preliminary mechanism for setting read options on a per-port basis. For instance, the case-insensitive read option is set (or unset) on the port when the reader encounters the #!fold-case or #!no-fold-case reader directives. Similarly, the #!curly-infix reader directive sets the curly-infix read option on the port,

and #!curly-infix-and-bracket-lists sets curly-infix and unsets square-brackets

on the port (see Section 7.5.42 [SRFI-105], page 621). There is currently no other way to access or set the per-port read options.

The boolean options may be toggled with read-enable and read-disable. The nonboolean keywords option must be set using read-set!.

[Scheme Procedure] read-enable option-name [Scheme Procedure] read-disable option-name

[Scheme Syntax] read-set! option-name value Modify the read options. read-enable should be used with boolean options and switches them on, read-disable switches them off.

read-set! can be used to set an option to a specific value. Due to historical oddities, it is a macro that expects an unquoted option name.

For example, to make read fold all symbols to their lower case (perhaps for compatibility with older Scheme code), you can enter:

(read-enable 'case-insensitive)

For more information on the effect of the r6rs-hex-escapes and hungry-eol-escapes options, see (see Section 6.6.5.1 [String Syntax], page 141).

For more information on the r7rs-symbols option, see (see Section 6.6.7.6 [Symbol Read Syntax], page 178).

6.17.3 Writing Scheme Values

Any scheme value may be written to a port. Not all values may be read back in (see Section 6.17.2 [Scheme Read], page 360), however.

[Scheme Procedure] write obj [port]

Send a representation of obj to port or to the current output port if not given.

The output is designed to be machine readable, and can be read back with read (see Section 6.17.2 [Scheme Read], page 360). Strings are printed in double quotes, with escapes if necessary, and characters are printed in ‘#\’ notation.

[Scheme Procedure] display obj [port]

Send a representation of obj to port or to the current output port if not given.

The output is designed for human readability, it differs from write in that strings are printed without double quotes and escapes, and characters are printed as per write-char, not in ‘#\’ form.

As was the case with the Scheme reader, there are a few options that affect the behavior of the Scheme printer.

[Scheme Procedure] print-options [setting]

Display the current settings of the read options. If setting is omitted, only a short form of the current read options is printed. Otherwise if setting is the symbol help, a complete options description is displayed.

The set of available options, and their default values, may be had by invoking printoptions at the prompt.

scheme@(guile-user)> (print-options)
(quote-keywordish-symbols reader highlight-suffix "}" highlight-prefix "{")
scheme@(guile-user)> (print-options 'help)
highlight-prefix          {       The string to print before highlighted values.
highlight-suffix          }       The string to print after highlighted values.
quote-keywordish-symbols  reader  How to print symbols that have a colon
                                  as their first or last character. The
                                  value '#f' does not quote the colons;
                                  '#t' quotes them; 'reader' quotes them
                                  when the reader option 'keywords' is
                                  not '#f'.
escape-newlines           yes     Render newlines as \n when printing
                                  using `write'.
r7rs-symbols              no      Escape symbols using R7RS |...| symbol
                                  notation.

These options may be modified with the print-set! syntax.

[Scheme Syntax] print-set! option-name value Modify the print options. Due to historical oddities, print-set! is a macro that expects an unquoted option name.

6.17.4 Procedures for On the Fly Evaluation

Scheme has the lovely property that its expressions may be represented as data. The eval procedure takes a Scheme datum and evaluates it as code.

[Scheme Procedure] eval exp module or state [C Function] scm_eval (exp, module or state)

Evaluate exp, a list representing a Scheme expression, in the top-level environment specified by module or state. While exp is evaluated (using primitive-eval), module or state is made the current module. The current module is reset to its previous value when eval returns. XXX - dynamic states. Example: (eval ’(+ 1 2) (interactionenvironment))

[Scheme Procedure] interaction-environment
[C Function] scm_interaction_environment ()

Return a specifier for the environment that contains implementation–defined bindings, typically a superset of those listed in the report. The intent is that this procedure will return the environment in which the implementation would evaluate expressions dynamically typed by the user.

See Section 6.19.11 [Environments], page 395, for other environments.

One does not always receive code as Scheme data, of course, and this is especially the case for Guile’s other language implementations (see Section 6.23 [Other Languages], page 432). For the case in which all you have is a string, we have eval-string. There is a legacy version of this procedure in the default environment, but you really want the one from

(ice-9 eval-string), so load it up:
(use-modules (ice-9 eval-string))

[Scheme Procedure] eval-string string [#:module=#f] [#:file=#f] [#:line=#f]

[#:column=#f] [#:lang=(current-language)] [#:compile?=#f] Parse string according to the current language, normally Scheme. Evaluate or compile the expressions it contains, in order, returning the last expression. If the module keyword argument is set, save a module excursion (see Section 6.19.8 [Module System Reflection], page 391) and set the current module to module before evaluation. The file, line, and column keyword arguments can be used to indicate that the source string begins at a particular source location. Finally, lang is a language, defaulting to the current language, and the expression is compiled if compile? is true or there is no evaluator for the given language.

[C Function] scm_eval_string (string) [C Function] scm_eval_string_in_module (string, module)

These C bindings call eval-string from (ice-9 eval-string), evaluating within module or the current module.

[C Function] SCM scm_c_eval_string (const char *string)

scm_eval_string, but taking a C string in locale encoding instead of an SCM.

[Scheme Procedure] apply proc arg … arglst [C Function] scm_apply_0 (proc, arglst) [C Function] scm_apply_1 (proc, arg1, arglst) [C Function] scm_apply_2 (proc, arg1, arg2, arglst) [C Function] scm_apply_3 (proc, arg1, arg2, arg3, arglst) [C Function] scm_apply (proc, arg, rest)

Call proc with arguments arg … and the elements of the arglst list. scm_apply takes parameters corresponding to a Scheme level (lambda (proc arg1 . rest) ...). So arg1 and all but the last element of the rest list make up arg …, and the last element of rest is the arglst list. Or if rest is the empty list SCM_EOL then there’s no arg …, and (arg1) is the arglst. arglst is not modified, but the rest list passed to scm_apply is modified.

[C Function] scm_call_0 (proc)

[C Function] scm_call_1 (proc, arg1) [C Function] scm_call_2 (proc, arg1, arg2) [C Function] scm_call_3 (proc, arg1, arg2, arg3) [C Function] scm_call_4 (proc, arg1, arg2, arg3, arg4) [C Function] scm_call_5 (proc, arg1, arg2, arg3, arg4, arg5) [C Function] scm_call_6 (proc, arg1, arg2, arg3, arg4, arg5, arg6) [C Function] scm_call_7 (proc, arg1, arg2, arg3, arg4, arg5, arg6, arg7) [C Function] scm_call_8 (proc, arg1, arg2, arg3, arg4, arg5, arg6, arg7, arg8) [C Function] scm_call_9 (proc, arg1, arg2, arg3, arg4, arg5, arg6, arg7, arg8, arg9)

Call proc with the given arguments.

[C Function] scm_call (proc, ...)

Call proc with any number of arguments. The argument list must be terminated by

SCM_UNDEFINED. For example:

          scm_call (scm_c_public_ref ("guile", "+"),
                    scm_from_int (1),
                    scm_from_int (2),
                    SCM_UNDEFINED);

[C Function] scm_call_n (proc, argv, nargs)

Call proc with the array of arguments argv, as a SCM*. The length of the arguments should be passed in nargs, as a size_t.

[Scheme Procedure] apply:nconc2last lst [C Function] scm_nconc2last (lst)

lst should be a list (arg1 … argN arglst), with arglst being a list. This function returns a list comprising arg1 to argN plus the elements of arglst. lst is modified to form the return. arglst is not modified, though the return does share structure with it. This operation collects up the arguments from a list which is apply style parameters.

[Scheme Procedure] primitive-eval exp

[C Function] scm_primitive_eval (exp)

Evaluate exp in the top-level environment specified by the current module.

6.17.5 Compiling Scheme Code

The eval procedure directly interprets the S-expression representation of Scheme. An alternate strategy for evaluation is to determine ahead of time what computations will be necessary to evaluate the expression, and then use that recipe to produce the desired results. This is known as compilation.

While it is possible to compile simple Scheme expressions such as (+ 2 2) or even "Hello world!", compilation is most interesting in the context of procedures. Compiling a lambda expression produces a compiled procedure, which is just like a normal procedure except typically much faster, because it can bypass the generic interpreter.

Functions from system modules in a Guile installation are normally compiled already, so they load and run quickly.

Note that well-written Scheme programs will not typically call the procedures in this section, for the same reason that it is often bad taste to use eval. By default, Guile automatically compiles any files it encounters that have not been compiled yet (see Section 4.2 [Invoking Guile], page 35). The compiler can also be invoked explicitly from the shell as

guild compile foo.scm.

(Why are calls to eval and compile usually in bad taste? Because they are limited, in that they can only really make sense for top-level expressions. Also, most needs for “compiletime” computation are fulfilled by macros and closures. Of course one good counterexample is the REPL itself, or any code that reads expressions from a port.)

Automatic compilation generally works transparently, without any need for user intervention. However Guile does not yet do proper dependency tracking, so that if file a.scm uses macros from b.scm, and b.scm changes, a.scm would not be automatically recompiled. To forcibly invalidate the auto-compilation cache, pass the --fresh-auto-compile option to Guile, or set the GUILE_AUTO_COMPILE environment variable to fresh (instead of to 0 or

For more information on the compiler itself, see Section 9.4 [Compiling to the Virtual Machine], page 791. For information on the virtual machine, see Section 9.3 [A Virtual Machine for Guile], page 772.

The command-line interface to Guile’s compiler is the guild compile command:

[Command] guild compile [option...] file...

Compile file, a source file, and store bytecode in the compilation cache or in the file specified by the -o option. The following options are available:

-L dir
--load-path=dir

Add dir to the front of the module load path.

-o ofile
--output=ofile

Write output bytecode to ofile. By convention, bytecode file names end in .go. When -o is omitted, the output file name is as for compile-file (see below).

-W warning
--warn=warning

Emit warnings of type warning; use --warn=help for a list of available warnings and their description. Currently recognized warnings in-

          clude unused-variable, unused-toplevel, unbound-variable, arity-
          mismatch, format, duplicate-case-datum, and bad-case-datum.
-f lang
--from=lang

Use lang as the source language of file. If this option is omitted, scheme is assumed.

-t lang
--to=lang

Use lang as the target language of file. If this option is omitted, objcode is assumed.

-T target
--target=target

Produce bytecode for target instead of %host-type (see Section 6.22.1 [Build Config], page 427). Target must be a valid GNU triplet, such

          as armv5tel-unknown-linux-gnueabi (see Section “Specifying Target

Triplets” in GNU Autoconf Manual).

Each file is assumed to be UTF-8-encoded, unless it contains a coding declaration as recognized by file-encoding (see Section 6.17.8 [Character Encoding of Source Files], page 370).

The compiler can also be invoked directly by Scheme code using the procedures below:

[Scheme Procedure] compile exp [#:env=#f] [#:from=(current-language)]

[#:to=value] [#:opts=()] Compile the expression exp in the environment env. If exp is a procedure, the result will be a compiled procedure; otherwise compile is mostly equivalent to eval.

For a discussion of languages and compiler options, See Section 9.4 [Compiling to the Virtual Machine], page 791.

[Scheme Procedure] compile-file file [#:output-file=#f] [#:from=(current-language)] [#:to=’objcode] [#:env=(default-environment from)] [#:opts=’()] [#:canonicalization=’relative]

Compile the file named file. Output will be written to a output-file. If you do not supply an output file name, output is written to a file in the cache directory, as computed by (compiled-file-

name file).

from and to specify the source and target languages. See Section 9.4 [Compiling to the Virtual Machine], page 791, for more information on these options, and on env and opts. As with guild compile, file is assumed to be UTF-8-encoded unless it contains a coding declaration.

[Scheme Procedure] compiled-file-name file

Compute a cached location for a compiled version of a Scheme file named file. This file will usually be below the $HOME/.cache/guile/ccache directory, depending on the value of the XDG_CACHE_HOME environment variable. The intention is that compiled-file-name provides a fallback location for caching auto-compiled files. If you want to place a compile file in the %load-compiled-path, you should pass the output-file option to compile-file, explicitly.

[Scheme Variable] %auto-compilation-options

This variable contains the options passed to the compile-file procedure when autocompiling source files. By default, it enables useful compilation warnings. It can be customized from ~/.guile.

6.17.6 Loading Scheme Code from File

[Scheme Procedure] load filename [reader] Load filename and evaluate its contents in the top-level environment. reader if provided should be either #f, or a procedure with the signature (lambda (port) ...) which reads the next expression from port. If reader is #f or absent, Guile’s built-in read procedure is used (see Section 6.17.2 [Scheme Read], page 360). The reader argument takes effect by setting the value of the current-reader fluid (see below) before loading the file, and restoring its previous value when loading is complete. The Scheme code inside filename can itself change the current reader procedure on the fly by setting current-reader fluid. If the variable %load-hook is defined, it should be bound to a procedure that will be called before any code is loaded. See documentation for %load-hook later in this section.

[Scheme Procedure] load-compiled filename

Load the compiled file named filename.

Compiling a source file (see Section 6.17 [Read/Load/Eval/Compile], page 357) and then calling load-compiled on the resulting file is equivalent to calling load on the source file.

[Scheme Procedure] primitive-load filename [C Function] scm_primitive_load (filename)

Load the file named filename and evaluate its contents in the top-level environment. filename must either be a full pathname or be a pathname relative to the current directory. If the variable %load-hook is defined, it should be bound to a procedure that will be called before any code is loaded. See the documentation for %load-hook later in this section.

[C Function] SCM scm_c_primitive_load (const char *filename)

scm_primitive_load, but taking a C string instead of an SCM.

[Variable] current-reader

current-reader holds the read procedure that is currently being used by the above loading procedures to read expressions (from the file that they are loading). currentreader is a fluid, so it has an independent value in each dynamic root and should be read and set using fluid-ref and fluid-set! (see Section 6.21.7 [Fluids and Dynamic States], page 419).

Changing current-reader is typically useful to introduce local syntactic changes, such that code following the fluid-set! call is read using the newly installed reader. The current-reader change should take place at evaluation time when the code is evaluated, or at compilation time when the code is compiled:

     (eval-when (compile eval)
       (fluid-set! current-reader my-own-reader))

The eval-when form above ensures that the current-reader change occurs at the right time.

[Variable] %load-hook

A procedure to be called (%load-hook filename) whenever a file is loaded, or #f for no such call. %load-hook is used by all of the loading functions (load and

primitive-load, and load-from-path and primitive-load-path documented in

the next section).

For example an application can set this to show what’s loaded,

     (set! %load-hook (lambda (filename)
                        (format #t "Loading ~a ...\n" filename)))
     (load-from-path "foo.scm")
     ⊣ Loading /usr/local/share/guile/site/foo.scm ...
[Scheme Procedure] current-load-port
[C Function] scm_current_load_port ()

Return the current-load-port. The load port is used internally by primitive-load.

6.17.7 Load Paths

The procedure in the previous section look for Scheme code in the file system at specific location. Guile also has some procedures to search the load path for code.

[Variable] %load-path

List of directories which should be searched for Scheme modules and libraries. When Guile starts up, %load-path is initialized to the default load path (list

(%library-dir) (%site-dir) (%global-site-dir) (%package-data-dir)). The

GUILE_LOAD_PATH environment variable can be used to prepend or append additional directories (see Section 4.2.2 [Environment Variables], page 38). See Section 6.22.1 [Build Config], page 427, for more on %site-dir and related procedures.

[Scheme Procedure] load-from-path filename

Similar to load, but searches for filename in the load paths. Preferentially loads a compiled version of the file, if it is available and up-to-date.

A user can extend the load path by calling add-to-load-path.

[Scheme Syntax] add-to-load-path dir

Add dir to the load path.

For example, a script might include this form to add the directory that it is in to the load path:

(add-to-load-path (dirname (current-filename)))

It’s better to use add-to-load-path than to modify %load-path directly, because addto-load-path takes care of modifying the path both at compile-time and at run-time.

[Scheme Procedure] primitive-load-path filename [exception-on-not-found]

[C Function] scm_primitive_load_path (filename)

Search %load-path for the file named filename and load it into the top-level environment. If filename is a relative pathname and is not found in the list of search paths, an error is signalled. Preferentially loads a compiled version of the file, if it is available and up-to-date. If filename is a relative pathname and is not found in the list of search paths, one of three things may happen, depending on the optional second argument, exceptionon-not-found. If it is #f, #f will be returned. If it is a procedure, it will be called with no arguments. (This allows a distinction to be made between exceptions raised by loading a file, and exceptions related to the loader itself.) Otherwise an error is signalled. For compatibility with Guile 1.8 and earlier, the C function takes only one argument, which can be either a string (the file name) or an argument list.

[Scheme Procedure] %search-load-path filename

[C Function] scm_sys_search_load_path (filename)

Search %load-path for the file named filename, which must be readable by the current user. If filename is found in the list of paths to search or is an absolute pathname, return its full pathname. Otherwise, return #f. Filenames may have any of the optional extensions in the %load-extensions list; %search-load-path will try each extension automatically.

[Variable] %load-extensions

A list of default file extensions for files containing Scheme code. %search-load-path tries each of these extensions when looking for a file to load. By default, %load-

extensions is bound to the list ("" ".scm").

As mentioned above, when Guile searches the %load-path for a source file, it will also search the %load-compiled-path for a corresponding compiled file. If the compiled file is as new or newer than the source file, it will be loaded instead of the source file, using

load-compiled.
[Variable] %load-compiled-path

Like %load-path, but for compiled files. By default, this path has two entries: one for compiled files from Guile itself, and one for site packages. The GUILE_LOAD_ COMPILED_PATH environment variable can be used to prepend or append additional directories (see Section 4.2.2 [Environment Variables], page 38).

When primitive-load-path searches the %load-compiled-path for a corresponding

compiled file for a relative path it does so by appending .go to the relative path. For exam-

ple, searching for ice-9/popen could find /usr/lib/guile/2.0/ccache/ice-9/popen.go,
and use it instead of /usr/share/guile/2.0/ice-9/popen.scm.
If primitive-load-path does not find a corresponding .go file in the %load-compiled-

path, or the .go file is out of date, it will search for a corresponding auto-compiled file in the fallback path, possibly creating one if one does not exist.

See Section 4.7 [Installing Site Packages], page 55, for more on how to correctly install site packages. See Section 6.19.4 [Modules and the File System], page 386, for more on the relationship between load paths and modules. See Section 6.17.5 [Compilation], page 364, for more on the fallback path and auto-compilation.

Finally, there are a couple of helper procedures for general path manipulation.

[Scheme Procedure] parse-path path [tail] [C Function] scm_parse_path (path, tail)

Parse path, which is expected to be a colon-separated string, into a list and return the resulting list with tail appended. If path is #f, tail is returned.

[Scheme Procedure] parse-path-with-ellipsis path base

[C Function] scm_parse_path_with_ellipsis (path, base)

Parse path, which is expected to be a colon-separated string, into a list and return the resulting list with base (a list) spliced in place of the ... path component, if present, or else base is added to the end. If path is #f, base is returned.

[Scheme Procedure] search-path path filename [extensions [require-exts?]] [C Function] scm_search_path (path, filename, rest)

Search path for a directory containing a file named filename. The file must be readable, and not a directory. If we find one, return its full filename; otherwise, return #f. If filename is absolute, return it unchanged. If given, extensions is a list of strings; for each directory in path, we search for filename concatenated with each extension. If require-exts? is true, require that the returned file name have one of the given extensions; if require-exts? is not given, it defaults to #f.

For compatibility with Guile 1.8 and earlier, the C function takes only three arguments.

6.17.8 Character Encoding of Source Files

Scheme source code files are usually encoded in ASCII or UTF-8, but the built-in reader can interpret other character encodings as well. When Guile loads Scheme source code, it uses the file-encoding procedure (described below) to try to guess the encoding of the file. In the absence of any hints, UTF-8 is assumed. One way to provide a hint about the encoding of a source file is to place a coding declaration in the top 500 characters of the file.

A coding declaration has the form coding: XXXXXX, where XXXXXX is the name of a character encoding in which the source code file has been encoded. The coding declaration must appear in a scheme comment. It can either be a semicolon-initiated comment, or the first block #! comment in the file.

The name of the character encoding in the coding declaration is typically lower case and containing only letters, numbers, and hyphens, as recognized by set-port-encoding! (see Section 6.14.1 [Ports], page 316). Common examples of character encoding names

are utf-8 and iso-8859-1, as defined by IANA (http://www.iana.org/assignments/

character-sets). Thus, the coding declaration is mostly compatible with Emacs.

However, there are some differences in encoding names recognized by Emacs and encoding names defined by IANA, the latter being essentially a subset of the former. For instance, latin-1 is a valid encoding name for Emacs, but it’s not according to the IANA standard, which Guile follows; instead, you should use iso-8859-1, which is both understood by Emacs and dubbed by IANA (IANA writes it uppercase but Emacs wants it lowercase and Guile is case insensitive.)

For source code, only a subset of all possible character encodings can be interpreted by the built-in source code reader. Only those character encodings in which ASCII text appears unmodified can be used. This includes UTF-8 and ISO-8859-1 through ISO-8859-15. The multi-byte character encodings UTF-16 and UTF-32 may not be used because they are not compatible with ASCII.

There might be a scenario in which one would want to read non-ASCII code from a port, such as with the function read, instead of with load. If the port’s character encoding is the same as the encoding of the code to be read by the port, not other special handling is necessary. The port will automatically do the character encoding conversion. The functions setlocale or by set-port-encoding! are used to set port encodings (see Section 6.14.1 [Ports], page 316).

If a port is used to read code of unknown character encoding, it can accomplish this in three steps. First, the character encoding of the port should be set to ISO-8859-1 using set-port-encoding!. Then, the procedure file-encoding, described below, is used to scan for a coding declaration when reading from the port. As a side effect, it rewinds the port after its scan is complete. After that, the port’s character encoding should be set to the encoding returned by file-encoding, if any, again by using set-port-encoding!. Then the code can be read as normal.

Alternatively, one can use the #:guess-encoding keyword argument of open-file and related procedures. See Section 6.14.9.1 [File Ports], page 326.

[Scheme Procedure] file-encoding port

[C Function] scm_file_encoding (port)

Attempt to scan the first few hundred bytes from the port for hints about its character encoding. Return a string containing the encoding name or #f if the encoding cannot be determined. The port is rewound.

Currently, the only supported method is to look for an Emacs-like character coding declaration (see Section “Recognize Coding” in The GNU Emacs Reference Manual). The coding declaration is of the form coding: XXXXX and must appear in a Scheme comment. Additional heuristics may be added in the future.

6.17.9 Delayed Evaluation

Promises are a convenient way to defer a calculation until its result is actually needed, and to run such a calculation only once. Also see Section 7.5.30 [SRFI-45], page 614.

[syntax] delay expr

Return a promise object which holds the given expr expression, ready to be evaluated by a later force.

[Scheme Procedure] promise? obj [C Function] scm_promise_p (obj)

Return true if obj is a promise.

[Scheme Procedure] force p [C Function] scm_force (p)

Return the value obtained from evaluating the expr in the given promise p. If p has previously been forced then its expr is not evaluated again, instead the value obtained at that time is simply returned.

During a force, an expr can call force again on its own promise, resulting in a recursive evaluation of that expr. The first evaluation to return gives the value for the promise. Higher evaluations run to completion in the normal way, but their results are ignored, force always returns the first value.

6.17.10 Local Evaluation

Guile includes a facility to capture a lexical environment, and later evaluate a new expression within that environment. This code is implemented in a module.

(use-modules (ice-9 local-eval))
[syntax] the-environment

Captures and returns a lexical environment for use with local-eval or local-

compile.

[Scheme Procedure] local-eval exp env [C Function] scm_local_eval (exp, env) [Scheme Procedure] local-compile exp env [opts=()]

Evaluate or compile the expression exp in the lexical environment env.

Here is a simple example, illustrating that it is the variable that gets captured, not just its value at one point in time.

     (define e (let ((x 100)) (the-environment)))
     (define fetch-x (local-eval '(lambda () x) e))
     (fetch-x)
     ⇒ 100
     (local-eval '(set! x 42) e)
     (fetch-x)
     ⇒ 42

While exp is evaluated within the lexical environment of (the-environment), it has the dynamic environment of the call to local-eval.

local-eval and local-compile can only evaluate expressions, not definitions.

     (local-eval '(define foo 42)
                 (let ((x 100)) (the-environment)))
     ⇒ syntax error: definition in expression context

Note that the current implementation of (the-environment) only captures “normal” lexical bindings, and pattern variables bound by syntax-case. It does not currently capture local syntax transformers bound by let-syntax, letrec-syntax or non-top-level definesyntax forms. Any attempt to reference such captured syntactic keywords via local-eval or local-compile produces an error.

6.17.11 Local Inclusion

This section has discussed various means of linking Scheme code together: fundamentally, loading up files at run-time using load and load-compiled. Guile provides another option to compose parts of programs together at expansion-time instead of at run-time.

[Scheme Syntax] include file-name Open file-name, at expansion-time, and read the Scheme forms that it contains, splicing them into the location of the include, within a begin. If file-name is a relative path, it is searched for relative to the path that contains the file that the include for appears in.

If you are a C programmer, if load in Scheme is like dlopen in C, consider include to be like the C preprocessor’s #include. When you use include, it is as if the contents of the included file were typed in instead of the include form.

Because the code is included at compile-time, it is available to the macroexpander. Syntax definitions in the included file are available to later code in the form in which the include appears, without the need for eval-when. (See Section 6.10.8 [Eval When], page 274.)

For the same reason, compiling a form that uses include results in one compilation unit, composed of multiple files. Loading the compiled file is one stat operation for the compilation unit, instead of 2*n in the case of load (once for each loaded source file, and once each corresponding compiled file, in the best case).

Unlike load, include also works within nested lexical contexts. It so happens that the optimizer works best within a lexical context, because all of the uses of bindings in a lexical context are visible, so composing files by including them within a (let () ...) can sometimes lead to important speed improvements.

On the other hand, include does have all the disadvantages of early binding: once the code with the include is compiled, no change to the included file is reflected in the future behavior of the including form.

Also, the particular form of include, which requires an absolute path, or a path relative to the current directory at compile-time, is not very amenable to compiling the source in one place, but then installing the source to another place. For this reason, Guile provides another form, include-from-path, which looks for the source file to include within a load path.

[Scheme Syntax] include-from-path file-name

Like include, but instead of expecting file-name to be an absolute file name, it is expected to be a relative path to search in the %load-path.

include-from-path is more useful when you want to install all of the source files for a package (as you should!). It makes it possible to evaluate an installed file from source, instead of relying on the .go file being up to date.

6.17.12 REPL Servers

The procedures in this section are provided by

     (use-modules (system repl server))

When an application is written in Guile, it is often convenient to allow the user to be able to interact with it by evaluating Scheme expressions in a REPL.

The procedures of this module allow you to spawn a REPL server, which permits interaction over a local or TCP connection. Guile itself uses them internally to implement the --listen switch, Section 4.2.1 [Command-line Options], page 35.

[Scheme Procedure] make-tcp-server-socket [#:host=#f] [#:addr]

[#:port=37146] Return a stream socket bound to a given address addr and port number port. If the host is given, and addr is not, then the host string is converted to an address. If neither is given, we use the loopback address.

[Scheme Procedure] make-unix-domain-server-socket

[#:path="/tmp/guile-socket"] Return a UNIX domain socket, bound to a given path.

[Scheme Procedure] run-server [server-socket]

[Scheme Procedure] spawn-server [server-socket]

Create and run a REPL, making it available over the given server-socket. If serversocket is not provided, it defaults to the socket created by calling make-tcp-server-socket with no arguments. run-server runs the server in the current thread, whereas spawn-server runs the server in a new thread.

[Scheme Procedure] stop-server-and-clients!

Closes the connection on all running server sockets.

Please note that in the current implementation, the REPL threads are cancelled without unwinding their stacks. If any of them are holding mutexes or are within a critical section, the results are unspecified.

6.17.13 Cooperative REPL Servers

The procedures in this section are provided by

     (use-modules (system repl coop-server))

Whereas ordinary REPL servers run in their own threads (see Section 6.17.12 [REPL Servers], page 373), sometimes it is more convenient to provide REPLs that run at specified times within an existing thread, for example in programs utilizing an event loop or in singlethreaded programs. This allows for safe access and mutation of a program’s data structures from the REPL, without concern for thread synchronization.

Although the REPLs are run in the thread that calls spawn-coop-repl-server and poll-coop-repl-server, dedicated threads are spawned so that the calling thread is not blocked. The spawned threads read input for the REPLs and to listen for new connections.

Cooperative REPL servers must be polled periodically to evaluate any pending expressions by calling poll-coop-repl-server with the object returned from spawn-coop-repl-server. The thread that calls poll-coop-repl-server will be blocked for as long as the expression takes to be evaluated or if the debugger is entered.

[Scheme Procedure] spawn-coop-repl-server [server-socket]

Create and return a new cooperative REPL server object, and spawn a new thread to listen for connections on server-socket. Proper functioning of the REPL server requires that poll-coop-repl-server be called periodically on the returned server object.

[Scheme Procedure] poll-coop-repl-server coop-server

Poll the cooperative REPL server coop-server and apply a pending operation if there is one, such as evaluating an expression typed at the REPL prompt. This procedure must be called from the same thread that called spawn-coop-repl-server.

6.18 Memory Management and Garbage Collection

Guile uses a garbage collector to manage most of its objects. While the garbage collector is designed to be mostly invisible, you sometimes need to interact with it explicitly.

See Section 5.4.2 [Garbage Collection], page 65 for a general discussion of how garbage collection relates to using Guile from C.

6.18.1 Function related to Garbage Collection

[Scheme Procedure] gc [C Function] scm_gc ()

Scans all of SCM objects and reclaims for further use those that are no longer accessible. You normally don’t need to call this function explicitly. It is called automatically when appropriate.

[C Function] SCM scm_gc_protect_object (SCM obj)

Protects obj from being freed by the garbage collector, when it otherwise might be. When you are done with the object, call scm_gc_unprotect_object on the object.

Calls to scm_gc_protect/scm_gc_unprotect_object can be nested, and the object

remains protected until it has been unprotected as many times as it was protected. It is an error to unprotect an object more times than it has been protected. Returns the SCM object it was passed. Note that storing obj in a C global variable has the same effect^{11}.

[C Function] SCM scm_gc_unprotect_object (SCM obj)

Unprotects an object from the garbage collector which was protected by scm_gc_ unprotect_object. Returns the SCM object it was passed.

[C Function] SCM scm_permanent_object (SCM obj)

Similar to scm_gc_protect_object in that it causes the collector to always mark the object, except that it should not be nested (only call scm_permanent_object on an object once), and it has no corresponding unpermanent function. Once an object is declared permanent, it will never be freed. Returns the SCM object it was passed.

[C Macro] void scm_remember_upto_here_1 (SCM obj)
[C Macro] void scm_remember_upto_here_2 (SCM obj1, SCM obj2)

Create a reference to the given object or objects, so they’re certain to be present on the stack or in a register and hence will not be freed by the garbage collector before this point. Note that these functions can only be applied to ordinary C local variables (ie. “automatics”). Objects held in global or static variables or some malloced block or the like cannot be protected with this mechanism.

[Scheme Procedure] gc-stats [C Function] scm_gc_stats ()

Return an association list of statistics about Guile’s current use of storage.

[Scheme Procedure] gc-live-object-stats
[C Function] scm_gc_live_object_stats ()

Return an alist of statistics of the current live objects.

[Function] void scm_gc_mark (SCM x)

Mark the object x, and recurse on any objects x refers to. If x’s mark bit is already set, return immediately. This function must only be called during the mark-phase of garbage collection, typically from a smob mark function.

11 In Guile up to version 1.8, C global variables were not scanned by the garbage collector; hence, scm_gc_

protect_object was the only way in C to prevent a Scheme object from being freed.

6.18.2 Memory Blocks

In C programs, dynamic management of memory blocks is normally done with the functions malloc, realloc, and free. Guile has additional functions for dynamic memory allocation that are integrated into the garbage collector and the error reporting system.

Memory blocks that are associated with Scheme objects (for example a smob) should be

allocated with scm_gc_malloc or scm_gc_malloc_pointerless. These two functions will

either return a valid pointer or signal an error. Memory blocks allocated this way can be freed with scm_gc_free; however, this is not strictly needed: memory allocated with scm_ gc_malloc or scm_gc_malloc_pointerless is automatically reclaimed when the garbage collector no longer sees any live reference to it^{12}.

Memory allocated with scm_gc_malloc is scanned for live pointers. This means that if scm_gc_malloc-allocated memory contains a pointer to some other part of the memory, the garbage collector notices it and prevents it from being reclaimed^{13}. Conversely, memory allocated with scm_gc_malloc_pointerless is assumed to be “pointer-less” and is not scanned.

For memory that is not associated with a Scheme object, you can use scm_malloc instead of malloc. Like scm_gc_malloc, it will either return a valid pointer or signal an error. However, it will not assume that the new memory block can be freed by a garbage collection. The memory must be explicitly freed with free.

There is also scm_gc_realloc and scm_realloc, to be used in place of realloc when appropriate, and scm_gc_calloc and scm_calloc, to be used in place of calloc when appropriate.

The function scm_dynwind_free can be useful when memory should be freed with libc’s free when leaving a dynwind context, See Section 6.13.10 [Dynamic Wind], page 309.

[C Function] void * scm_malloc (size t size)
[C Function] void * scm_calloc (size t size)

Allocate size bytes of memory and return a pointer to it. When size is 0, return NULL. When not enough memory is available, signal an error. This function runs the GC to free up some memory when it deems it appropriate.

The memory is allocated by the libc malloc function and can be freed with free. There is no scm_free function to go with scm_malloc to make it easier to pass memory back and forth between different modules.

The function scm_calloc is similar to scm_malloc, but initializes the block of memory to zero as well.

These functions will (indirectly) call scm_gc_register_allocation.

[C Function] void * scm_realloc (void *mem, size t new_size)

Change the size of the memory block at mem to new size and return its new location. When new size is 0, this is the same as calling free on mem and NULL is returned.

12 In Guile up to version 1.8, memory allocated with scm_gc_malloc had to be freed with scm_gc_free.

13 In Guile up to 1.8, memory allocated with scm_gc_malloc was not scanned. Consequently, the GC had

to be told explicitly about pointers to live objects contained in the memory block, e.g., via SMOB mark

functions (see Section 6.8 [Smobs], page 241)

When mem is NULL, this function behaves like scm_malloc and allocates a new block of size new size. When not enough memory is available, signal an error. This function runs the GC to free up some memory when it deems it appropriate.

This function will call scm_gc_register_allocation.

[C Function] void * scm_gc_malloc (size t size, const char *what)

[C Function] void * scm_gc_malloc_pointerless (size t size, const char
   *what)
[C Function] void * scm_gc_realloc (void *mem, size t old_size, size t

new_size, const char *what);

[C Function] void * scm_gc_calloc (size t size, const char *what)

Allocate size bytes of automatically-managed memory. The memory is automatically freed when no longer referenced from any live memory block. Memory allocated with scm_gc_malloc or scm_gc_calloc is scanned for pointers. Memory allocated by scm_gc_malloc_pointerless is not scanned. The scm_gc_realloc call preserves the “pointerlessness” of the memory area pointed to by mem. Note that you need to pass the old size of a reallocated memory block as well. See below for a motivation.

[C Function] void scm_gc_free (void *mem, size t size, const char *what)

Explicitly free the memory block pointed to by mem, which was previously allocated by one of the above scm_gc functions. Note that you need to explicitly pass the size parameter. This is done since it should normally be easy to provide this parameter (for memory that is associated with GC controlled objects) and help keep the memory management overhead very low. However, in Guile 2.x, size is always ignored.

[C Function] void scm_gc_register_allocation (size t size)

Informs the garbage collector that size bytes have been allocated, which the collector would otherwise not have known about. In general, Scheme will decide to collect garbage only after some amount of memory has been allocated. Calling this function will make the Scheme garbage collector know about more allocation, and thus run more often (as appropriate). It is especially important to call this function when large unmanaged allocations, like images, may be freed by small Scheme allocations, like SMOBs.

[C Function] void scm_dynwind_free (void *mem)

Equivalent  to  scm_dynwind_unwind_handler (free, mem, SCM_F_WIND_

EXPLICITLY). That is, the memory block at mem will be freed (using free from the C library) when the current dynwind is left.

[Scheme Procedure] malloc-stats

Return an alist ((what . n) ...) describing number of malloced objects. what is the second argument to scm_gc_malloc, n is the number of objects of that type currently allocated. This function is only available if the GUILE_DEBUG_MALLOC preprocessor macro was defined when Guile was compiled.

6.18.2.1 Upgrading from scm must malloc et al.

Version 1.6 of Guile and earlier did not have the functions from the previous section. In

their place, it had the functions scm_must_malloc, scm_must_realloc and scm_must_free.

This section explains why we want you to stop using them, and how to do this.

   The functions scm_must_malloc and scm_must_realloc behaved like scm_gc_malloc

and scm_gc_realloc do now, respectively. They would inform the GC about the newly allocated memory via the internal equivalent of scm_gc_register_allocation. However, scm_must_free did not unregister the memory it was about to free. The usual way to unregister memory was to return its size from a smob free function.

This disconnectedness of the actual freeing of memory and reporting this to the GC proved to be bad in practice. It was easy to make mistakes and report the wrong size because allocating and freeing was not done with symmetric code, and because it is cumbersome to compute the total size of nested data structures that were freed with multiple calls to scm_must_free. Additionally, there was no equivalent to scm_malloc, and it was tempting to just use scm_must_malloc and never to tell the GC that the memory has been freed.

The effect was that the internal statistics kept by the GC drifted out of sync with reality and could even overflow in long running programs. When this happened, the result was a dramatic increase in (senseless) GC activity which would effectively stop the program dead.

The functions scm_done_malloc and scm_done_free were introduced to help restore balance to the force, but existing bugs did not magically disappear, of course.

Therefore we decided to force everybody to review their code by deprecating the existing functions and introducing new ones in their place that are hopefully easier to use correctly.

For every use of scm_must_malloc you need to decide whether to use scm_malloc or scm_gc_malloc in its place. When the memory block is not part of a smob or some other Scheme object whose lifetime is ultimately managed by the garbage collector, use scm_ malloc and free. When it is part of a smob, use scm_gc_malloc and change the smob free function to use scm_gc_free instead of scm_must_free or free and make it return zero.

The important thing is to always pair scm_malloc with free; and to always pair scm_

gc_malloc with scm_gc_free.
   The same reasoning applies to scm_must_realloc and scm_realloc versus scm_gc_
realloc.

6.18.3 Weak References

[FIXME: This chapter is based on Mikael Djurfeldt’s answer to a question by Michael Livshin. Any mistakes are not theirs, of course. ]

Weak references let you attach bookkeeping information to data so that the additional information automatically disappears when the original data is no longer in use and gets garbage collected. In a weak key hash, the hash entry for that key disappears as soon as the key is no longer referenced from anywhere else. For weak value hashes, the same happens as soon as the value is no longer in use. Entries in a doubly weak hash disappear when either the key or the value are not used anywhere else anymore.

Object properties offer the same kind of functionality as weak key hashes in many situations. (see Section 6.11.2 [Object Properties], page 278)

Here’s an example (a little bit strained perhaps, but one of the examples is actually used in Guile):

Assume that you’re implementing a debugging system where you want to associate information about filename and position of source code expressions with the expressions themselves.

Hashtables can be used for that, but if you use ordinary hash tables it will be impossible for the scheme interpreter to "forget" old source when, for example, a file is reloaded.

To implement the mapping from source code expressions to positional information it is necessary to use weak-key tables since we don’t want the expressions to be remembered just because they are in our table.

To implement a mapping from source file line numbers to source code expressions you would use a weak-value table.

To implement a mapping from source code expressions to the procedures they constitute a doubly-weak table has to be used.

6.18.3.1 Weak hash tables

[Scheme Procedure] make-weak-key-hash-table [size]
[Scheme Procedure] make-weak-value-hash-table [size]
[Scheme Procedure] make-doubly-weak-hash-table [size]
[C Function] scm_make_weak_key_hash_table (size)
[C Function] scm_make_weak_value_hash_table (size)
[C Function] scm_make_doubly_weak_hash_table (size)

Return a weak hash table with size buckets. As with any hash table, choosing a good size for the table requires some caution.

You can modify weak hash tables in exactly the same way you would modify regular hash tables. (see Section 6.7.14 [Hash Tables], page 236)

[Scheme Procedure] weak-key-hash-table? obj
[Scheme Procedure] weak-value-hash-table? obj
[Scheme Procedure] doubly-weak-hash-table? obj
[C Function] scm_weak_key_hash_table_p (obj)
[C Function] scm_weak_value_hash_table_p (obj)
[C Function] scm_doubly_weak_hash_table_p (obj)

Return #t if obj is the specified weak hash table. Note that a doubly weak hash table is neither a weak key nor a weak value hash table.

6.18.3.2 Weak vectors

[Scheme Procedure] make-weak-vector size [fill] [C Function] scm_make_weak_vector (size, fill)

Return a weak vector with size elements. If the optional argument fill is given, all entries in the vector will be set to fill. The default value for fill is the empty list.

[Scheme Procedure] weak-vector elem … [Scheme Procedure] list->weak-vector l

[C Function] scm_weak_vector (l)

Construct a weak vector from a list: weak-vector uses the list of its arguments while list->weak-vector uses its only argument l (a list) to construct a weak vector the same way list->vector would.

[Scheme Procedure] weak-vector? obj

[C Function] scm_weak_vector_p (obj)

Return #t if obj is a weak vector.

[Scheme Procedure] weak-vector-ref wvect k [C Function] scm_weak_vector_ref (wvect, k)

Return the kth element of the weak vector wvect, or #f if that element has been collected.

[Scheme Procedure] weak-vector-set! wvect k elt [C Function] scm_weak_vector_set_x (wvect, k, elt)

Set the kth element of the weak vector wvect to elt.

6.18.4 Guardians

Guardians provide a way to be notified about objects that would otherwise be collected as garbage. Guarding them prevents the objects from being collected and cleanup actions can be performed on them, for example.

See R. Kent Dybvig, Carl Bruggeman, and David Eby (1993) "Guardians in a Generation-Based Garbage Collector". ACM SIGPLAN Conference on Programming Language Design and Implementation, June 1993.

[Scheme Procedure] make-guardian

[C Function] scm_make_guardian ()

Create a new guardian. A guardian protects a set of objects from garbage collection, allowing a program to apply cleanup or other actions. make-guardian returns a procedure representing the guardian. Calling the guardian procedure with an argument adds the argument to the guardian’s set of protected objects. Calling the guardian procedure without an argument returns one of the protected objects which are ready for garbage collection, or #f if no such object is available. Objects which are returned in this way are removed from the guardian. You can put a single object into a guardian more than once and you can put a single object into more than one guardian. The object will then be returned multiple times by the guardian procedures. An object is eligible to be returned from a guardian when it is no longer referenced from outside any guardian. There is no guarantee about the order in which objects are returned from a guardian. If you want to impose an order on finalization actions, for example, you can do that by keeping objects alive in some global data structure until they are no longer needed for finalizing other objects. Being an element in a weak vector, a key in a hash table with weak keys, or a value in a hash table with weak values does not prevent an object from being returned by a guardian. But as long as an object can be returned from a guardian it will not be removed from such a weak vector or hash table. In other words, a weak link does not prevent an object from being considered collectable, but being inside a guardian prevents a weak link from being broken. A key in a weak key hash table can be thought of as having a strong reference to its associated value as long as the key is accessible. Consequently, when the key is only accessible from within a guardian, the reference from the key to the value is also considered to be coming from within a guardian. Thus, if there is no other reference to the value, it is eligible to be returned from a guardian.

6.19 Modules

When programs become large, naming conflicts can occur when a function or global variable defined in one file has the same name as a function or global variable in another file. Even just a similarity between function names can cause hard-to-find bugs, since a programmer might type the wrong function name.

The approach used to tackle this problem is called information encapsulation, which consists of packaging functional units into a given name space that is clearly separated from other name spaces.

The language features that allow this are usually called the module system because programs are broken up into modules that are compiled separately (or loaded separately in an interpreter).

Older languages, like C, have limited support for name space manipulation and protection. In C a variable or function is public by default, and can be made local to a module with the static keyword. But you cannot reference public variables and functions from another module with different names.

More advanced module systems have become a common feature in recently designed languages: ML, Python, Perl, and Modula 3 all allow the renaming of objects from a foreign module, so they will not clutter the global name space.

In addition, Guile offers variables as first-class objects. They can be used for interacting with the module system.

6.19.1 General Information about Modules

A Guile module can be thought of as a collection of named procedures, variables and macros. More precisely, it is a set of bindings of symbols (names) to Scheme objects.

Within a module, all bindings are visible. Certain bindings can be declared public, in which case they are added to the module’s so-called export list; this set of public bindings is called the module’s public interface (see Section 6.19.3 [Creating Guile Modules], page 383).

A client module uses a providing module’s bindings by either accessing the providing module’s public interface, or by building a custom interface (and then accessing that). In a custom interface, the client module can select which bindings to access and can also algorithmically rename bindings. In contrast, when using the providing module’s public interface, the entire export list is available without renaming (see Section 6.19.2 [Using Guile Modules], page 382).

All Guile modules have a unique module name, for example (ice-9 popen) or (srfi srfi-11). Module names are lists of one or more symbols.

When Guile goes to use an interface from a module, for example (ice-9 popen), Guile first looks to see if it has loaded (ice-9 popen) for any reason. If the module has not been loaded yet, Guile searches a load path for a file that might define it, and loads that file.

The following subsections go into more detail on using, creating, installing, and otherwise manipulating modules and the module system.

6.19.2 Using Guile Modules

To use a Guile module is to access either its public interface or a custom interface (see Section 6.19.1 [General Information about Modules], page 381). Both types of access are handled by the syntactic form use-modules, which accepts one or more interface specifications and, upon evaluation, arranges for those interfaces to be available to the current module. This process may include locating and loading code for a given module if that code has not yet been loaded, following %load-path (see Section 6.19.4 [Modules and the File System], page 386).

An interface specification has one of two forms. The first variation is simply to name the module, in which case its public interface is the one accessed. For example:

     (use-modules (ice-9 popen))

Here, the interface specification is (ice-9 popen), and the result is that the current module now has access to open-pipe, close-pipe, open-input-pipe, and so on (see Section 7.2.10 [Pipes], page 500).

Note in the previous example that if the current module had already defined open-pipe, that definition would be overwritten by the definition in (ice-9 popen). For this reason (and others), there is a second variation of interface specification that not only names a module to be accessed, but also selects bindings from it and renames them to suit the current module’s needs. For example:

     (use-modules ((ice-9 popen)
                   #:select ((open-pipe . pipe-open) close-pipe)
                   #:renamer (symbol-prefix-proc 'unixy:)))

Here, the interface specification is more complex than before, and the result is that a custom interface with only two bindings is created and subsequently accessed by the current module. The mapping of old to new names is as follows:

      (ice-9 popen) sees:             current module sees:
      open-pipe                       unixy:pipe-open
      close-pipe                      unixy:close-pipe

This example also shows how to use the convenience procedure symbol-prefix-proc. You can also directly refer to bindings in a module by using the @ syntax. For example,

instead of using the use-modules statement from above and writing unixy:pipe-open to

refer to the pipe-open from the (ice-9 popen), you could also write (@ (ice-9 popen)

open-pipe). Thus an alternative to the complete use-modules statement would be

     (define unixy:pipe-open (@ (ice-9 popen) open-pipe))
     (define unixy:close-pipe (@ (ice-9 popen) close-pipe))

There is also @@, which can be used like @, but does not check whether the variable that is being accessed is actually exported. Thus, @@ can be thought of as the impolite version of @ and should only be used as a last resort or for debugging, for example.

Note that just as with a use-modules statement, any module that has not yet been loaded yet will be loaded when referenced by a @ or @@ form.

You can also use the @ and @@ syntaxes as the target of a set! when the binding refers to a variable.

[Scheme Procedure] symbol-prefix-proc prefix-sym

Return a procedure that prefixes its arg (a symbol) with prefix-sym.

[syntax] use-modules spec …

Resolve each interface specification spec into an interface and arrange for these to be accessible by the current module. The return value is unspecified. spec can be a list of symbols, in which case it names a module whose public interface is found and used. spec can also be of the form:

      (MODULE-NAME [#:select SELECTION] [#:renamer RENAMER])

in which case a custom interface is newly created and used. module-name is a list of symbols, as above; selection is a list of selection-specs; and renamer is a procedure that takes a symbol and returns its new name. A selection-spec is either a symbol or a pair of symbols (ORIG . SEEN), where orig is the name in the used module and seen is the name in the using module. Note that seen is also passed through renamer. The #:select and #:renamer clauses are optional. If both are omitted, the returned interface has no bindings. If the #:select clause is omitted, renamer operates on the used module’s public interface. In addition to the above, spec can also include a #:version clause, of the form:

      #:version VERSION-SPEC

where version-spec is an R6RS-compatible version reference. An error will be signaled in the case in which a module with the same name has already been loaded, if that module specifies a version and that version is not compatible with version-spec. See Section 6.19.5 [R6RS Version References], page 387, for more on version references. If the module name is not resolvable, use-modules will signal an error.

[syntax] @ module-name binding-name Refer to the binding named binding-name in module module-name. The binding must have been exported by the module.

[syntax] @@ module-name binding-name Refer to the binding named binding-name in module module-name. The binding must not have been exported by the module. This syntax is only intended for debugging purposes or as a last resort.

6.19.3 Creating Guile Modules

When you want to create your own modules, you have to take the following steps:

• Create a Scheme source file and add all variables and procedures you wish to export, or which are required by the exported procedures.

• Add a define-module form at the beginning.

• Export all bindings which should be in the public interface, either by using definepublic or export (both documented below).

[syntax] define-module module-name option …

module-name is a list of one or more symbols.

(define-module (ice-9 popen))

define-module makes this module available to Guile programs under the given module-name. option … are keyword/value pairs which specify more about the defined module. The recognized options and their meaning are shown in the following table.

#:use-module interface-specification
Equivalent to a (use-modules interface-specification) (see

Section 6.19.2 [Using Guile Modules], page 382).

#:autoload module symbol-list

Load module when any of symbol-list are accessed. For example,

     (define-module (my mod)
       #:autoload (srfi srfi-1) (partition delete-duplicates))
     ...
     (if something
         (set! foo (delete-duplicates ...)))

When a module is autoloaded, all its bindings become available. symbollist is just those that will first trigger the load. An autoload is a good way to put off loading a big module until it’s really needed, for instance for faster startup or if it will only be needed in certain circumstances. @ can do a similar thing (see Section 6.19.2 [Using Guile Modules], page 382), but in that case an @ form must be written every time a binding from the module is used.

#:export list

Export all identifiers in list which must be a list of symbols or pairs of symbols. This is equivalent to (export list) in the module body.

#:re-export list

Re-export all identifiers in list which must be a list of symbols or pairs of symbols. The symbols in list must be imported by the current module from other modules. This is equivalent to re-export below.

#:replace list

Export all identifiers in list (a list of symbols or pairs of symbols) and mark them as replacing bindings. In the module user’s name space, this will have the effect of replacing any binding with the same name that is not also “replacing”. Normally a replacement results in an “override” warning message, #:replace avoids that. In general, a module that exports a binding for which the (guile) module already has a definition should use #:replace instead of #:export.

#:replace, in a sense, lets Guile know that the module purposefully replaces a core binding. It is important to note, however, that this binding replacement is confined to the name space of the module user. In other words, the value of the core binding in question remains unchanged for other modules. Note that although it is often a good idea for the replaced binding to remain compatible with a binding in (guile), to avoid surprising the user, sometimes the bindings will be incompatible. For example, SRFI-19 exports its own version of current-time (see Section 7.5.16.2 [SRFI-19 Time], page 584) which is not compatible with the core current-time function (see Section 7.2.5 [Time], page 485). Guile assumes that a user importing a module knows what she is doing, and uses #:replace for this binding rather than #:export. A #:replace clause is equivalent to (export! list) in the module body. The #:duplicates (see below) provides fine-grain control about duplicate binding handling on the module-user side.

#:version list

Specify a version for the module in the form of list, a list of zero or more exact, nonnegative integers. The corresponding #:version option in the use-modules form allows callers to restrict the value of this option in various ways.

#:duplicates list

Tell Guile to handle duplicate bindings for the bindings imported by the current module according to the policy defined by list, a list of symbols. list must contain symbols representing a duplicate binding handling policy chosen among the following:

check Raises an error when a binding is imported from more than one place.

warn Issue a warning when a binding is imported from more than one place and leave the responsibility of actually handling the duplication to the next duplicate binding handler.

replace When a new binding is imported that has the same name as a previously imported binding, then do the following:

1. If the old binding was said to be replacing (via the #:replace option above) and the new binding is not replacing, the keep the old binding.

2. If the old binding was not said to be replacing and the new binding is replacing, then replace the old binding with the new one.

3. If neither the old nor the new binding is replacing, then keep the old one.

warn-override-core

Issue a warning when a core binding is being overwritten and actually override the core binding with the new one. first In case of duplicate bindings, the firstly imported binding is

always the one which is kept.

last In case of duplicate bindings, the lastly imported binding is always the one which is kept.

noop In case of duplicate bindings, leave the responsibility to the next duplicate handler.

If list contains more than one symbol, then the duplicate binding handlers which appear first will be used first when resolving a duplicate binding situation. As mentioned above, some resolution policies may explicitly leave the responsibility of handling the duplication to the next handler in list. If GOOPS has been loaded before the #:duplicates clause is processed, there are additional strategies available for dealing with generic functions. See Section 8.6.3 [Merging Generics], page 728, for more information. The default duplicate binding resolution policy is given by the default-

duplicate-binding-handler procedure, and is
     (replace warn-override-core warn last)

#:pure Create a pure module, that is a module which does not contain any of the standard procedure bindings except for the syntax forms. This is useful if you want to create safe modules, that is modules which do not know anything about dangerous procedures.

[syntax] export variable … Add all variables (which must be symbols or pairs of symbols) to the list of exported bindings of the current module. If variable is a pair, its car gives the name of the variable as seen by the current module and its cdr specifies a name for the binding in the current module’s public interface.

[syntax] define-public . . .
Equivalent to (begin (define foo ...) (export foo)).

[syntax] re-export variable …

Add all variables (which must be symbols or pairs of symbols) to the list of re-exported bindings of the current module. Pairs of symbols are handled as in export. Reexported bindings must be imported by the current module from some other module.

[syntax] export! variable … Like export, but marking the exported variables as replacing. Using a module with replacing bindings will cause any existing bindings to be replaced without issuing any warnings. See the discussion of #:replace above.

6.19.4 Modules and the File System

Typical programs only use a small subset of modules installed on a Guile system. In order to keep startup time down, Guile only loads modules when a program uses them, on demand.

When a program evaluates (use-modules (ice-9 popen)), and the module is not loaded, Guile searches for a conventionally-named file from in the load path.

In this case, loading (ice-9 popen) will eventually cause Guile to run (primitive-

load-path "ice-9/popen"). primitive-load-path will search for a file ice-9/popen

in the %load-path (see Section 6.17.7 [Load Paths], page 368). For each directory in %load-path, Guile will try to find the file name, concatenated with the extensions from %load-extensions. By default, this will cause Guile to stat ice-9/popen.scm, and then ice-9/popen. See Section 6.17.7 [Load Paths], page 368, for more on primitive-load-

path.

If a corresponding compiled .go file is found in the %load-compiled-path or in the fallback path, and is as fresh as the source file, it will be loaded instead of the source file. If no compiled file is found, Guile may try to compile the source file and cache away the resulting .go file. See Section 6.17.5 [Compilation], page 364, for more on compilation.

Once Guile finds a suitable source or compiled file is found, the file will be loaded. If, after loading the file, the module under consideration is still not defined, Guile will signal an error.

For more information on where and how to install Scheme modules, See Section 4.7 [Installing Site Packages], page 55.

6.19.5 R6RS Version References

Guile’s module system includes support for locating modules based on a declared version specifier of the same form as the one described in R6RS (see Section “Library form” in The Revised^6 Report on the Algorithmic Language Scheme). By using the #:version keyword in a define-module form, a module may specify a version as a list of zero or more exact, nonnegative integers.

This version can then be used to locate the module during the module search process. Client modules and callers of the use-modules function may specify constraints on the versions of target modules by providing a version reference, which has one of the following forms:

      (sub-version-reference ...)
      (and version-reference ...)
      (or version-reference ...)
      (not version-reference)

in which sub-version-reference is in turn one of:

      (sub-version)
      (>= sub-version)
      (<= sub-version)
      (and sub-version-reference ...)
      (or sub-version-reference ...)
      (not sub-version-reference)

in which sub-version is an exact, nonnegative integer as above. A version reference matches a declared module version if each element of the version reference matches a corresponding element of the module version, according to the following rules:

• The and sub-form matches a version or version element if every element in the tail of the sub-form matches the specified version or version element.

• The or sub-form matches a version or version element if any element in the tail of the sub-form matches the specified version or version element.

• The not sub-form matches a version or version element if the tail of the sub-form does not match the version or version element.

• The >= sub-form matches a version element if the element is greater than or equal to the sub-version in the tail of the sub-form.

• The <= sub-form matches a version element if the version is less than or equal to the sub-version in the tail of the sub-form.

• A sub-version matches a version element if one is eqv? to the other.

For example, a module declared as:

      (define-module (mylib mymodule) #:version (1 2 0))

would be successfully loaded by any of the following use-modules expressions:

      (use-modules ((mylib mymodule) #:version (1 2 (>= 0))))
      (use-modules ((mylib mymodule) #:version (or (1 2 0) (1 2 1))))
      (use-modules ((mylib mymodule) #:version ((and (>= 1) (not 2)) 2 0)))

6.19.6 R6RS Libraries

In addition to the API described in the previous sections, you also have the option to create modules using the portable library form described in R6RS (see Section “Library form” in The Revised^6 Report on the Algorithmic Language Scheme), and to import libraries created in this format by other programmers. Guile’s R6RS library implementation takes advantage of the flexibility built into the module system by expanding the R6RS library form into a corresponding Guile define-module form that specifies equivalent import and export requirements and includes the same body expressions. The library expression:

       (library (mylib (1 2))
         (import (otherlib (3)))
         (export mybinding))

is equivalent to the module definition:

       (define-module (mylib)
         #:version (1 2)
         #:use-module ((otherlib) #:version (3))
         #:export (mybinding))

Central to the mechanics of R6RS libraries is the concept of import and export levels, which control the visibility of bindings at various phases of a library’s lifecycle — macros necessary to expand forms in the library’s body need to be available at expand time; variables used in the body of a procedure exported by the library must be available at runtime. R6RS specifies the optional for sub-form of an import set specification (see below) as a mechanism by which a library author can indicate that a particular library import should take place at a particular phase with respect to the lifecycle of the importing library.

Guile’s library implementation uses a technique called implicit phasing (first described by Abdulaziz Ghuloum and R. Kent Dybvig), which allows the expander and compiler to automatically determine the necessary visibility of a binding imported from another library. As such, the for sub-form described below is ignored by Guile (but may be required by Schemes in which phasing is explicit).

[Scheme Syntax] library name (export export-spec ...) (import import-spec ...) body

... Defines a new library with the specified name, exports, and imports, and evaluates the specified body expressions in this library’s environment. The library name is a non-empty list of identifiers, optionally ending with a version specification of the form described above (see Section 6.19.3 [Creating Guile Modules], page 383). Each export-spec is the name of a variable defined or imported by the library, or must

take the form (rename (internal-name external-name) ...), where the identifier

internal-name names a variable defined or imported by the library and external-name is the name by which the variable is seen by importing libraries. Each import-spec must be either an import set (see below) or must be of the form

(for import-set import-level ...), where each import-level is one of:
       run
       expand
       (meta level)

where level is an integer. Note that since Guile does not require explicit phase specification, any import-sets found inside of for sub-forms will be “unwrapped” during expansion and processed as if they had been specified directly. Import sets in turn take one of the following forms:

       library-reference
       (library library-reference)
       (only import-set identifier ...)
       (except import-set identifier ...)
       (prefix import-set identifier)
       (rename import-set (internal-identifier external-identifier) ...)

where library-reference is a non-empty list of identifiers ending with an optional version reference (see Section 6.19.5 [R6RS Version References], page 387), and the other sub-forms have the following semantics, defined recursively on nested import-sets:

• The library sub-form is used to specify libraries for import whose names begin with the identifier “library.”

• The only sub-form imports only the specified identifiers from the given importset.

• The except sub-form imports all of the bindings exported by import-set except for those that appear in the specified list of identifiers.

• The prefix sub-form imports all of the bindings exported by import-set, first prefixing them with the specified identifier.

• The rename sub-form imports all of the identifiers exported by import-set. The binding for each internal-identifier among these identifiers is made visible to the importing library as the corresponding external-identifier; all other bindings are imported using the names provided by import-set.

Note that because Guile translates R6RS libraries into module definitions, an import specification may be used to declare a dependency on a native Guile module — although doing so may make your libraries less portable to other Schemes.

[Scheme Syntax] import import-spec ...

Import into the current environment the libraries specified by the given import specifications, where each import-spec takes the same form as in the library form described above.

6.19.7 Variables

Each module has its own hash table, sometimes known as an obarray, that maps the names defined in that module to their corresponding variable objects.

A variable is a box-like object that can hold any Scheme value. It is said to be undefined if its box holds a special Scheme value that denotes undefined-ness (which is different from all other Scheme values, including for example #f); otherwise the variable is defined.

On its own, a variable object is anonymous. A variable is said to be bound when it is associated with a name in some way, usually a symbol in a module obarray. When this happens, the name is said to be bound to the variable, in that module.

(That’s the theory, anyway. In practice, defined-ness and bound-ness sometimes get confused, because Lisp and Scheme implementations have often conflated — or deliberately drawn no distinction between — a name that is unbound and a name that is bound to a variable whose value is undefined. We will try to be clear about the difference and explain any confusion where it is unavoidable.)

Variables do not have a read syntax. Most commonly they are created and bound implicitly by define expressions: a top-level define expression of the form

     (define name value)

creates a variable with initial value value and binds it to the name name in the current module. But they can also be created dynamically by calling one of the constructor procedures

make-variable and make-undefined-variable.
[Scheme Procedure] make-undefined-variable
[C Function] scm_make_undefined_variable ()

Return a variable that is initially unbound.

[Scheme Procedure] make-variable init

[C Function] scm_make_variable (init)

Return a variable initialized to value init.

[Scheme Procedure] variable-bound? var

[C Function] scm_variable_bound_p (var)

Return #t if var is bound to a value, or #f otherwise. Throws an error if var is not a variable object.

[Scheme Procedure] variable-ref var

[C Function] scm_variable_ref (var)

Dereference var and return its value. var must be a variable object; see make-

     variable and make-undefined-variable.

[Scheme Procedure] variable-set! var val [C Function] scm_variable_set_x (var, val)

Set the value of the variable var to val. var must be a variable object, val can be any value. Return an unspecified value.

[Scheme Procedure] variable-unset! var

[C Function] scm_variable_unset_x (var)

Unset the value of the variable var, leaving var unbound.

[Scheme Procedure] variable? obj [C Function] scm_variable_p (obj)

Return #t if obj is a variable object, else return #f.

6.19.8 Module System Reflection

The previous sections have described a declarative view of the module system. You can also work with it programmatically by accessing and modifying various parts of the Scheme objects that Guile uses to implement the module system.

At any time, there is a current module. This module is the one where a top-level define and similar syntax will add new bindings. You can find other module objects with

resolve-module, for example.

These module objects can be used as the second argument to eval.

[Scheme Procedure] current-module

[C Function] scm_current_module ()

Return the current module object.

[Scheme Procedure] set-current-module module

[C Function] scm_set_current_module (module)

Set the current module to module and return the previous current module.

[Scheme Procedure] save-module-excursion thunk

Call thunk within a dynamic-wind such that the module that is current at invocation time is restored when thunk’s dynamic extent is left (see Section 6.13.10 [Dynamic Wind], page 309). More precisely, if thunk escapes non-locally, the current module (at the time of escape) is saved, and the original current module (at the time thunk’s dynamic extent was last entered) is restored. If thunk’s dynamic extent is re-entered, then the current module is saved, and the previously saved inner module is set current again.

[Scheme Procedure] resolve-module name [autoload=#t] [version=#f]

[#:ensure=#t]

[C Function] scm_resolve_module (name)

Find the module named name and return it. When it has not already been defined and autoload is true, try to auto-load it. When it can’t be found that way either, create an empty module if ensure is true, otherwise return #f. If version is true, ensure that the resulting module is compatible with the given version reference (see Section 6.19.5 [R6RS Version References], page 387). The name is a list of symbols.

[Scheme Procedure] resolve-interface name [#:select=#f] [#:hide=’()]

[#:prefix=#f] [#:renamer=#f] [#:version=#f] Find the module named name as with resolve-module and return its interface. The interface of a module is also a module object, but it contains only the exported bindings.

[Scheme Procedure] module-uses module

Return a list of the interfaces used by module.

[Scheme Procedure] module-use! module interface Add interface to the front of the use-list of module. Both arguments should be module objects, and interface should very likely be a module returned by resolve-interface.

[Scheme Procedure] reload-module module

Revisit the source file that corresponds to module. Raises an error if no source file is associated with the given module.

As mentioned in the previous section, modules contain a mapping between identifiers (as symbols) and storage locations (as variables). Guile defines a number of procedures to allow access to this mapping. If you are programming in C, Section 6.19.9 [Accessing Modules from C], page 392.

[Scheme Procedure] module-variable module name

Return the variable bound to name (a symbol) in module, or #f if name is unbound.

[Scheme Procedure] module-add! module name var Define a new binding between name (a symbol) and var (a variable) in module.

[Scheme Procedure] module-ref module name

Look up the value bound to name in module. Like module-variable, but also does a variable-ref on the resulting variable, raising an error if name is unbound.

[Scheme Procedure] module-define! module name value Locally bind name to value in module. If name was already locally bound in module, i.e., defined locally and not by an imported module, the value stored in the existing variable will be updated. Otherwise, a new variable will be added to the module, via

module-add!.

[Scheme Procedure] module-set! module name value Update the binding of name in module to value, raising an error if name is not already bound in module.

There are many other reflective procedures available in the default environment. If you find yourself using one of them, please contact the Guile developers so that we can commit to stability for that interface.

6.19.9 Accessing Modules from C

The last sections have described how modules are used in Scheme code, which is the recommended way of creating and accessing modules. You can also work with modules from C, but it is more cumbersome.

The following procedures are available.

[C Function] SCM scm_c_call_with_current_module (SCM module, SCM

(*func)(void *), void *data) Call func and make module the current module during the call. The argument data is passed to func. The return value of scm_c_call_with_current_module is the return value of func.

[C Function] SCM scm_public_variable (SCM module_name, SCM name)
[C Function] SCM scm_c_public_variable (const char *module_name, const char
   *name)

Find a the variable bound to the symbol name in the public interface of the module named module name.

module name should be a list of symbols, when represented as a Scheme object, or a space-separated string, in the const char * case. See scm_c_define_module below, for more examples.

Signals an error if no module was found with the given name. If name is not bound in the module, just returns #f.

[C Function] SCM scm_private_variable (SCM module_name, SCM name)
[C Function] SCM scm_c_private_variable (const char *module_name, const

char *name) Like scm_public_variable, but looks in the internals of the module named module name instead of the public interface. Logically, these procedures should only be called on modules you write.

[C Function] SCM scm_public_lookup (SCM module_name, SCM name)
[C Function] SCM scm_c_public_lookup (const char *module_name, const char
   *name)
[C Function] SCM scm_private_lookup (SCM module_name, SCM name)
[C Function] SCM scm_c_private_lookup (const char *module_name, const char
   *name)
Like scm_public_variable or scm_private_variable, but if the name is not bound

in the module, signals an error. Returns a variable, always.

     static SCM eval_string_var;
     /* NOTE: It is important that the call to 'my_init'
        happens-before all calls to 'my_eval_string'. */
     void my_init (void)
     {
       eval_string_var = scm_c_public_lookup ("ice-9 eval-string",
                                              "eval-string");
     }
     SCM my_eval_string (SCM str)
     {
       return scm_call_1 (scm_variable_ref (eval_string_var), str);
     }
[C Function] SCM scm_public_ref (SCM module_name, SCM name)
[C Function] SCM scm_c_public_ref (const char *module_name, const char *name)
[C Function] SCM scm_private_ref (SCM module_name, SCM name)

[C Function] SCM scm_c_private_ref (const char *module_name, const char

   *name)

Like scm_public_lookup or scm_private_lookup, but additionally dereferences the variable. If the variable object is unbound, signals an error. Returns the value bound to name in module name.

In addition, there are a number of other lookup-related procedures. We suggest that you use the scm_public_and scm_private_family of procedures instead, if possible.

[C Function] SCM scm_c_lookup (const char *name)

Return the variable bound to the symbol indicated by name in the current module. If there is no such binding or the symbol is not bound to a variable, signal an error.

[C Function] SCM scm_lookup (SCM name)

Like scm_c_lookup, but the symbol is specified directly.

[C Function] SCM scm_c_module_lookup (SCM module, const char *name)
[C Function] SCM scm_module_lookup (SCM module, SCM name)

Like scm_c_lookup and scm_lookup, but the specified module is used instead of the current one.

[C Function] SCM scm_module_variable (SCM module, SCM name)

Like scm_module_lookup, but if the binding does not exist, just returns #f instead of raising an error.

To define a value, use scm_define:

[C Function] SCM scm_c_define (const char *name, SCM val)

Bind the symbol indicated by name to a variable in the current module and set that variable to val. When name is already bound to a variable, use that. Else create a new variable.

[C Function] SCM scm_define (SCM name, SCM val)

Like scm_c_define, but the symbol is specified directly.

[C Function] SCM scm_c_module_define (SCM module, const char *name, SCM
   val)
[C Function] SCM scm_module_define (SCM module, SCM name, SCM val)

Like scm_c_define and scm_define, but the specified module is used instead of the current one.

In some rare cases, you may need to access the variable that scm_module_define would have accessed, without changing the binding of the existing variable, if one is present. In

that case, use scm_module_ensure_local_variable:
[C Function] SCM scm_module_ensure_local_variable (SCM module, SCM
   sym)

Like scm_module_define, but if the sym is already locally bound in that module, the variable’s existing binding is not reset. Returns a variable.

[C Function] SCM scm_module_reverse_lookup (SCM module, SCM variable)

Find the symbol that is bound to variable in module. When no such binding is found, return #f.

[C Function] SCM scm_c_define_module (const char *name, void (*init)(void *),

void *data) Define a new module named name and make it current while init is called, passing it data. Return the module. The parameter name is a string with the symbols that make up the module name, separated by spaces. For example, ‘"foo bar"’ names the module ‘(foo bar)’. When there already exists a module named name, it is used unchanged, otherwise, an empty module is created.

[C Function] SCM scm_c_resolve_module (const char *name)

Find the module name name and return it. When it has not already been defined, try to auto-load it. When it can’t be found that way either, create an empty module. The name is interpreted as for scm_c_define_module.

[C Function] SCM scm_c_use_module (const char *name)

Add the module named name to the uses list of the current module, as with (use-

modules name). The name is interpreted as for scm_c_define_module.

[C Function] SCM scm_c_export (const char *name, ...)

Add the bindings designated by name, ... to the public interface of the current module. The list of names is terminated by NULL.

6.19.10 provide and require

Aubrey Jaffer, mostly to support his portable Scheme library SLIB, implemented a provide/require mechanism for many Scheme implementations. Library files in SLIB provide a feature, and when user programs require that feature, the library file is loaded in.

For example, the file random.scm in the SLIB package contains the line

(provide 'random)

so to use its procedures, a user would type

(require 'random)

and they would magically become available, but still have the same names! So this method is nice, but not as good as a full-featured module system.

When SLIB is used with Guile, provide and require can be used to access its facilities.

6.19.11 Environments

Scheme, as defined in R5RS, does not have a full module system. However it does define the concept of a top-level environment. Such an environment maps identifiers (symbols) to Scheme objects such as procedures and lists: Section 3.4 [About Closure], page 26. In other words, it implements a set of bindings.

Environments in R5RS can be passed as the second argument to eval (see Section 6.17.4 [Fly Evaluation], page 362). Three procedures are defined to return environments:

scheme-report-environment, null-environment and interaction-environment (see

Section 6.17.4 [Fly Evaluation], page 362).

In addition, in Guile any module can be used as an R5RS environment, i.e., passed as the second argument to eval.

Note: the following two procedures are available only when the (ice-9 r5rs) module is loaded:

     (use-modules (ice-9 r5rs))
[Scheme Procedure] scheme-report-environment version

[Scheme Procedure] null-environment version

version must be the exact integer ‘5’, corresponding to revision 5 of the Scheme report (the Revised^5 Report on Scheme). scheme-report-environment returns a specifier for an environment that is empty except for all bindings defined in the report that are either required or both optional and supported by the implementation. nullenvironment returns a specifier for an environment that is empty except for the (syntactic) bindings for all syntactic keywords defined in the report that are either required or both optional and supported by the implementation. Currently Guile does not support values of version for other revisions of the report. The effect of assigning (through the use of eval) a variable bound in a schemereport-environment (for example car) is unspecified. Currently the environments specified by scheme-report-environment are not immutable in Guile.

6.20 Foreign Function Interface

The more one hacks in Scheme, the more one realizes that there are actually two computational worlds: one which is warm and alive, that land of parentheses, and one cold and dead, the land of C and its ilk.

But yet we as programmers live in both worlds, and Guile itself is half implemented in C. So it is that Guile’s living half pays respect to its dead counterpart, via a spectrum of interfaces to C ranging from dynamic loading of Scheme primitives to dynamic binding of stock C library procedures.

6.20.1 Foreign Libraries

Most modern Unices have something called shared libraries. This ordinarily means that they have the capability to share the executable image of a library between several running programs to save memory and disk space. But generally, shared libraries give a lot of additional flexibility compared to the traditional static libraries. In fact, calling them ‘dynamic’ libraries is as correct as calling them ‘shared’.

Shared libraries really give you a lot of flexibility in addition to the memory and disk space savings. When you link a program against a shared library, that library is not closely incorporated into the final executable. Instead, the executable of your program only contains enough information to find the needed shared libraries when the program is actually run. Only then, when the program is starting, is the final step of the linking process performed. This means that you need not recompile all programs when you install a new, only slightly modified version of a shared library. The programs will pick up the changes automatically the next time they are run.

Now, when all the necessary machinery is there to perform part of the linking at runtime, why not take the next step and allow the programmer to explicitly take advantage of it from within his program? Of course, many operating systems that support shared libraries do just that, and chances are that Guile will allow you to access this feature from within your Scheme programs. As you might have guessed already, this feature is called dynamic linking.^{14}

We titled this section “foreign libraries” because although the name “foreign” doesn’t leak into the API, the world of C really is foreign to Scheme – and that estrangement extends to components of foreign libraries as well, as we see in future sections.

[Scheme Procedure] dynamic-link [library] [C Function] scm_dynamic_link (library)

Find the shared library denoted by library (a string) and link it into the running Guile application. When everything works out, return a Scheme object suitable for representing the linked object file. Otherwise an error is thrown. How object files are searched is system dependent.

Normally, library is just the name of some shared library file that will be searched for in the places where shared libraries usually reside, such as in /usr/lib and

/usr/local/lib.

library should not contain an extension such as .so. The correct file name extension for the host operating system is provided automatically, according to libltdl’s rules (see Section “lt_dlopenext” in Shared Library Support for GNU).

When library is omitted, a global symbol handle is returned. This handle provides access to the symbols available to the program at run-time, including those exported by the program itself and the shared libraries already loaded.

[Scheme Procedure] dynamic-object? obj

[C Function] scm_dynamic_object_p (obj)

Return #t if obj is a dynamic library handle, or #f otherwise.

[Scheme Procedure] dynamic-unlink dobj

[C Function] scm_dynamic_unlink (dobj)

Unlink the indicated object file from the application. The argument dobj must have been obtained by a call to dynamic-link. After dynamic-unlink has been called on dobj, its content is no longer accessible.

(define libgl-obj (dynamic-link "libGL"))
libgl-obj
⇒ #<dynamic-object "libGL">
(dynamic-unlink libGL-obj)
libGL-obj
⇒ #<dynamic-object "libGL" (unlinked)>

As you can see, after calling dynamic-unlink on a dynamically linked library, it is marked as ‘(unlinked)’ and you are no longer able to use it with dynamic-call, etc.

14 Some people also refer to the final linking stage at program startup as ‘dynamic linking’, so if you want to

make yourself perfectly clear, it is probably best to use the more technical term dlopening, as suggested by Gordon Matzigkeit in his libtool documentation.

Whether the library is really removed from you program is system-dependent and will generally not happen when some other parts of your program still use it.

When dynamic linking is disabled or not supported on your system, the above functions throw errors, but they are still available.

6.20.2 Foreign Functions

The most natural thing to do with a dynamic library is to grovel around in it for a function pointer: a foreign function. dynamic-func exists for that purpose.

[Scheme Procedure] dynamic-func name dobj [C Function] scm_dynamic_func (name, dobj)

Return a “handle” for the func name in the shared object referred to by dobj. The handle can be passed to dynamic-call to actually call the function.

Regardless whether your C compiler prepends an underscore ‘_’ to the global names in a program, you should not include this underscore in name since it will be added automatically when necessary.

Guile has static support for calling functions with no arguments, dynamic-call.

[Scheme Procedure] dynamic-call func dobj [C Function] scm_dynamic_call (func, dobj)

Call the C function indicated by func and dobj. The function is passed no arguments and its return value is ignored. When function is something returned by dynamicfunc, call that function and ignore dobj. When func is a string , look it up in dynobj; this is equivalent to

            (dynamic-call (dynamic-func func dobj) #f)

dynamic-call is not very powerful. It is mostly intended to be used for calling specially written initialization functions that will then add new primitives to Guile. For example, we do not expect that you will dynamically link libX11 with dynamic-link and then construct a beautiful graphical user interface just by using dynamic-call. Instead, the usual way would be to write a special Guile-to-X11 glue library that has intimate knowledge about both Guile and X11 and does whatever is necessary to make them inter-operate smoothly. This glue library could then be dynamically linked into a vanilla Guile interpreter and activated by calling its initialization function. That function would add all the new types and primitives to the Guile interpreter that it has to offer.

(There is actually another, better option: simply to create a libX11 wrapper in Scheme via the dynamic FFI. See Section 6.20.6 [Dynamic FFI], page 407, for more information.)

Given some set of C extensions to Guile, the next logical step is to integrate these glue libraries into the module system of Guile so that you can load new primitives into a running system just as you can load new Scheme code.

[Scheme Procedure] load-extension lib init [C Function] scm_load_extension (lib, init)

Load and initialize the extension designated by LIB and INIT. When there is no pre-registered function for LIB/INIT, this is equivalent to

          (dynamic-call INIT (dynamic-link LIB))

When there is a pre-registered function, that function is called instead.

Normally, there is no pre-registered function. This option exists only for situations where dynamic linking is unavailable or unwanted. In that case, you would statically link your program with the desired library, and register its init function right after Guile has been initialized. As for dynamic-link, lib should not contain any suffix such as .so (see Section 6.20.1 [Foreign Libraries], page 396). It should also not contain any directory components. Libraries that implement Guile Extensions should be put into the normal locations for shared libraries. We recommend to use the naming convention libguile-bla-blum for a extension related to a module (bla blum). The normal way for a extension to be used is to write a small Scheme file that defines a module, and to load the extension into this module. When the module is auto-loaded, the extension is loaded as well. For example,

          (define-module (bla blum))
          (load-extension "libguile-bla-blum" "bla_init_blum")

6.20.3 C Extensions

The most interesting application of dynamically linked libraries is probably to use them for providing compiled code modules to Scheme programs. As much fun as programming in Scheme is, every now and then comes the need to write some low-level C stuff to make Scheme even more fun.

Not only can you put these new primitives into their own module (see the previous section), you can even put them into a shared library that is only then linked to your running Guile image when it is actually needed.

An example will hopefully make everything clear. Suppose we want to make the Bessel functions of the C library available to Scheme in the module ‘(math bessel)’. First we need to write the appropriate glue code to convert the arguments and return values of the functions from Scheme to C and back. Additionally, we need a function that will add them to the set of Guile primitives. Because this is just an example, we will only implement this for the j0 function.

      #include <math.h>
      #include <libguile.h>
      SCM
      j0_wrapper (SCM x)
      {
        return scm_from_double (j0 (scm_to_double (x, "j0")));
      }
      void
      init_math_bessel ()
      {
        scm_c_define_gsubr ("j0", 1, 0, 0, j0_wrapper);
      }

We can already try to bring this into action by manually calling the low level functions for performing dynamic linking. The C source file needs to be compiled into a shared library.

Here is how to do it on GNU/Linux, please refer to the libtool documentation for how to create dynamically linkable libraries portably.

      gcc -shared -o libbessel.so -fPIC bessel.c

Now fire up Guile:

     (define bessel-lib (dynamic-link "./libbessel.so"))
     (dynamic-call "init_math_bessel" bessel-lib)
     (j0 2)
     ⇒ 0.223890779141236

The filename ./libbessel.so should be pointing to the shared library produced with the gcc command above, of course. The second line of the Guile interaction will call the init_math_bessel function which in turn will register the C function j0_wrapper with the Guile interpreter under the name j0. This function becomes immediately available and we can call it from Scheme.

Fun, isn’t it? But we are only half way there. This is what apropos has to say about

j0:
      (apropos "j0")
       ⊣ (guile-user): j0     #<primitive-procedure j0>

As you can see, j0 is contained in the root module, where all the other Guile primitives like display, etc live. In general, a primitive is put into whatever module is the current module at the time scm_c_define_gsubr is called.

A compiled module should have a specially named module init function. Guile knows about this special name and will call that function automatically after having linked in the shared library. For our example, we replace init_math_bessel with the following code in

bessel.c:
      void
      init_math_bessel (void *unused)
      {
        scm_c_define_gsubr ("j0", 1, 0, 0, j0_wrapper);
        scm_c_export ("j0", NULL);
      }
      void
      scm_init_math_bessel_module ()
      {
        scm_c_define_module ("math bessel", init_math_bessel, NULL);
      }

The general pattern for the name of a module init function is: ‘scm_init_’, followed by the name of the module where the individual hierarchical components are concatenated with underscores, followed by ‘_module’.

After libbessel.so has been rebuilt, we need to place the shared library into the right place.

Once the module has been correctly installed, it should be possible to use it like this:

      guile> (load-extension "./libbessel.so" "scm_init_math_bessel_module")
      guile> (use-modules (math bessel))
      guile> (j0 2)
      0.223890779141236
      guile> (apropos "j0")
       ⊣ (math bessel): j0      #<primitive-procedure j0>

That’s it!

6.20.4 Modules and Extensions

The new primitives that you add to Guile with scm_c_define_gsubr (see Section 6.9.2 [Primitive Procedures], page 245) or with any of the other mechanisms are placed into the module that is current when the scm_c_define_gsubr is executed. Extensions loaded from the REPL, for example, will be placed into the (guile-user) module, if the REPL module was not changed.

To define C primitives within a specific module, the simplest way is:

     (define-module (foo bar))
     (load-extension "foobar-c-code" "foo_bar_init")
   When loaded with (use-modules (foo bar)), the load-extension call looks for the

foobar-c-code.so (etc) object file in Guile’s extensiondir, which is usually a subdirectory of the libdir. For example, if your libdir is /usr/lib, the extensiondir for the Guile

2.0.x series will be /usr/lib/guile/2.0/.

The extension path includes the major and minor version of Guile (the “effective version”), because Guile guarantees compatibility within a given effective version. This allows you to install different versions of the same extension for different versions of Guile.

If the extension is not found in the extensiondir, Guile will also search the standard system locations, such as /usr/lib or /usr/local/lib. It is preferable, however, to keep your extension out of the system library path, to prevent unintended interference with other dynamically-linked C libraries.

If someone installs your module to a non-standard location then the object file won’t be found. You can address this by inserting the install location in the foo/bar.scm file. This is convenient for the user and also guarantees the intended object is read, even if stray older or newer versions are in the loader’s path.

The usual way to specify an install location is with a prefix at the configure stage,

for instance ‘./configure prefix=/opt’ results in library files as say /opt/lib/foobar-

c-code.so. When using Autoconf (see Section “Introduction” in The GNU Autoconf Manual), the library location is in a libdir variable. Its value is intended to be expanded by make, and can by substituted into a source file like foo.scm.in

     (define-module (foo bar))
     (load-extension "XXextensiondirXX/foobar-c-code" "foo_bar_init")

with the following in a Makefile, using sed (see Section “Introduction” in SED),

     foo.scm: foo.scm.in
             sed 's|XXextensiondirXX|$(libdir)/guile/2.0|' <foo.scm.in >foo.scm

The actual pattern XXextensiondirXX is arbitrary, it’s only something which doesn’t otherwise occur. If several modules need the value, it can be easier to create one foo/config.scm with a define of the extensiondir location, and use that as required.

     (define-module (foo config))
     (define-public foo-config-extensiondir "XXextensiondirXX"")

Such a file might have other locations too, for instance a data directory for auxiliary files, or localedir if the module has its own gettext message catalogue (see Section 6.24 [Internationalization], page 436).

It will be noted all of the above requires that the Scheme code to be found in %load-path (see Section 6.17.7 [Load Paths], page 368). Presently it’s left up to the system administrator or each user to augment that path when installing Guile modules in non-default locations. But having reached the Scheme code, that code should take care of hitting any of its own private files etc.

6.20.5 Foreign Pointers

The previous sections have shown how Guile can be extended at runtime by loading compiled C extensions. This approach is all well and good, but wouldn’t it be nice if we didn’t have to write any C at all? This section takes up the problem of accessing C values from Scheme, and the next discusses C functions.

6.20.5.1 Foreign Types

The first impedance mismatch that one sees between C and Scheme is that in C, the storage locations (variables) are typed, but in Scheme types are associated with values, not variables. See Section 3.1.2 [Values and Variables], page 15.

So when describing a C function or a C structure so that it can be accessed from Scheme, the data types of the parameters or fields must be passed explicitly.

These “C type values” may be constructed using the constants and procedures from the (system foreign) module, which may be loaded like this:

     (use-modules (system foreign))

(system foreign) exports a number of values expressing the basic C types:

[Scheme Variable] int8 [Scheme Variable] uint8 [Scheme Variable] uint16 [Scheme Variable] int16 [Scheme Variable] uint32 [Scheme Variable] int32 [Scheme Variable] uint64 [Scheme Variable] int64 [Scheme Variable] float [Scheme Variable] double

These values represent the C numeric types of the specified sizes and signednesses.

In addition there are some convenience bindings for indicating types of platformdependent size:

[Scheme Variable] int [Scheme Variable] unsigned-int [Scheme Variable] long [Scheme Variable] unsigned-long [Scheme Variable] size_t [Scheme Variable] ssize_t [Scheme Variable] ptrdiff_t

Values exported by the (system foreign) module, representing C numeric types. For example, long may be equal? to int64 on a 64-bit platform.

[Scheme Variable] void

The void type. It can be used as the first argument to pointer->procedure to wrap a C function that returns nothing.

In addition, the symbol * is used by convention to denote pointer types. Procedures detailed in the following sections, such as pointer->procedure, accept it as a type descriptor.

6.20.5.2 Foreign Variables

Pointers to variables in the current address space may be looked up dynamically using

dynamic-pointer.

[Scheme Procedure] dynamic-pointer name dobj [C Function] scm_dynamic_pointer (name, dobj)

Return a “wrapped pointer” for the symbol name in the shared object referred to by dobj. The returned pointer points to a C object. Regardless whether your C compiler prepends an underscore ‘_’ to the global names in a program, you should not include this underscore in name since it will be added automatically when necessary.

For example, currently Guile has a variable, scm_numptob, as part of its API. It is declared as a C long. So, to create a handle pointing to that foreign value, we do:

(use-modules (system foreign))
(define numptob (dynamic-pointer "scm_numptob" (dynamic-link)))
numptob
⇒ #<pointer 0x7fb35b1b4688>

(The next section discusses ways to dereference pointers.) A value returned by dynamic-pointer is a Scheme wrapper for a C pointer.

[Scheme Procedure] pointer-address pointer [C Function] scm_pointer_address (pointer)

Return the numerical value of pointer.

     (pointer-address numptob)
     ⇒ 139984413364296 ; YMMV

[Scheme Procedure] make-pointer address [finalizer] Return a foreign pointer object pointing to address. If finalizer is passed, it should be a pointer to a one-argument C function that will be called when the pointer object becomes unreachable.

[Scheme Procedure] pointer? obj

Return #t if obj is a pointer object, #f otherwise.

[Scheme Variable] %null-pointer

A foreign pointer whose value is 0.

[Scheme Procedure] null-pointer? pointer

Return #t if pointer is the null pointer, #f otherwise.

For the purpose of passing SCM values directly to foreign functions, and allowing them to return SCM values, Guile also supports some unsafe casting operators.

[Scheme Procedure] scm->pointer scm

Return a foreign pointer object with the object-address of scm.

[Scheme Procedure] pointer->scm pointer

Unsafely cast pointer to a Scheme object. Cross your fingers!

Sometimes you want to give C extensions access to the dynamic FFI. At that point, the names get confusing, because “pointer” can refer to a SCM object that wraps a pointer, or to a void* value. We will try to use “pointer object” to refer to Scheme objects, and “pointer value” to refer to void * values.

[C Function] SCM scm_from_pointer (void *ptr, void (*finalizer) (void*))

Create a pointer object from a pointer value. If finalizer is non-null, Guile arranges to call it on the pointer value at some point after the pointer object becomes collectable.

[C Function] void* scm_to_pointer (SCM obj)

Unpack the pointer value from a pointer object.

6.20.5.3 Void Pointers and Byte Access

Wrapped pointers are untyped, so they are essentially equivalent to C void pointers. As in C, the memory region pointed to by a pointer can be accessed at the byte level. This is achieved using bytevectors (see Section 6.6.6 [Bytevectors], page 163). The (rnrs bytevectors) module contains procedures that can be used to convert byte sequences to Scheme objects such as strings, floating point numbers, or integers.

[Scheme Procedure] pointer->bytevector pointer len [offset [uvec type]]

[C Function] scm_pointer_to_bytevector (pointer, len, offset, uvec type)

Return a bytevector aliasing the len bytes pointed to by pointer. The user may specify an alternate default interpretation for the memory by passing the uvec type argument, to indicate that the memory is an array of elements of that type. uvec type should be something that array-type would return, like f32 or s16. When offset is passed, it specifies the offset in bytes relative to pointer of the memory region aliased by the returned bytevector. Mutating the returned bytevector mutates the memory pointed to by pointer, so buckle your seatbelts.

[Scheme Procedure] bytevector->pointer bv [offset]

[C Function] scm_bytevector_to_pointer (bv, offset)

Return a pointer pointer aliasing the memory pointed to by bv or offset bytes after bv when offset is passed.

In addition to these primitives, convenience procedures are available:

[Scheme Procedure] dereference-pointer pointer

Assuming pointer points to a memory region that holds a pointer, return this pointer.

[Scheme Procedure] string->pointer string [encoding]

Return a foreign pointer to a nul-terminated copy of string in the given encoding, defaulting to the current locale encoding. The C string is freed when the returned foreign pointer becomes unreachable. This is the Scheme equivalent of scm_to_stringn.

[Scheme Procedure] pointer->string pointer [length] [encoding] Return the string representing the C string pointed to by pointer. If length is omitted or -1, the string is assumed to be nul-terminated. Otherwise length is the number of bytes in memory pointed to by pointer. The C string is assumed to be in the given encoding, defaulting to the current locale encoding. This is the Scheme equivalent of scm_from_stringn.

Most object-oriented C libraries use pointers to specific data structures to identify objects. It is useful in such cases to reify the different pointer types as disjoint Scheme types.

The define-wrapped-pointer-type macro simplifies this.

[Scheme Syntax] define-wrapped-pointer-type type-name pred wrap unwrap

print Define helper procedures to wrap pointer objects into Scheme objects with a disjoint type. Specifically, this macro defines:

• pred, a predicate for the new Scheme type;

• wrap, a procedure that takes a pointer object and returns an object that satisfies pred;

• unwrap, which does the reverse.

wrap preserves pointer identity, for two pointer objects p1 and p2 that are equal?,

(eq? (wrap p1) (wrap p2)) ⇒ #t.

Finally, print should name a user-defined procedure to print such objects. The procedure is passed the wrapped object and a port to write to. For example, assume we are wrapping a C library that defines a type, bottle_t, and functions that can be passed bottle_t * pointers to manipulate them. We could write:

     (define-wrapped-pointer-type bottle
       bottle?
       wrap-bottle unwrap-bottle
       (lambda (b p)
         (format p "#<bottle of ~a ~x>"
                 (bottle-contents b)
                 (pointer-address (unwrap-bottle b)))))
     (define grab-bottle
       ;; Wrapper for `bottle_t *grab (void)'.
       (let ((grab (pointer->procedure '*
                                       (dynamic-func "grab_bottle" libbottle)
                                       '())))
         (lambda ()

                "Return a new bottle."
                (wrap-bottle (grab)))))
          (define bottle-contents
            ;; Wrapper for `const char *bottle_contents (bottle_t *)'.
            (let ((contents (pointer->procedure '*
                                                (dynamic-func "bottle_contents"
                                                               libbottle)
                                                '(*))))
              (lambda (b)
                "Return the contents of B."
                (pointer->string (contents (unwrap-bottle b))))))
          (write (grab-bottle))
          ⇒ #<bottle of Ch^ateau Haut-Brion 803d36>

In this example, grab-bottle is guaranteed to return a genuine bottle object satisfying bottle?. Likewise, bottle-contents errors out when its argument is not a genuine bottle object.

Going back to the scm_numptob example above, here is how we can read its value as a C long integer:

     (use-modules (rnrs bytevectors))
     (bytevector-uint-ref (pointer->bytevector numptob (sizeof long))
                          0 (native-endianness)
                          (sizeof long))
     ⇒ 8

If we wanted to corrupt Guile’s internal state, we could set scm_numptob to another value; but we shouldn’t, because that variable is not meant to be set. Indeed this point applies more widely: the C API is a dangerous place to be. Not only might setting a value crash your program, simply accessing the data pointed to by a dangling pointer or similar can prove equally disastrous.

6.20.5.4 Foreign Structs

Finally, one last note on foreign values before moving on to actually calling foreign functions. Sometimes you need to deal with C structs, which requires interpreting each element of the struct according to the its type, offset, and alignment. Guile has some primitives to support this.

[Scheme Procedure] sizeof type [C Function] scm_sizeof (type)

Return the size of type, in bytes.

type should be a valid C type, like int. Alternately type may be the symbol *, in which case the size of a pointer is returned. type may also be a list of types, in which case the size of a struct with ABI-conventional packing is returned.

[Scheme Procedure] alignof type [C Function] scm_alignof (type)

Return the alignment of type, in bytes. type should be a valid C type, like int. Alternately type may be the symbol *, in which case the alignment of a pointer is returned. type may also be a list of types, in which case the alignment of a struct with ABI-conventional packing is returned.

Guile also provides some convenience methods to pack and unpack foreign pointers wrapping C structs.

[Scheme Procedure] make-c-struct types vals

Create a foreign pointer to a C struct containing vals with types types. vals and types should be lists of the same length.

[Scheme Procedure] parse-c-struct foreign types

Parse a foreign pointer to a C struct, returning a list of values. types should be a list of C types.

For example, to create and parse the equivalent of a struct { int64_t a; uint8_t b;

}:
(parse-c-struct (make-c-struct (list int64 uint8)
                               (list 300 43))
                (list int64 uint8))
⇒ (300 43)

As yet, Guile only has convenience routines to support conventionally-packed structs.

But given the bytevector->pointer and pointer->bytevector routines, one can create

and parse tightly packed structs and unions by hand. See the code for (system foreign) for details.

6.20.6 Dynamic FFI

Of course, the land of C is not all nouns and no verbs: there are functions too, and Guile allows you to call them.

[Scheme Procedure] pointer->procedure return type func ptr arg types [C Procedure] scm_pointer_to_procedure (return type, func ptr, arg types)

Make a foreign function. Given the foreign void pointer func ptr, its argument and return types arg types and return type, return a procedure that will pass arguments to the foreign function and return appropriate values. arg types should be a list of foreign types. return_type should be a foreign type. See Section 6.20.5.1 [Foreign Types], page 402, for more information on foreign types.

Here is a better definition of (math bessel):

(define-module (math bessel)
  #:use-module (system foreign)
  #:export (j0))

(define libm (dynamic-link "libm"))
(define j0
  (pointer->procedure double
                      (dynamic-func "j0" libm)
                      (list double)))

That’s it! No C at all. Numeric arguments and return values from foreign functions are represented as Scheme

values. For example, j0 in the above example takes a Scheme number as its argument, and returns a Scheme number.

Pointers may be passed to and returned from foreign functions as well. In that case the type of the argument or return value should be the symbol *, indicating a pointer. For example, the following code makes memcpy available to Scheme:

(define memcpy
  (let ((this (dynamic-link)))
    (pointer->procedure '*
                        (dynamic-func "memcpy" this)
                        (list '* '* size_t))))

To invoke memcpy, one must pass it foreign pointers:

(use-modules (rnrs bytevectors))
(define src-bits
  (u8-list->bytevector '(0 1 2 3 4 5 6 7)))
(define src
  (bytevector->pointer src-bits))
(define dest
  (bytevector->pointer (make-bytevector 16 0)))
(memcpy dest src (bytevector-length src-bits))
(bytevector->u8-list (pointer->bytevector dest 16))
⇒ (0 1 2 3 4 5 6 7 0 0 0 0 0 0 0 0)

One may also pass structs as values, passing structs as foreign pointers. See Section 6.20.5.4 [Foreign Structs], page 406, for more information on how to express struct types and struct values.

“Out” arguments are passed as foreign pointers. The memory pointed to by the foreign pointer is mutated in place.

;; struct timeval {
;;      time_t      tv_sec;     /* seconds */
;;      suseconds_t tv_usec;    /* microseconds */
;; };
;; assuming fields are of type "long"
(define gettimeofday
  (let ((f (pointer->procedure

                 int
                 (dynamic-func "gettimeofday" (dynamic-link))
                 (list '* '*)))
             (tv-type (list long long)))
         (lambda ()
           (let* ((timeval (make-c-struct tv-type (list 0 0)))
                  (ret (f timeval %null-pointer)))
             (if (zero? ret)
                 (apply values (parse-c-struct timeval tv-type))
                 (error "gettimeofday returned an error" ret))))))
     (gettimeofday)
     ⇒ 1270587589
     ⇒ 499553

As you can see, this interface to foreign functions is at a very low, somewhat dangerous level^{15}.

The FFI can also work in the opposite direction: making Scheme procedures callable from C. This makes it possible to use Scheme procedures as “callbacks” expected by C function.

[Scheme Procedure] procedure->pointer return-type proc arg-types [C Function] scm_procedure_to_pointer (return type, proc, arg types)

Return a pointer to a C function of type return-type taking arguments of types argtypes (a list) and behaving as a proxy to procedure proc. Thus proc’s arity, supported argument types, and return type should match return-type and arg-types.

As an example, here’s how the C library’s qsort array sorting function can be made accessible to Scheme (see Section “Array Sort Function” in The GNU C Library Reference Manual):

     (define qsort!
       (let ((qsort (pointer->procedure void
                                        (dynamic-func "qsort"
                                                      (dynamic-link))
                                        (list '* size_t size_t '*))))
         (lambda (bv compare)
           ;; Sort bytevector BV in-place according to comparison
           ;; procedure COMPARE.
           (let ((ptr (procedure->pointer int
                                          (lambda (x y)
                                            ;; X and Y are pointers so,
                                            ;; for convenience, dereference
                                            ;; them before calling COMPARE.
                                            (compare (dereference-uint8* x)
                                                     (dereference-uint8* y)))
                                          (list '* '*))))

15 A contribution to Guile in the form of a high-level FFI would be most welcome.

        (qsort (bytevector->pointer bv)
               (bytevector-length bv) 1 ;; we're sorting bytes
               ptr)))))
(define (dereference-uint8* ptr)
  ;; Helper function: dereference the byte pointed to by PTR.
  (let ((b (pointer->bytevector ptr 1)))
    (bytevector-u8-ref b 0)))
(define bv
  ;; An unsorted array of bytes.
  (u8-list->bytevector '(7 1 127 3 5 4 77 2 9 0)))
;; Sort BV.
(qsort! bv (lambda (x y) (- x y)))
;; Let's see what the sorted array looks like:
(bytevector->u8-list bv)
⇒ (0 1 2 3 4 5 7 9 77 127)

And voilà!

Note that procedure->pointer is not supported (and not defined) on a few exotic architectures. Thus, user code may need to check (defined? 'procedure->pointer). Nevertheless, it is available on many architectures, including (as of libffi 3.0.9) x86, ia64, SPARC, PowerPC, ARM, and MIPS, to name a few.

6.21 Threads, Mutexes, Asyncs and Dynamic Roots

6.21.1 Arbiters

Arbiters are synchronization objects, they can be used by threads to control access to a shared resource. An arbiter can be locked to indicate a resource is in use, and unlocked when done.

An arbiter is like a light-weight mutex (see Section 6.21.4 [Mutexes and Condition Variables], page 415). It uses less memory and may be faster, but there’s no way for a thread to block waiting on an arbiter, it can only test and get the status returned.

[Scheme Procedure] make-arbiter name [C Function] scm_make_arbiter (name)

Return an object of type arbiter and name name. Its state is initially unlocked. Arbiters are a way to achieve process synchronization.

[Scheme Procedure] try-arbiter arb [C Function] scm_try_arbiter (arb)

If arb is unlocked, then lock it and return #t. If arb is already locked, then do nothing and return #f.

[Scheme Procedure] release-arbiter arb

[C Function] scm_release_arbiter (arb)

If arb is locked, then unlock it and return #t. If arb is already unlocked, then do nothing and return #f. Typical usage is for the thread which locked an arbiter to later release it, but that’s not required, any thread can release it.

6.21.2 Asyncs

Asyncs are a means of deferring the execution of Scheme code until it is safe to do so.

Guile provides two kinds of asyncs that share the basic concept but are otherwise quite different: system asyncs and user asyncs. System asyncs are integrated into the core of Guile and are executed automatically when the system is in a state to allow the execution of Scheme code. For example, it is not possible to execute Scheme code in a POSIX signal handler, but such a signal handler can queue a system async to be executed in the near future, when it is safe to do so.

System asyncs can also be queued for threads other than the current one. This way, you can cause threads to asynchronously execute arbitrary code.

User asyncs offer a convenient means of queuing procedures for future execution and triggering this execution. They will not be executed automatically.

6.21.2.1 System asyncs

To cause the future asynchronous execution of a procedure in a given thread, use system-

async-mark.

Automatic invocation of system asyncs can be temporarily disabled by calling callwith-blocked-asyncs. This function works by temporarily increasing the async blocking level of the current thread while a given procedure is running. The blocking level starts out at zero, and whenever a safe point is reached, a blocking level greater than zero will prevent the execution of queued asyncs.

Analogously, the procedure call-with-unblocked-asyncs will temporarily decrease the blocking level of the current thread. You can use it when you want to disable asyncs by default and only allow them temporarily.

   In addition to the C versions of call-with-blocked-asyncs and call-with-
unblocked-asyncs, C code can use scm_dynwind_block_asyncs and scm_dynwind_

unblock_asyncs inside a dynamic context (see Section 6.13.10 [Dynamic Wind], page 309) to block or unblock system asyncs temporarily.

[Scheme Procedure] system-async-mark proc [thread]

[C Function] scm_system_async_mark (proc)
[C Function] scm_system_async_mark_for_thread (proc, thread)

Mark proc (a procedure with zero arguments) for future execution in thread. When proc has already been marked for thread but has not been executed yet, this call has no effect. When thread is omitted, the thread that called system-async-mark is used. This procedure is not safe to be called from signal handlers. Use scm_sigaction or

     scm_sigaction_for_thread to install signal handlers.

[Scheme Procedure] call-with-blocked-asyncs proc
[C Function] scm_call_with_blocked_asyncs (proc)

Call proc and block the execution of system asyncs by one level for the current thread while it is running. Return the value returned by proc. For the first two variants, call proc with no arguments; for the third, call it with data.

[C Function] void * scm_c_call_with_blocked_asyncs (void * (*proc) (void

*data), void *data) The same but with a C function proc instead of a Scheme thunk.

[Scheme Procedure] call-with-unblocked-asyncs proc
[C Function] scm_call_with_unblocked_asyncs (proc)

Call proc and unblock the execution of system asyncs by one level for the current thread while it is running. Return the value returned by proc. For the first two variants, call proc with no arguments; for the third, call it with data.

[C Function] void * scm_c_call_with_unblocked_asyncs (void *(*proc) (void

*data), void *data) The same but with a C function proc instead of a Scheme thunk.

[C Function] void scm_dynwind_block_asyncs ()

During the current dynwind context, increase the blocking of asyncs by one level. This function must be used inside a pair of calls to scm_dynwind_begin and scm_ dynwind_end (see Section 6.13.10 [Dynamic Wind], page 309).

[C Function] void scm_dynwind_unblock_asyncs ()

During the current dynwind context, decrease the blocking of asyncs by one level. This function must be used inside a pair of calls to scm_dynwind_begin and scm_ dynwind_end (see Section 6.13.10 [Dynamic Wind], page 309).

6.21.2.2 User asyncs

A user async is a pair of a thunk (a parameterless procedure) and a mark. Setting the mark on a user async will cause the thunk to be executed when the user async is passed to run-asyncs. Setting the mark more than once is satisfied by one execution of the thunk.

User asyncs are created with async. They are marked with async-mark.

[Scheme Procedure] async thunk [C Function] scm_async (thunk)

Create a new user async for the procedure thunk.

[Scheme Procedure] async-mark a

[C Function] scm_async_mark (a)

Mark the user async a for future execution.

[Scheme Procedure] run-asyncs list of a [C Function] scm_run_asyncs (list of a)

Execute all thunks from the marked asyncs of the list list of a.

6.21.3 Threads

Guile supports POSIX threads, unless it was configured with --without-threads or the host lacks POSIX thread support. When thread support is available, the threads feature is provided (see Section 6.22.2.1 [Feature Manipulation], page 429).

The procedures below manipulate Guile threads, which are wrappers around the system’s POSIX threads. For application-level parallelism, using higher-level constructs, such as futures, is recommended (see Section 6.21.9 [Futures], page 424).

[Scheme Procedure] all-threads

[C Function] scm_all_threads ()

Return a list of all threads.

[Scheme Procedure] current-thread

[C Function] scm_current_thread ()

Return the thread that called this function.

[Scheme Procedure] call-with-new-thread thunk [handler]

Call thunk in a new thread and with a new dynamic state, returning the new thread. The procedure thunk is called via with-continuation-barrier. When handler is specified, then thunk is called from within a catch with tag #t that has handler as its handler. This catch is established inside the continuation barrier. Once thunk or handler returns, the return value is made the exit value of the thread and the thread is terminated.

[C Function] SCM scm_spawn_thread (scm t catch body body, void *body data,

scm t catch handler handler, void *handler data) Call body in a new thread, passing it body data, returning the new thread. The

function body is called via scm_c_with_continuation_barrier.

When handler is non-NULL, body is called via scm_internal_catch with tag SCM_ BOOL_T that has handler and handler data as the handler and its data. This catch is established inside the continuation barrier. Once body or handler returns, the return value is made the exit value of the thread and the thread is terminated.

[Scheme Procedure] thread? obj [C Function] scm_thread_p (obj)

Return #t ff obj is a thread; otherwise, return #f.

[Scheme Procedure] join-thread thread [timeout [timeoutval]]

[C Function] scm_join_thread (thread)

[C Function] scm_join_thread_timed (thread, timeout, timeoutval)

Wait for thread to terminate and return its exit value. Threads that have not been

created with call-with-new-thread or scm_spawn_thread have an exit value of

#f. When timeout is given, it specifies a point in time where the waiting should be aborted. It can be either an integer as returned by current-time or a pair as returned by gettimeofday. When the waiting is aborted, timeoutval is returned (if it is specified; #f is returned otherwise).

[Scheme Procedure] thread-exited? thread

[C Function] scm_thread_exited_p (thread)

Return #t if thread has exited, or #f otherwise.

[Scheme Procedure] yield

If one or more threads are waiting to execute, calling yield forces an immediate context switch to one of them. Otherwise, yield has no effect.

[Scheme Procedure] cancel-thread thread [C Function] scm_cancel_thread (thread)

Asynchronously notify thread to exit. Immediately after receiving this notification, thread will call its cleanup handler (if one has been set) and then terminate, aborting any evaluation that is in progress.

Because Guile threads are isomorphic with POSIX threads, thread will not receive its cancellation signal until it reaches a cancellation point. See your operating system’s POSIX threading documentation for more information on cancellation points; note that in Guile, unlike native POSIX threads, a thread can receive a cancellation notification while attempting to lock a mutex.

[Scheme Procedure] set-thread-cleanup! thread proc

[C Function] scm_set_thread_cleanup_x (thread, proc)

Set proc as the cleanup handler for the thread thread. proc, which must be a thunk, will be called when thread exits, either normally or by being canceled. Thread cleanup handlers can be used to perform useful tasks like releasing resources, such as locked mutexes, when thread exit cannot be predicted.

The return value of proc will be set as the exit value of thread.

To remove a cleanup handler, pass #f for proc.

[Scheme Procedure] thread-cleanup thread [C Function] scm_thread_cleanup (thread)

Return the cleanup handler currently installed for the thread thread. If no cleanup handler is currently installed, thread-cleanup returns #f.

Higher level thread procedures are available by loading the (ice-9 threads) module. These provide standardized thread creation.

[macro] make-thread proc arg …

Apply proc to arg … in a new thread formed by call-with-new-thread using a default error handler that display the error to the current error port. The arg … expressions are evaluated in the new thread.

[macro] begin-thread expr1 expr2 …

Evaluate forms expr1 expr2 … in a new thread formed by call-with-new-thread using a default error handler that display the error to the current error port.

6.21.4 Mutexes and Condition Variables

A mutex is a thread synchronization object, it can be used by threads to control access to a shared resource. A mutex can be locked to indicate a resource is in use, and other threads can then block on the mutex to wait for the resource (or can just test and do something else if not available). “Mutex” is short for “mutual exclusion”.

There are two types of mutexes in Guile, “standard” and “recursive”. They’re created by make-mutex and make-recursive-mutex respectively, the operation functions are then common to both.

Note that for both types of mutex there’s no protection against a “deadly embrace”. For instance if one thread has locked mutex A and is waiting on mutex B, but another thread owns B and is waiting on A, then an endless wait will occur (in the current implementation). Acquiring requisite mutexes in a fixed order (like always A before B) in all threads is one way to avoid such problems.

[Scheme Procedure] make-mutex flag …

[C Function] scm_make_mutex ()
[C Function] scm_make_mutex_with_flags (SCM flags)

Return a new mutex. It is initially unlocked. If flag … is specified, it must be a list of symbols specifying configuration flags for the newly-created mutex. The supported flags are:

unchecked-unlock

Unless this flag is present, a call to ‘unlock-mutex’ on the returned mutex when it is already unlocked will cause an error to be signalled.

allow-external-unlock

Allow the returned mutex to be unlocked by the calling thread even if it was originally locked by a different thread.

recursive

The returned mutex will be recursive.

[Scheme Procedure] mutex? obj [C Function] scm_mutex_p (obj)

Return #t if obj is a mutex; otherwise, return #f.

[Scheme Procedure] make-recursive-mutex
[C Function] scm_make_recursive_mutex ()

Create a new recursive mutex. It is initially unlocked. Calling this function is equivalent to calling ‘make-mutex’ and specifying the recursive flag.

[Scheme Procedure] lock-mutex mutex [timeout [owner]]

[C Function] scm_lock_mutex (mutex)

[C Function] scm_lock_mutex_timed (mutex, timeout, owner)

Lock mutex. If the mutex is already locked, then block and return only when mutex has been acquired. When timeout is given, it specifies a point in time where the waiting should be aborted. It can be either an integer as returned by current-time or a pair as returned by gettimeofday. When the waiting is aborted, #f is returned.

When owner is given, it specifies an owner for mutex other than the calling thread. owner may also be #f, indicating that the mutex should be locked but left unowned. For standard mutexes (make-mutex), and error is signalled if the thread has itself already locked mutex. For a recursive mutex (make-recursive-mutex), if the thread has itself already locked mutex, then a further lock-mutex call increments the lock count. An additional unlock-mutex will be required to finally release. If mutex was locked by a thread that exited before unlocking it, the next attempt to lock mutex will succeed, but abandoned-mutex-error will be signalled. When a system async (see Section 6.21.2.1 [System asyncs], page 411) is activated for a thread blocked in lock-mutex, the wait is interrupted and the async is executed. When the async returns, the wait resumes.

[C Function] void scm_dynwind_lock_mutex (SCM mutex)

Arrange for mutex to be locked whenever the current dynwind context is entered and to be unlocked when it is exited.

[Scheme Procedure] try-mutex mx [C Function] scm_try_mutex (mx)

Try to lock mutex as per lock-mutex. If mutex can be acquired immediately then this is done and the return is #t. If mutex is locked by some other thread then nothing is done and the return is #f.

[Scheme Procedure] unlock-mutex mutex [condvar [timeout]]

[C Function] scm_unlock_mutex (mutex)

[C Function] scm_unlock_mutex_timed (mutex, condvar, timeout)

Unlock mutex. An error is signalled if mutex is not locked and was not created with the unchecked-unlock flag set, or if mutex is locked by a thread other than the calling thread and was not created with the allow-external-unlock flag set. If condvar is given, it specifies a condition variable upon which the calling thread will wait to be signalled before returning. (This behavior is very similar to that of wait-condition-variable, except that the mutex is left in an unlocked state when the function returns.) When timeout is also given and not false, it specifies a point in time where the waiting should be aborted. It can be either an integer as returned by current-time or a pair as returned by gettimeofday. When the waiting is aborted, #f is returned. Otherwise the function returns #t.

[Scheme Procedure] mutex-owner mutex [C Function] scm_mutex_owner (mutex)

Return the current owner of mutex, in the form of a thread or #f (indicating no owner). Note that a mutex may be unowned but still locked.

[Scheme Procedure] mutex-level mutex [C Function] scm_mutex_level (mutex)

Return the current lock level of mutex. If mutex is currently unlocked, this value will be 0; otherwise, it will be the number of times mutex has been recursively locked by its current owner.

[Scheme Procedure] mutex-locked? mutex

[C Function] scm_mutex_locked_p (mutex)

Return #t if mutex is locked, regardless of ownership; otherwise, return #f.

[Scheme Procedure] make-condition-variable
[C Function] scm_make_condition_variable ()

Return a new condition variable.

[Scheme Procedure] condition-variable? obj

[C Function] scm_condition_variable_p (obj)

Return #t if obj is a condition variable; otherwise, return #f.

[Scheme Procedure] wait-condition-variable condvar mutex [time] [C Function] scm_wait_condition_variable (condvar, mutex, time)

Wait until condvar has been signalled. While waiting, mutex is atomically unlocked (as with unlock-mutex) and is locked again when this function returns. When time is given, it specifies a point in time where the waiting should be aborted. It can be either a integer as returned by current-time or a pair as returned by gettimeofday. When the waiting is aborted, #f is returned. When the condition variable has in fact been signalled, #t is returned. The mutex is re-locked in any case before wait-

condition-variable returns.

When a system async is activated for a thread that is blocked in a call to waitcondition-variable, the waiting is interrupted, the mutex is locked, and the async is executed. When the async returns, the mutex is unlocked again and the waiting is resumed. When the thread block while re-acquiring the mutex, execution of asyncs is blocked.

[Scheme Procedure] signal-condition-variable condvar
[C Function] scm_signal_condition_variable (condvar)

Wake up one thread that is waiting for condvar.

[Scheme Procedure] broadcast-condition-variable condvar
[C Function] scm_broadcast_condition_variable (condvar)

Wake up all threads that are waiting for condvar.

The following are higher level operations on mutexes. These are available from

(use-modules (ice-9 threads))

[macro] with-mutex mutex body1 body2 … Lock mutex, evaluate the body body1 body2 …, then unlock mutex. The return value is that returned by the last body form.

The lock, body and unlock form the branches of a dynamic-wind (see Section 6.13.10 [Dynamic Wind], page 309), so mutex is automatically unlocked if an error or new continuation exits the body, and is re-locked if the body is re-entered by a captured continuation.

[macro] monitor body1 body2 … Evaluate the body form body1 body2 … with a mutex locked so only one thread can execute that code at any one time. The return value is the return from the last body form. Each monitor form has its own private mutex and the locking and evaluation is as per with-mutex above. A standard mutex (make-mutex) is used, which means the body must not recursively re-enter the monitor form. The term “monitor” comes from operating system theory, where it means a particular bit of code managing access to some resource and which only ever executes on behalf of one process at any one time.

6.21.5 Blocking in Guile Mode

Up to Guile version 1.8, a thread blocked in guile mode would prevent the garbage collector from running. Thus threads had to explicitly leave guile mode with scm_without_guile () before making a potentially blocking call such as a mutex lock, a select () system call, etc. The following functions could be used to temporarily leave guile mode or to perform some common blocking operations in a supported way.

Starting from Guile 2.0, blocked threads no longer hinder garbage collection. Thus, the functions below are not needed anymore. They can still be used to inform the GC that a thread is about to block, giving it a (small) optimization opportunity for “stop the world” garbage collections, should they occur while the thread is blocked.

[C Function] void * scm_without_guile (void *(*func) (void *), void *data)

Leave guile mode, call func on data, enter guile mode and return the result of calling func. While a thread has left guile mode, it must not call any libguile functions except scm_with_guile or scm_without_guile and must not use any libguile macros. Also, local variables of type SCM that are allocated while not in guile mode are not protected from the garbage collector. When used from non-guile mode, calling scm_without_guile is still allowed: it simply calls func. In that way, you can leave guile mode without having to know whether the current thread is in guile mode or not.

[C Function] int scm_pthread_mutex_lock (pthread mutex t *mutex)

Like pthread_mutex_lock, but leaves guile mode while waiting for the mutex.

[C Function] int scm_pthread_cond_wait (pthread cond t *cond,

pthread mutex t *mutex)

[C Function] int scm_pthread_cond_timedwait (pthread cond t *cond,

pthread mutex t *mutex, struct timespec *abstime)

Like pthread_cond_wait and pthread_cond_timedwait, but leaves guile mode while

waiting for the condition variable.

[C Function] int scm_std_select (int nfds, fd set *readfds, fd set *writefds, fd set

*exceptfds, struct timeval *timeout) Like select but leaves guile mode while waiting. Also, the delivery of a system async causes this function to be interrupted with error code EINTR.

[C Function] unsigned int scm_std_sleep (unsigned int seconds)

Like sleep, but leaves guile mode while sleeping. Also, the delivery of a system async causes this function to be interrupted.

[C Function] unsigned long scm_std_usleep (unsigned long usecs)

Like usleep, but leaves guile mode while sleeping. Also, the delivery of a system async causes this function to be interrupted.

6.21.6 Critical Sections

[C Macro] SCM_CRITICAL_SECTION_START
[C Macro] SCM_CRITICAL_SECTION_END

These two macros can be used to delimit a critical section. Syntactically, they are both statements and need to be followed immediately by a semicolon. Executing SCM_CRITICAL_SECTION_START will lock a recursive mutex and block the executing of system asyncs. Executing SCM_CRITICAL_SECTION_END will unblock the execution of system asyncs and unlock the mutex. Thus, the code that executes between these two macros can only be executed in one thread at any one time and no system asyncs will run. However, because the mutex is a recursive one, the code might still be reentered by the same thread. You must either allow for this or avoid it, both by careful coding. On the other hand, critical sections delimited with these macros can be nested since the mutex is recursive. You must make sure that for each SCM_CRITICAL_SECTION_START, the corresponding SCM_CRITICAL_SECTION_END is always executed. This means that no non-local exit (such as a signalled error) might happen, for example.

[C Function] void scm_dynwind_critical_section (SCM mutex)
Call scm_dynwind_lock_mutex on mutex and call scm_dynwind_block_asyncs.

When mutex is false, a recursive mutex provided by Guile is used instead. The effect of a call to scm_dynwind_critical_section is that the current dynwind context (see Section 6.13.10 [Dynamic Wind], page 309) turns into a critical section. Because of the locked mutex, no second thread can enter it concurrently and because of the blocked asyncs, no system async can reenter it from the current thread. When the current thread reenters the critical section anyway, the kind of mutex determines what happens: When mutex is recursive, the reentry is allowed. When it is a normal mutex, an error is signalled.

6.21.7 Fluids and Dynamic States

A fluid is an object that can store one value per dynamic state. Each thread has a current dynamic state, and when accessing a fluid, this current dynamic state is used to provide the actual value. In this way, fluids can be used for thread local storage, but they are in fact more flexible: dynamic states are objects of their own and can be made current for more than one thread at the same time, or only be made current temporarily, for example.

Fluids can also be used to simulate the desirable effects of dynamically scoped variables. Dynamically scoped variables are useful when you want to set a variable to a value during some dynamic extent in the execution of your program and have them revert to their original value when the control flow is outside of this dynamic extent. See the description of with-fluids below for details.

New fluids are created with make-fluid and fluid? is used for testing whether an object is actually a fluid. The values stored in a fluid can be accessed with fluid-ref and

fluid-set!.

[Scheme Procedure] make-fluid [dflt]

[C Function] scm_make_fluid ()
[C Function] scm_make_fluid_with_default (dflt)

Return a newly created fluid, whose initial value is dflt, or #f if dflt is not given. Fluids are objects that can hold one value per dynamic state. That is, modifications to this value are only visible to code that executes with the same dynamic state as the modifying code. When a new dynamic state is constructed, it inherits the values from its parent. Because each thread normally executes with its own dynamic state, you can use fluids for thread local storage.

[Scheme Procedure] make-unbound-fluid
[C Function] scm_make_unbound_fluid ()

Return a new fluid that is initially unbound (instead of being implicitly bound to some definite value).

[Scheme Procedure] fluid? obj [C Function] scm_fluid_p (obj)

Return #t if obj is a fluid; otherwise, return #f.

[Scheme Procedure] fluid-ref fluid [C Function] scm_fluid_ref (fluid)

Return the value associated with fluid in the current dynamic root. If fluid has not been set, then return its default value. Calling fluid-ref on an unbound fluid produces a runtime error.

[Scheme Procedure] fluid-set! fluid value [C Function] scm_fluid_set_x (fluid, value)

Set the value associated with fluid in the current dynamic root.

[Scheme Procedure] fluid-unset! fluid

[C Function] scm_fluid_unset_x (fluid)

Disassociate the given fluid from any value, making it unbound.

[Scheme Procedure] fluid-bound? fluid

[C Function] scm_fluid_bound_p (fluid)

Returns #t if the given fluid is bound to a value, otherwise #f.

with-fluids* temporarily changes the values of one or more fluids, so that the given procedure and each procedure called by it access the given values. After the procedure returns, the old values are restored.

[Scheme Procedure] with-fluid* fluid value thunk [C Function] scm_with_fluid (fluid, value, thunk)

Set fluid to value temporarily, and call thunk. thunk must be a procedure with no argument.

[Scheme Procedure] with-fluids* fluids values thunk [C Function] scm_with_fluids (fluids, values, thunk)

Set fluids to values temporary, and call thunk. fluids must be a list of fluids and values must be the same number of their values to be applied. Each substitution is done in the order given. thunk must be a procedure with no argument. It is called inside a dynamic-wind and the fluids are set/restored when control enter or leaves the established dynamic extent.

[Scheme Macro] with-fluids ((fluid value) …) body1 body2 … Execute body body1 body2 … while each fluid is set to the corresponding value. Both fluid and value are evaluated and fluid must yield a fluid. The body is executed inside a dynamic-wind and the fluids are set/restored when control enter or leaves the established dynamic extent.

[C Function] SCM scm_c_with_fluids (SCM fluids, SCM vals, SCM (*cproc)(void

*), void *data)

[C Function] SCM scm_c_with_fluid (SCM fluid, SCM val, SCM (*cproc)(void *),

void *data) The function scm_c_with_fluids is like scm_with_fluids except that it takes a C function to call instead of a Scheme thunk.

The function scm_c_with_fluid is similar but only allows one fluid to be set instead of a list.

[C Function] void scm_dynwind_fluid (SCM fluid, SCM val)

This function must be used inside a pair of calls to scm_dynwind_begin and scm_ dynwind_end (see Section 6.13.10 [Dynamic Wind], page 309). During the dynwind context, the fluid fluid is set to val.

More precisely, the value of the fluid is swapped with a ‘backup’ value whenever the dynwind context is entered or left. The backup value is initialized with the val argument.

[Scheme Procedure] make-dynamic-state [parent]

[C Function] scm_make_dynamic_state (parent)

Return a copy of the dynamic state object parent or of the current dynamic state when parent is omitted.

[Scheme Procedure] dynamic-state? obj

[C Function] scm_dynamic_state_p (obj)

Return #t if obj is a dynamic state object; return #f otherwise.

[C Procedure] int scm_is_dynamic_state (SCM obj)

Return non-zero if obj is a dynamic state object; return zero otherwise.

[Scheme Procedure] current-dynamic-state
[C Function] scm_current_dynamic_state ()

Return the current dynamic state object.

[Scheme Procedure] set-current-dynamic-state state
[C Function] scm_set_current_dynamic_state (state)

Set the current dynamic state object to state and return the previous current dynamic state object.

[Scheme Procedure] with-dynamic-state state proc [C Function] scm_with_dynamic_state (state, proc)

Call proc while state is the current dynamic state object.

[C Procedure] void scm_dynwind_current_dynamic_state (SCM state)

Set the current dynamic state to state for the current dynwind context.

[C Procedure] void * scm_c_with_dynamic_state (SCM state, void

*(*func)(void *), void *data) Like scm_with_dynamic_state, but call func with data.

6.21.8 Parameters

A parameter object is a procedure. Calling it with no arguments returns its value. Calling it with one argument sets the value.

     (define my-param (make-parameter 123))
     (my-param) ⇒ 123
     (my-param 456)
     (my-param) ⇒ 456

The parameterize special form establishes new locations for parameters, those new locations having effect within the dynamic scope of the parameterize body. Leaving restores the previous locations. Re-entering (through a saved continuation) will again use the new locations.

     (parameterize ((my-param 789))
       (my-param)) ⇒ 789
     (my-param) ⇒ 456

Parameters are like dynamically bound variables in other Lisp dialects. They allow an application to establish parameter settings (as the name suggests) just for the execution of a particular bit of code, restoring when done. Examples of such parameters might be case-sensitivity for a search, or a prompt for user input.

Global variables are not as good as parameter objects for this sort of thing. Changes to them are visible to all threads, but in Guile parameter object locations are per-thread, thereby truly limiting the effect of parameterize to just its dynamic execution.

Passing arguments to functions is thread-safe, but that soon becomes tedious when there’s more than a few or when they need to pass down through several layers of calls before reaching the point they should affect. And introducing a new setting to existing code is often easier with a parameter object than adding arguments.

[Scheme Procedure] make-parameter init [converter]

Return a new parameter object, with initial value init. If a converter is given, then a call (converter val) is made for each value set, its return is the value stored. Such a call is made for the init initial value too. A converter allows values to be validated, or put into a canonical form. For example,

     (define my-param (make-parameter 123
                        (lambda (val)
                          (if (not (number? val))
                              (error "must be a number"))
                          (inexact->exact val))))
     (my-param 0.75)
     (my-param) ⇒ 3/4

[library syntax] parameterize ((param value) …) body1 body2 … Establish a new dynamic scope with the given params bound to new locations and set to the given values. body1 body2 … is evaluated in that environment. The value returned is that of last body form. Each param is an expression which is evaluated to get the parameter object. Often this will just be the name of a variable holding the object, but it can be anything that evaluates to a parameter. The param expressions and value expressions are all evaluated before establishing the new dynamic bindings, and they’re evaluated in an unspecified order. For example,

     (define prompt (make-parameter "Type something: "))
     (define (get-input)
       (display (prompt))
       ...)
     (parameterize ((prompt "Type a number: "))
       (get-input)
       ...)

Parameter objects are implemented using fluids (see Section 6.21.7 [Fluids and Dynamic States], page 419), so each dynamic state has its own parameter locations. That includes the separate locations when outside any parameterize form. When a parameter is created it gets a separate initial location in each dynamic state, all initialized to the given init value.

New code should probably just use parameters instead of fluids, because the interface is better. But for migrating old code or otherwise providing interoperability, Guile provides

the fluid->parameter procedure:

[Scheme Procedure] fluid->parameter fluid [conv]

Make a parameter that wraps a fluid. The value of the parameter will be the same as the value of the fluid. If the parameter is rebound in some dynamic extent, perhaps via parameterize, the new value will be run through the optional conv procedure, as with any parameter. Note that unlike make-parameter, conv is not applied to the initial value.

As alluded to above, because each thread usually has a separate dynamic state, each thread has its own locations behind parameter objects, and changes in one thread are not visible to any other. When a new dynamic state or thread is created, the values of parameters in the originating context are copied, into new locations.

Guile’s parameters conform to SRFI-39 (see Section 7.5.26 [SRFI-39], page 599).

6.21.9 Futures

The (ice-9 futures) module provides futures, a construct for fine-grain parallelism. A future is a wrapper around an expression whose computation may occur in parallel with the code of the calling thread, and possibly in parallel with other futures. Like promises, futures are essentially proxies that can be queried to obtain the value of the enclosed expression:

     (touch (future (+ 2 3)))
     ⇒ 5

However, unlike promises, the expression associated with a future may be evaluated on another CPU core, should one be available. This supports fine-grain parallelism, because even relatively small computations can be embedded in futures. Consider this sequential code:

     (define (find-prime lst1 lst2)
       (or (find prime? lst1)
           (find prime? lst2)))

The two arms of or are potentially computation-intensive. They are independent of one another, yet, they are evaluated sequentially when the first one returns #f. Using futures, one could rewrite it like this:

     (define (find-prime lst1 lst2)
       (let ((f (future (find prime? lst2))))
         (or (find prime? lst1)
             (touch f))))

This preserves the semantics of find-prime. On a multi-core machine, though, the computation of (find prime? lst2) may be done in parallel with that of the other find call, which can reduce the execution time of find-prime.

Futures may be nested: a future can itself spawn and then touch other futures, leading to a directed acyclic graph of futures. Using this facility, a parallel map procedure can be defined along these lines:

     (use-modules (ice-9 futures) (ice-9 match))
     (define (par-map proc lst)
       (match lst
         (()
          '())
         ((head tail ...)
          (let ((tail (future (par-map proc tail)))
                (head (proc head)))
            (cons head (touch tail))))))

Note that futures are intended for the evaluation of purely functional expressions. Expressions that have side-effects or rely on I/O may require additional care, such as explicit synchronization (see Section 6.21.4 [Mutexes and Condition Variables], page 415).

Guile’s futures are implemented on top of POSIX threads (see Section 6.21.3 [Threads], page 413). Internally, a fixed-size pool of threads is used to evaluate futures, such that offloading the evaluation of an expression to another thread doesn’t incur thread creation costs. By default, the pool contains one thread per available CPU core, minus one, to account for the main thread. The number of available CPU cores is determined using current-processor-count (see Section 7.2.7 [Processes], page 490).

When a thread touches a future that has not completed yet, it processes any pending future while waiting for it to complete, or just waits if there are no pending futures. When touch is called from within a future, the execution of the calling future is suspended, allowing its host thread to process other futures, and resumed when the touched future has completed. This suspend/resume is achieved by capturing the calling future’s continuation, and later reinstating it (see Section 6.13.5 [Prompts], page 296).

Note that par-map above is not tail-recursive. This could lead to stack overflows when lst is large compared to (current-processor-count). To address that, touch uses the suspend mechanism described above to limit the number of nested futures executing on the same stack. Thus, the above code should never run into stack overflows.

[Scheme Syntax] future exp

Return a future for expression exp. This is equivalent to:

          (make-future (lambda () exp))

[Scheme Procedure] make-future thunk

Return a future for thunk, a zero-argument procedure.

This procedure returns immediately. Execution of thunk may begin in parallel with the calling thread’s computations, if idle CPU cores are available, or it may start when touch is invoked on the returned future. If the execution of thunk throws an exception, that exception will be re-thrown when touch is invoked on the returned future.

[Scheme Procedure] future? obj

Return #t if obj is a future.

[Scheme Procedure] touch f

Return the result of the expression embedded in future f.

If the result was already computed in parallel, touch returns instantaneously. Otherwise, it waits for the computation to complete, if it already started, or initiates it. In the former case, the calling thread may process other futures in the meantime.

6.21.10 Parallel forms

The functions described in this section are available from

     (use-modules (ice-9 threads))

They provide high-level parallel constructs. The following functions are implemented in terms of futures (see Section 6.21.9 [Futures], page 424). Thus they are relatively cheap as they re-use existing threads, and portable, since they automatically use one thread per available CPU core.

[syntax] parallel expr …

Evaluate each expr expression in parallel, each in its own thread. Return the results of n expressions as a set of n multiple values (see Section 6.13.7 [Multiple Values], page 301).

[syntax] letpar ((var expr) …) body1 body2 … Evaluate each expr in parallel, each in its own thread, then bind the results to the corresponding var variables, and then evaluate body1 body2 … letpar is like let (see Section 6.12.2 [Local Bindings], page 287), but all the expressions for the bindings are evaluated in parallel.

[Scheme Procedure] par-map proc lst1 lst2 …

[Scheme Procedure] par-for-each proc lst1 lst2 …

Call proc on the elements of the given lists. par-map returns a list comprising the return values from proc. par-for-each returns an unspecified value, but waits for all calls to complete. The proc calls are (proc elem1 elem2 ...), where each elem is from the corresponding lst . Each lst must be the same length. The calls are potentially made in parallel, depending on the number of CPU cores available. These functions are like map and for-each (see Section 6.7.2.8 [List Mapping], page 193), but make their proc calls in parallel.

Unlike those above, the functions described below take a number of threads as an argument. This makes them inherently non-portable since the specified number of threads may differ from the number of available CPU cores as returned by current-processor-count (see Section 7.2.7 [Processes], page 490). In addition, these functions create the specified number of threads when they are called and terminate them upon completion, which makes them quite expensive.

Therefore, they should be avoided.

[Scheme Procedure] n-par-map n proc lst1 lst2 …

[Scheme Procedure] n-par-for-each n proc lst1 lst2 …

Call proc on the elements of the given lists, in the same way as par-map and parfor-each above, but use no more than n threads at any one time. The order in which calls are initiated within that threads limit is unspecified. These functions are good for controlling resource consumption if proc calls might be costly, or if there are many to be made. On a dual-CPU system for instance n = 4 might be enough to keep the CPUs utilized, and not consume too much memory.

[Scheme Procedure] n-for-each-par-map n sproc pproc lst1 lst2 … Apply pproc to the elements of the given lists, and apply sproc to each result returned by pproc. The final return value is unspecified, but all calls will have been completed before returning. The calls made are (sproc (pproc elem1 ... elemN)), where each elem is from the corresponding lst. Each lst must have the same number of elements.

The pproc calls are made in parallel, in separate threads. No more than n threads are used at any one time. The order in which pproc calls are initiated within that limit is unspecified. The sproc calls are made serially, in list element order, one at a time. pproc calls on later elements may execute in parallel with the sproc calls. Exactly which thread makes each sproc call is unspecified. This function is designed for individual calculations that can be done in parallel, but with results needing to be handled serially, for instance to write them to a file. The n limit on threads controls system resource usage when there are many calculations or when they might be costly. It will be seen that n-for-each-par-map is like a combination of n-par-map and

for-each,
     (for-each sproc (n-par-map n pproc lst1 ... lstN))

But the actual implementation is more efficient since each sproc call, in turn, can be initiated once the relevant pproc call has completed, it doesn’t need to wait for all to finish.

6.22 Configuration, Features and Runtime Options

Why is my Guile different from your Guile? There are three kinds of possible variation:

• build differences — different versions of the Guile source code, installation directories, configuration flags that control pieces of functionality being included or left out, etc.

• differences in dynamically loaded code — behaviour and features provided by modules that can be dynamically loaded into a running Guile

• different runtime options — some of the options that are provided for controlling Guile’s behaviour may be set differently.

Guile provides “introspective” variables and procedures to query all of these possible variations at runtime. For runtime options, it also provides procedures to change the settings of options and to obtain documentation on what the options mean.

6.22.1 Configuration, Build and Installation

The following procedures and variables provide information about how Guile was configured, built and installed on your system.

[Scheme Procedure] version

[Scheme Procedure] effective-version

[Scheme Procedure] major-version [Scheme Procedure] minor-version [Scheme Procedure] micro-version [C Function] scm_version ()

[C Function] scm_effective_version ()
[C Function] scm_major_version ()
[C Function] scm_minor_version ()
[C Function] scm_micro_version ()

Return a string describing Guile’s full version number, effective version number, major, minor or micro version number, respectively. The effective-version function returns the version name that should remain unchanged during a stable series. Currently that means that it omits the micro version. The effective version should be used for items like the versioned share directory name i.e. /usr/share/guile/2.0/

     (version) ⇒ "2.0.4"
     (effective-version) ⇒ "2.0"
     (major-version) ⇒ "2"
     (minor-version) ⇒ "0"
     (micro-version) ⇒ "4"
[Scheme Procedure] %package-data-dir
[C Function] scm_sys_package_data_dir ()

Return the name of the directory under which Guile Scheme files in general are stored. On Unix-like systems, this is usually /usr/local/share/guile or

/usr/share/guile.

[Scheme Procedure] %library-dir

[C Function] scm_sys_library_dir ()

Return the name of the directory where the Guile Scheme files that belong to the core Guile installation (as opposed to files from a 3rd party package) are installed.

On Unix-like systems this is usually /usr/local/share/guile/GUILE_EFFECTIVE_
VERSION or /usr/share/guile/GUILE_EFFECTIVE_VERSION;
for example /usr/local/share/guile/2.0.

[Scheme Procedure] %site-dir

[C Function] scm_sys_site_dir ()

Return the name of the directory where Guile Scheme files specific to your site should be installed. On Unix-like systems, this is usually /usr/local/share/guile/site

or /usr/share/guile/site.

[Scheme Procedure] %site-ccache-dir

[C Function] scm_sys_site_ccache_dir ()

Return the directory where users should install compiled .go files for use with this version of Guile. Might look something like /usr/lib/guile/2.0/site-ccache.

[Variable] %guile-build-info

Alist of information collected during the building of a particular Guile. Entries can be grouped into one of several categories: directories, env vars, and versioning info. Briefly, here are the keys in %guile-build-info, by group:

directories srcdir, top srcdir, prefix, exec prefix, bindir, sbindir, libexecdir, datadir, sysconfdir, sharedstatedir, localstatedir, libdir, infodir, mandir, includedir, pkgdatadir, pkglibdir, pkgincludedir

env vars LIBS

versioning info guileversion, libguileinterface, buildstamp

Values are all strings. The value for LIBS is typically found also as a part of pkgconfig --libs guile-2.0 output. The value for guileversion has form X.Y.Z, and should be the same as returned by (version). The value for libguileinterface is libtool compatible and has form CURRENT:REVISION:AGE (see Section “Library interface versions” in GNU Libtool). The value for buildstamp is the output of the

command ‘date -u +'%Y-%m-%d %T'’ (UTC).

In the source, %guile-build-info is initialized from libguile/libpath.h, which is completely generated, so deleting this file before a build guarantees up-to-date values for that build.

[Variable] %host-type

The canonical host type (GNU triplet) of the host Guile was configured for, e.g., "x86_64-unknown-linux-gnu" (see Section “Canonicalizing” in The GNU Autoconf Manual).

6.22.2 Feature Tracking

Guile has a Scheme level variable *features* that keeps track to some extent of the features that are available in a running Guile. *features* is a list of symbols, for example threads, each of which describes a feature of the running Guile process.

[Variable] *features*

A list of symbols describing available features of the Guile process.

You shouldn’t modify the *features* variable directly using set!. Instead, see the procedures that are provided for this purpose in the following subsection.

6.22.2.1 Feature Manipulation

To check whether a particular feature is available, use the provided? procedure:

[Scheme Procedure] provided? feature [Deprecated Scheme Procedure] feature? feature

Return #t if the specified feature is available, otherwise #f.

To advertise a feature from your own Scheme code, you can use the provide procedure:

[Scheme Procedure] provide feature

Add feature to the list of available features in this Guile process.

For C code, the equivalent function takes its feature name as a char * argument for convenience:

[C Function] void scm_add_feature (const char *str)

Add a symbol with name str to the list of available features in this Guile process.

6.22.2.2 Common Feature Symbols

In general, a particular feature may be available for one of two reasons. Either because the Guile library was configured and compiled with that feature enabled — i.e. the feature is built into the library on your system. Or because some C or Scheme code that was dynamically loaded by Guile has added that feature to the list.

In the first category, here are the features that the current version of Guile may define (depending on how it is built), and what they mean. array Indicates support for arrays (see Section 6.7.5 [Arrays], page 200).

array-for-each

Indicates availability of array-for-each and other array mapping procedures (see Section 6.7.5 [Arrays], page 200).

char-ready?

Indicates that the char-ready? function is available (see Section 6.14.2 [Reading], page 318).

complex Indicates support for complex numbers.

current-time

Indicates availability of time-related functions: times, get-internal-run-time and so on (see Section 7.2.5 [Time], page 485).

debug-extensions

Indicates that the debugging evaluator is available, together with the options for controlling it.

delay Indicates support for promises (see Section 6.17.9 [Delayed Evaluation], page 371).

EIDs Indicates that the geteuid and getegid really return effective user and group IDs (see Section 7.2.7 [Processes], page 490).

inexact Indicates support for inexact numbers.

i/o-extensions

Indicates availability of the following extended I/O procedures: ftell,

redirect-port, dup->fdes, dup2, fileno, isatty?, fdopen, primitive-

move->fdes and fdes->ports (see Section 7.2.2 [Ports and File Descriptors], page 469).

net-db Indicates availability of network database functions: scm_gethost,

scm_getnet, scm_getproto, scm_getserv, scm_sethost, scm_setnet,

scm_setproto, scm_setserv, and their ‘byXXX’ variants (see Section 7.2.11.2 [Network Databases], page 503).

posix Indicates support for POSIX functions: pipe, getgroups, kill, execl and so on (see Section 7.2 [POSIX], page 468).

fork Indicates support for the POSIX fork function (see Section 7.2.7 [Processes], page 490). This is a prerequisite for the (ice-9 popen) module (see Section 7.2.10 [Pipes], page 500).

random Indicates availability of random number generation functions: random, copyrandom-state, random-uniform and so on (see Section 6.6.2.14 [Random], page 127).

reckless Indicates that Guile was built with important checks omitted — you should never see this!

regex Indicates support for POSIX regular expressions using make-regexp, regexpexec and friends (see Section 6.15.1 [Regexp Functions], page 350). socket Indicates availability of socket-related functions: socket, bind, connect and

so on (see Section 7.2.11.4 [Network Sockets and Communication], page 511).

sort Indicates availability of sorting and merging functions (see Section 6.11.3 [Sorting], page 279).

system Indicates that the system function is available (see Section 7.2.7 [Processes], page 490).

threads Indicates support for multithreading (see Section 6.21.3 [Threads], page 413).

values Indicates support for multiple return values using values and call-with-values (see Section 6.13.7 [Multiple Values], page 301).

Available features in the second category depend, by definition, on what additional code your Guile process has loaded in. The following table lists features that you might encounter for this reason.

defmacro Indicates that the defmacro macro is available (see Section 6.10 [Macros], page 257).

describe Indicates that the (oop goops describe) module has been loaded, which provides a procedure for describing the contents of GOOPS instances.

readline Indicates that Guile has loaded in Readline support, for command line editing (see Section 7.8 [Readline Support], page 660).

record Indicates support for record definition using make-record-type and friends (see Section 6.7.9 [Records], page 219).

Although these tables may seem exhaustive, it is probably unwise in practice to rely on them, as the correspondences between feature symbols and available procedures/behaviour are not strictly defined. If you are writing code that needs to check for the existence of some procedure, it is probably safer to do so directly using the defined? procedure than to test for the corresponding feature using provided?.

6.22.3 Runtime Options

There are a number of runtime options available for paramaterizing built-in procedures, like read, and built-in behavior, like what happens on an uncaught error.

For more information on reader options, See Section 6.17.2 [Scheme Read], page 360. For more information on print options, See Section 6.17.3 [Scheme Write], page 361. Finally, for more information on debugger options, See Section 6.25.3.4 [Debug Options],

page 454.

6.22.3.1 Examples of option use

Here is an example of a session in which some read and debug option handling procedures are used. In this example, the user

1. Notices that the symbols abc and aBc are not the same

2. Examines the read-options, and sees that case-insensitive is set to “no”.

 3. Enables case-insensitive

4. Quits the recursive prompt 5. Verifies that now aBc and abc are the same

      scheme@(guile-user)> (define abc "hello")
      scheme@(guile-user)> abc
      $1 = "hello"
      scheme@(guile-user)> aBc
      <unknown-location>: warning: possibly unbound variable `aBc'
      ERROR: In procedure module-lookup:
      ERROR: Unbound variable: aBc
      Entering a new prompt.  Type `,bt' for a backtrace or `,q' to continue.
      scheme@(guile-user) [1]> (read-options 'help)
      copy              no    Copy source code expressions.
      positions         yes   Record positions of source code expressions.
      case-insensitive  no    Convert symbols to lower case.
      keywords          #f    Style of keyword recognition: #f, 'prefix or 'postfix.
      r6rs-hex-escapes  no    Use R6RS variable-length character and string hex escapes.
      square-brackets   yes   Treat `[' and `]' as parentheses, for R6RS compatibility.
      hungry-eol-escapes no   In strings, consume leading whitespace after an
                              escaped end-of-line.
      curly-infix       no    Support SRFI-105 curly infix expressions.
      scheme@(guile-user) [1]> (read-enable 'case-insensitive)
      $2 = (square-brackets keywords #f case-insensitive positions)
      scheme@(guile-user) [1]> ,q
      scheme@(guile-user)> aBc
      $3 = "hello"

6.23 Support for Other Languages

In addition to Scheme, a user may write a Guile program in an increasing number of other languages. Currently supported languages include Emacs Lisp and ECMAScript.

Guile is still fundamentally a Scheme, but it tries to support a wide variety of language building-blocks, so that other languages can be implemented on top of Guile. This allows users to write or extend applications in languages other than Scheme, too. This section describes the languages that have been implemented.

(For details on how to implement a language, See Section 9.4 [Compiling to the Virtual Machine], page 791.)

6.23.1 Using Other Languages

There are currently only two ways to access other languages from within Guile: at the

REPL, and programmatically, via compile, read-and-compile, and compile-file.

The REPL is Guile’s command prompt (see Section 4.4 [Using Guile Interactively], page 46). The REPL has a concept of the “current language”, which defaults to Scheme. The user may change that language, via the meta-command ,language.

For example, the following meta-command enables Emacs Lisp input:

     scheme@(guile-user)> ,language elisp
     Happy hacking with Emacs Lisp!  To switch back, type `,L scheme'.
     elisp@(guile-user)> (eq 1 2)
     $1 = #nil

Each language has its short name: for example, elisp, for Elisp. The same short name may be used to compile source code programmatically, via compile:

     elisp@(guile-user)> ,L scheme

     Happy hacking with Guile Scheme!  To switch back, type `,L elisp'.
     scheme@(guile-user)> (compile '(eq 1 2) #:from 'elisp)
     $2 = #nil

Granted, as the input to compile is a datum, this works best for Lispy languages, which have a straightforward datum representation. Other languages that need more parsing are better dealt with as strings.

The easiest way to deal with syntax-heavy language is with files, via compile-file and friends. However it is possible to invoke a language’s reader on a port, and then compile the resulting expression (which is a datum at that point). For more information, See Section 6.17.5 [Compilation], page 364.

For more details on introspecting aspects of different languages, See Section 9.4.1 [Compiler Tower], page 792.

6.23.2 Emacs Lisp

Emacs Lisp (Elisp) is a dynamically-scoped Lisp dialect used in the Emacs editor. See Section “Overview” in Emacs Lisp, for more information on Emacs Lisp.

We hope that eventually Guile’s implementation of Elisp will be good enough to replace Emacs’ own implementation of Elisp. For that reason, we have thought long and hard about how to support the various features of Elisp in a performant and compatible manner.

Readers familiar with Emacs Lisp might be curious about how exactly these various Elisp features are supported in Guile. The rest of this section focuses on addressing these concerns of the Elisp elect.

6.23.2.1 Nil

nil in ELisp is an amalgam of Scheme’s #f and '(). It is false, and it is the end-of-list; thus it is a boolean, and a list as well.

Guile has chosen to support nil as a separate value, distinct from #f and '(). This allows existing Scheme and Elisp code to maintain their current semantics. nil, which in Elisp would just be written and read as nil, in Scheme has the external representation

#nil.

This decision to have nil as a low-level distinct value facilitates interoperability between the two languages. Guile has chosen to have Scheme deal with nil as follows:

     (boolean? #nil) ⇒ #t
     (not #nil) ⇒ #t
     (null? #nil) ⇒ #t

And in C, one has:

     scm_is_bool (SCM_ELISP_NIL) ⇒ 1
     scm_is_false (SCM_ELISP_NIL) ⇒ 1
     scm_is_null (SCM_ELISP_NIL) ⇒ 1

In this way, a version of fold written in Scheme can correctly fold a function written in Elisp (or in fact any other language) over a nil-terminated list, as Elisp makes. The converse holds as well; a version of fold written in Elisp can fold over a '()-terminated list, as made by Scheme.

On a low level, the bit representations for #f, #t, nil, and '() are made in such a way that they differ by only one bit, and so a test for, for example, #f-or-nil may be made very efficiently. See libguile/boolean.h, for more information.

6.23.2.2 Equality

Since Scheme’s equal? must be transitive, and '() is not equal? to #f, to Scheme nil is

not equal? to #f or '().
(eq? #f '()) ⇒ #f
(eq? #nil '()) ⇒ #f
(eq? #nil #f) ⇒ #f
(eqv? #f '()) ⇒ #f
(eqv? #nil '()) ⇒ #f
(eqv? #nil #f) ⇒ #f
(equal? #f '()) ⇒ #f
(equal? #nil '()) ⇒ #f
(equal? #nil #f) ⇒ #f

However, in Elisp, '(), #f, and nil are all equal (though not eq).

(defvar f (make-scheme-false))
(defvar eol (make-scheme-null))
(eq f eol) ⇒ nil
(eq nil eol) ⇒ nil
(eq nil f) ⇒ nil
(equal f eol) ⇒ t
(equal nil eol) ⇒ t
(equal nil f) ⇒ t

These choices facilitate interoperability between Elisp and Scheme code, but they are not perfect. Some code that is correct standard Scheme is not correct in the presence of a second false and null value. For example:

(define (truthiness x)
  (if (eq? x #f)
      #f
      #t))

This code seems to be meant to test a value for truth, but now that there are two false values, #f and nil, it is no longer correct.

Similarly, there is the loop:

(define (my-length l)
  (let lp ((l l) (len 0))
    (if (eq? l '())
        len
        (lp (cdr l) (1+ len)))))

Here, my-length will raise an error if l is a nil-terminated list. Both of these examples are correct standard Scheme, but, depending on what they

really want to do, they are not correct Guile Scheme. Correctly written, they would test the properties of falsehood or nullity, not the individual members of that set. That is to say, they should use not or null? to test for falsehood or nullity, not eq? or memv or the like.

Fortunately, using not and null? is in good style, so all well-written standard Scheme programs are correct, in Guile Scheme.

Here are correct versions of the above examples:

     (define (truthiness* x)
       (if (not x)
           #f
           #t))
     ;; or: (define (t* x) (not (not x)))
     ;; or: (define (t** x) x)
     (define (my-length* l)
       (let lp ((l l) (len 0))
         (if (null? l)
             len
             (lp (cdr l) (1+ len)))))

This problem has a mirror-image case in Elisp:

     (deffn my-falsep (x)
       (if (eq x nil)
           t
           nil))

Guile can warn when compiling code that has equality comparisons with #f, '(), or nil. See Section 6.17.5 [Compilation], page 364, for details.

6.23.2.3 Dynamic Binding

In contrast to Scheme, which uses “lexical scoping”, Emacs Lisp scopes its variables dynamically. Guile supports dynamic scoping with its “fluids” facility. See Section 6.21.7 [Fluids and Dynamic States], page 419, for more information.

6.23.2.4 Other Elisp Features

Buffer-local and mode-local variables should be mentioned here, along with buckybits on characters, Emacs primitive data types, the Lisp-2-ness of Elisp, and other things. Contributions to the documentation are most welcome!

6.23.3 ECMAScript

ECMAScript (http://www.ecma-international.org/publications/files/ECMA-ST/

Ecma-262.pdf) was not the first non-Schemey language implemented by Guile, but it was the first implemented for Guile’s bytecode compiler. The goal was to support ECMAScript version 3.1, a relatively small language, but the implementor was completely irresponsible and got distracted by other things before finishing the standard library, and even some bits of the syntax. So, ECMAScript does deserve a mention in the manual, but it doesn’t deserve an endorsement until its implementation is completed, perhaps by some more responsible hacker.

In the meantime, the charitable user might investigate such invocations as ,L

ecmascript and cat test-suite/tests/ecmascript.test.

6.24 Support for Internationalization

Guile provides internationalization^{16} support for Scheme programs in two ways. First, procedures to manipulate text and data in a way that conforms to particular cultural conventions (i.e., in a “locale-dependent” way) are provided in the (ice-9 i18n). Second, Guile allows the use of GNU gettext to translate program message strings.

6.24.1 Internationalization with Guile

In order to make use of the functions described thereafter, the (ice-9 i18n) module must be imported in the usual way:

     (use-modules (ice-9 i18n))

The (ice-9 i18n) module provides procedures to manipulate text and other data in a way that conforms to the cultural conventions chosen by the user. Each region of the world or language has its own customs to, for instance, represent real numbers, classify characters, collate text, etc. All these aspects comprise the so-called “cultural conventions” of that region or language.

Computer systems typically refer to a set of cultural conventions as a locale. For each particular aspect that comprise those cultural conventions, a locale category is defined. For instance, the way characters are classified is defined by the LC_CTYPE category, while the language in which program messages are issued to the user is defined by the LC_MESSAGES category (see Section 7.2.13 [Locales], page 518 for details).

The procedures provided by this module allow the development of programs that adapt automatically to any locale setting. As we will see later, many of these procedures can optionally take a locale object argument. This additional argument defines the locale settings that must be followed by the invoked procedure. When it is omitted, then the current locale settings of the process are followed (see Section 7.2.13 [Locales], page 518).

The following procedures allow the manipulation of such locale objects.

[Scheme Procedure] make-locale category-list locale-name [base-locale] [C Function] scm_make_locale (category list, locale name, base locale)

Return a reference to a data structure representing a set of locale datasets. localename should be a string denoting a particular locale (e.g., "aa_DJ") and category-list should be either a list of locale categories or a single category as used with setlocale (see Section 7.2.13 [Locales], page 518). Optionally, if base-locale is passed, it should be a locale object denoting settings for categories not listed in category-list. The following invocation creates a locale object that combines the use of Swedish for messages and character classification with the default settings for the other categories (i.e., the settings of the default C locale which usually represents conventions in use in the USA):

          (make-locale (list LC_MESSAGES LC_CTYPE) "sv_SE")

The following example combines the use of Esperanto messages and conventions with monetary conventions from Croatia:

16 For concision and style, programmers often like to refer to internationalization as “i18n”.

     (make-locale LC_MONETARY "hr_HR"
                  (make-locale LC_ALL "eo_EO"))

A system-error exception (see Section 6.13.11 [Handling Errors], page 313) is raised by make-locale when locale-name does not match any of the locales compiled on the system. Note that on non-GNU systems, this error may be raised later, when the locale object is actually used.

[Scheme Procedure] locale? obj [C Function] scm_locale_p (obj)

Return true if obj is a locale object.

[Scheme Variable] %global-locale

[C Variable] scm_global_locale

This variable is bound to a locale object denoting the current process locale as installed using setlocale () (see Section 7.2.13 [Locales], page 518). It may be used like any other locale object, including as a third argument to make-locale, for instance.

6.24.2 Text Collation

The following procedures provide support for text collation, i.e., locale-dependent string and character sorting.

[Scheme Procedure] string-locale<? s1 s2 [locale] [C Function] scm_string_locale_lt (s1, s2, locale) [Scheme Procedure] string-locale>? s1 s2 [locale] [C Function] scm_string_locale_gt (s1, s2, locale) [Scheme Procedure] string-locale-ci<? s1 s2 [locale] [C Function] scm_string_locale_ci_lt (s1, s2, locale) [Scheme Procedure] string-locale-ci>? s1 s2 [locale] [C Function] scm_string_locale_ci_gt (s1, s2, locale)

Compare strings s1 and s2 in a locale-dependent way. If locale is provided, it should be locale object (as returned by make-locale) and will be used to perform the comparison; otherwise, the current system locale is used. For the -ci variants, the comparison is made in a case-insensitive way.

[Scheme Procedure] string-locale-ci=? s1 s2 [locale] [C Function] scm_string_locale_ci_eq (s1, s2, locale)

Compare strings s1 and s2 in a case-insensitive, and locale-dependent way. If locale is provided, it should be a locale object (as returned by make-locale) and will be used to perform the comparison; otherwise, the current system locale is used.

[Scheme Procedure] char-locale<? c1 c2 [locale] [C Function] scm_char_locale_lt (c1, c2, locale) [Scheme Procedure] char-locale>? c1 c2 [locale] [C Function] scm_char_locale_gt (c1, c2, locale) [Scheme Procedure] char-locale-ci<? c1 c2 [locale] [C Function] scm_char_locale_ci_lt (c1, c2, locale) [Scheme Procedure] char-locale-ci>? c1 c2 [locale] [C Function] scm_char_locale_ci_gt (c1, c2, locale)

Compare characters c1 and c2 according to either locale (a locale object as returned by make-locale) or the current locale. For the -ci variants, the comparison is made in a case-insensitive way.

[Scheme Procedure] char-locale-ci=? c1 c2 [locale] [C Function] scm_char_locale_ci_eq (c1, c2, locale)

Return true if character c1 is equal to c2, in a case insensitive way according to locale or to the current locale.

6.24.3 Character Case Mapping

The procedures below provide support for “character case mapping”, i.e., to convert characters or strings to their upper-case or lower-case equivalent. Note that SRFI-13 provides procedures that look similar (see Section 6.6.5.9 [Alphabetic Case Mapping], page 154). However, the SRFI-13 procedures are locale-independent. Therefore, they do not take into account specificities of the customs in use in a particular language or region of the world. For instance, while most languages using the Latin alphabet map lower-case letter “i” to upper-case letter “I”, Turkish maps lower-case “i” to “Latin capital letter I with dot above”. The following procedures allow programmers to provide idiomatic character mapping.

[Scheme Procedure] char-locale-downcase chr [locale] [C Function] scm_char_locale_upcase (chr, locale)

Return the lowercase character that corresponds to chr according to either locale or the current locale.

[Scheme Procedure] char-locale-upcase chr [locale]

[C Function] scm_char_locale_downcase (chr, locale)

Return the uppercase character that corresponds to chr according to either locale or the current locale.

[Scheme Procedure] char-locale-titlecase chr [locale]

[C Function] scm_char_locale_titlecase (chr, locale)

Return the titlecase character that corresponds to chr according to either locale or the current locale.

[Scheme Procedure] string-locale-upcase str [locale]

[C Function] scm_string_locale_upcase (str, locale)

Return a new string that is the uppercase version of str according to either locale or the current locale.

[Scheme Procedure] string-locale-downcase str [locale]

[C Function] scm_string_locale_downcase (str, locale)

Return a new string that is the down-case version of str according to either locale or the current locale.

[Scheme Procedure] string-locale-titlecase str [locale]

[C Function] scm_string_locale_titlecase (str, locale)

Return a new string that is the titlecase version of str according to either locale or the current locale.

6.24.4 Number Input and Output

The following procedures allow programs to read and write numbers written according to a particular locale. As an example, in English, “ten thousand and a half” is usually written 10,000.5 while in French it is written 10 000,5. These procedures allow such differences to be taken into account.

[Scheme Procedure] locale-string->integer str [base [locale]]

[C Function] scm_locale_string_to_integer (str, base, locale)

Convert string str into an integer according to either locale (a locale object as returned by make-locale) or the current process locale. If base is specified, then it determines the base of the integer being read (e.g., 16 for an hexadecimal number, 10 for a decimal number); by default, decimal numbers are read. Return two values (see Section 6.13.7 [Multiple Values], page 301): an integer (on success) or #f, and the number of characters read from str (0 on failure). This function is based on the C library’s strtol function (see Section “Parsing of Integers” in The GNU C Library Reference Manual).

[Scheme Procedure] locale-string->inexact str [locale]

[C Function] scm_locale_string_to_inexact (str, locale)

Convert string str into an inexact number according to either locale (a locale object as returned by make-locale) or the current process locale. Return two values (see Section 6.13.7 [Multiple Values], page 301): an inexact number (on success) or #f, and the number of characters read from str (0 on failure). This function is based on the C library’s strtod function (see Section “Parsing of Floats” in The GNU C Library Reference Manual).

[Scheme Procedure] number->locale-string number [fraction-digits [locale]] Convert number (an inexact) into a string according to the cultural conventions of either locale (a locale object) or the current locale. Optionally, fraction-digits may be bound to an integer specifying the number of fractional digits to be displayed.

[Scheme Procedure] monetary-amount->locale-string amount intl? [locale]

Convert amount (an inexact denoting a monetary amount) into a string according to the cultural conventions of either locale (a locale object) or the current locale. If intl? is true, then the international monetary format for the given locale is used (see Section “Currency Symbol” in The GNU C Library Reference Manual).

6.24.5 Accessing Locale Information

It is sometimes useful to obtain very specific information about a locale such as the word it uses for days or months, its format for representing floating-point figures, etc. The (ice-9 i18n) module provides support for this in a way that is similar to the libc functions nl_ langinfo () and localeconv () (see Section “Locale Information” in The GNU C Library Reference Manual). The available functions are listed below.

[Scheme Procedure] locale-encoding [locale]

Return the name of the encoding (a string whose interpretation is system-dependent) of either locale or the current locale.

The following functions deal with dates and times.

[Scheme Procedure] locale-day day [locale] [Scheme Procedure] locale-day-short day [locale] [Scheme Procedure] locale-month month [locale] [Scheme Procedure] locale-month-short month [locale]

Return the word (a string) used in either locale or the current locale to name the day (or month) denoted by day (or month), an integer between 1 and 7 (or 1 and 12). The -short variants provide an abbreviation instead of a full name.

[Scheme Procedure] locale-am-string [locale] [Scheme Procedure] locale-pm-string [locale]

Return a (potentially empty) string that is used to denote ante meridiem (or post meridiem) hours in 12-hour format.

[Scheme Procedure] locale-date+time-format [locale] [Scheme Procedure] locale-date-format [locale] [Scheme Procedure] locale-time-format [locale] [Scheme Procedure] locale-time+am/pm-format [locale] [Scheme Procedure] locale-era-date-format [locale]

[Scheme Procedure] locale-era-date+time-format [locale]

[Scheme Procedure] locale-era-time-format [locale]

These procedures return format strings suitable to strftime (see Section 7.2.5 [Time], page 485) that may be used to display (part of) a date/time according to certain constraints and to the conventions of either locale or the current locale (see Section “The Elegant and Fast Way” in The GNU C Library Reference Manual).

[Scheme Procedure] locale-era [locale] [Scheme Procedure] locale-era-year [locale]

These functions return, respectively, the era and the year of the relevant era used in locale or the current locale. Most locales do not define this value. In this case, the empty string is returned. An example of a locale that does define this value is the Japanese one.

The following procedures give information about number representation.

[Scheme Procedure] locale-decimal-point [locale]

[Scheme Procedure] locale-thousands-separator [locale]

These functions return a string denoting the representation of the decimal point or that of the thousand separator (respectively) for either locale or the current locale.

[Scheme Procedure] locale-digit-grouping [locale]

Return a (potentially circular) list of integers denoting how digits of the integer part of a number are to be grouped, starting at the decimal point and going to the left. The list contains integers indicating the size of the successive groups, from right to left. If the list is non-circular, then no grouping occurs for digits beyond the last group. For instance, if the returned list is a circular list that contains only 3 and the thousand separator is "," (as is the case with English locales), then the number 12345678 should be printed 12,345,678.

The following procedures deal with the representation of monetary amounts. Some of them take an additional intl? argument (a boolean) that tells whether the international or local monetary conventions for the given locale are to be used.

[Scheme Procedure] locale-monetary-decimal-point [locale]
[Scheme Procedure] locale-monetary-thousands-separator [locale]

[Scheme Procedure] locale-monetary-grouping [locale]

These are the monetary counterparts of the above procedures. These procedures apply to monetary amounts.

[Scheme Procedure] locale-currency-symbol intl? [locale]

Return the currency symbol (a string) of either locale or the current locale. The following example illustrates the difference between the local and international monetary formats:

     (define us (make-locale LC_MONETARY "en_US"))
     (locale-currency-symbol #f us)
     ⇒ "-$"
     (locale-currency-symbol #t us)
     ⇒ "USD "
[Scheme Procedure] locale-monetary-fractional-digits intl? [locale]

Return the number of fractional digits to be used when printing monetary amounts according to either locale or the current locale. If the locale does not specify it, then #f is returned.

[Scheme Procedure] locale-currency-symbol-precedes-positive? intl?

[locale]

[Scheme Procedure] locale-currency-symbol-precedes-negative? intl?

[locale]

[Scheme Procedure] locale-positive-separated-by-space? intl? [locale]
[Scheme Procedure] locale-negative-separated-by-space? intl? [locale]

These procedures return a boolean indicating whether the currency symbol should precede a positive/negative number, and whether a whitespace should be inserted between the currency symbol and a positive/negative amount.

[Scheme Procedure] locale-monetary-positive-sign [locale]
[Scheme Procedure] locale-monetary-negative-sign [locale]

Return a string denoting the positive (respectively negative) sign that should be used when printing a monetary amount.

[Scheme Procedure] locale-positive-sign-position
[Scheme Procedure] locale-negative-sign-position

These functions return a symbol telling where a sign of a positive/negative monetary amount is to appear when printing it. The possible values are:

parenthesize

The currency symbol and quantity should be surrounded by parentheses.

sign-before

Print the sign string before the quantity and currency symbol.

     sign-after

Print the sign string after the quantity and currency symbol.

     sign-before-currency-symbol

Print the sign string right before the currency symbol.

     sign-after-currency-symbol

Print the sign string right after the currency symbol.

     unspecified

Unspecified. We recommend you print the sign after the currency symbol.

Finally, the two following procedures may be helpful when programming user interfaces:

[Scheme Procedure] locale-yes-regexp [locale] [Scheme Procedure] locale-no-regexp [locale]

Return a string that can be used as a regular expression to recognize a positive (respectively, negative) response to a yes/no question. For the C locale, the default values are typically "^[yY]" and "^[nN]", respectively. Here is an example:

          (use-modules (ice-9 rdelim))
          (format #t "Does Guile rock?~%")
          (let lp ((answer (read-line)))
            (cond ((string-match (locale-yes-regexp) answer)
                   (format #t "High fives!~%"))
                  ((string-match (locale-no-regexp) answer)
                   (format #t "How about now? Does it rock yet?~%")
                   (lp (read-line)))
                  (else
                   (format #t "What do you mean?~%")
                   (lp (read-line)))))

For an internationalized yes/no string output, gettext should be used (see Section 6.24.6 [Gettext Support], page 442).

Example uses of some of these functions are the implementation of the number->locale-string and monetary-amount->locale-string procedures (see Section 6.24.4 [Number Input and Output], page 439), as well as that the SRFI-19 date and time conversion to/from strings (see Section 7.5.16 [SRFI-19], page 583).

6.24.6 Gettext Support

Guile provides an interface to GNU gettext for translating message strings (see Section “Introduction” in GNU gettext utilities).

Messages are collected in domains, so different libraries and programs maintain different message catalogues. The domain parameter in the functions below is a string (it becomes part of the message catalog filename).

When gettext is not available, or if Guile was configured ‘--without-nls’, dummy functions doing no translation are provided. When gettext support is available in Guile, the i18n feature is provided (see Section 6.22.2 [Feature Tracking], page 429).

[Scheme Procedure] gettext msg [domain [category]] [C Function] scm_gettext (msg, domain, category)

Return the translation of msg in domain. domain is optional and defaults to the domain set through textdomain below. category is optional and defaults to LC_ MESSAGES (see Section 7.2.13 [Locales], page 518).

Normal usage is for msg to be a literal string. xgettext can extract those from the source to form a message catalogue ready for translators (see Section “Invoking the xgettext Program” in GNU gettext utilities).

     (display (gettext "You are in a maze of twisty passages."))

_is a commonly used shorthand, an application can make that an alias for gettext. Or a library can make a definition that uses its specific domain (so an application can change the default without affecting the library).

     (define (_msg) (gettext msg "mylibrary"))
     (display (_"File not found."))

_is also a good place to perhaps strip disambiguating extra text from the message string, as for instance in Section “How to use gettext in GUI programs” in GNU gettext utilities.

[Scheme Procedure] ngettext msg msgplural n [domain [category]] [C Function] scm_ngettext (msg, msgplural, n, domain, category)

Return the translation of msg/msgplural in domain, with a plural form chosen appropriately for the number n. domain is optional and defaults to the domain set through textdomain below. category is optional and defaults to LC_MESSAGES (see Section 7.2.13 [Locales], page 518).

msg is the singular form, and msgplural the plural. When no translation is available, msg is used if n = 1, or msgplural otherwise. When translated, the message catalogue can have a different rule, and can have more than two possible forms.

As per gettext above, normal usage is for msg and msgplural to be literal strings, since xgettext can extract them from the source to build a message catalogue. For example,

     (define (done n)
       (format #t (ngettext "~a file processed\n"
                            "~a files processed\n" n)
                  n))
     (done 1) ⊣ 1 file processed
     (done 3) ⊣ 3 files processed

It’s important to use ngettext rather than plain gettext for plurals, since the rules for singular and plural forms in English are not the same in other languages. Only ngettext will allow translators to give correct forms (see Section “Additional functions for plural forms” in GNU gettext utilities).

[Scheme Procedure] textdomain [domain] [C Function] scm_textdomain (domain)

Get or set the default gettext domain. When called with no parameter the current domain is returned. When called with a parameter, domain is set as the current domain, and that new value returned. For example,

     (textdomain "myprog")
     ⇒ "myprog"

[Scheme Procedure] bindtextdomain domain [directory] [C Function] scm_bindtextdomain (domain, directory)

Get or set the directory under which to find message files for domain. When called without a directory the current setting is returned. When called with a directory, directory is set for domain and that new setting returned. For example,

     (bindtextdomain "myprog" "/my/tree/share/locale")
     ⇒ "/my/tree/share/locale"

When using Autoconf/Automake, an application should arrange for the configured localedir to get into the program (by substituting, or by generating a config file) and set that for its domain. This ensures the catalogue can be found even when installed in a non-standard location.

[Scheme Procedure] bind-textdomain-codeset domain [encoding] [C Function] scm_bind_textdomain_codeset (domain, encoding)

Get or set the text encoding to be used by gettext for messages from domain. encoding is a string, the name of a coding system, for instance "8859_1". (On a Unix/POSIX system the iconv program can list all available encodings.) When called without an encoding the current setting is returned, or #f if none yet set. When called with an encoding, it is set for domain and that new setting returned. For example,

     (bind-textdomain-codeset "myprog")
     ⇒ #f
     (bind-textdomain-codeset "myprog" "latin-9")
     ⇒ "latin-9"

The encoding requested can be different from the translated data file, messages will be recoded as necessary. But note that when there is no translation, gettext returns its msg unchanged, ie. without any recoding. For that reason source message strings are best as plain ASCII. Currently Guile has no understanding of multi-byte characters, and string functions won’t recognise character boundaries in multi-byte strings. An application will at least be able to pass such strings through to some output though. Perhaps this will change in the future.

6.25 Debugging Infrastructure

In order to understand Guile’s debugging facilities, you first need to understand a little about how Guile represent the Scheme control stack. With that in place we explain the low level trap calls that the virtual machine can be configured to make, and the trap and breakpoint infrastructure that builds on top of those calls.

6.25.1 Evaluation and the Scheme Stack

The idea of the Scheme stack is central to a lot of debugging. The Scheme stack is a reified representation of the pending function returns in an expression’s continuation. As Guile implements function calls using a stack, this reification takes the form of a number of nested stack frames, each of which corresponds to the application of a procedure to a set of arguments.

A Scheme stack always exists implicitly, and can be summoned into concrete existence as a first-class Scheme value by the make-stack call, so that an introspective Scheme program – such as a debugger – can present it in some way and allow the user to query its details. The first thing to understand, therefore, is how Guile’s function call convention creates the stack.

Broadly speaking, Guile represents all control flow on a stack. Calling a function involves pushing an empty frame on the stack, then evaluating the procedure and its arguments, then fixing up the new frame so that it points to the old one. Frames on the stack are thus linked together. A tail call is the same, except it reuses the existing frame instead of pushing on a new one.

In this way, the only frames that are on the stack are “active” frames, frames which need to do some work before the computation is complete. On the other hand, a function that has tail-called another function will not be on the stack, as it has no work left to do.

Therefore, when an error occurs in a running program, or the program hits a breakpoint, or in fact at any point that the programmer chooses, its state at that point can be represented by a stack of all the procedure applications that are logically in progress at that time, each of which is known as a frame. The programmer can learn more about the program’s state at that point by inspecting the stack and its frames.

6.25.1.1 Stack Capture

A Scheme program can use the make-stack primitive anywhere in its code, with first arg #t, to construct a Scheme value that describes the Scheme stack at that point.

     (make-stack #t)

⇒

     #<stack 25205a0>

Use start-stack to limit the stack extent captured by future make-stack calls.

[Scheme Procedure] make-stack obj arg … [C Function] scm_make_stack (obj, args)

Create a new stack. If obj is #t, the current evaluation stack is used for creating the stack frames, otherwise the frames are taken from obj (which must be a continuation or a frame object). arg … can be any combination of integer, procedure, prompt tag and #t values. These values specify various ways of cutting away uninteresting stack frames from the top and bottom of the stack that make-stack returns. They come in pairs like this:

     (inner_cut_1 outer_cut_1 inner_cut_2 outer_cut_2 ...).

Each inner cut i can be #t, an integer, a prompt tag, or a procedure. #t means to cut away all frames up to but excluding the first user module frame. An integer means to cut away exactly that number of frames. A prompt tag means to cut away all frames that are inside a prompt with the given tag. A procedure means to cut away all frames up to but excluding the application frame whose procedure matches the specified one. Each outer cut i can be an integer, a prompt tag, or a procedure. An integer means to cut away that number of frames. A prompt tag means to cut away all frames that are outside a prompt with the given tag. A procedure means to cut away frames down to but excluding the application frame whose procedure matches the specified one. If the outer cut i of the last pair is missing, it is taken as 0.

[Scheme Syntax] start-stack id exp

Evaluate exp on a new calling stack with identity id. If exp is interrupted during evaluation, backtraces will not display frames farther back than exp’s top-level form. This macro is a way of artificially limiting backtraces and stack procedures, largely as a convenience to the user.

6.25.1.2 Stacks

[Scheme Procedure] stack? obj [C Function] scm_stack_p (obj)

Return #t if obj is a calling stack.

[Scheme Procedure] stack-id stack [C Function] scm_stack_id (stack)

Return the identifier given to stack by start-stack.

[Scheme Procedure] stack-length stack [C Function] scm_stack_length (stack)

Return the length of stack.

[Scheme Procedure] stack-ref stack index

[C Function] scm_stack_ref (stack, index)

Return the index’th frame from stack.

[Scheme Procedure] display-backtrace stack port [first [depth [highlights]]]

[C Function] scm_display_backtrace_with_highlights (stack, port, first,

depth, highlights)

[C Function] scm_display_backtrace (stack, port, first, depth)

Display a backtrace to the output port port. stack is the stack to take the backtrace from, first specifies where in the stack to start and depth how many frames to display. first and depth can be #f, which means that default values will be used. If highlights is given it should be a list; the elements of this list will be highlighted wherever they appear in the backtrace.

6.25.1.3 Frames

[Scheme Procedure] frame? obj [C Function] scm_frame_p (obj)

Return #t if obj is a stack frame.

[Scheme Procedure] frame-previous frame

[C Function] scm_frame_previous (frame)

Return the previous frame of frame, or #f if frame is the first frame in its stack.

[Scheme Procedure] frame-procedure frame

[C Function] scm_frame_procedure (frame)

Return the procedure for frame, or #f if no procedure is associated with frame.

[Scheme Procedure] frame-arguments frame

[C Function] scm_frame_arguments (frame)

Return the arguments of frame.

[Scheme Procedure] frame-address frame

[Scheme Procedure] frame-instruction-pointer frame

[Scheme Procedure] frame-stack-pointer frame

Accessors for the three VM registers associated with this frame: the frame pointer (fp), instruction pointer (ip), and stack pointer (sp), respectively. See Section 9.3.2 [VM Concepts], page 773, for more information.

[Scheme Procedure] frame-dynamic-link frame [Scheme Procedure] frame-return-address frame

[Scheme Procedure] frame-mv-return-address frame

Accessors for the three saved VM registers in a frame: the previous frame pointer, the single-value return address, and the multiple-value return address. See Section 9.3.3 [Stack Layout], page 774, for more information.

[Scheme Procedure] frame-num-locals frame [Scheme Procedure] frame-local-ref frame i [Scheme Procedure] frame-local-set! frame i val

Accessors for the temporary values corresponding to frame’s procedure application. The first local is the first argument given to the procedure. After the arguments, there are the local variables, and after that temporary values. See Section 9.3.3 [Stack Layout], page 774, for more information.

[Scheme Procedure] display-application frame [port [indent]] [C Function] scm_display_application (frame, port, indent)

Display a procedure application frame to the output port port. indent specifies the indentation of the output.

Additionally, the (system vm frame) module defines a number of higher-level introspective procedures, for example to retrieve the names of local variables, and the source location to correspond to a frame. See its source code for more details.

6.25.2 Source Properties

As Guile reads in Scheme code from file or from standard input, it remembers the file name, line number and column number where each expression begins. These pieces of information are known as the source properties of the expression. Syntax expanders and the compiler propagate these source properties to compiled procedures, so that, if an error occurs when evaluating the transformed expression, Guile’s debugger can point back to the file and location where the expression originated.

The way that source properties are stored means that Guile cannot associate source properties with individual symbols, keywords, characters, booleans, or small integers. This can be seen by typing (xxx) and xxx at the Guile prompt (where the variable xxx has not been defined):

     scheme@(guile-user)> (xxx)
     <unnamed port>:4:1: In procedure module-lookup:
     <unnamed port>:4:1: Unbound variable: xxx
     scheme@(guile-user)> xxx
     ERROR: In procedure module-lookup:
     ERROR: Unbound variable: xxx

In the latter case, no source properties were stored, so the error doesn’t have any source information.

[Scheme Procedure] supports-source-properties? obj
[C Function] scm_supports_source_properties_p (obj)

Return #t if source properties can be associated with obj, otherwise return #f.

The recording of source properties is controlled by the read option named “positions” (see Section 6.17.2 [Scheme Read], page 360). This option is switched on by default.

The following procedures can be used to access and set the source properties of read expressions.

[Scheme Procedure] set-source-properties! obj alist

[C Function] scm_set_source_properties_x (obj, alist)

Install the association list alist as the source property list for obj.

[Scheme Procedure] set-source-property! obj key datum [C Function] scm_set_source_property_x (obj, key, datum)

Set the source property of object obj, which is specified by key to datum. Normally, the key will be a symbol.

[Scheme Procedure] source-properties obj

[C Function] scm_source_properties (obj)

Return the source property association list of obj.

[Scheme Procedure] source-property obj key [C Function] scm_source_property (obj, key)

Return the property specified by key from obj’s source properties.

If the positions reader option is enabled, supported expressions will have values set for

the filename, line and column properties.

Source properties are also associated with syntax objects. Procedural macros can get at the source location of their input using the syntax-source accessor. See Section 6.10.4 [Syntax Transformer Helpers], page 268, for more.

Guile also defines a couple of convenience macros built on syntax-source:

[Scheme Syntax] current-source-location

Expands to the source properties corresponding to the location of the (current-

     source-location) form.
[Scheme Syntax] current-filename

Expands to the current filename: the filename that the (current-filename) form appears in. Expands to #f if this information is unavailable.

If you’re stuck with defmacros (see Section 6.10.5 [Defmacros], page 270), and want to preserve source information, the following helper function might be useful to you:

[Scheme Procedure] cons-source xorig x y [C Function] scm_cons_source (xorig, x, y)

Create and return a new pair whose car and cdr are x and y. Any source properties associated with xorig are also associated with the new pair.

6.25.3 Programmatic Error Handling

For better or for worse, all programs have bugs, and dealing with bugs is part of programming. This section deals with that class of bugs that causes an exception to be raised – from your own code, from within a library, or from Guile itself.

6.25.3.1 Catching Exceptions

A common requirement is to be able to show as much useful context as possible when a Scheme program hits an error. The most immediate information about an error is the kind of error that it is – such as “division by zero” – and any parameters that the code which signalled the error chose explicitly to provide. This information originates with the error or throw call (or their C code equivalents, if the error is detected by C code) that signals the error, and is passed automatically to the handler procedure of the innermost applicable

catch or with-throw-handler expression.

Therefore, to catch errors that occur within a chunk of Scheme code, and to intercept basic information about those errors, you need to execute that code inside the dynamic context of a catch or with-throw-handler expression, or the equivalent in C. In Scheme, this means you need something like this:

     (catch #t
            (lambda ()
              ;; Execute the code in which
              ;; you want to catch errors here.
              ...)
            (lambda (key . parameters)
              ;; Put the code which you want
              ;; to handle an error here.
              ...))

The catch here can also be with-throw-handler; see Section 6.13.8.3 [Throw Handlers], page 306 for information on the when you might want to use with-throw-handler instead

of catch.

For example, to print out a message and return #f when an error occurs, you might use:

      (define (catch-all thunk)
        (catch #t
          thunk
          (lambda (key . parameters)
            (format (current-error-port)
                    "Uncaught throw to '~a: ~a\n" key parameters)
            #f)))
      (catch-all
       (lambda () (error "Not a vegetable: tomato")))
       ⊣ Uncaught throw to 'misc-error: (#f ~A (Not a vegetable: tomato) #f)
      ⇒ #f

The #t means that the catch is applicable to all kinds of error. If you want to restrict your catch to just one kind of error, you can put the symbol for that kind of error instead of #t. The equivalent to this in C would be something like this:

     SCM my_body_proc (void *body_data)
     {
       /* Execute the code in which
          you want to catch errors here. */
       ...
     }
     SCM my_handler_proc (void *handler_data,
                          SCM key,
                          SCM parameters)
     {
       /* Put the code which you want
          to handle an error here. */
       ...
     }
     {
       ...
       scm_c_catch (SCM_BOOL_T,
                    my_body_proc, body_data,
                    my_handler_proc, handler_data,
                    NULL, NULL);
       ...
     }

Again, as with the Scheme version, scm_c_catch could be replaced by scm_c_with_throw_ handler, and SCM_BOOL_T could instead be the symbol for a particular kind of error.

6.25.3.2 Capturing the full error stack

The other interesting information about an error is the full Scheme stack at the point where the error occurred; in other words what innermost expression was being evaluated, what was the expression that called that one, and so on. If you want to write your code so that it captures and can display this information as well, there are a couple important things to understand.

Firstly, the stack at the point of the error needs to be explicitly captured by a makestack call (or the C equivalent scm_make_stack). The Guile library does not do this “automatically” for you, so you will need to write code with a make-stack or scm_make_ stack call yourself. (We emphasise this point because some people are misled by the fact that the Guile interactive REPL code does capture and display the stack automatically. But the Guile interactive REPL is itself a Scheme program^{17} running on top of the Guile library, and which uses catch and make-stack in the way we are about to describe to capture the stack when an error occurs.)

And secondly, in order to capture the stack effectively at the point where the error occurred, the make-stack call must be made before Guile unwinds the stack back to the location of the prevailing catch expression. This means that the make-stack call must be made within the handler of a with-throw-handler expression, or the optional "pre-unwind" handler of a catch. (For the full story of how these alternatives differ from each other, see Section 6.13.8 [Exceptions], page 303. The main difference is that catch terminates the error, whereas with-throw-handler only intercepts it temporarily and then allow it to continue propagating up to the next innermost handler.)

So, here are some examples of how to do all this in Scheme and in C. For the purpose of these examples we assume that the captured stack should be stored in a variable, so that it can be displayed or arbitrarily processed later on. In Scheme:

     (let ((captured-stack #f))
       (catch #t
              (lambda ()
                ;; Execute the code in which
                ;; you want to catch errors here.
                ...)
              (lambda (key . parameters)
                ;; Put the code which you want
                ;; to handle an error after the
                ;; stack has been unwound here.
                ...)
              (lambda (key . parameters)
                ;; Capture the stack here:
                (set! captured-stack (make-stack #t))))
       ...
       (if captured-stack
           (begin
             ;; Display or process the captured stack.
             ...))
       ...)

And in C:

     SCM my_body_proc (void *body_data)
     {
       /* Execute the code in which

17 In effect, it is the default program which is run when no commands or script file are specified on the

Guile command line.

   you want to catch errors here. */
...
}
SCM my_handler_proc (void *handler_data,
                   SCM key,
                   SCM parameters)
{
/* Put the code which you want
   to handle an error after the
   stack has been unwound here. */
...
}
SCM my_preunwind_proc (void *handler_data,
                     SCM key,
                     SCM parameters)
{
/* Capture the stack here: */
*(SCM *)handler_data = scm_make_stack (SCM_BOOL_T, SCM_EOL);
}
{
SCM captured_stack = SCM_BOOL_F;
...
scm_c_catch (SCM_BOOL_T,
             my_body_proc, body_data,
             my_handler_proc, handler_data,
             my_preunwind_proc, &captured_stack);
...
if (captured_stack != SCM_BOOL_F)
{
  /* Display or process the captured stack. */
  ...
}
...
}

Once you have a captured stack, you can interrogate and display its details in any way that you want, using the stack-... and frame-... API described in Section 6.25.1.2 [Stacks], page 446 and Section 6.25.1.3 [Frames], page 446.

If you want to print out a backtrace in the same format that the Guile REPL does, you can use the display-backtrace procedure to do so. You can also use displayapplication to display an individual frame in the Guile REPL format.

6.25.3.3 Pre-Unwind Debugging

Instead of saving a stack away and waiting for the catch to return, you can handle errors directly, from within the pre-unwind handler.

For example, to show a backtrace when an error is thrown, you might want to use a procedure like this:

(define (with-backtrace thunk)
  (with-throw-handler #t
                      thunk
                      (lambda args (backtrace))))
(with-backtrace (lambda () (error "Not a vegetable: tomato")))

Since we used with-throw-handler here, we didn’t actually catch the error. See Section 6.13.8.3 [Throw Handlers], page 306, for more information. However, we did print out a context at the time of the error, using the built-in procedure, backtrace.

[Scheme Procedure] backtrace [highlights]

[C Function] scm_backtrace_with_highlights (highlights)
[C Function] scm_backtrace ()

Display a backtrace of the current stack to the current output port. If highlights is given it should be a list; the elements of this list will be highlighted wherever they appear in the backtrace.

The Guile REPL code (in system/repl/repl.scm and related files) uses a catch with a pre-unwind handler to capture the stack when an error occurs in an expression that was typed into the REPL, and debug that stack interactively in the context of the error.

These procedures are available for use by user programs, in the (system repl error-

handling) module.
(use-modules (system repl error-handling))

[Scheme Procedure] call-with-error-handling thunk [#:on-error

on-error=’debug] [#:post-error post-error=’catch] [#:pass-keys pass-keys=’(quit)] [#:trap-handler trap-handler=’debug]

Call a thunk in a context in which errors are handled. There are four keyword arguments:

on-error Specifies what to do before the stack is unwound.

Valid options are debug (the default), which will enter a debugger; pass, in which case nothing is done, and the exception is rethrown; or a procedure, which will be the pre-unwind handler.

post-error Specifies what to do after the stack is unwound.

Valid options are catch (the default), which will silently catch errors, returning the unspecified value; report, which prints out a description of the error (via display-error), and then returns the unspecified value; or a procedure, which will be the catch handler.

trap-handler Specifies a trap handler: what to do when a breakpoint is hit.

Valid options are debug, which will enter the debugger; pass, which does nothing; or disabled, which disables traps entirely. See Section 6.25.4 [Traps], page 455, for more information.

pass-keys A set of keys to ignore, as a list.

6.25.3.4 Debug options

The behavior of the backtrace procedure and of the default error handler can be parameterized via the debug options.

[Scheme Procedure] debug-options [setting]

Display the current settings of the debug options. If setting is omitted, only a short form of the current read options is printed. Otherwise if setting is the symbol help, a complete options description is displayed.

The set of available options, and their default values, may be had by invoking debugoptions at the prompt.

      scheme@(guile-user)>
      backwards       no      Display backtrace in anti-chronological order.
      width           79      Maximal width of backtrace.
      depth           20      Maximal length of printed backtrace.
      backtrace       yes     Show backtrace on error.
      stack           1048576 Stack size limit (measured in words;
                              0 = no check).
      show-file-name  #t      Show file names and line numbers in backtraces
                              when not `#f'.  A value of `base' displays only
                              base names, while `#t' displays full names.
      warn-deprecated no      Warn when deprecated features are used.

The boolean options may be toggled with debug-enable and debug-disable. The non-boolean keywords option must be set using debug-set!.

[Scheme Procedure] debug-enable option-name [Scheme Procedure] debug-disable option-name

[Scheme Syntax] debug-set! option-name value Modify the debug options. debug-enable should be used with boolean options and switches them on, debug-disable switches them off. debug-set! can be used to set an option to a specific value. Due to historical oddities, it is a macro that expects an unquoted option name.

Stack overflow

Stack overflow errors are caused by a computation trying to use more stack space than has been enabled by the stack option. There are actually two kinds of stack that can overflow, the C stack and the Scheme stack.

Scheme stack overflows can occur if Scheme procedures recurse too far deeply. An example would be the following recursive loop:

     scheme@(guile-user)> (let lp () (+ 1 (lp)))
     <unnamed port>:8:17: In procedure vm-run:
     <unnamed port>:8:17: VM: Stack overflow

The default stack size should allow for about 10000 frames or so, so one usually doesn’t hit this level of recursion. Unfortunately there is no way currently to make a VM with a bigger stack. If you are in this unfortunate situation, please file a bug, and in the meantime, rewrite your code to be tail-recursive (see Section 3.3.2 [Tail Calls], page 24).

The other limit you might hit would be C stack overflows. If you call a primitive procedure which then calls a Scheme procedure in a loop, you will consume C stack space. Guile tries to detect excessive consumption of C stack space, throwing an error when you have hit 80% of the process’ available stack (as allocated by the operating system), or 160 kilowords in the absence of a strict limit.

For example, looping through call-with-vm, a primitive that calls a thunk, gives us the following:

     scheme@(guile-user)> (use-modules (system vm vm))
     scheme@(guile-user)> (debug-set! stack 10000)
     scheme@(guile-user)> (let lp () (call-with-vm (the-vm) lp))
     ERROR: In procedure call-with-vm:
     ERROR: Stack overflow

If you get an error like this, you can either try rewriting your code to use less stack space, or increase the maximum stack size. To increase the maximum stack size, use debug-set!, for example:

     (debug-set! stack 200000)

But of course it’s better to have your code operate without so much resource consumption, avoiding loops through C trampolines.

6.25.4 Traps

Guile’s virtual machine can be configured to call out at key points to arbitrary user-specified procedures.

In principle, these hooks allow Scheme code to implement any model it chooses for examining the evaluation stack as program execution proceeds, and for suspending execution to be resumed later.

VM hooks are very low-level, though, and so Guile also has a library of higher-level traps on top of the VM hooks. A trap is an execution condition that, when fulfilled, will fire a handler. For example, Guile defines a trap that fires when control reaches a certain source location.

Finally, Guile also defines a third level of abstractions: per-thread trap states. A trap state exists to give names to traps, and to hold on to the set of traps so that they can be enabled, disabled, or removed. The trap state infrastructure defines the most useful abstractions for most cases. For example, Guile’s REPL uses trap state functions to set breakpoints and tracepoints.

The following subsections describe all this in detail, for both the user wanting to use traps, and the developer interested in understanding how the interface hangs together.

6.25.4.1 VM Hooks

Everything that runs in Guile runs on its virtual machine, a C program that defines a number of operations that Scheme programs can perform.

Note that there are multiple VM “engines” for Guile. Only some of them have support for hooks compiled in. Normally the deal is that you get hooks if you are running interactively, and otherwise they are disabled, as they do have some overhead (about 10 or 20 percent).

To ensure that you are running with hooks, pass --debug to Guile when running your program, or otherwise use the call-with-vm and set-vm-engine! procedures to ensure that you are running in a VM with the debug engine.

To digress, Guile’s VM has 6 different hooks (see Section 6.11.6 [Hooks], page 281) that can be fired at different times, which may be accessed with the following procedures.

All hooks are called with one argument, the frame in question. See Section 6.25.1.3 [Frames], page 446. Since these hooks may be fired very frequently, Guile does a terrible thing: it allocates the frames on the C stack instead of the garbage-collected heap.

The upshot here is that the frames are only valid within the dynamic extent of the call to the hook. If a hook procedure keeps a reference to the frame outside the extent of the hook, bad things will happen.

The interface to hooks is provided by the (system vm vm) module:

(use-modules (system vm vm))

The result of calling the-vm is usually passed as the vm argument to all of these procedures.

[Scheme Procedure] vm-next-hook vm

The hook that will be fired before an instruction is retired (and executed).

[Scheme Procedure] vm-push-continuation-hook vm

The hook that will be fired after preparing a new frame. Fires just before applying a procedure in a non-tail context, just before the corresponding apply-hook.

[Scheme Procedure] vm-pop-continuation-hook vm

The hook that will be fired before returning from a frame.

This hook is a bit trickier than the rest, in that there is a particular interpretation of the values on the stack. Specifically, the top value on the stack is the number of values being returned, and the next n values are the actual values being returned, with the last value highest on the stack.

[Scheme Procedure] vm-apply-hook vm

The hook that will be fired before a procedure is applied. The frame’s procedure will have already been set to the new procedure.

Note that procedure application is somewhat orthogonal to continuation pushes and pops. A non-tail call to a procedure will result first in a firing of the push-continuation hook, then this application hook, whereas a tail call will run without having fired a push-continuation hook.

[Scheme Procedure] vm-abort-continuation-hook vm

The hook that will be called after aborting to a prompt. See Section 6.13.5 [Prompts], page 296. The stack will be in the same state as for vm-pop-continuation-hook.

[Scheme Procedure] vm-restore-continuation-hook vm

The hook that will be called after restoring an undelimited continuation. Unfortunately it’s not currently possible to introspect on the values that were given to the continuation.

These hooks do impose a performance penalty, if they are on. Obviously, the vm-next-hook has quite an impact, performance-wise. Therefore Guile exposes a single, heavyhanded knob to turn hooks on or off, the VM trace level. If the trace level is positive, hooks run; otherwise they don’t.

For convenience, when the VM fires a hook, it does so with the trap level temporarily set to 0. That way the hooks don’t fire while you’re handling a hook. The trace level is restored to whatever it was once the hook procedure finishes.

[Scheme Procedure] vm-trace-level vm

Retrieve the “trace level” of the VM. If positive, the trace hooks associated with vm will be run. The initial trace level is 0.

[Scheme Procedure] set-vm-trace-level! vm level

Set the “trace level” of the VM.

See Section 9.3 [A Virtual Machine for Guile], page 772, for more information on Guile’s virtual machine.

6.25.4.2 Trap Interface

The capabilities provided by hooks are great, but hooks alone rarely correspond to what users want to do.

For example, if a user wants to break when and if control reaches a certain source location, how do you do it? If you install a “next” hook, you get unacceptable overhead for the execution of the entire program. It would be possible to install an “apply” hook, then if the procedure encompasses those source locations, install a “next” hook, but already you’re talking about one concept that might be implemented by a varying number of lower-level concepts.

It’s best to be clear about things and define one abstraction for all such conditions: the trap.

Considering the myriad capabilities offered by the hooks though, there is only a minimum of functionality shared by all traps. Guile’s current take is to reduce this to the absolute minimum, and have the only standard interface of a trap be “turn yourself on” or “turn yourself off”.

This interface sounds a bit strange, but it is useful to procedurally compose higher-level traps from lower-level building blocks. For example, Guile defines a trap that calls one handler when control enters a procedure, and another when control leaves the procedure. Given that trap, one can define a trap that adds to the next-hook only when within a given procedure. Building further, one can define a trap that fires when control reaches particular instructions within a procedure.

Or of course you can stop at any of these intermediate levels. For example, one might only be interested in calls to a given procedure. But the point is that a simple enable/disable interface is all the commonality that exists between the various kinds of traps, and furthermore that such an interface serves to allow “higher-level” traps to be composed from more primitive ones.

Specifically, a trap, in Guile, is a procedure. When a trap is created, by convention the trap is enabled; therefore, the procedure that is the trap will, when called, disable the trap, and return a procedure that will enable the trap, and so on.

Trap procedures take one optional argument: the current frame. (A trap may want to add to different sets of hooks depending on the frame that is current at enable-time.)

If this all sounds very complicated, it’s because it is. Some of it is essential, but probably most of it is not. The advantage of using this minimal interface is that composability is more lexically apparent than when, for example, using a stateful interface based on GOOPS. But perhaps this reflects the cognitive limitations of the programmer who made the current interface more than anything else.

6.25.4.3 Low-Level Traps

To summarize the last sections, traps are enabled or disabled, and when they are enabled, they add to various VM hooks.

Note, however, that traps do not increase the VM trace level. So if you create a trap, it will be enabled, but unless something else increases the VM’s trace level (see Section 6.25.4.1 [VM Hooks], page 455), the trap will not fire. It turns out that getting the VM trace level right is tricky without a global view of what traps are enabled. See Section 6.25.4.5 [Trap States], page 461, for Guile’s answer to this problem.

Traps are created by calling procedures. Most of these procedures share a set of common keyword arguments, so rather than document them separately, we discuss them all together here:

#:vm The VM to instrument. Defaults to the current thread’s VM.

#:closure?

For traps that depend on the current frame’s procedure, this argument specifies whether to trap on the only the specific procedure given, or on any closure that has the given procedure’s code. Defaults to #f.

#:current-frame

For traps that enable more hooks depending on their dynamic context, this argument gives the current frame that the trap is running in. Defaults to #f.

To have access to these procedures, you’ll need to have imported the (system vm traps) module:

     (use-modules (system vm traps))

[Scheme Procedure] trap-at-procedure-call proc handler [#:vm] [#:closure?]

A trap that calls handler when proc is applied.

[Scheme Procedure] trap-in-procedure proc enter-handler exit-handler

[#:current-frame] [#:vm] [#:closure?] A trap that calls enter-handler when control enters proc, and exit-handler when control leaves proc. Control can enter a procedure via:

• A procedure call.

• A return to a procedure’s frame on the stack.

• A continuation returning directly to an application of this procedure.

Control can leave a procedure via:

• A normal return from the procedure.

• An application of another procedure.

• An invocation of a continuation.

• An abort.

[Scheme Procedure] trap-instructions-in-procedure proc next-handler

exit-handler [#:current-frame] [#:vm] [#:closure?] A trap that calls next-handler for every instruction executed in proc, and exit-handler when execution leaves proc.

[Scheme Procedure] trap-at-procedure-ip-in-range proc range handler

[#:current-frame] [#:vm] [#:closure?] A trap that calls handler when execution enters a range of instructions in proc. range is a simple of pairs, ((start . end) ...). The start addresses are inclusive, and end addresses are exclusive.

[Scheme Procedure] trap-at-source-location file user-line handler

[#:current-frame] [#:vm] A trap that fires when control reaches a given source location. The user-line parameter is one-indexed, as a user counts lines, instead of zero-indexed, as Guile counts lines.

[Scheme Procedure] trap-frame-finish frame return-handler abort-handler [#:vm] A trap that fires when control leaves the given frame. frame should be a live frame in the current continuation. return-handler will be called on a normal return, and abort-handler on a nonlocal exit.

[Scheme Procedure] trap-in-dynamic-extent proc enter-handler return-handler

abort-handler [#:vm] [#:closure?] A more traditional dynamic-wind trap, which fires enter-handler when control enters proc, return-handler on a normal return, and abort-handler on a nonlocal exit. Note that rewinds are not handled, so there is no rewind handler.

[Scheme Procedure] trap-calls-in-dynamic-extent proc apply-handler

return-handler [#:current-frame] [#:vm] [#:closure?] A trap that calls apply-handler every time a procedure is applied, and return-handler for returns, but only during the dynamic extent of an application of proc.

[Scheme Procedure] trap-instructions-in-dynamic-extent proc next-handler

[#:current-frame] [#:vm] [#:closure?] A trap that calls next-handler for all retired instructions within the dynamic extent of a call to proc.

[Scheme Procedure] trap-calls-to-procedure proc apply-handler return-handler

[#:vm] A trap that calls apply-handler whenever proc is applied, and return-handler when it returns, but with an additional argument, the call depth. That is to say, the handlers will get two arguments: the frame in question, and the call depth (a non-negative integer).

[Scheme Procedure] trap-matching-instructions frame-pred handler [#:vm]

A trap that calls frame-pred at every instruction, and if frame-pred returns a true value, calls handler on the frame.

6.25.4.4 Tracing Traps

The (system vm trace) module defines a number of traps for tracing of procedure applications. When a procedure is traced, it means that every call to that procedure is reported to the user during a program run. The idea is that you can mark a collection of procedures for tracing, and Guile will subsequently print out a line of the form

|  |  (procedure args ...)

whenever a marked procedure is about to be applied to its arguments. This can help a programmer determine whether a function is being called at the wrong time or with the wrong set of arguments.

In addition, the indentation of the output is useful for demonstrating how the traced applications are or are not tail recursive with respect to each other. Thus, a trace of a non-tail recursive factorial implementation looks like this:

scheme@(guile-user)> (define (fact1 n)
                       (if (zero? n) 1
                           (* n (fact1 (1- n)))))
scheme@(guile-user)> ,trace (fact1 4)
trace: (fact1 4)
trace: |  (fact1 3)
trace: |  |  (fact1 2)
trace: |  |  |  (fact1 1)
trace: |  |  |  |  (fact1 0)
trace: |  |  |  |  1
trace: |  |  |  1
trace: |  |  2
trace: |  6
trace: 24

While a typical tail recursive implementation would look more like this:

scheme@(guile-user)> (define (facti acc n)
                       (if (zero? n) acc
                           (facti (* n acc) (1- n))))
scheme@(guile-user)> (define (fact2 n) (facti 1 n))
scheme@(guile-user)> ,trace (fact2 4)
trace: (fact2 4)
trace: (facti 1 4)
trace: (facti 4 3)
trace: (facti 12 2)
trace: (facti 24 1)
trace: (facti 24 0)
trace: 24

The low-level traps below (see Section 6.25.4.3 [Low-Level Traps], page 458) share some common options:

#:width The maximum width of trace output. Trace printouts will try not to exceed this column, but for highly nested procedure calls, it may be unavoidable. Defaults to 80.

#:vm The VM on which to add the traps. Defaults to the current thread’s VM.

#:prefix A string to print out before each trace line. As seen above in the examples, defaults to "trace: ".

To have access to these procedures, you’ll need to have imported the (system vm trace) module:

(use-modules (system vm trace))

[Scheme Procedure] trace-calls-to-procedure proc [#:width] [#:vm] [#:prefix]

Print a trace at applications of and returns from proc.

[Scheme Procedure] trace-calls-in-procedure proc [#:width] [#:vm] [#:prefix]

Print a trace at all applications and returns within the dynamic extent of calls to proc.

[Scheme Procedure] trace-instructions-in-procedure proc [#:width] [#:vm]

Print a trace at all instructions executed in the dynamic extent of calls to proc.

In addition, Guile defines a procedure to call a thunk, tracing all procedure calls and returns within the thunk.

[Scheme Procedure] call-with-trace thunk [#:calls?=#t] [#:instructions?=#f]

[#:width=80] [#:vm=(the-vm)] Call thunk, tracing all execution within its dynamic extent. If calls? is true, Guile will print a brief report at each procedure call and return, as given above. If instructions? is true, Guile will also print a message each time an instruction is executed. This is a lot of output, but it is sometimes useful when doing low-level optimization. Note that because this procedure manipulates the VM trace level directly, it doesn’t compose well with traps at the REPL.

See Section 4.4.4.5 [Profile Commands], page 49, for more information on tracing at the REPL.

6.25.4.5 Trap States

When multiple traps are present in a system, we begin to have a bookkeeping problem. How are they named? How does one disable, enable, or delete them?

Guile’s answer to this is to keep an implicit per-thread trap state. The trap state object is not exposed to the user; rather, API that works on trap states fetches the current trap state from the dynamic environment.

Traps are identified by integers. A trap can be enabled, disabled, or removed, and can have an associated user-visible name.

These procedures have their own module:

(use-modules (system vm trap-state))

[Scheme Procedure] add-trap! trap name

Add a trap to the current trap state, associating the given name with it. Returns a fresh trap identifier (an integer). Note that usually the more specific functions detailed in Section 6.25.4.6 [High-Level Traps], page 462 are used in preference to this one.

[Scheme Procedure] list-traps

List the current set of traps, both enabled and disabled. Returns a list of integers.

[Scheme Procedure] trap-name idx

Returns the name associated with trap idx, or #f if there is no such trap.

[Scheme Procedure] trap-enabled? idx

Returns #t if trap idx is present and enabled, or #f otherwise.

[Scheme Procedure] enable-trap! idx

Enables trap idx.

[Scheme Procedure] disable-trap! idx

Disables trap idx.

[Scheme Procedure] delete-trap! idx

Removes trap idx, disabling it first, if necessary.

6.25.4.6 High-Level Traps

The low-level trap API allows one to make traps that call procedures, and the trap state API allows one to keep track of what traps are there. But neither of these APIs directly helps you when you want to set a breakpoint, because it’s unclear what to do when the trap fires. Do you enter a debugger, or mail a summary of the situation to your great-aunt, or what?

So for the common case in which you just want to install breakpoints, and then have them all result in calls to one parameterizable procedure, we have the high-level trap interface.

Perhaps we should have started this section with this interface, as it’s clearly the one most people should use. But as its capabilities and limitations proceed from the lower layers, we felt that the character-building exercise of building a mental model might be helpful.

These procedures share a module with trap states:

(use-modules (system vm trap-state))

[Scheme Procedure] with-default-trap-handler handler thunk

Call thunk in a dynamic context in which handler is the current trap handler. Additionally, during the execution of thunk, the VM trace level (see Section 6.25.4.1 [VM Hooks], page 455) is set to the number of enabled traps. This ensures that traps will in fact fire. handler may be #f, in which case VM hooks are not enabled as they otherwise would be, as there is nothing to handle the traps.

The trace-level-setting behavior of with-default-trap-handler is one of its more useful aspects, but if you are willing to forgo that, and just want to install a global trap handler, there’s a function for that too:

[Scheme Procedure] install-trap-handler! handler

Set the current thread’s trap handler to handler.

Trap handlers are called when traps installed by procedures from this module fire. The current “consumer” of this API is Guile’s REPL, but one might easily imagine other trap handlers being used to integrate with other debugging tools.

[Scheme Procedure] add-trap-at-procedure-call! proc

Install a trap that will fire when proc is called. This is a breakpoint.

[Scheme Procedure] add-trace-at-procedure-call! proc

Install a trap that will print a tracing message when proc is called. See Section 6.25.4.4 [Tracing Traps], page 460, for more information. This is a tracepoint.

[Scheme Procedure] add-trap-at-source-location! file user-line

Install a trap that will fire when control reaches the given source location. user-line is one-indexed, as users count lines, instead of zero-indexed, as Guile counts lines. This is a source breakpoint.

[Scheme Procedure] add-ephemeral-trap-at-frame-finish! frame handler

Install a trap that will call handler when frame finishes executing. The trap will be removed from the trap state after firing, or on nonlocal exit. This is a finish trap, used to implement the “finish” REPL command.

[Scheme Procedure] add-ephemeral-stepping-trap! frame handler [#:into?]

[#:instruction?] Install a trap that will call handler after stepping to a different source line or instruction. The trap will be removed from the trap state after firing, or on nonlocal exit. If instruction? is false (the default), the trap will fire when control reaches a new source line. Otherwise it will fire when control reaches a new instruction. Additionally, if into? is false (not the default), the trap will only fire for frames at or prior to the given frame. If into? is true (the default), the trap may step into nested procedure invocations. This is a stepping trap, used to implement the “step”, “next”, “step-instruction”, and “next-instruction” REPL commands.

6.25.5 GDB Support

Sometimes, you may find it necessary to debug Guile applications at the C level. Doing so can be tedious, in particular because the debugger is oblivious to Guile’s SCM type, and thus unable to display SCM values in any meaningful way:

(gdb) frame
#0  scm_display (obj=0xf04310, port=0x6f9f30) at print.c:1437

To address that, Guile comes with an extension of the GNU Debugger (GDB) that contains a “pretty-printer” for SCM values. With this GDB extension, the C frame in the example above shows up like this:

(gdb) frame
#0  scm_display (obj=("hello" GDB!), port=#<port file 6f9f30>) at print.c:1437

Here GDB was able to decode the list pointed to by obj, and to print it using Scheme’s read syntax.

That extension is a .scm file installed alongside the libguile shared library. When GDB 7.8 or later is installed and compiled with support for extensions written in Guile, the extension is automatically loaded when debugging a program linked against libguile (see Section “Auto-loading” in Debugging with GDB). Note that the directory where libguile is installed must be among GDB’s auto-loading “safe directories” (see Section “Auto-loading safe path” in Debugging with GDB).

6.26 Code Coverage Reports

When writing a test suite for a program or library, it is desirable to know what part of the code is covered by the test suite. The (system vm coverage) module provides tools to gather code coverage data and to present them, as detailed below.

[Scheme Procedure] with-code-coverage vm thunk

Run thunk, a zero-argument procedure, using vm; instrument vm to collect code coverage data. Return code coverage data and the values returned by thunk.

[Scheme Procedure] coverage-data? obj

Return #t if obj is a coverage data object as returned by with-code-coverage.

[Scheme Procedure] coverage-data->lcov data port #:key modules Traverse code coverage information data, as obtained with with-code-coverage, and write coverage information to port in the .info format used by LCOV (http://ltp. sourceforge.net/coverage/lcov.php). The report will include all of modules (or, by default, all the currently loaded modules) even if their code was not executed. The generated data can be fed to LCOV’s genhtml command to produce an HTML report, which aids coverage data visualization.

Here’s an example use:

(use-modules (system vm coverage)
             (system vm vm))
(call-with-values (lambda ()
                    (with-code-coverage (the-vm)
                      (lambda ()
                        (do-something-tricky))))
  (lambda (data result)
    (let ((port (open-output-file "lcov.info")))

      (coverage-data->lcov data port)
      (close file))))

In addition, the module provides low-level procedures that would make it possible to write other user interfaces to the coverage data.

[Scheme Procedures] instrumented-source-files data

Return the list of “instrumented” source files, i.e., source files whose code was loaded at the time data was collected.

[Scheme Procedures] line-execution-counts data file

Return a list of line number/execution count pairs for file, or #f if file is not among the files covered by data. This includes lines with zero count.

[Scheme Procedures] instrumented/executed-lines data file

Return the number of instrumented and the number of executed source lines in file according to data.

[Scheme Procedures] procedure-execution-count data proc

Return the number of times proc’s code was executed, according to data, or #f if proc was not executed. When proc is a closure, the number of times its code was executed is returned, not the number of times this code associated with this particular closure was executed.

7 Guile Modules

7.1 SLIB

SLIB is a portable library of Scheme packages which can be used with Guile and other Scheme implementations. SLIB is not included in the Guile distribution, but can be installed separately (see Section 7.1.1 [SLIB installation], page 467). It is available from http://

people.csail.mit.edu/jaffer/SLIB.html.

After SLIB is installed, the following Scheme expression must be executed before the SLIB facilities can be used:

     (use-modules (ice-9 slib))

require can then be used in the usual way (see Section “Require” in The SLIB Manual). For example,

     (use-modules (ice-9 slib))
     (require 'primes)
     (prime? 13)
     ⇒ #t

A few Guile core functions are overridden by the SLIB setups; for example the SLIB version of delete-file returns a boolean indicating success or failure, whereas the Guile core version throws an error for failure. In general (and as might be expected) when SLIB is loaded it’s the SLIB specifications that are followed.

7.1.1 SLIB installation

The following procedure works, e.g., with SLIB version 3a3 (see Section “Installation” in The SLIB Portable Scheme Library):

1. Unpack SLIB and install it using make install from its directory. By default, this

    will install SLIB in /usr/local/lib/slib/. Running make install-info installs its

documentation, by default under /usr/local/info/.

2. Define the SCHEME_LIBRARY_PATH environment variable:

         $ SCHEME_LIBRARY_PATH=/usr/local/lib/slib/
         $ export SCHEME_LIBRARY_PATH

Alternatively, you can create a symlink in the Guile directory to SLIB, e.g.:

         ln -s /usr/local/lib/slib /usr/local/share/guile/2.0/slib

3. Use Guile to create the catalog file, e.g.,:

         # guile
         guile> (use-modules (ice-9 slib))
         guile> (require 'new-catalog)
         guile> (quit)
    The catalog data should now be in /usr/local/share/guile/2.0/slibcat.

If instead you get an error such as:

         Unbound variable: scheme-implementation-type

then a solution is to get a newer version of Guile, or to modify ice-9/slib.scm to use define-public for the offending variables.

7.1.2 JACAL

Jacal is a symbolic math package written in Scheme by Aubrey Jaffer. It is usually installed as an extra package in SLIB.

You can use Guile’s interface to SLIB to invoke Jacal:

     (use-modules (ice-9 slib))
     (slib:load "math")
     (math)

For complete documentation on Jacal, please read the Jacal manual. If it has been installed on line, you can look at Section “Jacal” in JACAL Symbolic Mathematics System. Oth-

erwise you can find it on the web at http://www-swiss.ai.mit.edu/~jaffer/JACAL.
html

7.2 POSIX System Calls and Networking

7.2.1 POSIX Interface Conventions

These interfaces provide access to operating system facilities. They provide a simple wrapping around the underlying C interfaces to make usage from Scheme more convenient. They are also used to implement the Guile port of scsh (see Section 7.17 [The Scheme shell (scsh)], page 693).

Generally there is a single procedure for each corresponding Unix facility. There are some exceptions, such as procedures implemented for speed and convenience in Scheme with no primitive Unix equivalent, e.g. copy-file.

The interfaces are intended as far as possible to be portable across different versions of Unix. In some cases procedures which can’t be implemented on particular systems may become no-ops, or perform limited actions. In other cases they may throw errors.

General naming conventions are as follows:

• The Scheme name is often identical to the name of the underlying Unix facility.

• Underscores in Unix procedure names are converted to hyphens.

• Procedures which destructively modify Scheme data have exclamation marks appended, e.g., recv!.

• Predicates (returning only #t or #f) have question marks appended, e.g., access?.

• Some names are changed to avoid conflict with dissimilar interfaces defined by scsh,

    e.g., primitive-fork.

• Unix preprocessor names such as EPERM or R_OK are converted to Scheme variables of the same name (underscores are not replaced with hyphens).

Unexpected conditions are generally handled by raising exceptions. There are a few procedures which return a special value if they don’t succeed, e.g., getenv returns #f if it the requested string is not found in the environment. These cases are noted in the documentation.

For ways to deal with exceptions, see Section 6.13.8 [Exceptions], page 303.

Errors which the C library would report by returning a null pointer or through some other means are reported by raising a system-error exception with scm-error (see Section 6.13.9 [Error Reporting], page 309). The data parameter is a list containing the Unix errno value (an integer). For example,

(define (my-handler key func fmt fmtargs data)
  (display key) (newline)
  (display func) (newline)
  (apply format #t fmt fmtargs) (newline)
  (display data) (newline))
(catch 'system-error
  (lambda () (dup2 -123 -456))
  my-handler)

⊣

system-error
dup2
Bad file descriptor
(9)
[Function] system-error-errno arglist

Return the errno value from a list which is the arguments to an exception handler. If the exception is not a system-error, then the return is #f. For example,

     (catch
      'system-error
      (lambda ()
        (mkdir "/this-ought-to-fail-if-I'm-not-root"))
      (lambda stuff
        (let ((errno (system-error-errno stuff)))
          (cond
           ((= errno EACCES)
            (display "You're not allowed to do that."))
           ((= errno EEXIST)
            (display "Already exists."))
           (#t
            (display (strerror errno))))
          (newline))))

7.2.2 Ports and File Descriptors

Conventions generally follow those of scsh, Section 7.17 [The Scheme shell (scsh)], page 693.

File ports are implemented using low-level operating system I/O facilities, with optional buffering to improve efficiency; see Section 6.14.9.1 [File Ports], page 326.

Note that some procedures (e.g., recv!) will accept ports as arguments, but will actually operate directly on the file descriptor underlying the port. Any port buffering is ignored, including the buffer which implements peek-char and unread-char.

The force-output and drain-input procedures can be used to clear the buffers.

Each open file port has an associated operating system file descriptor. File descriptors are generally not useful in Scheme programs; however they may be needed when interfacing with foreign code and the Unix environment.

A file descriptor can be extracted from a port and a new port can be created from a file descriptor. However a file descriptor is just an integer and the garbage collector doesn’t recognize it as a reference to the port. If all other references to the port were dropped, then it’s likely that the garbage collector would free the port, with the side-effect of closing the file descriptor prematurely.

To assist the programmer in avoiding this problem, each port has an associated revealed count which can be used to keep track of how many times the underlying file descriptor has been stored in other places. If a port’s revealed count is greater than zero, the file descriptor will not be closed when the port is garbage collected. A programmer can therefore ensure that the revealed count will be greater than zero if the file descriptor is needed elsewhere.

For the simple case where a file descriptor is “imported” once to become a port, it does not matter if the file descriptor is closed when the port is garbage collected. There is no need to maintain a revealed count. Likewise when “exporting” a file descriptor to the external environment, setting the revealed count is not required provided the port is kept open (i.e., is pointed to by a live Scheme binding) while the file descriptor is in use.

To correspond with traditional Unix behaviour, three file descriptors (0, 1, and 2) are automatically imported when a program starts up and assigned to the initial values of the current/standard input, output, and error ports, respectively. The revealed count for each is initially set to one, so that dropping references to one of these ports will not result in its garbage collection: it could be retrieved with fdopen or fdes->ports.

[Scheme Procedure] port-revealed port

[C Function] scm_port_revealed (port)

Return the revealed count for port.

[Scheme Procedure] set-port-revealed! port rcount [C Function] scm_set_port_revealed_x (port, rcount)

Sets the revealed count for a port to rcount. The return value is unspecified.

[Scheme Procedure] fileno port [C Function] scm_fileno (port)

Return the integer file descriptor underlying port. Does not change its revealed count.

[Scheme Procedure] port->fdes port

Returns the integer file descriptor underlying port. As a side effect the revealed count of port is incremented.

[Scheme Procedure] fdopen fdes modes [C Function] scm_fdopen (fdes, modes)

Return a new port based on the file descriptor fdes. Modes are given by the string modes. The revealed count of the port is initialized to zero. The modes string is the same as that accepted by open-file (see Section 6.14.9.1 [File Ports], page 326).

[Scheme Procedure] fdes->ports fdes

[C Function] scm_fdes_to_ports (fdes)

Return a list of existing ports which have fdes as an underlying file descriptor, without changing their revealed counts.

[Scheme Procedure] fdes->inport fdes

Returns an existing input port which has fdes as its underlying file descriptor, if one exists, and increments its revealed count. Otherwise, returns a new input port with a revealed count of 1.

[Scheme Procedure] fdes->outport fdes

Returns an existing output port which has fdes as its underlying file descriptor, if one exists, and increments its revealed count. Otherwise, returns a new output port with a revealed count of 1.

[Scheme Procedure] primitive-move->fdes port fdes

[C Function] scm_primitive_move_to_fdes (port, fdes)

Moves the underlying file descriptor for port to the integer value fdes without changing the revealed count of port. Any other ports already using this descriptor will be automatically shifted to new descriptors and their revealed counts reset to zero. The return value is #f if the file descriptor already had the required value or #t if it was moved.

[Scheme Procedure] move->fdes port fdes

Moves the underlying file descriptor for port to the integer value fdes and sets its revealed count to one. Any other ports already using this descriptor will be automatically shifted to new descriptors and their revealed counts reset to zero. The return value is unspecified.

[Scheme Procedure] release-port-handle port

Decrements the revealed count for a port.

[Scheme Procedure] fsync port or fd [C Function] scm_fsync (port or fd)

Copies any unwritten data for the specified output file descriptor to disk. If port or fd is a port, its buffer is flushed before the underlying file descriptor is fsync’d. The return value is unspecified.

[Scheme Procedure] open path flags [mode] [C Function] scm_open (path, flags, mode)

Open the file named by path for reading and/or writing. flags is an integer specifying how the file should be opened. mode is an integer specifying the permission bits of the file, if it needs to be created, before the umask (see Section 7.2.7 [Processes], page 490) is applied. The default is 666 (Unix itself has no default). flags can be constructed by combining variables using logior. Basic flags are:

[Variable] O_RDONLY

Open the file read-only.

[Variable] O_WRONLY

Open the file write-only.

[Variable] O_RDWR

Open the file read/write.

[Variable] O_APPEND

Append to the file instead of truncating.

[Variable] O_CREAT

Create the file if it does not already exist.

See Section “File Status Flags” in The GNU C Library Reference Manual, for additional flags.

[Scheme Procedure] open-fdes path flags [mode] [C Function] scm_open_fdes (path, flags, mode)

Similar to open but return a file descriptor instead of a port.

[Scheme Procedure] close fd or port [C Function] scm_close (fd or port)

Similar to close-port (see Section 6.14.4 [Closing], page 321), but also works on file descriptors. A side effect of closing a file descriptor is that any ports using that file descriptor are moved to a different file descriptor and have their revealed counts set to zero.

[Scheme Procedure] close-fdes fd [C Function] scm_close_fdes (fd)

A simple wrapper for the close system call. Close file descriptor fd, which must be an integer. Unlike close, the file descriptor will be closed even if a port is using it. The return value is unspecified.

[Scheme Procedure] unread-char char [port] [C Function] scm_unread_char (char, port)

Place char in port so that it will be read by the next read operation on that port. If called multiple times, the unread characters will be read again in “last-in, first-out” order (i.e. a stack). If port is not supplied, the current input port is used.

[Scheme Procedure] unread-string str port

Place the string str in port so that its characters will be read in subsequent read operations. If called multiple times, the unread characters will be read again in lastin first-out order. If port is not supplied, the current-input-port is used.

[Scheme Procedure] pipe [C Function] scm_pipe ()

Return a newly created pipe: a pair of ports which are linked together on the local machine. The CAR is the input port and the CDR is the output port. Data written (and flushed) to the output port can be read from the input port. Pipes are commonly used for communication with a newly forked child process. The need to flush the output port can be avoided by making it unbuffered using setvbuf.

[Variable] PIPE_BUF

A write of up to PIPE_BUF many bytes to a pipe is atomic, meaning when done it goes into the pipe instantaneously and as a contiguous block (see Section “Atomicity of Pipe I/O” in The GNU C Library Reference Manual).

Note that the output port is likely to block if too much data has been written but not yet read from the input port. Typically the capacity is PIPE_BUF bytes.

The next group of procedures perform a dup2 system call, if newfd (an integer) is supplied, otherwise a dup. The file descriptor to be duplicated can be supplied as an integer or contained in a port. The type of value returned varies depending on which procedure is used.

All procedures also have the side effect when performing dup2 that any ports using newfd are moved to a different file descriptor and have their revealed counts set to zero.

[Scheme Procedure] dup->fdes fd or port [fd] [C Function] scm_dup_to_fdes (fd or port, fd)

Return a new integer file descriptor referring to the open file designated by fd or port, which must be either an open file port or a file descriptor.

[Scheme Procedure] dup->inport port/fd [newfd]

Returns a new input port using the new file descriptor.

[Scheme Procedure] dup->outport port/fd [newfd]

Returns a new output port using the new file descriptor.

[Scheme Procedure] dup port/fd [newfd] Returns a new port if port/fd is a port, with the same mode as the supplied port, otherwise returns an integer file descriptor.

[Scheme Procedure] dup->port port/fd mode [newfd] Returns a new port using the new file descriptor. mode supplies a mode string for the port (see Section 6.14.9.1 [File Ports], page 326).

[Scheme Procedure] duplicate-port port modes

Returns a new port which is opened on a duplicate of the file descriptor underlying port, with mode string modes as for Section 6.14.9.1 [File Ports], page 326. The two ports will share a file position and file status flags. Unexpected behaviour can result if both ports are subsequently used and the original and/or duplicate ports are buffered. The mode string can include 0 to obtain an unbuffered duplicate port. This procedure is equivalent to (dup->port port modes).

[Scheme Procedure] redirect-port old port new port

[C Function] scm_redirect_port (old port, new port)

This procedure takes two ports and duplicates the underlying file descriptor from old port into new port. The current file descriptor in new port will be closed. After the redirection the two ports will share a file position and file status flags. The return value is unspecified.

Unexpected behaviour can result if both ports are subsequently used and the original and/or duplicate ports are buffered. This procedure does not have any side effects on other ports or revealed counts.

[Scheme Procedure] dup2 oldfd newfd [C Function] scm_dup2 (oldfd, newfd)

A simple wrapper for the dup2 system call. Copies the file descriptor oldfd to descriptor number newfd, replacing the previous meaning of newfd. Both oldfd and newfd must be integers. Unlike for dup->fdes or primitive-move->fdes, no attempt is made to move away ports which are using newfd. The return value is unspecified.

[Scheme Procedure] port-mode port

Return the port modes associated with the open port port. These will not necessarily be identical to the modes used when the port was opened, since modes such as “append” which are used only during port creation are not retained.

[Scheme Procedure] port-for-each proc [C Function] scm_port_for_each (SCM proc)

[C Function] scm_c_port_for_each (void (*proc)(void *, SCM), void *data)

Apply proc to each port in the Guile port table (FIXME: what is the Guile port table?) in turn. The return value is unspecified. More specifically, proc is applied exactly once to every port that exists in the system at the time port-for-each is invoked. Changes to the port table while port-for-each is running have no effect as far as port-for-each is concerned. The C function scm_port_for_each takes a Scheme procedure encoded as a SCM value, while scm_c_port_for_each takes a pointer to a C function and passes along a arbitrary data cookie.

[Scheme Procedure] setvbuf port mode [size] [C Function] scm_setvbuf (port, mode, size)

Set the buffering mode for port. mode can be:

[Variable] _IONBF

non-buffered

[Variable] _IOLBF

line buffered

[Variable] _IOFBF

block buffered, using a newly allocated buffer of size bytes. If size is omitted, a default size will be used.

Only certain types of ports are supported, most importantly file ports.

[Scheme Procedure] fcntl port/fd cmd [value] [C Function] scm_fcntl (object, cmd, value)

Apply cmd on port/fd, either a port or file descriptor. The value argument is used by the SET commands described below, it’s an integer value. Values for cmd are:

[Variable] F_DUPFD

Duplicate the file descriptor, the same as dup->fdes above does.

[Variable] F_GETFD

[Variable] F_SETFD

Get or set flags associated with the file descriptor. The only flag is the following,

[Variable] FD_CLOEXEC

“Close on exec”, meaning the file descriptor will be closed on an exec call (a successful such call). For example to set that flag,

          (fcntl port F_SETFD FD_CLOEXEC)

Or better, set it but leave any other possible future flags unchanged,

          (fcntl port F_SETFD (logior FD_CLOEXEC
                                      (fcntl port F_GETFD)))

[Variable] F_GETFL

[Variable] F_SETFL

Get or set flags associated with the open file. These flags are O_RDONLY etc described under open above. A common use is to set O_NONBLOCK on a network socket. The following sets that flag, and leaves other flags unchanged.

     (fcntl sock F_SETFL (logior O_NONBLOCK
                                 (fcntl sock F_GETFL)))

[Variable] F_GETOWN

[Variable] F_SETOWN

Get or set the process ID of a socket’s owner, for SIGIO signals.

[Scheme Procedure] flock file operation [C Function] scm_flock (file, operation)

Apply or remove an advisory lock on an open file. operation specifies the action to be done:

[Variable] LOCK_SH

Shared lock. More than one process may hold a shared lock for a given file at a given time.

[Variable] LOCK_EX

Exclusive lock. Only one process may hold an exclusive lock for a given file at a given time.

[Variable] LOCK_UN

Unlock the file.

[Variable] LOCK_NB

Don’t block when locking. This is combined with one of the other operations using logior (see Section 6.6.2.13 [Bitwise Operations], page 125). If flock would block an EWOULDBLOCK error is thrown (see Section 7.2.1 [Conventions], page 468).

The return value is not specified. file may be an open file descriptor or an open file descriptor port. Note that flock does not lock files across NFS.

[Scheme Procedure] select reads writes excepts [secs [usecs]] [C Function] scm_select (reads, writes, excepts, secs, usecs)

This procedure has a variety of uses: waiting for the ability to provide input, accept output, or the existence of exceptional conditions on a collection of ports or file descriptors, or waiting for a timeout to occur. It also returns if interrupted by a signal. reads, writes and excepts can be lists or vectors, with each member a port or a file descriptor. The value returned is a list of three corresponding lists or vectors containing only the members which meet the specified requirement. The ability of port buffers to provide input or accept output is taken into account. Ordering of the input lists or vectors is not preserved. The optional arguments secs and usecs specify the timeout. Either secs can be specified alone, as either an integer or a real number, or both secs and usecs can be specified as integers, in which case usecs is an additional timeout expressed in microseconds. If secs is omitted or is #f then select will wait for as long as it takes for one of the other conditions to be satisfied. The scsh version of select differs as follows: Only vectors are accepted for the first three arguments. The usecs argument is not supported. Multiple values are returned instead of a list. Duplicates in the input vectors appear only once in output. An additional select! interface is provided.

7.2.3 File System

These procedures allow querying and setting file system attributes (such as owner, permissions, sizes and types of files); deleting, copying, renaming and linking files; creating and removing directories and querying their contents; syncing the file system and creating special files.

[Scheme Procedure] access? path how [C Function] scm_access (path, how)

Test accessibility of a file under the real UID and GID of the calling process. The return is #t if path exists and the permissions requested by how are all allowed, or #f if not. how is an integer which is one of the following values, or a bitwise-OR (logior) of multiple values.

[Variable] R_OK

Test for read permission.

[Variable] W_OK

Test for write permission.

[Variable] X_OK

Test for execute permission.

[Variable] F_OK

Test for existence of the file. This is implied by each of the other tests, so there’s no need to combine it with them.

It’s important to note that access? does not simply indicate what will happen on attempting to read or write a file. In normal circumstances it does, but in a set-UID or set-GID program it doesn’t because access? tests the real ID, whereas an open or execute attempt uses the effective ID. A program which will never run set-UID/GID can ignore the difference between real and effective IDs, but for maximum generality, especially in library functions, it’s best not to use access? to predict the result of an open or execute, instead simply attempt that and catch any exception. The main use for access? is to let a set-UID/GID program determine what the invoking user would have been allowed to do, without the greater (or perhaps lesser) privileges afforded by the effective ID. For more on this, see Section “Testing File Access” in The GNU C Library Reference Manual.

[Scheme Procedure] stat object

[C Function] scm_stat (object)

Return an object containing various information about the file determined by object. object can be a string containing a file name or a port or integer file descriptor which is open on a file (in which case fstat is used as the underlying system call). The object returned by stat can be passed as a single parameter to the following procedures, all of which return integers:

[Scheme Procedure] stat:dev st

The device number containing the file.

[Scheme Procedure] stat:ino st

The file serial number, which distinguishes this file from all other files on the same device.

[Scheme Procedure] stat:mode st

The mode of the file. This is an integer which incorporates file type information and file permission bits. See also stat:type and stat:perms below.

[Scheme Procedure] stat:nlink st

The number of hard links to the file.

[Scheme Procedure] stat:uid st

The user ID of the file’s owner.

[Scheme Procedure] stat:gid st

The group ID of the file.

[Scheme Procedure] stat:rdev st

Device ID; this entry is defined only for character or block special files. On some systems this field is not available at all, in which case stat:rdev returns

     #f.

[Scheme Procedure] stat:size st

The size of a regular file in bytes.

[Scheme Procedure] stat:atime st

The last access time for the file, in seconds.

[Scheme Procedure] stat:mtime st

The last modification time for the file, in seconds.

[Scheme Procedure] stat:ctime st

The last modification time for the attributes of the file, in seconds.

[Scheme Procedure] stat:atimensec st [Scheme Procedure] stat:mtimensec st [Scheme Procedure] stat:ctimensec st

The fractional part of a file’s access, modification, or attribute modification time, in nanoseconds. Nanosecond timestamps are only available on some operating systems and file systems. If Guile cannot retrieve nanosecond-level timestamps for a file, these fields will be set to 0.

[Scheme Procedure] stat:blksize st

The optimal block size for reading or writing the file, in bytes. On some systems this field is not available, in which case stat:blksize returns a sensible suggested block size.

[Scheme Procedure] stat:blocks st

The amount of disk space that the file occupies measured in units of 512 byte blocks. On some systems this field is not available, in which case stat:blocks returns #f.

In addition, the following procedures return the information from stat:mode in a more convenient form:

[Scheme Procedure] stat:type st

A symbol representing the type of file. Possible values are ‘regular’,

‘directory’, ‘symlink’, ‘block-special’, ‘char-special’, ‘fifo’, ‘socket’,
and ‘unknown’.

[Scheme Procedure] stat:perms st

An integer representing the access permission bits.

[Scheme Procedure] lstat path [C Function] scm_lstat (path)

Similar to stat, but does not follow symbolic links, i.e., it will return information about a symbolic link itself, not the file it points to. path must be a string.

[Scheme Procedure] readlink path [C Function] scm_readlink (path)

Return the value of the symbolic link named by path (a string), i.e., the file that the link points to.

[Scheme Procedure] chown object owner group [C Function] scm_chown (object, owner, group)

Change the ownership and group of the file referred to by object to the integer values owner and group. object can be a string containing a file name or, if the platform supports fchown (see Section “File Owner” in The GNU C Library Reference Manual), a port or integer file descriptor which is open on the file. The return value is unspecified. If object is a symbolic link, either the ownership of the link or the ownership of the referenced file will be changed depending on the operating system (lchown is unsupported at present). If owner or group is specified as -1, then that ID is not changed.

[Scheme Procedure] chmod object mode [C Function] scm_chmod (object, mode)

Changes the permissions of the file referred to by object. object can be a string containing a file name or a port or integer file descriptor which is open on a file (in which case fchmod is used as the underlying system call). mode specifies the new permissions as a decimal number, e.g., (chmod "foo" #o755). The return value is unspecified.

[Scheme Procedure] utime pathname [actime [modtime [actimens [modtimens

[flags]]]]]

[C Function] scm_utime (pathname, actime, modtime, actimens, modtimens, flags)

utime sets the access and modification times for the file named by pathname. If actime or modtime is not supplied, then the current time is used. actime and modtime must be integer time values as returned by the current-time procedure. The optional actimens and modtimens are nanoseconds to add actime and modtime. Nanosecond precision is only supported on some combinations of file systems and operating systems.

     (utime "foo" (- (current-time) 3600))

will set the access time to one hour in the past and the modification time to the current time.

[Scheme Procedure] delete-file str [C Function] scm_delete_file (str)

Deletes (or “unlinks”) the file whose path is specified by str.

[Scheme Procedure] copy-file oldfile newfile [C Function] scm_copy_file (oldfile, newfile)

Copy the file specified by oldfile to newfile. The return value is unspecified.

[Scheme Procedure] sendfile out in count [offset] [C Function] scm_sendfile (out, in, count, offset)

Send count bytes from in to out, both of which must be either open file ports or file descriptors. When offset is omitted, start reading from in’s current position; otherwise, start reading at offset. Return the number of bytes actually sent. When in is a port, it is often preferable to specify offset, because in’s offset as a port may be different from the offset of its underlying file descriptor.

On systems that support it, such as GNU/Linux, this procedure uses the sendfile libc function, which usually corresponds to a system call. This is faster than doing a series of read and write system calls. A typical application is to send a file over a socket. In some cases, the sendfile libc function may return EINVAL or ENOSYS. In that case, Guile’s sendfile procedure automatically falls back to doing a series of read and write calls. In other cases, the libc function may send fewer bytes than count—for instance because out is a slow or limited device, such as a pipe. When that happens, Guile’s sendfile automatically retries until exactly count bytes were sent or an error occurs.

[Scheme Procedure] rename-file oldname newname [C Function] scm_rename (oldname, newname)

Renames the file specified by oldname to newname. The return value is unspecified.

[Scheme Procedure] link oldpath newpath [C Function] scm_link (oldpath, newpath)

Creates a new name newpath in the file system for the file named by oldpath. If oldpath is a symbolic link, the link may or may not be followed depending on the system.

[Scheme Procedure] symlink oldpath newpath [C Function] scm_symlink (oldpath, newpath)

Create a symbolic link named newpath with the value (i.e., pointing to) oldpath. The return value is unspecified.

[Scheme Procedure] mkdir path [mode] [C Function] scm_mkdir (path, mode)

Create a new directory named by path. If mode is omitted then the permissions of the directory file are set using the current umask (see Section 7.2.7 [Processes], page 490). Otherwise they are set to the decimal value specified with mode. The return value is unspecified.

[Scheme Procedure] rmdir path [C Function] scm_rmdir (path)

Remove the existing directory named by path. The directory must be empty for this to succeed. The return value is unspecified.

[Scheme Procedure] opendir dirname [C Function] scm_opendir (dirname)

Open the directory specified by dirname and return a directory stream. Before using this and the procedures below, make sure to see the higher-level procedures for directory traversal that are available (see Section 7.11 [File Tree Walk], page 676).

[Scheme Procedure] directory-stream? object

[C Function] scm_directory_stream_p (object)

Return a boolean indicating whether object is a directory stream as returned by

opendir.

[Scheme Procedure] readdir stream [C Function] scm_readdir (stream)

Return (as a string) the next directory entry from the directory stream stream. If there is no remaining entry to be read then the end of file object is returned.

[Scheme Procedure] rewinddir stream [C Function] scm_rewinddir (stream)

Reset the directory port stream so that the next call to readdir will return the first directory entry.

[Scheme Procedure] closedir stream [C Function] scm_closedir (stream)

Close the directory stream stream. The return value is unspecified.

Here is an example showing how to display all the entries in a directory:

(define dir (opendir "/usr/lib"))
(do ((entry (readdir dir) (readdir dir)))
    ((eof-object? entry))
  (display entry)(newline))
(closedir dir)

[Scheme Procedure] sync [C Function] scm_sync ()

Flush the operating system disk buffers. The return value is unspecified.

[Scheme Procedure] mknod path type perms dev [C Function] scm_mknod (path, type, perms, dev)

Creates a new special file, such as a file corresponding to a device. path specifies the name of the file. type should be one of the following symbols: ‘regular’, ‘directory’,

‘symlink’, ‘block-special’, ‘char-special’, ‘fifo’, or ‘socket’. perms (an integer)

specifies the file permissions. dev (an integer) specifies which device the special file refers to. Its exact interpretation depends on the kind of special file being created.

E.g.,

     (mknod "/dev/fd0" 'block-special #o660 (+ (* 2 256) 2))

The return value is unspecified.

[Scheme Procedure] tmpnam [C Function] scm_tmpnam ()

Return an auto-generated name of a temporary file, a file which doesn’t already exist. The name includes a path, it’s usually in /tmp but that’s system dependent.

Care must be taken when using tmpnam. In between choosing the name and creating the file another program might use that name, or an attacker might even make it a symlink pointing at something important and causing you to overwrite that.

The safe way is to create the file using open with O_EXCL to avoid any overwriting. A loop can try again with another name if the file exists (error EEXIST). mkstemp! below does that.

[Scheme Procedure] mkstemp! tmpl [C Function] scm_mkstemp (tmpl)

Create a new unique file in the file system and return a new buffered port open for reading and writing to the file. tmpl is a string specifying where the file should be created: it must end with ‘XXXXXX’ and those ‘X’s will be changed in the string to return the name of the file. (portfilename on the port also gives the name.) POSIX doesn’t specify the permissions mode of the file, on GNU and most systems it’s #o600. An application can use chmod to relax that if desired. For example #o666 less umask, which is usual for ordinary file creation,

     (let ((port (mkstemp! (string-copy "/tmp/myfile-XXXXXX"))))
       (chmod port (logand #o666 (lognot (umask))))
       ...)

[Scheme Procedure] tmpfile [C Function] scm_tmpfile ()

Return an input/output port to a unique temporary file named using the path prefix P_tmpdir defined in stdio.h. The file is automatically deleted when the port is closed or the program terminates.

[Scheme Procedure] dirname filename [C Function] scm_dirname (filename)

Return the directory name component of the file name filename. If filename does not contain a directory component, . is returned.

[Scheme Procedure] basename filename [suffix] [C Function] scm_basename (filename, suffix)

Return the base name of the file name filename. The base name is the file name without any directory components. If suffix is provided, and is equal to the end of basename, it is removed also.

     (basename "/tmp/test.xml" ".xml")
     ⇒ "test"

[Scheme Procedure] file-exists? filename

Return #t if the file named filename exists, #f if not.

Many operating systems, such as GNU, use / (forward slash) to separate the components of a file name; any file name starting with / is considered an absolute file name. These conventions are specified by the POSIX Base Definitions, which refer to conforming file names as “pathnames”. Some operating systems use a different convention; in particular, Windows uses \ (backslash) as the file name separator, and also has the notion of volume names like C:\ for absolute file names. The following procedures and variables provide support for portable file name manipulations.

[Scheme Procedure] system-file-name-convention

Return either posix or windows, depending on what kind of system this Guile is running on.

[Scheme Procedure] file-name-separator? c

Return true if character c is a file name separator on the host platform.

[Scheme Procedure] absolute-file-name? file-name

Return true if file-name denotes an absolute file name on the host platform.

[Scheme Variable] file-name-separator-string

The preferred file name separator. Note that on MinGW builds for Windows, both / and \ are valid separators. Thus, programs should not assume that file-name-separator-string is the only file name separator—e.g., when extracting the components of a file name.

7.2.4 User Information

The facilities in this section provide an interface to the user and group database. They should be used with care since they are not reentrant.

The following functions accept an object representing user information and return a selected component:

[Scheme Procedure] passwd:name pw

The name of the userid.

[Scheme Procedure] passwd:passwd pw

The encrypted passwd.

[Scheme Procedure] passwd:uid pw

The user id number.

[Scheme Procedure] passwd:gid pw

The group id number.

[Scheme Procedure] passwd:gecos pw

The full name.

[Scheme Procedure] passwd:dir pw

The home directory.

[Scheme Procedure] passwd:shell pw

The login shell.

[Scheme Procedure] getpwuid uid

Look up an integer userid in the user database.

[Scheme Procedure] getpwnam name

Look up a user name string in the user database.

[Scheme Procedure] setpwent

Initializes a stream used by getpwent to read from the user database. The next use of getpwent will return the first entry. The return value is unspecified.

[Scheme Procedure] getpwent

Read the next entry in the user database stream. The return is a passwd user object as above, or #f when no more entries.

[Scheme Procedure] endpwent

Closes the stream used by getpwent. The return value is unspecified.

[Scheme Procedure] setpw [arg] [C Function] scm_setpwent (arg)

If called with a true argument, initialize or reset the password data stream. Otherwise, close the stream. The setpwent and endpwent procedures are implemented on top of this.

[Scheme Procedure] getpw [user] [C Function] scm_getpwuid (user)

Look up an entry in the user database. user can be an integer, a string, or omitted, giving the behaviour of getpwuid, getpwnam or getpwent respectively.

The following functions accept an object representing group information and return a selected component:

[Scheme Procedure] group:name gr

The group name.

[Scheme Procedure] group:passwd gr

The encrypted group password.

[Scheme Procedure] group:gid gr

The group id number.

[Scheme Procedure] group:mem gr

A list of userids which have this group as a supplementary group.

[Scheme Procedure] getgrgid gid

Look up an integer group id in the group database.

[Scheme Procedure] getgrnam name

Look up a group name in the group database.

[Scheme Procedure] setgrent

Initializes a stream used by getgrent to read from the group database. The next use of getgrent will return the first entry. The return value is unspecified.

[Scheme Procedure] getgrent

Return the next entry in the group database, using the stream set by setgrent.

[Scheme Procedure] endgrent

Closes the stream used by getgrent. The return value is unspecified.

[Scheme Procedure] setgr [arg] [C Function] scm_setgrent (arg)

If called with a true argument, initialize or reset the group data stream. Otherwise, close the stream. The setgrent and endgrent procedures are implemented on top of this.

[Scheme Procedure] getgr [group] [C Function] scm_getgrgid (group)

Look up an entry in the group database. group can be an integer, a string, or omitted, giving the behaviour of getgrgid, getgrnam or getgrent respectively.

In addition to the accessor procedures for the user database, the following shortcut procedure is also available.

[Scheme Procedure] getlogin [C Function] scm_getlogin ()

Return a string containing the name of the user logged in on the controlling terminal of the process, or #f if this information cannot be obtained.

7.2.5 Time

[Scheme Procedure] current-time

[C Function] scm_current_time ()

Return the number of seconds since 1970-01-01 00:00:00 UTC, excluding leap seconds.

[Scheme Procedure] gettimeofday

[C Function] scm_gettimeofday ()

Return a pair containing the number of seconds and microseconds since 1970-01-01 00:00:00 UTC, excluding leap seconds. Note: whether true microsecond resolution is available depends on the operating system.

The following procedures either accept an object representing a broken down time and return a selected component, or accept an object representing a broken down time and a value and set the component to the value. The numbers in parentheses give the usual range.

[Scheme Procedure] tm:sec tm [Scheme Procedure] set-tm:sec tm val

Seconds (0-59).

[Scheme Procedure] tm:min tm [Scheme Procedure] set-tm:min tm val

Minutes (0-59).

[Scheme Procedure] tm:hour tm [Scheme Procedure] set-tm:hour tm val

Hours (0-23).

[Scheme Procedure] tm:mday tm [Scheme Procedure] set-tm:mday tm val

Day of the month (1-31).

[Scheme Procedure] tm:mon tm [Scheme Procedure] set-tm:mon tm val

Month (0-11).

[Scheme Procedure] tm:year tm [Scheme Procedure] set-tm:year tm val

Year (70-), the year minus 1900.

[Scheme Procedure] tm:wday tm [Scheme Procedure] set-tm:wday tm val

Day of the week (0-6) with Sunday represented as 0.

[Scheme Procedure] tm:yday tm [Scheme Procedure] set-tm:yday tm val

Day of the year (0-364, 365 in leap years).

[Scheme Procedure] tm:isdst tm [Scheme Procedure] set-tm:isdst tm val

Daylight saving indicator (0 for “no”, greater than 0 for “yes”, less than 0 for “unknown”).

[Scheme Procedure] tm:gmtoff tm [Scheme Procedure] set-tm:gmtoff tm val

Time zone offset in seconds west of UTC (-46800 to 43200). For example on East coast USA (zone ‘EST+5’) this would be 18000 (ie. 5 × 60 × 60) in winter, or 14400 (ie. 4 × 60 × 60) during daylight savings. Note tm:gmtoff is not the same as tm_gmtoff in the C tm structure. tm_gmtoff is seconds east and hence the negative of the value here.

[Scheme Procedure] tm:zone tm [Scheme Procedure] set-tm:zone tm val

Time zone label (a string), not necessarily unique.

[Scheme Procedure] localtime time [zone] [C Function] scm_localtime (time, zone)

Return an object representing the broken down components of time, an integer like the one returned by current-time. The time zone for the calculation is optionally specified by zone (a string), otherwise the TZ environment variable or the system default is used.

[Scheme Procedure] gmtime time [C Function] scm_gmtime (time)

Return an object representing the broken down components of time, an integer like the one returned by current-time. The values are calculated for UTC.

[Scheme Procedure] mktime sbd-time [zone] [C Function] scm_mktime (sbd time, zone)

For a broken down time object sbd-time, return a pair the car of which is an integer time like current-time, and the cdr of which is a new broken down time with normalized fields. zone is a timezone string, or the default is the TZ environment variable or the system default (see Section “Specifying the Time Zone with TZ” in GNU C Library Reference Manual). sbd-time is taken to be in that zone.

The following fields of sbd-time are used: tm:year, tm:mon, tm:mday, tm:hour, tm:min, tm:sec, tm:isdst. The values can be outside their usual ranges. For example tm:hour normally goes up to 23, but a value say 33 would mean 9 the following day.

tm:isdst in sbd-time says whether the time given is with daylight savings or not. This is ignored if zone doesn’t have any daylight savings adjustment amount.

The broken down time in the return normalizes the values of sbd-time by bringing them into their usual ranges, and using the actual daylight savings rule for that time in zone (which may differ from what sbd-time had). The easiest way to think of this is that sbd-time plus zone converts to the integer UTC time, then a localtime is applied to get the normal presentation of that time, in zone.

[Scheme Procedure] tzset [C Function] scm_tzset ()

Initialize the timezone from the TZ environment variable or the system default. It’s not usually necessary to call this procedure since it’s done automatically by other procedures that depend on the timezone.

[Scheme Procedure] strftime format tm [C Function] scm_strftime (format, tm)

Return a string which is broken-down time structure tm formatted according to the given format string.

format contains field specifications introduced by a ‘%’ character. See Section “Formatting Calendar Time” in The GNU C Library Reference Manual, or ‘man 3 strftime’, for the available formatting.

     (strftime "%c" (localtime (current-time)))
     ⇒ "Mon Mar 11 20:17:43 2002"

If setlocale has been called (see Section 7.2.13 [Locales], page 518), month and day names are from the current locale and in the locale character set.

[Scheme Procedure] strptime format string [C Function] scm_strptime (format, string)

Performs the reverse action to strftime, parsing string according to the specification supplied in format. The interpretation of month and day names is dependent on the current locale. The value returned is a pair. The CAR has an object with time components in the form returned by localtime or gmtime, but the time zone components are not usefully set. The CDR reports the number of characters from string which were used for the conversion.

[Variable] internal-time-units-per-second

The value of this variable is the number of time units per second reported by the following procedures.

[Scheme Procedure] times [C Function] scm_times ()

Return an object with information about real and processor time. The following procedures accept such an object as an argument and return a selected component:

[Scheme Procedure] tms:clock tms

The current real time, expressed as time units relative to an arbitrary base.

[Scheme Procedure] tms:utime tms

The CPU time units used by the calling process.

[Scheme Procedure] tms:stime tms

The CPU time units used by the system on behalf of the calling process.

[Scheme Procedure] tms:cutime tms

The CPU time units used by terminated child processes of the calling process, whose status has been collected (e.g., using waitpid).

[Scheme Procedure] tms:cstime tms

Similarly, the CPU times units used by the system on behalf of terminated child processes.

[Scheme Procedure] get-internal-real-time
[C Function] scm_get_internal_real_time ()

Return the number of time units since the interpreter was started.

[Scheme Procedure] get-internal-run-time
[C Function] scm_get_internal_run_time ()

Return the number of time units of processor time used by the interpreter. Both system and user time are included but subprocesses are not.

7.2.6 Runtime Environment

[Scheme Procedure] program-arguments

[Scheme Procedure] command-line

[Scheme Procedure] set-program-arguments
[C Function] scm_program_arguments ()
[C Function] scm_set_program_arguments_scm (lst)

Get the command line arguments passed to Guile, or set new arguments. The arguments are a list of strings, the first of which is the invoked program name. This is just "guile" (or the executable path) when run interactively, or it’s the script name when running a script with -s (see Section 4.2 [Invoking Guile], page 35).

     guile -L /my/extra/dir -s foo.scm abc def
     (program-arguments) ⇒ ("foo.scm" "abc" "def")

set-program-arguments allows a library module or similar to modify the arguments, for example to strip options it recognises, leaving the rest for the mainline. The argument list is held in a fluid, which means it’s separate for each thread. Neither the list nor the strings within it are copied at any point and normally should not be mutated.

The two names program-arguments and command-line are an historical accident, they both do exactly the same thing. The name scm_set_program_arguments_scm has an extra _scm on the end to avoid clashing with the C function below.

[C Function] void scm_set_program_arguments (int argc, char **argv, char *first)

Set the list of command line arguments for program-arguments and command-line above. argv is an array of null-terminated strings, as in a C main function. argc is the number of strings in argv, or if it’s negative then a NULL in argv marks its end. first is an extra string put at the start of the arguments, or NULL for no such extra. This is a convenient way to pass the program name after advancing argv to strip option arguments. Eg.

     {
       char *progname = argv[0];
       for (argv++; argv[0] != NULL && argv[0][0] == '-'; argv++)
         {
           /* munch option ... */
         }
       /* remaining args for scheme level use */
       scm_set_program_arguments (-1, argv, progname);
     }

This sort of thing is often done at startup under scm_boot_guile with options handled at the C level removed. The given strings are all copied, so the C data is not

accessed again once scm_set_program_arguments returns.

[Scheme Procedure] getenv name [C Function] scm_getenv (name)

Looks up the string name in the current environment. The return value is #f unless a string of the form NAME=VALUE is found, in which case the string VALUE is returned.

[Scheme Procedure] setenv name value Modifies the environment of the current process, which is also the default environment inherited by child processes. If value is #f, then name is removed from the environment. Otherwise, the string name=value is added to the environment, replacing any existing string with name matching name. The return value is unspecified.

[Scheme Procedure] unsetenv name

Remove variable name from the environment. The name can not contain a ‘=’ character.

[Scheme Procedure] environ [env] [C Function] scm_environ (env)

If env is omitted, return the current environment (in the Unix sense) as a list of strings. Otherwise set the current environment, which is also the default environment for child processes, to the supplied list of strings. Each member of env should be of the form name=value and values of name should not be duplicated. If env is supplied then the return value is unspecified.

[Scheme Procedure] putenv str [C Function] scm_putenv (str)

Modifies the environment of the current process, which is also the default environment inherited by child processes. If str is of the form NAME=VALUE then it will be written directly into the environment, replacing any existing environment string with name matching NAME. If str does not contain an equal sign, then any existing string with name matching str will be removed. The return value is unspecified.

7.2.7 Processes

[Scheme Procedure] chdir str [C Function] scm_chdir (str)

Change the current working directory to str. The return value is unspecified.

[Scheme Procedure] getcwd [C Function] scm_getcwd ()

Return the name of the current working directory.

[Scheme Procedure] umask [mode] [C Function] scm_umask (mode)

If mode is omitted, returns a decimal number representing the current file creation mask. Otherwise the file creation mask is set to mode and the previous value is returned. See Section “Assigning File Permissions” in The GNU C Library Reference Manual, for more on how to use umasks. E.g., (umask #o022) sets the mask to octal 22/decimal 18.

[Scheme Procedure] chroot path [C Function] scm_chroot (path)

Change the root directory to that specified in path. This directory will be used for path names beginning with /. The root directory is inherited by all children of the current process. Only the superuser may change the root directory.

[Scheme Procedure] getpid [C Function] scm_getpid ()

Return an integer representing the current process ID.

[Scheme Procedure] getgroups

[C Function] scm_getgroups ()

Return a vector of integers representing the current supplementary group IDs.

[Scheme Procedure] getppid [C Function] scm_getppid ()

Return an integer representing the process ID of the parent process.

[Scheme Procedure] getuid [C Function] scm_getuid ()

Return an integer representing the current real user ID.

[Scheme Procedure] getgid [C Function] scm_getgid ()

Return an integer representing the current real group ID.

[Scheme Procedure] geteuid [C Function] scm_geteuid ()

Return an integer representing the current effective user ID. If the system does not support effective IDs, then the real ID is returned. (provided? 'EIDs) reports whether the system supports effective IDs.

[Scheme Procedure] getegid [C Function] scm_getegid ()

Return an integer representing the current effective group ID. If the system does not support effective IDs, then the real ID is returned. (provided? 'EIDs) reports whether the system supports effective IDs.

[Scheme Procedure] setgroups vec [C Function] scm_setgroups (vec)

Set the current set of supplementary group IDs to the integers in the given vector vec. The return value is unspecified. Generally only the superuser can set the process group IDs (see Section “Setting Groups” in The GNU C Library Reference Manual).

[Scheme Procedure] setuid id [C Function] scm_setuid (id)

Sets both the real and effective user IDs to the integer id, provided the process has appropriate privileges. The return value is unspecified.

[Scheme Procedure] setgid id [C Function] scm_setgid (id)

Sets both the real and effective group IDs to the integer id, provided the process has appropriate privileges. The return value is unspecified.

[Scheme Procedure] seteuid id [C Function] scm_seteuid (id)

Sets the effective user ID to the integer id, provided the process has appropriate privileges. If effective IDs are not supported, the real ID is set instead—(provided? 'EIDs) reports whether the system supports effective IDs. The return value is unspecified.

[Scheme Procedure] setegid id [C Function] scm_setegid (id)

Sets the effective group ID to the integer id, provided the process has appropriate privileges. If effective IDs are not supported, the real ID is set instead—(provided? 'EIDs) reports whether the system supports effective IDs. The return value is unspecified.

[Scheme Procedure] getpgrp [C Function] scm_getpgrp ()

Return an integer representing the current process group ID. This is the POSIX definition, not BSD.

[Scheme Procedure] setpgid pid pgid [C Function] scm_setpgid (pid, pgid)

Move the process pid into the process group pgid. pid or pgid must be integers: they can be zero to indicate the ID of the current process. Fails on systems that do not support job control. The return value is unspecified.

[Scheme Procedure] setsid [C Function] scm_setsid ()

Creates a new session. The current process becomes the session leader and is put in a new process group. The process will be detached from its controlling terminal if it has one. The return value is an integer representing the new process group ID.

[Scheme Procedure] getsid pid [C Function] scm_getsid (pid)

Returns the session ID of process pid. (The session ID of a process is the process group ID of its session leader.)

[Scheme Procedure] waitpid pid [options] [C Function] scm_waitpid (pid, options)

This procedure collects status information from a child process which has terminated or (optionally) stopped. Normally it will suspend the calling process until this can be done. If more than one child process is eligible then one will be chosen by the operating system. The value of pid determines the behaviour:

pid greater than 0 Request status information from the specified child process.

pid equal to -1 or WAIT_ANY Request status information for any child process.

pid equal to 0 or WAIT_MYPGRP Request status information for any child process in the current process group.

pid less than -1 Request status information for any child process whose process group ID is −pid.

The options argument, if supplied, should be the bitwise OR of the values of zero or more of the following variables:

[Variable] WNOHANG

Return immediately even if there are no child processes to be collected.

[Variable] WUNTRACED

Report status information for stopped processes as well as terminated processes.

The return value is a pair containing:

1. The process ID of the child process, or 0 if WNOHANG was specified and no process was collected.

2. The integer status value.

The following three functions can be used to decode the process status code returned by

waitpid.

[Scheme Procedure] status:exit-val status

[C Function] scm_status_exit_val (status)

Return the exit status value, as would be set if a process ended normally through a call to exit or _exit, if any, otherwise #f.

[Scheme Procedure] status:term-sig status

[C Function] scm_status_term_sig (status)

Return the signal number which terminated the process, if any, otherwise #f.

[Scheme Procedure] status:stop-sig status

[C Function] scm_status_stop_sig (status)

Return the signal number which stopped the process, if any, otherwise #f.

[Scheme Procedure] system [cmd] [C Function] scm_system (cmd)

Execute cmd using the operating system’s “command processor”. Under Unix this is usually the default shell sh. The value returned is cmd’s exit status as returned by waitpid, which can be interpreted using the functions above. If system is called without arguments, return a boolean indicating whether the command processor is available.

[Scheme Procedure] system* arg1 arg2 …

[C Function] scm_system_star (args)

Execute the command indicated by arg1 arg2 . . .. The first element must be a string indicating the command to be executed, and the remaining items must be strings representing each of the arguments to that command. This function returns the exit status of the command as provided by waitpid. This value can be handled with status:exit-val and the related functions. system* is similar to system, but accepts only one string per-argument, and performs no shell interpretation. The command is executed using fork and execlp. Accordingly this function may be safer than system in situations where shell interpretation is not required. Example: (system* "echo" "foo" "bar")

[Scheme Procedure] quit [status] [Scheme Procedure] exit [status]

Terminate the current process with proper unwinding of the Scheme stack. The exit status zero if status is not supplied. If status is supplied, and it is an integer, that integer is used as the exit status. If status is #t or #f, the exit status is 0 or 1, respectively. The procedure exit is an alias of quit. They have the same functionality.

[Scheme Procedure] primitive-exit [status] [Scheme Procedure] primitive-_exit [status] [C Function] scm_primitive_exit (status)

[C Function] scm_primitive__exit (status)

Terminate the current process without unwinding the Scheme stack. The exit status is status if supplied, otherwise zero. primitive-exit uses the C exit function and hence runs usual C level cleanups (flush output streams, call atexit functions, etc, see Section “Normal Termination” in The GNU C Library Reference Manual)). primitive-_exit is the _exit system call (see Section “Termination Internals” in The GNU C Library Reference Manual). This terminates the program immediately, with neither Scheme-level nor C-level cleanups. The typical use for primitive-_exit is from a child process created with primitivefork. For example in a Gdk program the child process inherits the X server connection and a C-level atexit cleanup which will close that connection. But closing in the child would upset the protocol in the parent, so primitive-_exit should be used to exit without that.

[Scheme Procedure] execl filename arg … [C Function] scm_execl (filename, args)

Executes the file named by filename as a new process image. The remaining arguments are supplied to the process; from a C program they are accessible as the argv argument to main. Conventionally the first arg is the same as filename. All arguments must be strings. If arg is missing, filename is executed with a null argument list, which may have system-dependent side-effects. This procedure is currently implemented using the execv system call, but we call it execl because of its Scheme calling interface.

[Scheme Procedure] execlp filename arg … [C Function] scm_execlp (filename, args)

Similar to execl, however if filename does not contain a slash then the file to execute will be located by searching the directories listed in the PATH environment variable. This procedure is currently implemented using the execvp system call, but we call it execlp because of its Scheme calling interface.

[Scheme Procedure] execle filename env arg … [C Function] scm_execle (filename, env, args)

Similar to execl, but the environment of the new process is specified by env, which must be a list of strings as returned by the environ procedure. This procedure is currently implemented using the execve system call, but we call it execle because of its Scheme calling interface.

[Scheme Procedure] primitive-fork [C Function] scm_fork ()

Creates a new “child” process by duplicating the current “parent” process. In the child the return value is 0. In the parent the return value is the integer process ID of the child.

Note that it is unsafe to fork a process that has multiple threads running, as only the thread that calls primitive-fork will persist in the child. Any resources that other threads held, such as locked mutexes or open file descriptors, are lost. Indeed, POSIX specifies that only async-signal-safe procedures are safe to call after a multithreaded fork, which is a very limited set. Guile issues a warning if it detects a fork from a multi-threaded program.

If you are going to exec soon after forking, the procedures in (ice-9 popen) may be useful to you, as they fork and exec within an async-signal-safe function carefully written to ensure robust program behavior, even in the presence of threads. See Section 7.2.10 [Pipes], page 500, for more.

This procedure has been renamed from fork to avoid a naming conflict with the scsh fork.

[Scheme Procedure] nice incr [C Function] scm_nice (incr)

Increment the priority of the current process by incr. A higher priority value means that the process runs less often. The return value is unspecified.

[Scheme Procedure] setpriority which who prio [C Function] scm_setpriority (which, who, prio)

Set the scheduling priority of the process, process group or user, as indicated by which and who. which is one of the variables PRIO_PROCESS, PRIO_PGRP or PRIO_USER, and who is interpreted relative to which (a process identifier for PRIO_PROCESS, process group identifier for PRIO_PGRP, and a user identifier for PRIO_USER. A zero value of who denotes the current process, process group, or user. prio is a value in the range [−20,20]. The default priority is 0; lower priorities (in numerical terms) cause more favorable scheduling. Sets the priority of all of the specified processes. Only the super-user may lower priorities. The return value is not specified.

[Scheme Procedure] getpriority which who [C Function] scm_getpriority (which, who)

Return the scheduling priority of the process, process group or user, as indicated by which and who. which is one of the variables PRIO_PROCESS, PRIO_PGRP or PRIO_ USER, and who should be interpreted depending on which (a process identifier for PRIO_PROCESS, process group identifier for PRIO_PGRP, and a user identifier for PRIO_ USER). A zero value of who denotes the current process, process group, or user. Return the highest priority (lowest numerical value) of any of the specified processes.

[Scheme Procedure] getaffinity pid [C Function] scm_getaffinity (pid)

Return a bitvector representing the CPU affinity mask for process pid. Each CPU the process has affinity with has its corresponding bit set in the returned bitvector. The number of bits set is a good estimate of how many CPUs Guile can use without stepping on other processes’ toes.

Currently this procedure is only defined on GNU variants (see Section “CPU Affinity” in The GNU C Library Reference Manual).

[Scheme Procedure] setaffinity pid mask [C Function] scm_setaffinity (pid, mask)

Install the CPU affinity mask mask, a bitvector, for the process or thread with ID pid. The return value is unspecified. Currently this procedure is only defined on GNU variants (see Section “CPU Affinity” in The GNU C Library Reference Manual).

[Scheme Procedure] total-processor-count
[C Function] scm_total_processor_count ()

Return the total number of processors of the machine, which is guaranteed to be at least 1. A “processor” here is a thread execution unit, which can be either:

• an execution core in a (possibly multi-core) chip, in a (possibly multi- chip) module, in a single computer, or

• a thread execution unit inside a core in the case of hyper-threaded CPUs.

Which of the two definitions is used, is unspecified.

[Scheme Procedure] current-processor-count
[C Function] scm_current_processor_count ()

Like total-processor-count, but return the number of processors available to the current process. See setaffinity and getaffinity for more information.

7.2.8 Signals

The following procedures raise, handle and wait for signals.

Scheme code signal handlers are run via a system async (see Section 6.21.2.1 [System asyncs], page 411), so they’re called in the handler’s thread at the next safe opportunity. Generally this is after any currently executing primitive procedure finishes (which could be a long time for primitives that wait for an external event).

[Scheme Procedure] kill pid sig [C Function] scm_kill (pid, sig)

Sends a signal to the specified process or group of processes. pid specifies the processes to which the signal is sent:

pid greater than 0 The process whose identifier is pid.

pid equal to 0 All processes in the current process group.

pid less than -1 The process group whose identifier is -pid

pid equal to -1 If the process is privileged, all processes except for some special system processes. Otherwise, all processes with the current effective user ID.

sig should be specified using a variable corresponding to the Unix symbolic name, e.g.,

[Variable] SIGHUP

Hang-up signal.

[Variable] SIGINT

Interrupt signal.

A full list of signals on the GNU system may be found in Section “Standard Signals” in The GNU C Library Reference Manual.

[Scheme Procedure] raise sig [C Function] scm_raise (sig)

Sends a specified signal sig to the current process, where sig is as described for the kill procedure.

[Scheme Procedure] sigaction signum [handler [flags [thread]]] [C Function] scm_sigaction (signum, handler, flags) [C Function] scm_sigaction_for_thread (signum, handler, flags, thread)

Install or report the signal handler for a specified signal. signum is the signal number, which can be specified using the value of variables such

as SIGINT.

If handler is omitted, sigaction returns a pair: the CAR is the current signal hander, which will be either an integer with the value SIG_DFL (default action) or SIG_IGN (ignore), or the Scheme procedure which handles the signal, or #f if a non-Scheme procedure handles the signal. The CDR contains the current sigaction flags for the handler. If handler is provided, it is installed as the new handler for signum. handler can be a Scheme procedure taking one argument, or the value of SIG_DFL (default action) or SIG_IGN (ignore), or #f to restore whatever signal handler was installed before sigaction was first used. When a scheme procedure has been specified, that procedure will run in the given thread. When no thread has been given, the thread that made this call to sigaction is used. flags is a logior (see Section 6.6.2.13 [Bitwise Operations], page 125) of the following (where provided by the system), or 0 for none.

[Variable] SA_NOCLDSTOP

By default, SIGCHLD is signalled when a child process stops (ie. receives SIGSTOP), and when a child process terminates. With the SA_NOCLDSTOP flag, SIGCHLD is only signalled for termination, not stopping. SA_NOCLDSTOP has no effect on signals other than SIGCHLD.

[Variable] SA_RESTART

If a signal occurs while in a system call, deliver the signal then restart the system call (as opposed to returning an EINTR error from that call).

The return value is a pair with information about the old handler as described above. This interface does not provide access to the “signal blocking” facility. Maybe this is not needed, since the thread support may provide solutions to the problem of consistent access to data structures.

[Scheme Procedure] restore-signals

[C Function] scm_restore_signals ()

Return all signal handlers to the values they had before any call to sigaction was made. The return value is unspecified.

[Scheme Procedure] alarm i [C Function] scm_alarm (i)

Set a timer to raise a SIGALRM signal after the specified number of seconds (an integer). It’s advisable to install a signal handler for SIGALRM beforehand, since the default action is to terminate the process. The return value indicates the time remaining for the previous alarm, if any. The new value replaces the previous alarm. If there was no previous alarm, the return value is zero.

[Scheme Procedure] pause [C Function] scm_pause ()

Pause the current process (thread?) until a signal arrives whose action is to either terminate the current process or invoke a handler procedure. The return value is unspecified.

[Scheme Procedure] sleep secs [Scheme Procedure] usleep usecs [C Function] scm_sleep (secs) [C Function] scm_usleep (usecs)

Wait the given period secs seconds or usecs microseconds (both integers). If a signal arrives the wait stops and the return value is the time remaining, in seconds or microseconds respectively. If the period elapses with no signal the return is zero. On most systems the process scheduler is not microsecond accurate and the actual period slept by usleep might be rounded to a system clock tick boundary, which might be 10 milliseconds for instance. See scm_std_sleep and scm_std_usleep for equivalents at the C level (see Section 6.21.5 [Blocking], page 418).

[Scheme Procedure] getitimer which timer [Scheme Procedure] setitimer which timer interval seconds interval microseconds

periodic seconds periodic microseconds

[C Function] scm_getitimer (which timer)

[C Function] scm_setitimer (which timer, interval seconds, interval microseconds,

periodic seconds, periodic microseconds) Get or set the periods programmed in certain system timers. These timers have a current interval value which counts down and on reaching zero raises a signal. An optional periodic value can be set to restart from there each time, for periodic operation. which timer is one of the following values

[Variable] ITIMER_REAL

A real-time timer, counting down elapsed real time. At zero it raises SIGALRM. This is like alarm above, but with a higher resolution period.

[Variable] ITIMER_VIRTUAL

A virtual-time timer, counting down while the current process is actually using CPU. At zero it raises SIGVTALRM.

[Variable] ITIMER_PROF

A profiling timer, counting down while the process is running (like ITIMER_ VIRTUAL) and also while system calls are running on the process’s behalf. At zero it raises a SIGPROF.

This timer is intended for profiling where a program is spending its time (by looking where it is when the timer goes off).

getitimer returns the current timer value and its programmed restart value, as a list containing two pairs. Each pair is a time in seconds and microseconds: ((interval_

secs . interval_usecs) (periodic_secs . periodic_usecs)).

setitimer sets the timer values similarly, in seconds and microseconds (which must be integers). The periodic value can be zero to have the timer run down just once. The return value is the timer’s previous setting, in the same form as getitimer returns.

     (setitimer ITIMER_REAL
                5 500000     ;; first SIGALRM in 5.5 seconds time
                2 0)         ;; then repeat every 2 seconds

Although the timers are programmed in microseconds, the actual accuracy might not be that high.

7.2.9 Terminals and Ptys

[Scheme Procedure] isatty? port [C Function] scm_isatty_p (port)

Return #t if port is using a serial non–file device, otherwise #f.

[Scheme Procedure] ttyname port [C Function] scm_ttyname (port)

Return a string with the name of the serial terminal device underlying port.

[Scheme Procedure] ctermid [C Function] scm_ctermid ()

Return a string containing the file name of the controlling terminal for the current process.

[Scheme Procedure] tcgetpgrp port [C Function] scm_tcgetpgrp (port)

Return the process group ID of the foreground process group associated with the terminal open on the file descriptor underlying port.

If there is no foreground process group, the return value is a number greater than 1 that does not match the process group ID of any existing process group. This can happen if all of the processes in the job that was formerly the foreground job have terminated, and no other job has yet been moved into the foreground.

[Scheme Procedure] tcsetpgrp port pgid [C Function] scm_tcsetpgrp (port, pgid)

Set the foreground process group ID for the terminal used by the file descriptor underlying port to the integer pgid. The calling process must be a member of the same session as pgid and must have the same controlling terminal. The return value is unspecified.

7.2.10 Pipes

The following procedures are similar to the popen and pclose system routines. The code is in a separate “popen” module^{1}:

(use-modules (ice-9 popen))

[Scheme Procedure] open-pipe command mode [Scheme Procedure] open-pipe* mode prog [args...]

Execute a command in a subprocess, with a pipe to it or from it, or with pipes in both directions. open-pipe runs the shell command using ‘/bin/sh -c’. open-pipe* executes prog directly, with the optional args arguments (all strings). mode should be one of the following values. OPEN_READ is an input pipe, ie. to read from the subprocess. OPEN_WRITE is an output pipe, ie. to write to it.

[Variable] OPEN_READ

[Variable] OPEN_WRITE

[Variable] OPEN_BOTH

For an input pipe, the child’s standard output is the pipe and standard input is inherited from current-input-port. For an output pipe, the child’s standard input is the pipe and standard output is inherited from current-output-port. In all cases cases the child’s standard error is inherited from current-error-port (see Section 6.14.8 [Default Ports], page 325). If those current-X-ports are not files of some kind, and hence don’t have file descriptors for the child, then /dev/null is used instead. Care should be taken with OPEN_BOTH, a deadlock will occur if both parent and child are writing, and waiting until the write completes before doing any reading. Each direction has PIPE_BUF bytes of buffering (see Section 7.2.2 [Ports and File Descriptors], page 469), which will be enough for small writes, but not for say putting a big file through a filter.

[Scheme Procedure] open-input-pipe command

Equivalent to open-pipe with mode OPEN_READ.

     (let* ((port (open-input-pipe "date --utc"))
            (str  (read-line port)))
       (close-pipe port)
       str)
     ⇒ "Mon Mar 11 20:10:44 UTC 2002"

1 This module is only available on systems where the fork feature is provided (see Section 6.22.2.2 [Com-

mon Feature Symbols], page 429).

[Scheme Procedure] open-output-pipe command

Equivalent to open-pipe with mode OPEN_WRITE.

          (let ((port (open-output-pipe "lpr")))
            (display "Something for the line printer.\n" port)
            (if (not (eqv? 0 (status:exit-val (close-pipe port))))
                (error "Cannot print")))

[Scheme Procedure] open-input-output-pipe command

Equivalent to open-pipe with mode OPEN_BOTH.

[Scheme Procedure] close-pipe port

Close a pipe created by open-pipe, wait for the process to terminate, and return the wait status code. The status is as per waitpid and can be decoded with status:exitval etc (see Section 7.2.7 [Processes], page 490)

waitpid WAIT_ANY should not be used when pipes are open, since it can reap a pipe’s child process, causing an error from a subsequent close-pipe.

close-port (see Section 6.14.4 [Closing], page 321) can close a pipe, but it doesn’t reap the child process.

The garbage collector will close a pipe no longer in use, and reap the child process with waitpid. If the child hasn’t yet terminated the garbage collector doesn’t block, but instead checks again in the next GC.

Many systems have per-user and system-wide limits on the number of processes, and a system-wide limit on the number of pipes, so pipes should be closed explicitly when no longer needed, rather than letting the garbage collector pick them up at some later time.

7.2.11 Networking

7.2.11.1 Network Address Conversion

This section describes procedures which convert internet addresses between numeric and string formats.

IPv4 Address Conversion

An IPv4 Internet address is a 4-byte value, represented in Guile as an integer in host byte order, so that say “0.0.0.1” is 1, or “1.0.0.0” is 16777216.

Some underlying C functions use network byte order for addresses, Guile converts as necessary so that at the Scheme level its host byte order everywhere.

[Variable] INADDR_ANY

For a server, this can be used with bind (see Section 7.2.11.4 [Network Sockets and Communication], page 511) to allow connections from any interface on the machine.

[Variable] INADDR_BROADCAST

The broadcast address on the local network.

[Variable] INADDR_LOOPBACK

The address of the local host using the loopback device, ie. ‘127.0.0.1’.

[Scheme Procedure] inet-aton address [C Function] scm_inet_aton (address)

This function is deprecated in favor of inet-pton. Convert an IPv4 Internet address from printable string (dotted decimal notation) to an integer. E.g.,

     (inet-aton "127.0.0.1") ⇒ 2130706433

[Scheme Procedure] inet-ntoa inetid [C Function] scm_inet_ntoa (inetid)

This function is deprecated in favor of inet-ntop. Convert an IPv4 Internet address to a printable (dotted decimal notation) string. E.g.,

     (inet-ntoa 2130706433) ⇒ "127.0.0.1"

[Scheme Procedure] inet-netof address [C Function] scm_inet_netof (address)

Return the network number part of the given IPv4 Internet address. E.g.,

     (inet-netof 2130706433) ⇒ 127

[Scheme Procedure] inet-lnaof address [C Function] scm_lnaof (address)

Return the local-address-with-network part of the given IPv4 Internet address, using the obsolete class A/B/C system. E.g.,

     (inet-lnaof 2130706433) ⇒ 1

[Scheme Procedure] inet-makeaddr net lna [C Function] scm_inet_makeaddr (net, lna)

Make an IPv4 Internet address by combining the network number net with the localaddress-within-network number lna. E.g.,

     (inet-makeaddr 127 1) ⇒ 2130706433

IPv6 Address Conversion

An IPv6 Internet address is a 16-byte value, represented in Guile as an integer in host byte order, so that say “::1” is 1.

[Scheme Procedure] inet-ntop family address [C Function] scm_inet_ntop (family, address)

Convert a network address from an integer to a printable string. family can be AF_

INET or AF_INET6. E.g.,
     (inet-ntop AF_INET 2130706433) ⇒ "127.0.0.1"
     (inet-ntop AF_INET6 (- (expt 2 128) 1))
       ⇒ "ffff:ffff:ffff:ffff:ffff:ffff:ffff:ffff"

[Scheme Procedure] inet-pton family address [C Function] scm_inet_pton (family, address)

Convert a string containing a printable network address to an integer address. family

can be AF_INET or AF_INET6. E.g.,

     (inet-pton AF_INET "127.0.0.1") ⇒ 2130706433
     (inet-pton AF_INET6 "::1") ⇒ 1

7.2.11.2 Network Databases

This section describes procedures which query various network databases. Care should be taken when using the database routines since they are not reentrant.

getaddrinfo

The getaddrinfo procedure maps host and service names to socket addresses and associated information in a protocol-independent way.

[Scheme Procedure] getaddrinfo name service [hint flags [hint family [hint socktype

[hint protocol]]]]

[C Function] scm_getaddrinfo (name, service, hint flags, hint family, hint socktype,

hint protocol) Return a list of addrinfo structures containing a socket address and associated information for host name and/or service to be used in creating a socket with which to address the specified service.

     (let* ((ai (car (getaddrinfo "www.gnu.org" "http")))
            (s  (socket (addrinfo:fam ai) (addrinfo:socktype ai)
                        (addrinfo:protocol ai))))
       (connect s (addrinfo:addr ai))
       s)

When service is omitted or is #f, return network-level addresses for name. When name is #f service must be provided and service locations local to the caller are returned.

Additional hints can be provided. When specified, hint flags should be a bitwise-or of zero or more constants among the following:

AI_PASSIVE

Socket address is intended for bind.

AI_CANONNAME

Request for canonical host name, available via addrinfo:canonname. This makes sense mainly when DNS lookups are involved.

AI_NUMERICHOST

Specifies that name is a numeric host address string (e.g., "127.0.0.1"), meaning that name resolution will not be used.

AI_NUMERICSERV

Likewise, specifies that service is a numeric port string (e.g., "80").

AI_ADDRCONFIG

Return only addresses configured on the local system It is highly recommended to provide this flag when the returned socket addresses are to be used to make connections; otherwise, some of the returned addresses could be unreachable or use a protocol that is not supported.

AI_V4MAPPED

When looking up IPv6 addresses, return mapped IPv4 addresses if there is no IPv6 address available at all.

AI_ALL If this flag is set along with AI_V4MAPPED when looking up IPv6 addresses, return all IPv6 addresses as well as all IPv4 addresses, the latter mapped to IPv6 format.

When given, hint family should specify the requested address family, e.g., AF_INET6. Similarly, hint socktype should specify the requested socket type (e.g., SOCK_DGRAM), and hint protocol should specify the requested protocol (its value is interpreted as in calls to socket). On error, an exception with key getaddrinfo-error is thrown, with an error code (an integer) as its argument:

     (catch 'getaddrinfo-error
       (lambda ()
         (getaddrinfo "www.gnu.org" "gopher"))
       (lambda (key errcode)
         (cond ((= errcode EAI_SERVICE)
        (display "doesn't know about Gopher!\n"))
       ((= errcode EAI_NONAME)
        (display "www.gnu.org not found\\n"))
       (else
        (format #t "something wrong: ~a\n"
        (gai-strerror errcode))))))

Error codes are:

EAI_AGAIN

The name or service could not be resolved at this time. Future attempts may succeed.

EAI_BADFLAGS

hint flags contains an invalid value.

EAI_FAIL A non-recoverable error occurred when attempting to resolve the name.

EAI_FAMILY

hint family was not recognized.

EAI_NONAME

Either name does not resolve for the supplied parameters, or neither name nor service were supplied.

EAI_NODATA

This non-POSIX error code can be returned on some systems (GNU and Darwin, at least), for example when name is known but requests that were made turned out no data. Error handling code should be prepared to handle it when it is defined.

EAI_SERVICE

service was not recognized for the specified socket type.

EAI_SOCKTYPE

hint socktype was not recognized.

EAI_SYSTEM

A system error occurred. In C, the error code can be found in errno; this value is not accessible from Scheme, but in practice it provides little information about the actual error cause.

Users are encouraged to read the "POSIX specification (http://www.opengroup.

org/onlinepubs/9699919799/functions/getaddrinfo.html) for more details.

The following procedures take an addrinfo object as returned by getaddrinfo:

[Scheme Procedure] addrinfo:flags ai

Return flags for ai as a bitwise or of AI_values (see above).

[Scheme Procedure] addrinfo:fam ai

Return the address family of ai (a AF_value).

[Scheme Procedure] addrinfo:socktype ai

Return the socket type for ai (a SOCK_value).

[Scheme Procedure] addrinfo:protocol ai

Return the protocol of ai.

[Scheme Procedure] addrinfo:addr ai

Return the socket address associated with ai as a sockaddr object (see Section 7.2.11.3 [Network Socket Address], page 509).

[Scheme Procedure] addrinfo:canonname ai

Return a string for the canonical name associated with ai if the AI_CANONNAME flag was supplied.

The Host Database

A host object is a structure that represents what is known about a network host, and is the usual way of representing a system’s network identity inside software.

The following functions accept a host object and return a selected component:

[Scheme Procedure] hostent:name host

The “official” hostname for host.

[Scheme Procedure] hostent:aliases host

A list of aliases for host.

[Scheme Procedure] hostent:addrtype host

The host address type, one of the AF constants, such as AF_INET or AF_INET6.

[Scheme Procedure] hostent:length host

The length of each address for host, in bytes.

[Scheme Procedure] hostent:addr-list host

The list of network addresses associated with host. For AF_INET these are integer IPv4 address (see Section 7.2.11.1 [Network Address Conversion], page 501).

The following procedures can be used to search the host database. However, getaddrinfo should be preferred over them since it’s more generic and thread-safe.

[Scheme Procedure] gethost [host] [Scheme Procedure] gethostbyname hostname [Scheme Procedure] gethostbyaddr address [C Function] scm_gethost (host)

Look up a host by name or address, returning a host object. The gethost procedure will accept either a string name or an integer address; if given no arguments, it behaves like gethostent (see below). If a name or address is supplied but the address can not be found, an error will be thrown to one of the keys: host-not-found, try-again, no-recovery or no-data, corresponding to the equivalent h_error values. Unusual conditions may result in errors thrown to the system-error or misc_error keys.

     (gethost "www.gnu.org")
     ⇒ #("www.gnu.org" () 2 4 (3353880842))
     (gethostbyname "www.emacs.org")
     ⇒ #("emacs.org" ("www.emacs.org") 2 4 (1073448978))

The following procedures may be used to step through the host database from beginning to end.

[Scheme Procedure] sethostent [stayopen]

Initialize an internal stream from which host objects may be read. This procedure must be called before any calls to gethostent, and may also be called afterward to reset the host entry stream. If stayopen is supplied and is not #f, the database is not closed by subsequent gethostbyname or gethostbyaddr calls, possibly giving an efficiency gain.

[Scheme Procedure] gethostent

Return the next host object from the host database, or #f if there are no more hosts to be found (or an error has been encountered). This procedure may not be used before sethostent has been called.

[Scheme Procedure] endhostent

Close the stream used by gethostent. The return value is unspecified.

[Scheme Procedure] sethost [stayopen] [C Function] scm_sethost (stayopen)

If stayopen is omitted, this is equivalent to endhostent. Otherwise it is equivalent

to sethostent stayopen.

The Network Database

The following functions accept an object representing a network and return a selected component:

[Scheme Procedure] netent:name net

The “official” network name.

[Scheme Procedure] netent:aliases net

A list of aliases for the network.

[Scheme Procedure] netent:addrtype net

The type of the network number. Currently, this returns only AF_INET.

[Scheme Procedure] netent:net net

The network number.

The following procedures are used to search the network database:

[Scheme Procedure] getnet [net] [Scheme Procedure] getnetbyname net-name [Scheme Procedure] getnetbyaddr net-number [C Function] scm_getnet (net)

Look up a network by name or net number in the network database. The net-name argument must be a string, and the net-number argument must be an integer. getnet will accept either type of argument, behaving like getnetent (see below) if no arguments are given.

The following procedures may be used to step through the network database from beginning to end.

[Scheme Procedure] setnetent [stayopen]

Initialize an internal stream from which network objects may be read. This procedure must be called before any calls to getnetent, and may also be called afterward to reset the net entry stream. If stayopen is supplied and is not #f, the database is not closed by subsequent getnetbyname or getnetbyaddr calls, possibly giving an efficiency gain.

[Scheme Procedure] getnetent

Return the next entry from the network database.

[Scheme Procedure] endnetent

Close the stream used by getnetent. The return value is unspecified.

[Scheme Procedure] setnet [stayopen] [C Function] scm_setnet (stayopen)

If stayopen is omitted, this is equivalent to endnetent. Otherwise it is equivalent to

setnetent stayopen.

The Protocol Database

The following functions accept an object representing a protocol and return a selected component:

[Scheme Procedure] protoent:name protocol

The “official” protocol name.

[Scheme Procedure] protoent:aliases protocol

A list of aliases for the protocol.

[Scheme Procedure] protoent:proto protocol

The protocol number.

The following procedures are used to search the protocol database:

[Scheme Procedure] getproto [protocol] [Scheme Procedure] getprotobyname name [Scheme Procedure] getprotobynumber number [C Function] scm_getproto (protocol)

Look up a network protocol by name or by number. getprotobyname takes a string argument, and getprotobynumber takes an integer argument. getproto will accept either type, behaving like getprotoent (see below) if no arguments are supplied.

The following procedures may be used to step through the protocol database from beginning to end.

[Scheme Procedure] setprotoent [stayopen]

Initialize an internal stream from which protocol objects may be read. This procedure must be called before any calls to getprotoent, and may also be called afterward to reset the protocol entry stream. If stayopen is supplied and is not #f, the database is not closed by subsequent getprotobyname or getprotobynumber calls, possibly giving an efficiency gain.

[Scheme Procedure] getprotoent

Return the next entry from the protocol database.

[Scheme Procedure] endprotoent

Close the stream used by getprotoent. The return value is unspecified.

[Scheme Procedure] setproto [stayopen] [C Function] scm_setproto (stayopen)

If stayopen is omitted, this is equivalent to endprotoent. Otherwise it is equivalent

to setprotoent stayopen.

The Service Database

The following functions accept an object representing a service and return a selected component:

[Scheme Procedure] servent:name serv

The “official” name of the network service.

[Scheme Procedure] servent:aliases serv

A list of aliases for the network service.

[Scheme Procedure] servent:port serv

The Internet port used by the service.

[Scheme Procedure] servent:proto serv

The protocol used by the service. A service may be listed many times in the database under different protocol names.

The following procedures are used to search the service database:

[Scheme Procedure] getserv [name [protocol]]

[Scheme Procedure] getservbyname name protocol [Scheme Procedure] getservbyport port protocol

[C Function] scm_getserv (name, protocol)

Look up a network service by name or by service number, and return a network service object. The protocol argument specifies the name of the desired protocol; if the protocol found in the network service database does not match this name, a system error is signalled. The getserv procedure will take either a service name or number as its first argument; if given no arguments, it behaves like getservent (see below).

     (getserv "imap" "tcp")
     ⇒ #("imap2" ("imap") 143 "tcp")
     (getservbyport 88 "udp")
     ⇒ #("kerberos" ("kerberos5" "krb5") 88 "udp")

The following procedures may be used to step through the service database from beginning to end.

[Scheme Procedure] setservent [stayopen]

Initialize an internal stream from which service objects may be read. This procedure must be called before any calls to getservent, and may also be called afterward to reset the service entry stream. If stayopen is supplied and is not #f, the database is not closed by subsequent getservbyname or getservbyport calls, possibly giving an efficiency gain.

[Scheme Procedure] getservent

Return the next entry from the services database.

[Scheme Procedure] endservent

Close the stream used by getservent. The return value is unspecified.

[Scheme Procedure] setserv [stayopen] [C Function] scm_setserv (stayopen)

If stayopen is omitted, this is equivalent to endservent. Otherwise it is equivalent

to setservent stayopen.

7.2.11.3 Network Socket Address

A socket address object identifies a socket endpoint for communication. In the case of AF_INET for instance, the socket address object comprises the host address (or interface on the host) and a port number which specifies a particular open socket in a running client or server process. A socket address object can be created with,

[Scheme Procedure] make-socket-address AF INET ipv4addr port

[Scheme Procedure] make-socket-address AF INET6 ipv6addr port [flowinfo

[scopeid]]

[Scheme Procedure] make-socket-address AF UNIX path

[C Function] scm_make_socket_address (family, address, arglist)

Return a new socket address object. The first argument is the address family, one of the AF constants, then the arguments vary according to the family. For AF_INET the arguments are an IPv4 network address number (see Section 7.2.11.1 [Network Address Conversion], page 501), and a port number. For AF_INET6 the arguments are an IPv6 network address number and a port number. Optional flowinfo and scopeid arguments may be given (both integers, default 0). For AF_UNIX the argument is a filename (a string). The C function scm_make_socket_address takes the family and address arguments directly, then arglist is a list of further arguments, being the port for IPv4, port and optional flowinfo and scopeid for IPv6, or the empty list SCM_EOL for Unix domain.

The following functions access the fields of a socket address object,

[Scheme Procedure] sockaddr:fam sa

Return the address family from socket address object sa. This is one of the AF constants (e.g. AF_INET).

[Scheme Procedure] sockaddr:path sa

For an AF_UNIX socket address object sa, return the filename.

[Scheme Procedure] sockaddr:addr sa

For an AF_INET or AF_INET6 socket address object sa, return the network address number.

[Scheme Procedure] sockaddr:port sa

For an AF_INET or AF_INET6 socket address object sa, return the port number.

[Scheme Procedure] sockaddr:flowinfo sa

For an AF_INET6 socket address object sa, return the flowinfo value.

[Scheme Procedure] sockaddr:scopeid sa

For an AF_INET6 socket address object sa, return the scope ID value.

The functions below convert to and from the C struct sockaddr (see Section “Address Formats” in The GNU C Library Reference Manual). That structure is a generic type,

an application can cast to or from struct sockaddr_in, struct sockaddr_in6 or struct

sockaddr_un according to the address family.

In a struct sockaddr taken or returned, the byte ordering in the fields follows the C conventions (see Section “Byte Order Conversion” in The GNU C Library Reference Manual). This means network byte order for AF_INET host address (sin_addr.s_addr) and port number (sin_port), and AF_INET6 port number (sin6_port). But at the Scheme level these values are taken or returned in host byte order, so the port is an ordinary integer, and the host address likewise is an ordinary integer (as described in Section 7.2.11.1 [Network Address Conversion], page 501).

[C Function] struct sockaddr * scm_c_make_socket_address (SCM family,

SCM address, SCM args, size t *outsize) Return a newly-malloced struct sockaddr created from arguments like those taken

by scm_make_socket_address above.

The size (in bytes) of the struct sockaddr return is stored into *outsize. An application must call free to release the returned structure when no longer required.

[C Function] SCM scm_from_sockaddr (const struct sockaddr *address, unsigned

address size) Return a Scheme socket address object from the C address structure. address size is the size in bytes of address.

[C Function] struct sockaddr * scm_to_sockaddr (SCM address, size t

*address size) Return a newly-malloced struct sockaddr from a Scheme level socket address object. The size (in bytes) of the struct sockaddr return is stored into *outsize. An application must call free to release the returned structure when no longer required.

7.2.11.4 Network Sockets and Communication

Socket ports can be created using socket and socketpair. The ports are initially unbuffered, to make reading and writing to the same port more reliable. A buffer can be added to the port using setvbuf; see Section 7.2.2 [Ports and File Descriptors], page 469.

Most systems have limits on how many files and sockets can be open, so it’s strongly recommended that socket ports be closed explicitly when no longer required (see Section 6.14.1 [Ports], page 316).

Some of the underlying C functions take values in network byte order, but the convention in Guile is that at the Scheme level everything is ordinary host byte order and conversions are made automatically where necessary.

[Scheme Procedure] socket family style proto [C Function] scm_socket (family, style, proto)

Return a new socket port of the type specified by family, style and proto. All three parameters are integers. The possible values for family are as follows, where supported by the system,

[Variable] PF_UNIX

[Variable] PF_INET

[Variable] PF_INET6

The possible values for style are as follows, again where supported by the system,

[Variable] SOCK_STREAM
[Variable] SOCK_DGRAM

[Variable] SOCK_RAW

[Variable] SOCK_RDM

[Variable] SOCK_SEQPACKET

proto can be obtained from a protocol name using getprotobyname (see Section 7.2.11.2 [Network Databases], page 503). A value of zero means the default protocol, which is usually right. A socket cannot by used for communication until it has been connected somewhere, usually with either connect or accept below.

[Scheme Procedure] socketpair family style proto [C Function] scm_socketpair (family, style, proto)

Return a pair, the car and cdr of which are two unnamed socket ports connected to each other. The connection is full-duplex, so data can be transferred in either direction between the two. family, style and proto are as per socket above. But many systems only support socket pairs in the PF_UNIX family. Zero is likely to be the only meaningful value for proto.

[Scheme Procedure] getsockopt sock level optname [Scheme Procedure] setsockopt sock level optname value [C Function] scm_getsockopt (sock, level, optname) [C Function] scm_setsockopt (sock, level, optname, value)

Get or set an option on socket port sock. getsockopt returns the current value. setsockopt sets a value and the return is unspecified. level is an integer specifying a protocol layer, either SOL_SOCKET for socket level options, or a protocol number from the IPPROTO constants or getprotoent (see Section 7.2.11.2 [Network Databases], page 503).

[Variable] SOL_SOCKET
[Variable] IPPROTO_IP
[Variable] IPPROTO_TCP
[Variable] IPPROTO_UDP

optname is an integer specifying an option within the protocol layer. For SOL_SOCKET level the following optnames are defined (when provided by the system). For their meaning see Section “Socket-Level Options” in The GNU C Library Reference Manual, or man 7 socket.

[Variable] SO_DEBUG

[Variable] SO_REUSEADDR

[Variable] SO_STYLE

[Variable] SO_TYPE

[Variable] SO_ERROR

[Variable] SO_DONTROUTE
[Variable] SO_BROADCAST

[Variable] SO_SNDBUF

[Variable] SO_RCVBUF

[Variable] SO_KEEPALIVE
[Variable] SO_OOBINLINE
[Variable] SO_NO_CHECK
[Variable] SO_PRIORITY
[Variable] SO_REUSEPORT

The value taken or returned is an integer.

[Variable] SO_LINGER

The value taken or returned is a pair of integers (ENABLE . TIMEOUT). On old systems without timeout support (ie. without struct linger), only ENABLE has an effect but the value in Guile is always a pair.

For IP level (IPPROTO_IP) the following optnames are defined (when provided by the system). See man ip for what they mean.

[Variable] IP_MULTICAST_IF

This sets the source interface used by multicast traffic.

[Variable] IP_MULTICAST_TTL

This sets the default TTL for multicast traffic. This defaults to 1 and should be increased to allow traffic to pass beyond the local network.

[Variable] IP_ADD_MEMBERSHIP
[Variable] IP_DROP_MEMBERSHIP

These can be used only with setsockopt, not getsockopt. value is a pair (MULTIADDR . INTERFACEADDR) of integer IPv4 addresses (see Section 7.2.11.1 [Network Address Conversion], page 501). MULTIADDR is a multicast address to be added to or dropped from the interface INTERFACEADDR. IN-TERFACEADDR can be INADDR_ANY to have the system select the interface. INTERFACEADDR can also be an interface index number, on systems supporting that.

[Scheme Procedure] shutdown sock how [C Function] scm_shutdown (sock, how)

Sockets can be closed simply by using close-port. The shutdown procedure allows reception or transmission on a connection to be shut down individually, according to the parameter how:

0 Stop receiving data for this socket. If further data arrives, reject it.

1 Stop trying to transmit data from this socket. Discard any data waiting to be sent. Stop looking for acknowledgement of data already sent; don’t retransmit it if it is lost.

2 Stop both reception and transmission.

The return value is unspecified.

[Scheme Procedure] connect sock sockaddr [Scheme Procedure] connect sock AF INET ipv4addr port [Scheme Procedure] connect sock AF INET6 ipv6addr port [flowinfo [scopeid]] [Scheme Procedure] connect sock AF UNIX path [C Function] scm_connect (sock, fam, address, args)

Initiate a connection on socket port sock to a given address. The destination is either a socket address object, or arguments the same as make-socket-address would take to make such an object (see Section 7.2.11.3 [Network Socket Address], page 509). The return value is unspecified.

     (connect sock AF_INET INADDR_LOOPBACK 23)
     (connect sock (make-socket-address AF_INET INADDR_LOOPBACK 23))

[Scheme Procedure] bind sock sockaddr [Scheme Procedure] bind sock AF INET ipv4addr port [Scheme Procedure] bind sock AF INET6 ipv6addr port [flowinfo [scopeid]] [Scheme Procedure] bind sock AF UNIX path [C Function] scm_bind (sock, fam, address, args)

Bind socket port sock to the given address. The address is either a socket address object, or arguments the same as make-socket-address would take to make such an object (see Section 7.2.11.3 [Network Socket Address], page 509). The return value is unspecified. Generally a socket is only explicitly bound to a particular address when making a server, i.e. to listen on a particular port. For an outgoing connection the system will assign a local address automatically, if not already bound.

     (bind sock AF_INET INADDR_ANY 12345)
     (bind sock (make-socket-address AF_INET INADDR_ANY 12345))

[Scheme Procedure] listen sock backlog [C Function] scm_listen (sock, backlog)

Enable sock to accept connection requests. backlog is an integer specifying the maximum length of the queue for pending connections. If the queue fills, new clients will fail to connect until the server calls accept to accept a connection from the queue. The return value is unspecified.

[Scheme Procedure] accept sock [C Function] scm_accept (sock)

Accept a connection from socket port sock which has been enabled for listening with listen above. If there are no incoming connections in the queue, wait until one is available (unless O_NONBLOCK has been set on the socket, see Section 7.2.2 [Ports and File Descriptors], page 469). The return value is a pair. The car is a new socket port, connected and ready to communicate. The cdr is a socket address object (see Section 7.2.11.3 [Network Socket Address], page 509) which is where the remote connection is from (like getpeername below). All communication takes place using the new socket returned. The given sock remains bound and listening, and accept may be called on it again to get another incoming connection when desired.

[Scheme Procedure] getsockname sock [C Function] scm_getsockname (sock)

Return a socket address object which is the where sock is bound locally. sock may have obtained its local address from bind (above), or if a connect is done with an otherwise unbound socket (which is usual) then the system will have assigned an address. Note that on many systems the address of a socket in the AF_UNIX namespace cannot be read.

[Scheme Procedure] getpeername sock [C Function] scm_getpeername (sock)

Return a socket address object which is where sock is connected to, i.e. the remote endpoint. Note that on many systems the address of a socket in the AF_UNIX namespace cannot be read.

[Scheme Procedure] recv! sock buf [flags] [C Function] scm_recv (sock, buf, flags)

Receive data from a socket port. sock must already be bound to the address from which data is to be received. buf is a bytevector into which the data will be written. The size of buf limits the amount of data which can be received: in the case of packet protocols, if a packet larger than this limit is encountered then some data will be irrevocably lost. The optional flags argument is a value or bitwise OR of MSG_OOB, MSG_PEEK, MSG_

DONTROUTE etc.

The value returned is the number of bytes read from the socket. Note that the data is read directly from the socket file descriptor: any unread buffered port data is ignored.

[Scheme Procedure] send sock message [flags] [C Function] scm_send (sock, message, flags)

Transmit bytevector message on socket port sock. sock must already be bound to a destination address. The value returned is the number of bytes transmitted—it’s possible for this to be less than the length of message if the socket is set to be nonblocking. The optional flags argument is a value or bitwise OR of MSG_OOB, MSG_PEEK,

MSG_DONTROUTE etc.

Note that the data is written directly to the socket file descriptor: any unflushed buffered port data is ignored.

[Scheme Procedure] recvfrom! sock buf [flags [start [end]]] [C Function] scm_recvfrom (sock, buf, flags, start, end)

Receive data from socket port sock, returning the originating address as well as the data. This function is usually for datagram sockets, but can be used on streamoriented sockets too. The data received is stored in bytevector buf, using either the whole bytevector or just the region between the optional start and end positions. The size of buf limits the amount of data that can be received. For datagram protocols if a packet larger than this is received then excess bytes are irrevocably lost. The return value is a pair. The car is the number of bytes read. The cdr is a socket address object (see Section 7.2.11.3 [Network Socket Address], page 509) which is where the data came from, or #f if the origin is unknown. The optional flags argument is a or bitwise-OR (logior) of MSG_OOB, MSG_PEEK, MSG_

DONTROUTE etc.

Data is read directly from the socket file descriptor, any buffered port data is ignored. On a GNU/Linux system recvfrom! is not multi-threading, all threads stop while a recvfrom! call is in progress. An application may need to use select, O_NONBLOCK or MSG_DONTWAIT to avoid this.

[Scheme Procedure] sendto sock message sockaddr [flags] [Scheme Procedure] sendto sock message AF INET ipv4addr port [flags] [Scheme Procedure] sendto sock message AF INET6 ipv6addr port [flowinfo [scopeid

[flags]]]

[Scheme Procedure] sendto sock message AF UNIX path [flags] [C Function] scm_sendto (sock, message, fam, address, args and flags)

Transmit bytevector message as a datagram socket port sock. The destination is specified either as a socket address object, or as arguments the same as would be taken by make-socket-address to create such an object (see Section 7.2.11.3 [Network Socket Address], page 509). The destination address may be followed by an optional flags argument which is a logior (see Section 6.6.2.13 [Bitwise Operations], page 125) of MSG_OOB, MSG_PEEK,

MSG_DONTROUTE etc.

The value returned is the number of bytes transmitted – it’s possible for this to be less than the length of message if the socket is set to be non-blocking. Note that the data is written directly to the socket file descriptor: any unflushed buffered port data is ignored.

7.2.11.5 Network Socket Examples

The following give examples of how to use network sockets.

Internet Socket Client Example

The following example demonstrates an Internet socket client. It connects to the HTTP daemon running on the local machine and returns the contents of the root index URL.

(let ((s (socket PF_INET SOCK_STREAM 0)))
  (connect s AF_INET (inet-pton AF_INET "127.0.0.1") 80)
  (display "GET / HTTP/1.0\r\n\r\n" s)
  (do ((line (read-line s) (read-line s)))
      ((eof-object? line))
    (display line)
    (newline)))

Internet Socket Server Example

The following example shows a simple Internet server which listens on port 2904 for incoming connections and sends a greeting back to the client.

(let ((s (socket PF_INET SOCK_STREAM 0)))
  (setsockopt s SOL_SOCKET SO_REUSEADDR 1)

;; Specific address?

;; (bind s AF INET (inet-pton AF INET "127.0.0.1") 2904)

  (bind s AF_INET INADDR_ANY 2904)
  (listen s 5)
  (simple-format #t "Listening for clients in pid: ~S" (getpid))
  (newline)
  (while #t
    (let* ((client-connection (accept s))
           (client-details (cdr client-connection))
           (client (car client-connection)))

(simple-format #t "Got new client connection: ~S"
               client-details)
(newline)
(simple-format #t "Client address: ~S"
               (gethostbyaddr
                (sockaddr:addr client-details)))
(newline)

;; Send back the greeting to the client port

(display "Hello client\r\n" client)
(close client))))

7.2.12 System Identification

This section lists the various procedures Guile provides for accessing information about the system it runs on.

[Scheme Procedure] uname [C Function] scm_uname ()

Return an object with some information about the computer system the program is running on.

The following procedures accept an object as returned by uname and return a selected component (all of which are strings).

[Scheme Procedure] utsname:sysname un

The name of the operating system.

[Scheme Procedure] utsname:nodename un

The network name of the computer.

[Scheme Procedure] utsname:release un

The current release level of the operating system implementation.

[Scheme Procedure] utsname:version un

The current version level within the release of the operating system.

[Scheme Procedure] utsname:machine un

A description of the hardware.

[Scheme Procedure] gethostname

[C Function] scm_gethostname ()

Return the host name of the current processor.

[Scheme Procedure] sethostname name [C Function] scm_sethostname (name)

Set the host name of the current processor to name. May only be used by the superuser. The return value is not specified.

7.2.13 Locales

[Scheme Procedure] setlocale category [locale] [C Function] scm_setlocale (category, locale)

Get or set the current locale, used for various internationalizations. Locales are strings, such as ‘sv_SE’.

If locale is given then the locale for the given category is set and the new value returned. If locale is not given then the current value is returned. category should be one of the following values (see Section “Locale Categories” in The GNU C Library Reference Manual):

[Variable] LC_ALL

[Variable] LC_COLLATE

[Variable] LC_CTYPE

[Variable] LC_MESSAGES
[Variable] LC_MONETARY
[Variable] LC_NUMERIC

[Variable] LC_TIME

A common usage is ‘(setlocale LC_ALL "")’, which initializes all categories based on standard environment variables (LANG etc). For full details on categories and locale names see Section “Locales and Internationalization” in The GNU C Library Reference Manual.

Note that setlocale affects locale settings for the whole process. See Section 6.24.1 [i18n Introduction], page 436, for a thread-safe alternative.

7.2.14 Encryption

Please note that the procedures in this section are not suited for strong encryption, they are only interfaces to the well-known and common system library functions of the same name. They are just as good (or bad) as the underlying functions, so you should refer to your system documentation before using them (see Section “Encrypting Passwords” in The GNU C Library Reference Manual).

[Scheme Procedure] crypt key salt [C Function] scm_crypt (key, salt)

Encrypt key, with the addition of salt (both strings), using the crypt C library call.

Although getpass is not an encryption procedure per se, it appears here because it is often used in combination with crypt:

[Scheme Procedure] getpass prompt [C Function] scm_getpass (prompt)

Display prompt to the standard error output and read a password from /dev/tty. If this file is not accessible, it reads from standard input. The password may be up to 127 characters in length. Additional characters and the terminating newline character are discarded. While reading the password, echoing and the generation of signals by special characters is disabled.

7.3 HTTP, the Web, and All That

It has always been possible to connect computers together and share information between them, but the rise of the World Wide Web over the last couple of decades has made it much easier to do so. The result is a richly connected network of computation, in which Guile forms a part.

By “the web”, we mean the HTTP protocol^{2} as handled by servers, clients, proxies, caches, and the various kinds of messages and message components that can be sent and received by that protocol, notably HTML.

On one level, the web is text in motion: the protocols themselves are textual (though the payload may be binary), and it’s possible to create a socket and speak text to the web. But such an approach is obviously primitive. This section details the higher-level data types and operations provided by Guile: URIs, HTTP request and response records, and a conventional web server implementation.

The material in this section is arranged in ascending order, in which later concepts build on previous ones. If you prefer to start with the highest-level perspective, see Section 7.3.10 [Web Examples], page 542, and work your way back.

7.3.1 Types and the Web

It is a truth universally acknowledged, that a program with good use of data types, will be free from many common bugs. Unfortunately, the common practice in web programming seems to ignore this maxim. This subsection makes the case for expressive data types in web programming.

By “expressive data types”, we mean that the data types say something about how a program solves a problem. For example, if we choose to represent dates using SRFI 19 date records (see Section 7.5.16 [SRFI-19], page 583), this indicates that there is a part of the program that will always have valid dates. Error handling for a number of basic cases, like invalid dates, occurs on the boundary in which we produce a SRFI 19 date record from other types, like strings.

With regards to the web, data types are helpful in the two broad phases of HTTP messages: parsing and generation.

Consider a server, which has to parse a request, and produce a response. Guile will parse the request into an HTTP request object (see Section 7.3.6 [Requests], page 534), with each header parsed into an appropriate Scheme data type. This transition from an incoming stream of characters to typed data is a state change in a program—the strings might parse, or they might not, and something has to happen if they do not. (Guile throws an error in this case.) But after you have the parsed request, “client” code (code built on top of the Guile web framework) will not have to check for syntactic validity. The types already make this information manifest.

This state change on the parsing boundary makes programs more robust, as they themselves are freed from the need to do a number of common error checks, and they can use normal Scheme procedures to handle a request instead of ad-hoc string parsers.

2 Yes, the P is for protocol, but this phrase appears repeatedly in RFC 2616.

The need for types on the response generation side (in a server) is more subtle, though not less important. Consider the example of a POST handler, which prints out the text that a user submits from a form. Such a handler might include a procedure like this:

     ;; First, a helper procedure
     (define (para . contents)
       (string-append "<p>" (string-concatenate contents) "</p>"))
     ;; Now the meat of our simple web application
     (define (you-said text)
       (para "You said: " text))
     (display (you-said "Hi!"))
     ⊣ <p>You said: Hi!</p>

This is a perfectly valid implementation, provided that the incoming text does not contain the special HTML characters ‘<’, ‘>’, or ‘&’. But this provision of a restricted character set is not reflected anywhere in the program itself: we must assume that the programmer understands this, and performs the check elsewhere.

Unfortunately, the short history of the practice of programming does not bear out this assumption. A cross-site scripting (XSS) vulnerability is just such a common error in which unfiltered user input is allowed into the output. A user could submit a crafted comment to your web site which results in visitors running malicious Javascript, within the security context of your domain:

     (display (you-said "<script src=\"http://bad.com/nasty.js\" />"))
     ⊣ <p>You said: <script src="http://bad.com/nasty.js" /></p>

The fundamental problem here is that both user data and the program template are represented using strings. This identity means that types can’t help the programmer to make a distinction between these two, so they get confused.

There are a number of possible solutions, but perhaps the best is to treat HTML not as strings, but as native s-expressions: as SXML. The basic idea is that HTML is either text, represented by a string, or an element, represented as a tagged list. So ‘foo’ becomes ‘"foo"’, and ‘<b>foo</b>’ becomes ‘(b "foo")’. Attributes, if present, go in a tagged list

headed by ‘@’, like ‘(img (@ (src "http://example.com/foo.png")))’. See Section 7.22

[SXML], page 697, for more information.

The good thing about SXML is that HTML elements cannot be confused with text. Let’s make a new definition of para:

     (define (para . contents)
       `(p ,@contents))
     (use-modules (sxml simple))
     (sxml->xml (you-said "Hi!"))
     ⊣ <p>You said: Hi!</p>
     (sxml->xml (you-said "<i>Rats, foiled again!</i>"))
     ⊣ <p>You said: &lt;i&gt;Rats, foiled again!&lt;/i&gt;</p>

So we see in the second example that HTML elements cannot be unwittingly introduced into the output. However it is now perfectly acceptable to pass SXML to you-said; in fact, that is the big advantage of SXML over everything-as-a-string.

     (sxml->xml (you-said (you-said "<Hi!>")))
     ⊣ <p>You said: <p>You said: &lt;Hi!&gt;</p></p>

The SXML types allow procedures to compose. The types make manifest which parts are HTML elements, and which are text. So you needn’t worry about escaping user input; the type transition back to a string handles that for you. XSS vulnerabilities are a thing of the past.

Well. That’s all very nice and opinionated and such, but how do I use the thing? Read on!

7.3.2 Universal Resource Identifiers

Guile provides a standard data type for Universal Resource Identifiers (URIs), as defined in RFC 3986.

The generic URI syntax is as follows:

     URI := scheme ":" ["//" [userinfo "@"] host [":" port]] path \
            [ "?" query ] [ "#" fragment ]

For example, in the URI, ‘http://www.gnu.org/help/’, the scheme is http, the host is www.gnu.org, the path is /help/, and there is no userinfo, port, query, or fragment. All URIs have a scheme and a path (though the path might be empty). Some URIs have a host, and some of those have ports and userinfo. Any URI might have a query part or a fragment.

Userinfo is something of an abstraction, as some legacy URI schemes allowed userinfo of the form username:passwd. But since passwords do not belong in URIs, the RFC does not want to condone this practice, so it calls anything before the @ sign userinfo.

Properly speaking, a fragment is not part of a URI. For example, when a web browser fol-

lows a link to ‘http://example.com/#foo’, it sends a request for ‘http://example.com/’,

then looks in the resulting page for the fragment identified foo reference. A fragment identifies a part of a resource, not the resource itself. But it is useful to have a fragment field in the URI record itself, so we hope you will forgive the inconsistency.

     (use-modules (web uri))

The following procedures can be found in the (web uri) module. Load it into your Guile, using a form like the above, to have access to them.

[Scheme Procedure] build-uri scheme [#:userinfo=#f] [#:host=#f] [#:port=#f]

[#:path=""] [#:query=#f] [#:fragment=#f] [#:validate?=#t] Construct a URI object. scheme should be a symbol, port either a positive, exact integer or #f, and the rest of the fields are either strings or #f. If validate? is true, also run some consistency checks to make sure that the constructed URI is valid.

[Scheme Procedure] uri? obj [Scheme Procedure] uri-scheme uri [Scheme Procedure] uri-userinfo uri [Scheme Procedure] uri-host uri [Scheme Procedure] uri-port uri [Scheme Procedure] uri-path uri [Scheme Procedure] uri-query uri [Scheme Procedure] uri-fragment uri

A predicate and field accessors for the URI record type. The URI scheme will be a symbol, the port either a positive, exact integer or #f, and the rest either strings or #f if not present.

[Scheme Procedure] string->uri string

Parse string into a URI object. Return #f if the string could not be parsed.

[Scheme Procedure] uri->string uri

Serialize uri to a string. If the URI has a port that is the default port for its scheme, the port is not included in the serialization.

[Scheme Procedure] declare-default-port! scheme port

Declare a default port for the given URI scheme.

[Scheme Procedure] uri-decode str [#:encoding="utf-8"]

Percent-decode the given str, according to encoding, which should be the name of a character encoding. Note that this function should not generally be applied to a full URI string. For paths, use split-and-decode-uri-path instead. For query strings, split the query on & and = boundaries, and decode the components separately. Note also that percent-encoded strings encode bytes, not characters. There is no guarantee that a given byte sequence is a valid string encoding. Therefore this routine may signal an error if the decoded bytes are not valid for the given encoding. Pass #f for encoding if you want decoded bytes as a bytevector directly. See Section 6.14.1 [Ports], page 316, for more information on character encodings. Returns a string of the decoded characters, or a bytevector if encoding was #f.

Fixme: clarify return type. indicate default values. type of unescaped-chars.

[Scheme Procedure] uri-encode str [#:encoding="utf-8"] [#:unescaped-chars] Percent-encode any character not in the character set, unescaped-chars. The default character set includes alphanumerics from ASCII, as well as the special characters ‘-’, ‘.’, ‘_’, and ‘~’. Any other character will be percent-encoded, by writing out the character to a bytevector within the given encoding, then encoding each byte as %HH, where HH is the hexadecimal representation of the byte.

[Scheme Procedure] split-and-decode-uri-path path

Split path into its components, and decode each component, removing empty components. For example, "/foo/bar%20baz/" decodes to the two-element list, ("foo" "bar

baz").
[Scheme Procedure] encode-and-join-uri-path parts

URI-encode each element of parts, which should be a list of strings, and join the parts together with / as a delimiter.

For example, the list ("scrambled eggs" "biscuits&gravy") encodes as
"scrambled%20eggs/biscuits%26gravy".

7.3.3 The Hyper-Text Transfer Protocol

The initial motivation for including web functionality in Guile, rather than rely on an external package, was to establish a standard base on which people can share code. To that end, we continue the focus on data types by providing a number of low-level parsers and unparsers for elements of the HTTP protocol.

If you are want to skip the low-level details for now and move on to web pages, see Section 7.3.8 [Web Client], page 538, and see Section 7.3.9 [Web Server], page 540. Otherwise, load the HTTP module, and read on.

(use-modules (web http))

The focus of the (web http) module is to parse and unparse standard HTTP headers, representing them to Guile as native data structures. For example, a Date: header will be represented as a SRFI-19 date record (see Section 7.5.16 [SRFI-19], page 583), rather than as a string.

Guile tries to follow RFCs fairly strictly—the road to perdition being paved with compatibility hacks—though some allowances are made for not-too-divergent texts.

Header names are represented as lower-case symbols.

[Scheme Procedure] string->header name

Parse name to a symbolic header name.

[Scheme Procedure] header->string sym

Return the string form for the header named sym.

For example:

(string->header "Content-Length")
⇒ content-length
(header->string 'content-length)
⇒ "Content-Length"
(string->header "FOO")
⇒ foo
(header->string 'foo)
⇒ "Foo"

Guile keeps a registry of known headers, their string names, and some parsing and serialization procedures. If a header is unknown, its string name is simply its symbol name in title-case.

[Scheme Procedure] known-header? sym

Return #t if sym is a known header, with associated parsers and serialization procedures, or #f otherwise.

[Scheme Procedure] header-parser sym

Return the value parser for headers named sym. The result is a procedure that takes one argument, a string, and returns the parsed value. If the header isn’t known to Guile, a default parser is returned that passes through the string unchanged.

[Scheme Procedure] header-validator sym

Return a predicate which returns #t if the given value is valid for headers named sym. The default validator for unknown headers is string?.

[Scheme Procedure] header-writer sym

Return a procedure that writes values for headers named sym to a port. The resulting procedure takes two arguments: a value and a port. The default writer is display.

For more on the set of headers that Guile knows about out of the box, see Section 7.3.4 [HTTP Headers], page 525. To add your own, use the declare-header! procedure:

[Scheme Procedure] declare-header! name parser validator writer

[#:multiple?=#f] Declare a parser, validator, and writer for a given header.

For example, let’s say you are running a web server behind some sort of proxy, and your proxy adds an X-Client-Address header, indicating the IPv4 address of the original client. You would like for the HTTP request record to parse out this header to a Scheme value, instead of leaving it as a string. You could register this header with Guile’s HTTP stack like this:

(declare-header! "X-Client-Address"
  (lambda (str)
    (inet-aton str))
  (lambda (ip)
    (and (integer? ip) (exact? ip) (<= 0 ip #xffffffff)))
  (lambda (ip port)
    (display (inet-ntoa ip) port)))
[Scheme Procedure] declare-opaque-header! name

A specialised version of declare-header! for the case in which you want a header’s value to be returned/written “as-is”.

[Scheme Procedure] valid-header? sym val

Return a true value if val is a valid Scheme value for the header with name sym, or #f otherwise.

Now that we have a generic interface for reading and writing headers, we do just that.

[Scheme Procedure] read-header port

Read one HTTP header from port. Return two values: the header name and the parsed Scheme value. May raise an exception if the header was known but the value was invalid. Returns the end-of-file object for both values if the end of the message body was reached (i.e., a blank line).

[Scheme Procedure] parse-header name val

Parse val, a string, with the parser for the header named name. Returns the parsed value.

[Scheme Procedure] write-header name val port

Write the given header name and value to port, using the writer from header-writer.

[Scheme Procedure] read-headers port

Read the headers of an HTTP message from port, returning them as an ordered alist.

[Scheme Procedure] write-headers headers port

Write the given header alist to port. Doesn’t write the final ‘\r\n’, as the user might want to add another header.

The (web http) module also has some utility procedures to read and write request and response lines.

[Scheme Procedure] parse-http-method str [start] [end]

Parse an HTTP method from str. The result is an upper-case symbol, like GET.

[Scheme Procedure] parse-http-version str [start] [end]

Parse an HTTP version from str, returning it as a major–minor pair. For example, HTTP/1.1 parses as the pair of integers, (1 . 1).

[Scheme Procedure] parse-request-uri str [start] [end]

Parse a URI from an HTTP request line. Note that URIs in requests do not have to have a scheme or host name. The result is a URI object.

[Scheme Procedure] read-request-line port

Read the first line of an HTTP request from port, returning three values: the method, the URI, and the version.

[Scheme Procedure] write-request-line method uri version port Write the first line of an HTTP request to port.

[Scheme Procedure] read-response-line port

Read the first line of an HTTP response from port, returning three values: the HTTP version, the response code, and the “reason phrase”.

[Scheme Procedure] write-response-line version code reason-phrase port Write the first line of an HTTP response to port.

7.3.4 HTTP Headers

In addition to defining the infrastructure to parse headers, the (web http) module defines specific parsers and unparsers for all headers defined in the HTTP/1.1 standard.

For example, if you receive a header named ‘Accept-Language’ with a value ‘en, es;q=0.8’, Guile parses it as a quality list (defined below):

(parse-header 'accept-language "en, es;q=0.8")
⇒ ((1000 . "en") (800 . "es"))

The format of the value for ‘Accept-Language’ headers is defined below, along with all other headers defined in the HTTP standard. (If the header were unknown, the value would have been returned as a string.)

For brevity, the header definitions below are given in the form, Type name, indicating that values for the header name will be of the given Type. Since Guile internally treats header names in lower case, in this document we give types title-cased names. A short description of the each header’s purpose and an example follow.

For full details on the meanings of all of these headers, see the HTTP 1.1 standard, RFC 2616.

7.3.4.1 HTTP Header Types

Here we define the types that are used below, when defining headers.

[HTTP Header Type] Date

A SRFI-19 date.

[HTTP Header Type] KVList

A list whose elements are keys or key-value pairs. Keys are parsed to symbols. Values are strings by default. Non-string values are the exception, and are mentioned explicitly below, as appropriate.

[HTTP Header Type] SList

A list of strings.

[HTTP Header Type] Quality

An exact integer between 0 and 1000. Qualities are used to express preference, given multiple options. An option with a quality of 870, for example, is preferred over an option with quality 500. (Qualities are written out over the wire as numbers between 0.0 and 1.0, but since the standard only allows three digits after the decimal, it’s equivalent to integers between 0 and 1000, so that’s what Guile uses.)

[HTTP Header Type] QList

A quality list: a list of pairs, the car of which is a quality, and the cdr a string. Used to express a list of options, along with their qualities.

[HTTP Header Type] ETag

An entity tag, represented as a pair. The car of the pair is an opaque string, and the cdr is #t if the entity tag is a “strong” entity tag, and #f otherwise.

7.3.4.2 General Headers

General HTTP headers may be present in any HTTP message.

[HTTP Header] KVList cache-control

A key-value list of cache-control directives. See RFC 2616, for more details. If present, parameters to max-age, max-stale, min-fresh, and s-maxage are all parsed as non-negative integers. If present, parameters to private and no-cache are parsed as lists of header names, as symbols.

     (parse-header 'cache-control "no-cache,no-store"
     ⇒ (no-cache no-store)
     (parse-header 'cache-control "no-cache=\"Authorization,Date\",no-store"
     ⇒ ((no-cache . (authorization date)) no-store)
     (parse-header 'cache-control "no-cache=\"Authorization,Date\",max-age=10"
     ⇒ ((no-cache . (authorization date)) (max-age . 10))

[HTTP Header] List connection

A list of header names that apply only to this HTTP connection, as symbols. Additionally, the symbol ‘close’ may be present, to indicate that the server should close the connection after responding to the request.

     (parse-header 'connection "close")
     ⇒ (close)

[HTTP Header] Date date

The date that a given HTTP message was originated.

     (parse-header 'date "Tue, 15 Nov 1994 08:12:31 GMT")
     ⇒ #<date ...>
[HTTP Header] KVList pragma

A key-value list of implementation-specific directives.

     (parse-header 'pragma "no-cache, broccoli=tasty")
     ⇒ (no-cache (broccoli . "tasty"))

[HTTP Header] List trailer

A list of header names which will appear after the message body, instead of with the message headers.

     (parse-header 'trailer "ETag")
     ⇒ (etag)
[HTTP Header] List transfer-encoding

A list of transfer codings, expressed as key-value lists. The only transfer coding defined by the specification is chunked.

     (parse-header 'transfer-encoding "chunked")
     ⇒ ((chunked))

[HTTP Header] List upgrade

A list of strings, indicating additional protocols that a server could use in response to a request.

     (parse-header 'upgrade "WebSocket")
     ⇒ ("WebSocket")

FIXME: parse out more fully?

[HTTP Header] List via

A list of strings, indicating the protocol versions and hosts of intermediate servers and proxies. There may be multiple via headers in one message.

     (parse-header 'via "1.0 venus, 1.1 mars")
     ⇒ ("1.0 venus" "1.1 mars")

[HTTP Header] List warning

A list of warnings given by a server or intermediate proxy. Each warning is a itself a list of four elements: a code, as an exact integer between 0 and 1000, a host as a string, the warning text as a string, and either #f or a SRFI-19 date. There may be multiple warning headers in one message.

     (parse-header 'warning "123 foo \"core breach imminent\"")
     ⇒ ((123 "foo" "core-breach imminent" #f))

7.3.4.3 Entity Headers

Entity headers may be present in any HTTP message, and refer to the resource referenced in the HTTP request or response.

[HTTP Header] List allow

A list of allowed methods on a given resource, as symbols.

     (parse-header 'allow "GET, HEAD")
     ⇒ (GET HEAD)
[HTTP Header] List content-encoding

A list of content codings, as symbols.

     (parse-header 'content-encoding "gzip")
     ⇒ (gzip)
[HTTP Header] List content-language

The languages that a resource is in, as strings.

     (parse-header 'content-language "en")
     ⇒ ("en")
[HTTP Header] UInt content-length

The number of bytes in a resource, as an exact, non-negative integer.

     (parse-header 'content-length "300")
     ⇒ 300
[HTTP Header] URI content-location

The canonical URI for a resource, in the case that it is also accessible from a different URI.

     (parse-header 'content-location "http://example.com/foo")
     ⇒ #<<uri> ...>
[HTTP Header] String content-md5

The MD5 digest of a resource.

     (parse-header 'content-md5 "ffaea1a79810785575e29e2bd45e2fa5")
     ⇒ "ffaea1a79810785575e29e2bd45e2fa5"
[HTTP Header] List content-range

A range specification, as a list of three elements: the symbol bytes, either the symbol * or a pair of integers, indicating the byte rage, and either * or an integer, for the instance length. Used to indicate that a response only includes part of a resource.

     (parse-header 'content-range "bytes 10-20/*")
     ⇒ (bytes (10 . 20) *)
[HTTP Header] List content-type

The MIME type of a resource, as a symbol, along with any parameters.

     (parse-header 'content-length "text/plain")
     ⇒ (text/plain)
     (parse-header 'content-length "text/plain;charset=utf-8")
     ⇒ (text/plain (charset . "utf-8"))

Note that the charset parameter is something is a misnomer, and the HTTP specification admits this. It specifies the encoding of the characters, not the character set.

[HTTP Header] Date expires

The date/time after which the resource given in a response is considered stale.

     (parse-header 'expires "Tue, 15 Nov 1994 08:12:31 GMT")
     ⇒ #<date ...>
[HTTP Header] Date last-modified

The date/time on which the resource given in a response was last modified.

     (parse-header 'expires "Tue, 15 Nov 1994 08:12:31 GMT")
     ⇒ #<date ...>

7.3.4.4 Request Headers

Request headers may only appear in an HTTP request, not in a response.

[HTTP Header] List accept

A list of preferred media types for a response. Each element of the list is itself a list, in the same format as content-type.

     (parse-header 'accept "text/html,text/plain;charset=utf-8")
     ⇒ ((text/html) (text/plain (charset . "utf-8")))

Preference is expressed with quality values:

     (parse-header 'accept "text/html;q=0.8,text/plain;q=0.6")
     ⇒ ((text/html (q . 800)) (text/plain (q . 600)))
[HTTP Header] QList accept-charset

A quality list of acceptable charsets. Note again that what HTTP calls a “charset” is what Guile calls a “character encoding”.

     (parse-header 'accept-charset "iso-8859-5, unicode-1-1;q=0.8")
     ⇒ ((1000 . "iso-8859-5") (800 . "unicode-1-1"))
[HTTP Header] QList accept-encoding

A quality list of acceptable content codings.

     (parse-header 'accept-encoding "gzip,identity=0.8")
     ⇒ ((1000 . "gzip") (800 . "identity"))
[HTTP Header] QList accept-language

A quality list of acceptable languages.

     (parse-header 'accept-language "cn,en=0.75")
     ⇒ ((1000 . "cn") (750 . "en"))

[HTTP Header] Pair authorization

Authorization credentials. The car of the pair indicates the authentication scheme, like basic. For basic authentication, the cdr of the pair will be the base64-encoded ‘user:pass’ string. For other authentication schemes, like digest, the cdr will be a key-value list of credentials.

     (parse-header 'authorization "Basic QWxhZGRpbjpvcGVuIHNlc2FtZQ=="
     ⇒ (basic . "QWxhZGRpbjpvcGVuIHNlc2FtZQ==")

[HTTP Header] List expect

A list of expectations that a client has of a server. The expectations are key-value lists.

     (parse-header 'expect "100-continue")
     ⇒ ((100-continue))

[HTTP Header] String from

The email address of a user making an HTTP request.

     (parse-header 'from "bob@example.com")
     ⇒ "bob@example.com"

[HTTP Header] Pair host

The host for the resource being requested, as a hostname-port pair. If no port is given, the port is #f.

     (parse-header 'host "gnu.org:80")
     ⇒ ("gnu.org" . 80)
     (parse-header 'host "gnu.org")
     ⇒ ("gnu.org" . #f)
[HTTP Header] *|List if-match

A set of etags, indicating that the request should proceed if and only if the etag of the resource is in that set. Either the symbol *, indicating any etag, or a list of entity tags.

     (parse-header 'if-match "*")
     ⇒ *
     (parse-header 'if-match "asdfadf")
     ⇒ (("asdfadf" . #t))
     (parse-header 'if-match W/"asdfadf")
     ⇒ (("asdfadf" . #f))
[HTTP Header] Date if-modified-since

Indicates that a response should proceed if and only if the resource has been modified since the given date.

     (parse-header 'if-modified-since "Tue, 15 Nov 1994 08:12:31 GMT")
     ⇒ #<date ...>
[HTTP Header] *|List if-none-match

A set of etags, indicating that the request should proceed if and only if the etag of the resource is not in the set. Either the symbol *, indicating any etag, or a list of entity tags.

     (parse-header 'if-none-match "*")
     ⇒ *
[HTTP Header] ETag|Date if-range

Indicates that the range request should proceed if and only if the resource matches a modification date or an etag. Either an entity tag, or a SRFI-19 date.

     (parse-header 'if-range "\"original-etag\"")
     ⇒ ("original-etag" . #t)
[HTTP Header] Date if-unmodified-since

Indicates that a response should proceed if and only if the resource has not been modified since the given date.

     (parse-header 'if-not-modified-since "Tue, 15 Nov 1994 08:12:31 GMT")
     ⇒ #<date ...>
[HTTP Header] UInt max-forwards

The maximum number of proxy or gateway hops that a request should be subject to.

     (parse-header 'max-forwards "10")
     ⇒ 10
[HTTP Header] Pair proxy-authorization

Authorization credentials for a proxy connection. See the documentation for authorization above for more information on the format.

     (parse-header 'proxy-authorization "Digest foo=bar,baz=qux"
     ⇒ (digest (foo . "bar") (baz . "qux"))

[HTTP Header] Pair range

A range request, indicating that the client wants only part of a resource. The car of the pair is the symbol bytes, and the cdr is a list of pairs. Each element of the cdr indicates a range; the car is the first byte position and the cdr is the last byte position, as integers, or #f if not given.

     (parse-header 'range "bytes=10-30,50-")
     ⇒ (bytes (10 . 30) (50 . #f))

[HTTP Header] URI referer

The URI of the resource that referred the user to this resource. The name of the header is a misspelling, but we are stuck with it.

     (parse-header 'referer "http://www.gnu.org/")
     ⇒ #<uri ...>

[HTTP Header] List te

A list of transfer codings, expressed as key-value lists. A common transfer coding is

trailers.
     (parse-header 'te "trailers")
     ⇒ ((trailers))

[HTTP Header] String user-agent

A string indicating the user agent making the request. The specification defines a structured format for this header, but it is widely disregarded, so Guile does not attempt to parse strictly.

     (parse-header 'user-agent "Mozilla/5.0")
     ⇒ "Mozilla/5.0"

7.3.4.5 Response Headers

[HTTP Header] List accept-ranges

A list of range units that the server supports, as symbols.

     (parse-header 'accept-ranges "bytes")
     ⇒ (bytes)

[HTTP Header] UInt age

The age of a cached response, in seconds.

     (parse-header 'age "3600")
     ⇒ 3600

[HTTP Header] ETag etag

The entity-tag of the resource.

     (parse-header 'etag "\"foo\"")
     ⇒ ("foo" . #t)

[HTTP Header] URI location

A URI on which a request may be completed. Used in combination with a redirecting status code to perform client-side redirection.

     (parse-header 'location "http://example.com/other")
     ⇒ #<uri ...>
[HTTP Header] List proxy-authenticate

A list of challenges to a proxy, indicating the need for authentication.

     (parse-header 'proxy-authenticate "Basic realm=\"foo\"")
     ⇒ ((basic (realm . "foo")))
[HTTP Header] UInt|Date retry-after

Used in combination with a server-busy status code, like 503, to indicate that a client should retry later. Either a number of seconds, or a date.

     (parse-header 'retry-after "60")
     ⇒ 60
[HTTP Header] String server

A string identifying the server.

     (parse-header 'server "My first web server")
     ⇒ "My first web server"

[HTTP Header] *|List vary

A set of request headers that were used in computing this response. Used to indicate that server-side content negotiation was performed, for example in response to the accept-language header. Can also be the symbol *, indicating that all headers were considered.

     (parse-header 'vary "Accept-Language, Accept")
     ⇒ (accept-language accept)
[HTTP Header] List www-authenticate

A list of challenges to a user, indicating the need for authentication.

     (parse-header 'www-authenticate "Basic realm=\"foo\"")
     ⇒ ((basic (realm . "foo")))

7.3.5 Transfer Codings

HTTP 1.1 allows for various transfer codings to be applied to message bodies. These include various types of compression, and HTTP chunked encoding. Currently, only chunked encoding is supported by guile.

Chunked coding is an optional coding that may be applied to message bodies, to allow messages whose length is not known beforehand to be returned. Such messages can be split into chunks, terminated by a final zero length chunk.

In order to make dealing with encodings more simple, guile provides procedures to create ports that “wrap” existing ports, applying transformations transparently under the hood.

These procedures are in the (web http) module.

(use-modules (web http))

[Scheme Procedure] make-chunked-input-port port [#:keep-alive?=#f]

Returns a new port, that transparently reads and decodes chunk-encoded data from port. If no more chunk-encoded data is available, it returns the end-of-file object. When the port is closed, port will also be closed, unless keep-alive? is true.

(use-modules (ice-9 rdelim))
(define s "5\r\nFirst\r\nA\r\n line\n Sec\r\n8\r\nond line\r\n0\r\n")
(define p (make-chunked-input-port (open-input-string s)))
(read-line s)
⇒ "First line"
(read-line s)
⇒ "Second line"

[Scheme Procedure] make-chunked-output-port port [#:keep-alive?=#f]

Returns a new port, which transparently encodes data as chunk-encoded before writing it to port. Whenever a write occurs on this port, it buffers it, until the port is flushed, at which point it writes a chunk containing all the data written so far. When the port is closed, the data remaining is written to port, as is the terminating zero chunk. It also causes port to be closed, unless keep-alive? is true. Note. Forcing a chunked output port when there is no data is buffered does not write a zero chunk, as this would cause the data to be interpreted incorrectly by the client.

     (call-with-output-string
       (lambda (out)
         (define out* (make-chunked-output-port out #:keep-alive? #t))
         (display "first chunk" out*)
         (force-output out*)
         (force-output out*) ; note this does not write a zero chunk
         (display "second chunk" out*)
         (close-port out*)))
     ⇒ "b\r\nfirst chunk\r\nc\r\nsecond chunk\r\n0\r\n"

7.3.6 HTTP Requests

     (use-modules (web request))

The request module contains a data type for HTTP requests.

7.3.6.1 An Important Note on Character Sets

HTTP requests consist of two parts: the request proper, consisting of a request line and a set of headers, and (optionally) a body. The body might have a binary content-type, and even in the textual case its length is specified in bytes, not characters.

Therefore, HTTP is a fundamentally binary protocol. However the request line and headers are specified to be in a subset of ASCII, so they can be treated as text, provided that the port’s encoding is set to an ASCII-compatible one-byte-per-character encoding. ISO-8859-1 (latin-1) is just such an encoding, and happens to be very efficient for Guile.

So what Guile does when reading requests from the wire, or writing them out, is to set the port’s encoding to latin-1, and treating the request headers as text.

The request body is another issue. For binary data, the data is probably in a bytevector, so we use the R6RS binary output procedures to write out the binary payload. Textual data usually has to be written out to some character encoding, usually UTF-8, and then the resulting bytevector is written out to the port.

In summary, Guile reads and writes HTTP over latin-1 sockets, without any loss of generality.

7.3.6.2 Request API

[Scheme Procedure] request? obj [Scheme Procedure] request-method request [Scheme Procedure] request-uri request [Scheme Procedure] request-version request [Scheme Procedure] request-headers request [Scheme Procedure] request-meta request [Scheme Procedure] request-port request

A predicate and field accessors for the request type. The fields are as follows:

method The HTTP method, for example, GET.

uri The URI as a URI record.

version The HTTP version pair, like (1 . 1). headers The request headers, as an alist of parsed values.

meta An arbitrary alist of other data, for example information returned in the sockaddr from accept (see Section 7.2.11.4 [Network Sockets and Communication], page 511).

port The port on which to read or write a request body, if any.

[Scheme Procedure] read-request port [meta=’()]

Read an HTTP request from port, optionally attaching the given metadata, meta. As a side effect, sets the encoding on port to ISO-8859-1 (latin-1), so that reading one character reads one byte. See the discussion of character sets above, for more information. Note that the body is not part of the request. Once you have read a request, you may read the body separately, and likewise for writing requests.

[Scheme Procedure] build-request uri [#:method=’GET] [#:version=’(1 . 1)]

[#:headers=’()] [#:port=#f] [#:meta=’()] [#:validate-headers?=#t] Construct an HTTP request object. If validate-headers? is true, the headers are each run through their respective validators.

[Scheme Procedure] write-request r port

Write the given HTTP request to port. Return a new request, whose request-port will continue writing on port, perhaps using some transfer encoding.

[Scheme Procedure] read-request-body r

Reads the request body from r, as a bytevector. Return #f if there was no request body.

[Scheme Procedure] write-request-body r bv

Write bv, a bytevector, to the port corresponding to the HTTP request r.

The various headers that are typically associated with HTTP requests may be accessed with these dedicated accessors. See Section 7.3.4 [HTTP Headers], page 525, for more information on the format of parsed headers.

[Scheme Procedure] request-accept request [default=’()]

[Scheme Procedure] request-accept-charset request [default=’()] [Scheme Procedure] request-accept-encoding request [default=’()] [Scheme Procedure] request-accept-language request [default=’()]

[Scheme Procedure] request-allow request [default=’()]

[Scheme Procedure] request-authorization request [default=#f] [Scheme Procedure] request-cache-control request [default=’()] [Scheme Procedure] request-connection request [default=’()] [Scheme Procedure] request-content-encoding request [default=’()] [Scheme Procedure] request-content-language request [default=’()] [Scheme Procedure] request-content-length request [default=#f] [Scheme Procedure] request-content-location request [default=#f] [Scheme Procedure] request-content-md5 request [default=#f] [Scheme Procedure] request-content-range request [default=#f] [Scheme Procedure] request-content-type request [default=#f]

[Scheme Procedure] request-date request [default=#f] [Scheme Procedure] request-expect request [default=’()]

[Scheme Procedure] request-expires request [default=#f]

[Scheme Procedure] request-from request [default=#f] [Scheme Procedure] request-host request [default=#f]

[Scheme Procedure] request-if-match request [default=#f] [Scheme Procedure] request-if-modified-since request [default=#f] [Scheme Procedure] request-if-none-match request [default=#f] [Scheme Procedure] request-if-range request [default=#f] [Scheme Procedure] request-if-unmodified-since request [default=#f] [Scheme Procedure] request-last-modified request [default=#f] [Scheme Procedure] request-max-forwards request [default=#f]

[Scheme Procedure] request-pragma request [default=’()]

[Scheme Procedure] request-proxy-authorization request [default=#f]

[Scheme Procedure] request-range request [default=#f]

[Scheme Procedure] request-referer request [default=#f]

[Scheme Procedure] request-te request [default=#f] [Scheme Procedure] request-trailer request [default=’()]

[Scheme Procedure] request-transfer-encoding request [default=’()]

[Scheme Procedure] request-upgrade request [default=’()]

[Scheme Procedure] request-user-agent request [default=#f]

[Scheme Procedure] request-via request [default=’()] [Scheme Procedure] request-warning request [default=’()]

Return the given request header, or default if none was present.

[Scheme Procedure] request-absolute-uri r [default-host=#f] [default-port=#f] A helper routine to determine the absolute URI of a request, using the host header and the default host and port.

7.3.7 HTTP Responses

     (use-modules (web response))

As with requests (see Section 7.3.6 [Requests], page 534), Guile offers a data type for HTTP responses. Again, the body is represented separately from the request.

[Scheme Procedure] response? obj [Scheme Procedure] response-version response [Scheme Procedure] response-code response [Scheme Procedure] response-reason-phrase response [Scheme Procedure] response-headers response [Scheme Procedure] response-port response

A predicate and field accessors for the response type. The fields are as follows:

version The HTTP version pair, like (1 . 1).

code The HTTP response code, like 200.

     reason-phrase

The reason phrase, or the standard reason phrase for the response’s code. headers The response headers, as an alist of parsed values.

port The port on which to read or write a response body, if any.

[Scheme Procedure] read-response port

Read an HTTP response from port. As a side effect, sets the encoding on port to ISO-8859-1 (latin-1), so that reading one character reads one byte. See the discussion of character sets in Section 7.3.7 [Responses], page 536, for more information.

[Scheme Procedure] build-response [#:version=’(1 . 1)] [#:code=200]

[#:reason-phrase=#f] [#:headers=’()] [#:port=#f] [#:validate-headers?=#t] Construct an HTTP response object. If validate-headers? is true, the headers are each run through their respective validators.

[Scheme Procedure] adapt-response-version response version

Adapt the given response to a different HTTP version. Return a new HTTP response. The idea is that many applications might just build a response for the default HTTP version, and this method could handle a number of programmatic transformations to respond to older HTTP versions (0.9 and 1.0). But currently this function is a bit heavy-handed, just updating the version field.

[Scheme Procedure] write-response r port

Write the given HTTP response to port. Return a new response, whose response-port will continue writing on port, perhaps using some transfer encoding.

[Scheme Procedure] response-must-not-include-body? r

Some responses, like those with status code 304, are specified as never having bodies. This predicate returns #t for those responses. Note also, though, that responses to HEAD requests must also not have a body.

[Scheme Procedure] response-body-port r [#:decode?=#t] [#:keep-alive?=#t] Return an input port from which the body of r can be read. The encoding of the returned port is set according to r’s content-type header, when it’s textual, except if decode? is #f. Return #f when no body is available. When keep-alive? is #f, closing the returned port also closes r’s response port.

[Scheme Procedure] read-response-body r

Read the response body from r, as a bytevector. Returns #f if there was no response body.

[Scheme Procedure] write-response-body r bv

Write bv, a bytevector, to the port corresponding to the HTTP response r.

As with requests, the various headers that are typically associated with HTTP responses may be accessed with these dedicated accessors. See Section 7.3.4 [HTTP Headers], page 525, for more information on the format of parsed headers.

[Scheme Procedure] response-accept-ranges response [default=#f]

[Scheme Procedure] response-age response [default=’()] [Scheme Procedure] response-allow response [default=’()]

[Scheme Procedure] response-cache-control response [default=’()] [Scheme Procedure] response-connection response [default=’()] [Scheme Procedure] response-content-encoding response [default=’()] [Scheme Procedure] response-content-language response [default=’()] [Scheme Procedure] response-content-length response [default=#f] [Scheme Procedure] response-content-location response [default=#f] [Scheme Procedure] response-content-md5 response [default=#f] [Scheme Procedure] response-content-range response [default=#f] [Scheme Procedure] response-content-type response [default=#f]

[Scheme Procedure] response-date response [default=#f] [Scheme Procedure] response-etag response [default=#f]

[Scheme Procedure] response-expires response [default=#f] [Scheme Procedure] response-last-modified response [default=#f] [Scheme Procedure] response-location response [default=#f]

[Scheme Procedure] response-pragma response [default=’()]

[Scheme Procedure] response-proxy-authenticate response [default=#f] [Scheme Procedure] response-retry-after response [default=#f]

[Scheme Procedure] response-server response [default=#f] [Scheme Procedure] response-trailer response [default=’()]

[Scheme Procedure] response-transfer-encoding response [default=’()]

[Scheme Procedure] response-upgrade response [default=’()] [Scheme Procedure] response-vary response [default=’()] [Scheme Procedure] response-via response [default=’()] [Scheme Procedure] response-warning response [default=’()]

[Scheme Procedure] response-www-authenticate response [default=#f]

Return the given response header, or default if none was present.

[Scheme Procedure] text-content-type? type

Return #t if type, a symbol as returned by response-content-type, represents a textual type such as text/plain.

7.3.8 Web Client

(web client) provides a simple, synchronous HTTP client, built on the lower-level HTTP, request, and response modules.

     (use-modules (web client))

[Scheme Procedure] open-socket-for-uri uri

Return an open input/output port for a connection to URI.

[Scheme Procedure] http-get uri arg...

[Scheme Procedure] http-head uri arg... [Scheme Procedure] http-post uri arg...

[Scheme Procedure] http-put uri arg...

[Scheme Procedure] http-delete uri arg... [Scheme Procedure] http-trace uri arg...

[Scheme Procedure] http-options uri arg...

Connect to the server corresponding to uri and make a request over HTTP, using the appropriate method (GET, HEAD, etc.). All of these procedures have the same prototype: a URI followed by an optional sequence of keyword arguments. These keyword arguments allow you to modify the requests in various ways, for example attaching a body to the request, or setting specific headers. The following table lists the keyword arguments and their default values.

#:body #f
#:port (open-socket-for-uri uri)]
#:version '(1 . 1)
#:keep-alive? #f
#:headers '()
#:decode-body? #t
#:streaming? #f

If you already have a port open, pass it as port. Otherwise, a connection will be opened to the server corresponding to uri. Any extra headers in the alist headers will be added to the request. If body is not #f, a message body will also be sent with the HTTP request. If body is a string, it is encoded according to the content-type in headers, defaulting to UTF-8. Otherwise body should be a bytevector, or #f for no body. Although a message body may be sent with any request, usually only POST and PUT requests have bodies. If decode-body? is true, as is the default, the body of the response will be decoded to string, if it is a textual content-type. Otherwise it will be returned as a bytevector. However, if streaming? is true, instead of eagerly reading the response body from the server, this function only reads off the headers. The response body will be returned as a port on which the data may be read. Unless keep-alive? is true, the port will be closed after the full response body has been read. Returns two values: the response read from the server, and the response body as a string, bytevector, #f value, or as a port (if streaming? is true).

http-get is useful for making one-off requests to web sites. If you are writing a web spider or some other client that needs to handle a number of requests in parallel, it’s better to build an event-driven URL fetcher, similar in structure to the web server (see Section 7.3.9 [Web Server], page 540).

Another option, good but not as performant, would be to use threads, possibly via par-map or futures.

[Scheme Parameter] current-http-proxy

Either #f or a non-empty string containing the URL of the HTTP proxy server to be used by the procedures in the (web client) module, including open-socket-for-uri. Its initial value is based on the http_proxy environment variable.

     (current-http-proxy) ⇒ "http://localhost:8123/"
     (parameterize ((current-http-proxy #f))

        (http-get "http://example.com/"))  ; temporarily bypass proxy
      (current-http-proxy) ⇒ "http://localhost:8123/"

7.3.9 Web Server

(web server) is a generic web server interface, along with a main loop implementation for web servers controlled by Guile.

 (use-modules (web server))

The lowest layer is the <server-impl> object, which defines a set of hooks to open a server, read a request from a client, write a response to a client, and close a server. These hooks – open, read, write, and close, respectively – are bound together in a <serverimpl> object. Procedures in this module take a <server-impl> object, if needed.

A <server-impl> may also be looked up by name. If you pass the http symbol to runserver, Guile looks for a variable named http in the (web server http) module, which should be bound to a <server-impl> object. Such a binding is made by instantiation of the define-server-impl syntax. In this way the run-server loop can automatically load other backends if available.

The life cycle of a server goes as follows:

1. The open hook is called, to open the server. open takes zero or more arguments, depending on the backend, and returns an opaque server socket object, or signals an error.

2. The read hook is called, to read a request from a new client. The read hook takes one argument, the server socket. It should return three values: an opaque client socket, the request, and the request body. The request should be a <request> object, from (web request). The body should be a string or a bytevector, or #f if there is no body. If the read failed, the read hook may return #f for the client socket, request, and body.

3. A user-provided handler procedure is called, with the request and body as its arguments. The handler should return two values: the response, as a <response> record from (web response), and the response body as bytevector, or #f if not present. The respose and response body are run through sanitize-response, documented below. This allows the handler writer to take some convenient shortcuts: for example, instead of a <response>, the handler can simply return an alist of headers, in which case a default response object is constructed with those headers. Instead of a bytevector for the body, the handler can return a string, which will be serialized into an appropriate encoding; or it can return a procedure, which will be called on a port to write out the data. See the sanitize-response documentation, for more.

4. The write hook is called with three arguments: the client socket, the response, and the body. The write hook returns no values.

5. At this point the request handling is complete. For a loop, we loop back and try to read a new request.

6. If the user interrupts the loop, the close hook is called on the server socket.

A user may define a server implementation with the following form:

[Scheme Syntax] define-server-impl name open read write close Make a <server-impl> object with the hooks open, read, write, and close, and bind it to the symbol name in the current module.

[Scheme Procedure] lookup-server-impl impl

Look up a server implementation. If impl is a server implementation already, it is returned directly. If it is a symbol, the binding named impl in the (web server impl) module is looked up. Otherwise an error is signaled.

Currently a server implementation is a somewhat opaque type, useful only for passing to other procedures in this module, like read-client.

The (web server) module defines a number of routines that use <server-impl> objects to implement parts of a web server. Given that we don’t expose the accessors for the various fields of a <server-impl>, indeed these routines are the only procedures with any access to the impl objects.

[Scheme Procedure] open-server impl open-params Open a server for the given implementation. Return one value, the new server object. The implementation’s open procedure is applied to open-params, which should be a list.

[Scheme Procedure] read-client impl server

Read a new client from server, by applying the implementation’s read procedure to the server. If successful, return three values: an object corresponding to the client, a request object, and the request body. If any exception occurs, return #f for all three values.

[Scheme Procedure] handle-request handler request body state Handle a given request, returning the response and body.

The response and response body are produced by calling the given handler with request and body as arguments.

The elements of state are also passed to handler as arguments, and may be returned as additional values. The new state, collected from the handler’s return values, is then returned as a list. The idea is that a server loop receives a handler from the user, along with whatever state values the user is interested in, allowing the user’s handler to explicitly manage its state.

[Scheme Procedure] sanitize-response request response body “Sanitize” the given response and body, making them appropriate for the given request.

As a convenience to web handler authors, response may be given as an alist of headers, in which case it is used to construct a default response. Ensures that the response version corresponds to the request version. If body is a string, encodes the string to a bytevector, in an encoding appropriate for response. Adds a content-length and content-type header, as necessary.

If body is a procedure, it is called with a port as an argument, and the output collected as a bytevector. In the future we might try to instead use a compressing, chunkencoded port, and call this procedure later, in the write-client procedure. Authors are advised not to rely on the procedure being called at any particular time.

[Scheme Procedure] write-client impl server client response body Write an HTTP response and body to client. If the server and client support persistent connections, it is the implementation’s responsibility to keep track of the client thereafter, presumably by attaching it to the server argument somehow.

[Scheme Procedure] close-server impl server

Release resources allocated by a previous invocation of open-server.

Given the procedures above, it is a small matter to make a web server:

[Scheme Procedure] serve-one-client handler impl server state Read one request from server, call handler on the request and body, and write the response to the client. Return the new state produced by the handler procedure.

[Scheme Procedure] run-server handler [impl=’http] [open-params=’()] arg … Run Guile’s built-in web server.

handler should be a procedure that takes two or more arguments, the HTTP request and request body, and returns two or more values, the response and response body.

For examples, skip ahead to the next section, Section 7.3.10 [Web Examples], page 542.

The response and body will be run through sanitize-response before sending back to the client.

Additional arguments to handler are taken from arg . . .. These arguments comprise a state. Additional return values are accumulated into a new state, which will be used for subsequent requests. In this way a handler can explicitly manage its state.

The default web server implementation is http, which binds to a socket, listening for request on that port.

[HTTP Implementation] http [#:host=#f] [#:family=AF INET]

[#:addr=INADDR LOOPBACK] [#:port 8080] [#:socket] The default HTTP implementation. We document it as a function with keyword arguments, because that is precisely the way that it is – all of the open-params to run-server get passed to the implementation’s open function.

     ;; The defaults: localhost:8080
     (run-server handler)
     ;; Same thing
     (run-server handler 'http '())
     ;; On a different port
     (run-server handler 'http '(#:port 8081))
     ;; IPv6
     (run-server handler 'http '(#:family AF_INET6 #:port 8081))
     ;; Custom socket
     (run-server handler 'http `(#:socket ,(sudo-make-me-a-socket)))

7.3.10 Web Examples

Well, enough about the tedious internals. Let’s make a web application!

7.3.10.1 Hello, World!

The first program we have to write, of course, is “Hello, World!”. This means that we have to implement a web handler that does what we want.

Now we define a handler, a function of two arguments and two return values:

     (define (handler request request-body)
       (values response response-body))

In this first example, we take advantage of a short-cut, returning an alist of headers instead of a proper response object. The response body is our payload:

     (define (hello-world-handler request request-body)
       (values '((content-type . (text/plain)))
               "Hello World!"))

Now let’s test it, by running a server with this handler. Load up the web server module if you haven’t yet done so, and run a server with this handler:

     (use-modules (web server))
     (run-server hello-world-handler)

By default, the web server listens for requests on localhost:8080. Visit that address in your web browser to test. If you see the string, Hello World!, sweet!

7.3.10.2 Inspecting the Request

The Hello World program above is a general greeter, responding to all URIs. To make a more exclusive greeter, we need to inspect the request object, and conditionally produce different results. So let’s load up the request, response, and URI modules, and do just that.

     (use-modules (web server)) ; you probably did this already
     (use-modules (web request)
                  (web response)
                  (web uri))
     (define (request-path-components request)
       (split-and-decode-uri-path (uri-path (request-uri request))))
     (define (hello-hacker-handler request body)
       (if (equal? (request-path-components request)
                   '("hacker"))
           (values '((content-type . (text/plain)))
                   "Hello hacker!")
           (not-found request)))
     (run-server hello-hacker-handler)

Here we see that we have defined a helper to return the components of the URI path as a list of strings, and used that to check for a request to /hacker/. Then the success case is just as before – visit http://localhost:8080/hacker/ in your browser to check.

You should always match against URI path components as decoded by split-and-

decode-uri-path. The above example will work for /hacker/, //hacker///, and
/h%61ck%65r.

But we forgot to define not-found! If you are pasting these examples into a REPL, accessing any other URI in your web browser will drop your Guile console into the debugger:

<unnamed port>:38:7: In procedure module-lookup:
<unnamed port>:38:7: Unbound variable: not-found
Entering a new prompt.  Type `,bt' for a backtrace or `,q' to continue.
scheme@(guile-user) [1]>

So let’s define the function, right there in the debugger. As you probably know, we’ll want to return a 404 response.

;; Paste this in your REPL
(define (not-found request)
  (values (build-response #:code 404)
          (string-append "Resource not found: "
                         (uri->string (request-uri request)))))
;; Now paste this to let the web server keep going:
,continue

Now if you access http://localhost/foo/, you get this error message. (Note that some popular web browsers won’t show server-generated 404 messages, showing their own instead, unless the 404 message body is long enough.)

7.3.10.3 Higher-Level Interfaces

The web handler interface is a common baseline that all kinds of Guile web applications can use. You will usually want to build something on top of it, however, especially when producing HTML. Here is a simple example that builds up HTML output using SXML (see Section 7.22 [SXML], page 697).

First, load up the modules:

(use-modules (web server)
             (web request)
             (web response)
             (sxml simple))

Now we define a simple templating function that takes a list of HTML body elements, as SXML, and puts them in our super template:

(define (templatize title body)
  `(html (head (title ,title))
         (body ,@body)))

For example, the simplest Hello HTML can be produced like this:

(sxml->xml (templatize "Hello!" '((b "Hi!"))))

⊣

<html><head><title>Hello!</title></head><body><b>Hi!</b></body></html>

Much better to work with Scheme data types than to work with HTML as strings. Now we define a little response helper:

(define* (respond #:optional body #:key
                  (status 200)

                       (title "Hello hello!")
                       (doctype "<!DOCTYPE html>\n")
                       (content-type-params '((charset . "utf-8")))
                       (content-type 'text/html)
                       (extra-headers '())
                       (sxml (and body (templatize title body))))
       (values (build-response
                #:code status
                #:headers `((content-type
                             . (,content-type ,@content-type-params))
                            ,@extra-headers))
               (lambda (port)
                 (if sxml
                     (begin
                       (if doctype (display doctype port))
                       (sxml->xml sxml port))))))

Here we see the power of keyword arguments with default initializers. By the time the arguments are fully parsed, the sxml local variable will hold the templated SXML, ready for sending out to the client.

Also, instead of returning the body as a string, respond gives a procedure, which will be called by the web server to write out the response to the client.

Now, a simple example using this responder, which lays out the incoming headers in an HTML table.

     (define (debug-page request body)
       (respond
        `((h1 "hello world!")
          (table
           (tr (th "header") (th "value"))
           ,@(map (lambda (pair)
                    `(tr (td (tt ,(with-output-to-string
                                    (lambda () (display (car pair))))))
                         (td (tt ,(with-output-to-string
                                    (lambda ()
                                      (write (cdr pair))))))))
                  (request-headers request))))))
     (run-server debug-page)

Now if you visit any local address in your web browser, we actually see some HTML, finally.

7.3.10.4 Conclusion

Well, this is about as far as Guile’s built-in web support goes, for now. There are many ways to make a web application, but hopefully by standardizing the most fundamental data types, users will be able to choose the approach that suits them best, while also being able to switch between implementations of the server. This is a relatively new part of Guile, so if you have feedback, let us know, and we can take it into account. Happy hacking on the web!

7.4 The (ice-9 getopt-long) Module

The (ice-9 getopt-long) module exports two procedures: getopt-long and option-ref.

• getopt-long takes a list of strings — the command line arguments — an option specification, and some optional keyword parameters. It parses the command line arguments according to the option specification and keyword parameters, and returns a data structure that encapsulates the results of the parsing.

• option-ref then takes the parsed data structure and a specific option’s name, and returns information about that option in particular.

To make these procedures available to your Guile script, include the expression (usemodules (ice-9 getopt-long)) somewhere near the top, before the first usage of getopt-

long or option-ref.

7.4.1 A Short getopt-long Example

This section illustrates how getopt-long is used by presenting and dissecting a simple example. The first thing that we need is an option specification that tells getopt-long how to parse the command line. This specification is an association list with the long option name as the key. Here is how such a specification might look:

     (define option-spec
       '((version (single-char #\v) (value #f))
         (help    (single-char #\h) (value #f))))

This alist tells getopt-long that it should accept two long options, called version and help, and that these options can also be selected by the single-letter abbreviations v and h, respectively. The (value #f) clauses indicate that neither of the options accepts a value.

With this specification we can use getopt-long to parse a given command line:

     (define options (getopt-long (command-line) option-spec))

After this call, options contains the parsed command line and is ready to be examined

by option-ref. option-ref is called like this:
     (option-ref options 'help #f)

It expects the parsed command line, a symbol indicating the option to examine, and a default value. The default value is returned if the option was not present in the command line, or if the option was present but without a value; otherwise the value from the command line is returned. Usually option-ref is called once for each possible option that a script supports.

The following example shows a main program which puts all this together to parse its command line and figure out what the user wanted.

     (define (main args)
       (let* ((option-spec '((version (single-char #\v) (value #f))
                             (help    (single-char #\h) (value #f))))
              (options (getopt-long args option-spec))
              (help-wanted (option-ref options 'help #f))

              (version-wanted (option-ref options 'version #f)))
         (if (or version-wanted help-wanted)
             (begin
               (if version-wanted
                   (display "getopt-long-example version 0.3\n"))
               (if help-wanted
                   (display "\
     getopt-long-example [options]
       -v, --version    Display version
       -h, --help       Display this help
     ")))
             (begin
               (display "Hello, World!") (newline)))))

7.4.2 How to Write an Option Specification

An option specification is an association list (see Section 6.7.12 [Association Lists], page 228) with one list element for each supported option. The key of each list element is a symbol that names the option, while the value is a list of option properties:

     OPTION-SPEC ::=  '( (OPT-NAME1 (PROP-NAME PROP-VALUE) ...)
                         (OPT-NAME2 (PROP-NAME PROP-VALUE) ...)
                         (OPT-NAME3 (PROP-NAME PROP-VALUE) ...)
                         ...
                       )

Each opt-name specifies the long option name for that option. For example, a list element with opt-name background specifies an option that can be specified on the command line using the long option --background. Further information about the option — whether it takes a value, whether it is required to be present in the command line, and so on — is specified by the option properties.

In the example of the preceding section, we already saw that a long option name can have a equivalent short option character. The equivalent short option character can be set for an option by specifying a single-char property in that option’s property list. For example, a list element like '(output (single-char #\o) ...) specifies an option with long name --output that can also be specified by the equivalent short name -o.

The value property specifies whether an option requires or accepts a value. If the value property is set to #t, the option requires a value: getopt-long will signal an error if the option name is present without a corresponding value. If set to #f, the option does not take a value; in this case, a non-option word that follows the option name in the command line will be treated as a non-option argument. If set to the symbol optional, the option accepts a value but does not require one: a non-option word that follows the option name in the command line will be interpreted as that option’s value. If the option name for an option with '(value optional) is immediately followed in the command line by another option name, the value for the first option is implicitly #t.

The required? property indicates whether an option is required to be present in the command line. If the required? property is set to #t, getopt-long will signal an error if the option is not specified.

Finally, the predicate property can be used to constrain the possible values of an option. If used, the predicate property should be set to a procedure that takes one argument — the proposed option value as a string — and returns either #t or #f according as the proposed value is or is not acceptable. If the predicate procedure returns #f, getopt-long will signal an error.

By default, options do not have single-character equivalents, are not required, and do not take values. Where the list element for an option includes a value property but no predicate property, the option values are unconstrained.

7.4.3 Expected Command Line Format

In order for getopt-long to correctly parse a command line, that command line must conform to a standard set of rules for how command line options are specified. This section explains what those rules are.

getopt-long splits a given command line into several pieces. All elements of the argument list are classified to be either options or normal arguments. Options consist of two dashes and an option name (so-called long options), or of one dash followed by a single letter (short options).

Options can behave as switches, when they are given without a value, or they can be used to pass a value to the program. The value for an option may be specified using an equals sign, or else is simply the next word in the command line, so the following two invocations are equivalent:

     $ ./foo.scm --output=bar.txt
     $ ./foo.scm --output bar.txt

Short options can be used instead of their long equivalents and can be grouped together after a single dash. For example, the following commands are equivalent.

     $ ./foo.scm --version --help
     $ ./foo.scm -v --help
     $ ./foo.scm -vh

If an option requires a value, it can only be grouped together with other short options if it is the last option in the group; the value is the next argument. So, for example, with the following option specification —

     ((apples    (single-char #\a))
      (blimps    (single-char #\b) (value #t))
      (catalexis (single-char #\c) (value #t)))

— the following command lines would all be acceptable:

     $ ./foo.scm -a -b bang -c couth
     $ ./foo.scm -ab bang -c couth
     $ ./foo.scm -ac couth -b bang

But the next command line is an error, because -b is not the last option in its combination, and because a group of short options cannot include two options that both require values:

     $ ./foo.scm -abc couth bang

If an option’s value is optional, getopt-long decides whether the option has a value by looking at what follows it in the argument list. If the next element is a string, and it does not appear to be an option itself, then that string is the option’s value.

If the option -- appears in the argument list, argument parsing stops there and subsequent arguments are returned as ordinary arguments, even if they resemble options. So, with the command line

$ ./foo.scm --apples "Granny Smith" -- --blimp Goodyear

getopt-long will recognize the --apples option as having the value "Granny Smith", but will not treat --blimp as an option. The strings --blimp and Goodyear will be returned as ordinary argument strings.

7.4.4 Reference Documentation for getopt-long

[Scheme Procedure] getopt-long args grammar [#:stop-at-first-non-option #t] Parse the command line given in args (which must be a list of strings) according to the option specification grammar. The grammar argument is expected to be a list of this form:

((option (property value) ...) ...)

where each option is a symbol denoting the long option, but without the two leading dashes (e.g. version if the option is called --version). For each option, there may be list of arbitrarily many property/value pairs. The order of the pairs is not important, but every property may only appear once in the property list. The following table lists the possible properties:

(single-char char)

Accept -char as a single-character equivalent to --option. This is how to specify traditional Unix-style flags.

(required? bool)

If bool is true, the option is required. getopt-long will raise an error if it is not found in args.

(value bool)

If bool is #t, the option accepts a value; if it is #f, it does not; and if it is the symbol optional, the option may appear in args with or without a value.

(predicate func)

If the option accepts a value (i.e. you specified (value #t) for this option), then getopt-long will apply func to the value, and throw an exception if it returns #f. func should be a procedure which accepts a string and returns a boolean value; you may need to use quasiquotes to get it into grammar.

The #:stop-at-first-non-option keyword, if specified with any true value, tells getopt-long to stop when it gets to the first non-option in the command line. That is, at the first word which is neither an option itself, nor the value of an option. Everything in the command line from that word onwards will be returned as nonoption arguments.

getopt-long’s args parameter is expected to be a list of strings like the one returned by command-line, with the first element being the name of the command. Therefore getoptlong ignores the first element in args and starts argument interpretation with the second element.

getopt-long signals an error if any of the following conditions hold.

• The option grammar has an invalid syntax.

• One of the options in the argument list was not specified by the grammar.

• A required option is omitted.

• An option which requires an argument did not get one.

• An option that doesn’t accept an argument does get one (this can only happen using the long option --opt=value syntax).

• An option predicate fails.

#:stop-at-first-non-option is useful for command line invocations like guild [-

-help | --version] [script [script-options]] and cvs [general-options] command

[command-options], where there are options at two levels: some generic and understood by the outer command, and some that are specific to the particular script or command being invoked. To use getopt-long in such cases, you would call it twice: firstly with #:stopat-first-non-option #t, so as to parse any generic options and identify the wanted script or sub-command; secondly, and after trimming off the initial generic command words, with a script- or sub-command-specific option grammar, so as to process those specific options.

7.4.5 Reference Documentation for option-ref

[Scheme Procedure] option-ref options key default Search options for a command line option named key and return its value, if found. If the option has no value, but was given, return #t. If the option was not given, return default. options must be the result of a call to getopt-long.

option-ref always succeeds, either by returning the requested option value from the command line, or the default value.

The special key '() can be used to get a list of all non-option arguments.

7.5 SRFI Support Modules

SRFI is an acronym for Scheme Request For Implementation. The SRFI documents define a lot of syntactic and procedure extensions to standard Scheme as defined in R5RS.

Guile has support for a number of SRFIs. This chapter gives an overview over the available SRFIs and some usage hints. For complete documentation, design rationales and further examples, we advise you to get the relevant SRFI documents from the SRFI home

page http://srfi.schemers.org/.

7.5.1 About SRFI Usage

SRFI support in Guile is currently implemented partly in the core library, and partly as addon modules. That means that some SRFIs are automatically available when the interpreter is started, whereas the other SRFIs require you to use the appropriate support module explicitly.

There are several reasons for this inconsistency. First, the feature checking syntactic form cond-expand (see Section 7.5.2 [SRFI-0], page 551) must be available immediately, because it must be there when the user wants to check for the Scheme implementation, that is, before she can know that it is safe to use use-modules to load SRFI support modules. The second reason is that some features defined in SRFIs had been implemented in Guile before the developers started to add SRFI implementations as modules (for example SRFI-13 (see Section 7.5.11 [SRFI-13], page 578)). In the future, it is possible that SRFIs in the core library might be factored out into separate modules, requiring explicit module loading when they are needed. So you should be prepared to have to use use-modules someday in the future to access SRFI-13 bindings. If you want, you can do that already. We have included the module (srfi srfi-13) in the distribution, which currently does nothing, but ensures that you can write future-safe code.

Generally, support for a specific SRFI is made available by using modules named (srfi srfi-number), where number is the number of the SRFI needed. Another possibility is to use the command line option --use-srfi, which will load the necessary modules automatically (see Section 4.2 [Invoking Guile], page 35).

7.5.2 SRFI-0 - cond-expand

This SRFI lets a portable Scheme program test for the presence of certain features, and adapt itself by using different blocks of code, or fail if the necessary features are not available. There’s no module to load, this is in the Guile core.

A program designed only for Guile will generally not need this mechanism, such a program can of course directly use the various documented parts of Guile.

[syntax] cond-expand (feature body…) … Expand to the body of the first clause whose feature specification is satisfied. It is an error if no feature is satisfied. Features are symbols such as srfi-1, and a feature specification can use and, or and not forms to test combinations. The last clause can be an else, to be used if no other passes. For example, define a private version of alist-cons if SRFI-1 is not available.

          (cond-expand (srfi-1
                        )
                       (else
                        (define (alist-cons key val alist)
                          (cons (cons key val) alist))))

Or demand a certain set of SRFIs (list operations, string ports, receive and string operations), failing if they’re not available.

          (cond-expand ((and srfi-1 srfi-6 srfi-8 srfi-13)
                        ))

The Guile core has the following features,

     guile
     guile-2  ;; starting from Guile 2.x
     r5rs
     srfi-0
     srfi-4
     srfi-13

     srfi-14
     srfi-16
     srfi-23
     srfi-30
     srfi-39
     srfi-46
     srfi-55
     srfi-61
     srfi-62
     srfi-87
     srfi-105

Other SRFI feature symbols are defined once their code has been loaded with usemodules, since only then are their bindings available.

The ‘--use-srfi’ command line option (see Section 4.2 [Invoking Guile], page 35) is a good way to load SRFIs to satisfy cond-expand when running a portable program.

Testing the guile feature allows a program to adapt itself to the Guile module system, but still run on other Scheme systems. For example the following demands SRFI-8 (receive), but also knows how to load it with the Guile mechanism.

     (cond-expand (srfi-8
                   )
                  (guile
                   (use-modules (srfi srfi-8))))

Likewise, testing the guile-2 feature allows code to be portable between Guile 2.0 and previous versions of Guile. For instance, it makes it possible to write code that accounts for Guile 2.0’s compiler, yet be correctly interpreted on 1.8 and earlier versions:

     (cond-expand (guile-2 (eval-when (compile)
                             ;; This must be evaluated at compile time.
                             (fluid-set! current-reader my-reader)))
                  (guile
                           ;; Earlier versions of Guile do not have a
                           ;; separate compilation phase.
                           (fluid-set! current-reader my-reader)))

It should be noted that cond-expand is separate from the *features* mechanism (see Section 6.22.2 [Feature Tracking], page 429), feature symbols in one are unrelated to those in the other.

7.5.3 SRFI-1 - List library

The list library defined in SRFI-1 contains a lot of useful list processing procedures for construction, examining, destructuring and manipulating lists and pairs.

Since SRFI-1 also defines some procedures which are already contained in R5RS and thus are supported by the Guile core library, some list and pair procedures which appear in the SRFI-1 document may not appear in this section. So when looking for a particular list/pair processing procedure, you should also have a look at the sections Section 6.7.2 [Lists], page 188 and Section 6.7.1 [Pairs], page 185.

7.5.3.1 Constructors

New lists can be constructed by calling one of the following procedures.

[Scheme Procedure] xcons d a

Like cons, but with interchanged arguments. Useful mostly when passed to higherorder procedures.

[Scheme Procedure] list-tabulate n init-proc

Return an n-element list, where each list element is produced by applying the procedure init-proc to the corresponding list index. The order in which init-proc is applied to the indices is not specified.

[Scheme Procedure] list-copy lst

Return a new list containing the elements of the list lst. This function differs from the core list-copy (see Section 6.7.2.3 [List Constructors], page 189) in accepting improper lists too. And if lst is not a pair at all then it’s treated as the final tail of an improper list and simply returned.

[Scheme Procedure] circular-list elt1 elt2 …

Return a circular list containing the given arguments elt1 elt2 . . ..

[Scheme Procedure] iota count [start step] Return a list containing count numbers, starting from start and adding step each time. The default start is 0, the default step is 1. For example,

     (iota 6)  ⇒ (0 1 2 3 4 5)

     (iota 4 2.5 -2) ⇒ (2.5 0.5 -1.5 -3.5)

This function takes its name from the corresponding primitive in the APL language.

7.5.3.2 Predicates

The procedures in this section test specific properties of lists.

[Scheme Procedure] proper-list? obj

Return #t if obj is a proper list, or #f otherwise. This is the same as the core list? (see Section 6.7.2.2 [List Predicates], page 188). A proper list is a list which ends with the empty list () in the usual way. The empty list () itself is a proper list too.

     (proper-list? '(1 2 3)) ⇒ #t
     (proper-list? '())  ⇒ #t

[Scheme Procedure] circular-list? obj

Return #t if obj is a circular list, or #f otherwise. A circular list is a list where at some point the cdr refers back to a previous pair in the list (either the start or some later point), so that following the cdrs takes you around in a circle, with no end.

     (define x (list 1 2 3 4))
     (set-cdr! (last-pair x) (cddr x))
     x ⇒ (1 2 3 4 3 4 3 4 ...)
     (circular-list? x) ⇒ #t

[Scheme Procedure] dotted-list? obj

Return #t if obj is a dotted list, or #f otherwise. A dotted list is a list where the cdr of the last pair is not the empty list (). Any non-pair obj is also considered a dotted list, with length zero.

     (dotted-list? '(1 2 . 3)) ⇒ #t
     (dotted-list? 99)  ⇒ #t

It will be noted that any Scheme object passes exactly one of the above three tests

proper-list?, circular-list? and dotted-list?. Non-lists are dotted-list?, finite
lists are either proper-list? or dotted-list?, and infinite lists are circular-list?.

[Scheme Procedure] null-list? lst

Return #t if lst is the empty list (), #f otherwise. If something else than a proper or circular list is passed as lst, an error is signalled. This procedure is recommended for checking for the end of a list in contexts where dotted lists are not allowed.

[Scheme Procedure] not-pair? obj

Return #t is obj is not a pair, #f otherwise. This is shorthand notation (not (pair? obj)) and is supposed to be used for end-of-list checking in contexts where dotted lists are allowed.

[Scheme Procedure] list= elt= list1 … Return #t if all argument lists are equal, #f otherwise. List equality is determined by testing whether all lists have the same length and the corresponding elements are equal in the sense of the equality predicate elt=. If no or only one list is given, #t is returned.

7.5.3.3 Selectors

[Scheme Procedure] first pair [Scheme Procedure] second pair [Scheme Procedure] third pair [Scheme Procedure] fourth pair [Scheme Procedure] fifth pair [Scheme Procedure] sixth pair [Scheme Procedure] seventh pair [Scheme Procedure] eighth pair [Scheme Procedure] ninth pair [Scheme Procedure] tenth pair

These are synonyms for car, cadr, caddr, . . ..

[Scheme Procedure] car+cdr pair

Return two values, the car and the cdr of pair.

[Scheme Procedure] take lst i [Scheme Procedure] take! lst i

Return a list containing the first i elements of lst. take! may modify the structure of the argument list lst in order to produce the result.

[Scheme Procedure] drop lst i

Return a list containing all but the first i elements of lst.

[Scheme Procedure] take-right lst i

Return a list containing the i last elements of lst. The return shares a common tail with lst.

[Scheme Procedure] drop-right lst i [Scheme Procedure] drop-right! lst i

Return a list containing all but the i last elements of lst. drop-right always returns a new list, even when i is zero. drop-right! may modify the structure of the argument list lst in order to produce the result.

[Scheme Procedure] split-at lst i [Scheme Procedure] split-at! lst i

Return two values, a list containing the first i elements of the list lst and a list containing the remaining elements. split-at! may modify the structure of the argument list lst in order to produce the result.

[Scheme Procedure] last lst

Return the last element of the non-empty, finite list lst.

7.5.3.4 Length, Append, Concatenate, etc.

[Scheme Procedure] length+ lst

Return the length of the argument list lst. When lst is a circular list, #f is returned.

[Scheme Procedure] concatenate list-of-lists [Scheme Procedure] concatenate! list-of-lists

Construct a list by appending all lists in list-of-lists. concatenate! may modify the structure of the given lists in order to produce the result.

concatenate is the same as (apply append list-of-lists). It exists because some

Scheme implementations have a limit on the number of arguments a function takes, which the apply might exceed. In Guile there is no such limit.

[Scheme Procedure] append-reverse rev-head tail [Scheme Procedure] append-reverse! rev-head tail

Reverse rev-head, append tail to it, and return the result. This is equivalent to (append (reverse rev-head) tail), but its implementation is more efficient.

     (append-reverse '(1 2 3) '(4 5 6)) ⇒ (3 2 1 4 5 6)

append-reverse! may modify rev-head in order to produce the result.

[Scheme Procedure] zip lst1 lst2 … Return a list as long as the shortest of the argument lists, where each element is a list. The first list contains the first elements of the argument lists, the second list contains the second elements, and so on.

[Scheme Procedure] unzip1 lst [Scheme Procedure] unzip2 lst [Scheme Procedure] unzip3 lst [Scheme Procedure] unzip4 lst [Scheme Procedure] unzip5 lst

unzip1 takes a list of lists, and returns a list containing the first elements of each list, unzip2 returns two lists, the first containing the first elements of each lists and the second containing the second elements of each lists, and so on.

[Scheme Procedure] count pred lst1 lst2 … Return a count of the number of times pred returns true when called on elements from the given lists. pred is called with N parameters (pred elem1 ... elemN ), each element being from the corresponding list. The first call is with the first element of each list, the second with the second element from each, and so on. Counting stops when the end of the shortest list is reached. At least one list must be non-circular.

7.5.3.5 Fold, Unfold & Map

[Scheme Procedure] fold proc init lst1 lst2 … [Scheme Procedure] fold-right proc init lst1 lst2 …

Apply proc to the elements of lst1 lst2 … to build a result, and return that result. Each proc call is (proc elem1 elem2 ... previous), where elem1 is from lst1, elem2 is from lst2, and so on. previous is the return from the previous call to proc, or the given init for the first call. If any list is empty, just init is returned. fold works through the list elements from first to last. The following shows a list reversal and the calls it makes,

     (fold cons '() '(1 2 3))
     (cons 1 '())
     (cons 2 '(1))
     (cons 3 '(2 1)
     ⇒ (3 2 1)

fold-right works through the list elements from last to first, ie. from the right. So for example the following finds the longest string, and the last among equal longest,

     (fold-right (lambda (str prev)
                   (if (> (string-length str) (string-length prev))
                       str
                       prev))
                 ""
                 '("x" "abc" "xyz" "jk"))
     ⇒ "xyz"

If lst1 lst2 … have different lengths, fold stops when the end of the shortest is reached; fold-right commences at the last element of the shortest. Ie. elements past the length of the shortest are ignored in the other lsts. At least one lst must be non-circular.

fold should be preferred over fold-right if the order of processing doesn’t matter, or can be arranged either way, since fold is a little more efficient.

The way fold builds a result from iterating is quite general, it can do more than other iterations like say map or filter. The following for example removes adjacent duplicate elements from a list,

     (define (delete-adjacent-duplicates lst)
       (fold-right (lambda (elem ret)
                     (if (equal? elem (first ret))
                         ret
                         (cons elem ret)))
                   (list (last lst))
                   lst))
     (delete-adjacent-duplicates '(1 2 3 3 4 4 4 5))
     ⇒ (1 2 3 4 5)

Clearly the same sort of thing can be done with a for-each and a variable in which to build the result, but a self-contained proc can be re-used in multiple contexts, where a for-each would have to be written out each time.

[Scheme Procedure] pair-fold proc init lst1 lst2 … [Scheme Procedure] pair-fold-right proc init lst1 lst2 …

The same as fold and fold-right, but apply proc to the pairs of the lists instead of the list elements.

[Scheme Procedure] reduce proc default lst [Scheme Procedure] reduce-right proc default lst

reduce is a variant of fold, where the first call to proc is on two elements from lst, rather than one element and a given initial value.

If lst is empty, reduce returns default (this is the only use for default). If lst has just one element then that’s the return value. Otherwise proc is called on the elements of lst.

Each proc call is (proc elem previous), where elem is from lst (the second and subsequent elements of lst), and previous is the return from the previous call to proc. The first element of lst is the previous for the first call to proc.

For example, the following adds a list of numbers, the calls made to + are shown. (Of course + accepts multiple arguments and can add a list directly, with apply.)

     (reduce + 0 '(5 6 7)) ⇒ 18
     (+ 6 5) ⇒ 11
     (+ 7 11) ⇒ 18

reduce can be used instead of fold where the init value is an “identity”, meaning a value which under proc doesn’t change the result, in this case 0 is an identity since (+ 5 0) is just 5. reduce avoids that unnecessary call. reduce-right is a similar variation on fold-right, working from the end (ie. the right) of lst. The last element of lst is the previous for the first call to proc, and the elem values go from the second last. reduce should be preferred over reduce-right if the order of processing doesn’t matter, or can be arranged either way, since reduce is a little more efficient.

[Scheme Procedure] unfold p f g seed [tail-gen]

unfold is defined as follows:

     (unfold p f g seed) =
        (if (p seed) (tail-gen seed)
            (cons (f seed)
                  (unfold p f g (g seed))))

p Determines when to stop unfolding.

f Maps each seed value to the corresponding list element.

g Maps each seed value to next seed value.

seed The state value for the unfold.

tail-gen Creates the tail of the list; defaults to (lambda (x) '()).

g produces a series of seed values, which are mapped to list elements by f. These elements are put into a list in left-to-right order, and p tells when to stop unfolding.

[Scheme Procedure] unfold-right p f g seed [tail]

Construct a list with the following loop.

     (let lp ((seed seed) (lis tail))
        (if (p seed) lis
            (lp (g seed)
                (cons (f seed) lis))))

p Determines when to stop unfolding.

f Maps each seed value to the corresponding list element.

g Maps each seed value to next seed value.

seed The state value for the unfold.

tail The tail of the list; defaults to '().

[Scheme Procedure] map f lst1 lst2 … Map the procedure over the list(s) lst1, lst2, … and return a list containing the results of the procedure applications. This procedure is extended with respect to R5RS, because the argument lists may have different lengths. The result list will have the same length as the shortest argument lists. The order in which f will be applied to the list element(s) is not specified.

[Scheme Procedure] for-each f lst1 lst2 … Apply the procedure f to each pair of corresponding elements of the list(s) lst1, lst2, . . .. The return value is not specified. This procedure is extended with respect to R5RS, because the argument lists may have different lengths. The shortest argument list determines the number of times f is called. f will be applied to the list elements in left-to-right order.

[Scheme Procedure] append-map f lst1 lst2 … [Scheme Procedure] append-map! f lst1 lst2 …

Equivalent to

     (apply append (map f clist1 clist2 ...))

and

     (apply append! (map f clist1 clist2 ...))

Map f over the elements of the lists, just as in the map function. However, the results of the applications are appended together to make the final result. append-map uses append to append the results together; append-map! uses append!. The dynamic order in which the various applications of f are made is not specified.

[Scheme Procedure] map! f lst1 lst2 … Linear-update variant of map – map! is allowed, but not required, to alter the cons cells of lst1 to construct the result list. The dynamic order in which the various applications of f are made is not specified. In the n-ary case, lst2, lst3, … must have at least as many elements as lst1.

[Scheme Procedure] pair-for-each f lst1 lst2 …

Like for-each, but applies the procedure f to the pairs from which the argument lists are constructed, instead of the list elements. The return value is not specified.

[Scheme Procedure] filter-map f lst1 lst2 …

Like map, but only results from the applications of f which are true are saved in the result list.

7.5.3.6 Filtering and Partitioning

Filtering means to collect all elements from a list which satisfy a specific condition. Partitioning a list means to make two groups of list elements, one which contains the elements satisfying a condition, and the other for the elements which don’t.

The filter and filter! functions are implemented in the Guile core, See Section 6.7.2.6 [List Modification], page 191.

[Scheme Procedure] partition pred lst [Scheme Procedure] partition! pred lst

Split lst into those elements which do and don’t satisfy the predicate pred. The return is two values (see Section 6.13.7 [Multiple Values], page 301), the first being a list of all elements from lst which satisfy pred, the second a list of those which do not. The elements in the result lists are in the same order as in lst but the order in which the calls (pred elem) are made on the list elements is unspecified. partition does not change lst, but one of the returned lists may share a tail with it. partition! may modify lst to construct its return.

[Scheme Procedure] remove pred lst

[Scheme Procedure] remove! pred lst

Return a list containing all elements from lst which do not satisfy the predicate pred. The elements in the result list have the same order as in lst. The order in which pred is applied to the list elements is not specified.

remove! is allowed, but not required to modify the structure of the input list.

7.5.3.7 Searching

The procedures for searching elements in lists either accept a predicate or a comparison object for determining which elements are to be searched.

[Scheme Procedure] find pred lst Return the first element of lst which satisfies the predicate pred and #f if no such element is found.

[Scheme Procedure] find-tail pred lst

Return the first pair of lst whose car satisfies the predicate pred and #f if no such element is found.

[Scheme Procedure] take-while pred lst [Scheme Procedure] take-while! pred lst

Return the longest initial prefix of lst whose elements all satisfy the predicate pred.

take-while! is allowed, but not required to modify the input list while producing the result.

[Scheme Procedure] drop-while pred lst

Drop the longest initial prefix of lst whose elements all satisfy the predicate pred.

[Scheme Procedure] span pred lst [Scheme Procedure] span! pred lst [Scheme Procedure] break pred lst [Scheme Procedure] break! pred lst

span splits the list lst into the longest initial prefix whose elements all satisfy the predicate pred, and the remaining tail. break inverts the sense of the predicate.

span! and break! are allowed, but not required to modify the structure of the input list lst in order to produce the result.

Note that the name break conflicts with the break binding established by while (see Section 6.13.4 [while do], page 294). Applications wanting to use break from within a while loop will need to make a new define under a different name.

[Scheme Procedure] any pred lst1 lst2 … Test whether any set of elements from lst1 lst2 … satisfies pred. If so, the return value is the return value from the successful pred call, or if not, the return value is

#f.

If there are n list arguments, then pred must be a predicate taking n arguments. Each pred call is (pred elem1 elem2 ... ) taking an element from each lst. The calls are made successively for the first, second, etc. elements of the lists, stopping when pred returns non-#f, or when the end of the shortest list is reached.

The pred call on the last set of elements (i.e., when the end of the shortest list has been reached), if that point is reached, is a tail call.

[Scheme Procedure] every pred lst1 lst2 … Test whether every set of elements from lst1 lst2 … satisfies pred. If so, the return value is the return from the final pred call, or if not, the return value is #f. If there are n list arguments, then pred must be a predicate taking n arguments. Each pred call is (pred elem1 elem2 ...) taking an element from each lst. The calls are made successively for the first, second, etc. elements of the lists, stopping if pred returns #f, or when the end of any of the lists is reached. The pred call on the last set of elements (i.e., when the end of the shortest list has been reached) is a tail call. If one of lst1 lst2 …is empty then no calls to pred are made, and the return value is #t.

[Scheme Procedure] list-index pred lst1 lst2 … Return the index of the first set of elements, one from each of lst1 lst2 …, which satisfies pred. pred is called as (elem1 elem2 ...). Searching stops when the end of the shortest lst is reached. The return index starts from 0 for the first set of elements. If no set of elements pass, then the return value is #f.

     (list-index odd? '(2 4 6 9))  ⇒ 3
     (list-index = '(1 2 3) '(3 1 2)) ⇒ #f

[Scheme Procedure] member x lst [=]

Return the first sublist of lst whose car is equal to x. If x does not appear in lst, return #f. Equality is determined by equal?, or by the equality predicate = if given. = is called (= x elem), ie. with the given x first, so for example to find the first element greater than 5,

     (member 5 '(3 5 1 7 2 9) <) ⇒ (7 2 9)

This version of member extends the core member (see Section 6.7.2.7 [List Searching], page 192) by accepting an equality predicate.

7.5.3.8 Deleting

[Scheme Procedure] delete x lst [=] [Scheme Procedure] delete! x lst [=]

Return a list containing the elements of lst but with those equal to x deleted. The returned elements will be in the same order as they were in lst. Equality is determined by the = predicate, or equal? if not given. An equality call is made just once for each element, but the order in which the calls are made on the elements is unspecified. The equality calls are always (= x elem), ie. the given x is first. This means for instance elements greater than 5 can be deleted with (delete 5 lst <). delete does not modify lst, but the return might share a common tail with lst. delete! may modify the structure of lst to construct its return. These functions extend the core delete and delete! (see Section 6.7.2.6 [List Modification], page 191) in accepting an equality predicate. See also lset-difference (see Section 7.5.3.10 [SRFI-1 Set Operations], page 563) for deleting multiple elements from a list.

[Scheme Procedure] delete-duplicates lst [=] [Scheme Procedure] delete-duplicates! lst [=]

Return a list containing the elements of lst but without duplicates.

When elements are equal, only the first in lst is retained. Equal elements can be anywhere in lst, they don’t have to be adjacent. The returned list will have the retained elements in the same order as they were in lst.

Equality is determined by the = predicate, or equal? if not given. Calls (= x y) are made with element x being before y in lst. A call is made at most once for each combination, but the sequence of the calls across the elements is unspecified.

delete-duplicates does not modify lst, but the return might share a common tail with lst. delete-duplicates! may modify the structure of lst to construct its return.

In the worst case, this is an O(N^{2}) algorithm because it must check each element against all those preceding it. For long lists it is more efficient to sort and then compare only adjacent elements.

7.5.3.9 Association Lists

Association lists are described in detail in section Section 6.7.12 [Association Lists], page 228. The present section only documents the additional procedures for dealing with association lists defined by SRFI-1.

[Scheme Procedure] assoc key alist [=] Return the pair from alist which matches key. This extends the core assoc (see Section 6.7.12.3 [Retrieving Alist Entries], page 230) by taking an optional = comparison procedure.

The default comparison is equal?. If an = parameter is given it’s called (= key alistcar), i.e. the given target key is the first argument, and a car from alist is second.

For example a case-insensitive string lookup,

     (assoc "yy" '(("XX" . 1) ("YY" . 2)) string-ci=?)
     ⇒ ("YY" . 2)

[Scheme Procedure] alist-cons key datum alist Cons a new association key and datum onto alist and return the result. This is equivalent to

     (cons (cons key datum) alist)

acons (see Section 6.7.12.2 [Adding or Setting Alist Entries], page 228) in the Guile core does the same thing.

[Scheme Procedure] alist-copy alist

Return a newly allocated copy of alist, that means that the spine of the list as well as the pairs are copied.

[Scheme Procedure] alist-delete key alist [=] [Scheme Procedure] alist-delete! key alist [=]

Return a list containing the elements of alist but with those elements whose keys are equal to key deleted. The returned elements will be in the same order as they were in alist. Equality is determined by the = predicate, or equal? if not given. The order in which elements are tested is unspecified, but each equality call is made (= key alistkey), i.e. the given key parameter is first and the key from alist second. This means for instance all associations with a key greater than 5 can be removed with (alist-

delete 5 alist <).

alist-delete does not modify alist, but the return might share a common tail with alist. alist-delete! may modify the list structure of alist to construct its return.

7.5.3.10 Set Operations on Lists

Lists can be used to represent sets of objects. The procedures in this section operate on such lists as sets.

Note that lists are not an efficient way to implement large sets. The procedures here typically take time m × n when operating on m and n element lists. Other data structures like trees, bitsets (see Section 6.7.4 [Bit Vectors], page 198) or hash tables (see Section 6.7.14 [Hash Tables], page 236) are faster.

All these procedures take an equality predicate as the first argument. This predicate is used for testing the objects in the list sets for sameness. This predicate must be consistent with eq? (see Section 6.11.1 [Equality], page 276) in the sense that if two list elements are eq? then they must also be equal under the predicate. This simply means a given object must be equal to itself.

[Scheme Procedure] lset<= = list …

Return #t if each list is a subset of the one following it. I.e., list1 is a subset of list2, list2 is a subset of list3, etc., for as many lists as given. If only one list or no lists are given, the return value is #t. A list x is a subset of y if each element of x is equal to some element in y. Elements are compared using the given = procedure, called as (= xelem yelem).

     (lset<= eq?)  ⇒ #t
     (lset<= eqv? '(1 2 3) '(1))  ⇒ #f
     (lset<= eqv? '(1 3 2) '(4 3 1 2)) ⇒ #t

[Scheme Procedure] lset= = list …

Return #t if all argument lists are set-equal. list1 is compared to list2, list2 to list3, etc., for as many lists as given. If only one list or no lists are given, the return value is #t. Two lists x and y are set-equal if each element of x is equal to some element of y and conversely each element of y is equal to some element of x. The order of the elements in the lists doesn’t matter. Element equality is determined with the given = procedure, called as (= xelem yelem), but exactly which calls are made is unspecified.

     (lset= eq?)  ⇒ #t
     (lset= eqv? '(1 2 3) '(3 2 1)) ⇒ #t

     (lset= string-ci=? '("a" "A" "b") '("B" "b" "a")) ⇒ #t

[Scheme Procedure] lset-adjoin = list elem …

Add to list any of the given elems not already in the list. elems are consed onto the start of list (so the return value shares a common tail with list), but the order that the elems are added is unspecified. The given = procedure is used for comparing elements, called as (= listelem elem), i.e., the second argument is one of the given elem parameters.

     (lset-adjoin eqv? '(1 2 3) 4 1 5) ⇒ (5 4 1 2 3)

[Scheme Procedure] lset-union = list … [Scheme Procedure] lset-union! = list …

Return the union of the argument list sets. The result is built by taking the union of list1 and list2, then the union of that with list3, etc., for as many lists as given. For one list argument that list itself is the result, for no list arguments the result is the empty list. The union of two lists x and y is formed as follows. If x is empty then the result is y. Otherwise start with x as the result and consider each y element (from first to last). A y element not equal to something already in the result is consed onto the result. The given = procedure is used for comparing elements, called as (= relem yelem). The first argument is from the result accumulated so far, and the second is from the list being union-ed in. But exactly which calls are made is otherwise unspecified. Notice that duplicate elements in list1 (or the first non-empty list) are preserved, but that repeated elements in subsequent lists are only added once.

     (lset-union eqv?)  ⇒ ()
     (lset-union eqv? '(1 2 3))  ⇒ (1 2 3)
     (lset-union eqv? '(1 2 1 3) '(2 4 5) '(5)) ⇒ (5 4 1 2 1 3)

lset-union doesn’t change the given lists but the result may share a tail with the first non-empty list. lset-union! can modify all of the given lists to form the result.

[Scheme Procedure] lset-intersection = list1 list2 … [Scheme Procedure] lset-intersection! = list1 list2 …

Return the intersection of list1 with the other argument lists, meaning those elements of list1 which are also in all of list2 etc. For one list argument, just that list is returned. The test for an element of list1 to be in the return is simply that it’s equal to some element in each of list2 etc. Notice this means an element appearing twice in list1 but only once in each of list2 etc will go into the return twice. The return has its elements in the same order as they were in list1. The given = procedure is used for comparing elements, called as (= elem1 elemN). The first argument is from list1 and the second is from one of the subsequent lists. But exactly which calls are made and in what order is unspecified.

     (lset-intersection eqv? '(x y))  ⇒ (x y)
     (lset-intersection eqv? '(1 2 3) '(4 3 2))  ⇒ (2 3)
     (lset-intersection eqv? '(1 1 2 2) '(1 2) '(2 1) '(2)) ⇒ (2 2)

The return from lset-intersection may share a tail with list1. lsetintersection! may modify list1 to form its result.

[Scheme Procedure] lset-difference = list1 list2 … [Scheme Procedure] lset-difference! = list1 list2 …

Return list1 with any elements in list2, list3 etc removed (ie. subtracted). For one list argument, just that list is returned. The given = procedure is used for comparing elements, called as (= elem1 elemN). The first argument is from list1 and the second from one of the subsequent lists. But exactly which calls are made and in what order is unspecified.

     (lset-difference eqv? '(x y))  ⇒ (x y)
     (lset-difference eqv? '(1 2 3) '(3 1))  ⇒ (2)
     (lset-difference eqv? '(1 2 3) '(3) '(2)) ⇒ (1)

The return from lset-difference may share a tail with list1. lset-difference! may modify list1 to form its result.

[Scheme Procedure] lset-diff+intersection = list1 list2 … [Scheme Procedure] lset-diff+intersection! = list1 list2 …

Return two values (see Section 6.13.7 [Multiple Values], page 301), the difference and intersection of the argument lists as per lset-difference and lset-intersection above. For two list arguments this partitions list1 into those elements of list1 which are in list2 and not in list2. (But for more than two arguments there can be elements of list1 which are neither part of the difference nor the intersection.) One of the return values from lset-diff+intersection may share a tail with list1. lset-diff+intersection! may modify list1 to form its results.

[Scheme Procedure] lset-xor = list … [Scheme Procedure] lset-xor! = list …

Return an XOR of the argument lists. For two lists this means those elements which are in exactly one of the lists. For more than two lists it means those elements which appear in an odd number of the lists. To be precise, the XOR of two lists x and y is formed by taking those elements of x not equal to any element of y, plus those elements of y not equal to any element of x. Equality is determined with the given = procedure, called as (= e1 e2). One argument is from x and the other from y, but which way around is unspecified. Exactly which calls are made is also unspecified, as is the order of the elements in the result.

     (lset-xor eqv? '(x y))  ⇒ (x y)
     (lset-xor eqv? '(1 2 3) '(4 3 2)) ⇒ (4 1)

The return from lset-xor may share a tail with one of the list arguments. lset-xor! may modify list1 to form its result.

7.5.4 SRFI-2 - and-let*

The following syntax can be obtained with

(use-modules (srfi srfi-2))

or alternatively

(use-modules (ice-9 and-let-star))

[library syntax] and-let* (clause …) body … A combination of and and let*. Each clause is evaluated in turn, and if #f is obtained then evaluation stops and #f is returned. If all are non-#f then body is evaluated and the last form gives the return value, or if body is empty then the result is #t. Each clause should be one of the following,

(symbol expr)

Evaluate expr, check for #f, and bind it to symbol. Like let*, that binding is available to subsequent clauses.

(expr) Evaluate expr and check for #f.

symbol Get the value bound to symbol and check for #f.

Notice that (expr) has an “extra” pair of parentheses, for instance ((eq? x y)). One way to remember this is to imagine the symbol in (symbol expr) is omitted. and-let* is good for calculations where a #f value means termination, but where a non-#f value is going to be needed in subsequent expressions. The following illustrates this, it returns text between brackets ‘[...]’ in a string, or #f if there are no such brackets (ie. either string-index gives #f).

     (define (extract-brackets str)
       (and-let* ((start (string-index str #\[))
                  (end   (string-index str #\] start)))
         (substring str (1+ start) end)))

The following shows plain variables and expressions tested too. diagnostic-levels is taken to be an alist associating a diagnostic type with a level. str is printed only if the type is known and its level is high enough.

     (define (show-diagnostic type str)
       (and-let* (want-diagnostics
                  (level (assq-ref diagnostic-levels type))
                  ((>= level current-diagnostic-level)))
         (display str)))

The advantage of and-let* is that an extended sequence of expressions and tests doesn’t require lots of nesting as would arise from separate and and let*, or from

cond with =>.

7.5.5 SRFI-4 - Homogeneous numeric vector datatypes

SRFI-4 provides an interface to uniform numeric vectors: vectors whose elements are all of a single numeric type. Guile offers uniform numeric vectors for signed and unsigned 8-bit, 16-bit, 32-bit, and 64-bit integers, two sizes of floating point values, and, as an extension to SRFI-4, complex floating-point numbers of these two sizes.

The standard SRFI-4 procedures and data types may be included via loading the appropriate module:

(use-modules (srfi srfi-4))

This module is currently a part of the default Guile environment, but it is a good practice to explicitly import the module. In the future, using SRFI-4 procedures without importing the SRFI-4 module will cause a deprecation message to be printed. (Of course, one may call the C functions at any time. Would that C had modules!)

7.5.5.1 SRFI-4 - Overview

Uniform numeric vectors can be useful since they consume less memory than the nonuniform, general vectors. Also, since the types they can store correspond directly to C types, it is easier to work with them efficiently on a low level. Consider image processing as an example, where you want to apply a filter to some image. While you could store the pixels of an image in a general vector and write a general convolution function, things are much more efficient with uniform vectors: the convolution function knows that all pixels are unsigned 8-bit values (say), and can use a very tight inner loop.

This is implemented in Scheme by having the compiler notice calls to the SRFI-4 accessors, and inline them to appropriate compiled code. From C you have access to the raw array; functions for efficiently working with uniform numeric vectors from C are listed at the end of this section.

Uniform numeric vectors are the special case of one dimensional uniform numeric arrays.

There are 12 standard kinds of uniform numeric vectors, and they all have their own complement of constructors, accessors, and so on. Procedures that operate on a specific kind of uniform numeric vector have a “tag” in their name, indicating the element type.

u8 unsigned 8-bit integers

s8 signed 8-bit integers

u16 unsigned 16-bit integers

s16 signed 16-bit integers

u32 unsigned 32-bit integers

s32 signed 32-bit integers

u64 unsigned 64-bit integers

s64 signed 64-bit integers

f32  the C type float
f64  the C type double

In addition, Guile supports uniform arrays of complex numbers, with the nonstandard tags:

c32 complex numbers in rectangular form with the real and imaginary part being

          a float

c64 complex numbers in rectangular form with the real and imaginary part being

          a double

The external representation (ie. read syntax) for these vectors is similar to normal Scheme vectors, but with an additional tag from the tables above indicating the vector’s type. For example,

     #u16(1 2 3)
     #f64(3.1415 2.71)

Note that the read syntax for floating-point here conflicts with #f for false. In Standard Scheme one can write (1 #f3) for a three element list (1 #f 3), but for Guile (1 #f3) is invalid. (1 #f 3) is almost certainly what one should write anyway to make the intention clear, so this is rarely a problem.

7.5.5.2 SRFI-4 - API

Note that the c32 and c64 functions are only available from (srfi srfi-4 gnu).

[Scheme Procedure] u8vector? obj [Scheme Procedure] s8vector? obj [Scheme Procedure] u16vector? obj [Scheme Procedure] s16vector? obj [Scheme Procedure] u32vector? obj [Scheme Procedure] s32vector? obj [Scheme Procedure] u64vector? obj [Scheme Procedure] s64vector? obj [Scheme Procedure] f32vector? obj [Scheme Procedure] f64vector? obj [Scheme Procedure] c32vector? obj [Scheme Procedure] c64vector? obj [C Function] scm_u8vector_p (obj) [C Function] scm_s8vector_p (obj) [C Function] scm_u16vector_p (obj) [C Function] scm_s16vector_p (obj) [C Function] scm_u32vector_p (obj) [C Function] scm_s32vector_p (obj) [C Function] scm_u64vector_p (obj) [C Function] scm_s64vector_p (obj) [C Function] scm_f32vector_p (obj) [C Function] scm_f64vector_p (obj) [C Function] scm_c32vector_p (obj) [C Function] scm_c64vector_p (obj)

Return #t if obj is a homogeneous numeric vector of the indicated type.

[Scheme Procedure] make-u8vector n [value] [Scheme Procedure] make-s8vector n [value] [Scheme Procedure] make-u16vector n [value] [Scheme Procedure] make-s16vector n [value] [Scheme Procedure] make-u32vector n [value] [Scheme Procedure] make-s32vector n [value] [Scheme Procedure] make-u64vector n [value] [Scheme Procedure] make-s64vector n [value] [Scheme Procedure] make-f32vector n [value] [Scheme Procedure] make-f64vector n [value] [Scheme Procedure] make-c32vector n [value] [Scheme Procedure] make-c64vector n [value] [C Function] scm_make_u8vector (n, value) [C Function] scm_make_s8vector (n, value) [C Function] scm_make_u16vector (n, value) [C Function] scm_make_s16vector (n, value) [C Function] scm_make_u32vector (n, value) [C Function] scm_make_s32vector (n, value) [C Function] scm_make_u64vector (n, value) [C Function] scm_make_s64vector (n, value) [C Function] scm_make_f32vector (n, value) [C Function] scm_make_f64vector (n, value) [C Function] scm_make_c32vector (n, value) [C Function] scm_make_c64vector (n, value)

Return a newly allocated homogeneous numeric vector holding n elements of the indicated type. If value is given, the vector is initialized with that value, otherwise the contents are unspecified.

[Scheme Procedure] u8vector value … [Scheme Procedure] s8vector value … [Scheme Procedure] u16vector value … [Scheme Procedure] s16vector value … [Scheme Procedure] u32vector value … [Scheme Procedure] s32vector value … [Scheme Procedure] u64vector value … [Scheme Procedure] s64vector value … [Scheme Procedure] f32vector value … [Scheme Procedure] f64vector value … [Scheme Procedure] c32vector value … [Scheme Procedure] c64vector value … [C Function] scm_u8vector (values) [C Function] scm_s8vector (values) [C Function] scm_u16vector (values) [C Function] scm_s16vector (values) [C Function] scm_u32vector (values) [C Function] scm_s32vector (values) [C Function] scm_u64vector (values) [C Function] scm_s64vector (values) [C Function] scm_f32vector (values) [C Function] scm_f64vector (values) [C Function] scm_c32vector (values) [C Function] scm_c64vector (values)

Return a newly allocated homogeneous numeric vector of the indicated type, holding the given parameter values. The vector length is the number of parameters given.

[Scheme Procedure] u8vector-length vec [Scheme Procedure] s8vector-length vec [Scheme Procedure] u16vector-length vec [Scheme Procedure] s16vector-length vec [Scheme Procedure] u32vector-length vec [Scheme Procedure] s32vector-length vec [Scheme Procedure] u64vector-length vec [Scheme Procedure] s64vector-length vec [Scheme Procedure] f32vector-length vec [Scheme Procedure] f64vector-length vec [Scheme Procedure] c32vector-length vec [Scheme Procedure] c64vector-length vec

[C Function] scm_u8vector_length (vec)
[C Function] scm_s8vector_length (vec)
[C Function] scm_u16vector_length (vec)
[C Function] scm_s16vector_length (vec)
[C Function] scm_u32vector_length (vec)
[C Function] scm_s32vector_length (vec)
[C Function] scm_u64vector_length (vec)
[C Function] scm_s64vector_length (vec)
[C Function] scm_f32vector_length (vec)
[C Function] scm_f64vector_length (vec)
[C Function] scm_c32vector_length (vec)
[C Function] scm_c64vector_length (vec)

Return the number of elements in vec.

[Scheme Procedure] u8vector-ref vec i [Scheme Procedure] s8vector-ref vec i [Scheme Procedure] u16vector-ref vec i [Scheme Procedure] s16vector-ref vec i [Scheme Procedure] u32vector-ref vec i [Scheme Procedure] s32vector-ref vec i [Scheme Procedure] u64vector-ref vec i [Scheme Procedure] s64vector-ref vec i [Scheme Procedure] f32vector-ref vec i [Scheme Procedure] f64vector-ref vec i [Scheme Procedure] c32vector-ref vec i [Scheme Procedure] c64vector-ref vec i [C Function] scm_u8vector_ref (vec, i) [C Function] scm_s8vector_ref (vec, i) [C Function] scm_u16vector_ref (vec, i) [C Function] scm_s16vector_ref (vec, i) [C Function] scm_u32vector_ref (vec, i) [C Function] scm_s32vector_ref (vec, i) [C Function] scm_u64vector_ref (vec, i) [C Function] scm_s64vector_ref (vec, i) [C Function] scm_f32vector_ref (vec, i) [C Function] scm_f64vector_ref (vec, i) [C Function] scm_c32vector_ref (vec, i) [C Function] scm_c64vector_ref (vec, i)

Return the element at index i in vec. The first element in vec is index 0.

[Scheme Procedure] u8vector-set! vec i value [Scheme Procedure] s8vector-set! vec i value [Scheme Procedure] u16vector-set! vec i value [Scheme Procedure] s16vector-set! vec i value [Scheme Procedure] u32vector-set! vec i value [Scheme Procedure] s32vector-set! vec i value [Scheme Procedure] u64vector-set! vec i value [Scheme Procedure] s64vector-set! vec i value [Scheme Procedure] f32vector-set! vec i value [Scheme Procedure] f64vector-set! vec i value [Scheme Procedure] c32vector-set! vec i value [Scheme Procedure] c64vector-set! vec i value [C Function] scm_u8vector_set_x (vec, i, value) [C Function] scm_s8vector_set_x (vec, i, value) [C Function] scm_u16vector_set_x (vec, i, value) [C Function] scm_s16vector_set_x (vec, i, value) [C Function] scm_u32vector_set_x (vec, i, value) [C Function] scm_s32vector_set_x (vec, i, value) [C Function] scm_u64vector_set_x (vec, i, value) [C Function] scm_s64vector_set_x (vec, i, value) [C Function] scm_f32vector_set_x (vec, i, value) [C Function] scm_f64vector_set_x (vec, i, value) [C Function] scm_c32vector_set_x (vec, i, value) [C Function] scm_c64vector_set_x (vec, i, value)

Set the element at index i in vec to value. The first element in vec is index 0. The return value is unspecified.

[Scheme Procedure] u8vector->list vec [Scheme Procedure] s8vector->list vec [Scheme Procedure] u16vector->list vec [Scheme Procedure] s16vector->list vec [Scheme Procedure] u32vector->list vec [Scheme Procedure] s32vector->list vec [Scheme Procedure] u64vector->list vec [Scheme Procedure] s64vector->list vec [Scheme Procedure] f32vector->list vec [Scheme Procedure] f64vector->list vec [Scheme Procedure] c32vector->list vec [Scheme Procedure] c64vector->list vec

[C Function] scm_u8vector_to_list (vec)
[C Function] scm_s8vector_to_list (vec)
[C Function] scm_u16vector_to_list (vec)
[C Function] scm_s16vector_to_list (vec)
[C Function] scm_u32vector_to_list (vec)
[C Function] scm_s32vector_to_list (vec)
[C Function] scm_u64vector_to_list (vec)
[C Function] scm_s64vector_to_list (vec)

[C Function] scm_f32vector_to_list (vec)
[C Function] scm_f64vector_to_list (vec)
[C Function] scm_c32vector_to_list (vec)
[C Function] scm_c64vector_to_list (vec)

Return a newly allocated list holding all elements of vec.

[Scheme Procedure] list->u8vector lst [Scheme Procedure] list->s8vector lst [Scheme Procedure] list->u16vector lst [Scheme Procedure] list->s16vector lst [Scheme Procedure] list->u32vector lst [Scheme Procedure] list->s32vector lst [Scheme Procedure] list->u64vector lst [Scheme Procedure] list->s64vector lst [Scheme Procedure] list->f32vector lst [Scheme Procedure] list->f64vector lst [Scheme Procedure] list->c32vector lst [Scheme Procedure] list->c64vector lst

[C Function] scm_list_to_u8vector (lst)
[C Function] scm_list_to_s8vector (lst)
[C Function] scm_list_to_u16vector (lst)
[C Function] scm_list_to_s16vector (lst)
[C Function] scm_list_to_u32vector (lst)
[C Function] scm_list_to_s32vector (lst)
[C Function] scm_list_to_u64vector (lst)
[C Function] scm_list_to_s64vector (lst)
[C Function] scm_list_to_f32vector (lst)
[C Function] scm_list_to_f64vector (lst)
[C Function] scm_list_to_c32vector (lst)
[C Function] scm_list_to_c64vector (lst)

Return a newly allocated homogeneous numeric vector of the indicated type, initialized with the elements of the list lst.

[C Function] SCM scm_take_u8vector (const scm t uint8 *data, size t len) [C Function] SCM scm_take_s8vector (const scm t int8 *data, size t len) [C Function] SCM scm_take_u16vector (const scm t uint16 *data, size t len) [C Function] SCM scm_take_s16vector (const scm t int16 *data, size t len) [C Function] SCM scm_take_u32vector (const scm t uint32 *data, size t len) [C Function] SCM scm_take_s32vector (const scm t int32 *data, size t len) [C Function] SCM scm_take_u64vector (const scm t uint64 *data, size t len) [C Function] SCM scm_take_s64vector (const scm t int64 *data, size t len) [C Function] SCM scm_take_f32vector (const float *data, size t len) [C Function] SCM scm_take_f64vector (const double *data, size t len) [C Function] SCM scm_take_c32vector (const float *data, size t len) [C Function] SCM scm_take_c64vector (const double *data, size t len)

Return a new uniform numeric vector of the indicated type and length that uses the memory pointed to by data to store its elements. This memory will eventually be freed with free. The argument len specifies the number of elements in data, not its size in bytes.

The c32 and c64 variants take a pointer to a C array of floats or doubles. The real parts of the complex numbers are at even indices in that array, the corresponding imaginary parts are at the following odd index.

[C Function] const scm_t_uint8 * scm_u8vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_int8 * scm_s8vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_uint16 * scm_u16vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_int16 * scm_s16vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_uint32 * scm_u32vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_int32 * scm_s32vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_uint64 * scm_u64vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const scm_t_int64 * scm_s64vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const float * scm_f32vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const double * scm_f64vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const float * scm_c32vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] const double * scm_c64vector_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp) Like scm_vector_elements (see Section 6.7.3.4 [Vector Accessing from C], page 196), but returns a pointer to the elements of a uniform numeric vector of the indicated kind.

[C Function] scm_t_uint8 * scm_u8vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_int8 * scm_s8vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_uint16 * scm_u16vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_int16 * scm_s16vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_uint32 * scm_u32vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_int32 * scm_s32vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_uint64 * scm_u64vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] scm_t_int64 * scm_s64vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] float * scm_f32vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] double * scm_f64vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] float * scm_c32vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp)

[C Function] double * scm_c64vector_writable_elements (SCM vec,

scm t array handle *handle, size t *lenp, ssize t *incp) Like scm_vector_writable_elements (see Section 6.7.3.4 [Vector Accessing from C], page 196), but returns a pointer to the elements of a uniform numeric vector of the indicated kind.

7.5.5.3 SRFI-4 - Relation to bytevectors

Guile implements SRFI-4 vectors using bytevectors (see Section 6.6.6 [Bytevectors], page 163). Often when you have a numeric vector, you end up wanting to write its bytes somewhere, or have access to the underlying bytes, or read in bytes from somewhere else. Bytevectors are very good at this sort of thing. But the SRFI-4 APIs are nicer to use when doing number-crunching, because they are addressed by element and not by byte.

So as a compromise, Guile allows all bytevector functions to operate on numeric vectors. They address the underlying bytes in the native endianness, as one would expect.

Following the same reasoning, that it’s just bytes underneath, Guile also allows uniform vectors of a given type to be accessed as if they were of any type. One can fill a u32vector, and access its elements with u8vector-ref. One can use f64vector-ref on bytevectors. It’s all the same to Guile.

In this way, uniform numeric vectors may be written to and read from input/output ports using the procedures that operate on bytevectors.

See Section 6.6.6 [Bytevectors], page 163, for more information.

7.5.5.4 SRFI-4 - Guile extensions

Guile defines some useful extensions to SRFI-4, which are not available in the default Guile environment. They may be imported by loading the extensions module:

     (use-modules (srfi srfi-4 gnu))

[Scheme Procedure] any->u8vector obj [Scheme Procedure] any->s8vector obj [Scheme Procedure] any->u16vector obj [Scheme Procedure] any->s16vector obj [Scheme Procedure] any->u32vector obj [Scheme Procedure] any->s32vector obj [Scheme Procedure] any->u64vector obj [Scheme Procedure] any->s64vector obj [Scheme Procedure] any->f32vector obj [Scheme Procedure] any->f64vector obj [Scheme Procedure] any->c32vector obj [Scheme Procedure] any->c64vector obj

[C Function] scm_any_to_u8vector (obj)
[C Function] scm_any_to_s8vector (obj)
[C Function] scm_any_to_u16vector (obj)
[C Function] scm_any_to_s16vector (obj)
[C Function] scm_any_to_u32vector (obj)
[C Function] scm_any_to_s32vector (obj)
[C Function] scm_any_to_u64vector (obj)
[C Function] scm_any_to_s64vector (obj)
[C Function] scm_any_to_f32vector (obj)
[C Function] scm_any_to_f64vector (obj)
[C Function] scm_any_to_c32vector (obj)
[C Function] scm_any_to_c64vector (obj)

Return a (maybe newly allocated) uniform numeric vector of the indicated type, initialized with the elements of obj, which must be a list, a vector, or a uniform vector. When obj is already a suitable uniform numeric vector, it is returned unchanged.

7.5.6 SRFI-6 - Basic String Ports

SRFI-6 defines the procedures open-input-string, open-output-string and
get-output-string.

Note that although versions of these procedures are included in the Guile core, the core versions are not fully conformant with SRFI-6: attempts to read or write characters that are not supported by the current %default-port-encoding will fail.

We therefore recommend that you import this module, which supports all characters:

     (use-modules (srfi srfi-6))

7.5.7 SRFI-8 - receive

receive is a syntax for making the handling of multiple-value procedures easier. It is documented in See Section 6.13.7 [Multiple Values], page 301.

7.5.8 SRFI-9 - define-record-type

This SRFI is a syntax for defining new record types and creating predicate, constructor, and field getter and setter functions. It is documented in the “Compound Data Types” section of the manual (see Section 6.7.8 [SRFI-9 Records], page 216).

7.5.9 SRFI-10 - Hash-Comma Reader Extension

This SRFI implements a reader extension #,() called hash-comma. It allows the reader to give new kinds of objects, for use both in data and as constants or literals in source code. This feature is available with

     (use-modules (srfi srfi-10))

The new read syntax is of the form

     #,(tag arg...)

where tag is a symbol and the args are objects taken as parameters. tags are registered with the following procedure.

[Scheme Procedure] define-reader-ctor tag proc

Register proc as the constructor for a hash-comma read syntax starting with symbol tag, i.e. #,(tag arg...). proc is called with the given arguments (proc arg...) and the object it returns is the result of the read.

For example, a syntax giving a list of N copies of an object.

(define-reader-ctor 'repeat
  (lambda (obj reps)
    (make-list reps obj)))
(display '#,(repeat 99 3))
⊣ (99 99 99)

Notice the quote ' when the #,( ) is used. The repeat handler returns a list and the program must quote to use it literally, the same as any other list. Ie.

(display '#,(repeat 99 3))

⇒

(display '(99 99 99))

When a handler returns an object which is self-evaluating, like a number or a string, then there’s no need for quoting, just as there’s no need when giving those directly as literals. For example an addition,

(define-reader-ctor 'sum
  (lambda (x y)
    (+ x y)))
(display #,(sum 123 456)) ⊣ 579

A typical use for #,() is to get a read syntax for objects which don’t otherwise have one. For example, the following allows a hash table to be given literally, with tags and values, ready for fast lookup.

(define-reader-ctor 'hash
  (lambda elems
    (let ((table (make-hash-table)))
      (for-each (lambda (elem)
                  (apply hash-set! table elem))
                elems)
      table)))
(define (animal->family animal)
  (hash-ref '#,(hash ("tiger" "cat")
                     ("lion"  "cat")
                     ("wolf"  "dog"))
            animal))
(animal->family "lion") ⇒ "cat"

Or for example the following is a syntax for a compiled regular expression (see Section 6.15 [Regular Expressions], page 349).

     (use-modules (ice-9 regex))
     (define-reader-ctor 'regexp make-regexp)
     (define (extract-angs str)
       (let ((match (regexp-exec '#,(regexp "<([A-Z0-9]+)>") str)))
         (and match
              (match:substring match 1))))
     (extract-angs "foo <BAR> quux") ⇒ "BAR"

#,() is somewhat similar to define-macro (see Section 6.10 [Macros], page 257) in that handler code is run to produce a result, but #,() operates at the read stage, so it can appear in data for read (see Section 6.17.2 [Scheme Read], page 360), not just in code to be executed.

Because #,() is handled at read-time it has no direct access to variables etc. A symbol in the arguments is just a symbol, not a variable reference. The arguments are essentially constants, though the handler procedure can use them in any complicated way it might want.

Once (srfi srfi-10) has loaded, #,() is available globally, there’s no need to use (srfi srfi-10) in later modules. Similarly the tags registered are global and can be used anywhere once registered.

There’s no attempt to record what previous #,() forms have been seen, if two identical forms occur then two calls are made to the handler procedure. The handler might like to maintain a cache or similar to avoid making copies of large objects, depending on expected usage.

In code the best uses of #,() are generally when there’s a lot of objects of a particular kind as literals or constants. If there’s just a few then some local variables and initializers are fine, but that becomes tedious and error prone when there’s a lot, and the anonymous and compact syntax of #,() is much better.

7.5.10 SRFI-11 - let-values

This module implements the binding forms for multiple values let-values and let*- values. These forms are similar to let and let* (see Section 6.12.2 [Local Bindings], page 287), but they support binding of the values returned by multiple-valued expressions.

Write (use-modules (srfi srfi-11)) to make the bindings available.

     (let-values (((x y) (values 1 2))
                  ((z f) (values 3 4)))
        (+ x y z f))

⇒

     10

let-values performs all bindings simultaneously, which means that no expression in the binding clauses may refer to variables bound in the same clause list. let*-values, on the other hand, performs the bindings sequentially, just like let* does for single-valued expressions.

7.5.11 SRFI-13 - String Library

The SRFI-13 procedures are always available, See Section 6.6.5 [Strings], page 141.

7.5.12 SRFI-14 - Character-set Library

The SRFI-14 data type and procedures are always available, See Section 6.6.4 [Character Sets], page 133.

7.5.13 SRFI-16 - case-lambda

SRFI-16 defines a variable-arity lambda form, case-lambda. This form is available in the default Guile environment. See Section 6.9.5 [Case-lambda], page 251, for more information.

7.5.14 SRFI-17 - Generalized set!

This SRFI implements a generalized set!, allowing some “referencing” functions to be used as the target location of a set!. This feature is available from

     (use-modules (srfi srfi-17))

For example vector-ref is extended so that

     (set! (vector-ref vec idx) new-value)

is equivalent to

     (vector-set! vec idx new-value)

The idea is that a vector-ref expression identifies a location, which may be either fetched or stored. The same form is used for the location in both cases, encouraging visual clarity. This is similar to the idea of an “lvalue” in C.

The mechanism for this kind of set! is in the Guile core (see Section 6.9.8 [Procedures with Setters], page 255). This module adds definitions of the following functions as procedures with setters, allowing them to be targets of a set!,

     car, cdr, caar, cadr, cdar, cddr, caaar, caadr, cadar, caddr, cdaar, cdadr,
     cddar, cdddr, caaaar, caaadr, caadar, caaddr, cadaar, cadadr, caddar,
     cadddr, cdaaar, cdaadr, cdadar, cdaddr, cddaar, cddadr, cdddar, cddddr
     string-ref, vector-ref

The SRFI specifies setter (see Section 6.9.8 [Procedures with Setters], page 255) as a procedure with setter, allowing the setter for a procedure to be changed, eg. (set! (setter foo) my-new-setter-handler). Currently Guile does not implement this, a setter can only be specified on creation (getter-with-setter below).

[Function] getter-with-setter

The same as the Guile core make-procedure-with-setter (see Section 6.9.8 [Procedures with Setters], page 255).

7.5.15 SRFI-18 - Multithreading support

This is an implementation of the SRFI-18 threading and synchronization library. The functions and variables described here are provided by

(use-modules (srfi srfi-18))

As a general rule, the data types and functions in this SRFI-18 implementation are compatible with the types and functions in Guile’s core threading code. For example, mutexes created with the SRFI-18 make-mutex function can be passed to the built-in Guile function lock-mutex (see Section 6.21.4 [Mutexes and Condition Variables], page 415), and mutexes created with the built-in Guile function make-mutex can be passed to the SRFI-18 function mutex-lock!. Cases in which this does not hold true are noted in the following sections.

7.5.15.1 SRFI-18 Threads

Threads created by SRFI-18 differ in two ways from threads created by Guile’s built-in thread functions. First, a thread created by SRFI-18 make-thread begins in a blocked state and will not start execution until thread-start! is called on it. Second, SRFI-18 threads are constructed with a top-level exception handler that captures any exceptions that are thrown on thread exit. In all other regards, SRFI-18 threads are identical to normal Guile threads.

[Function] current-thread

Returns the thread that called this function. This is the same procedure as the samenamed built-in procedure current-thread (see Section 6.21.3 [Threads], page 413).

[Function] thread? obj

Returns #t if obj is a thread, #f otherwise. This is the same procedure as the samenamed built-in procedure thread? (see Section 6.21.3 [Threads], page 413).

[Function] make-thread thunk [name]

Call thunk in a new thread and with a new dynamic state, returning the new thread and optionally assigning it the object name name, which may be any Scheme object. Note that the name make-thread conflicts with the (ice-9 threads) function makethread. Applications wanting to use both of these functions will need to refer to them by different names.

[Function] thread-name thread

Returns the name assigned to thread at the time of its creation, or #f if it was not given a name.

[Function] thread-specific thread

[Function] thread-specific-set! thread obj

Get or set the “object-specific” property of thread. In Guile’s implementation of SRFI-18, this value is stored as an object property, and will be #f if not set.

[Function] thread-start! thread

Unblocks thread and allows it to begin execution if it has not done so already.

[Function] thread-yield!

If one or more threads are waiting to execute, calling thread-yield! forces an immediate context switch to one of them. Otherwise, thread-yield! has no effect. thread-yield! behaves identically to the Guile built-in function yield.

[Function] thread-sleep! timeout

The current thread waits until the point specified by the time object timeout is reached (see Section 7.5.15.4 [SRFI-18 Time], page 582). This blocks the thread only if timeout represents a point in the future. it is an error for timeout to be #f.

[Function] thread-terminate! thread

Causes an abnormal termination of thread. If thread is not already terminated, all mutexes owned by thread become unlocked/abandoned. If thread is the current thread, thread-terminate! does not return. Otherwise thread-terminate! returns an unspecified value; the termination of thread will occur before thread-terminate! returns. Subsequent attempts to join on thread will cause a “terminated thread exception” to be raised.

thread-terminate! is compatible with the thread cancellation procedures in the core threads API (see Section 6.21.3 [Threads], page 413) in that if a cleanup handler has been installed for the target thread, it will be called before the thread exits and its return value (or exception, if any) will be stored for later retrieval via a call to

thread-join!.

[Function] thread-join! thread [timeout [timeout-val]] Wait for thread to terminate and return its exit value. When a time value timeout is given, it specifies a point in time where the waiting should be aborted. When the waiting is aborted, timeout-val is returned if it is specified; otherwise, a jointimeout-exception exception is raised (see Section 7.5.15.5 [SRFI-18 Exceptions], page 582). Exceptions may also be raised if the thread was terminated by a call to

thread-terminate! (terminated-thread-exception will be raised) or if the thread

exited by raising an exception that was handled by the top-level exception handler (uncaught-exception will be raised; the original exception can be retrieved using

uncaught-exception-reason).

7.5.15.2 SRFI-18 Mutexes

The behavior of Guile’s built-in mutexes is parameterized via a set of flags passed to the make-mutex procedure in the core (see Section 6.21.4 [Mutexes and Condition Variables], page 415). To satisfy the requirements for mutexes specified by SRFI-18, the make-mutex procedure described below sets the following flags:

• recursive: the mutex can be locked recursively

• unchecked-unlock: attempts to unlock a mutex that is already unlocked will not raise an exception

• allow-external-unlock: the mutex can be unlocked by any thread, not just the thread that locked it originally

[Function] make-mutex [name]

Returns a new mutex, optionally assigning it the object name name, which may be any Scheme object. The returned mutex will be created with the configuration described above. Note that the name make-mutex conflicts with Guile core function make-mutex. Applications wanting to use both of these functions will need to refer to them by different names.

[Function] mutex-name mutex

Returns the name assigned to mutex at the time of its creation, or #f if it was not given a name.

[Function] mutex-specific mutex

[Function] mutex-specific-set! mutex obj

Get or set the “object-specific” property of mutex. In Guile’s implementation of SRFI-18, this value is stored as an object property, and will be #f if not set.

[Function] mutex-state mutex

Returns information about the state of mutex. Possible values are:

• thread T: the mutex is in the locked/owned state and thread T is the owner of the mutex

• symbol not-owned: the mutex is in the locked/not-owned state

• symbol abandoned: the mutex is in the unlocked/abandoned state

• symbol not-abandoned: the mutex is in the unlocked/not-abandoned state

[Function] mutex-lock! mutex [timeout [thread]] Lock mutex, optionally specifying a time object timeout after which to abort the lock attempt and a thread thread giving a new owner for mutex different than the current thread. This procedure has the same behavior as the lock-mutex procedure in the core library.

[Function] mutex-unlock! mutex [condition-variable [timeout]] Unlock mutex, optionally specifying a condition variable condition-variable on which to wait, either indefinitely or, optionally, until the time object timeout has passed, to be signalled. This procedure has the same behavior as the unlock-mutex procedure in the core library.

7.5.15.3 SRFI-18 Condition variables

SRFI-18 does not specify a “wait” function for condition variables. Waiting on a condition variable can be simulated using the SRFI-18 mutex-unlock! function described in the previous section, or Guile’s built-in wait-condition-variable procedure can be used.

[Function] condition-variable? obj

Returns #t if obj is a condition variable, #f otherwise. This is the same procedure as the same-named built-in procedure (see Section 6.21.4 [Mutexes and Condition Variables], page 415).

[Function] make-condition-variable [name]

Returns a new condition variable, optionally assigning it the object name name, which may be any Scheme object. This procedure replaces a procedure of the same name in the core library.

[Function] condition-variable-name condition-variable

Returns the name assigned to condition-variable at the time of its creation, or #f if it was not given a name.

[Function] condition-variable-specific condition-variable

[Function] condition-variable-specific-set! condition-variable obj

Get or set the “object-specific” property of condition-variable. In Guile’s implementation of SRFI-18, this value is stored as an object property, and will be #f if not set.

[Function] condition-variable-signal! condition-variable

[Function] condition-variable-broadcast! condition-variable

Wake up one thread that is waiting for condition-variable, in the case of conditionvariable-signal!, or all threads waiting for it, in the case of condition-variable-broadcast!. The behavior of these procedures is equivalent to that of the procedures

signal-condition-variable and broadcast-condition-variable in the core li-

brary.

7.5.15.4 SRFI-18 Time

The SRFI-18 time functions manipulate time in two formats: a “time object” type that represents an absolute point in time in some implementation-specific way; and the number of seconds since some unspecified “epoch”. In Guile’s implementation, the epoch is the Unix epoch, 00:00:00 UTC, January 1, 1970.

[Function] current-time

Return the current time as a time object. This procedure replaces the procedure of the same name in the core library, which returns the current time in seconds since the epoch.

[Function] time? obj

Returns #t if obj is a time object, #f otherwise.

[Function] time->seconds time

[Function] seconds->time seconds

Convert between time objects and numerical values representing the number of seconds since the epoch. When converting from a time object to seconds, the return value is the number of seconds between time and the epoch. When converting from seconds to a time object, the return value is a time object that represents a time seconds seconds after the epoch.

7.5.15.5 SRFI-18 Exceptions

SRFI-18 exceptions are identical to the exceptions provided by Guile’s implementation of SRFI-34. The behavior of exception handlers invoked to handle exceptions thrown from SRFI-18 functions, however, differs from the conventional behavior of SRFI-34 in that the continuation of the handler is the same as that of the call to the function. Handlers are called in a tail-recursive manner; the exceptions do not “bubble up”.

[Function] current-exception-handler

Returns the current exception handler.

[Function] with-exception-handler handler thunk

Installs handler as the current exception handler and calls the procedure thunk with no arguments, returning its value as the value of the exception. handler must be a procedure that accepts a single argument. The current exception handler at the time this procedure is called will be restored after the call returns.

[Function] raise obj

Raise obj as an exception. This is the same procedure as the same-named procedure defined in SRFI 34.

[Function] join-timeout-exception? obj

Returns #t if obj is an exception raised as the result of performing a timed join on a thread that does not exit within the specified timeout, #f otherwise.

[Function] abandoned-mutex-exception? obj

Returns #t if obj is an exception raised as the result of attempting to lock a mutex that has been abandoned by its owner thread, #f otherwise.

[Function] terminated-thread-exception? obj

Returns #t if obj is an exception raised as the result of joining on a thread that exited as the result of a call to thread-terminate!.

[Function] uncaught-exception? obj
[Function] uncaught-exception-reason exc

uncaught-exception? returns #t if obj is an exception thrown as the result of joining a thread that exited by raising an exception that was handled by the top-level exception handler installed by make-thread. When this occurs, the original exception is preserved as part of the exception thrown by thread-join! and can be accessed by calling uncaught-exception-reason on that exception. Note that because this exception-preservation mechanism is a side-effect of make-thread, joining on threads that exited as described above but were created by other means will not raise this

uncaught-exception error.

7.5.16 SRFI-19 - Time/Date Library

This is an implementation of the SRFI-19 time/date library. The functions and variables described here are provided by

(use-modules (srfi srfi-19))

Caution: The current code in this module incorrectly extends the Gregorian calendar leap year rule back prior to the introduction of those reforms in 1582 (or the appropriate year in various countries). The Julian calendar was used prior to 1582, and there were 10 days skipped for the reform, but the code doesn’t implement that.

This will be fixed some time. Until then calculations for 1583 onwards are correct, but prior to that any day/month/year and day of the week calculations are wrong.

7.5.16.1 SRFI-19 Introduction

This module implements time and date representations and calculations, in various time systems, including universal time (UTC) and atomic time (TAI).

For those not familiar with these time systems, TAI is based on a fixed length second derived from oscillations of certain atoms. UTC differs from TAI by an integral number of seconds, which is increased or decreased at announced times to keep UTC aligned to a mean solar day (the orbit and rotation of the earth are not quite constant).

So far, only increases in the TAI ↔ UTC difference have been needed. Such an increase is a “leap second”, an extra second of TAI introduced at the end of a UTC day. When working entirely within UTC this is never seen, every day simply has 86400 seconds. But when converting from TAI to a UTC date, an extra 23:59:60 is present, where normally a day would end at 23:59:59. Effectively the UTC second from 23:59:59 to 00:00:00 has taken two TAI seconds.

In the current implementation, the system clock is assumed to be UTC, and a table of leap seconds in the code converts to TAI. See comments in srfi-19.scm for how to update this table.

Also, for those not familiar with the terminology, a Julian Day is a real number which is a count of days and fraction of a day, in UTC, starting from -4713-01-01T12:00:00Z, ie. midday Monday 1 Jan 4713 B.C. A Modified Julian Day is the same, but starting from 1858-11-17T00:00:00Z, ie. midnight 17 November 1858 UTC. That time is julian day 2400000.5.

7.5.16.2 SRFI-19 Time

A time object has type, seconds and nanoseconds fields representing a point in time starting from some epoch. This is an arbitrary point in time, not just a time of day. Although times are represented in nanoseconds, the actual resolution may be lower.

The following variables hold the possible time types. For instance (current-time timeprocess) would give the current CPU process time.

[Variable] time-utc

Universal Coordinated Time (UTC).

[Variable] time-tai

International Atomic Time (TAI).

[Variable] time-monotonic

Monotonic time, meaning a monotonically increasing time starting from an unspecified epoch. Note that in the current implementation time-monotonic is the same as time-tai, and unfortunately is therefore affected by adjustments to the system clock. Perhaps this will change in the future.

[Variable] time-duration

A duration, meaning simply a difference between two times.

[Variable] time-process

CPU time spent in the current process, starting from when the process began.

[Variable] time-thread

CPU time spent in the current thread. Not currently implemented.

[Function] time? obj

Return #t if obj is a time object, or #f if not.

[Function] make-time type nanoseconds seconds Create a time object with the given type, seconds and nanoseconds.

[Function] time-type time

[Function] time-nanosecond time

[Function] time-second time

[Function] set-time-type! time type

[Function] set-time-nanosecond! time nsec

[Function] set-time-second! time sec

Get or set the type, seconds or nanoseconds fields of a time object. set-time-type! merely changes the field, it doesn’t convert the time value. For conversions, see Section 7.5.16.4 [SRFI-19 Time/Date conversions], page 587.

[Function] copy-time time

Return a new time object, which is a copy of the given time.

[Function] current-time [type]

Return the current time of the given type. The default type is time-utc. Note that the name current-time conflicts with the Guile core current-time function (see Section 7.2.5 [Time], page 485) as well as the SRFI-18 current-time function (see Section 7.5.15.4 [SRFI-18 Time], page 582). Applications wanting to use more than one of these functions will need to refer to them by different names.

[Function] time-resolution [type]

Return the resolution, in nanoseconds, of the given time type. The default type is

time-utc.

[Function] time<=? t1 t2

[Function] time<? t1 t2

[Function] time=? t1 t2

[Function] time>=? t1 t2

[Function] time>? t1 t2

Return #t or #f according to the respective relation between time objects t1 and t2. t1 and t2 must be the same time type.

[Function] time-difference t1 t2
[Function] time-difference! t1 t2

Return a time object of type time-duration representing the period between t1 and t2. t1 and t2 must be the same time type. time-difference returns a new time object, time-difference! may modify t1 to form its return.

[Function] add-duration time duration

[Function] add-duration! time duration

[Function] subtract-duration time duration

[Function] subtract-duration! time duration

Return a time object which is time with the given duration added or subtracted. duration must be a time object of type time-duration.

add-duration and subtract-duration return a new time object. add-duration!

and subtract-duration! may modify the given time to form their return.

7.5.16.3 SRFI-19 Date

A date object represents a date in the Gregorian calendar and a time of day on that date in some timezone.

The fields are year, month, day, hour, minute, second, nanoseconds and timezone. A date object is immutable, its fields can be read but they cannot be modified once the object is created.

[Function] date? obj

Return #t if obj is a date object, or #f if not.

[Function] make-date nsecs seconds minutes hours date month year zone-offset Create a new date object.

[Function] date-nanosecond date

Nanoseconds, 0 to 999999999.

[Function] date-second date

Seconds, 0 to 59, or 60 for a leap second. 60 is never seen when working entirely within UTC, it’s only when converting to or from TAI.

[Function] date-minute date

Minutes, 0 to 59.

[Function] date-hour date

Hour, 0 to 23.

[Function] date-day date

Day of the month, 1 to 31 (or less, according to the month).

[Function] date-month date

Month, 1 to 12.

[Function] date-year date

Year, eg. 2003. Dates B.C. are negative, eg. −46 is 46 B.C. There is no year 0, year −1 is followed by year 1.

[Function] date-zone-offset date

Time zone, an integer number of seconds east of Greenwich.

[Function] date-year-day date

Day of the year, starting from 1 for 1st January.

[Function] date-week-day date

Day of the week, starting from 0 for Sunday.

[Function] date-week-number date dstartw

Week of the year, ignoring a first partial week. dstartw is the day of the week which is taken to start a week, 0 for Sunday, 1 for Monday, etc.

[Function] current-date [tz-offset]

Return a date object representing the current date/time, in UTC offset by tz-offset. tz-offset is seconds east of Greenwich and defaults to the local timezone.

[Function] current-julian-day

Return the current Julian Day.

[Function] current-modified-julian-day

Return the current Modified Julian Day.

7.5.16.4 SRFI-19 Time/Date conversions

[Function] date->julian-day date
[Function] date->modified-julian-day date
[Function] date->time-monotonic date
[Function] date->time-tai date
[Function] date->time-utc date

[Function] julian-day->date jdn [tz-offset]

[Function] julian-day->time-monotonic jdn
[Function] julian-day->time-tai jdn
[Function] julian-day->time-utc jdn
[Function] modified-julian-day->date jdn [tz-offset]
[Function] modified-julian-day->time-monotonic jdn
[Function] modified-julian-day->time-tai jdn
[Function] modified-julian-day->time-utc jdn

[Function] time-monotonic->date time [tz-offset]

[Function] time-monotonic->time-tai time
[Function] time-monotonic->time-tai! time
[Function] time-monotonic->time-utc time
[Function] time-monotonic->time-utc! time

[Function] time-tai->date time [tz-offset]

[Function] time-tai->julian-day time
[Function] time-tai->modified-julian-day time
[Function] time-tai->time-monotonic time
[Function] time-tai->time-monotonic! time
[Function] time-tai->time-utc time
[Function] time-tai->time-utc! time

[Function] time-utc->date time [tz-offset]

[Function] time-utc->julian-day time
[Function] time-utc->modified-julian-day time
[Function] time-utc->time-monotonic time
[Function] time-utc->time-monotonic! time
[Function] time-utc->time-tai time
[Function] time-utc->time-tai! time

Convert between dates, times and days of the respective types. For instance timetai->time-utc accepts a time object of type time-tai and returns an object of type

     time-utc.

The ! variants may modify their time argument to form their return. The plain functions create a new object.

For conversions to dates, tz-offset is seconds east of Greenwich. The default is the local timezone, at the given time, as provided by the system, using localtime (see Section 7.2.5 [Time], page 485).

On 32-bit systems, localtime is limited to a 32-bit time_t, so a default tz-offset is only available for times between Dec 1901 and Jan 2038. For prior dates an application might like to use the value in 1902, though some locations have zone changes prior to that. For future dates an application might like to assume today’s rules extend indefinitely. But for correct daylight savings transitions it will be necessary to take an offset for the same day and time but a year in range and which has the same starting weekday and same leap/non-leap (to support rules like last Sunday in October).

7.5.16.5 SRFI-19 Date to string

[Function] date->string date [format]

Convert a date to a string under the control of a format. format should be a string containing ‘~’ escapes, which will be expanded as per the following conversion table. The default format is ‘~c’, a locale-dependent date and time.

Many of these conversion characters are the same as POSIX strftime (see Section 7.2.5 [Time], page 485), but there are some extras and some variations.

~~  literal ~
~a  locale abbreviated weekday, eg. ‘Sun’
~A  locale full weekday, eg. ‘Sunday’
~b  locale abbreviated month, eg. ‘Jan’
~B  locale full month, eg. ‘January’

~c

locale date and time, eg.

          ‘Fri Jul 14 20:28:42-0400 2000’

~d  day of month, zero padded, ‘01’ to ‘31’
~e  day of month, blank padded, ‘ 1’ to ‘31’
~f  seconds and fractional seconds, with locale decimal point, eg. ‘5.2’
~h  same as ~b
~H  hour, 24-hour clock, zero padded, ‘00’ to ‘23’
~I  hour, 12-hour clock, zero padded, ‘01’ to ‘12’
~j  day of year, zero padded, ‘001’ to ‘366’
~k  hour, 24-hour clock, blank padded, ‘ 0’ to ‘23’
~l  hour, 12-hour clock, blank padded, ‘ 1’ to ‘12’
~m  month, zero padded, ‘01’ to ‘12’
~M  minute, zero padded, ‘00’ to ‘59’
~n  newline
~N  nanosecond, zero padded, ‘000000000’ to ‘999999999’
~p  locale AM or PM
~r  time, 12 hour clock, ‘~I:~M:~S ~p’
~s  number of full seconds since “the epoch” in UTC

~S

second, zero padded ‘00’ to ‘60’

(usual limit is 59, 60 is a leap second)

~t  horizontal tab character

~T  time, 24 hour clock, ‘~H:~M:~S’
~U  week of year, Sunday first day of week, ‘00’ to ‘52’
~V  week of year, Monday first day of week, ‘01’ to ‘53’
~w  day of week, 0 for Sunday, ‘0’ to ‘6’
~W  week of year, Monday first day of week, ‘00’ to ‘52’
~y  year, two digits, ‘00’ to ‘99’
~Y  year, full, eg. ‘2003’
~z  time zone, RFC-822 style
~Z  time zone symbol (not currently implemented)
~1  ISO-8601 date, ‘~Y-~m-~d’
~2  ISO-8601 time+zone, ‘~H:~M:~S~z’
~3  ISO-8601 time, ‘~H:~M:~S’
~4  ISO-8601 date/time+zone, ‘~Y-~m-~dT~H:~M:~S~z’
~5  ISO-8601 date/time, ‘~Y-~m-~dT~H:~M:~S’

Conversions ‘~D’, ‘~x’ and ‘~X’ are not currently described here, since the specification and reference implementation differ.

Conversion is locale-dependent on systems that support it (see Section 6.24.5 [Accessing Locale Information], page 439). See Section 7.2.13 [Locales], page 518, for information on how to change the current locale.

7.5.16.6 SRFI-19 String to date

[Function] string->date input template Convert an input string to a date under the control of a template string. Return a newly created date object. Literal characters in template must match characters in input and ‘~’ escapes must match the input forms described in the table below. “Skip to” means characters up to one of the given type are ignored, or “no skip” for no skipping. “Read” is what’s then read, and “Set” is the field affected in the date object. For example ‘~Y’ skips input characters until a digit is reached, at which point it expects a year and stores that to the year field of the date.

Skip to  Read  Set
~~  no skip  literal ~  nothing
~a        char-alphabetic? locale abbreviated weekday name  nothing
~A        char-alphabetic? locale full weekday name  nothing

~b        char-alphabetic? locale abbreviated month name  date-month
~B        char-alphabetic? locale full month name  date-month
~d        char-numeric?  day of month  date-day

~e  no skip  day of month, blank padded  date-day
~h  same as ‘~b’

~H        char-numeric?  hour  date-hour

~k  no skip  hour, blank padded  date-hour

~m        char-numeric?  month  date-month
~M        char-numeric?  minute  date-minute
~S        char-numeric?  second  date-second

~y

no skip 2-digit year

                                                                           date-year within

50 years

~Y        char-numeric?  year  date-year

~z  no skip  time zone  date-zone-offset

Notice that the weekday matching forms don’t affect the date object returned, instead the weekday will be derived from the day, month and year. Conversion is locale-dependent on systems that support it (see Section 6.24.5 [Accessing Locale Information], page 439). See Section 7.2.13 [Locales], page 518, for information on how to change the current locale.

7.5.17 SRFI-23 - Error Reporting

The SRFI-23 error procedure is always available.

7.5.18 SRFI-26 - specializing parameters

This SRFI provides a syntax for conveniently specializing selected parameters of a function. It can be used with,

(use-modules (srfi srfi-26))

[library syntax] cut slot1 slot2 … [library syntax] cute slot1 slot2 …

Return a new procedure which will make a call (slot1 slot2 …) but with selected parameters specialized to given expressions. An example will illustrate the idea. The following is a specialization of write, sending

output to my-output-port,
     (cut write <> my-output-port)

⇒

     (lambda (obj) (write obj my-output-port))

The special symbol <> indicates a slot to be filled by an argument to the new procedure. my-output-port on the other hand is an expression to be evaluated and passed, ie. it specializes the behaviour of write.

<> A slot to be filled by an argument from the created procedure. Arguments are assigned to <> slots in the order they appear in the cut form, there’s no way to re-arrange arguments. The first argument to cut is usually a procedure (or expression giving a procedure), but <> is allowed there too. For example,

               (cut <> 1 2 3)

⇒

               (lambda (proc) (proc 1 2 3))

<...> A slot to be filled by all remaining arguments from the new procedure. This can only occur at the end of a cut form. For example, a procedure taking a variable number of arguments like max but in addition enforcing a lower bound,

               (define my-lower-bound 123)
               (cut max my-lower-bound <...>)

⇒

               (lambda arglist (apply max my-lower-bound arglist))

For cut the specializing expressions are evaluated each time the new procedure is called. For cute they’re evaluated just once, when the new procedure is created. The name cute stands for “cut with evaluated arguments”. In all cases the evaluations take place in an unspecified order.

The following illustrates the difference between cut and cute,

     (cut format <> "the time is ~s" (current-time))

⇒

     (lambda (port) (format port "the time is ~s" (current-time)))
     (cute format <> "the time is ~s" (current-time))

⇒

     (let ((val (current-time)))
       (lambda (port) (format port "the time is ~s" val))

(There’s no provision for a mixture of cut and cute where some expressions would be evaluated every time but others evaluated only once.)

cut is really just a shorthand for the sort of lambda forms shown in the above examples. But notice cut avoids the need to name unspecialized parameters, and is more compact. Use in functional programming style or just with map, for-each or similar is typical.

     (map (cut * 2 <>) '(1 2 3 4))
     (for-each (cut write <> my-port) my-list)

7.5.19 SRFI-27 - Sources of Random Bits

This subsection is based on the specification of SRFI-27 (http://srfi.schemers.org/ srfi-27/srfi-27.html) written by Sebastian Egner.

This SRFI provides access to a (pseudo) random number generator; for Guile’s builtin random number facilities, which SRFI-27 is implemented upon, See Section 6.6.2.14 [Random], page 127. With SRFI-27, random numbers are obtained from a random source, which encapsulates a random number generation algorithm and its state.

7.5.19.1 The Default Random Source

[Function] random-integer n

Return a random number between zero (inclusive) and n (exclusive), using the default random source. The numbers returned have a uniform distribution.

[Function] random-real

Return a random number in (0,1), using the default random source. The numbers returned have a uniform distribution.

[Function] default-random-source

A random source from which random-integer and random-real have been

derived using random-source-make-integers and random-source-make-reals

(see Section 7.5.19.3 [SRFI-27 Random Number Generators], page 593 for those procedures). Note that an assignment to default-random-source does not change random-integer or random-real; it is also strongly recommended not to assign a new value.

7.5.19.2 Random Sources

[Function] make-random-source

Create a new random source. The stream of random numbers obtained from each random source created by this procedure will be identical, unless its state is changed by one of the procedures below.

[Function] random-source? object

Tests whether object is a random source. Random sources are a disjoint type.

[Function] random-source-randomize! source

Attempt to set the state of the random source to a truly random value. The current implementation uses a seed based on the current system time.

[Function] random-source-pseudo-randomize! source i j

Changes the state of the random source s into the initial state of the (i, j)-th independent random source, where i and j are non-negative integers. This procedure provides a mechanism to obtain a large number of independent random sources (usually all derived from the same backbone generator), indexed by two integers. In contrast to random-source-randomize!, this procedure is entirely deterministic.

The state associated with a random state can be obtained an reinstated with the following procedures:

[Function] random-source-state-ref source
[Function] random-source-state-set! source state

Get and set the state of a random source. No assumptions should be made about the nature of the state object, besides it having an external representation (i.e. it can be passed to write and subsequently read back).

7.5.19.3 Obtaining random number generator procedures

[Function] random-source-make-integers source

Obtains a procedure to generate random integers using the random source source. The returned procedure takes a single argument n, which must be a positive integer, and returns the next uniformly distributed random integer from the interval {0, ..., n-1} by advancing the state of source. If an application obtains and uses several generators for the same random source source, a call to any of these generators advances the state of source. Hence, the generators do not produce the same sequence of random integers each but rather share a state. This also holds for all other types of generators derived from a fixed random sources. While the SRFI text specifies that “Implementations that support concurrency make sure that the state of a generator is properly advanced”, this is currently not the case in Guile’s implementation of SRFI-27, as it would cause a severe performance penalty. So in multi-threaded programs, you either must perform locking on random sources shared between threads yourself, or use different random sources for multiple threads.

[Function] random-source-make-reals source
[Function] random-source-make-reals source unit

Obtains a procedure to generate random real numbers 0 < x < 1 using the random source source. The procedure rand is called without arguments. The optional parameter unit determines the type of numbers being produced by the returned procedure and the quantization of the output. unit must be a number such that 0 < unit < 1. The numbers created by the returned procedure are of the same numerical type as unit and the potential output values are spaced by at most unit. One can imagine rand to create numbers as x * unit where x is a random integer in {1, ..., floor(1/unit)-1}. Note, however, that this need not be the way the values are actually created and that the actual resolution of rand can be much higher than unit. In case unit is absent it defaults to a reasonably small value (related to the width of the mantissa of an efficient number format).

7.5.20 SRFI-30 - Nested Multi-line Comments

Starting from version 2.0, Guile’s read supports SRFI-30/R6RS nested multi-line comments by default, Section 6.17.1.3 [Block Comments], page 359.

7.5.21 SRFI-31 - A special form ‘rec’ for recursive evaluation

SRFI-31 defines a special form that can be used to create self-referential expressions more conveniently. The syntax is as follows:

<rec expression> --> (rec <variable> <expression>)
<rec expression> --> (rec (<variable>+) <body>)

The first syntax can be used to create self-referential expressions, for example:

  guile> (define tmp (rec ones (cons 1 (delay ones))))

The second syntax can be used to create anonymous recursive functions:

  guile> (define tmp (rec (display-n item n)
                       (if (positive? n)
                           (begin (display n) (display-n (- n 1))))))
  guile> (tmp 42 3)
  424242
  guile>

7.5.22 SRFI-34 - Exception handling for programs

Guile provides an implementation of SRFI-34’s exception handling mechanisms (http://

srfi.schemers.org/srfi-34/srfi-34.html) as an alternative to its own built-in mecha-

nisms (see Section 6.13.8 [Exceptions], page 303). It can be made available as follows:

(use-modules (srfi srfi-34))

7.5.23 SRFI-35 - Conditions

SRFI-35 (http://srfi.schemers.org/srfi-35/srfi-35.html) implements conditions, a

data structure akin to records designed to convey information about exceptional conditions between parts of a program. It is normally used in conjunction with SRFI-34’s raise:

(raise (condition (&message
                    (message "An error occurred"))))

Users can define condition types containing arbitrary information. Condition types may inherit from one another. This allows the part of the program that handles (or “catches”) conditions to get accurate information about the exceptional condition that arose.

SRFI-35 conditions are made available using:

(use-modules (srfi srfi-35))

The procedures available to manipulate condition types are the following:

[Scheme Procedure] make-condition-type id parent field-names

Return a new condition type named id, inheriting from parent, and with the fields whose names are listed in field-names. field-names must be a list of symbols and must not contain names already used by parent or one of its supertypes.

[Scheme Procedure] condition-type? obj

Return true if obj is a condition type.

Conditions can be created and accessed with the following procedures:

[Scheme Procedure] make-condition type . field+value

Return a new condition of type type with fields initialized as specified by field+value, a sequence of field names (symbols) and values as in the following example:

     (let ((&ct (make-condition-type 'foo &condition '(a b c))))
       (make-condition &ct 'a 1 'b 2 'c 3))

Note that all fields of type and its supertypes must be specified.

[Scheme Procedure] make-compound-condition condition1 condition2 …

Return a new compound condition composed of condition1 condition2 . . .. The returned condition has the type of each condition of condition1 condition2 … (per

condition-has-type?).

[Scheme Procedure] condition-has-type? c type

Return true if condition c has type type.

[Scheme Procedure] condition-ref c field-name

Return the value of the field named field-name from condition c. If c is a compound condition and several underlying condition types contain a field named field-name, then the value of the first such field is returned, using the order in which conditions were passed to make-compound-condition.

[Scheme Procedure] extract-condition c type

Return a condition of condition type type with the field values specified by c. If c is a compound condition, extract the field values from the subcondition belonging to type that appeared first in the call to make-compound-condition that created the condition.

Convenience macros are also available to create condition types and conditions.

[library syntax] define-condition-type type supertype predicate field-spec...

Define a new condition type named type that inherits from supertype. In addition, bind predicate to a type predicate that returns true when passed a condition of type type or any of its subtypes. field-spec must have the form (field accessor) where field is the name of field of type and accessor is the name of a procedure to access field field in conditions of type type. The example below defines condition type &foo, inheriting from &condition with fields a, b and c:

     (define-condition-type &foo &condition
       foo-condition?
       (a  foo-a)
       (b  foo-b)
       (c  foo-c))

[library syntax] condition type-field-binding1 type-field-binding2 … Return a new condition or compound condition, initialized according to type-field-binding1 type-field-binding2 . . .. Each type-field-binding must have the form (type field-specs...), where type is the name of a variable bound to a condition type; each field-spec must have the form (field-name value) where field-name is a symbol denoting the field being initialized to value. As for make-condition, all fields must be specified. The following example returns a simple condition:

     (condition (&message (message "An error occurred")))

The one below returns a compound condition:

     (condition (&message (message "An error occurred"))
                (&serious))

Finally, SRFI-35 defines a several standard condition types.

[Variable] &condition

This condition type is the root of all condition types. It has no fields.

[Variable] &message

A condition type that carries a message describing the nature of the condition to humans.

[Scheme Procedure] message-condition? c

Return true if c is of type &message or one of its subtypes.

[Scheme Procedure] condition-message c

Return the message associated with message condition c.

[Variable] &serious

This type describes conditions serious enough that they cannot safely be ignored. It has no fields.

[Scheme Procedure] serious-condition? c

Return true if c is of type &serious or one of its subtypes.

[Variable] &error

This condition describes errors, typically caused by something that has gone wrong in the interaction of the program with the external world or the user.

[Scheme Procedure] error? c

Return true if c is of type &error or one of its subtypes.

7.5.24 SRFI-37 - args-fold

This is a processor for GNU getopt_long-style program arguments. It provides an alternative, less declarative interface than getopt-long in (ice-9 getopt-long) (see Section 7.4 [The (ice-9 getopt-long) Module], page 546). Unlike getopt-long, it supports repeated options and any number of short and long names per option. Access it with:

(use-modules (srfi srfi-37))

SRFI-37 principally provides an option type and the args-fold function. To use the library, create a set of options with option and use it as a specification for invoking args-

fold.

Here is an example of a simple argument processor for the typical ‘--version’ and ‘--help’ options, which returns a backwards list of files given on the command line:

(args-fold (cdr (program-arguments))
           (let ((display-and-exit-proc
                  (lambda (msg)
                    (lambda (opt name arg loads)
                      (display msg) (quit)))))
             (list (option '(#\v "version") #f #f
                           (display-and-exit-proc "Foo version 42.0\n"))
                   (option '(#\h "help") #f #f
                           (display-and-exit-proc
                            "Usage: foo scheme-file ..."))))
           (lambda (opt name arg loads)
             (error "Unrecognized option `~A'" name))
           (lambda (op loads) (cons op loads))
           '())

[Scheme Procedure] option names required-arg? optional-arg? processor Return an object that specifies a single kind of program option. names is a list of command-line option names, and should consist of characters for traditional getopt short options and strings for getopt_long-style long options. required-arg? and optional-arg? are mutually exclusive; one or both must be #f. If required-arg?, the option must be followed by an argument on the command line, such as ‘--opt=value’ for long options, or an error will be signalled. If optional-arg?, an argument will be taken if available. processor is a procedure that takes at least 3 arguments, called when args-fold encounters the option: the containing option object, the name used on the command line, and the argument given for the option (or #f if none). The rest of the arguments are args-fold “seeds”, and the processor should return seeds as well.

[Scheme Procedure] option-names opt

[Scheme Procedure] option-required-arg? opt
[Scheme Procedure] option-optional-arg? opt

[Scheme Procedure] option-processor opt

Return the specified field of opt, an option object, as described above for option.

[Scheme Procedure] args-fold args options unrecognized-option-proc operand-proc

seed … Process args, a list of program arguments such as that returned by (cdr (programarguments)), in order against options, a list of option objects as described above. All functions called take the “seeds”, or the last multiple-values as multiple arguments, starting with seed …, and must return the new seeds. Return the final seeds. Call unrecognized-option-proc, which is like an option object’s processor, for any options not found in options. Call operand-proc with any items on the command line that are not named options. This includes arguments after ‘--’. It is called with the argument in question, as well as the seeds.

7.5.25 SRFI-38 - External Representation for Data With Shared Structure

This subsection is based on the specification of SRFI-38 (http://srfi.schemers.org/ srfi-38/srfi-38.html) written by Ray Dillinger.

This SRFI creates an alternative external representation for data written and read using

write-with-shared-structure and read-with-shared-structure. It is identical to the

grammar for external representation for data written and read with write and read given in section 7 of R5RS, except that the single production

<datum> --> <simple datum> | <compound datum>

is replaced by the following five productions:

<datum> --> <defining datum> | <nondefining datum> | <defined datum>
<defining datum> -->  #<indexnum>=<nondefining datum>
<defined datum> --> #<indexnum>#
<nondefining datum> --> <simple datum> | <compound datum>
<indexnum> --> <digit 10>+

[Scheme procedure] write-with-shared-structure obj
[Scheme procedure] write-with-shared-structure obj port

[Scheme procedure] write-with-shared-structure obj port optarg

Writes an external representation of obj to the given port. Strings that appear in the written representation are enclosed in doublequotes, and within those strings backslash and doublequote characters are escaped by backslashes. Character objects are written using the #\ notation.

Objects which denote locations rather than values (cons cells, vectors, and non-zero-length strings in R5RS scheme; also Guile’s structs, bytevectors and ports and hashtables), if they appear at more than one point in the data being written, are preceded by ‘#N=’ the first time they are written and replaced by ‘#N#’ all subsequent times they are written, where N is a natural number used to identify that particular object. If objects which denote locations occur only once in the structure, then write-with-shared-structure must produce the same external representation for those objects

as write.

write-with-shared-structure terminates in finite time and produces a finite representation when writing finite data.

write-with-shared-structure returns an unspecified value. The port argument may be omitted, in which case it defaults to the value returned by (current-output-port). The optarg argument may also be omitted. If present, its effects on the output and return value are unspecified but write-with-shared-structure must still write a representation that can be read by read-with-shared-structure. Some implementations may wish to use optarg to specify formatting conventions, numeric radixes, or return values. Guile’s implementation ignores optarg.

For example, the code

     (begin (define a (cons 'val1 'val2))
            (set-cdr! a a)
            (write-with-shared-structure a))

should produce the output #1=(val1 . #1#). This shows a cons cell whose cdr contains itself.

[Scheme procedure] read-with-shared-structure
[Scheme procedure] read-with-shared-structure port

read-with-shared-structure converts the external representations of Scheme objects produced by write-with-shared-structure into Scheme objects. That is, it is a parser for the nonterminal ‘<datum>’ in the augmented external representation grammar defined above. read-with-shared-structure returns the next object parsable from the given input port, updating port to point to the first character past the end of the external representation of the object.

If an end-of-file is encountered in the input before any characters are found that can begin an object, then an end-of-file object is returned. The port remains open, and further attempts to read it (by read-with-shared-structure or read will also return an end-of-file object. If an end of file is encountered after the beginning of an object’s external representation, but the external representation is incomplete and therefore not parsable, an error is signalled.

The port argument may be omitted, in which case it defaults to the value returned by (current-input-port). It is an error to read from a closed port.

7.5.26 SRFI-39 - Parameters

This SRFI adds support for dynamically-scoped parameters. SRFI 39 is implemented in the Guile core; there’s no module needed to get SRFI-39 itself. Parameters are documented in Section 6.21.8 [Parameters], page 422.

This module does export one extra function: with-parameters*. This is a Guile-specific addition to the SRFI, similar to the core with-fluids* (see Section 6.21.7 [Fluids and Dynamic States], page 419).

[Function] with-parameters* param-list value-list thunk Establish a new dynamic scope, as per parameterize above, taking parameters from param-list and corresponding values from value-list. A call (thunk) is made in the new scope and the result from that thunk is the return from with-parameters*.

7.5.27 SRFI-41 - Streams

This subsection is based on the specification of SRFI-41 (http://srfi.schemers.org/

srfi-41/srfi-41.html) by Philip L. Bewig.

This SRFI implements streams, sometimes called lazy lists, a sequential data structure containing elements computed only on demand. A stream is either null or is a pair with a stream in its cdr. Since elements of a stream are computed only when accessed, streams can be infinite. Once computed, the value of a stream element is cached in case it is needed again. SRFI-41 can be made available with:

     (use-modules (srfi srfi-41))

7.5.27.1 SRFI-41 Stream Fundamentals

SRFI-41 Streams are based on two mutually-recursive abstract data types: An object of the stream abstract data type is a promise that, when forced, is either stream-null or is an object of type stream-pair. An object of the stream-pair abstract data type contains a stream-car and a stream-cdr, which must be a stream. The essential feature of streams is the systematic suspensions of the recursive promises between the two data types.

The object stored in the stream-car of a stream-pair is a promise that is forced the first time the stream-car is accessed; its value is cached in case it is needed again. The object may have any type, and different stream elements may have different types. If the stream-car is never accessed, the object stored there is never evaluated. Likewise, the stream-cdr is a promise to return a stream, and is only forced on demand.

7.5.27.2 SRFI-41 Stream Primitives

This library provides eight operators: constructors for stream-null and stream-pairs, type predicates for streams and the two kinds of streams, accessors for both fields of a stream-pair, and a lambda that creates procedures that return streams.

[Scheme Variable] stream-null

A promise that, when forced, is a single object, distinguishable from all other objects, that represents the null stream. stream-null is immutable and unique.

[Scheme Syntax] stream-cons object-expr stream-expr Creates a newly-allocated stream containing a promise that, when forced, is a streampair with object-expr in its stream-car and stream-expr in its stream-cdr. Neither object-expr nor stream-expr is evaluated when stream-cons is called. Once created, a stream-pair is immutable; there is no stream-set-car! or streamset-cdr! that modifies an existing stream-pair. There is no dotted-pair or improper stream as with lists.

[Scheme Procedure] stream? object

Returns true if object is a stream, otherwise returns false. If object is a stream, its promise will not be forced. If (stream? obj) returns true, then one of (stream-null? obj) or (stream-pair? obj) will return true and the other will return false.

[Scheme Procedure] stream-null? object

Returns true if object is the distinguished null stream, otherwise returns false. If object is a stream, its promise will be forced.

[Scheme Procedure] stream-pair? object

Returns true if object is a stream-pair constructed by stream-cons, otherwise returns false. If object is a stream, its promise will be forced.

[Scheme Procedure] stream-car stream

Returns the object stored in the stream-car of stream. An error is signalled if the argument is not a stream-pair. This causes the object-expr passed to stream-cons to be evaluated if it had not yet been; the value is cached in case it is needed again.

[Scheme Procedure] stream-cdr stream

Returns the stream stored in the stream-cdr of stream. An error is signalled if the argument is not a stream-pair.

[Scheme Syntax] stream-lambda formals body …

Creates a procedure that returns a promise to evaluate the body of the procedure. The last body expression to be evaluated must yield a stream. As with normal lambda, formals may be a single variable name, in which case all the formal arguments are collected into a single list, or a list of variable names, which may be null if there are no arguments, proper if there are an exact number of arguments, or dotted if a fixed number of arguments is to be followed by zero or more arguments collected into a list. Body must contain at least one expression, and may contain internal definitions preceding any expressions to be evaluated.

(define strm123
  (stream-cons 1
    (stream-cons 2
      (stream-cons 3
        stream-null))))
(stream-car strm123) ⇒ 1
(stream-car (stream-cdr strm123) ⇒ 2

(stream-pair?
  (stream-cdr
    (stream-cons (/ 1 0) stream-null))) ⇒ #f
(stream? (list 1 2 3)) ⇒ #f
(define iter
  (stream-lambda (f x)
    (stream-cons x (iter f (f x)))))
(define nats (iter (lambda (x) (+ x 1)) 0))
(stream-car (stream-cdr nats)) ⇒ 1
(define stream-add
  (stream-lambda (s1 s2)
    (stream-cons
      (+ (stream-car s1) (stream-car s2))
      (stream-add (stream-cdr s1)
                  (stream-cdr s2)))))
(define evens (stream-add nats nats))
(stream-car evens) ⇒ 0
(stream-car (stream-cdr evens)) ⇒ 2
(stream-car (stream-cdr (stream-cdr evens))) ⇒ 4

7.5.27.3 SRFI-41 Stream Library

[Scheme Syntax] define-stream (name args …) body … Creates a procedure that returns a stream, and may appear anywhere a normal define may appear, including as an internal definition. It may contain internal definitions of its own. The defined procedure takes arguments in the same way as stream-lambda.

define-stream is syntactic sugar on stream-lambda; see also stream-let, which is

also a sugaring of stream-lambda.

A simple version of stream-map that takes only a single input stream calls itself recursively:

     (define-stream (stream-map proc strm)
       (if (stream-null? strm)
           stream-null
           (stream-cons
             (proc (stream-car strm))
             (stream-map proc (stream-cdr strm))))))

[Scheme Procedure] list->stream list

Returns a newly-allocated stream containing the elements from list.

[Scheme Procedure] port->stream [port]

Returns a newly-allocated stream containing in its elements the characters on the port. If port is not given it defaults to the current input port. The returned stream has finite length and is terminated by stream-null. It looks like one use of port->stream would be this:

     (define s ;wrong!
       (with-input-from-file filename
         (lambda () (port->stream))))

But that fails, because with-input-from-file is eager, and closes the input port prematurely, before the first character is read. To read a file into a stream, say:

     (define-stream (file->stream filename)
       (let ((p (open-input-file filename)))
         (stream-let loop ((c (read-char p)))
           (if (eof-object? c)
               (begin (close-input-port p)
                      stream-null)
               (stream-cons c
                 (loop (read-char p)))))))

[Scheme Syntax] stream object-expr … Creates a newly-allocated stream containing in its elements the objects, in order. The object-exprs are evaluated when they are accessed, not when the stream is created. If no objects are given, as in (stream), the null stream is returned. See also list-

>stream.
     (define strm123 (stream 1 2 3))
     ; (/ 1 0) not evaluated when stream is created
     (define s (stream 1 (/ 1 0) -1))

[Scheme Procedure] stream->list [n] stream

Returns a newly-allocated list containing in its elements the first n items in stream. If stream has less than n items, all the items in the stream will be included in the returned list. If n is not given it defaults to infinity, which means that unless stream is finite stream->list will never return.

     (stream->list 10
       (stream-map (lambda (x) (* x x))
         (stream-from 0)))
       ⇒ (0 1 4 9 16 25 36 49 64 81)

[Scheme Procedure] stream-append stream …

Returns a newly-allocated stream containing in its elements those elements contained in its input streams, in order of input. If any of the input streams is infinite, no elements of any of the succeeding input streams will appear in the output stream.

See also stream-concat.

[Scheme Procedure] stream-concat stream

Takes a stream consisting of one or more streams and returns a newly-allocated stream containing all the elements of the input streams. If any of the streams in the input stream is infinite, any remaining streams in the input stream will never appear in the output stream. See also stream-append.

[Scheme Procedure] stream-constant object …

Returns a newly-allocated stream containing in its elements the objects, repeating in succession forever.

     (stream-constant 1) ⇒ 1 1 1 ...
     (stream-constant #t #f) ⇒ #t #f #t #f #t #f ...

[Scheme Procedure] stream-drop n stream

Returns the suffix of the input stream that starts at the next element after the first n elements. The output stream shares structure with the input stream; thus, promises forced in one instance of the stream are also forced in the other instance of the stream. If the input stream has less than n elements, stream-drop returns the null stream.

See also stream-take.

[Scheme Procedure] stream-drop-while pred stream

Returns the suffix of the input stream that starts at the first element x for which (pred x) returns false. The output stream shares structure with the input stream.

See also stream-take-while.

[Scheme Procedure] stream-filter pred stream

Returns a newly-allocated stream that contains only those elements x of the input stream which satisfy the predicate pred.

     (stream-filter odd? (stream-from 0))
        ⇒ 1 3 5 7 9 ...

[Scheme Procedure] stream-fold proc base stream Applies a binary procedure proc to base and the first element of stream to compute a new base, then applies the procedure to the new base and the next element of stream to compute a succeeding base, and so on, accumulating a value that is finally returned as the value of stream-fold when the end of the stream is reached. stream must be finite, or stream-fold will enter an infinite loop. See also stream-scan, which is similar to stream-fold, but useful for infinite streams. For readers familiar with other functional languages, this is a left-fold; there is no corresponding right-fold, since right-fold relies on finite streams that are fully-evaluated, in which case they may as well be converted to a list.

[Scheme Procedure] stream-for-each proc stream …

Applies proc element-wise to corresponding elements of the input streams for sideeffects; it returns nothing. stream-for-each stops as soon as any of its input streams is exhausted.

[Scheme Procedure] stream-from first [step]

Creates a newly-allocated stream that contains first as its first element and increments each succeeding element by step. If step is not given it defaults to 1. first and step may be of any numeric type. stream-from is frequently useful as a generator in stream-of expressions. See also stream-range for a similar procedure that creates finite streams.

[Scheme Procedure] stream-iterate proc base

Creates a newly-allocated stream containing base in its first element and applies proc to each element in turn to determine the succeeding element. See also stream-unfold

and stream-unfolds.

[Scheme Procedure] stream-length stream

Returns the number of elements in the stream; it does not evaluate its elements. stream-length may only be used on finite streams; it enters an infinite loop with infinite streams.

[Scheme Syntax] stream-let tag ((var expr) …) body … Creates a local scope that binds each variable to the value of its corresponding expression. It additionally binds tag to a procedure which takes the bound variables as arguments and body as its defining expressions, binding the tag with streamlambda. tag is in scope within body, and may be called recursively. When the expanded expression defined by the stream-let is evaluated, stream-let evaluates the expressions in its body in an environment containing the newly-bound variables, returning the value of the last expression evaluated, which must yield a stream. stream-let provides syntactic sugar on stream-lambda, in the same manner as normal let provides syntactic sugar on normal lambda. However, unlike normal let, the tag is required, not optional, because unnamed stream-let is meaningless. For example, stream-member returns the first stream-pair of the input strm with a stream-car x that satisfies (eql? obj x), or the null stream if x is not present in strm.

     (define-stream (stream-member eql? obj strm)
       (stream-let loop ((strm strm))
         (cond ((stream-null? strm) strm)
               ((eql? obj (stream-car strm)) strm)
               (else (loop (stream-cdr strm))))))

[Scheme Procedure] stream-map proc stream …

Applies proc element-wise to corresponding elements of the input streams, returning a newly-allocated stream containing elements that are the results of those procedure applications. The output stream has as many elements as the minimum-length input stream, and may be infinite.

[Scheme Syntax] stream-match stream clause …

Provides pattern-matching for streams. The input stream is an expression that evaluates to a stream. Clauses are of the form (pattern [fender] expression), consisting of a pattern that matches a stream of a particular shape, an optional fender that must succeed if the pattern is to match, and an expression that is evaluated if the pattern matches. There are four types of patterns:

• () matches the null stream.

• (pat0 pat1 …) matches a finite stream with length exactly equal to the number of pattern elements.

• (pat0 pat1 …. pat-rest) matches an infinite stream, or a finite stream with length at least as great as the number of pattern elements before the literal dot.

• pat matches an entire stream. Should always appear last in the list of clauses; it’s not an error to appear elsewhere, but subsequent clauses could never match.

Each pattern element may be either:

• An identifier, which matches any stream element. Additionally, the value of the stream element is bound to the variable named by the identifier, which is in scope in the fender and expression of the corresponding clause. Each identifier in a single pattern must be unique.

• A literal underscore (_), which matches any stream element but creates no bindings.

The patterns are tested in order, left-to-right, until a matching pattern is found; if fender is present, it must evaluate to a true value for the match to be successful. Pattern variables are bound in the corresponding fender and expression. Once the matching pattern is found, the corresponding expression is evaluated and returned as the result of the match. An error is signaled if no pattern matches the input stream.

stream-match is often used to distinguish null streams from non-null streams, binding head and tail:

     (define (len strm)
       (stream-match strm
         (() 0)
         ((head . tail) (+ 1 (len tail)))))

Fenders can test the common case where two stream elements must be identical; the else pattern is an identifier bound to the entire stream, not a keyword as in cond.

     (stream-match strm
       ((x y . _) (equal? x y) 'ok)
       (else 'error))

A more complex example uses two nested matchers to match two different stream arguments; (stream-merge lt? . strms) stably merges two or more streams ordered by the lt? predicate:

     (define-stream (stream-merge lt? . strms)
       (define-stream (merge xx yy)
         (stream-match xx (() yy) ((x . xs)
           (stream-match yy (() xx) ((y . ys)
             (if (lt? y x)
                 (stream-cons y (merge xx ys))
                 (stream-cons x (merge xs yy))))))))
       (stream-let loop ((strms strms))
         (cond ((null? strms) stream-null)
               ((null? (cdr strms)) (car strms))
               (else (merge (car strms)
                            (apply stream-merge lt?
                              (cdr strms)))))))

[Scheme Syntax] stream-of expr clause … Provides the syntax of stream comprehensions, which generate streams by means of looping expressions. The result is a stream of objects of the type returned by expr. There are four types of clauses:

• (var in stream-expr) loops over the elements of stream-expr, in order from the start of the stream, binding each element of the stream in turn to var. streamfrom and stream-range are frequently useful as generators for stream-expr.

• (var is expr) binds var to the value obtained by evaluating expr.

• (pred expr) includes in the output stream only those elements x which satisfy the predicate pred.

The scope of variables bound in the stream comprehension is the clauses to the right of the binding clause (but not the binding clause itself) plus the result expression. When two or more generators are present, the loops are processed as if they are nested from left to right; that is, the rightmost generator varies fastest. A consequence of this is that only the first generator may be infinite and all subsequent generators must be finite. If no generators are present, the result of a stream comprehension is a stream containing the result expression; thus, ‘(stream-of 1)’ produces a finite stream containing only the element 1.

     (stream-of (* x x)
       (x in (stream-range 0 10))
       (even? x))
       ⇒ 0 4 16 36 64
     (stream-of (list a b)
       (a in (stream-range 1 4))
       (b in (stream-range 1 3)))
       ⇒ (1 1) (1 2) (2 1) (2 2) (3 1) (3 2)
     (stream-of (list i j)
       (i in (stream-range 1 5))
       (j in (stream-range (+ i 1) 5)))
       ⇒ (1 2) (1 3) (1 4) (2 3) (2 4) (3 4)

[Scheme Procedure] stream-range first past [step]

Creates a newly-allocated stream that contains first as its first element and increments each succeeding element by step. The stream is finite and ends before past, which is not an element of the stream. If step is not given it defaults to 1 if first is less than past and -1 otherwise. first, past and step may be of any real numeric type. stream-range is frequently useful as a generator in stream-of expressions. See also stream-from for a similar procedure that creates infinite streams.

     (stream-range 0 10) ⇒ 0 1 2 3 4 5 6 7 8 9
     (stream-range 0 10 2) ⇒ 0 2 4 6 8

Successive elements of the stream are calculated by adding step to first, so if any of first, past or step are inexact, the length of the output stream may differ from

(ceiling (- (/ (- past first) step) 1).

[Scheme Procedure] stream-ref stream n

Returns the nth element of stream, counting from zero. An error is signaled if n is greater than or equal to the length of stream.

     (define (fact n)
       (stream-ref
         (stream-scan * 1 (stream-from 1))
         n))

[Scheme Procedure] stream-reverse stream

Returns a newly-allocated stream containing the elements of the input stream but in reverse order. stream-reverse may only be used with finite streams; it enters an infinite loop with infinite streams. stream-reverse does not force evaluation of the elements of the stream.

[Scheme Procedure] stream-scan proc base stream Accumulates the partial folds of an input stream into a newly-allocated output stream. The output stream is the base followed by (stream-fold proc base (stream-take i stream)) for each of the first i elements of stream.

     (stream-scan + 0 (stream-from 1))
       ⇒ (stream 0 1 3 6 10 15 ...)
     (stream-scan * 1 (stream-from 1))
       ⇒ (stream 1 1 2 6 24 120 ...)

[Scheme Procedure] stream-take n stream

Returns a newly-allocated stream containing the first n elements of the input stream. If the input stream has less than n elements, so does the output stream. See also

stream-drop.

[Scheme Procedure] stream-take-while pred stream

Takes a predicate and a stream and returns a newly-allocated stream containing those elements x that form the maximal prefix of the input stream which satisfy pred. See

also stream-drop-while.

[Scheme Procedure] stream-unfold map pred gen base The fundamental recursive stream constructor. It constructs a stream by repeatedly applying gen to successive values of base, in the manner of stream-iterate, then applying map to each of the values so generated, appending each of the mapped values to the output stream as long as (pred? base) returns a true value. See also

stream-iterate and stream-unfolds.

The expression below creates the finite stream ‘0 1 4 9 16 25 36 49 64 81’. Initially the base is 0, which is less than 10, so map squares the base and the mapped value becomes the first element of the output stream. Then gen increments the base by 1, so it becomes 1; this is less than 10, so map squares the new base and 1 becomes the second element of the output stream. And so on, until the base becomes 10, when pred stops the recursion and stream-null ends the output stream.

     (stream-unfold
       (lambda (x) (expt x 2)) ; map

       (lambda (x) (< x 10))   ; pred?
       (lambda (x) (+ x 1))    ; gen
       0)                      ; base

[Scheme Procedure] stream-unfolds proc seed

Returns n newly-allocated streams containing those elements produced by successive calls to the generator proc, which takes the current seed as its argument and returns n+1 values (proc seed) ⇒ seed result 0 … result n-1 where the returned seed is the input seed to the next call to the generator and result i indicates how to produce the next element of the ith result stream:

• (value): value is the next car of the result stream.

• #f: no value produced by this iteration of the generator proc for the result stream.

• (): the end of the result stream.

It may require multiple calls of proc to produce the next element of any particular

result stream. See also stream-iterate and stream-unfold.
     (define (stream-partition pred? strm)
       (stream-unfolds
         (lambda (s)
           (if (stream-null? s)
               (values s '() '())
               (let ((a (stream-car s))
                     (d (stream-cdr s)))
                 (if (pred? a)
                     (values d (list a) #f)
                     (values d #f (list a))))))
         strm))
     (call-with-values
       (lambda ()
         (stream-partition odd?
           (stream-range 1 6)))
       (lambda (odds evens)
         (list (stream->list odds)
               (stream->list evens))))
       ⇒ ((1 3 5) (2 4))

[Scheme Procedure] stream-zip stream …

Returns a newly-allocated stream in which each element is a list (not a stream) of the corresponding elements of the input streams. The output stream is as long as the shortest input stream, if any of the input streams is finite, or is infinite if all the input streams are infinite.

7.5.28 SRFI-42 - Eager Comprehensions

See the specification of SRFI-42 (http://srfi.schemers.org/srfi-42/srfi-42.html).

7.5.29 SRFI-43 - Vector Library

This subsection is based on the specification of SRFI-43 (http://srfi.schemers.org/

srfi-43/srfi-43.html) by Taylor Campbell.

SRFI-43 implements a comprehensive library of vector operations. It can be made available with:

(use-modules (srfi srfi-43))

7.5.29.1 SRFI-43 Constructors

[Scheme Procedure] make-vector size [fill]

Create and return a vector of size size, optionally filling it with fill. The default value of fill is unspecified.

     (make-vector 5 3) ⇒ #(3 3 3 3 3)

[Scheme Procedure] vector x …

Create and return a vector whose elements are x . . ..

     (vector 0 1 2 3 4) ⇒ #(0 1 2 3 4)

[Scheme Procedure] vector-unfold f length initial-seed … The fundamental vector constructor. Create a vector whose length is length and iterates across each index k from 0 up to length - 1, applying f at each iteration to the current index and current seeds, in that order, to receive n + 1 values: first, the element to put in the kth slot of the new vector and n new seeds for the next iteration. It is an error for the number of seeds to vary between iterations.

     (vector-unfold (lambda (i x) (values x (- x 1)))
                    10 0)
     ⇒ #(0 -1 -2 -3 -4 -5 -6 -7 -8 -9)
     (vector-unfold values 10)
     ⇒ #(0 1 2 3 4 5 6 7 8 9)

[Scheme Procedure] vector-unfold-right f length initial-seed …

Like vector-unfold, but it uses f to generate elements from right-to-left, rather than left-to-right.

     (vector-unfold-right (lambda (i x) (values x (+ x 1)))
                          10 0)
     ⇒ #(9 8 7 6 5 4 3 2 1 0)

[Scheme Procedure] vector-copy vec [start [end [fill]]]

Allocate a new vector whose length is end - start and fills it with elements from vec, taking elements from vec starting at index start and stopping at index end. start defaults to 0 and end defaults to the value of (vector-length vec). If end extends beyond the length of vec, the slots in the new vector that obviously cannot be filled by elements from vec are filled with fill, whose default value is unspecified.

     (vector-copy '#(a b c d e f g h i))
     ⇒ #(a b c d e f g h i)

     (vector-copy '#(a b c d e f g h i) 6)
     ⇒ #(g h i)
     (vector-copy '#(a b c d e f g h i) 3 6)
     ⇒ #(d e f)
     (vector-copy '#(a b c d e f g h i) 6 12 'x)
     ⇒ #(g h i x x x)

[Scheme Procedure] vector-reverse-copy vec [start [end]]

Like vector-copy, but it copies the elements in the reverse order from vec.

     (vector-reverse-copy '#(5 4 3 2 1 0) 1 5)
     ⇒ #(1 2 3 4)

[Scheme Procedure] vector-append vec …

Return a newly allocated vector that contains all elements in order from the subsequent locations in vec . . ..

     (vector-append '#(a) '#(b c d))
     ⇒ #(a b c d)

[Scheme Procedure] vector-concatenate list-of-vectors

Append each vector in list-of-vectors. Equivalent to (apply vector-append list-

of-vectors).
     (vector-concatenate '(#(a b) #(c d)))
     ⇒ #(a b c d)

7.5.29.2 SRFI-43 Predicates

[Scheme Procedure] vector? obj

Return true if obj is a vector, else return false.

[Scheme Procedure] vector-empty? vec

Return true if vec is empty, i.e. its length is 0, else return false.

[Scheme Procedure] vector= elt=? vec … Return true if the vectors vec … have equal lengths and equal elements according to elt=?. elt=? is always applied to two arguments. Element comparison must be consistent with eq? in the following sense: if (eq? a b) returns true, then (elt=? a b) must also return true. The order in which comparisons are performed is unspecified.

7.5.29.3 SRFI-43 Selectors

[Scheme Procedure] vector-ref vec i

Return the element at index i in vec. Indexing is based on zero.

[Scheme Procedure] vector-length vec

Return the length of vec.

7.5.29.4 SRFI-43 Iteration

[Scheme Procedure] vector-fold kons knil vec1 vec2 … The fundamental vector iterator. kons is iterated over each index in all of the vectors, stopping at the end of the shortest; kons is applied as

      (kons i state (vector-ref vec1 i) (vector-ref vec2 i) ...)

where state is the current state value, and i is the current index. The current state value begins with knil, and becomes whatever kons returned at the respective iteration. The iteration is strictly left-to-right.

[Scheme Procedure] vector-fold-right kons knil vec1 vec2 …

Similar to vector-fold, but it iterates right-to-left instead of left-to-right.

[Scheme Procedure] vector-map f vec1 vec2 …

Return a new vector of the shortest size of the vector arguments. Each element at index i of the new vector is mapped from the old vectors by

      (f i (vector-ref vec1 i) (vector-ref vec2 i) ...)

The dynamic order of application of f is unspecified.

[Scheme Procedure] vector-map! f vec1 vec2 …

Similar to vector-map, but rather than mapping the new elements into a new vector, the new mapped elements are destructively inserted into vec1. The dynamic order of application of f is unspecified.

[Scheme Procedure] vector-for-each f vec1 vec2 …

Call (f i (vector-ref vec1 i) (vector-ref vec2 i) ...) for each index i less than

the length of the shortest vector passed. The iteration is strictly left-to-right.

[Scheme Procedure] vector-count pred? vec1 vec2 … Count the number of parallel elements in the vectors that satisfy pred?, which is applied, for each index i less than the length of the smallest vector, to i and each parallel element in the vectors at that index, in order.

     (vector-count (lambda (i elt) (even? elt))
                   '#(3 1 4 1 5 9 2 5 6))
     ⇒ 3
     (vector-count (lambda (i x y) (< x y))
                   '#(1 3 6 9) '#(2 4 6 8 10 12))
     ⇒ 2

7.5.29.5 SRFI-43 Searching

[Scheme Procedure] vector-index pred? vec1 vec2 … Find and return the index of the first elements in vec1 vec2 … that satisfy pred?. If no matching element is found by the end of the shortest vector, return #f.

     (vector-index even? '#(3 1 4 1 5 9))
     ⇒ 2
     (vector-index < '#(3 1 4 1 5 9 2 5 6) '#(2 7 1 8 2))
     ⇒ 1
     (vector-index = '#(3 1 4 1 5 9 2 5 6) '#(2 7 1 8 2))
     ⇒ #f

[Scheme Procedure] vector-index-right pred? vec1 vec2 …

Like vector-index, but it searches right-to-left, rather than left-to-right. Note that the SRFI 43 specification requires that all the vectors must have the same length, but both the SRFI 43 reference implementation and Guile’s implementation allow vectors with unequal lengths, and start searching from the last index of the shortest vector.

[Scheme Procedure] vector-skip pred? vec1 vec2 … Find and return the index of the first elements in vec1 vec2 … that do not satisfy pred?. If no matching element is found by the end of the shortest vector, return #f. Equivalent to vector-index but with the predicate inverted.

     (vector-skip number? '#(1 2 a b 3 4 c d)) ⇒ 2

[Scheme Procedure] vector-skip-right pred? vec1 vec2 …

Like vector-skip, but it searches for a non-matching element right-to-left, rather than left-to-right. Note that the SRFI 43 specification requires that all the vectors must have the same length, but both the SRFI 43 reference implementation and Guile’s implementation allow vectors with unequal lengths, and start searching from the last index of the shortest vector.

[Scheme Procedure] vector-binary-search vec value cmp [start [end]]

Find and return an index of vec between start and end whose value is value using a binary search. If no matching element is found, return #f. The default start is 0 and the default end is the length of vec. cmp must be a procedure of two arguments such that (cmp a b) returns a negative integer if a < b, a positive integer if a > b, or zero if a = b. The elements of vec must be sorted in non-decreasing order according to cmp. Note that SRFI 43 does not document the start and end arguments, but both its reference implementation and Guile’s implementation support them.

     (define (char-cmp c1 c2)
       (cond ((char<? c1 c2) -1)
             ((char>? c1 c2) 1)
             (else 0)))
     (vector-binary-search '#(#\a #\b #\c #\d #\e #\f #\g #\h)
                           #\g
                           char-cmp)
     ⇒ 6

[Scheme Procedure] vector-any pred? vec1 vec2 … Find the first parallel set of elements from vec1 vec2 … for which pred? returns a true value. If such a parallel set of elements exists, vector-any returns the value that pred? returned for that set of elements. The iteration is strictly left-to-right.

[Scheme Procedure] vector-every pred? vec1 vec2 … If, for every index i between 0 and the length of the shortest vector argument, the set

of elements (vector-ref vec1 i)(vector-ref vec2 i) . . . satisfies pred?, vector-

every returns the value that pred? returned for the last set of elements, at the last index of the shortest vector. Otherwise it returns #f. The iteration is strictly left-to-right.

7.5.29.6 SRFI-43 Mutators

[Scheme Procedure] vector-set! vec i value

Assign the contents of the location at i in vec to value.

[Scheme Procedure] vector-swap! vec i j

Swap the values of the locations in vec at i and j.

[Scheme Procedure] vector-fill! vec fill [start [end]]

Assign the value of every location in vec between start and end to fill. start defaults to 0 and end defaults to the length of vec.

[Scheme Procedure] vector-reverse! vec [start [end]]

Destructively reverse the contents of vec between start and end. start defaults to 0 and end defaults to the length of vec.

[Scheme Procedure] vector-copy! target tstart source [sstart [send]] Copy a block of elements from source to target, both of which must be vectors, starting in target at tstart and starting in source at sstart, ending when (send - sstart) elements have been copied. It is an error for target to have a length less than (tstart + send - sstart). sstart defaults to 0 and send defaults to the length of source.

[Scheme Procedure] vector-reverse-copy! target tstart source [sstart [send]] Like vector-copy!, but this copies the elements in the reverse order. It is an error if target and source are identical vectors and the target and source ranges overlap;

however, if tstart = sstart, vector-reverse-copy! behaves as (vector-reverse!
target tstart send) would.

7.5.29.7 SRFI-43 Conversion

[Scheme Procedure] vector->list vec [start [end]]

Return a newly allocated list containing the elements in vec between start and end. start defaults to 0 and end defaults to the length of vec.

[Scheme Procedure] reverse-vector->list vec [start [end]]

Like vector->list, but the resulting list contains the specified range of elements of vec in reverse order.

[Scheme Procedure] list->vector proper-list [start [end]] Return a newly allocated vector of the elements from proper-list with indices between start and end. start defaults to 0 and end defaults to the length of proper-list. Note that SRFI 43 does not document the start and end arguments, but both its reference implementation and Guile’s implementation support them.

[Scheme Procedure] reverse-list->vector proper-list [start [end]]

Like list->vector, but the resulting vector contains the specified range of elements of proper-list in reverse order. Note that SRFI 43 does not document the start and end arguments, but both its reference implementation and Guile’s implementation support them.

7.5.30 SRFI-45 - Primitives for Expressing Iterative Lazy Algorithms

This subsection is based on the specification of SRFI-45 (http://srfi.schemers.org/ srfi-45/srfi-45.html) written by André van Tonder.

Lazy evaluation is traditionally simulated in Scheme using delay and force. However, these primitives are not powerful enough to express a large class of lazy algorithms that are iterative. Indeed, it is folklore in the Scheme community that typical iterative lazy algorithms written using delay and force will often require unbounded memory.

This SRFI provides set of three operations: {lazy, delay, force}, which allow the programmer to succinctly express lazy algorithms while retaining bounded space behavior in cases that are properly tail-recursive. A general recipe for using these primitives is provided. An additional procedure eager is provided for the construction of eager promises in cases where efficiency is a concern.

Although this SRFI redefines delay and force, the extension is conservative in the sense that the semantics of the subset {delay, force} in isolation (i.e., as long as the program does not use lazy) agrees with that in R5RS. In other words, no program that uses the R5RS definitions of delay and force will break if those definition are replaced by the SRFI-45 definitions of delay and force.

Guile also adds promise? to the list of exports, which is not part of the official SRFI-45.

[Scheme Procedure] promise? obj

Return true if obj is an SRFI-45 promise, otherwise return false.

[Scheme Syntax] delay expression Takes an expression of arbitrary type a and returns a promise of type (Promise a) which at some point in the future may be asked (by the force procedure) to evaluate the expression and deliver the resulting value.

[Scheme Syntax] lazy expression Takes an expression of type (Promise a) and returns a promise of type (Promise a) which at some point in the future may be asked (by the force procedure) to evaluate the expression and deliver the resulting promise.

[Scheme Procedure] force expression Takes an argument of type (Promise a) and returns a value of type a as follows: If a value of type a has been computed for the promise, this value is returned. Otherwise, the promise is first evaluated, then overwritten by the obtained promise or value, and then force is again applied (iteratively) to the promise.

[Scheme Procedure] eager expression Takes an argument of type a and returns a value of type (Promise a). As opposed to delay, the argument is evaluated eagerly. Semantically, writing (eager expression) is equivalent to writing

          (let ((value expression)) (delay value)).

However, the former is more efficient since it does not require unnecessary creation and evaluation of thunks. We also have the equivalence

          (delay expression) = (lazy (eager expression))

The following reduction rules may be helpful for reasoning about these primitives. However, they do not express the memoization and memory usage semantics specified above:

     (force (delay expression)) -> expression
     (force (lazy  expression)) -> (force expression)
     (force (eager value))      -> value

Correct usage

We now provide a general recipe for using the primitives {lazy, delay, force} to express lazy algorithms in Scheme. The transformation is best described by way of an example: Consider the stream-filter algorithm, expressed in a hypothetical lazy language as

     (define (stream-filter p? s)
       (if (null? s) '()
           (let ((h (car s))
                 (t (cdr s)))
             (if (p? h)
                 (cons h (stream-filter p? t))
                 (stream-filter p? t)))))

This algorithm can be expressed as follows in Scheme:

     (define (stream-filter p? s)
       (lazy
          (if (null? (force s)) (delay '())
              (let ((h (car (force s)))
                    (t (cdr (force s))))
                (if (p? h)
                    (delay (cons h (stream-filter p? t)))
                    (stream-filter p? t))))))

In other words, we

• wrap all constructors (e.g., '(), cons) with delay,

• apply force to arguments of deconstructors (e.g., car, cdr and null?),

• wrap procedure bodies with (lazy ...).

7.5.31 SRFI-46 Basic syntax-rules Extensions

Guile’s core syntax-rules supports the extensions specified by SRFI-46/R7RS. Tail patterns have been supported since at least Guile 2.0, and custom ellipsis identifiers have been supported since Guile 2.0.10. See Section 6.10.2 [Syntax Rules], page 258.

7.5.32 SRFI-55 - Requiring Features

SRFI-55 provides require-extension which is a portable mechanism to load selected SRFI modules. This is implemented in the Guile core, there’s no module needed to get SRFI-55 itself.

[library syntax] require-extension clause1 clause2 …

Require the features of clause1 clause2 …, throwing an error if any are unavailable. A clause is of the form (identifier arg...). The only identifier currently supported is srfi and the arguments are SRFI numbers. For example to get SRFI-1 and SRFI-6,

     (require-extension (srfi 1 6))

require-extension can only be used at the top-level.

A Guile-specific program can simply use-modules to load SRFIs not already in the core, require-extension is for programs designed to be portable to other Scheme implementations.

7.5.33 SRFI-60 - Integers as Bits

This SRFI provides various functions for treating integers as bits and for bitwise manipulations. These functions can be obtained with,

(use-modules (srfi srfi-60))

Integers are treated as infinite precision twos-complement, the same as in the core logical functions (see Section 6.6.2.13 [Bitwise Operations], page 125). And likewise bit indexes start from 0 for the least significant bit. The following functions in this SRFI are already in the Guile core,

logand, logior, logxor, lognot, logtest, logcount, integer-length,
logbit?, ash

[Function] bitwise-and n1 ...

[Function] bitwise-ior n1 ...

[Function] bitwise-xor n1 ...

[Function] bitwise-not n
[Function] any-bits-set? j k

[Function] bit-set? index n

[Function] arithmetic-shift n count

[Function] bit-field n start end

[Function] bit-count n

Aliases for logand, logior, logxor, lognot, logtest, logbit?, ash, bit-extract

and logcount respectively.

Note that the name bit-count conflicts with bit-count in the core (see Section 6.7.4 [Bit Vectors], page 198).

[Function] bitwise-if mask n1 n0

[Function] bitwise-merge mask n1 n0

Return an integer with bits selected from n1 and n0 according to mask. Those bits where mask has 1s are taken from n1, and those where mask has 0s are taken from n0.

     (bitwise-if 3 #b0101 #b1010) ⇒ 9
[Function] log2-binary-factors n
[Function] first-set-bit n

Return a count of how many factors of 2 are present in n. This is also the bit index of the lowest 1 bit in n. If n is 0, the return is −1.

     (log2-binary-factors 6) ⇒ 1
     (log2-binary-factors -8) ⇒ 3

[Function] copy-bit index n newbit Return n with the bit at index set according to newbit. newbit should be #t to set the bit to 1, or #f to set it to 0. Bits other than at index are unchanged in the return.

     (copy-bit 1 #b0101 #t) ⇒ 7

[Function] copy-bit-field n newbits start end Return n with the bits from start (inclusive) to end (exclusive) changed to the value newbits.

The least significant bit in newbits goes to start, the next to start + 1, etc. Anything in newbits past the end given is ignored.

     (copy-bit-field #b10000 #b11 1 3) ⇒ #b10110

[Function] rotate-bit-field n count start end

Return n with the bit field from start (inclusive) to end (exclusive) rotated upwards by count bits.

count can be positive or negative, and it can be more than the field width (it’ll be reduced modulo the width).

     (rotate-bit-field #b0110 2 1 4) ⇒ #b1010

[Function] reverse-bit-field n start end

Return n with the bits from start (inclusive) to end (exclusive) reversed.

     (reverse-bit-field #b101001 2 4) ⇒ #b100101

[Function] integer->list n [len]

Return bits from n in the form of a list of #t for 1 and #f for 0. The least significant len bits are returned, and the first list element is the most significant of those bits. If len is not given, the default is (integer-length n) (see Section 6.6.2.13 [Bitwise Operations], page 125).

     (integer->list 6) ⇒ (#t #t #f)
     (integer->list 1 4) ⇒ (#f #f #f #t)
[Function] list->integer lst
[Function] booleans->integer bool. . .

Return an integer formed bitwise from the given lst list of booleans, or for booleans- >integer from the bool arguments.

Each boolean is #t for a 1 and #f for a 0. The first element becomes the most significant bit in the return.

     (list->integer '(#t #f #t #f)) ⇒ 10

7.5.34 SRFI-61 - A more general cond clause

This SRFI extends RnRS cond to support test expressions that return multiple values, as well as arbitrary definitions of test success. SRFI 61 is implemented in the Guile core; there’s no module needed to get SRFI-61 itself. Extended cond is documented in Section 6.13.2 [Simple Conditional Evaluation], page 292.

7.5.35 SRFI-62 - S-expression comments.

Starting from version 2.0, Guile’s read supports SRFI-62/R7RS S-expression comments by default.

7.5.36 SRFI-64 - A Scheme API for test suites.

See the specification of SRFI-64 (http://srfi.schemers.org/srfi-64/srfi-64.html).

7.5.37 SRFI-67 - Compare procedures

See the specification of SRFI-67 (http://srfi.schemers.org/srfi-67/srfi-67.html).

7.5.38 SRFI-69 - Basic hash tables

This is a portable wrapper around Guile’s built-in hash table and weak table support. See Section 6.7.14 [Hash Tables], page 236, for information on that built-in support. Above that, this hash-table interface provides association of equality and hash functions with tables at creation time, so variants of each function are not required, as well as a procedure that takes care of most uses for Guile hash table handles, which this SRFI does not provide as such.

Access it with:

(use-modules (srfi srfi-69))

7.5.38.1 Creating hash tables

[Scheme Procedure] make-hash-table [equal-proc hash-proc #:weak weakness

start-size] Create and answer a new hash table with equal-proc as the equality function and hash-proc as the hashing function. By default, equal-proc is equal?. It can be any two-argument procedure, and should answer whether two keys are the same for this table’s purposes. My default hash-proc assumes that equal-proc is no coarser than equal? unless it is literally string-ci=?. If provided, hash-proc should be a two-argument procedure that takes a key and the current table size, and answers a reasonably good hash integer between 0 (inclusive) and the size (exclusive). weakness should be #f or a symbol indicating how “weak” the hash table is:

#f An ordinary non-weak hash table. This is the default.

key When the key has no more non-weak references at GC, remove that entry.

value When the value has no more non-weak references at GC, remove that entry.

key-or-value

When either has no more non-weak references at GC, remove the association.

As a legacy of the time when Guile couldn’t grow hash tables, start-size is an optional integer argument that specifies the approximate starting size for the hash table, which will be rounded to an algorithmically-sounder number.

By coarser than equal?, we mean that for all x and y values where (equal-proc x y), (equal? x y) as well. If that does not hold for your equal-proc, you must provide a hash-proc.

In the case of weak tables, remember that references above always refers to eq?-wise references. Just because you have a reference to some string "foo" doesn’t mean that an association with key "foo" in a weak-key table won’t be collected; it only counts as a reference if the two "foo"s are eq?, regardless of equal-proc. As such, it is usually only sensible to use eq? and hashq as the equivalence and hash functions for a weak table. See Section 6.18.3 [Weak References], page 378, for more information on Guile’s built-in weak table support.

[Scheme Procedure] alist->hash-table alist [equal-proc hash-proc #:weak

weakness start-size] As with make-hash-table, but initialize it with the associations in alist. Where keys are repeated in alist, the leftmost association takes precedence.

7.5.38.2 Accessing table items

[Scheme Procedure] hash-table-ref table key [default-thunk]

[Scheme Procedure] hash-table-ref/default table key default

Answer the value associated with key in table. If key is not present, answer the result of invoking the thunk default-thunk, which signals an error instead by default. hash-table-ref/default is a variant that requires a third argument, default, and answers default itself instead of invoking it.

[Scheme Procedure] hash-table-set! table key new-value Set key to new-value in table.

[Scheme Procedure] hash-table-delete! table key

Remove the association of key in table, if present. If absent, do nothing.

[Scheme Procedure] hash-table-exists? table key

Answer whether key has an association in table.

[Scheme Procedure] hash-table-update! table key modifier [default-thunk]

[Scheme Procedure] hash-table-update!/default table key modifier default

Replace key’s associated value in table by invoking modifier with one argument, the old value. If key is not present, and default-thunk is provided, invoke it with no arguments to get the “old value” to be passed to modifier as above. If default-thunk is not provided in such a case, signal an error. hash-table-update!/default is a variant that requires the fourth argument, which is used directly as the “old value” rather than as a thunk to be invoked to retrieve the “old value”.

7.5.38.3 Table properties

[Scheme Procedure] hash-table-size table

Answer the number of associations in table. This is guaranteed to run in constant time for non-weak tables.

[Scheme Procedure] hash-table-keys table

Answer an unordered list of the keys in table.

[Scheme Procedure] hash-table-values table

Answer an unordered list of the values in table.

[Scheme Procedure] hash-table-walk table proc

Invoke proc once for each association in table, passing the key and value as arguments.

[Scheme Procedure] hash-table-fold table proc init

Invoke (proc key value previous) for each key and value in table, where previous is the result of the previous invocation, using init as the first previous value. Answer the final proc result.

[Scheme Procedure] hash-table->alist table

Answer an alist where each association in table is an association in the result.

7.5.38.4 Hash table algorithms

Each hash table carries an equivalence function and a hash function, used to implement key lookups. Beginning users should follow the rules for consistency of the default hash-proc specified above. Advanced users can use these to implement their own equivalence and hash functions for specialized lookup semantics.

[Scheme Procedure] hash-table-equivalence-function hash-table

[Scheme Procedure] hash-table-hash-function hash-table

Answer the equivalence and hash function of hash-table, respectively.

[Scheme Procedure] hash obj [size]

[Scheme Procedure] string-hash obj [size] [Scheme Procedure] string-ci-hash obj [size] [Scheme Procedure] hash-by-identity obj [size]

Answer a hash value appropriate for equality predicate equal?, string=?, stringci=?, and eq?, respectively.

hash is a backwards-compatible replacement for Guile’s built-in hash.

7.5.39 SRFI-87 => in case clauses

Starting from version 2.0.6, Guile’s core case syntax supports => in clauses, as specified by SRFI-87/R7RS. See Section 6.13.2 [Conditionals], page 292.

7.5.40 SRFI-88 Keyword Objects

SRFI-88 (http://srfi.schemers.org/srfi-88/srfi-88.html) provides keyword objects,

which are equivalent to Guile’s keywords (see Section 6.6.8 [Keywords], page 180). SRFI-88 keywords can be entered using the postfix keyword syntax, which consists of an identifier followed by : (see Section 6.17.2 [Scheme Read], page 360). SRFI-88 can be made available with:

     (use-modules (srfi srfi-88))

Doing so installs the right reader option for keyword syntax, using (read-set! keywords 'postfix). It also provides the procedures described below.

[Scheme Procedure] keyword? obj

Return #t if obj is a keyword. This is the same procedure as the same-named built-in procedure (see Section 6.6.8.4 [Keyword Procedures], page 183).

     (keyword? foo:)  ⇒ #t
     (keyword? 'foo:)  ⇒ #t
     (keyword? "foo")  ⇒ #f

[Scheme Procedure] keyword->string kw

Return the name of kw as a string, i.e., without the trailing colon. The returned string may not be modified, e.g., with string-set!.

     (keyword->string foo:) ⇒ "foo"

[Scheme Procedure] string->keyword str

Return the keyword object whose name is str.

     (keyword->string (string->keyword "a b c"))  ⇒ "a b c"

7.5.41 SRFI-98 Accessing environment variables.

This is a portable wrapper around Guile’s built-in support for interacting with the current environment, See Section 7.2.6 [Runtime Environment], page 488.

[Scheme Procedure] get-environment-variable name

Returns a string containing the value of the environment variable given by the string name, or #f if the named environment variable is not found. This is equivalent to

(getenv name).
[Scheme Procedure] get-environment-variables

Returns the names and values of all the environment variables as an association list in which both the keys and the values are strings.

7.5.42 SRFI-105 Curly-infix expressions.

Guile’s built-in reader includes support for SRFI-105 curly-infix expressions. See the spec-

ification of SRFI-105 (http://srfi.schemers.org/srfi-105/srfi-105.html). Some

examples:

{n <= 5}  ⇒ (<= n 5)
{a + b + c}  ⇒ (+ a b c)
{a * {b + c}}  ⇒ (* a (+ b c))
{(- a) / b}  ⇒ (/ (- a) b)
{-(a) / b}  ⇒ (/ (- a) b) as well
{(f a b) + (g h)}  ⇒ (+ (f a b) (g h))
{f(a b) + g(h)}  ⇒ (+ (f a b) (g h)) as well
{f[a b] + g(h)}  ⇒ (+ ($bracket-apply$ f a b) (g h))
'{a + f(b) + x}  ⇒ '(+ a (f b) x)
{length(x) >= 6}  ⇒ (>= (length x) 6)
{n-1 + n-2}  ⇒ (+ n-1 n-2)
{n * factorial{n - 1}} ⇒ (* n (factorial (- n 1)))
{{a > 0} and {b >= 1}} ⇒ (and (> a 0) (>= b 1))
{f{n - 1}(x)}  ⇒ ((f (- n 1)) x)

     {a . z}  ⇒ ($nfx$ a . z)
     {a + b - c}  ⇒ ($nfx$ a + b - c)

To enable curly-infix expressions within a file, place the reader directive #!curly-infix before the first use of curly-infix notation. To globally enable curly-infix expressions in Guile’s reader, set the curly-infix read option.

Guile also implements the following non-standard extension to SRFI-105: if curly-infix is enabled and there is no other meaning assigned to square brackets (i.e. the squarebrackets read option is turned off), then lists within square brackets are read as normal lists but with the special symbol $bracket-list$ added to the front. To enable this combination of read options within a file, use the reader directive #!curly-infix-and-

bracket-lists. For example:
     [a b]  ⇒ ($bracket-list$ a b)
     [a . b] ⇒ ($bracket-list$ a . b)

For more information on reader options, See Section 6.17.2 [Scheme Read], page 360.

7.5.43 SRFI-111 Boxes.

SRFI-111 (http://srfi.schemers.org/srfi-111/srfi-111.html) provides boxes: ob-

jects with a single mutable cell.

[Scheme Procedure] box value Return a newly allocated box whose contents is initialized to value.

[Scheme Procedure] box? obj

Return true if obj is a box, otherwise return false.

[Scheme Procedure] unbox box

Return the current contents of box.

[Scheme Procedure] set-box! box value

Set the contents of box to value.

7.6 R6RS Support

See Section 6.19.6 [R6RS Libraries], page 388, for more information on how to define R6RS libraries, and their integration with Guile modules.

7.6.1 Incompatibilities with the R6RS

There are some incompatibilities between Guile and the R6RS. Some of them are intentional, some of them are bugs, and some are simply unimplemented features. Please let the Guile developers know if you find one that is not on this list.

• The R6RS specifies many situations in which a conforming implementation must signal a specific error. Guile doesn’t really care about that too much—if a correct R6RS program would not hit that error, we don’t bother checking for it.

• Multiple library forms in one file are not yet supported. This is because the expansion of library sets the current module, but does not restore it. This is a bug.

• R6RS unicode escapes within strings are disabled by default, because they conflict with Guile’s already-existing escapes. The same is the case for R6RS treatment of escaped newlines in strings. R6RS behavior can be turned on via a reader option. See Section 6.6.5.1 [String Syntax], page 141, for more information.

• A set! to a variable transformer may only expand to an expression, not a definition— even if the original set! expression was in definition context.

• Instead of using the algorithm detailed in chapter 10 of the R6RS, expansion of toplevel forms happens sequentially. For example, while the expansion of the following set of toplevel definitions does the correct thing:

     (begin
      (define even?
        (lambda (x)
          (or (= x 0) (odd? (- x 1)))))
      (define-syntax odd?
        (syntax-rules ()
          ((odd? x) (not (even? x)))))
      (even? 10))
     ⇒ #t

The same definitions outside of the begin wrapper do not:

     (define even?
       (lambda (x)
         (or (= x 0) (odd? (- x 1)))))
     (define-syntax odd?
       (syntax-rules ()
         ((odd? x) (not (even? x)))))
     (even? 10)
     <unnamed port>:4:18: In procedure even?:
     <unnamed port>:4:18: Wrong type to apply: #<syntax-transformer odd?>

This is because when expanding the right-hand-side of even?, the reference to odd? is not yet marked as a syntax transformer, so it is assumed to be a function. This bug will only affect top-level programs, not code in library forms. Fixing it for toplevel forms seems doable, but tricky to implement in a backward-compatible way. Suggestions and/or patches would be appreciated.

• The (rnrs io ports) module is incomplete. Work is ongoing to fix this.

• Guile does not prevent use of textual I/O procedures on binary ports. More generally, it does not make a sharp distinction between binary and textual ports (see Section 6.14.10.6 [R6RS Port Manipulation], page 337).

• Guile’s implementation of equal? may fail to terminate when applied to arguments containing cycles.

7.6.2 R6RS Standard Libraries

In contrast with earlier versions of the Revised Report, the R6RS organizes the procedures and syntactic forms required of conforming implementations into a set of “standard libraries” which can be imported as necessary by user programs and libraries. Here we briefly list the libraries that have been implemented for Guile.

We do not attempt to document these libraries fully here, as most of their functionality is already available in Guile itself. The expectation is that most Guile users will use the well-known and well-documented Guile modules. These R6RS libraries are mostly useful to users who want to port their code to other R6RS systems.

The documentation in the following sections reproduces some of the content of the library section of the Report, but is mostly intended to provide supplementary information about Guile’s implementation of the R6RS standard libraries. For complete documentation, design rationales and further examples, we advise you to consult the “Standard Libraries” section of the Report (see Section “Standard Libraries” in The Revised^6 Report on the Algorithmic Language Scheme).

7.6.2.1 Library Usage

Guile implements the R6RS ‘library’ form as a transformation to a native Guile module definition. As a consequence of this, all of the libraries described in the following subsections, in addition to being available for use by R6RS libraries and top-level programs, can also be imported as if they were normal Guile modules—via a use-modules form, say. For example, the R6RS “composite” library can be imported by:

       (import (rnrs (6)))
       (use-modules ((rnrs) :version (6)))

For more information on Guile’s library implementation, see (see Section 6.19.6 [R6RS Libraries], page 388).

7.6.2.2 rnrs base

The (rnrs base (6)) library exports the procedures and syntactic forms described in the main section of the Report (see Section “Base library” in The Revised^6 Report on the Algorithmic Language Scheme). They are grouped below by the existing manual sections to which they correspond.

[Scheme Procedure] boolean? obj [Scheme Procedure] not x

See Section 6.6.1 [Booleans], page 104, for documentation.

[Scheme Procedure] symbol? obj [Scheme Procedure] symbol->string sym [Scheme Procedure] string->symbol str

See Section 6.6.7.4 [Symbol Primitives], page 173, for documentation.

[Scheme Procedure] char? obj [Scheme Procedure] char=? [Scheme Procedure] char<? [Scheme Procedure] char>? [Scheme Procedure] char<=? [Scheme Procedure] char>=? [Scheme Procedure] integer->char n [Scheme Procedure] char->integer chr

See Section 6.6.3 [Characters], page 129, for documentation.

[Scheme Procedure] list? x [Scheme Procedure] null? x

See Section 6.7.2.2 [List Predicates], page 188, for documentation.

[Scheme Procedure] pair? x [Scheme Procedure] cons x y

[Scheme Procedure] car pair [Scheme Procedure] cdr pair

[Scheme Procedure] caar pair [Scheme Procedure] cadr pair [Scheme Procedure] cdar pair [Scheme Procedure] cddr pair [Scheme Procedure] caaar pair [Scheme Procedure] caadr pair [Scheme Procedure] cadar pair [Scheme Procedure] cdaar pair [Scheme Procedure] caddr pair [Scheme Procedure] cdadr pair [Scheme Procedure] cddar pair [Scheme Procedure] cdddr pair [Scheme Procedure] caaaar pair [Scheme Procedure] caaadr pair [Scheme Procedure] caadar pair [Scheme Procedure] cadaar pair [Scheme Procedure] cdaaar pair [Scheme Procedure] cddaar pair [Scheme Procedure] cdadar pair [Scheme Procedure] cdaadr pair [Scheme Procedure] cadadr pair [Scheme Procedure] caaddr pair [Scheme Procedure] caddar pair [Scheme Procedure] cadddr pair [Scheme Procedure] cdaddr pair [Scheme Procedure] cddadr pair [Scheme Procedure] cdddar pair [Scheme Procedure] cddddr pair

See Section 6.7.1 [Pairs], page 185, for documentation.

[Scheme Procedure] number? obj

See Section 6.6.2.1 [Numerical Tower], page 105, for documentation.

[Scheme Procedure] string? obj

See Section 6.6.5.2 [String Predicates], page 142, for documentation.

[Scheme Procedure] procedure? obj

See Section 6.9.7 [Procedure Properties], page 254, for documentation.

[Scheme Syntax] define name value [Scheme Syntax] set! variable-name value

See Section 3.1.3 [Definition], page 16, for documentation.

[Scheme Syntax] define-syntax keyword expression [Scheme Syntax] let-syntax ((keyword transformer) …) exp1 exp2 … [Scheme Syntax] letrec-syntax ((keyword transformer) …) exp1 exp2 …

See Section 6.10.1 [Defining Macros], page 257, for documentation.

[Scheme Syntax] identifier-syntax exp

See Section 6.10.6 [Identifier Macros], page 271, for documentation.

[Scheme Syntax] syntax-rules literals (pattern template) ...

See Section 6.10.2 [Syntax Rules], page 258, for documentation.

[Scheme Syntax] lambda formals body See Section 6.9.1 [Lambda], page 244, for documentation.

[Scheme Syntax] let bindings body [Scheme Syntax] let* bindings body [Scheme Syntax] letrec bindings body [Scheme Syntax] letrec* bindings body

See Section 6.12.2 [Local Bindings], page 287, for documentation.

[Scheme Syntax] let-values bindings body [Scheme Syntax] let*-values bindings body

See Section 7.5.10 [SRFI-11], page 577, for documentation.

[Scheme Syntax] begin expr1 expr2 ...

See Section 6.13.1 [begin], page 290, for documentation.

[Scheme Syntax] quote expr [Scheme Syntax] quasiquote expr [Scheme Syntax] unquote expr [Scheme Syntax] unquote-splicing expr

See Section 6.17.1.1 [Expression Syntax], page 357, for documentation.

[Scheme Syntax] if test consequence [alternate] [Scheme Syntax] cond clause1 clause2 ... [Scheme Syntax] case key clause1 clause2 ...

See Section 6.13.2 [Conditionals], page 292, for documentation.

[Scheme Syntax] and expr ... [Scheme Syntax] or expr ...

See Section 6.13.3 [and or], page 293, for documentation.

[Scheme Procedure] eq? x y [Scheme Procedure] eqv? x y [Scheme Procedure] equal? x y

[Scheme Procedure] symbol=? symbol1 symbol2 ...

See Section 6.11.1 [Equality], page 276, for documentation.

symbol=? is identical to eq?.

[Scheme Procedure] complex? z

See Section 6.6.2.4 [Complex Numbers], page 113, for documentation.

[Scheme Procedure] real-part z [Scheme Procedure] imag-part z

[Scheme Procedure] make-rectangular real part imaginary part

[Scheme Procedure] make-polar x y [Scheme Procedure] magnitude z [Scheme Procedure] angle z

See Section 6.6.2.10 [Complex], page 118, for documentation.

[Scheme Procedure] sqrt z [Scheme Procedure] exp z [Scheme Procedure] expt z1 z2 [Scheme Procedure] log z [Scheme Procedure] sin z [Scheme Procedure] cos z [Scheme Procedure] tan z [Scheme Procedure] asin z [Scheme Procedure] acos z [Scheme Procedure] atan z

See Section 6.6.2.12 [Scientific], page 123, for documentation.

[Scheme Procedure] real? x [Scheme Procedure] rational? x [Scheme Procedure] numerator x [Scheme Procedure] denominator x [Scheme Procedure] rationalize x eps

See Section 6.6.2.3 [Reals and Rationals], page 110, for documentation.

[Scheme Procedure] exact? x [Scheme Procedure] inexact? x [Scheme Procedure] exact z [Scheme Procedure] inexact z

See Section 6.6.2.5 [Exactness], page 113, for documentation. The exact and inexact procedures are identical to the inexact->exact and exact->inexact procedures provided by Guile’s code library.

[Scheme Procedure] integer? x

See Section 6.6.2.2 [Integers], page 106, for documentation.

[Scheme Procedure] odd? n [Scheme Procedure] even? n

[Scheme Procedure] gcd x ... [Scheme Procedure] lcm x ...

[Scheme Procedure] exact-integer-sqrt k

See Section 6.6.2.7 [Integer Operations], page 116, for documentation.

[Scheme Procedure] = [Scheme Procedure] < [Scheme Procedure] > [Scheme Procedure] <= [Scheme Procedure] >= [Scheme Procedure] zero? x [Scheme Procedure] positive? x [Scheme Procedure] negative? x

See Section 6.6.2.8 [Comparison], page 117, for documentation.

[Scheme Procedure] for-each f lst1 lst2 ...

See Section 7.5.3.5 [SRFI-1 Fold and Map], page 556, for documentation.

[Scheme Procedure] list elem …

See Section 6.7.2.3 [List Constructors], page 189, for documentation.

[Scheme Procedure] length lst [Scheme Procedure] list-ref lst k [Scheme Procedure] list-tail lst k

See Section 6.7.2.4 [List Selection], page 190, for documentation.

[Scheme Procedure] append lst … obj [Scheme Procedure] append [Scheme Procedure] reverse lst

See Section 6.7.2.5 [Append/Reverse], page 190, for documentation.

[Scheme Procedure] number->string n [radix] [Scheme Procedure] string->number str [radix]

See Section 6.6.2.9 [Conversion], page 118, for documentation.

[Scheme Procedure] string char ...

[Scheme Procedure] make-string k [chr] [Scheme Procedure] list->string lst

See Section 6.6.5.3 [String Constructors], page 143, for documentation.

[Scheme Procedure] string->list str [start [end]]

See Section 6.6.5.4 [List/String Conversion], page 144, for documentation.

[Scheme Procedure] string-length str [Scheme Procedure] string-ref str k [Scheme Procedure] string-copy str [start [end]] [Scheme Procedure] substring str start [end]

See Section 6.6.5.5 [String Selection], page 145, for documentation.

[Scheme Procedure] string=? s1 s2 s3 … [Scheme Procedure] string<? s1 s2 s3 … [Scheme Procedure] string>? s1 s2 s3 … [Scheme Procedure] string<=? s1 s2 s3 … [Scheme Procedure] string>=? s1 s2 s3 …

See Section 6.6.5.7 [String Comparison], page 148, for documentation.

[Scheme Procedure] string-append arg …

See Section 6.6.5.10 [Reversing and Appending Strings], page 155, for documentation.

[Scheme Procedure] string-for-each proc s [start [end]]

See Section 6.6.5.11 [Mapping Folding and Unfolding], page 156, for documentation.

[Scheme Procedure] + z1 ... [Scheme Procedure] - z1 z2 ... [Scheme Procedure] * z1 ... [Scheme Procedure] / z1 z2 ... [Scheme Procedure] max x1 x2 ... [Scheme Procedure] min x1 x2 ...

[Scheme Procedure] abs x [Scheme Procedure] truncate x [Scheme Procedure] floor x [Scheme Procedure] ceiling x [Scheme Procedure] round x

See Section 6.6.2.11 [Arithmetic], page 119, for documentation.

[Scheme Procedure] div x y [Scheme Procedure] mod x y [Scheme Procedure] div-and-mod x y

These procedures accept two real numbers x and y, where the divisor y must be non-zero. div returns the integer q and mod returns the real number r such that x = q ∗ y + r and 0 <= r < abs(y). div-and-mod returns both q and r, and is more efficient than computing each separately. Note that when y > 0, div returns floor(x/y), otherwise it returns ceiling(x/y).

     (div 123 10) ⇒ 12
     (mod 123 10) ⇒ 3
     (div-and-mod 123 10) ⇒ 12 and 3
     (div-and-mod 123 -10) ⇒ -12 and 3
     (div-and-mod -123 10) ⇒ -13 and 7
     (div-and-mod -123 -10) ⇒ 13 and 7
     (div-and-mod -123.2 -63.5) ⇒ 2.0 and 3.8
     (div-and-mod 16/3 -10/7) ⇒ -3 and 22/21

[Scheme Procedure] div0 x y [Scheme Procedure] mod0 x y [Scheme Procedure] div0-and-mod0 x y

These procedures accept two real numbers x and y, where the divisor y must be non-zero. div0 returns the integer q and mod0 returns the real number r such that x = q ∗ y + r and −abs(y/2) <= r < abs(y/2). div0-and-mod0 returns both q and r, and is more efficient than computing each separately. Note that div0 returns x/y rounded to the nearest integer. When x/y lies exactly half-way between two integers, the tie is broken according to the sign of y. If y > 0, ties are rounded toward positive infinity, otherwise they are rounded toward negative infinity. This is a consequence of the requirement that −abs(y/2) <= r < abs(y/2).

     (div0 123 10) ⇒ 12
     (mod0 123 10) ⇒ 3
     (div0-and-mod0 123 10) ⇒ 12 and 3

     (div0-and-mod0 123 -10) ⇒ -12 and 3
     (div0-and-mod0 -123 10) ⇒ -12 and -3
     (div0-and-mod0 -123 -10) ⇒ 12 and -3
     (div0-and-mod0 -123.2 -63.5) ⇒ 2.0 and 3.8
     (div0-and-mod0 16/3 -10/7) ⇒ -4 and -8/21

[Scheme Procedure] real-valued? obj [Scheme Procedure] rational-valued? obj [Scheme Procedure] integer-valued? obj

These procedures return #t if and only if their arguments can, respectively, be coerced to a real, rational, or integer value without a loss of numerical precision.

real-valued? will return #t for complex numbers whose imaginary parts are zero.

[Scheme Procedure] nan? x [Scheme Procedure] infinite? x [Scheme Procedure] finite? x

nan? returns #t if x is a NaN value, #f otherwise. infinite? returns #t if x is an infinite value, #f otherwise. finite? returns #t if x is neither infinite nor a NaN value, otherwise it returns #f. Every real number satisfies exactly one of these predicates. An exception is raised if x is not real.

[Scheme Syntax] assert expr

Raises an &assertion condition if expr evaluates to #f; otherwise evaluates to the value of expr.

[Scheme Procedure] error who message irritant1 ... [Scheme Procedure] assertion-violation who message irritant1 ...

These procedures raise compound conditions based on their arguments: If who is not #f, the condition will include a &who condition whose who field is set to who; a &message condition will be included with a message field equal to message; an &irritants condition will be included with its irritants list given by irritant1

....

error produces a compound condition with the simple conditions described above, as well as an &error condition; assertion-violation produces one that includes an

&assertion condition.

[Scheme Procedure] vector-map proc v [Scheme Procedure] vector-for-each proc v

These procedures implement the map and for-each contracts over vectors.

[Scheme Procedure] vector arg … [Scheme Procedure] vector? obj [Scheme Procedure] make-vector len [Scheme Procedure] make-vector len fill [Scheme Procedure] list->vector l [Scheme Procedure] vector->list v

See Section 6.7.3.2 [Vector Creation], page 194, for documentation.

[Scheme Procedure] vector-length vector [Scheme Procedure] vector-ref vector k [Scheme Procedure] vector-set! vector k obj [Scheme Procedure] vector-fill! v fill

See Section 6.7.3.3 [Vector Accessors], page 195, for documentation.

[Scheme Procedure] call-with-current-continuation proc

[Scheme Procedure] call/cc proc

See Section 6.13.6 [Continuations], page 300, for documentation.

[Scheme Procedure] values arg … [Scheme Procedure] call-with-values producer consumer

See Section 6.13.7 [Multiple Values], page 301, for documentation.

[Scheme Procedure] dynamic-wind in guard thunk out guard See Section 6.13.10 [Dynamic Wind], page 309, for documentation.

[Scheme Procedure] apply proc arg … arglst See Section 6.17.4 [Fly Evaluation], page 362, for documentation.

7.6.2.3 rnrs unicode

The (rnrs unicode (6)) library provides procedures for manipulating Unicode characters and strings.

[Scheme Procedure] char-upcase char [Scheme Procedure] char-downcase char [Scheme Procedure] char-titlecase char [Scheme Procedure] char-foldcase char

These procedures translate their arguments from one Unicode character set to an-

other. char-upcase, char-downcase, and char-titlecase are identical to their

counterparts in the Guile core library; See Section 6.6.3 [Characters], page 129, for documentation. char-foldcase returns the result of applying char-upcase to its argument, followed by char-downcase—except in the case of the Turkic characters U+0130 and U+0131, for which the procedure acts as the identity function.

[Scheme Procedure] char-ci=? char1 char2 char3 ... [Scheme Procedure] char-ci<? char1 char2 char3 ... [Scheme Procedure] char-ci>? char1 char2 char3 ... [Scheme Procedure] char-ci<=? char1 char2 char3 ... [Scheme Procedure] char-ci>=? char1 char2 char3 ...

These procedures facilitate case-insensitive comparison of Unicode characters. They are identical to the procedures provided by Guile’s core library. See Section 6.6.3 [Characters], page 129, for documentation.

[Scheme Procedure] char-alphabetic? char [Scheme Procedure] char-numeric? char [Scheme Procedure] char-whitespace? char [Scheme Procedure] char-upper-case? char [Scheme Procedure] char-lower-case? char [Scheme Procedure] char-title-case? char

These procedures implement various Unicode character set predicates. They are identical to the procedures provided by Guile’s core library. See Section 6.6.3 [Characters], page 129, for documentation.

[Scheme Procedure] char-general-category char

See Section 6.6.3 [Characters], page 129, for documentation.

[Scheme Procedure] string-upcase string [Scheme Procedure] string-downcase string [Scheme Procedure] string-titlecase string [Scheme Procedure] string-foldcase string

These procedures perform Unicode case folding operations on their input. See Section 6.6.5.9 [Alphabetic Case Mapping], page 154, for documentation.

[Scheme Procedure] string-ci=? string1 string2 string3 ... [Scheme Procedure] string-ci<? string1 string2 string3 ... [Scheme Procedure] string-ci>? string1 string2 string3 ... [Scheme Procedure] string-ci<=? string1 string2 string3 ... [Scheme Procedure] string-ci>=? string1 string2 string3 ...

These procedures perform case-insensitive comparison on their input. See Section 6.6.5.7 [String Comparison], page 148, for documentation.

[Scheme Procedure] string-normalize-nfd string [Scheme Procedure] string-normalize-nfkd string [Scheme Procedure] string-normalize-nfc string [Scheme Procedure] string-normalize-nfkc string

These procedures perform Unicode string normalization operations on their input. See Section 6.6.5.7 [String Comparison], page 148, for documentation.

7.6.2.4 rnrs bytevectors

The (rnrs bytevectors (6)) library provides procedures for working with blocks of binary data. This functionality is documented in its own section of the manual; See Section 6.6.6 [Bytevectors], page 163.

7.6.2.5 rnrs lists

The (rnrs lists (6)) library provides procedures additional procedures for working with lists.

[Scheme Procedure] find proc list This procedure is identical to the one defined in Guile’s SRFI-1 implementation. See Section 7.5.3.7 [SRFI-1 Searching], page 560, for documentation.

[Scheme Procedure] for-all proc list1 list2 ... [Scheme Procedure] exists proc list1 list2 ...

The for-all procedure is identical to the every procedure defined by SRFI-1; the exists procedure is identical to SRFI-1’s any. See Section 7.5.3.7 [SRFI-1 Searching], page 560, for documentation.

[Scheme Procedure] filter proc list

[Scheme Procedure] partition proc list

These procedures are identical to the ones provided by SRFI-1. See Section 6.7.2.6 [List Modification], page 191, for a description of filter; See Section 7.5.3.6 [SRFI-1 Filtering and Partitioning], page 559, for partition.

[Scheme Procedure] fold-left combine nil list1 list2 … [Scheme Procedure] fold-right combine nil list1 list2 …

These procedures are identical to the fold and fold-right procedures provided by SRFI-1. See Section 7.5.3.5 [SRFI-1 Fold and Map], page 556, for documentation.

[Scheme Procedure] remp proc list [Scheme Procedure] remove obj list [Scheme Procedure] remv obj list [Scheme Procedure] remq obj list

remove, remv, and remq are identical to the delete, delv, and delq procedures provided by Guile’s core library, (see Section 6.7.2.6 [List Modification], page 191). remp is identical to the alternate remove procedure provided by SRFI-1; See Section 7.5.3.8 [SRFI-1 Deleting], page 561.

[Scheme Procedure] memp proc list [Scheme Procedure] member obj list [Scheme Procedure] memv obj list [Scheme Procedure] memq obj list

member, memv, and memq are identical to the procedures provided by Guile’s core library; See Section 6.7.2.7 [List Searching], page 192, for their documentation. memp uses the specified predicate function proc to test elements of the list list—it behaves similarly to find, except that it returns the first sublist of list whose car satisfies proc.

[Scheme Procedure] assp proc alist [Scheme Procedure] assoc obj alist [Scheme Procedure] assv obj alist [Scheme Procedure] assq obj alist

assoc, assv, and assq are identical to the procedures provided by Guile’s core library; See Section 6.7.12.1 [Alist Key Equality], page 228, for their documentation. assp uses the specified predicate function proc to test keys in the association list alist.

[Scheme Procedure] cons* obj1 ... obj

[Scheme Procedure] cons* obj

This procedure is identical to the one exported by Guile’s core library. See Section 6.7.2.3 [List Constructors], page 189, for documentation.

7.6.2.6 rnrs sorting

The (rnrs sorting (6)) library provides procedures for sorting lists and vectors.

[Scheme Procedure] list-sort proc list [Scheme Procedure] vector-sort proc vector

These procedures return their input sorted in ascending order, without modifying the original data. proc must be a procedure that takes two elements from the input list or vector as arguments, and returns a true value if the first is “less” than the second, #f otherwise. list-sort returns a list; vector-sort returns a vector. Both list-sort and vector-sort are implemented in terms of the stable-sort procedure from Guile’s core library. See Section 6.11.3 [Sorting], page 279, for a discussion of the behavior of that procedure.

[Scheme Procedure] vector-sort! proc vector

Performs a destructive, “in-place” sort of vector, using proc as described above to determine an ascending ordering of elements. vector-sort! returns an unspecified value. This procedure is implemented in terms of the sort! procedure from Guile’s core library. See Section 6.11.3 [Sorting], page 279, for more information.

7.6.2.7 rnrs control

The (rnrs control (6)) library provides syntactic forms useful for constructing conditional expressions and controlling the flow of execution.

[Scheme Syntax] when test expression1 expression2 ... [Scheme Syntax] unless test expression1 expression2 ...

The when form is evaluated by evaluating the specified test expression; if the result is a true value, the expressions that follow it are evaluated in order, and the value of the final expression becomes the value of the entire when expression. The unless form behaves similarly, with the exception that the specified expressions are only evaluated if the value of test is false.

[Scheme Syntax] do ((variable init step) ...) (test expression ...) command ...

This form is identical to the one provided by Guile’s core library. See Section 6.13.4 [while do], page 294, for documentation.

[Scheme Syntax] case-lambda clause ...

This form is identical to the one provided by Guile’s core library. See Section 6.9.5 [Case-lambda], page 251, for documentation.

7.6.2.8 R6RS Records

The manual sections below describe Guile’s implementation of R6RS records, which provide support for user-defined data types. The R6RS records API provides a superset of the features provided by Guile’s “native” records, as well as those of the SRFI-9 records API; See Section 6.7.9 [Records], page 219, and Section 6.7.8 [SRFI-9 Records], page 216, for a description of those interfaces.

As with SRFI-9 and Guile’s native records, R6RS records are constructed using a recordtype descriptor that specifies attributes like the record’s name, its fields, and the mutability of those fields.

R6RS records extend this framework to support single inheritance via the specification of a “parent” type for a record type at definition time. Accessors and mutator procedures for the fields of a parent type may be applied to records of a subtype of this parent. A record type may be sealed, in which case it cannot be used as the parent of another record type.

The inheritance mechanism for record types also informs the process of initializing the fields of a record and its parents. Constructor procedures that generate new instances of a record type are obtained from a record constructor descriptor, which encapsulates the record-type descriptor of the record to be constructed along with a protocol procedure that defines how constructors for record subtypes delegate to the constructors of their parent types.

A protocol is a procedure used by the record system at construction time to bind arguments to the fields of the record being constructed. The protocol procedure is passed a procedure n that accepts the arguments required to construct the record’s parent type; this procedure, when invoked, will return a procedure p that accepts the arguments required to construct a new instance of the record type itself and returns a new instance of the record type.

The protocol should in turn return a procedure that uses n and p to initialize the fields of the record type and its parent type(s). This procedure will be the constructor returned by

As a trivial example, consider the hypothetical record type pixel, which encapsulates an x-y location on a screen, and voxel, which has pixel as its parent type and stores an additional coordinate. The following protocol produces a constructor procedure that accepts all three coordinates, uses the first two to initialize the fields of pixel, and binds the third to the single field of voxel.

       (lambda (n)
         (lambda (x y z)
           (let ((p (n x y)))
             (p z))))

It may be helpful to think of protocols as “constructor factories” that produce chains of delegating constructors glued together by the helper procedure n.

An R6RS record type may be declared to be nongenerative via the use of a unique generated or user-supplied symbol—or uid—such that subsequent record type declarations with the same uid and attributes will return the previously-declared record-type descriptor.

R6RS record types may also be declared to be opaque, in which case the various predicates and introspection procedures defined in (rnrs records introspection) will behave as if records of this type are not records at all.

Note that while the R6RS records API shares much of its namespace with both the SRFI-9 and native Guile records APIs, it is not currently compatible with either.

7.6.2.9 rnrs records syntactic

The (rnrs records syntactic (6)) library exports the syntactic API for working with R6RS records.

[Scheme Syntax] define-record-type name-spec record-clause … Defines a new record type, introducing bindings for a record-type descriptor, a record constructor descriptor, a constructor procedure, a record predicate, and accessor and mutator procedures for the new record type’s fields. name-spec must either be an identifier or must take the form (record-name constructor-name predicate-name), where record-name, constructor-name, and predicate-name are all identifiers and specify the names to which, respectively, the record-type descriptor, constructor, and predicate procedures will be bound. If name-spec is only an identifier, it specifies the name to which the generated record-type descriptor will be bound. Each record-clause must be one of the following:

• (fields field-spec*), where each field-spec specifies a field of the new record

type and takes one of the following forms:

  • (immutable field-name accessor-name), which specifies an immutable

field with the name field-name and binds an accessor procedure for it to the name given by accessor-name

  • (mutable field-name accessor-name mutator-name), which specifies a

mutable field with the name field-name and binds accessor and mutator procedures to accessor-name and mutator-name, respectively

• (immutable field-name), which specifies an immutable field with the name field-name; an accessor procedure for it will be created and named by appending record name and field-name with a hyphen separator

• (mutable field-name), which specifies a mutable field with the name fieldname; an accessor procedure for it will be created and named as described above; a mutator procedure will also be created and named by appending -set! to the accessor name

• field-name, which specifies an immutable field with the name field-name;

an access procedure for it will be created and named as described above • (parent parent-name), where parent-name is a symbol giving the name of the

record type to be used as the parent of the new record type

• (protocol expression), where expression evaluates to a protocol procedure which behaves as described above, and is used to create a record constructor descriptor for the new record type

• (sealed sealed?), where sealed? is a boolean value that specifies whether or not the new record type is sealed

• (opaque opaque?), where opaque? is a boolean value that specifies whether or not the new record type is opaque

• (nongenerative [uid]), which specifies that the record type is nongenerative via the optional uid uid. If uid is not specified, a unique uid will be generated at expansion time

• (parent-rtd parent-rtd parent-cd), a more explicit form of the parent form

above; parent-rtd and parent-cd should evaluate to a record-type descriptor and a record constructor descriptor, respectively

[Scheme Syntax] record-type-descriptor record-name

Evaluates to the record-type descriptor associated with the type specified by recordname.

[Scheme Syntax] record-constructor-descriptor record-name

Evaluates to the record-constructor descriptor associated with the type specified by record-name.

7.6.2.10 rnrs records procedural

The (rnrs records procedural (6)) library exports the procedural API for working with R6RS records.

[Scheme Procedure] make-record-type-descriptor name parent uid sealed?

opaque? fields Returns a new record-type descriptor with the specified characteristics: name must be a symbol giving the name of the new record type; parent must be either #f or a non-sealed record-type descriptor for the returned record type to extend; uid must be either #f, indicating that the record type is generative, or a symbol giving the type’s nongenerative uid; sealed? and opaque? must be boolean values that specify the sealedness and opaqueness of the record type; fields must be a vector of zero or more field specifiers of the form (mutable name) or (immutable name), where name is a symbol giving a name for the field. If uid is not #f, it must be a symbol

[Scheme Procedure] record-type-descriptor? obj

Returns #t if obj is a record-type descriptor, #f otherwise.

[Scheme Procedure] make-record-constructor-descriptor rtd

parent-constructor-descriptor protocol Returns a new record constructor descriptor that can be used to produce constructors for the record type specified by the record-type descriptor rtd and whose delegation and binding behavior are specified by the protocol procedure protocol. parent-constructor-descriptor specifies a record constructor descriptor for the parent type of rtd, if one exists. If rtd represents a base type, then parent-constructor-descriptor must be #f. If rtd is an extension of another type, parent-constructor-descriptor may still be #f, but protocol must also be #f in this case.

[Scheme Procedure] record-constructor rcd

Returns a record constructor procedure by invoking the protocol defined by the recordconstructor descriptor rcd.

[Scheme Procedure] record-predicate rtd

Returns the record predicate procedure for the record-type descriptor rtd.

[Scheme Procedure] record-accessor rtd k

Returns the record field accessor procedure for the kth field of the record-type descriptor rtd.

[Scheme Procedure] record-mutator rtd k

Returns the record field mutator procedure for the kth field of the record-type descriptor rtd. An &assertion condition will be raised if this field is not mutable.

7.6.2.11 rnrs records inspection

The (rnrs records inspection (6)) library provides procedures useful for accessing metadata about R6RS records.

[Scheme Procedure] record? obj

Return #t if the specified object is a non-opaque R6RS record, #f otherwise.

[Scheme Procedure] record-rtd record

Returns the record-type descriptor for record. An &assertion is raised if record is opaque.

[Scheme Procedure] record-type-name rtd

Returns the name of the record-type descriptor rtd.

[Scheme Procedure] record-type-parent rtd

Returns the parent of the record-type descriptor rtd, or #f if it has none.

[Scheme Procedure] record-type-uid rtd

Returns the uid of the record-type descriptor rtd, or #f if it has none.

[Scheme Procedure] record-type-generative? rtd

Returns #t if the record-type descriptor rtd is generative, #f otherwise.

[Scheme Procedure] record-type-sealed? rtd

Returns #t if the record-type descriptor rtd is sealed, #f otherwise.

[Scheme Procedure] record-type-opaque? rtd

Returns #t if the record-type descriptor rtd is opaque, #f otherwise.

[Scheme Procedure] record-type-field-names rtd

Returns a vector of symbols giving the names of the fields defined by the record-type descriptor rtd (and not any of its sub- or supertypes).

[Scheme Procedure] record-field-mutable? rtd k

Returns #t if the field at index k of the record-type descriptor rtd (and not any of its sub- or supertypes) is mutable.

7.6.2.12 rnrs exceptions

The (rnrs exceptions (6)) library provides functionality related to signaling and handling exceptional situations. This functionality is similar to the exception handling systems provided by Guile’s core library See Section 6.13.8 [Exceptions], page 303, and by the SRFI-18 and SRFI-34 modules—See Section 7.5.15.5 [SRFI-18 Exceptions], page 582, and Section 7.5.22 [SRFI-34], page 594, respectively—but there are some key differences in concepts and behavior.

A raised exception may be continuable or non-continuable. When an exception is raised non-continuably, another exception, with the condition type &non-continuable, will be raised when the exception handler returns locally. Raising an exception continuably captures the current continuation and invokes it after a local return from the exception handler.

Like SRFI-18 and SRFI-34, R6RS exceptions are implemented on top of Guile’s native throw and catch forms, and use custom “throw keys” to identify their exception types. As a consequence, Guile’s catch form can handle exceptions thrown by these APIs, but the reverse is not true: Handlers registered by the with-exception-handler procedure described below will only be called on exceptions thrown by the corresponding raise procedure.

[Scheme Procedure] with-exception-handler handler thunk

Installs handler, which must be a procedure taking one argument, as the current exception handler during the invocation of thunk, a procedure taking zero arguments. The handler in place at the time with-exception-handler is called is made current again once either thunk returns or handler is invoked after an exception is thrown from within thunk. This procedure is similar to the with-throw-handler procedure provided by Guile’s code library; (see Section 6.13.8.3 [Throw Handlers], page 306).

[Scheme Syntax] guard (variable clause1 clause2 ...) body Evaluates the expression given by body, first creating an ad hoc exception handler that binds a raised exception to variable and then evaluates the specified clauses as if they were part of a cond expression, with the value of the first matching clause becoming the value of the guard expression (see Section 6.13.2 [Conditionals], page 292). If none of the clause’s test expressions evaluates to #t, the exception is re-raised, with the exception handler that was current before the evaluation of the guard form. For example, the expression

     (guard (ex ((eq? ex 'foo) 'bar) ((eq? ex 'bar) 'baz))
       (raise 'bar))

evaluates to baz.

[Scheme Procedure] raise obj

Raises a non-continuable exception by invoking the currently-installed exception handler on obj. If the handler returns, a &non-continuable exception will be raised in the dynamic context in which the handler was installed.

[Scheme Procedure] raise-continuable obj

Raises a continuable exception by invoking currently-installed exception handler on obj.

7.6.2.13 rnrs conditions

The (rnrs condition (6)) library provides forms and procedures for constructing new condition types, as well as a library of pre-defined condition types that represent a variety of common exceptional situations. Conditions are records of a subtype of the &condition record type, which is neither sealed nor opaque. See Section 7.6.2.8 [R6RS Records], page 634.

Conditions may be manipulated singly, as simple conditions, or when composed with other conditions to form compound conditions. Compound conditions do not “nest”— constructing a new compound condition out of existing compound conditions will “flatten” them into their component simple conditions. For example, making a new condition out of a &message condition and a compound condition that contains an &assertion condition and another &message condition will produce a compound condition that contains two &message conditions and one &assertion condition.

The record type predicates and field accessors described below can operate on either simple or compound conditions. In the latter case, the predicate returns #t if the compound condition contains a component simple condition of the appropriate type; the field accessors return the requisite fields from the first component simple condition found to be of the appropriate type.

This library is quite similar to the SRFI-35 conditions module (see Section 7.5.23 [SRFI-35], page 594). Among other minor differences, the (rnrs conditions) library features slightly different semantics around condition field accessors, and comes with a larger number of pre-defined condition types. The two APIs are not currently compatible, however; the condition? predicate from one API will return #f when applied to a condition object created in the other.

[Condition Type] &condition [Scheme Procedure] condition? obj

The base record type for conditions.

[Scheme Procedure] condition condition1 ...

[Scheme Procedure] simple-conditions condition

The condition procedure creates a new compound condition out of its condition arguments, flattening any specified compound conditions into their component simple conditions as described above. simple-conditions returns a list of the component simple conditions of the compound condition condition, in the order in which they were specified at construction time.

[Scheme Procedure] condition-predicate rtd [Scheme Procedure] condition-accessor rtd proc

These procedures return condition predicate and accessor procedures for the specified condition record type rtd.

[Scheme Syntax] define-condition-type condition-type supertype constructor

predicate field-spec ... Evaluates to a new record type definition for a condition type with the name conditiontype that has the condition type supertype as its parent. A default constructor, which binds its arguments to the fields of this type and its parent types, will be bound to the identifier constructor; a condition predicate will be bound to predicate. The fields of the new type, which are immutable, are specified by the field-specs, each of which must be of the form:

     (field accessor)

where field gives the name of the field and accessor gives the name for a binding to an accessor procedure created for this field.

[Condition Type] &message [Scheme Procedure] make-message-condition message [Scheme Procedure] message-condition? obj [Scheme Procedure] condition-message condition

A type that includes a message describing the condition that occurred.

[Condition Type] &warning [Scheme Procedure] make-warning [Scheme Procedure] warning? obj

A base type for representing non-fatal conditions during execution.

[Condition Type] &serious

[Scheme Procedure] make-serious-condition

[Scheme Procedure] serious-condition? obj

A base type for conditions representing errors serious enough that cannot be ignored.

[Condition Type] &error [Scheme Procedure] make-error [Scheme Procedure] error? obj

A base type for conditions representing errors.

[Condition Type] &violation [Scheme Procedure] make-violation [Scheme Procedure] violation?

A subtype of &serious that can be used to represent violations of a language or library standard.

[Condition Type] &assertion

[Scheme Procedure] make-assertion-violation
[Scheme Procedure] assertion-violation? obj

A subtype of &violation that indicates an invalid call to a procedure.

[Condition Type] &irritants [Scheme Procedure] make-irritants-condition irritants

[Scheme Procedure] irritants-condition? obj

[Scheme Procedure] condition-irritants condition

A base type used for storing information about the causes of another condition in a compound condition.

[Condition Type] &who [Scheme Procedure] make-who-condition who [Scheme Procedure] who-condition? obj [Scheme Procedure] condition-who condition

A base type used for storing the identity, a string or symbol, of the entity responsible for another condition in a compound condition.

[Condition Type] &non-continuable
[Scheme Procedure] make-non-continuable-violation
[Scheme Procedure] non-continuable-violation? obj

A subtype of &violation used to indicate that an exception handler invoked by raise has returned locally.

[Condition Type] &implementation-restriction
[Scheme Procedure] make-implementation-restriction-violation
[Scheme Procedure] implementation-restriction-violation? obj

A subtype of &violation used to indicate a violation of an implementation restriction.

[Condition Type] &lexical

[Scheme Procedure] make-lexical-violation

[Scheme Procedure] lexical-violation? obj

A subtype of &violation used to indicate a syntax violation at the level of the datum syntax.

[Condition Type] &syntax [Scheme Procedure] make-syntax-violation form subform [Scheme Procedure] syntax-violation? obj [Scheme Procedure] syntax-violation-form condition [Scheme Procedure] syntax-violation-subform condition

A subtype of &violation that indicates a syntax violation. The form and subform fields, which must be datum values, indicate the syntactic form responsible for the condition.

[Condition Type] &undefined

[Scheme Procedure] make-undefined-violation
[Scheme Procedure] undefined-violation? obj

A subtype of &violation that indicates a reference to an unbound identifier.

7.6.2.14 I/O Conditions

These condition types are exported by both the (rnrs io ports (6)) and (rnrs io simple (6)) libraries.

[Condition Type] &i/o [Scheme Procedure] make-i/o-error [Scheme Procedure] i/o-error? obj

A condition supertype for more specific I/O errors.

[Condition Type] &i/o-read

[Scheme Procedure] make-i/o-read-error

[Scheme Procedure] i/o-read-error? obj

A subtype of &i/o; represents read-related I/O errors.

[Condition Type] &i/o-write

[Scheme Procedure] make-i/o-write-error

[Scheme Procedure] i/o-write-error? obj

A subtype of &i/o; represents write-related I/O errors.

[Condition Type] &i/o-invalid-position
[Scheme Procedure] make-i/o-invalid-position-error position
[Scheme Procedure] i/o-invalid-position-error? obj

[Scheme Procedure] i/o-error-position condition

A subtype of &i/o; represents an error related to an attempt to set the file position to an invalid position.

[Condition Type] &i/o-filename [Scheme Procedure] make-io-filename-error filename [Scheme Procedure] i/o-filename-error? obj [Scheme Procedure] i/o-error-filename condition

A subtype of &i/o; represents an error related to an operation on a named file.

[Condition Type] &i/o-file-protection
[Scheme Procedure] make-i/o-file-protection-error filename

[Scheme Procedure] i/o-file-protection-error? obj

A subtype of &i/o-filename; represents an error resulting from an attempt to access a named file for which the caller had insufficient permissions.

[Condition Type] &i/o-file-is-read-only
[Scheme Procedure] make-i/o-file-is-read-only-error filename
[Scheme Procedure] i/o-file-is-read-only-error? obj

A subtype of &i/o-file-protection; represents an error related to an attempt to write to a read-only file.

[Condition Type] &i/o-file-already-exists
[Scheme Procedure] make-i/o-file-already-exists-error filename
[Scheme Procedure] i/o-file-already-exists-error? obj

A subtype of &i/o-filename; represents an error related to an operation on an existing file that was assumed not to exist.

[Condition Type] &i/o-file-does-not-exist
[Scheme Procedure] make-i/o-file-does-not-exist-error
[Scheme Procedure] i/o-file-does-not-exist-error? obj

A subtype of &i/o-filename; represents an error related to an operation on a nonexistent file that was assumed to exist.

[Condition Type] &i/o-port [Scheme Procedure] make-i/o-port-error port [Scheme Procedure] i/o-port-error? obj [Scheme Procedure] i/o-error-port condition

A subtype of &i/o; represents an error related to an operation on the port port.

7.6.2.15 rnrs io ports

The (rnrs io ports (6)) library provides various procedures and syntactic forms for use in writing to and reading from ports. This functionality is documented in its own section of the manual; (see Section 6.14.10 [R6RS I/O Ports], page 332).

7.6.2.16 rnrs io simple

The (rnrs io simple (6)) library provides convenience functions for performing textual I/O on ports. This library also exports all of the condition types and associated procedures described in (see Section 7.6.2.14 [I/O Conditions], page 642). In the context of this section, when stating that a procedure behaves “identically” to the corresponding procedure in Guile’s core library, this is modulo the behavior wrt. conditions: such procedures raise the appropriate R6RS conditions in case of error, but otherwise behave identically.

Note: There are still known issues regarding condition-correctness; some errors may still be thrown as native Guile exceptions instead of the appropriate R6RS conditions.

[Scheme Procedure] eof-object [Scheme Procedure] eof-object? obj

These procedures are identical to the ones provided by the (rnrs io ports (6)) library. See Section 6.14.10 [R6RS I/O Ports], page 332, for documentation.

[Scheme Procedure] input-port? obj [Scheme Procedure] output-port? obj

These procedures are identical to the ones provided by Guile’s core library. See Section 6.14.1 [Ports], page 316, for documentation.

[Scheme Procedure] call-with-input-file filename proc [Scheme Procedure] call-with-output-file filename proc [Scheme Procedure] open-input-file filename [Scheme Procedure] open-output-file filename [Scheme Procedure] with-input-from-file filename thunk [Scheme Procedure] with-output-to-file filename thunk

These procedures are identical to the ones provided by Guile’s core library. See Section 6.14.9.1 [File Ports], page 326, for documentation.

[Scheme Procedure] close-input-port input-port [Scheme Procedure] close-output-port output-port

These procedures are identical to the ones provided by Guile’s core library. See Section 6.14.4 [Closing], page 321, for documentation.

[Scheme Procedure] peek-char

[Scheme Procedure] peek-char textual-input-port

[Scheme Procedure] read-char

[Scheme Procedure] read-char textual-input-port These procedures are identical to the ones provided by Guile’s core library. See Section 6.14.2 [Reading], page 318, for documentation.

[Scheme Procedure] read

[Scheme Procedure] read textual-input-port This procedure is identical to the one provided by Guile’s core library. See Section 6.17.2 [Scheme Read], page 360, for documentation.

[Scheme Procedure] display obj

[Scheme Procedure] display obj textual-output-port

[Scheme Procedure] newline

[Scheme Procedure] newline textual-output-port

[Scheme Procedure] write obj

[Scheme Procedure] write obj textual-output-port

[Scheme Procedure] write-char char

[Scheme Procedure] write-char char textual-output-port These procedures are identical to the ones provided by Guile’s core library. See Section 6.14.3 [Writing], page 320, for documentation.

7.6.2.17 rnrs files

The (rnrs files (6)) library provides the file-exists? and delete-file procedures,

which test for the existence of a file and allow the deletion of files from the file system, respectively.

These procedures are identical to the ones provided by Guile’s core library. See Section 7.2.3 [File System], page 476, for documentation.

7.6.2.18 rnrs programs

The (rnrs programs (6)) library provides procedures for process management and introspection.

[Scheme Procedure] command-line

This procedure is identical to the one provided by Guile’s core library. See Section 7.2.6 [Runtime Environment], page 488, for documentation.

[Scheme Procedure] exit [status]

This procedure is identical to the one provided by Guile’s core library. See Section 7.2.7 [Processes], page 490, for documentation.

7.6.2.19 rnrs arithmetic fixnums

The (rnrs arithmetic fixnums (6)) library provides procedures for performing arithmetic operations on an implementation-dependent range of exact integer values, which R6RS refers to as fixnums. In Guile, the size of a fixnum is determined by the size of the SCM type; a single SCM struct is guaranteed to be able to hold an entire fixnum, making fixnum computations particularly efficient—(see Section 6.3 [The SCM Type], page 100). On 32-bit systems, the most negative and most positive fixnum values are, respectively, -536870912 and 536870911.

Unless otherwise specified, all of the procedures below take fixnums as arguments, and will raise an &assertion condition if passed a non-fixnum argument or an &implementation-restriction condition if their result is not itself a fixnum.

[Scheme Procedure] fixnum? obj

Returns #t if obj is a fixnum, #f otherwise.

[Scheme Procedure] fixnum-width [Scheme Procedure] least-fixnum [Scheme Procedure] greatest-fixnum

These procedures return, respectively, the maximum number of bits necessary to represent a fixnum value in Guile, the minimum fixnum value, and the maximum fixnum value.

[Scheme Procedure] fx=? fx1 fx2 fx3 ... [Scheme Procedure] fx>? fx1 fx2 fx3 ... [Scheme Procedure] fx<? fx1 fx2 fx3 ... [Scheme Procedure] fx>=? fx1 fx2 fx3 ... [Scheme Procedure] fx<=? fx1 fx2 fx3 ...

These procedures return #t if their fixnum arguments are (respectively): equal, monotonically increasing, monotonically decreasing, monotonically nondecreasing, or monotonically nonincreasing; #f otherwise.

[Scheme Procedure] fxzero? fx [Scheme Procedure] fxpositive? fx [Scheme Procedure] fxnegative? fx [Scheme Procedure] fxodd? fx [Scheme Procedure] fxeven? fx

These numerical predicates return #t if fx is, respectively, zero, greater than zero, less than zero, odd, or even; #f otherwise.

[Scheme Procedure] fxmax fx1 fx2 ... [Scheme Procedure] fxmin fx1 fx2 ...

These procedures return the maximum or minimum of their arguments.

[Scheme Procedure] fx+ fx1 fx2 [Scheme Procedure] fx* fx1 fx2

These procedures return the sum or product of their arguments.

[Scheme Procedure] fx- fx1 fx2

[Scheme Procedure] fx- fx

Returns the difference of fx1 and fx2, or the negation of fx, if called with a single argument.

An &assertion condition is raised if the result is not itself a fixnum.

[Scheme Procedure] fxdiv-and-mod fx1 fx2

[Scheme Procedure] fxdiv fx1 fx2 [Scheme Procedure] fxmod fx1 fx2

[Scheme Procedure] fxdiv0-and-mod0 fx1 fx2 [Scheme Procedure] fxdiv0 fx1 fx2 [Scheme Procedure] fxmod0 fx1 fx2

These procedures implement number-theoretic division on fixnums; See ⟨undefined⟩ [(rnrs base)], page ⟨undefined⟩, for a description of their semantics.

[Scheme Procedure] fx+/carry fx1 fx2 fx3

Returns the two fixnum results of the following computation:

     (let* ((s (+ fx1 fx2 fx3))
            (s0 (mod0 s (expt 2 (fixnum-width))))
            (s1 (div0 s (expt 2 (fixnum-width)))))
       (values s0 s1))

[Scheme Procedure] fx-/carry fx1 fx2 fx3

Returns the two fixnum results of the following computation:

     (let* ((d (- fx1 fx2 fx3))
            (d0 (mod0 d (expt 2 (fixnum-width))))
            (d1 (div0 d (expt 2 (fixnum-width)))))
       (values d0 d1))

[Scheme Procedure] fx*/carry fx1 fx2 fx3

     Returns the two fixnum results of the following computation:
     (let* ((s (+ (* fx1 fx2) fx3))
            (s0 (mod0 s (expt 2 (fixnum-width))))
            (s1 (div0 s (expt 2 (fixnum-width)))))
       (values s0 s1))

[Scheme Procedure] fxnot fx

[Scheme Procedure] fxand fx1 ... [Scheme Procedure] fxior fx1 ... [Scheme Procedure] fxxor fx1 ...

These procedures are identical to the lognot, logand, logior, and logxor procedures provided by Guile’s core library. See Section 6.6.2.13 [Bitwise Operations], page 125, for documentation.

[Scheme Procedure] fxif fx1 fx2 fx3 Returns the bitwise “if” of its fixnum arguments. The bit at position i in the return value will be the ith bit from fx2 if the ith bit of fx1 is 1, the ith bit from fx3.

[Scheme Procedure] fxbit-count fx

Returns the number of 1 bits in the two’s complement representation of fx.

[Scheme Procedure] fxlength fx

Returns the number of bits necessary to represent fx.

[Scheme Procedure] fxfirst-bit-set fx

Returns the index of the least significant 1 bit in the two’s complement representation of fx.

[Scheme Procedure] fxbit-set? fx1 fx2

Returns #t if the fx2th bit in the two’s complement representation of fx1 is 1, #f otherwise.

[Scheme Procedure] fxcopy-bit fx1 fx2 fx3

Returns the result of setting the fx2th bit of fx1 to the fx2th bit of fx3.

[Scheme Procedure] fxbit-field fx1 fx2 fx3

Returns the integer representation of the contiguous sequence of bits in fx1 that starts at position fx2 (inclusive) and ends at position fx3 (exclusive).

[Scheme Procedure] fxcopy-bit-field fx1 fx2 fx3 fx4

Returns the result of replacing the bit field in fx1 with start and end positions fx2 and fx3 with the corresponding bit field from fx4.

[Scheme Procedure] fxarithmetic-shift fx1 fx2

[Scheme Procedure] fxarithmetic-shift-left fx1 fx2
[Scheme Procedure] fxarithmetic-shift-right fx1 fx2

Returns the result of shifting the bits of fx1 right or left by the fx2 positions.

fxarithmetic-shift is identical to fxarithmetic-shift-left.

[Scheme Procedure] fxrotate-bit-field fx1 fx2 fx3 fx4

Returns the result of cyclically permuting the bit field in fx1 with start and end positions fx2 and fx3 by fx4 bits in the direction of more significant bits.

[Scheme Procedure] fxreverse-bit-field fx1 fx2 fx3

Returns the result of reversing the order of the bits of fx1 between position fx2 (inclusive) and position fx3 (exclusive).

7.6.2.20 rnrs arithmetic flonums

The (rnrs arithmetic flonums (6)) library provides procedures for performing arithmetic operations on inexact representations of real numbers, which R6RS refers to as flonums.

Unless otherwise specified, all of the procedures below take flonums as arguments, and will raise an &assertion condition if passed a non-flonum argument.

[Scheme Procedure] flonum? obj

Returns #t if obj is a flonum, #f otherwise.

[Scheme Procedure] real->flonum x

Returns the flonum that is numerically closest to the real number x.

[Scheme Procedure] fl=? fl1 fl2 fl3 ... [Scheme Procedure] fl<? fl1 fl2 fl3 ... [Scheme Procedure] fl<=? fl1 fl2 fl3 ... [Scheme Procedure] fl>? fl1 fl2 fl3 ... [Scheme Procedure] fl>=? fl1 fl2 fl3 ...

These procedures return #t if their flonum arguments are (respectively): equal, monotonically increasing, monotonically decreasing, monotonically nondecreasing, or monotonically nonincreasing; #f otherwise.

[Scheme Procedure] flinteger? fl [Scheme Procedure] flzero? fl [Scheme Procedure] flpositive? fl [Scheme Procedure] flnegative? fl [Scheme Procedure] flodd? fl [Scheme Procedure] fleven? fl

These numerical predicates return #t if fl is, respectively, an integer, zero, greater than zero, less than zero, odd, even, #f otherwise. In the case of flodd? and fleven?, fl must be an integer-valued flonum.

[Scheme Procedure] flfinite? fl [Scheme Procedure] flinfinite? fl [Scheme Procedure] flnan? fl

These numerical predicates return #t if fl is, respectively, not infinite, infinite, or a NaN value.

[Scheme Procedure] flmax fl1 fl2 ... [Scheme Procedure] flmin fl1 fl2 ...

These procedures return the maximum or minimum of their arguments.

[Scheme Procedure] fl+ fl1 ... [Scheme Procedure] fl* fl ...

These procedures return the sum or product of their arguments.

[Scheme Procedure] fl- fl1 fl2 ...

[Scheme Procedure] fl- fl

[Scheme Procedure] fl/ fl1 fl2 ...

[Scheme Procedure] fl/ fl

These procedures return, respectively, the difference or quotient of their arguments when called with two arguments; when called with a single argument, they return the additive or multiplicative inverse of fl.

[Scheme Procedure] flabs fl

Returns the absolute value of fl.

[Scheme Procedure] fldiv-and-mod fl1 fl2 [Scheme Procedure] fldiv fl1 fl2 [Scheme Procedure] fldmod fl1 fl2 [Scheme Procedure] fldiv0-and-mod0 fl1 fl2 [Scheme Procedure] fldiv0 fl1 fl2 [Scheme Procedure] flmod0 fl1 fl2

These procedures implement number-theoretic division on flonums; See ⟨undefined⟩ [(rnrs base)], page ⟨undefined⟩, for a description for their semantics.

[Scheme Procedure] flnumerator fl [Scheme Procedure] fldenominator fl

These procedures return the numerator or denominator of fl as a flonum.

[Scheme Procedure] flfloor fl1 [Scheme Procedure] flceiling fl [Scheme Procedure] fltruncate fl [Scheme Procedure] flround fl

These procedures are identical to the floor, ceiling, truncate, and round procedures provided by Guile’s core library. See Section 6.6.2.11 [Arithmetic], page 119, for documentation.

[Scheme Procedure] flexp fl [Scheme Procedure] fllog fl [Scheme Procedure] fllog fl1 fl2 [Scheme Procedure] flsin fl [Scheme Procedure] flcos fl [Scheme Procedure] fltan fl [Scheme Procedure] flasin fl [Scheme Procedure] flacos fl [Scheme Procedure] flatan fl [Scheme Procedure] flatan fl1 fl2

These procedures, which compute the usual transcendental functions, are the flonum variants of the procedures provided by the R6RS base library (see ⟨undefined⟩ [(rnrs base)], page ⟨undefined⟩).

[Scheme Procedure] flsqrt fl

Returns the square root of fl. If fl is -0.0, -0.0 is returned; for other negative values, a NaN value is returned.

[Scheme Procedure] flexpt fl1 fl2

Returns the value of fl1 raised to the power of fl2.

The following condition types are provided to allow Scheme implementations that do not support infinities or NaN values to indicate that a computation resulted in such a value. Guile supports both of these, so these conditions will never be raised by Guile’s standard libraries implementation.

[Condition Type] &no-infinities

[Scheme Procedure] make-no-infinities-violation obj
[Scheme Procedure] no-infinities-violation?

A condition type indicating that a computation resulted in an infinite value on a Scheme implementation incapable of representing infinities.

[Condition Type] &no-nans

[Scheme Procedure] make-no-nans-violation obj

[Scheme Procedure] no-nans-violation? obj

A condition type indicating that a computation resulted in a NaN value on a Scheme implementation incapable of representing NaNs.

[Scheme Procedure] fixnum->flonum fx

Returns the flonum that is numerically closest to the fixnum fx.

7.6.2.21 rnrs arithmetic bitwise

The (rnrs arithmetic bitwise (6)) library provides procedures for performing bitwise arithmetic operations on the two’s complement representations of fixnums.

This library and the procedures it exports share functionality with SRFI-60, which provides support for bitwise manipulation of integers (see Section 7.5.33 [SRFI-60], page 616).

[Scheme Procedure] bitwise-not ei [Scheme Procedure] bitwise-and ei1 ... [Scheme Procedure] bitwise-ior ei1 ... [Scheme Procedure] bitwise-xor ei1 ...

These procedures are identical to the lognot, logand, logior, and logxor procedures provided by Guile’s core library. See Section 6.6.2.13 [Bitwise Operations], page 125, for documentation.

[Scheme Procedure] bitwise-if ei1 ei2 ei3

Returns the bitwise “if” of its arguments. The bit at position i in the return value will be the ith bit from ei2 if the ith bit of ei1 is 1, the ith bit from ei3.

[Scheme Procedure] bitwise-bit-count ei

Returns the number of 1 bits in the two’s complement representation of ei.

[Scheme Procedure] bitwise-length ei

Returns the number of bits necessary to represent ei.

[Scheme Procedure] bitwise-first-bit-set ei

Returns the index of the least significant 1 bit in the two’s complement representation of ei.

[Scheme Procedure] bitwise-bit-set? ei1 ei2

Returns #t if the ei2th bit in the two’s complement representation of ei1 is 1, #f otherwise.

[Scheme Procedure] bitwise-copy-bit ei1 ei2 ei3

Returns the result of setting the ei2th bit of ei1 to the ei2th bit of ei3.

[Scheme Procedure] bitwise-bit-field ei1 ei2 ei3

Returns the integer representation of the contiguous sequence of bits in ei1 that starts at position ei2 (inclusive) and ends at position ei3 (exclusive).

[Scheme Procedure] bitwise-copy-bit-field ei1 ei2 ei3 ei4

Returns the result of replacing the bit field in ei1 with start and end positions ei2 and ei3 with the corresponding bit field from ei4.

[Scheme Procedure] bitwise-arithmetic-shift ei1 ei2
[Scheme Procedure] bitwise-arithmetic-shift-left ei1 ei2
[Scheme Procedure] bitwise-arithmetic-shift-right ei1 ei2

Returns the result of shifting the bits of ei1 right or left by the ei2 positions. bitwise-

arithmetic-shift is identical to bitwise-arithmetic-shift-left.

[Scheme Procedure] bitwise-rotate-bit-field ei1 ei2 ei3 ei4

Returns the result of cyclically permuting the bit field in ei1 with start and end positions ei2 and ei3 by ei4 bits in the direction of more significant bits.

[Scheme Procedure] bitwise-reverse-bit-field ei1 ei2 ei3

Returns the result of reversing the order of the bits of ei1 between position ei2 (inclusive) and position ei3 (exclusive).

7.6.2.22 rnrs syntax-case

The (rnrs syntax-case (6)) library provides access to the syntax-case system for writing hygienic macros. With one exception, all of the forms and procedures exported by this library are “re-exports” of Guile’s native support for syntax-case; See Section 6.10.3 [Syntax Case], page 263, for documentation, examples, and rationale.

[Scheme Procedure] make-variable-transformer proc

Creates a new variable transformer out of proc, a procedure that takes a syntax object as input and returns a syntax object. If an identifier to which the result of this procedure is bound appears on the left-hand side of a set! expression, proc will be called with a syntax object representing the entire set! expression, and its return value will replace that set! expression.

[Scheme Syntax] syntax-case expression (literal ...) clause ...

The syntax-case pattern matching form.

[Scheme Syntax] syntax template

[Scheme Syntax] quasisyntax template [Scheme Syntax] unsyntax template [Scheme Syntax] unsyntax-splicing template

These forms allow references to be made in the body of a syntax-case output expression subform to datum and non-datum values. They are identical to the forms provided by Guile’s core library; See Section 6.10.3 [Syntax Case], page 263, for documentation.

[Scheme Procedure] identifier? obj [Scheme Procedure] bound-identifier=? id1 id2 [Scheme Procedure] free-identifier=? id1 id2

These predicate procedures operate on syntax objects representing Scheme identifiers. identifier? returns #t if obj represents an identifier, #f otherwise. boundidentifier=? returns #t if and only if a binding for id1 would capture a reference to id2 in the transformer’s output, or vice-versa. free-identifier=? returns #t if and only id1 and id2 would refer to the same binding in the output of the transformer, independent of any bindings introduced by the transformer.

[Scheme Procedure] generate-temporaries l

Returns a list, of the same length as l, which must be a list or a syntax object representing a list, of globally unique symbols.

[Scheme Procedure] syntax->datum syntax-object

[Scheme Procedure] datum->syntax template-id datum These procedures convert wrapped syntax objects to and from Scheme datum values. The syntax object returned by datum->syntax shares contextual information with the syntax object template-id.

[Scheme Procedure] syntax-violation whom message form

[Scheme Procedure] syntax-violation whom message form subform

Constructs a new compound condition that includes the following simple conditions: • If whom is not #f, a &who condition with the whom as its field

• A &message condition with the specified message

• A &syntax condition with the specified form and optional subform fields

7.6.2.23 rnrs hashtables

The (rnrs hashtables (6)) library provides structures and procedures for creating and accessing hash tables. The hash tables API defined by R6RS is substantially similar to both Guile’s native hash tables implementation as well as the one provided by SRFI-69; See Section 6.7.14 [Hash Tables], page 236, and Section 7.5.38 [SRFI-69], page 618, respectively. Note that you can write portable R6RS library code that manipulates SRFI-69 hash tables (by importing the (srfi :69) library); however, hash tables created by one API cannot be used by another.

Like SRFI-69 hash tables—and unlike Guile’s native ones—R6RS hash tables associate hash and equality functions with a hash table at the time of its creation. Additionally, R6RS allows for the creation (via hashtable-copy; see below) of immutable hash tables.

[Scheme Procedure] make-eq-hashtable

[Scheme Procedure] make-eq-hashtable k

Returns a new hash table that uses eq? to compare keys and Guile’s hashq procedure as a hash function. If k is given, it specifies the initial capacity of the hash table.

[Scheme Procedure] make-eqv-hashtable
[Scheme Procedure] make-eqv-hashtable k

Returns a new hash table that uses eqv? to compare keys and Guile’s hashv procedure as a hash function. If k is given, it specifies the initial capacity of the hash table.

[Scheme Procedure] make-hashtable hash-function equiv [Scheme Procedure] make-hashtable hash-function equiv k

Returns a new hash table that uses equiv to compare keys and hash-function as a hash function. equiv must be a procedure that accepts two arguments and returns a true value if they are equivalent, #f otherwise; hash-function must be a procedure that accepts one argument and returns a non-negative integer. If k is given, it specifies the initial capacity of the hash table.

[Scheme Procedure] hashtable? obj

Returns #t if obj is an R6RS hash table, #f otherwise.

[Scheme Procedure] hashtable-size hashtable

Returns the number of keys currently in the hash table hashtable.

[Scheme Procedure] hashtable-ref hashtable key default Returns the value associated with key in the hash table hashtable, or default if none is found.

[Scheme Procedure] hashtable-set! hashtable key obj Associates the key key with the value obj in the hash table hashtable, and returns an unspecified value. An &assertion condition is raised if hashtable is immutable.

[Scheme Procedure] hashtable-delete! hashtable key

Removes any association found for the key key in the hash table hashtable, and returns an unspecified value. An &assertion condition is raised if hashtable is immutable.

[Scheme Procedure] hashtable-contains? hashtable key

Returns #t if the hash table hashtable contains an association for the key key, #f otherwise.

[Scheme Procedure] hashtable-update! hashtable key proc default Associates with key in the hash table hashtable the result of calling proc, which must be a procedure that takes one argument, on the value currently associated key in hashtable—or on default if no such association exists. An &assertion condition is raised if hashtable is immutable.

[Scheme Procedure] hashtable-copy hashtable

[Scheme Procedure] hashtable-copy hashtable mutable Returns a copy of the hash table hashtable. If the optional argument mutable is provided and is a true value, the new hash table will be mutable.

[Scheme Procedure] hashtable-clear! hashtable [Scheme Procedure] hashtable-clear! hashtable k

Removes all of the associations from the hash table hashtable. The optional argument k, which specifies a new capacity for the hash table, is accepted by Guile’s (rnrs hashtables) implementation, but is ignored.

[Scheme Procedure] hashtable-keys hashtable

Returns a vector of the keys with associations in the hash table hashtable, in an unspecified order.

[Scheme Procedure] hashtable-entries hashtable

Return two values—a vector of the keys with associations in the hash table hashtable, and a vector of the values to which these keys are mapped, in corresponding but unspecified order.

[Scheme Procedure] hashtable-equivalence-function hashtable

Returns the equivalence predicated use by hashtable. This procedure returns eq? and eqv?, respectively, for hash tables created by make-eq-hashtable and make-

eqv-hashtable.

[Scheme Procedure] hashtable-hash-function hashtable

Returns the hash function used by hashtable. For hash tables created by make-eq-

hashtable or make-eqv-hashtable, #f is returned.

[Scheme Procedure] hashtable-mutable? hashtable

Returns #t if hashtable is mutable, #f otherwise.

A number of hash functions are provided for convenience:

[Scheme Procedure] equal-hash obj

Returns an integer hash value for obj, based on its structure and current contents. This hash function is suitable for use with equal? as an equivalence function.

[Scheme Procedure] string-hash string [Scheme Procedure] symbol-hash symbol

These procedures are identical to the ones provided by Guile’s core library. See Section 6.7.14.2 [Hash Table Reference], page 237, for documentation.

[Scheme Procedure] string-ci-hash string

Returns an integer hash value for string based on its contents, ignoring case. This hash function is suitable for use with string-ci=? as an equivalence function.

7.6.2.24 rnrs enums

The (rnrs enums (6)) library provides structures and procedures for working with enumerable sets of symbols. Guile’s implementation defines an enum-set record type that encapsulates a finite set of distinct symbols, the universe, and a subset of these symbols, which define the enumeration set.

The SRFI-1 list library provides a number of procedures for performing set operations on lists; Guile’s (rnrs enums) implementation makes use of several of them. See Section 7.5.3.10 [SRFI-1 Set Operations], page 563, for more information.

[Scheme Procedure] make-enumeration symbol-list

Returns a new enum-set whose universe and enumeration set are both equal to symbollist, a list of symbols.

[Scheme Procedure] enum-set-universe enum-set

Returns an enum-set representing the universe of enum-set, an enum-set.

[Scheme Procedure] enum-set-indexer enum-set

Returns a procedure that takes a single argument and returns the zero-indexed position of that argument in the universe of enum-set, or #f if its argument is not a member of that universe.

[Scheme Procedure] enum-set-constructor enum-set

Returns a procedure that takes a single argument, a list of symbols from the universe of enum-set, an enum-set, and returns a new enum-set with the same universe that represents a subset containing the specified symbols.

[Scheme Procedure] enum-set->list enum-set

Returns a list containing the symbols of the set represented by enum-set, an enum-set, in the order that they appear in the universe of enum-set.

[Scheme Procedure] enum-set-member? symbol enum-set

[Scheme Procedure] enum-set-subset? enum-set1 enum-set2 [Scheme Procedure] enum-set=? enum-set1 enum-set2

These procedures test for membership of symbols and enum-sets in other enum-sets. enum-set-member? returns #t if and only if symbol is a member of the subset specified by enum-set. enum-set-subset? returns #t if and only if the universe of enum-set1 is a subset of the universe of enum-set2 and every symbol in enum-set1 is present in enum-set2. enum-set=? returns #t if and only if enum-set1 is a subset, as per enum-set-subset? of enum-set2 and vice versa.

[Scheme Procedure] enum-set-union enum-set1 enum-set2

[Scheme Procedure] enum-set-intersection enum-set1 enum-set2 [Scheme Procedure] enum-set-difference enum-set1 enum-set2

These procedures return, respectively, the union, intersection, and difference of their enum-set arguments.

[Scheme Procedure] enum-set-complement enum-set

Returns enum-set’s complement (an enum-set), with regard to its universe.

[Scheme Procedure] enum-set-projection enum-set1 enum-set2

Returns the projection of the enum-set enum-set1 onto the universe of the enum-set enum-set2.

[Scheme Syntax] define-enumeration type-name (symbol ...) constructor-syntax Evaluates to two new definitions: A constructor bound to constructor-syntax that behaves similarly to constructors created by enum-set-constructor, above, and creates new enum-sets in the universe specified by (symbol ...); and a “predicate macro” bound to type-name, which has the following form:

     (type-name sym)

If sym is a member of the universe specified by the symbols above, this form evaluates to sym. Otherwise, a &syntax condition is raised.

7.6.2.25 rnrs

The (rnrs (6)) library is a composite of all of the other R6RS standard libraries—it imports and re-exports all of their exported procedures and syntactic forms—with the exception of the following libraries:

  • (rnrs eval (6))
  • (rnrs mutable-pairs (6))
  • (rnrs mutable-strings (6))
  • (rnrs r5rs (6))

7.6.2.26 rnrs eval

The (rnrs eval (6) library provides procedures for performing “on-the-fly” evaluation of expressions.

[Scheme Procedure] eval expression environment Evaluates expression, which must be a datum representation of a valid Scheme expression, in the environment specified by environment. This procedure is identical to the one provided by Guile’s code library; See Section 6.17.4 [Fly Evaluation], page 362, for documentation.

[Scheme Procedure] environment import-spec ...

Constructs and returns a new environment based on the specified import-specs, which must be datum representations of the import specifications used with the import form. See Section 6.19.6 [R6RS Libraries], page 388, for documentation.

7.6.2.27 rnrs mutable-pairs

The (rnrs mutable-pairs (6)) library provides the set-car! and set-cdr! procedures,

which allow the car and cdr fields of a pair to be modified.

These procedures are identical to the ones provide by Guile’s core library. See Section 6.7.1 [Pairs], page 185, for documentation. All pairs in Guile are mutable; consequently, these procedures will never throw the &assertion condition described in the R6RS libraries specification.

7.6.2.28 rnrs mutable-strings

The (rnrs mutable-strings (6)) library provides the string-set! and string-fill!

procedures, which allow the content of strings to be modified “in-place.”

These procedures are identical to the ones provided by Guile’s core library. See Section 6.6.5.6 [String Modification], page 147, for documentation. All strings in Guile are mutable; consequently, these procedures will never throw the &assertion condition described in the R6RS libraries specification.

7.6.2.29 rnrs r5rs

The (rnrs r5rs (6)) library exports bindings for some procedures present in R5RS but omitted from the R6RS base library specification.

[Scheme Procedure] exact->inexact z [Scheme Procedure] inexact->exact z

These procedures are identical to the ones provided by Guile’s core library. See Section 6.6.2.5 [Exactness], page 113, for documentation.

[Scheme Procedure] quotient n1 n2 [Scheme Procedure] remainder n1 n2 [Scheme Procedure] modulo n1 n2

These procedures are identical to the ones provided by Guile’s core library. See Section 6.6.2.7 [Integer Operations], page 116, for documentation.

[Scheme Syntax] delay expr

[Scheme Procedure] force promise The delay form and the force procedure are identical to their counterparts in Guile’s core library. See Section 6.17.9 [Delayed Evaluation], page 371, for documentation.

[Scheme Procedure] null-environment n

[Scheme Procedure] scheme-report-environment n

These procedures are identical to the ones provided by the (ice-9 r5rs) Guile module. See Section 6.19.11 [Environments], page 395, for documentation.

7.7 Pattern Matching

The (ice-9 match) module provides a pattern matcher, written by Alex Shinn, and compatible with Andrew K. Wright’s pattern matcher found in many Scheme implementations.

A pattern matcher can match an object against several patterns and extract the elements that make it up. Patterns can represent any Scheme object: lists, strings, symbols, records, etc. They can optionally contain pattern variables. When a matching pattern is found, an expression associated with the pattern is evaluated, optionally with all pattern variables bound to the corresponding elements of the object:

     (let ((l '(hello (world))))
       (match l           ;; <- the input object
         (('hello (who))  ;; <- the pattern
          who)))          ;; <- the expression evaluated upon matching
     ⇒ world

In this example, list l matches the pattern ('hello (who)), because it is a two-element list whose first element is the symbol hello and whose second element is a one-element list. Here who is a pattern variable. match, the pattern matcher, locally binds who to the value contained in this one-element list—i.e., the symbol world. An error would be raised if l did not match the pattern.

The same object can be matched against a simpler pattern:

     (let ((l '(hello (world))))
       (match l
         ((x y)
          (values x y))))
     ⇒ hello

     ⇒ (world)

Here pattern (x y) matches any two-element list, regardless of the types of these elements. Pattern variables x and y are bound to, respectively, the first and second element of l.

Patterns can be composed, and nested. For instance, ... (ellipsis) means that the previous pattern may be matched zero or more times in a list:

     (match lst
       (((heads tails ...) ...)
        heads))

This expression returns the first element of each list within lst. For proper lists of proper lists, it is equivalent to (map car lst). However, it performs additional checks to make sure that lst and the lists therein are proper lists, as prescribed by the pattern, raising an error if they are not.

Compared to hand-written code, pattern matching noticeably improves clarity and conciseness—no need to resort to series of car and cdr calls when matching lists, for instance. It also improves robustness, by making sure the input completely matches the pattern—conversely, hand-written code often trades robustness for conciseness. And of course, match is a macro, and the code it expands to is just as efficient as equivalent handwritten code.

The pattern matcher is defined as follows:

[Scheme Syntax] match exp clause1 clause2 … Match object exp against the patterns in clause1 clause2 … in the order in which they appear. Return the value produced by the first matching clause. If no clause matches, throw an exception with key match-error. Each clause has the form (pattern body1 body2 ...). Each pattern must follow the syntax described below. Each body is an arbitrary Scheme expression, possibly referring to pattern variables of pattern.

The syntax and interpretation of patterns is as follows:

        patterns:                       matches:
pat ::= identifier                      anything, and binds identifier
      | _anything
      | ()                              the empty list
      | #t                              #t
      | #f                              #f
      | string                          a string
      | number                          a number
      | character                       a character
      | 'sexp                           an s-expression
      | 'symbol                         a symbol (special case of s-expr)
      | (pat_1 ... pat_n)               list of n elements
      | (pat_1 ... pat_n . pat_{n+1})   list of n or more
      | (pat_1 ... pat_n pat_n+1 ooo)   list of n or more, each element
                                          of remainder must match pat_n+1
      | #(pat_1 ... pat_n)              vector of n elements

| #(pat_1 ... pat_n pat_n+1 ooo)  vector of n or more, each element
                                    of remainder must match pat_n+1
| #&pat                           box
| ($ record-name pat_1 ... pat_n) a record
| (= field pat)                   a ``field'' of an object
| (and pat_1 ... pat_n)           if all of pat_1 thru pat_n match
| (or pat_1 ... pat_n)            if any of pat_1 thru pat_n match
| (not pat_1 ... pat_n)           if all pat_1 thru pat_n don't match
| (? predicate pat_1 ... pat_n)   if predicate true and all of
                                    pat_1 thru pat_n match
| (set! identifier)               anything, and binds setter
| (get! identifier)               anything, and binds getter
| `qp                             a quasi-pattern
| (identifier *** pat)            matches pat in a tree and binds
                                  identifier to the path leading
                                  to the object that matches pat
ooo ::= ...                             zero or more
| ___zero or more
| ..1                             1 or more
  quasi-patterns:                 matches:
qp  ::= ()                              the empty list
| #t                              #t
| #f                              #f
| string                          a string
| number                          a number
| character                       a character
| identifier                      a symbol
| (qp_1 ... qp_n)                 list of n elements
| (qp_1 ... qp_n . qp_{n+1})      list of n or more
| (qp_1 ... qp_n qp_n+1 ooo)      list of n or more, each element
                                    of remainder must match qp_n+1
| #(qp_1 ... qp_n)                vector of n elements
| #(qp_1 ... qp_n qp_n+1 ooo)     vector of n or more, each element
                                    of remainder must match qp_n+1
| #&qp                            box
| ,pat                            a pattern
| ,@pat                           a pattern
The names quote, quasiquote, unquote, unquote-splicing, ?, _, $, and, or, not,

set!, get!, ..., and ___cannot be used as pattern variables.

Here is a more complex example:

(use-modules (srfi srfi-9))
(let ()

       (define-record-type person
         (make-person name friends)
         person?
         (name    person-name)
         (friends person-friends))
       (letrec ((alice (make-person "Alice" (delay (list bob))))
                (bob   (make-person "Bob" (delay (list alice)))))
         (match alice
           (($ person name (= force (($ person "Bob"))))
            (list 'friend-of-bob name))
           (_#f))))
     ⇒ (friend-of-bob "Alice")

Here the $ pattern is used to match a SRFI-9 record of type person containing two or more slots. The value of the first slot is bound to name. The = pattern is used to apply force on the second slot, and then checking that the result matches the given pattern. In other words, the complete pattern matches any person whose second slot is a promise that evaluates to a one-element list containing a person whose first slot is "Bob".

Please refer to the ice-9/match.upstream.scm file in your Guile installation for more details.

Guile also comes with a pattern matcher specifically tailored to SXML trees, See Section 7.16 [sxml-match], page 687.

7.8 Readline Support

Guile comes with an interface module to the readline library (see GNU Readline Library). This makes interactive use much more convenient, because of the command-line editing features of readline. Using (ice-9 readline), you can navigate through the current input line with the cursor keys, retrieve older command lines from the input history and even search through the history entries.

7.8.1 Loading Readline Support

The module is not loaded by default and so has to be loaded and activated explicitly. This is done with two simple lines of code:

     (use-modules (ice-9 readline))
     (activate-readline)

The first line will load the necessary code, and the second will activate readline’s features for the REPL. If you plan to use this module often, you should save these to lines to your .guile personal startup file.

You will notice that the REPL’s behaviour changes a bit when you have loaded the readline module. For example, when you press Enter before typing in the closing parentheses of a list, you will see the continuation prompt, three dots: ... This gives you a nice visual feedback when trying to match parentheses. To make this even easier, bouncing parentheses are implemented. That means that when you type in a closing parentheses, the cursor will jump to the corresponding opening parenthesis for a short time, making it trivial to make them match.

Once the readline module is activated, all lines entered interactively will be stored in a history and can be recalled later using the cursor-up and -down keys. Readline also understands the Emacs keys for navigating through the command line and history.

When you quit your Guile session by evaluating (quit) or pressing Ctrl-D, the history will be saved to the file .guile_history and read in when you start Guile for the next time. Thus you can start a new Guile session and still have the (probably long-winded) definition expressions available.

You can specify a different history file by setting the environment variable GUILE_ HISTORY. And you can make Guile specific customizations to your .inputrc by testing for application ‘Guile’ (see Section “Conditional Init Constructs” in GNU Readline Library). For instance to define a key inserting a matched pair of parentheses,

     $if Guile
       "\C-o": "()\C-b"
     $endif

7.8.2 Readline Options

The readline interface module can be tweaked in a few ways to better suit the user’s needs. Configuration is done via the readline module’s options interface, in a similar way to the evaluator and debugging options (see Section 6.22.3 [Runtime Options], page 431).

[Scheme Procedure] readline-options [Scheme Procedure] readline-enable option-name [Scheme Procedure] readline-disable option-name

[Scheme Syntax] readline-set! option-name value Accessors for the readline options. Note that unlike the enable/disable procedures, readline-set! is syntax, which expects an unquoted option name.

Here is the list of readline options generated by typing (readline-options 'help) in Guile. You can also see the default values.

      history-file    yes     Use history file.
      history-length  200     History length.
      bounce-parens   500     Time (ms) to show matching opening parenthesis
                              (0 = off).

The readline options interface can only be used after loading the readline module, because it is defined in that module.

7.8.3 Readline Functions

The following functions are provided by

     (use-modules (ice-9 readline))

There are two ways to use readline from Scheme code, either make calls to readline directly to get line by line input, or use the readline port below with all the usual reading functions.

[Function] readline [prompt]

Read a line of input from the user and return it as a string (without a newline at the end). prompt is the prompt to show, or the default is the string set in set-readline-

prompt! below.
     (readline "Type something: ") ⇒ "hello"
[Function] set-readline-input-port! port
[Function] set-readline-output-port! port

Set the input and output port the readline function should read from and write to. port must be a file port (see Section 6.14.9.1 [File Ports], page 326), and should usually be a terminal.

The default is the current-input-port and current-output-port (see

Section 6.14.8 [Default Ports], page 325) when (ice-9 readline) loads, which in an interactive user session means the Unix “standard input” and “standard output”.

7.8.3.1 Readline Port

[Function] readline-port

Return a buffered input port (see Section 7.14 [Buffered Input], page 684) which calls the readline function above to get input. This port can be used with all the usual reading functions (read, read-char, etc), and the user gets the interactive editing features of readline.

There’s only a single readline port created. readline-port creates it when first called, and on subsequent calls just returns what it previously made.

[Function] activate-readline

If the current-input-port is a terminal (see Section 7.2.9 [isatty?], page 499) then enable readline for all reading from current-input-port (see Section 6.14.8 [Default Ports], page 325) and enable readline features in the interactive REPL (see Section 3.3.3 [The REPL], page 25).

     (activate-readline)
     (read-char)
activate-readline enables readline on current-input-port simply by a
set-current-input-port to the readline-port above. An application can do that

directly if the extra REPL features that activate-readline adds are not wanted.

[Function] set-readline-prompt! prompt1 [prompt2]

Set the prompt string to print when reading input. This is used when reading through readline-port, and is also the default prompt for the readline function above.

prompt1 is the initial prompt shown. If a user might enter an expression across multiple lines, then prompt2 is a different prompt to show further input required. In the Guile REPL for instance this is an ellipsis (‘...’).

See set-buffered-input-continuation?! (see Section 7.14 [Buffered Input],

page 684) for an application to indicate the boundaries of logical expressions (assuming of course an application has such a notion).

7.8.3.2 Completion

[Function] with-readline-completion-function completer thunk

Call (thunk) with completer as the readline tab completion function to be used in any readline calls within that thunk. completer can be #f for no completion. completer will be called as (completer text state), as described in (see Section “How Completing Works” in GNU Readline Library). text is a partial word to be completed, and each completer call should return a possible completion string or #f when no more. state is #f for the first call asking about a new text then #t while getting further completions of that text. Here’s an example completer for user login names from the password file (see Section 7.2.4 [User Information], page 483), much like readline’s own rl_username_

completion_function,
     (define (username-completer-function text state)
       (if (not state)
           (setpwent))  ;; new, go to start of database
       (let more ((pw (getpwent)))
         (if pw
             (if (string-prefix? text (passwd:name pw))
                 (passwd:name pw)     ;; this name matches, return it
                 (more (getpwent)))   ;; doesn't match, look at next
             (begin
               ;; end of database, close it and return #f
               (endpwent)
               #f))))
[Function] apropos-completion-function text state

A completion function offering completions for Guile functions and variables (all defines). This is the default completion function.

[Function] filename-completion-function text state

A completion function offering filename completions. This is readline’s rl_filename_ completion_function (see Section “Completion Functions” in GNU Readline Library).

[Function] make-completion-function string-list

Return a completion function which offers completions from the possibilities in stringlist. Matching is case-sensitive.

7.9 Pretty Printing

The module (ice-9 pretty-print) provides the procedure pretty-print, which provides nicely formatted output of Scheme objects. This is especially useful for deeply nested or complex data structures, such as lists and vectors.

The module is loaded by entering the following:

(use-modules (ice-9 pretty-print))

This makes the procedure pretty-print available. As an example how pretty-print will format the output, see the following:

(pretty-print '(define (foo) (lambda (x)
(cond ((zero? x) #t) ((negative? x) -x) (else
(if (= x 1) 2 (* x x x)))))))

⊣

(define (foo)
  (lambda (x)
    (cond ((zero? x) #t)
          ((negative? x) -x)
          (else (if (= x 1) 2 (* x x x))))))

[Scheme Procedure] pretty-print obj [port] [keyword-options] Print the textual representation of the Scheme object obj to port. port defaults to the current output port, if not given.

The further keyword-options are keywords and parameters as follows,

#:display? flag

If flag is true then print using display. The default is #f which means use write style. (see Section 6.14.3 [Writing], page 320)

#:per-line-prefix string

Print the given string as a prefix on each line. The default is no prefix.

#:width columns

Print within the given columns. The default is 79.

Also exported by the (ice-9 pretty-print) module is truncated-print, a procedure to print Scheme datums, truncating the output to a certain number of characters. This is useful when you need to present an arbitrary datum to the user, but you only have one line in which to do so.

(define exp '(a b #(c d e) f . g))
(truncated-print exp #:width 10) (newline)
⊣ (a b . #)
(truncated-print exp #:width 15) (newline)
⊣ (a b # f . g)
(truncated-print exp #:width 18) (newline)
⊣ (a b #(c ...) . #)
(truncated-print exp #:width 20) (newline)
⊣ (a b #(c d e) f . g)
(truncated-print "The quick brown fox" #:width 20) (newline)
⊣ "The quick brown..."
(truncated-print (current-module) #:width 20) (newline)
⊣ #<directory (gui...>

truncated-print will not output a trailing newline. If an expression does not fit in the given width, it will be truncated – possibly ellipsized^{3}, or in the worst case, displayed as #.

3 On Unicode-capable ports, the ellipsis is represented by character ‘HORIZONTAL ELLIPSIS’ (U+2026),

otherwise it is represented by three dots.

[Scheme Procedure] truncated-print obj [port] [keyword-options] Print obj, truncating the output, if necessary, to make it fit into width characters. By default, obj will be printed using write, though that behavior can be overridden via the display? keyword argument. The default behaviour is to print depth-first, meaning that the entire remaining width will be available to each sub-expression of obj – e.g., if obj is a vector, each member of obj. One can attempt to “ration” the available width, trying to allocate it equally to each sub-expression, via the breadth-first? keyword argument. The further keyword-options are keywords and parameters as follows,

#:display? flag

If flag is true then print using display. The default is #f which means use write style. (see Section 6.14.3 [Writing], page 320)

#:width columns

Print within the given columns. The default is 79.

#:breadth-first? flag

If flag is true, then allocate the available width breadth-first among elements of a compound data structure (list, vector, pair, etc.). The default is #f which means that any element is allowed to consume all of the available width.

7.10 Formatted Output

The format function is a powerful way to print numbers, strings and other objects together with literal text under the control of a format string. This function is available from

(use-modules (ice-9 format))

A format string is generally more compact and easier than using just the standard procedures like display, write and newline. Parameters in the output string allow various output styles, and parameters can be taken from the arguments for runtime flexibility.

format is similar to the Common Lisp procedure of the same name, but it’s not identical and doesn’t have quite all the features found in Common Lisp.

C programmers will note the similarity between format and printf, though escape sequences are marked with ~ instead of %, and are more powerful.

[Scheme Procedure] format dest fmt arg … Write output specified by the fmt string to dest. dest can be an output port, #t for current-output-port (see Section 6.14.8 [Default Ports], page 325), or #f to return the output as a string. fmt can contain literal text to be output, and ~ escapes. Each escape has the form

     ~ [param [, param...] [:] [@] code

code is a character determining the escape sequence. The : and @ characters are optional modifiers, one or both of which change the way various codes operate. Optional parameters are accepted by some codes too. Parameters have the following forms,

[+/-]number

An integer, with optional + or -.

' (apostrophe)

The following character in the format string, for instance 'z for z.

v The next function argument as the parameter. v stands for “variable”, a parameter can be calculated at runtime and included in the arguments. Upper case V can be used too.

# The number of arguments remaining. (See ~* below for some usages.)

Parameters are separated by commas (,). A parameter can be left empty to keep its default value when supplying later parameters.

The following escapes are available. The code letters are not case-sensitive, upper and lower case are the same.

~a

~s Object output. Parameters: minwidth, padinc, minpad, padchar. ~a outputs an argument like display, ~s outputs an argument like write (see Section 6.14.3 [Writing], page 320).

     (format #t "~a" "foo") ⊣ foo
     (format #t "~s" "foo") ⊣ "foo"

~:a and ~:s put objects that don’t have an external representation in quotes like a string.

     (format #t "~:a" car) ⊣ "#<primitive-procedure car>"

If the output is less than minwidth characters (default 0), it’s padded on the right with padchar (default space). ~@a and ~@s put the padding on the left instead.

     (format #f "~5a" 'abc)  ⇒ "abc  "
     (format #f "~5,,,'-@a" 'abc) ⇒ "--abc"

minpad is a minimum for the padding then plus a multiple of padinc. Ie. the padding is minpad+N ∗padinc, where n is the smallest integer making the total object plus padding greater than or equal to minwidth. The default minpad is 0 and the default padinc is 1 (imposing no minimum or multiple).

     (format #f "~5,1,4a" 'abc) ⇒ "abc    "

~c Character. Parameter: charnum.

Output a character. The default is to simply output, as per write-char (see Section 6.14.3 [Writing], page 320). ~@c prints in write style. ~:c prints control characters (ASCII 0 to 31) in ^X form.

     (format #t "~c" #\z)  ⊣ z
     (format #t "~@c" #\z)  ⊣ #\z
     (format #t "~:c" #\newline) ⊣ ^J

If the charnum parameter is given then an argument is not taken but instead the character is (integer->char charnum) (see Section 6.6.3 [Characters], page 129). This can be used for instance to output characters given by their ASCII code.

     (format #t "~65c") ⊣ A
~d
~x
~o

~b Integer. Parameters: minwidth, padchar, commachar, commawidth. Output an integer argument as a decimal, hexadecimal, octal or binary integer (respectively), in a locale-independent way.

     (format #t "~d" 123) ⊣ 123

~@d etc shows a + sign is shown on positive numbers.

     (format #t "~@b" 12) ⊣ +1100

If the output is less than the minwidth parameter (default no minimum), it’s padded on the left with the padchar parameter (default space).

     (format #t "~5,'*d" 12) ⊣ ***12
     (format #t "~5,'0d" 12) ⊣ 00012
     (format #t "~3d"    1234) ⊣ 1234

~:d adds commas (or the commachar parameter) every three digits (or the commawidth parameter many). However, when your intent is to write numbers in a way that follows typographical conventions, using ~h is recommended.

     (format #t "~:d" 1234567)  ⊣ 1,234,567
     (format #t "~10,'*,'/,2:d" 12345) ⊣ ***1/23/45

Hexadecimal ~x output is in lower case, but the ~( and ~) case conversion directives described below can be used to get upper case.

     (format #t "~x"       65261) ⊣ feed
     (format #t "~:@(~x~)" 65261) ⊣ FEED

~r Integer in words, roman numerals, or a specified radix. Parameters: radix, minwidth, padchar, commachar, commawidth. With no parameters output is in words as a cardinal like “ten”, or ~:r prints an ordinal like “tenth”.

     (format #t "~r" 9) ⊣ nine        ;; cardinal
     (format #t "~r" -9) ⊣ minus nine  ;; cardinal
     (format #t "~:r" 9) ⊣ ninth       ;; ordinal

And also with no parameters, ~@r gives roman numerals and ~:@r gives old roman numerals. In old roman numerals there’s no “subtraction”, so 9 is VIIII instead of IX. In both cases only positive numbers can be output.

     (format #t "~@r" 89) ⊣ LXXXIX     ;; roman
     (format #t "~:@r" 89) ⊣ LXXXVIIII  ;; old roman

When a parameter is given it means numeric output in the specified radix. The modifiers and parameters following the radix are the same as described for ~d etc above.

     (format #f "~3r" 27) ⇒ "1000"    ;; base 3
     (format #f "~3,5r" 26) ⇒ "  222"   ;; base 3 width 5

~f Fixed-point float. Parameters: width, decimals, scale, overflowchar, padchar. Output a number or number string in fixed-point format, ie. with a decimal point.

     (format #t "~f" 5)  ⊣ 5.0
     (format #t "~f" "123") ⊣ 123.0
     (format #t "~f" "1e-1") ⊣ 0.1

~@f prints a + sign on positive numbers (including zero).

     (format #t "~@f" 0) ⊣ +0.0

If the output is less than width characters it’s padded on the left with padchar (space by default). If the output equals or exceeds width then there’s no padding. The default for width is no padding.

     (format #f "~6f" -1.5)  ⇒ "  -1.5"
     (format #f "~6,,,,'*f" 23) ⇒ "**23.0"
     (format #f "~6f" 1234567.0) ⇒ "1234567.0"

decimals is how many digits to print after the decimal point, with the value rounded or padded with zeros as necessary. (The default is to output as many decimals as required.)

     (format #t "~1,2f" 3.125) ⊣ 3.13
     (format #t "~1,2f" 1.5) ⊣ 1.50

scale is a power of 10 applied to the value, moving the decimal point that many places. A positive scale increases the value shown, a negative decreases it.

     (format #t "~,,2f" 1234) ⊣ 123400.0
     (format #t "~,,-2f" 1234) ⊣ 12.34

If overflowchar and width are both given and if the output would exceed width, then that many overflowchars are printed instead of the value.

     (format #t "~6,,,'xf" 12345) ⊣ 12345.
     (format #t "~5,,,'xf" 12345) ⊣ xxxxx

~h Localized number^{4}. Parameters: width, decimals, padchar.

Like ~f, output an exact or floating point number, but do so according to the current locale, or according to the given locale object when the : modifier is used (see Section 6.24.4 [Number Input and Output], page 439).

     (format #t "~h" 12345.5678)  ; with "C" as the current locale
     ⊣ 12345.5678
     (format #t "~14,,'*:h" 12345.5678
             (make-locale LC_ALL "en_US"))
     ⊣ ***12,345.5678
     (format #t "~,2:h" 12345.5678

4 The ~h format specifier first appeared in Guile version 2.0.6.

             (make-locale LC_NUMERIC "fr_FR"))
     ⊣ 12 345,56

~e Exponential float. Parameters: width, mantdigits, expdigits, intdigits, overflowchar, padchar, expchar. Output a number or number string in exponential notation.

     (format #t "~e" 5000.25) ⊣ 5.00025E+3
     (format #t "~e" "123.4") ⊣ 1.234E+2
     (format #t "~e" "1e4") ⊣ 1.0E+4

~@e prints a + sign on positive numbers (including zero). (This is for the mantissa, a + or - sign is always shown on the exponent.)

     (format #t "~@e" 5000.0) ⊣ +5.0E+3

If the output is less than width characters it’s padded on the left with padchar (space by default). The default for width is to output with no padding.

     (format #f "~10e" 1234.0)  ⇒ "  1.234E+3"
     (format #f "~10,,,,,'*e" 0.5) ⇒ "****5.0E-1"

mantdigits is the number of digits shown in the mantissa after the decimal point. The value is rounded or trailing zeros are added as necessary. The default mantdigits is to show as much as needed by the value.

     (format #f "~,3e" 11111.0) ⇒ "1.111E+4"
     (format #f "~,8e" 123.0) ⇒ "1.23000000E+2"

expdigits is the minimum number of digits shown for the exponent, with leading zeros added if necessary. The default for expdigits is to show only as many digits as required. At least 1 digit is always shown.

     (format #f "~,,1e" 1.0e99) ⇒ "1.0E+99"
     (format #f "~,,6e" 1.0e99) ⇒ "1.0E+000099"

intdigits (default 1) is the number of digits to show before the decimal point in the mantissa. intdigits can be zero, in which case the integer part is a single 0, or it can be negative, in which case leading zeros are shown after the decimal point.

     (format #t "~,,,3e" 12345.0) ⊣ 123.45E+2
     (format #t "~,,,0e" 12345.0) ⊣ 0.12345E+5
     (format #t "~,,,-3e" 12345.0) ⊣ 0.00012345E+8

If overflowchar is given then width is a hard limit. If the output would exceed width then instead that many overflowchars are printed.

     (format #f "~6,,,,'xe" 100.0) ⇒ "1.0E+2"
     (format #f "~3,,,,'xe" 100.0) ⇒ "xxx"

expchar is the exponent marker character (default E).

     (format #t "~,,,,,,'ee" 100.0) ⊣ 1.0e+2

~g General float. Parameters: width, mantdigits, expdigits, intdigits, overflowchar, padchar, expchar.

Output a number or number string in either exponential format the same as ~e, or fixed-point format like ~f but aligned where the mantissa would have been and followed by padding where the exponent would have been. Fixed-point is used when the absolute value is 0.1 or more and it takes no more space than the mantissa in exponential format, ie. basically up to mantdigits digits.

     (format #f "~12,4,2g" 999.0)  ⇒ "   999.0    "
     (format #f "~12,4,2g" "100000") ⇒ "  1.0000E+05"

The parameters are interpreted as per ~e above. When fixed-point is used, the decimals parameter to ~f is established from mantdigits, so as to give a total mantdigits + 1 figures.

~$ Monetary style fixed-point float. Parameters: decimals, intdigits, width, padchar. Output a number or number string in fixed-point format, ie. with a decimal point. decimals is the number of decimal places to show, default 2.

     (format #t "~$" 5)  ⊣ 5.00
     (format #t "~4$" "2.25") ⊣ 2.2500
     (format #t "~4$" "1e-2") ⊣ 0.0100

~@$ prints a + sign on positive numbers (including zero).

     (format #t "~@$" 0) ⊣ +0.00

intdigits is a minimum number of digits to show in the integer part of the value (default 1).

     (format #t "~,3$" 9.5) ⊣ 009.50

     (format #t "~,0$" 0.125) ⊣ .13

If the output is less than width characters (default 0), it’s padded on the left with padchar (default space). ~:$ puts the padding after the sign.

     (format #f "~,,8$" -1.5) ⇒ "   -1.50"

     (format #f "~,,8:$" -1.5) ⇒ "-   1.50"
     (format #f "~,,8,'.:@$" 3) ⇒ "+...3.00"

Note that floating point for dollar amounts is generally not a good idea, because a cent 0.01 cannot be represented exactly in the binary floating point Guile uses, which leads to slowly accumulating rounding errors. Keeping values as cents (or fractions of a cent) in integers then printing with the scale option in ~f may be a better approach.

~i Complex fixed-point float. Parameters: width, decimals, scale, overflowchar, padchar. Output the argument as a complex number, with both real and imaginary part shown (even if one or both are zero). The parameters and modifiers are the same as for fixed-point ~f described above. The real and imaginary parts are both output with the same given parameters and modifiers, except that for the imaginary part the @ modifier is always enabled, so as to print a + sign between the real and imaginary parts.

     (format #t "~i" 1) ⊣ 1.0+0.0i

~p Plural. No parameters.

Output nothing if the argument is 1, or ‘s’ for any other value.

     (format #t "enter name~p" 1) ⊣ enter name
     (format #t "enter name~p" 2) ⊣ enter names

~@p prints ‘y’ for 1 or ‘ies’ otherwise.

     (format #t "pupp~@p" 1) ⊣ puppy
     (format #t "pupp~@p" 2) ⊣ puppies

~:p re-uses the preceding argument instead of taking a new one, which can be convenient when printing some sort of count.

     (format #t "~d cat~:p" 9) ⊣ 9 cats
     (format #t "~d pupp~:@p" 5) ⊣ 5 puppies

~p is designed for English plurals and there’s no attempt to support other languages. ~[ conditionals (below) may be able to help. When using gettext to translate messages ngettext is probably best though (see Section 6.24 [Internationalization], page 436).

~y Structured printing. Parameters: width. ~y outputs an argument using pretty-print (see Section 7.9 [Pretty Printing], page 663). The result will be formatted to fit within width columns (79 by default), consuming multiple lines if necessary. ~@y outputs an argument using truncated-print (see Section 7.9 [Pretty Printing], page 663). The resulting code will be formatted to fit within width columns (79 by default), on a single line. The output will be truncated if necessary. ~:@y is like ~@y, except the width parameter is interpreted to be the maximum column to which to output. That is to say, if you are at column 10, and ~60:@y is seen, the datum will be truncated to 50 columns.

~?

~k Sub-format. No parameters.

Take a format string argument and a second argument which is a list of arguments for that string, and output the result.

     (format #t "~?" "~d ~d" '(1 2))  ⊣ 1 2

~@? takes arguments for the sub-format directly rather than in a list.

     (format #t "~@? ~s" "~d ~d" 1 2 "foo") ⊣ 1 2 "foo"

~? and ~k are the same, ~k is provided for T-Scheme compatibility.

~* Argument jumping. Parameter: N. Move forward N arguments (default 1) in the argument list. ~:* moves backwards. (N cannot be negative.)

     (format #f "~d ~2*~d" 1 2 3 4) ⇒ "1 4"
     (format #f "~d ~:*~d" 6)  ⇒ "6 6"

~@* moves to argument number N. The first argument is number 0 (and that’s the default for N).

     (format #f "~d~d again ~@*~d~d" 1 2) ⇒ "12 again 12"
     (format #f "~d~d~d ~1@*~d~d" 1 2 3) ⇒ "123 23"

A # move to the end followed by a : modifier move back can be used for an absolute position relative to the end of the argument list, a reverse of what the @ modifier does.

     (format #t "~#*~2:*~a" 'a 'b 'c 'd) ⊣ c

At the end of the format string the current argument position doesn’t matter, any further arguments are ignored.

~t Advance to a column position. Parameters: colnum, colinc, padchar. Output padchar (space by default) to move to the given colnum column. The start of the line is column 0, the default for colnum is 1.

     (format #f "~tX") ⇒ " X"
     (format #f "~3tX") ⇒ "   X"

If the current column is already past colnum, then the move is to there plus a multiple of colinc, ie. column colnum + N ∗ colinc for the smallest N which makes that value greater than or equal to the current column. The default colinc is 1 (which means no further move).

     (format #f "abcd~2,5,'.tx") ⇒ "abcd...x"

~@t takes colnum as an offset from the current column. colnum many pad characters are output, then further padding to make the current column a multiple of colinc, if it isn’t already so.

     (format #f "a~3,5'*@tx") ⇒ "a****x"

~t is implemented using port-column (see Section 6.14.2 [Reading], page 318), so it works even there has been other output before format.

~~ Tilde character. Parameter: n.

Output a tilde character ~, or n many if a parameter is given. Normally ~ introduces an escape sequence, ~~ is the way to output a literal tilde.

~% Newline. Parameter: n.

Output a newline character, or n many if a parameter is given. A newline (or a few newlines) can of course be output just by including them in the format string.

~& Start a new line. Parameter: n.

Output a newline if not already at the start of a line. With a parameter, output that many newlines, but with the first only if not already at the start of a line. So for instance 3 would be a newline if not already at the start of a line, and 2 further newlines.

~_Space character. Parameter: n.

Output a space character, or n many if a parameter is given. With a variable parameter this is one way to insert runtime calculated padding (~t or the various field widths can do similar things).

     (format #f "~v_foo" 4) ⇒ "    foo"

~/ Tab character. Parameter: n.

Output a tab character, or n many if a parameter is given.

~| Formfeed character. Parameter: n. Output a formfeed character, or n many if a parameter is given.

~! Force output. No parameters.

At the end of output, call force-output to flush any buffers on the destination (see Section 6.14.3 [Writing], page 320). ~! can occur anywhere in the format string, but the force is done at the end of output. When output is to a string (destination #f), ~! does nothing.

~newline (ie. newline character)

Continuation line. No parameters. Skip this newline and any following whitespace in the format string, ie. don’t send it to the output. This can be used to break up a long format string for readability, but not print the extra whitespace.

     (format #f "abc~
                 ~d def~
                 ~d" 1 2) ⇒ "abc1 def2"

~:newline skips the newline but leaves any further whitespace to be printed normally. ~@newline prints the newline then skips following whitespace.

~( ~) Case conversion. No parameters. Between ~( and ~) the case of all output is changed. The modifiers on ~( control the conversion.

~( — lower case. ~:@( — upper case.

For example,

     (format #t "~(Hello~)") ⊣ hello
     (format #t "~:@(Hello~)") ⊣ HELLO

In the future it’s intended the modifiers : and @ alone will capitalize the first letters of words, as per Common Lisp format, but the current implementation of this is flawed and not recommended for use. Case conversions do not nest, currently. This might change in the future, but if it does then it will be to Common Lisp style where the outermost conversion has priority, overriding inner ones (making those fairly pointless).

~{ ~} Iteration. Parameter: maxreps (for ~{). The format between ~{ and ~} is iterated. The modifiers to ~{ determine how arguments are taken. The default is a list argument with each iteration successively consuming elements from it. This is a convenient way to output a whole list.

     (format #t "~{~d~}"     '(1 2 3))  ⊣ 123
     (format #t "~{~s=~d ~}" '("x" 1 "y" 2)) ⊣ "x"=1 "y"=2

~:{ takes a single argument which is a list of lists, each of those contained lists gives the arguments for the iterated format.

     (format #t "~:{~dx~d ~}" '((1 2) (3 4) (5 6)))
     ⊣ 1x2 3x4 5x6

~@{ takes arguments directly, with each iteration successively consuming arguments.

     (format #t "~@{~d~}"     1 2 3)  ⊣ 123
     (format #t "~@{~s=~d ~}" "x" 1 "y" 2) ⊣ "x"=1 "y"=2

~:@{ takes list arguments, one argument for each iteration, using that list for the format.

     (format #t "~:@{~dx~d ~}" '(1 2) '(3 4) '(5 6))
     ⊣ 1x2 3x4 5x6

Iterating stops when there are no more arguments or when the maxreps parameter to ~{ is reached (default no maximum).

     (format #t "~2{~d~}" '(1 2 3 4)) ⊣ 12

If the format between ~{ and ~} is empty, then a format string argument is taken (before iteration argument(s)) and used instead. This allows a sub-format (like ~? above) to be iterated.

     (format #t "~{~}" "~d" '(1 2 3)) ⊣ 123

Iterations can be nested, an inner iteration operates in the same way as described, but of course on the arguments the outer iteration provides it. This can be used to work into nested list structures. For example in the following the inner ~{~d~}x is applied to (1 2) then (3 4 5) etc.

     (format #t "~{~{~d~}x~}" '((1 2) (3 4 5))) ⊣ 12x345x

See also ~^below for escaping from iteration.

~[ ~; ~] Conditional. Parameter: selector. A conditional block is delimited by ~[ and ~], and ~; separates clauses within the block. ~[ takes an integer argument and that number clause is used. The first clause is number 0.

     (format #f "~[peach~;banana~;mango~]" 1) ⇒ "banana"

The selector parameter can be used for the clause number, instead of taking an argument.

     (format #f "~2[peach~;banana~;mango~]") ⇒ "mango"

If the clause number is out of range then nothing is output. Or the last clause can be ~:; to use that for a number out of range.

     (format #f "~[banana~;mango~]"         99) ⇒ ""
     (format #f "~[banana~;mango~:;fruit~]" 99) ⇒ "fruit"

~:[ treats the argument as a flag, and expects two clauses. The first is used if the argument is #f or the second otherwise.

     (format #f "~:[false~;not false~]" #f) ⇒ "false"
     (format #f "~:[false~;not false~]" 'abc) ⇒ "not false"
     (let ((n 3))
       (format #t "~d gnu~:[s are~; is~] here" n (= 1 n)))
     ⊣ 3 gnus are here

~@[ also treats the argument as a flag, and expects one clause. If the argument is #f then no output is produced and the argument is consumed, otherwise the clause is used and the argument is not consumed, it’s left for the clause. This can be used for instance to suppress output if #f means something not available.

     (format #f "~@[temperature=~d~]" 27) ⇒ "temperature=27"
     (format #f "~@[temperature=~d~]" #f) ⇒ ""

~^Escape. Parameters: val1, val2, val3.

Stop formatting if there are no more arguments. This can be used for instance to have a format string adapt to a variable number of arguments.

     (format #t "~d~^~d" 1) ⊣ 1
     (format #t "~d~^~d" 1 2) ⊣ 1 2

Within a ~{ ~} iteration, ~^stops the current iteration step if there are no more arguments to that step, but continuing with possible further steps and the rest of the format. This can be used for instance to avoid a separator on the last iteration, or to adapt to variable length argument lists.

     (format #f "~{~d~^/~} go"    '(1 2 3))  ⇒ "1/2/3 go"
     (format #f "~:{ ~d~^~d~} go" '((1) (2 3))) ⇒ " 1 23 go"

Within a ~? sub-format, ~^operates just on that sub-format. If it terminates the sub-format then the originating format will still continue.

     (format #t "~? items" "~d~^~d" '(1)) ⊣ 1 items
     (format #t "~? items" "~d~^~d" '(1 2)) ⊣ 1 2 items

The parameters to ~^(which are numbers) change the condition used to terminate. For a single parameter, termination is when that value is zero (notice this makes plain ~^equivalent to ~#^). For two parameters, termination is when those two are equal. For three parameters, termination is when val1 ≤ val2 and val2 ≤ val3.

~q Inquiry message. Insert a copyright message into the output. ~:q inserts the format implementation version.

It’s an error if there are not enough arguments for the escapes in the format string, but any excess arguments are ignored.

Iterations ~{~} and conditionals ~[~;~] can be nested, but must be properly nested, meaning the inner form must be entirely within the outer form. So it’s not possible, for instance, to try to conditionalize the endpoint of an iteration.

     (format #t "~{ ~[ ... ~] ~}" ...)       ;; good
     (format #t "~{ ~[ ... ~} ... ~]" ...)   ;; bad

The same applies to case conversions ~( ~), they must properly nest with respect to iterations and conditionals (though currently a case conversion cannot nest within another case conversion). When a sub-format (~?) is used, that sub-format string must be self-contained. It cannot for instance give a ~{ to begin an iteration form and have the ~} up in the originating format, or similar.

Guile contains a format procedure even when the module (ice-9 format) is not loaded. The default format is simple-format (see Section 6.14.3 [Writing], page 320), it doesn’t support all escape sequences documented in this section, and will signal an error if you try to use one of them. The reason for two versions is that the full format is fairly large and requires some time to load. simple-format is often adequate too.

7.11 File Tree Walk

The functions in this section traverse a tree of files and directories. They come in two flavors: the first one is a high-level functional interface, and the second one is similar to the C ftw and nftw routines (see Section “Working with Directory Trees” in GNU C Library Reference Manual).

(use-modules (ice-9 ftw))

[Scheme Procedure] file-system-tree file-name [enter? [stat]]

Return a tree of the form (file-name stat children ...) where stat is the result of (stat file-name) and children are similar structures for each file contained in file-name when it designates a directory. The optional enter? predicate is invoked as (enter? name stat) and should return true to allow recursion into directory name; the default value is a procedure that always returns #t. When a directory does not match enter?, it nonetheless appears in the resulting tree, only with zero children. The stat argument is optional and defaults to lstat, as for file-system-fold (see below.) The example below shows how to obtain a hierarchical listing of the files under the module/language directory in the Guile source tree, discarding their stat info:

     (use-modules (ice-9 match))
     (define remove-stat
       ;; Remove the `stat' object the `file-system-tree' provides
       ;; for each file in the tree.
       (match-lambda
         ((name stat)              ; flat file
          name)
         ((name stat children ...) ; directory
          (list name (map remove-stat children)))))

(let ((dir (string-append (assq-ref %guile-build-info 'top_srcdir)
                      "/module/language")))
(remove-stat (file-system-tree dir)))

⇒

("language"
(("value" ("spec.go" "spec.scm"))
("scheme"
("spec.go"
"spec.scm"
"compile-tree-il.scm"
"decompile-tree-il.scm"
"decompile-tree-il.go"
"compile-tree-il.go"))
("tree-il"
("spec.go"
"fix-letrec.go"
"inline.go"
"fix-letrec.scm"
"compile-glil.go"
"spec.scm"
"optimize.scm"
"primitives.scm"
...))
...))

It is often desirable to process directories entries directly, rather than building up a tree of entries in memory, like file-system-tree does. The following procedure, a combinator, is designed to allow directory entries to be processed directly as a directory tree is traversed; in fact, file-system-tree is implemented in terms of it.

[Scheme Procedure] file-system-fold enter? leaf down up skip error init file-name

[stat] Traverse the directory at file-name, recursively, and return the result of the successive applications of the leaf, down, up, and skip procedures as described below. Enter sub-directories only when (enter? path stat result) returns true. When a sub-directory is entered, call (down path stat result), where path is the path of the sub-directory and stat the result of (false-if-exception (stat path)); when

it is left, call (up path stat result).

For each file in a directory, call (leaf path stat result). When enter? returns #f, or when an unreadable directory is encountered, call (skip

path stat result).

When file-name names a flat file, (leaf path stat init) is returned. When an opendir or stat call fails, call (error path stat errno result), with errno being the operating system error number that was raised—e.g., EACCES—and stat either #f or the result of the stat call for that entry, when available.

The special . and .. entries are not passed to these procedures. The path argument to the procedures is a full file name—e.g., "../foo/bar/gnu"; if file-name is an absolute file name, then path is also an absolute file name. Files and directories, as identified by their device/inode number pair, are traversed only once. The optional stat argument defaults to lstat, which means that symbolic links are not followed; the stat procedure can be used instead when symbolic links are to be followed (see Section 7.2.3 [File System], page 476). The example below illustrates the use of file-system-fold:

(define (total-file-size file-name)
"Return the size in bytes of the files under FILE-NAME (similar
to `du --apparent-size' with GNU Coreutils.)"
(define (enter? name stat result)
  ;; Skip version control directories.
  (not (member (basename name) '(".git" ".svn" "CVS"))))
(define (leaf name stat result)
  ;; Return RESULT plus the size of the file at NAME.
  (+ result (stat:size stat)))
;; Count zero bytes for directories.
(define (down name stat result) result)
(define (up name stat result) result)
;; Likewise for skipped directories.
(define (skip name stat result) result)
;; Ignore unreadable files/directories but warn the user.
(define (error name stat errno result)
  (format (current-error-port) "warning: ~a: ~a~%"
          name (strerror errno))
  result)
(file-system-fold enter? leaf down up skip error
                         0  ; initial counter is zero bytes
                         file-name))
(total-file-size ".")
⇒ 8217554
(total-file-size "/dev/null")
⇒ 0

The alternative C-like functions are described below.

[Scheme Procedure] scandir name [select? [entry<?]] Return the list of the names of files contained in directory name that match predicate select? (by default, all files). The returned list of file names is sorted according to entry<?, which defaults to string-locale<? such that file names are sorted in the locale’s alphabetical order (see Section 6.24.2 [Text Collation], page 437). Return #f when name is unreadable or is not a directory. This procedure is modeled after the C library function of the same name (see Section “Scanning Directory Content” in GNU C Library Reference Manual).

[Scheme Procedure] ftw startname proc [’hash-size n] Walk the file system tree descending from startname, calling proc for each file and directory. Hard links and symbolic links are followed. A file or directory is reported to proc only once, and skipped if seen again in another place. One consequence of this is that ftw is safe against circularly linked directory structures. Each proc call is (proc filename statinfo flag) and it should return #t to continue, or any other value to stop. filename is the item visited, being startname plus a further path and the name of the item. statinfo is the return from stat (see Section 7.2.3 [File System], page 476) on filename. flag is one of the following symbols,

regular filename is a file, this includes special files like devices, named pipes, etc.

directory

filename is a directory.

invalid-stat

An error occurred when calling stat, so nothing is known. statinfo is #f in this case.

directory-not-readable

filename is a directory, but one which cannot be read and hence won’t be recursed into.

symlink filename is a dangling symbolic link. Symbolic links are normally followed and their target reported, the link itself is reported if the target does not exist.

The return value from ftw is #t if it ran to completion, or otherwise the non-#t value from proc which caused the stop. Optional argument symbol hash-size and an integer can be given to set the size of the hash table used to track items already visited. (see Section 6.7.14.2 [Hash Table Reference], page 237) In the current implementation, returning non-#t from proc is the only valid way to terminate ftw. proc must not use throw or similar to escape.

[Scheme Procedure] nftw startname proc [’chdir] [’depth] [’hash-size n] [’mount]

[’physical] Walk the file system tree starting at startname, calling proc for each file and directory. nftw has extra features over the basic ftw described above. Like ftw, hard links and symbolic links are followed. A file or directory is reported to proc only once, and skipped if seen again in another place. One consequence of this is that nftw is safe against circular linked directory structures.

Each proc call is (proc filename statinfo flag base level) and it should return

#t to continue, or any other value to stop. filename is the item visited, being startname plus a further path and the name of the item. statinfo is the return from stat on filename (see Section 7.2.3 [File System], page 476). base is an integer offset into filename which is where the basename for this item begins. level is an integer giving the directory nesting level, starting from 0 for the contents of startname (or that item itself if it’s a file). flag is one of the following symbols,

regular filename is a file, including special files like devices, named pipes, etc.

directory

filename is a directory.

directory-processed

filename is a directory, and its contents have all been visited. This flag is given instead of directory when the depth option below is used.

invalid-stat

An error occurred when applying stat to filename, so nothing is known about it. statinfo is #f in this case.

directory-not-readable

filename is a directory, but one which cannot be read and hence won’t be recursed into.

stale-symlink

filename is a dangling symbolic link. Links are normally followed and their target reported, the link itself is reported if its target does not exist.

symlink When the physical option described below is used, this indicates filename is a symbolic link whose target exists (and is not being followed).

The following optional arguments can be given to modify the way nftw works. Each is passed as a symbol (and hash-size takes a following integer value).

chdir Change to the directory containing the item before calling proc. When nftw returns the original current directory is restored. Under this option, generally the base parameter to each proc call should be used to pick out the base part of the filename. The filename is still a path but with a changed directory it won’t be valid (unless the startname directory was absolute).

depth Visit files “depth first”, meaning proc is called for the contents of each directory before it’s called for the directory itself. Normally a directory is reported first, then its contents. Under this option, the flag to proc for a directory is directory-

processed instead of directory.
hash-size n

Set the size of the hash table used to track items already visited. (see Section 6.7.14.2 [Hash Table Reference], page 237) mount Don’t cross a mount point, meaning only visit items on the same file

system as startname (ie. the same stat:dev).

physical Don’t follow symbolic links, instead report them to proc as symlink. Dangling links (those whose target doesn’t exist) are still reported as

          stale-symlink.

The return value from nftw is #t if it ran to completion, or otherwise the non-#t value from proc which caused the stop. In the current implementation, returning non-#t from proc is the only valid way to terminate ftw. proc must not use throw or similar to escape.

7.12 Queues

The functions in this section are provided by

(use-modules (ice-9 q))

This module implements queues holding arbitrary scheme objects and designed for efficient first-in / first-out operations.

make-q creates a queue, and objects are entered and removed with enq! and deq!. q-push! and q-pop! can be used too, treating the front of the queue like a stack.

[Scheme Procedure] make-q

Return a new queue.

[Scheme Procedure] q? obj Return #t if obj is a queue, or #f if not. Note that queues are not a distinct class of objects but are implemented with cons cells. For that reason certain list structures can get #t from q?.

[Scheme Procedure] enq! q obj

Add obj to the rear of q, and return q.

[Scheme Procedure] deq! q [Scheme Procedure] q-pop! q

Remove and return the front element from q. If q is empty, a q-empty exception is thrown. deq! and q-pop! are the same operation, the two names just let an application match

enq! with deq!, or q-push! with q-pop!.

[Scheme Procedure] q-push! q obj

Add obj to the front of q, and return q.

[Scheme Procedure] q-length q

Return the number of elements in q.

[Scheme Procedure] q-empty? q

Return true if q is empty.

[Scheme Procedure] q-empty-check q

Throw a q-empty exception if q is empty.

[Scheme Procedure] q-front q

Return the first element of q (without removing it). If q is empty, a q-empty exception is thrown.

[Scheme Procedure] q-rear q

Return the last element of q (without removing it). If q is empty, a q-empty exception is thrown.

[Scheme Procedure] q-remove! q obj

Remove all occurrences of obj from q, and return q. obj is compared to queue elements using eq?.

The q-empty exceptions described above are thrown just as (throw 'q-empty), there’s no message etc like an error throw.

A queue is implemented as a cons cell, the car containing a list of queued elements, and the cdr being the last cell in that list (for ease of enqueuing).

(list . last-cell)

If the queue is empty, list is the empty list and last-cell is #f.

An application can directly access the queue list if desired, for instance to search the elements or to insert at a specific point.

[Scheme Procedure] sync-q! q

Recompute the last-cell field in q. All the operations above maintain last-cell as described, so normally there’s no need for sync-q!. But if an application modifies the queue list then it must either maintain last-cell similarly, or call sync-q! to recompute it.

7.13 Streams

This section documents Guile’s legacy stream module. For a more complete and portable stream library, see Section 7.5.27 [SRFI-41], page 599.

A stream represents a sequence of values, each of which is calculated only when required. This allows large or even infinite sequences to be represented and manipulated with familiar operations like “car”, “cdr”, “map” or “fold”. In such manipulations only as much as needed is actually held in memory at any one time. The functions in this section are available from

(use-modules (ice-9 streams))

Streams are implemented using promises (see Section 6.17.9 [Delayed Evaluation], page 371), which is how the underlying calculation of values is made only when needed, and the values then retained so the calculation is not repeated. Here is a simple example producing a stream of all odd numbers,

(define odds (make-stream (lambda (state)
                            (cons state (+ state 2)))
                          1))
(stream-car odds)  ⇒ 1
(stream-car (stream-cdr odds)) ⇒ 3

stream-map could be used to derive a stream of odd squares,

(define (square n) (* n n))
(define oddsquares (stream-map square odds))

These are infinite sequences, so it’s not possible to convert them to a list, but they could be printed (infinitely) with for example

(stream-for-each (lambda (n sq)
                   (format #t "~a squared is ~a\n" n sq))
                 odds oddsquares)

⊣

1 squared is 1
3 squared is 9
5 squared is 25
7 squared is 49
...

[Scheme Procedure] make-stream proc initial-state Return a new stream, formed by calling proc successively. Each call is (proc state), it should return a pair, the car being the value for the stream, and the cdr being the new state for the next call. For the first call state is the given initial-state. At the end of the stream, proc should return some non-pair object.

[Scheme Procedure] stream-car stream

Return the first element from stream. stream must not be empty.

[Scheme Procedure] stream-cdr stream

Return a stream which is the second and subsequent elements of stream. stream must not be empty.

[Scheme Procedure] stream-null? stream

Return true if stream is empty.

[Scheme Procedure] list->stream list [Scheme Procedure] vector->stream vector

Return a stream with the contents of list or vector. list or vector should not be modified subsequently, since it’s unspecified whether changes there will be reflected in the stream returned.

[Scheme Procedure] port->stream port readproc

Return a stream which is the values obtained by reading from port using readproc. Each read call is (readproc port), and it should return an EOF object (see Section 6.14.2 [Reading], page 318) at the end of input. For example a stream of characters from a file,

     (port->stream (open-input-file "/foo/bar.txt") read-char)

[Scheme Procedure] stream->list stream

Return a list which is the entire contents of stream.

[Scheme Procedure] stream->reversed-list stream

Return a list which is the entire contents of stream, but in reverse order.

[Scheme Procedure] stream->list&length stream

Return two values (see Section 6.13.7 [Multiple Values], page 301), being firstly a list which is the entire contents of stream, and secondly the number of elements in that list.

[Scheme Procedure] stream->reversed-list&length stream

Return two values (see Section 6.13.7 [Multiple Values], page 301) being firstly a list which is the entire contents of stream, but in reverse order, and secondly the number of elements in that list.

[Scheme Procedure] stream->vector stream

Return a vector which is the entire contents of stream.

[Function] stream-fold proc init stream1 stream2 … Apply proc successively over the elements of the given streams, from first to last until the end of the shortest stream is reached. Return the result from the last proc call. Each call is (proc elem1 elem2 ... prev), where each elem is from the corresponding stream. prev is the return from the previous proc call, or the given init for the first call.

[Function] stream-for-each proc stream1 stream2 … Call proc on the elements from the given streams. The return value is unspecified. Each call is (proc elem1 elem2 ...), where each elem is from the corresponding stream. stream-for-each stops when it reaches the end of the shortest stream.

[Function] stream-map proc stream1 stream2 … Return a new stream which is the results of applying proc to the elements of the given streams. Each call is (proc elem1 elem2 ...), where each elem is from the corresponding stream. The new stream ends when the end of the shortest given stream is reached.

7.14 Buffered Input

The following functions are provided by

(use-modules (ice-9 buffered-input))

A buffered input port allows a reader function to return chunks of characters which are to be handed out on reading the port. A notion of further input for an application level logical expression is maintained too, and passed through to the reader.

[Scheme Procedure] make-buffered-input-port reader

Create an input port which returns characters obtained from the given reader function. reader is called (reader cont), and should return a string or an EOF object. The new port gives precisely the characters returned by reader, nothing is added, so if any newline characters or other separators are desired they must come from the reader function.

The cont parameter to reader is #f for initial input, or #t when continuing an expression. This is an application level notion, set with set-buffered-input-continuation?! below. If the user has entered a partial expression then it allows reader for instance to give a different prompt to show more is required.

[Scheme Procedure] make-line-buffered-input-port reader

Create an input port which returns characters obtained from the specified reader function, similar to make-buffered-input-port above, but where reader is expected to be a line-oriented. reader is called (reader cont), and should return a string or an EOF object as above. Each string is a line of input without a newline character, the port code inserts a newline after each string.

[Scheme Procedure] set-buffered-input-continuation?! port cont

Set the input continuation flag for a given buffered input port. An application uses this by calling with a cont flag of #f when beginning to read a new logical expression. For example with the Scheme read function (see Section 6.17.2 [Scheme Read], page 360),

     (define my-port (make-buffered-input-port my-reader))
     (set-buffered-input-continuation?! my-port #f)
     (let ((obj (read my-port)))
       ...

7.15 Expect

The macros in this section are made available with:

(use-modules (ice-9 expect))

expect is a macro for selecting actions based on the output from a port. The name comes from a tool of similar functionality by Don Libes. Actions can be taken when a particular string is matched, when a timeout occurs, or when end-of-file is seen on the port. The expect macro is described below; expect-strings is a front-end to expect based on regexec (see the regular expression documentation).

[Macro] expect-strings clause …

By default, expect-strings will read from the current input port. The first term in each clause consists of an expression evaluating to a string pattern (regular expression). As characters are read one-by-one from the port, they are accumulated in a buffer string which is matched against each of the patterns. When a pattern matches, the remaining expression(s) in the clause are evaluated and the value of the last is returned. For example:

     (with-input-from-file "/etc/passwd"
       (lambda ()
         (expect-strings
           ("^nobody" (display "Got a nobody user.\n")
                      (display "That's no problem.\n"))
           ("^daemon" (display "Got a daemon user.\n")))))

The regular expression is compiled with the REG_NEWLINE flag, so that the ^and $ anchors will match at any newline, not just at the start and end of the string. There are two other ways to write a clause: The expression(s) to evaluate can be omitted, in which case the result of the regular expression match (converted to strings, as obtained from regexec with match-pick set to "") will be returned if the pattern matches. The symbol => can be used to indicate that the expression is a procedure which will accept the result of a successful regular expression match. E.g.,

     ("^daemon" => write)
     ("^d(aemon)" => (lambda args (for-each write args)))
     ("^da(em)on" => (lambda (all sub)
                       (write all) (newline)
                       (write sub) (newline)))

The order of the substrings corresponds to the order in which the opening brackets occur. A number of variables can be used to control the behaviour of expect (and expectstrings). Most have default top-level bindings to the value #f, which produces the default behaviour. They can be redefined at the top level or locally bound in a form enclosing the expect expression.

expect-port

A port to read characters from, instead of the current input port.

expect-timeout

expect will terminate after this number of seconds, returning #f or the value returned by expect-timeout-proc.

expect-timeout-proc

A procedure called if timeout occurs. The procedure takes a single argument: the accumulated string.

expect-eof-proc

A procedure called if end-of-file is detected on the input port. The procedure takes a single argument: the accumulated string.

expect-char-proc

A procedure to be called every time a character is read from the port. The procedure takes a single argument: the character which was read.

expect-strings-compile-flags

Flags to be used when compiling a regular expression, which are passed to make-regexp See Section 6.15.1 [Regexp Functions], page 350. The default value is regexp/newline.

expect-strings-exec-flags

Flags to be used when executing a regular expression, which are passed to regexp-exec See Section 6.15.1 [Regexp Functions], page 350. The default value is regexp/noteol, which prevents $ from matching the end of the string while it is still accumulating, but still allows it to match after a line break or at the end of file.

Here’s an example using all of the variables:

      (let ((expect-port (open-input-file "/etc/passwd"))
            (expect-timeout 1)
            (expect-timeout-proc
              (lambda (s) (display "Times up!\n")))
            (expect-eof-proc
              (lambda (s) (display "Reached the end of the file!\n")))
            (expect-char-proc display)
            (expect-strings-compile-flags (logior regexp/newline regexp/icase))
            (expect-strings-exec-flags 0))
         (expect-strings
           ("^nobody"  (display "Got a nobody user\n"))))

[Macro] expect clause …

expect is used in the same way as expect-strings, but tests are specified not as patterns, but as procedures. The procedures are called in turn after each character is read from the port, with two arguments: the value of the accumulated string and a flag to indicate whether end-of-file has been reached. The flag will usually be #f, but if end-of-file is reached, the procedures are called an additional time with the final accumulated string and #t. The test is successful if the procedure returns a non-false value. If the => syntax is used, then if the test succeeds it must return a list containing the arguments to be provided to the corresponding expression. In the following example, a string will only be matched at the beginning of the file:

     (let ((expect-port (open-input-file "/etc/passwd")))
       (expect
          ((lambda (s eof?) (string=? s "fnord!"))
             (display "Got a nobody user!\n"))))

The control variables described for expect-strings also influence the behaviour of expect, with the exception of variables whose names begin with expect-strings-.

7.16 sxml-match: Pattern Matching of SXML

The (sxml match) module provides syntactic forms for pattern matching of SXML trees, in a “by example” style reminiscent of the pattern matching of the syntax-rules and syntax-case macro systems. See Section 7.22 [SXML], page 697, for more information on SXML.

The following example^{5} provides a brief illustration, transforming a music album catalog language into HTML.

(define (album->html x)
  (sxml-match x
    [(album (@ (title ,t)) (catalog (num ,n) (fmt ,f)) ...)
     `(ul (li ,t)
          (li (b ,n) (i ,f)) ...)]))
Three macros are provided: sxml-match, sxml-match-let, and sxml-match-let*.

5 This example is taken from a paper by Krishnamurthi et al. Their paper was the first to show the

usefulness of the syntax-rules style of pattern matching for transformation of XML, though the language described, XT3D, is an XML language.

Compared to a standard s-expression pattern matcher (see Section 7.7 [Pattern Matching], page 657), sxml-match provides the following benefits:

• matching of SXML elements does not depend on any degree of normalization of the SXML;

• matching of SXML attributes (within an element) is under-ordered; the order of the attributes specified within the pattern need not match the ordering with the element being matched;

• all attributes specified in the pattern must be present in the element being matched; in the spirit that XML is ’extensible’, the element being matched may include additional attributes not specified in the pattern.

The present module is a descendant of WebIt!, and was inspired by an s-expression pattern matcher developed by Erik Hilsdale, Dan Friedman, and Kent Dybvig at Indiana University.

Syntax

sxml-match provides case-like form for pattern matching of XML nodes.

[Scheme Syntax] sxml-match input-expression clause1 clause2 … Match input-expression, an SXML tree, according to the given clauses (one or more), each consisting of a pattern and one or more expressions to be evaluated if the pattern match succeeds. Optionally, each clause within sxml-match may include a guard expression.

The pattern notation is based on that of Scheme’s syntax-rules and syntax-case macro systems. The grammar for the sxml-match syntax is given below:

match-form ::= (sxml-match input-expression
                 clause+)
clause ::= [node-pattern action-expression+]
         | [node-pattern (guard expression*) action-expression+]
node-pattern ::= literal-pattern
               | pat-var-or-cata
               | element-pattern
               | list-pattern
literal-pattern ::= string
                  | character
                  | number
                  | #t
                  | #f
attr-list-pattern ::= (@ attribute-pattern*)
                    | (@ attribute-pattern* . pat-var-or-cata)
attribute-pattern ::= (tag-symbol attr-val-pattern)

attr-val-pattern ::= literal-pattern
                   | pat-var-or-cata
                   | (pat-var-or-cata default-value-expr)
element-pattern ::= (tag-symbol attr-list-pattern?)
                  | (tag-symbol attr-list-pattern? nodeset-pattern)
                  | (tag-symbol attr-list-pattern?
                                nodeset-pattern? . pat-var-or-cata)
list-pattern ::= (list nodeset-pattern)
               | (list nodeset-pattern? . pat-var-or-cata)
               | (list)
nodeset-pattern ::= node-pattern
                  | node-pattern ...
                  | node-pattern nodeset-pattern
                  | node-pattern ... nodeset-pattern
pat-var-or-cata ::= (unquote var-symbol)
                  | (unquote [var-symbol*])
                  | (unquote [cata-expression -> var-symbol*])

Within a list or element body pattern, ellipses may appear only once, but may be followed by zero or more node patterns.

Guard expressions cannot refer to the return values of catamorphisms.

Ellipses in the output expressions must appear only in an expression context; ellipses are not allowed in a syntactic form.

The sections below illustrate specific aspects of the sxml-match pattern matcher.

Matching XML Elements

The example below illustrates the pattern matching of an XML element:

     (sxml-match '(e (@ (i 1)) 3 4 5)
       [(e (@ (i ,d)) ,a ,b ,c) (list d a b c)]
       [,otherwise #f])

Each clause in sxml-match contains two parts: a pattern and one or more expressions which are evaluated if the pattern is successfully match. The example above matches an element e with an attribute i and three children.

Pattern variables are must be “unquoted” in the pattern. The above expression binds d to 1, a to 3, b to 4, and c to 5.

Ellipses in Patterns

As in syntax-rules, ellipses may be used to specify a repeated pattern. Note that the pattern item ... specifies zero-or-more matches of the pattern item.

The use of ellipses in a pattern is illustrated in the code fragment below, where nested ellipses are used to match the children of repeated instances of an a element, within an element d.

     (define x '(d (a 1 2 3) (a 4 5) (a 6 7 8) (a 9 10)))
     (sxml-match x
       [(d (a ,b ...) ...)
        (list (list b ...) ...)])

The above expression returns a value of ((1 2 3) (4 5) (6 7 8) (9 10)).

Ellipses in Quasiquote’d Output

Within the body of an sxml-match form, a slightly extended version of quasiquote is provided, which allows the use of ellipses. This is illustrated in the example below.

     (sxml-match '(e 3 4 5 6 7)
       [(e ,i ... 6 7) `("start" ,(list 'wrap i) ... "end")]
       [,otherwise #f])

The general pattern is that `(something ,i ...) is rewritten as `(something ,@i).

Matching Nodesets

A nodeset pattern is designated by a list in the pattern, beginning the identifier list. The example below illustrates matching a nodeset.

     (sxml-match '("i" "j" "k" "l" "m")
       [(list ,a ,b ,c ,d ,e)
        `((p ,a) (p ,b) (p ,c) (p ,d) (p ,e))])

This example wraps each nodeset item in an HTML paragraph element. This example can be rewritten and simplified through using ellipsis:

     (sxml-match '("i" "j" "k" "l" "m")
       [(list ,i ...)
        `((p ,i) ...)])

This version will match nodesets of any length, and wrap each item in the nodeset in an HTML paragraph element.

Matching the “Rest” of a Nodeset

Matching the “rest” of a nodeset is achieved by using a . rest) pattern at the end of an element or nodeset pattern.

This is illustrated in the example below:

     (sxml-match '(e 3 (f 4 5 6) 7)
       [(e ,a (f . ,y) ,d)
        (list a y d)])

The above expression returns (3 (4 5 6) 7).

Matching the Unmatched Attributes

Sometimes it is useful to bind a list of attributes present in the element being matched, but which do not appear in the pattern. This is achieved by using a . rest) pattern at the end of the attribute list pattern. This is illustrated in the example below:

     (sxml-match '(a (@ (z 1) (y 2) (x 3)) 4 5 6)
       [(a (@ (y ,www) . ,qqq) ,t ,u ,v)
        (list www qqq t u v)])

The above expression matches the attribute y and binds a list of the remaining attributes to the variable qqq. The result of the above expression is (2 ((z 1) (x 3)) 4 5 6).

This type of pattern also allows the binding of all attributes:

     (sxml-match '(a (@ (z 1) (y 2) (x 3)))
       [(a (@ . ,qqq))
        qqq])

Default Values in Attribute Patterns

It is possible to specify a default value for an attribute which is used if the attribute is not present in the element being matched. This is illustrated in the following example:

     (sxml-match '(e 3 4 5)
       [(e (@ (z (,d 1))) ,a ,b ,c) (list d a b c)])

The value 1 is used when the attribute z is absent from the element e.

Guards in Patterns

Guards may be added to a pattern clause via the guard keyword. A guard expression may include zero or more expressions which are evaluated only if the pattern is matched. The body of the clause is only evaluated if the guard expressions evaluate to #t.

The use of guard expressions is illustrated below:

     (sxml-match '(a 2 3)
       ((a ,n) (guard (number? n)) n)
       ((a ,m ,n) (guard (number? m) (number? n)) (+ m n)))

Catamorphisms

The example below illustrates the use of explicit recursion within an sxml-match form. This example implements a simple calculator for the basic arithmetic operations, which are represented by the XML elements plus, minus, times, and div.

     (define simple-eval
       (lambda (x)
         (sxml-match x
           [,i (guard (integer? i)) i]
           [(plus ,x ,y) (+ (simple-eval x) (simple-eval y))]
           [(times ,x ,y) (* (simple-eval x) (simple-eval y))]
           [(minus ,x ,y) (- (simple-eval x) (simple-eval y))]
           [(div ,x ,y) (/ (simple-eval x) (simple-eval y))]
           [,otherwise (error "simple-eval: invalid expression" x)])))

Using the catamorphism feature of sxml-match, a more concise version of simple-eval can be written. The pattern ,[x] recursively invokes the pattern matcher on the value bound in this position.

     (define simple-eval
       (lambda (x)

(sxml-match x
[,i (guard (integer? i)) i]
[(plus ,[x] ,[y]) (+ x y)]
[(times ,[x] ,[y]) (* x y)]
[(minus ,[x] ,[y]) (- x y)]
[(div ,[x] ,[y]) (/ x y)]
[,otherwise (error "simple-eval: invalid expression" x)])))

Named-Catamorphisms

It is also possible to explicitly name the operator in the “cata” position. Where ,[id*] recurs to the top of the current sxml-match, ,[cata -> id*] recurs to cata. cata must evaluate to a procedure which takes one argument, and returns as many values as there are identifiers following ->.

Named catamorphism patterns allow processing to be split into multiple, mutually recursive procedures. This is illustrated in the example below: a transformation that formats a “TV Guide” into HTML.

(define (tv-guide->html g)
(define (cast-list cl)
(sxml-match cl
[(CastList (CastMember (Character (Name ,ch)) (Actor (Name ,a))) ...)
 `(div (ul (li ,ch ": " ,a) ...))]))
(define (prog p)
(sxml-match p
[(Program (Start ,start-time) (Duration ,dur) (Series ,series-title)
          (Description ,desc ...))
 `(div (p ,start-time
          (br) ,series-title
          (br) ,desc ...))]
[(Program (Start ,start-time) (Duration ,dur) (Series ,series-title)
          (Description ,desc ...)
          ,[cast-list -> cl])
 `(div (p ,start-time
          (br) ,series-title
          (br) ,desc ...)
       ,cl)]))
(sxml-match g
[(TVGuide (@ (start ,start-date)
           (end ,end-date))
        (Channel (Name ,nm) ,[prog -> p] ...) ...)
`(html (head (title "TV Guide"))
      (body (h1 "TV Guide")
            (div (h2 ,nm) ,p ...) ...))]))
sxml-match-let and sxml-match-let*

[Scheme Syntax] sxml-match-let ((pat expr) ...) expression0 expression ...

[Scheme Syntax] sxml-match-let* ((pat expr) ...) expression0 expression ...

These forms generalize the let and let* forms of Scheme to allow an XML pattern in the binding position, rather than a simple variable.

For example, the expression below:

     (sxml-match-let ([(a ,i ,j) '(a 1 2)])
       (+ i j))

binds the variables i and j to 1 and 2 in the XML value given.

7.17 The Scheme shell (scsh)

An incomplete port of the Scheme shell (scsh) was once available for Guile as a separate package. However this code has bitrotten somewhat. The pieces are available in Guile’s legacy CVS repository, which may be browsed at http://cvs.savannah.gnu.org/viewvc/

guile/guile-scsh/?root=guile.

For information about scsh see http://www.scsh.net/.

This bitrotting is a bit of a shame, as there is a good deal of well-written Scheme code in scsh. Adopting this code and porting it to current Guile should be an educational experience, in addition to providing something of value to Guile folks.

7.18 Curried Definitions

The macros in this section are provided by

     (use-modules (ice-9 curried-definitions))

and replace those provided by default.

Prior to Guile 2.0, Guile provided a type of definition known colloquially as a “curried definition”. The idea is to extend the syntax of define so that you can conveniently define procedures that return procedures, up to any desired depth.

For example,

     (define ((foo x) y)
       (list x y))

is a convenience form of

     (define foo
       (lambda (x)
         (lambda (y)
           (list x y))))

[Scheme Syntax] define (… (name args …) …) body … [Scheme Syntax] define* (… (name args …) …) body … [Scheme Syntax] define-public (… (name args …) …) body …

Create a top level variable name bound to the procedure with parameter list args. If name is itself a formal parameter list, then a higher order procedure is created using that formal-parameter list, and returning a procedure that has parameter list args. This nesting may occur to arbitrary depth. define* is similar but the formal parameter lists take additional options as described in Section 6.9.4.1 [lambda* and define*], page 248. For example,

          (define* ((foo #:keys (bar 'baz) (quux 'zot)) frotz #:rest rest)
            (list bar quux frotz rest))
          ((foo #:quux 'foo) 1 2 3 4 5)
          ⇒ (baz foo 1 (2 3 4 5))

define-public is similar to define but it also adds name to the list of exported bindings of the current module.

7.19 Statprof

(statprof) is a fairly simple statistical profiler for Guile.

A simple use of statprof would look like this:

     (statprof-reset 0 50000 #t)
     (statprof-start)
     (do-something)
     (statprof-stop)
     (statprof-display)

This would reset statprof, clearing all accumulated statistics, then start profiling, run some code, stop profiling, and finally display a gprof flat-style table of statistics which will look something like this:

       %   cumulative      self              self    total
      time    seconds   seconds    calls  ms/call  ms/call  name
      35.29      0.23      0.23     2002     0.11     0.11  -
      23.53      0.15      0.15     2001     0.08     0.08  positive?
      23.53      0.15      0.15     2000     0.08     0.08  +
      11.76      0.23      0.08     2000     0.04     0.11  do-nothing
       5.88      0.64      0.04     2001     0.02     0.32  loop
       0.00      0.15      0.00        1     0.00   150.59  do-something
      ...

All of the numerical data with the exception of the calls column is statistically approximate. In the following column descriptions, and in all of statprof, "time" refers to execution time (both user and system), not wall clock time.

% time The percent of the time spent inside the procedure itself (not counting children).

cumulative seconds The total number of seconds spent in the procedure, including children.

self seconds The total number of seconds spent in the procedure itself (not counting children).

calls The total number of times the procedure was called.

self ms/call The average time taken by the procedure itself on each call, in ms.

total ms/call The average time taken by each call to the procedure, including time spent in child functions. name The name of the procedure.

The profiler uses eq? and the procedure object itself to identify the procedures, so it won’t confuse different procedures with the same name. They will show up as two different rows in the output.

Right now the profiler is quite simplistic. I cannot provide call-graphs or other higher level information. What you see in the table is pretty much all there is. Patches are welcome :-)

7.20 Implementation notes

The profiler works by setting the unix profiling signal ITIMER_PROF to go off after the interval you define in the call to statprof-reset. When the signal fires, a sampling routine is run which looks at the current procedure that’s executing, and then crawls up the stack, and for each procedure encountered, increments that procedure’s sample count. Note that if a procedure is encountered multiple times on a given stack, it is only counted once. After the sampling is complete, the profiler resets profiling timer to fire again after the appropriate interval.

Meanwhile, the profiler keeps track, via get-internal-run-time, how much CPU time (system and user – which is also what ITIMER_PROF tracks), has elapsed while code has been executing within a statprof-start/stop block.

The profiler also tries to avoid counting or timing its own code as much as possible.

7.21 Usage

[Function] statprof-active?

Returns #t if statprof-start has been called more times than statprof-stop, #f otherwise.

[Function] statprof-start

Start the profiler.

[Function] statprof-stop

Stop the profiler.

[Function] statprof-reset sample-seconds sample-microseconds count-calls?

[full-stacks?] Reset the statprof sampler interval to sample-seconds and sample-microseconds. If count-calls? is true, arrange to instrument procedure calls as well as collecting statistical profiling data. If full-stacks? is true, collect all sampled stacks into a list for later analysis. Enables traps and debugging as necessary.

[Function] statprof-accumulated-time

Returns the time accumulated during the last statprof run.

[Function] statprof-sample-count

Returns the number of samples taken during the last statprof run.

[Function] statprof-fold-call-data proc init

Fold proc over the call-data accumulated by statprof. Cannot be called while statprof is active. proc should take two arguments, (call-data prior-result). Note that a given proc-name may appear multiple times, but if it does, it represents different functions with the same name.

[Function] statprof-proc-call-data proc

Returns the call-data associated with proc, or #f if none is available.

[Function] statprof-call-data-name cd
[Function] statprof-call-data-calls cd
[Function] statprof-call-data-cum-samples cd
[Function] statprof-call-data-self-samples cd
[Function] statprof-call-data->stats call-data

Returns an object of type statprof-stats.

[Function] statprof-stats-proc-name stats
[Function] statprof-stats-%-time-in-proc stats
[Function] statprof-stats-cum-secs-in-proc stats
[Function] statprof-stats-self-secs-in-proc stats
[Function] statprof-stats-calls stats
[Function] statprof-stats-self-secs-per-call stats
[Function] statprof-stats-cum-secs-per-call stats
[Function] statprof-display .

Displays a gprof-like summary of the statistics collected. Unless an optional port argument is passed, uses the current output port.

[Function] statprof-display-anomolies

A sanity check that attempts to detect anomolies in statprof’s statistics.

[Function] statprof-fetch-stacks

Returns a list of stacks, as they were captured since the last call to statprof-reset. Note that stacks are only collected if the full-stacks? argument to statprof-reset is true.

[Function] statprof-fetch-call-tree
Return a call tree for the previous statprof run.
The return value is a list of nodes, each of which is of the type:
@@code
 node ::= (@@var@{proc@} @@var@{count@} . @@var@{nodes@})
@@end code

[Function] statprof thunk [#:loop] [#:hz] [#:count-calls?] [#:full-stacks?] Profiles the execution of thunk. The stack will be sampled hz times per second, and the thunk itself will be called loop times. If count-calls? is true, all procedure calls will be recorded. This operation is somewhat expensive. If full-stacks? is true, at each sample, statprof will store away the whole call tree,

for later analysis. Use statprof-fetch-stacks or statprof-fetch-call-tree to

retrieve the last-stored stacks.

[Special Form] with-statprof args

Profiles the expressions in its body. Keyword arguments:

#:loop Execute the body loop number of times, or #f for no looping default: #f

#:hz Sampling rate default: 20

#:count-calls?

Whether to instrument each function call (expensive) default: #f

#:full-stacks?

Whether to collect away all sampled stacks into a list default: #f

[Function] gcprof thunk [#:loop] [#:full-stacks?] Do an allocation profile of the execution of thunk. The stack will be sampled soon after every garbage collection, yielding an approximate idea of what is causing allocation in your program. Since GC does not occur very frequently, you may need to use the loop parameter, to cause thunk to be called loop times. If full-stacks? is true, at each sample, statprof will store away the whole call tree,

for later analysis. Use statprof-fetch-stacks or statprof-fetch-call-tree to

retrieve the last-stored stacks.

7.22 SXML

SXML is a native representation of XML in terms of standard Scheme data types: lists, symbols, and strings. For example, the simple XML fragment:

<parrot type="African Grey"><name>Alfie</name></parrot>

may be represented with the following SXML:

(parrot (@ (type "African Grey)) (name "Alfie"))

SXML is very general, and is capable of representing all of XML. Formally, this means that SXML is a conforming implementation of the http://www.w3.org/TR/xml-infoset/

(XML Information Set) standard.

Guile includes several facilities for working with XML and SXML: parsers, serializers, and transformers.

7.22.1 SXML Overview

(This section needs to be written; volunteers welcome.)

7.22.2 Reading and Writing XML

The (sxml simple) module presents a basic interface for parsing XML from a port into the Scheme SXML format, and for serializing it back to text.

     (use-modules (sxml simple))

[Scheme Procedure] xml->sxml [string-or-port] [#:namespaces=’()] [#:declare-namespaces?=#t] [#:trim-whitespace?=#f] [#:entities=’()] [#:default-entity-handler=#f] [#:doctype-handler=#f]

Use SSAX to parse an XML document into SXML. Takes one optional argument, string-or-port, which defaults to the current input port. Returns the resulting SXML document. If string-or-port is a port, it will be left pointing at the next available character in the port.

As is normal in SXML, XML elements parse as tagged lists. Attributes, if any, are placed after the tag, within an @ element. The root of the resulting XML will be contained in a special tag, *TOP*. This tag will contain the root element of the XML, but also any prior processing instructions.

     (xml->sxml "<foo/>")
     ⇒ (*TOP* (foo))
     (xml->sxml "<foo>text</foo>")
     ⇒ (*TOP* (foo "text"))
     (xml->sxml "<foo kind=\"bar\">text</foo>")
     ⇒ (*TOP* (foo (@ (kind "bar")) "text"))
     (xml->sxml "<?xml version=\"1.0\"?><foo/>")
     ⇒ (*TOP* (*PI* xml "version=\"1.0\"") (foo))

All namespaces in the XML document must be declared, via xmlns attributes. SXML elements built from non-default namespaces will have their tags prefixed with their URI. Users can specify custom prefixes for certain namespaces with the #:namespaces keyword argument to xml->sxml.

     (xml->sxml "<foo xmlns=\"http://example.org/ns1\">text</foo>")
     ⇒ (*TOP* (http://example.org/ns1:foo "text"))
     (xml->sxml "<foo xmlns=\"http://example.org/ns1\">text</foo>"
                #:namespaces '((ns1 . "http://example.org/ns1")))
     ⇒ (*TOP* (ns1:foo "text"))
     (xml->sxml "<foo xmlns:bar=\"http://example.org/ns2\"><bar:baz/></foo>"
                #:namespaces '((ns2 . "http://example.org/ns2")))
     ⇒ (*TOP* (foo (ns2:baz)))

By default, namespaces passed to xml->sxml are treated as if they were declared on the root element. Passing a false #:declare-namespaces? argument will disable this behavior, requiring in-document declarations of namespaces before use..

     (xml->sxml "<foo><ns2:baz/></foo>"
                #:namespaces '((ns2 . "http://example.org/ns2")))
     ⇒ (*TOP* (foo (ns2:baz)))
     (xml->sxml "<foo><ns2:baz/></foo>"
                #:namespaces '((ns2 . "http://example.org/ns2"))
                #:declare-namespaces? #f)
     ⇒ error: undeclared namespace: `bar'

By default, all whitespace in XML is significant. Passing the #:trim-whitespace? keyword argument to xml->sxml will trim whitespace in front, behind and between elements, treating it as “unsignificant”. Whitespace in text fragments is left alone.

     (xml->sxml "<foo>\n<bar> Alfie the parrot! </bar>\n</foo>")
     ⇒ (*TOP* (foo "\n" (bar " Alfie the parrot! ") "\n"))
     (xml->sxml "<foo>\n<bar> Alfie the parrot! </bar>\n</foo>"
                #:trim-whitespace? #t)
     ⇒ (*TOP* (foo (bar " Alfie the parrot! ")))

Parsed entities may be declared with the #:entities keyword argument, or handled with

the #:default-entity-handler. By default, only the standard &lt;, &gt;, &amp;, &apos;

and &quot; entities are defined, as well as the &#N; and &#xN; (decimal and hexadecimal) numeric character entities.

     (xml->sxml "<foo>&amp;</foo>")
     ⇒ (*TOP* (foo "&"))
     (xml->sxml "<foo>&nbsp;</foo>")
     ⇒ error: undefined entity: nbsp
     (xml->sxml "<foo>&#xA0;</foo>")
     ⇒ (*TOP* (foo "\xa0"))
     (xml->sxml "<foo>&nbsp;</foo>"
                #:entities '((nbsp . "\xa0")))
     ⇒ (*TOP* (foo "\xa0"))
     (xml->sxml "<foo>&nbsp; &foo;</foo>"
                #:default-entity-handler
                (lambda (port name)
                  (case name
                    ((nbsp) "\xa0")
                    (else
                     (format (current-warning-port)
                             "~a:~a:~a: undefined entitity: ~a\n"
                             (or (port-filename port) "<unknown file>")
                             (port-line port) (port-column port)
                             name)
                     (symbol->string name)))))
     ⊣ <unknown file>:0:17: undefined entitity: foo
     ⇒ (*TOP* (foo "\xa0 foo"))

By default, xml->sxml skips over the <!DOCTYPE> declaration, if any. This behavior can be overridden with the #:doctype-handler argument, which should be a procedure of three arguments: the docname (a symbol), systemid (a string), and the internal doctype subset (as a string or #f if not present).

The handler should return keyword arguments as multiple values, as if it were calling its continuation with keyword arguments. The continuation accepts the #:entities and #:namespaces keyword arguments, in the same format that xml->sxml itself takes. These entities and namespaces will be prepended to those given to the xml->sxml invocation.

     (define (handle-foo docname systemid internal-subset)
       (case docname
         ((foo)
          (values #:entities '((greets . "<i>Hello, world!</i>"))))
         (else
          (values))))
     (xml->sxml "<!DOCTYPE foo><p>&greets;</p>"
                #:doctype-handler handle-foo)
     ⇒ (*TOP* (p (i "Hello, world!")))

If the document has no doctype declaration, the doctype-handler is invoked with #f for the three arguments.

In the future, the continuation may accept other keyword arguments, for example to validate the parsed SXML against the doctype.

[Scheme Procedure] sxml->xml tree [port]

Serialize the SXML tree tree as XML. The output will be written to the current output port, unless the optional argument port is present.

[Scheme Procedure] sxml->string sxml

Detag an sxml tree sxml into a string. Does not perform any formatting.

7.22.3 SSAX: A Functional XML Parsing Toolkit

Guile’s XML parser is based on Oleg Kiselyov’s powerful XML parsing toolkit, SSAX.

7.22.3.1 History

Back in the 1990s, when the world was young again and XML was the solution to all of its problems, there were basically two kinds of XML parsers out there: DOM parsers and SAX parsers.

A DOM parser reads through an entire XML document, building up a tree of “DOM objects” representing the document structure. They are very easy to use, but sometimes you don’t actually want all of the information in a document; building an object tree is not necessary if all you want to do is to count word frequencies in a document, for example.

SAX parsers were created to give the programmer more control on the parsing process. A programmer gives the SAX parser a number of “callbacks”: functions that will be called on various features of the XML stream as they are encountered. SAX parsers are more efficient, but much harder to user, as users typically have to manually maintain a stack of open elements.

Kiselyov realized that the SAX programming model could be made much simpler if the callbacks were formulated not as a linear fold across the features of the XML stream, but as a tree fold over the structure implicit in the XML. In this way, the user has a very convenient, functional-style interface that can still generate optimal parsers.

The xml->sxml interface from the (sxml simple) module is a DOM-style parser built using SSAX, though it returns SXML instead of DOM objects.

7.22.3.2 Implementation

(sxml ssax) is a package of low-to-high level lexing and parsing procedures that can be combined to yield a SAX, a DOM, a validating parser, or a parser intended for a particular document type. The procedures in the package can be used separately to tokenize or parse various pieces of XML documents. The package supports XML Namespaces, internal and external parsed entities, user-controlled handling of whitespace, and validation. This module therefore is intended to be a framework, a set of “Lego blocks” you can use to build a parser following any discipline and performing validation to any degree. As an example of the parser construction, this file includes a semi-validating SXML parser.

SSAX has a “sequential” feel of SAX yet a “functional style” of DOM. Like a SAX parser, the framework scans the document only once and permits incremental processing. An application that handles document elements in order can run as efficiently as possible. Unlike a SAX parser, the framework does not require an application register stateful callbacks and surrender control to the parser. Rather, it is the application that can drive the framework – calling its functions to get the current lexical or syntax element. These functions do not maintain or mutate any state save the input port. Therefore, the framework permits parsing of XML in a pure functional style, with the input port being a monad (or a linear, read-once parameter).

Besides the port, there is another monad – seed. Most of the middle- and high-level parsers are single-threaded through the seed. The functions of this framework do not process or affect the seed in any way: they simply pass it around as an instance of an opaque datatype. User functions, on the other hand, can use the seed to maintain user’s state, to accumulate parsing results, etc. A user can freely mix his own functions with those of the framework. On the other hand, the user may wish to instantiate a high-level parser:

SSAX:make-elem-parser or SSAX:make-parser. In the latter case, the user must provide

functions of specific signatures, which are called at predictable moments during the parsing: to handle character data, element data, or processing instructions (PI). The functions are always given the seed, among other parameters, and must return the new seed.

From a functional point of view, XML parsing is a combined pre-post-order traversal of a “tree” that is the XML document itself. This down-and-up traversal tells the user about an element when its start tag is encountered. The user is notified about the element once more, after all element’s children have been handled. The process of XML parsing therefore is a fold over the raw XML document. Unlike a fold over trees defined in [1], the parser is necessarily single-threaded – obviously as elements in a text XML document are laid down sequentially. The parser therefore is a tree fold that has been transformed to accept an accumulating parameter [1,2].

Formally, the denotational semantics of the parser can be expressed as

       parser:: (Start-tag -> Seed -> Seed) ->
         (Start-tag -> Seed -> Seed -> Seed) ->
         (Char-Data -> Seed -> Seed) ->
         XML-text-fragment -> Seed -> Seed
       parser fdown fup fchar "<elem attrs> content </elem>" seed
        = fup "<elem attrs>" seed
      (parser fdown fup fchar "content" (fdown "<elem attrs>" seed))

       parser fdown fup fchar "char-data content" seed
        = parser fdown fup fchar "content" (fchar "char-data" seed)
       parser fdown fup fchar "elem-content content" seed
        = parser fdown fup fchar "content" (
      parser fdown fup fchar "elem-content" seed)

Compare the last two equations with the left fold

       fold-left kons elem:list seed = fold-left kons list (kons elem seed)

The real parser created by SSAX:make-parser is slightly more complicated, to account for processing instructions, entity references, namespaces, processing of document type declaration, etc.

The XML standard document referred to in this module is http://www.w3.org/TR/

1998/REC-xml-19980210.html

The present file also defines a procedure that parses the text of an XML document or of a separate element into SXML, an S-expression-based model of an XML Information Set. SXML is also an Abstract Syntax Tree of an XML document. SXML is similar but not identical to DOM; SXML is particularly suitable for Scheme-based XML/HTML authoring, SXPath queries, and tree transformations. See SXML.html for more details. SXML is a term implementation of evaluation of the XML document [3]. The other implementation is context-passing.

The present frameworks fully supports the XML Namespaces Recommendation: http://

www.w3.org/TR/REC-xml-names/.

Other links:

[1] Jeremy Gibbons, Geraint Jones, "The Under-appreciated Unfold," Proc. ICFP’98, 1998, pp. 273-279.

[2] Richard S. Bird, The promotion and accumulation strategies in transformational programming, ACM Trans. Progr. Lang. Systems, 6(4):487-504, October 1984.

[3] Ralf Hinze, "Deriving Backtracking Monad Transformers," Functional Pearl. Proc ICFP’00, pp. 186-197.

7.22.3.3 Usage

[Scheme Procedure] current-ssax-error-port

[Scheme Procedure] with-ssax-error-to-port port thunk [Scheme Procedure] xml-token?

      -- Scheme Procedure: pair? x
          Return `#t' if X is a pair; otherwise return `#f'.

[Scheme Syntax] xml-token-kind token [Scheme Syntax] xml-token-head token

[Scheme Procedure] make-empty-attlist

[Scheme Procedure] attlist-add attlist name-value

[Scheme Procedure] attlist-null? x

Return #t if x is the empty list, else #f.

[Scheme Procedure] attlist-remove-top attlist [Scheme Procedure] attlist->alist attlist [Scheme Procedure] attlist-fold kons knil lis1 [Scheme Procedure] define-parsed-entity! entity str

Define a new parsed entity. entity should be a symbol.

Instances of &entity; in XML text will be replaced with the string str, which will then be parsed.

[Scheme Procedure] reset-parsed-entity-definitions!

Restore the set of parsed entity definitions to its initial state.

[Scheme Procedure] ssax:uri-string->symbol uri-str

[Scheme Procedure] ssax:skip-internal-dtd port
[Scheme Procedure] ssax:read-pi-body-as-string port
[Scheme Procedure] ssax:reverse-collect-str-drop-ws fragments
[Scheme Procedure] ssax:read-markup-token port

[Scheme Procedure] ssax:read-cdata-body port str-handler seed [Scheme Procedure] ssax:read-char-ref port [Scheme Procedure] ssax:read-attributes port entities

[Scheme Procedure] ssax:complete-start-tag tag-head port elems entities

namespaces

[Scheme Procedure] ssax:read-external-id port

[Scheme Procedure] ssax:read-char-data port expect-eof? str-handler seed [Scheme Procedure] ssax:xml->sxml port namespace-prefix-assig

[Scheme Syntax] ssax:make-parser . kw-val-pairs [Scheme Syntax] ssax:make-pi-parser orig-handlers

[Scheme Syntax] ssax:make-elem-parser my-new-level-seed my-finish-element

my-char-data-handler my-pi-handlers

7.22.4 Transforming SXML 7.22.4.1 Overview

SXML expression tree transformers

Pre-Post-order traversal of a tree and creation of a new tree

      pre-post-order:: <tree> x <bindings> -> <new-tree>

where

       <bindings> ::= (<binding> ...)
       <binding> ::= (<trigger-symbol> *preorder* . <handler>) |
                     (<trigger-symbol> *macro* . <handler>) |
      (<trigger-symbol> <new-bindings> . <handler>) |
      (<trigger-symbol> . <handler>)
       <trigger-symbol> ::= XMLname | *text* | *default*
       <handler> :: <trigger-symbol> x [<tree>] -> <new-tree>

The pre-post-order function visits the nodes and nodelists pre-post-order (depth-first). For each <Node> of the form (name <Node> ...), it looks up an association with the given name among its <bindings>. If failed, pre-post-order tries to locate a *default* binding. It’s an error if the latter attempt fails as well. Having found a binding, the pre-post-order function first checks to see if the binding is of the form

      (<trigger-symbol> *preorder* . <handler>)

If it is, the handler is ’applied’ to the current node. Otherwise, the pre-post-order function first calls itself recursively for each child of the current node, with <new-bindings> prepended to the <bindings> in effect. The result of these calls is passed to the <handler> (along with the head of the current <Node>). To be more precise, the handler is applied to the head of the current node and its processed children. The result of the handler, which should also be a <tree>, replaces the current <Node>. If the current <Node> is a text string or other atom, a special binding with a symbol *text* is looked up.

A binding can also be of a form

      (<trigger-symbol> *macro* . <handler>)

This is equivalent to *preorder* described above. However, the result is re-processed again, with the current stylesheet.

7.22.4.2 Usage

[Scheme Procedure] SRV:send-reply . fragments

Output the fragments to the current output port.

The fragments are a list of strings, characters, numbers, thunks, #f, #t – and other fragments. The function traverses the tree depth-first, writes out strings and characters, executes thunks, and ignores #f and '(). The function returns #t if anything was written at all; otherwise the result is #f If #t occurs among the fragments, it is not written out but causes the result of SRV:send-reply to be #t.

[Scheme Procedure] foldts fdown fup fhere seed tree

[Scheme Procedure] post-order tree bindings

[Scheme Procedure] pre-post-order tree bindings

[Scheme Procedure] replace-range beg-pred end-pred forest

7.22.5 SXML Tree Fold 7.22.5.1 Overview

(sxml fold) defines a number of variants of the fold algorithm for use in transforming SXML trees. Additionally it defines the layout operator, fold-layout, which might be described as a context-passing variant of SSAX’s pre-post-order.

7.22.5.2 Usage

[Scheme Procedure] foldt fup fhere tree The standard multithreaded tree fold. fup is of type [a] -> a. fhere is of type object -> a.

[Scheme Procedure] foldts fdown fup fhere seed tree The single-threaded tree fold originally defined in SSAX. See Section 7.22.3 [SSAX], page 700, for more information.

[Scheme Procedure] foldts* fdown fup fhere seed tree A variant of foldts that allows pre-order tree rewrites. Originally defined in Andy Wingo’s 2007 paper, Applications of fold to XML transformation.

[Scheme Procedure] fold-values proc list . seeds A variant of fold that allows multi-valued seeds. Note that the order of the arguments differs from that of fold. See Section 7.5.3.5 [SRFI-1 Fold and Map], page 556.

[Scheme Procedure] foldts*-values fdown fup fhere tree . seeds A variant of foldts* that allows multi-valued seeds. Originally defined in Andy Wingo’s 2007 paper, Applications of fold to XML transformation.

[Scheme Procedure] fold-layout tree bindings params layout stylesheet A traversal combinator in the spirit of pre-post-order. See Section 7.22.4 [Transforming SXML], page 703. fold-layout was originally presented in Andy Wingo’s 2007 paper, Applications of fold to XML transformation.

     bindings := (<binding>...)
     binding  := (<tag> <bandler-pair>...)
               | (*default* . <post-handler>)
               | (*text* . <text-handler>)
     tag      := <symbol>
     handler-pair := (pre-layout . <pre-layout-handler>)
               | (post . <post-handler>)
               | (bindings . <bindings>)
               | (pre . <pre-handler>)
               | (macro . <macro-handler>)

pre-layout-handler A function of three arguments:

kids the kids of the current node, before traversal

params the params of the current node layout the layout coming into this node

pre-layout-handler is expected to use this information to return a layout to pass to the kids. The default implementation returns the layout given in the arguments.

post-handler A function of five arguments:

tag the current tag being processed

params the params of the current node

layout the layout coming into the current node, before any kids were processed

klayout the layout after processing all of the children

kids the already-processed child nodes

post-handler should return two values, the layout to pass to the next node and the final tree.

text-handler text-handler is a function of three arguments:

text the string

params the current params

layout the current layout

text-handler should return two values, the layout to pass to the next node and the value to which the string should transform.

7.22.6 SXPath

7.22.6.1 Overview

SXPath: SXML Query Language

SXPath is a query language for SXML, an instance of XML Information set (Infoset) in the form of s-expressions. See (sxml ssax) for the definition of SXML and more details. SXPath is also a translation into Scheme of an XML Path Language, XPath (http://www. w3.org/TR/xpath). XPath and SXPath describe means of selecting a set of Infoset’s items or their properties.

To facilitate queries, XPath maps the XML Infoset into an explicit tree, and introduces important notions of a location path and a current, context node. A location path denotes a selection of a set of nodes relative to a context node. Any XPath tree has a distinguished, root node – which serves as the context node for absolute location paths. Location path is recursively defined as a location step joined with a location path. A location step is a simple query of the database relative to a context node. A step may include expressions that further filter the selected set. Each node in the resulting set is used as a context node for the adjoining location path. The result of the step is a union of the sets returned by the latter location paths.

The SXML representation of the XML Infoset (see SSAX.scm) is rather suitable for querying as it is. Bowing to the XPath specification, we will refer to SXML information items as ’Nodes’:

       <Node> ::= <Element> | <attributes-coll> | <attrib>
          | "text string" | <PI>

This production can also be described as

     <Node> ::= (name . <Nodeset>) | "text string"

An (ordered) set of nodes is just a list of the constituent nodes:

       <Nodeset> ::= (<Node> ...)

Nodesets, and Nodes other than text strings are both lists. A <Nodeset> however is either an empty list, or a list whose head is not a symbol. A symbol at the head of a node is either an XML name (in which case it’s a tag of an XML element), or an administrative name such as ’@’. This uniform list representation makes processing rather simple and elegant, while avoiding confusion. The multi-branch tree structure formed by the mutually-recursive datatypes <Node> and <Nodeset> lends itself well to processing by functional languages.

A location path is in fact a composite query over an XPath tree or its branch. A singe step is a combination of a projection, selection or a transitive closure. Multiple steps are combined via join and union operations. This insight allows us to elegantly implement XPath as a sequence of projection and filtering primitives – converters – joined by combinators. Each converter takes a node and returns a nodeset which is the result of the corresponding query relative to that node. A converter can also be called on a set of nodes. In that case it returns a union of the corresponding queries over each node in the set. The union is easily implemented as a list append operation as all nodes in a SXML tree are considered distinct, by XPath conventions. We also preserve the order of the members in the union. Query combinators are high-order functions: they take converter(s) (which is a Node|Nodeset -> Nodeset function) and compose or otherwise combine them. We will be concerned with only relative location paths [XPath]: an absolute location path is a relative path applied to the root node.

Similarly to XPath, SXPath defines full and abbreviated notations for location paths. In both cases, the abbreviated notation can be mechanically expanded into the full form by simple rewriting rules. In case of SXPath the corresponding rules are given as comments to a sxpath function, below. The regression test suite at the end of this file shows a representative sample of SXPaths in both notations, juxtaposed with the corresponding XPath expressions. Most of the samples are borrowed literally from the XPath specification, while the others are adjusted for our running example, tree1.

7.22.6.2 Usage

[Scheme Procedure] nodeset? x

[Scheme Procedure] node-typeof? crit

[Scheme Procedure] node-eq? other

[Scheme Procedure] node-equal? other

[Scheme Procedure] node-pos n [Scheme Procedure] filter pred?

-- Scheme Procedure: filter pred list
Return all the elements of 2nd arg LIST that satisfy predicate
PRED.  The list is not disordered - elements that appear in the
result list occur in the same order as they occur in the argument
list.  The returned list may share a common tail with the argument
list.  The dynamic order in which the various applications of pred
are made is not specified.
     (filter even? '(0 7 8 8 43 -4)) => (0 8 8 -4)

[Scheme Procedure] take-until pred?

[Scheme Procedure] take-after pred?

[Scheme Procedure] map-union proc lst

[Scheme Procedure] node-reverse node-or-nodeset

[Scheme Procedure] node-trace title

[Scheme Procedure] select-kids test-pred?

[Scheme Procedure] node-self pred?

-- Scheme Procedure: filter pred list
Return all the elements of 2nd arg LIST that satisfy predicate
PRED.  The list is not disordered - elements that appear in the
result list occur in the same order as they occur in the argument
list.  The returned list may share a common tail with the argument
list.  The dynamic order in which the various applications of pred
are made is not specified.
     (filter even? '(0 7 8 8 43 -4)) => (0 8 8 -4)

[Scheme Procedure] node-join . selectors

[Scheme Procedure] node-reduce . converters

[Scheme Procedure] node-or . converters

[Scheme Procedure] node-closure test-pred?

[Scheme Procedure] node-parent rootnode

[Scheme Procedure] sxpath path

7.22.7 (sxml ssax input-parse)

7.22.7.1 Overview

A simple lexer.

The procedures in this module surprisingly often suffice to parse an input stream. They either skip, or build and return tokens, according to inclusion or delimiting semantics. The list of characters to expect, include, or to break at may vary from one invocation of a function to another. This allows the functions to easily parse even context-sensitive languages.

EOF is generally frowned on, and thrown up upon if encountered. Exceptions are mentioned specifically. The list of expected characters (characters to skip until, or breakcharacters) may include an EOF "character", which is to be coded as the symbol, *eof*.

The input stream to parse is specified as a port, which is usually the last (and optional) argument. It defaults to the current input port if omitted.

If the parser encounters an error, it will throw an exception to the key parser-error.

The arguments will be of the form (port message specialising-msg*).

The first argument is a port, which typically points to the offending character or its neighborhood. You can then use port-column and port-line to query the current position. message is the description of the error. Other arguments supply more details about the problem.

7.22.7.2 Usage

[Scheme Procedure] peek-next-char [port]

[Scheme Procedure] assert-curr-char expected-chars comment [port]

[Scheme Procedure] skip-until arg [port]

[Scheme Procedure] skip-while skip-chars [port]

[Scheme Procedure] next-token prefix-skipped-chars break-chars [comment] [port]

[Scheme Procedure] next-token-of incl-list/pred [port]

[Scheme Procedure] read-text-line [port] [Scheme Procedure] read-string n [port]

[Scheme Procedure] find-string-from-port?  .

Looks for str in <input-port>, optionally within the first max-no-char characters.

7.22.8 (sxml apply-templates)

7.22.8.1 Overview

Pre-order traversal of a tree and creation of a new tree:

      apply-templates:: tree x <templates> -> <new-tree>

where

       <templates> ::= (<template> ...)
       <template>  ::= (<node-test> <node-test> ... <node-test> . <handler>)
       <node-test> ::= an argument to node-typeof? above
       <handler>   ::= <tree> -> <new-tree>

This procedure does a normal, pre-order traversal of an SXML tree. It walks the tree, checking at each node against the list of matching templates.

If the match is found (which must be unique, i.e., unambiguous), the corresponding handler is invoked and given the current node as an argument. The result from the handler, which must be a <tree>, takes place of the current node in the resulting tree. The name of the function is not accidental: it resembles rather closely an apply-templates function of XSLT.

7.22.8.2 Usage

[Scheme Procedure] apply-templates tree templates

7.23 Texinfo Processing

7.23.1 (texinfo)

7.23.1.1 Overview

Texinfo processing in scheme

This module parses texinfo into SXML. TeX will always be the processor of choice for print output, of course. However, although makeinfo works well for info, its output in other formats is not very customizable, and the program is not extensible as a whole. This module aims to provide an extensible framework for texinfo processing that integrates texinfo into the constellation of SXML processing tools.

Notes on the SXML vocabulary

Consider the following texinfo fragment:

      @deffn Primitive set-car! pair value
      This function...
      @end deffn

Logically, the category (Primitive), name (set-car!), and arguments (pair value) are “attributes” of the deffn, with the description as the content. However, texinfo allows for @-commands within the arguments to an environment, like @deffn, which means that texinfo “attributes” are PCDATA. XML attributes, on the other hand, are CDATA. For this reason, “attributes” of texinfo @-commands are called “arguments”, and are grouped under the special element, ‘%’.

Because ‘%’ is not a valid NCName, stexinfo is a superset of SXML. In the interests of interoperability, this module provides a conversion function to replace the ‘%’ with ‘texinfoarguments’.

7.23.1.2 Usage

[Function] call-with-file-and-dir filename proc

Call the one-argument procedure proc with an input port that reads from filename. During the dynamic extent of proc’s execution, the current directory will be (dirname filename). This is useful for parsing documents that can include files by relative path name.

[Variable] texi-command-specs

[Function] texi-command-depth command max-depth

Given the texinfo command command, return its nesting level, or #f if it nests too deep for max-depth. Examples:

      (texi-command-depth 'chapter 4)  ⇒ 1
      (texi-command-depth 'top 4)  ⇒ 0
      (texi-command-depth 'subsection 4)  ⇒ 3
      (texi-command-depth 'appendixsubsec 4) ⇒ 3
      (texi-command-depth 'subsection 2)  ⇒ #f

[Function] texi-fragment->stexi string-or-port

Parse the texinfo commands in string-or-port, and return the resultant stexi tree. The head of the tree will be the special command, *fragment*.

[Function] texi->stexi port

Read a full texinfo document from port and return the parsed stexi tree. The parsing will start at the @settitle and end at @bye or EOF.

[Function] stexi->sxml tree

Transform the stexi tree tree into sxml. This involves replacing the % element that keeps the texinfo arguments with an element for each argument. FIXME: right now it just changes % to texinfo-arguments – that doesn’t hang with the idea of making a dtd at some point

7.23.2 (texinfo docbook)

7.23.2.1 Overview

This module exports procedures for transforming a limited subset of the SXML representation of docbook into stexi. It is not complete by any means. The intention is to gather a number of routines and stylesheets so that external modules can parse specific subsets of docbook, for example that set generated by certain tools.

7.23.2.2 Usage

[Variable] *sdocbook->stexi-rules*
[Variable] *sdocbook-block-commands*

[Function] sdocbook-flatten sdocbook

"Flatten" a fragment of sdocbook so that block elements do not nest inside each other. Docbook is a nested format, where e.g. a refsect2 normally appears inside a refsect1. Logical divisions in the document are represented via the tree topology; a refsect2 element contains all of the elements in its section. On the contrary, texinfo is a flat format, in which sections are marked off by standalone section headers like @subsection, and block elements do not nest inside each other. This function takes a nested sdocbook fragment sdocbook and flattens all of the sections, such that e.g.

      (refsect1 (refsect2 (para "Hello")))

becomes

      ((refsect1) (refsect2) (para "Hello"))

Oftentimes (always?) sectioning elements have <title> as their first element child; users interested in processing the refsect* elements into proper sectioning elements

like chapter might be interested in replace-titles and filter-empty-elements.

See [replace-titles], page 712, and [filter-empty-elements], page 712. Returns a nodeset; that is to say, an untagged list of stexi elements. See Section 7.22.6 [SXPath], page 706, for the definition of a nodeset.

[Function] filter-empty-elements sdocbook

Filters out empty elements in an sdocbook nodeset. Mostly useful after running

sdocbook-flatten.

[Function] replace-titles sdocbook-fragment Iterate over the sdocbook nodeset sdocbook-fragment, transforming contiguous refsect and title elements into the appropriate texinfo sectioning command. Most useful after having run sdocbook-flatten. For example:

      (replace-titles '((refsect1) (title "Foo") (para "Bar.")))
         ⇒ '((chapter "Foo") (para "Bar."))

7.23.3 (texinfo html)

7.23.3.1 Overview

This module implements transformation from stexi to HTML. Note that the output of stexi->shtml is actually SXML with the HTML vocabulary. This means that the output can be further processed, and that it must eventually be serialized by sxml->xml. See Section 7.22.2 [Reading and Writing XML], page 698.

References (i.e., the @ref family of commands) are resolved by a ref-resolver. See [texinfo html add-ref-resolver!], page 712.

7.23.3.2 Usage

[Function] add-ref-resolver! proc

Add proc to the head of the list of ref-resolvers. proc will be expected to take the name of a node and the name of a manual and return the URL of the referent, or #f to pass control to the next ref-resolver in the list. The default ref-resolver will return the concatenation of the manual name, #, and the node name.

[Function] stexi->shtml tree

Transform the stexi tree into shtml, resolving references via ref-resolvers. See the module commentary for more details.

[Function] urlify str

7.23.4 (texinfo indexing)

7.23.4.1 Overview

Given a piece of stexi, return an index of a specified variety.

Note that currently, stexi-extract-index doesn’t differentiate between different kinds of index entries. That’s a bug ;)

7.23.4.2 Usage

[Function] stexi-extract-index tree manual-name kind

Given an stexi tree tree, index all of the entries of type kind. kind can be one of the

predefined texinfo indices (concept, variable, function, key, program, type) or

one of the special symbols auto or all. auto will scan the stext for a (printindex) statement, and all will generate an index from all entries, regardless of type. The returned index is a list of pairs, the car of which is the entry (a string) and the cdr of which is a node name (a string).

7.23.5 (texinfo string-utils)

7.23.5.1 Overview

Module ‘(texinfo string-utils)’ provides various string-related functions useful to Guile’s texinfo support.

7.23.5.2 Usage

[Function] escape-special-chars str special-chars escape-char Returns a copy of str with all given special characters preceded by the given escapechar. special-chars can either be a single character, or a string consisting of all the special characters.

     ;; make a string regexp-safe...
      (escape-special-chars "***(Example String)***"
                           "[]()/*."
                           #\\)
     => "\\*\\*\\*\\(Example String\\)\\*\\*\\*"
     ;; also can escape a singe char...
      (escape-special-chars "richardt@vzavenue.net"
                           #\@
                           #\@)
     => "richardt@@vzavenue.net"

[Function] transform-string str match? replace [start] [end] Uses match? against each character in str, and performs a replacement on each character for which matches are found. match? may either be a function, a character, a string, or #t. If match? is a function, then it takes a single character as input, and should return ‘#t’ for matches. match? is a character, it is compared to each string character using char=?. If match? is a string, then any character in that string will be considered a match. #t will cause every character to be a match. If replace is a function, it is called with the matched character as an argument, and the returned value is sent to the output string via ‘display’. If replace is anything else, it is sent through the output string via ‘display’. Note that te replacement for the matched characters does not need to be a single character. That is what differentiates this function from ‘string-map’, and what makes it useful for applications such as converting ‘#\&’ to ‘"&amp;"’ in web page text. Some other functions in this module are just wrappers around common uses of ‘transform-string’. Transformations not possible with this function should probably be done with regular expressions. If start and end are given, they control which portion of the string undergoes transformation. The entire input string is still output, though. So, if start is ‘5’, then the first five characters of str will still appear in the returned string.

     ; these two are equivalent...
      (transform-string str #\space #\-) ; change all spaces to -'s
      (transform-string str (lambda (c) (char=? #\space c)) #\-)

[Function] expand-tabs str [tab-size]

Returns a copy of str with all tabs expanded to spaces. tab-size defaults to 8. Assuming tab size of 8, this is equivalent to:

      (transform-string str #\tab "        ")

[Function] center-string str [width] [chr] [rchr]

Returns a copy of str centered in a field of width characters. Any needed padding is done by character chr, which defaults to ‘#\space’. If rchr is provided, then the padding to the right will use it instead. See the examples below. left and rchr on the right. The default width is 80. The default chr and rchr is ‘#\space’. The string is never truncated.

      (center-string "Richard Todd" 24)
     => "      Richard Todd      "
      (center-string " Richard Todd " 24 #\=)
     => "===== Richard Todd ====="
      (center-string " Richard Todd " 24 #\< #\>)
     => "<<<<< Richard Todd >>>>>"

[Function] left-justify-string str [width] [chr]

left-justify-string str [width chr]. Returns a copy of str padded with chr such that it is left justified in a field of width characters. The default width is 80. Unlike ‘string-pad’ from srfi-13, the string is never truncated.

[Function] right-justify-string str [width] [chr]

Returns a copy of str padded with chr such that it is right justified in a field of width characters. The default width is 80. The default chr is ‘#\space’. Unlike ‘string-pad’ from srfi-13, the string is never truncated.

[Function] collapse-repeated-chars str [chr] [num]

Returns a copy of str with all repeated instances of chr collapsed down to at most num instances. The default value for chr is ‘#\space’, and the default value for num is 1.

(collapse-repeated-chars "H  e  l  l  o")
=> "H e l l o"
(collapse-repeated-chars "H--e--l--l--o" #\-)
=> "H-e-l-l-o"
(collapse-repeated-chars "H-e--l---l----o" #\- 2)
=> "H-e--l--l--o"

[Function] make-text-wrapper [#:line-width] [#:expand-tabs?] [#:tab-width] [#:collapse-whitespace?] [#:subsequent-indent] [#:initial-indent] [#:break-long-words?]

Returns a procedure that will split a string into lines according to the given parameters.

#:line-width

This is the target length used when deciding where to wrap lines. Default is 80.

#:expand-tabs?

Boolean describing whether tabs in the input should be expanded. Default is #t.

#:tab-width

If tabs are expanded, this will be the number of spaces to which they expand. Default is 8.

#:collapse-whitespace?

Boolean describing whether the whitespace inside the existing text should be removed or not. Default is #t. If text is already well-formatted, and is just being wrapped to fit in a different width, then set this to ‘#f’. This way, many common text conventions (such as two spaces between sentences) can be preserved if in the original text. If the input text spacing cannot be trusted, then leave this setting at the default, and all repeated whitespace will be collapsed down to a single space.

#:initial-indent

Defines a string that will be put in front of the first line of wrapped text. Default is the empty string, “”.

#:subsequent-indent

Defines a string that will be put in front of all lines of wrapped text, except the first one. Default is the empty string, “”.

#:break-long-words?

If a single word is too big to fit on a line, this setting tells the wrapper what to do. Defaults to #t, which will break up long words. When set to #f, the line will be allowed, even though it is longer than the defined

          #:line-width.

The return value is a procedure of one argument, the input string, which returns a list of strings, where each element of the list is one line.

[Function] fill-string str . kwargs

Wraps the text given in string str according to the parameters provided in kwargs, or the default setting if they are not given. Returns a single string with the wrapped text. Valid keyword arguments are discussed in make-text-wrapper.

[Function] string->wrapped-lines str . kwargs

string->wrapped-lines str keywds .... Wraps the text given in string str according to the parameters provided in keywds, or the default setting if they are not given. Returns a list of strings representing the formatted lines. Valid keyword arguments

are discussed in make-text-wrapper.

7.23.6 (texinfo plain-text)

7.23.6.1 Overview

Transformation from stexi to plain-text. Strives to re-create the output from info; comes pretty damn close.

7.23.6.2 Usage

[Function] stexi->plain-text tree

Transform tree into plain text. Returns a string.

7.23.7 (texinfo serialize)

7.23.7.1 Overview

Serialization of stexi to plain texinfo.

7.23.7.2 Usage

[Function] stexi->texi tree

Serialize the stexi tree into plain texinfo.

7.23.8 (texinfo reflection)

7.23.8.1 Overview

Routines to generare stexi documentation for objects and modules.

Note that in this context, an object is just a value associated with a location. It has nothing to do with GOOPS.

7.23.8.2 Usage

[Function] module-stexi-documentation sym-name [%docs-resolver]

[#:docs-resolver] Return documentation for the module named sym-name. The documentation will be formatted as stexi (see Section 7.23.1 [texinfo], page 710).

[Function] script-stexi-documentation scriptpath

Return documentation for given script. The documentation will be taken from the script’s commentary, and will be returned in the stexi format (see Section 7.23.1 [texinfo], page 710).

[Function] object-stexi-documentation [ ] [#:force]

[Function] package-stexi-standard-copying name version updated years

copyright-holder permissions Create a standard texinfo copying section.

years is a list of years (as integers) in which the modules being documented were released. All other arguments are strings.

[Function] package-stexi-standard-titlepage name version updated authors

Create a standard GNU title page.

authors is a list of (name . email) pairs. All other arguments are strings.

Here is an example of the usage of this procedure:

       (package-stexi-standard-titlepage
        "Foolib"
        "3.2"
        "26 September 2006"
        '(("Alyssa P Hacker" . "alyssa@example.com"))
        '(2004 2005 2006)
        "Free Software Foundation, Inc."
        "Standard GPL permissions blurb goes here")
[Function] package-stexi-generic-menu name entries

Create a menu from a generic alist of entries, the car of which should be the node name, and the cdr the description. As an exception, an entry of #f will produce a separator.

[Function] package-stexi-standard-menu name modules module-descriptions

extra-entries Create a standard top node and menu, suitable for processing by makeinfo.

[Function] package-stexi-extended-menu name module-pairs script-pairs

extra-entries Create an "extended" menu, like the standard menu but with a section for scripts.

[Function] package-stexi-standard-prologue name filename category

description copying titlepage menu Create a standard prologue, suitable for later serialization to texinfo and .info creation with makeinfo.

Returns a list of stexinfo forms suitable for passing to package-stexi-documentation as the prologue. See [texinfo reflection package-stexi-documentation], page 718, [texinfo reflection package-stexi-standard-titlepage], page 717, [texinfo reflection package-stexi-standard-copying], page 717, and [texinfo reflection package-stexi-standard-menu], page 717.

[Function] package-stexi-documentation modules name filename prologue

epilogue [#:module-stexi-documentation-args] [#:scripts] Create stexi documentation for a package, where a package is a set of modules that is released together. modules is expected to be a list of module names, where a module name is a list of symbols. The stexi that is returned will be titled name and a texinfo filename of filename. prologue and epilogue are lists of stexi forms that will be spliced into the output document before and after the generated modules documentation, respectively. See [texinfo reflection package-stexi-standard-prologue], page 717, to create a conventional GNU texinfo prologue. module-stexi-documentation-args is an optional argument that, if given, will be added to the argument list when module-texi-documentation is called. For example, it might be useful to define a #:docs-resolver argument.

[Function] package-stexi-documentation-for-include modules

module-descriptions [#:module-stexi-documentation-args] Create stexi documentation for a package, where a package is a set of modules that is released together. modules is expected to be a list of module names, where a module name is a list of symbols. Returns an stexinfo fragment. Unlike package-stexi-documentation, this function simply produces a menu and the module documentations instead of producing a full texinfo document. This can be useful if you write part of your manual by hand, and just use @include to pull in the automatically generated parts. module-stexi-documentation-args is an optional argument that, if given, will be added to the argument list when module-texi-documentation is called. For example, it might be useful to define a #:docs-resolver argument.

8 GOOPS

GOOPS is the object oriented extension to Guile. Its implementation is derived from STk-3.99.3 by Erick Gallesio and version 1.3 of Gregor Kiczales’ Tiny-Clos. It is very close in spirit to CLOS, the Common Lisp Object System, but is adapted for the Scheme language.

GOOPS is a full object oriented system, with classes, objects, multiple inheritance, and generic functions with multi-method dispatch. Furthermore its implementation relies on a meta object protocol — which means that GOOPS’s core operations are themselves defined as methods on relevant classes, and can be customised by overriding or redefining those methods.

To start using GOOPS you first need to import the (oop goops) module. You can do this at the Guile REPL by evaluating:

     (use-modules (oop goops))

8.1 Copyright Notice

The material in this chapter is partly derived from the STk Reference Manual written by Erick Gallesio, whose copyright notice is as follows.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-1999 Erick Gallesio - I3S-CNRS/ESSI <eg@unice.fr> Permission to use, copy, modify, distribute,and license this software and its documentation for any purpose is hereby granted, provided that existing copyright notices are retained in all copies and that this notice is included verbatim in any distributions. No written agreement, license, or royalty fee is required for any of the authorized uses. This software is provided “AS IS” without express or implied warranty.

The material has been adapted for use in Guile, with the author’s permission.

8.2 Class Definition

A new class is defined with the define-class syntax:

     (define-class class (superclass ...)
        slot-description ...
        class-option ...)

class is the class being defined. The list of superclasses specifies which existing classes, if any, to inherit slots and properties from. Slots hold per-instance^{1} data, for instances of that class — like “fields” or “member variables” in other object oriented systems. Each slot-description gives the name of a slot and optionally some “properties” of this slot; for example its initial value, the name of a function which will access its value, and so on. Class options, slot descriptions and inheritance are discussed more below.

[syntax] define-class name (super …) slot-definition … class-option … Define a class called name that inherits from supers, with direct slots defined by slot-definitions and class-options. The newly created class is bound to the variable name name in the current environment. Each slot-definition is either a symbol that names the slot or a list,

1 Usually — but see also the #:allocation slot option.

     (slot-name-symbol . slot-options)

where slot-name-symbol is a symbol and slot-options is a list with an even number of elements. The even-numbered elements of slot-options (counting from zero) are slot option keywords; the odd-numbered elements are the corresponding values for those keywords. Each class-option is an option keyword and corresponding value.

As an example, let us define a type for representing a complex number in terms of two real numbers.^{2} This can be done with the following class definition:

(define-class <my-complex> (<number>)
   r i)

This binds the variable <my-complex> to a new class whose instances will contain two slots. These slots are called r and i and will hold the real and imaginary parts of a complex number. Note that this class inherits from <number>, which is a predefined class.^{3}

Slot options are described in the next section. The possible class options are as follows.

[class option] #:metaclass metaclass

The #:metaclass class option specifies the metaclass of the class being defined. metaclass must be a class that inherits from <class>. For the use of metaclasses, see Section 8.11.1 [Metaobjects and the Metaobject Protocol], page 743 and Section 8.11.2 [Metaclasses], page 745. If the #:metaclass option is absent, GOOPS reuses or constructs a metaclass for the new class by calling ensure-metaclass (see Section 8.11.5 [ensure-metaclass], page 747).

[class option] #:name name

The #:name class option specifies the new class’s name. This name is used to identify the class whenever related objects - the class itself, its instances and its subclasses - are printed. If the #:name option is absent, GOOPS uses the first argument to define-class as the class name.

8.3 Instance Creation and Slot Access

An instance (or object) of a defined class can be created with make. make takes one mandatory parameter, which is the class of the instance to create, and a list of optional arguments that will be used to initialize the slots of the new instance. For instance the following form

(define c (make <my-complex>))

creates a new <my-complex> object and binds it to the Scheme variable c.

[generic] make

[method] make (class <class>) initarg … Create and return a new instance of class class, initialized using initarg . . ..

2 Of course Guile already provides complex numbers, and <complex> is in fact a predefined class in GOOPS;

but the definition here is still useful as an example.

3 <number> is the direct superclass of the predefined class <complex>; <complex> is the superclass of

<real>, and <real> is the superclass of <integer>.

In theory, initarg … can have any structure that is understood by whatever methods get applied when the initialize generic function is applied to the newly allocated instance. In practice, specialized initialize methods would normally call (next-method), and so eventually the standard GOOPS initialize methods are applied. These methods expect initargs to be a list with an even number of elements, where evennumbered elements (counting from zero) are keywords and odd-numbered elements are the corresponding values. GOOPS processes initialization argument keywords automatically for slots whose definition includes the #:init-keyword option (see Section 8.4 [init-keyword], page 721). Other keyword value pairs can only be processed by an initialize method that is specialized for the new instance’s class. Any unprocessed keyword value pairs are ignored.

[generic] make-instance

[method] make-instance (class <class>) initarg …

make-instance is an alias for make.

The slots of the new complex number can be accessed using slot-ref and slot-set!. slot-set! sets the value of an object slot and slot-ref retrieves it.

(slot-set! c 'r 10)
(slot-set! c 'i 3)
(slot-ref c 'r) ⇒ 10
(slot-ref c 'i) ⇒ 3

The (oop goops describe) module provides a describe function that is useful for seeing all the slots of an object; it prints the slots and their values to standard output.

(describe c)

⊣

#<<my-complex> 401d8638> is an instance of class <my-complex>
Slots are:
     r = 10
     i = 3

8.4 Slot Options

When specifying a slot (in a (define-class ...) form), various options can be specified in addition to the slot’s name. Each option is specified by a keyword. The list of possible keywords is as follows.

[slot option] #:init-value init-value [slot option] #:init-form init-form [slot option] #:init-thunk init-thunk [slot option] #:init-keyword init-keyword

These options provide various ways to specify how to initialize the slot’s value at instance creation time. init-value specifies a fixed initial slot value (shared across all new instances of the class). init-thunk specifies a thunk that will provide a default value for the slot. The thunk is called when a new instance is created and should return the desired initial slot value. init-form specifies a form that, when evaluated, will return an initial value for the slot. The form is evaluated each time that an instance of the class is created, in the lexical environment of the containing define-class expression. init-keyword specifies a keyword that can be used to pass an initial slot value to make when creating a new instance. Note that, since an init-value value is shared across all instances of a class, you should only use it when the initial value is an immutable value, like a constant. If you want to initialize a slot with a fresh, independently mutable value, you should use init-thunk or init-form instead. Consider the following example.

     (define-class <chbouib> ()
       (hashtab #:init-value (make-hash-table)))

Here only one hash table is created and all instances of <chbouib> have their hashtab slot refer to it. In order to have each instance of <chbouib> refer to a new hash table, you should instead write:

     (define-class <chbouib> ()
       (hashtab #:init-thunk make-hash-table))

or:

     (define-class <chbouib> ()
       (hashtab #:init-form (make-hash-table)))

If more than one of these options is specified for the same slot, the order of precedence, highest first is

• #:init-keyword, if init-keyword is present in the options passed to make

  • #:init-thunk, #:init-form or #:init-value.

If the slot definition contains more than one initialization option of the same precedence, the later ones are ignored. If a slot is not initialized at all, its value is unbound. In general, slots that are shared between more than one instance are only initialized at new instance creation time if the slot value is unbound at that time. However, if the new instance creation specifies a valid init keyword and value for a shared slot, the slot is re-initialized regardless of its previous value. Note, however, that the power of GOOPS’ metaobject protocol means that everything written here may be customized or overridden for particular classes! The slot initializations described here are performed by the least specialized method of the generic function initialize, whose signature is

     (define-method (initialize (object <object>) initargs) ...)

The initialization of instances of any given class can be customized by defining a initialize method that is specialized for that class, and the author of the specialized method may decide to call next-method - which will result in a call to the next less specialized initialize method - at any point within the specialized code, or maybe not at all. In general, therefore, the initialization mechanisms described here may be modified or overridden by more specialized code, or may not be supported at all for particular classes.

[slot option] #:getter getter [slot option] #:setter setter [slot option] #:accessor accessor

Given an object obj with slots named foo and bar, it is always possible to read and write those slots by calling slot-ref and slot-set! with the relevant slot name; for example:

     (slot-ref obj 'foo)
     (slot-set! obj 'bar 25)

The #:getter, #:setter and #:accessor options, if present, tell GOOPS to create generic function and method definitions that can be used to get and set the slot value more conveniently. getter specifies a generic function to which GOOPS will add a method for getting the slot value. setter specifies a generic function to which GOOPS will add a method for setting the slot value. accessor specifies an accessor to which GOOPS will add methods for both getting and setting the slot value. So if a class includes a slot definition like this:

     (c #:getter get-count #:setter set-count #:accessor count)

GOOPS defines generic function methods such that the slot value can be referenced using either the getter or the accessor -

     (let ((current-count (get-count obj))) ...)
     (let ((current-count (count obj))) ...)

- and set using either the setter or the accessor -

     (set-count obj (+ 1 current-count))
     (set! (count obj) (+ 1 current-count))

Note that • with an accessor, the slot value is set using the generalized set! syntax

• in practice, it is unusual for a slot to use all three of these options: read-only, write-only and read-write slots would typically use only #:getter, #:setter and #:accessor options respectively.

The binding of the specified names is done in the environment of the define-class expression. If the names are already bound (in that environment) to values that cannot be upgraded to generic functions, those values are overwritten when the defineclass expression is evaluated. For more detail, see Section 8.11.9 [ensure-generic], page 753.

[slot option] #:allocation allocation

The #:allocation option tells GOOPS how to allocate storage for the slot. Possible values for allocation are

  • #:instance

Indicates that GOOPS should create separate storage for this slot in each new instance of the containing class (and its subclasses). This is the default.

  • #:class

Indicates that GOOPS should create storage for this slot that is shared by all instances of the containing class (and its subclasses). In other words, a slot in class C with allocation #:class is shared by all instances for which (isa? instance c). This permits defining a kind of global variable which can be accessed only by (in)direct instances of the class which defines the slot.

  • #:each-subclass

Indicates that GOOPS should create storage for this slot that is shared by all direct instances of the containing class, and that whenever a subclass of the containing class is defined, GOOPS should create a new storage for the slot that is shared by all direct instances of the subclass. In other words, a slot with allocation #:each-subclass is shared by all instances with the same class-of.

  • #:virtual

Indicates that GOOPS should not allocate storage for this slot. The slot definition must also include the #:slot-ref and #:slot-set! options to specify how to reference and set the value for this slot. See the example below.

Slot allocation options are processed when defining a new class by the generic function compute-get-n-set, which is specialized by the class’s metaclass. Hence new types of slot allocation can be implemented by defining a new metaclass and a method for compute-get-n-set that is specialized for the new metaclass. For an example of how to do this, see Section 8.11.6 [Customizing Class Definition], page 750.

[slot option] #:slot-ref getter [slot option] #:slot-set! setter

The #:slot-ref and #:slot-set! options must be specified if the slot allocation is #:virtual, and are ignored otherwise. getter should be a closure taking a single instance parameter that returns the current slot value. setter should be a closure taking two parameters - instance and new-val - that sets the slot value to new-val.

8.5 Illustrating Slot Description

To illustrate slot description, we can redefine the <my-complex> class seen before. A definition could be:

(define-class <my-complex> (<number>)
   (r #:init-value 0 #:getter get-r #:setter set-r! #:init-keyword #:r)
   (i #:init-value 0 #:getter get-i #:setter set-i! #:init-keyword #:i))

With this definition, the r and i slots are set to 0 by default, and can be initialised to other values by calling make with the #:r and #:i keywords. Also the generic functions get-r, set-r!, get-i and set-i! are automatically defined to read and write the slots.

(define c1 (make <my-complex> #:r 1 #:i 2))
(get-r c1) ⇒ 1
(set-r! c1 12)
(get-r c1) ⇒ 12
(define c2 (make <my-complex> #:r 2))
(get-r c2) ⇒ 2
(get-i c2) ⇒ 0

Accessors can both read and write a slot. So, another definition of the <my-complex> class, using the #:accessor option, could be:

     (define-class <my-complex> (<number>)
        (r #:init-value 0 #:accessor real-part #:init-keyword #:r)
        (i #:init-value 0 #:accessor imag-part #:init-keyword #:i))

With this definition, the r slot can be read with:

     (real-part c)

and set with:

     (set! (real-part c) new-value)

Suppose now that we want to manipulate complex numbers with both rectangular and polar coordinates. One solution could be to have a definition of complex numbers which uses one particular representation and some conversion functions to pass from one representation to the other. A better solution is to use virtual slots, like this:

     (define-class <my-complex> (<number>)
        ;; True slots use rectangular coordinates
        (r #:init-value 0 #:accessor real-part #:init-keyword #:r)
        (i #:init-value 0 #:accessor imag-part #:init-keyword #:i)
        ;; Virtual slots access do the conversion
        (m #:accessor magnitude #:init-keyword #:magn
           #:allocation #:virtual
           #:slot-ref (lambda (o)
                       (let ((r (slot-ref o 'r)) (i (slot-ref o 'i)))
                         (sqrt (+ (* r r) (* i i)))))
           #:slot-set! (lambda (o m)
                         (let ((a (slot-ref o 'a)))
                           (slot-set! o 'r (* m (cos a)))
                           (slot-set! o 'i (* m (sin a))))))
        (a #:accessor angle #:init-keyword #:angle
           #:allocation #:virtual
           #:slot-ref (lambda (o)
                       (atan (slot-ref o 'i) (slot-ref o 'r)))
           #:slot-set! (lambda(o a)
                        (let ((m (slot-ref o 'm)))
                           (slot-set! o 'r (* m (cos a)))
                           (slot-set! o 'i (* m (sin a)))))))

In this class definition, the magnitude m and angle a slots are virtual, and are calculated, when referenced, from the normal (i.e. #:allocation #:instance) slots r and i, by calling the function defined in the relevant #:slot-ref option. Correspondingly, writing m or a leads to calling the function defined in the #:slot-set! option. Thus the following expression

     (slot-set! c 'a 3)

permits to set the angle of the c complex number.

     (define c (make <my-complex> #:r 12 #:i 20))
     (real-part c) ⇒ 12
     (angle c) ⇒ 1.03037682652431
     (slot-set! c 'i 10)

     (set! (real-part c) 1)
     (describe c)

⊣

     #<<my-complex> 401e9b58> is an instance of class <my-complex>
     Slots are:
          r = 1
          i = 10
          m = 10.0498756211209
          a = 1.47112767430373

Since initialization keywords have been defined for the four slots, we can now define the standard Scheme primitives make-rectangular and make-polar.

     (define make-rectangular
        (lambda (x y) (make <my-complex> #:r x #:i y)))
     (define make-polar
        (lambda (x y) (make <my-complex> #:magn x #:angle y)))

8.6 Methods and Generic Functions

A GOOPS method is like a Scheme procedure except that it is specialized for a particular set of argument classes, and will only be used when the actual arguments in a call match the classes in the method definition.

     (define-method (+ (x <string>) (y <string>))
       (string-append x y))
     (+ "abc" "de") ⇒ "abcde"

A method is not formally associated with any single class (as it is in many other object oriented languages), because a method can be specialized for a combination of several classes. If you’ve studied object orientation in non-Lispy languages, you may remember discussions such as whether a method to stretch a graphical image around a surface should be a method of the image class, with a surface as a parameter, or a method of the surface class, with an image as a parameter. In GOOPS you’d just write

     (define-method (stretch (im <image>) (sf <surface>))
       ...)

and the question of which class the method is more associated with does not need answering.

There can simultaneously be several methods with the same name but different sets of specializing argument classes; for example:

     (define-method (+ (x <string>) (y <string)) ...)
     (define-method (+ (x <matrix>) (y <matrix>)) ...)
     (define-method (+ (f <fish>) (b <bicycle>)) ...)
     (define-method (+ (a <foo>) (b <bar>) (c <baz>)) ...)

A generic function is a container for the set of such methods that a program intends to use.

If you look at a program’s source code, and see (+ x y) somewhere in it, conceptually what is happening is that the program at that point calls a generic function (in this case, the generic function bound to the identifier +). When that happens, Guile works out which of the generic function’s methods is the most appropriate for the arguments that the function is being called with; then it evaluates the method’s code with the arguments as formal parameters. This happens every time that a generic function call is evaluated — it isn’t assumed that a given source code call will end up invoking the same method every time.

Defining an identifier as a generic function is done with the define-generic macro. Definition of a new method is done with the define-method macro. Note that definemethod automatically does a define-generic if the identifier concerned is not already a generic function, so often an explicit define-generic call is not needed.

[syntax] define-generic symbol

Create a generic function with name symbol and bind it to the variable symbol. If symbol was previously bound to a Scheme procedure (or procedure-with-setter), the old procedure (and setter) is incorporated into the new generic function as its default procedure (and setter). Any other previous value, including an existing generic function, is discarded and replaced by a new, empty generic function.

[syntax] define-method (generic parameter …) body … Define a method for the generic function or accessor generic with parameters parameters and body body . . .. generic is a generic function. If generic is a variable which is not yet bound to a generic function object, the expansion of define-method will include a call to definegeneric. If generic is (setter generic-with-setter), where generic-with-setter is a variable which is not yet bound to a generic-with-setter object, the expansion will

     include a call to define-accessor.

Each parameter must be either a symbol or a two-element list (symbol class). The symbols refer to variables in the body forms that will be bound to the parameters supplied by the caller when calling this method. The classes, if present, specify the possible combinations of parameters to which this method can be applied. body … are the bodies of the method definition.

define-method expressions look a little like Scheme procedure definitions of the form

     (define (name formals ...) . body)

The important difference is that each formal parameter, apart from the possible “rest” argument, can be qualified by a class name: formal becomes (formal class). The meaning of this qualification is that the method being defined will only be applicable in a particular generic function invocation if the corresponding argument is an instance of class (or one of its subclasses). If more than one of the formal parameters is qualified in this way, then the method will only be applicable if each of the corresponding arguments is an instance of its respective qualifying class.

Note that unqualified formal parameters act as though they are qualified by the class <top>, which GOOPS uses to mean the superclass of all valid Scheme types, including both primitive types and GOOPS classes.

For example, if a generic function method is defined with parameters (s1 <square>) and (n <number>), that method is only applicable to invocations of its generic function that have two parameters where the first parameter is an instance of the <square> class and the second parameter is a number.

8.6.1 Accessors

An accessor is a generic function that can also be used with the generalized set! syntax (see Section 6.9.8 [Procedures with Setters], page 255). Guile will handle a call like

     (set! (accessor args...) value)

by calling the most specialized method of accessor that matches the classes of args and value. define-accessor is used to bind an identifier to an accessor.

[syntax] define-accessor symbol

Create an accessor with name symbol and bind it to the variable symbol. If symbol was previously bound to a Scheme procedure (or procedure-with-setter), the old procedure (and setter) is incorporated into the new accessor as its default procedure (and setter). Any other previous value, including an existing generic function or accessor, is discarded and replaced by a new, empty accessor.

8.6.2 Extending Primitives

Many of Guile’s primitive procedures can be extended by giving them a generic function definition that operates in conjunction with their normal C-coded implementation. When a primitive is extended in this way, it behaves like a generic function with the C-coded implementation as its default method.

This extension happens automatically if a method is defined (by a define-method call) for a variable whose current value is a primitive. But it can also be forced by calling

enable-primitive-generic!.

[primitive procedure] enable-primitive-generic! primitive

Force the creation of a generic function definition for primitive.

Once the generic function definition for a primitive has been created, it can be retrieved

using primitive-generic-generic.

[primitive procedure] primitive-generic-generic primitive

Return the generic function definition of primitive.

primitive-generic-generic raises an error if primitive is not a primitive with generic capability.

8.6.3 Merging Generics

GOOPS generic functions and accessors often have short, generic names. For example, if a vector package provides an accessor for the X coordinate of a vector, that accessor may just be called x. It doesn’t need to be called, for example, vector:x, because GOOPS will work out, when it sees code like (x obj), that the vector-specific method of x should be called if obj is a vector.

That raises the question, though, of what happens when different packages define a generic function with the same name. Suppose we work with a graphical package which needs to use two independent vector packages for 2D and 3D vectors respectively. If both packages export x, what does the code using those packages end up with?

Section 6.19.3 [duplicate binding handlers], page 383 explains how this is resolved for conflicting bindings in general. For generics, there is a special duplicates handler, mergegenerics, which tells the module system to merge generic functions with the same name. Here is an example:

     (define-module (math 2D-vectors)
       #:use-module (oop goops)
       #:export (x y ...))
     (define-module (math 3D-vectors)
       #:use-module (oop goops)
       #:export (x y z ...))
     (define-module (my-module)
       #:use-module (oop goops)
       #:use-module (math 2D-vectors)
       #:use-module (math 3D-vectors)
       #:duplicates (merge-generics))

The generic function x in (my-module) will now incorporate all of the methods of x from both imported modules.

To be precise, there will now be three distinct generic functions named x: x in (math

2D-vectors), x in (math 3D-vectors), and x in (my-module); and these functions share

their methods in an interesting and dynamic way.

To explain, let’s call the imported generic functions (in (math 2D-vectors) and (math 3D-vectors)) the ancestors, and the merged generic function (in (my-module)), the descendant. The general rule is that for any generic function G, the applicable methods are selected from the union of the methods of G’s descendant functions, the methods of G itself and the methods of G’s ancestor functions.

Thus ancestor functions effectively share methods with their descendants, and vice versa. In the example above, x in (math 2D-vectors) will share the methods of x in (my-module) and vice versa.^{4} Sharing is dynamic, so adding another new method to a descendant implies adding it to that descendant’s ancestors too.

8.6.4 Next-method

When you call a generic function, with a particular set of arguments, GOOPS builds a list of all the methods that are applicable to those arguments and orders them by how closely the method definitions match the actual argument types. It then calls the method at the top of this list. If the selected method’s code wants to call on to the next method in this list, it can do so by using next-method.

     (define-method (Test (a <integer>)) (cons 'integer (next-method)))
     (define-method (Test (a <number>))  (cons 'number  (next-method)))
     (define-method (Test a)             (list 'top))

With these definitions,

     (Test 1) ⇒ (integer number top)
     (Test 1.0) ⇒ (number top)

4 But note that x in (math 2D-vectors) doesn’t share methods with x in (math 3D-vectors), so modu-

larity is still preserved.

(Test #t) ⇒ (top)

next-method is always called as just (next-method). The arguments for the next method call are always implicit, and always the same as for the original method call.

If you want to call on to a method with the same name but with a different set of arguments (as you might with overloaded methods in C++, for example), you do not use next-method, but instead simply write the new call as usual:

(define-method (Test (a <number>) min max)
  (if (and (>= a min) (<= a max))
      (display "Number is in range\n"))
  (Test a))
(Test 2 1 10)

⊣

Number is in range

⇒

(integer number top)

(You should be careful in this case that the Test calls do not lead to an infinite recursion, but this consideration is just the same as in Scheme code in general.)

8.6.5 Generic Function and Method Examples

Consider the following definitions:

(define-generic G)
(define-method (G (a <integer>) b) 'integer)
(define-method (G (a <real>) b) 'real)
(define-method (G a b) 'top)

The define-generic call defines G as a generic function. The three next lines define methods for G. Each method uses a sequence of parameter specializers that specify when the given method is applicable. A specializer permits to indicate the class a parameter must belong to (directly or indirectly) to be applicable. If no specializer is given, the system defaults it to <top>. Thus, the first method definition is equivalent to

(define-method (G (a <integer>) (b <top>)) 'integer)

Now, let’s look at some possible calls to the generic function G:

(G 2 3)  ⇒ integer
(G 2 #t) ⇒ integer
(G 1.2 'a) ⇒ real
(G #t #f) ⇒ top
(G 1 2 3) ⇒ error (since no method exists for 3 parameters)

The methods above use only one specializer per parameter list. But in general, any or all of a method’s parameters may be specialized. Suppose we define now:

(define-method (G (a <integer>) (b <number>))  'integer-number)
(define-method (G (a <integer>) (b <real>))    'integer-real)
(define-method (G (a <integer>) (b <integer>)) 'integer-integer)
(define-method (G a (b <number>))              'top-number)

With these definitions:

(G 1 2) ⇒ integer-integer
(G 1 1.0) ⇒ integer-real
(G 1 #t) ⇒ integer
(G 'a 1) ⇒ top-number

As a further example we shall continue to define operations on the <my-complex> class. Suppose that we want to use it to implement complex numbers completely. For instance a definition for the addition of two complex numbers could be

(define-method (new-+ (a <my-complex>) (b <my-complex>))
  (make-rectangular (+ (real-part a) (real-part b))
                    (+ (imag-part a) (imag-part b))))

To be sure that the + used in the method new-+ is the standard addition we can do:

(define-generic new-+)
(let ((+ +))
  (define-method (new-+ (a <my-complex>) (b <my-complex>))
    (make-rectangular (+ (real-part a) (real-part b))
                      (+ (imag-part a) (imag-part b)))))

The define-generic ensures here that new-+ will be defined in the global environment. Once this is done, we can add methods to the generic function new-+ which make a closure on the + symbol. A complete writing of the new-+ methods is shown in Figure 8.1.

     (define-generic new-+)
     (let ((+ +))
       (define-method (new-+ (a <real>) (b <real>)) (+ a b))
       (define-method (new-+ (a <real>) (b <my-complex>))
         (make-rectangular (+ a (real-part b)) (imag-part b)))
       (define-method (new-+ (a <my-complex>) (b <real>))
         (make-rectangular (+ (real-part a) b) (imag-part a)))
       (define-method (new-+ (a <my-complex>) (b <my-complex>))
         (make-rectangular (+ (real-part a) (real-part b))
                           (+ (imag-part a) (imag-part b))))
       (define-method (new-+ (a <number>))  a)
       (define-method (new-+) 0)
       (define-method (new-+ . args)
         (new-+ (car args)
           (apply new-+ (cdr args)))))
     (set! + new-+)

Figure 8.1: Extending + to handle complex numbers

We take advantage here of the fact that generic function are not obliged to have a fixed number of parameters. The four first methods implement dyadic addition. The fifth method says that the addition of a single element is this element itself. The sixth method says that using the addition with no parameter always return 0 (as is also true for the primitive +). The last method takes an arbitrary number of parameters^{5}. This method acts as a kind of reduce: it calls the dyadic addition on the car of the list and on the result of applying it on its rest. To finish, the set! permits to redefine the + symbol to our extended addition.

To conclude our implementation (integration?) of complex numbers, we could redefine standard Scheme predicates in the following manner:

     (define-method (complex? c <my-complex>) #t)
     (define-method (complex? c)           #f)
     (define-method (number? n <number>) #t)
     (define-method (number? n)          #f)
     ...

5 The parameter list for a define-method follows the conventions used for Scheme procedures. In particular

it can use the dot notation or a symbol to denote an arbitrary number of parameters

Standard primitives in which complex numbers are involved could also be redefined in the same manner.

8.6.6 Handling Invocation Errors

If a generic function is invoked with a combination of parameters for which there is no applicable method, GOOPS raises an error.

[generic] no-method

[method] no-method (gf <generic>) args When an application invokes a generic function, and no methods at all have been defined for that generic function, GOOPS calls the no-method generic function. The default method calls goops-error with an appropriate message.

[generic] no-applicable-method

[method] no-applicable-method (gf <generic>) args

When an application applies a generic function to a set of arguments, and no methods have been defined for those argument types, GOOPS calls the no-applicable-method generic function. The default method calls goops-error with an appropriate message.

[generic] no-next-method

[method] no-next-method (gf <generic>) args

When a generic function method calls (next-method) to invoke the next less specialized method for that generic function, and no less specialized methods have been defined for the current generic function arguments, GOOPS calls the no-next-method generic function. The default method calls goops-error with an appropriate message.

8.7 Inheritance

Here are some class definitions to help illustrate inheritance:

(define-class A () a)
(define-class B () b)
(define-class C () c)
(define-class D (A B) d a)
(define-class E (A C) e c)
(define-class F (D E) f)

A, B, C have a null list of superclasses. In this case, the system will replace the null list by a list which only contains <object>, the root of all the classes defined by define-class. D, E, F use multiple inheritance: each class inherits from two previously defined classes. Those class definitions define a hierarchy which is shown in Figure 8.2. In this figure, the class <top> is also shown; this class is the superclass of all Scheme objects. In particular, <top> is the superclass of all standard Scheme types.

Figure 8.2: A class hierarchy.

When a class has superclasses, its set of slots is calculated by taking the union of its own slots and those of all its superclasses. Thus each instance of D will have three slots, a, b and d). The slots of a class can be discovered using the class-slots primitive. For instance,

     (class-slots A) ⇒ ((a))
     (class-slots E) ⇒ ((a) (e) (c))
     (class-slots F) ⇒ ((e) (c) (b) (d) (a) (f))

The ordering of the returned slots is not significant.

8.7.1 Class Precedence List

What happens when a class inherits from two or more superclasses that have a slot with the same name but incompatible definitions — for example, different init values or slot allocations? We need a rule for deciding which slot definition the derived class ends up with, and this rule is provided by the class’s Class Precedence List.^{6}

Another problem arises when invoking a generic function, and there is more than one method that could apply to the call arguments. Here we need a way of ordering the applicable methods, so that Guile knows which method to use first, which to use next if that method calls next-method, and so on. One of the ingredients for this ordering is determining, for each given call argument, which of the specializing classes, from each applicable method’s definition, is the most specific for that argument; and here again the class precedence list helps.

6 This section is an adaptation of material from Jeff Dalton’s (J.Dalton@ed.ac.uk) Brief introduction to

CLOS

If inheritance was restricted such that each class could only have one superclass — which is known as single inheritance — class ordering would be easy. The rule would be simply that a subclass is considered more specific than its superclass.

With multiple inheritance, ordering is less obvious, and we have to impose an arbitrary rule to determine precedence. Suppose we have

     (define-class X ()
        (x #:init-value 1))
     (define-class Y ()
        (x #:init-value 2))
     (define-class Z (X Y)
        (...))

Clearly the Z class is more specific than X or Y, for instances of Z. But which is more specific out of X and Y — and hence, for the definitions above, which #:init-value will take effect when creating an instance of Z? The rule in GOOPS is that the superclasses listed earlier are more specific than those listed later. Hence X is more specific than Y, and the #:init-value for slot x in instances of Z will be 1.

Hence there is a linear ordering for a class and all its superclasses, from most specific to least specific, and this ordering is called the Class Precedence List of the class.

In fact the rules above are not quite enough to always determine a unique order, but they give an idea of how things work. For example, for the F class shown in Figure 8.2, the class precedence list is

     (f d e a c b <object> <top>)

In cases where there is any ambiguity (like this one), it is a bad idea for programmers to rely on exactly what the order is. If the order for some superclasses is important, it can be expressed directly in the class definition.

The precedence list of a class can be obtained by calling class-precedence-list. This function returns a ordered list whose first element is the most specific class. For instance:

     (class-precedence-list B) ⇒ (#<<class> B 401b97c8>
                                          #<<class> <object> 401e4a10>
                                          #<<class> <top> 4026a9d8>)

Or for a more immediately readable result:

     (map class-name (class-precedence-list B)) ⇒ (B <object> <top>)

8.7.2 Sorting Methods

Now, with the idea of the class precedence list, we can state precisely how the possible methods are sorted when more than one of the methods of a generic function are applicable to the call arguments.

The rules are that • the applicable methods are sorted in order of specificity, and the most specific method

is used first, then the next if that method calls next-method, and so on

• a method M1 is more specific than another method M2 if the first specializing class that differs, between the definitions of M1 and M2, is more specific, in M1’s definition, for the corresponding actual call argument, than the specializing class in M2’s definition

• a class C1 is more specific than another class C2, for an object of actual class C, if C1 comes before C2 in C’s class precedence list.

8.8 Introspection

Introspection, or reflection, means being able to obtain information dynamically about GOOPS objects. It is perhaps best illustrated by considering an object oriented language that does not provide any introspection, namely C++.

Nothing in C++ allows a running program to obtain answers to the following types of question:

• What are the data members of this object or class?

• What classes does this class inherit from?

• Is this method call virtual or non-virtual?

  • If I invoke Employee::adjustHoliday(), what class contains the adjustHoliday()

method that will be applied?

In C++, answers to such questions can only be determined by looking at the source code, if you have access to it. GOOPS, on the other hand, includes procedures that allow answers to these questions — or their GOOPS equivalents — to be obtained dynamically, at run time.

8.8.1 Classes

A GOOPS class is itself an instance of the <class> class, or of a subclass of <class>. The definition of the <class> class has slots that are used to describe the properties of a class, including the following.

[primitive procedure] class-name class

Return the name of class class. This is the value of class’s name slot.

[primitive procedure] class-direct-supers class

Return a list containing the direct superclasses of class. This is the value of class’s

     direct-supers slot.

[primitive procedure] class-direct-slots class

Return a list containing the slot definitions of the direct slots of class. This is the value of class’s direct-slots slot.

[primitive procedure] class-direct-subclasses class

Return a list containing the direct subclasses of class. This is the value of class’s

     direct-subclasses slot.

[primitive procedure] class-direct-methods class

Return a list of all the generic function methods that use class as a formal parameter specializer. This is the value of class’s direct-methods slot.

[primitive procedure] class-precedence-list class

Return the class precedence list for class class (see Section 8.7.1 [Class Precedence List], page 734). This is the value of class’s cpl slot.

[primitive procedure] class-slots class

Return a list containing the slot definitions for all class’s slots, including any slots that are inherited from superclasses. This is the value of class’s slots slot.

[procedure] class-subclasses class

Return a list of all subclasses of class.

[procedure] class-methods class

Return a list of all methods that use class or a subclass of class as one of its formal parameter specializers.

8.8.2 Instances

[primitive procedure] class-of value

Return the GOOPS class of any Scheme value.

[primitive procedure] instance? object

Return #t if object is any GOOPS instance, otherwise #f.

[procedure] is-a? object class Return #t if object is an instance of class or one of its subclasses.

You can use the is-a? predicate to ask whether any given value belongs to a given class, or class-of to discover the class of a given value. Note that when GOOPS is loaded (by code using the (oop goops) module) built-in classes like <string>, <list> and <number> are automatically set up, corresponding to all Guile Scheme types.

(is-a? 2.3 <number>) ⇒ #t
(is-a? 2.3 <real>) ⇒ #t
(is-a? 2.3 <string>) ⇒ #f
(is-a? '("a" "b") <string>) ⇒ #f
(is-a? '("a" "b") <list>) ⇒ #t
(is-a? (car '("a" "b")) <string>) ⇒ #t
(is-a? <string> <class>) ⇒ #t
(is-a? <class> <string>) ⇒ #f
(class-of 2.3) ⇒ #<<class> <real> 908c708>
(class-of #(1 2 3)) ⇒ #<<class> <vector> 908cd20>
(class-of <string>) ⇒ #<<class> <class> 8bd3e10>
(class-of <class>) ⇒ #<<class> <class> 8bd3e10>

8.8.3 Slots

[procedure] class-slot-definition class slot-name

Return the slot definition for the slot named slot-name in class class. slot-name should be a symbol.

[procedure] slot-definition-name slot-def

Extract and return the slot name from slot-def.

[procedure] slot-definition-options slot-def

Extract and return the slot options from slot-def.

[procedure] slot-definition-allocation slot-def

Extract and return the slot allocation option from slot-def. This is the value of the #:allocation keyword (see Section 8.4 [allocation], page 721), or #:instance if the #:allocation keyword is absent.

[procedure] slot-definition-getter slot-def

Extract and return the slot getter option from slot-def. This is the value of the #:getter keyword (see Section 8.4 [getter], page 721), or #f if the #:getter keyword is absent.

[procedure] slot-definition-setter slot-def

Extract and return the slot setter option from slot-def. This is the value of the #:setter keyword (see Section 8.4 [setter], page 721), or #f if the #:setter keyword is absent.

[procedure] slot-definition-accessor slot-def

Extract and return the slot accessor option from slot-def. This is the value of the #:accessor keyword (see Section 8.4 [accessor], page 721), or #f if the #:accessor keyword is absent.

[procedure] slot-definition-init-value slot-def

Extract and return the slot init-value option from slot-def. This is the value of the #:init-value keyword (see Section 8.4 [init-value], page 721), or the unbound value if the #:init-value keyword is absent.

[procedure] slot-definition-init-form slot-def

Extract and return the slot init-form option from slot-def. This is the value of the #:init-form keyword (see Section 8.4 [init-form], page 721), or the unbound value if the #:init-form keyword is absent.

[procedure] slot-definition-init-thunk slot-def

Extract and return the slot init-thunk option from slot-def. This is the value of the #:init-thunk keyword (see Section 8.4 [init-thunk], page 721), or #f if the #:initthunk keyword is absent.

[procedure] slot-definition-init-keyword slot-def

Extract and return the slot init-keyword option from slot-def. This is the value of the #:init-keyword keyword (see Section 8.4 [init-keyword], page 721), or #f if the #:init-keyword keyword is absent.

[procedure] slot-init-function class slot-name

Return the initialization function for the slot named slot-name in class class. slotname should be a symbol.

The returned initialization function incorporates the effects of the standard #:init-thunk, #:init-form and #:init-value slot options. These initializations can be overridden by the #:init-keyword slot option or by a specialized initialize method, so, in general, the function returned by slot-init-function may be irrelevant. For a fuller discussion, see Section 8.4 [init-value], page 721.

8.8.4 Generic Functions

A generic function is an instance of the <generic> class, or of a subclass of <generic>. The definition of the <generic> class has slots that are used to describe the properties of a generic function.

[primitive procedure] generic-function-name gf

Return the name of generic function gf.

[primitive procedure] generic-function-methods gf

Return a list of the methods of generic function gf. This is the value of gf ’s methods slot.

Similarly, a method is an instance of the <method> class, or of a subclass of <method>; and the definition of the <method> class has slots that are used to describe the properties of a method.

[primitive procedure] method-generic-function method

Return the generic function that method belongs to. This is the value of method’s

generic-function slot.

[primitive procedure] method-specializers method

Return a list of method’s formal parameter specializers . This is the value of method’s

specializers slot.

[primitive procedure] method-procedure method

Return the procedure that implements method. This is the value of method’s

procedure slot.
[generic] method-source

[method] method-source (m <method>)

Return an expression that prints to show the definition of method m.

     (define-generic cube)
     (define-method (cube (n <number>))
       (* n n n))
     (map method-source (generic-function-methods cube))

⇒

     ((method ((n <number>)) (* n n n)))

8.8.5 Accessing Slots

Any slot, regardless of its allocation, can be queried, referenced and set using the following four primitive procedures.

[primitive procedure] slot-exists? obj slot-name

Return #t if obj has a slot with name slot-name, otherwise #f.

[primitive procedure] slot-bound? obj slot-name Return #t if the slot named slot-name in obj has a value, otherwise #f.

slot-bound? calls the generic function slot-missing if obj does not have a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740).

[primitive procedure] slot-ref obj slot-name Return the value of the slot named slot-name in obj.

slot-ref calls the generic function slot-missing if obj does not have a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740).

slot-ref calls the generic function slot-unbound if the named slot in obj does not have a value (see Section 8.8.5 [Accessing Slots], page 740).

[primitive procedure] slot-set! obj slot-name value Set the value of the slot named slot-name in obj to value.

slot-set! calls the generic function slot-missing if obj does not have a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740).

GOOPS stores information about slots in classes. Internally, all of these procedures work by looking up the slot definition for the slot named slot-name in the class (class-of obj), and then using the slot definition’s “getter” and “setter” closures to get and set the slot value.

The next four procedures differ from the previous ones in that they take the class as an explicit argument, rather than assuming (class-of obj). Therefore they allow you to apply the “getter” and “setter” closures of a slot definition in one class to an instance of a different class.

[primitive procedure] slot-exists-using-class? class obj slot-name

Return #t if class has a slot definition for a slot with name slot-name, otherwise #f.

[primitive procedure] slot-bound-using-class? class obj slot-name

Return #t if applying slot-ref-using-class to the same arguments would call the generic function slot-unbound, otherwise #f.

slot-bound-using-class? calls the generic function slot-missing if class does not have a slot definition for a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740).

[primitive procedure] slot-ref-using-class class obj slot-name

Apply the “getter” closure for the slot named slot-name in class to obj, and return its result. slot-ref-using-class calls the generic function slot-missing if class does not have a slot definition for a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740). slot-ref-using-class calls the generic function slot-unbound if the application of the “getter” closure to obj returns an unbound value (see Section 8.8.5 [Accessing Slots], page 740).

[primitive procedure] slot-set-using-class! class obj slot-name value Apply the “setter” closure for the slot named slot-name in class to obj and value. slot-set-using-class! calls the generic function slot-missing if class does not have a slot definition for a slot called slot-name (see Section 8.8.5 [Accessing Slots], page 740).

Slots whose allocation is per-class rather than per-instance can be referenced and set without needing to specify any particular instance.

[procedure] class-slot-ref class slot-name

Return the value of the slot named slot-name in class class. The named slot must have #:class or #:each-subclass allocation (see Section 8.4 [allocation], page 721). If there is no such slot with #:class or #:each-subclass allocation, class-slot-ref calls the slot-missing generic function with arguments class and slot-name. Otherwise, if the slot value is unbound, class-slot-ref calls the slot-unbound generic function, with the same arguments.

[procedure] class-slot-set! class slot-name value Set the value of the slot named slot-name in class class to value. The named slot must have #:class or #:each-subclass allocation (see Section 8.4 [allocation], page 721).

If there is no such slot with #:class or #:each-subclass allocation, class-slot-ref

calls the slot-missing generic function with arguments class and slot-name.

When a slot-ref or slot-set! call specifies a non-existent slot name, or tries to reference a slot whose value is unbound, GOOPS calls one of the following generic functions.

[generic] slot-missing

[method] slot-missing (class <class>) slot-name

[method] slot-missing (class <class>) (object <object>) slot-name

[method] slot-missing (class <class>) (object <object>) slot-name value When an application attempts to reference or set a class or instance slot by name, and the slot name is invalid for the specified class or object, GOOPS calls the slotmissing generic function. The default methods all call goops-error with an appropriate message.

[generic] slot-unbound

[method] slot-unbound (object <object>)

[method] slot-unbound (class <class>) slot-name

[method] slot-unbound (class <class>) (object <object>) slot-name When an application attempts to reference a class or instance slot, and the slot’s value is unbound, GOOPS calls the slot-unbound generic function. The default methods all call goops-error with an appropriate message.

8.9 Error Handling

The procedure goops-error is called to raise an appropriate error by the default methods of the following generic functions:

• slot-missing (see Section 8.8.5 [slot-missing], page 740)

• slot-unbound (see Section 8.8.5 [slot-unbound], page 740)

• no-method (see Section 8.6.6 [no-method], page 733)

• no-applicable-method (see Section 8.6.6 [no-applicable-method], page 733)

• no-next-method (see Section 8.6.6 [no-next-method], page 733)

If you customize these functions for particular classes or metaclasses, you may still want to use goops-error to signal any error conditions that you detect.

[procedure] goops-error format-string arg … Raise an error with key goops-error and error message constructed from formatstring and arg . . .. Error message formatting is as done by scm-error.

8.10 GOOPS Object Miscellany

Here we cover some points about GOOPS objects that aren’t substantial enough to merit sections on their own.

Object Equality

When GOOPS is loaded, eqv?, equal? and = become generic functions, and you can define methods for them, specialized for your own classes, so as to control what the various kinds of equality mean for your classes.

For example, the assoc procedure, for looking up an entry in an alist, is specified as using equal? to determine when the car of an entry in the alist is the same as the key parameter that assoc is called with. Hence, if you had defined a new class, and wanted to use instances of that class as the keys in an alist, you could define a method for equal?, for your class, to control assoc’s lookup precisely.

Cloning Objects

[generic] shallow-clone

[method] shallow-clone (self <object>)

Return a “shallow” clone of self. The default method makes a shallow clone by allocating a new instance and copying slot values from self to the new instance. Each slot value is copied either as an immediate value or by reference.

[generic] deep-clone

[method] deep-clone (self <object>)

Return a “deep” clone of self. The default method makes a deep clone by allocating a new instance and copying or cloning slot values from self to the new instance. If a slot value is an instance (satisfies instance?), it is cloned by calling deep-clone on that value. Other slot values are copied either as immediate values or by reference.

Write and Display

[primitive generic] write object port [primitive generic] display object port

When GOOPS is loaded, write and display become generic functions with special methods for printing

• objects - instances of the class <object>

• foreign objects - instances of the class <foreign-object>

• classes - instances of the class <class>

• generic functions - instances of the class <generic>

• methods - instances of the class <method>.

write and display print non-GOOPS values in the same way as the Guile primitive write and display functions.

In addition to the cases mentioned, you can of course define write and display methods for your own classes, to customize how instances of those classes are printed.

8.11 The Metaobject Protocol

At this point, we’ve said about as much as can be said about GOOPS without having to confront the idea of the metaobject protocol. There are a couple more topics that could be discussed in isolation first — class redefinition, and changing the class of existing instances — but in practice developers using them will be advanced enough to want to understand the metaobject protocol too, and will probably be using the protocol to customize exactly what happens during these events.

So let’s plunge in. GOOPS is based on a “metaobject protocol” (aka “MOP”) derived from the ones used in CLOS (the Common Lisp Object System), tiny-clos (a small Scheme implementation of a subset of CLOS functionality) and STKlos.

The MOP underlies many possible GOOPS customizations — such as defining an initialize method to customize the initialization of instances of an application-defined class — and an understanding of the MOP makes it much easier to explain such customizations in a precise way. And at a deeper level, understanding the MOP is a key part of understanding GOOPS, and of taking full advantage of GOOPS’ power, by customizing the behaviour of GOOPS itself.

8.11.1 Metaobjects and the Metaobject Protocol

The building blocks of GOOPS are classes, slot definitions, instances, generic functions and methods. A class is a grouping of inheritance relations and slot definitions. An instance is an object with slots that are allocated following the rules implied by its class’s superclasses and slot definitions. A generic function is a collection of methods and rules for determining which of those methods to apply when the generic function is invoked. A method is a procedure and a set of specializers that specify the type of arguments to which the procedure is applicable.

Of these entities, GOOPS represents classes, generic functions and methods as “metaobjects”. In other words, the values in a GOOPS program that describe classes, generic functions and methods, are themselves instances (or “objects”) of special GOOPS classes that encapsulate the behaviour, respectively, of classes, generic functions, and methods.

(The other two entities are slot definitions and instances. Slot definitions are not strictly instances, but every slot definition is associated with a GOOPS class that specifies the behaviour of the slot as regards accessibility and protection from garbage collection. Instances are of course objects in the usual sense, and there is no benefit from thinking of them as metaobjects.)

The “metaobject protocol” (or “MOP”) is the specification of the generic functions which determine the behaviour of these metaobjects and the circumstances in which these generic functions are invoked.

For a concrete example of what this means, consider how GOOPS calculates the set of slots for a class that is being defined using define-class. The desired set of slots is the union of the new class’s direct slots and the slots of all its superclasses. But define-class itself does not perform this calculation. Instead, there is a method of the initialize generic function that is specialized for instances of type <class>, and it is this method that performs the slot calculation.

initialize is a generic function which GOOPS calls whenever a new instance is created, immediately after allocating memory for a new instance, in order to initialize the new instance’s slots. The sequence of steps is as follows.

• define-class uses make to make a new instance of the <class> class, passing as initialization arguments the superclasses, slot definitions and class options that were specified in the define-class form.

• make allocates memory for the new instance, and invokes the initialize generic function to initialize the new instance’s slots.

• The initialize generic function applies the method that is specialized for instances of type <class>, and this method performs the slot calculation.

In other words, rather than being hardcoded in define-class, the default behaviour of class definition is encapsulated by generic function methods that are specialized for the

class <class>.

It is possible to create a new class that inherits from <class>, which is called a “metaclass”, and to write a new initialize method that is specialized for instances of the new metaclass. Then, if the define-class form includes a #:metaclass class option whose value is the new metaclass, the class that is defined by the define-class form will be an instance of the new metaclass rather than of the default <class>, and will be defined in accordance with the new initialize method. Thus the default slot calculation, as well as any other aspect of the new class’s relationship with its superclasses, can be modified or overridden.

In a similar way, the behaviour of generic functions can be modified or overridden by creating a new class that inherits from the standard generic function class <generic>, writing appropriate methods that are specialized to the new class, and creating new generic functions that are instances of the new class.

The same is true for method metaobjects. And the same basic mechanism allows the application class author to write an initialize method that is specialized to their application class, to initialize instances of that class.

Such is the power of the MOP. Note that initialize is just one of a large number of generic functions that can be customized to modify the behaviour of application objects and classes and of GOOPS itself. Each following section covers a particular area of GOOPS functionality, and describes the generic functions that are relevant for customization of that area.

8.11.2 Metaclasses

A metaclass is the class of an object which represents a GOOPS class. Put more succinctly, a metaclass is a class’s class.

Most GOOPS classes have the metaclass <class> and, by default, any new class that is created using define-class has the metaclass <class>.

But what does this really mean? To find out, let’s look in more detail at what happens when a new class is created using define-class:

     (define-class <my-class> (<object>) . slots)

Guile expands this to something like:

     (define <my-class> (class (<object>) . slots))

which in turn expands to:

     (define <my-class>
       (make <class> #:dsupers (list <object>) #:slots slots))

As this expansion makes clear, the resulting value of <my-class> is an instance of the class <class> with slot values specifying the superclasses and slot definitions for the class <my-class>. (#:dsupers and #:slots are initialization keywords for the dsupers and dslots slots of the <class> class.)

Now suppose that you want to define a new class with a metaclass other than the default <class>. This is done by writing:

     (define-class <my-class2> (<object>)
        slot ...
        #:metaclass <my-metaclass>)

and Guile expands this to something like:

     (define <my-class2>
       (make <my-metaclass> #:dsupers (list <object>) #:slots slots))

In this case, the value of <my-class2> is an instance of the more specialized class <mymetaclass>. Note that <my-metaclass> itself must previously have been defined as a subclass of <class>. For a full discussion of when and how it is useful to define new metaclasses, see Section 8.11.3 [MOP Specification], page 746.

Now let’s make an instance of <my-class2>:

     (define my-object (make <my-class2> ...))

All of the following statements are correct expressions of the relationships between my-

object, <my-class2>, <my-metaclass> and <class>.

• my-object is an instance of the class <my-class2>.

  • <my-class2> is an instance of the class <my-metaclass>.

• <my-metaclass> is an instance of the class <class>.

  • The class of my-object is <my-class2>.
  • The class of <my-class2> is <my-metaclass>.
  • The class of <my-metaclass> is <class>.

8.11.3 MOP Specification

The aim of the MOP specification in this chapter is to specify all the customizable generic function invocations that can be made by the standard GOOPS syntax, procedures and methods, and to explain the protocol for customizing such invocations.

A generic function invocation is customizable if the types of the arguments to which it is applied are not completely determined by the lexical context in which the invocation appears. For example, the (initialize instance initargs) invocation in the default make-instance method is customizable, because the type of the instance argument is determined by the class that was passed to make-instance.

(Whereas — to give a counter-example — the (make <generic> #:name ',name) invocation in define-generic is not customizable, because all of its arguments have lexically determined types.)

When using this rule to decide whether a given generic function invocation is customizable, we ignore arguments that are expected to be handled in method definitions as a single “rest” list argument.

For each customizable generic function invocation, the invocation protocol is explained by specifying

• what, conceptually, the applied method is intended to do

• what assumptions, if any, the caller makes about the applied method’s side effects • what the caller expects to get as the applied method’s return value.

8.11.4 Instance Creation Protocol

make <class> . initargs (method)
  • allocate-instance class initargs (generic)

The applied allocate-instance method should allocate storage for a new instance of class class and return the uninitialized instance.

  • initialize instance initargs (generic)

instance is the uninitialized instance returned by allocate-instance. The applied method should initialize the new instance in whatever sense is appropriate for its class. The method’s return value is ignored.

make itself is a generic function. Hence the make invocation itself can be customized in the case where the new instance’s metaclass is more specialized than the default <class>, by defining a make method that is specialized to that metaclass.

Normally, however, the method for classes with metaclass <class> will be applied. This method calls two generic functions:

• (allocate-instance class . initargs)

• (initialize instance . initargs)

allocate-instance allocates storage for and returns the new instance, uninitialized. You might customize allocate-instance, for example, if you wanted to provide a GOOPS wrapper around some other object programming system.

To do this, you would create a specialized metaclass, which would act as the metaclass for all classes and instances from the other system. Then define an allocate-instance method, specialized to that metaclass, which calls a Guile primitive C function (or FFI code), which in turn allocates the new instance using the interface of the other object system.

In this case, for a complete system, you would also need to customize a number of other generic functions like make and initialize, so that GOOPS knows how to make classes from the other system, access instance slots, and so on.

initialize initializes the instance that is returned by allocate-instance. The standard GOOPS methods perform initializations appropriate to the instance class.

• At the least specialized level, the method for instances of type <object> performs internal GOOPS instance initialization, and initializes the instance’s slots according to the slot definitions and any slot initialization keywords that appear in initargs.

• The method for instances of type <class> calls (next-method), then performs the class initializations described in Section 8.11.5 [Class Definition Protocol], page 747.

• and so on for generic functions, methods, operator classes …

Similarly, you can customize the initialization of instances of any application-defined class by defining an initialize method specialized to that class.

Imagine a class whose instances’ slots need to be initialized at instance creation time by querying a database. Although it might be possible to achieve this a combination of #:init-thunk keywords and closures in the slot definitions, it may be neater to write an initialize method for the class that queries the database once and initializes all the dependent slot values according to the results.

8.11.5 Class Definition Protocol

Here is a summary diagram of the syntax, procedures and generic functions that may be involved in class definition.

define-class (syntax)

• class (syntax)

• make-class (procedure)

         • ensure-metaclass (procedure)
         • make metaclass ... (generic)
             • allocate-instance (generic)
             • initialize (generic)
                 • compute-cpl (generic)
                    • compute-std-cpl (procedure)
                 • compute-slots (generic)
                 • compute-get-n-set (generic)
                 • compute-getter-method (generic)

               • compute-setter-method (generic)
• class-redefinition (generic)
    • remove-class-accessors (generic)
    • update-direct-method! (generic)
    • update-direct-subclass! (generic)

Wherever a step above is marked as “generic”, it can be customized, and the detail shown below it is only “correct” insofar as it describes what the default method of that generic function does. For example, if you write an initialize method, for some metaclass, that does not call next-method and does not call compute-cpl, then compute-cpl will not be called when a class is defined with that metaclass.

A (define-class ...) form (see Section 8.2 [Class Definition], page 719) expands to an expression which

• checks that it is being evaluated only at top level

• defines any accessors that are implied by the slot-definitions

• uses class to create the new class

• checks for a previous class definition for name and, if found, handles the redefinition by invoking class-redefinition (see Section 8.12 [Redefining a Class], page 755).

[syntax] class name (super …) slot-definition … class-option … Return a newly created class that inherits from supers, with direct slots defined by slot-definitions and class-options. For the format of slot-definitions and class-options, see Section 8.2 [define-class], page 719.

class expands to an expression which

• processes the class and slot definition options to check that they are well-formed, to convert the #:init-form option to an #:init-thunk option, to supply a default environment parameter (the current top-level environment) and to evaluate all the bits that need to be evaluated

• calls make-class to create the class with the processed and evaluated parameters.

[procedure] make-class supers slots class-option … Return a newly created class that inherits from supers, with direct slots defined by slots and class-options. For the format of slots and class-options, see Section 8.2 [define-class], page 719, except note that for make-class, slots is a separate list of slot definitions.

make-class

• adds <object> to the supers list if supers is empty or if none of the classes in supers have <object> in their class precedence list

• defaults the #:environment, #:name and #:metaclass options, if they are not specified by options, to the current top-level environment, the unbound value, and (ensure-

  metaclass supers) respectively

• checks for duplicate classes in supers and duplicate slot names in slots, and signals an error if there are any duplicates

• calls make, passing the metaclass as the first parameter and all other parameters as option keywords with values.

[procedure] ensure-metaclass supers env

Return a metaclass suitable for a class that inherits from the list of classes in supers. The returned metaclass is the union by inheritance of the metaclasses of the classes in supers. In the simplest case, where all the supers are straightforward classes with metaclass <class>, the returned metaclass is just <class>. For a more complex example, suppose that supers contained one class with metaclass

<operator-class> and one with metaclass <foreign-object-class>. Then the

returned metaclass would be a class that inherits from both <operator-class> and

<foreign-object-class>.

If supers is the empty list, ensure-metaclass returns the default GOOPS metaclass

<class>.

GOOPS keeps a list of the metaclasses created by ensure-metaclass, so that each required type of metaclass only has to be created once. The env parameter is ignored.

[generic] make metaclass initarg … metaclass is the metaclass of the class being defined, either taken from the #:metaclass class option or computed by ensure-metaclass. The applied method must create and return the fully initialized class metaobject for the new class definition.

The (make metaclass initarg ...) invocation is a particular case of the instance creation protocol covered in the previous section. It will create an class metaobject with metaclass metaclass. By default, this metaobject will be initialized by the initialize method that is specialized for instances of type <class>.

The initialize method for classes (signature (initialize <class> initargs)) calls

the following generic functions.

• compute-cpl class (generic)

The applied method should compute and return the class precedence list for class as a list of class metaobjects. When compute-cpl is called, the following class metaob-

ject slots have all been initialized: name, direct-supers, direct-slots, direct-
subclasses (empty), direct-methods. The value returned by compute-cpl will be

stored in the cpl slot.

• compute-slots class (generic)

The applied method should compute and return the slots (union of direct and inherited) for class as a list of slot definitions. When compute-slots is called, all the class metaobject slots mentioned for compute-cpl have been initialized, plus the following:

cpl, redefined (#f), environment. The value returned by compute-slots will be

stored in the slots slot.

• compute-get-n-set class slot-def (generic)
initialize calls compute-get-n-set for each slot computed by compute-slots. The

applied method should compute and return a pair of closures that, respectively, get and set the value of the specified slot. The get closure should have arity 1 and expect a single argument that is the instance whose slot value is to be retrieved. The set closure should have arity 2 and expect two arguments, where the first argument is the instance whose slot value is to be set and the second argument is the new value for that slot. The closures should be returned in a two element list: (list get set). The closures returned by compute-get-n-set are stored as part of the value of the class metaobject’s getters-n-setters slot. Specifically, the value of this slot is a list with the same number of elements as there are slots in the class, and each element looks either like

     (slot-name-symbol init-function . index)

or like

     (slot-name-symbol init-function get set)

Where the get and set closures are replaced by index, the slot is an instance slot and index is the slot’s index in the underlying structure: GOOPS knows how to get and set the value of such slots and so does not need specially constructed get and set closures. Otherwise, get and set are the closures returned by compute-get-n-set. The structure of the getters-n-setters slot value is important when understanding the next customizable generic functions that initialize calls…

• compute-getter-method class gns (generic)

initialize calls compute-getter-method for each of the class’s slots (as determined by compute-slots) that includes a #:getter or #:accessor slot option. gns is the element of the class metaobject’s getters-n-setters slot that specifies how the slot in question is referenced and set, as described above under compute-get-n-set. The applied method should create and return a method that is specialized for instances of type class and uses the get closure to retrieve the slot’s value. initialize uses add-method! to add the returned method to the generic function named by the slot definition’s #:getter or #:accessor option.

• compute-setter-method class gns (generic)
compute-setter-method is invoked with the same arguments as compute-getter-

method, for each of the class’s slots that includes a #:setter or #:accessor slot option. The applied method should create and return a method that is specialized for instances of type class and uses the set closure to set the slot’s value. initialize then uses add-method! to add the returned method to the generic function named by the slot definition’s #:setter or #:accessor option.

8.11.6 Customizing Class Definition

If the metaclass of the new class is something more specialized than the default <class>, then the type of class in the calls above is more specialized than <class>, and hence it becomes possible to define generic function methods, specialized for the new class’s metaclass, that can modify or override the default behaviour of initialize, compute-cpl or

compute-get-n-set.

compute-cpl computes the class precedence list (“CPL”) for the new class (see Section 8.7.1 [Class Precedence List], page 734), and returns it as a list of class objects. The CPL is important because it defines a superclass ordering that is used, when a generic function is invoked upon an instance of the class, to decide which of the available generic function methods is the most specific. Hence compute-cpl could be customized in order to modify the CPL ordering algorithm for all classes with a special metaclass.

The default CPL algorithm is encapsulated by the compute-std-cpl procedure, which is called by the default compute-cpl method.

[procedure] compute-std-cpl class

Compute and return the class precedence list for class according to the algorithm described in Section 8.7.1 [Class Precedence List], page 734.

compute-slots computes and returns a list of all slot definitions for the new class. By default, this list includes the direct slot definitions from the define-class form, plus the slot definitions that are inherited from the new class’s superclasses. The default computeslots method uses the CPL computed by compute-cpl to calculate this union of slot definitions, with the rule that slots inherited from superclasses are shadowed by direct slots with the same name. One possible reason for customizing compute-slots would be to implement an alternative resolution strategy for slot name conflicts.

compute-get-n-set computes the low-level closures that will be used to get and set the value of a particular slot, and returns them in a list with two elements.

The closures returned depend on how storage for that slot is allocated. The standard compute-get-n-set method, specialized for classes of type <class>, handles the standard GOOPS values for the #:allocation slot option (see Section 8.4 [allocation], page 721). By defining a new compute-get-n-set method for a more specialized metaclass, it is possible to support new types of slot allocation.

Suppose you wanted to create a large number of instances of some class with a slot that should be shared between some but not all instances of that class - say every 10 instances should share the same slot storage. The following example shows how to implement and use a new type of slot allocation to do this.

     (define-class <batched-allocation-metaclass> (<class>))
     (let ((batch-allocation-count 0)
           (batch-get-n-set #f))
       (define-method (compute-get-n-set
                          (class <batched-allocation-metaclass>) s)
         (case (slot-definition-allocation s)
           ((#:batched)
            ;; If we've already used the same slot storage for 10 instances,
            ;; reset variables.
            (if (= batch-allocation-count 10)
                (begin
                  (set! batch-allocation-count 0)
                  (set! batch-get-n-set #f)))
            ;; If we don't have a current pair of get and set closures,
            ;; create one.  make-closure-variable returns a pair of closures
            ;; around a single Scheme variable - see goops.scm for details.
            (or batch-get-n-set
                (set! batch-get-n-set (make-closure-variable)))
            ;; Increment the batch allocation count.
            (set! batch-allocation-count (+ batch-allocation-count 1))
            batch-get-n-set)

           ;; Call next-method to handle standard allocation types.
           (else (next-method)))))
     (define-class <class-using-batched-slot> ()
       ...
       (c #:allocation #:batched)
       ...
       #:metaclass <batched-allocation-metaclass>)
   The usage of compute-getter-method and compute-setter-method is described in Sec-

tion 8.11.5 [Class Definition Protocol], page 747.

   compute-cpl and compute-get-n-set are called by the standard initialize method

for classes whose metaclass is <class>. But initialize itself can also be modified, by defining an initialize method specialized to the new class’s metaclass. Such a method could complete override the standard behaviour, by not calling (next-method) at all, but more typically it would perform additional class initialization steps before and/or after calling (next-method) for the standard behaviour.

8.11.7 Method Definition

define-method (syntax)
  • add-method! target method (generic)

define-method invokes the add-method! generic function to handle adding the new method to a variety of possible targets. GOOPS includes methods to handle target as

• a generic function (the most common case)

• a procedure

• a primitive generic (see Section 8.6.2 [Extending Primitives], page 728)

By defining further methods for add-method!, you can theoretically handle adding methods to further types of target.

8.11.8 Method Definition Internals

define-method:

• checks the form of the first parameter, and applies the following steps to the accessor’s setter if it has the (setter ...) form

• interpolates a call to define-generic or define-accessor if a generic function is not already defined with the supplied name

• calls method with the parameters and body, to make a new method instance

• calls add-method! to add this method to the relevant generic function.

[syntax] method (parameter …) body … Make a method whose specializers are defined by the classes in parameters and whose procedure definition is constructed from the parameter symbols and body forms. The parameter and body parameters should be as for define-method (see Section 8.6 [define-method], page 726).

method:

• extracts formals and specializing classes from the parameters, defaulting the class for unspecialized parameters to <top>

• creates a closure using the formals and the body forms

• calls make with metaclass <method> and the specializers and closure using the

#:specializers and #:procedure keywords.

[procedure] make-method specializers procedure Make a method using specializers and procedure. specializers should be a list of classes that specifies the parameter combinations to which this method will be applicable. procedure should be the closure that will applied to the generic function parameters when this method is invoked.

make-method is a simple wrapper around make with metaclass <method>.

[generic] add-method! target method Generic function for adding method method to target.

[method] add-method! (generic <generic>) (method <method>)

Add method method to the generic function generic.

[method] add-method! (proc <procedure>) (method <method>)

If proc is a procedure with generic capability (see Section 8.6.2 [generic-capability?], page 728), upgrade it to a primitive generic and add method to its generic function definition.

[method] add-method! (pg <primitive-generic>) (method <method>)

Add method method to the generic function definition of pg.

Implementation: (add-method! (primitive-generic-generic pg) method).

[method] add-method! (whatever <top>) (method <method>)

Raise an error indicating that whatever is not a valid generic function.

8.11.9 Generic Function Internals

define-generic calls ensure-generic to upgrade a pre-existing procedure value, or make with metaclass <generic> to create a new generic function.

define-accessor calls ensure-accessor to upgrade a pre-existing procedure value, or make-accessor to create a new accessor.

[procedure] ensure-generic old-definition [name] Return a generic function with name name, if possible by using or upgrading olddefinition. If unspecified, name defaults to #f. If old-definition is already a generic function, it is returned unchanged. If old-definition is a Scheme procedure or procedure-with-setter, ensure-generic returns a new generic function that uses old-definition for its default procedure and setter. Otherwise ensure-generic returns a new generic function with no defaults and no methods.

[procedure] make-generic [name]

Return a new generic function with name (car name). If unspecified, name defaults to #f.

   ensure-generic calls make with metaclasses <generic> and <generic-with-setter>,

depending on the previous value of the variable that it is trying to upgrade.

make-generic is a simple wrapper for make with metaclass <generic>.

[procedure] ensure-accessor proc [name]

Return an accessor with name name, if possible by using or upgrading proc. If unspecified, name defaults to #f.

If proc is already an accessor, it is returned unchanged.

If proc is a Scheme procedure, procedure-with-setter or generic function, ensureaccessor returns an accessor that reuses the reusable elements of proc.

Otherwise ensure-accessor returns a new accessor with no defaults and no methods.

[procedure] make-accessor [name]

Return a new accessor with name (car name). If unspecified, name defaults to #f.

   ensure-accessor calls make with metaclass <generic-with-setter>, as well as calls
to ensure-generic, make-accessor and (tail recursively) ensure-accessor.

make-accessor calls make twice, first with metaclass <generic> to create a generic function for the setter, then with metaclass <generic-with-setter> to create the accessor, passing the setter generic function as the value of the #:setter keyword.

8.11.10 Generic Function Invocation

There is a detailed and customizable protocol involved in the process of invoking a generic function — i.e., in the process of deciding which of the generic function’s methods are applicable to the current arguments, and which one of those to apply. Here is a summary diagram of the generic functions involved.

apply-generic (generic)

• no-method (generic)

  • compute-applicable-methods (generic)
  • sort-applicable-methods (generic)
     • method-more-specific? (generic)
  • apply-methods (generic)
     • apply-method (generic)
     • no-next-method (generic)
  • no-applicable-method

We do not yet have full documentation for these. Please refer to the code

(oop/goops.scm) for details.

8.12 Redefining a Class

Suppose that a class <my-class> is defined using define-class (see Section 8.2 [defineclass], page 719), with slots that have accessor functions, and that an application has created several instances of <my-class> using make (see Section 8.3 [make], page 720). What then happens if <my-class> is redefined by calling define-class again?

8.12.1 Default Class Redefinition Behaviour

GOOPS’ default answer to this question is as follows.

• All existing direct instances of <my-class> are converted to be instances of the new class. This is achieved by preserving the values of slots that exist in both the old and new definitions, and initializing the values of new slots in the usual way (see Section 8.3 [make], page 720).

• All existing subclasses of <my-class> are redefined, as though the define-class expressions that defined them were re-evaluated following the redefinition of <my-class>, and the class redefinition process described here is applied recursively to the redefined subclasses.

• Once all of its instances and subclasses have been updated, the class metaobject previously bound to the variable <my-class> is no longer needed and so can be allowed to be garbage collected.

To keep things tidy, GOOPS also needs to do a little housekeeping on methods that are associated with the redefined class.

• Slot accessor methods for slots in the old definition should be removed from their generic functions. They will be replaced by accessor methods for the slots of the new class definition.

• Any generic function method that uses the old <my-class> metaobject as one of its formal parameter specializers must be updated to refer to the new <my-class> metaobject. (Whenever a new generic function method is defined, define-method adds the method to a list stored in the class metaobject for each class used as a formal parameter specializer, so it is easy to identify all the methods that must be updated when a class is redefined.)

If this class redefinition strategy strikes you as rather counter-intuitive, bear in mind that it is derived from similar behaviour in other object systems such as CLOS, and that experience in those systems has shown it to be very useful in practice.

Also bear in mind that, like most of GOOPS’ default behaviour, it can be customized…

8.12.2 Customizing Class Redefinition

When define-class notices that a class is being redefined, it constructs the new class metaobject as usual, then invokes the class-redefinition generic function with the old and new classes as arguments. Therefore, if the old or new classes have metaclasses other than the default <class>, class redefinition behaviour can be customized by defining a class-redefinition method that is specialized for the relevant metaclasses.

[generic] class-redefinition

Handle the class redefinition from old-class to new-class, and return the new class metaobject that should be bound to the variable specified by define-class’s first argument.

[method] class-redefinition (old-class <class>) (new-class <class>)

Implements GOOPS’ default class redefinition behaviour, as described in Section 8.12.1 [Default Class Redefinition Behaviour], page 755. Returns the metaobject for the new class definition.

The default class-redefinition method, for classes with the default metaclass <class>, calls the following generic functions, which could of course be individually customized.

[generic] remove-class-accessors! old

The default remove-class-accessors! method removes the accessor methods of the old class from all classes which they specialize.

[generic] update-direct-method! method old new

The default update-direct-method! method substitutes the new class for the old in all methods specialized to the old class.

[generic] update-direct-subclass! subclass old new
     The default update-direct-subclass! method invokes class-redefinition recur-

sively to handle the redefinition of subclasses.

An alternative class redefinition strategy could be to leave all existing instances as instances of the old class, but accepting that the old class is now “nameless”, since its name has been taken over by the new definition. In this strategy, any existing subclasses could also be left as they are, on the understanding that they inherit from a nameless superclass.

This strategy is easily implemented in GOOPS, by defining a new metaclass, that will be used as the metaclass for all classes to which the strategy should apply, and then defining a class-redefinition method that is specialized for this metaclass:

     (define-class <can-be-nameless> (<class>))
     (define-method (class-redefinition (old <can-be-nameless>)
                                        (new <class>))
       new)

When customization can be as easy as this, aren’t you glad that GOOPS implements the far more difficult strategy as its default!

8.13 Changing the Class of an Instance

When a class is redefined, any existing instance of the redefined class will be modified for the new class definition before the next time that any of the instance’s slots is referenced or set. GOOPS modifies each instance by calling the generic function change-class.

More generally, you can change the class of an existing instance at any time by invoking the generic function change-class with two arguments: the instance and the new class.

The default method for change-class decides how to implement the change of class by looking at the slot definitions for the instance’s existing class and for the new class. If the new class has slots with the same name as slots in the existing class, the values for those slots are preserved. Slots that are present only in the existing class are discarded. Slots that are present only in the new class are initialized using the corresponding slot definition’s init function (see Section 8.8.1 [slot-init-function], page 736).

[generic] change-class instance new-class

[method] change-class (obj <object>) (new <class>)

Modify instance obj to make it an instance of class new.

The value of each of obj’s slots is preserved only if a similarly named slot exists in new; any other slot values are discarded. The slots in new that do not correspond to any of obj’s pre-existing slots are initialized according to new’s slot definitions’ init functions.

The default change-class method also invokes another generic function, updateinstance-for-different-class, as the last thing that it does before returning.

The applied update-instance-for-different-class method can make any further

adjustments to new-instance that are required to complete or modify the change of class. The return value from the applied method is ignored.

[generic] update-instance-for-different-class old-instance new-instance

A generic function that can be customized to put finishing touches to an instance whose class has just been changed. The default update-instance-for-different-class method does nothing.

Customized change of class behaviour can be implemented by defining change-class methods that are specialized either by the class of the instances to be modified or by the metaclass of the new class.

9 Guile Implementation

At some point, after one has been programming in Scheme for some time, another level of Scheme comes into view: its implementation. Knowledge of how Scheme can be implemented turns out to be necessary to become an expert hacker. As Peter Norvig notes in his retrospective on PAIP^{1}, “The expert Lisp programmer eventually develops a good ‘efficiency model’.”

By this Norvig means that over time, the Lisp hacker eventually develops an understanding of how much her code “costs” in terms of space and time.

This chapter describes Guile as an implementation of Scheme: its history, how it represents and evaluates its data, and its compiler. This knowledge can help you to make that step from being one who is merely familiar with Scheme to being a real hacker.

9.1 A Brief History of Guile

Guile is an artifact of historical processes, both as code and as a community of hackers. It is sometimes useful to know this history when hacking the source code, to know about past decisions and future directions.

Of course, the real history of Guile is written by the hackers hacking and not the writers writing, so we round up the section with a note on current status and future directions.

9.1.1 The Emacs Thesis

The story of Guile is the story of bringing the development experience of Emacs to the mass of programs on a GNU system.

Emacs, when it was first created in its GNU form in 1984, was a new take on the problem of “how to make a program”. The Emacs thesis is that it is delightful to create composite programs based on an orthogonal kernel written in a low-level language together with a powerful, high-level extension language.

Extension languages foster extensible programs, programs which adapt readily to different users and to changing times. Proof of this can be seen in Emacs’ current and continued existence, spanning more than a quarter-century.

Besides providing for modification of a program by others, extension languages are good for intension as well. Programs built in “the Emacs way” are pleasurable and easy for their authors to flesh out with the features that they need.

After the Emacs experience was appreciated more widely, a number of hackers started to consider how to spread this experience to the rest of the GNU system. It was clear that the easiest way to Emacsify a program would be to embed a shared language implementation into it.

9.1.2 Early Days

Tom Lord was the first to fully concentrate his efforts on an embeddable language runtime, which he named “GEL”, the GNU Extension Language.

1 PAIP is the common abbreviation for Paradigms of Artificial Intelligence Programming, an old but still

useful text on Lisp. Norvig’s retrospective sums up the lessons of PAIP, and can be found at http://

norvig.com/Lisp-retro.html.

GEL was the product of converting SCM, Aubrey Jaffer’s implementation of Scheme, into something more appropriate to embedding as a library. (SCM was itself based on an implementation by George Carrette, SIOD.)

Lord managed to convince Richard Stallman to dub GEL the official extension language for the GNU project. It was a natural fit, given that Scheme was a cleaner, more modern Lisp than Emacs Lisp. Part of the argument was that eventually when GEL became more capable, it could gain the ability to execute other languages, especially Emacs Lisp.

Due to a naming conflict with another programming language, Jim Blandy suggested a new name for GEL: “Guile”. Besides being a recursive acronym, “Guile” craftily follows the naming of its ancestors, “Planner”, “Conniver”, and “Schemer”. (The latter was truncated to “Scheme” due to a 6-character file name limit on an old operating system.) Finally, “Guile” suggests “guy-ell”, or “Guy L. Steele”, who, together with Gerald Sussman, originally discovered Scheme.

Around the same time that Guile (then GEL) was readying itself for public release, another extension language was gaining in popularity, Tcl. Many developers found advantages in Tcl because of its shell-like syntax and its well-developed graphical widgets library, Tk. Also, at the time there was a large marketing push promoting Tcl as a “universal extension language”.

Richard Stallman, as the primary author of GNU Emacs, had a particular vision of what extension languages should be, and Tcl did not seem to him to be as capable as Emacs Lisp. He posted a criticism to the comp.lang.tcl newsgroup, sparking one of the internet’s legendary flamewars. As part of these discussions, retrospectively dubbed the “Tcl Wars”, he announced the Free Software Foundation’s intent to promote Guile as the extension language for the GNU project.

It is a common misconception that Guile was created as a reaction to Tcl. While it is true that the public announcement of Guile happened at the same time as the “Tcl wars”, Guile was created out of a condition that existed outside the polemic. Indeed, the need for a powerful language to bridge the gap between extension of existing applications and a more fully dynamic programming environment is still with us today.

9.1.3 A Scheme of Many Maintainers

Surveying the field, it seems that Scheme implementations correspond with their maintainers on an N-to-1 relationship. That is to say, that those people that implement Schemes might do so on a number of occasions, but that the lifetime of a given Scheme is tied to the maintainership of one individual.

Guile is atypical in this regard.

Tom Lord maintained Guile for its first year and a half or so, corresponding to the end of 1994 through the middle of 1996. The releases made in this time constitute an arc from SCM as a standalone program to Guile as a reusable, embeddable library, but passing through a explosion of features: embedded Tcl and Tk, a toolchain for compiling and disassembling Java, addition of a C-like syntax, creation of a module system, and a start at a rich POSIX interface.

Only some of those features remain in Guile. There were ongoing tensions between providing a small, embeddable language, and one which had all of the features (e.g. a graphical toolkit) that a modern Emacs might need. In the end, as Guile gained in uptake, the development team decided to focus on depth, documentation and orthogonality rather than on breadth. This has been the focus of Guile ever since, although there is a wide range of third-party libraries for Guile.

Jim Blandy presided over that period of stabilization, in the three years until the end of 1999, when he too moved on to other projects. Since then, Guile has had a group maintainership. The first group was Maciej Stachowiak, Mikael Djurfeldt, and Marius Vollmer, with Vollmer staying on the longest. By late 2007, Vollmer had mostly moved on to other things, so Neil Jerram and Ludovic Courtès stepped up to take on the primary maintenance responsibility. Jerram and Courtès were joined by Andy Wingo in late 2009.

Of course, a large part of the actual work on Guile has come from other contributors too numerous to mention, but without whom the world would be a poorer place.

9.1.4 A Timeline of Selected Guile Releases

guile-i — 4 February 1995 SCM, turned into a library.

guile-ii — 6 April 1995 A low-level module system was added. Tcl/Tk support was added, allowing extension of Scheme by Tcl or vice versa. POSIX support was improved, and there was an experimental stab at Java integration.

guile-iii — 18 August 1995 The C-like syntax, ctax, was improved, but mostly this release featured a start at the task of breaking Guile into pieces.

1.0 — 5 January 1997 #f was distinguished from '(). User-level, cooperative multi-threading was added. Source-level debugging became more useful, and programmer’s and user’s manuals were begun. The module system gained a high-level interface, which is still used today in more or less the same form.

1.1 — 16 May 1997 1.2 — 24 June 1997

Support for Tcl/Tk and ctax were split off as separate packages, and have remained there since. Guile became more compatible with SCSH, and more useful as a UNIX scripting language. Libguile could now be built as a shared library, and third-party extensions written in C became loadable via dynamic linking.

1.3.0 — 19 October 1998 Command-line editing became much more pleasant through the use of the readline library. The initial support for internationalization via multi-byte strings was removed; 10 years were to pass before proper internationalization would land again. Initial Emacs Lisp support landed, ports gained better support for file descriptors, and fluids were added.

1.3.2 — 20 August 1999 1.3.4 — 25 September 1999 1.4 — 21 June 2000

A long list of lispy features were added: hooks, Common Lisp’s format, optional and keyword procedure arguments, getopt-long, sorting, random numbers, and many other fixes and enhancements. Guile also gained an interactive debugger, interactive help, and better backtraces.

1.6 — 6 September 2002 Guile gained support for the R5RS standard, and added a number of SRFI modules. The module system was expanded with programmatic support for identifier selection and renaming. The GOOPS object system was merged into Guile core.

1.8 — 20 February 2006 Guile’s arbitrary-precision arithmetic switched to use the GMP library, and added support for exact rationals. Guile’s embedded user-space threading was removed in favor of POSIX pre-emptive threads, providing true multiprocessing. Gettext support was added, and Guile’s C API was cleaned up and orthogonalized in a massive way.

2.0 — 16 February 2010 A virtual machine was added to Guile, along with the associated compiler and toolchain. Support for internationalization was finally reimplemented, in terms of unicode, locales, and libunistring. Running Guile instances became controllable and debuggable from within Emacs, via Geiser. Guile caught up to features found in a number of other Schemes: SRFI-18 threads, module-hygienic macros, a profiler, tracer, and debugger, SSAX XML integration, bytevectors, a dynamic FFI, delimited continuations, module versions, and partial support for R6RS.

9.1.5 Status, or: Your Help Needed

Guile has achieved much of what it set out to achieve, but there is much remaining to do.

There is still the old problem of bringing existing applications into a more Emacs-like experience. Guile has had some successes in this respect, but still most applications in the GNU system are without Guile integration.

Getting Guile to those applications takes an investment, the “hacktivation energy” needed to wire Guile into a program that only pays off once it is good enough to enable new kinds of behavior. This would be a great way for new hackers to contribute: take an application that you use and that you know well, think of something that it can’t yet do, and figure out a way to integrate Guile and implement that task in Guile.

With time, perhaps this exposure can reverse itself, whereby programs can run under Guile instead of vice versa, eventually resulting in the Emacsification of the entire GNU system. Indeed, this is the reason for the naming of the many Guile modules that live in the ice-9 namespace, a nod to the fictional substance in Kurt Vonnegut’s novel, Cat’s Cradle, capable of acting as a seed crystal to crystallize the mass of software.

Implicit to this whole discussion is the idea that dynamic languages are somehow better than languages like C. While languages like C have their place, Guile’s take on this question is that yes, Scheme is more expressive than C, and more fun to write. This realization carries an imperative with it to write as much code in Scheme as possible rather than in other languages.

These days it is possible to write extensible applications almost entirely from highlevel languages, through byte-code and native compilation, speed gains in the underlying hardware, and foreign call interfaces in the high-level language. Smalltalk systems are like this, as are Common Lisp-based systems. While there already are a number of pure-Guile applications out there, users still need to drop down to C for some tasks: interfacing to system libraries that don’t have prebuilt Guile interfaces, and for some tasks requiring high performance.

The addition of the virtual machine in Guile 2.0, together with the compiler infrastructure, should go a long way to addressing the speed issues. But there is much optimization to be done. Interested contributors will find lots of delightful low-hanging fruit, from simple profile-driven optimization to hacking a just-in-time compiler from VM bytecode to native code.

Still, even with an all-Guile application, sometimes you want to provide an opportunity for users to extend your program from a language with a syntax that is closer to C, or to Python. Another interesting idea to consider is compiling e.g. Python to Guile. It’s not that far-fetched of an idea: see for example IronPython or JRuby.

And then there’s Emacs itself. Though there is a somewhat-working Emacs Lisp language frontend for Guile, it cannot yet execute all of Emacs Lisp. A serious integration of Guile with Emacs would replace the Elisp virtual machine with Guile, and provide the necessary C shims so that Guile could emulate Emacs’ C API. This would give lots of exciting things to Emacs: native threads, a real object system, more sophisticated types, cleaner syntax, and access to all of the Guile extensions.

Finally, there is another axis of crystallization, the axis between different Scheme implementations. Guile does not yet support the latest Scheme standard, R6RS, and should do so. Like all standards, R6RS is imperfect, but supporting it will allow more code to run on Guile without modification, and will allow Guile hackers to produce code compatible with other schemes. Help in this regard would be much appreciated.

9.2 Data Representation

Scheme is a latently-typed language; this means that the system cannot, in general, determine the type of a given expression at compile time. Types only become apparent at run time. Variables do not have fixed types; a variable may hold a pair at one point, an integer at the next, and a thousand-element vector later. Instead, values, not variables, have fixed types.

In order to implement standard Scheme functions like pair? and string? and provide garbage collection, the representation of every value must contain enough information to accurately determine its type at run time. Often, Scheme systems also use this information to determine whether a program has attempted to apply an operation to an inappropriately typed value (such as taking the car of a string).

Because variables, pairs, and vectors may hold values of any type, Scheme implementations use a uniform representation for values — a single type large enough to hold either a complete value or a pointer to a complete value, along with the necessary typing information.

The following sections will present a simple typing system, and then make some refinements to correct its major weaknesses. We then conclude with a discussion of specific choices that Guile has made regarding garbage collection and data representation.

9.2.1 A Simple Representation

The simplest way to represent Scheme values in C would be to represent each value as a pointer to a structure containing a type indicator, followed by a union carrying the real value. Assuming that SCM is the name of our universal type, we can write:

     enum type { integer, pair, string, vector, ... };
     typedef struct value *SCM;
     struct value {
       enum type type;
       union {
         int integer;
         struct { SCM car, cdr; } pair;
         struct { int length; char *elts; } string;
         struct { int length; SCM  *elts; } vector;
         ...
       } value;
     };

with the ellipses replaced with code for the remaining Scheme types.

This representation is sufficient to implement all of Scheme’s semantics. If x is an SCM value:

• To test if x is an integer, we can write x->type == integer.

• To find its value, we can write x->value.integer.

• To test if x is a vector, we can write x->type == vector.

• If we know x is a vector, we can write x->value.vector.elts[0] to refer to its first element.

• If we know x is a pair, we can write x->value.pair.car to extract its car.

9.2.2 Faster Integers

Unfortunately, the above representation has a serious disadvantage. In order to return an integer, an expression must allocate a struct value, initialize it to represent that integer, and return a pointer to it. Furthermore, fetching an integer’s value requires a memory reference, which is much slower than a register reference on most processors. Since integers are extremely common, this representation is too costly, in both time and space. Integers should be very cheap to create and manipulate.

One possible solution comes from the observation that, on many architectures, heapallocated data (i.e., what you get when you call malloc) must be aligned on an eight-byte boundary. (Whether or not the machine actually requires it, we can write our own allocator for struct value objects that assures this is true.) In this case, the lower three bits of the structure’s address are known to be zero.

This gives us the room we need to provide an improved representation for integers. We make the following rules:

• If the lower three bits of an SCM value are zero, then the SCM value is a pointer to a struct value, and everything proceeds as before.

• Otherwise, the SCM value represents an integer, whose value appears in its upper bits.

Here is C code implementing this convention:

     enum type { pair, string, vector, ... };
     typedef struct value *SCM;
     struct value {
       enum type type;
       union {
         struct { SCM car, cdr; } pair;
         struct { int length; char *elts; } string;
         struct { int length; SCM  *elts; } vector;
         ...
       } value;
     };
     #define POINTER_P(x) (((int) (x) & 7) == 0)
     #define INTEGER_P(x) (! POINTER_P (x))
     #define GET_INTEGER(x)  ((int) (x) >> 3)
     #define MAKE_INTEGER(x) ((SCM) (((x) << 3) | 1))

Notice that integer no longer appears as an element of enum type, and the union has lost its integer member. Instead, we use the POINTER_P and INTEGER_P macros to make a coarse classification of values into integers and non-integers, and do further type testing as before.

Here’s how we would answer the questions posed above (again, assume x is an SCM value):

• To test if x is an integer, we can write INTEGER_P (x).

• To find its value, we can write GET_INTEGER (x).

• To test if x is a vector, we can write:

           POINTER_P (x) && x->type == vector

Given the new representation, we must make sure x is truly a pointer before we dereference it to determine its complete type.

• If we know x is a vector, we can write x->value.vector.elts[0] to refer to its first element, as before.

• If we know x is a pair, we can write x->value.pair.car to extract its car, just as before.

This representation allows us to operate more efficiently on integers than the first. For example, if x and y are known to be integers, we can compute their sum as follows:

MAKE_INTEGER (GET_INTEGER (x) + GET_INTEGER (y))

Now, integer math requires no allocation or memory references. Most real Scheme systems actually implement addition and other operations using an even more efficient algorithm, but this essay isn’t about bit-twiddling. (Hint: how do you decide when to overflow to a bignum? How would you do it in assembly?)

9.2.3 Cheaper Pairs

However, there is yet another issue to confront. Most Scheme heaps contain more pairs than any other type of object; Jonathan Rees said at one point that pairs occupy 45% of the heap in his Scheme implementation, Scheme 48. However, our representation above spends three SCM-sized words per pair — one for the type, and two for the car and cdr. Is there any way to represent pairs using only two words?

Let us refine the convention we established earlier. Let us assert that:

• If the bottom three bits of an SCM value are #b000, then it is a pointer, as before.

• If the bottom three bits are #b001, then the upper bits are an integer. This is a bit more restrictive than before.

• If the bottom two bits are #b010, then the value, with the bottom three bits masked out, is the address of a pair.

Here is the new C code:

enum type { string, vector, ... };
typedef struct value *SCM;
struct value {
  enum type type;
  union {
    struct { int length; char *elts; } string;
    struct { int length; SCM  *elts; } vector;
    ...
  } value;
};
struct pair {
  SCM car, cdr;
};
#define POINTER_P(x) (((int) (x) & 7) == 0)
#define INTEGER_P(x)  (((int) (x) & 7) == 1)
#define GET_INTEGER(x)  ((int) (x) >> 3)
#define MAKE_INTEGER(x) ((SCM) (((x) << 3) | 1))
#define PAIR_P(x) (((int) (x) & 7) == 2)
#define GET_PAIR(x) ((struct pair *) ((int) (x) & ~7))

Notice that enum type and struct value now only contain provisions for vectors and strings; both integers and pairs have become special cases. The code above also assumes that an int is large enough to hold a pointer, which isn’t generally true.

Our list of examples is now as follows:

• To test if x is an integer, we can write INTEGER_P (x); this is as before.

• To find its value, we can write GET_INTEGER (x), as before.

• To test if x is a vector, we can write:

           POINTER_P (x) && x->type == vector

We must still make sure that x is a pointer to a struct value before dereferencing it to find its type.

• If we know x is a vector, we can write x->value.vector.elts[0] to refer to its first element, as before.

• We can write PAIR_P (x) to determine if x is a pair, and then write GET_PAIR (x)->car to refer to its car.

This change in representation reduces our heap size by 15%. It also makes it cheaper to decide if a value is a pair, because no memory references are necessary; it suffices to check the bottom two bits of the SCM value. This may be significant when traversing lists, a common activity in a Scheme system.

Again, most real Scheme systems use a slightly different implementation; for example, if GET PAIR subtracts off the low bits of x, instead of masking them off, the optimizer will often be able to combine that subtraction with the addition of the offset of the structure member we are referencing, making a modified pointer as fast to use as an unmodified pointer.

9.2.4 Conservative Garbage Collection

Aside from the latent typing, the major source of constraints on a Scheme implementation’s data representation is the garbage collector. The collector must be able to traverse every live object in the heap, to determine which objects are not live, and thus collectable.

There are many ways to implement this. Guile’s garbage collection is built on a library, the Boehm-Demers-Weiser conservative garbage collector (BDW-GC). The BDW-GC “just works”, for the most part. But since it is interesting to know how these things work, we include here a high-level description of what the BDW-GC does.

Garbage collection has two logical phases: a mark phase, in which the set of live objects is enumerated, and a sweep phase, in which objects not traversed in the mark phase are collected. Correct functioning of the collector depends on being able to traverse the entire set of live objects.

In the mark phase, the collector scans the system’s global variables and the local variables on the stack to determine which objects are immediately accessible by the C code. It then scans those objects to find the objects they point to, and so on. The collector logically sets a mark bit on each object it finds, so each object is traversed only once.

When the collector can find no unmarked objects pointed to by marked objects, it assumes that any objects that are still unmarked will never be used by the program (since there is no path of dereferences from any global or local variable that reaches them) and deallocates them.

In the above paragraphs, we did not specify how the garbage collector finds the global and local variables; as usual, there are many different approaches. Frequently, the programmer must maintain a list of pointers to all global variables that refer to the heap, and another list (adjusted upon entry to and exit from each function) of local variables, for the collector’s benefit.

The list of global variables is usually not too difficult to maintain, since global variables are relatively rare. However, an explicitly maintained list of local variables (in the author’s personal experience) is a nightmare to maintain. Thus, the BDW-GC uses a technique called conservative garbage collection, to make the local variable list unnecessary.

The trick to conservative collection is to treat the stack as an ordinary range of memory, and assume that every word on the stack is a pointer into the heap. Thus, the collector marks all objects whose addresses appear anywhere in the stack, without knowing for sure how that word is meant to be interpreted.

In addition to the stack, the BDW-GC will also scan static data sections. This means that global variables are also scanned when looking for live Scheme objects.

Obviously, such a system will occasionally retain objects that are actually garbage, and should be freed. In practice, this is not a problem. The alternative, an explicitly maintained list of local variable addresses, is effectively much less reliable, due to programmer error. Interested readers should see the BDW-GC web page at http://www.hpl.hp.com/

personal/Hans_Boehm/gc, for more information.

9.2.5 The SCM Type in Guile

Guile classifies Scheme objects into two kinds: those that fit entirely within an SCM, and those that require heap storage.

The former class are called immediates. The class of immediates includes small integers, characters, boolean values, the empty list, the mysterious end-of-file object, and some others.

The remaining types are called, not surprisingly, non-immediates. They include pairs, procedures, strings, vectors, and all other data types in Guile. For non-immediates, the SCM word contains a pointer to data on the heap, with further information about the object in question is stored in that data.

This section describes how the SCM type is actually represented and used at the C level. Interested readers should see libguile/tags.h for an exposition of how Guile stores type information.

In fact, there are two basic C data types to represent objects in Guile: SCM and scm_t_

bits.

9.2.5.1 Relationship between SCM and scm_t_bits

A variable of type SCM is guaranteed to hold a valid Scheme object. A variable of type scm_t_bits, on the other hand, may hold a representation of a SCM value as a C integral type, but may also hold any C value, even if it does not correspond to a valid Scheme object.

For a variable x of type SCM, the Scheme object’s type information is stored in a form that is not directly usable. To be able to work on the type encoding of the scheme value, the SCM variable has to be transformed into the corresponding representation as a scm_t_ bits variable y by using the SCM_UNPACK macro. Once this has been done, the type of the scheme object x can be derived from the content of the bits of the scm_t_bits value y, in the way illustrated by the example earlier in this chapter (see Section 9.2.3 [Cheaper Pairs], page 766). Conversely, a valid bit encoding of a Scheme value as a scm_t_bits variable can be transformed into the corresponding SCM value using the SCM_PACK macro.

9.2.5.2 Immediate objects

A Scheme object may either be an immediate, i.e. carrying all necessary information by itself, or it may contain a reference to a cell with additional information on the heap. Although in general it should be irrelevant for user code whether an object is an immediate or not, within Guile’s own code the distinction is sometimes of importance. Thus, the following low level macro is provided:

[Macro] int SCM_IMP (SCM x)

A Scheme object is an immediate if it fulfills the SCM_IMP predicate, otherwise it holds an encoded reference to a heap cell. The result of the predicate is delivered as a C style boolean value. User code and code that extends Guile should normally not be required to use this macro.

Summary:

• Given a Scheme object x of unknown type, check first with SCM_IMP (x) if it is an immediate object.

• If so, all of the type and value information can be determined from the scm_t_bits value that is delivered by SCM_UNPACK (x).

There are a number of special values in Scheme, most of them documented elsewhere in this manual. It’s not quite the right place to put them, but for now, here’s a list of the C names given to some of these values:

[Macro] SCM SCM_EOL

The Scheme empty list object, or “End Of List” object, usually written in Scheme as

'().
[Macro] SCM SCM_EOF_VAL

The Scheme end-of-file value. It has no standard written representation, for obvious reasons.

[Macro] SCM SCM_UNSPECIFIED

The value returned by some (but not all) expressions that the Scheme standard says return an “unspecified” value. This is sort of a weirdly literal way to take things, but the standard read-eval-print loop prints nothing when the expression returns this value, so it’s not a bad idea to return this when you can’t think of anything else helpful.

[Macro] SCM SCM_UNDEFINED

The “undefined” value. Its most important property is that is not equal to any valid Scheme value. This is put to various internal uses by C code interacting with Guile. For example, when you write a C function that is callable from Scheme and which takes optional arguments, the interpreter passes SCM_UNDEFINED for any arguments you did not receive. We also use this to mark unbound variables.

[Macro] int SCM_UNBNDP (SCM x)

Return true if x is SCM_UNDEFINED. Note that this is not a check to see if x is SCM_UNBOUND. History will not be kind to us.

9.2.5.3 Non-immediate objects

A Scheme object of type SCM that does not fulfill the SCM_IMP predicate holds an encoded reference to a heap cell. This reference can be decoded to a C pointer to a heap cell using the SCM2PTR macro. The encoding of a pointer to a heap cell into a SCM value is done using the PTR2SCM macro.

[Macro] scm_t_cell * SCM2PTR (SCM x)

Extract and return the heap cell pointer from a non-immediate SCM object x.

[Macro] SCM PTR2SCM (scm t cell * x)

Return a SCM value that encodes a reference to the heap cell pointer x.

Note that it is also possible to transform a non-immediate SCM value by using SCM_ UNPACK into a scm_t_bits variable. However, the result of SCM_UNPACK may not be used as a pointer to a scm_t_cell: only SCM2PTR is guaranteed to transform a SCM object into a valid pointer to a heap cell. Also, it is not allowed to apply PTR2SCM to anything that is not a valid pointer to a heap cell.

Summary:

• Only use SCM2PTR on SCM values for which SCM_IMP is false!

  • Don’t use (scm_t_cell *) SCM_UNPACK (x)! Use SCM2PTR (x) instead!

• Don’t use PTR2SCM for anything but a cell pointer!

9.2.5.4 Allocating Cells

Guile provides both ordinary cells with two slots, and double cells with four slots. The following two function are the most primitive way to allocate such cells.

If the caller intends to use it as a header for some other type, she must pass an appropriate magic value in word 0, to mark it as a member of that type, and pass whatever value as word 1, etc that the type expects. You should generally not need these functions, unless you are implementing a new datatype, and thoroughly understand the code in

<libguile/tags.h>.

If you just want to allocate pairs, use scm_cons.

[Function] SCM scm_cell (scm t bits word 0, scm t bits word 1)

Allocate a new cell, initialize the two slots with word 0 and word 1, and return it.

Note that word 0 and word 1 are of type scm_t_bits. If you want to pass a SCM object, you need to use SCM_UNPACK.

[Function] SCM scm_double_cell (scm t bits word 0, scm t bits word 1, scm t bits

word 2, scm t bits word 3) Like scm_cell, but allocates a double cell with four slots.

9.2.5.5 Heap Cell Type Information

Heap cells contain a number of entries, each of which is either a scheme object of type SCM or a raw C value of type scm_t_bits. Which of the cell entries contain Scheme objects and which contain raw C values is determined by the first entry of the cell, which holds the cell type information.

[Macro] scm_t_bits SCM_CELL_TYPE (SCM x)

For a non-immediate Scheme object x, deliver the content of the first entry of the heap cell referenced by x. This value holds the information about the cell type.

[Macro] void SCM_SET_CELL_TYPE (SCM x, scm t bits t)

For a non-immediate Scheme object x, write the value t into the first entry of the heap cell referenced by x. The value t must hold a valid cell type.

9.2.5.6 Accessing Cell Entries

For a non-immediate Scheme object x, the object type can be determined by reading the cell type entry using the SCM_CELL_TYPE macro. For each different type of cell it is known which cell entries hold Scheme objects and which cell entries hold raw C data. To access the different cell entries appropriately, the following macros are provided.

[Macro] scm_t_bits SCM_CELL_WORD (SCM x, unsigned int n)

Deliver the cell entry n of the heap cell referenced by the non-immediate Scheme object x as raw data. It is illegal, to access cell entries that hold Scheme objects by using these macros. For convenience, the following macros are also provided.

• SCM CELL WORD 0 (x) ⇒ SCM CELL WORD (x, 0)

• SCM CELL WORD 1 (x) ⇒ SCM CELL WORD (x, 1)

• …

• SCM CELL WORD n (x) ⇒ SCM CELL WORD (x, n)

[Macro] SCM SCM_CELL_OBJECT (SCM x, unsigned int n)

Deliver the cell entry n of the heap cell referenced by the non-immediate Scheme object x as a Scheme object. It is illegal, to access cell entries that do not hold Scheme objects by using these macros. For convenience, the following macros are also provided.

• SCM CELL OBJECT 0 (x) ⇒ SCM CELL OBJECT (x, 0)

• SCM CELL OBJECT 1 (x) ⇒ SCM CELL OBJECT (x, 1)

• …

• SCM CELL OBJECT n (x) ⇒ SCM CELL OBJECT (x, n)

[Macro] void SCM_SET_CELL_WORD (SCM x, unsigned int n, scm t bits w)

Write the raw C value w into entry number n of the heap cell referenced by the non-immediate Scheme value x. Values that are written into cells this way may only be read from the cells using the SCM_CELL_WORD macros or, in case cell entry 0 is written, using the SCM_CELL_TYPE macro. For the special case of cell entry 0 it has to be made sure that w contains a cell type information which does not describe a Scheme object. For convenience, the following macros are also provided.

• SCM SET CELL WORD 0 (x, w) ⇒ SCM SET CELL WORD (x, 0, w)

• SCM SET CELL WORD 1 (x, w) ⇒ SCM SET CELL WORD (x, 1, w)

• …

• SCM SET CELL WORD n (x, w) ⇒ SCM SET CELL WORD (x, n, w)

[Macro] void SCM_SET_CELL_OBJECT (SCM x, unsigned int n, SCM o)

Write the Scheme object o into entry number n of the heap cell referenced by the non-immediate Scheme value x. Values that are written into cells this way may only be read from the cells using the SCM_CELL_OBJECT macros or, in case cell entry 0 is written, using the SCM_CELL_TYPE macro. For the special case of cell entry 0 the writing of a Scheme object into this cell is only allowed if the cell forms a Scheme pair. For convenience, the following macros are also provided.

• SCM SET CELL OBJECT 0 (x, o) ⇒ SCM SET CELL OBJECT (x, 0, o)

• SCM SET CELL OBJECT 1 (x, o) ⇒ SCM SET CELL OBJECT (x, 1, o)

• …

• SCM SET CELL OBJECT n (x, o) ⇒ SCM SET CELL OBJECT (x, n, o)

Summary:

• For a non-immediate Scheme object x of unknown type, get the type information by

    using SCM_CELL_TYPE (x).

• As soon as the cell type information is available, only use the appropriate access methods to read and write data to the different cell entries.

9.3 A Virtual Machine for Guile

Guile has both an interpreter and a compiler. To a user, the difference is transparent— interpreted and compiled procedures can call each other as they please.

The difference is that the compiler creates and interprets bytecode for a custom virtual machine, instead of interpreting the S-expressions directly. Loading and running compiled code is faster than loading and running source code.

The virtual machine that does the bytecode interpretation is a part of Guile itself. This section describes the nature of Guile’s virtual machine.

9.3.1 Why a VM?

For a long time, Guile only had an interpreter. Guile’s interpreter operated directly on the S-expression representation of Scheme source code.

But while the interpreter was highly optimized and hand-tuned, it still performs many needless computations during the course of evaluating an expression. For example, application of a function to arguments needlessly consed up the arguments in a list. Evaluation of an expression always had to figure out what the car of the expression is – a procedure, a memoized form, or something else. All values have to be allocated on the heap. Et cetera.

The solution to this problem was to compile the higher-level language, Scheme, into a lower-level language for which all of the checks and dispatching have already been done—the code is instead stripped to the bare minimum needed to “do the job”.

The question becomes then, what low-level language to choose? There are many options. We could compile to native code directly, but that poses portability problems for Guile, as it is a highly cross-platform project.

So we want the performance gains that compilation provides, but we also want to maintain the portability benefits of a single code path. The obvious solution is to compile to a virtual machine that is present on all Guile installations.

The easiest (and most fun) way to depend on a virtual machine is to implement the virtual machine within Guile itself. This way the virtual machine provides what Scheme needs (tail calls, multiple values, call/cc) and can provide optimized inline instructions for Guile (cons, struct-ref, etc.).

So this is what Guile does. The rest of this section describes that VM that Guile implements, and the compiled procedures that run on it.

Before moving on, though, we should note that though we spoke of the interpreter in the past tense, Guile still has an interpreter. The difference is that before, it was Guile’s main evaluator, and so was implemented in highly optimized C; now, it is actually implemented in Scheme, and compiled down to VM bytecode, just like any other program. (There is still a C interpreter around, used to bootstrap the compiler, but it is not normally used at runtime.)

The upside of implementing the interpreter in Scheme is that we preserve tail calls and multiple-value handling between interpreted and compiled code. The downside is that the interpreter in Guile 2.0 is slower than the interpreter in 1.8. We hope the that the compiler’s speed makes up for the loss!

Also note that this decision to implement a bytecode compiler does not preclude native compilation. We can compile from bytecode to native code at runtime, or even do ahead of time compilation. More possibilities are discussed in Section 9.4.8 [Extending the Compiler], page 804.

9.3.2 VM Concepts

Compiled code is run by a virtual machine (VM). Each thread has its own VM. When a compiled procedure is run, Guile looks up the virtual machine for the current thread and executes the procedure using that VM.

Guile’s virtual machine is a stack machine—that is, it has few registers, and the instructions defined in the VM operate by pushing and popping values from a stack.

Stack memory is exclusive to the virtual machine that owns it. In addition to their stacks, virtual machines also have access to the global memory (modules, global bindings, etc) that is shared among other parts of Guile, including other VMs.

A VM has generic instructions, such as those to reference local variables, and instructions designed to support Guile’s languages – mathematical instructions that support the entire numerical tower, an inlined implementation of cons, etc.

The registers that a VM has are as follows:

• ip - Instruction pointer

• sp - Stack pointer

• fp - Frame pointer

In other architectures, the instruction pointer is sometimes called the “program counter” (pc). This set of registers is pretty typical for stack machines; their exact meanings in the context of Guile’s VM are described in the next section.

9.3.3 Stack Layout

While not strictly necessary to understand how to work with the VM, it is instructive and sometimes entertaining to consider the structure of the VM stack.

Logically speaking, a VM stack is composed of “frames”. Each frame corresponds to the application of one compiled procedure, and contains storage space for arguments, local variables, intermediate values, and some bookkeeping information (such as what to do after the frame computes its value).

While the compiler is free to do whatever it wants to, as long as the semantics of a computation are preserved, in practice every time you call a function, a new frame is created. (The notable exception of course is the tail call case, see Section 3.3.2 [Tail Calls], page 24.)

Within a frame, you have the data associated with the function application itself, which is of a fixed size, and the stack space for intermediate values. Sometimes only the former is referred to as the “frame”, and the latter is the “stack”, although all pending application frames can have some intermediate computations interleaved on the stack.

The structure of the fixed part of an application frame is as follows:

                  Stack
        | ...              |
        | Intermed. val. 0 | <- fp + bp->nargs + bp->nlocs = SCM_FRAME_UPPER_ADDRESS (fp)
        +==================+
        | Local variable 1 |
        | Local variable 0 | <- fp + bp->nargs
        | Argument 1       |
        | Argument 0       | <- fp
        | Program          | <- fp - 1
        +------------------+
        | Return address   |
        | MV return address|
        | Dynamic link     | <- fp - 4 = SCM_FRAME_DATA_ADDRESS (fp) = SCM_FRAME_LOWER_ADDRESS (fp)
        +==================+
        |                  |

In the above drawing, the stack grows upward. The intermediate values stored in the application of this frame are stored above SCM_FRAME_UPPER_ADDRESS (fp). bp refers to the struct scm_objcode data associated with the program at fp - 1. nargs and nlocs are properties of the compiled procedure, which will be discussed later.

The individual fields of the frame are as follows:

Return address The ip that was in effect before this program was applied. When we return from this activation frame, we will jump back to this ip.

MV return address The ip to return to if this application returns multiple values. For continuations that only accept one value, this value will be NULL; for others, it will be an ip that points to a multiple-value return address in the calling code. That code will expect the top value on the stack to be an integer—the number of values being returned—and that below that integer there are the values being returned.

Dynamic link This is the fp in effect before this program was applied. In effect, this and the return address are the registers that are always “saved”. The dynamic link links the current frame to the previous frame; computing a stack trace involves traversing these frames.

Local variable n Lambda-local variables that are all allocated as part of the frame. This makes access to variables very cheap.

Argument n The calling convention of the VM requires arguments of a function application to be pushed on the stack, and here they are. References to arguments dispatch to these locations on the stack.

Program This is the program being applied. For more information on how programs are implemented, See Section 9.3.5 [VM Programs], page 776.

9.3.4 Variables and the VM

Consider the following Scheme code as an example:

       (define (foo a)
         (lambda (b) (list foo a b)))

Within the lambda expression, foo is a top-level variable, a is a lexically captured variable, and b is a local variable.

Another way to refer to a and b is to say that a is a “free” variable, since it is not defined within the lambda, and b is a “bound” variable. These are the terms used in the lambda calculus, a mathematical notation for describing functions. The lambda calculus is useful because it allows one to prove statements about functions. It is especially good at describing scope relations, and it is for that reason that we mention it here.

Guile allocates all variables on the stack. When a lexically enclosed procedure with free variables—a closure—is created, it copies those variables into its free variable vector. References to free variables are then redirected through the free variable vector.

If a variable is ever set!, however, it will need to be heap-allocated instead of stackallocated, so that different closures that capture the same variable can see the same value. Also, this allows continuations to capture a reference to the variable, instead of to its value at one point in time. For these reasons, set! variables are allocated in “boxes”—actually, in variable cells. See Section 6.19.7 [Variables], page 390, for more information. References to set! variables are indirected through the boxes.

Thus perhaps counterintuitively, what would seem “closer to the metal”, viz set!, actually forces an extra memory allocation and indirection.

Going back to our example, b may be allocated on the stack, as it is never mutated. a may also be allocated on the stack, as it too is never mutated. Within the enclosed

lambda, its value will be copied into (and referenced from) the free variables vector. foo is a top-level variable, because foo is not lexically bound in this example.

9.3.5 Compiled Procedures are VM Programs

By default, when you enter in expressions at Guile’s REPL, they are first compiled to VM object code, then that VM object code is executed to produce a value. If the expression evaluates to a procedure, the result of this process is a compiled procedure.

A compiled procedure is a compound object, consisting of its bytecode, a reference to any captured lexical variables, an object array, and some metadata such as the procedure’s arity, name, and documentation. You can pick apart these pieces with the accessors in (system vm program). See Section 6.9.3 [Compiled Procedures], page 245, for a full API reference.

The object array of a compiled procedure, also known as the object table, holds all Scheme objects whose values are known not to change across invocations of the procedure: constant strings, symbols, etc. The object table of a program is initialized right before a program is loaded with load-program. See Section 9.3.6.8 [Loading Instructions], page 786, for more information.

Variable objects are one such type of constant object: when a global binding is defined, a variable object is associated to it and that object will remain constant over time, even if the value bound to it changes. Therefore, toplevel bindings only need to be looked up once. Thereafter, references to the corresponding toplevel variables from within the program are then performed via the toplevel-ref instruction, which uses the object vector, and are almost as fast as local variable references.

We can see how these concepts tie together by disassembling the foo function we defined earlier to see what is going on:

      scheme@(guile-user)> (define (foo a) (lambda (b) (list foo a b)))
      scheme@(guile-user)> ,x foo
         0    (assert-nargs-ee/locals 1)
         2    (object-ref 1)                  ;; #<procedure 8ebec20 at <current input>:0:17 (b)>
         4    (local-ref 0)                   ;; `a'
         6    (make-closure 0 1)
         9    (return)
      ----------------------------------------
      Disassembly of #<procedure 8ebec20 at <current input>:0:17 (b)>:
         0    (assert-nargs-ee/locals 1)
         2    (toplevel-ref 1)                ;; `foo'
         4    (free-ref 0)                    ;; (closure variable)
         6    (local-ref 0)                   ;; `b'
         8    (list 0 3)                      ;; 3 elements         at (unknown file):0:29
        11    (return)

First there’s some prelude, where foo checks that it was called with only 1 argument. Then at ip 2, we load up the compiled lambda. Ip 4 loads up ‘a’, so that it can be captured into a closure by at ip 6—binding code (from the compiled lambda) with data (the free-variable vector). Finally we return the closure.

The second stanza disassembles the compiled lambda. After the prelude, we note that toplevel variables are resolved relative to the module that was current when the procedure was created. This lookup occurs lazily, at the first time the variable is actually referenced, and the location of the lookup is cached so that future references are very cheap. See Section 9.3.6.2 [Top-Level Environment Instructions], page 779, for more details.

Then we see a reference to a free variable, corresponding to a. The disassembler doesn’t have enough information to give a name to that variable, so it just marks it as being a “closure variable”. Finally we see the reference to b, then the list opcode, an inline implementation of the list scheme routine.

9.3.6 Instruction Set

There are about 180 instructions in Guile’s virtual machine. These instructions represent atomic units of a program’s execution. Ideally, they perform one task without conditional branches, then dispatch to the next instruction in the stream.

Instructions themselves are one byte long. Some instructions take parameters, which follow the instruction byte in the instruction stream.

Sometimes the compiler can figure out that it is compiling a special case that can be run more efficiently. So, for example, while Guile offers a generic test-and-branch instruction, it also offers specific instructions for special cases, so that the following cases all have their own test-and-branch instructions:

     (if pred then else)
     (if (not pred) then else)
     (if (null? l) then else)
     (if (not (null? l)) then else)

In addition, some Scheme primitives have their own inline implementations, e.g. cons, and list, as we saw in the previous section.

So Guile’s instruction set is a complete instruction set, in that it provides the instructions that are suited to the problem, and is not concerned with making a minimal, orthogonal set of instructions. More instructions may be added over time.

9.3.6.1 Lexical Environment Instructions

These instructions access and mutate the lexical environment of a compiled procedure—its free and bound variables.

Some of these instructions have long- variants, the difference being that they take 16-bit arguments, encoded in big-endianness, instead of the normal 8-bit range.

See Section 9.3.3 [Stack Layout], page 774, for more information on the format of stack frames.

[Instruction] local-ref index [Instruction] long-local-ref index

Push onto the stack the value of the local variable located at index within the current stack frame. Note that arguments and local variables are all in one block. Thus the first argument, if any, is at index 0, and local bindings follow the arguments.

[Instruction] local-set index [Instruction] long-local-set index

Pop the Scheme object located on top of the stack and make it the new value of the local variable located at index within the current stack frame.

[Instruction] box index Pop a value off the stack, and set the indexnth local variable to a box containing that value. A shortcut for make-variable then local-set, used when binding boxed variables.

[Instruction] empty-box index

Set the indexth local variable to a box containing a variable whose value is unbound. Used when compiling some letrec expressions.

[Instruction] local-boxed-ref index [Instruction] local-boxed-set index

Get or set the value of the variable located at index within the current stack frame.

A shortcut for local-ref then variable-ref or variable-set, respectively.

[Instruction] free-ref index

Push the value of the captured variable located at position index within the program’s vector of captured variables.

[Instruction] free-boxed-ref index [Instruction] free-boxed-set index

Get or set a boxed free variable. A shortcut for free-ref then variable-ref or variable-set, respectively. Note that there is no free-set instruction, as variables that are set! must be boxed.

[Instruction] make-closure num-free-vars Pop num-free-vars values and a program object off the stack in that order, and push a new program object closing over the given free variables. num-free-vars is encoded as a two-byte big-endian value. The free variables are stored in an array, inline to the new program object, in the order that they were on the stack (not the order they are popped off). The new closure shares state with the original program. At the time of this writing, the space overhead of closures is 3 words, plus one word for each free variable.

[Instruction] fix-closure index

Fix up the free variables array of the closure stored in the indexth local variable. index is a two-byte big-endian integer. This instruction will pop as many values from the stack as are in the corresponding closure’s free variables array. The topmost value on the stack will be stored as the closure’s last free variable, with other values filling in free variable slots in order. fix-closure is part of a hack for allocating mutually recursive procedures. The hack is to store the procedures in their corresponding local variable slots, with space already allocated for free variables. Then once they are all in place, this instruction fixes up their procedures’ free variable bindings in place. This allows most letrec-bound procedures to be allocated unboxed on the stack.

[Instruction] local-bound? index [Instruction] long-local-bound? index

Push #t on the stack if the indexth local variable has been assigned, or #f otherwise. Mostly useful for handling optional arguments in procedure prologues.

9.3.6.2 Top-Level Environment Instructions

These instructions access values in the top-level environment: bindings that were not lexically apparent at the time that the code in question was compiled.

The location in which a toplevel binding is stored can be looked up once and cached for later. The binding itself may change over time, but its location will stay constant.

Currently only toplevel references within procedures are cached, as only procedures have a place to cache them, in their object tables.

[Instruction] toplevel-ref index [Instruction] long-toplevel-ref index

Push the value of the toplevel binding whose location is stored in at position index in the current procedure’s object table. The long- variant encodes the index over two bytes. Initially, a cell in a procedure’s object table that is used by toplevel-ref is initialized to one of two forms. The normal case is that the cell holds a symbol, whose binding will be looked up relative to the module that was current when the current program was created. Alternately, the lookup may be performed relative to a particular module, determined at compile-time (e.g. via @ or @@). In that case, the cell in the object table holds a list: (modname sym public?). The symbol sym will be looked up in the module named modname (a list of symbols). The lookup will be performed against the module’s public interface, unless public? is #f, which it is for example when compiling @@. In any case, if the symbol is unbound, an error is signalled. Otherwise the initial form is replaced with the looked-up variable, an in-place mutation of the object table. This mechanism provides for lazy variable resolution, and an important cached fast-path once the variable has been successfully resolved. This instruction pushes the value of the variable onto the stack.

[Instruction] toplevel-set index [Instruction] long-toplevel-set index

Pop a value off the stack, and set it as the value of the toplevel variable stored at index in the object table. If the variable has not yet been looked up, we do the lookup

as in toplevel-ref.

[Instruction] define

Pop a symbol and a value from the stack, in that order. Look up its binding in the current toplevel environment, creating the binding if necessary. Set the variable to the value.

[Instruction] link-now

Pop a value, x, from the stack. Look up the binding for x, according to the rules for toplevel-ref, and push that variable on the stack. If the lookup fails, an error will be signalled.

This instruction is mostly used when loading programs, because it can do toplevel variable lookups without an object table.

[Instruction] variable-ref

Dereference the variable object which is on top of the stack and replace it by the value of the variable it represents.

[Instruction] variable-set

Pop off two objects from the stack, a variable and a value, and set the variable to the value.

[Instruction] variable-bound?

Pop off the variable object from top of the stack and push #t if it is bound, or #f otherwise. Mostly useful in procedure prologues for defining default values for boxed optional variables.

[Instruction] make-variable

Replace the top object on the stack with a variable containing it. Used in some circumstances when compiling letrec expressions.

9.3.6.3 Procedure Call and Return Instructions

[Instruction] new-frame

Push a new frame on the stack, reserving space for the dynamic link, return address, and the multiple-values return address. The frame pointer is not yet updated, because the frame is not yet active – it has to be patched by a call instruction to get the return address.

[Instruction] call nargs Call the procedure located at sp[-nargs] with the nargs arguments located from

sp[-nargs + 1] to sp[0].

This instruction requires that a new frame be pushed on the stack before the procedure, via new-frame. See Section 9.3.3 [Stack Layout], page 774, for more information. It patches up that frame with the current ip as the return address, then dispatches to the first instruction in the called procedure, relying on the called procedure to return one value to the newly-created continuation. Because the new frame pointer will point to sp[-nargs + 1], the arguments don’t have to be shuffled around – they are already in place.

[Instruction] tail-call nargs

Transfer control to the procedure located at sp[-nargs] with the nargs arguments

located from sp[-nargs + 1] to sp[0].

Unlike call, which requires a new frame to be pushed onto the stack, tail-call simply shuffles down the procedure and arguments to the current stack frame. This instruction implements tail calls as required by RnRS.

[Instruction] apply nargs [Instruction] tail-apply nargs

Like call and tail-call, except that the top item on the stack must be a list. The elements of that list are then pushed on the stack and treated as additional arguments, replacing the list itself, then the procedure is invoked as usual.

[Instruction] call/nargs

[Instruction] tail-call/nargs

These are like call and tail-call, except they take the number of arguments from the stack instead of the instruction stream. These instructions are used in the implementation of multiple value returns, where the actual number of values is pushed on the stack.

[Instruction] mv-call nargs offset Like call, except that a multiple-value continuation is created in addition to a singlevalue continuation. The offset (a three-byte value) is an offset within the instruction stream; the multiplevalue return address in the new frame (see Section 9.3.3 [Stack Layout], page 774) will be set to the normal return address plus this offset. Instructions at that offset will expect the top value of the stack to be the number of values, and below that values themselves, pushed separately.

[Instruction] return

Free the program’s frame, returning the top value from the stack to the current continuation. (The stack should have exactly one value on it.) Specifically, the sp is decremented to one below the current fp, the ip is reset to the current return address, the fp is reset to the value of the current dynamic link, and then the returned value is pushed on the stack.

[Instruction] return/values nvalues

[Instruction] return/nvalues

Return the top nvalues to the current continuation. In the case of return/nvalues, nvalues itself is first popped from the top of the stack. If the current continuation is a multiple-value continuation, return/values pushes the number of values on the stack, then returns as in return, but to the multiple-value return address. Otherwise if the current continuation accepts only one value, i.e. the multiple-value return address is NULL, then we assume the user only wants one value, and we give them the first one. If there are no values, an error is signaled.

[Instruction] return/values* nvalues

Like a combination of apply and return/values, in which the top value on the stack is interpreted as a list of additional values. This is an optimization for the common

(apply values ...) case.

[Instruction] truncate-values nbinds nrest

Used in multiple-value continuations, this instruction takes the values that are on the stack (including the number-of-values marker) and truncates them for a binding construct. For example, a call to (receive (x y . z) (foo) ...) would, logically speaking, pop off the values returned from (foo) and push them as three values, corresponding to x, y, and z. In that case, nbinds would be 3, and nrest would be 1 (to indicate that one of the bindings was a rest argument). Signals an error if there is an insufficient number of values.

[Instruction] call/cc [Instruction] tail-call/cc

Capture the current continuation, and then call (or tail-call) the procedure on the top of the stack, with the continuation as the argument. call/cc does not require a new-frame to be pushed on the stack, as call does, because it needs to capture the stack before the frame is pushed. Both the VM continuation and the C continuation are captured.

9.3.6.4 Function Prologue Instructions

A function call in Guile is very cheap: the VM simply hands control to the procedure. The procedure itself is responsible for asserting that it has been passed an appropriate number of arguments. This strategy allows arbitrarily complex argument parsing idioms to be developed, without harming the common case.

For example, only calls to keyword-argument procedures “pay” for the cost of parsing keyword arguments. (At the time of this writing, calling procedures with keyword arguments is typically two to four times as costly as calling procedures with a fixed set of arguments.)

[Instruction] assert-nargs-ee n
[Instruction] assert-nargs-ge n

Assert that the current procedure has been passed exactly n arguments, for the -ee case, or n or more arguments, for the -ge case. n is encoded over two bytes. The number of arguments is determined by subtracting the frame pointer from the stack pointer (sp - (fp -1)). See Section 9.3.3 [Stack Layout], page 774, for more details on stack frames.

[Instruction] br-if-nargs-ne n offset [Instruction] br-if-nargs-gt n offset [Instruction] br-if-nargs-lt n offset

Jump to offset if the number of arguments is not equal to, greater than, or less than n. n is encoded over two bytes, and offset has the normal three-byte encoding. These instructions are used to implement multiple arities, as in case-lambda. See Section 6.9.5 [Case-lambda], page 251, for more information.

[Instruction] bind-optionals n

If the procedure has been called with fewer than n arguments, fill in the remaining arguments with an unbound value (SCM_UNDEFINED). n is encoded over two bytes. The optionals can be later initialized conditionally via the local-bound? instruction.

[Instruction] push-rest n

Pop off excess arguments (more than n), collecting them into a list, and push that list. Used to bind a rest argument, if the procedure has no keyword arguments. Procedures with keyword arguments use bind-rest instead.

[Instruction] bind-rest n idx

Pop off excess arguments (more than n), collecting them into a list. The list is then assigned to the idxth local variable.

[Instruction] bind-optionals/shuffle nreq nreq-and-opt ntotal [Instruction] bind-optionals/shuffle-or-br nreq nreq-and-opt ntotal offset

Shuffle keyword arguments to the top of the stack, filling in the holes with SCM_ UNDEFINED. Each argument is encoded over two bytes.

This instruction is used by procedures with keyword arguments. nreq is the number of required arguments to the procedure, and nreq-and-opt is the total number of positional arguments (required plus optional). bind-optionals/shuffle will scan the stack from the nreqth argument up to the nreq-and-optth, and start shuffling when it sees the first keyword argument or runs out of positional arguments.

bind-optionals/shuffle-or-br does the same, except that it checks if there are too many positional arguments before shuffling. If this is the case, it jumps to offset, encoded using the normal three-byte encoding.

Shuffling simply moves the keyword arguments past the total number of arguments, ntotal, which includes keyword and rest arguments. The free slots created by the shuffle are filled in with SCM_UNDEFINED, so they may be conditionally initialized later in the function’s prologue.

[Instruction] bind-kwargs idx ntotal flags Parse keyword arguments, assigning their values to the corresponding local variables. The keyword arguments should already have been shuffled above the ntotalth stack

slot by bind-optionals/shuffle.

The parsing is driven by a keyword arguments association list, looked up from the idxth element of the procedures object array. The alist is a list of pairs of the form (kw . index), mapping keyword arguments to their local variable indices.

There are two bitflags that affect the parser, allow-other-keys? (0x1) and rest? (0x2). Unless allow-other-keys? is set, the parser will signal an error if an unknown key is found. If rest? is set, errors parsing the keyword arguments will be ignored, as a later bind-rest instruction will collect all of the tail arguments, including the keywords, into a list. Otherwise if the keyword arguments are invalid, an error is signalled.

idx and ntotal are encoded over two bytes each, and flags is encoded over one byte.

[Instruction] reserve-locals n

Resets the stack pointer to have space for n local variables, including the arguments. If this operation increments the stack pointer, as in a push, the new slots are filled with SCM_UNBOUND. If this operation decrements the stack pointer, any excess values are dropped.

reserve-locals is typically used after argument parsing to reserve space for local variables.

[Instruction] assert-nargs-ee/locals n
[Instruction] assert-nargs-ge/locals n

A combination of assert-nargs-ee and reserve-locals. The number of arguments is encoded in the lower three bits of n, a one-byte value. The number of additional local variables is take from the upper 5 bits of n.

9.3.6.5 Trampoline Instructions

Though most applicable objects in Guile are procedures implemented in bytecode, not all are. There are primitives, continuations, and other procedure-like objects that have their own calling convention. Instead of adding special cases to the call instruction, Guile wraps these other applicable objects in VM trampoline procedures, then provides special support for these objects in bytecode.

Trampoline procedures are typically generated by Guile at runtime, for example in response to a call to scm_c_make_gsubr. As such, a compiler probably shouldn’t emit code with these instructions. However, it’s still interesting to know how these things work, so we document these trampoline instructions here.

[Instruction] subr-call nargs

Pop off a foreign pointer (which should have been pushed on by the trampoline), and call it directly, with the nargs arguments from the stack. Return the resulting value or values to the calling procedure.

[Instruction] foreign-call nargs

Pop off an internal foreign object (which should have been pushed on by the trampoline), and call that foreign function with the nargs arguments from the stack. Return the resulting value to the calling procedure.

[Instruction] continuation-call

Pop off an internal continuation object (which should have been pushed on by the trampoline), and reinstate that continuation. All of the procedure’s arguments are passed to the continuation. Does not return.

[Instruction] partial-cont-call

Pop off two objects from the stack: the dynamic winds associated with the partial continuation, and the VM continuation object. Unroll the continuation onto the stack, rewinding the dynamic environment and overwriting the current frame, and pass all arguments to the continuation. Control flow proceeds where the continuation was captured.

9.3.6.6 Branch Instructions

All the conditional branch instructions described below work in the same way: • They pop off Scheme object(s) located on the stack for use in the branch condition • If the condition is true, then the instruction pointer is increased by the offset passed

as an argument to the branch instruction;

• Program execution proceeds with the next instruction (that is, the one to which the instruction pointer points).

Note that the offset passed to the instruction is encoded as three 8-bit integers, in bigendian order, effectively giving Guile a 24-bit relative address space.

[Instruction] br offset Jump to offset. No values are popped.

[Instruction] br-if offset

Jump to offset if the object on the stack is not false.

[Instruction] br-if-not offset

Jump to offset if the object on the stack is false.

[Instruction] br-if-eq offset

Jump to offset if the two objects located on the stack are equal in the sense of eq?. Note that, for this instruction, the stack pointer is decremented by two Scheme objects instead of only one.

[Instruction] br-if-not-eq offset

Same as br-if-eq for non-eq? objects.

[Instruction] br-if-null offset

Jump to offset if the object on the stack is '().

[Instruction] br-if-not-null offset

Jump to offset if the object on the stack is not '().

9.3.6.7 Data Constructor Instructions

These instructions push simple immediate values onto the stack, or construct compound data structures from values on the stack.

[Instruction] make-int8 value

Push value, an 8-bit integer, onto the stack.

[Instruction] make-int8:0

Push the immediate value 0 onto the stack.

[Instruction] make-int8:1

Push the immediate value 1 onto the stack.

[Instruction] make-int16 value

Push value, a 16-bit integer, onto the stack.

[Instruction] make-uint64 value

Push value, an unsigned 64-bit integer, onto the stack. The value is encoded in 8 bytes, most significant byte first (big-endian).

[Instruction] make-int64 value

Push value, a signed 64-bit integer, onto the stack. The value is encoded in 8 bytes, most significant byte first (big-endian), in twos-complement arithmetic.

[Instruction] make-false

Push #f onto the stack.

[Instruction] make-true

Push #t onto the stack.

[Instruction] make-nil

Push #nil onto the stack.

[Instruction] make-eol

Push '() onto the stack.

[Instruction] make-char8 value

Push value, an 8-bit character, onto the stack.

[Instruction] make-char32 value

Push value, an 32-bit character, onto the stack. The value is encoded in big-endian order.

[Instruction] make-symbol

Pops a string off the stack, and pushes a symbol.

[Instruction] make-keyword value

Pops a symbol off the stack, and pushes a keyword.

[Instruction] list n

Pops off the top n values off of the stack, consing them up into a list, then pushes that list on the stack. What was the topmost value will be the last element in the list. n is a two-byte value, most significant byte first.

[Instruction] vector n

Create and fill a vector with the top n values from the stack, popping off those values and pushing on the resulting vector. n is a two-byte value, like in vector.

[Instruction] make-struct n

Make a new struct from the top n values on the stack. The values are popped, and the new struct is pushed. The deepest value is used as the vtable for the struct, and the rest are used in order as the field initializers. Tail arrays are not supported by this instruction.

[Instruction] make-array n

Pop an array shape from the stack, then pop the remaining n values, pushing a new array. n is encoded over three bytes. The array shape should be appropriate to store n values. See Section 6.7.5.2 [Array Procedures], page 202, for more information on array shapes.

Many of these data structures are constant, never changing over the course of the different invocations of the procedure. In that case it is often advantageous to make them once when the procedure is created, and just reference them from the object table thereafter. See Section 9.3.4 [Variables and the VM], page 775, for more information on the object table.

[Instruction] object-ref n

[Instruction] long-object-ref n

Push nth value from the current program’s object vector. The “long” variant has a 16-bit index instead of an 8-bit index.

9.3.6.8 Loading Instructions

In addition to VM instructions, an instruction stream may contain variable-length data embedded within it. This data is always preceded by special loading instructions, which interpret the data and advance the instruction pointer to the next VM instruction.

All of these loading instructions have a length parameter, indicating the size of the embedded data, in bytes. The length itself is encoded in 3 bytes.

[Instruction] load-number length

Load an arbitrary number from the instruction stream. The number is embedded in the stream as a string.

[Instruction] load-string length

Load a string from the instruction stream. The string is assumed to be encoded in the “latin1” locale.

[Instruction] load-wide-string length

Load a UTF-32 string from the instruction stream. length is the length in bytes, not in codepoints.

[Instruction] load-symbol length

Load a symbol from the instruction stream. The symbol is assumed to be encoded in the “latin1” locale. Symbols backed by wide strings may be loaded via load-wide-

string then make-symbol.

[Instruction] load-array length

Load a uniform array from the instruction stream. The shape and type of the array are popped off the stack, in that order.

[Instruction] load-program

Load bytecode from the instruction stream, and push a compiled procedure. This instruction pops one value from the stack: the program’s object table, as a vector, or #f in the case that the program has no object table. A program that does not reference toplevel bindings and does not use object-ref does not need an object table. This instruction is unlike the rest of the loading instructions, because instead of parsing its data, it directly maps the instruction stream onto a C structure, struct scm_objcode. See Section 9.4.6 [Bytecode and Objcode], page 803, for more information. The resulting compiled procedure will not have any free variables captured, so it may be loaded only once but used many times to create closures.

9.3.6.9 Dynamic Environment Instructions

Guile’s virtual machine has low-level support for dynamic-wind, dynamic binding, and composable prompts and aborts.

[Instruction] wind

Pop an unwind thunk and a wind thunk from the stack, in that order, and push them onto the “dynamic stack”. The unwind thunk will be called on nonlocal exits, and the wind thunk on reentries. Used to implement dynamic-wind. Note that neither thunk is actually called; the compiler should emit calls to wind and unwind for the normal dynamic-wind control flow. See Section 6.13.10 [Dynamic Wind], page 309.

[Instruction] unwind

Pop off the top entry from the “dynamic stack”, for example, a wind/unwind thunk pair. unwind instructions should be properly paired with their winding instructions, like wind.

[Instruction] wind-fluids n

Pop off n values and n fluids from the stack, in that order. Set the fluids to the values by creating a with-fluids object and pushing that object on the dynamic stack. See Section 6.21.7 [Fluids and Dynamic States], page 419.

[Instruction] unwind-fluids

Pop a with-fluids object from the dynamic stack, and swap the current values of its fluids with the saved values of its fluids. In this way, the dynamic environment is left as it was before the corresponding wind-fluids instruction was processed.

[Instruction] fluid-ref

Pop a fluid from the stack, and push its current value.

[Instruction] fluid-set

Pop a value and a fluid from the stack, in that order, and set the fluid to the value.

[Instruction] prompt escape-only? offset Establish a dynamic prompt. See Section 6.13.5 [Prompts], page 296, for more information on prompts.

The prompt will be pushed on the dynamic stack. The normal control flow should ensure that the prompt is popped off at the end, via unwind.

If an abort is made to this prompt, control will jump to offset, a three-byte relative address. The continuation and all arguments to the abort will be pushed on the stack, along with the total number of arguments (including the continuation. If control returns to the handler, the prompt is already popped off by the abort mechanism. (Guile’s prompt implements Felleisen’s –F– operator.)

If escape-only? is nonzero, the prompt will be marked as escape-only, which allows an abort to this prompt to avoid reifying the continuation.

[Instruction] abort n

Abort to a dynamic prompt.

This instruction pops one tail argument list, n arguments, and a prompt tag from the stack. The dynamic environment is then searched for a prompt having the given tag. If none is found, an error is signalled. Otherwise all arguments are passed to the prompt’s handler, along with the captured continuation, if necessary.

If the prompt’s handler can be proven to not reference the captured continuation, no continuation is allocated. This decision happens dynamically, at run-time; the general case is that the continuation may be captured, and thus resumed. A reinstated continuation will have its arguments pushed on the stack, along with the number of arguments, as in the multiple-value return convention. Therefore an abort instruction should be followed by code ready to handle the equivalent of a multiply-valued return.

9.3.6.10 Miscellaneous Instructions

[Instruction] nop

Does nothing! Used for padding other instructions to certain alignments.

[Instruction] halt

Exits the VM, returning a SCM value. Normally, this instruction is only part of the “bootstrap program”, a program run when a virtual machine is first entered; compiled Scheme procedures will not contain this instruction. If multiple values have been returned, the SCM value will be a multiple-values object (see Section 6.13.7 [Multiple Values], page 301).

[Instruction] break

Does nothing, but invokes the break hook.

[Instruction] drop

Pops off the top value from the stack, throwing it away.

[Instruction] dup

Re-pushes the top value onto the stack.

[Instruction] void

Pushes “the unspecified value” onto the stack.

9.3.6.11 Inlined Scheme Instructions

The Scheme compiler can recognize the application of standard Scheme procedures. It tries to inline these small operations to avoid the overhead of creating new stack frames.

Since most of these operations are historically implemented as C primitives, not inlining them would entail constantly calling out from the VM to the interpreter, which has some costs—registers must be saved, the interpreter has to dispatch, called procedures have to do much type checking, etc. It’s much more efficient to inline these operations in the virtual machine itself.

All of these instructions pop their arguments from the stack and push their results, and take no parameters from the instruction stream. Thus, unlike in the previous sections, these instruction definitions show stack parameters instead of parameters from the instruction stream.

[Instruction] not x [Instruction] not-not x [Instruction] eq? x y [Instruction] not-eq? x y [Instruction] null? [Instruction] not-null? [Instruction] eqv? x y [Instruction] equal? x y [Instruction] pair? x y [Instruction] list? x [Instruction] set-car! pair x [Instruction] set-cdr! pair x [Instruction] cons x y [Instruction] car x [Instruction] cdr x [Instruction] vector-ref x y [Instruction] vector-set x n y [Instruction] struct? x [Instruction] struct-ref x n [Instruction] struct-set x n v [Instruction] struct-vtable x [Instruction] class-of x [Instruction] slot-ref struct n [Instruction] slot-set struct n x

Inlined implementations of their Scheme equivalents.

Note that caddr and friends compile to a series of car and cdr instructions.

9.3.6.12 Inlined Mathematical Instructions

Inlining mathematical operations has the obvious advantage of handling fixnums without function calls or allocations. The trick, of course, is knowing when the result of an operation will be a fixnum, and there might be a couple bugs here.

More instructions could be added here over time.

As in the previous section, the definitions below show stack parameters instead of instruction stream parameters.

[Instruction] add x y [Instruction] add1 x [Instruction] sub x y [Instruction] sub1 x [Instruction] mul x y [Instruction] div x y [Instruction] quo x y [Instruction] rem x y [Instruction] mod x y [Instruction] ee? x y [Instruction] lt? x y [Instruction] gt? x y [Instruction] le? x y [Instruction] ge? x y [Instruction] ash x n [Instruction] logand x y [Instruction] logior x y [Instruction] logxor x y

Inlined implementations of the corresponding mathematical operations.

9.3.6.13 Inlined Bytevector Instructions

Bytevector operations correspond closely to what the current hardware can do, so it makes sense to inline them to VM instructions, providing a clear path for eventual native compilation. Without this, Scheme programs would need other primitives for accessing raw bytes – but these primitives are as good as any.

As in the previous section, the definitions below show stack parameters instead of instruction stream parameters.

The multibyte formats (u16, f64, etc) take an extra endianness argument. Only aligned native accesses are currently fast-pathed in Guile’s VM.

[Instruction] bv-u8-ref bv n [Instruction] bv-s8-ref bv n

[Instruction] bv-u16-native-ref bv n
[Instruction] bv-s16-native-ref bv n
[Instruction] bv-u32-native-ref bv n
[Instruction] bv-s32-native-ref bv n
[Instruction] bv-u64-native-ref bv n
[Instruction] bv-s64-native-ref bv n
[Instruction] bv-f32-native-ref bv n
[Instruction] bv-f64-native-ref bv n

[Instruction] bv-u16-ref bv n endianness [Instruction] bv-s16-ref bv n endianness [Instruction] bv-u32-ref bv n endianness [Instruction] bv-s32-ref bv n endianness [Instruction] bv-u64-ref bv n endianness [Instruction] bv-s64-ref bv n endianness [Instruction] bv-f32-ref bv n endianness [Instruction] bv-f64-ref bv n endianness

[Instruction] bv-u8-set bv n val [Instruction] bv-s8-set bv n val [Instruction] bv-u16-native-set bv n val [Instruction] bv-s16-native-set bv n val [Instruction] bv-u32-native-set bv n val [Instruction] bv-s32-native-set bv n val [Instruction] bv-u64-native-set bv n val [Instruction] bv-s64-native-set bv n val [Instruction] bv-f32-native-set bv n val [Instruction] bv-f64-native-set bv n val

[Instruction] bv-u16-set bv n val endianness [Instruction] bv-s16-set bv n val endianness [Instruction] bv-u32-set bv n val endianness [Instruction] bv-s32-set bv n val endianness [Instruction] bv-u64-set bv n val endianness [Instruction] bv-s64-set bv n val endianness [Instruction] bv-f32-set bv n val endianness [Instruction] bv-f64-set bv n val endianness

Inlined implementations of the corresponding bytevector operations.

9.4 Compiling to the Virtual Machine

Compilers have a mystique about them that is attractive and off-putting at the same time. They are attractive because they are magical – they transform inert text into live results, like throwing the switch on Frankenstein’s monster. However, this magic is perceived by many to be impenetrable.

This section aims to pay attention to the small man behind the curtain. See Section 6.17 [Read/Load/Eval/Compile], page 357, if you’re lost and you just wanted

to know how to compile your .scm file.

9.4.1 Compiler Tower

Guile’s compiler is quite simple, actually – its compilers, to put it more accurately. Guile defines a tower of languages, starting at Scheme and progressively simplifying down to languages that resemble the VM instruction set (see Section 9.3.6 [Instruction Set], page 777).

Each language knows how to compile to the next, so each step is simple and understandable. Furthermore, this set of languages is not hardcoded into Guile, so it is possible for the user to add new high-level languages, new passes, or even different compilation targets.

Languages are registered in the module, (system base language):

(use-modules (system base language))

They are registered with the define-language form.

[Scheme Syntax] define-language [#:name] [#:title] [#:reader] [#:printer] [#:parser=#f] [#:compilers=’()] [#:decompilers=’()] [#:evaluator=#f] [#:joiner=#f] [#:for-humans?=#t] [#:make-default-environment=make-fresh-user-module]

Define a language. This syntax defines a #<language> object, bound to name in the current environment. In addition, the language will be added to the global language set. For example, this is the language definition for Scheme:

     (define-language scheme
       #:title "Scheme"
       #:reader      (lambda (port env) ...)
       #:compilers   `((tree-il . ,compile-tree-il))
       #:decompilers `((tree-il . ,decompile-tree-il))
       #:evaluator (lambda (x module) (primitive-eval x))
       #:printer write
       #:make-default-environment (lambda () ...))

The interesting thing about having languages defined this way is that they present a uniform interface to the read-eval-print loop. This allows the user to change the current language of the REPL:

scheme@(guile-user)> ,language tree-il
Happy hacking with Tree Intermediate Language!  To switch back, type `,L scheme'.
tree-il@(guile-user)> ,L scheme
Happy hacking with Scheme!  To switch back, type `,L tree-il'.
scheme@(guile-user)>

Languages can be looked up by name, as they were above.

[Scheme Procedure] lookup-language name

Looks up a language named name, autoloading it if necessary. Languages are autoloaded by looking for a variable named name in a module named

(language name spec).

The language object will be returned, or #f if there does not exist a language with that name.

Defining languages this way allows us to programmatically determine the necessary steps for compiling code from one language to another.

[Scheme Procedure] lookup-compilation-order from to

Recursively traverses the set of languages to which from can compile, depth-first, and return the first path that can transform from to to. Returns #f if no path is found. This function memoizes its results in a cache that is invalidated by subsequent calls to define-language, so it should be quite fast.

There is a notion of a “current language”, which is maintained in the current-language parameter, defined in the core (guile) module. This language is normally Scheme, and may be rebound by the user. The run-time compilation interfaces (see Section 6.17 [Read/Load/Eval/Compile], page 357) also allow you to choose other source and target languages.

The normal tower of languages when compiling Scheme goes like this:

• Scheme

• Tree Intermediate Language (Tree-IL)

• Guile Lowlevel Intermediate Language (GLIL)

• Assembly

• Bytecode

• Objcode

Object code may be serialized to disk directly, though it has a cookie and version prepended to the front. But when compiling Scheme at run time, you want a Scheme value: for example, a compiled procedure. For this reason, so as not to break the abstraction, Guile defines a fake language at the bottom of the tower:

• Value

Compiling to value loads the object code into a procedure, and wakes the sleeping giant. Perhaps this strangeness can be explained by example: compile-file defaults to com-

piling to object code, because it produces object code that has to live in the barren world outside the Guile runtime; but compile defaults to compiling to value, as its product re-enters the Guile world.

Indeed, the process of compilation can circulate through these different worlds indefinitely, as shown by the following quine:

     ((lambda (x) ((compile x) x)) '(lambda (x) ((compile x) x)))

9.4.2 The Scheme Compiler

The job of the Scheme compiler is to expand all macros and all of Scheme to its most primitive expressions. The definition of “primitive” is given by the inventory of constructs provided by Tree-IL, the target language of the Scheme compiler: procedure applications, conditionals, lexical references, etc. This is described more fully in the next section.

The tricky and amusing thing about the Scheme-to-Tree-IL compiler is that it is completely implemented by the macro expander. Since the macro expander has to run over all of the source code already in order to expand macros, it might as well do the analysis at the same time, producing Tree-IL expressions directly.

Because this compiler is actually the macro expander, it is extensible. Any macro which the user writes becomes part of the compiler.

The Scheme-to-Tree-IL expander may be invoked using the generic compile procedure:

     (compile '(+ 1 2) #:from 'scheme #:to 'tree-il)

⇒

      #<<application> src: #f
                      proc: #<<toplevel-ref> src: #f name: +>
                      args: (#<<const> src: #f exp: 1>
                             #<<const> src: #f exp: 2>)>

Or, since Tree-IL is so close to Scheme, it is often useful to expand Scheme to Tree-IL, then translate back to Scheme. For that reason the expander provides two interfaces. The former is equivalent to calling (macroexpand '(+ 1 2) 'c), where the 'c is for “compile”. With 'e (the default), the result is translated back to Scheme:

     (macroexpand '(+ 1 2))
     ⇒ (+ 1 2)
     (macroexpand '(let ((x 10)) (* x x)))
     ⇒ (let ((x84 10)) (* x84 x84))

The second example shows that as part of its job, the macro expander renames lexicallybound variables. The original names are preserved when compiling to Tree-IL, but can’t be represented in Scheme: a lexical binding only has one name. It is for this reason that the native output of the expander is not Scheme. There’s too much information we would lose if we translated to Scheme directly: lexical variable names, source locations, and module hygiene.

Note however that macroexpand does not have the same signature as compile-tree-il. compile-tree-il is a small wrapper around macroexpand, to make it conform to the general form of compiler procedures in Guile’s language tower.

Compiler procedures take three arguments: an expression, an environment, and a keyword list of options. They return three values: the compiled expression, the corresponding environment for the target language, and a “continuation environment”. The compiled expression and environment will serve as input to the next language’s compiler. The “continuation environment” can be used to compile another expression from the same source language within the same module.

For example, you might compile the expression, (define-module (foo)). This will result in a Tree-IL expression and environment. But if you compiled a second expression, you would want to take into account the compile-time effect of compiling the previous expression, which puts the user in the (foo) module. That is purpose of the “continuation environment”; you would pass it as the environment when compiling the subsequent expression.

For Scheme, an environment is a module. By default, the compile and compile-file procedures compile in a fresh module, such that bindings and macros introduced by the expression being compiled are isolated:

     (eq? (current-module) (compile '(current-module)))
     ⇒ #f
     (compile '(define hello 'world))
     (defined? 'hello)
     ⇒ #f
     (define / *)
     (eq? (compile '/) /)
     ⇒ #f

Similarly, changes to the current-reader fluid (see Section 6.17.6 [Loading], page 366) are isolated:

     (compile '(fluid-set! current-reader (lambda args 'fail)))
     (fluid-ref current-reader)
     ⇒ #f

Nevertheless, having the compiler and compilee share the same name space can be achieved by explicitly passing (current-module) as the compilation environment:

     (define hello 'world)
     (compile 'hello #:env (current-module))
     ⇒ world

9.4.3 Tree-IL

Tree Intermediate Language (Tree-IL) is a structured intermediate language that is close in expressive power to Scheme. It is an expanded, pre-analyzed Scheme.

Tree-IL is “structured” in the sense that its representation is based on records, not S-expressions. This gives a rigidity to the language that ensures that compiling to a lowerlevel language only requires a limited set of transformations. For example, the Tree-IL type <const> is a record type with two fields, src and exp. Instances of this type are created via make-const. Fields of this type are accessed via the const-src and const-exp procedures. There is also a predicate, const?. See Section 6.7.9 [Records], page 219, for more information on records.

All Tree-IL types have a src slot, which holds source location information for the expression. This information, if present, will be residualized into the compiled object code, allowing backtraces to show source information. The format of src is the same as that returned by Guile’s source-properties function. See Section 6.25.2 [Source Properties], page 447, for more information.

Although Tree-IL objects are represented internally using records, there is also an equivalent S-expression external representation for each kind of Tree-IL. For example, the Sexpression representation of #<const src: #f exp: 3> expression would be:

     (const 3)

Users may program with this format directly at the REPL:

     scheme@(guile-user)> ,language tree-il
     Happy hacking with Tree Intermediate Language!  To switch back, type `,L scheme'.
     tree-il@(guile-user)> (apply (primitive +) (const 32) (const 10))

⇒ 42

The src fields are left out of the external representation. One may create Tree-IL objects from their external representations via calling parse-

tree-il, the reader for Tree-IL. If any source information is attached to the input Sexpression, it will be propagated to the resulting Tree-IL expressions. This is probably the easiest way to compile to Tree-IL: just make the appropriate external representations in S-expression format, and let parse-tree-il take care of the rest.

[Scheme Variable] <void> src [External Representation] (void)

An empty expression. In practice, equivalent to Scheme’s (if #f #f).

[Scheme Variable] <const> src exp [External Representation] (const exp)

A constant.

[Scheme Variable] <primitive-ref> src name [External Representation] (primitive name)

A reference to a “primitive”. A primitive is a procedure that, when compiled, may be open-coded. For example, cons is usually recognized as a primitive, so that it compiles down to a single instruction. Compilation of Tree-IL usually begins with a pass that resolves some <module-ref> and <toplevel-ref> expressions to <primitive-ref> expressions. The actual compilation pass has special cases for applications of certain primitives, like apply or

cons.

[Scheme Variable] <lexical-ref> src name gensym [External Representation] (lexical name gensym)

A reference to a lexically-bound variable. The name is the original name of the variable in the source program. gensym is a unique identifier for this variable.

[Scheme Variable] <lexical-set> src name gensym exp

[External Representation] (set! (lexical name gensym) exp)

Sets a lexically-bound variable.

[Scheme Variable] <module-ref> src mod name public?

[External Representation] (@ mod name) [External Representation] (@@ mod name)

A reference to a variable in a specific module. mod should be the name of the module,

e.g. (guile-user).

If public? is true, the variable named name will be looked up in mod’s public interface, and serialized with @; otherwise it will be looked up among the module’s private bindings, and is serialized with @@.

[Scheme Variable] <module-set> src mod name public? exp

[External Representation] (set! (@ mod name) exp) [External Representation] (set! (@@ mod name) exp)

Sets a variable in a specific module.

[Scheme Variable] <toplevel-ref> src name [External Representation] (toplevel name)

References a variable from the current procedure’s module.

[Scheme Variable] <toplevel-set> src name exp [External Representation] (set! (toplevel name) exp)

Sets a variable in the current procedure’s module.

[Scheme Variable] <toplevel-define> src name exp [External Representation] (define (toplevel name) exp)

Defines a new top-level variable in the current procedure’s module.

[Scheme Variable] <conditional> src test then else

[External Representation] (if test then else)

A conditional. Note that else is not optional.

[Scheme Variable] <application> src proc args [External Representation] (apply proc . args)

A procedure call.

[Scheme Variable] <sequence> src exps [External Representation] (begin . exps)

Like Scheme’s begin.

[Scheme Variable] <lambda> src meta body

[External Representation] (lambda meta body)

A closure. meta is an association list of properties for the procedure. body is a single Tree-IL expression of type <lambda-case>. As the <lambda-case> clause can chain to an alternate clause, this makes Tree-IL’s <lambda> have the expressiveness

of Scheme’s case-lambda.

[Scheme Variable] <lambda-case> req opt rest kw inits gensyms body alternate

[External Representation] (lambda-case ((req opt rest kw inits gensyms)
   body) [alternate])

One clause of a case-lambda. A lambda expression in Scheme is treated as a caselambda with one clause. req is a list of the procedure’s required arguments, as symbols. opt is a list of the optional arguments, or #f if there are no optional arguments. rest is the name of the rest argument, or #f.

kw is a list of the form, (allow-other-keys? (keyword name var) ...), where key-

word is the keyword corresponding to the argument named name, and whose corresponding gensym is var. inits are tree-il expressions corresponding to all of the optional and keyword arguments, evaluated to bind variables whose value is not supplied by the procedure caller. Each init expression is evaluated in the lexical context of previously bound variables, from left to right. gensyms is a list of gensyms corresponding to all arguments: first all of the required arguments, then the optional arguments if any, then the rest argument if any, then all of the keyword arguments. body is the body of the clause. If the procedure is called with an appropriate number of arguments, body is evaluated in tail position. Otherwise, if there is an alternate, it should be a <lambda-case> expression, representing the next clause to try. If there is no alternate, a wrong-number-of-arguments error is signaled.

[Scheme Variable] <let> src names gensyms vals exp

[External Representation] (let names gensyms vals exp)

Lexical binding, like Scheme’s let. names are the original binding names, gensyms are gensyms corresponding to the names, and vals are Tree-IL expressions for the values. exp is a single Tree-IL expression.

[Scheme Variable] <letrec> in-order? src names gensyms vals exp

[External Representation] (letrec names gensyms vals exp)
[External Representation] (letrec* names gensyms vals exp)

A version of <let> that creates recursive bindings, like Scheme’s letrec, or letrec* if in-order? is true.

[Scheme Variable] <dynlet> fluids vals body

[External Representation] (dynlet fluids vals body)

Dynamic binding; the equivalent of Scheme’s with-fluids. fluids should be a list of Tree-IL expressions that will evaluate to fluids, and vals a corresponding list of expressions to bind to the fluids during the dynamic extent of the evaluation of body.

[Scheme Variable] <dynref> fluid [External Representation] (dynref fluid)

A dynamic variable reference. fluid should be a Tree-IL expression evaluating to a fluid.

[Scheme Variable] <dynset> fluid exp [External Representation] (dynset fluid exp)

A dynamic variable set. fluid, a Tree-IL expression evaluating to a fluid, will be set to the result of evaluating exp.

[Scheme Variable] <dynwind> winder body unwinder

[External Representation] (dynwind winder body unwinder)

A dynamic-wind. winder and unwinder should both evaluate to thunks. Ensure that the winder and the unwinder are called before entering and after leaving body. Note that body is an expression, without a thunk wrapper.

[Scheme Variable] <prompt> tag body handler

[External Representation] (prompt tag body handler)

A dynamic prompt. Instates a prompt named tag, an expression, during the dynamic extent of the execution of body, also an expression. If an abort occurs to this prompt, control will be passed to handler, a <lambda-case> expression with no optional or keyword arguments, and no alternate. The first argument to the <lambda-case> will be the captured continuation, and then all of the values passed to the abort. See Section 6.13.5 [Prompts], page 296, for more information.

[Scheme Variable] <abort> tag args tail

[External Representation] (abort tag args tail)

An abort to the nearest prompt with the name tag, an expression. args should be a list of expressions to pass to the prompt’s handler, and tail should be an expression that will evaluate to a list of additional arguments. An abort will save the partial continuation, which may later be reinstated, resulting in the <abort> expression evaluating to some number of values.

There are two Tree-IL constructs that are not normally produced by higher-level compilers, but instead are generated during the source-to-source optimization and analysis passes that the Tree-IL compiler does. Users should not generate these expressions directly, unless they feel very clever, as the default analysis pass will generate them as necessary.

[Scheme Variable] <let-values> src names gensyms exp body

[External Representation] (let-values names gensyms exp body)

Like Scheme’s receive – binds the values returned by evaluating exp to the lambdalike bindings described by gensyms. That is to say, gensyms may be an improper list.

     <let-values> is an optimization of <application> of the primitive, call-with-
     values.

[Scheme Variable] <fix> src names gensyms vals body

[External Representation] (fix names gensyms vals body)

Like <letrec>, but only for vals that are unset lambda expressions.

fix is an optimization of letrec (and let).

Tree-IL implements a compiler to GLIL that recursively traverses Tree-IL expressions, writing out GLIL expressions into a linear list. The compiler also keeps some state as to whether the current expression is in tail context, and whether its value will be used in future computations. This state allows the compiler not to emit code for constant expressions that will not be used (e.g. docstrings), and to perform tail calls when in tail position.

Most optimization, such as it currently is, is performed on Tree-IL expressions as sourceto-source transformations. There will be more optimizations added in the future.

Interested readers are encouraged to read the implementation in (language tree-il compile-glil) for more details.

9.4.4 GLIL

Guile Lowlevel Intermediate Language (GLIL) is a structured intermediate language whose expressions more closely approximate Guile’s VM instruction set. Its expression types are

defined in (language glil).

[Scheme Variable] <glil-program> meta . body

A unit of code that at run-time will correspond to a compiled procedure. meta should be an alist of properties, as in Tree-IL’s <lambda>. body is an ordered list of GLIL expressions.

[Scheme Variable] <glil-std-prelude> nreq nlocs else-label A prologue for a function with no optional, keyword, or rest arguments. nreq is the number of required arguments. nlocs the total number of local variables, including the arguments. If the procedure was not given exactly nreq arguments, control will jump to else-label, if given, or otherwise signal an error.

[Scheme Variable] <glil-opt-prelude> nreq nopt rest nlocs else-label A prologue for a function with optional or rest arguments. Like <glil-std-prelude>, with the addition that nopt is the number of optional arguments (possibly zero) and rest is an index of a local variable at which to bind a rest argument, or #f if there is no rest argument.

[Scheme Variable] <glil-kw-prelude> nreq nopt rest kw allow-other-keys? nlocs

else-label A prologue for a function with keyword arguments. Like <glil-opt-prelude>, with the addition that kw is a list of keyword arguments, and allow-other-keys? is a flag indicating whether to allow unknown keys. See Section 9.3.6.4 [Function Prologue Instructions], page 782, for details on the format of kw.

[Scheme Variable] <glil-bind> . vars

An advisory expression that notes a liveness extent for a set of variables. vars is a list

of (name type index), where type should be either argument, local, or external.

<glil-bind> expressions end up being serialized as part of a program’s metadata and do not form part of a program’s code path.

[Scheme Variable] <glil-mv-bind> vars rest

A multiple-value binding of the values on the stack to vars. If rest is true, the last element of vars will be treated as a rest argument. In addition to pushing a binding annotation on the stack, like <glil-bind>, an expression is emitted at compilation time to make sure that there are enough values available to bind. See the notes on truncate-values in Section 9.3.6.3 [Procedure Call and Return Instructions], page 780, for more information.

[Scheme Variable] <glil-unbind>

Closes the liveness extent of the most recently encountered <glil-bind> or <glilmv-bind> expression. As GLIL expressions are compiled, a parallel stack of live bindings is maintained; this expression pops off the top element from that stack. Bindings are written into the program’s metadata so that debuggers and other tools can determine the set of live local variables at a given offset within a VM program.

[Scheme Variable] <glil-source> loc

Records source information for the preceding expression. loc should be an association list of containing line column, and filename keys, e.g. as returned by source-

properties.

[Scheme Variable] <glil-void>

Pushes “the unspecified value” on the stack.

[Scheme Variable] <glil-const> obj

Pushes a constant value onto the stack. obj must be a number, string, symbol, keyword, boolean, character, uniform array, the empty list, or a pair or vector of constants.

[Scheme Variable] <glil-lexical> local? boxed? op index Accesses a lexically bound variable. If the variable is not local? it is free. All variables may have ref, set, and bound? as their op. Boxed variables may also have the ops box, empty-box, and fix, which correspond in semantics to the VM instructions box, empty-box, and fix-closure. See Section 9.3.3 [Stack Layout], page 774, for more information.

[Scheme Variable] <glil-toplevel> op name

Accesses a toplevel variable. op may be ref, set, or define.

[Scheme Variable] <glil-module> op mod name public?

Accesses a variable within a specific module. See Tree-IL’s <module-ref>, for more information.

[Scheme Variable] <glil-label> label

Creates a new label. label can be any Scheme value, and should be unique.

[Scheme Variable] <glil-branch> inst label

Branch to a label. label should be a <ghil-label>. inst is a branching instruction:

br-if, br, etc.

[Scheme Variable] <glil-call> inst nargs

This expression is probably misnamed, as it does not correspond to function calls. <glil-call> invokes the VM instruction named inst, noting that it is called with nargs stack arguments. The arguments should be pushed on the stack already. What happens to the stack afterwards depends on the instruction.

[Scheme Variable] <glil-mv-call> nargs ra

Performs a multiple-value call. ra is a <glil-label> corresponding to the multiplevalue return address for the call. See the notes on mv-call in Section 9.3.6.3 [Procedure Call and Return Instructions], page 780, for more information.

[Scheme Variable] <glil-prompt> label escape-only?

Push a dynamic prompt into the stack, with a handler at label. escape-only? is a flag that is propagated to the prompt, allowing an abort to avoid capturing a continuation in some cases. See Section 6.13.5 [Prompts], page 296, for more information.

Users may enter in GLIL at the REPL as well, though there is a bit more bookkeeping to do:

scheme@(guile-user)> ,language glil
Happy hacking with Guile Lowlevel Intermediate Language (GLIL)!
To switch back, type `,L scheme'.
glil@(guile-user)> (program () (std-prelude 0 0 #f)
                       (const 3) (call return 1))
⇒ 3

Just as in all of Guile’s compilers, an environment is passed to the GLIL-to-object code compiler, and one is returned as well, along with the object code.

9.4.5 Assembly

Assembly is an S-expression-based, human-readable representation of the actual bytecodes that will be emitted for the VM. As such, it is a useful intermediate language both for compilation and for decompilation.

Besides the fact that it is not a record-based language, assembly differs from GLIL in four main ways:

• Labels have been resolved to byte offsets in the program.

• Constants inside procedures have either been expressed as inline instructions or cached in object arrays.

• Procedures with metadata (source location information, liveness extents, procedure names, generic properties, etc) have had their metadata serialized out to thunks.

• All expressions correspond directly to VM instructions – i.e., there is no <glillexical> which can be a ref or a set.

Assembly is isomorphic to the bytecode that it compiles to. You can compile to bytecode, then decompile back to assembly, and you have the same assembly code.

The general form of assembly instructions is the following:

(inst arg ...)

The inst names a VM instruction, and its args will be embedded in the instruction stream. The easiest way to see assembly is to play around with it at the REPL, as can be seen in this annotated example:

scheme@(guile-user)> ,pp (compile '(+ 32 10) #:to 'assembly)
(load-program
((:LCASE16 . 2))  ; Labels, unused in this case.
8                 ; Length of the thunk that was compiled.
(load-program     ; Metadata thunk.
()
17
#f              ; No metadata thunk for the metadata thunk.
(make-eol)
(make-eol)
(make-int8 2)   ; Liveness extents, source info, and arities,
(make-int8 8)   ; in a format that Guile knows how to parse.
(make-int8:0)
(list 0 3)
(list 0 1)
(list 0 3)
(return))
(assert-nargs-ee/locals 0)  ; Prologue.
(make-int8 32)    ; Actual code starts here.
(make-int8 10)
(add)
(return))

Of course you can switch the REPL to assembly and enter in assembly S-expressions directly, like with other languages, though it is more difficult, given that the length fields have to be correct.

9.4.6 Bytecode and Objcode

Finally, the raw bytes. There are actually two different “languages” here, corresponding to two different ways to represent the bytes.

“Bytecode” represents code as uniform byte vectors, useful for structuring and destructuring code on the Scheme level. Bytecode is the next step down from assembly:

     scheme@(guile-user)> (compile '(+ 32 10) #:to 'bytecode)
     ⇒ #vu8(8 0 0 0 25 0 0 0            ; Header.
            95 0                            ; Prologue.
            10 32 10 10 148 66 17           ; Actual code.
            0 0 0 0 0 0 0 9                 ; Metadata thunk.
            9 10 2 10 8 11 18 0 3 18 0 1 18 0 3 66)

“Objcode” is bytecode, but mapped directly to a C structure, struct scm_objcode:

     struct scm_objcode {
       scm_t_uint32 len;
       scm_t_uint32 metalen;
       scm_t_uint8 base[0];
     };

As one might imagine, objcode imposes a minimum length on the bytecode. Also, the len and metalen fields are in native endianness, which makes objcode (and bytecode) system-dependent.

Objcode also has a couple of important efficiency hacks. First, objcode may be mapped directly from disk, allowing compiled code to be loaded quickly, often from the system’s disk cache, and shared among multiple processes. Secondly, objcode may be embedded in other objcode, allowing procedures to have the text of other procedures inlined into their bodies, without the need for separate allocation of the code. Of course, the objcode object itself does need to be allocated.

Procedures related to objcode are defined in the (system vm objcode) module.

[Scheme Procedure] objcode? obj [C Function] scm_objcode_p (obj)

Returns #f if obj is object code, #f otherwise.

[Scheme Procedure] bytecode->objcode bytecode

[C Function] scm_bytecode_to_objcode (bytecode)

Makes a bytecode object from bytecode, which should be a bytevector. See Section 6.6.6 [Bytevectors], page 163.

[Scheme Variable] load-objcode file

[C Function] scm_load_objcode (file)

Load object code from a file named file. The file will be mapped into memory via mmap, so this is a very fast operation. On disk, object code has an sixteen-byte cookie prepended to it, to prevent accidental loading of arbitrary garbage.

[Scheme Variable] write-objcode objcode file [C Function] scm_write_objcode (objcode)

Write object code out to a file, prepending the sixteen-byte cookie.

[Scheme Variable] objcode->bytecode objcode

[C Function] scm_objcode_to_bytecode (objcode)

Copy object code out to a bytevector for analysis by Scheme.

The following procedure is actually in (system vm program), but we’ll mention it here:

[Scheme Variable] make-program objcode objtable [free-vars=#f] [C Function] scm_make_program (objcode, objtable, free vars)

Load up object code into a Scheme program. The resulting program will have objtable as its object table, which should be a vector or #f, and will capture the free variables from free-vars.

Object code from a file may be disassembled at the REPL via the meta-command ,disassemble-file, abbreviated as ,xx. Programs may be disassembled via ,disassemble, abbreviated as ,x.

Compiling object code to the fake language, value, is performed via loading objcode into a program, then executing that thunk with respect to the compilation environment. Normally the environment propagates through the compiler transparently, but users may specify the compilation environment manually as well, as a module.

9.4.7 Writing New High-Level Languages

In order to integrate a new language lang into Guile’s compiler system, one has to create the module (language lang spec) containing the language definition and referencing the parser, compiler and other routines processing it. The module hierarchy in (language brainfuck) defines a very basic Brainfuck implementation meant to serve as easy-to-understand example on how to do this. See for instance http://en.wikipedia.org/ wiki/Brainfuck for more information about the Brainfuck language itself.

9.4.8 Extending the Compiler

At this point we take a detour from the impersonal tone of the rest of the manual. Admit it: if you’ve read this far into the compiler internals manual, you are a junkie. Perhaps a course at your university left you unsated, or perhaps you’ve always harbored a desire to hack the holy of computer science holies: a compiler. Well you’re in good company, and in a good position. Guile’s compiler needs your help.

There are many possible avenues for improving Guile’s compiler. Probably the most important improvement, speed-wise, will be some form of native compilation, both just-in-time and ahead-of-time. This could be done in many ways. Probably the easiest strategy would be to extend the compiled procedure structure to include a pointer to a native code vector, and compile from bytecode to native code at run-time after a procedure is called a certain number of times.

The name of the game is a profiling-based harvest of the low-hanging fruit, running programs of interest under a system-level profiler and determining which improvements would give the most bang for the buck. It’s really getting to the point though that native compilation is the next step.

The compiler also needs help at the top end, enhancing the Scheme that it knows to also understand R6RS, and adding new high-level compilers. We have JavaScript and Emacs Lisp mostly complete, but they could use some love; Lua would be nice as well, but whatever language it is that strikes your fancy would be welcome too.

Compilers are for hacking, not for admiring or for complaining about. Get to it!

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 815

Concept Index

This index contains concepts, keywords and non-Schemey names for several features, to make it easier to locate the desired sections.

!

!# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359

#

#! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 #,(). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575

(

(ice-9 match) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657

.

.guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 .guile file, not loading . . . . . . . . . . . . . . . . . . . . . . . . 37 .guile_history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 .inputrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661

/

/etc/hosts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 /etc/protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 /etc/services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508

A

absolute file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 addrinfo object type . . . . . . . . . . . . . . . . . . . . . . . . . 503 affinity, CPU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228, 562 arbiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 argument specialize. . . . . . . . . . . . . . . . . . . . . . . . . . . 590 arguments (command line). . . . . . . . . . . . . . . . . . . . . 35 arity, variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 251, 578 association list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 association List. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 asyncs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 atomic time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 autoload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 automatic compilation . . . . . . . . . . . . . . . . . . . . . . . . 364 automatically-managed memory . . . . . . . . . . . . . . 376

B

begin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 binary input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339, 340 binary output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 binary port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 binding renamer . . . . . . . . . . . . . . . . . . . . . . . . . 382, 383 bindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 bitwise logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 block comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 Block read/write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 BOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455, 463 Buffered input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 buildstamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 byte order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 byte order mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 bytevector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

C

816 Guile Reference Manual

C port interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 callbacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 409 canonical host type . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 case folding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 chaining environments . . . . . . . . . . . . . . . . . . . . . . . . . 27 character encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 child processes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 Closing ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 code coverage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 Code coverage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 code point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 code point, designated . . . . . . . . . . . . . . . . . . . . . . . . 129 code point, reserved . . . . . . . . . . . . . . . . . . . . . . . . . . 129 codec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 codeset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 command line . . . . . . . . . . . . . . . . . . . . . . . . . . . 488, 489 command line history. . . . . . . . . . . . . . . . . . . . . . . . . 660 command-line arguments . . . . . . . . . . . . . . . . . . . . . . 35 Command-line Options . . . . . . . . . . . . . . . . . . . . . . . . 35 commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 composable continuations . . . . . . . . . . . . . . . . . . . . . 296 cond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 condition variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 conditional evaluation . . . . . . . . . . . . . . . . . . . . . . . . 292 conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 conservative garbage collection. . . . . . . . . . . . . . . . 376 continuation, escape . . . . . . . . . . . . . . . . . . . . . . . . . . 297 continuations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 Cooperative REPL server. . . . . . . . . . . . . . . . . . . . . 374 copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 coverage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 cultural conventions . . . . . . . . . . . . . . . . . . . . . . . . . . 436 curly-infix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 curly-infix-and-bracket-lists . . . . . . . . . . . . . . . . . . . 621 current directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 custom binary input ports . . . . . . . . . . . . . . . . . . . . 339 custom binary output ports. . . . . . . . . . . . . . . . . . . 344

D

database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 datadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583, 586 date conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 date to string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 date, from string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 debug options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 debugging virtual machine (command line). . . . . 37 Default ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 default slot value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 definition splicing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 delayed evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 delimited continuations . . . . . . . . . . . . . . . . . . . . . . . 296 designated code point . . . . . . . . . . . . . . . . . . . . . . . . 129 device file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 directory contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 directory traversal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 domain-specific language . . . . . . . . . . . . . . . . . . . . . 257 DSL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 duplicate binding . . . . . . . . . . . . . . . . . . . . . . . . 384, 385 duplicate binding handlers . . . . . . . . . . . . . . . . . . . . 385

E

EDSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 effective version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Emacs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 emacs regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 embedded domain-specific language . . . . . . . . . . . 257 encapsulation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 encryption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 End of file object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 end-of-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 end-of-line style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 endianness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 environment . . . . . . . . . . . . . . . . . . . . . . . . . 26, 395, 489 environment variables . . . . . . . . . . . . . . . . . . . . . 38, 621 environment, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 environment, top level . . . . . . . . . . . . . . . . . . . . . . . . . 26 EOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 errno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468, 469 error handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 error-signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 evaluate expression, command-line argument … 35 exception handling . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 exec_prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 export. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 expression sequencing. . . . . . . . . . . . . . . . . . . . . . . . . 290 extensiondir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 extensiondir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401

F

ffi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 file descriptor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 file locking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 file name separator . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 file options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 File port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 file system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 file system combinator . . . . . . . . . . . . . . . . . . . . . . . . 677 file system traversal . . . . . . . . . . . . . . . . . . . . . . . . . . 676 file tree walk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 finalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 finalizer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 fine-grain parallelism . . . . . . . . . . . . . . . . . . . . . . . . . 424 fluids. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 fold-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 foreign function interface . . . . . . . . . . . . . . . . . . . . . 396 formatted output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 functional setters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 futures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 424

G

GC-managed memory . . . . . . . . . . . . . . . . . . . . . . . . 376 GDB support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 Geiser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 general cond clause . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 GNU triplet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 GPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 group file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 guardians, testing for GC’d objects . . . . . . . . . . . 286 guild . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Guile threads. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 guile-2 SRFI-0 feature . . . . . . . . . . . . . . . . . . . . . . 552 guile-snarf deprecated macros. . . . . . . . . . . . . . . . . 102 guile-snarf example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 guile-snarf invocation . . . . . . . . . . . . . . . . . . . . . . . . . . 82 guile-snarf recognized macros . . . . . . . . . . . . . . . . . 102 guile-tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 GUILE_HISTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 guileversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428

H

hash-comma. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 higher-order functions . . . . . . . . . . . . . . . . . . . . . . . . 253 host name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 host name lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 HTTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519

I

i18n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 iconv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 IEEE-754 floating point numbers . . . . . . . . . . . . . 168 if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 includedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 infodir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 information encapsulation . . . . . . . . . . . . . . . . . . . . 381 init file, not loading. . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Initializing Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 inlining. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 720 integers as bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 772 invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 invocation (command-line arguments) . . . . . . . . . . 35 IPv4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 IPv6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 iteration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294

J

JACAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 Jaffer, Aubrey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 julian day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584, 587

K

keyword objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620

L

lambda. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 LANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 leap second. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 LGPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 libexecdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 libguileinterface . . . . . . . . . . . . . . . . . . . . . . . . . . 428 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 license . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Line buffered input . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 Line continuation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 Line input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 list constructor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 list delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 list filter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 list fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 list map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 list partition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 list predicate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 list search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 list selector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 list set operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 load path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 loading srfi modules (command line) . . . . . . . . . . . 37 local bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 local environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 local time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 local variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139, 436, 518 locale category . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 locale object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 localstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 looping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 low-level locale information . . . . . . . . . . . . . . . . . . . 439

M

macro expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 mandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 match structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 math – symbolic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468

                                                memory-allocation-error . . . . . . . . . . . . . . . . . . . 314

misc-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 modified julian day . . . . . . . . . . . . . . . . . . . . . . 584, 587 module version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 multiline comments. . . . . . . . . . . . . . . . . . . . . . . . . . . 359 multiple values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 multiple values and cond . . . . . . . . . . . . . . . . . . . . . 292 mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415

N

name space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 name space - private. . . . . . . . . . . . . . . . . . . . . . . . . . 381 named let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294, 295 network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 network address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 network database . . . . . . . . . . . . . . . . . . . 503, 505, 506 network examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 network protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 network services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 network socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 network socket address . . . . . . . . . . . . . . . . . . . . . . . 509 no-fold-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 non-local exit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 numerical-overflow . . . . . . . . . . . . . . . . . . . . . . . . . 314

O

object array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 776 object table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 776 options (command line). . . . . . . . . . . . . . . . . . . . . . . . 35 options - debug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 options - print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 options - read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 out-of-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 overflow, stack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 overriding binding . . . . . . . . . . . . . . . . . . . . . . . 384, 385

P

parallel forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 parallelism . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 424 parameter object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 parameter specialize . . . . . . . . . . . . . . . . . . . . . . . . . . 590 parameter specializers . . . . . . . . . . . . . . . . . . . . . . . . 730 Paredit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 partial evaluator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 password file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 pattern matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657 pattern matching (SXML) . . . . . . . . . . . . . . . . . . . . 687 pattern variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472, 500 pkg-config . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 pkgdatadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 pkgincludedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 pkglibdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 polar form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 118 Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 port buffering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 port encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 Port implementation. . . . . . . . . . . . . . . . . . . . . . . . . . 346 Port, block read/write . . . . . . . . . . . . . . . . . . . . . . . . 324 Port, C interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 Port, close. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 Port, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 Port, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 Port, line input/output . . . . . . . . . . . . . . . . . . . . . . . 322 Port, random access . . . . . . . . . . . . . . . . . . . . . . . . . . 321 Port, soft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 Port, string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 Port, types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 Port, void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 portability between 2.0 and older versions. . . . . 552 POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 POSIX threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 precise marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 pretty printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 primitive procedures . . . . . . . . . . . . . . . . . . . . . . . . . . 245 primitive-load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 print options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 procedure documentation . . . . . . . . . . . . . . . . . . . . . 255 procedure inlining . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 procedure properties. . . . . . . . . . . . . . . . . . . . . . . . . . 254 procedure with setter . . . . . . . . . . . . . . . . . . . . . . . . . 255 process group . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499, 500 process priority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 process time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 Profiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 program arguments . . . . . . . . . . . . . . . . . . . . . . 488, 489 promises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 prompts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 ptob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 pure module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386

Q

q-empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681

R

R6RS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 332, 622 R6RS block comments . . . . . . . . . . . . . . . . . . . . . . . . 359 R6RS ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 r7rs-symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 Random access, ports. . . . . . . . . . . . . . . . . . . . . . . . . 321 re-export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 read options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 Reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 readline options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 receive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215, 216 recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 recursive expression . . . . . . . . . . . . . . . . . . . . . . . . . . 593 regex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

                                                regular-expression-syntax . . . . . . . . . . . . . . . . . 314

remembering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 REPL server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 replacing binding . . . . . . . . . . . . . . . . . . . . . . . . 384, 385 reserved code point . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

S

sameness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 sbindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 Scheme Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 SCM data type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 script mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 SCSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 search and replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 sequencing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 service name lookup . . . . . . . . . . . . . . . . . . . . . . . . . . 503 services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 setter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 Setting breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 Setting tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 shadowing an imported variable binding . . . . . . . 27 sharedstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 site. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 site path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 sitedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 SLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 slot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 smob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 socket address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 socket client example . . . . . . . . . . . . . . . . . . . . . . . . . 516 socket examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 socket server example. . . . . . . . . . . . . . . . . . . . . . . . . 516 Soft port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 sorting lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 sorting vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 source file encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 source properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 specialize parameter . . . . . . . . . . . . . . . . . . . . . . . . . . 590 splicing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 SRFI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550 SRFI-0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 SRFI-1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 SRFI-10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 SRFI-105. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 SRFI-11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 SRFI-111. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 SRFI-13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 SRFI-14 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 SRFI-16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251, 578 SRFI-17 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 SRFI-18 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 SRFI-19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 SRFI-2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 SRFI-23 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 SRFI-26 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 SRFI-27 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591, 592, 593 SRFI-30 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 SRFI-30 block comments . . . . . . . . . . . . . . . . . . . . . 359 SRFI-31 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 SRFI-34 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 SRFI-35 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 SRFI-37 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 SRFI-38 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 SRFI-39 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422, 424, 599 SRFI-4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 SRFI-41 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599 SRFI-42 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 608 SRFI-43 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 SRFI-45 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 SRFI-46 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 SRFI-55 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 SRFI-6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 SRFI-60 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 SRFI-61 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 SRFI-62 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 SRFI-64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 SRFI-67 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 SRFI-69 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 SRFI-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 SRFI-87 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 SRFI-88 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 SRFI-88 keyword syntax. . . . . . . . . . . . . . . . . . . . . . 182 SRFI-9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 SRFI-98 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 stack overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 stack-overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 standard error output . . . . . . . . . . . . . . . . . . . . . . . . 326 standard input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 standard output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 startup (command-line arguments) . . . . . . . . . . . . . 35 streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 String port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 string to date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 string, from date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 switches (command line) . . . . . . . . . . . . . . . . . . . . . . . 35 SXML pattern matching . . . . . . . . . . . . . . . . . . . . . . 687 symbolic math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 sysconfdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 system asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 system clock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 system name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 system-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

T

TAI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583, 584 tail calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 temporary file . . . . . . . . . . . . . . . . . . . . . . . . . . . 481, 482 terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 textual port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 thread time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 threads. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485, 583, 584 time conversion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 time formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 time parsing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 top level environment . . . . . . . . . . . . . . . . . . . . . . . . . . 26 top_srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 Trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 Tracepoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 Tracing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 transcoder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 transformation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 Traps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 truncated printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 Types of ports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326

U

Unicode code point . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 Unicode string encoding . . . . . . . . . . . . . . . . . . . . . . 169 universal time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 unless . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 user asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 user information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 UTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583, 584

V

variable arity . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251, 578 variable definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 variable, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 vcell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 VHash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 vlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 VList-based hash lists . . . . . . . . . . . . . . . . . . . . . . . . 234 VM hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 VM trace level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457 Void port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332

W

warnings, compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 Web. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 when . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 wizards. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 word order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 wrapped pointer types . . . . . . . . . . . . . . . . . . . . . . . . 405 Writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 wrong-number-of-args . . . . . . . . . . . . . . . . . . . . . . . 314 wrong-type-arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 WWW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519

Procedure Index 821

Procedure Index

This is an alphabetical list of all the procedures and macros in Guile. It also includes Guile’s Autoconf macros.

When looking for a particular procedure, please look under its Scheme name as well as under its C name. The C name can be constructed from the Scheme names by a simple transformation described in the section See Section 6.1 [API Overview], page 99.

#

#:accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:each-subclass. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724 #:getter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:init-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 #:init-keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 #:init-thunk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 #:init-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 #:instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:metaclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 720 #:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 720 #:setter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 723 #:slot-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724, 725 #:slot-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724, 725 #:virtual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 %char-set-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

%default-port-conversion-strategy . . . . . . . . 318

%library-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 %make-void-port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 %package-data-dir . . . . . . . . . . . . . . . . . . . . . . . . . . 428 %read-delimited!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 %read-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 %search-load-path . . . . . . . . . . . . . . . . . . . . . . . . . . 368 %site-ccache-dir . . . . . . . . . . . . . . . . . . . . . . . . 55, 428 %site-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 428 %string-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

&

&assertion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 &error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &i/o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 &i/o-decoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 &i/o-encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335

&i/o-file-already-exists . . . . . . . . . . . . . . . . . . 643
&i/o-file-does-not-exist . . . . . . . . . . . . . . . . . . 643

&i/o-file-is-read-only. . . . . . . . . . . . . . . . . . . . . 643 &i/o-file-protection . . . . . . . . . . . . . . . . . . . . . . . 642 &i/o-filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 &i/o-invalid-position . . . . . . . . . . . . . . . . . . . . . . 642 &i/o-port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 &i/o-read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 &i/o-write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642

                                                &implementation-restriction . . . . . . . . . . . . . . . 641

&irritants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &lexical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 &no-infinities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 &no-nans . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 &non-continuable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &serious . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 &undefined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 &violation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 &warning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 &who . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641

’

' . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358

(

(oop goops). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 629 *scm_to_latin1_stringn. . . . . . . . . . . . . . . . . . . . . 162 *scm_to_stringn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 *scm_to_utf32_stringn . . . . . . . . . . . . . . . . . . . . . . 162 *scm_to_utf8_stringn . . . . . . . . . . . . . . . . . . . . . . . 162

                              +

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 629

,

, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358 ,@. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358

-

- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 629 ->char-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 629

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 627 <= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 628

=

= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 627 ==. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 627 >= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 628

@

@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 @@. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383

‘

` . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358

1

1+. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 1-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

A

abandoned-mutex-exception? . . . . . . . . . . . . . . . . 583

abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299, 788 abort-to-prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119, 629 absolute-file-name? . . . . . . . . . . . . . . . . . . . . . . . . 483 accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 access? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 acons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 acos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 acosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 activate-readline . . . . . . . . . . . . . . . . . . . . . . . . . . 662 adapt-response-version. . . . . . . . . . . . . . . . . . . . . 537 add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 add-duration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 add-duration! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585

add-ephemeral-stepping-trap! . . . . . . . . . . . . . . 463
add-ephemeral-trap-at-frame-finish! . . . . . . 463

add-hook!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 add-method! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753 add-ref-resolver! . . . . . . . . . . . . . . . . . . . . . . . . . . 712 add-to-load-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . 368

add-trace-at-procedure-call! . . . . . . . . . . . . . . 463

add-trap!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462

add-trap-at-procedure-call! . . . . . . . . . . . . . . . 463

add-trap-at-source-location! . . . . . . . . . . . . . . 463

add1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 addrinfo:addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 addrinfo:canonname . . . . . . . . . . . . . . . . . . . . . . . . . 505 addrinfo:fam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 addrinfo:flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 addrinfo:protocol . . . . . . . . . . . . . . . . . . . . . . . . . . 505 addrinfo:socktype . . . . . . . . . . . . . . . . . . . . . . . . . . 505 alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 alignof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 alist->hash-table. . . . . . . . . . . . . . . . . . . . . . 238, 619 alist->hashq-table . . . . . . . . . . . . . . . . . . . . . . . . . 238 alist->hashv-table . . . . . . . . . . . . . . . . . . . . . . . . . 238 alist->hashx-table . . . . . . . . . . . . . . . . . . . . . . . . . 238 alist->vhash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 alist-cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 alist-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 alist-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 alist-delete! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 all-threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293, 626 and-let* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 and=> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 627 any . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 any->c32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 any->c64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 any->f32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->f64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->s16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->s32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->s64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->u16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->u32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->u64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any->u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 any-bits-set? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190, 628 append! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 append-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 append-map! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 append-reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 append-reverse!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 apply. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363, 631, 780 apply-templates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 apply:nconc2last. . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

apropos-completion-function . . . . . . . . . . . . . . . 663

args-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 arithmetic-shift. . . . . . . . . . . . . . . . . . . . . . . . . . . . 616

arity:allow-other-keys? . . . . . . . . . . . . . . . . . . . 247

arity:end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 arity:kw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 arity:nopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 arity:nreq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 arity:rest? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 arity:start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 array->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 array-copy! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-copy-in-order! . . . . . . . . . . . . . . . . . . . . . . . 204 array-dimensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-equal? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 array-in-bounds?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 array-index-map!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 array-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-map! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-map-in-order! . . . . . . . . . . . . . . . . . . . . . . . . 204 array-rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 array-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 array-shape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 array-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 array? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 ash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126, 790 asin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 asinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 assert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 assert-curr-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 assert-nargs-ee. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 assert-nargs-ee/locals. . . . . . . . . . . . . . . . . . . . . 783 assert-nargs-ge. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 assert-nargs-ge/locals. . . . . . . . . . . . . . . . . . . . . 783 assertion-violation . . . . . . . . . . . . . . . . . . . . . . . . 630 assertion-violation? . . . . . . . . . . . . . . . . . . . . . . . 641 assoc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230, 562, 633 assoc-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assoc-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 assoc-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 assq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230, 633 assq-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assq-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 assq-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230, 633 assv-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assv-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 assv-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 async . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 async-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 atan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 atanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 attlist->alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 attlist-add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 attlist-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 attlist-null? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 attlist-remove-top . . . . . . . . . . . . . . . . . . . . . . . . . 703

B

backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 453 basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291, 626 begin-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 binary-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 bind-kwargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 bind-optionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 bind-optionals/shuffle. . . . . . . . . . . . . . . . . . . . . 783

bind-optionals/shuffle-or-br . . . . . . . . . . . . . . 783

bind-rest. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782

bind-textdomain-codeset . . . . . . . . . . . . . . . . . . . 444

binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 binding:boxed? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 binding:end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 binding:index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 binding:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 binding:start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 bindtextdomain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 bit-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199, 616 bit-count* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 bit-extract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 bit-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 bit-invert! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 bit-position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 bit-set*!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 bit-set? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 bitvector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 bitvector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 bitvector-fill!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 bitvector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 bitvector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 bitvector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 bitvector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 bitwise-and. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616, 650

bitwise-arithmetic-shift . . . . . . . . . . . . . . . . . . 651
bitwise-arithmetic-shift-left. . . . . . . . . . . . . 651
bitwise-arithmetic-shift-right . . . . . . . . . . . 651

bitwise-bit-count . . . . . . . . . . . . . . . . . . . . . . . . . . 650 bitwise-bit-field . . . . . . . . . . . . . . . . . . . . . . . . . . 651 bitwise-bit-set?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 bitwise-copy-bit. . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 bitwise-copy-bit-field. . . . . . . . . . . . . . . . . . . . . 651 bitwise-first-bit-set . . . . . . . . . . . . . . . . . . . . . . 650 bitwise-if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616, 650 bitwise-ior. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616, 650 bitwise-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 bitwise-merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 bitwise-not. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616, 650

bitwise-reverse-bit-field . . . . . . . . . . . . . . . . . 651
bitwise-rotate-bit-field . . . . . . . . . . . . . . . . . . 651

bitwise-xor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616, 650 boolean? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105, 624 booleans->integer . . . . . . . . . . . . . . . . . . . . . . . . . . 617 bound-identifier=? . . . . . . . . . . . . . . . . . . . . 268, 652 box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622, 778 box? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 br. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 br-if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 br-if-eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 br-if-nargs-gt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 br-if-nargs-lt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 br-if-nargs-ne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 br-if-not. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 br-if-not-eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 br-if-not-null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 br-if-null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 break . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 295, 560, 789 break! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 break-at-source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

broadcast-condition-variable . . . . . . . . . . . . . . 417

buffer-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 buffer-mode? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 build-request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 build-response . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 build-uri. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 bv-f32-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f32-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f32-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f32-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f64-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f64-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f64-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-f64-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s16-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s16-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s16-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s16-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s32-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s32-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s32-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s32-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s64-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s64-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s64-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s64-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s8-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-s8-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u16-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u16-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u16-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u16-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u32-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u32-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u32-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u32-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u64-native-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u64-native-set . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u64-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u64-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u8-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bv-u8-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 791 bytecode->objcode . . . . . . . . . . . . . . . . . . . . . . . . . . 803 bytevector->pointer . . . . . . . . . . . . . . . . . . . . . . . . 404 bytevector->sint-list . . . . . . . . . . . . . . . . . . . . . . 168 bytevector->string . . . . . . . . . . . . . . . . . . . . 159, 336 bytevector->u8-list . . . . . . . . . . . . . . . . . . . . . . . . 168 bytevector->uint-list . . . . . . . . . . . . . . . . . . . . . . 168 bytevector-copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 bytevector-copy!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 bytevector-fill!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

                                                bytevector-ieee-double-native-ref . . . . . . . . 169
                                                bytevector-ieee-double-native-set! . . . . . . . 169
                                                bytevector-ieee-double-ref . . . . . . . . . . . . . . . . 168
                                                bytevector-ieee-double-set! . . . . . . . . . . . . . . . 169
                                                bytevector-ieee-single-native-ref . . . . . . . . 169
                                                bytevector-ieee-single-native-set! . . . . . . . 169
                                                bytevector-ieee-single-ref . . . . . . . . . . . . . . . . 168
                                                bytevector-ieee-single-set! . . . . . . . . . . . . . . . 169

bytevector-length . . . . . . . . . . . . . . . . . . . . . . . . . . 164

                                                bytevector-s16-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-s16-native-set! . . . . . . . . . . . . . . . . 167

bytevector-s16-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-s16-set! . . . . . . . . . . . . . . . . . . . . . . . . 167

                                                bytevector-s32-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-s32-native-set! . . . . . . . . . . . . . . . . 167

bytevector-s32-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-s32-set! . . . . . . . . . . . . . . . . . . . . . . . . 167

                                                bytevector-s64-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-s64-native-set! . . . . . . . . . . . . . . . . 167

bytevector-s64-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-s64-set! . . . . . . . . . . . . . . . . . . . . . . . . 167 bytevector-s8-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-s8-set! . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-sint-ref . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-sint-set! . . . . . . . . . . . . . . . . . . . . . . . 166

                                                bytevector-u16-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-u16-native-set! . . . . . . . . . . . . . . . . 167

bytevector-u16-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-u16-set! . . . . . . . . . . . . . . . . . . . . . . . . 166

                                                bytevector-u32-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-u32-native-set! . . . . . . . . . . . . . . . . 167

bytevector-u32-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-u32-set! . . . . . . . . . . . . . . . . . . . . . . . . 167

                                                bytevector-u64-native-ref . . . . . . . . . . . . . . . . . 167
                                                bytevector-u64-native-set! . . . . . . . . . . . . . . . . 167

bytevector-u64-ref . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-u64-set! . . . . . . . . . . . . . . . . . . . . . . . . 167 bytevector-u8-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-u8-set! . . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-uint-ref . . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector-uint-set! . . . . . . . . . . . . . . . . . . . . . . . 166 bytevector=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 bytevector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

C

c32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 c32vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 c32vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 c32vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 c32vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 c32vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 c64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 c64vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 c64vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 c64vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 c64vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 c64vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 caaaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 caaadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 caaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 caadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 caaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 caadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 caar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cadaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cadadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 caddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cadddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 caddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 780

call-with-blocked-asyncs . . . . . . . . . . . . . . . . . . 412
call-with-current-continuation. . . . . . . 300, 631
call-with-error-handling . . . . . . . . . . . . . . . . . . 453
call-with-escape-continuation. . . . . . . . . . . . . 298

call-with-file-and-dir. . . . . . . . . . . . . . . . . . . . . 710 call-with-input-file . . . . . . . . . . . . . . . . . . 328, 644 call-with-input-string. . . . . . . . . . . . . . . . . . . . . 330 call-with-new-thread . . . . . . . . . . . . . . . . . . . . . . . 413

call-with-output-encoded-string . . . . . . . . . . 159

call-with-output-file . . . . . . . . . . . . . . . . . 328, 644

call-with-output-string . . . . . . . . . . . . . . . . . . . 330

call-with-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 call-with-prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 call-with-trace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461

call-with-unblocked-asyncs . . . . . . . . . . . . . . . . 412

call-with-values . . . . . . . . . . . . . . . . . . . . . . . 302, 631 call/cc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300, 631, 782 call/ec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 call/nargs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 cancel-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625, 789 car+cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293, 626 case-lambda. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251, 634 case-lambda* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 cd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 cdaaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cdaadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cdaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cdadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cdaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cdadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cdar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cddaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cddadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187, 625 cddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cdddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cddddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cdddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625 cddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625

cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625, 789
ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 629
ceiling-quotient. . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
ceiling-remainder . . . . . . . . . . . . . . . . . . . . . . . . . . 121
ceiling/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
center-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714
centered-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . 122
centered-remainder . . . . . . . . . . . . . . . . . . . . . . . . . 122
centered/. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122
change-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757
char->integer . . . . . . . . . . . . . . . . . . . . . . . . . . 133, 624
char-alphabetic? . . . . . . . . . . . . . . . . . . . . . . . 132, 631
char-ci<=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 631
char-ci<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 631
char-ci=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 631
char-ci>=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132, 631
char-ci>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 631
char-downcase . . . . . . . . . . . . . . . . . . . . . . . . . . 133, 631
char-foldcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631
char-general-category . . . . . . . . . . . . . . . . . 132, 632
char-is-both? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132
char-locale-ci<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
char-locale-ci=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 438
char-locale-ci>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
char-locale-downcase . . . . . . . . . . . . . . . . . . . . . . . 438
char-locale-titlecase . . . . . . . . . . . . . . . . . . . . . . 438
char-locale-upcase . . . . . . . . . . . . . . . . . . . . . . . . . 438
char-locale<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
char-locale>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
char-lower-case? . . . . . . . . . . . . . . . . . . . . . . . 132, 631
char-numeric? . . . . . . . . . . . . . . . . . . . . . . . . . . 132, 631
char-ready? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318
char-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135
char-set->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set->string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set-adjoin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-adjoin!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-any . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set-complement . . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-complement! . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-contains? . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135
char-set-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set-cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
char-set-cursor-next . . . . . . . . . . . . . . . . . . . . . . . 134
char-set-delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-delete!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

                                                char-set-diff+intersection . . . . . . . . . . . . . . . . 138
                                                char-set-diff+intersection! . . . . . . . . . . . . . . . 139

char-set-difference . . . . . . . . . . . . . . . . . . . . . . . . 138
char-set-difference! . . . . . . . . . . . . . . . . . . . . . . . 139
char-set-every . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
char-set-filter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136
char-set-filter!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 136
char-set-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
char-set-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . 135
char-set-hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
char-set-intersection . . . . . . . . . . . . . . . . . . . . . . 138

char-set-intersection!. . . . . . . . . . . . . . . . . . . . . 139 char-set-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 char-set-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 char-set-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 char-set-unfold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 char-set-unfold!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 char-set-union . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 char-set-union!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set-xor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 char-set-xor! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set<= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 char-set=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 char-set?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 char-title-case?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 char-titlecase . . . . . . . . . . . . . . . . . . . . . . . . . 133, 631 char-upcase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133, 631 char-upper-case? . . . . . . . . . . . . . . . . . . . . . . . 132, 631 char-whitespace? . . . . . . . . . . . . . . . . . . . . . . . 132, 631 char<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 char<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 char=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 char>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 char>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 char? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 624 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 chroot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 circular-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 circular-list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 748 class-direct-methods . . . . . . . . . . . . . . . . . . . . . . . 736 class-direct-slots . . . . . . . . . . . . . . . . . . . . . . . . . 736

class-direct-subclasses . . . . . . . . . . . . . . . . . . . 736

class-direct-supers . . . . . . . . . . . . . . . . . . . . . . . . 736 class-methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 class-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 736 class-of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737, 790 class-precedence-list . . . . . . . . . . . . . . . . . . . . . . 737 class-redefinition . . . . . . . . . . . . . . . . . . . . . . . . . 756 class-slot-definition . . . . . . . . . . . . . . . . . . . . . . 737 class-slot-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 741 class-slot-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 741 class-slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 class-subclasses. . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 clear-value-history! . . . . . . . . . . . . . . . . . . . . . . . . 47 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 close-fdes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 close-input-port . . . . . . . . . . . . . . . . . . . . . . . 321, 644 close-output-port. . . . . . . . . . . . . . . . . . . . . . 321, 644 close-pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 close-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 close-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 closedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481

collapse-repeated-chars . . . . . . . . . . . . . . . . . . . 715

command-line . . . . . . . . . . . . . . . . . . . . . . . . . . . 488, 645 compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 365 compile-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 366 compiled-file-name . . . . . . . . . . . . . . . . . . . . . . . . . 366 complex? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 627 compose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 compute-std-cpl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 concatenate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 concatenate! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 cond. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292, 626 cond-expand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 595, 640 condition-accessor . . . . . . . . . . . . . . . . . . . . . . . . . 640 condition-has-type? . . . . . . . . . . . . . . . . . . . . . . . . 595 condition-irritants . . . . . . . . . . . . . . . . . . . . . . . . 641 condition-message. . . . . . . . . . . . . . . . . . . . . . 596, 640 condition-predicate . . . . . . . . . . . . . . . . . . . . . . . . 640 condition-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 595 condition-type?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594

                                                condition-variable-broadcast!. . . . . . . . . . . . . 582
                                                condition-variable-name . . . . . . . . . . . . . . . . . . . 581
                                                condition-variable-signal! . . . . . . . . . . . . . . . . 582
                                                condition-variable-specific . . . . . . . . . . . . . . . 582
                                                condition-variable-specific-set! . . . . . . . . . 582

condition-variable? . . . . . . . . . . . . . . . . . . . 417, 581 condition-who . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 condition? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625, 789 cons* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189, 633 cons-source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 const . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 continuation-call . . . . . . . . . . . . . . . . . . . . . . . . . . 784 continue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 copy-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 copy-bit-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 copy-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 copy-random-state . . . . . . . . . . . . . . . . . . . . . . . . . . 127 copy-time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 copy-tree. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 cosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 coverage-data->lcov . . . . . . . . . . . . . . . . . . . . . . . . 464 coverage-data? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 crypt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 ctermid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 current-date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 current-dynamic-state . . . . . . . . . . . . . . . . . . . . . . 422 current-error-port . . . . . . . . . . . . . . . . . . . . 326, 343

                                                current-exception-handler . . . . . . . . . . . . . . . . . 582

current-filename. . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 current-http-proxy . . . . . . . . . . . . . . . . . . . . . . . . . 539 current-input-port . . . . . . . . . . . . . . . . . . . . 325, 339 current-julian-day . . . . . . . . . . . . . . . . . . . . . . . . . 587 current-load-port . . . . . . . . . . . . . . . . . . . . . . . . . . 367

                                                current-modified-julian-day . . . . . . . . . . . . . . . 587

current-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 current-output-port . . . . . . . . . . . . . . . . . . . 325, 343

                                                current-processor-count . . . . . . . . . . . . . . . . . . . 496
                                                current-source-location . . . . . . . . . . . . . . . . . . . 449

current-ssax-error-port . . . . . . . . . . . . . . . . . . . 702

current-thread . . . . . . . . . . . . . . . . . . . . . . . . . 413, 579 current-time. . . . . . . . . . . . . . . . . . . . . . . 485, 582, 585 cut . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 cute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590

D

date->julian-day. . . . . . . . . . . . . . . . . . . . . . . . . . . . 587

date->modified-julian-day . . . . . . . . . . . . . . . . . 587

date->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 date->time-monotonic . . . . . . . . . . . . . . . . . . . . . . . 587 date->time-tai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 date->time-utc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 date-day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-hour. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-nanosecond. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-second . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-week-day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-week-number. . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-year-day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date-zone-offset. . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 date? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 datum->random-state . . . . . . . . . . . . . . . . . . . . . . . . 128 datum->syntax . . . . . . . . . . . . . . . . . . . . . . . . . . 266, 652 debug-disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 debug-enable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 debug-options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 debug-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 declare-default-port! . . . . . . . . . . . . . . . . . . . . . . 522 declare-header!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 declare-opaque-header!. . . . . . . . . . . . . . . . . . . . . 524 deep-clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742

default-duplicate-binding-handler . . . . . . . . 386

default-prompt-tag . . . . . . . . . . . . . . . . . . . . . . . . . 296 default-random-source . . . . . . . . . . . . . . . . . . . . . . 592 define . . . . . . . . . . . . . . . . . . . . . . . . . 287, 626, 693, 779 define* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249, 693 define*-public . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 define-accessor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 define-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 define-condition-type . . . . . . . . . . . . . . . . . 595, 640 define-enumeration . . . . . . . . . . . . . . . . . . . . . . . . . 655 define-generic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727

define-immutable-record-type . . . . . . . . . . . . . . 217

define-inlinable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 define-language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 792 define-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 define-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 define-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 define-once . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 define-parsed-entity! . . . . . . . . . . . . . . . . . . . . . . 703 define-public . . . . . . . . . . . . . . . . . . . . . . . . . . 386, 693 define-reader-ctor . . . . . . . . . . . . . . . . . . . . . . . . . 576 define-record-type . . . . . . . . . . . . . . . . . . . . 216, 635 define-server-impl . . . . . . . . . . . . . . . . . . . . . . . . . 540 define-stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 601 define-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . 257, 626

define-syntax-parameter . . . . . . . . . . . . . . . . . . . 273

define-syntax-rule . . . . . . . . . . . . . . . . . . . . . . . . . 262 define-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290

define-wrapped-pointer-type . . . . . . . . . . . . . . . 405

defined? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 defmacro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 defmacro*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 defmacro*-public. . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 defvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 delay. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371, 614, 657 delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192, 561 delete! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192, 561 delete-duplicates . . . . . . . . . . . . . . . . . . . . . . . . . . 562 delete-duplicates! . . . . . . . . . . . . . . . . . . . . . . . . . 562 delete-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 delete-trap! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 delete1! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 delq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 delq! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 delq1! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 delv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 delv! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 delv1! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 denominator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 627 deq! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 dereference-pointer . . . . . . . . . . . . . . . . . . . . . . . . 404 describe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 directory-stream? . . . . . . . . . . . . . . . . . . . . . . . . . . 480 dirname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 disable-trap! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 disable-value-history! . . . . . . . . . . . . . . . . . . . . . . 47 disassemble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 disassemble-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 display . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361, 644, 743 display-application . . . . . . . . . . . . . . . . . . . . . . . . 447 display-backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . 446 display-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 div . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629, 790 div-and-mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 div0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 div0-and-mod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294, 634 dotted-list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

doubly-weak-hash-table? . . . . . . . . . . . . . . . . . . . 379

down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 drain-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 drop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555, 789 drop-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 drop-right! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 drop-while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 dup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473, 789 dup->fdes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 dup->inport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 dup->outport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 dup->port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 dup2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 duplicate-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 dynamic-call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 dynamic-func . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 dynamic-link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 dynamic-object?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 dynamic-pointer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 dynamic-state? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 dynamic-unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 dynamic-wind . . . . . . . . . . . . . . . . . . . . . . . . . . . 311, 631

E

eager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 ee? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 effective-version . . . . . . . . . . . . . . . . . . . . . . . . . . 427 eighth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 empty-box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778

enable-primitive-generic! . . . . . . . . . . . . . . . . . 728

enable-trap! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 enable-value-history! . . . . . . . . . . . . . . . . . . . . . . . 47

encode-and-join-uri-path . . . . . . . . . . . . . . . . . . 522

end-of-char-set?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 endgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 endhostent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 endianness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 endnetent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 endprotoent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 endpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 endservent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 enq! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 ensure-accessor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 754 ensure-generic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753 ensure-metaclass. . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 enum-set->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-complement . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-constructor . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-difference . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-indexer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-intersection . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-member?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-projection . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-subset?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-union . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 enum-set-universe . . . . . . . . . . . . . . . . . . . . . . . . . . 654 enum-set=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 environ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 eof-object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337, 643 eof-object? . . . . . . . . . . . . . . . . . . . . . . . . 318, 336, 643 eol-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 eq? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276, 626, 789 equal-hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 654 equal?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277, 626, 789 eqv? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277, 626, 789 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 309, 630 error-handling-mode . . . . . . . . . . . . . . . . . . . . . . . . 335 error-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 error? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596, 641 escape-special-chars . . . . . . . . . . . . . . . . . . . . . . . 713 euclidean-quotient . . . . . . . . . . . . . . . . . . . . . . . . . 120 euclidean-remainder . . . . . . . . . . . . . . . . . . . . . . . . 120 euclidean/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 eval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362, 656 eval-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 eval-when . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274, 367 even? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 627 every . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 exact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 exact->inexact . . . . . . . . . . . . . . . . . . . . . . . . . 114, 657 exact-integer-sqrt . . . . . . . . . . . . . . . . . . . . 117, 627 exact-integer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 exact? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 627 execl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 execle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 execlp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 exit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493, 645 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 expand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 expand-tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 expect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 687 expect-strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 export! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 expt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123, 627 extract-condition . . . . . . . . . . . . . . . . . . . . . . . . . . 595

F

f32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 f32vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 f32vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 f32vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 f32vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 f32vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 f64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 f64vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 f64vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 f64vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 f64vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 f64vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 false-if-exception . . . . . . . . . . . . . . . . . . . . . . . . . 309 fchmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 fchown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 fcntl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 fdes->inport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 fdes->outport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 fdes->ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 fdopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 feature? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 fflush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 fifth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 file-encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 file-exists? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 file-name-separator? . . . . . . . . . . . . . . . . . . . . . . . 483 file-options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 file-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 file-system-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . 677 file-system-tree. . . . . . . . . . . . . . . . . . . . . . . . . . . . 676

filename-completion-function . . . . . . . . . . . . . . 663

fileno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 fill-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 filter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192, 633, 708 filter! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 filter-empty-elements . . . . . . . . . . . . . . . . . . . . . . 712 filter-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 find. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560, 632 find-string-from-port?. . . . . . . . . . . . . . . . . . . . . 709 find-tail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 finish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 finite? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 630 first . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 first-set-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 fix-closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778 fixnum->flonum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 fixnum-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fixnum? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fl* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fl>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flacos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flasin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flatan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flceiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flcos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldenominator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldiv-and-mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldiv0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldiv0-and-mod0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fldmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 fleven? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flexpt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flfinite?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flfloor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flinfinite? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flinteger? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 fllog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flmin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flmod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flnan? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flnegative? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 flnumerator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 flock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475

flodd? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648
flonum? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648
floor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 629
floor-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
floor-remainder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
floor/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
flpositive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648
flround . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649
flsin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649
flsqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649
fltan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649
fltruncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649
fluid->parameter. . . . . . . . . . . . . . . . . . . . . . . . . . . . 423
fluid-bound? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420
fluid-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420, 788
fluid-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 788
fluid-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420
fluid-unset! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420
fluid? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420
flush-all-ports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321
flush-output-port . . . . . . . . . . . . . . . . . . . . . . . . . . 343
flzero? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648
fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556
fold-layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705
fold-left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633
fold-matches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352
fold-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556, 633
fold-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705
foldt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705
foldts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704, 705
foldts* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705
foldts*-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705
for-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632
for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . 193, 558, 628
force. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371, 614, 657
force-output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321
foreign-call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784
format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665
fourth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554
frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
frame-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-dynamic-link . . . . . . . . . . . . . . . . . . . . . . . . . 447

                                                frame-instruction-pointer . . . . . . . . . . . . . . . . . 447

frame-local-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-local-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 447

                                                frame-mv-return-address . . . . . . . . . . . . . . . . . . . 447

frame-num-locals. . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-previous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-procedure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447
frame-return-address . . . . . . . . . . . . . . . . . . . . . . . 447
frame-stack-pointer . . . . . . . . . . . . . . . . . . . . . . . . 447
frame? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446
free-boxed-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778
free-boxed-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778
free-identifier=?. . . . . . . . . . . . . . . . . . . . . . 268, 652
free-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778
fstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477

fsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 ftell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 ftruncate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 ftw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679 future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 future? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 fx* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx*/carry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx+/carry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx-/carry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fx<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fx<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fx=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fx>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fx>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fxand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxarithmetic-shift . . . . . . . . . . . . . . . . . . . . . . . . . 647

fxarithmetic-shift-left . . . . . . . . . . . . . . . . . . . 647
fxarithmetic-shift-right . . . . . . . . . . . . . . . . . . 647

fxbit-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxbit-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxbit-set? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxcopy-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxcopy-bit-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxdiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxdiv-and-mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxdiv0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxdiv0-and-mod0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxeven? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fxfirst-bit-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxlength . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxmin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxmod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 fxnegative? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fxnot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxodd? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fxpositive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 fxreverse-bit-field . . . . . . . . . . . . . . . . . . . . . . . . 647 fxrotate-bit-field . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxxor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 fxzero? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645

G

gc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 374 gc-live-object-stats . . . . . . . . . . . . . . . . . . . . . . . 375 gc-stats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 375 gcd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 627 gcprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 ge? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 generate-temporaries . . . . . . . . . . . . . . . . . . 268, 652

generic-function-methods . . . . . . . . . . . . . . . . . . 739

generic-function-name . . . . . . . . . . . . . . . . . . . . . . 739 gensym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 get-bytevector-all . . . . . . . . . . . . . . . . . . . . . . . . . 341 get-bytevector-n. . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 get-bytevector-n! . . . . . . . . . . . . . . . . . . . . . . . . . . 340 get-bytevector-some . . . . . . . . . . . . . . . . . . . . . . . . 341 get-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 get-datum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342

get-environment-variable . . . . . . . . . . . . . . . . . . 621
get-environment-variables . . . . . . . . . . . . . . . . . 621

get-internal-real-time. . . . . . . . . . . . . . . . . . . . . 488 get-internal-run-time . . . . . . . . . . . . . . . . . . . . . . 488 get-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 get-output-string . . . . . . . . . . . . . . . . . . . . . . . . . . 331 get-print-state. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 get-string-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 get-string-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 get-string-n! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 get-u8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 getaddrinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 getaffinity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 getcwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 getegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 geteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 getgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 getgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 getgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 getgrgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 getgrnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 getgroups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 gethost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 gethostbyaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 gethostbyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 gethostent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 gethostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 getitimer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 getlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 getnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 getnetbyaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 getnetbyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 getnetent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 getopt-long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 549 getpass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 getpeername . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 getpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 getpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 getppid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 getpriority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 getproto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 getprotobyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 getprotobynumber. . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 getprotoent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 getpw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 getpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 getpwnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 getpwuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 getserv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 getservbyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 getservbyport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 getservent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 getsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 getsockname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 getsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 getter-with-setter . . . . . . . . . . . . . . . . . . . . . . . . . 578 gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 gettimeofday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 getuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 gmtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 goops-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 greatest-fixnum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 group:gid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 group:mem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 group:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 group:passwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 gt? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 guard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 639 guild compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 GUILE_CHECK_RETVAL . . . . . . . . . . . . . . . . . . . . . . . . . . 95 GUILE_FLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 GUILE_MODULE_AVAILABLE . . . . . . . . . . . . . . . . . . . . . . 95 GUILE_MODULE_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . 95 GUILE_MODULE_EXPORTS . . . . . . . . . . . . . . . . . . . . . . . . 96 GUILE_MODULE_REQUIRED . . . . . . . . . . . . . . . . . . . . . . . 96

GUILE_MODULE_REQUIRED_EXPORT . . . . . . . . . . . . . . . 96

GUILE_PKG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 GUILE_PROGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 GUILE_SITE_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

H

halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789 handle-request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 hash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239, 620 hash-by-identity. . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 hash-clear! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hash-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 hash-create-handle! . . . . . . . . . . . . . . . . . . . . . . . . 240 hash-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 hash-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hash-for-each-handle . . . . . . . . . . . . . . . . . . . . . . . 241 hash-get-handle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hash-map->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hash-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hash-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hash-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hash-table->alist . . . . . . . . . . . . . . . . . . . . . . . . . . 620 hash-table-delete! . . . . . . . . . . . . . . . . . . . . . . . . . 619

hash-table-equivalence-function . . . . . . . . . . 620

hash-table-exists? . . . . . . . . . . . . . . . . . . . . . . . . . 619 hash-table-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620

hash-table-hash-function . . . . . . . . . . . . . . . . . . 620

hash-table-keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 hash-table-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 hash-table-ref/default. . . . . . . . . . . . . . . . . . . . . 619 hash-table-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 hash-table-size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 hash-table-update! . . . . . . . . . . . . . . . . . . . . . . . . . 619

hash-table-update!/default . . . . . . . . . . . . . . . . 619

hash-table-values . . . . . . . . . . . . . . . . . . . . . . . . . . 620 hash-table-walk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 hash-table? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashq-create-handle! . . . . . . . . . . . . . . . . . . . . . . . 240 hashq-get-handle. . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hashq-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashq-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashq-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashtable-clear!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-contains? . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-delete! . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-entries . . . . . . . . . . . . . . . . . . . . . . . . . . 654

hashtable-equivalence-function . . . . . . . . . . . 654
hashtable-hash-function . . . . . . . . . . . . . . . . . . . 654

hashtable-keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-mutable? . . . . . . . . . . . . . . . . . . . . . . . . . 654 hashtable-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable-update! . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashtable? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 hashv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashv-create-handle! . . . . . . . . . . . . . . . . . . . . . . . 240 hashv-get-handle. . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hashv-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashv-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashv-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashx-create-handle! . . . . . . . . . . . . . . . . . . . . . . . 240 hashx-get-handle. . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hashx-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashx-remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 hashx-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 header->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 header-parser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 header-validator. . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 header-writer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 hook->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 hook-empty? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 hook? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 hostent:addr-list . . . . . . . . . . . . . . . . . . . . . . . . . . 505 hostent:addrtype. . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 hostent:aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 hostent:length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 hostent:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 http . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 http-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-head. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-post. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 http-trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538

I

i/o-decoding-error? . . . . . . . . . . . . . . . . . . . . . . . . 335

i/o-encoding-error-char . . . . . . . . . . . . . . . . . . . 335

i/o-encoding-error? . . . . . . . . . . . . . . . . . . . . . . . . 335 i/o-error-filename . . . . . . . . . . . . . . . . . . . . . . . . . 642 i/o-error-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 i/o-error-position . . . . . . . . . . . . . . . . . . . . . . . . . 642 i/o-error? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642

i/o-file-already-exists-error? . . . . . . . . . . . 643
i/o-file-does-not-exist-error? . . . . . . . . . . . 643
i/o-file-is-read-only-error? . . . . . . . . . . . . . . 643
i/o-file-protection-error? . . . . . . . . . . . . . . . . 642

i/o-filename-error? . . . . . . . . . . . . . . . . . . . . . . . . 642

i/o-invalid-position-error? . . . . . . . . . . . . . . . 642

i/o-port-error?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 i/o-read-error?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 i/o-write-error?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 identifier-syntax . . . . . . . . . . . . . . . . . 271, 272, 626 identifier?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265, 652 identity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292, 626 imag-part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 627

implementation-restriction-violation? . . . 641

import . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 390 in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 include-from-path . . . . . . . . . . . . . . . . . . . . . . . . . . 373 inet-aton. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-lnaof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-makeaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-netof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-ntoa. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-ntop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inet-pton. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 inexact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 inexact->exact . . . . . . . . . . . . . . . . . . . . . . . . . 114, 657 inexact? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114, 627 inf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 inf? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 infinite?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 input-port? . . . . . . . . . . . . . . . . . . . . . . . . 317, 338, 644 inspect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 install-trap-handler! . . . . . . . . . . . . . . . . . . . . . . 463 instance?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737

instrumented-source-files . . . . . . . . . . . . . . . . . 465
instrumented/executed-lines . . . . . . . . . . . . . . . 465

integer->char . . . . . . . . . . . . . . . . . . . . . . . . . . 133, 624 integer->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 integer-expt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 integer-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 integer-valued?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 integer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107, 627

interaction-environment . . . . . . . . . . . . . . . . . . . 362

iota . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 irritants-condition? . . . . . . . . . . . . . . . . . . . . . . . 641 is-a? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 isatty? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499

J

join-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413

                                                join-timeout-exception? . . . . . . . . . . . . . . . . . . . 583

julian-day->date. . . . . . . . . . . . . . . . . . . . . . . . . . . . 587

                                                julian-day->time-monotonic . . . . . . . . . . . . . . . . 587

julian-day->time-tai . . . . . . . . . . . . . . . . . . . . . . . 587 julian-day->time-utc . . . . . . . . . . . . . . . . . . . . . . . 587

K

keyword->string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 keyword->symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 keyword? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183, 621 kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 known-header? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523

L

lalr-parser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244, 626 lambda* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 last . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 last-pair. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 latin-1-codec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 lazy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 lchown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 lcm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 627 le? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 least-fixnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 left-justify-string . . . . . . . . . . . . . . . . . . . . . . . . 714 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190, 628 length+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287, 295, 626 let*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288, 626 let*-values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577, 626

                                                let-escape-continuation . . . . . . . . . . . . . . . . . . . 298

let-keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 let-keywords* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 let-optional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 let-optional* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 let-syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258, 626 let-values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577, 626 let/ec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 letpar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 letrec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288, 626 letrec* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289, 626 letrec-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . 258, 626 lexical-violation? . . . . . . . . . . . . . . . . . . . . . . . . . 641 library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 line-execution-counts . . . . . . . . . . . . . . . . . . . . . . 465 link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 link-now . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 779 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189, 628, 786 list->array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 list->bitvector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 list->c32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->c64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->char-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 list->char-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 list->f32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->f64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 list->s16vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->s32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->s64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->stream . . . . . . . . . . . . . . . . . . . . . . . . . . . 601, 683 list->string . . . . . . . . . . . . . . . . . . . . . . . . . . . 144, 628 list->symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 list->typed-array . . . . . . . . . . . . . . . . . . . . . . . . . . 203 list->u16vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->u32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->u64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 list->vector. . . . . . . . . . . . . . . . . . . . . . . 194, 613, 630 list->vlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 list->weak-vector . . . . . . . . . . . . . . . . . . . . . . . . . . 379 list-cdr-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 list-cdr-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 list-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189, 553 list-head. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 list-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 list-matches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352 list-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190, 628 list-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 list-sort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 list-tabulate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 list-tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190, 628 list-traps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 list= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 list?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189, 625, 789 listen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 366 load-array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 load-compiled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 366 load-extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 load-from-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 load-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 load-objcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 803 load-program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 load-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 load-symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 load-wide-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 local-bound? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 779 local-boxed-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778 local-boxed-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778 local-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 local-eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 local-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 777 local-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778 locale-am-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-currency-symbol. . . . . . . . . . . . . . . . . . . . . 441

locale-currency-symbol-precedes-negative?

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441

locale-currency-symbol-precedes-positive?

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441

locale-date+time-format . . . . . . . . . . . . . . . . . . . 440

locale-date-format . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-day-short. . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-decimal-point . . . . . . . . . . . . . . . . . . . . . . . 440 locale-digit-grouping . . . . . . . . . . . . . . . . . . . . . . 440 locale-encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 locale-era . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440

locale-era-date+time-format . . . . . . . . . . . . . . . 440

locale-era-date-format. . . . . . . . . . . . . . . . . . . . . 440 locale-era-time-format. . . . . . . . . . . . . . . . . . . . . 440 locale-era-year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440

locale-monetary-decimal-point. . . . . . . . . . . . . 441
locale-monetary-fractional-digits . . . . . . . . 441
locale-monetary-grouping . . . . . . . . . . . . . . . . . . 441
locale-monetary-negative-sign. . . . . . . . . . . . . 441
locale-monetary-positive-sign. . . . . . . . . . . . . 441
locale-monetary-thousands-separator . . . . . . 441

locale-month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-month-short . . . . . . . . . . . . . . . . . . . . . . . . . 440

locale-negative-separated-by-space? . . . . . . 441
locale-negative-sign-position. . . . . . . . . . . . . 441

locale-no-regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 locale-pm-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 440

locale-positive-separated-by-space? . . . . . . 441
locale-positive-sign-position. . . . . . . . . . . . . 441

locale-string->inexact. . . . . . . . . . . . . . . . . . . . . 439 locale-string->integer. . . . . . . . . . . . . . . . . . . . . 439

locale-thousands-separator . . . . . . . . . . . . . . . . 440
locale-time+am/pm-format . . . . . . . . . . . . . . . . . . 440

locale-time-format . . . . . . . . . . . . . . . . . . . . . . . . . 440 locale-yes-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 442 locale? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 locals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 localtime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 lock-mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 log10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 log2-binary-factors . . . . . . . . . . . . . . . . . . . . . . . . 616 logand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 790 logbit? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 logcount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 logior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 790 lognot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 logtest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 logxor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 790 long-local-bound? . . . . . . . . . . . . . . . . . . . . . . . . . . 779 long-local-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 777 long-local-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778 long-object-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786 long-toplevel-ref . . . . . . . . . . . . . . . . . . . . . . . . . . 779 long-toplevel-set . . . . . . . . . . . . . . . . . . . . . . . . . . 779 lookahead-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 lookahead-u8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340

lookup-compilation-order . . . . . . . . . . . . . . . . . . 793

lookup-language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 792 lookup-server-impl . . . . . . . . . . . . . . . . . . . . . . . . . 541 lset-adjoin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 lset-diff+intersection. . . . . . . . . . . . . . . . . . . . . 565

lset-diff+intersection! . . . . . . . . . . . . . . . . . . . 565

lset-difference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 lset-difference!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 lset-intersection . . . . . . . . . . . . . . . . . . . . . . . . . . 564 lset-intersection! . . . . . . . . . . . . . . . . . . . . . . . . . 564 lset-union . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 lset-union! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 lset-xor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 lset-xor!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 lset<= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 lset= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 lstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 lt? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790

M

macro-binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 macro-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 macro-transformer . . . . . . . . . . . . . . . . . . . . . . . . . . 275 macro-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 macro? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 magnitude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 627 major-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 720, 749 make-accessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 754 make-arbiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 make-array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202, 786

make-assertion-violation . . . . . . . . . . . . . . . . . . 641

make-binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 make-bitvector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

make-buffered-input-port . . . . . . . . . . . . . . . . . . 684

make-bytevector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 make-c-struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 make-c32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-c64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-char32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786 make-char8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786

make-chunked-input-port . . . . . . . . . . . . . . . . . . . 533
make-chunked-output-port . . . . . . . . . . . . . . . . . . 533

make-class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 748 make-closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 778

make-completion-function . . . . . . . . . . . . . . . . . . 663
make-compound-condition . . . . . . . . . . . . . . . . . . . 594

make-condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 594 make-condition-type . . . . . . . . . . . . . . . . . . . . . . . . 594

make-condition-variable . . . . . . . . . . . . . . . 417, 581
make-custom-binary-input-port. . . . . . . . . . . . . 339
make-custom-binary-output-port . . . . . . . . . . . 344

make-date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586

make-doubly-weak-hash-table . . . . . . . . . . . . . . . 379

make-dynamic-state . . . . . . . . . . . . . . . . . . . . . . . . . 421 make-empty-attlist . . . . . . . . . . . . . . . . . . . . . . . . . 702 make-enumeration. . . . . . . . . . . . . . . . . . . . . . . . . . . . 654 make-eol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-eq-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . . 652 make-eqv-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . 652 make-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 make-f32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-f64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-fluid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 make-future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 make-generic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 754 make-guardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 make-hash-table . . . . . . . . . . . . . . . . . . . . . . . . 238, 618 make-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 make-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282

make-i/o-decoding-error . . . . . . . . . . . . . . . . . . . 335
make-i/o-encoding-error . . . . . . . . . . . . . . . . . . . 335

make-i/o-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642

make-i/o-file-already-exists-error . . . . . . . 643
make-i/o-file-does-not-exist-error . . . . . . . 643
make-i/o-file-is-read-only-error . . . . . . . . . 643
make-i/o-file-protection-error . . . . . . . . . . . 642
make-i/o-invalid-position-error . . . . . . . . . . 642

make-i/o-port-error . . . . . . . . . . . . . . . . . . . . . . . . 643 make-i/o-read-error . . . . . . . . . . . . . . . . . . . . . . . . 642 make-i/o-write-error . . . . . . . . . . . . . . . . . . . . . . . 642

make-implementation-restriction-violation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 make-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 make-int16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-int64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-int8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-int8:0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-int8:1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-io-filename-error. . . . . . . . . . . . . . . . . . . . . 642

make-irritants-condition . . . . . . . . . . . . . . . . . . 641

make-keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786 make-lexical-violation. . . . . . . . . . . . . . . . . . . . . 641

make-line-buffered-input-port. . . . . . . . . . . . . 685

make-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 make-locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 make-message-condition. . . . . . . . . . . . . . . . . . . . . 640 make-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753 make-mutex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415, 580 make-nil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785

make-no-infinities-violation . . . . . . . . . . . . . . 650

make-no-nans-violation. . . . . . . . . . . . . . . . . . . . . 650

make-non-continuable-violation . . . . . . . . . . . 641

make-object-property . . . . . . . . . . . . . . . . . . . . . . . 278 make-parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 make-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 make-polar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 627

make-procedure-with-setter . . . . . . . . . . . . . . . . 256

make-program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 804 make-prompt-tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 make-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 make-random-source . . . . . . . . . . . . . . . . . . . . . . . . . 592

make-record-constructor-descriptor . . . . . . . 637

make-record-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

make-record-type-descriptor . . . . . . . . . . . . . . . 637

make-rectangular . . . . . . . . . . . . . . . . . . . . . . . 118, 627 make-recursive-mutex . . . . . . . . . . . . . . . . . . . . . . . 415 make-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 350 make-s16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-s32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-s64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-serious-condition. . . . . . . . . . . . . . . . . . . . . 641 make-shared-array . . . . . . . . . . . . . . . . . . . . . . . . . . 206 make-socket-address . . . . . . . . . . . . . . . . . . . . . . . . 509 make-soft-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 make-stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 make-stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 make-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144, 628 make-struct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221, 786 make-struct-layout . . . . . . . . . . . . . . . . . . . . . . . . . 225 make-struct/no-tail . . . . . . . . . . . . . . . . . . . . . . . . 221 make-symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179, 786

make-syntax-transformer . . . . . . . . . . . . . . . . . . . 275

make-syntax-violation . . . . . . . . . . . . . . . . . . . . . . 642 make-tcp-server-socket. . . . . . . . . . . . . . . . . . . . . 373 make-text-wrapper . . . . . . . . . . . . . . . . . . . . . . . . . . 715 make-thread. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414, 579 make-time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 make-transcoder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 make-true. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-typed-array. . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 make-u16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-u32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-u64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 make-uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 make-unbound-fluid . . . . . . . . . . . . . . . . . . . . . . . . . 420

make-undefined-variable . . . . . . . . . . . . . . . . . . . 390
make-undefined-violation . . . . . . . . . . . . . . . . . . 642
make-unix-domain-server-socket . . . . . . . . . . . 373

make-variable . . . . . . . . . . . . . . . . . . . . . . . . . . 390, 780

make-variable-transformer . . . . . . . . . . . . 272, 651

make-vector . . . . . . . . . . . . . . . . . . . . . . . . 194, 609, 630 make-violation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 make-vtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 make-warning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640

make-weak-key-hash-table . . . . . . . . . . . . . . . . . . 379
make-weak-value-hash-table . . . . . . . . . . . . . . . . 379

make-weak-vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 make-who-condition . . . . . . . . . . . . . . . . . . . . . . . . . 641 malloc-stats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193, 558 map! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 map-in-order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 map-union. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 658 match:count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 match:end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 match:prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 match:start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 match:string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 match:substring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 match:suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 629 member. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193, 561, 633 memp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 memq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193, 633 memv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193, 633 merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 merge! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 message-condition? . . . . . . . . . . . . . . . . . . . . 596, 640 method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 752

                                                method-generic-function . . . . . . . . . . . . . . . . . . . 739

method-procedure. . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 method-source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 method-specializers . . . . . . . . . . . . . . . . . . . . . . . . 739 micro-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 629 minor-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 mkstemp! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 mktime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629, 790 mod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629

                                                modified-julian-day->date . . . . . . . . . . . . . . . . . 587
                                                modified-julian-day->time-monotonic . . . . . . 587
                                                modified-julian-day->time-tai. . . . . . . . . . . . . 587
                                                modified-julian-day->time-utc. . . . . . . . . . . . . 587

module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 module-add! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 module-define! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 module-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 module-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392

                                                module-stexi-documentation . . . . . . . . . . . . . . . . 716

module-use! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 module-uses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 module-variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 modulo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 657 modulo-expt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

                                                monetary-amount->locale-string . . . . . . . . . . . 439

monitor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 move->fdes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 mul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 mutex-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 mutex-lock! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex-locked? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 mutex-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex-owner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 mutex-specific . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex-specific-set! . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex-state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex-unlock! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mutex? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 mv-call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781

N

n-for-each-par-map . . . . . . . . . . . . . . . . . . . . . . . . . 426 n-par-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 n-par-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 nan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 nan?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 630 native-endianness . . . . . . . . . . . . . . . . . . . . . . . . . . 164 native-eol-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 native-transcoder . . . . . . . . . . . . . . . . . . . . . . . . . . 336 negate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 negative? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 628 netent:addrtype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 netent:aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 netent:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 netent:net . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 new-frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 780 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320, 644 next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 next-token . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 next-token-of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 nftw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679 ngettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 nice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 ninth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 nl_langinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 no-applicable-method . . . . . . . . . . . . . . . . . . . . . . . 733

no-infinities-violation? . . . . . . . . . . . . . . . . . . 650

no-method. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733 no-nans-violation? . . . . . . . . . . . . . . . . . . . . . . . . . 650 no-next-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733 node-closure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-eq? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 node-equal? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 node-join. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-pos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 node-reduce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-self. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 node-typeof? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 nodeset? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707

non-continuable-violation? . . . . . . . . . . . . . . . . 641

nop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 788 not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105, 624, 789 not-eq? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789 not-not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789 not-null?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789 not-pair?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 null-environment . . . . . . . . . . . . . . . . . . . . . . . 396, 657 null-list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 null-pointer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 null?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189, 625, 789 number->locale-string . . . . . . . . . . . . . . . . . . . . . . 439 number->string . . . . . . . . . . . . . . . . . . . . . . . . . 118, 628 number? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106, 625 numerator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 627

O

objcode->bytecode . . . . . . . . . . . . . . . . . . . . . . . . . . 804 objcode? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 803 object->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 object-properties . . . . . . . . . . . . . . . . . . . . . . . . . . 278 object-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 object-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 786

                                                object-stexi-documentation . . . . . . . . . . . . . . . . 717

odd?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 627 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471

                                                open-bytevector-input-port . . . . . . . . . . . . . . . . 339
                                                open-bytevector-output-port . . . . . . . . . . . . . . . 344

open-fdes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 open-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 open-file-input-port . . . . . . . . . . . . . . . . . . . . . . . 338 open-file-output-port . . . . . . . . . . . . . . . . . . . . . . 343 open-input-file . . . . . . . . . . . . . . . . . . . . . . . . 328, 644 open-input-output-pipe. . . . . . . . . . . . . . . . . . . . . 501 open-input-pipe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 open-input-string . . . . . . . . . . . . . . . . . . . . . . . . . . 330 open-output-file . . . . . . . . . . . . . . . . . . . . . . . 328, 644 open-output-pipe. . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 open-output-string . . . . . . . . . . . . . . . . . . . . . . . . . 330 open-pipe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 open-pipe* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 open-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 open-socket-for-uri . . . . . . . . . . . . . . . . . . . . . . . . 538 opendir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 optimize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 597 option-names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 option-optional-arg? . . . . . . . . . . . . . . . . . . . . . . . 597 option-processor. . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 option-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550 option-required-arg? . . . . . . . . . . . . . . . . . . . . . . . 597 or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293, 626 output-port?. . . . . . . . . . . . . . . . . . . . . . . 317, 343, 644

P

                                                package-stexi-documentation . . . . . . . . . . . . . . . 718
                                                package-stexi-documentation-for-include

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718

                                                package-stexi-extended-menu . . . . . . . . . . . . . . . 717
                                                package-stexi-generic-menu . . . . . . . . . . . . . . . . 717
                                                package-stexi-standard-copying . . . . . . . . . . . 717
                                                package-stexi-standard-menu . . . . . . . . . . . . . . . 717
                                                package-stexi-standard-prologue . . . . . . . . . . 717
                                                package-stexi-standard-titlepage . . . . . . . . . 717

pair-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 pair-fold-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 pair-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 pair?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186, 625, 789 par-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 par-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 parallel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 parameterize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 parse-c-struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 parse-header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 parse-http-method . . . . . . . . . . . . . . . . . . . . . . . . . . 525 parse-http-version . . . . . . . . . . . . . . . . . . . . . . . . . 525 parse-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369

parse-path-with-ellipsis . . . . . . . . . . . . . . . . . . 369

parse-request-uri . . . . . . . . . . . . . . . . . . . . . . . . . . 525 partial-cont-call . . . . . . . . . . . . . . . . . . . . . . . . . . 784 partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559, 633 partition! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 passwd:dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:gecos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:gid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:passwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 passwd:uid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 pclose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 peek-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319, 644 peek-next-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 PKG_CHECK_MODULES. . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 pointer->bytevector . . . . . . . . . . . . . . . . . . . . . . . . 404 pointer->procedure . . . . . . . . . . . . . . . . . . . . . . . . . 407 pointer->scm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 pointer->string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 pointer-address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 pointer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 poll-coop-repl-server . . . . . . . . . . . . . . . . . . . . . . 374 popen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 port->fdes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 port->stream . . . . . . . . . . . . . . . . . . . . . . . . . . . 602, 683 port-closed? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 port-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

port-conversion-strategy . . . . . . . . . . . . . . . . . . 318

port-encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 port-eof?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 port-filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 port-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474

port-has-port-position? . . . . . . . . . . . . . . . . . . . 338
port-has-set-port-position!? . . . . . . . . . . . . . . 338

port-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 port-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329, 474 port-position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 338 port-revealed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 port-transcoder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 port-with-print-state . . . . . . . . . . . . . . . . . . . . . . 320 port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317, 337 positive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 628 post-order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 pre-post-order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 pretty-print. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 664 primitive-_exit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 primitive-eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 primitive-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 primitive-fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494

primitive-generic-generic . . . . . . . . . . . . . . . . . 728

primitive-load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 367 primitive-load-path . . . . . . . . . . . . . . . . . . . . . . . . 368 primitive-move->fdes . . . . . . . . . . . . . . . . . . . . . . . 471 print-options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 print-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 256 procedure->pointer . . . . . . . . . . . . . . . . . . . . . . . . . 409

                                                procedure-documentation . . . . . . . . . . . . . . . . . . . 255
                                                procedure-execution-count . . . . . . . . . . . . . . . . . 465

procedure-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 procedure-properties . . . . . . . . . . . . . . . . . . . . . . . 255 procedure-property . . . . . . . . . . . . . . . . . . . . . . . . . 255 procedure-source. . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 procedure-with-setter?. . . . . . . . . . . . . . . . . . . . . 256 procedure?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254, 625 profile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 program-arguments . . . . . . . . . . . . . . . . . . . . . . . . . . 488

                                                program-arguments-alist . . . . . . . . . . . . . . . . . . . 248

program-arities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 program-arity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 program-bindings. . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 program-free-variables. . . . . . . . . . . . . . . . . . . . . 246 program-lambda-list . . . . . . . . . . . . . . . . . . . . . . . . 248 program-meta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 program-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 program-objcode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 program-objects. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 program-sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 program? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 promise? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371, 614 prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 788 proper-list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 protoent:aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 protoent:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 protoent:proto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 provide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 provided?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 PTR2SCM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 push-rest. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 put-bytevector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 put-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 put-datum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 put-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 put-u8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 putenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 pwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490

Q

q-empty-check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 q-empty? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 q-front . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 q-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 q-pop! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 q-push! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 q-rear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 q-remove!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 q?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 quasiquote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358, 626 quasisyntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 493 quo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358, 626 quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 657

R

raise. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497, 583, 639 raise-continuable . . . . . . . . . . . . . . . . . . . . . . . . . . 639 random . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 random-integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 random-real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592

random-source-make-integers . . . . . . . . . . . . . . . 593
random-source-make-reals . . . . . . . . . . . . . . . . . . 593
random-source-pseudo-randomize! . . . . . . . . . . 592
random-source-randomize! . . . . . . . . . . . . . . . . . . 592
random-source-state-ref . . . . . . . . . . . . . . . . . . . 592
random-source-state-set! . . . . . . . . . . . . . . . . . . 592

random-source? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 random-state->datum . . . . . . . . . . . . . . . . . . . . . . . . 128

random-state-from-platform . . . . . . . . . . . . . . . . 128

random:exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 random:hollow-sphere! . . . . . . . . . . . . . . . . . . . . . . 128 random:normal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 random:normal-vector! . . . . . . . . . . . . . . . . . . . . . . 128 random:solid-sphere! . . . . . . . . . . . . . . . . . . . . . . . 128 random:uniform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 rational-valued?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 rational? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 627 rationalize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 627 re-export. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360, 644 read-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318, 644 read-client . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 read-delimited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 read-delimited!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 read-disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 read-enable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 read-hash-extend. . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 read-header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 read-headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 read-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 read-line! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 read-options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 read-request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 read-request-body . . . . . . . . . . . . . . . . . . . . . . . . . . 535 read-request-line . . . . . . . . . . . . . . . . . . . . . . . . . . 525 read-response . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 read-response-body . . . . . . . . . . . . . . . . . . . . . . . . . 537 read-response-line . . . . . . . . . . . . . . . . . . . . . . . . . 525 read-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 read-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323, 709 read-string! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 read-string!/partial . . . . . . . . . . . . . . . . . . . . . . . 324 read-text-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709

read-with-shared-structure . . . . . . . . . . . . . . . . 598

readdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 662 readline-disable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 readline-enable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 readline-options. . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 readline-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 662 readline-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 readlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 real->flonum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 real-part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 627 real-valued? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 real? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 627 rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 receive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 record-accessor . . . . . . . . . . . . . . . . . . . . . . . . 219, 637 record-constructor . . . . . . . . . . . . . . . . . . . . 219, 637

record-constructor-descriptor. . . . . . . . . . . . . 636

record-field-mutable? . . . . . . . . . . . . . . . . . . . . . . 638 record-modifier. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 record-mutator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 record-predicate . . . . . . . . . . . . . . . . . . . . . . . 219, 637 record-rtd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 638 record-type-descriptor . . . . . . . . . . . . . . . . 220, 636

record-type-descriptor? . . . . . . . . . . . . . . . . . . . 637
record-type-field-names . . . . . . . . . . . . . . . . . . . 638

record-type-fields . . . . . . . . . . . . . . . . . . . . . . . . . 220

record-type-generative? . . . . . . . . . . . . . . . . . . . 638

record-type-name . . . . . . . . . . . . . . . . . . . . . . . 220, 638 record-type-opaque? . . . . . . . . . . . . . . . . . . . . . . . . 638 record-type-parent . . . . . . . . . . . . . . . . . . . . . . . . . 638 record-type-sealed? . . . . . . . . . . . . . . . . . . . . . . . . 638 record-type-uid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 638 record? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219, 638 recv! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 recvfrom!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 redirect-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 reduce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 reduce-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 regexp-exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp-match? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 regexp-quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 regexp-substitute . . . . . . . . . . . . . . . . . . . . . . . . . . 352

regexp-substitute/global . . . . . . . . . . . . . . . . . . 353

regexp? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 release-arbiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 release-port-handle . . . . . . . . . . . . . . . . . . . . . . . . 471 reload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 reload-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 rem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116, 657 remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559, 633 remove! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559

remove-class-accessors! . . . . . . . . . . . . . . . . . . . 756

remove-hook! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 remp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 remq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 remv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 rename-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480

repl-default-option-set! . . . . . . . . . . . . . . . . . . . 52

replace-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 replace-titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 request-absolute-uri . . . . . . . . . . . . . . . . . . . . . . . 536 request-accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 request-accept-charset. . . . . . . . . . . . . . . . . . . . . 535

request-accept-encoding . . . . . . . . . . . . . . . . . . . 535
request-accept-language . . . . . . . . . . . . . . . . . . . 535

request-allow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 request-authorization . . . . . . . . . . . . . . . . . . . . . . 535 request-cache-control . . . . . . . . . . . . . . . . . . . . . . 535 request-connection . . . . . . . . . . . . . . . . . . . . . . . . . 535

request-content-encoding . . . . . . . . . . . . . . . . . . 535
request-content-language . . . . . . . . . . . . . . . . . . 535

request-content-length. . . . . . . . . . . . . . . . . . . . . 535

request-content-location . . . . . . . . . . . . . . . . . . 535

request-content-md5 . . . . . . . . . . . . . . . . . . . . . . . . 535 request-content-range . . . . . . . . . . . . . . . . . . . . . . 535 request-content-type . . . . . . . . . . . . . . . . . . . . . . . 536 request-date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-expect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-expires. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-headers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-if-match. . . . . . . . . . . . . . . . . . . . . . . . . . . . 536

request-if-modified-since . . . . . . . . . . . . . . . . . 536

request-if-none-match . . . . . . . . . . . . . . . . . . . . . . 536 request-if-range. . . . . . . . . . . . . . . . . . . . . . . . . . . . 536

request-if-unmodified-since . . . . . . . . . . . . . . . 536

request-last-modified . . . . . . . . . . . . . . . . . . . . . . 536 request-max-forwards . . . . . . . . . . . . . . . . . . . . . . . 536 request-meta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-pragma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536

request-proxy-authorization . . . . . . . . . . . . . . . 536

request-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-referer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-te . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-trailer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536

request-transfer-encoding . . . . . . . . . . . . . . . . . 536

request-upgrade. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-uri . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-user-agent . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 request-via . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request-warning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 request? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 require-extension . . . . . . . . . . . . . . . . . . . . . . . . . . 615 reserve-locals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 reset-hook! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283

reset-parsed-entity-definitions! . . . . . . . . . 703

resolve-interface . . . . . . . . . . . . . . . . . . . . . . . . . . 391 resolve-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 response-accept-ranges. . . . . . . . . . . . . . . . . . . . . 538 response-age . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-allow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-body-port . . . . . . . . . . . . . . . . . . . . . . . . . 537 response-cache-control. . . . . . . . . . . . . . . . . . . . . 538 response-code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 response-connection . . . . . . . . . . . . . . . . . . . . . . . . 538

                                                response-content-encoding . . . . . . . . . . . . . . . . . 538
                                                response-content-language . . . . . . . . . . . . . . . . . 538
                                                response-content-length . . . . . . . . . . . . . . . . . . . 538
                                                response-content-location . . . . . . . . . . . . . . . . . 538

response-content-md5 . . . . . . . . . . . . . . . . . . . . . . . 538 response-content-range. . . . . . . . . . . . . . . . . . . . . 538 response-content-type . . . . . . . . . . . . . . . . . . . . . . 538 response-date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-etag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-expires. . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-headers. . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 response-last-modified. . . . . . . . . . . . . . . . . . . . . 538 response-location . . . . . . . . . . . . . . . . . . . . . . . . . . 538

                                                response-must-not-include-body? . . . . . . . . . . 537

response-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 response-pragma. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538

                                                response-proxy-authenticate . . . . . . . . . . . . . . . 538

response-reason-phrase. . . . . . . . . . . . . . . . . . . . . 536 response-retry-after . . . . . . . . . . . . . . . . . . . . . . . 538 response-server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-trailer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 538

                                                response-transfer-encoding . . . . . . . . . . . . . . . . 538

response-upgrade. . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-vary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-version. . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 response-via . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 response-warning. . . . . . . . . . . . . . . . . . . . . . . . . . . . 538

                                                response-www-authenticate . . . . . . . . . . . . . . . . . 538

response?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 restore-signals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498

                                                restricted-vector-sort! . . . . . . . . . . . . . . . . . . . 280

return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 return/nvalues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 return/values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 return/values* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 628 reverse! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 reverse-bit-field . . . . . . . . . . . . . . . . . . . . . . . . . . 617 reverse-list->string . . . . . . . . . . . . . . . . . . . . . . . 144 reverse-list->vector . . . . . . . . . . . . . . . . . . . . . . . 613 reverse-vector->list . . . . . . . . . . . . . . . . . . . . . . . 613 rewinddir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 right-justify-string . . . . . . . . . . . . . . . . . . . . . . . 714 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 rotate-bit-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 629 round-ash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 round-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 round-remainder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 round/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 run-asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 run-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 run-server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373, 542

S

s16vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s16vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s16vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s16vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s16vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s16vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 s32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s32vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s32vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s32vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s32vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s32vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 s64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s64vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s64vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s64vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s64vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s64vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s8vector->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s8vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 s8vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 s8vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 s8vector?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 sanitize-response . . . . . . . . . . . . . . . . . . . . . . . . . . 541 save-module-excursion . . . . . . . . . . . . . . . . . . . . . . 391 scandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 678

scheme-report-environment . . . . . . . . . . . . 396, 657

scm->pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 scm-error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 scm_abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 scm_access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 scm_acons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_add_feature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 scm_add_hook_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_alarm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 scm_alignof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 scm_all_threads. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 scm_angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 scm_any_to_c32vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_c64vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_f32vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_f64vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_s16vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_s32vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_s64vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_s8vector . . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_u16vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_u32vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_u64vector . . . . . . . . . . . . . . . . . . . . . . . 575 scm_any_to_u8vector . . . . . . . . . . . . . . . . . . . . . . . . 575 scm_append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 scm_append_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 scm_apply. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_apply_0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_apply_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_apply_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_apply_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 SCM_ARG1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARG7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 SCM_ARGn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 scm_array_contents . . . . . . . . . . . . . . . . . . . . . . . . . 207 scm_array_copy_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_dimensions . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_fill_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_for_each . . . . . . . . . . . . . . . . . . . . . . . . . 205 scm_array_get_handle . . . . . . . . . . . . . . . . . . . . . . . 209

                                                scm_array_handle_bit_elements. . . . . . . . . . . . . 212
                                                scm_array_handle_bit_writable_elements . . 213
                                                scm_array_handle_c32_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_c32_writable_elements . . 212
                                                scm_array_handle_c64_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_c64_writable_elements . . 212

scm_array_handle_dims . . . . . . . . . . . . . . . . . . . . . . 209

                                                scm_array_handle_elements . . . . . . . . . . . . . . . . . 210
                                                scm_array_handle_f32_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_f32_writable_elements . . 212
                                                scm_array_handle_f64_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_f64_writable_elements . . 212

scm_array_handle_pos . . . . . . . . . . . . . . . . . . . . . . . 210 scm_array_handle_rank . . . . . . . . . . . . . . . . . . . . . . 209 scm_array_handle_ref . . . . . . . . . . . . . . . . . . . . . . . 210

                                                scm_array_handle_release . . . . . . . . . . . . . . . . . . 209
                                                scm_array_handle_s16_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_s16_writable_elements . . 211
                                                scm_array_handle_s32_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_s32_writable_elements . . 212
                                                scm_array_handle_s64_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_s64_writable_elements . . 212
                                                scm_array_handle_s8_elements . . . . . . . . . . . . . . 211
                                                scm_array_handle_s8_writable_elements . . . 211

scm_array_handle_set . . . . . . . . . . . . . . . . . . . . . . . 210

                                                scm_array_handle_u16_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_u16_writable_elements . . 211
                                                scm_array_handle_u32_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_u32_writable_elements . . 211
                                                scm_array_handle_u64_elements. . . . . . . . . . . . . 211
                                                scm_array_handle_u64_writable_elements . . 212
                                                scm_array_handle_u8_elements . . . . . . . . . . . . . . 211
                                                scm_array_handle_u8_writable_elements . . . 211
                                                scm_array_handle_uniform_element_size . . . 211
                                                scm_array_handle_uniform_elements . . . . . . . . 210
                                                scm_array_handle_uniform_writable_elements

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

                                                scm_array_handle_writable_elements . . . . . . . 210

scm_array_in_bounds_p . . . . . . . . . . . . . . . . . . . . . . 203 scm_array_index_map_x . . . . . . . . . . . . . . . . . . . . . . 205 scm_array_length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_map_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 scm_array_rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 scm_array_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 scm_array_to_list . . . . . . . . . . . . . . . . . . . . . . . . . . 204 scm_array_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 scm_ash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 SCM_ASSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 scm_assert_smob_type . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_ASSERT_TYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 scm_assoc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_assoc_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 scm_assoc_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . 232 scm_assoc_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_assq_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 scm_assq_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . . 232 scm_assq_set_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_assv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_assv_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 scm_assv_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . . 232 scm_assv_set_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 scm_async. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 scm_async_mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 scm_backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453

scm_backtrace_with_highlights. . . . . . . . . . . . . 453

scm_basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 scm_bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514

scm_bind_textdomain_codeset . . . . . . . . . . . . . . . 444

scm_bindtextdomain . . . . . . . . . . . . . . . . . . . . . . . . . 444 scm_bit_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 scm_bit_count_star . . . . . . . . . . . . . . . . . . . . . . . . . 200 scm_bit_extract. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 scm_bit_invert_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 scm_bit_position. . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 scm_bit_set_star_x . . . . . . . . . . . . . . . . . . . . . . . . . 199 scm_bitvector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_bitvector_elements. . . . . . . . . . . . . . . . . . . . . 200 scm_bitvector_fill_x . . . . . . . . . . . . . . . . . . . . . . . 199 scm_bitvector_length . . . . . . . . . . . . . . . . . . . . . . . 198 scm_bitvector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_bitvector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_bitvector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_bitvector_to_list . . . . . . . . . . . . . . . . . . . . . . 199

scm_bitvector_writable_elements . . . . . . . . . . 200

scm_boolean_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 scm_boot_guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

scm_broadcast_condition_variable . . . . . . . . . 417
scm_bytecode_to_objcode . . . . . . . . . . . . . . . . . . . 803
SCM_BYTEVECTOR_CONTENTS . . . . . . . . . . . . . . . . . . . 165

scm_bytevector_copy . . . . . . . . . . . . . . . . . . . . . . . . 165 scm_bytevector_copy_x . . . . . . . . . . . . . . . . . . . . . . 165 scm_bytevector_eq_p . . . . . . . . . . . . . . . . . . . . . . . . 165 scm_bytevector_fill_x . . . . . . . . . . . . . . . . . . . . . . 165

scm_bytevector_ieee_double_native_ref . . . 169
scm_bytevector_ieee_double_native_set_x

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

scm_bytevector_ieee_double_ref . . . . . . . . . . . 168

                                                scm_bytevector_ieee_double_set_x . . . . . . . . . 169
                                                scm_bytevector_ieee_single_native_ref . . . 169
                                                scm_bytevector_ieee_single_native_set_x

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

                                                scm_bytevector_ieee_single_ref . . . . . . . . . . . 168
                                                scm_bytevector_ieee_single_set_x . . . . . . . . . 169

scm_bytevector_length . . . . . . . . . . . . . . . . . . . . . . 164 SCM_BYTEVECTOR_LENGTH . . . . . . . . . . . . . . . . . . . . . . 165 scm_bytevector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

                                                scm_bytevector_s16_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_s16_native_set_x . . . . . . . . . . 167

scm_bytevector_s16_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_s16_set_x . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_s32_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_s32_native_set_x . . . . . . . . . . 167

scm_bytevector_s32_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_s32_set_x . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_s64_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_s64_native_set_x . . . . . . . . . . 167

scm_bytevector_s64_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_s64_set_x . . . . . . . . . . . . . . . . . . 167

scm_bytevector_s8_ref . . . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_s8_set_x . . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_sint_ref . . . . . . . . . . . . . . . . . . . 166
                                                scm_bytevector_sint_set_x . . . . . . . . . . . . . . . . . 166
                                                scm_bytevector_to_pointer . . . . . . . . . . . . . . . . . 404
                                                scm_bytevector_to_sint_list . . . . . . . . . . . . . . . 168
                                                scm_bytevector_to_u8_list . . . . . . . . . . . . . . . . . 168
                                                scm_bytevector_to_uint_list . . . . . . . . . . . . . . . 168
                                                scm_bytevector_u16_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_u16_native_set_x . . . . . . . . . . 167

scm_bytevector_u16_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_u16_set_x . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_u32_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_u32_native_set_x . . . . . . . . . . 167

scm_bytevector_u32_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_u32_set_x . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_u64_native_ref. . . . . . . . . . . . . 167
                                                scm_bytevector_u64_native_set_x . . . . . . . . . . 167

scm_bytevector_u64_ref. . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_u64_set_x . . . . . . . . . . . . . . . . . . 167

scm_bytevector_u8_ref . . . . . . . . . . . . . . . . . . . . . . 166

                                                scm_bytevector_u8_set_x . . . . . . . . . . . . . . . . . . . 167
                                                scm_bytevector_uint_ref . . . . . . . . . . . . . . . . . . . 166
                                                scm_bytevector_uint_set_x . . . . . . . . . . . . . . . . . 166

scm_c_angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_c_array_rank. . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

                                                scm_c_bind_keyword_arguments . . . . . . . . . . . . . . 183

scm_c_bitvector_length. . . . . . . . . . . . . . . . . . . . . 198 scm_c_bitvector_ref . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_c_bitvector_set_x . . . . . . . . . . . . . . . . . . . . . . 199

                                                scm_c_bytevector_length . . . . . . . . . . . . . . . . . . . 164

scm_c_bytevector_ref . . . . . . . . . . . . . . . . . . . . . . . 165 scm_c_bytevector_set_x. . . . . . . . . . . . . . . . . . . . . 165

                                                scm_c_call_with_blocked_asyncs . . . . . . . . . . . 412
                                                scm_c_call_with_current_module . . . . . . . . . . . 392
                                                scm_c_call_with_unblocked_asyncs . . . . . . . . . 412

scm_c_catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 scm_c_define . . . . . . . . . . . . . . . . . . . . . . . . . . . 287, 394 scm_c_define_gsubr . . . . . . . . . . . . . . . . . . . . . . . . . 245 scm_c_define_module . . . . . . . . . . . . . . . . . . . . . . . . 395 scm_c_downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 scm_c_eval_string . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_c_export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 scm_c_hook_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 scm_c_hook_init. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 scm_c_hook_remove . . . . . . . . . . . . . . . . . . . . . . . . . . 285 scm_c_hook_run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 scm_c_imag_part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

scm_c_locale_stringn_to_number . . . . . . . . . . . 118

scm_c_lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 scm_c_magnitude. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_c_make_bitvector . . . . . . . . . . . . . . . . . . . . . . . 198 scm_c_make_bytevector . . . . . . . . . . . . . . . . . . . . . . 164 scm_c_make_gsubr. . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 scm_c_make_polar. . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_c_make_rectangular. . . . . . . . . . . . . . . . . . . . . 119

scm_c_make_socket_address . . . . . . . . . . . . . . . . . 510

scm_c_make_string . . . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_c_make_struct . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_c_make_structv . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_c_make_vector . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_c_module_define . . . . . . . . . . . . . . . . . . . . . . . . 394 scm_c_module_lookup . . . . . . . . . . . . . . . . . . . . . . . . 394 scm_c_nvalues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 scm_c_port_for_each . . . . . . . . . . . . . . . . . . . . . . . . 474 scm_c_primitive_load . . . . . . . . . . . . . . . . . . . . . . . 367 scm_c_private_lookup . . . . . . . . . . . . . . . . . . . . . . . 393 scm_c_private_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_c_private_variable. . . . . . . . . . . . . . . . . . . . . 393 scm_c_public_lookup . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_c_public_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_c_public_variable . . . . . . . . . . . . . . . . . . . . . . 393 scm_c_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 scm_c_real_part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_c_resolve_module . . . . . . . . . . . . . . . . . . . . . . . 395 scm_c_round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_c_run_hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_c_string_length . . . . . . . . . . . . . . . . . . . . . . . . 145 scm_c_string_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 scm_c_string_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 147 scm_c_substring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 scm_c_substring_copy . . . . . . . . . . . . . . . . . . . . . . . 146

scm_c_substring_read_only . . . . . . . . . . . . . . . . . 146

scm_c_substring_shared. . . . . . . . . . . . . . . . . . . . . 146 scm_c_symbol_length . . . . . . . . . . . . . . . . . . . . . . . . 176 scm_c_titlecase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 scm_c_truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_c_upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 scm_c_use_module. . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 scm_c_value_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 scm_c_values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 scm_c_vector_length . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_c_vector_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_c_vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 196

scm_c_with_continuation_barrier . . . . . . . . . . 316

                                                scm_c_with_dynamic_state . . . . . . . . . . . . . . . . . . 422

scm_c_with_fluid. . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 scm_c_with_fluids . . . . . . . . . . . . . . . . . . . . . . . . . . 421

                                                scm_c_with_throw_handler . . . . . . . . . . . . . . . . . . 307

scm_c_write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 scm_c32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_c32vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_c32vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_c32vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_c32vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_c32vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_c32vector_to_list . . . . . . . . . . . . . . . . . . . . . . 572

                                                scm_c32vector_writable_elements . . . . . . . . . . 574

scm_c64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_c64vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_c64vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_c64vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_c64vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_c64vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_c64vector_to_list . . . . . . . . . . . . . . . . . . . . . . 572

                                                scm_c64vector_writable_elements . . . . . . . . . . 574

scm_caaaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caaadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cadaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cadadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cadar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cadddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_caddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 scm_call_n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364

                                                scm_call_with_blocked_asyncs . . . . . . . . . . . . . . 412
                                                scm_call_with_input_string . . . . . . . . . . . . . . . . 330
                                                scm_call_with_output_string . . . . . . . . . . . . . . . 330
                                                scm_call_with_unblocked_asyncs . . . . . . . . . . . 412

scm_calloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 376 scm_cancel_thread . . . . . . . . . . . . . . . . . . . . . . . . . . 414 scm_car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 SCM_CAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 scm_catch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

                                                scm_catch_with_pre_unwind_handler . . . . . . . . 305

scm_cdaaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdaadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cddaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cddadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cddar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cddddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scm_cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 SCM_CDR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 scm_ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_ceiling_divide . . . . . . . . . . . . . . . . . . . . . . . . . 121 scm_ceiling_quotient . . . . . . . . . . . . . . . . . . . . . . . 121 scm_ceiling_remainder . . . . . . . . . . . . . . . . . . . . . . 121 scm_cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 SCM_CELL_OBJECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771 SCM_CELL_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771 SCM_CELL_WORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771 scm_centered_divide . . . . . . . . . . . . . . . . . . . . . . . . 122 scm_centered_quotient . . . . . . . . . . . . . . . . . . . . . . 122 scm_centered_remainder. . . . . . . . . . . . . . . . . . . . . 122 scm_char_alphabetic_p . . . . . . . . . . . . . . . . . . . . . . 132 scm_char_downcase . . . . . . . . . . . . . . . . . . . . . . . . . . 133

scm_char_general_category . . . . . . . . . . . . . . . . . 132

scm_char_is_both_p . . . . . . . . . . . . . . . . . . . . . . . . . 132 scm_char_locale_ci_eq . . . . . . . . . . . . . . . . . . . . . . 438 scm_char_locale_ci_gt . . . . . . . . . . . . . . . . . . . . . . 437 scm_char_locale_ci_lt . . . . . . . . . . . . . . . . . . . . . . 437

scm_char_locale_downcase . . . . . . . . . . . . . . . . . . 438

scm_char_locale_gt . . . . . . . . . . . . . . . . . . . . . . . . . 437 scm_char_locale_lt . . . . . . . . . . . . . . . . . . . . . . . . . 437

scm_char_locale_titlecase . . . . . . . . . . . . . . . . . 438

scm_char_locale_upcase. . . . . . . . . . . . . . . . . . . . . 438 scm_char_lower_case_p . . . . . . . . . . . . . . . . . . . . . . 132 scm_char_numeric_p . . . . . . . . . . . . . . . . . . . . . . . . . 132 scm_char_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 scm_char_ready_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 scm_char_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 scm_char_set_adjoin . . . . . . . . . . . . . . . . . . . . . . . . 138 scm_char_set_adjoin_x . . . . . . . . . . . . . . . . . . . . . . 138 scm_char_set_any. . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

scm_char_set_complement . . . . . . . . . . . . . . . . . . . 138
scm_char_set_complement_x . . . . . . . . . . . . . . . . . 138
scm_char_set_contains_p . . . . . . . . . . . . . . . . . . . 137

scm_char_set_copy . . . . . . . . . . . . . . . . . . . . . . . . . . 135 scm_char_set_count . . . . . . . . . . . . . . . . . . . . . . . . . 137 scm_char_set_cursor . . . . . . . . . . . . . . . . . . . . . . . . 134

scm_char_set_cursor_next . . . . . . . . . . . . . . . . . . 134

scm_char_set_delete . . . . . . . . . . . . . . . . . . . . . . . . 138 scm_char_set_delete_x . . . . . . . . . . . . . . . . . . . . . . 138

scm_char_set_diff_plus_intersection . . . . . . 138
scm_char_set_diff_plus_intersection_x . . . 139
scm_char_set_difference . . . . . . . . . . . . . . . . . . . 138

                                                scm_char_set_difference_x . . . . . . . . . . . . . . . . . 139

scm_char_set_eq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_every . . . . . . . . . . . . . . . . . . . . . . . . . 137
scm_char_set_filter . . . . . . . . . . . . . . . . . . . . . . . . 136
scm_char_set_filter_x . . . . . . . . . . . . . . . . . . . . . . 136
scm_char_set_fold . . . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_for_each . . . . . . . . . . . . . . . . . . . . . . 135
scm_char_set_hash . . . . . . . . . . . . . . . . . . . . . . . . . . 134

                                                scm_char_set_intersection . . . . . . . . . . . . . . . . . 138
                                                scm_char_set_intersection_x . . . . . . . . . . . . . . . 139

scm_char_set_leq. . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_map. . . . . . . . . . . . . . . . . . . . . . . . . . . . 135
scm_char_set_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_size . . . . . . . . . . . . . . . . . . . . . . . . . . 137
scm_char_set_to_list . . . . . . . . . . . . . . . . . . . . . . . 137
scm_char_set_to_string. . . . . . . . . . . . . . . . . . . . . 137
scm_char_set_unfold . . . . . . . . . . . . . . . . . . . . . . . . 134
scm_char_set_unfold_x . . . . . . . . . . . . . . . . . . . . . . 135
scm_char_set_union . . . . . . . . . . . . . . . . . . . . . . . . . 138
scm_char_set_union_x . . . . . . . . . . . . . . . . . . . . . . . 139
scm_char_set_xor. . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
scm_char_set_xor_x . . . . . . . . . . . . . . . . . . . . . . . . . 139
scm_char_titlecase . . . . . . . . . . . . . . . . . . . . . . . . . 133
scm_char_to_integer . . . . . . . . . . . . . . . . . . . . . . . . 133
scm_char_upcase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133
scm_char_upper_case_p . . . . . . . . . . . . . . . . . . . . . . 132
scm_char_whitespace_p . . . . . . . . . . . . . . . . . . . . . . 132
scm_chdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_chmod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479
scm_chown. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479
scm_chroot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_close. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472
scm_close_fdes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472
scm_close_input_port . . . . . . . . . . . . . . . . . . . . . . . 321
scm_close_output_port . . . . . . . . . . . . . . . . . . . . . . 321
scm_close_port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321
scm_closedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481
scm_complex_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

                                                scm_condition_variable_p . . . . . . . . . . . . . . . . . . 417

scm_connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513
scm_cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186
scm_cons_source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449
scm_copy_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479
scm_copy_random_state . . . . . . . . . . . . . . . . . . . . . . 127
scm_copy_tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280

                                                SCM_CRITICAL_SECTION_END . . . . . . . . . . . . . . . . . . 419
                                                SCM_CRITICAL_SECTION_START . . . . . . . . . . . . . . . . 419

scm_crypt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518
scm_ctermid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499

                                                scm_current_dynamic_state . . . . . . . . . . . . . . . . . 422

scm_current_error_port. . . . . . . . . . . . . . . . . . . . . 326
scm_current_input_port. . . . . . . . . . . . . . . . . . . . . 325
scm_current_load_port . . . . . . . . . . . . . . . . . . . . . . 367
scm_current_module . . . . . . . . . . . . . . . . . . . . . . . . . 391

                                                scm_current_output_port . . . . . . . . . . . . . . . . . . . 325
                                                scm_current_processor_count . . . . . . . . . . . . . . . 496

scm_current_thread . . . . . . . . . . . . . . . . . . . . . . . . . 413

scm_current_time. . . . . . . . . . . . . . . . . . . . . . . . . . . . 485

scm_datum_to_random_state . . . . . . . . . . . . . . . . . 128

scm_define. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287, 394 SCM_DEFINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 scm_defined_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 scm_delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delete_file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 scm_delete_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delete1_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 scm_delq_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delq1_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 scm_delv_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_delv1_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 scm_denominator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_difference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_directory_stream_p. . . . . . . . . . . . . . . . . . . . . 480 scm_dirname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482

scm_display_application . . . . . . . . . . . . . . . . . . . 447

scm_display_backtrace . . . . . . . . . . . . . . . . . . . . . . 446

scm_display_backtrace_with_highlights . . . 446

scm_display_error . . . . . . . . . . . . . . . . . . . . . . . . . . 313 scm_divide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_done_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_done_malloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_double_cell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770

scm_doubly_weak_hash_table_p . . . . . . . . . . . . . . 379

scm_drain_input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 scm_dup_to_fdes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 scm_dup2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 scm_dynamic_call. . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 scm_dynamic_func. . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 scm_dynamic_link. . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 scm_dynamic_object_p . . . . . . . . . . . . . . . . . . . . . . . 397 scm_dynamic_pointer . . . . . . . . . . . . . . . . . . . . . . . . 403 scm_dynamic_state_p . . . . . . . . . . . . . . . . . . . . . . . . 421 scm_dynamic_unlink . . . . . . . . . . . . . . . . . . . . . . . . . 397 scm_dynamic_wind. . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 scm_dynwind_begin . . . . . . . . . . . . . . . . . . . . . . . . . . 312

scm_dynwind_block_asyncs . . . . . . . . . . . . . . . . . . 412
scm_dynwind_critical_section . . . . . . . . . . . . . . 419
scm_dynwind_current_dynamic_state . . . . . . . . 422
scm_dynwind_current_error_port . . . . . . . . . . . 326
scm_dynwind_current_input_port . . . . . . . . . . . 326
scm_dynwind_current_output_port . . . . . . . . . . 326

scm_dynwind_end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 scm_dynwind_fluid . . . . . . . . . . . . . . . . . . . . . . . . . . 421 scm_dynwind_free . . . . . . . . . . . . . . . . . . . . . . . 313, 377 scm_dynwind_lock_mutex. . . . . . . . . . . . . . . . . . . . . 416

scm_dynwind_rewind_handler . . . . . . . . . . . . . . . . 313
scm_dynwind_rewind_handler_with_scm . . . . . . 313
scm_dynwind_unblock_asyncs . . . . . . . . . . . . . . . . 412
scm_dynwind_unwind_handler . . . . . . . . . . . . . . . . 312
scm_dynwind_unwind_handler_with_scm . . . . . . 312

scm_effective_version . . . . . . . . . . . . . . . . . . . . . . 427 scm_end_of_char_set_p . . . . . . . . . . . . . . . . . . . . . . 134 scm_environ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 scm_eof_object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 scm_eof_object_p . . . . . . . . . . . . . . . . . . . . . . . 318, 336 SCM_EOF_VAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 SCM_EOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 scm_eq_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 scm_equal_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 scm_eqv_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 scm_error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_error_scm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 scm_euclidean_divide . . . . . . . . . . . . . . . . . . . . . . . 120 scm_euclidean_quotient. . . . . . . . . . . . . . . . . . . . . 120

                                                scm_euclidean_remainder . . . . . . . . . . . . . . . . . . . 120

scm_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 scm_eval_string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363

                                                scm_eval_string_in_module . . . . . . . . . . . . . . . . . 363

scm_even_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 scm_exact_integer_p . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_exact_integer_sqrt. . . . . . . . . . . . . . . . . . . . . 117 scm_exact_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 scm_exact_to_inexact . . . . . . . . . . . . . . . . . . . . . . . 114 scm_execl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_execle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_execlp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_f32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_f32vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_f32vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_f32vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_f32vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_f32vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_f32vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_f32vector_writable_elements . . . . . . . . . . 574

scm_f64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_f64vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_f64vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_f64vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_f64vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_f64vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_f64vector_to_list . . . . . . . . . . . . . . . . . . . . . . 572

                                                scm_f64vector_writable_elements . . . . . . . . . . 574

scm_fcntl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 scm_fdes_to_ports . . . . . . . . . . . . . . . . . . . . . . . . . . 471 scm_fdopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 scm_file_encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 371 scm_file_port_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 scm_fileno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 scm_finite_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_flock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 scm_floor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_floor_divide. . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 scm_floor_quotient . . . . . . . . . . . . . . . . . . . . . . . . . 121 scm_floor_remainder . . . . . . . . . . . . . . . . . . . . . . . . 121 scm_fluid_bound_p . . . . . . . . . . . . . . . . . . . . . . . . . . 420 scm_fluid_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 scm_fluid_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 scm_fluid_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 scm_fluid_unset_x . . . . . . . . . . . . . . . . . . . . . . . . . . 420 scm_flush_all_ports . . . . . . . . . . . . . . . . . . . . . . . . 321 scm_force. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 scm_force_output. . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 scm_fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_frame_arguments . . . . . . . . . . . . . . . . . . . . . . . . 447 scm_frame_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 scm_frame_previous . . . . . . . . . . . . . . . . . . . . . . . . . 447 scm_frame_procedure . . . . . . . . . . . . . . . . . . . . . . . . 447 scm_from_bool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 scm_from_char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_from_double. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_from_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_from_int16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_int32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_int64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_int8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_intmax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

scm_from_latin1_keyword . . . . . . . . . . . . . . . . . . . 183

scm_from_latin1_string. . . . . . . . . . . . . . . . . . . . . 162

scm_from_latin1_stringn . . . . . . . . . . . . . . . . . . . 162

scm_from_latin1_symbol. . . . . . . . . . . . . . . . . . . . . 176

scm_from_locale_keyword . . . . . . . . . . . . . . . . . . . 183
scm_from_locale_keywordn . . . . . . . . . . . . . . . . . . 183

scm_from_locale_string. . . . . . . . . . . . . . . . . . . . . 160

scm_from_locale_stringn . . . . . . . . . . . . . . . . . . . 160

scm_from_locale_symbol. . . . . . . . . . . . . . . . . . . . . 176

scm_from_locale_symboln . . . . . . . . . . . . . . . . . . . 176

scm_from_long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_long_long . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_mpz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_pointer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 scm_from_ptrdiff_t . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_schar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_from_short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

scm_from_signed_integer . . . . . . . . . . . . . . . . . . . 109

scm_from_size_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_sockaddr . . . . . . . . . . . . . . . . . . . . . . . . . . 511 scm_from_ssize_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_stringn. . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 scm_from_uchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_from_uint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_uint16. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_uint32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_uint64. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_uint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_uintmax. . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_ulong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_from_ulong_long . . . . . . . . . . . . . . . . . . . . . . . . 110

scm_from_unsigned_integer . . . . . . . . . . . . . . . . . 109

scm_from_ushort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_from_utf32_string . . . . . . . . . . . . . . . . . . . . . . 162 scm_from_utf32_stringn. . . . . . . . . . . . . . . . . . . . . 162 scm_from_utf8_keyword . . . . . . . . . . . . . . . . . . . . . . 183 scm_from_utf8_string . . . . . . . . . . . . . . . . . . . . . . . 162 scm_from_utf8_stringn . . . . . . . . . . . . . . . . . . . . . . 162 scm_from_utf8_symbol . . . . . . . . . . . . . . . . . . . . . . . 176 scm_fsync. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 scm_ftell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 scm_gc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 scm_gc_calloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377

scm_gc_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377

                                                scm_gc_live_object_stats . . . . . . . . . . . . . . . . . . 375

scm_gc_malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377

                                                scm_gc_malloc_pointerless . . . . . . . . . . . . . . . . . 377

scm_gc_mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 375
scm_gc_protect_object . . . . . . . . . . . . . . . . . . . . . . 375
scm_gc_realloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377

                                                scm_gc_register_allocation . . . . . . . . . . . . . . . . 377

scm_gc_stats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 375

                                                scm_gc_unprotect_object . . . . . . . . . . . . . . . . . . . 375

scm_gcd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116
scm_gensym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176
scm_geq_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117
scm_get_bytevector_all. . . . . . . . . . . . . . . . . . . . . 341
scm_get_bytevector_n . . . . . . . . . . . . . . . . . . . . . . . 340
scm_get_bytevector_n_x. . . . . . . . . . . . . . . . . . . . . 340

                                                scm_get_bytevector_some . . . . . . . . . . . . . . . . . . . 341
                                                scm_get_internal_real_time . . . . . . . . . . . . . . . . 488
                                                scm_get_internal_run_time . . . . . . . . . . . . . . . . . 488

scm_get_output_string . . . . . . . . . . . . . . . . . . . . . . 331
scm_get_print_state . . . . . . . . . . . . . . . . . . . . . . . . 320
scm_get_u8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340
scm_getaddrinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503
scm_getaffinity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495
scm_getcwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_getegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491
scm_getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489
scm_geteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491
scm_getgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491
scm_getgrgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485
scm_getgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_gethost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506
scm_gethostname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517
scm_getitimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498
scm_getlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485
scm_getnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507
scm_getpass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518
scm_getpeername. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514
scm_getpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492
scm_getpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_getppid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490
scm_getpriority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495
scm_getproto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508
scm_getpwuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484
scm_getserv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509
scm_getsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492
scm_getsockname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514
scm_getsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512
scm_gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443
scm_gettimeofday. . . . . . . . . . . . . . . . . . . . . . . . . . . . 485
scm_getuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491
SCM_GLOBAL_KEYWORD . . . . . . . . . . . . . . . . . . . . . . . . . 103
SCM_GLOBAL_SYMBOL . . . . . . . . . . . . . . . . . . . . . . . . . . 103
SCM_GLOBAL_VARIABLE . . . . . . . . . . . . . . . . . . . . . . . . 103

                                                SCM_GLOBAL_VARIABLE_INIT . . . . . . . . . . . . . . . . . . 103

scm_gmtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486
scm_gr_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117
scm_hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

scm_hash_clear_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hash_count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

scm_hash_create_handle_x . . . . . . . . . . . . . . . . . . 240

scm_hash_fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 scm_hash_for_each . . . . . . . . . . . . . . . . . . . . . . . . . . 240

scm_hash_for_each_handle . . . . . . . . . . . . . . . . . . 241

scm_hash_get_handle . . . . . . . . . . . . . . . . . . . . . . . . 240 scm_hash_map_to_list . . . . . . . . . . . . . . . . . . . . . . . 240 scm_hash_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hash_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hash_set_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hash_table_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

scm_hashq_create_handle_x . . . . . . . . . . . . . . . . . 240

scm_hashq_get_handle . . . . . . . . . . . . . . . . . . . . . . . 240 scm_hashq_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashq_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashq_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

scm_hashv_create_handle_x . . . . . . . . . . . . . . . . . 240

scm_hashv_get_handle . . . . . . . . . . . . . . . . . . . . . . . 240 scm_hashv_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashv_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashv_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239

scm_hashx_create_handle_x . . . . . . . . . . . . . . . . . 240

scm_hashx_get_handle . . . . . . . . . . . . . . . . . . . . . . . 240 scm_hashx_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashx_remove_x . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hashx_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 scm_hook_empty_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_hook_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_hook_to_list. . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 SCM_HOOKP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_imag_part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 SCM_IMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 scm_inet_aton . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inet_makeaddr . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inet_netof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inet_ntoa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inet_ntop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inet_pton . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_inexact_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 scm_inexact_to_exact . . . . . . . . . . . . . . . . . . . . . . . 114 scm_inf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_inf_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_init_guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 scm_input_port_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 scm_integer_expt. . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 scm_integer_length . . . . . . . . . . . . . . . . . . . . . . . . . 126 scm_integer_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 scm_integer_to_char . . . . . . . . . . . . . . . . . . . . . . . . 133

scm_interaction_environment . . . . . . . . . . . . . . . 362

scm_internal_catch . . . . . . . . . . . . . . . . . . . . . . . . . 305 scm_is_array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 scm_is_bitvector. . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_is_bool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 scm_is_bytevector . . . . . . . . . . . . . . . . . . . . . . . . . . 164 scm_is_complex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

scm_is_dynamic_state . . . . . . . . . . . . . . . . . . . . . . . 421
scm_is_eq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277
scm_is_exact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114
scm_is_exact_integer . . . . . . . . . . . . . . . . . . . . . . . 108
scm_is_false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105
scm_is_inexact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114
scm_is_integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108
scm_is_keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
scm_is_null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_is_number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106
scm_is_pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186
scm_is_rational. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112
scm_is_real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112
scm_is_signed_integer . . . . . . . . . . . . . . . . . . . . . . 108
scm_is_simple_vector . . . . . . . . . . . . . . . . . . . . . . . 196
scm_is_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143
scm_is_symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174
scm_is_true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105
scm_is_typed_array . . . . . . . . . . . . . . . . . . . . . . . . . 202

                                                scm_is_unsigned_integer . . . . . . . . . . . . . . . . . . . 108

scm_is_vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195
scm_isatty_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499
scm_join_thread. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413
scm_join_thread_timed . . . . . . . . . . . . . . . . . . . . . . 413
SCM_KEYWORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103
scm_keyword_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
scm_keyword_to_symbol . . . . . . . . . . . . . . . . . . . . . . 183
scm_kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496
scm_last_pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190
scm_lcm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116
scm_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190
scm_leq_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117
scm_less_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117
scm_link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480
scm_list_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_cdr_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 191
scm_list_copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190
scm_list_n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189
scm_list_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190
scm_list_set_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191
scm_list_tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190
scm_list_to_bitvector . . . . . . . . . . . . . . . . . . . . . . 199
scm_list_to_c32vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_c64vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_char_set . . . . . . . . . . . . . . . . . . . . . . . 135
scm_list_to_char_set_x. . . . . . . . . . . . . . . . . . . . . 135
scm_list_to_f32vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_f64vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_s16vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_s32vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_s64vector . . . . . . . . . . . . . . . . . . . . . . 572
scm_list_to_s8vector . . . . . . . . . . . . . . . . . . . . . . . 572

scm_list_to_typed_array . . . . . . . . . . . . . . . . . . . 203

scm_list_to_u16vector . . . . . . . . . . . . . . . . . . . . . . 572 scm_list_to_u32vector . . . . . . . . . . . . . . . . . . . . . . 572 scm_list_to_u64vector . . . . . . . . . . . . . . . . . . . . . . 572 scm_list_to_u8vector . . . . . . . . . . . . . . . . . . . . . . . 572 scm_listen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 scm_lnaof. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 scm_load_extension . . . . . . . . . . . . . . . . . . . . . . . . . 398 scm_load_objcode. . . . . . . . . . . . . . . . . . . . . . . . . . . . 803 scm_local_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 scm_locale_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437

scm_locale_string_to_inexact . . . . . . . . . . . . . . 439
scm_locale_string_to_integer . . . . . . . . . . . . . . 439

scm_localtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 scm_lock_mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 scm_lock_mutex_timed . . . . . . . . . . . . . . . . . . . . . . . 415 scm_logand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_logbit_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_logcount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 scm_logior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_lognot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_logtest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_lookahead_u8. . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 scm_lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 scm_loxor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 scm_lstat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 scm_macro_binding . . . . . . . . . . . . . . . . . . . . . . . . . . 275 scm_macro_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 scm_macro_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 scm_macro_transformer . . . . . . . . . . . . . . . . . . . . . . 275 scm_macro_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 scm_magnitude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 scm_major_version . . . . . . . . . . . . . . . . . . . . . . . . . . 427 scm_make_arbiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 scm_make_array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 scm_make_bitvector . . . . . . . . . . . . . . . . . . . . . . . . . 198 scm_make_bytevector . . . . . . . . . . . . . . . . . . . . . . . . 164 scm_make_c32vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_c64vector . . . . . . . . . . . . . . . . . . . . . . . . . 569

scm_make_condition_variable . . . . . . . . . . . . . . . 417
scm_make_custom_binary_input_port . . . . . . . . 339
scm_make_custom_binary_output_port . . . . . . . 344
scm_make_doubly_weak_hash_table . . . . . . . . . . 379

scm_make_dynamic_state. . . . . . . . . . . . . . . . . . . . . 421 scm_make_f32vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_f64vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_fluid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420

scm_make_fluid_with_default . . . . . . . . . . . . . . . 420

scm_make_guardian . . . . . . . . . . . . . . . . . . . . . . . . . . 380 scm_make_hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 scm_make_locale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 scm_make_mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415

scm_make_mutex_with_flags . . . . . . . . . . . . . . . . . 415

scm_make_polar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 scm_make_port_type . . . . . . . . . . . . . . . . . . . . . . . . . 347

scm_make_procedure_with_setter . . . . . . . . . . . 256

scm_make_program. . . . . . . . . . . . . . . . . . . . . . . . . . . . 804 scm_make_rectangular . . . . . . . . . . . . . . . . . . . . . . . 118

                                                scm_make_recursive_mutex . . . . . . . . . . . . . . . . . . 415

scm_make_regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 350 scm_make_s16vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_s32vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_s64vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_shared_array . . . . . . . . . . . . . . . . . . . . . . 206 scm_make_smob_type . . . . . . . . . . . . . . . . . . . . . . . . . 241

                                                scm_make_socket_address . . . . . . . . . . . . . . . . . . . 509

scm_make_soft_port . . . . . . . . . . . . . . . . . . . . . . . . . 331 scm_make_stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 scm_make_string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_make_struct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_make_struct_layout. . . . . . . . . . . . . . . . . . . . . 225 scm_make_symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 scm_make_typed_array . . . . . . . . . . . . . . . . . . . . . . . 202 scm_make_u16vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_u32vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_u64vector . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_make_unbound_fluid. . . . . . . . . . . . . . . . . . . . . 420

                                                scm_make_undefined_variable . . . . . . . . . . . . . . . 390

scm_make_variable . . . . . . . . . . . . . . . . . . . . . . . . . . 390 scm_make_vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194

                                                scm_make_weak_key_hash_table . . . . . . . . . . . . . . 379
                                                scm_make_weak_value_hash_table . . . . . . . . . . . 379

scm_make_weak_vector . . . . . . . . . . . . . . . . . . . . . . . 379 scm_malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 376 scm_map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 scm_markcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 scm_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 scm_memory_error. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_memq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 scm_memv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 scm_merge. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 scm_merge_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 scm_micro_version . . . . . . . . . . . . . . . . . . . . . . . . . . 427 scm_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_minor_version . . . . . . . . . . . . . . . . . . . . . . . . . . 427 scm_misc_error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_mkdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 scm_mknod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 scm_mkstemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 scm_mktime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 scm_module_define . . . . . . . . . . . . . . . . . . . . . . . . . . 394

                                                scm_module_ensure_local_variable . . . . . . . . . 394

scm_module_lookup . . . . . . . . . . . . . . . . . . . . . . . . . . 394

                                                scm_module_reverse_lookup . . . . . . . . . . . . . . . . . 395

scm_module_variable . . . . . . . . . . . . . . . . . . . . . . . . 394 scm_modulo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 scm_modulo_expt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 scm_must_calloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_must_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_must_malloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_must_realloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 scm_mutex_level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 scm_mutex_locked_p . . . . . . . . . . . . . . . . . . . . . . . . . 417 scm_mutex_owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 scm_mutex_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 scm_nan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_nan_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_native_endianness . . . . . . . . . . . . . . . . . . . . . . 164 scm_nconc2last . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 scm_negative_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 scm_new_double_smob . . . . . . . . . . . . . . . . . . . . . . . . 243 scm_new_smob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 scm_newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 scm_ngettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 scm_nice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 scm_not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 scm_null_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 scm_num_eq_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 scm_num_overflow. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_number_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 scm_number_to_string . . . . . . . . . . . . . . . . . . . . . . . 118 scm_numerator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_objcode_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 803

scm_objcode_to_bytecode . . . . . . . . . . . . . . . . . . . 804

scm_object_properties . . . . . . . . . . . . . . . . . . . . . . 278 scm_object_property . . . . . . . . . . . . . . . . . . . . . . . . 279 scm_object_to_string . . . . . . . . . . . . . . . . . . . . . . . 281 scm_odd_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 scm_oneminus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_oneplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471

scm_open_bytevector_input_port . . . . . . . . . . . 339
scm_open_bytevector_output_port . . . . . . . . . . 344

scm_open_fdes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 scm_open_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326

scm_open_file_with_encoding . . . . . . . . . . . . . . . 326

scm_open_input_string . . . . . . . . . . . . . . . . . . . . . . 330 scm_open_output_string. . . . . . . . . . . . . . . . . . . . . 330 scm_opendir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 scm_out_of_range. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_output_port_p . . . . . . . . . . . . . . . . . . . . . . . . . . 317 SCM_PACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 scm_pair_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 scm_parse_path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369

scm_parse_path_with_ellipsis . . . . . . . . . . . . . . 369

scm_pause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 scm_peek_char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 scm_permanent_object . . . . . . . . . . . . . . . . . . . . . . . 375 scm_pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 scm_pointer_address . . . . . . . . . . . . . . . . . . . . . . . . 403

scm_pointer_to_bytevector . . . . . . . . . . . . . . . . . 404
scm_pointer_to_procedure . . . . . . . . . . . . . . . . . . 407

scm_port_closed_p . . . . . . . . . . . . . . . . . . . . . . . . . . 321 scm_port_column. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

scm_port_conversion_strategy . . . . . . . . . . . . . . 318

scm_port_encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 317 scm_port_filename . . . . . . . . . . . . . . . . . . . . . . . . . . 329 scm_port_for_each . . . . . . . . . . . . . . . . . . . . . . . . . . 474 scm_port_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 scm_port_mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 scm_port_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 scm_port_revealed . . . . . . . . . . . . . . . . . . . . . . . . . . 470

                                                scm_port_with_print_state . . . . . . . . . . . . . . . . . 320

scm_positive_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 scm_primitive__exit . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_primitive_eval . . . . . . . . . . . . . . . . . . . . . . . . . 364 scm_primitive_exit . . . . . . . . . . . . . . . . . . . . . . . . . 494 scm_primitive_load . . . . . . . . . . . . . . . . . . . . . . . . . 367

                                                scm_primitive_load_path . . . . . . . . . . . . . . . . . . . 368
                                                scm_primitive_move_to_fdes . . . . . . . . . . . . . . . . 471

scm_private_lookup . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_private_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_private_variable . . . . . . . . . . . . . . . . . . . . . . . 393 scm_procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256

                                                scm_procedure_documentation . . . . . . . . . . . . . . . 255

scm_procedure_name . . . . . . . . . . . . . . . . . . . . . . . . . 254 scm_procedure_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

                                                scm_procedure_properties . . . . . . . . . . . . . . . . . . 255

scm_procedure_property. . . . . . . . . . . . . . . . . . . . . 255 scm_procedure_source . . . . . . . . . . . . . . . . . . . . . . . 254

                                                scm_procedure_to_pointer . . . . . . . . . . . . . . . . . . 409
                                                scm_procedure_with_setter_p . . . . . . . . . . . . . . . 256

scm_product . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 scm_program_arguments . . . . . . . . . . . . . . . . . . . . . . 488 scm_program_arities . . . . . . . . . . . . . . . . . . . . . . . . 247

                                                scm_program_free_variables . . . . . . . . . . . . . . . . 246

scm_program_meta. . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 scm_program_module . . . . . . . . . . . . . . . . . . . . . . . . . 246 scm_program_objcode . . . . . . . . . . . . . . . . . . . . . . . . 246 scm_program_objects . . . . . . . . . . . . . . . . . . . . . . . . 246 scm_program_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 scm_promise_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 SCM_PTAB_ENTRY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345

                                                scm_pthread_cond_timedwait . . . . . . . . . . . . . . . . 418

scm_pthread_cond_wait . . . . . . . . . . . . . . . . . . . . . . 418 scm_pthread_mutex_lock. . . . . . . . . . . . . . . . . . . . . 418 SCM_PTOBNUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 scm_public_lookup . . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_public_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_public_variable . . . . . . . . . . . . . . . . . . . . . . . . 393 scm_put_bytevector . . . . . . . . . . . . . . . . . . . . . . . . . 344 scm_put_u8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 scm_putenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 scm_quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 scm_raise. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 scm_random . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 scm_random_exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

                                                scm_random_hollow_sphere_x . . . . . . . . . . . . . . . . 128

scm_random_normal . . . . . . . . . . . . . . . . . . . . . . . . . . 128

                                                scm_random_normal_vector_x . . . . . . . . . . . . . . . . 128
                                                scm_random_solid_sphere_x . . . . . . . . . . . . . . . . . 128
                                                scm_random_state_from_platform . . . . . . . . . . . 128
                                                scm_random_state_to_datum . . . . . . . . . . . . . . . . . 128

scm_random_uniform . . . . . . . . . . . . . . . . . . . . . . . . . 128 scm_rational_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 scm_rationalize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 scm_read_char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 scm_read_delimited_x . . . . . . . . . . . . . . . . . . . . . . . 324 scm_read_hash_extend . . . . . . . . . . . . . . . . . . . . . . . 360 scm_read_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324

scm_read_string_x_partial . . . . . . . . . . . . . . . . . 324

scm_readdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 scm_readlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 scm_real_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 scm_real_part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 scm_realloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 376 scm_recv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 scm_recvfrom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 scm_redirect_port . . . . . . . . . . . . . . . . . . . . . . . . . . 473 scm_regexp_exec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 scm_regexp_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352 scm_release_arbiter . . . . . . . . . . . . . . . . . . . . . . . . 411 scm_remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

scm_remember_upto_here_1 . . . . . . . . . . . . . . . . . . 375
scm_remember_upto_here_2 . . . . . . . . . . . . . . . . . . 375

scm_remove_hook_x . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 scm_reset_hook_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_resolve_module . . . . . . . . . . . . . . . . . . . . . . . . . 391 scm_restore_signals . . . . . . . . . . . . . . . . . . . . . . . . 498

scm_restricted_vector_sort_x . . . . . . . . . . . . . . 280

scm_reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

scm_reverse_list_to_string . . . . . . . . . . . . . . . . 144

scm_reverse_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 scm_rewinddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 scm_rmdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 scm_round_ash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 scm_round_divide. . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 scm_round_number. . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_round_quotient . . . . . . . . . . . . . . . . . . . . . . . . . 123 scm_round_remainder . . . . . . . . . . . . . . . . . . . . . . . . 123 scm_run_asyncs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 scm_run_hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 scm_s16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_s16vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_s16vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s16vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_s16vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s16vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_s16vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

scm_s16vector_writable_elements . . . . . . . . . . 573

scm_s32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_s32vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_s32vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s32vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_s32vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s32vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_s32vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

scm_s32vector_writable_elements . . . . . . . . . . 573

scm_s64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_s64vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_s64vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s64vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_s64vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s64vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_s64vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_s64vector_writable_elements . . . . . . . . . . 574

scm_s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_s8vector_elements . . . . . . . . . . . . . . . . . . . . . . 573 scm_s8vector_length . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s8vector_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_s8vector_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_s8vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_s8vector_to_list . . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_s8vector_writable_elements . . . . . . . . . . . 573

scm_search_path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369

                                                scm_seed_to_random_state . . . . . . . . . . . . . . . . . . 128

scm_seek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 scm_select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 scm_send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 scm_sendfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 scm_sendto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 scm_set_car_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 scm_set_cdr_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 SCM_SET_CELL_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 772 SCM_SET_CELL_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . 771 SCM_SET_CELL_WORD . . . . . . . . . . . . . . . . . . . . . . . . . . 771

                                                scm_set_current_dynamic_state. . . . . . . . . . . . . 422
                                                scm_set_current_error_port . . . . . . . . . . . . . . . . 326
                                                scm_set_current_input_port . . . . . . . . . . . . . . . . 326

scm_set_current_module. . . . . . . . . . . . . . . . . . . . . 391

                                                scm_set_current_output_port . . . . . . . . . . . . . . . 326
                                                scm_set_object_properties_x . . . . . . . . . . . . . . . 278
                                                scm_set_object_property_x . . . . . . . . . . . . . . . . . 279

scm_set_port_close . . . . . . . . . . . . . . . . . . . . . . . . . 347 scm_set_port_column_x . . . . . . . . . . . . . . . . . . . . . . 320

                                                scm_set_port_conversion_strategy_x . . . . . . . 317
                                                scm_set_port_encoding_x . . . . . . . . . . . . . . . . . . . 317

scm_set_port_end_input. . . . . . . . . . . . . . . . . . . . . 348 scm_set_port_equalp . . . . . . . . . . . . . . . . . . . . . . . . 347

                                                scm_set_port_filename_x . . . . . . . . . . . . . . . . . . . 329

scm_set_port_flush . . . . . . . . . . . . . . . . . . . . . . . . . 348 scm_set_port_free . . . . . . . . . . . . . . . . . . . . . . . . . . 347

                                                scm_set_port_input_waiting . . . . . . . . . . . . . . . . 348

scm_set_port_line_x . . . . . . . . . . . . . . . . . . . . . . . . 320 scm_set_port_mark . . . . . . . . . . . . . . . . . . . . . . . . . . 347 scm_set_port_print . . . . . . . . . . . . . . . . . . . . . . . . . 347

                                                scm_set_port_revealed_x . . . . . . . . . . . . . . . . . . . 470

scm_set_port_seek . . . . . . . . . . . . . . . . . . . . . . . . . . 348 scm_set_port_truncate . . . . . . . . . . . . . . . . . . . . . . 348

                                                scm_set_procedure_properties_x . . . . . . . . . . . 255
                                                scm_set_procedure_property_x . . . . . . . . . . . . . . 255
                                                scm_set_program_arguments . . . . . . . . . . . . . . . . . 489
                                                scm_set_program_arguments_scm. . . . . . . . . . . . . 488

SCM_SET_SMOB_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SET_SMOB_DATA_2 . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SET_SMOB_DATA_3 . . . . . . . . . . . . . . . . . . . . . . . . 243 scm_set_smob_equalp . . . . . . . . . . . . . . . . . . . . . . . . 242 SCM_SET_SMOB_FLAGS . . . . . . . . . . . . . . . . . . . . . . . . . 243 scm_set_smob_free . . . . . . . . . . . . . . . . . . . . . . . . . . 242 scm_set_smob_mark . . . . . . . . . . . . . . . . . . . . . . . . . . 242 SCM_SET_SMOB_OBJECT . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SET_SMOB_OBJECT_2 . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SET_SMOB_OBJECT_3 . . . . . . . . . . . . . . . . . . . . . . 243 scm_set_smob_print . . . . . . . . . . . . . . . . . . . . . . . . . 242

scm_set_source_properties_x . . . . . . . . . . . . . . . 448
scm_set_source_property_x . . . . . . . . . . . . . . . . . 448
scm_set_struct_vtable_name_x . . . . . . . . . . . . . . 223
scm_set_thread_cleanup_x . . . . . . . . . . . . . . . . . . 414

scm_setaffinity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 scm_setegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 scm_seteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 scm_setgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 scm_setgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 scm_setgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 scm_sethost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 scm_sethostname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 scm_setitimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 scm_setlocale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 scm_setnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 scm_setpgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 scm_setpriority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 scm_setproto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 scm_setpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 scm_setserv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 scm_setsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 scm_setsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 scm_setuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 scm_setvbuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474

scm_shared_array_increments . . . . . . . . . . . . . . . 207
scm_shared_array_offset . . . . . . . . . . . . . . . . . . . 207

scm_shared_array_root . . . . . . . . . . . . . . . . . . . . . . 207 scm_shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 scm_shutdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 scm_sigaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497

scm_sigaction_for_thread . . . . . . . . . . . . . . . . . . 497
scm_signal_condition_variable. . . . . . . . . . . . . 417

scm_simple_format . . . . . . . . . . . . . . . . . . . . . . . . . . 320

SCM_SIMPLE_VECTOR_LENGTH . . . . . . . . . . . . . . . . . . 196

SCM_SIMPLE_VECTOR_REF . . . . . . . . . . . . . . . . . . . . . . 196 SCM_SIMPLE_VECTOR_SET . . . . . . . . . . . . . . . . . . . . . . 197

scm_sint_list_to_bytevector . . . . . . . . . . . . . . . 168

scm_sizeof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 scm_sleep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 scm_sloppy_assoc. . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 scm_sloppy_assq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 scm_sloppy_assv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 SCM_SMOB_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_DATA_2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_DATA_3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_FLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_OBJECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_OBJECT_2 . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_OBJECT_2_LOC . . . . . . . . . . . . . . . . . . . . . . 244 SCM_SMOB_OBJECT_3 . . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SMOB_OBJECT_3_LOC . . . . . . . . . . . . . . . . . . . . . . 244 SCM_SMOB_OBJECT_LOC . . . . . . . . . . . . . . . . . . . . . . . . 244 SCM_SMOB_PREDICATE . . . . . . . . . . . . . . . . . . . . . . . . . 243 SCM_SNARF_INIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 scm_socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 scm_socketpair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 scm_sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 scm_sort_list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 scm_sort_list_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 scm_sort_x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 scm_sorted_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 scm_source_properties . . . . . . . . . . . . . . . . . . . . . . 448 scm_source_property . . . . . . . . . . . . . . . . . . . . . . . . 448 scm_spawn_thread. . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 scm_stable_sort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 scm_stable_sort_x . . . . . . . . . . . . . . . . . . . . . . . . . . 280 scm_stack_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 scm_stack_length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 scm_stack_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 scm_stack_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 scm_stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 scm_status_exit_val . . . . . . . . . . . . . . . . . . . . . . . . 493 scm_status_stop_sig . . . . . . . . . . . . . . . . . . . . . . . . 493 scm_status_term_sig . . . . . . . . . . . . . . . . . . . . . . . . 493 scm_std_select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 scm_std_sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 scm_std_usleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 scm_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 scm_strftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 scm_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_string_any . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 scm_string_append . . . . . . . . . . . . . . . . . . . . . . . . . . 155

                                                scm_string_append_shared . . . . . . . . . . . . . . . . . . 155
                                                scm_string_bytes_per_char . . . . . . . . . . . . . . . . . 163

scm_string_capitalize . . . . . . . . . . . . . . . . . . . . . . 154

                                                scm_string_capitalize_x . . . . . . . . . . . . . . . . . . . 154

scm_string_ci_eq. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_ci_ge. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_ci_gt. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_ci_le. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_ci_lt. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_ci_neq . . . . . . . . . . . . . . . . . . . . . . . . . . 150

                                                scm_string_ci_to_symbol . . . . . . . . . . . . . . . . . . . 175

scm_string_compare . . . . . . . . . . . . . . . . . . . . . . . . . 149 scm_string_compare_ci . . . . . . . . . . . . . . . . . . . . . . 149 scm_string_concatenate. . . . . . . . . . . . . . . . . . . . . 155

                                                scm_string_concatenate_reverse . . . . . . . . . . . 155
                                                scm_string_concatenate_reverse_shared . . . 156
                                                scm_string_concatenate_shared. . . . . . . . . . . . . 156

scm_string_contains . . . . . . . . . . . . . . . . . . . . . . . . 153 scm_string_contains_ci. . . . . . . . . . . . . . . . . . . . . 153 scm_string_copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 scm_string_copy_x . . . . . . . . . . . . . . . . . . . . . . . . . . 148 scm_string_count. . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 scm_string_delete . . . . . . . . . . . . . . . . . . . . . . . . . . 158 scm_string_downcase . . . . . . . . . . . . . . . . . . . . . . . . 154 scm_string_downcase_x . . . . . . . . . . . . . . . . . . . . . . 154 scm_string_drop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_drop_right . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 scm_string_every. . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 scm_string_fill_x . . . . . . . . . . . . . . . . . . . . . . . . . . 147 scm_string_filter . . . . . . . . . . . . . . . . . . . . . . . . . . 158 scm_string_fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 scm_string_fold_right . . . . . . . . . . . . . . . . . . . . . . 157 scm_string_for_each . . . . . . . . . . . . . . . . . . . . . . . . 156

scm_string_for_each_index . . . . . . . . . . . . . . . . . 156

scm_string_ge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_gt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_index. . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 scm_string_index_right. . . . . . . . . . . . . . . . . . . . . 153 scm_string_join. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_string_le . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_length . . . . . . . . . . . . . . . . . . . . . . . . . . 145

scm_string_locale_ci_eq . . . . . . . . . . . . . . . . . . . 437
scm_string_locale_ci_gt . . . . . . . . . . . . . . . . . . . 437
scm_string_locale_ci_lt . . . . . . . . . . . . . . . . . . . 437
scm_string_locale_downcase . . . . . . . . . . . . . . . . 438

scm_string_locale_gt . . . . . . . . . . . . . . . . . . . . . . . 437 scm_string_locale_lt . . . . . . . . . . . . . . . . . . . . . . . 437

scm_string_locale_titlecase . . . . . . . . . . . . . . . 438
scm_string_locale_upcase . . . . . . . . . . . . . . . . . . 438

scm_string_lt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_string_map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 scm_string_map_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 scm_string_neq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

scm_string_normalize_nfc . . . . . . . . . . . . . . . . . . 151
scm_string_normalize_nfd . . . . . . . . . . . . . . . . . . 151
scm_string_normalize_nfkc . . . . . . . . . . . . . . . . . 151
scm_string_normalize_nfkd . . . . . . . . . . . . . . . . . 151

scm_string_null_p . . . . . . . . . . . . . . . . . . . . . . . . . . 143 scm_string_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 scm_string_pad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_pad_right . . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_prefix_ci_p. . . . . . . . . . . . . . . . . . . . . 152

scm_string_prefix_length . . . . . . . . . . . . . . . . . . 152
scm_string_prefix_length_ci . . . . . . . . . . . . . . . 152

scm_string_prefix_p . . . . . . . . . . . . . . . . . . . . . . . . 152 scm_string_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 scm_string_replace . . . . . . . . . . . . . . . . . . . . . . . . . 158 scm_string_reverse . . . . . . . . . . . . . . . . . . . . . . . . . 155 scm_string_reverse_x . . . . . . . . . . . . . . . . . . . . . . . 155 scm_string_rindex . . . . . . . . . . . . . . . . . . . . . . . . . . 152 scm_string_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 scm_string_skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 scm_string_skip_right . . . . . . . . . . . . . . . . . . . . . . 153 scm_string_split. . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 scm_string_suffix_ci_p. . . . . . . . . . . . . . . . . . . . . 153

scm_string_suffix_length . . . . . . . . . . . . . . . . . . 152
scm_string_suffix_length_ci . . . . . . . . . . . . . . . 152

scm_string_suffix_p . . . . . . . . . . . . . . . . . . . . . . . . 152 scm_string_tabulate . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_string_take. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_take_right . . . . . . . . . . . . . . . . . . . . . . 146 scm_string_titlecase . . . . . . . . . . . . . . . . . . . . . . . 155 scm_string_titlecase_x. . . . . . . . . . . . . . . . . . . . . 155 scm_string_to_char_set. . . . . . . . . . . . . . . . . . . . . 136

scm_string_to_char_set_x . . . . . . . . . . . . . . . . . . 136

scm_string_to_list . . . . . . . . . . . . . . . . . . . . . . . . . 144 scm_string_to_number . . . . . . . . . . . . . . . . . . . . . . . 118 scm_string_to_symbol . . . . . . . . . . . . . . . . . . . . . . . 174 scm_string_to_utf16 . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_string_to_utf32 . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_string_to_utf8 . . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_string_tokenize . . . . . . . . . . . . . . . . . . . . . . . . 158 scm_string_trim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 scm_string_trim_both . . . . . . . . . . . . . . . . . . . . . . . 147 scm_string_trim_right . . . . . . . . . . . . . . . . . . . . . . 147 scm_string_unfold . . . . . . . . . . . . . . . . . . . . . . . . . . 157

                                                scm_string_unfold_right . . . . . . . . . . . . . . . . . . . 157

scm_string_upcase . . . . . . . . . . . . . . . . . . . . . . . . . . 154 scm_string_upcase_x . . . . . . . . . . . . . . . . . . . . . . . . 154 scm_string_xcopy_x . . . . . . . . . . . . . . . . . . . . . . . . . 158 scm_strptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 scm_struct_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_struct_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_struct_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_struct_vtable . . . . . . . . . . . . . . . . . . . . . . . . . . 222 scm_struct_vtable_name. . . . . . . . . . . . . . . . . . . . . 223 scm_struct_vtable_p . . . . . . . . . . . . . . . . . . . . . . . . 224 scm_substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 scm_substring_copy . . . . . . . . . . . . . . . . . . . . 145, 146 scm_substring_downcase. . . . . . . . . . . . . . . . . . . . . 154

                                                scm_substring_downcase_x . . . . . . . . . . . . . . . . . . 154

scm_substring_fill_x . . . . . . . . . . . . . . . . . . . . . . . 147 scm_substring_hash . . . . . . . . . . . . . . . . . . . . . . . . . 150 scm_substring_hash_ci . . . . . . . . . . . . . . . . . . . . . . 151 scm_substring_move_x . . . . . . . . . . . . . . . . . . . . . . . 148

                                                scm_substring_read_only . . . . . . . . . . . . . . . . . . . 146

scm_substring_shared . . . . . . . . . . . . . . . . . . . . . . . 146 scm_substring_to_list . . . . . . . . . . . . . . . . . . . . . . 144 scm_substring_upcase . . . . . . . . . . . . . . . . . . . . . . . 154 scm_substring_upcase_x. . . . . . . . . . . . . . . . . . . . . 154 scm_sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

                                                scm_supports_source_properties_p . . . . . . . . . 448

SCM_SYMBOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 scm_symbol_fref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 scm_symbol_fset_x . . . . . . . . . . . . . . . . . . . . . . . . . . 177 scm_symbol_hash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 scm_symbol_interned_p . . . . . . . . . . . . . . . . . . . . . . 179 scm_symbol_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 scm_symbol_pref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 scm_symbol_pset_x . . . . . . . . . . . . . . . . . . . . . . . . . . 177 scm_symbol_to_keyword . . . . . . . . . . . . . . . . . . . . . . 183 scm_symbol_to_string . . . . . . . . . . . . . . . . . . . . . . . 174 scm_symlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 scm_sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 scm_sys_library_dir . . . . . . . . . . . . . . . . . . . . . . . . 428 scm_sys_make_void_port. . . . . . . . . . . . . . . . . . . . . 332

                                                scm_sys_package_data_dir . . . . . . . . . . . . . . . . . . 428
                                                scm_sys_search_load_path . . . . . . . . . . . . . . . . . . 368
                                                scm_sys_site_ccache_dir . . . . . . . . . . . . . . . . . . . 428

scm_sys_site_dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 scm_sys_string_dump . . . . . . . . . . . . . . . . . . . . . . . . 163 scm_syserror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_syserror_msg. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 scm_system_async_mark . . . . . . . . . . . . . . . . . . . . . . 411

                                                scm_system_async_mark_for_thread . . . . . . . . . 411

scm_system_star. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 scm_take_c32vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_c64vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_f32vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_f64vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_locale_string. . . . . . . . . . . . . . . . . . . . . 160

scm_take_locale_stringn . . . . . . . . . . . . . . . . . . . 160

scm_take_locale_symbol. . . . . . . . . . . . . . . . . . . . . 176

scm_take_locale_symboln . . . . . . . . . . . . . . . . . . . 176

scm_take_s16vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_s32vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_s64vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_s8vector . . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_u16vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_u32vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_u64vector . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_take_u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . 572 scm_tcgetpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 scm_tcsetpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 scm_textdomain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 scm_thread_cleanup . . . . . . . . . . . . . . . . . . . . . . . . . 414 scm_thread_exited_p . . . . . . . . . . . . . . . . . . . . . . . . 414 scm_thread_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 scm_throw. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 scm_thunk_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 scm_times. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 scm_tmpfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 scm_tmpnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 scm_to_bool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 scm_to_char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_char_set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 scm_to_double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 scm_to_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_int16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_int32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_int64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_int8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_intmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_locale_string . . . . . . . . . . . . . . . . . . . . . . . 160

scm_to_locale_stringbuf . . . . . . . . . . . . . . . . . . . 161

scm_to_locale_stringn . . . . . . . . . . . . . . . . . . . . . . 160 scm_to_long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_long_long. . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_mpz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 scm_to_pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 scm_to_ptrdiff_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_schar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_signed_integer . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_sockaddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 scm_to_ssize_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_uintmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_ulong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 scm_to_ulong_long . . . . . . . . . . . . . . . . . . . . . . . . . . 109

                                                scm_to_unsigned_integer . . . . . . . . . . . . . . . . . . . 109

scm_to_ushort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

                                                scm_total_processor_count . . . . . . . . . . . . . . . . . 496

scm_transpose_array . . . . . . . . . . . . . . . . . . . . . . . . 207 scm_truncate_divide . . . . . . . . . . . . . . . . . . . . . . . . 122 scm_truncate_file . . . . . . . . . . . . . . . . . . . . . . . . . . 322 scm_truncate_number . . . . . . . . . . . . . . . . . . . . . . . . 120 scm_truncate_quotient . . . . . . . . . . . . . . . . . . . . . . 122 scm_truncate_remainder. . . . . . . . . . . . . . . . . . . . . 122 scm_try_arbiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 scm_try_mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 scm_ttyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 scm_typed_array_p . . . . . . . . . . . . . . . . . . . . . . . . . . 202 scm_tzset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 scm_u16vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_u16vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_u16vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u16vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_u16vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u16vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_u16vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_u16vector_writable_elements . . . . . . . . . . 573

scm_u32vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_u32vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_u32vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u32vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_u32vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u32vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_u32vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_u32vector_writable_elements . . . . . . . . . . 573

scm_u64vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_u64vector_elements. . . . . . . . . . . . . . . . . . . . . 573 scm_u64vector_length . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u64vector_p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_u64vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u64vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_u64vector_to_list . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_u64vector_writable_elements . . . . . . . . . . 573
                                                scm_u8_list_to_bytevector . . . . . . . . . . . . . . . . . 168

scm_u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 scm_u8vector_elements . . . . . . . . . . . . . . . . . . . . . . 573 scm_u8vector_length . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u8vector_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 scm_u8vector_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 scm_u8vector_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 571 scm_u8vector_to_list . . . . . . . . . . . . . . . . . . . . . . . 571

                                                scm_u8vector_writable_elements . . . . . . . . . . . 573
                                                scm_ucs_range_to_char_set . . . . . . . . . . . . . . . . . 136
                                                scm_ucs_range_to_char_set_x . . . . . . . . . . . . . . . 136
                                                scm_uint_list_to_bytevector . . . . . . . . . . . . . . . 168

scm_umask. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 scm_uname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 SCM_UNBNDP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 SCM_UNDEFINED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 scm_unget_bytevector . . . . . . . . . . . . . . . . . . . . . . . 341

                                                scm_uniform_array_read_x . . . . . . . . . . . . . . . . . . 205
                                                scm_uniform_array_write . . . . . . . . . . . . . . . . . . . 205

scm_unlock_mutex. . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 scm_unlock_mutex_timed. . . . . . . . . . . . . . . . . . . . . 416 SCM_UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 scm_unread_char . . . . . . . . . . . . . . . . . . . . . . . . 319, 472 scm_unread_string . . . . . . . . . . . . . . . . . . . . . . . . . . 319 SCM_UNSPECIFIED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 scm_usleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 scm_utf16_to_string . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_utf32_to_string . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_utf8_to_string . . . . . . . . . . . . . . . . . . . . . . . . . 169 scm_utime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 scm_values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 SCM_VARIABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 scm_variable_bound_p . . . . . . . . . . . . . . . . . . . . . . . 390 SCM_VARIABLE_INIT . . . . . . . . . . . . . . . . . . . . . . . . . . 103 scm_variable_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 scm_variable_ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 scm_variable_set_x . . . . . . . . . . . . . . . . . . . . . . . . . 390 scm_variable_unset_x . . . . . . . . . . . . . . . . . . . . . . . 391 scm_vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 scm_vector_copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 scm_vector_elements . . . . . . . . . . . . . . . . . . . . . . . . 197 scm_vector_fill_x . . . . . . . . . . . . . . . . . . . . . . . . . . 196 scm_vector_length . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_vector_move_left_x. . . . . . . . . . . . . . . . . . . . . 196

scm_vector_move_right_x . . . . . . . . . . . . . . . . . . . 196

scm_vector_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_vector_ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_vector_set_x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 scm_vector_to_list . . . . . . . . . . . . . . . . . . . . . . . . . 194

scm_vector_writable_elements . . . . . . . . . . . . . . 197

scm_version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427

scm_wait_condition_variable . . . . . . . . . . . . . . . 417

scm_waitpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492

scm_weak_key_hash_table_p . . . . . . . . . . . . . . . . . 379
scm_weak_value_hash_table_p . . . . . . . . . . . . . . . 379

scm_weak_vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 scm_weak_vector_p . . . . . . . . . . . . . . . . . . . . . . . . . . 380 scm_weak_vector_ref . . . . . . . . . . . . . . . . . . . . . . . . 380 scm_weak_vector_set_x . . . . . . . . . . . . . . . . . . . . . . 380

scm_with_continuation_barrier. . . . . . . . . . . . . 315

scm_with_dynamic_state. . . . . . . . . . . . . . . . . . . . . 422 scm_with_fluid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 scm_with_fluids. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 scm_with_guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 scm_with_throw_handler. . . . . . . . . . . . . . . . . . . . . 306 scm_without_guile . . . . . . . . . . . . . . . . . . . . . . . . . . 418 scm_write_char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 scm_write_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 scm_write_objcode . . . . . . . . . . . . . . . . . . . . . . . . . . 804

scm_write_string_partial . . . . . . . . . . . . . . . . . . 325

scm_wrong_num_args . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_wrong_type_arg . . . . . . . . . . . . . . . . . . . . . . . . . 314 scm_wrong_type_arg_msg. . . . . . . . . . . . . . . . . . . . . 314 scm_xsubstring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 scm_zero_p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SCM2PTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770

script-stexi-documentation . . . . . . . . . . . . . . . . 717

sdocbook-flatten. . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 search-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 second . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 seconds->time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 seed->random-state . . . . . . . . . . . . . . . . . . . . . . . . . 128 seek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 select-kids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 sendfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 sendto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 serious-condition? . . . . . . . . . . . . . . . . . . . . 596, 641 serve-one-client. . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 servent:aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 servent:name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 servent:port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 servent:proto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626, 724 set-box! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622

                                                set-buffered-input-continuation?! . . . . . . . . 685

set-car! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 789 set-cdr! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 789

                                                set-current-dynamic-state . . . . . . . . . . . . . . . . . 422

set-current-error-port. . . . . . . . . . . . . . . . . . . . . 326 set-current-input-port. . . . . . . . . . . . . . . . . . . . . 326 set-current-module . . . . . . . . . . . . . . . . . . . . . . . . . 391

                                                set-current-output-port . . . . . . . . . . . . . . . . . . . 326

set-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 set-fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 set-object-properties!. . . . . . . . . . . . . . . . . . . . . 278 set-object-property! . . . . . . . . . . . . . . . . . . . . . . . 279 set-port-column!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

                                                set-port-conversion-strategy!. . . . . . . . . . . . . 317

set-port-encoding! . . . . . . . . . . . . . . . . . . . . 317, 370 set-port-filename! . . . . . . . . . . . . . . . . . . . . . . . . . 329 set-port-line! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 set-port-position! . . . . . . . . . . . . . . . . . . . . . . . . . 338 set-port-revealed! . . . . . . . . . . . . . . . . . . . . . . . . . 470

                                                set-procedure-properties! . . . . . . . . . . . . . . . . . 255
                                                set-procedure-property! . . . . . . . . . . . . . . . . . . . 255

set-program-arguments . . . . . . . . . . . . . . . . . . . . . . 488

                                                set-readline-input-port! . . . . . . . . . . . . . . . . . . 662
                                                set-readline-output-port! . . . . . . . . . . . . . . . . . 662

set-readline-prompt! . . . . . . . . . . . . . . . . . . . . . . . 662

                                                set-record-type-printer! . . . . . . . . . . . . . . . . . . 217

set-source-properties!. . . . . . . . . . . . . . . . . . . . . 448 set-source-property! . . . . . . . . . . . . . . . . . . . . . . . 448

                                                set-struct-vtable-name! . . . . . . . . . . . . . . . . . . . 223

set-symbol-property! . . . . . . . . . . . . . . . . . . . . . . . 177 set-thread-cleanup! . . . . . . . . . . . . . . . . . . . . . . . . 414 set-time-nanosecond! . . . . . . . . . . . . . . . . . . . . . . . 585 set-time-second!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 set-time-type! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 set-tm:gmtoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 set-tm:isdst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:mday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 set-tm:min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 set-tm:mon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 set-tm:wday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:yday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-tm:zone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 set-vm-trace-level! . . . . . . . . . . . . . . . . . . . . . . . . 457 setaffinity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 setegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 seteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 setgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 setgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 setgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 setgroups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 sethost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 sethostent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 sethostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 setitimer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 setlocale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 setnet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 setnetent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 setpgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 setpriority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 setproto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 setprotoent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 setpw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 setpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 setserv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 setservent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 setsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 setsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 setter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 setuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 setvbuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 seventh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 shallow-clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742

shared-array-increments . . . . . . . . . . . . . . . . . . . 207

shared-array-offset . . . . . . . . . . . . . . . . . . . . . . . . 207 shared-array-root . . . . . . . . . . . . . . . . . . . . . . . . . . 207 shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 shutdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 sigaction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497

signal-condition-variable . . . . . . . . . . . . . . . . . 417

simple-conditions . . . . . . . . . . . . . . . . . . . . . . . . . . 640 simple-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 sinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 sint-list->bytevector . . . . . . . . . . . . . . . . . . . . . . 168 sixth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 sizeof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 skip-until . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 skip-while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 sloppy-assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 sloppy-assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

sloppy-assv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

                                                slot-bound-using-class? . . . . . . . . . . . . . . . . . . . 740

slot-bound? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740

                                                slot-definition-accessor . . . . . . . . . . . . . . . . . . 738
                                                slot-definition-allocation . . . . . . . . . . . . . . . . 738

slot-definition-getter. . . . . . . . . . . . . . . . . . . . . 738

                                                slot-definition-init-form . . . . . . . . . . . . . . . . . 738
                                                slot-definition-init-keyword . . . . . . . . . . . . . . 738
                                                slot-definition-init-thunk . . . . . . . . . . . . . . . . 738
                                                slot-definition-init-value . . . . . . . . . . . . . . . . 738

slot-definition-name . . . . . . . . . . . . . . . . . . . . . . . 738

                                                slot-definition-options . . . . . . . . . . . . . . . . . . . 738

slot-definition-setter. . . . . . . . . . . . . . . . . . . . . 738

                                                slot-exists-using-class? . . . . . . . . . . . . . . . . . . 740

slot-exists? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740
slot-init-function . . . . . . . . . . . . . . . . . . . . . . . . . 738
slot-missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 741
slot-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . 721, 740, 790
slot-ref-using-class . . . . . . . . . . . . . . . . . . . . . . . 740
slot-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790
slot-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721, 740
slot-set-using-class! . . . . . . . . . . . . . . . . . . . . . . 741
slot-unbound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 741
sockaddr:addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510
sockaddr:fam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510
sockaddr:flowinfo . . . . . . . . . . . . . . . . . . . . . . . . . . 510
sockaddr:path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510
sockaddr:port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510
sockaddr:scopeid. . . . . . . . . . . . . . . . . . . . . . . . . . . . 510
socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511
socketpair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512
sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279
sort! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280
sort-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280
sort-list! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280
sorted? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279
source-properties . . . . . . . . . . . . . . . . . . . . . . . . . . 448
source-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448
source:addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247
source:column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247
source:file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247
source:line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247
span . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560
span! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560
spawn-coop-repl-server. . . . . . . . . . . . . . . . . . . . . 374
spawn-server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373

                                                split-and-decode-uri-path . . . . . . . . . . . . . . . . . 522

split-at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555
split-at!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555
sqrt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123, 627
SRV:send-reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704

                                                ssax:complete-start-tag . . . . . . . . . . . . . . . . . . . 703

ssax:make-elem-parser . . . . . . . . . . . . . . . . . . . . . . 703
ssax:make-parser. . . . . . . . . . . . . . . . . . . . . . . . . . . . 703
ssax:make-pi-parser . . . . . . . . . . . . . . . . . . . . . . . . 703
ssax:read-attributes . . . . . . . . . . . . . . . . . . . . . . . 703
ssax:read-cdata-body . . . . . . . . . . . . . . . . . . . . . . . 703
ssax:read-char-data . . . . . . . . . . . . . . . . . . . . . . . . 703

ssax:read-char-ref . . . . . . . . . . . . . . . . . . . . . . . . . 703 ssax:read-external-id . . . . . . . . . . . . . . . . . . . . . . 703 ssax:read-markup-token. . . . . . . . . . . . . . . . . . . . . 703

ssax:read-pi-body-as-string . . . . . . . . . . . . . . . 703
ssax:reverse-collect-str-drop-ws . . . . . . . . . 703

ssax:skip-internal-dtd. . . . . . . . . . . . . . . . . . . . . 703

ssax:uri-string->symbol . . . . . . . . . . . . . . . . . . . 703

ssax:xml->sxml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 stable-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 stable-sort! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 stack-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 stack-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 stack-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 stack? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 standard-error-port . . . . . . . . . . . . . . . . . . . . . . . . 343 standard-input-port . . . . . . . . . . . . . . . . . . . . . . . . 339 standard-output-port . . . . . . . . . . . . . . . . . . . . . . . 343 start-stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:atime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:atimensec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:blksize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:blocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:ctime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:ctimensec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:dev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:gid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:ino . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:mtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:mtimensec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:nlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:perms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:rdev. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 stat:size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 stat:uid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 statistics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 statprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697

statprof-accumulated-time . . . . . . . . . . . . . . . . . 695

statprof-active?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 695

statprof-call-data->stats . . . . . . . . . . . . . . . . . 696
statprof-call-data-calls . . . . . . . . . . . . . . . . . . 696
statprof-call-data-cum-samples . . . . . . . . . . . 696
statprof-call-data-name . . . . . . . . . . . . . . . . . . . 696
statprof-call-data-self-samples . . . . . . . . . . 696

statprof-display. . . . . . . . . . . . . . . . . . . . . . . . . . . . 696

statprof-display-anomolies . . . . . . . . . . . . . . . . 696
statprof-fetch-call-tree . . . . . . . . . . . . . . . . . . 696

statprof-fetch-stacks . . . . . . . . . . . . . . . . . . . . . . 696

statprof-fold-call-data . . . . . . . . . . . . . . . . . . . 696
statprof-proc-call-data . . . . . . . . . . . . . . . . . . . 696

statprof-reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 statprof-sample-count . . . . . . . . . . . . . . . . . . . . . . 695 statprof-start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695

statprof-stats-%-time-in-proc. . . . . . . . . . . . . 696

statprof-stats-calls . . . . . . . . . . . . . . . . . . . . . . . 696

statprof-stats-cum-secs-in-proc . . . . . . . . . . 696

                                                statprof-stats-cum-secs-per-call . . . . . . . . . 696
                                                statprof-stats-proc-name . . . . . . . . . . . . . . . . . . 696
                                                statprof-stats-self-secs-in-proc . . . . . . . . . 696
                                                statprof-stats-self-secs-per-call . . . . . . . . 696

statprof-stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 status:exit-val. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 status:stop-sig. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 status:term-sig. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 stexi->plain-text . . . . . . . . . . . . . . . . . . . . . . . . . . 716 stexi->shtml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 stexi->sxml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 stexi->texi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 stexi-extract-index . . . . . . . . . . . . . . . . . . . . . . . . 713

                                                stop-server-and-clients! . . . . . . . . . . . . . . . . . . 374

stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 stream->list . . . . . . . . . . . . . . . . . . . . . . . . . . . 602, 683 stream->list&length . . . . . . . . . . . . . . . . . . . . . . . . 684 stream->reversed-list . . . . . . . . . . . . . . . . . . . . . . 684

                                                stream->reversed-list&length . . . . . . . . . . . . . . 684

stream->vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 stream-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 stream-car. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600, 683 stream-cdr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600, 683 stream-concat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 stream-cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 stream-constant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 stream-drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 stream-drop-while . . . . . . . . . . . . . . . . . . . . . . . . . . 603 stream-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 stream-fold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603, 684 stream-for-each . . . . . . . . . . . . . . . . . . . . . . . . 603, 684 stream-from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 stream-iterate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604 stream-lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 stream-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604 stream-let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604 stream-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604, 684 stream-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 604 stream-null? . . . . . . . . . . . . . . . . . . . . . . . . . . . 600, 683 stream-of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 606 stream-pair? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 stream-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 606 stream-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-scan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-take . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-take-while . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-unfold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 stream-unfolds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 608 stream-zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 608 stream? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 strftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143, 628 string->bytevector . . . . . . . . . . . . . . . . . . . . 159, 336 string->char-set. . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 string->char-set! . . . . . . . . . . . . . . . . . . . . . . . . . . 136 string->date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 string->header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 string->keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 string->list . . . . . . . . . . . . . . . . . . . . . . . . . . . 144, 628 string->number . . . . . . . . . . . . . . . . . . . . . . . . . 118, 628 string->pointer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 string->symbol . . . . . . . . . . . . . . . . . . . . . . . . . 174, 624 string->uri . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 string->utf16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 string->utf32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 string->utf8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 string->wrapped-lines . . . . . . . . . . . . . . . . . . . . . . 716 string-any . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 string-append . . . . . . . . . . . . . . . . . . . . . . . . . . 155, 628 string-append/shared . . . . . . . . . . . . . . . . . . . . . . . 155 string-bytes-per-char . . . . . . . . . . . . . . . . . . . . . . 163 string-capitalize . . . . . . . . . . . . . . . . . . . . . . . . . . 154 string-capitalize! . . . . . . . . . . . . . . . . . . . . . . . . . 154 string-ci->symbol . . . . . . . . . . . . . . . . . . . . . . . . . . 175 string-ci-hash . . . . . . . . . . . . . . . . . . . . . . . . . 620, 654 string-ci< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci<= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 632 string-ci<> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 632 string-ci= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 632 string-ci> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci>= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string-ci>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 632 string-ci>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 632 string-compare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-compare-ci . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-concatenate . . . . . . . . . . . . . . . . . . . . . . . . . 155

string-concatenate-reverse . . . . . . . . . . . . . . . . 155
string-concatenate-reverse/shared . . . . . . . . 156
string-concatenate/shared . . . . . . . . . . . . . . . . . 156

string-contains. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 string-contains-ci . . . . . . . . . . . . . . . . . . . . . . . . . 153 string-copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 628 string-copy! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 string-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 string-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 string-downcase . . . . . . . . . . . . . . . . . . . . . . . . 154, 632 string-downcase!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 string-drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 string-drop-right . . . . . . . . . . . . . . . . . . . . . . . . . . 146 string-every . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 string-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 string-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 string-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 string-fold-right . . . . . . . . . . . . . . . . . . . . . . . . . . 157 string-foldcase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 string-for-each . . . . . . . . . . . . . . . . . . . . . . . . 156, 629 string-for-each-index . . . . . . . . . . . . . . . . . . . . . . 156 string-hash . . . . . . . . . . . . . . . . . . . . . . . . 150, 620, 654 string-hash-ci . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 string-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

string-index-right . . . . . . . . . . . . . . . . . . . . . . . . . 153
string-join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
string-length . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 628
string-locale-ci<? . . . . . . . . . . . . . . . . . . . . . . . . . 437
string-locale-ci=? . . . . . . . . . . . . . . . . . . . . . . . . . 437
string-locale-ci>? . . . . . . . . . . . . . . . . . . . . . . . . . 437
string-locale-downcase. . . . . . . . . . . . . . . . . . . . . 438

                                                string-locale-titlecase . . . . . . . . . . . . . . . . . . . 438

string-locale-upcase . . . . . . . . . . . . . . . . . . . . . . . 438
string-locale<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
string-locale>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437
string-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156
string-map! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156
string-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 350
string-normalize-nfc . . . . . . . . . . . . . . . . . . 151, 632
string-normalize-nfd . . . . . . . . . . . . . . . . . . 151, 632
string-normalize-nfkc . . . . . . . . . . . . . . . . . 151, 632
string-normalize-nfkd . . . . . . . . . . . . . . . . . 151, 632
string-null? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143
string-pad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146
string-pad-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . 146
string-prefix-ci? . . . . . . . . . . . . . . . . . . . . . . . . . . 152
string-prefix-length . . . . . . . . . . . . . . . . . . . . . . . 152

                                                string-prefix-length-ci . . . . . . . . . . . . . . . . . . . 152

string-prefix? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152
string-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145, 628
string-replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
string-reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155
string-reverse!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155
string-rindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152
string-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147
string-skip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153
string-skip-right . . . . . . . . . . . . . . . . . . . . . . . . . . 153
string-split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145
string-suffix-ci? . . . . . . . . . . . . . . . . . . . . . . . . . . 153
string-suffix-length . . . . . . . . . . . . . . . . . . . . . . . 152

                                                string-suffix-length-ci . . . . . . . . . . . . . . . . . . . 152

string-suffix? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152
string-tabulate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
string-take . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146
string-take-right . . . . . . . . . . . . . . . . . . . . . . . . . . 146
string-titlecase . . . . . . . . . . . . . . . . . . . . . . . 155, 632
string-titlecase! . . . . . . . . . . . . . . . . . . . . . . . . . . 155
string-tokenize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
string-trim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147
string-trim-both. . . . . . . . . . . . . . . . . . . . . . . . . . . . 147
string-trim-right . . . . . . . . . . . . . . . . . . . . . . . . . . 147
string-unfold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157
string-unfold-right . . . . . . . . . . . . . . . . . . . . . . . . 157
string-upcase . . . . . . . . . . . . . . . . . . . . . . . . . . 154, 632
string-upcase! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154
string-xcopy! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
string< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150
string<= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150
string<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 628
string<> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149
string<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 628
string= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

string=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 628 string> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string>= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 string>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 628 string>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 628 string? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143, 625 strptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 strtod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 struct-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222, 790 struct-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 struct-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 struct-vtable . . . . . . . . . . . . . . . . . . . . . . . . . . 222, 790 struct-vtable-name . . . . . . . . . . . . . . . . . . . . . . . . . 223 struct-vtable? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 struct? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222, 790 sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 sub1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 790 subr-call. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146, 628 substring-fill!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 substring-move!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 substring/copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 substring/read-only . . . . . . . . . . . . . . . . . . . . . . . . 146 substring/shared. . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 subtract-duration . . . . . . . . . . . . . . . . . . . . . . . . . . 585 subtract-duration! . . . . . . . . . . . . . . . . . . . . . . . . . 585

supports-source-properties? . . . . . . . . . . . . . . . 448

sxml->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 sxml->xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 sxml-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 sxml-match-let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 692 sxml-match-let*. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 692 sxpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 symbol->keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 symbol->string . . . . . . . . . . . . . . . . . . . . . . . . . 174, 624 symbol-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 symbol-fref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 symbol-fset! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 symbol-hash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 654 symbol-interned?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 symbol-pref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 symbol-prefix-proc . . . . . . . . . . . . . . . . . . . . . . . . . 383 symbol-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

symbol-property-remove! . . . . . . . . . . . . . . . . . . . 177

symbol-pset! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 symbol=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 symbol? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 624 symlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 sync-q! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264, 651 syntax->datum . . . . . . . . . . . . . . . . . . . . . . . . . . 266, 652 syntax-case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263, 651 syntax-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 syntax-local-binding . . . . . . . . . . . . . . . . . . . . . . . 269

syntax-locally-bound-identifiers . . . . . . . . . 270

syntax-module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 syntax-parameterize . . . . . . . . . . . . . . . . . . . . . . . . 273 syntax-rules . . . . . . . . . . . . . . . . . . . . . . . . . . . 258, 626 syntax-source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 syntax-violation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 syntax-violation-form . . . . . . . . . . . . . . . . . . . . . . 642

                                                syntax-violation-subform . . . . . . . . . . . . . . . . . . 642

syntax-violation? . . . . . . . . . . . . . . . . . . . . . . . . . . 642 system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 system* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 system-async-mark . . . . . . . . . . . . . . . . . . . . . . . . . . 411 system-error-errno . . . . . . . . . . . . . . . . . . . . . . . . . 469

                                                system-file-name-convention . . . . . . . . . . . . . . . 482

T

tail-apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 780 tail-call. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 780 tail-call/cc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 782 tail-call/nargs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 take . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 take! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 take-after . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 take-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 take-until . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 take-while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 take-while! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 627 tanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 tcgetpgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 tcsetpgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 tenth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

                                                terminated-thread-exception? . . . . . . . . . . . . . . 583

texi->stexi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 texi-command-depth . . . . . . . . . . . . . . . . . . . . . . . . . 711 texi-fragment->stexi . . . . . . . . . . . . . . . . . . . . . . . 711 text-content-type? . . . . . . . . . . . . . . . . . . . . . . . . . 538 textdomain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 textual-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 the-environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 third . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 thread-cleanup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 thread-exited? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414 thread-join! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 580 thread-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 thread-sleep! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 580 thread-specific. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 thread-specific-set! . . . . . . . . . . . . . . . . . . . . . . . 579 thread-start! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 thread-terminate! . . . . . . . . . . . . . . . . . . . . . . . . . . 580 thread-yield! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 thread? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413, 579 throw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 thunk? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 time->seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 time-difference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-difference!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-monotonic->date . . . . . . . . . . . . . . . . . . . . . . . 587

time-monotonic->time-tai . . . . . . . . . . . . . . . . . . 587
time-monotonic->time-tai! . . . . . . . . . . . . . . . . . 587
time-monotonic->time-utc . . . . . . . . . . . . . . . . . . 587
time-monotonic->time-utc! . . . . . . . . . . . . . . . . . 587

time-nanosecond. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-resolution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-second . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-tai->date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 time-tai->julian-day . . . . . . . . . . . . . . . . . . . . . . . 587

time-tai->modified-julian-day. . . . . . . . . . . . . 587
time-tai->time-monotonic . . . . . . . . . . . . . . . . . . 587
time-tai->time-monotonic! . . . . . . . . . . . . . . . . . 587

time-tai->time-utc . . . . . . . . . . . . . . . . . . . . . . . . . 587 time-tai->time-utc! . . . . . . . . . . . . . . . . . . . . . . . . 587 time-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time-utc->date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 time-utc->julian-day . . . . . . . . . . . . . . . . . . . . . . . 587

time-utc->modified-julian-day. . . . . . . . . . . . . 587
time-utc->time-monotonic . . . . . . . . . . . . . . . . . . 587
time-utc->time-monotonic! . . . . . . . . . . . . . . . . . 587

time-utc->time-tai . . . . . . . . . . . . . . . . . . . . . . . . . 587 time-utc->time-tai! . . . . . . . . . . . . . . . . . . . . . . . . 587 time<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 time? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582, 584 times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 tm:gmtoff. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 tm:isdst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:mday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 tm:min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 tm:mon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 tm:wday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:yday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tm:zone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 tmpfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 tmpnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 tms:clock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 tms:cstime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 tms:cutime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 tms:stime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 tms:utime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 toplevel-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 779 toplevel-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 779 total-processor-count . . . . . . . . . . . . . . . . . . . . . . 496 touch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

trace-calls-in-procedure . . . . . . . . . . . . . . . . . . 461
trace-calls-to-procedure . . . . . . . . . . . . . . . . . . 461
trace-instructions-in-procedure . . . . . . . . . . 461

tracepoint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 transcoded-port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 transcoder-codec. . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 transcoder-eol-style . . . . . . . . . . . . . . . . . . . . . . . 336

                                                transcoder-error-handling-mode . . . . . . . . . . . 336

transform-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 transpose-array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 trap-at-procedure-call. . . . . . . . . . . . . . . . . . . . . 458

                                                trap-at-procedure-ip-in-range. . . . . . . . . . . . . 459
                                                trap-at-source-location . . . . . . . . . . . . . . . . . . . 459
                                                trap-calls-in-dynamic-extent . . . . . . . . . . . . . . 459
                                                trap-calls-to-procedure . . . . . . . . . . . . . . . . . . . 459

trap-enabled? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 trap-frame-finish . . . . . . . . . . . . . . . . . . . . . . . . . . 459 trap-in-dynamic-extent. . . . . . . . . . . . . . . . . . . . . 459 trap-in-procedure . . . . . . . . . . . . . . . . . . . . . . . . . . 458

                                                trap-instructions-in-dynamic-extent . . . . . . 459
                                                trap-instructions-in-procedure . . . . . . . . . . . 459
                                                trap-matching-instructions . . . . . . . . . . . . . . . . 460

trap-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 322, 629 truncate-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 truncate-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . 121 truncate-remainder . . . . . . . . . . . . . . . . . . . . . . . . . 121 truncate-values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 781 truncate/. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 truncated-print. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 try-arbiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 try-mutex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 ttyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 typed-array? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 tzset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487

U

u16vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u16vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u16vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u16vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 u16vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u16vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 u32vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u32vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u32vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u32vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 u32vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u32vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 u64vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u64vector->list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u64vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 u64vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 u64vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u64vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 u8-list->bytevector . . . . . . . . . . . . . . . . . . . . . . . . 168 u8vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u8vector->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u8vector-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 u8vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570 u8vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 u8vector?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 ucs-range->char-set . . . . . . . . . . . . . . . . . . . . . . . . 136 ucs-range->char-set! . . . . . . . . . . . . . . . . . . . . . . . 136 uint-list->bytevector . . . . . . . . . . . . . . . . . . . . . . 168 umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 uname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 unbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622

uncaught-exception-reason . . . . . . . . . . . . . . . . . 583

uncaught-exception? . . . . . . . . . . . . . . . . . . . . . . . . 583 undefined-violation? . . . . . . . . . . . . . . . . . . . . . . . 642 unfold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 unfold-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 unget-bytevector. . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 uniform-array-read! . . . . . . . . . . . . . . . . . . . . . . . . 205 uniform-array-write . . . . . . . . . . . . . . . . . . . . . . . . 205 unless . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292, 634 unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 unlock-mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 unquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358, 626 unquote-splicing . . . . . . . . . . . . . . . . . . . . . . . 358, 626 unread-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319, 472 unread-string . . . . . . . . . . . . . . . . . . . . . . . . . . 319, 472 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 unsyntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 unsyntax-splicing . . . . . . . . . . . . . . . . . . . . . . . . . . 651 unwind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 unwind-fluids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 788 unzip1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 unzip2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 unzip3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 unzip4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 unzip5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 update-direct-method! . . . . . . . . . . . . . . . . . . . . . . 756

update-direct-subclass! . . . . . . . . . . . . . . . . . . . 756
update-instance-for-different-class . . . . . . 757

uri->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-decode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-encode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-fragment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 uri-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 uri-query. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 uri-scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 uri-userinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 uri? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 urlify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 use-modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 usleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 utf-16-codec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 utf-8-codec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 utf16->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 utf32->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 utf8->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 utime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 utsname:machine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 utsname:nodename. . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 utsname:release. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 utsname:sysname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 utsname:version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517

V

valid-header? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 value-history-enabled? . . . . . . . . . . . . . . . . . . . . . . 47 values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302, 631 variable-bound? . . . . . . . . . . . . . . . . . . . . . . . . 390, 780 variable-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . 390, 780 variable-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 780 variable-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 variable-unset!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 variable?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 vector . . . . . . . . . . . . . . . . . . . . . . . . . 194, 609, 630, 786 vector->list. . . . . . . . . . . . . . . . . . . . . . . 194, 613, 630 vector->stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 vector-any . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 612 vector-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 vector-binary-search . . . . . . . . . . . . . . . . . . . . . . . 612 vector-concatenate . . . . . . . . . . . . . . . . . . . . . . . . . 610 vector-copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 609 vector-copy! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 vector-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 vector-empty? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 vector-every . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 612 vector-fill!. . . . . . . . . . . . . . . . . . . . . . . 196, 613, 631 vector-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 vector-fold-right . . . . . . . . . . . . . . . . . . . . . . . . . . 611 vector-for-each . . . . . . . . . . . . . . . . . . . . . . . . 611, 630 vector-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 vector-index-right . . . . . . . . . . . . . . . . . . . . . . . . . 612 vector-length . . . . . . . . . . . . . . . . . . . . . 195, 610, 631 vector-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611, 630 vector-map! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 vector-move-left! . . . . . . . . . . . . . . . . . . . . . . . . . . 196 vector-move-right! . . . . . . . . . . . . . . . . . . . . . . . . . 196 vector-ref . . . . . . . . . . . . . . . . . . . . 195, 610, 631, 789 vector-reverse!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 vector-reverse-copy . . . . . . . . . . . . . . . . . . . . . . . . 610 vector-reverse-copy! . . . . . . . . . . . . . . . . . . . . . . . 613 vector-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789 vector-set! . . . . . . . . . . . . . . . . . . . . . . . . 195, 613, 631 vector-skip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 612 vector-skip-right . . . . . . . . . . . . . . . . . . . . . . . . . . 612 vector-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 633 vector-sort! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 634 vector-swap! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 vector-unfold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 vector-unfold-right . . . . . . . . . . . . . . . . . . . . . . . . 609 vector= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 610, 630 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 vhash-assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-assv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 vhash-consq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 vhash-consv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 vhash-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-delq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-delv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-fold* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-fold-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-foldq* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash-foldv* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 vhash? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 violation? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 vlist->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist-cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist-drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist-fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-fold-right. . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-null? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-take . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 vlist-unfold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist-unfold-right . . . . . . . . . . . . . . . . . . . . . . . . . 215 vlist? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214

vm-abort-continuation-hook . . . . . . . . . . . . . . . . 456

vm-apply-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 vm-next-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456

vm-pop-continuation-hook . . . . . . . . . . . . . . . . . . 456
vm-push-continuation-hook . . . . . . . . . . . . . . . . . 456
vm-restore-continuation-hook . . . . . . . . . . . . . . 456

vm-trace-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457 void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 789

W

wait-condition-variable . . . . . . . . . . . . . . . . . . . 417

waitpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 warning? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 weak-key-hash-table? . . . . . . . . . . . . . . . . . . . . . . . 379 weak-value-hash-table?. . . . . . . . . . . . . . . . . . . . . 379 weak-vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 weak-vector-ref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 weak-vector-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 weak-vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 when. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292, 634 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 who-condition? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 wind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 787 wind-fluids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 788 with-code-coverage . . . . . . . . . . . . . . . . . . . . . . . . . 464

                                                with-continuation-barrier . . . . . . . . . . . . . . . . . 315
                                                with-default-trap-handler . . . . . . . . . . . . . . . . . 462

with-dynamic-state . . . . . . . . . . . . . . . . . . . . . . . . . 422 with-ellipsis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 with-error-to-file . . . . . . . . . . . . . . . . . . . . . . . . . 328 with-exception-handler . . . . . . . . . . . . . . . . 582, 639 with-fluid* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 with-fluids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 with-fluids* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 with-input-from-file . . . . . . . . . . . . . . . . . . 328, 644 with-input-from-string. . . . . . . . . . . . . . . . . . . . . 330 with-mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 with-output-to-file . . . . . . . . . . . . . . . . . . . 328, 644 with-output-to-string . . . . . . . . . . . . . . . . . . . . . . 330 with-parameters*. . . . . . . . . . . . . . . . . . . . . . . . . . . . 599

                                                with-readline-completion-function . . . . . . . . 663
                                                with-ssax-error-to-port . . . . . . . . . . . . . . . . . . . 702

with-statprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 with-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 with-throw-handler . . . . . . . . . . . . . . . . . . . . . . . . . 306 write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361, 644, 743 write-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320, 644 write-client . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 write-header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 write-headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 write-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 write-objcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 804 write-request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 write-request-body . . . . . . . . . . . . . . . . . . . . . . . . . 535 write-request-line . . . . . . . . . . . . . . . . . . . . . . . . . 525 write-response . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 write-response-body . . . . . . . . . . . . . . . . . . . . . . . . 537 write-response-line . . . . . . . . . . . . . . . . . . . . . . . . 525 write-string/partial . . . . . . . . . . . . . . . . . . . . . . . 325

                                                write-with-shared-structure . . . . . . . . . . . . . . . 598

X

xcons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 xml->sxml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 xml-token-head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 xml-token-kind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 xml-token? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 xsubstring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

Y

yield . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 414

Z

zero? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 628 zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555

Variable Index 861

Variable Index

This is an alphabetical list of all the important variables and constants in Guile.

When looking for a particular variable or constant, please look under its Scheme name as well as under its C name. The C name can be constructed from the Scheme names by a simple transformation described in the section See Section 6.1 [API Overview], page 99.

%

%auto-compilation-options . . . . . . . . . . . . . . . . . 366

%default-port-encoding. . . . . . . . . . . . . . . . . . . . . 317 %global-locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 %guile-build-info . . . . . . . . . . . . . . . . . . . . . . . . . . 428 %host-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 %load-compiled-path . . . . . . . . . . . . . . . . . . . . . . . . 369 %load-extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 %load-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 367 %load-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 %null-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403

&

&condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 595 &error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 &message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 &serious . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596

*

*features* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 *random-state* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

*sdocbook->stexi-rules* . . . . . . . . . . . . . . . . . . . 711
*sdocbook-block-commands* . . . . . . . . . . . . . . . . . 711

<

<standard-vtable> . . . . . . . . . . . . . . . . . . . . . . . . . . 224

_IOFBF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 _IOLBF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 _IONBF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474

A

accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 accept-charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 accept-encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 accept-language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 accept-ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 after-gc-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 age . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 allow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 authorization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530

B

block-growth-factor . . . . . . . . . . . . . . . . . . . . . . . . 214

C

cache-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 char-set:ascii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:blank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:designated . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:digit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set:empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:full . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 char-set:graphic. . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:hex-digit . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:iso-control . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:letter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set:letter+digit . . . . . . . . . . . . . . . . . . . . . . 140 char-set:lower-case . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set:printing . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:punctuation . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 char-set:title-case . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set:upper-case . . . . . . . . . . . . . . . . . . . . . . . . 139 char-set:whitespace . . . . . . . . . . . . . . . . . . . . . . . . 140 connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 content-encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-language. . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-location. . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-md5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 content-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 current-reader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 367

D

date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402

E

etag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 expect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 expires . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529

F

F_DUPFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_GETFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_GETFL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_GETOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 F_SETFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_SETFL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 F_SETOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 FD_CLOEXEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475

file-name-separator-string . . . . . . . . . . . . . . . . 483

float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530

G

GUILE_AUTO_COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . 39 GUILE_HISTORY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 GUILE_INSTALL_LOCALE . . . . . . . . . . . . . . . . . . . . . . . . 39

GUILE_LOAD_COMPILED_PATH . . . . . . . . . . . . . . . . . . . 40

GUILE_LOAD_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 GUILE_STACK_SIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 GUILE_WARN_DEPRECATED . . . . . . . . . . . . . . . . . . . . . . . 40

H

HOME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530

I

if-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 if-modified-since . . . . . . . . . . . . . . . . . . . . . . . . . . 530 if-none-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 if-range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 if-unmodified-since . . . . . . . . . . . . . . . . . . . . . . . . 531 INADDR_ANY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 INADDR_BROADCAST. . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 INADDR_LOOPBACK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 int16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 int32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 int64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 int8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402

internal-time-units-per-second . . . . . . . . . . . 487

IP_ADD_MEMBERSHIP . . . . . . . . . . . . . . . . . . . . . . . . . . 513 IP_DROP_MEMBERSHIP . . . . . . . . . . . . . . . . . . . . . . . . . 513 IP_MULTICAST_IF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 IP_MULTICAST_TTL. . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 IPPROTO_IP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 IPPROTO_TCP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 IPPROTO_UDP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 ITIMER_PROF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 ITIMER_REAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 ITIMER_VIRTUAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499

L

last-modified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 LC_ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_COLLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_CTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_MESSAGES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_MONETARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_NUMERIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 LC_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 LOCK_EX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 LOCK_NB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 LOCK_SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 LOCK_UN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402

M

max-forwards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 MSG_DONTROUTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 MSG_OOB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 MSG_PEEK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515

O

O_APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 O_CREAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 O_RDONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 O_RDWR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 O_WRONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 472 OPEN_BOTH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 OPEN_READ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 OPEN_WRITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500

P

PF_INET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 PF_INET6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 PF_UNIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 PIPE_BUF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 pragma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 PRIO_PGRP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 PRIO_PROCESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 PRIO_USER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 proxy-authenticate . . . . . . . . . . . . . . . . . . . . . . . . . 532 proxy-authorization . . . . . . . . . . . . . . . . . . . . . . . . 531 ptrdiff_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402

R

R_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 referer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 regexp/basic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp/extended. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp/icase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp/newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp/notbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 regexp/noteol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 retry-after . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532

Variable Index 863

S

SA_NOCLDSTOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 SA_RESTART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 scm_after_gc_c_hook . . . . . . . . . . . . . . . . . . . . . . . . 286 scm_after_gc_hook . . . . . . . . . . . . . . . . . . . . . . . . . . 286 scm_after_sweep_c_hook. . . . . . . . . . . . . . . . . . . . . 286 scm_before_gc_c_hook . . . . . . . . . . . . . . . . . . . . . . . 285 scm_before_mark_c_hook. . . . . . . . . . . . . . . . . . . . . 285

scm_before_sweep_c_hook . . . . . . . . . . . . . . . . . . . 285

SCM_BOOL_F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 SCM_BOOL_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 SCM_C_HOOK_AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 SCM_C_HOOK_NORMAL . . . . . . . . . . . . . . . . . . . . . . . . . . 284 SCM_C_HOOK_OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 scm_char_set_ascii . . . . . . . . . . . . . . . . . . . . . . . . . 140 scm_char_set_blank . . . . . . . . . . . . . . . . . . . . . . . . . 140

scm_char_set_designated . . . . . . . . . . . . . . . . . . . 140

scm_char_set_digit . . . . . . . . . . . . . . . . . . . . . . . . . 139 scm_char_set_empty . . . . . . . . . . . . . . . . . . . . . . . . . 140 scm_char_set_full . . . . . . . . . . . . . . . . . . . . . . . . . . 141 scm_char_set_graphic . . . . . . . . . . . . . . . . . . . . . . . 140 scm_char_set_hex_digit. . . . . . . . . . . . . . . . . . . . . 140

scm_char_set_iso_control . . . . . . . . . . . . . . . . . . 140

scm_char_set_letter . . . . . . . . . . . . . . . . . . . . . . . . 139

scm_char_set_letter_and_digit. . . . . . . . . . . . . 140
scm_char_set_lower_case . . . . . . . . . . . . . . . . . . . 139

scm_char_set_printing . . . . . . . . . . . . . . . . . . . . . . 140

scm_char_set_punctuation . . . . . . . . . . . . . . . . . . 140

scm_char_set_symbol . . . . . . . . . . . . . . . . . . . . . . . . 140

scm_char_set_title_case . . . . . . . . . . . . . . . . . . . 139
scm_char_set_upper_case . . . . . . . . . . . . . . . . . . . 139
scm_char_set_whitespace . . . . . . . . . . . . . . . . . . . 140

scm_endianness_big . . . . . . . . . . . . . . . . . . . . . . . . . 164 scm_endianness_little . . . . . . . . . . . . . . . . . . . . . . 164 SCM_F_WIND_EXPLICITLY . . . . . . . . . . . . . . . . . . . . . . 312 scm_global_locale . . . . . . . . . . . . . . . . . . . . . . . . . . 437 scm_ptobs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 scm_t_int16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_int32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_int64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_int8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_intmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_uint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_uint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_uint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 scm_t_uintmax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

scm_vtable_index_layout . . . . . . . . . . . . . . . . . . . 223
scm_vtable_index_printer . . . . . . . . . . . . . . . . . . 223

SEEK_CUR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 SEEK_END . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 SEEK_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 SIGHUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 SIGINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SO_BROADCAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_DONTROUTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_KEEPALIVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_LINGER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_NO_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_OOBINLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_PRIORITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_RCVBUF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_REUSEADDR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_REUSEPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_SNDBUF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SO_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 SOCK_DGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 SOCK_RAW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 SOCK_RDM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 SOCK_SEQPACKET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 SOCK_STREAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 SOL_SOCKET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 ssize_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 standard-vtable-fields. . . . . . . . . . . . . . . . . . . . . 224 stream-null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599

T

te. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 texi-command-specs . . . . . . . . . . . . . . . . . . . . . . . . . 711 time-duration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 time-monotonic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 time-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 time-tai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 time-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 time-utc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 trailer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 transfer-encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 527

U

uint16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 uint32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 uint64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 uint8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 unsigned-int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 unsigned-long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 upgrade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 user-agent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532

V

vary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 via . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 vlist-null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 vtable-index-layout . . . . . . . . . . . . . . . . . . . . . . . . 223 vtable-index-printer . . . . . . . . . . . . . . . . . . . . . . . 223 vtable-offset-user . . . . . . . . . . . . . . . . . . . . . . . . . 225

W

W_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 WAIT_ANY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 WAIT_MYPGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 warning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 WNOHANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 WUNTRACED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 www-authenticate. . . . . . . . . . . . . . . . . . . . . . . . . . . . 533

X

X_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476

Type Index 865

Type Index

This is an alphabetical list of all the important data types defined in the Guile Programmers Manual.

(

(@. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 (@@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 (abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 (apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (const . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 (define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (dynlet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (dynref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (dynset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (dynwind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (fix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 (if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (lambda-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (let-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 (letrec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (letrec* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (lexical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 (primitive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 (prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 (set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796, 797 (toplevel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 (void) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796

<

<abort> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 <application> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <conditional> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <const> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <dynlet> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <dynref> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <dynset> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <dynwind>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <fix> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 <glil-bind> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-branch> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-call> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-const> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-kw-prelude> . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-label> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-lexical> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-module> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-mv-bind> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-mv-call> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-opt-prelude> . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-program> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 <glil-prompt> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-source> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-std-prelude> . . . . . . . . . . . . . . . . . . . . . . . . . 799 <glil-toplevel>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 801 <glil-unbind> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <glil-void> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 800 <lambda-case> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <lambda> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <let-values> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 799 <let> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <letrec> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <lexical-ref> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <lexical-set> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <module-ref> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <module-set> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <primitive-ref>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796 <prompt> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 798 <sequence> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <toplevel-define> . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <toplevel-ref> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <toplevel-set> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 797 <void> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 796

A

Alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Association Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

B

Booleans . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

C

Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 Complex numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

D

Date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526

E

ETag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 Exact numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

H

Hash Tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281

I

Inexact numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Integer numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

K

Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 KVList . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526

L

Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

N

Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

P

Pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 Parameter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422

Q

QList . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 Quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 Queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681

R

Rational numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 Real numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 Regular expressions . . . . . . . . . . . . . . . . . . . . . . . . 349

S

SCM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 scm_port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 scm_ptob_descriptor . . . . . . . . . . . . . . . . . . . . . . . . 345 scm_t_array_dim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 scm_t_array_handle . . . . . . . . . . . . . . . . . . . . . . . . . 209 scm_t_bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 scm_t_c_hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 scm_t_c_hook_function . . . . . . . . . . . . . . . . . . . . . . 285 scm_t_c_hook_type . . . . . . . . . . . . . . . . . . . . . . . . . . 284 scm_t_catch_body. . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 scm_t_catch_handler . . . . . . . . . . . . . . . . . . . . . . . . 306 scm_t_dynwind_flags . . . . . . . . . . . . . . . . . . . . . . . . 312 scm_t_signed_bits . . . . . . . . . . . . . . . . . . . . . . . . . . 100

scm_t_string_failed_conversion_handler . . 161

scm_t_wchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 scm_t_wind_flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 SList . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 sockaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 Socket address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 struct sockaddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

V

Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

R5RS Index 867

R5RS Index

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

-

- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

A

abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 acos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 asin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 assv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 atan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

B

boolean? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

C

call-with-current-continuation . . . . . . . . . . . 300

call-with-input-file . . . . . . . . . . . . . . . . . . . . . . . 328 call-with-output-file . . . . . . . . . . . . . . . . . . . . . . 328 call-with-values. . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ceiling-quotient. . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ceiling-remainder . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ceiling/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 centered-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . 119 centered-remainder . . . . . . . . . . . . . . . . . . . . . . . . . 119 centered/. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 char->integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 char-alphabetic?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char-ci<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char-ci<?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char-ci=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char-ci>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char-ci>?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char-downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 char-lower-case?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char-numeric? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char-ready? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 char-titlecase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 char-upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 char-upper-case?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char-whitespace?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 char<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 char? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 close-input-port. . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 close-output-port . . . . . . . . . . . . . . . . . . . . . . . . . . 321 complex? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 current-input-port . . . . . . . . . . . . . . . . . . . . . . . . . 325 current-output-port . . . . . . . . . . . . . . . . . . . . . . . . 325

D

delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 div . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 div-and-mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 div0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 div0-and-mod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 dynamic-wind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311

E

eof-object? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 eq? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 equal? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 eqv? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 euclidean-quotient . . . . . . . . . . . . . . . . . . . . . . . . . 119 euclidean-remainder . . . . . . . . . . . . . . . . . . . . . . . . 119 euclidean/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 even? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 exact->inexact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 exact? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 expt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

F

floor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 floor-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 floor-remainder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 floor/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 for-each . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371

G

gcd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

I

imag-part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 inexact->exact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 inexact? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 input-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 integer->char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 integer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

interaction-environment . . . . . . . . . . . . . . . . . . . 362

L

lcm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 list->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 list->symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 list->vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 list-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 list-tail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 list? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 366 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

M

magnitude. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 make-polar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 make-rectangular. . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 make-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 make-vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 memq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 memv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 mod0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 modulo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

N

negative?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 null? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 number->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 number? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

O

odd? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 open-input-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 open-output-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 output-port? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317

P

pair? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 peek-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 positive?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 procedure? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

Q

quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

R

rational?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 read-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 real-part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 real? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 remainder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 round-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 round-remainder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 round/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

S

set-car! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 set-cdr! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 string->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 string->number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 string->symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 string-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 string-ci<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-ci=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-ci>=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-ci>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 string-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 string-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 string-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 string-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 string-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 string<=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 149 string<? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 string=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 string>=?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 string>? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

R5RS Index 869

string? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 substring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 symbol->string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 symbol-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 symbol? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

T

tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 truncate-quotient . . . . . . . . . . . . . . . . . . . . . . . . . . 119 truncate-remainder . . . . . . . . . . . . . . . . . . . . . . . . . 119 truncate/. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

V

values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 vector->list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 vector-fill! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 vector-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 vector-ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 vector-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 vector? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

W

with-input-from-file . . . . . . . . . . . . . . . . . . . . . . . 329 with-output-to-file . . . . . . . . . . . . . . . . . . . . . . . . 329 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 write-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

Z

zero? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117